A Complete Bibliography of Publications in Computers and Mathematics with Applications (2010–2019)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

08 August 2024 Version 1.09

Title word cross-reference

(1 + 1) [UKI11, Ade16]. (1 + 3) [LL10a]. (1 + α) [BMA11]. (1, 1) [CBS18]. (2 + 1) [Abd18b, AA15, GMZ15, GTZ19, KKK15, KTK18, KT18a, LM19a, LCW19, LGZ19, LTJ^{+}16, zLYmL18, zLZ19, LWW19a, MYZ18, MCF18, ML19a, MGS^{+}14, MPfTX18, Osm18, RKSA18, RS18a, Ray18, RMY19, WZMY18, YTD^{+}18, YGS^{+}16, ZYSY17, ZTSC16, AK16, Ade16, DYH11, KSG11, Kuo18, LK15, LMMF17, MZM18, MK17, Ray17, RRO17, WWD18, XX10, YYLW19, YTS^{+}17, ZA15, ZZ11b, ZDM11]. (3 + 1) [AGT19, BM18a, BB08, CTSX16, CZ15, CZ18, DTYZ18, GZD18, GABC16, HLLM19, HTM18, Her19, HTY^{+}19, HTWS15, HYC18, KK19b, KK19c, LDL11, LM18a, LZL19, Liu18a, LZM18, LYC^{+}19, LWW19b, LXY19, Man18, Men18, PTZ19, QTW^{+}18, SR17a, SG10b, TZ15, TTX^{+}16, Üns18, YM17, YMLL18, YTL^{+}18, YHC18, ZSAN19, ZDZY17, ZLT18, Asl11, LD11b, LL16d, SR15, WTYZ17, ZM17a]. (3 + α) [LHD18]. (4 + 1) [AAEG17, TDXQ18]. (A)(C, α) [EÇ10]. (α, β) [ADA11, SJN10a, DSA09, Jun10]. (α, m) [ÖAK11]. (AXB, CXD) = (E, F) [ZWLZ18]. (C) [KT11a]. (C, 1) [Çan11a]. (ϵ_{γ}, ϵ_{γ} ∨ q_{δ}) [MZJ11]. (F, α, ρ, d) [GKS10]. (\frac{G′}{G}) [MCKM12]. (\frac{G′}{G}) [JKB11, GMZ15, KBA11, SG10b]. (H(·, ·), η) [WD10]. (h, q) [Sim10]. (∈, ∈ ∨q_{k}) [SM11]. (∈, ∈ ∨q) [ÖI12]. (∈, ∈ ∨q_{k}) [SJN10b].

(L, M) [Abb10]. (m, n) [Dav10]. (n + 1) [FYYT11]. (p, 6) [BFG11]. (p, 8) [BFG11]. (p, q) [BD11c]. (ϕ_{1}, ϕ_{2}) [WZF16]. (ψ, ϕ) [ĆSCD11, SS11c, AKS11, APS12]. (q, h) [Rah11a, Rah11b]. (Q, R) [MM11]. (R, S) [DH10b, XC16]. (t, n) [ZPWY12]. (T, S) [SR12]. (w/g) [Gep16]. + [Laz10]. 1 [DPM15, GDM13, KPG18, OVV^{+}16, Oru17, WFY17, Was13, WY15]. 1024 [CSCM13]. 13 [DDD10]. 14 [ZSW15]. 18 [ZSW15]. 1 < α < 2 [SW12]. 2 [ATO19, ALHZMC^{+}19, AY18, BMRA10, BPC17, BP18, BR12d, BC17, CKMR11, CLL19, CL15, DN18b, DA18c, DFP^{+}13, EAEH18, FST19, FVVS16, FF15, FES^{+}19, Gha17, GR19a, HHS^{+}17, HHGA19, KYR15, KYW^{+}18, KKVS19, KPG18, KV18, LXZ18, LL19b, MLL16, Ma18b, MDG19, Oru17, PLMS14, RCG15, SJL^{+}19, SHH16, SEM13, SAU11, Wan18, WZC^{+}19b, Wen18b, WFC16, Ye16, Ye17b, YKKS10, YT13, YWT18, YQ18, ZN18, Zha11b, ZZLB18, ZZG19, ZWH^{+}19, ZZL15, dVLV18]. 2D + t [GV11a]. 2\frac{1}{2}D [LL16c]. 2n [Yan11d]. 3 [AY18, AT17, BOY12, BGF15, BGM19b, CYP16, CCHG17, CL12b, CM19b, DWZ13, DLZ19, DFP^{+}13, Ers16, FZ17, FSTN18, FSZ18, FES^{+}19, FMPR15, GGLP15, GVSP12, Gha17, GIMZ14, JCF19, KLK15, KN11, KK19a, LZ18a, Li16b, LL19b, Li19b, LZG13, Liu18c, LB18, LRBA15, MK18, MM18b, MJWD19, MKS13, MSZG17, PAE^{+}12, PDHL12, QCG15, RCM11, SPT17, SWL16, SLKK19, Tod13, TM10, WFY17, WD16, Wei12b, Wen18a, WH11b, WCH13, XZ17, YSW16, Yan19, YZS18, ZB19, Zbo19, ZZ16b, ZaY17, ZZ18b, ZC11a, ZS18]. 3^{∗} [TKH10]. 4 [CS14b, DNZ^{+}13, QHT16, Wu18a]. 512 [CSCM13]. 6 [YC12]. 8 [SISH12]. 9 [FVVS16]. ^{+} [MGY11, ZMG10]. _{2} [ZAK18b]. 3 [ZAK18b]. _{4} [GSI19]. _{6} [GSI19]. A [DD10, Zha11a]. a, b ∈ R [dS15]. A_{T,S}^{(2)} [SC13]. A_{1}XA_{1}^{∗} = B_{1} [ZLZ11]. A_{2}XA_{2}^{∗} = B_{2} [ZLZ11]. A_{i}XB_{i} = F_{i} [DLD10, SS17]. α [BL17, BNP18, LY17, TMMASG10, YY12, YL18a]. antitridiag_{n}(a, 0, b) [dS15]. ARCH(∞) [AD12]. AX + XB = C [WLM13]. AX = B [LZGZ11, XC17, TTZW18]. AXAH + CY CH = F [XCS18]. AXB + CXD = E [ZMLZ16]. AXB = C [LGG12, WLD13, ZT13]. B [CZ15, LZL19, HS11b]. B(r, s, t) [FBB10]. B[X; \frac{1}{2^{2}} Z_{0}] [SKdA11]. BDM_{k} [Erv12]. BL [ZJ10a]. bv_{p}, (1 < p < ∞) [FBB10]. C [HLS11, LKCN19, SKST10, AES11, JW19b, SCK11]. C^{0} [SCGW18, WZH18]. C^{1} [HH10b, LV11, PGF18, ST16]. C^{2} [KV17a, KV17b]. C^{H}A^{−1}B [DFS14]. CD4^{+} [EOM11]. CI [SR12]. d [CKN11, SMBY10, WWW14, YZ19]. Δ [Çak11b, KA11, Bor11, SPL19]. Δ_{a,b} [AES11]. Ḃ_{∞,∞}^{−1} [GLR13]. e_{2} [Mah11]. ℓ^{1} [Gos10]. ℓ_{p} [FBB10]. ϵ [YjH18]. η [Ant10]. F [MS11a, Dra10]. F W C [LZ12b]. G [Ant14, ASV11, APS12, Çak11c, MAK12, PT11, WY11b, ZLLF12]. Γ [AMD10, DZK10, DZY11, JMLF11, MZLF10, SDM10]. γ_{×2} [WS10a]. GF(p^{m}) [CKK^{+}10]. H [DFG19, TM12, Tom13, Ver08, Ver12, ZFZQ10, DSA10, DHMU16, MYZ12, PGF18, Pas14, PPC13, PPC15, RZ17, VAK^{+}19]. H(α, β) [CH11b]. H^{1} [BS17, Hou15, SW16a, ZS13, SZL^{+}17, WLL^{+}18]. H^{1}(R^{N}) [LC10d]. H^{2} [JZL18]. H_{∞} [Ahn12, LZT11, LCYC12, WS12]. H_{v} [DZY11]. hk [SD18]. hp [BS14a, BS15b, BS15a, BGGCGRSP16, BS15c, CDG15, CW15b, CHH14, CX18, FWW14, FMPR15, GSZ14, Hak14, HW16, JZ13, KM15, ZD15, BPS19, BDM^{+}19b, CGGM19, CNSV17, DESV18, DMRS18, EGG16, FM18, JZE^{+}18, LCW17, PS18]. hpq [Zbo19]. I [BMC13, CT10a, DG10a, QRMH18, Sea15, TL10a, YMHL18, ZZC13a].

I − A [BT11]. ILU [DM19a]. J [SB19]. K [ĆSCD11, HCL11, AJS19, AKS10, HpD11, HL10, JSEM13, JL15b, KP10b, MM16, Pes13, SD12b, SWW11, WYY11, XC17, YL18a, YT12, sHC11]. K(m, n) [Bis10]. K^{2} [SXB^{+}12]. κ [KPG13]. L [ABFGZ11, LS10b, LZH12, MP19c, Ram11, Sim10, YZXW10]. L^{1} [LCH19, PSS18, Zha15b]. L^{2} [CM16c, EM14, GH18, HZM11, JMHF13, MW14, Ye17a]. L^{∞} [He16]. l^{p} [GH10]. L^{p}(log L)^{α} [BX10]. l_{0} [DCKY15]. L_{2} [CCKP15]. L_{∞} [zSdZ10]. L_{p} [AD10b]. λ [SH12a]. λ_{i} [CsH10]. l ∈ {1, 2, 4} [YH11b]. LL^{∗} [Lee15]. LU [Abd18a, ASA16, CLMM18, PPC13]. M [PXT10, SSR11, Bai11b, LZ11e, LLL12, MKR12, Noo10, ÖŞ11, Wu16, Yas12, ZY15c]. M + 1 [PXT10]. {m, n} [MTN19]. C [HD14b]. F_{2}[u]/(u^{4} − 1) [YS12]. R^{3} [CS16, Lü14]. R^{N} [CF16b, AEO15, CCX13, LL16a, LX15a, LL14b, WZF16, WW15, ZWW13, ZCY11, ZZ16c]. R^{n} \ {0} [SYY13]. R^{4} [WX18b]. A [LWBW13]. D [YL14]. L [qJhY12, SP10]. N (β) [UBF11]. P_{n}^{∗} [Wu10]. N [LMMF17, LM19a, LTSW16, LWW19a, RC17b, YHC18, ZZ11b, ABR10, ACC18, Che16, Dra11a, Gao12, GZ14, HLWX11, HL11a, HMWZ16, LCW10, ŞGY11, Sat11, gShYL10, SS14b, TZG10, Wu10, YZGW10]. n = 2 [Dai14]. NM [Wu10]. ω [MM16, YL18a]. ⊗_{B}^{k} [Sat11]. P [HpD11, PS12b, Amb19, AX11, BLS18, CNSV17, CHZ19, CP15c, CP16c, Chu18, DZ21, DLS18b, DHMU16, EAA10, GALO18, HY18, HGW18, KS11, KLL19, KSKK11, LCW10, LD18, LJX12, NT17, PC11b, PA12, RZ17, SD18, SHH16, Sim10, SYY13, Tod18, Tru19, VAK^{+}19, WZH10, WW19c, XZR16, YY10b, YL10a, Yan11c]. p(·) [BRR16]. p(t) [CTA12]. p(x) [AEO15, ZZ15b, KSZ18, Zha18a, ZG18a]. {P, k + 1} [YqS16]. P_{3} [Zha18c]. P_{4} [Zha18c]. ϕ [GKM11, QHW11, YL18a, ZC10]. ϕ_{0} [WW11a]. P I^{λ}D^{μ} [EK13]. Q [MMRN12, CsH10, CS11c, Fer11, Gal11a, Kim10, Liu11a, Mah10a, Mah10b, Ost11, OÖ11, Yan11b, ZP10, dCM12]. q > 1 [Gal11a, OÖ11]. Q_{n} [KP10b]. QR [LYS12a, LYS12b]. R [BR18, Mar11a]. R^{2} [dSSV17]. R^{3} [Ben17, CTC17]. ρ − (η, θ) [PN10]. RT_{k} [Erv12]. S [DMPV10, SSO10, Set12]. S_{8} [HSMG12]. σ [qJhY12, KD10, KD11, MM10b]. σ → 0+ [WY19a]. SIMS [ZY13]. T [Cia12, GWL11]. τ [ZC10]. θ [DMZ10]. V [Ant10, WN18]. φ [CL16b]. W [HCL11]. W_{p}^{1} [JLL18]. X [ZCZ17]. x^{2} − kxy + y^{2} + lx = 0 [YH11b]. x_{n+1} = (α − βx_{n−k})/g(x_{n}, x_{n−1}, …, x_{n−k+1}) [Ham10]. x_{n+1} = \frac{x_{n−1}}{β+γx_{n−2}^{2}x_{n−4}+γx_{n−2}x_{n−4}^{2}}

[ECY11]. xC′_{ν}(x) + γC_{ν}(x) = 0 [GS12]. X ± A^{∗}X^{−q}A = Q(q ≥ 1) [YL10b]. y [QRMH18]. Y A = D [XC17]. |A|_{k} [Sav10].

-accretive [WD10]. -adapted [PPC15]. -adaptive [BS14a, CGGM19, JZ13, JZE^{+}18, PGF18, PPC13, RZ17, BS15b, BS15a, CDG15, CHH14, FWW14, GSZ14, Pas14]. -adaptivity [VAK^{+}19]. -adic [Sim10]. -AFEM [CNSV17]. -algebras [CKN11, qJhY12, SR12, ZJ10a]. -analogues [Rah11b]. -approximation [MTN19, AD10b, Ant10]. -approximations [Zbo19]. -ary [YZGW10]. -based [CLMM18]. -BEM [FMPR15]. -Bernoulli [Kim10]. -Bernstein [OÖ11]. -biharmonic [BRR16]. -bit [CSCM13, FVVS16]. -block [CBS18]. -boundary [Hak14]. -bounded [KS11]. -boundedness [WW11a]. -breaking [KTK18]. -cages [BFG11]. -Cauchy [DG10a]. -Cebysev [Yan11b]. -chromatic [KP10b]. -coercivity [Cia12]. -concave [ZC10]. -conditions [KT11a]. -conforming [BS17, DFG19]. -conjugate [BR18]. -connected [TKH10]. -conservative [MM10b]. -constant [SISH12]. -continuity [Çak11c]. -contraction [ASV11]. -contractive [HLS11]. -control [LCH19]. -convergence [KD10, KA11, PS12b]. -convex [SSO10, PT11, Set12]. -convexity [GKS10, ÖAK11]. -coupled [LTSW16]. -critical [WS10a, Ye17a]. -curve [DMPV10]. -cuts [TMMASG10]. -cycle [WN18]. -cycles [BR12d]. -D [KK19a, SLKK19, YZ19, CCHG17, FF15, GDM13, Tod13, Wei12b]. -difference [Liu11a]. -differences [Fer11]. -differentiability [ZFZQ10]. -dimensional [AGT19, Abd18b, AK16, Ade16, AA15, AAEG17, Asl11, BM18a, BB08, CTSX16, CZ15, CZ18, DYH11, DTYZ18, EGAA19, FYYT11, Gao12, GZ14, GZD18, GABC16, GMZ15, GTZ19, HLLM19, HTM18, Her19, HTY^{+}19, HTWS15, HYC18, KSG11, KKK15, KT18a, KK19b, KK19c, Kuo18, LHD18, LK15, LD11b, LDL11, LMMF17, LM18a, LZL19, LM19a, LCW19, LGZ19, LL10a, LTJ^{+}16, LL16d, Liu18a, zLYmL18, LZM18, LYC^{+}19, LWW19b, zLZ19, LWW19a, LXY19, MLL16, MYZ18, MCF18, ML19a, MR17, MKR12, Man18, MZM18, MGS^{+}14, MPfTX18, Men18, MK17, Osm18, PTZ19, QTW^{+}18, Ray17, Ray18, RMY19, SR15, Sat11, gShYL10, SS14b, SMBY10, SG10b, TDXQ18, TZ15, TTX^{+}16, UKI11, WWW14, WTYZ17, WWD18, Was13, WZMY18, XX10, YTD^{+}18, YGS^{+}16, YMLL18, YYLW19, YTS^{+}17, YTL^{+}18, YHC18, ZSAN19, ZA15, ZZ11b, ZM17a, ZDZY17, ZLT18, ZDM11, ZTSC16]. -dimensions [RRO17, YM17]. -distance [SCK11]. -distribution [MS11a]. -divergence [YY12]. -eigenvalue [LKCN19]. -estimates [CM16c]. -expansion [Gep16, GMZ15, JKB11, KBA11, MCKM12, SG10b]. -FEM [BPS19, FM18]. -filled [RSL^{+}18]. -finite [PS18, BS15c, CW15b, HW16]. -fold [Noo10]. -fractional [Amb19, BL17, YjH18]. -functions [CL16b, Sim10]. -fuzzy [DSA09, Jun10, SJN10a, Abb10, ABFGZ11, LS10b, LZH12, MZJ11, ÖI12, Ram11, SP10, SR12, SJN10b, SM11, YZXW10]. -Galerkin [LCW17, SZL^{+}17, WLL^{+}18, Hou15]. -Gini [Dra10]. -harmonic [LWBW13]. -Hessenberg [SB19]. -Horadam [YT12]. -hypergroupoids [SKST10]. -hypernear-rings [DZK10]. -hyperrings [Dav10, MZLF10]. -ideals [DZY11, DSA10]. -interior [ZD15]. -intra- [MYZ12]. -intuitionistic [ADA11, qJhY12]. -invex [Ant10]. -invexity [Ant14, PN10, ZLLF12]. -involutory [XC17]. -irregular [ACC18]. -Kirchhoff [XZR16]. -Laplace [Chu18, ZY15c]. -Laplacian [DZ21, BLS18, CTA12, CHZ19, CP15c, CP16c, DLS18b, GALO18, GKM11, HY18, HGW18, KSZ18, KLL19, LCW10, LD18, LJX12, NT17, PC11b, SD18, SYY13, Tod18, Tru19, WZH10, WZF16, WW19c, Wu16, YY10b, Yan11c, Zha18a, ZZ15b, ZG18a, AEO15]. -linear [QHT16]. -logic [Wu10]. -loop [LMMF17]. -matrix [SH12a]. -maximal [Ver12, Ver08]. -means [JW19b, HCL11]. -mesh [BDM^{+}19b, DMRS18]. -method [DMZ10]. -methods [KM15]. -metric [ASV11, APS12, ĆSCD11, HS11b, MAK12]. -monotone [Zha11a]. -nonexpansive [QHW11]. -nonlocal [QRMH18]. -norm [GWL11, KSKK11, He16]. -normed [ŞGY11]. -norms [BX10]. -order [BMA11]. -packing [CKMR11]. -phase [MK18]. -ply [ABR10]. -point [Bai11b, LCW10, LZ11e, LLL12, Yas12]. -potential [SPL19]. -problems [Zha15b]. -projection [HZM11, JMHF13, MW14]. -pullback [YL14]. -quasi-hemiregular [MYZ12]. -quasi-monotone [Bor11]. -quasi-slowly [Çak11b]. -queue [WY11b]. -rectangular [HMWZ16]. -refinement [DESV18]. -rings [DZY11]. -robust [CNSV17]. -Salié [ZP10]. -scheme [MP19c]. -semigroups [JMLF11, SDM10]. -semihypergroups [AMD10]. -skew [DH10b]. -smooth [KV17a, KV17b].

-soliton [LWW19a, ZZ11b]. -solitons [YHC18]. -spaces [LZ12b]. -sparsity [PSS18]. -species [Che16, TZG10]. -spotty [ÖŞ11]. -stability [JZL18, WW11a]. -Stancu [Gal11a, Mah10b]. -starlike [SWW11]. -statistical [DD10]. -strict [CsH10]. -strictly [sHC11]. -summability [Sav10]. -superlinear [Wu18a]. -symmetric [dCM12, DH10b, XC16]. -Szász [Mah10a]. -tensor [MMRN12]. -th-order [LM19a]. -theory [GH10]. -time [Dra11a]. -times [AJS19]. -top- [HpD11]. -transform [Ost11]. -trees [Mar11a]. -tridiagonal [JSEM13, JL15b]. -tuple [SD12b, Zha11b]. -type [BD11c, CZ15, LZL19]. -uniformly [AKS10, CsH10, CS11c]. -valent [AX11, EAA10, YL10a]. -valently [PA12]. -valued [Wu10]. -variable [KYR15]. -velocity [ZSW15]. -version [SHH16]. -weak [ĆSCD11]. -weakly [AKS11, APS12, SS11c].

0-1 [RFP11, RF12].

1 [BA11, Cia16, LXYT11, WY11b, Xu11b]. 1-soliton [Bis10]. 12 [FSM19]. 1227 [EM19]. 17th [MV10]. 17th-order [MV10]. 1D [He16].

2 [CFS17]. 2010 [VBW10b]. 2014 [Bho14a]. 2015 [Ano16a]. 2016 [YRDR18]. 2d [ABB17, He16, HD14b]. 2nd [HPV^{+}18].

3.0 [Váz16]. 35 [Li10c]. 3rd [VBW10a, VBW10b].

4- [WS10a]. 49 [ID10]. 4th [CX18, VBW10a, VBW10b].

50 [WWW11a, WWW11b]. 55 [Ver12]. 56 [Asl11, fDxZ11]. 57 [Pen11]. 58 [AR10b, JY11]. 59 [AD11a, ÇE10a, Def10a, GXZ11, IIHuI10, KPS10a, KK11b, SK12].

60 [Jia11, VBW10a, WYG12, Wan13a]. 61 [KYR11a, LLH11a, MP16, Pan17, WWW11a, XLD11a]. 62 [WLDL11a, Yaz11]. 63 [Ant14, LYS12a, LW12a, McN12]. 64 [Ran15, Wan13a]. 65 [Her14]. 66 [SCBCB^{+}17, XWH16]. 67 [JFS20]. 69 [LK15, SKTC19]. 6th [Sle13].

71 [Her19]. 75 [DZ21]. 76 [EM19]. 78 [JL20].

802.11 [TXZ^{+}10].

a-posteriori [GSZ14]. Abaqus [LZL^{+}18]. ABC [AT19]. abduction [BMS12]. Abel [Dar11, EKE18, XH11b]. Abel-type [EKE18]. Abelian [LLL11]. abortion [KX12]. Abound [LYC^{+}19]. above [RY10]. absence [JKS19]. absolute [Bor10, KM12, ZCH14, ZGD14]. absorbing [MDG19, RHD18, SDH15]. absorption [CP15c, CP16c, MM18a, ZT19]. Abstract [PRR10, AKA11, BK12a, KR11, LJ11, Sal10]. Abundant [YM17, EGAA19]. accelerate [RH15]. Accelerated [DTR19, ED12, WWXW19, BGGCGRSP16, BKMT14, BKNR19, CM19a, DLZ19, GEZ14, HWXC19, KWAS16, KLK15, QCG15, WMW13, YCW^{+}14]. Accelerating [LYC15, PPD10, LZZ11a, ZWG11]. acceleration [BBO10, Bre14, DCRL13, MO14, WCZ^{+}19]. accelerators [MSZG17]. access [BMS12, HSMY12, WZCC10]. according [ABRL18]. account [CD12, SST19]. accounting [Kia19]. accretive [WD10]. Accumulation [XLK11, IL13]. Accuracy [Rua19, TMCM19, CLTA11, DWZ13, DH11b, DFG19, FDG^{+}17, HZL17, LH16, MS15, SZL^{+}17, SSL14, WLZ^{+}18a]. Accurate [BJPT16, Gem16, HM10, JLCS10, AHF16, BPC17, BDM19a, BL14, BZ18, CPP15, DBS12, FZL^{+}18, GGGR17, HMY15, HIS19, HD14b, JW15, KO11a, KL16b, KPG18, KVR11, MS18, SMF17, SST12]. Acknowledgements [Ano10a]. ACOMEN [SVP^{+}19]. acoustic [ADGG13, BK15, DL11, EK16, HMY15, HNPS13, HKW15, JPK17, JPK18, LZ18a, LCP16, MPMTV15, Ray16, Sea14, Sea16, SSH15]. acoustics [BPC17, VAK^{+}19]. acquisition [GMAM12]. across [BOT14, EPP18, JSGP16]. across-time [BOT14]. ACSP [QCYL12]. actin [ABH12]. action [Kup14]. Active [FQLC18, JW19b, BSK11, CChL14, DB10, Fer12, HSK11, HT18, LSW16, RSL^{+}18, RKA^{+}18, SRDD17, WHD14, XJYL17, yYsZyYL13, jZsQdLmG19]. actively [BG10b]. activities [YY10c]. activity [BCF10, Laz10, RIW12]. actual [ABRL18, BCF^{+}14, WSS10]. actuated [LL12c]. actuating [CL12b]. actuator [JCWZ16, LLZ12]. actuators [NGG12, RMA10]. ACW [MN12]. Aczél [HX10, Yan10b]. Ad [SSM12, CHY12, DA12, LLWZ11]. ad-hoc [DA12]. Adams [Gar13, LLML15, PIAH10]. adaptation [BDM^{+}19b, ILV^{+}19, KH18, Li18e, MPS18]. Adapted [DMP18, LHL14a, PPC15]. adapting [AD19b, MSTB17]. Adaptive [AsNAd10, Amo15, BKDM13, CY14a, CKRW19, EFK15, GN19, HW16, LCCC10, PZJ^{+}16, SKTD13, YLC12, AA10a, AKL18b, BS14a, BS15b, BS15a, Bar17, BGGCGRSP16, BC17, CDG15, CGGM19, CNV14, CCR16, CGH14, Che11c, CCH^{+}12, CL12b, CHY19, CHH14, DESV18, DNP15, DVY14, DLQ16, FWW14, FOX11, GSZ14, GM14a, HM19, HD19, JSGP16, JZ13, JW19b, JZE^{+}18, KWAS16, KS12a, KKJ15, LMR19, LY13, LZC13, LCK17, LTLL12, MMRN12, MA16, MSTB17, MKA^{+}10, OZF19, eOS18, PGF18, Pas14, PPC13, PD17, PDHL12, RZ17, SPS^{+}13, SWOF19, SZGG11, SS19, SW16b, Sul16, TSB16, VFM19, VHPVNXW18, WZY13, YZM^{+}19, Zbo19, ZY17c, Zha17a, ZLY^{+}13, Zhe11, ZOZZ12, ZG19].

Adaptive-expectation [LCCC10]. Adaptively [TS14, JVMF19, JK18]. adaptivity [CFPP14, DPM15, GGT14, KdLK19, PS18, VAK^{+}19]. additive [BR16, CG14, DSB19, FMPR15, KC11, KRAS19, RMK19, SCvdV^{+}19, SM17, SNEP19, WvDRG19, YL14, ZTY^{+}19, ZLS13]. additive-quadratic [KC11]. additively [KMS19]. addressing [FBL11]. adhering [DVM12]. Adhesion [CFLM19, YLF19]. ADI [AD15, ADZ19, WLZ18b, BK14, CL17b, DZW16, HP17, HCLL18, QX19, WW11c, WV14, WW18b, WCLD18, XZ18, ZLJ^{+}18, ZC11c]. adiabatic [BG15, NNK13]. adic [Sim10]. adjacent [LKL^{+}15]. Adjoint [CHBTD14, KFTT13, KTH13, LWR16, MCL^{+}13, LGL^{+}14, Zha17b]. Adjoint-based [KTH13]. adjointness [Zha17b]. adjustable [LLD10]. adjustment [ASSV18, KALAS11, LTX^{+}13]. ADM [NGG12]. ADM-Padé [NGG12]. admission [BPKM10]. Adomian [Aba10a, AO10a, AH11a, AER12, BPG10, Dua11, DR12, DCRL13, ESBR10, HSWZ11, ZSD10]. Adopting [cFpClC13]. adoption [SSS^{+}11a]. adsorption [MMS^{+}18]. adult [ADY12, BP13]. Advanced [DGOZ13, GIM15, KSD^{+}19, LY12a, SPCS13, GHCZ18, Sag10, SSS^{+}11a, TSB16, Yüz12c, SVP^{+}19]. Advances [AER12, PHWM10, PL10a, Zho11, Zho12, ZFLM18, ZFLM19, LXK11]. Advection [RAW^{+}16, AJAR18, AM14a, AV14, Auc18, CBBE16, DMP18, EO15, ESBR10, HLL^{+}15, KW14a, KGM11, LPLR19, LZB12, LHL15, MHHC18, MPY16, MV17, SOJC10, Sou12, Tam16, TM18, VAS^{+}18, WCB13, YYK16, ZLPM13, ZLJ^{+}18, ZLZ10, ZDF^{+}14]. advection-diffusion [AM14a, KGM11, MV17]. advection-diffusion-reaction [AV14, KW14a, LHL15, MPY16, WCB13]. advection-dispersion [ESBR10, LZB12, ZLPM13]. advection-reaction [CBBE16, DMP18, SOJC10]. adversary [ZPWY12]. aerodynamic [GC19b]. AES [CSCM13]. AFC [JJ19]. AFEM [CNSV17]. affected [Pes13]. affine [Liu15a]. affordable [BCD^{+}16]. afforestation [TZ18]. AFM [WY11a]. against [JS12a, KLL10, QCYL12]. age [Aki17, Hu19, MCL15, SCSF19, ZK16]. age-structure [SCSF19]. age-structured [Aki17, ZK16]. agent [CS11a, MB17, SBKS12, WWH12]. agent-based [SBKS12, WWH12]. agents [SSS^{+}11a]. ages [DYQM14]. aggregation [AZB13, GEZ14]. aggressive [SD15b]. aggressive-invasive [SD15b]. aging [FdOdSS17, RR11, YY10c]. aging/deteriorating [YY10c]. agreement [HPC12, NCL13]. aid [Lep11b]. aided [DSB19, WMSH11]. air [DGOZ13, IC12, JS12b, LWC13, ODAZ15, ySGL^{+}10]. air-water [ySGL^{+}10]. aircraft [YGH11]. airfoil [PLR15]. Airy [CJCV10, CJRR11]. AKNS [RKSA18]. Al [GSI19, ZAK18b]. al. [SPLHCB14]. alcohol [Laz10, ZZ17]. ALE [KKVS19, MR19, ZHV19, ZLS19]. ALE/embedded [ZLS19]. aleatory [WLT13a]. Algebra [GGAVRC^{+}19, AS11a, LY17, MBH16]. Algebraic [ASN11, AvdW13, JZJ18, BJS15, CTP10, GL10, GEZ14, GV11b, GZW^{+}18, GH12b, HVR18, IB11, LBW11, LYC15, Liu15a, MO14, Pap15, SBEB10, Tom13, WW10b, YLJ12, ZMA10]. algebraic-differential [ZMA10]. algebraically [HS11a]. algebras [CKN11, qJhY12, JLP10a, JLP10b, JLK11, JK11b, JK11c, MZJ11, SR12, YZ10b, ZJ10a, JKK10]. Algorithm [BGRS11, GACMO13, LEP11a, SLMZ12, tWqLzGkP11, AAR11, AIA13, AHF16, Ade16, AM15, ALLH11, AWJH19, ACC18, BDS15, BMM12a, BA16, Bog10, BBM18, BM18c, CB11a, CD14, CGY11, CC11, Che15b, CX16, CR18b, dSCM12, DCG^{+}12, DVY14, DLWW12, DWI^{+}12, DJZ18, EMR10, EHO^{+}12, FJB19, Far11, FVVS16, FOX11, FLWJ11, qGpWhL11, GHR10, GALO18, GIMZ14, GHC^{+}15b, GN11, GGO16, Haj18a, Haj18b, HDHL11, HG18a, HCHH12, HY11, HL18b, Hua10b, HXS^{+}15, HCL11, ILS13, JKK11, JA11, JS12a, Jav11, JMB10, JKS12, JL17a, JW10, JW19b, KV10, KYO10, KCL14a, KMS15, KS10b, KCL12, Laz10, LLG^{+}11, LLD10, LCQF19, LTLL12, LYC12, LLCG16, LLH10, LJJ17, LJLY19, LSD10, LXF11, LTL^{+}12, LRH13, Liu15a, LLLW18, LRCG16, LDHH13, MSQ^{+}11, MJB18, MN12, Mor13, MW10, NZ16, NZ14, fNS11, OMS10, OVV^{+}16, OPDC12, Oru17, PH13, Pas14, Ped18, PG10]. algorithm [PSD^{+}13, RT11, RFP11, SS17, SSS^{+}11a, TH19, TL12, TZMZ12, TW18, TTC14, Ver08, Ver12, WYLZ10, WCD10, WZ16, WLGL10, WQNF12, WY19b, yXpYxZT11, qXjH11, XG10, XY14, XJYL17, XDL12, YY14, YXYH10, YJ19, YH19, YD12, YZ12, YDK^{+}12, Yüz12a, ZQ11a, Zha11a, ZDL11, ZLC^{+}14, ZY17a, ZY17b, Zha19a, ZGZ13, ZH15b, ZXW13, ZLGL11]. algorithm-based [FLWJ11]. Algorithmic [KPG13, MV12, GD11, TCM15]. Algorithms [JL16, AEF15, ABCR10, AAP12, AK11, AGK15, Bac14c, BSN13, BE18, Bur13, CGY10b, Che12a, CDY11, CLMM18, Chu11c, DM12a, DZS10, DGOZ13, Dua11, DR12, FHZ10, FGHZ14, FRSC16, cFpClC13, HSZ15, HM15, HM18b, JL15b, JJ15, JL19, JL20, Lit13, Liu15b, LW18c, Loh16, MVB^{+}12, OO10, ODAZ15, PDN19, PBK19, QHW11, RdSSS11, RF12, SH10, SLM16, ST12, WLW16, WLW17, WWXW19, YY12, YNLK10, YW14, ZN18, Zha15a]. alignment [KH18, SSK13]. all-speed [QCS^{+}19]. Allee [GOGYL^{+}11, PQB^{+}16, ZY13]. allelopathy [MKG13]. Allen [BMH19, CHY19, GGT14, LL14a, Lee16, LLJK10, LJK^{+}19, LCWZ18, LLYL19, MSW18]. Allmaras [PLR15]. allocating [RZL11]. allocation [Chu12a, ES11, RZL11, RR11, WL11a, WZWS11, ZLG^{+}10]. Alloy [WRW^{+}19]. Almost [BK11b, KPR13, LFJ11, XY10, Abb11, Bor11, CZN11, CZN12, CCD10, EO14, HP10, HGW11, KS11, MN10b, WYN12, Xue13, ZL11]. ALOHA [TZMZ12]. along [MZM19, RBB12, Ran15]. alternate [CB19a]. Alternating [ADZ19, EKZ17, SHM13, AD15, BKMT14, BCJ19, BXKZ11, BX14, CLA19, hGzS15, GL16, KST10, LHY18, LDHH13, WH16, WLZ18b, WS11b]. alternating-direction [CLA19]. Alternative [Tol12, ZYZ^{+}16, AEH19, Git14, HB19, HZ16, Uze10, ZZX^{+}14a, ZYZC18]. AM [HGJP19]. ambiguity [BBC^{+}11]. Ambrosetti [She18b, ZZ15b]. Ambrosio [MPS18]. American [BS10b, BC16, CW14, CWDL17, CWY19, CEJV16, ECJ16, HMV18, KCC^{+}13, LZL16, MZC17, MVKK14, YKA18, ZC11c]. American-style [LZL16]. AMMOS ProtLig [JPP12]. among [MLY18]. Ampère [Dai14]. amplitude [NJ16, YASK10]. anaerobic [DK14]. analogue [HmZ11, Par11b]. analogues [Kli10, Rah11b]. analyses [GGS16, GS18]. Analysing [EGGS^{+}12]. Analysis [AD19a, AM11, ALMLM14, APwS18, BM18d, BCF10, CCN14, CSU13, EG18, FP19, FSCG11, GS15a, GK16, GHCZ18, GJX18, HKS19a, HLL13, HSJ15, HT16a, HM17d, HMZ18, JLWX18, KTDT17, KO13, LZWC16, LXY19, MR17, Moo18, MNT15, NB11, PLKC16, RSS10, SL16c, TH19, WSL10, WL11a, WHS11, XSYL19, Yaz11, YAS^{+}11, ZLW10, ZHY14, ZLL17, lZxLhY12, ABM11, Abd18b, AO10a, AHF16, ADY12, Ade17, ASFM15, ABDKD12, AGU14, ATH18, ATH19, ASSV18, ABN18, AYY12, AJS14, BTEM19, BBDS11, BDS17, BCFQ19, BC15, BCD^{+}16, BCC14, BGRV15, BPS18, BO18a, BN16, BKM11, BDB12, BBM18, BMSS18, BGL^{+}15, Bra16, BK12b, CH11a, CLT^{+}13, CPP10, CCJP11, CCRS17, CJ12, Cha18, CW10a, Che12c, CCCW16, CYZZ18, Cho17, ĆCM10, CJ15, CGO19, CJPB10, DdSF13, DGT18, DS18b, DSZ18, DWS19, DFW^{+}18, DH18]. analysis [DYWL19, Dun18, EAAS18, EM14, FSM19, FGL10, cFpClC13, FXCC18, FOS19, GBG11, GMP18, GRBT16, GOS18, GGVRB19, GD11, GY11, GGGR13, GS11b, GTZ19, HLLM19, HH16, HHY^{+}11, HW19b, HWH^{+}15, HH10a, HLY16, HCLL18, Ikh11, JGSS10, JLD19, JW15, JN14, JO19, KEHB18, KWPK13, KLCD16, KVJB15, KNIF13, KP19a, KC12, KC19, Kia18, KLTS11, KT15, KHF^{+}19, KPG18, KM13, KPG13, KKSK18, KS10b, KS12b, KKJ15, KCK19, Kuo16, LZL^{+}18, LMMT15, LL12a, LLZ11, LDW11, LYS12a, LYS12b, LL13, LY15, Li16b, LS16, LM17, LPML19, LBJ10, Lit13, LZB12, LTL^{+}12, LGL^{+}14, LZ12c, Ma10a, MCQ11, Mac12b, MA10b, MC10b, MAN^{+}15, MH11, MG15, MP19a, MMFT^{+}19, MW16, MPGW19, MVB^{+}12, NXHN14, OBCG19, Ogi12, OTiSY16, ODAZ15, OF16, PXXZ16, Per18, Pet11, PLT^{+}19, QXLL11, QGGL13, RSL^{+}18, Ray17, RS18a, Ray18, RS15, RÖ10, RDE^{+}17]. analysis [RTT17, SG11a, SR17a, Sea14, SISH12, SWL15, SW16a, SZL^{+}17, SW17, SY18, Sin16, ST19, SJS^{+}11, SLYY13, SJC14, SSC19, SM10, Tan18, TL18, TZ13, Tim13, TT14, Tol12, Tom11, TAA14, UKA15, Váz16, WYLZ10, WY11a, WC11a, Wan12, WLT13b, WLHZ14, WZ16, WHTZ16, WY16, WLZ^{+}18a, WZ18b, WY18a, WKBR18, WLXZ18, WZH18, WYK10, WH14, WO10, WL17b, WGY^{+}18, XKH10, XH11a, XHH^{+}19, YG17, YWL^{+}11a, YXX11, YCHW18, YQWZ19, YWHC11, YXWL14, YASK10, YZAX10, YW11b, YS19b, ZAK18b, ZR18a, Zed10, ZW11a, ZTH11, ZWL11, ZHQG12, ZHZ14, ZZ16a, ZJB19, ZY19, ZC11a, ZZ14, ZD15, ZZW15, ZMA10].

Analytic [Bog11, CCNT16, JQSS12, LMDL11, LTJ^{+}16, MPfTX18, Odi10, OSZP13, SLCC12, Alo11, AS10a, AS11c, ANR11, BGRS11, BBD10, CSSW12, CDN14, DM12b, EZM12, HHS^{+}10, MB17, Noo10, NNAS11a, NNAS11b, NNAS11c, NM11b, NODA11, ODR10, PA12, SKPW14, SRM11b, Swa10, TO11]. Analytical [AJAR18, CB11d, Che11b, GS11a, GZ14, JLTB12, KKAM11, KDG11, MLZ^{+}16, PDM11, Tha19a, XCS18, YMDZ10, ZZLB18, ZMA10, AÖ10b, CS13, FKKS11, FL13a, Gup11, HM10, JRZK11, KLTS11, LL14a, Lee17, Lu11, MSG11, MPMTV15, NB17, PLW^{+}18, SSH15, TA11, UKI11, XHA13, XWL18]. Analytically [CXZ15b]. Analyticity [LLY18b]. Analyzing [WO18, ZJZ^{+}11, SMM19]. ancient [ZQ11a]. and/or [HLY17b]. Anderson [BKNR19]. aneurysm [ZAK18b]. aneurysms [OHK^{+}19, WSC16]. angle [WHQ^{+}18]. angles [KYW^{+}18, WL11d]. angular [LSM11, Ma18a]. animals [OMS10]. Anisotropic [DMRS18, Li18e, LWL14, MPS18, SRDD17, WRW13, ZWH^{+}19, BDM^{+}19b, BM18b, BS17, CHM18, CST14, DLZ17, DB15, FIS18b, GKS17, GHCZ18, HHY13, KMS15, MTN19, MSTB17, MM19, NB11, ORR16, TSB16, VGC^{+}15, WLZ^{+}18a, ZSQ^{+}18, ZY19, Li18e]. anisotropically [GSZ14]. anisotropy [BM19]. Annealing [dCMdSGTdC^{+}16, Che12a, VB10a, ZLY^{+}13]. annotation [ALLH11, ALLQ13]. Annual [HPV^{+}18]. annular [ATH18, BKE18, MG15]. annulus [HO19]. Anomalous [CSZK10, LZCL18, AM13a, ADZ19, CJ18b, FIM18, LCA^{+}17, MZL13, MD18, MA17, QLT^{+}18, WW16, Yil19, ZJZ18]. anopheles [ADL12]. ant [qGpWhL11, HB12a, HY11]. antennas [HB12a, LHZ^{+}11]. Anti [AN11a, YZY10, AA11, AsNAd10, CYP16, LM18c, XHM14, Yan12a, YT11, YqS16, dS15, dS16]. anti- [YqS16]. anti-centrosymmetric [LM18c]. anti-pentadiagonal [dS16]. Anti-periodic [AN11a, YZY10, AA11]. anti-reflexive [XHM14]. anti-synchronization [AsNAd10]. anti-trapping [CYP16]. anti-tridiagonal [dS15]. anti-waves [YT11]. anti-windup [Yan12a]. antiperiodic [SLL12a]. antiring [ZL12b]. antisymmetric [Kli10]. any [BC17, KLL10]. AOR [CM16b]. AP [DG10a]. Apostol [KYR15, LS11d, Öza11]. appearing [Was13]. Appell [OCNG12]. Appl [AR10b, AD11a, Asl11, Bho14a, ÇE10a, Def10a, fDxZ11, DZ21, GXZ11, Her14, IIHuI10, ID10, JL20, Jia11, JFS20, JY11, KPS10a, KK11b, KYR11a, Li10c, LYS12a, LLH11a, LW12a, MP16, McN12, Pen11, Ran15, SK12, Ver12, WYG12, Wan13a, WLDL11a, XLD11a, XWH16]. Appl. [Her19, Pan17, SCBCB^{+}17, SKTC19]. Apple [GMAM12]. Applicability [PKD19]. Application [ATUC15, ASB12, Asl11, AYY12, BC15, BSZ16, BB08, CL11, CCY18, CFdM^{+}18, Cia12, CCM14, DFGG13, EAEH18, ED11b, EE18, FLLF10, GWR^{+}18, Gha17, HMP^{+}15, HPR19, HHS^{+}10, IIHI10, IIHuI10, JYF^{+}11, Jum10, KBA11, KKT13, KP18, KB10b, KLK15, KTA12, KLTS11, LX10a, LX12a, Liu11a, MN11a, NKM16, Pal12, PdlF10, PH13, QY13, RAD13, RCM11, Sae11, SKPW14, SPST18, SG10b, TL18, WY11a, WLW^{+}11, yXpYxZT11, YP10, YKKS10, ZN18, ZDF^{+}14, Ahm10, AD16, BR12a, BP11a, BL14, Bis14, Bor11, BCF10, CZY13, DWI^{+}12, DBH^{+}14, Elb11, FWFL11, GGLP15, Ge10, HBK^{+}19, HY11, HCL11, JW18a, JCZZ13, JBBL17, KEHB18, KLCD16, KBDC12, KK13a, KYAA10, KKBR19, KFYW11, Li10a, LZYW13, LZWC16, LLY18b, LZ19b, LLH11a, LLH11b, Liu17, LT11, Ma19, MPS18, MLZ^{+}16, Mok11, NHH13, OBAAD10, Özd18, PA15, PATA11, RY10, RDE^{+}17]. application [Rua19, SPL19, Sca11, SZ12a, ST16, SLMZ12, SK10b, SK12, SAR18, TDXQ18, TM12, VMO10, Von19, WD10, WPH11, Wu18b, WY19b, YMSL11, ZZ15a, ZQ11a, ZHJD13, ZHW^{+}11]. Applications [AKMS18, Ant14, CH11b, EKE18, GGAVGG19, GIM15, HNK13, LK15, MV10, MN10c, NNAS11a, RT10, SIL19, Tri11, Yaz11, ZM17b, AP19a, AJJAD^{+}10, ADK10, Amb12, AS15a, AD12, BMRA10, BJLZ12, BMC13, BS12b, BR16, Cao19, CDG16, CCHG17, CBM10, CCFV12, CP10, CWQJ12, CGS12, DV10, DLWW12, DL11, Dun11, EGG16, FIMV18, FGHZ13, FES17, FF14, FMGR19, FRZ15, GTG11, GHMN16, Hak14, HJ13, HRMS12, HLvS18, HLSN16, HSS^{+}12, HSMG12, JMST11, KM15, KA10b, KÖC^{+}18, LZL^{+}18, LjHO10, LB12, LZD17, LKCN19, LWZG10, LNKU12, LY11d, LL15, LZ12b, LLW15, Mat19, MI16, MS11b, MN10b, NA11, OMT12, PZL^{+}18, PHWM10, RRC11, RA12, Sal10, Sal11, Sar11, SKM11, TXZ^{+}10, TZG10, TNF11, Tha19a, Tia11, THH12, TMCM19, Ver08, Ver12, VBCJ10, WHD14, WZH18, XHH^{+}19, Xue13, YB13, YS12, YS16a, YjH18]. applications [ZT16a, ZM18, ZC10, ZX11, ZJ12, Zha15a, ZC17, CTS19, EM19]. applied [Ala10, AA18, BDHR18, Boy16, dSCM12, DFP^{+}13, Gur13, JLP10a, JLK11, KKL16, KQ11, LZG19, MBS17, dCMdSGTdC^{+}16, MP19b, NH15, SH12a, SPP18, VAK^{+}19, YL18a, YZMA18, ZLY14]. Applying [GWZ11, EZRR10, HCHH12, JM15, MF11]. appraisal [PDN19]. Approach [YLK10, AD15, Abd18a, Ahn12, AEF15, And12, ÁBÁPM11, ATH18, AB16, Asl10, BQ15, BQ17, Bai19, BS10b, BLS18, BB15, BZZ^{+}10, BM13a, BE12, BK18, BWZ16, BPKM10, CCY10, CW10a, CHL18, CZ11b, dSCM12, CGJ^{+}14, DA16, DA18b, DA18c, DFS11, DMPV10, Dub13, Elb15, ES11, EEBM10, FPW^{+}11, FL11b, Gal11c, GD11, GS11b, HDS11, HH18b, HC18, Hon12, HSC17, ILP14, JSGP16, JNBK13, JRZK11, JK10, JM15, JMST11, KB10a, KG14, KLK15, KC12, KAJ11, KMRN12, KS12a, KSKK11, KT18b, KK19b, KK19c, KdLK19, LZL16, LR14, LTX^{+}13, LZC13, LW17, LL12d, Lin12, LMZ18, LKL^{+}15, Lu11, LZS12, uHS12, Mai16, MBH11, MO14, Mil18, Moh15, MHH11, NH15, NGG12, OKTR11, Özu15, PZAR19, PTP14, PM10, PMA17, PATA11, PDHL12, SD11a, SYG11, SDH13, ST16, SRS11]. approach [SH12b, SY18, TL10b, TYY^{+}12, TAPA^{+}17, TT14, TMO13, UCK16, VP11, Von19, WH11a, WAG^{+}14, WY16, WLZ^{+}18a, WKBR18, WD12, Wu11b, Xu11a, XC11b, Yan12a, YKC11, Yaz11, YAS^{+}11, Yüz11, ZB19, ZZ19, gZnZpZbD12, ZHJZ11, ZNWG11, ZL12a, Zha18i, ZC11b, ZG18a, ZHW^{+}11, ZPGW16]. approach-based [HDS11]. approaches [AJJAD^{+}10, BTEM19, Che11e, LG17, PLKCC12, RS13, ZLL18]. Approximate [ASY^{+}11, BAO^{+}12, BKZ17, DM12b, DDK11, GBG11, Gup11, HP19b, HLZD11, HLW19, Lam12, LJ10a, MB10b, SSA12, ARESH18, BF16, Dun18, EOM11, GHR10, GSR14, GN11, LZ11c, NKM15, Oru17, QLT^{+}18, SRM11a, SKPW14, TAS11, TC10, VA11, WZW10, WWXW19, WZC^{+}19a, XCS18, YY15, YKC11, Yao16, ZTC14, ZG14, ZMA10]. approximated [LCHZ19, YCLY15]. Approximating [Dra11a, FJP18, CL16b, EZM12, HDT11]. Approximation [Ahn10, BM11b, Büy10, CFN11, CS12, CAP10, CS11d, Gal11a, GW15a, LD13b, Mah10a, MCN10, PAT12, AH10a, AJT19, AS15a, AM10c, AD10b, Ana11a, Ana12, AK11, Ant10, APwS18, BM12a, BHM19, BGIN13, BGGCGRSP16, BDGS13, BL14, BKY10, Boy16, BV17, BK15, BMP15, CM10a, CB11c, CDG15, CY14a, CCNT16, CHXL18, CP15b, CCBSRFRM11, CC11, CS14a, CR19, Cia16, CFS17, CFLX18, CJRR11, CDD12, DE11, DO11, DE10, DCG^{+}12, fDxZ11, DW18a, DK12, DCKY15, DFS14, Dun11, EG18, EHL^{+}14, FRSC16, FGHZ17, GMS18, Ge10, GM14b, GJX18, GHL18, HS11a, HVR18, HKS14, HC14, KPS10a, KPS10b, KD11, KP10a, KKS10, KLRW12, LT13, LA11, LZC11, LR15, LL16b, Lin14, LC13, LWL14, LZLL18, Lu12, LvdVX18, LS10d, Mai10, MTN19, Mur08, NMR15, NSYY13, PW18, RMS10, RSB14]. approximation [SLK12, SZW11, SI17, Tha19a, Wan16a, WD16, WLL^{+}18, XFL16, XC11a, Yil19, Yil10, Yüz12b, Yüz12c, ZHB11, ZCLW19, ZWH^{+}19, ZG16]. approximation-preserving [AH10a]. Approximations [BBC^{+}11, LF11a, Akm15, AMD10, ACE17, AS19, Bay19, CS14b, DFG19, DWZ16, FDG^{+}17, FN14, GY15, Gue13, KN12, LMR19, LL12b, LRH13, LPP15, MCP13, MV12, Mor10b, RA11b, Sou11, TMDTTC16, TMCM19, VGK^{+}16, Vel15, WNTW19, YSX^{+}19, Zbo19, ZW11c, ZZX16]. April [Ano18-62, Ano18-65, Ano19-52, Ano19-54]. aquifer [LS17b]. aquifers [LWL14]. arbitrage [CCJP11, Sin16]. arbitrage-free [Sin16]. Arbitrary [Par15, Wei14, AZ15, AHOP18, Gue13, HL18a, HGSL18, LPML19, Liu16a, LRCG16, MJWD19, Ned12, SS19, TS14, XLK11, ZZ11b]. arc [Gao11, KSMT11, Par17a, Par17b, TMMASG10, ZD12]. arc-roof [ZD12]. Archimedian [SP10]. architecture [AK12, LWR16, LCT12, SSESG12, WY19b, YLC16, YGH11]. architectures [ZMM18]. arcs [Lu12]. area [CSCM13, HR14, LCT12]. areas [NM11a]. areca [Hua12]. argument [Bic11, XY10, ZL11]. arguments [Can11d, Don10a, KMT10]. arising [AB16, BN14a, BNR10, CM10a, CL16a, CM18c, CG15, CL16b, CDN19, Du12, ES18, FJP18, GV18, KK15, KFYW11, KS12b, LM18a, LB11, Oh15, RY11, RS12b, SLKK19, Tan18, YYK16, YP10, YYYH19]. arithmetic [ÇA14, Ikh11, WLC11]. arm [KBGC12, SJS^{+}10]. Armstrong [CTS19].

Armstrong-Frederick [CTS19]. Arnoldi [AEH19, JW10, WTM17]. array [ĆCM10, HMF16]. arrays [CDW11, Kia18, Wan10b]. arrival [CTD10]. arterial [SRDD17]. arteries [MAN^{+}15]. artery [LYY12, SR10b, ZAK18b]. Artificial [KP18, OAY11, GH14, HBE14, HBE15, HCHH12, KAG11, MSH10, Naz13, OAKR16, PH13, TLR17, TL12, WLC11, WQNF12, yXpYxZT11]. ARX [LSD10]. ARX-like [LSD10]. ary [ES11, YZGW10]. ASCAM [Ano16a]. ASCAM-2015 [Ano16a]. Asian [DNS16, JYK16, ZM16b]. aspect [BSK11, CY14b, KYAA10]. aspect-ratio [CY14b]. aspects [BP11a, GHT^{+}15, TC18, TCM15, ZK16]. assembling [RPTD10]. assembly [BCG17]. Assessment [FBL11, CANA19, HH10a, KC19]. asset [BM12a, ES11, GLW18, MVKK14]. assignment [LMRS10, MAPS10, RES10]. assimilation [CS14b, DNZ^{+}13, RMB^{+}14]. assistance [KPL11]. assisted [GGVRB19, ZHQG12, FMP19]. associated [APRM11, ASN11, ASMEE11a, CS10b, Cvi10, FLDZ12, GW12a, Las10, MP10b, NM11c, NM11b, ORR16, Pov19, Sok10, ZZT11]. ASST [VGC^{+}15]. assumed [SA16]. assumption [Tsi11]. Asymmetric [BKE18, Eba14, KV10, MA10b, MC11, MCR11, WL11b]. asymmetry [WSC16]. Asymptotic [AIIZ10, AJ11, BMA11, BCC14, BG19, CKR10, EPP18, He11, IIHI10, JFC14, KLP17, KJ11, KM11, LPK15, Li14, Mah11, MMMG12, MDL18, Pen11, SISH12, SCBCB^{+}13, SCBCB^{+}17, WW19c, AKL18a, BM10b, BPR18, BM13a, BD19, CC19, Chu10, DD19b, EG10, GÇK10, GMB12, HKI12, HM10, IIHuI10, JKN10, JLL19, KPG13, MWL18, MP11b, MP16, MAB19, MMT18, MY10, Mor10a, Mor11a, NSYY13, PR11, PP12, Rey12, SAIZ15, TTX^{+}16, WY19a, WS10b, Xue13, YY10a, ZMWH18, ZD15].

Asymptotical [WYL19, LCC13]. Asymptotically [CCKP15, DSR10, BR12d, CT17a, CS10c, GL17b, GDZ11, KKS10, KA10b, KT11b, LT15b, LLT16a, LLT16b, QHW11, QHT16, SC16, XLT17, ZS11a, ZRC11, ZTZ16a, ZZM17, sHC11]. asymptotics [FLDZ12, MZL13]. asynchronous [CDFP12]. Atanassov [ZW11c]. atmosphere [GZW^{+}18, LZ19c]. atmospheric [GGM^{+}13]. attached [DDK11, Gen11b]. Attack [CJP12, CJP15, QCYL12]. attacks [JS12a, KLL10, LKLP12, MKA^{+}10, ZHJZ11]. attainable [GN11]. attention [WZWS11]. attitude [KW12]. attractant [QMW18]. attraction [SG10a, Wan16b, ZMH16]. attraction-repulsion [Wan16b, ZMH16]. Attractivity [CZ11a, LC10b, YW10]. attractor [DHGF17, FNZ13, PP14, WZ18a, YL14]. attractors [MX15, ML19b, MN17, SM17, WL17a, XY11, YLG17, YL18b, ZZ16c]. attribute [HKHK13, LCCC10, LS10c, Med12, PCK13, WL13a]. attribute-based [HKHK13]. attributes [NGL10]. attrition [GV11a]. augmentation [Che11d, CDY11, HH11, SBJ15]. Augmented [Liu16e, Bay19, CGO19, DH11a, DBH^{+}14, GRBT16, GOS18, PW10, SWS19, WLW17, WNTW19]. August [Ano18-71, Ano18-72, Ano19-64]. Australia [VBW10a, VBW10b]. authenticated [GLM^{+}11, HPC12, HPY10, NCL13]. Authentication [CJP12, KRM^{+}10, CJP15, SPLHCB14, SCKH10, YWK^{+}10]. auto [CZ15, MSTB17, WF17]. auto-adapting [MSTB17]. auto-Bäcklund [CZ15]. autocatalysis [GLL14, Zho13]. autocorrelation [NWZ11]. autoimmune [DDLM13]. automata [CM12b, FNZ13, SSM12, SAU11, WPH11, ZLC^{+}11a]. automata-driven [CM12b]. automated [DRT^{+}15, KBDC12, LL12a, RT11, SSS^{+}11a]. Automatic [CCL^{+}12, WNC12, jZsQdLmG19, ZKBE16, ALLH11, KJK18, LWJ10, MBKK10, PAE^{+}12, TNT12]. Automatically [LYX11, NCC13]. automaton [HBS^{+}10]. automorphic [Abb11, CZN11, CZN12, CCD10, MN10b]. automotive [CI18, CCKY12]. autonomous [CZN11, CZL17, LLP19, PP14, PP12, WL17a, WZ18a, WW19c, YL14, YLG17, Zha18h, ZZ16c]. autoregressive [YD12]. Auxiliary [AK18, WCD10, CH19, CWWY15, GH15, HCHH12]. availability [DLT12, MMR11, Tod12]. avascular [Amo18]. AVC [CUK12]. average [CCL^{+}12, Ili12, Mok11, Tod12, WC11b, YD12, ZDL11, ZSQ^{+}18]. averaged [DJD18, Izs15]. avoid [KALAS11]. avoidance [YMM12]. Avoiding [MBJ16]. Awad [Pan17]. aware [AVV18, AGK15, BMS19, SSM12, VFM19]. axially [HKK^{+}16, HL11c, SA16]. axiomatic [SH12b]. Axioms [CFPP14]. Axis [TLR17, CY14b, LZY12]. Axis-symmetrical [TLR17]. Axisymmetric [LL12b, BGRV15, CZF10, EF14, HLNZ19, LH16, Oh15, OY19, ZYZ^{+}16, ZaY17, ZYZ^{+}17, ZYZC18, Zha18g].

B [FZBF10, FAIV10, JKZ11, LZZ11b, CQRW11, GW15a, GMI11, GMI12, HH15, KB19, KPG13, LXL14, MRS15, MD15, Moh15, PZJ^{+}16, PZAR19, PLT^{+}19, TWLYÖ10, TTX^{+}16, YZ19, ZJB19]. B-spline [GW15a, GMI11, GMI12, LXL14, MD15, Moh15, PLT^{+}19]. B-splines [CQRW11, HH15, KB19, MRS15, PZJ^{+}16, PZAR19, TWLYÖ10]. B-type [TTX^{+}16]. Babuska [DFG^{+}18]. Back [KRM^{+}10, MS13, MTV13, CEF^{+}13, OKTR13a, Ste16]. Back-propagating [MTV13]. backbone [LLWZ11]. background [FGY^{+}17]. Bäcklund [CZ15, zLYmL18, LSX13, WF17, WF18]. backorders [CB10]. Backward [AGU14, AJY13, ACTB19, DE11, DH16, HXX19, LX12a, LLML15, LHTL19, Oan13, PH19, SCSF19, Tha19b, THD19, TNV19, WL12b, WZ18c, ZZWG16]. Backwards [YLG17]. bacterial [IL13]. baffled [WFC16]. Balakrishnan [HH18a, HW19a, KLP17, LPK15]. balance [AP19a, EZRR10, FZBF10, KAA19, Yaz11, YAS^{+}11, YASK10]. Balanced [HLC11, Chu12a, FM18, FR16, LX18, MDBCF16, XL11, Yan10a]. balancing [YZ15]. Balkan [ZGD13]. ball [BJLZ12, XLD11a, XLD11b]. ball-joint-like [BJLZ12]. ballast [CCKY12]. BAM [BV11, LHW11]. Banach [KK11b, KYR11a, BKT11, BHM12, CsH10, CB11b, CS11d, CS11b, CS11c, CSN11, FL14, HLWX11, HRMS12, KK10a, KKS10, KYR11b, KT11b, Lad16, LjHO10, LLX11, Nil11, PCM12, RGdSRLAJ10, SK11a, SCC12b, WD10, XLK11, ZLL11]. band [DIS19, Gos10, Sön11]. band-limited [Gos10]. banded [BDS15, LLCG16, MY13]. Bandwidth [WZ17a, QZY11, WL11a]. bandwidths [LLCG16]. Bankruptcy [Che11e]. Bapat [LX10a]. barotropic [eMA18]. barrier [CXZ15b, CKM12, DH17, FSRB15, FSB17]. Barycentric [Wac11, Gue13]. basal [GB18]. baseball [ALHZMC^{+}19]. based [AM10a, AFGL10, AK12, AMA14, ALLH11, ALLQ13, AdSSS19, AD11c, AOW18, ATH18, ABK10, AuIK11, BJLZ12, BPC17, BEAA11, BPM12, BC15, BBR10b, BJS15, BM18b, BMM12b, BLS17, BH14, BMS12, BR16, BM18c, BMP15, CB11b, CBBE16, CD14, CDP16, CCHG17, CR18a, CB19a, Che12a, CQLX11, CDY11, CC11, CLW11, CCH^{+}12, CL12c, CXZ15b, CM19a, CHY19, CHT11, CTZ17, CLMM18, CPT15, CJP12, CJP15, dSCM12, Cos18, DHQ11, DH11b, DA16, DM16, DA18c, Dej11, DB15, DYX11, DVY14, DK14, DGOZ13, DBEE11, DMRS18, DDM^{+}18, DWI^{+}12, DCKY15, DGLU18, DM19b, DH16, DNZ^{+}13, DLQ16, DNR13, Eba11, ES10, EE10, ESL11, FSH10, FJC16, FXC18, FNZ13, FGY^{+}17, FZ19, FWW14, FG18, FLLF10, FPW^{+}11, FL11b, FN14, FLP13, FLWJ11, GGAV18, GACMO13, GVJ13, GCE18, GGS16, GUK13, Gha17, GF16, GOT19]. based [GJX18, HDS11, HNK13, HZ10, HHY^{+}11, HC18, HAESLB14, HLB14, HSMT19, HD14b, HSC17, HCHH12, HCL12b, HYS^{+}14, HWW13, HXS^{+}15, HM18b, HD19, HPY10, HKHK13, IK12, JPCY13, JCWZ16, JW19b, JMNZ19, JK11c, KSPP11, KEHB18, KWPK13, KJK18, KK13b, KRP12, KK14a, KTDT17, KPP13, KH18, KX12, KSO16, KTH13, Ku15, KV18, KPL11, KORR10, LZ11a, LHL12a, LWWY12, LLH14, LG17, LLG^{+}11, LXZ13, LJJ11, LD11a, LHHZ12, LZH12, xLlFwWL12, LY12b, LY13, LTX^{+}13, LLML15, LSZ16, LZ15a, LSC17, LTLL12, LLH10, LW11c, LL12d, LYL12, LNKU12, LCCC10, LW10, LLX^{+}10, LS10c, LLG10, LXF11, LGH^{+}11, LHL12b, LGL^{+}14, LHL14a, Liu15b, Liu16e, LYSZ19, MSH10, MJWD19, MWWL11, MM18c, MHL11, Mok11, MR15, NKM15, Naz13, NM11a, NCL13, fNS11, OHMAK18, Oru17, PGQ16, PC11a, PXXZ16, PGDL18, PMM17, PZJ^{+}16, PZAR19, PGW19, PBK19, PKTH13, PATA11, Pov12b, PB13, QKR19]. based [QXLL11, RSL^{+}18, RKA^{+}18, RH15, RKW12, RWTW19, RFP11, RF12, RC18, SPLHCB14, Sag10, SH11, SWOF19, SBKS12, SBB^{+}18, SR16, gShYL10, Sha12a, SLXC11, SC13, SS17, SCC^{+}12a, SGK18, SSM12, SCKH10, SGZW18, uIAH10, SJS^{+}10, Sul16, SYL10, SLZ11, SCGW18, ST12, THZ^{+}11, TL10b, TH19, TNT12, TZXP11, TMDTTC16, TMLF19, TT12, TRL19, UKAL10, VC12, Ver08, Ver12, VAK^{+}19, VKJ13, VZM14, VHPVNXW18, WSL10, WCD10, WZW10, WYD10, WSS10, WZCC10, WXYW11, tWqLzGkP11, WMZW11, WWZ12, WRW13, WLT13a, WL13a, Wan16a, Wan19a, WZWS11, WFL11a, WZKY12, WRW^{+}19, Wei14, Wei17, WMW13, WD12, WBA^{+}18, WBN18, WWH12, WPH11, WTC^{+}12, WY19b, XY16, XCM12, XJYL17, XJ18, XDL12, YT18a, YXYH10, YY12, YC12, Yan12a, yYqWqZC13, YH15, YD12, YLC12, YGH11, Yoo17, YZMA18, hYlLL11, YNS^{+}14, Yüz12b, ZB19, Zbo19, ZJ10a, ZDL11, gZnZpZbD12, ZHY14]. based [ZL14b, Zha15a, ZYZ^{+}16, ZCY16, ZS16, ZY17b, ZYZC18, ZFY^{+}19, Zha19a, Zha19b, ZHL12, ZcHS18, ZJZ^{+}11, ZWJ^{+}11, ZLY12, ZGZ13, ZHW^{+}11, ZC11c, ZXW13, ZD12, jZsQdLmG19]. Bases [DM18, Ben17, Erv12, ID10, JW05, MP10b, Wac10, XSLS11, Yil10, YSS11b]. Bashforth [LLML15, PIAH10]. basic [Das15, Kir10b, RF12]. basin [SG10a]. Basis [MvS18, PGQ16, BM11a, BS15a, BASW18, Ber16, BL14, Boy10, BM13b, Che15b, CJN19, DHQ11, DDMQ19, DO11, DAM14, DM16, Dej11, GK18, HT12a, HT13, HT16b, IQR16, KKT13, KRP12, KY11, LW19a, MFSL19, NMR15, NB17, PT15, RKP12, RFK16, RKF18, RA12, Saj14, SL16b, TMDTTC16, Wei14, ZC16, Zha18d, ZcHS18, ZZL18a]. basis-finite [MFSL19]. basket [SL16b]. batch [CTD10, KPP13, KX12]. Bates [BC16]. BATMAN [KHIB12]. batteries [CD19]. battery [JLD19, YLH12]. Baxter [DT16, DD16, RWW18]. Baxter-like [DT16, DD16, RWW18]. Bayesian [ASB12, DGLS19, DNS18, GACMO13, TZMZ12, VJM15, YC10a, YMSL11]. Bazilevic [ANR11]. BBM [RS18a]. BCC [AS11a]. BCC-algebra [AS11a]. BCI [JLK11, MZJ11]. BCI-algebras [JLP10a, JLP10b, JLK11, MZJ11]. BCK [JLP10a, JLP10b, JK11c]. BCK-algebras [JK11c]. BCK/BCI [JLP10a, JLP10b]. BCK/BCI-algebras [JLP10a, JLP10b]. BCR [LM18c]. BDDC [KCL14a, TW18]. BDF [QX19, Yan15]. BDMk [Ben17]. be [RS12a, Swa10]. beaches [FER15]. Beam [WRW^{+}19, AML^{+}14, BMAR18, BDO11, KLTS11, LZ12a, MR15, RSP18, SA16, SM19, TAS11, YW11b]. beam-columns [YW11b]. beams [HL11c, LMR14, RYK13, TAA14, WY18a]. bearing [CJ12, MH11, Wan12]. Beavers [HQ19]. bed [IL13, PLMS14, RMK19]. Beddington [LY11c, WWG10, YX16]. beds [ZLC^{+}11a]. bee [HCHH12, PH13, TL12, WQNF12]. behaved [TTC14]. behavior [AKL18a, CFRS10, CC19, CO19, Don10a, GZZ^{+}16, GÇK10, GMB12, GF16, HBS^{+}10, JFC14, JLL19, Kar10b, KJ11, KM11, Li12a, Li14, Li19a, MWL18, MY10, PR11, Pen11, PSD^{+}13, PP12, RMA10, SK14a, ST14, TM18, WW19c, WY19a, WS10b, Ye17a, YZAX10, YCS19, ZYG10, ZW16a]. behaviors [AÇT11, Chu10, GZD18, TTX^{+}16, ZTZJ14, ZC11b]. behaviour [AJ11, HAESLB14, Sha18, YZ15]. behaviours [HMP^{+}15]. Beketov [Dos10]. Bell [Ikh11]. Bellman [YYYH19, CC11, HX10, Yan10b]. Belousov [LV11]. Beltrami [PL17]. BEM [BPC17, BS15b, BN16, Bol16, FMPR15, FH17, FP18, Kat11, MT19a, OP14, Wei14, Wei17, Zha17a]. BEM-based [Wei14, Wei17]. BEMD [TZXP11]. Bénard [RGVR17, FST19, Ma18b]. Benchmark [HAESLB14, HLB14, SPT17]. Benchmarking [OVV^{+}16, RHC15, PTH^{+}16]. bending [CCN14, GF16, LMR14, MM13, MG15, SSL14, VĆC10]. benign [CB19a]. Benjamin [CdR18, DAM14, HGHA19, HS12, KB19, NNWAS11, Oru17, YL18b]. Benney [Gup11, WY16]. benzenoid [DYX11]. Bernoulli [MP16, Kim10, MP11b, Öza11, OSS10, SPST18, TT14, YK18, ZB19]. Bernstein [CMGR11, JLCS10, OÖ11, Tac11, WFL11a]. Bernstein-type [CMGR11]. Berselli [ZWY19]. Besov [MP11c, QY17, QaY18, RA19, Via15, Zha18e, Zha18f]. Bessel [Sat11, She18b, YSS11a, YSS11b, Yüz12b, Yüz12c, Zhu10]. Best [DE10, XCS18]. beta [OCNG12, YYLW19]. Bethe [RT10]. better [Liu14]. between [Ala10, AJ12, BTEM19, BCF10, CM18a, CP16a, mCfX10, CFB11, DT11, Din10, EAMA19, FAIV10, GK16, HVA10, KRCJ11, LJK^{+}19, LGH^{+}11, ML19a, MNJ^{+}13, RGdSRLAJ10, SPH10, Tia11, xYsHjL11, ZJB19, ZLZG11]. beyond [LZS12]. Bézier [Ahn10, AM17, UMY11, Dej11]. Bézier-based [Dej11]. Bezout [WW10b]. BGK [CZ11b, LF11b, LLG10, LGH^{+}11, PGW19]. BGK-based [LLG10, LGH^{+}11]. bi [CT10a, CD16, DNS15, LBH^{+}12, RFP11, RF12, XZR16, XS10]. bi-fuzzy [XS10]. bi-iteration [DNS15]. bi-Newton [DNS15]. bi-nonlocal [XZR16]. bi-objective [LBH^{+}12, RFP11, RF12]. bi-projection [CD16]. bi-quasi-variational [CT10a]. biaxial [LL12c]. BiCG [XW19b]. BiCGStab [SL16c]. biconjugate [Haj18a]. BiCOR [ZHJD13]. BiCORSTAB [GHC^{+}15b, SJHC14]. BiCR [GHM^{+}14, XW19b]. BiCR-type [GHM^{+}14]. bicyclic [ZYS10]. bidirectional [BHZJ19, GH12a]. bidomain [LRBA15]. bifunction [NN12]. bifurcate [CD10]. Bifurcation [ABN18, BF11, CWW19, CJ12, DGA18, KL19, MW17, Wan12, ARK13, AGU14, CW15a, CM13b, CO19, CHH14, DWY15, GLL14, LYZ11, LW15, LWN15, Li17a, Li18a, LS19b, LPY16, MG16, RYK13, SCSF19, SZ14, TZ13, WY11a, ZLC11b, ZZW15, ZZC13b]. Bifurcations [AZ10, BBDS11, CSW11a, CS13]. biharmonic [BPS19, BRR16, DN18a, DK18, Gha17, GNP14, HHG14, Kon16, Moo18, MWY17, ST19, WW14a, ZC17]. biholomorphic [XL10, XL11]. bijective [GXZ10]. bilateral [BDS17]. bilayer [FG18, KHWK10]. bilevel [LLW10, WZW10]. Bilinear [WF18, CLH13, DTYZ18, DFS14, FSH10, FW18, FGHZ17, GZZ^{+}16, GS19, LC12a, MF11, NSYY13, ÜM16, WLZ^{+}18a]. billiards [CCF13]. bimolecular [Zho13]. binary [DLF^{+}11, FOX11, GT15, GT16, LLSW10, MMHGM17, SR10a, SKST10, TC18, WK13, XDL12]. Bingham [CD16, FKF13]. bio [Sen12]. bio-mathematical [Sen12]. biobjective [LLW10]. biochemical [NCV^{+}18]. bioconvection [MMFT^{+}19, UABK16]. biodegradation [BDB12, DK14]. biofilm [CAC14]. biofilms [DEFP11]. biogas [DK14]. biogeography [LY12b]. biological [LB11, Mag10, SD11b]. biology [CFB11, HY13, YP10]. biomass [WHLC11]. BIOMATH [AM12b, AM13d, AD14]. biomechanical [DRT^{+}15]. biomechanics [GDF12]. biomedical [CGS12]. bioreactor [AM12a, BDB12]. bioreactors [IL13]. Biorthogonal [BS15a, VT11]. biosciences [CLM14]. Biot [BKR^{+}19, CR18a, Fu19, HMY18, RDE^{+}17, RTT17, RTL19]. Bipanconnectivity [Yan10a]. bipancyclic [MW11]. biparametric [GK11a]. bipartite [YA11]. bipartition [ABK10]. biped [SCC^{+}12a]. birth [AF13, AGD10]. birth-death [AGD10]. births [GCG12]. bistable [CS16, She16]. bisymmetric [HA18a]. bit [CKK^{+}10, CSCM13, ĆCM10, FVVS16]. bit-plane [ĆCM10]. bit-serial [CKK^{+}10]. bivariate [Dej11, LW11d, OCNG12, ST16]. BKP [CZ17, Elb11, TZ15, bZM17]. BL [YZ10b]. BL-algebras [YZ10b]. Black [Bho14a, Bho14b, CHXL18, CXZ15b, CWY19, Dav17, DH17, HG18b, Jum10, LC13, Moh15, WY19a, ZLTY16, ZH18b]. blackjack [Coo10]. Blaschke [CQRW11]. blast [XWL18]. blending [LXL14]. blind [DM19b, FSH10, PXT10, WNC12, YXX11]. BLMP [KT18b]. blob [YKKS10]. Bloch [AET19, BDGBM11, BDGBM12, CLA19, LFAL19, Pet11, QLT^{+}18]. Block [CBS18, CL16b, DFS14, MBH11, AEH19, AJY13, AT18a, CDP16, DDM^{+}18, DBS12, HM17d, HMZ18, JK10, JW10, LS11b, LR17b, LS17b, NKA18, SBJ15, WmN13, YXYH10, ZZ15a, ZYW17]. block-based [CDP16]. block-centered [LR17b, LS17b]. block-grid [DBS12]. block-structured [JK10]. block-triangular [ZYW17]. blocked [AM11]. blocking [LA11]. blocks [HH11]. blogs [FPW^{+}11]. Blood [SR10b, MAN^{+}15, SRV10, WH11a, ZAK18b]. blossoming [ST16]. Blow [DG10b, DS18b, DZ21, HGW18, KB15, Li18c, Liu13, LLY13, LWSL19, LW19b, MF18a, Sam19, SD19, Tan17, TF17, Bag17, CC17, CCP19, Din13, DL18, FIW17, HL18a, HGSL18, LX15a, MF18b, MTAS17, NT17, PQBK17, SNSK19, SRG16, SLW18, Tia17, Wu16, XZ19, Zha18e, ZSL19, ZZ10b, ZY15c, Zho16]. Blow-up [DG10b, DS18b, DZ21, HGW18, KB15, Li18c, Liu13, LLY13, LWSL19, LW19b, MF18a, Sam19, SD19, Tan17, TF17, Bag17, CC17, CCP19, Din13, DL18, FIW17, HL18a, HGSL18, MF18b, MTAS17, NT17, PQBK17, SNSK19, SRG16, SLW18, Tia17, Wu16, XZ19, Zha18e, ZSL19, ZZ10b, ZY15c, Zho16]. blowing [Nab19, UABK16]. blowing-up [Nab19]. Blowup [Yu17, BS16a, ZY17d]. BLP [KKK15]. blue [HMSC10]. bluff [Özd18]. bluff-body [Özd18]. blurred [LHLH15, YZM^{+}19]. Board [Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o, Ano13p]. Board [Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano14x, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h, Ano17i]. Board [Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano18a, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18-32, Ano18-33, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano18-40, Ano18-41, Ano18-42, Ano18-43, Ano18-44, Ano18-45, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano19a, Ano19-27, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano19-36, Ano19-37, Ano19-38, Ano19-39, Ano19-40, Ano19-41, Ano19-42, Ano19-43, Ano19-44, Ano19-45, Ano19-46, Ano19-47, Ano19-48, Ano19b, Ano19c]. Board [Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z]. bodies [BR13a, KP19b, MGN^{+}16, NB11]. body [DTR19, HMY15, LZZC12, LCT12, Özd18, SEM13, WZC^{+}19b, YGR11]. Bogoyavlenskii [AKMUH17, CM18b, zLYmL18, MCKM12, WF17, WF18, XX10, YS16a]. Bogoyavlensky [Abd18b, LCW19, Ray17]. Bogoyavlensky-Konopelchenko [Ray17]. Bohemia [RB19]. boiling [BRFH16]. Boiti [ES17, LM18a, Liu18a, LZM18, MR17, PTZ19]. Boiti-Leon-Manna-Pempinelli [MR17]. Boltzmann [GGGR17, SKTC19, AvB16, ASFM15, AM10b, AML^{+}14, ABR^{+}14, ADGG13, ADGL14, AYY12, AJS14, BRFH16, BV10, BR13a, BS18b, BFS15, CYP16, CZY11, CWH13, CCSZ14, CCY18, CS18, CZ11b, CLL11, DL16, DSK^{+}14, Del13, DZW^{+}15, DLT10, DL11, Dub13, DJD18, EN11, EAAS18, EAMA19, FIVM17, FIMV18, FVVS16, FES^{+}19, GWR^{+}18, GLW13, GZR^{+}13, GSPK15, GUK13, GR10, GD10b, Gol13, GM14b, GSZ11, HP19a, HZ10, HBK^{+}19, Hei10b, HG16, HS18, HMSC10, HBE14, HBE15, HTL10, HSC17, HYS^{+}14, HLSN15, HLSN16, HL18b, HLNZ19, HWyL11, Ima17, JK10, JLD19, JQG14, KS15a, KTH13, KVR11, Kup11, KGM11, LMPE18, LD13a, LWR16, LDS10, LZZC12, LCS15, LCS18, LSC17, LLML10, LYN11, LG13, LGL^{+}14, Lóp19, LBL14, MMHGM17, MPGW19, MJ14, OKTR11, OKTR13b, OAY11, OHK^{+}19, OF16, PLR15, PTH^{+}16, PD11]. Boltzmann [QKR19, QWGJ15, RSL^{+}18, RKA^{+}18, Rhe10, RJGS^{+}19, SSHH^{+}18, SK14a, ST15, SPT17, Sag10, SKG^{+}11, SL18a, SG11b, ST14, ST18, Sle13, Ste16, SKFG11, SKTC15, SZP^{+}11, TR14, TTT10, THC^{+}18, TDM13, Tro13, UMLF13, VMC^{+}14, WXYW11, WH11a, WWLL13, WAG^{+}14, WSC16, WLA18, WMP^{+}19, Wan19a, WL13b, WTSS10, WHQ^{+}18, WZHW13, WGY^{+}18, WCZ^{+}19, XZ11, YY11a, YLY12, Yan10c, YB13, Yan13, YSW16, YKKS10, YZMA18, YCW^{+}14, YNS^{+}14, YCS19, ZCSG13, ZZX^{+}14a, ZZX^{+}14b, ZY15a, ZY15b, ZSW15, ZYZ^{+}16, ZYZ^{+}17, ZYZC18, ZHW^{+}18, Zha13, ZM16b, ZHW^{+}11, ZZC13a, ZZC13b, ZPGW16]. Boltzmann-level [ST15, SPT17]. Bona [CdR18, DAM14, HGHA19, KB19, NNWAS11, Oru17, YL18b]. bond [EHO^{+}12]. bonds [ZLL18]. bone [CSU13, GGL13, NCV^{+}18, SS13]. Boolean [LZH12]. Booleans [BGF15]. Boosting [TNT12]. Boosting-based [TNT12]. bootstrap [HLC11]. bordered [Abd18a, JL15a, JL19, JL20, Mar16]. borders [AY18]. Borel [ET12]. Born [KY10]. borne [LZ11b]. borrow [CM18c]. Bose [ZHS^{+}19]. bot [KWPK13]. both [Gen11a, He16, YZMZ16]. bottom [SZA^{+}18]. Boubaker [CM11a, YMDZ10]. bounce [OKTR13a, Ste16]. bounce-back [OKTR13a, Ste16]. bound [AJ12, CChL14, DESV18, DBH^{+}14, DP15, KP10b, WZ10, YLB16, ZY15c, ZSY19]. boundaries [ANP19, BP18, DL16, FGL10, Gol13, SDH13, SG10a, Yan10c]. Boundary [AJRWS12, AAH^{+}18, FHS18, MY10, SbX19, WH11b, Ye19, ZMM18, ZYT^{+}16, ZKWW17, ZHC17, AEG18, AA11, AN11a, ANP11, AW11, AN12, AH11a, AIIZ10, AA18, AER12, Arq18, BZ10, Bai11b, Bai12, BS12a, BNR10, BCD^{+}16, BG19, BCJ19, BJPT16, Bic11, Bog11, Bol16, BT14, BRR16, Boy16, BHH16, CHS11, ÇT12, Cha11a, CHBTD14, Che11d, Che12b, CRXL15, CCCW16, CLL19, CR19, CLJ11, CH17, CHS18, CX18, DPBL16, DGB10b, DL10, DGL12, DN10, DH11b, DA18c, DG10b, Din13, DSM18, DGK10, DBEE11, DB12, DZO^{+}19, DMV11, DBS12, DLT10, ESN10, EF14, EE18, FIVM17, FIMV18, FMS19, FWFL11, FP19, FPB17, FBTS19, FH16, Fer11, FM12b, FXCC18, FES^{+}19, GG18, GWZ11, GCE18, GB16, GH14, GGGR17, GMI12, GKM11, GR13b, GHCZ18, GW12c, GR15, GF19, HSK11, Hak14, HZ10, HLWX11, HP13]. boundary [HH17, HKS19a, HYCP11, HL11a, HL11b, HSWZ11, HBE14, HBE15, HD14b, HW19b, HXL11, HYS^{+}14, HLSN15, Ibr16, Jan10, JM15, JPK17, JPK18, JL11b, JZ11, JK18, KAG11, KA10a, Kar17, KP10a, Kaw15, KKBR19, KFYW11, KS12a, KJ11, Kim11, Kim15, KZ16, KPK18a, KSF14, KHF^{+}19, KVR11, KYA15, LS19a, LC11a, LVF^{+}16, LN98, Li10c, LLZ10, LCW10, LYZ11, LW11a, LW11b, LS11a, LL11, LZ12a, LJSK13, LCK13, LY15, LH16, LFC16, LCP16, LS16, Li19b, LPML19, LJ10a, LZ11e, LZ11d, LZ14, LJ10b, LLML10, LHF11, LCZ11, LLL12, LS12a, LS12b, Liu12, Liu13, LLY13, Luc10, LWC13, LDHH13, Ma10a, MM10a, MDVM17, MCR11, MA17, MT10, MDG19, MW14, MV11, MN10a, MS12b, NKM15, NNAS11b, OKTR13a, PC11b, Par18, Pit12, Pov12a, QWL19, QCG15, RBB12, Ran15, REHA11, RCH19, RR14, SC19a, Saj14, SMF17, SDH15]. boundary [SDH13, SD19, SKPW14, SPST18, uIAA15, SS18b, SG16b, SKM11, Sta11, Ste16, SZ11, Su12, SZW11, SLL12a, SZ12b, Tan17, THC^{+}18, TB10, TAA14, Tur10, WZH10, WWA11, Wan11, WAZ11a, WZF12b, WZM^{+}16, WS17, WW18b, WCH18, WZC^{+}19b, WCCS15, WCZ^{+}19, XD17, XWY11, XZC12, XLD11a, XLD11b, YK18, YY15, YY10b, Yan11c, Yan11d, Yan12b, YZY10, YC10b, Yas12, YDL11, YNS^{+}14, YK17, ZLS19, ZSH11, ZYWZ17, yZjM10, ZJ10b, Zha10, ZBF11, Zha11c, ZS11b, ZD11, ZSQ^{+}18, Zha19c, ZSS10, ZL10b, ZHW^{+}11, lZxLhY12, ZS18, dPLM18, uRK11]. boundary-domain [BCJ19]. boundary-lattice [FES^{+}19, HYS^{+}14, HLSN15, WCZ^{+}19, YNS^{+}14].

Boundary-layer [Ye19].

boundary-layer-resolving [Boy16].

boundary-type [FXCC18].

boundary-value [AEG18, Pov12a, SKM11]. Bounded [Bog10, LXK11, TD10a, AaC19, AA10b, AA10c, BCB11, BR13b, CDM12, DG13a, Das15, EAEH18, FSTN18, FST19, HY18, HWW13, KS11, LKU10, LWKK10, LZKU11, MDRRV11, MD18, NNAS11b, NNAS11c, NUH12, PB12, Wan16b, WZ18c, XL11, ZWH^{+}19, ZH18b]. boundedly [MM10b]. Boundedness [Gao12, MM12, QMW18, WPL16, WHZL17, Wan18, WS10b, ZT19, ZW16a, ZMH16, ZW16b, LMY19, Sed13, WW11a, YÇG12, YDW15]. bounding [Kaw15, fNS11]. Bounds [BS16a, CWQJ12, MF18b, Tia17, ARK13, AJ10, Bag17, DLS18a, DWZ16, HGSL18, Ipe12, LMR19, LD13b, LA11, LY10a, MP11a, Nie10, Pul16, SRG16, ZP18a]. Boussinesq [CN16a, CGO19, JZL18, KKD13, LB18, QY17, SMYK19, WD16, Ye16, Ye17b]. Boussinesq-like [SMYK19]. Box [WYD10]. boxes [HSMG12]. BPES [CM11a, YMDZ10]. BPS [LXZ18]. brain [CB19a]. branch [KL19]. branching [ASB12]. Bratu [EM19, SMK18, AP10, Moh14, RSM17, VAB12]. Bratu-type [AP10, VAB12]. breakdown [CCCW10]. Breaking [CDG16, KTK18, ML19a, Osm18, XX10, YTD^{+}18, ZDM11]. breakwater [LFC16, VHPVNXW18]. breakwaters [LGH^{+}11]. breast [FIS18b, LW12c]. Breather [LZ19c, PTZ19, ZCZ17, CZ18, DTYZ18, HTY^{+}19, LD11b, WTYZ17, YTD^{+}18, ZDM11]. breather-type [ZDM11]. breathers [YHC18]. Breeding [KALAS11]. Bregman [CX16, SCC12b]. Brenke [VSI12]. bridge [MX15]. bridges [TL18]. Bright [HTWS15, CDW11, QTW^{+}18]. bright-dark [QTW^{+}18]. Brinkman [AB10b, GS18, GOS18, ILS13, MR14, RTL19, ZMFL18].

Brittle [Yoo17, MPS18]. broadband [HKW15]. broadcast [IC12]. broadcasting [HB12a]. Broer [DYH11, ZA15]. Brownian [BH11, Dun11, MP11c, TC16, WL11d, XaZH19]. Bruijn [SD12b, XCXW10]. Bruno [XW10, XC11b]. Brute [CJP12, CJP15]. Brute-force [CJP12, CJP15]. bubble [ABR^{+}14, CZY11, HL12, PC14, QWGJ15, ST15, Zha18c]. bubbling [ZLC^{+}11a]. Bubnov [CKSL^{+}14]. Buckling [YW11b, ATH18, ATH19, BKE18, BL14, GBG11, HL11c, RYK13, VLFS12, ZW11a]. budget [Lin10a, Lin11, WL11a]. budgeting [ZHT11]. Bulgaria [GGM^{+}13, Sen12]. Bullen [THH12]. Bullen-type [THH12]. bump [ALHZMC^{+}19]. Burger [EK16, GM14a]. Burgers [BHJ14, aZW17, AB10a, AJAR18, Bra10, DGA18, DS18a, DAM14, EMQ18, HGHA19, HA10, KB19, KSMN11, LCS18, LGZ19, LL10a, LMDL11, LZZ11a, LWZ16, MDRRV11, MK17, Oru17, RC17b, Sah17, YGS17, YYLW19, ZTC14, ZYZ11, ZSD10]. Burgers’-type [BHJ14, YGS17]. Burr [WL11b]. business [CSSW12]. Butcher [CM14]. butterfly [BBM18]. buyer [BO10, KGJ11]. BVP [HY10, HL11a]. BVPs [DC10, FRSW11].

C [LWD15, KRD16, XA13, AM11]. C-M [LWD15]. C/O [XA13]. CA [JL12]. cable [ABH12, KSMT11]. CAD [LZL^{+}18]. cage [DGGBTRJF12]. cages [BFG11]. Cahn [AEH18, BN14b, BMH19, CHY19, DFJS10, GR19a, GGT14, LL14a, Lee16, LS19a, LLJK10, LJSK13, LCK17, LQMW18, LJK^{+}19, LCWZ18, LLYL19, MSW18]. Calaogero [JPB11]. calculate [HD14b]. calculating [BGRS11, JCF19]. Calculation [HK10, SST19, FLP13, GN11, JPCY13, YZMA18, ZKBE16]. Calculations [LHZ^{+}11, BO18b, WY19b]. Calculus [DLS14, AT11, Ana10, AE12b, CPT15, DSVS15, EKE18, GS11b, Hol11, Kir10a, Kir10b, Mag10, MT12, OMT12, PAT13, Sal10, Uze10, VBCJ10, dCM12]. Calderón [DLS14, DSVS15, GGR19, SBB^{+}18]. calibrate [HZ16]. calibration [SYZ19]. Call [KRM^{+}10]. Call-Back [KRM^{+}10]. calls [CCL^{+}12, LZL16]. Calogero [AKMUH17, CM18b]. Camassa [LXY19]. Camellia [Rob14]. Camera [GMAM12]. Campanato [LZG13]. CAMWA [VBW10a, WWW11a]. cancellous [GGL13]. cancer [Amo15, FIS18b, GL17a, KC19, SSM^{+}17, SNSK19, SD15b, VJM15]. cannibalism [BP13, LLY18a]. cantilever [GB18, KLTS11]. Cantor [ZZT11]. cap [LWJ10]. capacitance [LLG^{+}11]. capacitors [ZLG^{+}10]. capacity [CCH^{+}12, RZL11]. capillary [BMSS18, CP15a, TG14, Tao18]. capillary-gravity [TG14, Tao18]. capital [ZHT11]. Capped [LX10b]. capture [IS14, ZOZZ12]. capturing [LX12b]. Caputo [Abd11, AVR17, AJT19, AR17, MT10, MV17]. Caputo-Fabrizio [AVR17]. Carathéodory [KC10, MG11, SOS11]. carbon [Kia16, Kia18, PvdM13, RYK13, TZ18]. cardiac [CT17c]. cardinality [KST10]. cardiovascular [RKD18]. Carleman [ZYY10]. Carlo [AEF15, AGK15, DGLS19, HC18, Li12b, Mil18, MDW13, PA15, SWOF19, VMC^{+}14]. Carlson [CS10a]. CARMA [YD12]. Carreau [GK16]. carrier [LCT12]. carrying [ZWL11]. CART [FNZ13]. Cartesian [Bac14b, FGY^{+}17, LZZC12, LNW19, QCLC17]. Cartesian-mesh [LZZC12]. cascade [FLWJ11, KYY12]. Cascaded [GUK13, FIVM17, FIMV18, HP19a, LBL14]. Cascadic [Bac14c, HRHP17]. case [CM11a, CNV14, CST14, DG13a, FIW17, Izs15, LMW10, RPTD10, SXB^{+}12, ZMG10]. casing [EPP18]. CAT [AE12a, KA11]. cataloging [LEP11a]. catalyst [SCA14]. catalytic [GRW14, Kei13]. catalyzed [ABN18]. Categories [XS10]. categorization [MLY11]. catheterized [SR10b]. cathode [AaC19]. Catmull [PXXZ16]. Cattaneo [XWH16, QXG13]. Cattaneo-type [XWH16, QXG13]. Cauchy [AM10c, AKA11, BKMT14, BKMT17, BCJ19, DG10a, FPB17, Fur13, GKK11, HKK^{+}16, KTDT17, KKD13, LZH16, LR17a, Liu16a, LW18a, LW18b, LLX11, Nab19, SMBY10, YZM^{+}19, ZSL19, ZY10b]. Cauchy-type [Fur13]. causal [ZSS10]. cavitating [CZY11]. cavities [SSPL10]. cavity [AJS14, BK18, HSK11, KG14, MMFT^{+}19, PD11, RSL^{+}18, SK14a, SRRP18, ZZC13a, ZZC13b]. Cayley [LLL11]. CBA [KRM^{+}10]. CBS [Pul16]. CCA [LWHY10, yYqWqZC13]. CCA-secure [yYqWqZC13]. CCD [HP17]. CD4 [MGY11]. CE [RAZ19]. CE/SE [RAZ19]. Cebysev [Sar10, Yan11b]. Cell [CIN^{+}18, GVJ13, WKBR18, ZKR^{+}12, AaC19, BCCZ18, BMS19, BGM19b, CD14, CM19b, CCM14, EOM11, HJD15, HD19, JZE^{+}18, KC19, NGL10, PLKCC12, SRGL13, SW19, TPHD18, YLF19, YZMZ16]. Cell-based [GVJ13, CD14]. cell-centered [BMS19, BGM19b, CM19b, SRGL13]. Cell-element [CIN^{+}18]. cell-to-cell [SW19, YZMZ16]. cell-vertex [CCM14]. cells [AZB13, Amo15, BMS19, Ben12, BZZ^{+}10, CANA19, CS13, DDLM13, EOM11, GZR^{+}13, IL13, MGY11, SCA14, SKH12, WYL19]. cellular [BKR11, Bur13, CCH^{+}12, FNZ13, HBS^{+}10, KRM^{+}10, MAPS10, SAU11, SI10, XGH17a]. Center [hYlLL11, HLT12, WZ11b, YH12, Zha18i]. Center-of-gravity [hYlLL11]. centered [BMS19, BGM19b, CM19b, LR17b, LS17b, SRGL13]. centers [CD10]. Central [DL16, HPV^{+}18, SND19, ZZG19, HP10, JYYL16].

Central-moment [DL16]. central-upwind [JYYL16]. Centralized [VBK13]. centres [WD12]. centroid [CG13, HVA10]. centroidal [FJWW16, WJWW12]. centrosymmetric [LM18c]. ceramics [WvDRG19]. cerebral [OHK^{+}19]. Certain [DRS11, KYR15, Par17a, SRM11b, AS10a, AS10b, ASMEE11a, AX11, Arq18, BK10, EAA10, GCG12, KHUO12, Las10, Li11a, LY11a, LZH16, LR17a, LXK11, MP10a, MP11a, MPZ11, MS11b, NNAS11a, NNAS11c, PA12, RS12a, Sok11, SWW11, Swa10, TÇA12, TO11, WCCS15]. Certificate [LHHZ12]. Certificate-based [LHHZ12]. certificateless [CPL11, HPC12]. Certification [MS13]. certified [Che15b]. Cesàro [Çan11c, ET12]. CESE [SZA^{+}18]. CFD [AVV18]. CFS [MSFS18]. CGMY [HLvS18]. CGMY-process [HLvS18]. CGMYe [GLW18]. Chaikin [WZXL11]. chain [Bae10, CSSW12, Chu11b, Chu11c, Chu12b, ELS11, GTZ19, KGJ11, LM19a, MW17, PQB^{+}16, PQBK17, TZ13, ZSY14]. chains [LKCN19, WHW11, XL11]. challenges [FP19]. chamber [XWL18]. Change [VĆC10, Boy16, CPP10, CWW15, HD14a, HH10a, JFS14, JFS20, Lin14, MHH11]. change-points [HH10a]. changes [KM12, KÖC^{+}18, SKK12, ZGD13]. changing [Bao16, CTC17, CTG17, CT17a, CHZ19, CT17b, DL14, DLS18b, HY10, HY18, LL19a, LL11, LY14, Li18b, LC10d, LC16, Su12, ZT18a]. channel [ASFM15, AM10b, ABV11, CDFP12, CLC16, CV14, DdSF13, Eba14, GD10b, HBK^{+}19, KP18, KLL10, LG13, MC10a, MMA12, MS10c, NPR10, QWGJ15, XZ11, YZMA18, ZJZ^{+}11]. channelling [WXYW11]. channels [AM18, Kup14, LWZG10, Sha12a, ST14]. Chaos [AEG11, DGB10a, Dan12, DZS10, JW11, LYY12, UKAL10, And12, BDGBM12, Bra13, CJ12, CZMZ11, HA18b, Kuo11, PC12, PSD^{+}13, SSAM11, TL12, WYLZ10, WHS11, WY15, Zhe11]. Chaos-based [UKAL10]. Chaotic [TTC14, YMSL11, AsNAd10, Alo11, CZY13, Che11c, dSCM12, DGTC13, DWI^{+}12, FDB13, FDXW11, GGGR17, HGN^{+}10, JW11, Kun12, LW10, MHM11, OBAAD10, OSZP13, SZDO10, yS10, TL12, WY11a, Wei12b, ZTSC16]. Characteristic [ELS11, LR17b, BSS18, LPLR19, Li17b, LZ18c, RT10, ySW10, SLW14, XFH19]. Characteristics [YTD^{+}18, HCZ16, Ikh11, LHZ^{+}11, Luk11, PHM^{+}19, TZXP11, YS19a, ZWL11, ZCSG13]. Characterization [Sar11, Sha12a, ZW11c, CP16c, LSM10, PS12b, RS14a, ŞGY11]. Characterizations [DSA10, MYZ12, SK10a, SJN10a, SM11, LN10, LY10b, RdSSS11]. characterized [SJN10b, SGK18]. Characterizing [HDT11]. charge [Yaz11, YAS^{+}11]. charged [Yaz11, YAS^{+}11]. charging [YLH12]. Charney [Boy16]. chattering [CZMZ11]. cheaters [MVB^{+}12]. Chebysev [Gav12]. Chebyshev [BE12, Boy16, DBEE11, GTL16, Jav11, KB10a, SK14b, WMW13]. chemical [BG13, CYM13, HTGSH13, LZP^{+}19, MN17, ZWC10]. chemically [BKR10, RSV11, VPR11]. chemically-reacting [BKR10]. chemoattractant [WHZL17]. chemostat [WWG10]. chemotactic [DGZ13]. Chemotaxis [MGTH16, CST14, CKRW19, FHA16, HLZM16, IS14, JLL19, LMY19, LMZ18, QMW18, SST12, WPL16, Wan16b, WHZL17, Wan18, XFH19, XBHN16, ZT19, ZW16a, ZMH16, ZW16b, ZMWH18]. Chemotaxis-driven [MGTH16]. chemotaxis-fluid [CST14]. chemotaxis-growth [QMW18, ZMH16, ZMWH18]. chemotaxis-haptotaxis [LMY19, ZW16b]. chemotaxis-Stokes [WPL16, Wan18]. chemotherapy [SD15b, WHLC11]. Chen [KM10]. Cherub [JCZZ13]. chest [LWWY12]. chi [Cvi10]. Chinese [FPW^{+}11, ZQ11a]. chip [SSESG12]. Chlodowsky [Büy10]. Cho [KSJ12, SPLHCB14]. Choice [YC11, Nes10, Rus16]. choices [ADGG13]. cholesteric [HMSC10]. Choquard [Amb19, FZ18, LT18, Luo19, ZXZ17, ZT18a]. Chosen [HPY10, GH12a]. chosen-ciphertext [GH12a]. chromatic [IB10, KP10b]. chromosomes [MCQ11]. Chua [RMM11]. CIM [LEP11a]. ciphers [QGGL13]. ciphertext [GH12a, HPY10, HKHK13]. circle [BCB11, DWI^{+}12, LTX^{+}13]. circuit [RMM11, SST19]. Circulant [LLLC14, IB10, PB11]. circular [AVR17, Far11, FAIV10, Hal13, LPML19, Lu12, LLL16, QZY11, SB14, THGG14]. circular-linear [QZY11]. clamped [KTA12]. Clark [PXXZ16]. Clarke [LMZ17]. class [ABJ11, AEO15, AJT19, And12, AB16, Ara18, ANR11, AGK15, APTZ19, Bai11b, Bai12, BM10a, BR12b, BM10b, BMTV12, BNTT14, BX14, BKK12, Bre14, CHS11, Cao19, CFRS10, CM16b, CTC17, CM17, CO19, CN11, CN16b, DWY15, Dan12, DT16, DSZ18, DZ21, DC10, DSL18, DRS11, Eba11, Ehr18, EAAED10, FWZ16, Fer11, GYH11, HHY13, HY18, HKT11, HGW18, HP19b, HR15, HXL11, HM17b, HGW11, HM14, HM15, HWXC18a, JMLF11, Jal14, JW11, JL19, JL20, JZ11, JKS18, KCL14a, KCL16, Kir10a, KK10c, KV18, KHUO12, LCS15, LZD17, LM19b, LMRS10, LLH11a, LLH11b, LZB12, LX15a, LLT16a, LLT16b, LYZ17, LL14b, MP10b, MN11a, MS11a, Mar11b, MM18d, Mor10b, NT17, NM11b, Ouy11, PRR18, PA12, PLT17, RS12a, RZ16, RWW18, SD12a, Sed13, SBS12, Smo17, Sok11, SWW11, SGQ12, SLW18, Swa10]. class [TJ10, TM17, TÇA12, TY16, THY^{+}10, UBF11, Vel15, WC10b, WSCL11, WKS13, Wan15, WZF16, XY10, XY11, XZ19, YY11b, YM13, ZA15, ZM16a, ZD11, ZLY14, ZYW15, ZZWG16, Zha19a, Zha14, Zha15b, ZJZ18, ZZT11, ZT18a, lZxLhY12, ZLW19]. Classes [Sok10, AE12a, AKS10, ASMEE11a, Arq18, BBD10, Dra11b, EAA10, Liu11b, MN10b, NNAS11a, TTG16]. classic [EGSHR10]. classical [CTS19, Iom18, Li16a, Uze10]. Classification [jC11, ALLQ13, BSN13, CLW11, GAVOF11, Hua12, LXF11, LW12c, MWWL11, PTL13, SYO12, SS16a, TJQS13, WLYX13, Xia11]. classifier [FNZ13, Pes13, VKJ13]. classifiers [KYY12]. classify [LCA^{+}17]. classifying [Tol12]. clause [LWJ10]. Clenshaw [Som13]. Clifford [FRZ15]. Clifford-valued [FRZ15]. climate [eMA18, CWW15]. climatic [ZGD13]. clinical [PATA11]. Closed [PMM17, SMDI18, SBM13, WFDW10a, JMST11, KTK18, KW12, LCN10, LC13, THY^{+}10, XSLS11]. Closed-form [PMM17, SBM13, WFDW10a, LC13]. closed-loop [KW12, THY^{+}10]. closure [BMS19, CRG16, GJX18]. closures [BS11a, NN13, YG11]. cloud [HSS^{+}12, QXLL11]. Cluster [RKW12, CW10a]. Cluster-based [RKW12]. Clustering [Che12a, FPW^{+}11, HYL10, fLcJ10, NGL10, NHIN16, SRS11, THZ^{+}11, YYL11]. clusters [KCC^{+}13]. CMA [LWHY10]. CMRH [AT18a]. CNPE [LK14]. CNTRC [ATH18, ATH19]. co [KLTS11]. co-planar [KLTS11]. coagulation [WL11d]. coal [RB19]. coalesced [ST18]. coalescence [BKL14, DFJS10]. coalition [XKH10]. coalitions [Bur13]. coarse [DWZ13, Jum10, KHWK10]. coarse-grained [Jum10]. Coast [VBW10a, VBW10b]. coating [Lin14, SISH12]. coaxial [ZJB19, ZLZG11]. cobweb [MM11]. Code [CCHG17, BDM19a, BB15, HMP^{+}15, MCQ11, PG10, ZKBE16]. codec [YZ12]. codes [BKL^{+}19, ÖŞ11, SKdA11, TBP19, YS12]. coding [EE10, Sha12b, WLC11, WHC12, ZPS^{+}12]. coefficient [CL19, CTSX16, CRG16, CYL17, CTM^{+}13, Dem10b, ES17, FL11b, GR19b, HJ13, IMS19, KK19a, Lam12, Laz10, LWL11, zLYLQ19, LGVS19, MMOJ14, MAS11, MG11, NHIN16, NM11c, RSS16, RS14b, SMC10, Smo17, WLZ^{+}18a, WRY18, WZC^{+}19a, WLXZ18, WTLS18, YTC^{+}18, ZWH^{+}19]. coefficient-based [FL11b]. coefficients [AJRWS12, AEDL14, BH10, BNTT14, BBO10, CHLY15, CUK12, CNP14, DWZ16, DGLU18, ES18, Erg19, FZL^{+}18, GC19b, HH18a, HP17, HLI14, HLY17b, Ili10, JS11, KN11, KÖ10, LCS15, LXZ18, LN19, LZ19c, MDW11, MHHC18, Mar11b, MNT15, NLA19, Oan13, Osm18, RAD13, SD10b, SK11b, SYG11, SRGL13, SLM16, TF17, TUT11, WW18a, WS10b, WK13, XZ18, YH12, ZYG10, ZW11d, ZZL15]. coercive [BR16]. coercivity [Cia12]. Coexistence [MCL15, WWG10, ZLZ18, MY16]. cognitive [Ogi12, OYXK12]. Cohen [Hua11]. coherence [Mah14, ZLY14]. cohesive [JQG14]. Coincidence [Che12b, ĆSCD11, YBC11, Ber12, CM10b, LZ12b]. coincidences [CYM13]. coli [XBHN16]. collapse [MLSLM15]. collars [MTV13]. Collision [BSS18, OVV^{+}16, OHK^{+}19, TDM13, WL11d]. collision-stream [OVV^{+}16]. collisions [CM13a, DLF^{+}11, HTWS15, OVV^{+}16, Zha13]. collocated [PBM19]. Collocation [KAG11, AuIA17, BMAR18, BPF13, CCNT16, CDY11, CN11, DGLU18, ESL11, EE18, GCE18, JSGP16, JL18, KS12a, KBK19, MN11a, MRS15, Moh15, eOS18, PLT^{+}19, QX19, RFK16, SYG11, SYW11, YZX18, YCHW18, YQWZ19, Yüz12c, YK17, ZZL18a, ZT18b]. colloid [GQF^{+}10, ySGL^{+}10]. colloidal [NWZ11]. colony [qGpWhL11, HB12a, HCHH12, PH13, TL12, WQNF12]. Color [LW10, HCL11, JMNZ19, LHLH15, TT12]. colored [LD11a]. colorization [JMNZ19]. colour [BSL11, BLS17]. Colpitts [MMMG12]. column [GBG11, Liu14, MLSLM15, Tsi11]. columns [VLFS12, YW11b]. comb [LZCL18]. Combination [JY11, Mil18, SWS19, YLY^{+}09, AD15, CL16b, DA18b, JWX^{+}13, LSV18, RS12a, ZAK18b]. combinations [BM11b, PJ17]. combined [BN14a, DFGG13, DNS18, FHZ13, GWR^{+}18, HA18b, KVR11, LZLL18, WLL^{+}18, WN18, WO18, ZGD14]. combustion [XX19]. coming [BDM19a]. commensurate [CB11d]. Comment [Abd18b, Asl11, EM19, LK15, Pan17]. Comments [Ant14, DZ21, Her14, Her19, NTR15, Gro19]. commodity [LNKU12]. Common [Ahm10, AIB10, CSW11b, KPR10, LEP11a, SCK11, ZLZ11, Cha11b, Che11b, CAP10, CS11d, CM10b, KKS10, KSM12, LY11d, MAK12, NS11, Pir11, Qiu12, RI12, VBK13, WW10a, YBC11, YLK10]. common-opponent [KSM12]. communication [HTWS15, HPY10, KRM^{+}10, PKTH13, WD13]. commutators [Gao12]. Commuting [RWW18, DD16, NS11]. Compact [RZ16, AM13a, AZ17, CCBSRFRM11, CT17a, CT10a, Fis18a, Gal11a, Lee11b, LS19a, Li16b, LLYL19, MKHC11, Mah10a, Mah10b, PCO16, PC17, QKR19, QZF19, TMDTTC16, WH10, WWA11, WW11c, WV16, WW16, WZM^{+}16, WW18b, WH18, WX18a, WCLD18, YLG17, YT13, YTZ17, YT18b, THGG14]. compactness [LS10b]. company [HHS^{+}10]. comparative [CGK14, CJRR11, EAMA19, FMGR19, JRB15, KRP12, LL10a, uIAH10, ZZX^{+}14b, ZYZ^{+}17]. comparing [BPG10, Bol16]. Comparison [Ala10, BKY10, CEF^{+}13, CHK^{+}18, CCFV12, CFdM^{+}18, GSY10, GT15, GT16, HSWZ11, LG17, LJK^{+}19, UKI11, VMC^{+}14, WZHW13, AO10a, AVV18, Che11e, CSCM13, HB19, IB11, KHIB12, PMM17, PKD19, Yak11, YB13]. Comparisons [WLW16]. compartmental [dPRVRB13]. compatibility [JK18]. compensation [LCT12]. compensator [Yan12a]. competing [MY16, TM18]. competition [ACTB19, BF11, BD11d, CS16, CP16a, GCG12, JWX14, TT12, XZZ16, YW10]. competition-diffusion [CS16, YW10]. Competitive [BK16, MY16, AIA13, ZC11b]. Competitive-cooperative [BK16]. complementarity [CM11c, FZ19, Kha10, LLH11a, LLH11b, RWTW19, TM12, ZL14b, Zha14]. complementarity-type [Kha10]. complementary [QZ11, YGR11]. Complete [DD16, RTB14, CWQJ12, CP16c, HZL17, KZ16, LZH12, LC11b, NS11, Raf12, RT11, YWL17, pZ10, ZYSY17]. completed [DZW16, DLZ17]. completely [SKJ10, SGQ12]. completion [AA10c, ŞGY11, WWXW19, YY10c]. Complex [LK14, WCZ13, BR12b, CM16b, CW18, FCZ12, GKLR11, HD14b, HWXC18a, HWXC19, IQR16, KK13a, LBZL11, LZT11, LJSK13, LM19b, LZ16c, LLL16, LRTV10, MZC17, Mag10, Mah10a, NZ16, NNAS11b, PZ11a, RTRR18, RI12, SPS^{+}13, SS18a, TPHD18, TN11, TMCM19, Uze10, XY17, XWY17, XWY18, YKA18, ZZ15a, ZM16a, ZSZ17, ZH11, ZDLC14, ZY15a, ZYT^{+}16, ZWLZ18]. complexes [Oh15]. Complexities [Kei13]. Complexiton [Üns18, WZMY18, LZM18]. complexitons [ZM17b]. Complexity [GY13, PL10b, SS11a, AAR11, Bur13, Dej11, HTGSH13, LCT12, Lit13, LW11d, XA13, YZ12, ZZF18, ZM13]. compliant [HZLM10, LL12a]. component [BN14b, CLW11, DB10, Elb11, MM19, RZZ19, WhJxLwW11, WCW13, YB13, YYC11, ZaY17]. components [Che12c, CSS10, HGJP19, PLKC16, SNEP19, Yan19, ZC11a]. composite [BI12, BL14, BML11, CGY10b, CGY11, CHS18, DD19b, DB10, DCN^{+}18, FVZ18, HD14a, Kia16, MM18c, PHM^{+}19, RYK13, STC18, Sza15, Tim13, ZTW^{+}19]. composite-type [CGY10b]. composites [BCSCB^{+}15, Ryl15a, Ryl15b]. composition [GGM^{+}13, LG12, Özd18]. compositional [DVY14, SSS16]. compound [CGK14, GMZ15, HSBL11, YLS12]. compounds [DB10]. Compressed [BMP15]. compressibility [OAKR16, OAY11, KP18]. Compressible [HMF16, AP19b, BSZ16, FSTN18, FL13a, GLP18, GGO16, HKP17, HG16, JVMF19, MA16, RQ18, SSS16, Ser19, YSW16, YM13, ZR18b, ZR18c]. compression [Che15b, GA10, LW11c, TZXP11]. compressive [WLW17]. Comput [AR10b, AD11a, Asl11, Bho14a, ÇE10a, Def10a, fDxZ11, DZ21, GXZ11, Her14, IIHuI10, ID10, JL20, Jia11, JFS20, JY11, KPS10a, KK11b, KYR11a, Li10c, LYS12a, LLH11a, LW12a, MP16, McN12, Pen11, Ran15, SK12, Ver12, WWW11a, WWW11b, WYG12, Wan13a, WLDL11a, XLD11a, XWH16]. Comput. [Her19, Pan17, SCBCB^{+}17, SKTC19]. Computable [RSH18, DESV18, Kim14, ZP18a]. Computation [Bri10, ÇA14, DVMS13, GLP18, Hei10a, LZL^{+}13, NKM15, Pap15, YH12, YWL^{+}11b, Ade17, ARK13, ADD^{+}15, CDP16, Che11e, DLZ19, DC15, FKKS11, LSX13, LMZK16, MMRN12, MGN^{+}16, MNT15, PW18, Pet14, SSP13, SM14, TXL19, XX10, ZHW^{+}18, Zha18i, lZxLhY12, dPRVRB13]. COmputational [SVP^{+}19, BCHS18, BP11a, Erv12, EAAS18, HD14a, MAN^{+}15, NPR10, RF12, WKP^{+}14, ABL15, AM12a, ABK^{+}13, Ben17, DCN^{+}18, DHMU16, FT15, FMGR19, GGK18, HHG14, KLK15, KÖC^{+}18, LKL^{+}15, OC14, PRS18, SP12, WCB13]. Computationally [LCK17, AGK15]. Computations [AEF15, DD19b, GIM15, BCF^{+}14, FGHZ14, GUK13, GD10b, NHH13, RA18, ST15, SPT17]. compute [SC13]. computed [CX16, DRT^{+}15]. Computer [Ben12, GGAVRC^{+}19, LY12a, SK12, DSB19, JK12, PHWM10, RMS12, SK10b, SJS^{+}11, Zha11b]. Computers [Ant14, LK15, LZL^{+}13, Yaz11, YWL^{+}11b, CBM10, Kan15, PHWM10, RPTD10, EM19]. Computing [BK14, BRS11, BDK^{+}11, CJPR10, CEJV16, GS12, Haj18a, Ji14, SPCS13, TMMASG10, Als10, BHJ14, CCJP11, CS14b, DNS15, Dos18, GPV11, KTDT17, KL16b, KORR10, LCQF19, LXF11, MR15, Ned12, fNS11, OYXK12, PL10a, PV12, TL10b, XSLS11, YWK^{+}10, YWW^{+}12]. concave [LS17a, ZC10]. concave-convex [LS17a]. concentrated [DVMS13]. concentrating [ANP19]. concentration [Amb19, CHT11, MC10b, YL16, Ye17a, YW19]. concentric [BTEM19, WL13b]. Concept [HBE15, WL11c, ABFGZ11, BS12b, Med12, MHH11, SGK18]. concepts [FF14]. Conceptual [HGJP19]. concerned [KHUO12]. concerning [Bai11a, DFG19, QYL10, SKJ10]. concrete [LHM11]. concurrent [GPV11, PV12]. condensates [ZHS^{+}19]. condensation [PÁAP^{+}15]. Condition [WH11b, AM14a, AEDL14, AKS11, BK11a, BKT11, BRR16, CT17a, CLJ11, CC17, DG10a, DA18c, Dia17, DSM18, DBS12, GG18, GB16, Gon13, HBE15, HQ19, HXL11, Ipe12, JPK17, LS19a, Liu13, LWC13, MC10a, MKL11, MDG19, MGW11, OKTR13a, RR14, San12, SS11c, She18b, SS18b, Ste16, THC^{+}18, TNHK19, Wu18a, YY15, Yan10c, ZZ15b, ZWY19]. conditionally [LW19a, MDRRV11]. conditioned [BT14, HT16b]. conditioning [Jaw13]. Conditions [SCSF19, AN12, AER12, AYH17, AHOP18, AKRT14, jASzZ12, AD12, Arq18, BKL14, BM18c, Boy16, CM18a, CP15b, CHBTD14, CDN19, DPBL16, DGB10b, DG10b, Din13, FPB17, GH14, GGGR17, GKLR11, GN12, GV11a, Goo11a, GKM11, GW12c, GR15, GH12b, HSK11, HKS19a, HL11a, HBE14, HZ11, JM15, JPK17, JPK18, JL11a, JL11b, KT11a, KC10, KJ11, Kim15, KZ16, KP19c, KSF14, KHF^{+}19, KH18, LC10b, LZ12a, LJSK13, LCK13, LPML19, LH10a, LJ10b, LLML10, LHF11, LLY13, MCR11, MA17, MT10, Mar12, MV11, OC10, RMS10, RNB11, SMF10, Saj12, Saj14, SMF17, SDH15, SD19, SOS11, SW12, SPST18, uIAA15, SRM11b, SZW11, SLL12a, Swa10, TTMJ12, Tan17, eT10, TAA14, TÇ11, TÇ12, WC10b, WAZ11a, WL12a, WZM^{+}16, WW18b, XD17, YL10a, Yan12b, Yan13, YDL11, YK17, Zha11c, ZS11b]. conditions [ZD11, ZZM17, ZZL15]. conducting [SL12, STS19]. conduction [CZF10, DCN^{+}18, FXCC18, GHCZ18, Hei10b, KK14a, LX12a, PK19, QXG13, QWL19, SWL19, WCCS15, XWH16, ZSQ^{+}18, ZSLZ19]. conductive [EPP18, GRS18, Ye19]. conductivity [BD19, KWFY11, MM18a, WCCS15]. conduits [AHF16]. cone [AR10d, AHF10, ASMM11, CM11c, CSW11b, CM10b, GJ10, HS11b, IN10, KPR10, Kar10a, KLMV12, MCB10, Pan17, REHA11, Sha10, SCK11, YBC11, Zha11d, ZLL12]. cone-invex [GJ10]. cones [AH10b, TM12]. Conference [BCHS18, VBW10a, VBW10b]. configuration [GGGR17, Ögü13]. configurations [DCN^{+}18, HR14, Mic17, SBA10]. confined [XWL18, ZCSG13]. confining [YCS19]. conflict [JTCC11]. confluent [Res16]. conformable [AK18, CJ18a, EGAA19]. conformal [LLLC14]. Conforming [GM19, BS17, DFG19, RC18, SDH15, WCSW18, WBZY18]. congestion [BBO10]. congruent [KK14b]. conic [Ahn10, NNAS11a, NM11c, NM11b, ODR10]. conical [GZ14, MG11, WZ18b]. conjecture [HL10, LEN10, SKJ10]. Conjectures [Deu10]. conjugacy [Dos12]. Conjugate [RJGS^{+}19, ZY17a, ABR10, BR18, Don10b, DFS14, Haj18b, Haj18c, HP13, HLSN15, HM17b, HM14, HM17c, HM18b, Ima17, KLK15, KM14, LZD17, LFZ19b, OP14, TM17, TNP17, WZ17b, WFDW10a, WFDW10b, WLDL11a, WLDL11b, XM15, ZT13, ZY17b, ZXW13]. conjugation [SKPW14]. Connected [Bog10, DRS11, EAEH18, JMLF11, LW18b, MW11, TKH10, VDV13]. Connection [GMAM12, GJX18, LLL11]. connectivity [DT11, LLWZ11]. conquer [FNW18, LLCG16]. Consequences [GOGYL^{+}11]. Conservation [LHD18, MK17, Zha17b, Abd18b, BKZ17, CG13, Col14, EK16, FM19, Kim15, KPS17, Kim18, LXY19, LDG19, Ma19, MW13, RR18, Ray17, RS18a, Ray18, WLXZ18, YÖ10]. conservative [ADGS18, hCTM11, CL17b, DZ16, EDC14, JW19a, Lee16, LJSK13, Li16b, MM10b, PS12b, Was13]. Conserved [AK16, JQSS12, MSTB17]. conserving [LD13a, Por18]. Consideration [CJP12, CJP15, HD14a, CZ10]. considered [ZS11b]. considering [GGGR17, LYJ15, MAPS10, SSIP19, SJS^{+}10, TTMJ12]. consistency [Gro19]. consistent [CTM^{+}13, YMHL18, Yu11b, ZHV19]. consolidation [BKR^{+}19, BKNR19, Fu19, HMY18]. constancy [Rey12]. constant [AEDL14, BRS11, CM12a, IBB10, Kim14, LZH16, MLL16, Mor10c, SISH12, Wei10a, WHG11, ZL11]. constants [Pul16]. constitutive [HNPS13, LZCL18, NB11, Pov19, XJ18, YLDL11]. Constrained [BMS19, KKL16, SPL19, AG10a, AO18, AD11d, CLH13, Che15a, CChL14, DBH^{+}14, GYTD12, GHMN16, Haj18c, HR15, HM17d, KFTT13, LS11c, LH12a, LL15, LW18c, MP19a, Ped18, PW10, PSS18, SA16, WZ17b, ZLGL11, KM18]. constraining [Loh16]. constrains [KKBR19]. constraint [Dia17, GN11, GWL11, LZ15a, Lin11, LC10d, MMR11, ZCY16, ZZX16]. constraints [BB18, DPBL16, DM12a, DP15, FGPP17, Gao15, HA18a, HC14, KLY16, MN11b, PZ11b, PSS18, TCM18, Yoo17]. constricted [GD10b]. construct [KCL12]. constructed [ABRL18, ID10, JW05, KBS11]. Constructing [HWW13, MZ10, MI16, gZnZpZbD12, Zha18i]. Construction [BGIN13, GK13, HSMG12, KAS11b, LSX13, NPD17, NB17, Ben17, Che14, Gur13, LSW16, LW11d, ST16, Som13, WZH18, WY19b, XG10]. Constructions [SKdA11]. Constructive [DC10, WKBR18]. Consumer [LY12a]. consumption [WHZL17]. Contact [WHQ^{+}18, BS15b, BHM19, BDS17, BZKR15, Bri10, CLM12, Che12c, CXMO19, EKZ17, HZL17, JO19, LRV13, RZ17, SH12a, SbX19, XHH^{+}19, ZR16]. container [HLT12, RT11]. containing [EZM12, PK19]. containment [GIMZ14]. contaminant [YYK16]. contaminated [ID16]. contamination [DA18b, LHL15]. content [xLlFwWL12, Zho16]. content-based [xLlFwWL12]. context [AFGL10, BCD^{+}16, KK12, OYXK12]. contexts [LML11]. contiguous [RRC11]. continua [Kim17]. Continuation [AS15b, Che19a, SSC19]. Continued [SVY16, ZZT11]. continuities [Çak11a]. Continuity [Goo10, SSL11, Zha18h, ZZ16c, Çak11c, Sca11, ZYWZ17]. Continuous [AJY13, XY11, AF13, Ahn10, AHOP18, CDG15, CEF^{+}13, CTM^{+}13, CT10b, Ciu11, Col14, Dan12, DFGG13, DD10, DW18a, DMRS18, Dra11b, FWW14, Hof18, KVJB15, KV17a, KV17b, KdLK19, Lad16, LHY18, Mar12, Nie10, OBCG19, Ols10, PP10, PGF18, PLKCC12, PLKCC13, SDH13, TWLYÖ10, WLM13, Wan19b, WLGL10, ZM18, ZZ16a, ZLL17]. continuous-discontinuous [DFGG13]. continuous-installment [Ciu11]. continuous-time [CTM^{+}13, Mar12, WLGL10]. continuum [BKK12, KBGC12, LWL14, NEB14, ZSY14]. contour [CCSZ14, FQLC18, HT18, JW19b, WHD14, XJYL17, Yan18b, yYsZyYL13, jZsQdLmG19]. contours [KBDC12]. Contraction [CNV14, AKT12, ABH12, ASV11, BBR10a, ĆSCD11]. contractions [CsH10, KPR10, KR11, KS10a, Kar10a, KL12a, RK10]. contractive [AKS11, APS12, Cha11b, GDZ11, HLS11, KK10b, LC11b, SS11c, SH10, Zha11d]. contrast [GGR19, YXX11, ZLW18]. contributions [Per18, WLHZ14]. contributon [YZMA18]. Control [CL12b, Dav17, HL12, JMB10, LY12a, VBW10a, VBW10b, AKT12, AMA14, Akm15, Ala10, AR10c, AHO16, AEG11, AAA12, AT19, ADL12, Ant14, ADS14, AM17, BKT12, BZK12, BR12b, BMJ10, BMM12b, BMS12, BKDM13, BL12, BPKM10, BS15c, BHH16, CY14a, CN13, CZMZ11, CL12c, CLH13, CHLY15, CRXL15, CS11a, CCK18, DZS10, Def10a, Def10b, DNP15, DGGBTRJF12, DLWW12, DMV11, DD13, ES10, EF14, EE18, FGHZ17, FOS19, GG18, Gal11e, GRW14, GS15b, GS19, GNP14, GN11, HSMY12, HM18a, HMM12, HLL^{+}15, Hou15, HLY17a, HSC17, HLCY12, HC14, HLZ15, HZP18, IS14, JS12b, JW11, JCWZ16, KBGC12, KP19b, KSMT11, KKL16, KKL^{+}13, KTH13, KYA15, KRBS18, Kun12, KLY16, LZ11a, Lan12, LCH19, Li10b, LZT11, LZY12, LLC13, LZC13, LZZ18, LCYC12, LYY12, Lin12, LL12c, LXP11, LH12b, LMZ17, LDY11, LLZ12]. control [Mac12a, MM16, MK18, Mop11, MN11b, MM19, NMR15, eOS18, PS18, PLT^{+}19, PM13, PSS18, RES10, RGVR17, SSAM11, SZGG11, SCC^{+}12a, SSM12, Śmi11, SbX19, SJS^{+}10, SJS^{+}11, TC10, UKYK17, Vel15, WCB13, WZCC10, WL11a, WS12, Wei12a, XL15, XLK11, YHZY11, Yan12a, YWHC11, YLH12, YLC12, Yil19, YSB15, ZZHF12, ZLLF12, ZZ10a, Zhe11, ZFC11, ZG16, ZZX16, ZP18c, FIS18b]. Controllability [DB11, Guo12, TG11, BKT12, BZK12, DGL12, Gao15, HP19b, HLW19, LMZ17, MB10b, SRM11a, SSA12, Wei12a, YT18a, ZG14]. controllable [Zha18i]. controlled [LC12b, RKP12, YD12, YZAX10, Zha18a]. controller [Ala10, Che11c, CL12b, CL12c, dSCM12, EK13, HHY^{+}11, Kuo11, LC12b, OSZP13, Zhe11]. controllers [CFRS10, HK10, PC12]. Controlling [Che11c, BFF^{+}11]. controls [GN11, KW12, WZWX11]. Convection [ASMM11, Pan17, AHF16, AM10a, Akm15, AM10b, ABV11, AG11, AJS14, Bac14a, BDM^{+}19b, BSK11, BP18, BS14b, CKSL^{+}14, CGGM19, CEQ14, CHBTD14, CM14, CCM14, DWZ13, DZW16, Das12, DN18b, GSZ14, HSK11, HO19, HLSN16, HLNZ19, HHM12, Ima17, JJ19, JLD19, KNT12, KW14b, LCM14, LCS15, LZ19a, LZ15b, LZ18c, MD18, MCB10, MSV18, Nes10, Oru19, QZ16, QZF19, RSL^{+}18, RKA^{+}18, RBB12, Ran15, REHA11, RKF18, RS14b, RRP16, SPH10, SZZ11, SRRP18, ySW10, SG11b, SLW14, SSPL10, SZP^{+}11, Tro13, WV14, WLL^{+}18, WZY13, XFL16, YSB15, ZHZ14, ZHJ14, ZLL17]. convection-dominated [CHBTD14, Oru19, ySW10, SLW14, ZLL17]. Convective [Yan13, GR19a, MA10b, MC11, MCR11, MS10c, NPR10, PDN19, SSSB11, UKA15, UKYK17, WWLL13]. Convergence [Ara18, BNTT14, BGH14, CH11a, DGR18, DMZ10, DSZ18, DWS19, DFW^{+}18, DYWL19, FIS18b, HH16, Kar17, KA10b, KYO10, KT11b, Lad16, LBJ10, LSD10, Ma10a, Mah10b, MKL11, MP11c, MVB^{+}12, Naz13, PRS18, RS15, SL16a, Śmi11, SOJC10, SJC14, TT14, VGK^{+}16, WKS13, WHG11, WH14, YXS10, YX11b, Yan15, ZS11a, ZRC11, ZZL^{+}18b, sHC11, AE12a, AT17, AA13, Ban13, BK11b, Bra16, BC17, CsH10, CB11b, Çan11b, CM16a, CCG18, CQ13, CAH11, CS10c, CS11c, CJ15, CdR18, DN18b, DD19a, Deb12, DD10, DH18, DBH^{+}14, DCRL13, FHZ13, FM18, FT15, GY11, GGGR13, Gon13, GDZ11, GH18, GD16, HP19a, He16, HRHP17, Hua10b, KS10a, KBS11, KD10, KK10a, KK11b, KA11, Lee11a, LjHO10, LM17, LM19b, LBW11, LR13, Mia18, MS12a, Mor13, MT19b, MME10, Ngo18, PS12b, Pir11, QHW11]. convergence [ŞGY11, She11, SL16c, SV11, SYW11, zSdZ10, TÇ11, XZ10, XY15, Yan17, YCHW18, YQWZ19, YH19, YLLN16, pZ10, ZSZ17, ZLL11, ZSY14, ZL14b, ZH19, ZL16, ZZ14]. convergent [AHP^{+}14, ABLS15, CG14, Dun18, LT13, LN98, Li10c]. converging [CM12a, Mor10c]. Conversion [ID10, JW05]. converter [Ala10, GRW14]. convertible [ZLL18]. converting [GR13a]. Convex [SLL17, Ahn12, AKS10, ADK10, CHLY15, CyL11, CCK18, Dra10, DRS11, Gue13, Hal13, jHlXZ11, IBB10, JW18b, KA10b, Lee11a, LS17a, LY10b, LL15, LFAL19, MLG17, MSFS18, Noo11, PT11, RS12a, SSO10, Set12, SK11d, THD11, Wac10, ZJ12]. Convexification [KK19a]. Convexificators [GN12]. Convexity [BBD10, GKS10, MB11, Mos10, ÖAK11]. conveying [XXG10]. convolution [AD12, SG16b, Sok10]. Convolutional [WRW13]. cooling [DLWW12, KDG11, MA10b, MC11]. cooperative [BK16, CL16a, YWL17, ZC11b]. coordinate [Boy16, GS11a]. Coordinated [ES10, LZY12]. coordinates [BJPT16, BE18, GV11b, Gue13, Wac11, WV16, ZY15a]. coordination [KGJ11].

Córdoba [ZWY19]. core [BC11, DÇ12, Fia15, ZMM18]. Coreduction [MW10]. cores [AH10a]. Coriolis [HBK^{+}19, ZYSY17]. Corner [FM18, CK15, Cho17, CCK18, DBS12, KL16b, SR10a, WZXL11]. corner-cutting [WZXL11]. corners [dS16]. corona [YKRV11]. coronary [LYY12]. corporate [ZLY12]. correct [ZPGW16]. correcting [KM12]. correction [AZ17, ACC18, EE10, FIVM17, FIMV18, GSD^{+}19, PBM19, XJLX10, YH15, ZHJ14]. corrector [AGPCC10, KBAF18, WFL11b, ZC11c]. correlated [LLZ11]. Correlation [LMS13, Sha12a, SS19]. correspondences [IN10]. corresponding [fLcJ10]. Corrigendum [AD11a, Bho14a, ÇE10a, Def10a, GXZ11, IIHuI10, Jia11, KPS10a, KK11b, Li10c, LLH11a, LW12a, MP16, McN12, Ran15, SCBCB^{+}17, SKTC19, VBW10a, Ver12, Wan13a, WLDL11a, XLD11a, Yaz11]. corrugated [AHF16, AB10b, GK16, GK18, HHM12]. cortical [GGK18, TH13]. cosh [FYYT11, SH11]. cosmological [MLL16]. Cost [CJP12, CJP15, Chu12b, DB15, DWZ16, DH10c, LCH19, WCB13, WKP^{+}14]. costs [CJPR10, LYLX11, LW17, LC13]. Couette [FL13a, GSZ11, HAESLB14, HLB14, JNJ^{+}11, LZZ11b, MC11]. Coulomb [HKS19a]. countable [CGHY11, CS11b, CS11c]. counteracting [CZ11b]. Counting [DDD10, QCYL12, SS11a]. Couple [Kar10a]. Coupled [AKS11, APS12, Ber12, GYTD12, LHL15, LT11, NS11, SSIP19, AB10a, Ade17, AAA12, AS19, BR18, BS11b, BT15, BTB18, BO18a, BS12b, BK15, CST14, Cha18, CC19, CM17, Che18, Che19b, DM16, DSM18, Elb15, FIVM17, FIMV18, FWFL11, FJP18, GRS18, HB19, Haj18b, HC16, HBK^{+}19, HQ19, HM14, HTWS15, HM17c, HM18b, JKB11, JKMS12, KTDT17, KRD16, KKSK18, Ku18, KCK19, LG10, LCS18, LZPZ19, LWL14, LTSW16, LM18c, MKHC11, MM16, MRS^{+}12, MSW18, MP19d, MK17, PMA17, Por18, ST15, SPT17, SSH15, SMDI18, Sha10, SBM13, SSC19, zSdZ10, SLL12b, SS13, TM17, WJ11, WZKY12, WY15, WL10, WFDW10b, XZZ16, XHM14, Yan12b, YQWZ19, ZR18c, Zed10, ZBF11, Zha15a, ZY17a, Zha18d, ZFY^{+}19, ZZ10b, ZW11d, uRK11]. Coupling [BDPM12, CH19, JM16, JK18, KW14a, SC19a, AKSW19, BBL19, DPBL16, DTR19, ED12, FHS18, FH17, HLB14, JM15, MUB^{+}16, NEB14, RRAK19, SKCL19, SR17c, WC11a, WLA18, WCZ^{+}19, ZHY14, ZD18]. couplings [MZ10, Yu11b]. Courcelle [AK16]. Cournot [DH10a]. covalent [EHO^{+}12]. covalent-bond-driven [EHO^{+}12]. covering [CQ11, Ge10, LBZL11]. covert [CDFP12]. CPU [Bre14, WWZ12, YW14]. CPUs [SKG^{+}11]. crack [GB18, PK19]. cracked [TAA14]. cracks [AB18]. Cramer [SWC11]. Crank [AY12, BN14a, CY19, DM15, DH18, GH18, HA10, ZLJ^{+}18]. crash [MVB^{+}12]. crash-faults [MVB^{+}12]. CRE [WF17]. creating [OO10]. credentials [BMS12]. credit [Chu11c, Chu12b, HC18]. Creep [XJ18, XC13, YLC18]. criteria [AAZ10, Ant10, CYL17, DVY14, Dos12, FAHZ17, FZ17, FL10, GFZ16, Hon10, Hua10a, KÖ10, KC12, LLY10, LXK11, LS12c, LZG13, MZ11, RM17, SK11c, TZG10, YÇG12, YX11a, Yan19, Ye15, Ye16, ZFZ10b, ZZ16b, Zha18f, Zha18g]. Criterion [Kup10, FSZ17, GLR13, GGR15, LB18, Wen18a, WY18b, Wen18b, XZ17, Ye17b, Zha18e]. critical [ACD^{+}11, AJ12, CCP19, D’A18a, DSL18, FIW17, HL11c, JZ11, KLL19, LeT10, LSCG16, LY14, LL16a, LT18, Li18b, LS19b, LRZ18, LLZ18, LLT16b, MPLR18, QaY18, RC17b, SC16, SC19b, WS10a, WhJxLwW11, Wan15, WS17, WZ15, XW18, Ye17a, YW19, ZJ17, ZLG18, ZT16b, ZT18a, dSSV17]. criticism [MN12]. Crocco [YT18a]. cross [ABLS15, ASMM11, BGGCGRSP16, CW15a, CLM14, FMSV17, Gal12, GS15a, GV18, JWX14, LD11b, Li13, Li16a, Li17a, Li18a, LD19, Pan17, WY18a, Yan18a, ZY15b, Zha18b, ZBFC19, ZTW^{+}19, ZLZ18]. cross-diffusion [ABLS15, ASMM11, CW15a, CLM14, Gal12, GS15a, GV18, JWX14, Li13, Li16a, Li17a, Li18a, LD19, Pan17, Yan18a, ZY15b, Zha18b]. cross-diffusive [ZLZ18]. cross-kink [LD11b, ZBFC19]. cross-ply [ZTW^{+}19]. cross-sections [WY18a]. crowd [EAEH18]. Crowley [Li14, YXWL14]. Crowley-Martin [Li14]. CRS [CM17]. cruise [DNP15]. Cryptanalysis [Lee11b, SPLHCB14, You11]. cryptography [KK10c, PHWM10]. cryptosystem [Lee11b, UKAL10, WH10, yYqWqZC13, You11]. cryptosystems [WMZW11]. crystal [CYP16, DFW^{+}18, EGG16, GLR13, KMS19, KSF14, LZG13, MSW18, YWT18, ZDB19]. crystalline [SKH12]. crystallization [EF14]. crystals [GFZ16, MMRN12]. CSPs [FOX11]. CTL [YX16, ZLC11b]. CTL-response [ZLC11b]. Cu [ZAK18b]. cubature [Pap15]. cube [WWW14]. cubes [CKMR11]. Cubic [JLK11, CT17a, Che18, Gao17, GMI11, KB19, KG11, MD15, SYW11, VFM19]. cubic-order [KG11]. cubic-quintic-septimal [Che18]. cubical [DW18a]. cuboid [RKA^{+}18, WMP^{+}19]. cumulant [GSPK15]. cumulative [LCA^{+}17]. CuO [RKA^{+}18]. curl [SCGW18, Tom13]. current [AR10c, BGRV15, BD19, Cao19, CYP16, CLN^{+}19, NKM16, Pal13]. current-hole [NKM16]. curse [Pes13]. Curtis [Som13]. curvature [AM15, Ahn10, HWW13, JCF19, Ma19, TMZ^{+}15]. curvature-preserving [TMZ^{+}15]. Curve [PU10, Boy16, DMPV10, DRS11, KK10c]. curved [CS16, Ibr16, PLW^{+}18, Rua19, THC^{+}18, WY18a, Yan10c]. curves [Ahn10, AD11d, HH18b, JRSZ12, WW10b, Wu11b]. curvilinear [PC17, SYI12, ZY15a]. CUSP [HH17]. cuspidal [AZ10]. cutoff [GOT19]. Cuts [SLKK19, TMMASG10]. cutting [SR10a, WZXL11]. cycle [CL12a, Dol11, Ebr11, GCG12, GM18b, WN18]. cycle-separated [Dol11]. cycles [AZ10, BR12d, CD10, CSS10, DDD10, GOGYL^{+}11, SS11a, ZZHF12]. Cyclic [YS12, CM19a, EMR10, JL17b, JL17a]. cyclic-reduction-based [CM19a]. cylinder [ALHZMC^{+}19, CNH17, GS12, LWC13, SEM13, Tia19, WL13b]. cylinders [FAIV10, HMF16, UMLF13, ZJB19, ZLZG11]. cylindrical [ES18, GSZ11, LPML19, MC10b, MDL18, ZTW^{+}19].

D [KK19a, Pan17, SLKK19, YZ19, ATO19, ALHZMC^{+}19, AY18, AT17, BMRA10, BPC17, BOY12, BGF15, BP18, BGM19b, BC17, CYP16, CCHG17, CL12b, CLL19, CL15, CM19b, CS14b, DWZ13, DLZ19, DPM15, DN18b, DA18c, DFP^{+}13, DNZ^{+}13, EAEH18, Ers16, FZ17, FSTN18, FSZ18, FST19, FVVS16, FF15, FES^{+}19, FMPR15, GGLP15, GVSP12, Gha17, GR19a, GIMZ14, GDM13, HHS^{+}17, HHGA19, JCF19, KLK15, KN11, KYW^{+}18, KKVS19, KPG18, KV18, LZ18a, Li16b, LXZ18, LL19b, Li19b, LZG13, Liu18c, LB18, LRBA15, Ma18b, MM18b, MDG19, MJWD19, MKS13, MSZG17, OVV^{+}16, Oru17, PAE^{+}12, PLMS14, PDHL12, QCG15, RCM11, RCG15, SPT17, SJL^{+}19, SHH16, SEM13, SWL16, SAU11, Tod13, WFY17, WD16, Wan18, WZC^{+}19b, Wei12b, WY15, Wen18a, Wen18b, WH11b, WCH13, WFC16, XZ17, YC12, YSW16, Yan19, Ye16, Ye17b, YKKS10, YT13]. D [YWT18, YZS18, YQ18, ZN18, ZB19, Zbo19, ZZ16b, ZaY17, ZZLB18, ZZ18b, ZZG19, ZWH^{+}19, ZC11a, ZZL15, ZS18, dVLV18]. D-acoustics [BPC17]. D-based [Zbo19].

D-cylinder [ALHZMC^{+}19]. D-Helmholtz [GVSP12]. D-partial [KV18]. D-shell [MJWD19]. D-Var [CS14b]. D-vorticity [RCG15]. D2Q5 [EAMA19]. D2Q9 [EAMA19]. D3Q27 [SKTC19, SKTC15]. daily [CHT11]. dam [yXpYxZT11]. damage [BBL19, CFdM^{+}18, LKL^{+}15, WhJxLwW11]. Damped [CT10b, KP18, AJS19, DM10, FIW17, LL12a, Liu11a, MSV18, ML19b, MTAS17, MAST18, MB10b, RZ16, WL17a, WZ18a, YLG17]. damper [CJ12]. damping [Bag17, BG19, BTB18, BDS10, Bis10, CCRS17, Cve11, FL14, HKJ14, HH18a, HW19a, HEP10, HLY17b, KLP17, KPK18b, KYA15, LPK15, LLY10, LZY11, QZM17b, QM19, SR18, yZjM10, ZQ11b]. Darbo [SZ12a]. Darboux [LM19a, WW18a]. Darcy [RTL19, AS19, AGPR19, BO18a, HQ19, JJH16, LR14, LHL18, LWL14, Ngo18, OY19, SZ17]. Dark [HTY^{+}19, CDW11, QTW^{+}18, ZLW18]. Data [JPP12, LZL^{+}13, MBT^{+}13, SSK13, ST12, WL12b, ZM13, AY18, ACD^{+}11, ASB12, BWL18, Che15b, CP10, CH17, CS14b, DNZ^{+}13, FSZS18, FKC12, Fer12, FMGR19, FRSC16, HP13, Hot13, HLT12, HYL10, HH10b, IC12, JZL18, KBAF18, KLH^{+}12, KOPS13, LW19a, LZ19a, LCYC12, LGVS19, MHH11, ODAZ15, PMM17, PQBK17, RMB^{+}14, RKW12, SYO12, Tol12, THY^{+}10, VP11, WYY11, WYK10, WK13, WCW13, XW19a, YT18a, Ye15, ZKWW17, ZL13]. database [CCY10]. databases [Cha13]. Date [YTS^{+}17]. dates [LYLX11, WW10a]. datum [FIM18, HKP17]. Davey [HLTL17, KSG11, QRMH18, THfL17, WLXZ18]. DCT [CUK12]. DD [Jav11]. DDoS [LKLP12]. de-noising [BPM12, MSTB17]. DEA [WO10]. deadlock [TYY^{+}12]. dealing [YF10]. DeAngelis [LY11c, WWG10, YX16]. death [AGD10]. deaths [GCG12]. deblurring [DM19b, LX12a, Liu15b]. debonding [CFLM19]. debugging [CCDL10]. DEC [MSFS18]. Decay [GZN19, MAST18, RNQ13, RNQ16, CCRS17, DZ21, Dua18, FH16, HKJ14, HC16, HW19a, HGW18, JPK17, JLF17, KKLJ11, KPK18a, Kup14, NWZ11, Par15, Par18, yZjM10, ZW16a]. decaying [AD12, BR12d, KNIF13, WHZL17]. December [Ano18-64, Ano19-61, Ano19-62, VBW10a, VBW10b]. Decentralized [CMS10, CS11a, LWC13, LLC13, THY^{+}10]. decision [ÇE10a, ÇE10b, DH10a, FLLF10, KC12, LML11, LS10c, PdlF10, PCK13, SLCC12, TZ18, WL13a, XKH10]. decomposed [GK11a]. Decomposing [LMW10]. decomposition [Aba10a, AO10a, ASA16, AH11a, AER12, BOY12, BA16, BPG10, BR16, BM18c, CGH14, Che15b, CY19, CNP14, CJK17, DHQ11, DA18b, DR12, DCRL13, ESBR10, FSZS18, FRZ15, GM18a, GVSP12, HSWZ11, IQR16, LW19a, LLH14, LY10a, LYN11, LW18c, LZG19, fNS11, SK14b, YZ15, ZLC^{+}14, ZGZ13, ZSD10]. decomposition-discrete [DA18b]. deconvolution [Dun18, SM10, ZHJD13]. Decoupled [ZH15a, ZH15b, CR18b, LWZ16, PLKC16]. decoupling [JJH16, THY^{+}10]. deduced [Gal12]. deduction [BMS12]. deduction/abduction [BMS12]. deductive [JKK10]. deep [ZZC13a, ZZC13b]. deep-cavity [ZZC13a, ZZC13b]. default [HC18, LPP15]. defeating [Boy10]. defect [AZ17, ĆCM10, YH15, ZHJ14]. defect-correction [ZHJ14]. defective [KHF^{+}19]. Defense [JS12a]. deficient [EHO^{+}12]. defined [EAAED10, Haz11, IN10, MS11a, Mar11b, MG11, ODR10, ORD11, PB12, PA12, RS12a]. definite [BKL^{+}19, HM15, LW19a, WWB13, WLM13, YL10b]. Definition [LZYW13, hYxLL10]. deflated [ZG19]. deflection [FG18, KLTS11, YLY12]. deformable [FIMV18, KVR11, RTL19, TRL19]. Deformation [DVM12, ATH18, FGY^{+}17, KMS15, LHZ^{+}11, MJWD19, RZ17]. deformations [CCNT16, VĆV11]. Degasperis [JPB11, YT11]. degenerate [AAD17, CST14, CLS19, FIS18b, GJ19, KPK18b, LSW10, MP19a, SSS16, SNDK18, Wan16b, XWY11, ZW11b, ZZ10b]. Degradation [LTLL12]. Degree [Bog10, ABRL18, AD11d, BGH14, Che12b, DYX11, LZ10, LZYW13, LHH10, Pap15, Wu10, YMM12]. Degree-Bounded [Bog10]. degrees [BS17, Rus16]. DEIM [DA18b]. delay [AR09, AR10b, AAZ10, AK10, AdAS11, AM12c, AA13, BKR11, BP11b, BCF13, BB10b, BDS10, BB12, BDGBM11, BKDM13, CFN11, CANK11, CZL17, CM18c, DRK10, DB11, DQ10, EBENF10, EKS10, FH16, GGEB12, HK10, HW19a, HF10, JPK18, KLP17, KPK18a, KCL12, KK14c, LPK15, Li10d, LJ11, LB12, LCYC12, LLY10, LKU10, LWKK10, LZKU11, LZ15b, MP11a, MG16, MLY18, MM11, Ouy11, PR11, Pen11, SRS11, SYW11, SW19, TZ13, VB10b, WHS11, WS12, Wei10a, Wei12a, WL10, YG17, YX11b, YX11a, YLZ17, YZAX10, YW11a, ZGW11, ZQ11b, ZW11b, ZL11, ZPS^{+}12, ZLC11b, ZLW19, dAS18, dSAC11]. delay-dependent [LCYC12]. Delayed [Wei12b, Ahn12, Bai11a, CSW11a, HY16, HM17a, Hua11, LZT11, LZC13, LW15, Li18c, LLY18a, LRV13, MY16, SSAM11, TZG10, XZZ16, XZ18, XMW10, XGH17a, XGH17b, YX16, Zen11, ZZ16a, ZZF18, ZSW19, ZZW15]. delays [BKT12, BZK12, BKR11, CW10b, DW18b, DHGF17, HY15, HGW11, LC12a, LJJ11, LHW11, LXK11, LZ12c, MSA12, MM12, SI10, Tia19, WYN12, Xu11b, XH11a, YY10a, YW10]. delivery [FJP18]. Delta [Kuo16, RA12, MT11]. demand [BO10, BMM12a, DRK10, DH10c, JCZZ13, San11]. Dempster [KBDC12]. dendritic [SZP^{+}11]. dengue [ZLZ18]. denial [MKA^{+}10]. denoising [BSL11, GV18, JRB15, Liu15b, RSS16, SGZW18, SLZ11, WRW13, Yua18, ZY17c, ZFY^{+}19]. densities [MD10]. density [BO18b, CZY11, Che16, CYL17, FZ17, FST19, GHC15a, GOT19, KH13, LLML15, LX17, Özu15, QY17, RTRR18, SSHH^{+}18, ST15, SGK18, SKTD13, VMFF18, WZXS17, WX18b, XZ17, YZS18, ZZ18b]. density-dependent [CYL17, FZ17, FST19, QY17, WZXS17, XZ17, YZS18, ZZ18b]. dependable [OPDC12]. dependence [LX17, MS12a, WX18b, YLS12]. dependent [ABT19, AdAS11, ÁBÁPM11, ATH19, BKE18, BO18a, BQS16, Bis10, CHBTD14, Che16, CYL17, CXMO19, CT10b, CM14, DNS16, DRK10, FZ17, FSZ18, FST19, FSRB15, FZBF10, HG18a, HP13, Her14, HZLM10, HLI14, HLY17b, JLWX18, JLL19, KVW18, KYA15, Lee11c, LR14, LLZ11, LY11b, LL13, LLML15, Li17a, LZ19a, LCYC12, LFJ11, LZ12c, MWL18, Oan13, OY19, OSA13, Pen11, QZ16, QY17, RZZ19, SD15a, SLM18, SG16b, SG16a, SMH18, SZ14, TF17, TCHW19, THGG14, Tro13, VV14, WWW11a, WWW11b, WL11c, WQRZ14, Wan14, WD16, WHTZ16, WZXS17, XZ17, XHH^{+}19, YF10, YS19a, YZS18, ZJ10b, ZZ11a, ZZ18b, ZL19a, ZG18b, dAS18, dSAC11]. depending [KYAA10]. deposited [PvdM13]. deposition [ÖKJR19]. depreciable [ZZXY12]. depression [GGK18]. Depth [DJD18, Dem10a, Dem10b, GZW^{+}18, WFL11b]. Depth-averaged [DJD18]. depth-integrated [WFL11b]. Derivation [BBL19, EMR10, FES17, Jum10, Lóp19, OF16, PZL^{+}18, BPR18, Del13]. Derivative [WZ17b, AK18, BC10, CDW11, CJ18a, DB15, FKeT12, Fur13, GGR15, JRZK11, JLCS10, LEN10, LZZ11b, LDL^{+}15b, Liu18c, LLLW18, Lup11, MT10, MV17, Ned12, NA11, OC10, PC17, STC18, SHM13, WL11c, WLT13a, WCH18, YLC18, ZY10a, aZW17]. Derivative-free [WZ17b, Ned12]. Derivatives [MS11a, AVR17, AS11b, AT19, AR17, CCJP11, CSZK10, EZM12, FM11, Gal11e, GS11a, GZ10, HS11a, JKZ11, KCC^{+}13, Kli10, KY10, Li18d, MNJ^{+}13, NCV^{+}18, PAT12, Set12, SOK19, Tha19a, XJ18, Yan11d]. derive [CB10]. derived [FL13a, KB10c, zLYLQ19, RTV17, Sön11]. Deriving [HNPS13]. describe [ABL10, Laz10]. described [NB11]. describing [BPR18, De 10, HY16, OHMAK18]. description [JTC^{+}10b]. descriptor [Lin12]. descriptors [SID15]. Design [CFRS10, CLM12, Kuo11, LZY12, MK18, PC12, AM14b, BM12b, CI18, CZY13, Cha13, DSB19, EK13, ES10, HGJP19, HT12b, HWH^{+}15, HLCY12, KH13, LLWZ11, MPGW19, PM10, PBS12, SAR18, Váz16, Ver08, Ver12, WSL10, WKBR18, Yan12a, YWHC11]. Designing [FDXW11, MAPS10, ZPGW16]. desktops [Fia15]. destruction [AEG11]. desynchronization [HPR19]. detailed [SST19]. Detecting [ACD^{+}11, Liu18b, XLZW11]. Detection [Hua12, LKLP12, fLcJ10, CCSZ14, DWI^{+}12, ES11, EE10, GWZ11, KWPK13, KPL11, KYY12, LCW12, MBKK10, MHH11, Ögü13, RIW12, SLXC11, SCC^{+}12a, TXZ^{+}10, TT12, WLC11]. deteriorated [LZ18b]. deteriorating [Chu11c, DRK10, RR11, VCM11, YY10c]. deterioration [LLW11, MM11, WWW11a, WWW11b, YX11c]. determinant [CZ17, JL15a, JL17a, PW18]. determinants [JYL16, JL17b]. Determination [SG16b, XW19a, XXG10, CHL18, CSS10, HLI14, Ikh11, Mom11]. determine [HL11c, KLTS11, Pal12, SCvdV^{+}19]. determined [TNP17, ZL14a]. Determining [EUTS18, PA15, Mok11]. deterministic [Bra13, TKBMT17, VMC^{+}14]. developed [SPH10]. Development [DNP15, FJB19, SWL16, TLR17, FIS18b, LYL12]. developments [CC15, WZF12b]. deviating [Bic11, Can11d, Don10a, XY10]. deviation [ZZG19]. deviations [YLS12]. deviatoric [KKVS19]. device [MDW13, WXYW11]. devices [AdSSS19, CGS12, LNKU12, MM18b, SJS^{+}11]. DFT [JMNZ19, LL12d]. DFT-based [LL12d]. DG [FLdS14, GSZ14, Kim18, NDC^{+}19, VMC^{+}14]. DGLM [Kim18]. DGPMHSS [CW18]. diagonal [GSR14, HM17d, HMZ18, JKS12, Wu18b, Xue13]. diagonalizable [DD16]. diagonalization [CLMM18, Lam12]. diagonally [FHZ13]. diagrams [BF11]. dichotomies [ZFZ10a]. Dichotomy [Chu18, ZFZ10b]. dictionary [LYSZ19]. dielectric [Hei10a]. dielectricity [Zbo19]. dielectrics [Kup14]. Difference [BR13b, THGG14, AVZ15, Ara18, AG10b, AS10c, AKA11, AY12, AA13, BT11, BC12, BK11b, BR12d, BK13, CDM12, CLM12, CLM14, CTZ17, CV14, CG14, Cos18, DL19, DO11, DA18d, DÇ12, DFS11, DSR10, DT17, fDxZ11, Dos12, Dra10, FNW18, hGzS15, GCE18, GÇK10, GMB12, Goo11a, GL16, Gur13, GH10, GH12b, GH13b, HmZ11, He16, Hei10b, HG16, HS18, HTL10, Jan10, JM16, JW18b, JAJ18, KP13, KO11a, Kar10b, KÖ10, KPR13, KNZ19, KO11b, KB10c, LCM14, Li12a, LLML15, Li16b, LR17b, LX18, LWKK10, Liu11a, LXK11, LZKU11, LDL^{+}15a, LS17b, LLLW18, LLYL19, LR13, MX10, MZC17, Mil18, MFSL19, MMH11, MHM11, MM18d, MM12, Mur08, MKPS11, Pal12, PS16, PS12a, Par11a, Pit12, QYL10, RMM11, RS12b, RZ16, RGHZ15, RS18b, SD10b, Saj12, SMM19]. difference [SZ12a, Sed13, SDH13, SB14, Sou12, SK19, zSdZ10, SR18, SI17, TTT10, TUT11, WW11a, WWA11, WV16, WZM^{+}16, WH16, WH18, WS10b, WX18a, XGH17b, Xue13, YT18a, YZMZ16, YTZ17, YT18b, Yüz11, Yüz12b, ZCLW19, dVLV18]. difference-finite [YT18a]. difference/Adams [LLML15]. difference/finite [DA18d, LDL^{+}15a, LLLW18]. difference/spectral [ZCLW19]. Differences [MvS18, Abd11, Fer11, HL11a, OAY11, XW10]. differencing [GP11, LK18]. Different [GABC16, Her19, AE12a, BTEM19, CL19, CHK^{+}18, CP16a, DCN^{+}18, GD10a, GAVOF11, HSK11, HL18b, Ima17, KHF^{+}19, KM13, Kuo11, LG17, LJK^{+}19, RNQ16, SZC^{+}18, WL11d, WZHW13]. Differentiability [Smo17, TM12, ZFZQ10]. differentiable [Dra11a]. Differential [BC15, GGAVGG19, GGAVRC^{+}19, KSJ12, QGGL13, TS11a, Zho11, Zho12, ZFLM19, AB10a, Abb11, AZH10, AK10, AA11, AdAS11, ANP11, AW11, AN12, AuIK17, AK18, AJY13, AMA14, AÇT11, ArEM10, AJ10, AJ11, AJ12, Alo11, ABL10, AKRT14, ASMEE11b, AM12c, Arq18, AS10c, AuIA17, BD11a, BD11b, BM10a, BMA11, BH10, BP11b, BMTV12, BCC14, BB10a, BHM12, BCK11, BDS10, BB12, BE11, BKY10, Bis14, BY11, BCF^{+}14, BOT14, BH11, BK12b, BK10, CCJ10, CCJV11, CP15b, Cha11a, CANK11, CZN11, CZN12, CW10b, CZ11a, CM18c, CJ18a, CY19, CS14a, CJCV10, CJRR11, CMT12, DE11, DCL17, DN10, DB11, DQ10, DMD10, DRD12, DPZ13, DMZ10, fDxZ12, DHY19, Don10a, DM19b, DW18b, DCRL13, EBENF10, ED11a, Els10, EOM11, ESL11, FID14, FT10, FSHZ11, FN14, FOX11, FNW18, GGAV18]. differential [GY15, GYTD12, hGzS15, GLLC19, GK13, Gen10, Gep16, GR13a, GTL16, GY11, GS11a, GHR10, Goo11b, GR13b, GJ12, GW12c, GMZ15, Gup11, GGO16, HG18a, HY16, HO10, HB12b, HLvS18, HY11, HF10, HLZD11, HLY12b, HT16b, IB11, Ibr11, JYF^{+}11, JNBK13, JL11a, JL11b, JW18b, JS11, Jia12, JD12, Kat11, Kaw15, KRCJ11, KY11, KK14c, KV18, KMT10, KM11, Lam12, LeT10, Lee15, LC11a, LLZ10, Li10e, LF11c, LJ11, LZC11, LS11b, LY11a, LB12, Li18d, LZ11e, LH10a, LLY10, LKU10, LJ10b, Liu11a, LHF11, LXP11, LZY11, LCZ11, LS12c, LS12a, LS12b, LJX12, Liu12, LLLC14, Liu16d, Lup11, LLX11, MP10a, MPZ11, MBH11, Mar11b, MLZ^{+}16, MP11c, MS12a, MD15, MI16, MM18d, MS15, MN10b, MGW11, MT18, NLA19, Naw11, NB11, Odi10, OBAAD10, ODR10, ORD11, Oua12, Ouy11, ÖZ11, PC17, PIAH10, Pen11]. differential [QX19, RSM17, RM17, RS12b, RSDR11, RS15, RÖ10, SD10a, SD10b, SRM11a, SCSF19, SBEB10, SAIZ15, SKTH11, SZDO10, SS14a, SXM11, SW12, SGZW18, SYW11, Su12, SLL12a, SLL12b, TN11, TZWM10, TAS11, TNF11, TB10, TMDTTC16, Tom11, TMSO12, TNHK19, VA10, VA11, VBCJ10, WC10b, WY11a, WS11a, Wan11, WG11, WAZ11a, WZF12a, WZF12b, Wei10a, WAW15, cW11, WZ11b, XD10, XY10, XH11b, Yak11, YÇG12, YXS10, YX11b, YX11a, Yan12b, Yan17, YZY10, YW11b, YSS11a, Yüz11, YSS11b, Yüz12b, Yüz12c, ZA10, ZFZQ10, ZJ10b, ZYG10, Zha10, ZW11a, Zha11c, ZS11b, ZTH11, ZD11, ZW11b, ZL11, ZY17c, Zha17b, ZSHL11, ZCH12, ZLY^{+}13, ZZL18a, ZL10b, ZFLM18, ZLGL11, ZMA10, dSAC11, uRK11]. differential-algebraic [IB11, SBEB10]. differential-difference [RS12b, Yüz11]. Differentiating [FCZ12]. differentiation [AJY13, BMC13, CBM10, GS15a, HCL12a, SK14b, TXL19]. Differently [TL10a]. differintegral [PA12]. diffraction [LHL12a, WZ18b, XD17]. Diffuse [KY10]. diffused [ST15]. Diffusion [FRAK15, SOK19, VPR11, AM13a, ADZ19, AD19a, AM10a, Akm15, AJRWS12, AJAR18, AM14a, AKL18a, AO18, AK19, AAA12, AJT19, Amo18, ABLS15, ADS14, AB16, Ara18, AV14, ABN18, Auc18, ASMM11, Bac14a, BDM^{+}19b, BP18, BL17, BK16, BS14b, BSY^{+}19, CW15a, CGGM19, CS16, CM18a, CM15, CST14, CEQ14, CHBTD14, CLM14, CSZK10, CYS10, Che16, CLC16, CJ18b, CL17b, CJK17, CM14, CG14, CFLX18, CRRS11, CCM14, DMP18, DG13a, DPBL16, DWZ13, DZW16, DNS16, DN18b, DA18d, fDxZ11, DS18b, DMV11, DH16, DHGF17, DJZ18, FJC16, FP19, FM18, FZL^{+}18, FMSV17, FIM18, GD10a, Gal12, GS15a, GV18, GZR^{+}13, Gao15, Gao17, GJ19, GGVRB19, GH14, GSZ14, GS11b, GM14a, GM14c, GL16, GML17b, GLW18, HMV18, HHY13, HXX19, HY13, HY15, HO19, HPR19, HHNLGC18, Her14, HM17a]. diffusion [HLSN16, IW18, JGSS10, JA11, JRB15, JKN10, JMB10, JJ19, JWX14, JW19a, JLTB12, JAJ18, KBAF18, KW14a, KW14b, KP19c, KAA19, KCK19, KBK19, KGM11, LGHR16, LCM14, LHD18, LN98, Li10c, LDW11, Li13, LZC13, LCS15, Li16a, LLFT17, Li17a, LLY18b, LXZ18, Li18c, Li18a, LD19, LCA^{+}17, LN19, LHL15, LDL^{+}15a, LZ15b, LDL^{+}15b, Liu17, LZCL18, LLLW18, LZ18c, LWSL19, LW19b, LZG19, LHTL19, Luc10, LMP13, LGVS19, MZL13, MGTH16, MY16, MW17, MF18b, MF18a, Ma18a, MLY18, MBS17, MD18, Mah14, MA17, MHHC18, MCL^{+}13, MSTB17, MPY16, MZB10, MAS11, MV17, MP19c, MD15, Mop11, MN11b, MY13, MAH18, Mur08, Nab19, Nes10, OSA13, Özu15, PS16, Pan17, PH19, Pov12a, Pov19, QZF19, QLT^{+}18, RSS16, RKF18, RS14b, RZZ19, SD11a, SD15a, SC19a, SCA14, STC18, SHM13, SZZ11, She16, SG11b, SLW14]. diffusion [SZL^{+}17, SY18, SS16c, Sou11, Sou12, SLZ11, SFM15, SND19, SI17, Tam16, TC16, TM18, TTM19, TF17, Tim13, THD19, TKBMT17, TKHL18, VAS^{+}18, WCB13, WFY17, WW11c, WL12b, WRW13, WV14, WL15, WW16, WZ17a, WZXS17, Wan18, WW18b, WLZ^{+}18a, WLL^{+}18, WHS18, WRY18, WYL19, WZC^{+}19a, WZ18c, WZY13, WCLD18, XFL16, XA13, XW19a, XL15, XZZ16, XZ18, XHA13, XX19, YY14, YY15, YW10, Yan18a, YZX18, YCHW18, YP10, YXWL14, YLL^{+}14, YKA18, YDW15, YSB15, YSX^{+}19, ZN18, ZLY14, ZTZ15, ZY15b, ZWZ16, ZTZ16a, ZZWG16, ZLJ^{+}18, Zha18b, ZSL19, ZSW19, ZLA17, ZZL^{+}18b, ZcHS18, ZWH^{+}19, ZLZ10, ZJZ18, ZWJ^{+}11, ZG16, ZZW15, ZZ16c, ZZ17, ZLW19]. diffusion-chemotaxis [MGTH16]. diffusion-convection [RS14b]. diffusion-convection-reaction [Akm15]. diffusion-reaction [AO18, BDM^{+}19b, CCM14, KW14b, LCM14, LZ15b, QZF19]. diffusion-wave [BL17, CLC16, DA18d, GS11b, GML17b, JGSS10, JA11, JLTB12, LMP13, Pov12a, SC19a, WZ18c, XW19a, YCHW18, ZWH^{+}19]. diffusion-wave/diffusion [JLTB12]. diffusions [Lee14, LMS13, MN17]. diffusive [ABV11, CWW19, CB19a, CP16a, CSW11a, DM10, HHNLGC18, Li14, LWN15, LX17, LLY18a, LPY16, MG16, MCB10, MW16, Pov19, RRP16, SZ14, SW19, TZ13, WW19a, Xu14, XGH17a, XGH17b, YL13, YX16, YZMZ16, YLZ17, YZ18, ZL14a, ZHW14, ZZ16a, ZZF18, Zho13, ZLZ18, ZG18b]. diffusivity [WHZL17]. digital [LWZG10, LTL^{+}12]. digraphs [LH10b, XCXW10]. Dimension [KV17a, Bac14a, BS12b, IBB10, LCW17, MBS17, WZ18a, YKRV11, ZY11]. dimensional [AM13a, ADZ19, AGT19, Abd18b, AK16, Ade16, AHHM19, AuIK17, AA15, AAEG17, AJRWS12, Ali15, AdSSS19, AK19, Asl11, AuIA17, BMRA10, BHZJ19, BM18a, BBBM16, BZ18, BQS16, BGP13, BL17, BB08, BSY^{+}19, BPF13, CNV14, CTSX16, CB19a, CWH13, CLB14, Che14, CCCW16, CS18, CZ15, CR13, CL17b, CZ18, DYH11, DM18, Def10a, Def10b, DAM14, DM15, DA16, DD10, DLS14, DSVS15, DDM^{+}18, DFW^{+}18, DH18, DTYZ18, Dos12, Dua18, DHMU16, EGAA19, FR15, FYYT11, FLZ14a, FDG^{+}17, Fio14, GGR15, GQF^{+}10, Gao12, hGzS15, GH14, Gha17, Gha18, GZ14, GHCZ18, GZD18, GABC16, GMZ15, GL16, GML17b, GTZ19, GC19a, HLLM19, HG18a, HGHA19, HTM18, Her19, HA16a, HSZ15, HBE15, HVO17, Hu18, HTY^{+}19, HA10, HTWS15, HYC18, Ima17, JL18, JK12, KG14, KNT12, KSG11, KK14a, KCL14b, KPG18, KKK15, KT18a]. dimensional [KK19b, KK19c, Kuo18, Kup14, LCM14, LHD18, LK15, LK13, Lep11b, LMW10, LD11b, LDL11, LZH16, LS16, LR17a, LMMF17, LCS18, LM18a, LZL19, LMY19, LM19a, LCW19, LGZ19, LSC17, LL10a, LCLL16, LTJ^{+}16, LL16d, LWZ16, Liu18a, zLYmL18, LZM18, LYC^{+}19, LWW19b, zLZ19, LWW19a, LFAL19, LLYL19, LXY19, LGVS19, LDHH13, MP10a, MP11a, MPZ11, MLL16, MYZ18, Ma18a, MCF18, ML19a, MR17, MMRN12, MKR12, Man18, MZM18, MGS^{+}14, MPfTX18, Men18, MS15, MM19, MK17, MSFS18, MMA12, MM10b, Oru19, Osm18, Pas14, PTZ19, PD11, Por18, QTW^{+}18, QMW18, RKA^{+}18, Ray17, RS18a, Ray18, RG18, RMY19, RNQ13, SSHH^{+}18, SK14a, SH18, SR15, SR17a, Saj12, SMF17, Sat11, Sea16, gShYL10, SS14b, SZL^{+}17, SY18, SMBY10, SPST18, uIAA15, SG10b, SZC^{+}18, SK14b, TDXQ18, TZ15, TS11a, TCHW19, TT14, TTX^{+}16, UMY11, UKI11, Üns18, VC12]. dimensional [VFM19, VMFF18, WZH10, WV14, WWW14, WPL16, WV16, WW16, WY16, WHS17, WTYZ17, WWD18, WCH18, WHS18, WY18a, WMP^{+}19, WZC^{+}19a, Was13, WLZ18b, WZMY18, WCLD18, XX10, XW14, XZ18, XZ11, XX19, YMM12, YTD^{+}18, YGS^{+}16, YLZ17, YGS17, YJ19, YCLY15, Yil19, YMLL18, YYLW19, YTS^{+}17, YTL^{+}18, YHC18, YSX^{+}19, YK17, Zak18a, ZSAN19, ZA15, ZZ11b, ZCSG13, ZZX^{+}14a, ZDLC14, ZTC14, ZSW15, ZC16, ZHC17, ZYSY17, ZM17a, ZDZY17, ZLJ^{+}18, ZLT18, Zha18f, Zha18d, ZDM11, ZYZ11, ZLA17, ZTSC16, ZW16a, ZMH16, ZJZ18, ZSD10, ZDF^{+}14, ZT18b]. Dimensionality [HKS19b, Pes13]. dimensionally [FW18]. dimensions [BZT16, Cal19, CLA19, CX18, CDS15, FGB19, GGAV18, GSPK15, HP17, KO11b, KLP10, LP12, LNW19, NWZ11, Ped18, Pes13, RZ16, RRO17, RC17b, Rus16, SBM13, WWLL13, WH16, XG10, YK18, YM17, YQWZ19, YL18b, ZCLW19, ZHW^{+}11]. Diophantine [Kes10, YH11b]. dipole [RCRV14]. Dirac [FRZ15, OVV^{+}16]. Dirac-generated [OVV^{+}16]. Direct [PPC15, TTT10, WH11b, YLL^{+}14, BCD^{+}16, Che12c, Cos18, DB12, ES18, Fia15, GH12a, HNK13, KP18, LCP10, PPC13, Ped18, PAT13, WKBR18, ZWZ16]. direct-forcing [BCD^{+}16]. directed [SLZ11]. Direction [SHM13, VFM19, AD15, ADZ19, BXKZ11, BX14, CLA19, EKZ17, hGzS15, Haj18b, Haj18c, KKC^{+}10, SLM16, TM17, WH16, WLZ18b]. Direction-aware [VFM19]. direction-splitting [SLM16]. directional [Bis14, CB19a, GGR15, Liu18c, PCS13]. directions [KOPS13]. Dirichlet [ABB17, CLJ11, CH17, DGB10b, DM12b, DMV11, GG18, GR15, HP13, HT13, Ibr16, KN11, KSF14, LJSK13, LWBW13, LCZ11, ORR16, Sal10, SPST18, Tod18, WZM^{+}16, WCH18, WHG11, ZJ10b, ZZ15b]. Dirichlet-to-Neumann [KSF14]. Dirichlet-type [Sal10]. disc [Amb12, AM13c, LSM11]. discharge [MC10b, RZL11]. disconjugacy [MZ11]. discontinuities [HM10]. Discontinuity [BC11, BSS18, HKP17, ZLW10]. Discontinuous [BDF16, CDL17, CN16b, FHS18, HT16a, Moo18, PGQ16, RRAK19, TCM15, AvB16, AEH18, AMGC19, AYH17, Bac14b, BR12b, BQS16, BKZ17, CCN14, CGGM19, CW17, CWHW17, CD10, Che19b, CFdM^{+}18, CHH14, DFGG13, EDC14, FRAK15, GM19, HKS14, Hof18, Izs15, JMHF13, JRA^{+}18, KWAS16, KCL14a, KW14a, KCL14b, KW14b, Kim15, KCL16, KPS17, KS15a, KRBS18, LPLR19, LSS17, LXZ18, LQMW18, LMLB19, LZLL18, MNPD15, MLG17, NCC13, OBCG19, Pet15, QY13, RAW^{+}16, RTT17, Rob14, SRGL13, SL18a, SND19, TAPA^{+}17, UMLF13, WZKY12, WSM^{+}19, Yil19, YSB15, ZSY14, ZP18a, ZZL15, ZD15]. discount [KGJ11]. discovery [ZHQG12, YWL^{+}11b]. Discrete [Bab14, BCPS15, EAEH18, LGL^{+}14, PAT13, PLKCC12, WSCW16, AD12, AE12b, BP19, BKR11, BM10b, BMS19, Bis14, BN14b, BS18b, CTP10, CGH14, CEF^{+}13, CC11, CS18, CH17, CC17, Chu18, CCP19, DA18b, DSVS15, FSRB15, FSB17, GGH18, Goo10, HHY13, HmZ11, HBE15, Hol11, HL18b, HHM12, KK15, Krn12, LMPE18, LP10, LX16, LGC^{+}17, LZJ12, MZ10, MM10a, Ma19, Mah14, MCP13, MT12, MSA12, MAS11, MZ11, yN11, OSZP13, Pal12, PSP10, Par11b, PC11b, PCM12, PLKCC13, RKA^{+}18, Rah11a, Rah11b, Rey12, Sal16, SS11b, SOJC10, TZG10, Wan13b, WLA18, WN18, WZKY12, ZHZ14, ZZL^{+}18b, ZSD10, ZY13]. discrete-time [MSA12, yN11, PCM12, ZY13]. discretisation [AMGC19, BS14b, Von19]. discretisations [SBvdV13]. discretization [Cia12, DH11b, DFM15, FIW13, HB19, JHW15, KO11b, KPG18, KK14c, MR19, MS15, MT18, RCH19, Tam16, Tar17, WBZY18, XZZ16, YYYH19, ZN18, ZH15a]. Discretizations [ÖKJR19, AVZ15, Cal19, DM15, EM14, FF15, JJ19, JK18, KW14a, LY13, MW13, RGHZ15, TAPA^{+}17, TW18, YH15, ZLS13]. discretized [BO18a, GM14b, HKS19a, JW19a]. discriminant [LLSW10, SLYY13]. Discussion [WAW15]. disease [ASB12, Li11b, MW16, SG11a, ZHW14]. Diseases [FdOdSS17, ABLS15, HY16, LZ11b]. disjunctive [GXZ11, XGXZ10]. disk [Ikh11, KM13, LYZ11]. disks [KDG11, Mah10a, Mah10b]. disorder [Iom18]. dispatch [BGRS11]. dispersal [BL18, DD13, ZLY17]. dispersion [BDPM12, Bis10, CTZ17, CV14, ESBR10, GJX18, HLY16, JFS14, JFS20, LZB12, MC10b, SIL19, YYK16, YZ15, ZLPM13]. dispersive [CJK18, CWH13, HS12, KB10b, MDL18, NZ16, RMY19, UKI11, WWD18]. displacement [HCZ16, LR15, LCHZ19, yXpYxZT11]. dissipation [CSCM13, DD19b, KKLJ11, Ma18a, NPR10, NDC^{+}19, QY17, SIL19, Ye16, YWT18, YQ18, YZ19]. dissipative [D’A18a, Fer12, LSS17, SM17, yZjM10]. dissolution [CM15]. dissolution-diffusion [CM15]. distance [AJ12, CSW11b, SCK11, yYsZyYL13]. distorted [ZSY19]. distortion [XL10]. distress [ZLY12]. Distributed [FM12b, HB12a, AF13, ALLH11, ALLQ13, BZK12, BKR11, BB10b, BB12, BGR14, Can11d, CLT^{+}13, CL19, CCRS17, FWFL11, hGzS15, GNP14, JW18b, JCWZ16, LLFT17, LLY18b, LLY10, LLH10, MKA^{+}10, MM19, OPDC12, STC18, SLMZ12, WCZ13, Xu11b, YZAX10, ZLJ^{+}18, ZL19a, ZGL14]. distributed-order [hGzS15, JW18b, LLFT17, ZLJ^{+}18, ZL19a]. distribution [ABT19, BY11, CDG15, HSBL11, HW11, HLT12, JL12, LL12d, LXYT11, MDVM17, MS11a, PFBL10, RMK19, SID15, VBK13, WLHZ14, WD12, YW14, ZGZ13]. distributional [LSJ12]. Distributions [CG15, FF15, Lam13, OCNG12, RKW12, VDV13]. disturbance [LZY12]. disystems [BFSO12]. div [CCKP15, DFG19, Tom13]. Divergence [Kia19, DFG19, FDG^{+}17, Tre18, YY12, Zha18c]. divergence-free [Zha18c]. divergent [Ehr18]. diverse [DCN^{+}18]. Diversity [MYZ18, YMLL18]. divide [FNW18, LLCG16]. divide-and-conquer [FNW18, LLCG16]. divided [XW10]. divisibility [ZP10]. divisor [VBK13]. DNA [MCQ11, YS12]. does [RNB11]. DOF [BJLZ12]. DoGIP [Von19]. Domain [CDW11, CNP14, CJK17, HMSC10, LW19a, BA16, BCJ19, BM18c, BPZ19, BO18b, BSY^{+}19, CLT^{+}13, CFRS10, CL17a, CCNT16, CY19, De 10, DSWB18, Deu10, EAEH18, EAAS18, FIVM17, FSTN18, FST19, FIW17, GM18a, GGGR17, HHS^{+}17, IQR16, Ibr16, JLTB12, KSD^{+}19, KB10c, LLH14, LYJ15, LSM11, Liu16a, LW18c, LW18b, LZS12, LLL16, MLG17, MPMTV15, NM11b, OY19, PLR15, Sam19, SCKH10, SP12, Sön11, WJ11, WZ18c, WH11b, YWK^{+}10, YZ15, YLG17, ZHY14, ZWH^{+}19]. domains [AD19a, BT15, BBBM16, Chu10, DG13a, Das15, DM12b, DC15, DHMU16, FID14, FMS19, FM18, Fis18a, GVSP12, GH14, GH15, GHCZ18, Hal13, HY18, HKK^{+}16, JW18b, JLL18, LL12b, LJSK13, LFAL19, LRCG16, MG11, MLG17, MSFS18, NNAS11a, NM11c, Oh15, ODR10, QLT^{+}18, Rua19, SDH13, TS14, TH13, THGG14, WL17a, XL11, Yan10c, YL14, YL18b, YKKS10]. dome [PBS12]. dominated [CKSL^{+}14, CHBTD14, Oru19, ySW10, SLW14, ZLL17]. domination [CMMP11, CKMR11, CLM11, SD12b]. Double [ABV11, CNH17, EKS10, Liu18a, MCB10, Von19, AD10b, AZ10, ÇA10a, CXZ15b, CH11b, mCfX10, Ciz12, DH17, FZ19, FSRB15, FSB17, GOGYL^{+}11, GMZ15, KZ16, LFZ19a, Liu15a, MGS^{+}14, MME10, OAKR16, RTL19, RRP16, TD10a]. Double-delay [EKS10]. Double-diffusive [ABV11, MCB10, RRP16]. Double-grid [Von19]. Double-periodic [Liu18a]. double-population [OAKR16]. doubling [HM15]. Doubly [MP19d, LSW10, LW18b, PLW^{+}18]. doubly-connected [LW18b]. doubly-curved [PLW^{+}18]. down [LZL16, LLSW10]. down-and-out [LZL16]. downdating [ZLC^{+}14]. DP [AD11d, KLP10, KCL14b]. DPG [BGH14, CDG16, CHBTD14, DG13b, DGH17, FH17, Füh18, HKS14, HK15, NPD17, PD17, RBTD14, RC17a, CGHW14]. DQ [DA18b]. DQEM [TAA14]. DQM [Gha17]. drag [CT10b, CLL11]. drainage [NNL13]. drains [MNJ^{+}13]. Drazin [JW18a]. DRBEM [STS19]. drift [BK12b, CUK12, FP19, SS16a]. drift-diffusion [FP19]. drilling [CL12b, MTV13]. drive [LC12b]. driven [BK18, BH11, BM19, CCKY12, CM12b, DZS10, EHO^{+}12, FQLC18, JCWZ16, MGTH16, MP11c, MT19b, PD11, PSD^{+}13, QWGJ15, SK14a, Sah11, SSPL10, VAK^{+}19, XSYL19, XaZH19, YLC12, ZZC13a, ZZC13b, aZW17]. driver [KPL11]. dromion [KSG11, RKSA18]. dromion-like [KSG11]. Drone [RKSA18]. drop [BDGS13, BPS18, GK19, MJ14, TTT10]. droplet [DLF^{+}11, Mom11, ST18, TY13, YCS19]. droplets [MSW18]. Dropping [WhJxLwW11]. drops [BKL14]. Drude [LSS17, SWL16]. drug [EE18, FJP18]. drum [KH13]. dry [PLMS14]. DSI [KAK^{+}12]. DSP [WHC12]. Dual [CPP15, CN16a, TL18, BCFQ19, BB15, CEQ14, CKLL10, DGA18, DS18a, Dub13, Hof18, LP12, LZCL18, SC19a, XJYL17, ZWL11, ZLC^{+}11a, RKA^{+}18]. dual-duct [ZWL11]. Dual-mixed [CN16a]. Dual-MRT [RKA^{+}18]. dual-phase-lag [BCFQ19, LZCL18]. dual-primal [Hof18]. dual-scale [ZLC^{+}11a]. duality [AH10b, Ant14, GKS10, GJ10, PN10, ZLLF12]. Dubrovsky [KKK16, WZMY18]. duct [FBTS19, SB14, ZWL11]. ductile [Yoo17]. ducts [STS19]. due [ANN10, DVMS13, FZBF10, HSK11, LYLX11, LZZ11a, WW10a, WCH13, ZWG11]. Duffing [ASY^{+}11, Gen11a, RKP12, YASK10]. Duhem [LL12c]. durability [SJS^{+}10]. during [GCG12, RIW12, ZOZZ12]. dust [EK16]. dustiness [RB19]. dusty [ES18, EEBM10, MC10a]. Dynamic [ATZ11, DEFP11, GZD18, KBGC12, yN11, RSP18, SJS^{+}11, YCS19, ZZHF12, ZTZJ14, AR09, AR10b, AB18, AAZ10, Bab14, BDS17, BMM12a, BMS13, BMSS18, BFF^{+}11, Can11d, CP15a, DFS11, DLT12, FH16, FL10, GGEB12, GLM^{+}11, GA10, HDS11, HEP10, Hon10, Hua10a, JZR15, KSPP11, KJA10, Kia18, KAK^{+}12, LZ11c, LL12a, LCW10, LCP10, Li11a, MP11a, MAPS10, MH11, MR10, MS17, MG15, OPDC12, PR11, PH13, PCK13, PZAR19, RFP11, RF12, SD12a, Sca11, SA16, SSK13, TZMZ12, VAS^{+}18, WSL10, WSG10, WZ11a, Wan12, WLXG18, XC16, ZFZ10a, ZHJ11, ZGW11, ZQ11b, ZHQG12, ZLW18, ZHJZ11]. Dynamical [ABM11, DS18a, NJV13, TM18, XZL^{+}11, ZK16, Alo11, ALMLM14, BKT12, BZK12, LX10a, LZT11, LB11, Ma10a, SIL19, SS18b, SLYY13, WL10, YZAX10]. dynamically [SW16b]. Dynamics [CLS19, Che16, IL13, KHWK10, Sah17, SW19, SMYK19, Tao18, WZXS17, WFC16, Xu14, YL13, YZ18, ZL14a, ZZ16a, ZWZ16, ZG18b, ABT19, ABL15, ABL10, AB16, BMS19, BPKM10, CM13b, CL16a, CM19b, DGTC13, DFJS10, ECY11, FDB13, GM14c, GPV11, HPR19, KW12, LZ11b, LDS10, LJK^{+}19, LMPG13, MLY18, MCQ11, Mag10, MB17, NNR14, Özu15, PC14, PS12a, PQB^{+}16, PV12, RKSA18, RCG15, SP12, SZ14, TTM19, TY13, Tre18, TTX^{+}16, VV14, VJM15, WHS11, WSM^{+}19, Xu11b, XGH17a, YX16, YZMZ16, YGR11, YKKS10, YWL17, ZLMZ18, Zha18h].

e-services [KC12]. E-UBIAS [Ogi12]. E.A [AIB10, MAK12]. eagle [GYTD12]. earliness [LYLX11, MMR10, MMR11]. earliness-tardiness [LYLX11]. early [Amo15]. earthquake [LMPG13]. earthquakes [LKL^{+}15]. easy [CB10]. EBM [GSI19]. eccentric [DT11]. eccentricity [Ili12]. ECG [JWX^{+}13]. Echocardiographic [RMB^{+}14]. echocardiography [AY18]. eco [ZSW19]. eco-epidemiological [ZSW19]. ecological [JLWX18, YZAX10]. ecological-mathematical [JLWX18]. ecology [WLW^{+}11]. economic [BGRS11, Chu11b, DRK11]. economy [PG10, ZBL12]. eddies [BK18, MS17]. eddy [ADD^{+}15, BGRV15, BD19, Cao19, CRG16, CLN^{+}19, DSWB18, DNR13, HBK^{+}19, Sag10, SKFG11]. eddy-resolving [DSWB18]. edema [RdSRL19]. Edge [Ryl15a, BDHR18, Cia16, CFS17, EG18, FJWW16, LW11c, WJWW12, YA11, ZLS13]. edge-weighted [FJWW16, WJWW12]. edges [MSTB17, YA11]. Editorial [BBH^{+}19, Ano10b, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano10j, Ano10k, Ano10l, Ano10m, Ano10n, Ano10o, Ano10p, Ano10q, Ano10r, Ano10s, Ano10t, Ano10u, Ano10v, Ano10w, Ano10x, Ano10y, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano11g, Ano11h, Ano11i, Ano11j, Ano11k, Ano11l, Ano11m, Ano11n, Ano11o, Ano11p, Ano11q, Ano11r, Ano11s, Ano11t, Ano11u, Ano11v, Ano11w, Ano11x, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano12k, Ano12l, Ano12m, Ano12n, Ano12o, Ano12p, Ano12q, Ano12r, Ano12s, Ano12t, Ano12u, Ano12v, Ano12w, Ano12x, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano13g, Ano13h, Ano13i, Ano13j, Ano13k, Ano13l, Ano13m, Ano13n, Ano13o]. Editorial [Ano13p, Ano13q, Ano13r, Ano13s, Ano13t, Ano13u, Ano13v, Ano13w, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano14l, Ano14m, Ano14n, Ano14o, Ano14p, Ano14q, Ano14r, Ano14s, Ano14t, Ano14u, Ano14v, Ano14w, Ano14x, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano15m, Ano15n, Ano15o, Ano15p, Ano15q, Ano15r, Ano15s, Ano15t, Ano15u, Ano15v, Ano15w, Ano15x, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano16k, Ano16l, Ano16m, Ano16n, Ano16o, Ano16p, Ano16q, Ano16r, Ano16s, Ano16t, Ano16u, Ano16v, Ano16w, Ano16x, Ano16y, Ano17a, Ano17b, Ano17c, Ano17d, Ano17e, Ano17f, Ano17g, Ano17h]. Editorial [Ano17i, Ano17j, Ano17k, Ano17l, Ano17m, Ano17n, Ano17o, Ano17p, Ano17q, Ano17r, Ano18a, Ano18-27, Ano18-28, Ano18-29, Ano18-30, Ano18-31, Ano18-32, Ano18-33, Ano18-34, Ano18-35, Ano18-36, Ano18-37, Ano18-38, Ano18-39, Ano18-40, Ano18-41, Ano18-42, Ano18-43, Ano18-44, Ano18-45, Ano18-46, Ano18-47, Ano18-48, Ano18-49, Ano18-50, Ano18-51, Ano18-52, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano18j, Ano18k, Ano18l, Ano18m, Ano18n, Ano18o, Ano18p, Ano18q, Ano18r, Ano18s, Ano18t, Ano18u, Ano18v, Ano18w, Ano18x, Ano18y, Ano18z, Ano19a, Ano19-27, Ano19-28, Ano19-29, Ano19-30, Ano19-31, Ano19-32, Ano19-33, Ano19-34, Ano19-35, Ano19-36, Ano19-37, Ano19-38, Ano19-39, Ano19-40, Ano19-41, Ano19-42, Ano19-43, Ano19-44, Ano19-45, Ano19-46, Ano19-47, Ano19-48, Ano19b]. Editorial [Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano19j, Ano19k, Ano19l, Ano19m, Ano19n, Ano19o, Ano19p, Ano19q, Ano19r, Ano19s, Ano19t, Ano19u, Ano19v, Ano19w, Ano19x, Ano19y, Ano19z]. Effect [AD19b, BMJ10, BSK11, GM14c, Her14, OSA13, PHM^{+}19, SEM13, SSPL10, Yan18a, AJS14, COR18, Def10a, Def10b, EUTS18, GOGYL^{+}11, Lee11c, LL13, LZZ11a, MGTH16, Mac12a, MTV13, PQB^{+}16, PFBL10, RR11, SRRP18, TBP19, WSS10, WH11a, WSC16, YG17, YYLW19, ZFC11]. Effective [sCYhX18, Gos10, MWY17, BCSCB^{+}15, JM15, LFZ19a, LZS12]. Effects [BKL14, GSD^{+}19, LZP^{+}19, Sah11, SR10b, TS11b, Amo18, ASMM11, CP15a, CD12, Das12, FKKS11, Far11, GM18b, HLB14, HHM12, KWFY11, Kia19, KM13, LSW16, Lin14, MCB10, Pal13, Pan17, Ryl15a, SNSK19, UABK16, WW10a, WWW11a, WWW11b, WL15, Xu14, YY10c, YX11c, ZY13]. efficacy [IS14]. efficiency [ABK^{+}13, Cer18, EGGS^{+}12, ILP14, McN12, MP12, SKCL19, WL11d]. Efficient [BN14b, CDP16, CCK12, EGG16, EHL^{+}14, FGY^{+}17, FHZ10, FGHZ13, FGHZ14, FSB17, GH12a, HSMY12, HKW15, KVR11, LKK12, LZ16a, LWHY10, Liu15b, LX15b, LW18c, Loh16, LHL^{+}14b, McN12, MP12, OKTR13a, PAE^{+}12, RSS16, ST15, SZC^{+}18, SST12, TJQS13, XY17, XWY17, XWY18, YY14, AWJH19, ALMLM14, BS10b, BGF15, BSL11, BLS17, Bho14a, Bho14b, CPP15, CGK14, DLZ17, DB12, ECJ16, FR15, GS19, GLM^{+}11, GML17b, HMY15, HG18a, HPC12, HHG14, HCL12b, HWXC18a, IC12, JL17a, KM12, KAJ11, KSMT11, Li12b, LCK17, LSC17, LK18, LCHZ19, LRCG16, MSQ^{+}11, MD18, OP14, Res16, SD15a, SRGL13, SG14, TYY^{+}12, TSB16, WW14a, WZM^{+}16, WH18, XG10, XY15, yYqWqZC13, YZ12, Yüz12a, ZWLZ18, ZPS^{+}12, ZC11a]. EFG [AD15, DA16]. EFS [LL19b]. eIF4F [NJV13]. eigenanalysis [LCP16]. eigenproblem [Boy16]. eigenproblems [Gem16, JW10]. eigenspace [GN19]. eigenstructure [GH16]. Eigenvalue [Bai12, NHIN16, WS11a, AS15a, BMJ19, CM19a, CKLL10, De 10, EGG16, HA18a, Haj18c, HDT11, KSS13, LY13, LKCN19, LC10d, LvdVX18, MJB18, MRR18, RES10, RA11b, SWL15, TCM18, WmN13, WTM17, WFZ12, YH15]. eigenvalues [BL14, HT18, ORR16, TL10b, WBZY18, WZ10, YLB16]. eigenvectors [TL10b]. eighth [GK11a, SV11]. eighth-order [GK11a]. Einstein [WL13a, ZHS^{+}19]. El-Sayed [DR12]. elastic [BKE18, BDO11, BCSCB^{+}15, CFLM19, Cve11, DSVS15, DDK11, FL14, GHC15a, Gen11b, GF16, Her14, HL11c, JHW15, KP19b, LZ12a, LTT13, LMLB19, MTV13, MDL18, MNT15, OSA13, SCBCB^{+}13, SCBCB^{+}17, YLY12, Zha18d, ZTW^{+}19]. elastic-plastic [LMLB19]. elasticity [BS15b, BCG17, BS18a, BRROP19, CGJ^{+}14, DB15, DDM^{+}18, EFK15, FG18, FLdS14, FPB17, GGS16, GF16, Hes18, HVO17, HMWZ16, Kar17, LH16, LCQL17, Liu16a, QCLC17, RDE^{+}17, RS18b, WZC^{+}19b, ZR18b]. elasticity-type [DDM^{+}18]. elasto [CTS19, HAESLB14]. elasto-plastic [CTS19]. elastodynamic [MDG19]. elastodynamics [CDL17, FMS19, HPS18, TMLF19]. Elastoplastic [Bol16, ZW11a]. elastoplasticity [CHK^{+}18]. elastostatics [EKZ17]. Electric [WRW13, BN16, EPP18, KSMT11, Kup14, NZ16, TG14, Tao18, Yaz11, YAS^{+}11]. Electrical [dCMdSGTdC^{+}16, GKS17, LLG^{+}11, MH11, ZTZJ14]. electrically [MRS^{+}12, SL12, YLC12]. electricity [BBO10]. electro [BCSCB^{+}15, EKZ17, HLNZ19, SCBCB^{+}13, SCBCB^{+}17, YWHC11]. electro-elastostatics [EKZ17]. electro-thermo-convection [HLNZ19]. electrocardiology [LRBA15]. electrochemical [CS13, HNK13]. Electromagnetic [GGR19, BT14, CM10a, CJK18, CWH13, Cia16, CFS17, CHS18, CDS15, HMZ18, JN14, LZ18a, LRZ18, NJ16, OTiSY16, SBB^{+}18, SST19, ZKWW17, dBD17]. electromagnetism [AG10a]. electromagnetism-like [AG10a]. electromechanical [LLZ12]. electron [AML^{+}14]. electronic [CCKY12]. electroosmotic [AVR17]. electrophysiology [CT17c]. electrospinning [HL12, XWN11, XLF12]. electrostatic [BI12, NGG12, RMA10]. electrostatics [RCRV14]. electrothermal [MDW13]. Element [BDGG14, CM16c, CDN19, DA18c, DJD18, GCDG17, GD16, KDU15, MRR18, PGQ16, RCM11, Rus16, Vac18, dVDR18, AD15, ADZ19, AD19a, AvB16, ADGS18, ABRL18, AHHM19, AAB^{+}13, AJRWS12, AAD17, AEDL14, AJT19, AOW18, ACAS11, ABSV18, ACE17, AS19, AP19b, AO13, AKL18b, BCCZ18, BCH^{+}18, BN14a, BP19, BS10b, BGGCGRSP16, BB15, BDGS13, BWL18, BS18a, BO18b, BK15, BC17, BSY^{+}19, BS17, CKL18, CDG15, CIN^{+}18, Cal19, CPP10, CGH14, CR18a, CEQ14, CW15b, Cha18, Che12c, CLH13, Che14, CHLY15, CCZ18, CLL19, CHY19, CS14a, CFdM^{+}18, CK15, Cho17, CCK18, CELY18, CB19b, CN16a, CFLX18, CX18, DRT^{+}15, DH11b, DA16, DA18d, DFGG13, DFG19, DSZ18, DWS19, DZO^{+}19, DNS18, DGLU18, DZ17, Dun18, DHMU16, EG18, FKF13, FLdS14, FDG^{+}17, FN14, Fia15, FIW13, FGHZ17, GY15]. element [GL17a, GML17a, GLLC19, GLP18, GP19, GMS18, GGH18, GHCZ18, GS15b, GLZ18, HLL13, HMP^{+}15, HR19, HVO17, HH15, Hou15, HD16, HLY17a, HW16, Hu15, HMWZ16, HCZ16, Hu18, HK17, IK16, IS14, Ibr16, JMHF13, JZ13, JM16, JPS14, JPCY13, JHW15, JW15, JLL18, KB19, KK15, KL16b, KKL^{+}13, KVR11, KRBS18, KCK19, KdLK19, LT13, LC10a, LL12b, LL12a, LMR14, LN98, Li10c, Li10b, LY13, LR15, LY15, LLML15, LZWC16, LL16b, LS16, Li18e, LXZ18, LD18, LL19b, LNW19, LZ19a, LRH13, LWL14, LHL15, LDL^{+}15a, LDL^{+}15b, LLLW18, LZ18c, LCHZ19, LFAL19, MBS17, MMHGM17, MP19b, MLGY16, MWY17, MT18, NJ16, NNK13, Naz13, NHH13, NH15, NSYY13, OBCG19, OTiSY16, Oh15, OY19, PÁAP^{+}15, PFDG17, PGF18, PPC13, QY13, QCLC17, RZ17, RMB^{+}14, RGHZ15, RHC15, Rua19, SD15a, SD18, SBvdV13, SDH15, SD11b, ySW10, SLW14, SWL15, SW16a, SZL^{+}17]. element [SJC14, SST12, SR17c, TH19, TCHW19, TS16, UMLF13, VHPVNXW18, WSW11, WW14a, WC15, Wan16a, WHTZ16, WHS17, WC17, WCSW18, WLA18, WLZ^{+}18a, WY18a, WW19b, WZC^{+}19b, WL13b, WH14, WFZ12, WV15, XFL16, XD17, XZZ16, XFY18, XFH19, XCZQ16, YT18a, YH15, YLB16, YSX^{+}19, ZMM18, ZTR11, Zbo19, ZRE16, ZR18b, ZYWZ17, ZS13, ZHY14, ZHZ14, ZDLC14, ZH15a, ZT15, ZHC17, ZL18, Zha18c, ZY19, ZH19, ZWH^{+}19, ZLZ10, ZC11a, ZLHF19, ZLS13, ZH15b, ZG16, ZMFL18, ZZL15, ZC17, aZ18, ZKBE16, FIS18b]. element-free [CLL19, DSZ18, DWS19, LY15, LZWC16, LL16b, LD18, LL19b, ZDLC14, ZL18]. element-level [PÁAP^{+}15]. Elementary [MI16, DT16, Ji14, LSW16]. Elements [FHS18, AOW18, BPM14, BBBM16, BZKR15, BDHR18, BS15c, CKSL^{+}14, CH19, CFS17, DM18, DDD10, DGH17, EDC14, FVZ18, FMSV17, FKDN15, GLLC19, HD14b, JN14, KW14a, KB13, Li15, LCQL17, LBvB^{+}16, MNPD15, PS18, SDH13, SRDD17, SWL19, WBZY18, WN18, WBA^{+}18, YLK10]. elements.Part [Cia16]. elevated [CV14]. elevation [AD11d]. elimination [AR10c, Pas14, SC13]. Elliott [VKJ13]. ellipse [WCQ^{+}19]. ellipsoidal [BKMT16]. elliptic [ARK13, AKLS19, AVZ15, ANP19, APwS18, AG10b, AKL18b, AuIA17, BS15a, Ber16, BCJ19, BBBM16, BWL18, BMP15, CY14a, CGHW14, CCNT16, CCX13, CM13b, CCZ18, CT17b, CNP14, CCK18, CWQJ12, Cov13, DCL17, DWS19, DB12, DGLU18, DHMU16, ES18, Elb11, EZRR10, FDG^{+}17, FN14, FOS19, Gha18, GTL16, GS15b, GLZ18, HJ13, HZL17, HT16a, Hou15, HLY17a, HRHP17, HC14, HX14, HM17d, HT19, HK17, IQR16, Jal14, JK12, KK15, KN11, Kaw15, KM18, KTDT17, KK10c, Ku15, Ku18, Lee15, LCH19, LN98, Li10c, Li10b, LY14, Li19a, LC10d, LRH13, LYZ17, LW18b, MCL15, MV11, Mil18, MP19b, MY10, MP19d, NXHN14, Oru19, eOS18, PT15, RCH19, Saj14, uIAA15, SJC14, Tod15, Tod18, WZF16, WS17, WHS17, WFZ12, WH11b, XW14, XLD11a, XLD11b, Yao10, Yao16, ZYWZ17, ZHL12, Zho19, ZCY11, lZxLhY12]. elliptic-parabolic [AG10b]. elliptical [HZL17, ZH19]. elongation [DRZ10]. ELS [WCD10]. embedded [AB18, FKDN15, GF16, MM18b, PC11a, RBB12, Ran15, REHA11, WHC12, WL17b, ZLS19]. embeddings [CMT12, FWW14]. Emden [AR10a, KM11, VA10]. emergency [ZC11b]. Empirical [SCvdV^{+}19, ZLY12, DA18b, RKW12, WCH18, YWW^{+}12]. employing [FT15, JS12a]. Enabling [ADD^{+}15]. encasing [DRZ10]. enclosure [AG11, BSK11, HHM12, Pop14, RKA^{+}18, SSH15]. enclosures [PDN19, ZD12]. encoder [CUK12]. encoding [SKdA11]. encryption [GH12a, HKHK13, LW10, RWZ13, yYqWqZC13]. end [WL11a]. end-to-end [WL11a]. endemic [Sun10, ZK16]. endoscopy [Tri11]. endothelial [YLF19]. endpoint [Boy16]. ends [ZTW^{+}19]. energetic [FZBF10]. energization [GSY10]. Energy [AGK15, Chu12a, Izs15, JPK17, KKLJ11, KPK18a, LS19a, STDLM19, BT15, BM19, CT18, CT17b, DD19b, FH16, FR16, Gal10a, GOT19, HL18a, HGSL18, HB12a, HW16, KB15, LD13a, LL12d, LW18a, LZS12, Par15, Por18, SRS11, Ste16, VĆC10, VĆV11, WW15, Wu16, Yaz11, YAS^{+}11, YASK10, ZWW13, ZY15c]. Energy-balanced [Chu12a]. energy-based [GOT19]. energy-conserving [LD13a].

Energy-stable [STDLM19]. enforcement [CM12b]. engine [Ebr11]. ENgineering [SVP^{+}19, GGAVGG19, KLK13, LAM^{+}16, YLLK14, CCFV12, DELK13, FCZ12, HLK10, LLR^{+}19, LCQF19, LLW15, RS13, SKM11, ZLG^{+}11]. engines [JS12b]. Enhanced [AG11, DSWB18, GP19, JS12b, ABR^{+}14, FDG^{+}17, FJP18, JGK13, KMS19, LHHZ12, RC18, WZC^{+}19b]. enhanced-strain [RC18]. enhancement [CCH^{+}12, PvdM13, Rua19, ZLY14, ZLW18]. Enhancing [CDD12, GACMO13, Mah14]. enriched [WCSW18, Zha18c]. enrichment [GD16]. ensemble [ALLQ13, TNT12]. Enterprise [XDL12]. entire [ZY10a]. Entropically [KP18]. entropies [Gor13]. Entropy [BTEM19, MCQ11, RSL^{+}18, ZDV13, AvB16, BPM12, BDK^{+}11, Dub13, LCA^{+}17, MBKK10, NN13, RKA^{+}18, ZD12, jZsQdLmG19]. Enumeration [CJMS10, BR12c, TM10]. enumerators [ÖŞ11]. enveloping [Wri13]. envelopment [Tol12, WYK10]. environment [DH10a, DRK11, JXZ^{+}10, KHIB12, LBH^{+}12, MM18b, YWK^{+}10, ZLZ18]. environmental [DD13, IK12, PTP14, VMAVGCM19, YYYH19]. environments [AFGL10, LLL13, XBHN16]. enzyme [ABN18, CLM12]. enzyme-catalyzed [ABN18]. EOQ [CB10, San11]. epidemic [Aki17, AGPCC10, BL18, CW15a, DYQM14, GCR^{+}18, MMOJ14, PSP10, XL15, XMW10, ZY13]. epidemiological [AGU14, CLM12, ZSW19]. epileptic [HPR19]. epistemic [WLT13a, WLT13b]. epitaxial [FAHZ17, Zho19]. EPQ [CB10]. Epstein [MAB19]. equal [HR14, LFJ12, Ray18, SIL19]. equality [DBH^{+}14, WZ17b]. Equation [ZHV19, AM13a, ADZ19, AD19a, eMA18, AGT19, AvB16, Abd18b, AZ17, AO10a, AEH18, Ade17, AJS19, AK18, Akm15, AB10b, AA15, AAEG17, ARK13, Ali15, AJ12, AS11b, Amb19, ASY^{+}11, Asl10, AO13, AET19, AKMUH17, Bac14b, Bag17, BCH^{+}18, BM18a, BAO^{+}12, BCK11, BKMT14, BKMT17, BN16, BDGBM11, BC10, BJPT16, Bis10, BP18, BT14, BM18c, BL17, Bra10, BS18b, BPX11, BM19, BSY^{+}19, BK10, CCJV11, CM11a, CNR10, CHXL18, CP15b, CTSX16, CEQ14, CLS19, CBKR10, CF16a, CW10b, CDW11, CC11, CLB14, CXZ15b, CLC16, CWDL17, CM18b, CM18c, CJ18a, CHY19, Che19a, CZ15, CTZ17, CR19, CK15, CJK17, CJRR11, CCM14, CGJ^{+}14, CZ18, D’A18a, DE11, DWZ13, DLZ17, DLZ19, DN18a, DPM15, Dar11, DGA18, DS18a, DL19, Dav17, DO11, DD19a, DAM14, DA18c, DA18d]. equation [Dem10b, DZ18, DM12b, DT16, DZ21, DSM18, DZO^{+}19, DSVS15, DD16, DFW^{+}18, DTYZ18, DMV11, DBS12, Dos18, DW18b, DZ16, DJZ18, EMQ18, EGAA19, EK16, EM19, EN11, ESBR10, Erg19, EOM11, EM14, FID14, FMS19, FYYT11, FLZ14b, FW18, FJ19, FZL^{+}18, FZ18, FIW17, Fio14, FNW18, GL10, GGLP15, GZZ^{+}16, Gao17, GVSP12, GMP18, GJ19, GSPK15, GÇK10, GH14, GR13a, GSY10, GR19a, GS12, GMI11, Gol13, GL17b, GOT19, Goo11a, GKK11, GM14a, GM14b, GZD18, GS19, GABC16, GW15b, GL16, GML17b, GZW^{+}18, GTZ19, Gup11, GS11c, GHL18, GWL11, HKJ14, Hal13, HGSL18, HLTL17, HH18a, HW19a, HLLM19, HH16, He16, HGW18, HTM18, HB12b, Her19, HA16a, HHG14, HLvS18, HHGA19, HLSN16, HL18b, Hu18, HTY^{+}19, HLY12b, HX14, HXS^{+}15, HM15, HLX18, HYC18, HT19, HLI14, HLY17b, IMS19, JKB11, JKK12]. equation [Jav11, JPK17, JPK18, JQSS12, JFC14, JHW15, JLF17, JAJ18, KBA11, KB10a, KK11a, KO11a, KMS10, KB19, KSZ18, Kar10b, KK13a, KN11, KM14, KPR13, KC11, KK14a, KKAM11, KAJ11, KB15, KO11b, KBCS16, KJ11, KKLJ11, Kim14, KPK18b, KS15a, KQ11, KSMN11, KKK14, KS15b, KTK17, KT18b, KBK19, KK19b, KK19c, Kuo18, Kup11, KKD13, LMPE18, LS10a, LCM14, LHD18, LK15, LZL16, LL14a, Lee16, Lee17, LS19a, LSCG16, LSW10, LLJK10, LYZ11, LD11b, LDL11, LDW11, LZGZ11, LZ12a, LGG12, Li12a, LJSK13, LTT13, Li16b, LSZ16, LCK17, LMMF17, LLFT17, LLY18b, LD18, LT18, LM18a, LZL19, Li19b, LM19a, LCW19, LGZ19, LWL11, LZ14, LZ16c, LZZ18, LL10a, LT15b, LN19, LWKK10, LMDL11, LCZ11, LW12a, LS12c, LJX12, LW12b, Liu13, LTJ^{+}16, LL16d, Liu18a, LW18a, LZLL18, LW18b, zLYmL18, LZM18, LYC^{+}19, LWSL19, LZ19c, LLP19, LWW19b, zLZ19, zLYLQ19]. equation [LWW19a, Loh16, LLYL19, Lóp19, LSX13, LMZK16, LXY19, Luc10, LMP13, LGVS19, Luo18, MDW11, MX15, MF18b, MF18a, MYZ18, MCF18, MR17, MDRRV11, MCKM12, MA17, MZM18, MHHC18, MBHV10, MPY16, Men18, ML19b, MTAS17, MAST18, MV17, Moh15, Moo18, Mop11, MN11b, MY13, MNT15, MWY17, MT18, MCN10, Nab19, NLA19, NNWAS11, Oan13, OZF19, Oru17, Osm18, Pal12, PDM11, PC17, PP14, Par15, Par18, Pen11, PTZ19, PJ17, PL17, Pit12, Pov12a, Pov19, Pu12, QXG13, QRMH18, QX19, QTW^{+}18, QCG15, RKSA18, RNB11, RS12b, RAD13, RS14b, Ray17, RS18a, Ray18, RG18, RQ18, RWW18, RMY19, Ros12, RC17b, hRWH18, RZZ19, SCV10, SD10b, SD11a, SD18, Sah17, SR15, SR17a, Saj12, Saj14, SEY12, SH11, SMK18, Sam19, SCSF19, SBvdV13, SHM13, Sea11, Sea14, Sea15, Sea16, SIL19, SZDO10, Sha18]. equation [SA16, SS14a, SD19, SS14b, SY18, SBM10, SMBY10, SGZW18, SBM13, SS18b, SS16c, SZC^{+}18, Sou12, SC19b, SLZ11, SRG16, SND19, SMYK19, Tam16, TDXQ18, TUT11, Tha19b, TCHW19, Tod18, TS16, Tru19, TTX^{+}16, TKHL18, Üns18, UKYK17, VB10b, VRD11, WG11, WLD13, WLM13, WW14a, WQRZ14, WV16, WZM^{+}16, WHTZ16, WH16, WS17, WL17a, WF17, WTYZ17, WF18, WW18a, WZ18a, WLL^{+}18, WHS18, WRY18, WY19a, WMP^{+}19, Wan19a, Wei10b, WLL12, WZ18c, WLXZ18, WY18b, WL16, WW14b, WZ15, WW15, WZMY18, WX18a, WTLS18, XX10, XH11b, XHA13, XWH16, XX17, XCS18, XSYL19, XW19b, XX19, XaZH19, XLT17, YT18a, YYK16, YY15, YTD^{+}18, YC11, YL16, YM17, YCHW18, YÖ10, YMDZ10, Yil19, YLG10, YTF10, YT11, YL14, YMLL18, YL18b, YTC^{+}18, YYLW19, YW19, YYC11, YMHL18, YZMA18, Yu11a, YqS16, YS16a, YT18b, YH11b, YTS^{+}17, YTL^{+}18, YHC18, Yüz12a, YK17, ZT13, ZSAN18]. equation [ZSAN19, Zha18a, yZjM10, ZWC10, Zha10, Zha11c, ZS11b, ZS13, ZDLC14, ZY15a, ZTZ16b, ZC16, ZMLZ16, ZY17c, ZCZ17, ZXZ17, ZZM17, bZM17, ZY17b, ZM17a, ZDZY17, ZLJ^{+}18, ZH18a, ZWLZ18, ZLT18, ZCT18, ZCLW19, ZL19a, ZBFC19, ZDM11, ZCH12, ZL12a, qZM17a, ZTZJ14, ZLZ10, ZKW15, ZY17d, ZG14, ZY15c, ZG16, ZZX16, Zho19, ZSY19, ZCY11, ZD15, ZZ16c, ZLL18, ZS18, aZW17, aZ18, dSSV17, Li18e]. Equations [GGAVGG19, GGAVRC^{+}19, Zho11, Zho12, ZFLM19, AB10a, Abb11, AEH19, ARESH18, AIA13, ABT19, ABL15, AM10a, Ade16, AR09, AR10b, AZH10, AAZ10, AK10, AA11, AdAS11, ABJ11, AÖ10b, Agr10, ANP11, AW11, AN12, ATUC15, AuIK17, AK18, AJY13, AMA14, AP10, AÇT11, AA15, AJAR18, AAH^{+}18, AM14a, AHO16, AJ10, AJ11, AK19, AJT19, ABL10, AHOP18, AKRT14, AT17, AD12, AM12c, Arq18, ANP19, APwS18, AS10c, AY12, AA13, Asl11, APTZ19, AuIA17, BMRA10, BM11a, BD11a, BD11b, BJQS18, BN14a, BK11a, BKT11, BK12a, BM10a, BMA11, BC12, BR12a, BH10, BG15, BP11b, BMTV12, BI12, BCF13, BR18, BS11b, BB10a, BT15, BK14, BB10b, BDS10, BB12, BS16b, Ber16, BO18a, BDGBM12, BBBM16, BZ18, BQS16, BGPP11, BE11, BG11, BE12]. equations [BKY10, BML11, BMS13, BY11, BCF^{+}14, BDO11, BDHR18, BOT14, BH11, BHJ14, BB08, Bra16, BR12d, BR13b, BC17, BS14b, BMH19, BK12b, BPF13, CCJ10, Can11d, CBBE16, CH11a, CM18a, CDG16, CRA19, CDM12, Cha11a, CANK11, CZN11, CZN12, CLM14, CYS10, CZ11a, CDY11, CM11b, CCX13, CLCF14, CM16b, CM17, CZL17, CJ18a, CY19, CLA19, jC11, CZ17, CT17b, CS14a, CFdM^{+}18, CS12, CL17b, CH17, Chu11a, Chu10, CP15c, CP16c, CC17, Chu18, CELY18, CCP19, CFS17, CHS18, CdR18, CJCV10, CDS15, CN16b, CFF15, Dai14, DZW16, DLS18a, DDMQ19, DCL17, DH10b, DM15, DM16, DQ10, DM19a, DMD10, DRD12, DPZ13, DSR10, DG10b, DMZ10, DLD10, fDxZ11, fDxZ12, Din13, DHY19, Don10a, DC15, DK18, DNZ^{+}13, DCRL13, DLS18b, Dua18, EO14, EBENEA10, EBENF10, ED11a, Elb11, Els10, EG18]. equations [Ers16, ESL11, EKE18, FT10, FSHZ11, FH11, FN14, FH16, Fer12, Fia15, Fis18a, FM12b, FIW13, FRZ15, GGR15, GGAV18, GY15, hGzS15, GML17a, GH16, GRBT16, GK13, Gen10, Gen11a, Gep16, GB16, GH15, Gha17, GTL16, GY11, GMB12, GS11a, GS11b, GHR10, Gol13, Gon13, Goo11b, GGEB12, Gro19, GR13b, GJ12, GW12c, GMZ15, GLW18, GH18, GZN19, GM11, GGO16, GH10, GH12b, GH13b, HMF^{+}19, Haj18a, Haj18b, HL18a, HC16, HG18a, HGHA19, HKI12, HEP10, HY13, HY16, HLL13, HKK^{+}16, HP17, HP19b, HVR18, HIS19, HS13, Hes14, HSJ15, HSMT19, HA16b, HNPS13, Hon10, HXL11, HM17b, HF10, HA10, Hua10a, HLZD11, HZM11, HLY12b, HM14, HM17c, HM18b, HCLL18, HLW19, HM19, HT16b, IB11, Ibr11, ILS13, Iom18, JKK11, JGSS10, JYF^{+}11, JA11, JKK12, JNBK13, Jan10, JPB11, JJ19]. equations [JW18a, JL11b, JW18b, JW19a, JL19, JL20, JS11, JLTB12, Jia12, JD12, JYYL16, JZJ18, JL18, JZL18, JKS18, JKS19, Jum10, JKMS12, KP13, KVW18, Kar18, KÖ10, Kat11, KSG11, Kaw15, Kes10, KM12, KKBR19, KBAF18, KW11, KTDT17, KMRN12, Kig10, KG11, KL16b, KAA19, KPG18, Kon16, Ku18, KY11, KK14c, KT18a, KV18, KMT10, KM11, LK15, LT13, LC10a, LL12b, Lee15, LC11a, LLZ10, Li10b, LCW10, Li10e, LF11c, LJ11, LZC11, LS11b, LW11b, Li11a, LWBW13, LCS15, LZWC16, LL16a, LH16, Li17b, LSS17, LCQL17, Li18c, LQMW18, LCS18, LHY18, Li18d, LX18, LJK^{+}19, LZ11e, LK18, LRZ18, LBJ10, LLY10, LBW11, LL16c, LKU10, LJ10b, Liu11a, LXK11, LLH11a, LY11d, LZKU11, LHF11, LXP11, LZY11, LLH11b, LS12b, Liu12, LHL12b, LRH13, LLLC14, Liu15a, LX15a, LL15, LZ15b, LLT16a, LLT16b, LWZ16, Liu16d, Liu16c, LYZ17, Liu17, Liu18c, LCLL18, LB18]. equations [LFAL19, LW13, Lu11, LHTL19, Luk11, Luo19, LT11, LLX11, LM18c, MX10, MKHC11, MF11, MZL13, MLL16, MWL18, Ma19, MBS17, MN11a, MTM11, MBH11, MKR12, Man18, MTN19, Mar16, MR10, MPfTX18, MP19a, MDBCF16, MO14, MLZ^{+}16, MP11c, MS12a, MD15, MPLR18, MY10, MI16, MM18d, MS15, MSG11, MN10b, Mor13, MGW11, MAH18, Mur08, MB10b, Naw11, Ned12, NMR15, NZ16, NT17, Odi10, Ols10, Oru19, OF16, Ouy11, ÖZ11, Pan11, PLKC16, PS16, PR11, PS12a, Par11a, PIAH10, PGW19, PH19, PZA19, Per18, PPD10, Pet11, PT15, Por18, Pov12b, Pov19, PP12, PLT17, QAA^{+}16, QZF19, QHT16, QLT^{+}18, QY17, QCS^{+}19, RM17, RHMA18, RSDR11, RZ16, RKF18, Ray16, RS15, Rey12, RÖ10, RTT17, RTV17, RCM11, RA12, RA19, SP10, SD10a, SNH10, SNMA12, SD15a, SC19a]. equations [SYG11, SR15, SSA12, SH11, SZA^{+}18, SD12a, Sal11, STC18, San12, SBEB10, SAIZ15, SMM19, SKTH11, SZ12a, SBB^{+}18, Sed13, SYI12, Sha14, SHH16, SBS12, SG14, SXM11, SZZ11, SM17, She16, SS17, SWL16, SG11b, SC16, SW16a, SZL^{+}17, SW17, SLM18, SW12, SOK19, SPST18, SG16b, SG16a, SG10b, SWC11, SJC14, Sou11, SSC19, SLM16, SYW11, Su12, zSdZ10, SLL12b, SK14b, SLW18, SR18, TZWM10, TZ15, TTG16, TM17, TG11, TS11a, TNP17, TAS11, TNF11, TB10, Tia19, TMDTTC16, TT14, Tom11, TY16, TDN19, TNHK19, TTC14, UKI11, ÜM16, VA10, VA11, WC10b, WSG10, WKG10, WW11a, WS11a, WSW11, WW11c, Wan11, WZWX11, WAZ11a, WZF12a, WZF12b, WKS13, WV14, Wan15, WW16, Wan16a, WC17, WW18b, WH18, WLZ^{+}18a, WCH18, WW19c, WW19b, Wan19b, WZC^{+}19a, WZG19, Waz11b, WS10b, Wen18a, Wen18b, WFDW10a, WFDW10b, WLDL11a]. equations [WH11b, WFL11b, cW11, WLDL11b, Wu11a, Wu16, Wu18b, WCLD18, XWY11, XZ10, XY15, XYXZ16, XHM14, XM15, XZ18, XY10, XC17, XZ19, Yak11, YK18, YXS10, YX11b, YX11a, Yan12b, Yan17, YGS17, Yan18b, Yan18c, YQWZ19, YS19a, YZY10, Ye16, Ye17b, YP10, YL10b, YLG17, YN16, YDL11, Yu11b, YT13, YS16b, YS17, YTZ17, YWT18, YZS18, YQ18, YSX^{+}19, Yus09, YSS11a, Yüz11, YSS11b, Yüz12b, Yüz12c, ZA10, ZTR11, ZM16a, ZFZ10a, ZFZQ10, ZJ10b, ZYG10, ZLZ11, ZTH11, ZD11, ZGW11, ZQ11b, ZL11, ZLY14, ZTC14, ZLL14, ZCH14, ZH15a, ZYW15, Zha15a, ZZ16b, ZY17a, Zha17b, Zha18e, Zha18g, Zha18d, ZZL19, Zha19a, ZY19, ZZG19, Zha19c, ZYZ11, ZSHL11, ZHL12, ZWMD16, ZLL17, ZLA17, ZZL^{+}18b, ZP18a, ZZL18a, ZWH^{+}19, ZÖXL^{+}19, ZTSC16, ZJZ18, ZL10b, ZLS13, ZT18a, ZJ10c, ZW11d, ZWW13, ZFLM18, ZP18b, ZP18c, ZSD10, ZQ14a, ZLW19, ZC17]. equations [ZDF^{+}14, ZT18b, ZMA10, dVLV18, dAS18, dSAC11, uRK11]. Equi [KD10]. Equi-statistical [KD10]. Equidistributed [Mic17]. equidistribution [KH18, MN10a]. equilibrated [CLN^{+}19, NB17]. equilibria [BBDS11, BL12, CGY10b, NKM15, VGK^{+}16]. Equilibrium [SL16c, AK11, CB11a, CB11c, CGHY11, CyL11, CAP10, CS11b, DGM14, Gro19, GDM13, JL16, JQG14, KJA10, KPR13, KM10, Laz10, NN12, Pal12, Pir11, PT11, SK11a, Sca11, She11, She12, SW11, Sun10, YNLK10, ZLL11, sHC11]. equipment [ZZXY12]. Equivalence [WS16, PKK12]. Equivalent [AAB^{+}13, LCH19, SNEP19, Che12c, RHD18, SST19]. Ericksen [MSW18, Zha18f]. Ericksen/Allen [MSW18]. Erratum [ID10, KYR11a, LYS12a, WWW11a, WWW11b]. Error [CKL18, CXMO19, DHMU16, FT10, FM11, HR19, HLY17a, HC14, HZP18, LA11, LZH16, LR17a, RTT17, SZ17, WLC11, XFY18, ZCW15, ZHZ14, AV14, Bac14a, BP19, BPC17, BO18a, BN16, BQS16, BWL18, BV17, BS15c, CCKP15, CCZ18, CJN19, CHY19, CUK12, Cho17, CH17, CGO19, CLN^{+}19, CGJ^{+}14, DESV18, DPM15, DA18d, DSZ18, DWS19, DMRS18, EO15, EE10, FR16, FOS19, GG18, GRBT16, GGS16, GS18, GOS18, GSZ14, GH18, HM18a, HLL^{+}15, Hou15, Ibr16, Ikh11, Izs15, Kim14, KT15, KKJ15, LMR19, LHL12a, LX16, LCH19, LSZ11, LPP15, MMT18, MRR18, OC14, OY19, PS18, PLT^{+}19, PL17, QZ16, RDE^{+}17, RCRV14, RA11b, RC18, VP11, VAK^{+}19, WZ18b, WNTW19, Wei17, YS19a, ZDL11, ZY11, ZD15]. error-control [PLT^{+}19]. error-resilient [CUK12]. errors [CTM^{+}13, KT11b, TTMJ12, XC11a, XLK11]. errors-in-variables [CTM^{+}13]. Escherichia [XBHN16]. escrow [HKHK13]. Escrowable [NCL13]. essential [JM15]. Essentially [GDM13, LX18]. Establishing [HHS^{+}10]. Estimate [WZ10, BQS16, BV17, CKL18, CXMO19, DA18d, DLS18b, GSZ14, LWSL19, MPZ11, VP11, Wei17, Wu18b, ZHJZ11, ZY11, ZY15c]. estimates [Bac14a, BP19, BWL18, CCRS17, CCZ18, CM16c, CH17, DSZ18, DWS19, DHMU16, EGG16, GH18, HR19, Hou15, HLY17a, HC14, HZP18, HT19, Ibr16, Izs15, JLL18, Kar17, Kim14, KT15, LCH19, LWBW13, LW19b, LPP15, MMT18, McN12, MP12, MRR18, OY19, OÖ11, QZ16, RCRV14, RA11b, SZ17, WKP^{+}14, XFY18, YS19a, ZCW15]. Estimating [WL12a, LWC13, WL11b]. Estimation [Luk11, OC14, AHV10, AY18, AEG11, ASB12, BKR11, CJ18b, CHY19, CMS10, CTM^{+}13, FSM19, FJC16, FM11, FR16, HW11, KSMT11, KPP13, LK14, LX16, LD13b, LD11a, LL13, Lin10b, LCT12, LSD10, LB11, MM13, PL17, Pop13, SMC10, SLYY13, TWLYÖ10, TZMZ12, WYD10, ZDL11, ZGZ13]. estimator [AV14, CJN19, RC18, WL11b, WZ17a]. estimators [CCKP15, CLN^{+}19, KKJ15, VAK^{+}19]. eTCR [HKKK13]. Euclidean [PDHL12, tWqLzGkP11]. Euler [ABT19, ABL15, Agr10, AM12c, BM12a, Bri10, CM12a, Gol13, JYYL16, LX18, MLL16, Mor10c, Mor10d, MT18, OF16, Öza11, OSS10, PTP14, QCS^{+}19, Sle13, WFL11b, XZ10, YDL11]. Eulerian [AP19a]. European [ASSV18, Ciu11, KDU15, LW17, ZLTY16, ZBL12, ZH18b]. evacuation [EAEH18, GW12b, ZC11b]. evaluating [DA12, RZL11, SD15a, Zha11b]. Evaluation [FL13a, HMF16, HWyL11, MMHGM17, WTC^{+}12, BS10b, BA11, BSN13, CM10a, CP16b, CKM12, GP19, Hot13, HSBL11, Ima17, ID10, JW05, JLCS10, KRGS^{+}10, LLSS13, dCMdSGTdC^{+}16, RFK16, SSS^{+}11a, WhJxLwW11, WWH12]. evanescent [SKCL19]. Even [YZMA18, JW18a, LLY10, LZY11, ZYG10]. even-order [JW18a, LLY10, LZY11, ZYG10]. Even-parity [YZMA18]. Event [JCWZ16, WCZ13, dlPBTW13]. Event-driven [JCWZ16]. events [AM11]. Eventually [BC12]. evidential [KBDC12]. Evolution [SD15b, AA15, Asl11, APTZ19, Bis10, BB08, CLCF14, CZL17, Chu11a, CP15c, CP16c, CFF15, DGA18, DB11, EO14, EBENEA10, FOX11, GYTD12, GM11, HY11, JKK12, JKS18, KS12b, LK15, LL16d, LMZ17, MB10b, NT17, PRR18, RSB14, SSA12, SZDO10, SG10b, TTG16, eT10, WFC16, XYXZ16, YMLL18, YS17, YHC18, ZLL14, ZM17a, ZLT18, ZLY^{+}13, ZJ10c, ZY15c, ZLGL11]. Evolutionary [CG15, GACMO13, LYC12, SLMZ12, ZM18, AZB13, BDK^{+}11, Bra13, Bur13, Che11e, CQLX11, CH17, DZS10, DSZ18, HDHL11, HM18a, KOPS13, LHL18, MVB^{+}12, OSZP13, SZ17, THZ^{+}11, YC10a, Zha18a, dAS18]. evolutions [ZTY^{+}19]. EVSS [Wan13b]. Exact [AA15, BEAA11, BGF15, DGL12, Gep16, HBE14, JKB11, KSG11, LCM14, LK15, LD11b, LDL11, LZL19, MDW11, MP11b, MP16, MCKM12, Tar17, XX10, YGS17, ZLZG11, Ade17, AAEG17, BCSCB^{+}15, CCKP15, Cie13, CJCV10, DGA18, DGZ13, DPZ13, DLS18b, EK16, EM19, FKDN15, GK13, HLTL17, JRZK11, JL12, KNZ19, Kuo18, zLZ19, zLYLQ19, LSX13, uHS12, MM18a, MI16, PMA17, Ray16, SR17a, SH11, SMK18, Sea11, SK11d, TZ15, Wei10b, YMLL18, YYLW19, ZTW^{+}19, ZBL12]. example [AZB13, BDHR18]. exchange [Col14, GLM^{+}11, KMT19, SCA14, TL18, YNS^{+}14]. exchange-based [YNS^{+}14]. exchanged [HYS^{+}14]. exchanged-based [HYS^{+}14]. excitation [GB18, SP12, WCH13]. excitations [BDGS13, BPS18, LKL^{+}15]. excited [ZHS^{+}19]. exclusion [MY16]. Exclusive [GXZ11, XGXZ10]. EXCMG [DLZ19]. Executing [HSS^{+}12]. exemplified [Zbo19]. exhaustive [WY11b]. Existence [AZH10, AA11, ABJ11, AEO15, ALI11b, BS12a, BK11a, BK12a, BH10, CW19, CM15, CP15b, Cha11a, CC19, CCX13, CTC17, CTG17, CO19, CHZ19, CyL11, CS11b, DN18a, FSHZ11, FKeT12, GTC18, GRS18, GL17b, Goo11b, Goo11a, GH13b, HYCP11, HVR18, HB12b, HX14, JL11a, JPS10, Jia12, JZ11, Kim11, KMT10, LLZ10, LF11c, LJ11, LZ12a, LL16a, LT18, Li18b, LS19b, LZ11e, LZ11d, LLZ18, LH10a, LKU10, LWKK10, LZKU11, LLL12, Lü14, Luo18, MX10, MTAS17, MS12a, MP19d, MAH18, Ouy11, PT11, PLT17, QCT17, Qiu12, SLK12, SM17, She18a, She18b, SLL12a, SZ12b, SLL12b, TJ10, eT10, TB10, TDN19, TNHK19, WZF16, WS17, WL17a, WW14b, WZ15, Wu18a, XZR16, XC18, XLD11a, XLD11b, Yan10d, YW19, ZFZ10a, ZBF11, Zha11c, ZZ15b, ZXZ17, ZH18a, ZDB19, ZCH12, ZJ10c, ZZ18c, ZCY11, ZLW19, dSAC11, AKL18a, AKRT14, BZ10]. existence [BM10a, BB12, BK10, CT11b, CHS19, DG10b, DC10, FH16, HKJ14, HL18a, Jal14, JLWX18, JS11, JLL19, JLF17, KKD13, Li13, Li16a, Li17b, LX15a, Luc10, Ma18a, RHMA18, SNDK18, SW12, SW11, Sta11, WAZ11a, WAW15, XWY11, XZ19, Ye15, Ye17a, YLG10, YM13, YS19b, ZSH11, ZFZQ10, yZjM10, ZFZ10b, ZS11b, ZJ17, ZSL19, Zho16, AaC19]. exotic [KKT13]. Exp [Asl11, GM11, NTR15, AKMUH17, YP10, Ade16, Asl10, BB08, Cha11c, Chu11a, RNB11, ZZ11b]. Exp-function [Asl11, GM11, NTR15, AKMUH17, YP10, Ade16, Asl10, BB08, Cha11c, Chu11a, RNB11, ZZ11b]. expandable [HD16]. Expanded [WW19b, IK16, LRH13, SJC14]. expanding [QWGJ15]. Expansion [CM11a, WH11b, YMDZ10, CK15, CRRS11, Dem10a, Dra11a, Gep16, GMZ15, HA18b, HLZD11, JKB11, KBA11, KY10, LZH16, LR17a, MCKM12, MBHV10, NSYY13, PZ11b, SG10b, YH12]. expansion-iterative [MBHV10]. expansions [MD10]. expansive [WC11a]. expectation [LCCC10]. expedited [JFS14, JFS20]. expensive [QAA^{+}16]. experiment [CI18]. Experimental [GSI19, CB19b, MLSLM15, YWL^{+}11a]. experiments [Ehr18, HNK13, KV17a, XC13]. Explicit [AM17, Ben17, CY19, DWZ16, Kaw15, LL12a, PBK19, AVV18, AKSW19, BDHR18, DGM14, HMF^{+}19, HLL^{+}15, KAS11b, MMT18, MVKK14, MM18d, OC14, PRS18, Sou12, XZZ16, ZGD14]. Explicit/implicit [CY19]. exploitation [SLCC12]. Exploring [DM18]. exponent [D’A18a, DL18, DSL18, FRZ15, KSZ18, LeT10, LY14, LL16a, LT18, LS19b, LLZ18, LLT16b, MAST18, MPLR18, Sha18, She18a, SC19b, WLW^{+}11]. Exponential [Ahn12, Cvi10, IW18, KP13, KO13, ML19b, SS14b, WL10, Bai11a, BEAA11, BTB18, BK13, CL16b, FM18, FXCC18, GP11, Gar13, HA18b, JHW15, Kli10, LK18, LZZ11a, LO16, MF18a, MSV18, MSA12, MT18, PS12a, PCM12, SI10, Tam16, UMLF13, WSS10, WV14, XH11a, YT18a, ZFZ10a, ZFZ10b, ZJ12]. Exponentially [ED11a, Chu11c, CFLX18, TH13, WHZL17, XFL16]. exponentiation [KBDC12]. exponents [MV11, RA19, SNDK18, Wan15, WS17]. expression [JL12, LY11d, MBH16, PC11a]. expressions [Tia11, TO11]. Extended [CAH11, Kim17, AEH19, Ade17, AM10b, ABCR10, Bho14a, Bho14b, CL12a, EAAED10, EAA10, FVZ18, He16, KO11a, KO11b, KK19c, KM10, LYC^{+}19, LMZK16, Man18, MZM18, NUNAS11, PXXZ16, Sea15, SJPS11, ZSAN18, ZSAN19, GD16]. Extending [CRG16, JTC^{+}10b, XW19b, Boy16]. extensible [MX15, YW11b]. Extension [AZ15, PB12, PCK13, RR14, YSW16, BS12b, BGM19b, CM14, DDD10, Dos18, GY11, HRMS12, KV10, LCW12, PC11a, RZ17, TÇA11, XKH10, ZWY19, ZL13]. extensions [Jic10, LSM10, XW10]. extensive [LZF18, WZG19]. Exterior [Dai14, BT15, BBBM16, CLL19, DGR18, FIW17, GVSP12, LDHH13, Sam19]. exterior-like [BT15]. external [COR18, DS18a, PK19, Sah17]. Extinction [CP15c, GCG12]. extinctive [CP16c]. Extra [Boy16, MHL11]. extracted [AY18]. Extracting [BPG10, LYX11]. extraction [HpD11, JPP12, PDHL12, ZWJ^{+}11]. extragradient [CB11b, CAY12, YLK10]. Extrapolation [FHZ10, ZGD14, DZW16, DLZ17, DH18, FHZ13, hGzS15, Gos10, GH18, HRHP17, WW16, WW18b, ZDF^{+}14]. Extremal [LZ10, QZ11, Ili12]. extreme [LLD10]. extruder [HHS^{+}17]. eye [FdOdSS17]. eyed [CNH17].

F [Pan17]. Faà [XW10, XC11b]. Faber [Gal11a, SKPW14]. Fabrizio [AVR17, MV17]. face [HCL12b]. face-structured [HCL12b]. Facebook [HCHH12]. facial [PC11a]. factor [CKL18, DESV18, KL16b, WCH18, YZGW10]. Factored [LZ11c, BK14]. factorial [CI18, Mor10b]. Factorization [Raf12, RTB14, JJ13, LYS12a, LYS12b, LZD17, YXYH10, ZHB11]. factorizations [PPC13]. factors [ASSV18, BS16a, Bor10, Dos18, DD13, NJV13, PTP14, Sar11, Sav10, SS19]. FADE [CLJ11]. failure [DRK11, FSM19, LL13, VC12, WLT13a, Yoo17, ZC11a]. failures [ABDKD12, YWW^{+}12]. fairing [HH18b]. fairness [CZ10]. falling [JK11c]. families [GHM^{+}14, IBB10, KYO10, RR19, ZR18b]. family [AS11b, BG10a, CsH10, CGHY11, CS10c, CS11b, CS11c, CSN11, Fer12, Gar13, GK11a, GK11b, GDZ11, KS10a, KBS11, KKS10, KA10b, KT11b, Li15, LCC12, Mor13, Pir11, She11, SOJC10, sHC11]. fan [LSM10]. Fast [Bho14a, Bho14b, BCG17, BE18, BG13, Che11d, CL17b, DWZ13, DFM15, FF15, HIS19, KQ11, KYY12, LCLL16, PJ17, QCG15, Som13, TXL19, VMFF18, ZZL19, AP19a, AML^{+}14, BC16, BDS15, CDY11, Che15b, CWDL17, CUK12, Du12, DJZ18, FSB17, FZL^{+}18, FNW18, GIMZ14, GM14a, GA10, HLT12, JW18b, JW19a, Li19b, LN19, Liu16a, LCW17, dCMdSGTdC^{+}16, MR19, YZWW14, ZN18, ZLC^{+}14, CC15]. fatigue [LYJ15]. fault [LTT13, PA15, PFBL10, XCM12]. faulted [PFBL10]. faults [MVB^{+}12]. FCMBP [THZ^{+}11]. FD [MFSL19, Bay19, FLP13, GV11a, HMV18, JM15]. FDAEs [IB11]. FDM [SS19]. FDTD [HCLL18, NZ16, SWL16]. FE [ABB17, JM15]. feasibility [CAY12]. feasible [Zha15b]. Feature [PTL13, JWX^{+}13, MLY11, PDHL12]. features [CLW11, FHH13, IK12, JMNZ19, LHHZ12, NJV13, SST19, ZLY12]. February [Ano18-73, Ano18-55, Ano19-63, Ano19-65]. Feedback [BHH16, KLY16, XCXW10, ATZ11, Che11c, HSC17, HLZ15, LZT11, LCYC12, LMZ17, SSAM11, Wei12b, Zhe11]. feedforward [SYO12]. Feir [HS12]. Fejér [BCB11, Som13]. Fekete [ODR10]. FEM [AHP^{+}14, BS14a, BS15a, BLS18, BPS19, Bol16, BV17, CGHW14, CO15, EGG16, FWW14, FM18, FP18, GM18a, GOS18, Git14, GM19, GH18, PZJ^{+}16, QZM17b, SW17, Von19, Wei14, Wei17, Yan15, ZR18c]. FEM-based [PZJ^{+}16]. FEMs [WY15, YS19a, YWH14]. Fermat [KV10]. ferroelectric [MM18b]. ferromagnetic [AHP^{+}14, GTZ19, LM19a, MGS^{+}14]. ferromagnetism [LTJ^{+}16]. FESTUNG [FRAK15, JRA^{+}18, RAW^{+}16]. FETI [CHK^{+}18, KLP10, KCL14b]. fever [ZLZ18]. Feynman [DZ18]. FFT [BM18b, Cos18, VZM14]. FFT-based [BM18b, Cos18, VZM14]. FG [ATH18, ATH19]. FGM [BKE18]. fiber [HTWS15, LTSW16, Sza15, YTC^{+}18]. fiber-reinforced [Sza15]. fibering [ZG18a]. fibers [AGH^{+}15, PvdM13]. Fibonacci [CKMR11, CM10c, DRS11, EE10]. Fibonacci-polynomial [EE10]. fibre [ZR18b]. fibrous [BCSCB^{+}15, Ryl15a]. Fick [BHZJ19]. Fickian [KRP12]. fictitious [BZKR15, FT15, HHS^{+}17]. Field [WRW13, BPS19, BTEM19, BN16, BKMT16, CPP10, CYP16, CFdM^{+}18, COR18, DFW^{+}18, FGB19, HHM12, ILP^{+}11, KVW18, LK11, LTT13, LH19, LSC17, LCWZ18, LC12b, MCB10, PFDG17, PGF18, SLL17, SWC11, STDLM19, Sul16, TG14, Tao18, WCW13, Yaz11, YAS^{+}11, ZRE16, ZKWW17]. fields [AS11d, Cia16, CFS17, Kia18, Kup14, LMW10, LX18, LRZ18, OTiSY16, ÖŞ11, Ryl15b, TMLF19]. fifth [BQS16, RR18, SG14]. fifth-order [RR18]. filament [YNS^{+}14]. file [JXZ^{+}10]. filled [AG11, GYH11, RSL^{+}18, RKA^{+}18, SRRP18]. Filling [FGPP17, DC15, HLZM16, IS14, MGTH16, ZOZZ12]. film [BRFH16, CJ12, GHL18, JK12, KWFY11, KS12b, MJ14, Mom11]. films [Lin14]. Filter [TKBMT17, ZZC13a, ZZC13b, HDHL11, ILP14, KPP13, PTL13, ZLL12]. Filter-matrix [ZZC13a, ZZC13b]. filtering [BKP11, CFN11, CCR16, CCSZ14, DNR13, KNIF13, KPP13, TSB16, TMZ^{+}15]. filtering-based [DNR13]. filters [Ali11a, JK11b, Pet15, YZ10b]. filtration [Zho16]. final [CLJ11, TKBMT17, TDN19, TNHK19, XW19a]. finance [PG10]. Financial [Baa13, Lee14, DGTC13, SS19, Tha19a, WHS11, ZLY12]. financing [Chu11c]. find [FHH13]. Finding [SBS12, ZM14, Boy16, Gem16, KG11, PZ11a, PZ11b, Pas14, PPD10, XYXZ16]. Fine [JCZZ13, AES11, FBB10]. Fine-grained [JCZZ13]. fines [IL13]. fines/bacterial [IL13]. fingering [ST14]. fingerprinting [TXZ^{+}10]. fingers [LL12a]. Finite [AG10b, BR18, BDGS13, BDGG14, BK15, CPP10, Cha18, CFLX18, CDN19, DK18, FXC18, FN14, FIS18b, GMS18, GCDG17, GD16, KO11b, KCK19, LBvB^{+}16, LDL^{+}15a, MvS18, RCM11, SNSK19, SLW18, SI17, THGG14, WZ18a, WL13b, XFL16, ZKR^{+}12, ZH15a, ZCLW19, ZY17d, ZG16, ZQ14a, AD19a, AvB16, ADGS18, ABRL18, AHHM19, ABCR10, AAD17, AEDL14, AJT19, AOW18, ABLS15, ACAS11, AVZ15, Ara18, ACE17, AS19, AP19b, AO13, AKL18b, BCCZ18, BCH^{+}18, BN14a, BP19, BS10b, BPM14, BGGCGRSP16, BB15, BMM12a, BBBM16, BWL18, BS18a, BZKR15, BDHR18, BMSS18, Boy10, BO18b, BC17, BSY^{+}19, BS15c, BS17, CsH10, CB11c, CKSL^{+}14, CKL18, CDG15, CGH14, CH19, CR18a, CEQ14, CW15b, CG13, CWW15, CLM12, CLM14, CLH13, CHLY15, CS18, CCZ18, Che19b, CHY19, CTZ17, CV14, CS14a]. finite [CFdM^{+}18, CCCW10, CK15, Cho17, CCK18, CS10c, CELY18, Cia16, CFS17, CB19b, CM14, CG14, CN16a, CX18, CCM14, Cos18, CSU13, CT17c, DRT^{+}15, DL19, DO11, DA18d, DFGG13, DFG19, DT17, fDxZ11, DNS18, DGLU18, DZ17, Dun18, EFK15, EDC14, ED12, EG18, FKF13, FLdS14, FDG^{+}17, Fia15, FVZ18, FIW13, FMSV17, FGHZ17, FNW18, GY15, GL17a, GLW13, GML17a, GLP18, GP19, GCE18, GGH18, GD11, GR10, GS15b, GDZ11, GLZ18, GZW^{+}18, Gur13, GGO16, HMF^{+}19, HHY13, HLL13, He16, Hei10b, HJD15, HS18, HR19, HTL10, HVO17, HH15, Hou15, HD16, HLY17a, HW16, Hu15, HMWZ16, HCZ16, Hu18, HZM11, HD19, IK16, IS14, Ibr16, JSGP16, JMHF13, JZ13, JM16, JPS14, JPCY13, JW18b, JW19a, JLTB12, JHW15, JYYL16, JLL18, JN14, JZE^{+}18, JAJ18, KS10a, KB19, KK15, KKS10, KA10b]. finite [KW14a, KL16b, KT11b, KKL^{+}13, KVR11, KB13, KRBS18, KdLK19, LT13, LC10a, LL12b, LL12a, LMR14, LN98, Li10c, LMW10, Li10b, LY13, LR15, LLML15, Li15, LR17b, LLFT17, LW17, LCQL17, Li18e, LXZ18, LHL18, LX18, LNW19, LZ19a, LH12b, LRH13, LWL14, LHL15, LX15a, LDL^{+}15b, LZJY16, LS17b, LLLW18, LZ18c, LCHZ19, LFAL19, LLYL19, LR13, MZC17, MBS17, MZM19, MS11a, MMHGM17, Mil18, MP19b, MNPD15, MFSL19, MMH11, MHM11, MM18d, MLGY16, MWY17, MT18, Mur08, NJ16, NNK13, Naz13, NHH13, NH15, NSYY13, OBCG19, OTiSY16, Oh15, OY19, ÖŞ11, PS16, PPC13, PS18, QY13, QCLC17, RZ17, RMM11, RMB^{+}14, RQ18, RGHZ15, RHC15, Rua19, RS18b, Saj12, SBvdV13, SMM19, SDH15, SDH13, SRDD17, SD11b, gShYL10, SH12a, ySW10, SLW14, SWL15, SW16a, SZL^{+}17, SB14, SJC14, Sou12, SST12, SK19, SR17c, SR18, TPHD18, TTT10]. finite [Tam16, TH19, TCHW19, TS16, TCM15, VHPVNXW18, WSW11, WWA11, WW14a, WAG^{+}14, WC15, WZM^{+}16, Wan16a, WHTZ16, WHS17, WC17, WCSW18, WH18, WLZ^{+}18a, WBZY18, WW19b, WH14, WFZ12, WBA^{+}18, WV15, Wu18b, WX18a, XD17, XZZ16, XFY18, XFH19, XGH17b, XCZQ16, YT18a, YH15, YZMZ16, YWH14, YSX^{+}19, ZTR11, Zbo19, ZRE16, ZR18b, ZR18c, ZYWZ17, ZS13, ZHY14, ZHZ14, ZT15, Zha18c, ZL19b, ZY19, ZH19, ZWH^{+}19, ZLZ10, ZC11a, ZH15b, ZMFL18, ZSY19, ZZL15, aZ18, ZKBE16, dVLV18, sHC11, DJD18, WKBR18]. finite-difference [Cos18, Gur13, Hei10b, HS18, Saj12]. finite-element [AvB16, AHHM19, Rua19]. finite-size [GLW13, WAG^{+}14]. finite-source [GD11]. Finite-time [SNSK19]. finite-volume [GR10, GGO16, RQ18]. finitely [Lam13]. finned [Ima17]. fins [RSL^{+}18]. Firefly [dSCM12]. firms [Che11e]. First [GGAVRC^{+}19, Hes14, HZ11, Jan10, QZ16, VLFS12, BS15c, CCKP15, CW17, CLTA11, DG13b, DGT18, HM18a, JRZK11, JZL18, KPG18, Ku15, LKU10, LHF11, MDVM17, MBHV10, MJWD19, SL18b, TZWM10, TZZ11, TT14, Tsi11, Waz11b, YJ19, Yüz12b]. First-order [Jan10, CCKP15, DG13b, DGT18, HM18a, KPG18, Ku15, LKU10, SL18b, TZZ11]. Fisher [AÖ10b, He16, KO11a, KO11b, LLSW10].

FISM [FKC12]. fission [DYH11, YYC11]. fit [Che12c]. fitness [SSM12]. Fitted [Arq18, BP11b, CWW15, LZZC12]. fitting [HH18b, HD19, JCF19, RKW12, yS10, ZL13]. FitzHugh [AAA12, UKYK17, Zha18h, ZS15]. Five [Boy16, RQ18, ZRE16]. five-equation [RQ18]. five-field [ZRE16]. Fixed [AE12a, AHF10, AH11b, HLS11, IN10, KK10b, MTM11, Mar11a, XLY10, ZC10, Zha11d, Ahm10, Aki17, AIB10, AR10d, AKS11, APS12, BS12b, BR12c, BKR^{+}19, BKNR19, CB11a, CB11b, CB11c, CFN11, CAY12, Cha11b, CAP10, CSW11b, CS11d, CM10b, ĆSCD11, DL16, EE18, GRS12, KPR10, Kar10a, KT11a, KL12a, KKS10, LJYS18, LC11b, LT11, MP11b, MP16, MAK12, NS11, NKM15, Pir11, Pop11, Qiu12, RH15, RA11a, RI12, SK11a, SLK12, SZ12a, Sha10, She12, SCK11, YBC11, YLK10, Zha14, dBD17, sHC11]. fixed-point [Zha14]. Fixed-points [XLY10]. fixed-stress [BKR^{+}19, BKNR19]. FL [LXF11]. flame [Özd18]. flames [AM13b, AM18, XA13]. flash [WY19b]. flat [DVM12, FZBF10, IC12, PvdM13, TMDTTC16]. flatness [VMO10]. flavored [DLS14]. flaw [SR16]. Flexible [SSESG12, AT18a, DM12a, DRK11, KSMT11, KRD16, Kuo16, NKA18, SST12, YLC12, YNS^{+}14, ZG19]. flexible-joint [YLC12]. flexural [CRA19]. flicker [AR10c]. flight [YGH11]. floating [BR13a, ÇA14, DTR19]. flocculation [JQG14]. flocking [BHK16]. flooding [JS12a]. Flow [BHZJ19, ZZC13b, AVR17, ASFM15, ALHZMC^{+}19, Ala10, AR10c, AKSW19, AdSSS19, AR17, AGPR19, AP19b, AJS14, BP19, BASW18, BPR18, BV10, BGP13, BMSS18, BPZ19, BKR10, BFS15, CCR16, CCY18, CZ11b, CV14, CP10, Col14, Das12, DSWB18, DNP15, DZW^{+}15, DSL11, DSM18, DGM14, DFP^{+}13, DKM17, Eba14, EG10, EUTS18, EEBM10, FKKS11, FBTS19, FZBF10, GQF^{+}10, GLW13, GZR^{+}13, GHT^{+}15, GD10b, GOT19, GK16, GK18, GDM13, GSZ11, HMF16, HLB14, HMM12, IIHI10, IIHuI10, IBSS11, JNJ^{+}11, JK10, JK11a, JGK13, JVMF19, JQG14, KMT19, KP18, KG14, KSD^{+}19, KRCJ11, KWFY11, KPG18, KTH13, KCL12, KRBS18, KORR10, LDS10, LR14, LF11b, LBZL11, Lin11, Lin14, LBvB^{+}16, LZZ11a, LZZ11b, LGL^{+}14, LZP^{+}19, MMS^{+}18, MA10b, MC11, MCR11, MAN^{+}15, MGN^{+}16, MM18a, MDL18, MEAMHHV18, MM19, MMA12, NMR15]. flow [NNL13, OHMAK18, PZZ18, PTH^{+}16, QaY18, RHMA18, RSV11, RBB12, Ran15, RA18, RQ18, RTL19, Rua19, SSS16, SK14a, Sah11, SL12, SPH10, Sal16, SCA14, Sea15, STS19, ySGL^{+}10, SB14, SKTD13, SR10b, SRV10, TR14, TH19, TC18, Tod12, TDM13, UKA15, UMLF13, WH11a, WL13b, WCQ^{+}19, WBN18, WS16, YZ10a, YB13, YNS^{+}14, ZAK18b, ZLS19, ZCSG13, ZZX^{+}14a, ZZX^{+}14b, ZZLB18, ZDB19, ZJB19, ZNWG11, ZLC^{+}11a, ZWG11, ZD18]. flow/Darcy [LWL14]. flowmeter [HBK^{+}19]. flows [AD19b, BTEM19, BB18, BSZ16, Bis14, CP15a, CYL17, CL15, CD16, CHH14, Cos18, DdSF13, DVY14, DJD18, Ers16, EAAS18, FLZ14a, FL13a, GLR13, GLP18, GJX18, GGO16, HBK^{+}19, HG16, HS18, HWyL11, IK16, JKZ11, KRP12, KSO16, KLRW12, LMPE18, LZZC12, LLML15, LZPZ19, LH19, LSC17, LMLB19, LG13, LZG13, LZB15, MSV18, MA16, MC10a, MR14, MMHGM17, MS17, MLZ^{+}16, NNK13, NDC^{+}19, NN13, OHK^{+}19, PDN19, PLR15, PD11, PM10, PLMS14, RAZ19, RY10, RRAK19, RTRR18, SSHH^{+}18, ST15, SPT17, SMF17, SS16a, Ser19, SZ17, SL18a, SGK18, SPP18, SKFG11, SKTC15, SKTC19, Sul16, TY13, VMP15, WWLL13, Wan13b, Wan14, WSCW16, XZ11, YY11a, YSW16, YL18a, ZSW15, ZYZ^{+}16, ZLZG11, ZZC13a, ZZC13b, ZQ14b]. flowshop [AA10b, AA10c, LLW11, LBH^{+}12]. flowtime [LCC12]. fluctuating [DH10c]. fluctuations [De 10, VV14]. Fluid [HMF16, MMS^{+}18, AGT19, ABT19, ABV11, AGD10, AYY12, Bab14, BHK16, BHZJ19, BP19, BPR18, BMS13, BV10, BGP13, BPZ19, BKR10, BHH16, CCR16, CTSX16, CST14, CLTA11, CRXL15, CYZZ18, CL15, Das12, DSK^{+}14, DTR19, DFP^{+}13, DKM17, DNR13, Eba14, Ebr11, EDC14, EEBM10, FKKS11, FZBF10, FAIV10, GK16, GK18, GF19, HB19, HZ10, HHS^{+}17, HLSN15, HTY^{+}19, IBSS11, JNJ^{+}11, JRZK11, JKZ11, KEHB18, KNT12, KRCJ11, KDG11, KPG18, KSO16, KTH13, KVR11, LX16, LM18a, LL16c, LZZ11a, LZZ11b, LGL^{+}14, Liu18a, zLYLQ19, Ma18b, uHS12, MZM19, MSV18, MC10a, MC11, MCR11, MM18a, MMT18, MMHGM17, MR19, MUB^{+}16, MLZ^{+}16, MMH11, MEAMHHV18, MLSLM15, MM19, OHMAK18, PFDG17, RSV11, REHA11, RQ18, RHC15, Rua19, Sah11, SL12, SPH10, SCA14, SISH12, SM10, Sul16, TTX^{+}16, TRL19, UKA15, VPR11, VHPVNXW18]. fluid [WSW11, Wan14, WSCW16, WTLS18, XXG10, XBHN16, YZ10a, YJ19, Yoo17, YTL^{+}18, Zak18a, ZYT^{+}16, ZZLB18, ZLZG11, ZWG11, ZLHF19, ZH15b, ZPGW16, ZD18]. fluid-conveying [XXG10]. fluid-dynamic [BMS13]. fluid-particle [ZYT^{+}16]. fluid-solid [BPZ19, HLSN15, MMT18]. fluid-structure [HB19, HZ10, LX16, MZM19, MUB^{+}16, RHC15, Yoo17]. fluid-submerged [VHPVNXW18]. fluid-thin [KEHB18]. fluidic [HWY14]. fluidized [ZLC^{+}11a]. fluids [ATO19, AVR17, BB18, HLB14, HZLM10, KL16a, Kup11, RY10, Tri11, Yan19, Ye19, YM13, ZR16, Zha19c]. flutter [Kia19]. Flux [SL16c, AP19b, BRR16, CS18, DFG19, EA10, FDG^{+}17, GWR^{+}18, HLSN16, LWSL19, LvS15, MN10a, PK19, QCS^{+}19, RA18, RDE^{+}17, Ryl15a, SS16a, XWY11, YSW16, YZMA18, ZQ14b]. flux-vector [ZQ14b]. foams [HJD15]. foci [Wei12b]. Fokas [AAEG17, EGAA19, TDXQ18]. Fokker [AAH^{+}18, AHOP18, Loh16, SK14b, ZL12a]. fold [Noo10]. folding [TH13]. foldon [DYH11]. follower [yN11]. following [HR15]. food [Bae10, MW17, PQB^{+}16, PQBK17, TZ13]. food-chain [TZ13]. force [BCD^{+}16, CJP12, CJP15, FIVM17, FIMV18, HM10, KNZ19, LX10a, TBP19, WSCW16, WGY^{+}18, ZYSY17, BDR19]. force-correction [FIVM17, FIMV18]. Forced [HF10, AAZ10, AM10b, KNIF13, LS12c, PHM^{+}19, SSH15, SK11c, SZP^{+}11, WAG^{+}14, WL16]. forces [Cve11, SEM13, WZC^{+}19b, WCH13]. Forchheimer [AGPR19, IK16]. forcing [BCD^{+}16, FLZ14b, Gen11a, Ima17, KYA15, MPGW19]. forecasting [LCCC10, SLCC12]. foreign [PLKCC12]. forensic [GMAM12]. Foreword [Mai12]. form [BD19, CL17b, DLD10, DFS14, HZ16, HZ18, JLCS10, JW15, KTK18, Kup11, LTX10, LC13, uHS12, MTN19, MV12, Men18, PMM17, SNH10, SMDI18, SBM13, TC10, WFDW10a, dS15]. formal [LML11, Med12]. formation [AGT19, FL13b, GZW^{+}18, IW18, MGTH16, NGL10, RdSRL19, SD11b, SFM15, YGH11, ZY15b, ZS15, Zho13]. formations [CM10a]. formats [EHL^{+}14, LYM12]. forms [CDG16, CP10, ID16, Lam13, LNP^{+}12, XD10]. formula [AJY13, ADK10, AM13c, BM10b, Dos10, GTG11, GZ14, IS12, KB10b, Mor10a, Mor11b, Pal12, WCH18, XW10, XC11b, ZH18b]. formulae [KAS11a, LCN10, SH12b, Wei10a]. formulas [Mor11a, PMM17, Tha19a]. Formulation [GGVRB19, BPS19, BGR14, Bog10, BM18d, CL11, CGO19, FKF13, Fio14, FIW13, Füh18, GGS16, GV11a, HMF16, HMWZ16, JRA^{+}18, KPS10a, KPS10b, Kar17, Kar18, KLP10, KCL14b, KCL16, LMR14, LVF^{+}16, LPML19, LSV18, MMT18, MSFS18, NJ16, RS14b, RG11, RTT17, SH12a, SP12, SAR18, TLR17, VGC^{+}15, WHS17, WHS18, XW14, YASK10, YTZ17, YT18b, ZRE16, ZR18b]. formulations [AHHM19, AJRWS12, BQ17, BT14, CW17, CR19, CLN^{+}19, Sag10]. Fornberg [AO10a, GS11c, Lu11, SEY12]. Fortin [Che14, NPD17]. forum [LYX11]. forward [DE11, HLvS18, Tha19b]. forward-backward [DE11, Tha19b]. foundation [BKE18]. Foundations [Ana10, TWLYÖ10]. Four [DD10, ZL10a, AW11, LW11a, MM10b, SR10a, YGH11]. Four-dimensional [DD10, MM10b]. four-point [AW11, LW11a]. four-tiered [YGH11]. Fourier [AK19, BHZJ19, BM13b, BCF10, CNR10, CK15, DZ16, EO15, GGGR13, HLX18, ILP^{+}11, LL14a, Lee17, LK18, Oh15, RA19, Sat11, THGG14, ZJZ18]. Fourth [PS16, WGY^{+}18, AM13a, AZ17, BCK11, BDO11, Bra10, CY14a, ÇT12, CLTA11, CT17b, CM16c, DL10, DL19, DC10, Fis18a, FR16, GK11b, GS11b, GL16, HY10, HLL13, HK17, IBSS11, KS12a, LCN10, LW11a, Li16b, LDL^{+}15a, LDL^{+}15b, LLLW18, LLYL19, MX10, NLA19, Naw11, PCO16, PR11, SXM11, SGZW18, WWA11, WRW13, WZM^{+}16, XY10, YZX18, YC10b, ZSH11, ZY17c, ZZX16, Zho19]. fourth-grade [IBSS11]. Fourth-order [WGY^{+}18, AM13a, AZ17, BDO11, Bra10, CY14a, CLTA11, CT17b, DL19, DC10, Fis18a, GS11b, GL16, HLL13, KS12a, LCN10, Li16b, LDL^{+}15a, LDL^{+}15b, LLLW18, LLYL19, MX10, NLA19, Naw11, SXM11, SGZW18, WWA11, WZM^{+}16, XY10, YZX18, YC10b, ZSH11, ZY17c, Zho19]. fourth-step [GK11b]. Fowler [KM11]. FP [YÖ10]. FR [LZD17]. Fractal [Hot13, PU10, Ryl15b, SG10a, WZXL11, YL18b, BI12, CSZK10, FJC16, KLMV12, Lin10b, LW11c, WZ18a, ZDV13]. fraction [GK11a, GK11b]. Fractional [AT11, Ana12, BMC13, BI12, BDGBM11, CYS10, DV10, DPZ13, DB10, EK13, GS11b, IMD11, JKMS12, KB10a, KFYW11, KA13, LMPG13, LRTV10, Mag10, MBJ16, Oua12, PM13, RRGTV10, Sea15, SC19b, YjH18, Zho11, Zho12, ZFLM19, ABM11, AM13a, AD15, AD19a, ARESH18, Abd11, AEH18, ABB17, AHV10, AZH10, AA11, AdAS11, AJS19, AN11a, ANP11, AW11, AN12, AK18, ArEM10, AAH^{+}18, AM14a, Alo11, AKL18a, AK19, Amb19, AT18b, AJT19, AT19, Ana10, AM10c, AHOP18, AHJM18, AKRT14, AR17, Arq18, AE12b, APTZ19, AET19, BZ10, Bai11b, Bai12, BS12a, BHZJ19, BK11a, BKT11, BK12a, BKT12, BZK12, BM10a, BMA11, BR12a, BMJ10, BB18, BG19, BB10a, BHM12, BCK11, BDGBM12, BZ18, BH11, BL17, BMH19, BSY^{+}19, BK10, CHS11, CLT^{+}13, CFRS10, CHXL18]. fractional [CP15b, Cha11a, CANK11, CZN12, CB11d, CC19, CCD10, CSZK10, CBM10, CZ11a, Che12b, CLCF14, CLB14, CW14, CXZ15b, CLC16, CWDL17, CZL17, CYZZ18, CJ18a, CHZ19, CLA19, CWY19, CTM^{+}13, CL17b, DG13a, DGB10a, DGB10b, DGTC13, DN10, Dar11, DGA18, DN18b, DH17, DB11, DA18c, DA18d, DQ10, DZW^{+}15, DZ18, DMD10, DRD12, fDxZ11, fDxZ12, DBEE11, DM19b, DMV11, DH16, DCRL13, Dun11, DZ16, DJZ18, EBENEA10, EBENF10, EGAA19, EKS10, EN11, ESBR10, ESN10, Els10, EOM11, EZM12, ESL11, EKE18, FDB13, FLZ14b, FJC16, FJ19, FES17, FSHZ11, FZL^{+}18, FZ18, Fer11, FNW18, FKeT12, GD10a, GTC18, GP11, Gar13, GMP18, GK13, GH14, GR13a, GY11, GS11a, GL17b, Goo10, Goo11b, Goo11a, GZ10, GABC16, GJX18, Guo12, GJ12, GW12c, GML17b, GLW18, GF19, Gup11, GS11c, HK10, HXX19, HS11a]. fractional [HP19b, HO10, HM10, HB12b, Her14, Her19, HHG14, HA16b, Hol11, HNPS13, HLvS18, HTV13, HA18b, HHGA19, HGN^{+}10, HXL11, HLZD11, HLX18, HLW19, HT19, IB11, Ibr11, ILP^{+}11, IW18, JGSS10, JYF^{+}11, JA11, JNBK13, JRZK11, JKZ11, JMB10, JL11b, JW18b, JW19a, JLTB12, Jia12, JD12, JL18, JLL19, JZ11, JZ12, JC12, JLF17, Jum10, KK11a, Kar18, Kat11, KK14a, KKBR19, KBAF18, KKAM11, KAJ11, Kir10a, Kir10b, Kli10, KCK19, KBK19, Kun12, LZ11a, Lan12, LHD18, LeT10, LEN10, LMS13, LLZ10, LCP10, LJ11, LZC11, LDW11, LS11b, LX12a, LLC13, LZ16a, Li17b, LLY18b, LS19b, LGZ19, LD19, LZ11e, LZ11d, LK18, LRZ18, LN19, LJ10b, LB11, LZZ11b, LS12a, LS12b, LZB12, LJX12, Liu12, LC13, LDL^{+}15a, LDL^{+}15b, Liu17, LZCL18, LZLL18, LCWZ18, LLLW18, LZP^{+}19, LZG19, LFAL19, LDY11, LCC13, LXY19, LHTL19, Luc10]. fractional [LMP13, Luk11, Luo19, LLX11, MP10a, MPZ11, Mac12a, Mac12b, MT10, MAB19, MHHC18, MCL^{+}13, MZB10, MNJ^{+}13, MLZ^{+}16, MB10a, MV17, MP11c, MMH11, MHM11, MRS^{+}12, MM18d, MN10b, Mop11, MN11b, MY13, MGW11, MAH18, Mur08, MCN10, Nab19, Naw11, NB11, Odi10, OMT12, OSA13, Ouy11, Pad18, PCS13, PS16, PH19, PZA19, Pet11, PAT12, PAT13, Pov12a, Pov12b, Pov19, PK19, PA12, Pu12, PLT17, QXG13, QX19, QZF19, QLT^{+}18, RMS10, RM17, RZ16, RA18, Ray16, RS18a, Ray18, RG18, RSS10, RA19, RZZ19, SD10a, SD11a, SMF10, SSAM11, SC19a, SR15, SEY12, SRM11a, SSA12, Sal10, Sam19, STC18, SAIZ15, SMM19, SKTH11, Set12, She18b, SZL^{+}17, SY18, SW12, SS18b, SS16c, SZC^{+}18, Sou11, Sou12, Sta11, SZ11, Su12, SLL12a, SZ12b, SLL12b, SND19, SI17, Tar17]. fractional [eT10, Tha19b, TB10, TS11b, TC10, TMSO12, Tri11, THD19, Tru19, TKBMT17, TKHL18, TDN19, TNHK19, VA11, VMO10, WHS11, Wan11, WZWX11, WAZ11a, WL11c, WL12b, WZF12a, WZF12b, WV14, WV16, WW16, WH16, WW18b, WH18, WLZ^{+}18a, WLL^{+}18, WRY18, WW19c, WZC^{+}19a, Wei10a, Wei12a, WZKY12, WLL12, WZ18c, WLGL10, cW11, WW14b, XW19a, XZR16, XH11b, XC13, XHA13, XWH16, XJ18, XSYL19, XX19, XaZH19, Yak11, YY14, YY15, YXS10, YZ10a, xYsHjL11, Yan12b, YL16, YGS17, YCHW18, YJ19, YQWZ19, YDL11, Yu11a, YW11a, YDW15, YQ18, YZ19, YSX^{+}19, Zak18a, Zha10, ZBF11, Zha11c, ZS11b, ZTH11, ZW11b, ZLPM13, ZCH14, ZTZ16b, ZLTY16, ZZM17, ZZLB18, ZH18a, Zha18b, ZCLW19, ZZL19, ZSL19, ZL19a, ZJB19, ZHS^{+}19, ZSLZ19, ZY19, ZSHL11, ZCH12, ZL12a, ZLA17, ZZL^{+}18b, ZWH^{+}19, ZLZ10, ZJZ18, ZL10b, ZJ10c, ZG16, ZFLM18, ZZ18c, ZP18b, ZP18c, ZLW19, aZW17]. fractional [aZ18, ZMA10, dSAC11, uRK11]. fractional-delay [YW11a]. fractional-in-space [RZ16]. Fractional-order [EK13, IMD11, ABM11, AHV10, DGTC13, DM19b, FDB13, HK10, JZ12, JC12, Kun12, LZ11a, LCP10, LLC13, LCC13, MHM11, MRS^{+}12, Pet11, SD10a, XC13, xYsHjL11]. fractional-power [HM10]. fractional-step [DN18b, DZW^{+}15]. fractionally [DM10]. Fractions [SVY16, ZZT11]. Fracture [BCPS15, CR18b, Che19b, LS17b, MPS18]. fractured [AGPR19]. fractures [DJD18]. fragmentation [CHY12]. frame [FNZ13, LGHR16]. framed [TZMZ12]. Framework [HSS^{+}12, BMS19, BM18b, BMS12, BP18, CMMP11, DNS18, DA12, HpD11, IBG12, IL13, KSPP11, KK12, Kim17, KRD16, LYM12, MS17, Rob14, RGHZ15, WTSS10, ZHV19]. frameworks [SBB^{+}18]. FRAP [PMM17]. Fréchet [CANK11, KY10, Ols10]. Frederick [CTS19]. Fredholm [BMRA10, BGPP11, BKY10, KMRN12, MN11a, MBH11, MSG11, SD10b, Waz11b, YSS11b]. Free [ABR^{+}14, EF14, JK11a, Kia16, WL17b, AD15, AML^{+}14, BC15, CB19a, CQ13, CCY18, CLL19, CJN19, DA16, DA18c, DSZ18, DWS19, EE18, FF15, FRSC16, Gal10d, GOT19, HW19b, JK10, JGK13, JW15, Kia18, LD13b, LHY11, LMR14, LSW16, LY15, LZWC16, LL16b, LD18, LL19b, LPML19, Ned12, PLW^{+}18, PHM^{+}19, RSL^{+}18, RZL11, RRAK19, RSS10, RS18b, SH18, SSH15, Sin16, Ste16, UKA15, VCM11, VAS^{+}18, WmN13, WZ17b, WL13b, YK18, YDL11, ZDLC14, ZHC17, ZL18, Zha18c, Zha19a, ZTW^{+}19]. free-surface-flow [JK10]. freedom [Rus16, YMM12]. frequencies [XXG10]. Frequency [MMT18, YASK10, BD19, BT14, CL11, CFRS10, CCKY12, CJP12, CJP15, GM18a, LYJ15, LCT12, MR15, MPMTV15, OTiSY16, PD17, QCG15, RG11, dBD17]. frequency-amplitude [YASK10]. Frequency-explicit [MMT18]. frequent [FKC12]. friction [BDS17, DKM17, SH12a, TS11b, YGR11]. frictional [BS15b]. frictionless [SbX19]. Friedrichs [GMS15]. friendly [GVJ13]. Fritz [CS10a]. Frobenius [BR12c]. frog [CCL^{+}12]. front [BP11a, HBS^{+}10, LGVS19, PC14]. front-tracking [PC14]. frontal [WKP^{+}14]. frontier [He11]. fronts [CS16, CDW11, JSGP16, She16]. frustration [YA11]. FSI [RKD18, WCZ^{+}19]. FSKD [LZL^{+}13, YWL^{+}11b]. FTS [LCCC10]. fuel [AaC19, BGRS11, CS13, GZR^{+}13, HMP^{+}15, JS12b, SCA14]. Fujita [CCP19, DL18]. Fukuyama [SR17a]. full [CI18, KVW18, WSH12]. FullSWOF2D [WBN18]. Fully [DM19a, GGH18, HM18a, Kim14, LX16, SPH10, TG14, Tao18, WL16, ZP18a, AsNAd10, BP19, BO18a, CGO19, DWY15, DVY14, DSVS15, FN14, GGVRB19, Li10e, LWZ16, MM18b, Pap15, WZKY12, ZR18c, ZHZ14, ZZL^{+}18b]. fully-discrete [ZZL^{+}18b]. fully-mixed [CGO19, GGVRB19]. fully-symmetric [Pap15]. Function [DDM^{+}18, MvS18, PW11, Ade16, AH11a, ASMEE11a, Asl10, Asl11, AKMUH17, BASW18, BL14, BB08, Cha11c, CP16b, CZ10, CHL18, Chu11a, DHQ11, DO11, DCG^{+}12, Elb11, GK11a, GK11b, GM11, KKT13, KC12, KRGS^{+}10, KH13, KYA15, KY11, Lam13, LLW11, LLWZ11, LLSS13, LCLL18, Luh12, LLW15, dCMdSGTdC^{+}16, MSTB17, MB10a, Mor10b, Mor11a, NTR15, NWZ11, OCNG12, PA12, RNB11, RRC11, RFK16, RKF18, RHD18, Saj14, SL16b, SKJ10, SJPS11, SW19, TSB16, ÜM16, WWG10, WL11b, WLHZ14, Wei10b, YX16, YP10, YTZ17, ZY10a, ZZ11b, ZX11, ZJ12]. Function-based [DDM^{+}18]. function-vorticity [YTZ17]. Functional [BH11, AZH10, AJ11, Bae10, BH10, BO18b, BPX11, CANK11, CP16a, CW10b, CZ11a, Chu10, DE11, DN10, GP19, GKK11, GKM11, GJ12, Jan10, JLWX18, JS11, Jia12, JD12, KP13, KC11, KH18, LG10, Li13, Li14, LWD15, LH10a, LZY11, LY11c, LS12b, MPS18, MRR11, Par17a, SP10, SCV10, Su12, TNF11, WSG10, YXWL14, ZL14a, ZCY16, ZCH12]. functional-differential [CW10b]. functionally [AB18, FXCC18, Kia16, KM13]. functionals [DB15, LG17, RSB14]. Functions [BT10, AKT12, AD10a, AD11a, ABR10, AT11, ADK10, Ant10, AS10a, AS10b, ASMEE11a, AX11, AS11c, ASMEE11b, ANR11, AuIK11, BS10a, BMRA10, BM11a, BS15a, Ber16, BL14, Boy10, BM13b, BBD10, CCJV11, CCN14, CGH14, CEF^{+}13, CKN11, Che11a, CS10b, CH11b, CL16b, Cvi10, DE10, DAM14, DM16, DMPV10, DD10, DW18a, DP15, Dra10, Dra11a, Dra11b, DRS11, EAAED10, EAA10, ELS11, FRZ15, FKDN15, Fur13, GYH11, GMP18, GS12, GK18, Gue13, GJ10, HS11a, Haz11, HT12a, HT13, HK15, HKKK13, HT16b, JW19b, KVJB15, KV17a, KV17b, Kaw15, KRP12, KC10, KLTS11, Kir10a, Kir10b, Kli10, KHUO12, LS10a, Las10, LW19a, LY10b, LY11a, LSM11, MBH11, Mar11b, MJ10, MG11, MI16, MN10b, Mos10, NP12, Noo10, NNAS11a, NNAS11b, NNAS11c, NM11c, Noo11, NM11b, NB17, NODA11, Ols10]. functions [ODR10, ORD11, OSS10, PT15, PG10, QYL10, RKP12, RS12a, RRGTV10, RA12, Sae11, SK11b, SSO10, SID15, SKPW14, SOS11, Sim10, uIAH10, SRM11b, Sok10, Sok11, SWW11, Som13, SGQ12, SW16b, Swa10, SK11d, Tha19a, TMDTTC16, THD11, THH12, TO11, Wan08, Was13, Wei14, Wri13, XC11a, XLY10, XZL10, YL10a, YH12, YY11b, YH11a, YLLN16, Yüz12b, ZZ11b, ZX11, ZC16, Zha18d, ZcHS18, ZZL18a, Zhu10, ZS18]. functions-based [ZcHS18]. Fundamental [DH16, MV17, AA18, BML11, BL17, Deu10, Erg19, GK18, LCLL16, LW18a, LMP13, OZF19, SVY16]. fundus [MAN^{+}15]. furnaces [KSMT11]. Further [HSBL11, KÖ10, LC12a, LZ19b, PB11, DG10a]. fusiform [WSC16]. fusion [DYH11, HLT12, JMNZ19, LZS12, RMK19, ST12, YYC11]. Future [WZCC10, BEAA11, Eba11, PHWM10]. Fuzzifications [JK11b]. fuzzified [MM11]. Fuzzy [AN11b, AD11c, BL11, BKM11, BKK11, DZK10, Dav10, DZY11, DLWW12, HX11, JKK10, JLP10a, JK11c, LKS10, NGL10, ÖI12, PdlF10, PATA11, PB13, SYL10, SKH12, WO10, WPH11, Yan11a, ZZ10a, Abb10, AH10a, ADA11, Ahn12, AD11b, ABFGZ11, BV11, BKR11, BBC^{+}11, BC11, BMM12a, CS10a, ÇA10a, CA10b, CKN11, CZY13, CCBSRFRM11, CL12c, CHT11, DH10a, DRK11, Deb12, DMPV10, DSA09, FJWW16, FLLF10, GKK11, GSS11, AB11, GWL11, HVA10, Haz11, jHlXZ11, Hua11, HYL10, JS12b, JTC^{+}10a, Jia11, qJhY12, JMST11, JW19b, Jun10, JLP10b, JY11, KD12, KOPS13, KSKK11, Kuo11, LF11a, LN10, LS10b, Li10a, LZH12, LH12a, LZYW13, Lin12, LXF11, Liu11b, LC12b, LWC13, MZLF10, MZJ11, MS10a, MSA12, MZQ11, MBKK10, Mok11, MSG11, MME10, NM11a, PC12, PB12, PCK13, Ram11, SP10, SR12, Sae11, SDM10]. fuzzy [SK10a, SJN10a, SJN10b, SM11, gShYL10, SMC10, SZGG11, SLYY13, SSL11, TMMASG10, THZ^{+}11, TK11, TD10a, VP11, VB10a, WSL10, WC11b, WYG12, Wan13a, WL13a, WZWS11, WYK10, WD12, XS10, YÇG12, YLY^{+}09, YWHC11, Yil10, YZXW10, YZ10b, YZGW10, YLJ12, hYxLL10, hYlLL11, ZJ10a, ZFZQ10, ZJZ^{+}11, ZW11c]. fuzzy-stochastic [DRK11]. FVM [MLSLM15]. FWA [Mok11].

G [BA11, CyL11, Jic10, LXYT11, Pan17, WY11b, BA11, Yu11b]. G-convex [CyL11]. G-WKI [Yu11b]. G. [EMRS12]. G.729 [YZ12]. G/G/1 [BA11]. G/M/1 [BA11]. GA [DWI^{+}12, KX12]. GA-based [KX12]. GA/PSO [DWI^{+}12]. Gaddum [AH10c]. gain [IK12, KJA10]. Gait [SCC^{+}12a]. Galerkin [AD15, AvB16, AEG18, AEH18, AMGC19, AB10b, AT18b, AYH17, AO13, Bac14b, BP19, BP11a, BNTT14, BS16b, BDF16, BQS16, BS14b, BKZ17, CKSL^{+}14, CCN14, CGGM19, CNV14, Cer18, CEQ14, CW17, CWHW17, CWWY15, CFdM^{+}18, CDL17, CHH14, CN16b, DA16, DA18c, DFGG13, DWZ16, DSZ18, DWS19, DB12, DHY19, EDC14, EHL^{+}14, FRAK15, FHS18, GL17a, GML17a, Git14, GM19, GML17b, GHL18, HA16a, HT16a, Hof18, Hou15, HMY18, Izs15, JMHF13, JRA^{+}18, JZL18, Kar17, KB19, KWAS16, KCL14a, KW14a, KCL14b, KW14b, Kim15, KCL16, KPS17, KS15a, KSMN11, LPLR19, LZWC16, LL16b, LSS17, LXZ18, LQMW18, LD18, LNW19, LMLB19, LZ15b, LWZ16, LCW17, LZLL18, LZ18c, LFAL19, Loh16, MBS17, MT19a, MLG17, Moo18, MWY17, NCC13, NN13, OBCG19, PGQ16, Pet15, Pul16, QY13, RAD13, RCH19, RAW^{+}16, RRAK19, RTT17, Rob14, RCM11]. Galerkin [Ros12, RC18, SNMA12, SDH15, SL18a, SXM11, SW16a, SZL^{+}17, SW17, SLM18, SSS11b, SR17c, SND19, TAPA^{+}17, TCM15, TW18, UMLF13, VZM14, WW14a, WLL^{+}18, WZKY12, WSM^{+}19, XW14, Yil19, YSB15, YK17, ZYWZ17, ZS13, ZDLC14, ZSY14, ZLJ^{+}18, ZL18, ZLL17, ZP18a, ZY11, ZZX16, ZD15, ZC17, aZ18]. Galerkin-like [YK17]. game [CQLX11, DH10a, KWPK13, SLXC11, ZC11b]. game-theoretical [ZC11b]. games [CA10b, KJA10, LH12a, yN11]. Gamma [Bai19, Che11a, LNP^{+}12, Mor11a, PG10]. gap [LGC^{+}17]. gaps [HL11a]. Gardner [PDM11]. Gas [LBZL11, ABL15, BMS13, CM19b, GZR^{+}13, GT15, GT16, HBS^{+}10, HL18b, KSO16, Kup14, LF11b, LZPZ19, LZZ19, LG13, LZB15, Mok11, PZZ18, PLMS14, SCA14, Wan12, WSCW16, WSM^{+}19, ZLC^{+}11a]. Gas-kinetic [LBZL11, HL18b, LF11b, LZPZ19, LZZ19, LZB15, PZZ18, PLMS14]. gas-liquid [KSO16]. gas-lubricated [Wan12]. gas-mixtures [GT15, GT16]. gas-solid [ZLC^{+}11a]. gas-vapor [Kup14]. gas-water [LF11b]. Gate [CM12b]. gauge [CDN19]. gauged [Luo18]. Gause [LPY16, WX18b]. Gause-type [LPY16]. Gauss [AD10b, BDM19a, ED11a, Gon13, PCS13, RRC11]. Gaussian [Boy10, BM13b, DO11, FLP13, LFJ11, LSJ12, Luh12, MDVM17, MR15, MT19b, PLT^{+}19, RFK16, RKF18, SC13]. Gaussian-based [FLP13]. Gaussians [Xia11]. Gaussons [YS17]. GBH [Jav11]. GCD [Ipe12]. gear [CJ12]. gear-bearing [CJ12]. gel [GGH18]. Gelfand [EM19, SMK18]. gels [LCA^{+}17]. General [CCRS17, HC16, HW19a, LTX10, NN12, Par18, Pet15, AB16, Ara18, BM10b, BGF15, BS11b, BPG10, BKY10, BBD10, CsH10, Cao19, CL17a, CGY10a, CGY10b, CGY11, CKW13, CT18, DYX11, FN14, Fra11, Gal11c, GTC18, Haj18a, Haj18b, HVO17, JJ13, JL15a, JL15b, JJ15, JYL16, KA10b, KTDT17, KL19, Lee11c, LY11d, Luo19, MB11, uHS12, Mar16, NUNAS11, PC17, PW18, RWTW19, SCSF19, Sar11, SOK19, Som13, SW19, TM18, TZG10, TNF11, Tod15, TAA14, TÇ12, TC10, Wan08, Wan18, WL10, Xu14, XGH17b, YH11a, YX11c, Zha11a, ZHY14, ZSQ^{+}18]. Generalised [MS10a]. Generalization [TD10b, VSI12, AZ15, BMY13, GSR14, Liu10, Noo11, SS11b, XZL10]. Generalizations [QH11, LZJ12, MB10a, Pop11, Yak11]. Generalized [Agr10, ABDKD12, AHOP18, CT10a, CELY18, DMD10, GP11, Gue13, jHlXZ11, KT11a, Kha10, KLMV12, MC11, MT10, MKS13, MD10, OMT12, OCNG12, Pet14, Pov19, PB13, RK10, STC18, SS16b, SC16, TM17, WJWW12, YDL11, ZWMD16, ZLWL11, ADZ19, AD19a, AKT12, AE12a, Ade16, AA15, ABCR10, AH11b, ASMEE11a, ASMEE11b, ANR11, ASV11, BEAA11, BK11b, BR18, BC10, Bis10, CB11a, CM16a, CMR17, CGY10b, CGHY11, CTSX16, Cha11b, Cha11c, CLF10, CDW11, CLTA11, CTC17, CM18b, CyL11, CZ15, CZ17, CS12, CAP10, CSW11b, CS10c, CBB15, Cie11, ĆSCD11, Ciz12, CM10c, DAM14, DM16, Dos18, Dua18, DKM17, ES17, FWZ16, FSZ17, FM12a, FHA16, FSCG11, GML17a, GS11b, GKK11, GJ10, HLLM19, HGHA19, HY15, HKK^{+}16, HP17, Her14, HZ11, HM17b, HTY^{+}19, HM14, HM17c, HLX18, HM18b, HYC18]. generalized [HWXC18b, HWXC19, IK16, JPS10, JBBL17, JK11b, KV10, KSG11, KM14, KAS11a, KK10b, KB10c, Kir10b, KSMN11, KK19b, LX10a, LK15, LCP10, LS11b, LXL12, LZWC16, LZL19, LCW19, LWL11, LC11b, LL12c, LMRS10, LMDL11, LZZ11a, LSZ11, LZZ11b, LFJ12, LGL^{+}14, Liu16e, Liu16c, LZ19c, LWW19b, LWW19a, LS11d, LXY19, Luc10, Lup11, LM18c, ML19a, MDRRV11, MRS15, Mar11a, MPfTX18, Men18, MPS18, MLZ^{+}16, Moh15, Mor10a, MK17, MN11c, NTR15, NNAS11a, NUH12, Oru17, OSA13, OSS10, PC12, PT11, PMA17, PA12, PL10b, QTW^{+}18, QL10, RMM11, RG18, RA11a, RTV17, Ros12, SK11a, Sah17, SR15, SR17b, SPL19, Sav10, SIL19, SSR11, She11, SC13, SEM13, SKPW14, SMBY10, Śmi11, SK19, TM12, TZ15, TCM18, TWLYÖ10, TBP19, TTX^{+}16, ÜM16, WZXL11, WmN13, WY16, WTYZ17, Wei10b, Wu10, WZ11b, qXjH11, XHM14, XHA13, XCS18, YTD^{+}18]. generalized [Yan10d, Yan10b, YJ19, YT12, Ye15, YT11, YYLW19, YN16, YASK10, Yu11a, Yu17, Zak18a, ZS11a, ZSZ17, ZFZQ10, Zha11d, Zha15a, ZY17a, ZXZ17, ZDZY17, ZL18, ZZ18a, ZL12a, ZHJD13, ZLZG11, ZWG11, ZJZ18, yZjH12, ZYW17, GD16, SD12b, Yan18c]. generalized-PSS [Yan18c]. Generalized/Extended [GD16]. Generalizing [MT11]. generate [LRCG16, Tso13]. generated [ÇHK11, DSWB18, FLDZ12, MvS18, OVV^{+}16, Seg19, TTT10, Tia11, YGS^{+}16]. Generating [AM14b, DMP18, Ma19, OSS10]. generation [BTEM19, BH14, FF15, GC19a, HTL10, KYW^{+}18, KH18, MM18a, RSL^{+}18, RKA^{+}18, Sin16, SW10, VMFF18, ZN18, ZD12]. generation/absorption [MM18a]. generations [SLMZ12]. generator [Çan11b, DGGBTRJF12, FSCG11, HMP^{+}15]. generators [PPD10]. generic [GK11a, MBH16]. genesis [AAP12]. genetic [AAP12, BMM12a, BD11d, Bog10, cFpClC13, HDHL11, KS10b, SBKS12, SKK12, XDL12, YC10a]. Genocchi [Öza11, OSS10]. genus [WZF16]. Geo [AM11, AM11, LXYT11]. Geo/G/1 [LXYT11]. Geo/Geo/c [AM11]. geodesics [HWW13]. geodetic [CHM^{+}10]. geological [CM10a]. geomechanics [AKSW19, KLK15]. Geometric [AGPR19, CS11a, FGPP17, Hai10, JYK16, RKD18, RC17a, WL13a, WY16, WKBR18, ZWY19]. geometrically [KVJB15, KV17a, KV17b]. geometries [BV10, IQR16, KVJB15, KV17a, KV17b, TPHD18]. Geometry [XY16, DNS18, HN18, HGJP19, Hot13, HWH^{+}15, WJWW12, ZS18, WKBR18]. GeoPDEs [Váz16]. geostrophic [FIW13, WY18b, Zha18e]. gesture [LNKU12]. GF [YXYH10]. Gilbert [BM19, LT13, Pu12]. Gini [Dra10]. Ginzburg [FSZ18, RSB14, WH18, XaZH19, ZY15a, ZZL19]. Girkmann [DFGG13]. girth [ZYS10]. given [CS11a, LZ10, LHH10, SK11c, WZ10, ZYS10]. Givens [Abd18a, BMY13]. giving [EZM12]. glaciology [MLGY16]. GLHSS [FZ14, Mia18]. Global [AKL18a, AET19, CM13b, CQ13, CL15, DWY15, Din13, DYQM14, FAHZ17, FT10, FH16, HKJ14, HY13, JPK18, JLL19, KG14, Kar10b, KKD13, LZ11b, Li13, Li16a, LMY19, LZ14, LL16c, LX15a, Ma18a, Ma18b, NT17, RHMA18, RC17b, RA19, Sun10, TTM19, VBW10a, VBW10b, Wan16b, WW19a, XWY11, XMW10, Xu11b, XGH17a, XZ19, YX16, YLG10, YWT18, YZS18, YQ18, YZ19, ZFZQ10, yZjM10, ZZ18b, ZSL19, ZMWH18, ZZ10b, Zho16, AJS19, AG10a, BM10a, BG15, BS11b, BSN13, BPF13, CCKP15, CANK11, DG10b, FSTN18, FST19, FQLC18, GYTD12, GYH11, HL18a, HSMT19, HXL11, JLF17, JKS18, JKS19, Kim14, LC10b, Li12a, LY12b, LW15, Li18a, LPY16, MN10a, Mor13, PG10, SI10, SKK12, WHD14, WQNF12, WLYX13, XJYL17, YW10, Ye15, YN16, YS19b, ZH18a, ZFY^{+}19, ZLW19]. global-model [BSN13]. Gmeiner [SLW11]. GMLS [DM16]. GMRES [MZES12, WNTW19, YLC16, ZG19]. GMSSOR [ZZ14]. GNFS [YXYH10]. GNU [FRAK15, JRA^{+}18, RAW^{+}16]. Goal [DPM15, SWOF19, BDM^{+}19b, BWZ16]. Goal-oriented [DPM15, SWOF19, BDM^{+}19b, BWZ16]. Godunov [Ban13, CS18, Ma10a]. Godunov-type [CS18]. Gołab [CBKR10].

Gold [VBW10a, VBW10b]. golden [RG11]. goodwill [GM18b]. Gordon [CT18, DM16, DL14, GML17a, MS18, Wan19a, CM11a, FYYT11, HHGA19, KTDT17, LZWC16, MPfTX18, SH11, WV16, WL17a, Wei10b, YMDZ10]. Gould [KAS11a]. governed [Akm15, GS19, GNP14, HLZ15, Li10b, YSB15, ZG16, ZZX16]. governing [CW14, CWDL17, GGO16, JKMS12, Moh15, ZLTY16]. Gower [Li14, LWN15, YL13, YXWL14]. GPA [Bra13]. GPBiCOR [GHC^{+}15b]. GPGPU [DM19a]. GPGPUs [JK11a]. GPL [PHM^{+}19]. GPU [AVV18, BF16, BKL^{+}19, Bre14, GEZ14, KWAS16, KLK15, KCC^{+}13, KORR10, LWR16, OKTR13a, OKTR13b, OAKR16, WFL11a, WCZ^{+}19, YLC16, YCW^{+}14, YW14]. GPU-accelerated [KLK15]. GPU-aware [AVV18]. GPU-based [WFL11a]. GPUs [BPM14, LYC15, SKG^{+}11]. grade [CLTA11, FKKS11, IBSS11, KRCJ11, uHS12, MC11, MM18a, RHMA18, SL12, YJ19, Zak18a]. graded [AB18, BN16, BZT16, CFLX18, FXCC18, Kia16, KM13, Osm18, XFL16]. graded-index [Osm18]. Gradient [AD16, HM18b, LSW16, WYD10, AM18, Bra16, Don10b, DFS14, FLH10, FL11a, HP13, HR15, HM14, KLK15, KM14, KMS19, KHF^{+}19, LD11a, LZD17, LGC^{+}17, LNW19, LSD10, LLY13, LFZ19b, MF18b, Mai16, SS17, Tan17, TNP17, WZ17b, XM15, XBHN16, ZT13, ZDL11, Zha15a, ZY17a, ZY17b, ZZLB18, ZFY^{+}19, Zha19a, ZLZG11, ZXW13]. Gradient-based [HM18b, WYD10, Zha15a, ZY17b, Zha19a]. gradient-enhanced [KMS19]. gradient-type [FLH10, FL11a]. gradients [CNH17, RSB14, SZA^{+}18]. grading [LMMT15]. gradostat [DKG14, MKG13]. grained [JCZZ13, Jum10, KHWK10]. Grammian [CZ17, Men18, YTS^{+}17]. Grammian-type [CZ17]. granular [CBB15, LMPE18, LXF11]. Granulation [MWWL11]. Granulation-based [MWWL11]. graph [BBR10a, FPW^{+}11, KYAA10, KSO16]. graph-based [KSO16]. graphene [FG18, GF16, KHF^{+}19, WL17b]. graphical [YW11a]. graphics [DSK^{+}14, MSZG17, OKTR11]. graphs [AD11b, AH10c, BFG11, CHM^{+}10, CLM11, CP10, CSS10, Dol11, IB10, LZ10, LLL11, LSM10, PB11, WFY17, WS10a, Wu18b, YA11, YKRV11, ZYS10, ZM13]. Grasping [GY13, LL12a, YC12]. grating [LHL12a, XD17]. gravitactic [MMFT^{+}19]. gravitation [Gal11c, Gal11d]. gravitational [LX18]. gravity [GZW^{+}18, RS18a, TG14, Tao18, hYlLL11]. gray [YZWW14]. GrCCA [LXF11]. greatest [VBK13]. green [CSSW12, AH11a, BX10, LG12, RHD18]. grey [WO10]. grid [Amo15, AD19b, BN14a, BWL18, BH14, CCZ18, CR18b, CR19, DWZ13, DBS12, HD16, HCZ16, Hu18, JJH16, KX12, LPLR19, LZZ19, LZ19a, LRH13, LDL^{+}15b, LCHZ19, LSV18, MN10a, Nes10, PGW19, QCLC17, QM19, TCHW19, Von19, WC17, WBZY18, WFZ12, WY19b, ZPGW16, ZD18]. gridless [HH17]. grids [ATZ11, Bac14b, BG15, BE18, CG13, FVZ18, GGO16, HG16, HMWZ16, HLX18, JK10, KNT12, LZJY16, Nes10, Pas14, PPC13, PPC15, RS18b, SRGL13, SKG^{+}11, SKFG11, SW16b, SR18, VFM19, YT13, Zha18c, ZZL18a, ZÖXL^{+}19]. GRLW [Ros12]. Gronwall [AE12b]. Gross [SSC19]. Grossberg [Hua11]. Ground [CT17a, LLP19, Luo19, QHT16, ZTZ15, ZTZ16a, ZT18a, Zho19, CC19, CTC17, GTC18, LSCG16, LT18, Li18b, LLT16a, LLT16b, WZ15, Wu18a, XLT17, YW19, ZHS^{+}19]. groundwater [BFS15, DA18b, ID16, LHL15]. Group [SS16a, eMA18, Bre14, CYZZ18, Chu10, Deu10, GZ10, HPY10, Ji14, KKK16, LL12d, MR17, OMS10, PKTH13, Ray18, RÖ10, RRP16, UKA15, XKH10]. group-harmonic [LL12d]. groups [BSZ16, LLL11, MCL15, SA11b, WC10a]. growing [GDF12, LSM11, TH13]. growth [Amo18, AM12c, BGL^{+}15, CYP16, CZY11, CSU13, CAC14, EE18, FAHZ17, HLZM16, HMSC10, KRP12, KFTT13, LSCG16, Özu15, QMW18, SC16, SZP^{+}11, WZ15, XL10, ZH11, ZHC17, ZMH16, ZMWH18, Zho19, dSSV17]. Grünwald [SKTH11]. Grüss [FM12a, Gav12, SS11b, TD10b, Yan11b]. Guaranteed [KT15, LMR19, dPRVRB13, CFRS10, CLN^{+}19, DESV18]. guest [JXZ^{+}10]. guest-transparent [JXZ^{+}10]. guide [WD13]. guided [HB19, RT11]. guides [KSF14]. Gurland [Mor11b]. Guseman [PRR10]. Guseman-type [PRR10]. gyroscopic [CM19a]. gyrotactic [UABK16].

H [KKAM11, BC17, Jic10]. h-adaptive [BC17]. H-G-A [Jic10]. H.264 [CUK12]. H.264/AVC [CUK12]. Haar [AuIK11, AuIA17, HGHA19, Lep11b, Oru19, uIAH10]. habitat [AEG11]. habitats [BL18]. Hadamard [ADK10, Krn12, ÖAK11, THD11, THH12]. Hadamard-type [ADK10, ÖAK11, THD11, THH12]. haematopoietic [Ben12]. Hagen [RSV11]. Hahn [HRMS12]. half [BMTV12, DVMS13, GF19, Kim11, KMT10, LLY10, LZY11, MR10, Pov12a, RHD18, Sha14, SZ11, Ye19, ZQ11b]. half-line [Kim11, SZ11]. half-linear [KMT10, LLY10, LZY11, MR10, ZQ11b]. half-plane [Pov12a, RHD18]. half-space [DVMS13]. halftoning [ZN18]. Hall [EUTS18, FKKS11, FSZ17, Pal13, Wen18a, Ye15]. Hall-magnetohydrodynamics [Ye15]. Halpern [Nil11, SCC12b]. Halpern-type [Nil11]. halved [AJS19]. Hamel [EG10, MGB^{+}11]. Hami [XC13].

Hamilton [CC11, Kim17, SD12a, Tre18, YYYH19, ZQ14a]. Hamiltonian [Yaz11, AZ10, CT11b, CTA12, Gal11b, HDHW11, HM17b, HM17c, MZ10, MZ11, MP19a, TJ10, TZZ11, TM17, TCM18, Xu11a, XCS18, YAS^{+}11, Yu11a, ZZ18c, dVLV18]. hammer [CRXL15, MZM19]. Hammerstein [CS12, LD11a, WCD10]. hamstring [GZ10]. hand [AT18a, KM12]. handed [LNKU12]. handshake [KK13b]. hanging [BS17, CDG15]. Hankel [AHV10, dS16]. Hankel-norm [AHV10]. Hanusse [MN17]. Hanusse-type [MN17]. haptic [BJLZ12]. haptotaxis [LMY19, SNSK19, ZW16b]. Haragus [AK16]. hard [CM18c, GT15, GT16, Kaw15, Par17a, Par17b]. hard-to-borrow [CM18c]. Hardness [RdSSS11]. Hardy [Gao12, LZJ12, NP12, She18a, SC19b, WS17, ZYY10]. Harmonic [BNR10, CLN^{+}19, KNIF13, AKS10, AM14b, BJQS18, BP11a, BAO^{+}12, BDHR18, Cao19, Ciz12, DLS14, Dos18, GVSP12, GH16, LWBW13, LCK13, LL12d, LDHH13, MMT18, MDG19, PD17, SBM13, VAK^{+}19, Wan10b, YY11b, ZLS13]. harmonics [AR10c, BOY12]. harmony [WQNF12]. Hartmann [RA18, STS19]. harvesting [BP13, LW15, ZZ16a, ZL10a]. Hash [CJP12, CJP15, HKKK13, SPLHCB14]. Hash-based [CJP12, CJP15, SPLHCB14]. Hausdorff [MN10c, SSR11, WCH18]. having [GKLR11, KAG11]. HBV [HY15]. HCRF [Hon12]. HDG [JRA^{+}18, Fu19, GS18, HPS18]. healthcare [KK12]. heart [KEHB18]. Heat [CK15, EA10, BHZJ19, BB15, CL19, CZF10, CR19, CC17, DSM18, DZO^{+}19, DLWW12, DCN^{+}18, Ebr11, EUTS18, FXCC18, GWR^{+}18, GH15, GRS18, GMI11, GD10b, GK18, GHCZ18, HP13, HLSN15, HHM12, HLI14, Ima17, JMB10, JFC14, KMT19, KK14a, KKSK18, KQ11, LY11b, LX12a, Li17b, LZZ11a, MSV18, MC10a, MM18a, MS10c, NPR10, OP14, Pal13, PCO16, PvdM13, PK19, QXG13, QWL19, RA18, RJGS^{+}19, Ryl15a, Sah11, SL12, Sam19, SAIZ15, SWL19, SRRP18, TL18, TY16, UKA15, Wan19b, WCCS15, Wu18b, XWH16, YF10, YL18a, Ye19, ZZHF12, ZCH14, ZYT^{+}16, ZSQ^{+}18, ZJB19, ZSLZ19, ZWG11, ZKW15, ZG14, aZ18]. heat- [SAIZ15]. heat-conductive [Ye19]. heat-like [GH15]. heated [CLTA11, LWC13, NPR10, YJ19, Zak18a]. heating [PKD19, QXG13, SSPL10, XWH16]. heatline [RSL^{+}18]. heaving [DTR19]. heavy [WL12a]. Heisenberg [GTZ19, LM19a, LTJ^{+}16, MGS^{+}14]. helices [Lu12]. helicopter [SG10a]. Helmholtz [ASY^{+}11, BKMT14, BKMT17, BML11, BM18c, CGH14, CC15, CF16a, CLL19, CTZ17, Cia12, DPM15, DL19, Du12, Erg19, EM14, GVSP12, Li19b, LZ16c, LCLL18, QCG15, RHD18, RCM11, SJL^{+}19, SS14b, WX18a, ZD15]. Helmholtz-like [Cia12]. Helmholtz-type [BML11, LCLL18]. helper [RWZ13]. hematocrit [SR10b]. hematopoiesis [LC10b]. hemiregular [MYZ12]. hemirings [ADA11, DSA09, DSA10, Jun10, MYZ12, SM11, YJZ11]. hemivariational [BHM19, HLZ15, Mat19, MP19d, SMH18, XHH^{+}19]. hemivariational-variational [Mat19]. hemodynamics [WSC16]. Hencky [RSH18]. Henon [LYZ11]. herbivore [Li11b, ZL14a]. herbivores [Li11b]. HerEOS [ZHV19]. Hermite [Ali15, BCB11, CRRS11, CDD12, KAS11a, Krn12, ÖAK11, THD11, ZL13, ZQ14a, ZT18b]. Hermitian [BMJ19, BR18, Cao19, FZ14, FWZ16, GHC^{+}15b, LGG12, LM18b, MJB18, Mia18, WLD13, WWB13, WLM13, XCS18, XCZQ16, YqS16, ZLZ11, ZYW15, ZS16]. Hermitian/skew [ZS16]. Hermitian/skew-Hermitian [ZS16]. Heronian [Ciz12]. Hertz [SH12a]. Hertzian [HZL17]. Herz [Gao12].

Hessenberg [SB19, GH13a, HSMT19]. Hessian [DP15, HT18]. Heston [CGK14, HZ16, SS19, ZZ19]. heteroclinic [WX18b]. heterogeneous [CF16a, CELY18, CFS17, GM18a, Kei13, LYM12, LZP^{+}19, MM19, SSS16, SWOF19, SCKH10, SCBCB^{+}13, SCBCB^{+}17, SM19, SLM16, VMP15, Vel15, ZLZ18]. Heuristic [HCL12a, ÁBÁPM11, DM12a, LML11, LTX^{+}13, LCC12, LMRS10, OMS10]. Heuristics [AA10b, KX12]. hexagonal [ATZ11, KST10, SAU11]. hexahedral [WZH18]. HID [CCKY12]. Hidden [ML19a, SYO12]. hierarchic [SBKS12, SRDD17]. Hierarchical [KLCD16, Pul16, ZDL11, BSN13, CCK12, KÖC^{+}18, PZJ^{+}16, PZAR19, TAPA^{+}17, WZH18, Zbo19]. hierarchy [Elb11, Gep16, NN13, YGS^{+}16, Yu11a, Yu11b]. Higgs [JKB11]. High [ABT19, AK19, CO15, CT17b, CS14a, CN16b, GM18a, Git14, HZL17, HG18b, KYW^{+}18, KPS17, Kim18, KK14c, Lee16, LH16, LHL14a, LZLL18, MKHC11, NHH13, RQ18, SZL^{+}17, WBN18, YWH14, ZWW13, dVDR18, AEH18, ADGS18, AS11b, BGIN13, BS16b, BF16, BZKR15, BDHR18, BD19, BT14, BS18b, CKSL^{+}14, CF16a, CR13, CFdM^{+}18, CL17b, CM14, CT17c, DWZ13, DM18, DH11b, DAM14, DZW^{+}15, DT17, DHGF17, DLF^{+}11, Fu19, FKDN15, Gha18, GGR19, GGO16, HS11a, HBK^{+}19, KM15, KS15a, LZJY16, MBH16, MDG19, MNPD15, MS15, MM19, MSZG17, MR15, NH15, OTiSY16, OSZP13, ODAZ15, Pes13, PD17, RES10, Res16, RTRR18, RR19, RR14, SSHH^{+}18, ST15, SCA14, SS14b, SGK18, SSP13, SK14b, SSL14, VMFF18, WV14, WLZ^{+}18a, WZC^{+}19a, Wei17, XY15, XZ18, YZM^{+}19, YTZ17, YWW^{+}12, YSS11a, Yüz11, ZCY16]. high [ZC11a, ZW11d]. High-accuracy [LH16]. high-conductivity [BD19]. high-dimensional [CR13, VMFF18]. high-frequency [BT14, GM18a]. high-level [MBH16]. High-order [ABT19, AK19, CS14a, CN16b, GM18a, Git14, KPS17, Lee16, LHL14a, LZLL18, MKHC11, NHH13, YWH14, dVDR18, AEH18, ADGS18, AS11b, CM14, DM18, DH11b, DT17, DHGF17, Fu19, HS11a, KM15, LZJY16, MDG19, MNPD15, MM19, NH15, SSP13, WV14, WZC^{+}19a, XZ18, YZM^{+}19, YTZ17, YSS11a, Yüz11, ZCY16]. high-performance [GGO16, ODAZ15]. High-quality [KYW^{+}18]. Higher [ABRL18, ATUC15, BLS18, Cvi11, DZW16, DN18b, Dem10a, DGLS19, DGK10, DR12, GTG11, GKM11, GKS10, HKKK13, KJK18, KS15b, MT12, MP19b, Tre18, Yas12, BPM14, BS12b, BC10, BPX11, BS17, CM19b, DYH11, GJ10, GM11, GH10, GH12b, GH13b, HL11b, Jaw13, Kar10b, Kig10, KHF^{+}19, KB13, LF11c, Li12a, LKCN19, LMLB19, LYSZ19, ÖKJR19, PC17, QMW18, RZ17, SD10b, SDH15, Sea11, Sea15, Sed13, WWA11, WW11c, XG10, ZSAN18, ZR18c, ZA15, ZTH11, ZW16a, ZMH16]. higher-dimensional [QMW18, ZW16a]. Higher-order [Cvi11, DN18b, DR12, GKS10, KJK18, KS15b, MT12, Tre18, BC10, BS17, CM19b, GJ10, HL11b, LF11c, LKCN19, LMLB19, LYSZ19, SD10b, Sea15, WWA11, WW11c, ZSAN18, ZA15, ZTH11]. highlight [BB15]. Highly [BZ18, DBS12, EPP18, HMY15, LL12a, LZZC12, SSSB11, SLM16]. hijacker [WD13]. Hilbert [AD11a, AD10a, Arq18, Buo11, CGY11, CZN11, Deu10, Dra11a, Dra11b, FLDZ12, HKK^{+}16, JH10, KK13a, KN12, Krn12, Mai10, QL10, SH10, Sle13]. Hilbert-type [JH10]. Hilfer [AM14a, FKeT12]. Hilliard [AEH18, BN14b, BMH19, DFJS10, GR19a, GGT14, LS19a, LJSK13, LCK17, LQMW18, LJK^{+}19, LCWZ18]. Hindmarsh [CZMZ11]. Hirota [DTYZ18, FW18, GZZ^{+}16, LTSW16, LWW19a, MF11]. histogram [QXLL11]. history [CXMO19, SMH18, XHH^{+}19]. history-dependent [CXMO19, SMH18, XHH^{+}19]. hitting [MDVM17, TTT10]. HIV [MGY11, Xu11b, YWL^{+}11a, ZLC11b]. HIV-1 [Xu11b]. HJB [CC11, CWDL17, XX17]. HMM [KSPP11]. HMM-based [KSPP11]. Hoc [SSM12, CHY12, DA12, LLWZ11, BK18, KG14]. Hodge [FRZ15]. HOFEIM [DKRY15, YRDR18]. Hohenberg [KKAM11, Lee17, PP14]. Hohlov [SRM11b]. hold [RNB11]. Hölder [QH11, Wan19b]. Hölderian [Sca11]. hole [DC15, NKM16, QWGJ15]. hole-filling [DC15]. holes [FGPP17, TM19]. Holling [Bae10, CP16a, LG10, Li13, YL13, Zha18b]. Holling-type [YL13]. hollow [KM13]. Holm [LXY19]. holomorphic [Ibr11, MKS13]. Homann [SL12]. homeomorphism [ÇT12, HY10]. homoclinic [CT11b, CTA12, DTYZ18, YTD^{+}18, ZZ18c]. homogeneous [Aba10b, ABB17, AYH17, CLB14, GKLR11, GDM13, GLL14, JYL16, KNIF13, LC12a, LRV13, MEAMHHV18, RSDR11, San11, Wan16a, Zha18e, Zha18f, ZZL15]. homogenisation [HD14a]. Homogenization [Sza15, Tim14, TRL19, AGDP19, BM18b, CO15, HJD15, Li10b, NXHN14, OC14, RTL19, Sal16, SCBCB^{+}13, SCBCB^{+}17, VZM14]. homologies [CEF^{+}13]. homology [DW18a, MW10]. homomorphism [ÇT12, HY10]. homotopic [XSLS11]. Homotopy [AIIZ10, GS11c, IIHI10, JGSS10, JNJ^{+}11, KW11, MGB^{+}11, SISH12, ZTH11, AO10a, AÖ10b, BD11c, BGPP11, BE11, BG11, CH11a, Eba14, EG10, FH11, Gup11, HKI12, HM10, IIHuI10, KLTS11, Liu16d, MH11, Naw11, PP10, RMA10, RY10, SAIZ15, XXG10, Yus09, ZYSY17, ZMA10]. honoring [BCHS18]. Hood [Che14]. hop [AZB13, DA12, YDK^{+}12]. Hopf [CW15a, CSW11a, GLL14, LW15, LWN15, LPY16, SZ14, TZ13, ZLC11b]. Hopfield [DHGF17, KS10b]. Hopper [KAS11a]. Horadam [YT12]. horizon [BMM12a, MT12]. horizontal [BTEM19, LZ11b, PC12, SZA^{+}18, ST18, WCH13]. host [LZ11b, WZCC10, YZMZ16]. hosts [NNR14]. hot [Che12c]. hot-fit [Che12c]. Householder [SB19]. HPM [RY11]. HPM-Padé [RY11]. HSL [TT12]. HSS [CCG18, LHY18, LZ16b, LZ17, SS18a, XY17, XWY17, Yan17]. HT [SCA14]. HT-PEM [SCA14]. HTLV [EOM11]. HTLV-I [EOM11]. Huber [Zha19b]. Hull [SS19, CHM^{+}10]. human [EOM11, GW12b, LNP^{+}12, MCQ11]. Humbert [CH11b]. humoral [TTM19]. hunting [OMS10]. Hurst [WLW^{+}11]. Hurwitz [SJPS11]. Huxley [KSMN11, MDRRV11]. HVDC [ES10]. Hybrid [Buo11, CB11a, CGHY11, FSM19, HMSC10, JYYL16, LGC^{+}17, MZC17, MBH11, NEB14, SWL19, She12, TL12, VB10a, WQNF12, Ahm10, AEF15, AD16, AuIK11, CW10a, CAC14, DCG^{+}12, DWI^{+}12, FLdS14, GACMO13, HJD15, HY11, JGK13, JW11, KLK15, KMS15, LLJK10, Liu15b, MFSL19, OP14, PZZ18, PS18, QHW11, SK11a, SH10, uIAH10, SJC14, WL13a, WHD14, WN18, qXjH11, XY14, YHZY11, YW14, ZHJZ11, ZSHL11, ZLY^{+}13]. hybrid-based [GACMO13]. Hybridizable [FHS18]. hybridization [CDL17]. Hybridized [JRA^{+}18, GHC^{+}15b, JPS14]. hydraulic [CN13, GHT^{+}15]. hydro [EEBM10, WXYW11]. hydro-magnetic [EEBM10]. Hydrodynamic [WXYW11, WCH13, CM19b, LMLB19, QY13, WLA18, WSM^{+}19]. Hydrodynamics [ATO19, WRW^{+}19, BMS19, FM11, GR10, ILV^{+}19, SPP18, ZHV19, ZPGW16]. hydrogen [YZ15]. hydrological [YMSL11]. hydromagnetic [MC11, NPR10]. hydrostatic [RRAK19]. Hyers [Cie11]. hygro [KHF^{+}19]. hyper [AS11a, JKK10]. hyper-MV-algebras [JKK10]. hyper-MV-deductive [JKK10]. hyperalgebras [AN11b]. Hyperbolic [Col14, AEDL14, Ana11a, Bag17, BHM12, BDR19, BM12b, BKZ17, CW17, DHY19, EE18, FM19, FJP18, GY15, HA16a, JJC11, Kig10, Kim15, KPS17, Kim18, LYS12a, LYS12b, LZ14, LMZ17, MD15, RR18, SW17, SPST18, TT14, ÜM16, Vac18, WC17, WW19b, ZWG11, ZY17d, ZT18b]. hyperbolic-parabolic [FJP18]. hyperbolic/trigonometric [BM12b]. hyperbolicity [BRS11]. hyperchaos [Zhe11]. hyperchaotic [Che11c, MHM11]. hypercube [KP10b]. hypercubes [MW11, Yan10a]. hyperelastic [WY15]. hyperelasticity [YWH14, ZRE16, ZR18c]. hypergeometric [ASMEE11a, CH11b, Mos10, OCNG12, RRC11, Swa10]. hypergroupoids [HM12, SKST10]. hypergroups [CJMS10, LF11a, SYL10, TM10]. hyperinterpolation [WWW14]. hyperlattices [HX11]. hypermodules [YZXW10]. hypernear [DZK10]. hyperparameter [SW10]. hyperrings [Dav10, MZLF10]. hypersingular [FMPR15, HT12a, LTT13]. hyperstructure [YKD11]. hypersubstitutions [PL10b]. hypervector [AD11c]. hypoxia [SSM^{+}17]. hysteresis [CP15a, ZKW15].

i-Math [WNC12]. IB [GC19b]. ICA [WYLZ10]. ICNC [LZL^{+}13, YWL^{+}11b]. icosahedral [Mic17]. Ideal [ŞGY11, BSS18, Deb12, MME10, QaY18, XZ17, yYqWqZC13]. ideals [ADA11, Ali11a, CKN11, DZY11, DSA09, DSA10, Jun10, JLP10b, JLK11, JK11c, MZJ11, SDM10, SK10a, SJN10a, SJN10b, SM11, Yan11a]. Idempotent [Lit13]. Identification [AZB13, FPB17, HP13, TKHL18, ZL19a, CD12, CZF10, CJP12, CJP15, ES10, FWFL11, Hon12, KP19c, LKK12, LL12d, Liu16d, MMOJ14, MCL^{+}13, MAS11, SR16, SS18b, SKH12, TL12, WLXG18, WRY18, WV15, XCZQ16, YD12, YS19b, ZKWW17]. identifiers [WZCC10]. identifying [LY11b, SBM10, TMLF19, WQRZ14]. Identities [Kim10, GLM^{+}11, LS11d, SBB^{+}18, Wan10b]. identity [HPY10, NCL13, ÖŞ11, yYqWqZC13]. identity-based [HPY10, NCL13, yYqWqZC13]. idle [JL12]. IEEE [TXZ^{+}10]. IgA [LMR19]. ignition [JS12b]. II [THD11, CHBTD14, CFB11, DNZ^{+}13, Far11, GCDG17, HAESLB14, Ikh11, LLW10, Li13, MBKK10, RAW^{+}16, Tao18, ZH15a, Zho11, ZQ14a, ZZC13b]. IIA [MV10]. III [DHMU16, HLTL17, JRA^{+}18, THfL17, WL11b, YL13, Zho12]. Il’ichev [AK16]. Ill [HT16b, CH11a, CDY11, Jaw13, LP10, Liu14, MZES12, ZT16a]. Ill-conditioned [HT16b]. ill-conditioning [Jaw13]. ill-posed [CH11a, CDY11, LP10, Liu14, MZES12, ZT16a]. illiquid [CJPB10, GW15b]. ILP [Bog10]. ILW [GZW^{+}18]. Image [JMNZ19, PGDL18, QXLL11, ALLH11, ALLQ13, Bai19, Bar17, BSL11, BLS17, CLW11, CX16, DM19b, FJWW16, FQLC18, GV18, HLT12, HCL11, JRB15, JZR15, KJK18, KSZ18, LGHR16, LKK12, LLG^{+}11, LXZ13, LK11, xLlFwWL12, LZ17, LW11c, LW10, Liu15b, Liu16e, LZG19, LW12c, LZS12, LYSZ19, MAN^{+}15, dCMdSGTdC^{+}16, MSTB17, PZJ^{+}16, PZAR19, RSS16, SGZW18, SLZ11, TSB16, TZXP11, TMZ^{+}15, WRW13, WPH11, XJYL17, YZWW14, yYsZyYL13, ZT16a, ZN18, ZCY16, ZY17c, ZLW18, ZFY^{+}19, Zha19b, ZWJ^{+}11, jZsQdLmG19]. imagery [UKAL10]. images [CCSZ14, HT18, LTLL12, LHL14a, LHLH15, UCK16, WHD14, YZWW14, YZM^{+}19, Yua18]. imaging [Amb12, CHM18, CGJ^{+}14, ILP^{+}11, Par17a, Par17b, RMB^{+}14, XY14]. Immersed [ZYT^{+}16, ZHL12, ABRL18, BCCZ18, BCH^{+}18, BCD^{+}16, CAC14, DL16, FIVM17, FIMV18, FES^{+}19, GLZ18, HZ10, HR19, HYS^{+}14, HLSN15, KVR11, QCLC17, RHC15, SMF17, WCZ^{+}19, YNS^{+}14, ZHW^{+}11, ZZL15]. immersed-boundary [FIVM17, FIMV18]. immersogeometric [KEHB18]. immiscible [KRBS18, TDM13]. immobile [ZLPM13]. immune [BCC14, BF11, BD11d, DDLM13, FLWJ11, HDHL11, KLY16, TL12, yXpYxZT11, XDL12, YX16, ZLY^{+}13]. immunity [AGU14, TTM19, XMW10]. Impact [DD13, PÁAP^{+}15, RB19, CS14b, GK19, Li11b, MJ14, RS13, ZZ17]. impacted [RSP18]. impaired [WNC12]. Impedance [dCMdSGTdC^{+}16, GKS17, PFBL10, RHD18]. imperfect [BBL19, CCDL10, DRK11, SG11a, TTMJ12]. imperfections [RKD18]. Imperialist [AIA13]. impinging [YL18a, ZCSG13]. implant [DRT^{+}15]. Implementation [CDG15, PTH^{+}16, QKR19, ABB17, BDHR18, CPP15, DSK^{+}14, DFW^{+}18, EMR10, FHZ10, HKS19a, Kan15, KO13, KB13, MJWD19, MPGW19, MSZG17, MWY17, OKTR13a, OKTR13b, OAKR16, PC12, PLR15, RPTD10, SCSF19, Váz16, WZM^{+}16, WHC12, ZH15a, ZYZC18, ZKBE16]. implementational [TCM15]. implementations [KSS13, MZES12, SKG^{+}11]. implication [NUH12]. implicational [TL10a]. implications [Liu11b, hYlLL11]. implicative [JK11c]. Implicit [fDxZ11, DGM14, HH17, Mur08, SHM13, XZZ16, AD15, ADZ19, AVV18, CB19a, CY19, CLA19, DFM15, DGT18, FHZ13, hGzS15, HMF^{+}19, HZ10, He16, KBAF18, KYO10, MS15, PZZ18, PT11, PRS18, SD15a, SL16c, WH16, WZKY12, WLZ18b, XZ10, YH19]. Implicit-explicit [DGM14, XZZ16, PRS18]. implicit-spectral [ADZ19]. implicit-symplectic [DGT18]. implies [And12]. Importance [LLZ11, WLT13b, BM13b, LL13, WC11b]. important [Kir10a]. imprecise [FSM19]. improve [ZLG^{+}10]. Improved [Aba10a, HWXC16, KKC^{+}10, Mor11a, Pop14, SR15, SLMZ12, SR10a, tWqLzGkP11, WLA18, ACC18, BCCZ18, Chu11c, FNZ13, FL11a, GLR13, qGpWhL11, HYS^{+}14, JW19b, KP10b, LCC12, LLCG16, LZG13, SC19a, SLZ11, ST12, THZ^{+}11, WXF10, WHS17, Wen18a, WY18b, XJLX10, XDL12, YL18a, Yao16, Ye17b, Yua18, Zak18a, ZDLC14, ZÖXL^{+}19, ZY11]. Improvement [DH11b, Gav12, XW18, BG11, sCYhX18, JH10, Yus09]. improvements [Akm15, BCB11, LZJ12]. Improving [AGH^{+}15, DRK10, HP19a, KSPP11]. impulse [IMD11, Jia12, LHLH15]. impulses [Bai11a, BV11, Che12b, HF10, Hua10a]. Impulsive [BB10a, GJ12, GW12c, HO10, Wan11, AK10, AW11, AKRT14, ANN10, Bae10, BKT11, DB11, Guo12, JL11a, LCW10, LHW11, LB12, LH10a, LHF11, LCZ11, LW12a, LS12c, LW12b, ÖZ11, RM17, SI10, TNF11, TB10, WS11a, WAZ11a, WZF12a, WYN12, WZF12b, XZC12, XH11a, YHZY11, YZY10, ZD11]. impulsively [YZAX10]. In-plane [ATH18]. in-series [dPRVRB13]. in/on [FG18]. Incentive [CQLX11, RZL11, LLX^{+}10]. incidence [GCR^{+}18, Sun10, SW19, TTM19, XMW10, XGH17b, XXH18, YX16, YZ18]. incident [HCT12, LLL13, LFC16]. inclined [RBB12, Ran15]. include [CRG16, MT11]. included [RKA^{+}18]. including [CDG16, RSL^{+}18]. Inclusion [ASMEE11a, AX11, KC19, QL10, Zha11a]. inclusions [AA11, ANP11, ABCR10, AKLS19, BHM12, DN10, HO10, JL11a, KP19b, LKS10, LH10a, LMZ17, MW10, Oua12, PRR18, SWL19, WD10, ZM18]. inclusions/voids [SWL19]. Incomplete [Raf12, RTB14, RT11, YT18a]. incompressible [BHK16, BS18a, BMSS18, CCR16, CYL17, DH18, DLQ16, Ers16, FZ17, FST19, GLP18, GHT^{+}15, GD10b, GZN19, HHS^{+}17, HS18, HLB14, KLRW12, LVF^{+}16, LR15, LLML15, LR17b, LM18a, LH19, LL16c, Liu16c, Liu18a, Ma18a, MGN^{+}16, NN13, PLKC16, QaY18, QM19, RA19, SL18a, TPHD18, TH19, Wan14, Wu11a, WSCW16, XSYL19, YY11a, Yan18c, YH19, YS19a, Ye19, YT13, YZS18, YT18b, YNS^{+}14, ZRE16, ZZX^{+}14a, ZZX^{+}14b, ZT15, ZSW15, ZYZ^{+}16, ZYZ^{+}17, ZZ18b, ZDB19]. Incorporating [CCY10, KNZ19, SZA^{+}18]. incorporation [SGK18]. increase [Mor13]. increasing [ARK13, Bor11, ÇT12, HY10]. Incremental [FSZS18, Res16, DLWW12]. incubation [Aki17, NNR14]. indefinite [HH11, LT15b, MT11, SS18a]. independent [AGD10, CLW11, Dol11, Hon12, KX12, Liu10, PZ11b, XZL10]. index [DT11, KLTS11, Kir10a, LH10b, LHH10, MBKK10, Osm18]. indexing [IC12]. indicated [AM15]. indicator [BPC17, CJN19]. indicators [LMR19]. indices [DT11, DYX11, KMT10]. Indirect [DGGBTRJF12, JPCY13, MCP13, MR19, QMW18, ZWZ16]. indirectly [ABLS15]. individual [GCG12]. induced [AGT19, ABV11, Amo18, BG10b, MLY18, WCH13, ZZLB18, ZTSC16, ZY13]. inductance [KSMT11]. induction [DGGBTRJF12, LLSW10, LC12b, SST19]. industrial [Hot13, LZL^{+}18]. industry [Cha13]. inequalities [AD10a, AD11a, ADK10, Ana10, Ana11b, BX10, BD11c, BXKZ11, BX14, BS15c, CGY10a, CKW13, CW15b, CAP10, CT10a, mCfX10, DV10, DSZ18, DWS19, Dra11b, FM12a, GW12a, Gav12, GALO18, HmZ11, HKT11, HR15, HX10, HLZ15, HN10, JJC11, JPS10, Li10d, Li11a, LG12, Liu10, LSZ11, LZ12b, MPZ11, Mai16, MP19d, NDT11, NZ14, NM11c, NUNAS11, NN12, eOS18, ÖAK11, QH11, Rah11b, SSO10, Sar10, SS11b, Set12, SMH18, SJPS11, TZZ11, TÇA12, THD11, THH12, TD10b, Von11, Wan10a, qXjH11, XD10, XJLX10, XHH^{+}19, Yan11b, ZCW15, ZHJ11, Zhu10]. inequality [AE12b, BLS18, BHM19, CS10a, CB11a, CB11b, CB11c, CY14a, CLF10, CW14, Ciz12, CTS19, Dia17, JH10, Jic10, Krn12, KLY16, LGG12, LB12, LZJ12, MJ10, MG11, QZM17b, QM19, RGdSRLAJ10, She11, She12, SLW11, WLS10, XZL10, Yan10b, YLK10, ZLZ11, ZCY16, ZYY10, YYYH19]. inertia [HLB14, PSD^{+}13]. inertial [BKL14]. Inexact [TAPA^{+}17, BX14, DBH^{+}14, HM17d, Śmi11, THD19, qXjH11]. inextensible [ZRE16]. inf [Kim14]. inf-sup [Kim14]. infection [Aki17, EOM11, MGY11, TTM19, WZXS17, WYL19, Xu11b, XGH17a, XGH17b, ZLC11b]. infectious [ASB12, HY16, LJYS18]. inference [LWC13, WSL10]. Inferential [WWH12]. infinite [BR12a, BB10b, CANK11, CLC16, De 10, DQ10, FAIV10, LSZ16, LZ11e, LLL12, MT12, Mor13, Ols10, Pir11, PK19, She11, She12, TMO13, YY10a]. Infinitely [Bao16, CTA12, CF16b, CT18, DL14, WW15, ZD11, ZTZ16b, FLZ14a, KK10a, KK11b]. infinitesimal [VĆC10, VĆV11]. infinity [JFC14, WZ11b]. inflammatory [RdSRL19]. inflationary [TTMJ12]. inflow [HMF^{+}19, HKP17]. inflow-implicit [HMF^{+}19]. Influence [AM18, RMK19, SJS^{+}11, WvDRG19, ZGD13, AM13b, GC19a, RKD18, WL11b, YLF19, YLL^{+}14, ZG18b]. influenza [KKL16]. Information [Gal10a, Gal11c, Gal11d, LEP11a, SPCS13, FQLC18, Gal11b, HpD11, JPCY13, KAK^{+}12, OO12, Ögü13, SBA10, XDL12, Zha11b, ZWJ^{+}11]. infrared [HLT12]. inherent [FT15, SCvdV^{+}19]. inhomogeneity [LLL16]. Inhomogeneous [GG18, ZR18a, ABK10, CH17, GLL14, HL11c, LHTL19, NB11, YTC^{+}18, ZY17d]. initial [Aba10b, AM14a, AHOP18, Arq18, BKL14, CL19, CM18a, CP16b, CCD10, CLCF14, CN11, DN10, DBEE11, Fer12, GCE18, Goo10, GR13b, GF19, HL18a, HGSL18, JZL18, KBAF18, KFYW11, KB15, KP19c, KAS11b, Lan12, LZ14, LZ19a, LH10a, Luc10, OC10, RMS10, RNB11, RS15, VAB12, WC10b, WG11, WQRZ14, WHLC11, Wu16, XW19a, YC11, YK17, Zha18a, ZY15c]. initial-boundary [GCE18, KFYW11]. initial-boundary-value [GR13b, Luc10]. initialization [MGW11, Pan11]. initiation [NJV13]. initiatives [San11]. injected [DVM12]. injection [CCY18, DSWB18, SRRP18, ZOZZ12]. injective [Gon13]. injective-overdetermined [Gon13]. injector [CI18]. inlet [ZCSG13]. Inner [Pop13, FWFL11]. Innovation [SC13]. innovative [KLH^{+}12]. input [GAVOF11, LLZ11, LL13, LCYC12, LXYT11, ODAZ15]. inputs [CZY13, LLH14, Tol12]. insect [ADL12, DD13]. insight [Bay19]. inspection [RMS12, TTMJ12]. inspired [KBGC12, OMS10]. instabilities [GD10a, Kia19]. Instability [FHA16, SPP18, GTZ19, HS12, KPR13, Kup10, LK13, Lin14, NGG12, Pal12, SJL^{+}19]. installment [Ciu11]. instance [xLlFwWL12]. insulated [RWZ13]. insulation [RJGS^{+}19]. Integer [Koj10, dS15, AR17, CKR10, EKS10, HTGSH13, YXYH10, xYsHjL11]. integer-order [EKS10]. Integrability [GTZ19, LV11]. integrable [Fur13, MZ10, MCF18, Yu11b, Zed10]. Integrablization [AHJM18]. Integral [LHF11, MPZ11, RS12a, RS14b, ZSS10, ABJ11, AD10b, AKRT14, AS10b, AX11, AS11c, BMRA10, BM11a, BR12a, BN16, BGPP11, BE12, BT14, BBD10, CDY11, CR19, CHS18, CDS15, CMT12, Dar11, DV10, EMQ18, EKE18, FMS19, FM12a, Fra11, FMPR15, Gen11a, Gep16, GHMN16, GN11, HY10, HKI12, HC14, IB10, JL11b, Kar17, KKBR19, KMRN12, KQ11, Las10, LZL16, Li10d, LL11, Li11a, LTT13, LH16, LSZ16, LS12c, Luk11, LT11, MB11, MN11a, MTM11, MKR12, MT11, MBHV10, MSG11, Mor10b, MCN10, Ols10, PB11, RMS10, Rah11b, RA12, SNH10, SNMA12, SYG11, Saj12, Sal11, SPL19, SBB^{+}18, UBF11, Waz11b, WLL12, YK18, Yan12b, Yan18b, ZL19a, ZZX16, ZLL18]. integral-preserving [AX11]. integrals [BCJ19, CS10a, CR13, CWQJ12, HZL17, HD14b, KRGS^{+}10, LZH16, LS16, LR17a, PCS13, PAT12, SD15a, Set12, SK11c, WHS12]. Integrated [BP18, BO10, Chu12b, HH18b, HY11, Kia16, QZF19, TMDTTC16, WFL11b, YXYH10]. Integrating [LZL^{+}18, HK10, LYM12, PTP14, WC11b, WLYX13]. Integration [KX12, SVY16, AuIK11, BMA11, BPM14, CP10, CN11, DGLS19, Hai10, HD19, KMS10, KLH^{+}12, KB13, RC18, SBvdV13, SL18a, uIAH10, TS14, WLA18, WZC^{+}19b]. integrator [AHP^{+}14, ALMLM14, HA18b, IW18, JHW15, MS18, Tam16, UMLF13, YT18a]. integrators [CL16b, Gar13]. integrity [JXZ^{+}10]. integro [Abb11, AdAS11, AKRT14, BMTV12, BCC14, BKY10, DB11, FID14, GY15, HLZD11, HLY12b, KV18, LH10a, LHF11, LXP11, LS12b, MBH11, Naw11, QX19, RÖ10, SD10b, SS14a, TS11a, TAS11, Tom11, WC10b, WG11, WAW15, YSS11a, YSS11b, ZA10, ZTH11, dSAC11]. integro-differential [Abb11, AdAS11, AKRT14, BMTV12, BCC14, BKY10, DB11, FID14, GY15, HLZD11, HLY12b, LH10a, LHF11, LXP11, LS12b, MBH11, Naw11, QX19, RÖ10, SS14a, TS11a, TAS11, Tom11, WC10b, WG11, WAW15, YSS11a, YSS11b, ZA10, ZTH11, dSAC11]. integro-differential-difference [SD10b]. integrodifferential [BK11a, BK12a, WZWX11, WW14b]. Intel [CCHG17]. intelligence [DCG^{+}12, HSZ15]. Intelligent [XCM12, Che11e, HDHL11, YC12]. intense [Bis14]. intensity [CKL18, Dos18, FQLC18, GC19b, KL16b, VC12, WCH18]. intensive [AGK15]. inter [PKTH13, SCKH10, YWK^{+}10]. inter-domain [SCKH10, YWK^{+}10]. inter-group [PKTH13]. Interacting [THfL17, BKMT16, MV12, MW13, PP10, PLKCC13, VDV13]. Interaction [FW18, WWD18, BHK16, BM18a, GZD18, HB19, HZ10, HTM18, HLY16, HYC18, IMS19, KEHB18, KSD^{+}19, LX16, LWW19b, LWW19a, MYZ18, MZM19, MGN^{+}16, MMT18, MUB^{+}16, RMY19, RHC15, TTG16, VHPVNXW18, WLA18, XYXZ16, ZYT^{+}16, ZBFC19]. Interactions [ML19a, GSY10, HTY^{+}19, LNKU12, LGH^{+}11, Yoo17]. interactive [AM14b, DMPV10, HWH^{+}15, ZLHF19]. interconnected [LLC13, THY^{+}10]. interconnecting [Hof18]. interdependent [BKK11]. interest [CXZ15a, HZ18, LW11d]. Interface [ZHL12, ABRL18, AYH17, BCH^{+}18, BMS19, BGR14, BBL19, BPZ19, CPT15, CAC14, Du12, FGB19, GLZ18, HR19, Hes14, HSJ15, HQ19, JM15, JCF19, JK12, KEHB18, MA16, PBK19, QCLC17, ST15, ySGL^{+}10, WC15, WHS17, WCSW18, WHS18, WLZ18b, XW14, Yan15]. interface-aware [BMS19]. interface-level [CAC14]. interfaces [HT16a, LSZ16, NZ16, RTRR18, WD13]. interference [MLY18]. interferon [LNP^{+}12]. interferon-gamma [LNP^{+}12]. interharmonic [LL12d]. Interior [HKP17, ZG14, CLL19, GNP14, Ram11, SDM10, SCGW18, ZCW15, ZD15]. interlacing [MCP13]. intermediate [ESBR10, GDZ11, ZRC11, sHC11]. internal [BP18, JPCY13, RSL^{+}18, Rus16]. internally [GK18]. international [Che11b, WD12, BCHS18]. Internet [WZCC10, XCM12, gZnZpZbD12]. interplay [CFB11, KP19b]. interpolant [GW15a, ST16]. interpolants [BGIN13, BL14, CDP16]. interpolated [OKTR13a]. Interpolating [BD16, AD15, LS10a, Wu11b]. interpolation [BCB11, CD14, Cia16, DA18b, GWZ11, HH10b, LW19a, LXZ13, LZH16, LL19b, PU10, QKR19, RFK16, Seg19, SW16b, Tha19a, VP11, Von19, WXF10, Wei17, Zha18d]. interpolations [Bic11]. interpretation [Del13, XA13]. interruption [CTD10]. intersection [BE18]. Intersections [CYM13]. Interspecies [CP16a]. Interval [AAZ10, AD11b, JTC^{+}10a, Jia11, LS12c, dCMdSGTdC^{+}16, SK11c, WYG12, Wan13a, BCB11, BKM11, Boy10, FLLF10, HDT11, Ikh11, JY11, LZYW13, LZ11e, Lit13, LZ12c, WL13a, YLY^{+}09, YW11a]. Interval-valued [AD11b, JTC^{+}10a, Jia11, WYG12, Wan13a, FLLF10, JY11, WL13a, YLY^{+}09]. intervals [Bai12, BEAA11, Li12a, LLL12]. intra [MYZ12, PKTH13, SRS11]. intra-clustering [SRS11]. intra-group [PKTH13]. intracellular [Xu11b, XBHN16]. intrinsic [COR18, GV11b, GL16]. Introducing [HMP^{+}15, KNZ19]. Introduction [FGHZ13, FGHZ14]. intruders [LX12b]. intrusion [SLXC11, TXZ^{+}10]. Intuitionistic [AB11, SR12, ADA11, Deb12, JTC^{+}10a, Jia11, qJhY12, Li10a, MME10, PCK13, WYG12, Wan13a, WL13a, YLJ12, hYxLL10, ZW11c]. invariance [ADGL14, Mah14]. Invariant [Li12a, MN17, PDHL12, RS18a, Ray18, RÖ10, YÖ10, BG13, GH13a, GK13, Guo12, JC12, KKVS19, KT18a, PP12]. invariants [KYAA10, KK10c]. invasion [Amo15, GL17a, HLZM16, HBS^{+}10, SNSK19, ZZF18]. invasive [BFF^{+}11, SD15b]. Inventory [BMM12a, CB10, Chu12b, TTMJ12, WW10c]. Inver [tWqLzGkP11]. Inver-over [tWqLzGkP11]. Inverse [BCHS18, BM13b, DNS15, FRSC16, GHC15a, MMOJ14, XC16, AZ15, APRM11, AKMS18, AS15a, BT10, BF16, CZF10, Col18, CHS19, CGJ^{+}14, Ehr18, FIM18, GR19b, GSR14, GN19, HA18a, Haj18c, HSZ15, JKB11, JW18a, JL15a, JWX^{+}13, KK19a, LP10, LZ11c, LLY18b, Lin10b, Liu16a, Liu17, LW18a, LW18b, LRBA15, LGVS19, MA17, MM13, MPGW19, ORR16, Par17a, SYZ19, SC13, SBM10, SS16c, Smo17, SSP13, TCM18, WCB13, WmN13, WCW13, WW14b, XL15, YY14, YF10, YS19b, dBD17]. Inverse-free [FRSC16, WmN13]. Inverse-Quartic [BM13b]. inverses [Ji14, JL15b, JBBL17, LZ19b, Pet14, WSH12]. Inversion [Amb12, DGLS19, Hal13, JSEM13, AM13c, GSZ11, Res16, YLDL11, Zha19a]. inversion-free [Zha19a]. inversions [CLJ11]. inverted [ASMM11, BBBM16, Pan17]. inverting [EMR10]. investigate [MMS^{+}18]. Investigating [DLT12]. Investigation [HL18b, ST18, TR14, WTSS10, CB19b, CLL11, GC19b, KDG11, LLL13, LFC16, LGH^{+}11, LCWZ18, MR17, PN16, PvdM13, SHM13, SZ12a]. Investigations [WFZ12, MM18b]. investment [LZF18]. invex [Ant10, GJ10, NP12]. invexity [Ant14, PN10, ZLLF12]. inviscid [HG16, Ye17b]. involutory [XC17]. involving [AH10b, AHO16, AS10a, AS11c, ASMEE11b, BT10, BDGS13, CHZ19, CS10b, DSL18, EAA10, FKeT12, Fur13, Gen11a, GALO18, GJ10, HMY15, LS17a, LJX12, LYZ17, LS11d, MV11, NUH12, OCNG12, PC11b, RSS10, She18b, SRM11b, TB10, VSI12, XZL10, Yan11d, Zha11a, ZZ16b, ZLG18, ZT16b, ZY15c, ZG18a, Zhu10, dSSV17]. Ion [ABK10, GSY10, CD19, EK16, JLD19, Sea14, Sea16, YLH12]. ion-acoustic [Sea14, Sea16]. ionic [Tim14, TRL19]. Ionkin [BCK11]. ions [ABK10]. iontophoresis [FdOP17]. IOT [gZnZpZbD12, ST12]. IP [CHY12]. iPad [GMAM12]. IPDG [BS14a]. IPDG/TDG [BS14a]. irrational [LX10a]. irreducible [CKK^{+}10]. irregular [ACC18, FID14, Fis18a, GHL18, LFAL19, QLT^{+}18, WY18a, WH11b]. irreversible [CYM13]. irrotational [FLZ14a]. Isentropic [PBM19]. Ishikawa [KK11b, Buo11, KK10a, XLK11]. islanding [LCW12]. isoflux [HHM12]. Isogeometric [HT16a, KVJB15, KP19a, MRS15, Moo18, PXXZ16, BMAR18, GR15, Hof18, HWH^{+}15, KLCD16, KV17a, KV17b, KT15, KKJ15, LZL^{+}18, LMMT15, NXHN14, ST19, Tan18, Váz16, WZH18, WKP^{+}14]. isolated [BG10c]. isolation [SG11a]. isomorphism [MZLF10]. isoparametric [GLLC19, Rua19]. isosceles [AG11]. isothermal [GFZ16, MLL16, PFDG17, SS16a]. Isotropic [GKS17, BM13a, FR15, FWW14, KNIF13, KQ11, LX15b, XBHN16, ZR18c, ZY17d]. isotropy [ADGG13]. ISPH [RTRR18, TY13]. Issue [LK15, LZL^{+}13, SPCS13, VBW10a, YWL^{+}11b, ZFLM19, CBM10, PHWM10, VBW10b]. issues [AJJAD^{+}10, HSS^{+}12, PTH^{+}16]. Itô [Ade16, KP13, MKR12, Pal12, VB10b]. Itô-type [KP13]. item [DRK10]. items [Chu11c, TTMJ12]. itemsets [FKC12]. iterated [HKK^{+}16, Lam13]. Iteration [AMA14, YMDZ10, Aba10b, AP10, AM14a, BMM12b, BPG10, BKY10, CM16a, CMR17, CCG18, CW10b, CWDL17, CS11b, DNS15, Els10, FZ14, FWZ16, FZ19, Far11, Gen10, Gen11b, GH15, HSWZ11, Ikh11, KA10a, KFYW11, Kim11, KT11b, LM17, LHY18, LM18b, LZ17, LHL12b, Lu11, Mia18, NLA19, Naw11, RWTW19, SEY12, SS10, SCC12b, TTZW18, WLD13, WLM13, WMW13, cW11, XY17, XWY17, XWY18, YXS10, YC11, YX11b, Yan17, ZS11a, ZM16a, ZYW15, ZY10b]. iterations [BMJ19, Mah10b, MM18c, Nil11, RH15, ZHY14]. Iterative [BKMT17, BOT14, CGY10b, DL10, DLD10, DSM18, HMF^{+}19, JRB15, Lan12, NUNAS11, RSM17, RH15, SR16, Wan16a, WFDW10b, XHM14, AJAR18, ABCR10, BJQS18, BGGCGRSP16, BR18, BPG10, Buo11, CsH10, CGY10a, CGY11, CC11, CM16b, CX16, CM17, DGB10b, DN18a, DM19a, FIVM17, FIMV18, FL14, GHC^{+}15b, HH16, HYS^{+}14, HWXC16, HWXC18a, HM18b, KS10a, KK10a, KK11b, KM14, KA10b, KYR11a, KYR11b, KA11, KV18, LP10, Lee11a, LP12, LY15, LFZ19a, LBW11, Liu15a, MBHV10, MO14, NHIN16, Ned12, NKA18, OTiSY16, PÁAP^{+}15, PPD10, Pet14, RS15, She12, SS17, SZ12b, TCM18, TA11, WYD10, WW11c, WKS13, XM15, XLK11, YD12, ZT16a, ZDL11, ZY17b, ZZL19, Zha19a, ZL16, ZGL14]. iteratively [WWB13]. iterators [Mor13]. Ito [LWW19a, MYZ18, Zed10]. IV [ZFLM18, Bae10, CP16a]. IVFSs [WW11b]. IVIFSs [WW11b]. Ivo [DFG^{+}18]. IVPs [AR10a, Eba11].

J [McN12, HSMG12, SSR11]. Jacobi [CC11, YYYH19, CLMM18, DHY19, ES18, Elb11, LPML19, MJB18, PCS13, SD12a, TÇA11, Tre18, VT11, ZQ14a]. Jacobi-like [CLMM18]. Jacobi-type [PCS13]. Jacobian [CW18, ZKBE16]. Janous [SLW11]. January [Ano18-54, Ano18-67, Ano19-50, Ano19-59]. Jarratt [ATUC15]. Jarratt-like [ATUC15]. Jaulent [YGS^{+}16]. Jeffery [EG10, MGB^{+}11]. Jeffrey [Eba14]. Jeffreys [BB18, GF19]. Jenkins [WYD10]. Jensen [AD11a, AD10a, GKK11, KC11, MRR11]. jet [GUK13, YL18a]. jets [CY14b, ZCSG13]. Jimbo [YTS^{+}17, BM18a, HTM18, LDL11, LYC^{+}19, LSX13, Man18, YM17, ZBFC19].

Job [YWW^{+}12, AK12, DM12a]. job-shop [AK12, DM12a]. jobs [MMR10, MMR11, WSS10]. Joint [PZAR19, BJLZ12, CLMM18, DH10c, KGJ11, YLC12]. Jordan [BFSO12, Zhu10]. Joseph [HQ19]. July [Ano18-53, Ano18-61, Ano19-49]. jump [AYH17, BM19, DNS16, HMV18, Lee14, LZ12c, RR14, TC16, VB10b, WZ17a, YKA18, ZZL15]. jump-diffusion [DNS16, HMV18, TC16, WZ17a, YKA18]. jump-diffusions [Lee14]. jumping [BV11, ST18]. jumps [BS10b, Che11b, CNP14]. June [Ano18-69, Ano18-70, Ano19-58]. justification [CRA19]. juvenile [ADY12]. juvenile-adult [ADY12].

Kac [DZ18]. Kadomtsev [LXY19, WF17, WF18, Ade17, CTSX16, CZ15, Gep16, HTY^{+}19, IMS19, JPB11, KTK17, LZL19, LWL11, LMZK16, MCF18, QTW^{+}18, SIL19, SMYK19, TTX^{+}16, WTYZ17, WW18a, WL16, WTLS18, YMHL18, YTL^{+}18]. Kansa [CYS10, CHL18, FJB19, YY14]. Kansa-type [YY14]. Kantorovich [SNH10, WFL11a]. Kantorovich-quadrature [SNH10]. Karman [Par18, KPK18a, Par15]. Karst [LWL14, LS17b]. Kashiwara [YTS^{+}17]. Kassem [EM19]. Kaup [DYH11, ZA15]. Kautz [SD12b]. Kawarada [Pad18]. KD [KK17]. KdV [Her19, YYC11, GABC16, zLYLQ19, MDW11, RG18, SR15, Sea11, Sea15, Üns18]. KdV-type [Üns18]. KdV-Zakharov [Her19]. Keller [FL13b, FHA16]. keloid [BF11]. keloid-immune [BF11]. keloids [FBL11]. Kelvin [BN14a, BP19, SJL^{+}19]. Kerberos [SCKH10]. Kernel [PB13, Arq18, CC15, DA18c, DGLU18, DH16, KTDT17, LLSW10, MKS13, SG16b, TS11b, WSCL11, WLHZ14]. kernel-based [DGLU18, DH16, KTDT17]. Kernels [YG11, DM15, MFSL19, QX19, WBN18, Wu18b]. Kernighan [tWqLzGkP11]. key [CM13a, CSCM13, GLM^{+}11, HPC12, Lee11b, LD11a, NCL13, PKTH13, RWZ13, VBK13, WH10, WMZW11]. key-insulated [RWZ13]. keys [LW10, RWZ13]. KGS [DM16]. Khokhlov [KKK14]. kind [BM11a, BS15c, CWQJ12, CHS18, DWS19, MBHV10, MSG11, OCNG12, SNMA12, WXF10, Waz11b, XX17, Yüz12b, pZ10]. kinds [Çak11a, MZJ11, WFZ12]. kinetic [Bab14, BF11, BMS13, BKK12, Bre14, EAEH18, HL18b, KNZ19, KGM11, LF11b, LBZL11, LZPZ19, LZZ19, LZB15, PZZ18, PLMS14, WSCW16, ZYZ^{+}16, ZYZC18, ZQ14b]. kinetics [BG13, CLM12, CYM13, Lóp19, SL18b]. King [CM11b, DK12]. King-type [DK12]. kink [LD11b, WF18, bZM17, ZBFC19, qZM17a]. Kirchhoff [CTG17, CC19, CO19, CHZ19, DLS18a, DN18a, DSL18, HY18, HL18a, HGSL18, HCF16, KLL19, KJ11, LSCG16, LL16a, Li18b, LRZ18, LYZ17, LLP19, Lü14, MM13, MPLR18, QHT16, She18a, SC19b, TNV19, WW15, Wu18a, XZR16, XC18, ZJ17, ZT16b, ZWW13]. Kirchhoff-type [CC19, CO19, DSL18, HCF16, LL16a, Lü14, MPLR18, SC19b, WW15, ZWW13]. Kit [GMAM12]. KKL [RSL^{+}18]. KKM [HS11b]. Klein [CM11a, CT18, DM16, DL14, GML17a, HHGA19, MS18, WV16, Wan19a, YMDZ10]. KN [Yu11a]. knapsack [Lee11b, RFP11, RF12, WH10, You11]. KNC [CCHG17]. KNL [CCHG17]. knot [JRSZ12, LD13b]. knots [SLKK19]. Knowledge [YWL^{+}11b, HMP^{+}15, LML11, LTX^{+}13, NB17]. knowledge-based [LTX^{+}13]. knowledge-reduction [LML11]. known [Kup11, fLcJ10, LZ12c]. Koch [PU10]. Kolmogorov [He16, HLvS18, KO11a, KO11b, MP19a, TZG10, XLZW11]. Konopelchenko [Abd18b, KKK16, LCW19, Ray17, WZMY18]. Kontorovich [Ehr18].

Korteweg [Asl10, Dem10b, ES18, KT18a, LGZ19, YYC11]. KP [AGT19, CZ17, HLLM19, HYC18, LWW19b, MZM18, Sah17, TZ15, YS16b, ZCZ17, ZCT18, qZM17a]. KP-like [YS16b]. KPI [CZ18]. Krasner [Dav10]. Krätzel [KRGS^{+}10]. Kriging [DA16, LLW15]. Krylov [BJRF19, BS11b, BJS15, BOT14, CL16b, DGR18, Liu15a, SL18b, WmN13]. Kuhn [GN12]. Kukles [PL10a]. Kummer [MB10a]. Kutta [KMS10, KAS11b, MBJ16, FHZ13, MVKK14, SLL17, TFS11, Tsi11, ZGD14]. Kuznetsov [Che19a, DGA18, DS18a, Her19, KK19c, Kuo18, zLZ19, Ray18, GABC16, PMA17, SR15, Sea14, Sea16, ZTZJ14]. Kwon [KSJ12].

L [KKT13]. L-stable [KKT13]. L1 [ZJZ18]. labels [BSN13]. Lacunary [Deb12, KS11]. laden [AJS14, GLW13]. lag [AS11b, BCFQ19, FDXW11, KM13, LZCL18]. Lagrange [Agr10, BS16b, BGR14, EO15, KEHB18, Kim15, KPS17, LP10, PBM19, PBK19, WLW17, YÇG12, YDL11]. Lagrangian [AP19a, BMS19, BG15, BGM19b, CM19b, CMT12, ILV^{+}19, KSD^{+}19, LMLB19, LMRS10, Liu16e, PW10, RÖ10, SWS19, TC16, WSM^{+}19]. Lagrangians [DBH^{+}14]. Laguerre [MB10a, MD10, WJ11]. lakes [LZ19c]. Lamé [BG19, BTB18, ZS18]. laminae [Sza15]. Laminar [GK18, HLB14, GD10b, LWC13]. laminated [MG15, SSH15, ZTW^{+}19]. laminates [DD19b]. lamps [CCKY12]. Lanczos [Haj18a, YXYH10]. Landau [LT13, AET19, BM19, FSZ18, Pu12, RSB14, WH18, XaZH19, ZY15a, ZZL19, ZY17d]. Landweber [HXX19]. Lane [AR10a, VA10]. language [GA10]. Laplace [BG10c, Chu18, DM12b, DBS12, Dos18, Hol11, LW18a, MZC17, MZM19, NLA19, OZF19, ORR16, Par11b, PL17, Rah11a, WCH18, Zbo19, ZY15c]. Laplacian [DZ21, ABB17, AEO15, BLS18, CTP10, CDM12, CTA12, CHZ19, CP15c, CP16c, CCP19, DG13a, DLS18b, GALO18, GKM11, GJX18, HY18, HGW18, HT19, Ili10, KSZ18, KLL19, LCW10, LD18, LSM10, LL10b, LJX12, NT17, PC11b, SD18, SYY13, Tar17, Tod18, Tru19, VLJH18, WZH10, WZF16, WW19c, Wu16, YY10b, Yan11c, YZ19, ZYS10, Zha18a, ZZ15b, ZG18a]. Laplacian-based [GJX18]. Large [FG18, GW12b, HBK^{+}19, JKN10, Sag10, ZXW13, ADD^{+}15, BJS15, CCY10, CZY11, CChL14, DGOZ13, FGHZ14, FLH10, FL11a, GH15, GOT19, HIS19, HAESLB14, JW10, KLK15, KM14, KLTS11, LC12a, LHY11, LCQF19, LLH10, LFZ19b, MY10, OPDC12, SKFG11, SM14, Sun11, THY^{+}10, WLYX13, XY11, YLS12, YWW^{+}12, Zha14, Zha15b, ZD15]. Large-eddy [HBK^{+}19, Sag10]. large-scale [BJS15, CCY10, CChL14, DGOZ13, FLH10, LC12a, LHY11, LCQF19, LFZ19b, SM14, THY^{+}10, YWW^{+}12, Zha15b]. Laser [BGM19a, WRW^{+}19, BG10b, KÖC^{+}18, QXG13, SSIP19, XWH16, ZTY^{+}19, ZTSC16, FSM19]. laser-induced [ZTSC16]. Lasota [HGW11, WYN12]. latency [DYQM14]. lateness [AA10b]. Latent [xLlFwWL12]. Lateral [WCQ^{+}19]. Lattice [ASFM15, AJS14, BV10, BFS15, CYP16, ÇA10a, CZY11, CCSZ14, DSK^{+}14, FVVS16, GLW13, GZR^{+}13, GUK13, GR10, GD10b, Hei10b, HLSN16, JMST11, LWR16, LCS15, MJ14, OF16, PD11, SK14a, Sag10, SG11b, SKFG11, SZP^{+}11, Tro13, WWLL13, YLY12, YB13, Yan13, YKKS10, ZCSG13, ZY15a, ZY15b, AM10b, AML^{+}14, ABR^{+}14, ADGG13, ADGL14, AYY12, BRFH16, BR13a, Bre14, CWH13, CCY18, CZ11b, CLL11, DL16, Del13, DZW^{+}15, DLT10, DL11, Dub13, EAAS18, EAMA19, FIVM17, FIMV18, FES^{+}19, GWR^{+}18, GSPK15, GGGR17, GSZ11, HP19a, HZ10, HBS^{+}10, HBK^{+}19, HG16, HS18, HMSC10, HBE14, HBE15, HTL10, HSC17, HYS^{+}14, HLSN15, HWyL11, Ima17, JK10, JLD19, JQG14, KNZ19, KTH13, KVR11, Kup11, KGM11, LD13a, LDS10, LZZC12, LSC17, LLML10, LYN11, LG13, LGL^{+}14, LBL14, LW11d, MMHGM17, MPGW19, OKTR11]. lattice [OKTR13b, OVV^{+}16, OAY11, OHK^{+}19, PLR15, PTH^{+}16, PGW19, QWGJ15, RSL^{+}18, Rhe10, RJGS^{+}19, SSHH^{+}18, ST15, SPT17, SKG^{+}11, ST14, ST18, Ste16, SKTC15, SKTC19, TR14, TTT10, THC^{+}18, TDM13, UMLF13, WXYW11, WH11a, WAG^{+}14, WSC16, WLA18, WMP^{+}19, Wan19a, WL13b, WTSS10, WHQ^{+}18, WZHW13, WGY^{+}18, WCZ^{+}19, XZ11, YY11a, Yan10c, YSW16, YCW^{+}14, YNS^{+}14, YCS19, ZZX^{+}14a, ZZX^{+}14b, ZSW15, ZYZ^{+}16, ZYZ^{+}17, ZYZC18, ZHW^{+}18, Zha13, ZLC^{+}11a, ZM16b, ZHW^{+}11, ZZC13a, ZZC13b, ZPGW16, DJD18, GGGR17, HLNZ19, LCS18, QKR19, RKA^{+}18]. lattice-BGK [PGW19]. Lattice-Boltzmann [BV10, Hei10b, YKKS10, CLL11, WMP^{+}19]. Lattice-Boltzmann-based [Sag10]. lattice-gas [HBS^{+}10]. Lattice-valued [JMST11]. lattices [BM18b, BFS15, KD12, LZH12, LXF11, Med12, PB12, yYqWqZC13]. LAVIR [Bar17]. law [AM14a, BHZJ19, BDS17, BPZ19, DKM17, Gal11c, KAA19, MA10b, SMBY10, Sun11, TS11b, VPR11, WLXZ18, Zha19c, Zho13]. laws [Abd18b, AP19a, BKZ17, CG13, Col14, EK16, FM19, Kim15, KPS17, Kim18, LHD18, LXY19, LDG19, Ma19, MW13, MK17, RR18, Ray17, RS18a, Ray18, YÖ10, Zha17b, ZKBE16]. Lax [BKZ17, LM19a, MR17]. Layer [SNEP19, AGT19, Bog11, Boy16, Hak14, HTL10, HW19b, KS12a, KYY12, LYL12, zLYLQ19, MM18b, RBB12, Ran15, REHA11, SYO12, Tur10, WO18, Ye19, Zha19c]. layer-based [LYL12]. layered [CJ18b, QWL19, SRV10, WL17b]. layers [AGH^{+}15, FP19, GZR^{+}13, HN18, KAG11, Kia16, LTT13, LLG10, LZ18c, Ma10a, MLG17, SCA14]. layout [HGJP19]. LBE [Kup10, Kup14]. LBM [GOT19, GC19b, JK11a, JGK13, KORR10, YLF19]. LCP [YGR11]. LDG [Bac14a]. lead [BMM12a]. lead-time [BMM12a]. leader [yN11]. leaders [yN11]. leading [GS15a]. leakage [BKR11, Cha13]. leaks [MEAMHHV18]. leapfrog [HCLL18]. learned [LYSZ19]. Learning [LS10d, PB13, Ahn12, FHH13, HCHH12, KSPP11, Lan12, Lee11c, LLD10, xLlFwWL12, LZC13, LLW15, MKA^{+}10, SSM12, SLYY13, TNT12, Wan08, WSS10, WW10a, WWW11a, WWW11b, WWH12, WLYX13, YH11a, YX11c]. leasing [ZZXY12]. Least [BDGG14, GCDG17, HA18a, HK17, KDU15, LR14, RCM11, TCM18, BQ15, BQ17, BJQS18, BJ19, BG14, BC17, CCKP15, CT18, DM16, Dia17, GB16, HS13, Hes14, HSJ15, Hes18, HM17c, Ku18, LC10a, LGG12, LL16b, LZ17, LJJ17, LJLY19, LY11d, Liu14, MLGY16, PFDG17, PGF18, PATA11, RMB^{+}14, TJQS13, TBP19, WLT13b, YC10b, YD12, ZLZ11, ZMLZ16, ZY17a, ZWLZ18, ZL13]. Least-Squares [KDU15, RCM11, HK17, BG14, BC17, CCKP15, GB16, HS13, Ku18, LC10a, LZ17, LY11d, MLGY16, PFDG17, PGF18]. leaves [LHH10]. Lebedev [Ehr18]. Lebesgue [Li17b]. Leffler [GMP18, Kir10a, LCP10]. Left [Gal11e, Mok11, NM11a, RTB14]. left-looking [RTB14]. Legendre [AR10a, BKY10, CCJV11, CNV14, Cvi10, DB12, ED11a, JYF^{+}11, KK14a, MKL11, MBH11, RG18, SD12a, SXM11, VAB12, ZLJ^{+}18, ZC17]. Legendre-spectral [AR10a]. length [CL12a, KRP12, LXYT11, MP11b, MP16]. lengths [Don10b, Gao11, TMMASG10]. Lengyel [MAB19]. Lengyel-Epstein [MAB19]. Leon [ES17, LM18a, Liu18a, LZM18, PTZ19, MR17]. Leray [SM10].

Lerch [Cvi10, SJPS11]. Leslie

[Li14, LWN15, YL13, YXWL14, Zha18f]. Letnikov [SKTH11]. level [AMGC19, AVZ15, BDF16, BS17, CCY10, Chu12b, CAC14, DLQ16, FLLF10, FIW13, HMF^{+}19, He16, JZE^{+}18, KWAS16, LHL12b, MBH16, MM18d, MS15, PÁAP^{+}15, PZAR19, ST15, SPT17, SHH16, TS16, WCLD18, ZT15, ZLS13]. level-set [BDF16, HMF^{+}19]. leveled [IC12]. leveling [qGpWhL11]. levels [CZ10, ZGD13]. Levenberg [HM19]. Lévy [Li18c, LPP15]. Lévy-type [LPP15]. LFSRs [QGGL13]. Li [YLH12]. Li-ion [YLH12]. liability [LS11c]. LIBOR [LSV18]. lid [PD11, SK14a, SSPL10, ZZC13a, ZZC13b]. lid-driven [PD11, SK14a, SSPL10, ZZC13a, ZZC13b]. Lidstone [Yan10d]. Lie [Abd18b, eMA18, Ade17, BSZ16, CYZZ18, KKK16, KT18b, KK19b, KK19c, LXY19, Ray17, Ray18, RÖ10, RRP16, SR17a, Sin16, UKA15, WLXZ18]. Lie-group [CYZZ18]. Life [Nab19, CL12a, GM18b, KAK^{+}12, LYJ15, MSH10]. lifetime [BEAA11]. Lifshitz [AET19, BM19, LT13, Pu12, ZY17d]. lifting [Pan11]. light [HCT12, HLY16]. like [ATUC15, AG10a, BJLZ12, BMAR18, BM18a, BT15, BA16, CNR10, CLMM18, Cia12, DT16, DD16, DRS11, GH15, GC19b, HM15, HM19, HN10, KSG11, LSD10, LMZK16, QL10, RKSA18, RWW18, RGdSRLAJ10, SAIZ15, SMYK19, Von11, WKG10, Wei14, YS16b, YK17, ZDZY17, ZBFC19, Zha19b]. likelihood [LD11a]. Limit [CD10, AZ10, Bab14, GOGYL^{+}11, HP10, KLTS11, LFJ11, LSJ12, MS18, Pit12, RSH18, TDXQ18, ZYZ^{+}17, ZZG19]. limitation [ZY13]. limitation-induced [ZY13]. limitations [CDG15]. limited [Gos10]. limiter [VFM19]. limiters [HLL^{+}15, LvS15]. limiting [KdLK19, RAW^{+}16]. limits [Gal10b]. Lin [Gup11, tWqLzGkP11]. Linbox [YXYH10]. line [Amb12, AM13c, ANN10, BMTV12, BB15, CQ11, HSD10, Kim11, MCF18, Mai16, MS11a, SZ11, WZ17b, WZC^{+}19b, ZZXY12, ZTZJ14, ZLG^{+}10]. line-search [Mai16]. Linear [ADGG13, CYM13, CDM10, HM17a, MF11, RSV11, ÜM16, vdW14, Abd18a, AJ10, AJ12, AT18a, AD12, BM11a, Bag17, BS15b, BDM19a, BM11b, BDM^{+}19b, BMTV12, BGRV15, BD16, BM13a, BQS16, BGPP11, BG11, BKY10, BCG17, BOT14, BRROP19, BKP11, BPX11, BK12b, CCJ10, CB11d, CM16b, CDL17, CL16b, DDD10, Deb12, DDM^{+}18, DBH^{+}14, Dra10, ED11a, FMS19, FZ19, FLdS14, FPB17, Fia15, Fio14, FL10, FLDZ12, FP18, Gal10d, GY15, Gar13, GGS16, GK11a, GK11b, GR13b, GHM^{+}14, GHC^{+}15b, Guo12, GH12b, GH13b, HKJ14, HH11, Hes18, HVO17, HMWZ16, HWXC18a, HWXC19, IW18, JGSS10, JGK13, JKS12, JJ13, JJ15, JL19, JL20, JC12, JCF19, JBBL17, KP13, KD10, KKLJ11, Koj10, Kun12, KMT10, LZ11a, LLW11, LHL12a, Lee15, LZGZ11, LLC13, LM19b, LZ16c, LLY10, LT15b, LLH11a, LZY11, LLH11b, LRH13, Liu14]. linear [LCW17, LCC13, LHL^{+}14b, LS10d, MWL18, Mah11, MBH16, MGN^{+}16, MZES12, MR10, MM18c, MMH11, MS15, MSG11, MS12b, NHIN16, OBCG19, Odi10, ÖŞ11, Pal12, Pan11, PR11, PJ17, PCM12, Pop14, QZY11, QHT16, RSDR11, RWTW19, Rey12, RS18b, SD10b, SYG11, ŞGY11, Saj14, SBEB10, SS18a, SKPW14, SL16c, SST19, SP12, STDLM19, SJHC14, SS13, Tan18, TWLYÖ10, TAPA^{+}17, Tia11, TTZW18, THY^{+}10, VMO10, VRD11, WW10b, WLD13, WZ17b, WNTW19, WLGL10, WMW13, XY17, XWY17, XWY18, XHM14, XM15, XC13, XC17, Yan17, Yan18c, YGR11, YqS16, Yus09, YSS11a, Yüz11, YSS11b, Yüz12c, ZKR^{+}12, Zbo19, ZZ15a, ZM16a, ZSZ17, ZFZ10a, ZQ11b, ZL14b, ZYW15, ZS16, ZM17a, ZZ10a, ZZL18a, ZWH^{+}19, ZG19, ZM17b]. linearisation [MS12b]. Linearizability [GKLR11]. linearization [TH19]. linearized [AEH18, CX16, HP17, OF16, SW17, WSW11, WCLD18, Yan15]. Linearly [KBAF18, He16, OKTR13a]. LinearOperator [MBH16]. lines [DM15, GZW^{+}18, HWW13, ZR16]. Linguistic [OO12, ABFGZ11, CHT11, OO10, Zha11b]. link [OAKR16]. link-wise [OAKR16]. links [ES10, Kuo16]. Liouville [EM19, jASzZ12, CKLL10, KA13, Liu12, RNB11, SMK18, YY10b]. lipid [KHWK10]. Lipschitz [BH10, JS11, Lad16, SH10]. Lipschitz-continuous [Lad16]. Lipschitzian [THH12]. liquid [GLR13, GFZ16, KSO16, Kup10, Kup11, Kup14, Laz10, LZG13, MMRN12, MJ14, MSW18, NNL13, WO18, YWT18, ZDB19]. liquidity [CCJP11]. lithium [CD19, JLD19]. lithium-ion [CD19, JLD19]. Littlewood [NP12, SC19b]. Liu [AS10a, AS11c, DGB10a, HSMG12]. living [CFB11]. LMI [LZ11a, SMF10, Yan12a]. LMI-based [LZ11a]. load [AR10c, DVMS13, RSH18, WLXG18, XWL18, XCZQ16, ZWL11]. load-carrying [ZWL11]. loaded [FG18]. loading [AB18, KLTS11, PK19]. loads [HL11c, KÖC^{+}18, VLFS12, ZWL11]. loan [CXZ15a, FXC18, LWJ10]. loans [LX10b]. Lobachevsky [COR18]. Local [AuIK17, Akm15, AYH17, DFS11, DZ17, Gon13, HA10, KM15, Via15, WD16, WLYX13, YM13, ZTC14, ARK13, Bac14b, BS16b, BQS16, CCKP15, CQ11, CQ13, CDD12, DPBL16, DA18b, FWZ16, FQLC18, FIM18, GGGR13, GV11b, HN18, HKS19a, HD16, HSC17, HLY16, ID16, KP10a, KBK19, LMR19, Li17b, LG12, LZLL18, LL17, LPP15, LZS12, Pu12, Ryl15b, SYZ19, SDH15, yS10, SND19, Tha19a, TDM13, tWqLzGkP11, WHS18, WZKY12, XJYL17, YY15, YH15, YGS17, YLC18, gZnZpZbD12, ZWMD16, jZsQdLmG19]. local-volatility [LL17, SYZ19]. local-world [gZnZpZbD12]. locality [Hu19, LRTV10]. localization [HVR18, LKK12, zLYmL18, Sza15, YC12]. Localized [CPP15, HYC18, ZcHS18, KSG11, KKLJ11, xLlFwWL12, LWW19b, LWW19a, WL16, YKC11, YCLY15, Yao16, ZZ10b]. Locally [Bar17, Cie13, EDC14, BG15, CCRS17, FMPR15, JW19a, Lad16, Wu18b]. locate [Chu11c]. location [APT11, CGJ^{+}14, KV10, LGVS19, Mok11, MP19d, RG11, SSPL10]. locked [BG10b]. Locking [LMR14, BHM19, RS18b]. Locking-free [LMR14, RS18b]. locomotion [SCC^{+}12a]. LOD [WW16]. Log [FXC18, Boy16, KWPK13, PW18, ZJ12]. log-and-polynomials [Boy16]. log-convex [ZJ12]. log-determinant [PW18]. Log-Stable [FXC18]. logarithmic [AAR11, mCfX10, DZ21, HGW18, LTLL12, NT17, Tru19, YS17, ZH18a, Zho19]. Logarithmically [GLR13, LZG13, WY18b]. logging [MTV13, MPMTV15, dlPBTW13]. logging-while-drilling [MTV13]. logic [HHY^{+}11, PB12, SH12b, Wu10, ZJZ^{+}11]. logical [CL12c]. logics [JTC^{+}10b]. Logistic [SZDO10, Xia11, ABM11, BR13b, FL13b, HLZM16, LMZ18, LW12a, LW12b, PATA11, ZW16a]. logistically [TH13]. logistics [WWH12]. Long [Li19a, Ngo18, NWZ11, PQB^{+}16, FLZ14a, KBA11, LG13, MS12a, NNR14, Ros12, SIL19, SEM13, UKI11, WWD18]. long-range [MS12a]. Long-time [NWZ11]. long-wave [SIL19]. longitudinal [HHM12, Kia18]. looking [Raf12, RTB14]. Loop [KLCD16, AZ10, DLWW12, KW12, LMMF17, THY^{+}10]. loops [EKS10]. loss [WL12a, ZLG^{+}10]. lossy [Che15b, HCLL18, ZR18a]. lost [EMRS12]. lot [CCY10, DRK11, DH10c, WW10c]. lot-sizing [DRK11, DH10c]. Lotka [ACTB19, CS16, CSW11a, GKLR11, JWX14, MG16, Via15]. Love [MM13]. Low [CGHW14, CS14b, LCT12, AEH19, BK14, BD19, CY14b, sCYhX18, CZ11b, CCKY12, Cia16, CJK17, EG18, EHL^{+}14, GSD^{+}19, HP19a, HVO17, LCQL17, NDC^{+}19, PDN19, QCG15, ySW10, WCB13, WWXW19, ZLW18]. Low-complexity [LCT12]. low-contrast [ZLW18]. low-dissipation [NDC^{+}19]. low-frequency [BD19, CCKY12]. low-Mach [sCYhX18]. Low-order [CGHW14, HVO17, LCQL17]. Low-rank [CS14b, AEH19, BK14, EHL^{+}14]. low-regularity [Cia16, CJK17]. Lower [Bag17, SRG16, HGSL18, HL10, KP10b, MMA12, San12, WZ10, Wu18b, YLB16, ZS11b]. lowest [SLW14]. LQP [BX14]. LRBF [ZcHS18]. LSQR [LJJ17, LJLY19]. lubricated [Wan12]. Lucas [CKMR11, Oru17]. Lump [CM18b, IMS19, LCW19, MCF18, MZM18, TTG16, YMHL18, BM18a, HLLM19, HTY^{+}19, LM18a, SMYK19, WF18, YM17, ZCZ17, bZM17, ZDZY17, ZLT18, qZM17a]. lump-kink [bZM17, qZM17a]. Lump-type [LCW19, BM18a, HLLM19, YM17]. Lumped [FMSV17, XFY18]. lumps [YHC18]. Lupas [DK12, Ost11]. Lyapunov [DRD12, HmZ11, LCP10, TZZ11, TÇA12, ZHJ11]. Lyapunov-type [HmZ11, TZZ11, TÇA12, ZHJ11]. lymphotropic [EOM11].

M [BA11, EM19, ES11, LWD15, McN12, WL11b, ZLT18, WY11b]. M-ary [ES11]. M-estimator [WL11b]. M-lump [ZLT18]. M/G/1 [WY11b]. Mabrouk [EM19]. MAC [SR18]. Maccari [JKB11, SMDI18]. MacCormack [Ngo18]. Mach [sCYhX18, GSD^{+}19, HP19a, NDC^{+}19, PDN19]. Mach-number [GSD^{+}19]. Machine [WWH12, AA10b, AA10c, CQ11, CLW11, CCCW10, DM12a, DRK11, Hua12, Lee11c, LLW11, LLD10, LYLX11, LCC12, LW12c, MH11, MMR10, MMR11, RMS12, Wan08, WW10a, WWW11a, WWW11b, YY10c, YH11a, YX11c]. machinery [MSH10]. machines [CQ11, GAVOF11, TJQS13]. macromolecular [BNR10]. Macroscopic [DNP15]. MacWilliams [ÖŞ11]. MADM [Li10a]. magnet [HLCY12]. magnetic [BTEM19, BDGS13, BPS18, CL15, COR18, EEBM10, FST19, GGR19, HHM12, ILP^{+}11, Kia18, Ma18a, Ma18b, MCB10, SLKK19, XWN11, Yua18]. magnetized [Sea16]. magneto [BCSCB^{+}15, Her14, KHF^{+}19, LL16c, Ma18a, OSA13, SCBCB^{+}13, SCBCB^{+}17, YQ18, ZZ18b]. magneto-hygro-thermal [KHF^{+}19]. magneto-micropolar [LL16c, Ma18a, YQ18, ZZ18b]. magneto-thermoelastic [Her14, OSA13]. magnetoelectroelastodynamics [ZLWL11]. magnetogasdynamics [BSS18]. Magnetohydrodynamic [AVR17, CYL17, DM15, FVVS16, KLRW12, RY11, UKA15, Wu11a]. magnetohydrodynamics [AHHM19, DH18, FKKS11, YH19, Ye15, ZHY14, ZT15]. magnetosonic [MR19]. magnetostatic [CDN19]. magnification [GOT19]. Mahony [DAM14, HGHA19, KB19, Oru17, CdR18, NNWAS11, YL18b]. main [MJ10]. maintenance [ÁBÁPM11, CL12a, DM12a, YY10c]. majorant [Gon13]. majorants [RSH18]. majority [DDD10]. Majorization [PA12]. Makinde [Pan17]. making [ÇE10a, ÇE10b, FLLF10, LS10c, PCK13, WL13a]. Malaria [NNR14, PM13]. malware [DW18b, ZZW15]. mammalian [KC19]. mammals [KALAS11]. mammograms [LTL^{+}12]. management [BBO10, BPKM10, CSSW12, CCK12, Chu11b, Chu12b, cFpClC13, HDS11, HMP^{+}15, HMM12, HHS^{+}10, LS11c, OO12, OPDC12, PKTH13, VMAVGCM19]. Manakov [OVV^{+}16]. MANET [KHIB12]. manifold [QCT17, Tru19]. manifolds [BG13, PP12]. manipulator [YMM12]. manipulators [LH12b]. Mann [ZS11a]. Manna [LM18a, Liu18a, LZM18, MR17, PTZ19]. mannequin [JMDA13]. manufactured [KMS19, SNEP19]. manufacturing [DSB19, HDS11, KRAS19, MAPS10, RMK19, SCvdV^{+}19, WvDRG19, ZTY^{+}19]. many [Bao16, CTA12, CF16b, CT18, DL14, KK10a, KK11b, Lam13, LWKK10, LZKU11, WW15, ZMM18, ZD11, ZTZ16b]. many-core [ZMM18]. map [dSCM12, ZG18a]. MAPLE [PG10, XYXZ16]. mapped [WBA^{+}18]. mapping [ASV11, LLLC14, Zha11d, yZjH12]. mappings [AE12a, AKS10, APS12, BBR10a, CGY11, CGHY11, Cha11b, CS10c, CS11b, CM10b, Cie11, GDZ11, HLS11, HS11b, JPS10, KK10a, KK11b, KKS10, KA10b, KYO10, KYR11a, KYR11b, KT11b, LN10, LjHO10, LT11, Mai10, MS10b, Mar11a, NS11, Nil11, Pir11, QHW11, SLK12, Set12, She11, SH10, SCC12b, SSL11, XL10, XL11, YG11, ZS11a, ZRC11, ZLL11, sHC11]. MapReduce [ALLH11, ALLQ13]. MapReduce-based [ALLH11, ALLQ13]. maps [Ahm10, AH11b, And12, AvdW13, Bur13, CS11d, CGM10, JWX^{+}13, KK10b, LC11b, LW10, vdW14]. Marangoni [RGVR17]. March [Ano18-57, Ano19-51, LK15]. marching [KKT13]. Marcinkiewicz [HT19]. margin [LWJ10]. markers [WLC11]. market [BBO10, ELS11, LSV18]. markets [CJPB10, GW15b]. Markov [BY11, ELS11, LS11c, LKCN19, WHW11]. Markov-modulated [BY11]. Markovian [BV11, LJYS18, LZ12c, PdlF10, VB10b]. Marquardt [HM19]. Martin [Li14, YXWL14]. martingale [DFS11]. Maruyama [BM12a]. Mascheroni [CM12a, Mor10c, Mor10d]. masked [KLL10]. masonry [KKSK18]. Mass [DZ16, BHZJ19, Col14, DDK11, EAMA19, Gal10a, GBG11, Gen11b, HBK^{+}19, Lee16, LZP^{+}19, MS10c, NPR10, Sul16, XFY18]. Mass-conservative [DZ16]. mass-energy [Gal10a]. massively [Cos18, Kan15]. massively-parallel [Cos18]. master [AHOP18, ZWC10, ZWMD16]. matched [MLG17]. matches [KSM12]. matching [BL11, BE18, HL11a, HT16a, LG17, fLcJ10, WS10a]. material [BZKR15, Che18, FPB17, FMGR19, HMP^{+}15, HCHH12, JFS14, JFS20, TMLF19, XZL^{+}11, ZLS19, ZKBE16]. materials [BHM19, BML11, FXCC18, KMS19, QWL19, XLF12, XJ18]. Matern [LCLL18]. Math [AR10b, AD11a, Asl11, Bho14a, ÇE10a, Def10a, fDxZ11, DZ21, GXZ11, Her14, Her19, IIHuI10, ID10, JL20, Jia11, JFS20, JY11, KPS10a, KK11b, KYR11a, Li10c, LYS12a, LLH11a, LW12a, MP16, McN12, Pan17, Pen11, Ran15, SK12, SCBCB^{+}17, SKTC19, Ver12, WWW11a, WWW11b, WYG12, Wan13a, WLDL11a, XLD11a, XWH16, WNC12, WNC12]. mathematic [LHL12a]. Mathematical [ADL12, BPS18, CCDL10, CJK18, DKG14, GD10a, HHY^{+}11, JK12, Kuo16, SG11a, Ser19, SS13, AaC19, CFB11, DDLM13, DNR13, GL17a, JLWX18, MAPS10, MNJ^{+}13, OHMAK18, OO10, PBS12, PLKCC13, RMS12, RdSRL19, SSM^{+}17, SR17a, Sen12, SNSK19, SD15b, TSB16, WHLC11, YP10, ZQ11a, ASSV18]. Mathematics [Ant14, BCHS18, EM19, FdOdSS17, LK15, LZL^{+}13, Yaz11, YWL^{+}11b, Baa13, CBM10, CFB11, Gal11e, Lit13, PHWM10, SK10b, SK12]. Mathieu [CS10b]. mating [ZY13]. MATLAB [FRAK15, BCG17, JRA^{+}18, OBCG19, RAW^{+}16, Váz16]. MATLAB/GNU [FRAK15, JRA^{+}18, RAW^{+}16]. matrices [BKL^{+}19, CW18, DH10b, Din10, DT16, DD16, DIS19, EMR10, GH13a, HA18a, HDT11, HT16b, Ipe12, JSEM13, JL15a, JL15b, JYL16, JL17b, KKBR19, LFZ19a, LLCG16, Mis14, MM10b, NHIN16, PW18, RWW18, Res16, RC17a, SSR11, TL10b, TM17, WSH12, XC16, ZL12b, dS15, dS16]. Matrix [AEF15, jASzZ12, DH10b, LJLY19, XKH10, AZ15, ADD^{+}15, BT11, BK11b, BR18, BS11b, ÇE10a, ÇE10b, CA10b, CP16b, CM17, CP10, ÇA14, DLD10, DT16, DD10, DBEE11, DD16, ED11b, EHL^{+}14, FZ19, FG18, GMP18, Gem16, Haj18a, Haj18b, HT18, HH16, HSMT19, HM17b, HM14, HM15, HM17c, HM18b, JJ13, JL17a, KPS10a, KPS10b, KBDC12, KMRN12, KB10c, LKLP12, LHY11, LS11b, LZGZ11, LGG12, LH12a, LZD17, LY11d, LLLC14, LM18c, MKR12, Mar16, MKPS11, PS12b, PW10, RWW18, RWTW19, SD10a, SB19, SH12a, SS17, SPST18, SSP13, SWC11, Sön11, SI17, TM17, TCM18, TÇA11, Tia11, TT14, VT11, WLD13, WLW16, WLW17, WWXW19, WFDW10a, WZ10, WFDW10b, WLDL11a, WLDL11b, XHM14, XM15, XC17, XCS18, YL10b, YqS16, ZT13, ZLZ11, ZYW15, Zha15a, ZMLZ16, ZY17a, ZY17b, ZWLZ18, ZZ18a, Zha19a, ZZC13a, ZZC13b]. matrix-free [LHY11]. matroids [XS10]. maturation [YLZ17]. maturity [SXB^{+}12]. max [GWL11, KD11, YXP^{+}13, FKC12]. max- [GWL11]. max-product [KD11]. max/min [YXP^{+}13]. Maxallent [Gor13]. maxima [CHT11, LFJ11, LSJ12, GHR10]. maximal [CGHY11, FKC12, KST10, LG12, Liu14, Ver08, Ver12]. Maximizers [Gor13]. Maximum [Bog10, KST10, LD11a, LZS12, YXX11, YLH12, AA10b, BZZ^{+}10, GÇK10, KK15, KCL12, LHH10, Liu17, LMP13, Mah14, MMR10, MMR11, NN13, SH18, Tod12, XFY18]. Maxwell [AVR17, BJQS18, BHZJ19, BDHR18, BD19, CLT^{+}13, CDG16, CTC17, CT18, CFS17, DGL12, DGR18, DL14, EG18, FZ17, GT15, GT16, HLY12b, HCLL18, JRZK11, KVW18, LT13, LSS17, MSV18, NZ16, Per18, RÖ10, SBvdV13, SWL16, SLM18, SG16b, SG16a, Tri11, YZ10a, ZZLB18, ZSLZ19, ZLZG11, ZWG11, ZLS13]. May [Ano18-66, Ano18-68, Ano19-55, Ano19-57]. MBS [KKG10]. MCC [LR15]. McKean [MB10b]. MDS [LMPG13]. Mean [BKMT16, CCJ10, CCJV11, CJCV10, XH11a, ASB12, BM12a, BKM11, BR13b, ÇT11a, CZN11, mCfX10, Dra10, LYLX11, RG11, TC16, TÇ12, WHS12]. mean-reverting [TC16]. mean-reverting-theta [BM12a]. mean-variance-skewness [BKM11]. means [ADK10, Çan11a, Ciz12, DM10, GK18, GDM13, Hal13, HM10, HCL11, JW19b, KVW18, MN10c, MN11c, MKPS11, PAT12, ZKWW17]. Measurable [AKLS19]. Measure [MKPS11, Fer12, FOS19, MT19b, MN10c, WLT13a, WLT13b, XD10]. measured [ZM13]. measurement [BKP11, KVW18, SG16b, SG16a, ST12, WHQ^{+}18]. measurements [CL19, MMS^{+}18, MPMTV15, YLC12]. Measures [LS10b, SH12b, VP11, WZ11a]. MEBDF [DIJ12]. Mechanical [KP19b, SNEP19, AGH^{+}15, DM10, DPZ13, FG18, GLLC19, HMP^{+}15, HNPS13, MV10, SSIP19, YWHC11]. mechanics [BDM19a, CTSX16, Gal11e, JZJ18, LL19b, LZF18, MMH11, WZG19, WTLS18]. Mechanism [CY14b, CW17, CQLX11, CYM13, GZW^{+}18, HT12b, LHL12a, LLX^{+}10, RZL11]. mechanized [BCF^{+}14]. media [AGPR19, AP19b, ABK10, BASW18, BV10, BKNR19, CHM18, CP15a, CLB14, CJ18b, DVY14, EAAS18, EAMA19, GM18a, GQF^{+}10, GGGR17, GGR19, HWyL11, HCLL18, IK16, KMT19, KB10b, KRP12, KRBS18, LR15, MCB10, MM19, OHMAK18, PZL^{+}18, PCO16, RTL19, RJGS^{+}19, SSS16, SSHH^{+}18, SWOF19, SIL19, SCBCB^{+}13, SCBCB^{+}17, SP12, TR14, Tim13, Tim14, TMCM19, TRL19, UABK16, VZM14, XY16, YYK16, YB13, YLL^{+}14, ZSQ^{+}18, ZD18]. mediated [DDLM13, NJV13]. Medical [LXZ13, HT18]. Medium [Li18e, AKSW19, ASMM11, CWH13, CBB15, EEBM10, FLZ14b, GF16, GDM13, Her14, KNT12, KBCS16, LHL12a, Liu13, LZP^{+}19, MCR11, MS10c, NPR10, OSA13, Pan17, RHMA18, RBB12, Ran15, REHA11, SD19, SWL16, WO18, ZSLZ19, dBD17]. Meeting [HPV^{+}18]. Melnikov [YH12]. Melon [XC13]. Melting [WRW^{+}19, AML^{+}14, JFS14, JFS20, KÖC^{+}18, SSIP19, SM19, BGM19a]. melts [ZH11]. member [Gep16]. membership [DMPV10]. membrane [CNR10, GHC15a, SCA14]. membranes [BM18d, CIN^{+}18, VĆC10]. Memetic [tWqLzGkP11, EHO^{+}12]. memory [Def10a, Def10b, JKN10, JPK17, MWL18, MSV18, OHMAK18, Par15, Par18, SM17, TG11, TS11b, WL11c, WKP^{+}14, WY19b, XH11b, ZG14]. memory-based [OHMAK18]. memory-dependent [WL11c]. MEMS [LZ14]. Menger [AIB10]. Menten [CLM12]. Meromorphic [LSM11, AS10b, ASMEE11a, AX11, Las10, LY11a, Mar11b, QYL10, XLY10, ZX11]. meromorphically [EAA10, ORD11]. Merton [FES17, Jum10]. Mesh [CB19a, LMMT15, ABL10, Amo18, AKL18b, BGGCGRSP16, BDM^{+}19b, BP18, CBBE16, CHBTD14, CKRW19, CFLX18, DN18b, DVY14, DMRS18, DFP^{+}13, DLQ16, FGY^{+}17, FF15, Hak14, ILV^{+}19, JVMF19, KYW^{+}18, KMS15, KH18, KCL12, LZZC12, Li18e, LFAL19, MPS18, MKA^{+}10, PZZ18, SB14, Sul16, THGG14, VFM19, XFL16]. mesh-dependent [CHBTD14]. mesh-free [FF15]. meshes [AVV18, ABSV18, ACC18, BGF15, BN16, BDR19, BZT16, CBBE16, CWHW17, FMPR15, GSZ14, GGGR13, HMF^{+}19, HVO17, JW19a, JCF19, KSD^{+}19, PC17, SYI12, Tod13, Vac18, WW14a, WZH18, Wei14, Wei17, ZL19b, ZY19, ZMFL18, ZSY19, ZQ14a, dVDR18]. Meshfree [MR14, RC18, SL16a, SK19].

Meshless [LY15, MJWD19, AD15, BHJ14, BPF13, CD14, DAM14, DA16, DM16, DA18b, EFK15, FR15, FXCC18, ID16, JM15, JM16, KRP12, KBK19, LW18b, MHHC18, Mil18, MP19b, QWL19, ZN18, ZC16, ZSQ^{+}18]. mesoscale [MS17]. Mesoscopic [DELK13, HLK10, KLK13, LAM^{+}16, LLR^{+}19, YLLK14, ZLG^{+}11, MPGW19]. message [KC12]. meta [DM12a, KX12, OMS10]. meta-heuristic [DM12a, OMS10]. meta-heuristics [KX12]. Metadynamics [LNP^{+}12]. metal [BGL^{+}15, HJD15, ÖKJR19]. metallic [HLY16]. metamaterial [HLY12b]. metamaterials [HKW15, LSS17]. Metamodel [SJS^{+}10]. Metamodel-based [SJS^{+}10]. metapopulation [GMS18]. metering [LC10c]. Method [AMA14, AIIZ10, CPP15, DSK^{+}14, DK18, IIHI10, KDU15, PD11, Rus16, SISH12, SNEP19, SL16c, WH11b, YMDZ10, ZKR^{+}12, ZYT^{+}16, ZHL12, Aba10a, Aba10b, AB10a, AD15, ADZ19, AD19a, eMA18, AEH19, AvB16, AEG18, AO10a, AEH18, AM10a, AHV10, AMGC19, Ade17, AR10a, AÖ10b, ATUC15, AuIK17, Ahn12, AA10a, AK18, AB10b, AA15, AH11a, AJAR18, ArEM10, AM10b, AG10a, AM14a, Ali15, AGDP19, AHO16, AAD17, AEDL14, Alo11, AA18, AER12, AK19, AHF10, AT18a, AML^{+}14, AT18b, AOW18, ABLS15, ACAS11, Ant10, AVZ15, ABSV18, Arq18, AG10b, Asl11, AO13, AYY12, AuIA17, AJS14, BM11a, Bac14b, Bac14a, BCCZ18, BCH^{+}18, BN14a, BP19, BC16, BDM19a, BGGCGRSP16, BDM^{+}19b, BP11b, BMTV12, BRFH16, BC15, BASW18, BZZ^{+}10, BK14, BMM12b, BS16b, BCJ19, BLS17, BDF16]. method [BBBM16, BWL18, BPG10, BGPP11, BE11, BG11, Bic11, BKY10, BML11, BXKZ11, BX14, BG14, BR13a, BKR^{+}19, BHJ14, BGH14, Boy16, BB08, BRROP19, BSY^{+}19, BFS15, BPF13, CB11b, CKL18, CCN14, Cal19, Çan11c, CH11a, CM16a, CB10, CDG16, CGY10a, CGHY11, CAY12, Cer18, CR18a, CEQ14, CHBTD14, CB19a, Cha11c, CWW15, Cha18, CYS10, CW10b, CM11c, CM11b, CQ13, CCSZ14, CW14, CWWY15, CHLY15, CM16b, CCCW16, CTG17, CWDL17, CM17, CCY18, CS18, CW18, CJ18a, CHL18, CM19a, CLL19, Che19b, CLA19, CWY19, CHY19, CHT11, CChL14, CPT15, CM19b, CAH11, CK15, Cho17, CCK18, CL17b, Chu11a, CD16, CFS17, CM14, CX18, CRRS11, CSU13, CLL11, CAC14, DGB10b, DRT^{+}15, DWZ13, DZW16, DLZ17, DL10, DN18a, DH11a, DN18b, DSWB18, DH11b, DAM14, DM15, DA16, DM16, DA18b, DA18c, Del13]. method [DG13b, DRD12, DFGG13, DM12b, DT17, DMZ10, DSZ18, DWS19, DBEE11, DB12, DHY19, DZO^{+}19, DC15, DCN^{+}18, DGLU18, DBS12, Dos18, DH16, DLC19, DCRL13, DLQ16, Eba14, EFK15, EMQ18, ECJ16, EM19, ESBR10, ES18, Elb11, EZRR10, Els10, ET12, EZM12, EE10, ESL11, EG10, EE18, EAAS18, EKZ17, EGGS^{+}12, FID14, FJB19, FZ14, FZ19, FLH10, FL11a, FSRB15, FSB17, FT10, FH11, FN14, FZL^{+}18, Fia15, FRAK15, FNW18, FHS18, Fu19, FES^{+}19, Füh18, GL10, GWR^{+}18, GEZ14, GL17a, GGLP15, hGzS15, GLLC19, GLP18, GP19, GCE18, GRBT16, GS18, GK13, GGH18, Gen10, Gen11b, Gep16, GB16, GR13a, GH15, GY11, GMI11, Gon13, GR19b, GM14a, GK18, GR13b, GHCZ18, GNP14, GMZ15, GML17b, Gup11, GS11c, GM11, HMV18, HP19a, Haj18c, HSD10, HHY13, HXX19, HZ10, HP13]. method [HS11a, HH17, HKI12, HY16, HBK^{+}19, HJ13, He16, HHS^{+}17, HP17, HO19, HJD15, HG16, HS18, HM10, HA16a, HS13, Hes14, HSJ15, Hes18, HSWZ11, HBE14, HBE15, HT13, HKS14, HK15, HHG14, HA16b, HTL10, HVO17, HHGA19, HD14b, HLY17a, HYS^{+}14, HLSN15, HMWZ16, HCZ16, HRHP17, HG18b, Hu18, HMY18, HA10, HLZD11, HL11c, HZM11, HZP18, HWXC18a, HLX18, HCLL18, HWXC19, HM19, HD19, HLC11, IQR16, IK16, IS14, IIHuI10, JKB11, JGSS10, JKK12, JNJ^{+}11, JZ13, JM16, JRA^{+}18, Jaw13, JPS14, JJH16, JW18b, JW19a, JL18, JXZ^{+}10, JW15, JZL18, JZE^{+}18, KBA11, KAG11, KSD^{+}19, Kan15, KVW18, KBDC12, KA10a, KBS11, KWAS16, KK10a, KK11b, KRP12, KM14, KRCJ11, KM12, KKS10, KK14a, KFYW11, KW11, KNZ19, KL12b, KG11, KLH^{+}12, KW14b, Kim15, KL16b, KZ16, Kim18, KLTS11, KN12]. method [KS15a, KFTT13, KHF^{+}19, KAS11b, KQ11, KVR11, KSMN11, KK17, KTK18, KBK19, Kup11, LCW12, LP10, LK15, LDS10, LC10a, LP12, LWWY12, LLH14, LL14a, Lee15, Lee17, LC11a, LHY11, LMR14, Li10b, Li10a, LCP10, LLJK10, LYZ11, LML11, LLZ11, LW11a, LY11b, xLlFwWL12, Li12b, LZZC12, LY13, LJSK13, LZYW13, LY15, LZWC16, LFC16, LSZ16, LCP16, LL16b, LS16, LR17b, LCK17, LZD17, LGC^{+}17, LLFT17, LXZ18, LQMW18, LD18, LHY18, LHL18, LL19b, LFZ19a, Li19b, LPML19, LH19, LM19b, LK18, LZ19a, LZ17, LMLB19, LBJ10, LBW11, LYL12, LNKU12, LCLL16, LS10c, Liu11a, LL12e, LHL12b, LGL^{+}14, LHL15, LX15b, LDL^{+}15a, LL15, LZ15b, LDL^{+}15b, Liu16e, LZJY16, LWZ16, Liu16d, LS17b, LCW17, LW18a, LZLL18, LW18b, Liu18b, LZ18c, LCHZ19, LFZ19b, LFAL19, LW13, Lu11, Luk11, LWC13, LM18c, LBL14, MKHC11, MS18, MR17, MBS17, MDRRV11, MDVM17, Mai16, MN11a, MTM11, MKR12, MCKM12]. method [MTN19, MHHC18, MH11, MV10, MCL^{+}13, MBHV10, MMHGM17, MG15, MLY11, Mia18, MP19b, MFSL19, MHL11, MGB^{+}11, MN10a, Mok11, MSG11, MLGY16, MR15, MS12b, MT18, MDW13, NLA19, NNK13, NKM15, NTR15, Naw11, Naz13, Ned12, Ngo18, NDC^{+}19, Oan13, OKTR11, OKTR13b, OAKR16, OBAAD10, OTiSY16, OZF19, OAY11, OC14, OP14, OY19, Oru19, OHK^{+}19, eOS18, Özd18, PGQ16, PP10, PC14, PLKC16, PLW^{+}18, Par11b, PCK13, PFDG17, PGF18, PLR15, PXT10, PW10, PTH^{+}16, PD17, Pul16, QY13, QZ16, QZF19, QCLC17, QWGJ15, QWL19, QCG15, RZ17, RMM11, RMA10, RY10, RNB11, RSL^{+}18, RKA^{+}18, RAD13, RFK16, RKF18, RG18, RMS12, RWTW19, RAW^{+}16, RRAK19, RBTD14, RRP16, Ros12, RHC15, RS18b, SNH10, SD15a, SD18, SSAM11, SK11a, SC19a, SR15, Saj14, SEY12, SMF17, SMK18, SRGL13]. method [SR17b, SBEB10, SAIZ15, SMM19, SSIP19, SKTH11, SKG^{+}11, Sea11, SDH13, SG10a, ST16, SVY16, SD11b, SZ17, SHH16, SL18a, SA16, SG14, SS14a, SXM11, SZZ11, ST14, SW16a, SZL^{+}17, ST18, SGK18, SPST18, Śmi11, SST19, SMH18, SSP13, SS10, SG10b, SLYY13, SJC14, Sou12, SKM11, SW16b, SM14, SYW11, SSS11b, SK19, SKTC15, SKTC19, Sul16, SZ12b, SJHC14, SCGW18, SND19, SI17, TLR17, TR14, TPHD18, TTT10, THZ^{+}11, TDXQ18, TL10a, TC16, TM17, TCM18, TS11a, TNP17, TZXP11, TMZ^{+}15, TTZW18, TXL19, TCHW19, TDM13, TÇ12, TMLF19, TC10, TCM15, TT12, VAS^{+}18, VAB12, VBK13, VZM14, VHPVNXW18, WZW10, WKG10, Wan10a, WZCC10, WY11a, WSW11, WWA11, WXYW11, WW11c, WKS13, WL13a, Wan13b, WmN13, WQRZ14, WC15, WW16, WHTZ16, WZ17b, WTM17, WC17, WW18b, WCSW18, WH18, WLXG18, WCH18, WLL^{+}18]. method [WY18a, WW19b, WZC^{+}19b, Waz11b, Wei10b, WZKY12, WCCS15, WLZ18b, WYK10, WHW11, WH14, WHQ^{+}18, cW11, WZY13, WL17b, WSM^{+}19, XW14, XD17, XY15, XFY18, XFH19, XHM14, XM15, XXG10, XJLX10, XX17, XWL18, XCZQ16, YT18a, YY15, YXS10, YXYH10, YF10, YMSL11, YC11, YX11b, Yan11d, Yan17, YZX18, YCHW18, YLK10, YC10b, YKC11, YCLY15, YZ15, Yao16, Yaz11, YP10, YAS^{+}11, YKKS10, Yoo17, YASK10, YS16a, YTZ17, YW11b, YNS^{+}14, YSB15, YLLN16, Yun13, YS19b, Yus09, ZT13, ZT16a, Zak18a, ZTR11, ZSAN18, ZSAN19, Zbo19, ZS11a, ZCW15, ZM16a, ZSZ17, ZLS19, ZWC10, ZW11a, ZZ11b, ZTH11, ZLL12, ZLPM13, ZS13, ZHY14, ZHZ14, ZDLC14, ZTC14, ZHJ14, ZYW15, ZC16, ZKWW17, ZHC17, ZYSY17, ZLJ^{+}18, ZWLZ18, ZL18, ZSQ^{+}18, ZLW18, ZL19b, Zha19b, ZY19, ZH19, ZYZ11, ZHJD13, Zha14, Zha15b, ZLL17, ZP18a, ZcHS18, ZLZ10, ZC11a, ZOZZ12, ZL16, ZWJ^{+}11]. method [ZY10b, ZY11, ZL13, ZH15b, ZSD10, ZHW^{+}11, ZC11c, ZZL15, ZC17, ZQ14b, ZT18b, aZ18, ZMA10, dPLM18, CC15, CM16c, GD16, MRR18, WKBR18, dVDR18]. methodologies [CCFV12]. Methodology [HNK13, LLL13, MA16, MBT^{+}13, SCvdV^{+}19]. Methods [BDGG14, CN16b, GIM15, GCDG17, KLK13, LAM^{+}16, RCM11, SVP^{+}19, YLLK14, ADGS18, AAB^{+}13, AD16, AS11b, AO18, AYH17, AGPR19, AP19b, APTZ19, AS15b, AKMUH17, BQ15, BJ19, BJRF19, Ban13, BP11a, BNTT14, BS11b, BCD^{+}16, BJS15, Ber16, Bho14a, Bho14b, BQS16, BH14, Boy16, BLyS18, BC17, BK12b, Buo11, BKZ17, BS17, CCKP15, ÇT11a, CGGM19, CNV14, CMR17, CTP10, CWHW17, CCG18, CF16a, Che11d, CDY11, CLTA11, CLH13, Che15a, CLC16, CM18c, CCZ18, CY19, CJN19, CTM^{+}13, CGK14, CFdM^{+}18, CAP10, CSN11, CDL17, CELY18, CJK17, CJ15, CHH14, CN16a, CL16b, CN11, CT17c, CKR10, DIJ12, DMP18, DNS15, DELK13, fDxZ12, DSM18, DGM14, Don10b, DYWL19, DGH17, DFS14, DZ16, DHMU16, Eba11, EO15, ED11a, FMS19, FWZ16, FHZ13, FGHZ13, FLdS14, FT15, FMPR15, GML17a, GP11, GGVRB19, GTG11]. methods [Gem16, GK11a, GK11b, GSZ14, GR19a, Git14, GHM^{+}14, GHC^{+}15b, GDF12, GABC16, HN18, HCL12a, HH16, He11, HLL13, HR19, Her19, HSMT19, HR15, HLK10, Hof18, HH15, Hou15, HW16, Hu15, HCZ16, HM14, HWXC16, HPS18, IB11, Ikh11, IW18, Izs15, JMHF13, JM15, JLL18, JAJ18, KMS10, KM15, KBAF18, KKAM11, KTA12, KKC^{+}10, KCL14a, KCL16, KPS17, KKVS19, KO13, KTH13, KQ11, Ku18, KY11, KV18, KdLK19, LW19a, Lee16, LLR^{+}19, LN98, Li10c, LLML15, LH16, LSS17, LM17, LM18b, LNW19, LGH^{+}11, LCZ11, LZB12, LWQZ18, LCHZ19, Loh16, LR13, LO16, Ma10a, MZC17, MRS15, MVKK14, MR14, MO14, MD15, MBJ16, MM19, MWY17, MT19b, NEB14, NCC13, NKA18, NUNAS11, OBCG19, OO10, Oh15, PIAH10, PPD10, Pet14, PS18, PAT13, RMB^{+}14, RH15, RIW12, RTT17, Rob14, RJGS^{+}19, RCG15, RC18, SBvdV13]. methods [SHM13, SWOF19, SDH15, SL18a, SBS12, SL16b, SL18b, ySW10, SLM18, SLL17, SV11, Ste16, SR17c, SCGW18, TH19, Tha19a, TS16, TFS11, UKI11, VMP15, WJ11, WLD13, WLM13, WW14a, WLA18, WMW13, WFZ12, WV15, WZHW13, XL15, XZ10, XZZ16, XY17, XWY17, XWY18, XW19b, Yan10c, Yan18b, YQWZ19, YYL11, ZSY14, ZL14b, ZT15, ZZWG16, Zha17a, ZTY^{+}19, Zha13, ZLA17, ZLG^{+}11, ZJZ18, ZZ14, ZM16b, ZSY19, ZD15, ZGD14, dVLV18, dPRVRB13, GGGR17, Vac18, DJD18]. metric [AKT12, AR10d, AHF10, AKS11, ASV11, APS12, BBR10a, CSW11b, CM10b, ĆSCD11, HLS11, Haz11, HS11b, IBB10, IN10, KPR10, KR11, Kar10a, KYAA10, KA10b, KLMV12, Lee11a, LC11b, MAK12, NS11, PRR10, Qiu12, RK10, RA11a, RI12, Sha10, SS11c, SCK11, WFY17, YBC11, YKRV11, Zha11d]. metrics [GAVOF11, PDHL12]. MEW [Sah17]. MFE [WLL^{+}18]. MFEAs [QM19]. MFEM [AD19a]. mfront [HMP^{+}15]. MFS [LCK13, NKM16, YY14]. MHD [DM15, BGP13, CYZZ18, Das12, EUTS18, FSZ17, FSTN18, FBTS19, GGR15, GZN19, LZZ11a, LZZ11b, Liu18c, LZP^{+}19, MC10a, MM18a, MI16, MGB^{+}11, Pal13, RY10, RHMA18, RRP16, STS19, SB14, TH19, Wen18a, Wen18b, XZ17, YS19a, YZS18, ZH15a, ZaY17, ZZLB18, ZJB19, Zha19c]. MHSS [ZSZ17]. MI [WCD10]. MIB [XW14]. micellar [KL16a]. Michaelis [CLM12]. micro [ASFM15, BM13a, CCY18, GZR^{+}13, GSZ11, LG13, RMA10, RR14, YLF19, YWHC11]. micro-actuators [RMA10]. micro-channel [LG13]. micro-electro-mechanical [YWHC11]. micro-particles [ASFM15]. micro-polar [BM13a]. micro-tomography [GZR^{+}13]. micro-vessels [YLF19]. micro/nanoflows [RR14]. microcalcification [MBKK10]. microchannel [AYY12, HCT12, LZB15, PC14, ySGL^{+}10]. microfluidic [CGS12, PN16]. microgrid [HMM12]. micromagnetics [PRS18]. micromorphic [ZLWL11]. micron [GK19]. micron-scale [GK19]. micropolar [ABV11, BPR18, BGP13, BPZ19, CL15, Das12, LL16c, Ma18a, UKA15, YQ18, ZZ18b]. microRNA [NJV13]. microRNA-mediated [NJV13]. microscale [CD19, GQF^{+}10]. microscopic [BKK12]. microstructural [ZTY^{+}19]. microstructures [AGDP19]. microtissues [PN16]. middleware [AJJAD^{+}10]. midpoint [PRS18]. migration [Ben12, CQ11, Par17b, WCQ^{+}19]. mild [BH10, CZN11, CZN12, CCD10, HB12b, JS11, LJ11, SW12, TNHK19, ZDB19, ZJ10c, ZLW19]. Miller [Ebr11, NODA11]. Milstein [CFN11, KPR13]. Milstein-type [KPR13]. Mimetic [dVLV18, AD16, AVZ15, BG14, GB16, JAJ18, RGHZ15]. MIMO [Sha12a, SZGG11, VMO10]. min [YXP^{+}13]. Mindlin [CCN14]. mine [RB19]. MINI [CB19b]. miniature [KW12]. minimal [AM15, BWL18, CWW19, Ili10, KM12, SJHC14]. minimax [LZ12b]. minimisation [HW16]. Minimization [FT10, KSS13, Pir11, WHD14, ZD12]. minimize [AA10b, AA10c, LYLX11, ZLG^{+}10]. minimized [PGDL18]. Minimizing [DP15, MMR10, MMR11, YY10c, ZGL14, CTZ17, HWY14]. Minimum [BDGG14, GCDG17, HM17b, HB12a, KCL12, YS19b]. Minimum-norm [HM17b]. Mining [FKC12, LZL^{+}13, KOPS13]. Minkowski [AZ15, UMY11]. Minty [RGdSRLAJ10]. Miodek [YGS^{+}16]. Mirakjan [Mah10a]. Mirror [JRSZ12]. Mirror-curves [JRSZ12]. miscible [HCZ16, LR15, LCHZ19]. MISTY [LLSS13]. Mitigation [TBP19]. mitochondria [KC19]. Mittag [GMP18, Kir10a, LCP10]. Mittag-Leffler [GMP18, Kir10a, LCP10]. Miwa [BM18a, HTM18, LDL11, LYC^{+}19, LSX13, Man18, YM17, YTS^{+}17, ZBFC19]. Miwa-like [BM18a, ZBFC19]. Mixed [BGR14, BM12b, JN14, KRBS18, LLML15, LTSW16, MS12a, MS10c, SRRP18, bZM17, qZM17a, AD19a, AAZ10, AH10b, ABCR10, AD16, AS19, AP19b, APwS18, AJS14, BQ15, BQ17, BJ19, BL10, BNR10, Ber12, BGPP11, BM18d, BHH16, CB11a, CB11c, CGY10b, CLH13, CHLY15, CL15, CS11b, CB19b, CN16a, CGO19, DPBL16, DN10, Das12, DW18b, FKF13, FDG^{+}17, FGHZ17, GRBT16, GGS16, GOS18, GGVRB19, GS15b, HTGSH13, HLL13, HT13, HSBL11, Hou15, HLY17a, HMWZ16, HCZ16, Hu18, IK16, JJH16, KC11, LCQL17, LL16c, LRH13, LDL^{+}15b, LCHZ19, LT11, Ma18a, Ma18b, MP10b, Mat19, NPR10, NUNAS11, ÖZ11, PS18, RBB12, Ran15, RTT17, SK11a, SC19a, SPH10, SHM13, SLW14, SW16a, SZL^{+}17, SSPL10, SJC14, SR17c, TMMASG10, WAZ11a, WC17, WW19b, WFZ12, qXjH11, XH11a, YNLK10, ZR18b, ZR18c, ZS13, ZWH^{+}19, ZMFL18, QZM17b]. mixed-FEM [GOS18, QZM17b]. mixed-hybrid [PS18]. mixed-integer [HTGSH13]. Mixed-type [LTSW16]. mixer [MMS^{+}18]. Mixing [ATO19, GSD^{+}19, HTL10, KRP12, LLG10, ZCSG13]. mixture [BKR10, DB10, FM17, Gro19, PFDG17, RNQ13, RNQ16, TC18]. mixtures [GT15, GT16, Laz10, PXT10, VMP15]. MLAIN [IC12]. MNC [HHS^{+}10]. mobile [DA12, HSS^{+}12, JCWZ16, LX12b, ZLPM13]. mobile-immobile [ZLPM13]. mobility [BDPM12, CCK12, DFJS10]. Mocanu [NUH12, NODA11]. Modal [ZT18b, Che12c]. mode [BJLZ12, BRFH16, BG10b, CZMZ11, CL12b, FDB13, HGN^{+}10, HLCY12, Kuo11, LZY12, LZ12c, PC12, YWHC11, ZR18a]. mode-dependent [LZ12c]. mode-locked [BG10b]. Model [ÇA14, DK14, HTGSH13, JJH16, LEP11a, WJ11, WRW^{+}19, ABM11, AD15, ADY12, AK16, AFGL10, AZB13, Aki17, AaC19, AdSSS19, AWJH19, AEG11, ABLS15, ABR^{+}14, AAP12, AGU14, ALMLM14, ADS14, ASSV18, ABK10, BM12a, BN14a, BP19, BMAR18, BS10b, BC16, BMS19, BBDS11, BP11a, BG15, BPR18, BO10, BMM12a, BGRV15, BPS18, Bho14a, Bho14b, BF11, BSZ16, BMS13, Bis14, BD19, BSN13, BDB12, BKR^{+}19, BGL^{+}15, BNP18, BK15, BFF^{+}11, CW15a, CCN14, CLT^{+}13, CCR16, CYP16, CCJP11, CD19, CM15, CN13, CST14, CB19a, CWH13, CXZ15a, CL16a, CM18c, CJ18b, Che19b, CZ11b, CXMO19, CV14, CKRW19, CGK14, CCCW10, Chu12b, CT10b, CTS19, CFLM19, CGO19, CHS19, CS13, CAC14, DD19a, DH17, DDLM13, DZW^{+}15, DNZ^{+}13, DYQM14, DNR13, ECJ16, ES10, EOM11, EZM12]. model [FGB19, FJWW16, FJC16, FSZ18, FQLC18, FKC12, FIS18b, FL13b, FHA16, Fu19, GWR^{+}18, GSI19, Gal12, GS15a, GL17a, GCR^{+}18, GMS18, GGH18, GB18, GGK18, GGL13, GOGYL^{+}11, GM18b, GDM13, GJX18, HLZM16, HT18, HNK13, HY15, HBS^{+}10, HZ16, Hei10b, HHNLGC18, HMM12, HW19b, HQ19, HL18b, HMY18, HLNZ19, Hu19, HGW11, HLY16, HLC11, HKW15, IS14, JKN10, JS12b, JGK13, JWX14, JLWX18, JW19b, JQG14, JK12, KP18, KGJ11, KL16a, KK12, KSM12, Koj10, KÖC^{+}18, KSO16, KS15b, KM10, LMPE18, LHL12a, LPLR19, LLG^{+}11, LC10b, LK11, Li11b, LW15, LCS15, LWN15, LWD15, Li16a, LJYS18, LMY19, LZPZ19, LSC17, LL12c, LMZ18, LCCC10, LY11c, LJX12, LC13, LRV13, LWL14, LS17b, LZCL18, LRBA15, LLL16, MGTH16, MW17, MLY18, MMRN12, MAPS10, MM16, MNJ^{+}13, MGY11, MW16, MM11, MSW18, MSZG17, OHMAK18, OAKR16]. model [ODAZ15, Özu15, PZL^{+}18, PSP10, PQB^{+}16, PQBK17, PHPK12, PLR15, PGW19, PM13, PBS12, PLMS14, PDHL12, QXLL11, QZY11, RPTD10, RSL^{+}18, RKA^{+}18, RA18, RQ18, RTV17, RTL19, RRO17, SSM^{+}17, SG11a, SH18, Sal16, San11, SCSF19, SWOF19, SBKS12, SS16a, SNSK19, SMC10, SGZW18, SP12, SS19, SD15b, SSK13, SZ14, SSS^{+}11a, SW10, SLZ11, SFM15, SW19, TTMJ12, TM18, TZ18, TTM19, TZ13, TNV19, VGK^{+}16, VGC^{+}15, WCD10, WWG10, WC11a, WYN12, WHD14, WHZL17, WYL19, WW19a, Wan19a, WZWS11, WHLC11, WTSS10, WCW13, WGY^{+}18, yXpYxZT11, XL15, XY14, XZ11, XMW10, Xu11b, XLF12, XC13, Xu14, XC16, XBHN16, XGH17a, XJYL17, XGH17b, XWL18, XXH18, YY11a, YLY12, YG17, YZ10a, YWL^{+}11a, YL13, YZWW14, YX16, YZMZ16, YGS^{+}16, YLZ17, YZ18, YL18a, YLC18, YZM^{+}19, YXWL14, YKA18, YLDL11, yYsZyYL13, YDW15, YWL17, Yua18, ZK16, ZZ19]. model [ZH11, Zha11b, ZZ11a, ZLPM13, ZZX^{+}14a, ZL14a, ZHW14, ZY15a, ZYZ^{+}16, ZWZ16, ZCY16, ZLTY16, ZYZC18, ZHW^{+}18, Zha18b, ZZF18, ZSW19, ZSLZ19, ZLC^{+}11a, ZS15, ZW16b, Zho13, ZLC11b, ZZW15, ZZ17, ZLY17, ZLZ18, ZY13, jZsQdLmG19].

Model-based [DK14, RSL^{+}18, SWOF19, WCD10]. modeled [ZS18]. Modeling [ABH12, BCH^{+}18, CN13, COR18, GPV11, JVMF19, KMS19, MGN^{+}16, PV12, Pet11, SPS^{+}13, SL18b, SKK12, THY^{+}10, YGR11, ZC11b, ADL12, Baa13, BNR10, BHM19, BCC14, BZZ^{+}10, BWZ16, BKR10, CCDL10, CANA19, CP16b, CSZK10, Che12c, CJPB10, DEFP11, DNP15, DYX11, DFP^{+}13, DNZ^{+}13, DW18b, DA12, Ebr11, EK16, EAEH18, EE18, Fio14, GD10a, GQF^{+}10, GACMO13, GR10, GV11a, GGR19, HWH^{+}15, IBG12, IL13, JLD19, KMT19, KEHB18, KBGC12, Kei13, Kuo16, LZYW13, MCL15, MY16, MZM19, MM18b, MRS^{+}12, MEAMHHV18, ÖKJR19, PKD19, PLKCC12, RS12b, RdSRL19, SCA14, SWOF19, Ser19, Sha12a, SZP^{+}11, WA19, YC10a, YKKS10, ZZHF12, ZSY14, ZHC17, Zho19]. Modelling [CD12, CGS12, DLF^{+}11, GZ10, KKK12, QLT^{+}18, WWW11a, WWW11b, Amo15, BASW18, Ben12, BD11d, CI18, GHT^{+}15, HMP^{+}15, LDS10, RTL19, SM19, TRL19]. models [eMA18, AM12a, AGPCC10, AD19b, Bab14, BL18, BGM19a, BCFQ19, BDPM12, BCPS15, BKM11, BBL19, BKK12, BKMT16, BK16, Bre14, BS18b, Cao19, CB10, CJK18, CLM12, Che16, CR18b, CWY19, CTM^{+}13, CJPR10, Coo10, CRRS11, DPBL16, DNS16, DGZ13, DPZ13, DGOZ13, DLT12, DMRS18, DGM14, DKG14, Dun18, ES18, EPP18, FCZ12, FAHZ17, GM18a, GUK13, GFZ16, GGT14, GLL14, Gur13, HMV18, HNK13, HKS19a, HBK^{+}19, HpD11, HPR19, HH10a, HWyL11, IW18, JRB15, JL16, JBBL17, KNZ19, LD13a, Laz10, LLZ11, LZ16a, LCS18, LB11, LZB12, LCW17, LCWZ18, LL17, LSV18, LPP15, LRTV10, LPY16, LW11d, MZC17, Mag10, MV12, MMOJ14, MPY16, MW13, MPGW19, MBT^{+}13, NB11, NCV^{+}18, OHK^{+}19, PP10, PFDG17, PGF18, PLKCC13, QY13, RSS10, RR11, Sag10, SYZ19, SRDD17, SD11b, SLL17, Sin16]. models [SM10, STDLM19, Sul16, Sun10, TWLYÖ10, TH13, VMP15, WRW13, Wan14, WZ17a, WZXS17, Wri13, WS16, XY16, XFH19, XJ18, YB13, YW14, YZ19, Zbo19, ZZXY12, ZZX^{+}14b, ZSW15, ZYZ^{+}17, ZZ10a, ZLY12, dPRVRB13]. moderate [Pap15, ZZG19]. moderate-degree [Pap15]. modernization [CCHG17]. modernized [WBA^{+}18]. modes [MTV13, SA16]. modification [Bra13, DK12, KKC^{+}10, Özd18, PSP10, SR17b, SS10]. modifications [Cie13, EGGS^{+}12]. Modified [Aba10b, AY12, FH11, HA10, NLA19, SH10, YMDZ10, YS16a, ABM11, AM13a, ADZ19, AR10a, AA15, BGIN13, BM18c, Bra10, CL11, CB11b, CFN11, CM11c, CM11b, CTG17, CW18, CT17b, CdR18, DH11a, DMPV10, DYWL19, DR12, EK16, FM19, Gen10, Gep16, GABC16, HD14a, Her19, HM14, Jav11, JJH16, JW10, KMS10, KK10a, KK11b, KNZ19, KBCS16, LK15, LG10, LZD17, LCQF19, LHL15, zLYLQ19, uHS12, MM18a, MJWD19, MGS^{+}14, MP19c, MD15, NNWAS11, OHMAK18, PP14, RR18, Ray18, RMY19, SK11a, Sea16, SIL19, TFS11, WKG10, WW16, WTM17, WLW17, WW18a, WW15, YL13, YS19a, Yil19, YXWL14, ZWC10, ZTZJ14, ZYW17, Zhu10, ZH18b, ZLGL11]. modular [AH11b, Deu10]. modulated [AGDP19, AGD10, BY11, Che18]. modulation [GTZ19, WAG^{+}14]. Module [AK12, JPP12, WD13]. Module-based [AK12]. modules [AS11d, AB11]. modulus [FZ19, RWTW19, ZL14b]. modulus-based [FZ19, RWTW19, ZL14b]. Moffatt [BK18]. moisture [KKSK18, Zho16]. MOL [DM15]. mold [ZOZZ12]. molding [CCY18]. Molecular [KHWK10, SID15, EHO^{+}12, NEB14]. molecular-continuum [NEB14]. molecules [BO18b, EHO^{+}12, GT15, GT16]. mollification [AM10a, MAS11]. MOMCMC [Li12b]. Moment [FXC18, AvB16, BM11b, DL16, FN14, HD19, LMR14, pZ10]. moments [BF11, KYA15, MD10, Nie10]. momentum [HYS^{+}14, KS15a, YNS^{+}14]. Monge [Dai14]. monitoring [BCF10, HLT12, JXZ^{+}10]. Monk [BCHS18, Col18, Per18]. monochromatic [HCT12]. Monocular [YC12]. monocyclic [AZB13]. monodimensional [Dub13]. monodromy [ÇA14]. monolithic [NH15]. monopolies [ATZ11]. Monotone [FL14, Ber12, Bor11, CGHY11, CM11c, CT10a, LZ12a, LL15, LT11, Mai16, Tia19, WW11c, XJLX10, Zha11a]. monotonic [ABL15, DBH^{+}14, SKJ10, SGQ12]. monotonicity [Ver08, Ver12]. Monte [AEF15, AGK15, DGLS19, HC18, Li12b, Mil18, MDW13, PA15, SWOF19, VMC^{+}14]. Monte-Carlo [HC18, VMC^{+}14]. Montgomery [YXYH10]. Moore [BT10, LZ19b, SSP13]. Morley [CGH14, YLB16]. morphochemical [BGL^{+}15]. Morrey [Gao12, LZG13]. Morris [LSW16]. Mortar [KCL16, APwS18, KW14a, RZ17, SD18]. mosaics [JRSZ12]. MOSFETs [VMC^{+}14]. mosquito [ADL12, DD13]. motif [fLcJ10]. motifs [fLcJ10]. motility [XBHN16]. motion [BJLZ12, BN14a, BH11, DTR19, Dun11, FAIV10, Gal11e, GC19a, JNJ^{+}11, MP11c, SPL19, SM10, TC16, WSW11, XaZH19]. motions [ZTSC16]. motivated [GMS15]. motor [BJLZ12, HLCY12, LC12b, MK18, SST19]. Motsa [Pan17]. Moulton [PIAH10]. Mountain [SZ12b]. mounted [CJ12]. movable [FWFL11, KHF^{+}19]. movement [PLKCC12, RIW12]. Moving [ABL10, Amo18, BDR19, CKRW19, CDM10, DL16, DVMS13, DA16, DM16, EAEH18, FIMV18, FGL10, HH17, HSWZ11, JSGP16, Kia19, KÖC^{+}18, LHL12a, LL16b, LX12b, LGVS19, Ma10a, MDVM17, MT19a, MJ14, PTH^{+}16, SA16, ST18, TBP19, UKA15, WLT13b, YD12, ZDL11, ZWL11, ZR16]. MPS [NKM16, SGK18]. MPS-based [SGK18]. MR [LW12c, XY14]. MRF [LYSZ19]. MRT [GUK13, MPGW19, RKA^{+}18, ZPGW16]. muffin [BO18b]. Multi [BMS13, BL17, BO18b, DDMQ19, GACMO13, GWL11, IC12, LY12b, MGS^{+}14, OKTR13b, SH18, SKG^{+}11, VKJ13, ADZ19, ARESH18, AFGL10, ATUC15, AuIK17, AK19, BZ10, BRFH16, BMM12a, BZ18, BKK12, BN14b, BSY^{+}19, BS17, CCY10, CCL^{+}12, CLA19, CS11a, CCFV12, CS11d, Cie11, DM12a, DMPV10, DGK10, DA12, ESN10, FIVM17, FL11a, Fia15, FMGR19, FHA16, Gha17, Gha18, GY11, GN19, GLW18, HMY15, HG18a, HL11b, HW19b, JGSS10, JA11, JLTB12, JZE^{+}18, JK18, KGJ11, KSD^{+}19, KV17a, KC12, KRCJ11, Kim11, KK10b, Kir10a, KSO16, Kuo18, KM10, LGHR16, LWR16, LXZ13, LW11b, LZY12, xLlFwWL12, Li12b, LZZC12, LY13, LWL11, LSM10, LCLL16, LCCC10, LMDL11, Liu12, LL16d, Liu17, zLZ19, LFAL19, MKL11, MGN^{+}16, MVKK14, MLZ^{+}16, MB17, MM19, NB11, yN11, OBAAD10, Osm18, PLR15, RPTD10, SC19a]. multi [SLL12b, ST12, TZXP11, VCM11, WZH10, WWA11, WL13a, WZ16, WYL19, WO18, WKP^{+}14, XBHN16, YXS10, YX11b, YB13, YKA18, Yüz12a, Zak18a, ZMM18, ZL14b, ZLA17, ZZL^{+}18b, ZWH^{+}19, ZDF^{+}14, ZT18b, ZD18]. multi- [ZMM18]. multi-agent [CS11a, MB17]. multi-asset [GLW18, MVKK14]. multi-attribute [LCCC10, WL13a]. multi-axis [LZY12]. multi-body [HMY15]. Multi-classifier [VKJ13]. multi-component [BN14b, MM19, YB13]. multi-context [AFGL10]. multi-core [Fia15]. multi-criteria [KC12]. multi-delay [YX11b]. Multi-dimensional [BL17, SH18, ADZ19, AuIK17, AK19, BZ18, Gha17, Gha18, HG18a, LCLL16, Zak18a, ZDF^{+}14, ZT18b]. Multi-domain [BO18b, FIVM17, KSD^{+}19, PLR15]. multi-fan [LSM10]. multi-fluid [KSO16]. multi-frame [LGHR16]. multi-frontal [WKP^{+}14]. Multi-GPU [OKTR13b, LWR16]. multi-grid [ZD18]. multi-hop [DA12]. multi-index [Kir10a]. multi-instance [xLlFwWL12]. multi-layer [HW19b]. multi-leader-follower [yN11]. multi-level [BS17, CCY10, JZE^{+}18]. Multi-leveled [IC12]. multi-material [FMGR19]. multi-mode [BRFH16]. Multi-Objective [GACMO13, GWL11, BMM12a, CCFV12, DM12a, DMPV10, KGJ11, Li12b, RPTD10]. Multi-operator [LY12b]. multi-order [GY11, YXS10]. multi-pantograph [LW11b, Yüz12a]. multi-parameter [GN19, NB11]. multi-parameters [ZL14b]. multi-patch [JK18, KV17a]. multi-phase [BKK12, MGN^{+}16, MM19]. multi-point [BZ10, DGK10, ESN10, HL11b, Kim11, Liu12, WZH10, WWA11]. multi-projection [MKL11]. multi-quadratic [Cie11]. multi-relaxation [KM10]. multi-relaxation-time [LZZC12]. multi-resolution [LXZ13, TZXP11, WZ16]. multi-scale [LY13, XBHN16]. multi-sensor [ST12]. Multi-soliton [MGS^{+}14, Kuo18, LMDL11, LL16d, zLZ19, Osm18]. multi-solitonic [LWL11]. multi-species [FHA16]. multi-stage [CCL^{+}12]. multi-state [VCM11, YKA18]. multi-step [ATUC15, FL11a, KRCJ11, OBAAD10]. multi-target [WYL19]. Multi-temperature [BMS13]. multi-term [ARESH18, BSY^{+}19, CLA19, JGSS10, JA11, JLTB12, Liu17, LFAL19, MLZ^{+}16, SC19a, SLL12b, ZLA17, ZZL^{+}18b, ZWH^{+}19]. Multi-thread [SKG^{+}11]. multi-valued [CS11d, KK10b]. multi-vendor [KGJ11]. multi-waves [WO18]. multiattribute [WO10]. multicast [KCL12]. multicomponent [KNT12, WS16]. multicore [WWZ12]. Multidimensional [Izs15, Mac12b, She16, ACD^{+}11, DGOZ13, DB12, Erg19, HA16b, Koj10, LN19, MD18, RS15, Saj14, SW16b, ZY17d]. multiexponent [HL10]. multifield [KRD16]. multiformalism [IBG12]. multifrontal [YW14]. Multigrid [AZ17, BLyS18, Che15a, SCGW18, AGPR19, ACC18, BA16, CTP10, CO15, CWWY15, DWZ13, DLZ17, DLZ19, DDM^{+}18, GEZ14, GGGR13, GALO18, HRHP17, Kan15, LYC15, RC17a, SRGL13, ST19, WN18, YSX^{+}19]. multigrid-convergence [GGGR13]. multigrid-like [BA16]. multigroup [Sun10]. multiharmonic [KKL^{+}13]. Multilayer [KY11, RCH19]. multilayered [AGH^{+}15, LLL16]. multilayers [AHP^{+}14]. Multilevel [CDY11, LT15a, YH15, ZZL18a, Bac14c, Che11d, DGLU18, JZR15, LCQF19, PA15, SWOF19, Sha12b, TL12, Tom13]. multimaterial [BMS19]. multimedia [CL12c]. multinational [ZHT11]. Multiobjective [AH10b, GJ10, Ant10, Ant14, GKS10, KMS15, ZLLF12, ZGZ13]. multipath [YDK^{+}12]. Multiphase [CI18, LK11, LBL14, AWJH19, DVY14, DGM14, HWyL11, KSD^{+}19, LSC17, MA16, RTRR18, SGK18, WGY^{+}18, YZWW14]. multiphysics [GGH18]. Multiple [Ade16, AD11d, CZ15, CZ18, HLWX11, LL19a, LCW10, LY14, LM18a, LJ10b, LC16, PB13, San12, WSG10, ZJ10b, ZT16b, AT18a, BKT12, BGRS11, BG10a, CTG17, CHZ19, CMS10, DWZ13, ES10, FNZ13, FVVS16, GGAV18, GZZ^{+}16, HH10a, HGW11, JTCC11, JC12, KBGC12, KG11, LCN10, LYZ11, LSC17, LWZG10, Lin11, LKL^{+}15, LS10c, LC10d, LTL^{+}12, Liu12, Liu16a, MCKM12, NGL10, OHK^{+}19, PCK13, PPD10, RWZ13, SKTC15, SKTC19, TAA14, UABK16, WW10a, WYN12, WL17b, WZMY18, YLC16, ZSW15, ZM17a, ZHW^{+}18, ZL10b, ZCY11]. multiple-attractor [FNZ13]. multiple-attribute [LS10c]. multiple-GPU [YLC16]. Multiple-lump [LM18a]. multiple-point [ZL10b]. multiple-relaxation [OHK^{+}19]. multiple-relaxation-time [LSC17, SKTC15, SKTC19, ZSW15, ZHW^{+}18]. multiple-scale [Liu16a]. Multiplicative [Uze10, LYN11, LX15b, UCK16, aZW17]. multiplicativity [Tac11]. Multiplicity [DSL18, GJ19, LL14b, MV11, MPLR18, PD11, SYY13, YL16, Amb19, BS12a, CW19, HYCP11, HX14, Jal14, Li14, LL16a, LLZ18, Lü14, TJ10, WZF16, WS17, XC18, YWL17, ZXZ17, ZZ18c]. multiplier [BGR14, CKK^{+}10, EAAED10, EAA10, EKZ17, Kim15, KPS17, WLW17]. multipliers [KEHB18, PBK19]. Multipoint [AP19b, AIIZ10, AG10b, GK11b, Jaw13]. Multipole [CC15, Che18, QCG15, TXL19]. Multiquadric [CHL18, GWZ11, GW15a, WXF10]. Multirange [KM10]. multirate [AKSW19]. multiresolution [FRSC16]. Multiscale [DCN^{+}18, IL13, LDS10, SLM16, Tim13, WA19, APwS18, CHLY15, CELY18, CFS17, DWZ13, DLZ17, DLZ19, DC15, DNR13, GGK18, Grm13, ILP14, KW14b, Li10b, Liu16d, MS18, WH14, ZHZ14]. multiscale-homotopy [Liu16d]. multiselection [Als10]. multisensor [LZS12]. multislope [ZÖXL^{+}19]. multisplitting [ZL14b]. Multistability [YY12]. multistep [AS11b, BMJ19, Eba11, XZ18]. multisymplectic [MKHC11]. Multivalent [ABR10, AS10a, ASMEE11b, EAAED10, Mar11b, ORD11, SK11b, Sok10, YY11b]. multivalued [And12, DKM17, KYR11a, KYR11b, LC11b, Qiu12, RA11a, XZC12].

Multivariable [RMS10, dSCM12, GD12, LSD10, TÇA11, ZDL11, ZZHF12]. Multivariate [Ana11a, BGIN13, LKCN19, MD10, QZY11, Seg19, yS10, WMZW11, XC11a]. Müntz [ESL11]. MUSCL [KS15b, ZÖXL^{+}19]. muscle [GZ10, LTL^{+}12]. MUSIC [Par17a]. MUSIC-type [Par17a]. Musubi [QKR19]. mutant [LNP^{+}12]. mutation [Gal12, LY12b]. mutations [BD11d]. Mutual [CJP12, CJP15, MLY18, SPLHCB14]. MV [JKK10]. MVIM [YMDZ10]. Myoelectric [RIW12].

nabla [Ana10]. Nagumo

[AAA12, UKYK17, Zha18h, ZS15, ZSY19]. naive [And12]. nano [Eba14, PvdM13, RJGS^{+}19]. nano-particles [Eba14]. nanobeams [BC15, Kia19]. nanocantilever [NGG12]. nanochannel [LZB15]. nanoflows [RR14]. nanofluid [AG11, AR17, BTEM19, PZL^{+}18, RSL^{+}18, RKA^{+}18, RBB12, Ran15, REHA11, SRRP18, UABK16, ZJB19, ZSLZ19]. nanofluid-saturated [UABK16]. nanoindentation [CD12]. nanoinks [WA19]. Nanoparticles [ZAK18b, JFS14, JFS20, WL11d]. nanoplates [WL17b]. nanoporous [XLF12]. nanoscale [Hei10b]. nanoscience [KYAA10]. nanostructures [HLY16, PHM^{+}19]. nanotube [Kia16, RYK13]. nanotubes [Kia18]. nanowires [SKCL19]. narrow [LLCG16]. Nash [KJA10]. native [BCG17]. native-MATLAB [BCG17]. Natural [HSK11, LZL^{+}13, MSV18, REHA11, YWL^{+}11b, AG11, CKSL^{+}14, CCSZ14, GA10, HHM12, Ima17, JZ13, LZ19a, MT10, QZ16, RKA^{+}18, SZP^{+}11, Tro13, XXG10, ZHZ14, ZHJ14]. nature [GÇK10]. Navier [FZ17, AT17, BT15, BS16b, Bis14, BNP18, BPF13, DM19a, DNZ^{+}13, DLQ16, Dua18, Ers16, Fis18a, GRBT16, GOS18, GH18, HS13, HSJ15, HZM11, JJH16, JYYL16, KLRW12, LHL12b, Liu16c, MC10a, MM16, PLKC16, PGW19, QZM17b, RTV17, RA19, SMM19, SYI12, SHH16, Wan16a, WMP^{+}19, WH14, XSYL19, Yan18c, YT13, YTZ17, YT18b, ZZ16b, Zha18g, ZWY19, ZP18b, ZP18c, ZS18]. Navigation [LLWZ11]. Near [AM15, Por18, HR14, Oru19, ySGL^{+}10, SZ14, YH12, YL18a, YCLY15, ZLMZ18]. near-singular [YCLY15]. Near-stability [AM15]. near-wall [YL18a]. Nearly [Pan11, BS18a, JKS12, LS16, TPHD18, ZRE16]. nearness [LZ19b]. Necessary [AD12, ZFZ10b, MGW11]. necessity [TMO13]. negative [EM14, LZD17, Mar11b, SK11b, WS10b, ZCH14, Zha19c]. negotiation [KC12]. Nehari [GTC18, QCT17, Tru19]. neighborhood [LW11c]. Neighborhoods [YY11b, SK11b]. neighbors [Pas14]. nematic [ABL10, GLR13, GFZ16, LZG13, MMRN12, ZDB19]. Nernst [YQWZ19]. Nested [ZT13, Hu15, KM14, ZT16a]. net [AF13, AM17]. Network [BCPS15, CZ10, SLCC12, Ana11a, Ana12, BPKM10, CSSW12, CHY12, DHQ11, FCZ12, Ge10, HHS^{+}10, KS10b, KY11, LYM12, LJYS18, LCT12, LLX^{+}10, MSH10, ORR16, RPTD10, SSESG12, SXB^{+}12, SCKH10, SKH12, TMMASG10, TL10b, VMAVGCM19, YC10a, YLC12, ZH11, Zha11b, ZPS^{+}12]. network-on-chip [SSESG12]. networked [CMS10, YHZY11]. Networks [GACMO13, Ahn12, Bai11a, BV11, BKR11, BR12b, BDPM12, CQLX11, CHY12, CL12c, CCK12, Chu12a, CP15c, CC17, CCP19, Das15, DW18b, DHGF17, DA12, HTGSH13, HB12a, HIS19, HMM12, HCL12b, Hua11, IQR16, KCL12, LLWZ11, LZT11, LHW11, LZC13, Lin11, LRV13, MSA12, MKA^{+}10, PHPK12, PFBL10, SPS^{+}13, SYO12, SD12b, SLXC11, SSM12, SI10, TKH10, Tod12, WL11a, WL10, XC11a, XY11, YYK16, YXP^{+}13, YDK^{+}12, gZnZpZbD12, ZSY14, ZPS^{+}12, dlPBTW13]. Neumann [AER12, Arq18, DG10b, HCF16, KSF14, KPG18, Pov12a, RSWZ10, WW18b]. Neural [ORR16, Ahn12, Ana11a, Ana12, Bai11a, BV11, BKR11, DHGF17, HPR19, Hua11, HLC11, KS10b, KY11, LHW11, LZC13, MSH10, MSA12, SYO12, SI10, TL10b, XC11a, XY11, YYK16, YLC12]. neuron [MRS^{+}12]. neuronal [KS15b, RS12b]. neurons [CZMZ11, XY11]. neutral [AZH10, BD11a, BD11b, BV11, BH10, Can11d, CP15b, CW10b, Don10a, GGEB12, JS11, LW11b, LH10a, LKU10, LWKK10, LZKU11, LY11c, PR11, YX11a, ZYG10, ZZ11a, ZL11, ZJ10c, dSAC11]. neutral-delay [ZL11]. neutral-type [BV11]. Newmark [DM10]. Newmark-diffusive [DM10]. Newton [Lee15, AS15b, Cha18, CM11c, CW18, CChL14, DNS15, Gal11c, Gon13, KG11, LL19a, LHY11, LBJ10, LHL12b, NZ14, SNH10, SG10a, Śmi11, SL16c, XX17, ZLG18, Zha19b, Zho16]. Newton- [Lee15]. Newton-type [NZ14]. Newtonian [ATO19, DKM17, HHS^{+}17, HZLM10, HWY14, KRCJ11, KDG11, LX16, uHS12, MCR11, Mom11, PM10, REHA11, Sah11, ST14, WH11a, YM13]. next [He11]. nexuses [THB12]. NF [KPG13]. NFEM [GS19]. NHSS [XWY18]. Ni [WRW^{+}19]. Ni-based [WRW^{+}19]. Nicolson [AY12, BN14a, CY19, DM15, DH18, GH18, HA10, ZLJ^{+}18]. nilpotent [YH12, ZL12b]. Nitsche [WC15, dPLM18]. NLEQZK [ZSAN19]. NLmZK [ZSAN19]. NLP [GR13a]. NLS [ZSAN18]. NMF [IK12, YY12]. NMF-based [IK12]. NN [Pes13]. No [CZMZ11, BRR16, GB16, Liu16b]. No-chattering [CZMZ11]. no-flux [BRR16]. no-slip [GB16]. no-stationary [Liu16b]. nodal [AEH18, CWHW17, Rus16, RC18]. node [FF15, FES^{+}19, VMFF18, Wei12b, ZN18, ZSQ^{+}18]. node-foci [Wei12b]. nodes [BS17, CQLX11]. Noise [BG10b, Bai19, BR12d, BM19, DCKY15, GB18, KTDT17, LD11a, Li18c, LX15b, LHLH15, MX15, MT19b, SM17, TKBMT17, UCK16, XSYL19, YZM^{+}19, YL14, aZW17]. Noise-induced [BG10b]. noising [BPM12, MSTB17]. Noisy [SSK13, ZL13]. Non [ABL15, BMAR18, BDO11, BM18c, BK12b, Ers16, GDM13, KPP13, KHWK10, LZ18a, LZZC12, MHM11, Par11a, SP10, SDH15, WF17, Zho16, ATO19, ARK13, AYH17, AR17, AM12c, BH10, BMTV12, BBR10b, BGRV15, BGPP11, BSS18, BE18, BFS15, CGGM19, Cao19, CCJP11, CZN11, CLB14, CL16a, CZL17, CLMM18, CCK18, CT10a, CKR10, DSL11, DHMU16, EAEH18, FZ14, FWZ16, Fio14, FHS18, Gen11a, GHC^{+}15b, GFZ16, GMZ15, HHS^{+}17, HT16a, Hu15, Hu19, HLY16, HZLM10, HWY14, IC12, IW18, JK10, JGK13, JS11, JZL18, JQG14, KP10a, KRP12, KRCJ11, KBAF18, KDG11, KNZ19, LLW11, LA11, LCK13, LZD17, LSS17, LJYS18, LHY18, LM18b, LX18, LZ19a, LZF18, LLP19, LV11, LRTV10, uHS12, MCR11, MGN^{+}16, Mia18, MPGW19, MMH11, Mom11, MS12b, MSFS18, NPR10, OBCG19, OBAAD10, Oru19, PR11, PP14, PT15, PM10]. non [QCT17, RMM11, REHA11, RHD18, Rua19, Sah11, SMM19, SKG^{+}11, SDH13, She18a, ST14, SL16c, SST19, SKFG11, SR18, SKH12, Tia19, TAA14, WH11a, WWB13, Wan16a, WL17a, WHS17, WZ18a, WBZY18, WW19c, WZG19, WL10, WY11b, XGH17b, YY15, YJ19, YTF10, YL14, YLG17, YM13, ZCSG13, ZZ16a, ZTZ16b, Zha18h, ZZL15, ZZ16c]. non-absorbing [RHD18]. non-alternating [LHY18]. non-arbitrage [CCJP11]. Non-Archimedian [SP10]. Non-auto [WF17]. non-autonomous [CZN11, CZL17, LLP19, PP14, WL17a, WZ18a, WW19c, YL14, YLG17, Zha18h, ZZ16c]. Non-body-fitted [LZZC12]. non-chaotic [OBAAD10]. non-compact [CT10a]. Non-conforming [SDH15, WBZY18]. non-connected [EAEH18]. non-convex [CCK18, MSFS18]. non-cooperative [CL16a]. non-crystalline [SKH12]. non-delay [WL10]. non-dissipative [LSS17]. non-equilibrium [JQG14]. non-exhaustive [WY11b]. non-existence [QCT17, She18a]. non-extensive [LZF18, WZG19]. non-Fickian [KRP12]. non-flat [IC12]. non-harmonic [LCK13]. non-Hermitian [Cao19, FZ14, FWZ16, GHC^{+}15b, LM18b, Mia18, WWB13]. non-homogeneous [AYH17, CLB14, Wan16a, ZZL15]. non-ideal [BSS18]. non-integer [AR17, CKR10]. non-integrability [LV11]. non-integral [Gen11a]. non-isoparametric [Rua19]. non-isothermal [GFZ16]. non-linear [BMTV12, BGRV15, BGPP11, Fio14, IW18, JGK13, LLW11, MGN^{+}16, MS12b, OBCG19, PR11, SL16c, SST19]. non-Lipschitz [BH10, JS11]. non-local [ARK13, HLY16, KP10a, YY15]. non-locality [Hu19, LRTV10]. non-Markovian [LJYS18]. non-matching [BE18, HT16a]. non-monotone [Tia19]. Non-monotonic [ABL15]. non-negative [LZD17]. non-nested [Hu15]. Non-Newton [Zho16]. non-Newtonian [ATO19, HHS^{+}17, HZLM10, HWY14, KRCJ11, KDG11, uHS12, MCR11, Mom11, PM10, REHA11, Sah11, ST14, WH11a, YM13]. Non-normal [BK12b]. non-orthogonal [BFS15, CLMM18]. Non-oscillation [Par11a]. non-oscillatory [LX18]. Non-overlapping [BM18c]. non-parametric [Fio14]. Non-periodic [KHWK10, LZ18a]. non-periodicity [BBR10b]. non-preemptive [LA11]. Non-prismatic [BMAR18]. non-radial [ZTZ16b]. Non-recursive [KPP13]. non-smooth [DSL11, DHMU16, JZL18, KBAF18, LZ19a, SDH13, YJ19, ZZ16a].

Non-standard [MHM11, MMH11, RMM11, SMM19, XGH17b]. Non-stationary [Ers16, CGGM19]. non-strongly [PT15]. non-symmetric [FHS18, YTF10]. non-traditional [WHS17]. non-traveling [GMZ15]. Non-trivial [BDO11]. non-uniform [AM12c, JK10, KNZ19, MPGW19, Oru19, SKG^{+}11, SKFG11, SR18, TAA14, ZCSG13]. non-uniformly [NPR10]. nonanalytic [DBS12]. nonautonomous [BCF13, CTA12, TZWM10, TZG10, WW18a]. nonbreaking [FER15]. nonclassic [Dua11]. noncommensurate [JC12]. noncompact [CLCF14]. noncompactness [MN10c, MKPS11]. Nonconforming [GLZ18, KDU15, SWL15, ZMFL18, GS15b, HMWZ16, Li15, QCLC17, RA11b, SSL14, ZS13, ZY19]. nonconservative [YÖ10]. nonconvex [ALI11b, Bai19]. nondecreasing [LZ11d]. nondifferentiable [AT11, Ant10]. Nonexistence [AJS19, JKS18, JPK18, XWY11]. nonexpansive [AE12a, Buo11, CGY11, CGHY11, CS10c, CS11d, CS11b, KK10a, KK11b, KKS10, KA10b, KYO10, KYR11a, KYR11b, KT11b, LjHO10, Mai10, Mar11a, Nil11, Pir11, QHW11, She11, SCC12b, ZS11a, ZLL11]. nonhomogeneous [CHZ19, DMV11, LWBW13, MR14, WW18b, ZZ18b, dSSV17]. noninteger [MCP13]. Nonlinear [CFN11, FP19, GF16, LXP11, LS12a, LS12b, MBS17, MH11, MG15, MTAS17, SW10, WZF12a, YK18, YWHC11, Aba10b, Abb11, AM13a, AIA13, ABJ11, AW11, ATUC15, AÇT11, AA15, AJAR18, ABCR10, AER12, APwS18, Asl11, APTZ19, BMRA10, BM11a, BD11b, BR12a, BAO^{+}12, BCC14, Ber12, BDGS13, Ber16, BC10, BG10a, BE12, BDO11, BWZ16, BB08, BGL^{+}15, CL11, Can11d, CH11a, CCR16, CCRS17, CGY10b, Cha11a, CN13, Cha11b, CC19, CDW11, Che11d, hCTM11, CLTA11, CM11b, CQ13, CW18, CCZ18, Che18, CAP10, Chu11a, CP16c, CCP19, ĆSCD11, CRRS11, CFF15, Cve11, DWY15, DL10, DN18a, DAM14, Dem10b, DG10b, Din13, DL14, DSM18, Don10a, DLC19, DC10, DL18, DCRL13, Dub13, DNR13, DZ16, DDK11, EO14, ESN10, EZRR10, FM19, FT10, FSHZ11, FN14, FH16, FIW17, FIS18b, FP18, GCR^{+}18]. nonlinear [GRBT16, qGpWhL11, Gen11b, Gep16, GMZ15, GW15b, GML17b, GTZ19, GM11, Gur13, GN11, HH16, He11, HmZ11, HYCP11, HP17, HS12, HB12b, HA18b, HHGA19, HXL11, Hu18, Hua10a, HTWS15, HM15, HM19, HD19, Ibr11, Iom18, JKK11, JGSS10, JKK12, JKN10, Jan10, JPB11, JJC11, JPK18, JMB10, JJ19, JKS18, KA10a, Kar10a, KL12a, KK13a, KK15, KFYW11, KW11, Kig10, KJ11, KG11, KPK18a, Ku15, KS12b, KCK19, KY10, LX10a, Lan12, LK15, LKS10, LC10a, LK14, LR14, LHY11, LLZ10, LCN10, LCP10, LZC11, LW11a, LS11a, LY11a, LZ12a, LZWC16, Li17b, LX17, LD18, LZ11e, LK18, LBJ10, LC11b, LKU10, LWKK10, LJ10b, Liu11a, LXK11, LZKU11, Liu13, LLY13, LHL15, Liu15a, LDL^{+}15a, LL15, LZ15b, LDL^{+}15b, LTJ^{+}16, LC16, LL16d, LW18b, LLLW18, LWSL19, LW19b, LGVS19, Luo18, MZ10, MX10, MM10a, MKHC11, MPZ11, MC10b, MN11a, MTM11]. nonlinear [MBH11, MAS11, MPfTX18, MAST18, MV11, MP19c, MDL18, MI16, MM18d, MGB^{+}11, MLGY16, Mor13, MY13, Ned12, NDT11, Oru17, Ouy11, ÖZ11, PZA19, PPD10, Pit12, QAA^{+}16, QZM17b, RKP12, RMA10, RY11, RZ16, RA18, RKF18, Ray16, RG11, RTV17, SNH10, SD15a, SLK12, SD12a, Sam19, SCSF19, San12, Sea14, Sea15, Sea16, SIL19, Sed13, Sha18, SMDI18, SNDK18, SBS12, SG14, SS14a, SD19, SG11b, SZGG11, SW17, SGZW18, SG16b, SG10b, SLYY13, SM14, zSdZ10, Sun10, SLL12b, SRG16, TN11, TZZ11, TTG16, Tan17, TTM19, TG11, TG14, TF17, Tao18, TS11a, TAS11, TA11, TÇA12, TY16, THY^{+}10, TTC14, Tur10, WKG10, WHS11, WhJxLwW11, WWW11a, WWW11b, WWA11, Wan11, WAZ11a, Wan12, WKS13, WL15, WV16, WC17, Wan18, WH18, WLL^{+}18, WZG19, WS10b, WFL11a, WAW15, cW11, WCLD18, XWY11]. nonlinear [XY15, XYXZ16, XC18, XZ18, XY10, XGH17a, XGH17b, XXH18, Yan12b, Yan15, YGS^{+}16, YZ18, YQWZ19, YC10b, Yaz11, YP10, YAS^{+}11, YMLL18, YTC^{+}18, YASK10, YLDL11, YSB15, YHC18, ZA10, ZLW10, ZM18, ZYG10, Zha10, ZQ11a, ZBF11, Zha11c, ZS11b, ZD11, ZHJ11, ZGW11, ZLL12, ZLL14, Zha17b, ZM17a, ZLT18, Zha19a, ZY19, ZCH12, ZL12a, Zha14, Zha18i, ZW16a, ZJZ18, ZWJ^{+}11, ZW11d, ZLW19]. nonlinearities [AAZ10, BS16a, Gao17, GKLR11, LS17a, Luo19, MAST18, Sha18, SK11c, ZJ10b]. nonlinearity [ARK13, CT18, D’A18a, DZ21, GM11, HY10, HGW18, LRZ18, LLY13, MF18b, MF18a, ML19b, NT17, QAA^{+}16, QHT16, Sea11, SMBY10, Su12, Tan17, Tru19, WS12, ZTZ16a, ZLG18, ZT16b, ZT18a, ZW11d, Zho19]. nonlinearly [AG10a]. Nonlocal [Kia18, Tod15, WZWX11, ZL10b, AW11, AAD17, AT18b, AT19, ABLS15, AG10b, BK11a, BKT11, BS16a, BL18, BC15, CW19, CP15b, Cha18, DG13a, DPBL16, DB11, DS18b, DSL18, FG18, GJ19, GF16, Goo11a, HLWX11, JL11a, JKS18, KSZ18, KB15, KHF^{+}19, LGHR16, LSW10, LS19b, LZ14, LLZ18, LH10a, Liu13, LYZ17, LWSL19, zLYLQ19, LDG19, MA17, PH19, Pov19, QRMH18, Saj12, Saj14, SW12, SOK19, uIAA15, SZW11, eT10, Tia19, Tod18, TNHK19, VDB13, WC10b, WF17, Wan19b, XZR16, XSYL19, XZ19, YW10, ZLY14, ZZWG16, ZCZ17, ZLG18, ZL19a, ZSW19, ZWMD16, ZcHS18, ZG18a, ZLY17]. Nonmonotone [AA10a, HSD10, BDS17, BWL18, DKM17, TNP17]. Nonnegative [JD12, Mis14, BM13b, JL11b, YXX11]. Nonnegativity [DGZ13]. nonnormal [KSKK11]. Nonoscillation [MR10, FL10]. nonoscillatory [LKU10]. nonparametric [LWWY12, QZY11, WZ17a]. nonparametric-based [LWWY12]. nonpolynomial [SKM11]. nonrectangular [AD19a]. nonrigid [PZJ^{+}16]. nonself [KT11b]. nonseparable [Ali15]. nonsingular [HWXC16, LM18b, SS16b]. nonsmooth [LW13, SbX19]. nonsmoothness [FP19]. nonstability [BPX11]. Nonstandard [YZMZ16, AGPCC10, CLM12, CLM14, Gur13, Lóp19]. nonstationary [GH18, Ngo18]. nonsteady [BPR18]. nonsymmetric [AT18a, BJS15, CM16a, CBS18, HH11, HWXC18b, LZ15a, LBW11, SJHC14, Tan18, YH15]. Nontrivial [ESN10, CCX13, QCT17, ZJ17]. nonuniform [LNW19, PC17, WCQ^{+}19, YT13]. nonzero [KMT10]. Noor [Lee11a]. Noor-type [Lee11a]. Nordhaus [AH10c]. Nörlund [Bor10]. norm [AHV10, CCKP15, DCKY15, GWL11, He16, HM17b, Ipe12, Izs15, KSKK11, NDT11, NCC13, OÖ11, WWW14, WZ10]. Normal [BZKR15, BS18b, BK12b, LWC13, MSQ^{+}11, SLK12, XHH^{+}19, hYlLL11]. normalistic [SA11b]. normalization [WC11b]. normalized [HP10, MN10a, XC18, Ye17a, pZ10]. normed [Deb12, GKK11, MME10, SP10, ŞGY11, Yil10]. norms [BX10, CHBTD14, EM14, FM18, FR16, HR19, Liu11b]. North [RB19]. Note [BPX11, HKS14, Jun10, KK11a, ZYY10, AR10b, Ban13, BMY13, BY11, CTP10, CSW11a, CHS19, FDB13, Hai10, JS11, JY11, Lee11c, LFJ12, Min11, MT18, Özu15, RR11, SAIZ15, Tom11, Von11, WC10a, XZ10, YT12, YH11a, YS17, ZLZ10, dPLM18, uRK11]. Notes [BG11, CM11a, hCTM11, fDxZ11, GDZ11, HLY12a, zLZ19, NKA18, PV12, WWW11a, WYG12, Wan13a, XWH16]. notice [JL20, JFS20, Pen11, SK12]. Novel [CM13a, CP10, LB11, LWW19b, Man18, SYO12, CCKY12, CAC14, DCG^{+}12, FPW^{+}11, GLLC19, HT18, HT12b, JZR15, KSPP11, KCL12, LW11a, LLFT17, MS17, MBKK10, OMS10, Par17b, QCYL12, Sha12b, TT12, Von19, WC11a, WLZ^{+}18a, WZC^{+}19a, XFH19, yYsZyYL13, ZA15, ZLPM13, ZCY16, ZLGL11]. November [Ano18-60, Ano18-63, Ano19-60]. nozzles [FLZ14a]. NSFD [MW13, Was13]. NSGA [LLW10]. NTP [ID10, JW05]. nuclear [GIMZ14]. Null [Gao15, YT18a, HKK^{+}16, ZG14]. null-solutions [HKK^{+}16]. nullcline [HNK13]. nullcline-based [HNK13]. number [AH10a, BC11, BR12c, Bri10, CKMR11, sCYhX18, DZW^{+}15, Dol11, Far11, GCG12, GOGYL^{+}11, GSD^{+}19, HVA10, HP19a, HL12, IB10, Koj10, KP10b, LHH10, LYN11, MKL11, MMR10, MMR11, Ned12, NM11a, NDC^{+}19, PDN19, QZ11, WW10b, Wan10b, XY11, ZM14, ZD15]. number-valued [BC11]. numbers [BBC^{+}11, BS16b, CHM^{+}10, Cvi11, Dia17, DRS11, EMRS12, GSS11, HBK^{+}19, Haz11, Ipe12, KSKK11, Li10a, LLL11, LH12a, NM11a, Sae11, Sun11, TD10a, VP11, XCXW10, ZP10]. numeric [DR12, EZM12]. numeric-analytic [EZM12]. Numerical [AB10a, AM13b, ABK^{+}13, BMAR18, BPM14, BDS17, BCFQ19, BRFH16, BGRV15, BSL11, BGP13, BMSS18, CM10a, CYP16, CCJP11, CHXL18, CLTA11, CLB14, CM18c, CR13, CLJ11, CJPB10, CFF15, CS13, CLL11, CM10c, DNS16, DA16, DZ18, DGT18, DSL11, Dun18, ESL11, FSRB15, GGM^{+}13, Gen11a, GIM15, GTL16, GMI11, GDF12, HB19, HGHA19, HKI12, HHS^{+}17, HJD15, HPR19, HTL10, HLL^{+}15, HCLL18, HD19, HZLM10, IB11, JO19, KMT19, KMRN12, KSF14, KKSK18, Ku15, KLRW12, LS10a, LK13, LZC11, LDW11, LS11b, Li16b, LFC16, LGH^{+}11, LHM11, LZB12, LvS15, Mai16, MC11, MTN19, MM18b, MAS11, MP19a, MD15, MEAMHHV18, OHK^{+}19, PDN19, PN16, PvdM13, RS12b, RBB12, Ran15, RA18, RSB14, RCG15, RRO17, RZZ19, SSS16, SD10b, Sal16, SHM13, SYI12, uIAA15, Sou11, SM10, SM14, SK14b]. Numerical [TY13, UABK16, Vel15, Wan19a, XFH19, XWN11, XHA13, XWL18, XHH^{+}19, YYK16, Yan18b, YJ19, YZ15, YTC^{+}18, YSS11b, ZA10, ZZWG16, ZLTY16, ZTY^{+}19, ZLA17, ZSD10, dBD17, AD15, AJRWS12, AJAR18, AM14a, AHO16, AS11b, AWJH19, ATH18, ABSV18, AB16, AGPCC10, ASSV18, AuIK11, BM11a, BC16, BHM19, BA11, BBDS11, BP11b, BMTV12, BPS18, Bho14a, Bho14b, BZ18, Bic11, BR12c, BM18d, BHJ14, BN14b, Boy16, Bra10, BMP15, BK12b, CCR16, CD19, CCRS17, CC11, Cie13, CJ15, CJCV10, CJRR11, Cos18, CN11, DMP18, DGL12, DGZ13, DD19a, Def10a, Def10b, DAM14, DM16, DNP15, DTR19, fDxZ12, DFW^{+}18, DNR13, DHMU16, EMQ18, ECJ16, FJC16, FGHZ13, FER15, FSB17, FGL10, GSI19, GMP18, GHT^{+}15, GR13a, Gha18, GY11, GALO18, GDM13, GW15b, HCL12a, HG18a, HY16, HO19]. numerical [HHNLGC18, HA16b, HLY16, HT16b, HKW15, ILS13, JKN10, JL17a, KP18, KMS10, KV17a, Kat11, KHF^{+}19, KAS11b, KB13, KSMN11, Kup10, LLJK10, LBZL11, LY12b, LJSK13, LZ16a, LSZ16, LNW19, Lin14, LLG10, LCWZ18, LRBA15, LSV18, LDY11, LDL10, MDRRV11, MKR12, Mar16, MGY11, MMFT^{+}19, MN10a, MT19b, OTiSY16, Özu15, Pan11, PMM17, Per18, Pet11, Por18, RKP12, Res16, RTRR18, RTL19, SSM^{+}17, SCSF19, SS14a, uIAH10, SKTD13, TS14, TMDTTC16, Tom11, UKA15, VGK^{+}16, VA10, WW14a, WRY18, WKBR18, WZKY12, WHW11, XG10, YT18a, YYYH19, Yüz11, Yüz12b, ZTR11, ZLPM13, ZH15a, ZZLB18, ZHS^{+}19, ZYZ11, ZL12a, ZD12]. Numerically [BDM19a, DH17, JL19, JL20, MBHV10, Oru19, TC10]. numerics [MUB^{+}16]. NURBS [HH18b, Kar17, SH18, WZ16, WZC^{+}19b]. NURBS-enhanced [WZC^{+}19b]. Nutrient [MKG13, DKG14, XBHN16]. nuts [Hua12]. NVIDIA [LYC15]. Nyström [AOW18, DLS14, KMS10, KAS11b, MBJ16]. Nyström-based [AOW18].

O [Pan17, ZAK18b, XA13]. object [HH18b, HCL12b, LKK12, Med12, YC12]. object-oriented [Med12]. Objective [GACMO13, BMM12a, CCFV12, DM12a, DMPV10, GWL11, KGJ11, Li12b, LBH^{+}12, dCMdSGTdC^{+}16, RPTD10, RFP11, RF12, Smo17]. objects [CHS18, KLMV12, WYY11]. oblique [LHM11, WL11d, WCH13]. obliquely [LFC16]. oblivious [HSMY12]. observability [Guo12]. Observable [GD12, CCJP11, Gal10b]. observation [CS14b, ZL19a]. observations [CLJ11]. Observer [Yan12a, CZY13, JCWZ16, LZY12, THY^{+}10]. Observer-based [Yan12a, JCWZ16]. obstacle [AD16, BS14a, BS15a, BLS18, BV17, GG18, GM19, LK14, eOS18, ZSZ18]. obstacles [DJD18]. obstructed [ASFM15]. obtained [AHF10, Krn12, SKST10]. Obtaining [GH14, ED11b, HM15, TCM18]. obtuse [KYW^{+}18]. ocean [eMA18, FIW13, LPLR19, Liu18b]. oceans [LZ19c]. octahedral [HR14]. Octave [FRAK15, JRA^{+}18, RAW^{+}16, Váz16]. October [Ano18-58, Ano18-59, Ano19-56]. octopus [KBGC12]. octree [QKR19]. odd [HL11a]. ODE [CP16b, HL11a, RY11]. ODEs [ATUC15, FRSW11]. OEMA [WCD10]. off [Ikh11, MS15, SRS11]. off-step [MS15]. offset [LCT12]. offspring [ASB12]. oil [Mok11, WHD14]. old [PBS12]. Oldroyd [FZBF10, FAIV10, JKZ11, LZZ11b, SISH12, WSW11, YZ19, ZJB19]. Oldroyd-B [FZBF10, FAIV10, JKZ11, LZZ11b, YZ19, ZJB19]. OLSR [KHIB12]. on-demand [JCZZ13]. On-diagonal [Wu18b]. On-line [CQ11, ZZXY12]. On-site [LLL13]. One [THC^{+}18, ADS14, Bac14a, BQS16, CNV14, CWH13, GGR15, GYH11, GK11a, HBE15, HW11, HA10, KYR11a, KYR11b, KL12b, KPG18, LW10, LCW17, Liu18c, MY16, MMRN12, MBS17, Mai16, MHL11, Mor13, RNQ13, SD15a, TT14, WZH10, WCW13, ZY10a, ZY11, ZD12]. one- [HA10]. one-dimensional [BQS16, CNV14, CWH13, HBE15, KPG18, MMRN12, TT14, WZH10]. One-point [THC^{+}18]. one-predator [ADS14]. one-sided [ZD12]. one-step [KYR11a, KYR11b, Mai16, Mor13]. one-time [LW10]. one-variable [GK11a]. Online [KWPK13, KSMT11, LLD10]. only [HW11, Kup11, Tsi11, Wei12b]. onto [DVM12]. Open [KW12, ACE17, CV14, CHH14, HMP^{+}15, HT12a, LVF^{+}16, LWBW13, LSZ11, Par17a, YZ15]. open-channel [CV14]. Open-loop [KW12]. open-source [HMP^{+}15]. OpenCL [KB13, MSZG17]. operating [PTL13, SKH12]. operation [ZLY^{+}13]. operational [DBEE11, ED11b, KKBR19, KMRN12, LS11b, MKR12, SD10a, VBCJ10]. operations [ASN11, GXZ10, Ji14, JTCC11, SA11a]. Operator [BCF13, GR19a, AM10a, AES11, AS10a, AS10b, AS11c, ASMEE11b, AKA11, BCK11, BX10, BD11c, CH11a, CGHY11, Che14, CWY19, CH11b, DG13a, FRAK15, Fra11, FMPR15, FBB10, GTG11, GSD^{+}19, GNP14, Hu15, KSJ12, Lad16, LY12b, LG12, LS10c, LJX12, Lup11, Mar11b, MPY16, ODR10, ORD11, ORR16, PA12, QYL10, RAW^{+}16, SPL19, San12, Sat11, She18b, SRM11b, Sok10, Tac11, Tod15, TDM13, VLJH18, tWqLzGkP11, WL13a, WWW14]. operator-difference [AKA11].

Operator-splitting [GR19a]. operators [AD10a, AD11a, AM10c, AD10b, BK11a, BC11, BM10b, BM11b, Ber12, BBD10, CMGR11, CGY10a, CT10a, CDD12, DK12, Dra11a, Dra11b, Gao12, HT12a, KD10, KD11, Kir10b, Lad16, Lam12, Las10, LS10d, MB11, Mah10a, Mah11, MCP13, MKL11, MKPS11, NPD17, NNAS11a, OÖ11, PB12, Qiu12, RA11a, RRGTV10, SOJC10, UBF11, VSI12, WD10, WXF10, WZWX11, XJLX10, Yao10, ZM18, ZC10, Zha11a, ZSS10]. opinion [MB17]. opponent [KSM12]. opposite [Abd18a]. opposite-bordered [Abd18a]. opposition [DWI^{+}12]. opposition-based [DWI^{+}12]. Optical [KY10, Hei10b, HTWS15, KB10b, LTSW16, YTC^{+}18]. optical-fiber [HTWS15]. optics [MPfTX18]. Optimal [AMA14, AR10c, AIIZ10, AT19, BO10, BMM12b, BBO10, CL12a, CRXL15, DH18, DMV11, FMPR15, GRW14, HLZ15, IIHI10, KYA15, LS11c, LSS17, LMZ17, LZF18, MP10b, MM16, Mop11, MN11b, MT19b, PM10, SISH12, SV11, SBA10, Sul16, TZ18, UKYK17, VMAVGCM19, Yil19, ZHT11, ZSY14, Zha13, ZLG^{+}10, Akm15, Ala10, Als10, AEG11, ADS14, BNTT14, BQS16, BC17, BPKM10, CCN14, CY14a, CLH13, CHLY15, CHL18, CCK18, mCfX10, Chu11c, DL19, DO11, Def10a, Def10b, DRZ10, EF14, EG10, EE18, EUTS18, FHH13, FES17, FGHZ17, FES^{+}19, FOS19, GK11b, GSR14, GM18b, GS15b, GS19, GNP14, GH18, GD16, HKI12, HM10, HK15, HR14, Hou15, HLY17a, Ibr16, IIHuI10, IBSS11, JLL18, Jum10, KBS11, KM12, KP19b, KKL16, KKL^{+}13, KRBS18, LCH19, Li10b, LZZ18, LXP11, Liu15a, LDY11, MH11, MO14]. optimal [NMR15, NCC13, eOS18, Pan11, PA15, RGVR17, SAIZ15, SBS12, SbX19, SW16b, Tod13, TC10, VCM11, Vel15, WZWX11, WX18a, XL15, YMSL11, YC11, YLH12, YSB15, ZZ10a, ZG16, ZZX16, ZP18c].

Optimality [Mar12, CNV14, CGH14]. optimally [MS18]. optimisation [KTH13, Tod12]. optimistic [HCL12b]. Optimization [BR16, CCCW10, GGGR13, KP19c, Laz10, MCP13, TM19, VBW10a, VBW10b, APT11, Ahn12, AA10a, AG10a, AGDP19, AdSSS19, ABSV18, BMS19, BBO10, CI18, CChL14, CCFV12, DCG^{+}12, DK14, DMRS18, DNS18, DLQ16, DH10c, EHO^{+}12, FJB19, FLH10, FL11a, FBL11, FMP19, FLWJ11, GYTD12, GYH11, qGpWhL11, GALO18, GWL11, HSD10, HGJP19, HRMS10, HRMS12, HB12a, HM17d, HWY14, JPCY13, KGJ11, KM18, KL12b, KMS15, KFTT13, LWR16, LLH14, LYJ15, LG17, LY12b, LTX^{+}13, LH19, LLH10, LGL^{+}14, LW18c, LFZ19b, Mar12, MHL11, OMS10, PH13, PW10, PG10, PM10, PSS18, RR11, RGdSRLAJ10, SLMZ12, SJS^{+}10, SW10, TWLYÖ10, TTC14, VHPVNXW18, WXYW11, WWZ12, WQRZ14, WZ17b, WMW13, WQNF12, XLZW11, XDL12, Yoo17, ZB19, ZLGL11]. Optimization-based [BR16]. optimize [CIN^{+}18, LCW12, RJGS^{+}19]. Optimized [DSK^{+}14, KZ16, KAS11b, LKLP12, Mah14, SB14]. Optimizing [SXB^{+}12, YS19b, LBH^{+}12, YW14]. Optimum [GGGR17, HT12b, SKH12]. Option [GSS11, CWW15, CW14, CM18c, CKM12, CJPR10, CJPB10, CEJV16, DT17, ECJ16, ELS11, FSRB15, FSB17, GW15b, HZ16, HLvS18, KKT13, KCC^{+}13, LvS15, MvS18, Moh15, Tha19a, YKA18, ZM16b]. Options [KDU15, ACAS11, ASSV18, BS10b, BC16, CXZ15b, CWDL17, CWY19, CGK14, Ciu11, DNS16, DH17, GLW18, HMV18, JYK16, Lee14, LW17, MZC17, MVKK14, SL16b, TC16, ZLTY16, ZBL12, ZH18b]. Orbit [Lam13, WX18b]. orbits [CT11b]. Order [GGAVRC^{+}19, ABM11, AM13a, ABT19, AZ17, AEH18, ADGS18, AHV10, AAZ10, ATUC15, Akm15, AÇT11, AH11a, AJ10, AJ11, AJ12, AS11b, AK19, ABL10, ATH18, ALI11b, Bac14b, BD11a, BD11b, BQ17, BMA11, BPM14, BLS18, BEAA11, BGIN13, BCK11, BDS10, BS16b, BBBM16, BZ18, BC10, BQS16, BG10a, BZKR15, BDO11, BDHR18, BWZ16, Bra16, Bra10, BS18b, BGM19b, BC17, BPX11, BMH19, BS17, CKSL^{+}14, CCKP15, Can11d, CLT^{+}13, CY14a, CGHW14, CO15, ÇT12, CW17, CB11d, CF16a, CM11c, CT11b, CLTA11, CLB14, CW14, CT17b, CS14a, CTM^{+}13, CFdM^{+}18, CM16c, CM19b, CSN11, CM14, CX18, CDD12, CT17c, CN16b, CKR10, Cvi11, DYH11, DZW16, DLZ17, DGTC13, DL10, DGL12, DN10, DN18b, DM18, DL19, DH11b, Dem10a, DG13b, DFG19, DGLS19, DGT18, DT17, DGK10, DBEE11]. order [DLS14, Don10a, DFW^{+}18, DM19b, DC10, DFP^{+}13, DNZ^{+}13, DR12, DHGF17, ED11a, EKS10, EK13, ESN10, EOM11, FDB13, FM19, FN14, FZL^{+}18, Fis18a, FM12b, FR16, Fu19, FKDN15, GY15, GM18a, hGzS15, GP11, GK13, GTG11, GK11a, GK11b, Gha18, GY11, Git14, GMI12, GS11b, Goo11b, GGEB12, GKM11, GGT14, GL16, GKS10, GJ10, GM11, GH10, GH12b, GH13b, HY10, HK10, HSD10, HLWX11, HS11a, HH17, HEP10, HLL13, HO10, HL11a, HL11b, Hes14, HZ11, HVO17, HM18a, HKKK13, HTV13, HHGA19, HGN^{+}10, HLL^{+}15, HG18b, HF10, Hua10a, HK17, IMD11, JGSS10, JA11, Jan10, Jaw13, JMB10, JW18a, JW18b, Jia12, JL18, JZ12, JC12, JZL18, KO11a, KJK18, KBS11, Kar10b, KM15, KKBR19, KS12a, Kig10, KG11, KPS17, Kim18, KS15a, KHF^{+}19, KPG18, KB13, Ku15, KS12b, KK14c]. order [KS15b, Kun12, LZ11a, Lan12, Lee15, Lee16, LZ10, Li10e, LCN10, LCP10, LF11c, LW11a, LL11, Li12a, LLC13, Li16b, LLFT17, LCQL17, LLY18b, LKCN19, LM19a, LZZ19, LJ10a, LMLB19, LLY10, LKU10, LWKK10, LXK11, LZKU11, LHF11, LZY11, LCZ11, LL12e, LHL14a, LX15b, LDL^{+}15a, LDL^{+}15b, LZJY16, LZLL18, LLLW18, LLYL19, Lóp19, LCC13, LYSZ19, LLX11, MX10, MKHC11, MB11, Mac12a, MCP13, MV10, MT12, MDG19, MPY16, MJWD19, MP19b, MNPD15, MHM11, MRS^{+}12, MM18d, Mor13, MBJ16, MM19, MB10b, NLA19, NPD17, Naw11, NCV^{+}18, NHH13, NH15, NZ16, NNAS11b, NB17, OSZP13, Ouy11, ÖZ11, ÖKJR19, PN10, PCO16, PC17, PS16, PR11, Par11a, Pet11, QZ16, RZ17, RSDR11, RES10, RR18, RG18, RCH19, RQ18, RR19, RR14, SD10a, SD10b, SMF10, Sal16, Sal10, STC18, SBvdV13, SAIZ15, SDH15, Sea11, Sea15]. order [Sed13, SG14, SL18b, SXM11, ySW10, SLW14, SW12, SGZW18, SSP13, SV11, SW11, Sou12, SKM11, SZ11, SK11d, TJ10, TZWM10, TZZ11, THC^{+}18, TMDTTC16, Tod15, TT14, Tre18, TC10, TMSO12, TM10, Tsi11, VLFS12, VRD11, WHS11, WS11a, WWA11, WW11c, Wan11, WJ11, WRW13, WV14, WZM^{+}16, WTM17, WLL^{+}18, WW19b, WZC^{+}19a, WLZ18b, Wei14, Wei17, WGY^{+}18, WX18a, XZC12, XY15, XZ18, XY10, XH11b, XC13, XW18, YXS10, YY10b, YX11a, Yan11d, xYsHjL11, Yan15, YZX18, YZM^{+}19, YZY10, YC10b, Yas12, YWH14, YTZ17, YSS11a, Yüz11, ZA10, ZSAN18, ZR18c, ZA15, ZSH11, ZYWZ17, ZFZQ10, ZJ10b, ZYG10, ZS11b, ZTH11, ZGW11, ZQ11b, ZL11, ZLL12, ZLPM13, ZCY16, ZY17c, ZLJ^{+}18, ZL19a, ZHL12, ZZX16, ZM16b, Zho19, dVDR18]. Ordered [AR10d, Ali11a, AKS11, APS12, CSW11b, ĆSCD11, DGB10a, DDD10, GRS12, HLS11, KPR10, NS11, Qiu12, RK10, SK10a, Sha10, SS11c, SCK11]. ordering [Ste16]. orderings [BS11a]. orders [BR12a, BMJ10, CDG15, GW15a, JC12, KAJ11, THD19]. ordinal [LLH10]. ordinary [AJY13, LeT10, Li10e, LF11c, RSDR11, SBEB10]. organ [LWHL12]. organic [ZZHF12]. organizing [Bur13, JWX^{+}13]. Orientation [FKDN15, Loh16, LC12b]. orientations [AGH^{+}15]. Oriented [Ste16, AGDP19, BDM^{+}19b, BWZ16, DPM15, Med12, SWOF19, TYY^{+}12, XSLS11]. orifice [YCS19]. Origin [TMO13, WCH18]. originating [KS15b]. Orlicz [Haz11, HCF16, Yao10]. Ornstein [Li18d]. Orthogonal [APRM11, YZX18, BT10, BFS15, CLMM18, DA18b, ED11b, FSZS18, MS11a, QX19, XSLS11]. orthogonality [RES10]. orthotropic [FG18, LTT13]. oscillating [Çak11b, CHLY15, FAIV10, ZLZG11]. Oscillation [BD11a, BD11b, BB10b, Can11d, Don10a, Dos12, FL10, HEP10, Hua10a, LLY10, LZY11, ÖZ11, RM17, YX11a, ZYG10, ZGW11, ZQ11b, AAZ10, AK10, CDM12, DDK11, GGEB12, KÖ10, LXK11, LS12c, Par11a, SH18, SK11c]. oscillations [BCD^{+}16, BG10b, FT15, HF10, LvS15, MNPD15, OSZP13, Yaz11, YAS^{+}11]. oscillator [ArEM10, CL11, hCTM11, CKR10, FT10, Gen11b, MMMG12, RKP12, RSS10, ZLW10]. Oscillators [Cve11, CT10b, EZRR10, HM10, RG11, YASK10, ZQ11a]. Oscillatory [GDM13, ANP19, KAS11b, LX18, PR11]. Oseen [BLyS18]. osmosis [YTF10]. osmotic [CIN^{+}18]. Ostrowski [Ana11b, Dra11b, FM12a, HN10, KBS11, Sar10, SS11b, Set12, TD10b, Von11, XZL10]. Ostrowski-Grüss [FM12a, TD10b]. Ostrowski-like [HN10, Von11]. Ostrowski-type [XZL10]. other [ZSAN18]. Otsu [LTL^{+}12]. outcome [KSM12]. Outer [AK11, Ji14, Pet14]. outflow [HMF^{+}19, Yan13]. outflow-explicit [HMF^{+}19]. outliers [WL11b]. output [LZT11, ZDL11]. outputs [Tol12]. over-penalized [ZCW15]. over-relaxed [Hua10b, Ver08, Ver12]. over-represented [fLcJ10]. overcoming [Cho17]. overdetermined [Gon13, hRWH18]. overland [WBN18]. overlap [DHQ11]. Overlapping [BL14, BM18c, Cal19, SRV10]. overrelaxation [DH11a, HWXC19]. overspecified [HP13]. overview [CJK18, RS13]. Owa [AS10a, AS11c]. oxygen [IS12]. ozone [CHT11].

P [GRS12, Pan17, WY15]. p-FEMs [WY15]. P2P [LLX^{+}10]. PA [FSM19]. PA-12 [FSM19]. package [JPP12, MV10, WBA^{+}18, XYXZ16]. packed [KPS10a, KPS10b]. packing [CKMR11, LTX^{+}13]. packs [LRCG16]. Padé [NGG12, RY11, KKT13, KPS10a, KPS10b]. Pages [Ano18-62, Ano18-65, Ano18-71, Ano18-72, Ano18-64, Ano18-73, Ano18-55, Ano18-54, Ano18-67, Ano18-69, Ano18-70, Ano18-53, Ano18-61, Ano18-57, Ano18-66, Ano18-68, Ano18-60, Ano18-63, Ano18-58, Ano18-59, Ano18-74, Ano18-56, Ano19-52, Ano19-54, Ano19-64, Ano19-61, Ano19-62, Ano19-63, Ano19-65, Ano19-50, Ano19-59, Ano19-58, Ano19-49, Ano19-51, Ano19-55, Ano19-57, Ano19-60, Ano19-56, Ano19-53, LK15]. Painlevé [HLLM19]. paintings [PAE^{+}12]. pair [CM10b, GJX18, GC19a, LM19a, LY11d, MR17, NS11, ZCT18]. pairing [FSH10]. pairs [MCF18, RKSA18, San12, Tsi11]. Pan [McN12]. panels [RJGS^{+}19]. pantograph [LW11b, XZ10, Yüz12a]. paper [Abd18b, JY11, Pan17]. parabolic [AMGC19, AAD17, AYH17, AG10b, AS10c, AY12, AA13, BS14a, BS16a, BCK11, Bra16, CL17a, CLS19, Cha18, CW14, CY19, Chu18, CELY18, CCP19, CG14, CJ15, Dai14, DN18b, DG10b, Din13, DSZ18, DZ21, DL18, EE18, FID14, FJP18, GHMN16, GM19, HGSL18, HGW18, HZP18, HLW19, JLF17, Kar18, KB15, KKL^{+}13, KK14c, LMR19, LSW10, Li19a, LZZ18, LLY13, LW18c, MT19a, MS15, MT18, MT19b, NLA19, NSYY13, Oan13, Saj12, SHM13, SNDK18, SBM10, SLM16, SLW18, Tan17, TG11, Tha19b, Tia17, TNV19, VB10b, WQRZ14, WLZ18b, XWY11, XFY18, XDH16, XZ19, Yan15, ZS13, ZY19, ZZ10b]. parabolic-hyperbolic [EE18]. parabolized [Bis14, DNZ^{+}13]. paraboloid [JCF19]. paradigm [KKD13]. paraelectric [MM18b]. Parallel [BH14, CP16b, Fia15, KBCS16, KCC^{+}13, KSS13, LZ19a, LO16, MUB^{+}16, YSX^{+}19, AML^{+}14, ADD^{+}15, BKR^{+}19, Cos18, DdSF13, DM12a, DCG^{+}12, DZO^{+}19, DZ17, HD16, Kan15, KRCJ11, LCC12, MGN^{+}16, MNJ^{+}13, PW10, RWZ13, SPH10, SKCL19, TH19, VMC^{+}14, WFL11a, WBN18, WKP^{+}14, WY19b, YXYH10, ZS16]. Parallel-in-time [YSX^{+}19, BKR^{+}19]. paralleled [San12]. parallelism [GL16]. Parallelization [Hof18, JZE^{+}18]. parallelotopes [CR13]. Parameter [ADGG13, FJC16, FWFL11, MCL^{+}13, MN10a, TDXQ18, ARK13, CLT^{+}13, CD12, CHL18, CEJV16, DGTC13, DO11, FJB19, FIM18, GYH11, GH15, GN19, HYCP11, HO19, Her14, JCWZ16, KJK18, KP19c, KK14c, KGM11, LK14, Li12b, LL13, Lin10b, LB11, LHL14a, Liu16d, Luh12, MSQ^{+}11, MA17, MP11c, NB11, OSA13, SLYY13, SYW11, TWLYÖ10, TL12, VAS^{+}18, WYD10, YLDL11, ZDL11, Zha19c, dPLM18]. parameter-free [VAS^{+}18]. Parameter-uniform [MN10a, HO19, KK14c]. Parameterized [HWXC18b, ZZ15a, ZYW15, DWZ16, LM17, LM19b, LJ10a, XY17, ZM16a, ZSZ17]. parameterizing [Boy16]. Parameters [CJ18b, AZB13, AsNAd10, AEG11, AAP12, BV11, BZT16, CCJP11, CZMZ11, DL11, EUTS18, FSRB15, KBDC12, LLZ11, Luk11, PA15, RSS10, Sha14, Ste16, TMLF19, WL11b, WLHZ14, WH11b, WZY13, YS19b, ZL14b, ZJZ^{+}11]. Parametric [WLDL11a, WLDL11b, AKV11, CLF10, Fio14, HWH^{+}15, ID10, JW05, Pop13, Pop14, WY18a, yZjH12]. parametrization [MS17]. parametrized [DDMQ19, IQR16, NMR15, PGQ16]. paranormed [DÇ12]. parasitical [ZL10a]. paraxial [SS14b]. Pareto [SLMZ12]. parity [YZMA18]. ParMooN [WBA^{+}18]. Part [Ikh11, Sea15, ZZC13a, ZZC13b, HHNLGC18, LT15b, CFS17, DHMU16, FRAK15, JRA^{+}18, RAW^{+}16, Tao18]. Partial [DW18b, GGAV18, GGAVGG19, GGAVRC^{+}19, RES10, AKT12, AuIK17, AMA14, AM13b, ABL10, AGU14, Arq18, AKS11, AuIA17, BKE18, BH10, BP11b, BHM12, BE11, BCF^{+}14, CM18c, CJ18a, CY19, CL15, CS14a, ĆCM10, DCL17, fDxZ12, DHY19, Els10, FID14, FN14, FNW18, GR13a, GTL16, GR13b, GMZ15, GGO16, HG18a, HY16, HLvS18, HP10, IBSS11, JNBK13, JW18b, JS11, KÖ10, Kat11, Kaw15, KK14c, KV18, Lee15, LC11a, Li11a, Li18d, LL16c, LLLC14, Liu16d, Ma18a, Ma18b, MI16, MM18d, MS15, MT18, NLA19, Ouy11, PC17, PW10, RM17, RS15, SCSF19, SAIZ15, SS14a, SGZW18, TN11, TWLYÖ10, VA11, Ye16, YQ18, ZY17c, Zha17b, ZZL18a]. Partially [Sha10, AM13b, BSK11, BKR^{+}19, DDD10, HA18a, LZ12c, NS11, RK10, SS11c, XWL18, YZ15]. Particle [ATO19, JKK11, MZB10, NNL13, WRW^{+}19, AJS14, BCCZ18, BBO10, CV14, COR18, DA18c, DH10c, FM11, FLWJ11, GSY10, HDHL11, JNJ^{+}11, KPP13, LCW12, Laz10, LTX^{+}13, PTH^{+}16, RMB^{+}14, RCG15, SGK18, SPP18, TTC14, WLA18, WvDRG19, XLZW11, ZYT^{+}16]. particle-in-cell [BCCZ18]. particles [ASFM15, BHK16, BKMT16, CDM10, Eba14, GLW13, KVR11, NNL13, NWZ11, SK14a, THC^{+}18, WAG^{+}14, ZLHF19]. Particular [LCLL18, BPF13, DCL17, DLC19, NKM15, YY15, YKC11, YCLY15, Yao16, ZTC14]. Particulate [ZYT^{+}16]. partition [BASW18, CDP16, CHT11, GGLP15, GCE18, HR14, KdLK19, MBT^{+}13, SL16b]. partitioned [KRD16, MUB^{+}16, PBK19, SZ17]. partitioning [Koj10]. partitions [EMRS12, LNW19]. parts [FSM19]. party [HPC12, KWPK13]. party-play [KWPK13]. Pass [SZ12b, WHC12]. passenger [ZNWG11]. passification [LZ12c]. passing [CHY12, LFC16]. passive [SRDD17]. passive-active [SRDD17]. Passivity [LZ12c]. past [Lee11c, MS13, Sah11, UMLF13, VPR11]. past-sequence-dependent [Lee11c]. patch [JK18, KVJB15, KV17a, KV17b, YLZ17]. patches [Dej11]. path [BSN13, Gao11, HR15, KYAA10, LJJ11, Lin10a, TMMASG10, VMO10, WZ17b]. path-following [HR15]. pathogen [ZWZ16]. paths [CSS10, DA12, Lin11]. pathway [XBHN16]. patients [ZJZ^{+}11]. Pattern [KS10b, SFM15, YXWL14, ZS15, AAR11, CS11a, FL13b, IW18, MGTH16, SD11b, TJQS13, VB10a, ZY15b, Zho13]. patterns [BGL^{+}15, DA16, Hu19, IS14, LZ11c, LD19, Liu16b, TG14, Zha18b]. Pawlak [AMD10]. payment [MM11]. payments [DRK10]. payoffs [LH12a]. PCM [JFS14, JFS20]. PCO’2010 [VBW10a]. PD [HK10]. PDE [AM17, BLS17, BH14, CGHW14, DGLS19, FSZS18, FF15, GMS18, HG18b, HWW13, HM17d, KM18, KFTT13, LG17, LW18c, LSV18, LR13, PMM17, PSS18, TMZ^{+}15, Tre18, YLC18]. PDE-based [BLS17, HWW13]. PDE-constrained [KM18, HM17d, KFTT13, LW18c, PSS18]. PDES [BNTT14, ATUC15, AHJM18, BQ17, BMP15, CCNT16, DWZ16, Gha17, Gha18, HJ13, HA18b, KMS15, LLH14, Lep11b, MMH11, MBJ16, PLT^{+}19, uIAA15, SK19, Yao16, YSB15, ZM16b, ZFLM18, ZFLM19, ZZL15]. pdf [Özd18]. Pearson [LMS13]. Peclet [DZW^{+}15]. pectoral [LTL^{+}12]. pedestrian [KYY12]. peer [SBA10]. peer-to-peer [SBA10]. PEM [SCA14]. PEMFCs [DD19a]. Pempinelli [LM18a, Liu18a, LZM18, MR17, PTZ19]. Pempinlli [ES17]. penalized [Ibr16, ZCW15]. penalty [CW17, CQ13, CW14, GP19, GNP14, LP12, LW17, SMH18, SCGW18, TNT12, WSW11, ZCW15, ZT15, ZD15]. penalty-free [CQ13]. pendulum [BAO^{+}12]. penetrable [ZSZ18]. penetration [LHM11, XHH^{+}19]. Peninsula [ZGD13]. Penrose [BT10, LZ19b, SSP13]. penta [JKS12]. penta-diagonal [JKS12]. pentadiagonal [EMR10, JJ13, JJ15, JYL16, JL17b, JL17a, dS16]. pepper [DCKY15]. perceptrons [KY11]. perceptual [WHC12]. percolation [CLB14, JL18]. perfect [PB11]. Perfectly [MLG17]. perforated [DC15]. perforation [LHZ^{+}11]. Performance [DdSF13, KHIB12, PLT^{+}19, YW14, AD19b, AGK15, BF16, CIN^{+}18, DA12, Ebr11, GGO16, KSPP11, KL12b, MSZG17, ODAZ15, PÁAP^{+}15, PPC15, PSD^{+}13, RF12, SJS^{+}10, WBN18, WTC^{+}12, YWW^{+}12]. performances [AGH^{+}15]. Peridynamic [MSZG17, LCW17, SL16a]. perimeter [Bri10]. Period [And12, Aki17, JL12, LJYS18, NNR14]. Periodic [DHGF17, ES17, FRSW11, Haj18b, LSW10, LW11b, PRR18, TZWM10, WYN12, Wri13, Zen11, ZKW15, AGT19, AA11, AN11a, AK12, AGDP19, ÁBÁPM11, AD12, ASY^{+}11, AKMUH17, BL18, BDGS13, DMP18, Dan12, DS18a, DSR10, DCN^{+}18, FJ19, Gao17, GÇK10, GL17b, GH13b, HN18, HMF16, HB12b, HXL11, HGW11, JJ13, JJ15, Kia18, KHWK10, KK19b, LZ18a, LS19a, Li10e, LD11b, LF11c, LZZ18, LT15b, LY11c, LLT16a, LLT16b, Liu18a, LDL10, MM10a, PLT17, Sah17, She16, SC16, SCBCB^{+}13, SCBCB^{+}17, SM14, TJ10, TZ15, TUT11, TNF11, Tim14, TTX^{+}16, VZM14, WF18, XY10, XLT17, YBC11, YZY10, YLG17, ZD11, ZZ11a, ZL11, ZTZ16a, ZZM17, ZSZ18, ZBFC19, ZL10a, ZDM11]. periodic-type [ZDM11]. periodically [WHLC11]. Periodicity [YY10a, AR09, AR10b, BBR10b, GH12b, QYL10, SZ12a, Sed13, TUT11]. Peristaltic [Tri11, Eba14]. Permanence [TZG10, ZHW14, YZAX10]. permanent [HLCY12]. permeability [XY16]. permeable [Pal13, REHA11, RRP16, SPH10]. permissible [DRK10, MM11]. permitted [RZL11]. permutation [Als10]. persistent [CEF^{+}13, DW18a, MW10]. personal [RPTD10]. personalized [HCHH12]. perspective [FCZ12, KJA10, SYZ19, VMAVGCM19]. persymmetric [dS16]. Perturbation [LYS12a, LYS12b, RSDR11, AÖ10b, AP10, BGPP11, BE11, BG11, CH11a, CTG17, CG15, DS18a, Dem10a, DM12b, Eba14, FH11, Gup11, GS11c, HN18, JNJ^{+}11, KW11, LX10a, LY10a, MH11, MGB^{+}11, Naw11, PP10, RMA10, RY10, Sah17, XXG10, YL16, Yus09, ZYSY17]. perturbation-iteration [AP10]. Perturbations [GH13a, Bae10, BR13b, CHM18, CRRS11, HDHW11, LW12a, LW12b, LCC13, ÖZ11]. Perturbed [DIJ12, HCF16, ABCR10, AHO16, ALI11b, BR12d, BZT16, CG14, CJ15, CX18, Dar11, DN18b, FM18, FBTS19, FR16, FH17, HO19, KAG11, KSMN11, KK14c, LN98, Li10c, LLC13, Lü14, LGVS19, MN10a, RS12b, Tur10, WSCL11, WZY13, dS16]. pervasive [YWK^{+}10]. Peter [BCHS18, Col18, Per18]. Petrov [BS14b, CCN14, CEQ14, EDC14, NCC13, RCH19, Rob14, Ros12, SXM11]. Petviashvili [SMYK19, Ade17, CTSX16, CZ15, Gep16, HTY^{+}19, IMS19, JPB11, KTK17, LZL19, LWL11, LMZK16, LXY19, MCF18, QTW^{+}18, SIL19, TTX^{+}16, WF17, WTYZ17, WF18, WW18a, WL16, WTLS18, YMHL18, YTL^{+}18]. Petviashvili-like [LMZK16]. Pexiderized [CBKR10]. PGSS [DYWL19]. Phan [ZH15b]. Phan-Thien [ZH15b]. Phase [KS12b, ATO19, AS11b, ABK^{+}13, BTEM19, BGM19a, BCFQ19, BSZ16, BKK12, BMSS18, BV17, CP15a, CPP10, CYP16, CFdM^{+}18, Col14, DTR19, DFW^{+}18, DJD18, FGB19, GOT19, GDM13, JFS14, JFS20, KÖC^{+}18, KM13, LMPE18, LK11, LH19, LSC17, Lin14, LZCL18, LCWZ18, MGN^{+}16, MK18, MM19, PFDG17, PGF18, RKW12, SSS16, SSHH^{+}18, SPT17, SSIP19, SS16a, Ser19, SLL17, STDLM19, Sul16, TR14, TY13, TDM13, Wan14, WCW13, WS16, ZR16, ZLC^{+}11a, ZQ14b]. phase-field [CYP16, FGB19, LK11, LH19, LCWZ18, PFDG17, PGF18, SLL17, Sul16]. phase-field-based [LSC17]. phase-lag [AS11b, KM13]. phase-type [RKW12]. phases [CTD10, HMSC10, Kup11]. phenomena [CHH14, EAMA19, GT15, GT16, KB15, Liu13, LLY13, LWW19b, LMPG13, MF18a, Sam19, SD19, TTG16, Tan17, TF17, Ye19]. Phenomenon [Boy10, IMS19, MM11, ZBFC19]. phenylenes [DYX11]. phi [CDM12, CCHG17]. phi-Laplacian [CDM12]. phone [KRM^{+}10]. phonons [Hei10b]. photobleaching [SL18b]. photonic [EGG16, KSF14]. photovoltaic [LCW12, YLH12]. PHT [VAK^{+}19]. PHT-splines [VAK^{+}19]. Physical [ZSY19, LDW11, MC10a]. Physical-bound-preserving [ZSY19]. physics [FMGR19, Grm13, SR17a, Xu11a]. phytoplankton [YLZ17]. pi [Mor11b]. PIC [TBP19]. Picard [Zha19b]. PID [dSCM12, DZS10]. PIDE [YKA18]. piece [Dan12]. piece-wise [Dan12]. piecewise [Gen11b, JCF19, LZH16, WW10b, ZL11, ZHL12]. piecewise-linear [JCF19]. PIES [Bol16, ZS18]. piezo [CL12b, LL12c]. piezo-actuated [LL12c]. piezo-actuating [CL12b]. piezoelectric [AB18, Kia16, KYA15, RYK13, SSH15]. piezoelectricity [FR15]. pile [LGH^{+}11]. Pilot [WZWS11]. PIM [YLLN16]. pinning [ZFC11]. pioneer [Sen12]. pipe [AB10b, BPR18, DLWW12, LWL14, MC10b, MEAMHHV18, YZ10a]. pipe-flow [LWL14]. pipe-flow/Darcy [LWL14]. pipeline [KSO16, WWZ12]. pipelines [CRXL15]. pipes [BTEM19, XXG10]. Pitaevskii [SSC19]. pitching [GMS15]. pivoting [BDM19a, Raf12, RTB14]. planar [DM12b, GKLR11, Hal13, KV17a, KLTS11, Luo18, QCLC17, YMM12]. Planck [AAH^{+}18, AHOP18, Loh16, SK14b, YQWZ19, ZL12a]. plane [ATH18, ĆCM10, Her14, JNJ^{+}11, Kar17, KS12b, LCQL17, Liu16a, MA10b, OSA13, PJ17, Pov12a, PK19, RHD18, Sha14, XCZQ16, ZM14]. planktonic [ZLMZ18]. planning [HHS^{+}10, MAPS10, VB10a]. plant [Li11b, ZL14a, ZZF18]. plant-herbivore [Li11b, ZL14a]. plants [AZB13, GIMZ14]. plasma [BCCZ18, MR19, Sea14, Sea16, Xu11a, YTL^{+}18, ZTSC16]. plasmas [ES18]. plastic [CTS19, HAESLB14, LMLB19]. plasticity [CTS19, KMS19, RSH18, TPHD18]. plate [AGH^{+}15, BHZJ19, BM13a, CCN14, D’A18a, FZBF10, KTA12, KYA15, LZZ11a, Par18, PvdM13, RBB12, Ran15, SSH15, SM17, SSL14, TM19, UKA15, YLY12, ZWL11, ZWG11]. plates [ATH18, ATH19, BKE18, BL14, GK16, HSK11, HD14a, JW15, KRCJ11, Kia16, MM13, MG15, SPH10, ZW11a]. platform [HMP^{+}15, HWH^{+}15, LZL^{+}18, PC12, WD13]. platforms [BF16]. play [KWPK13]. Playing [Bur13]. PLC [ZMG10]. PLEIADES [HMP^{+}15]. Plücker [BE18]. plume [JMDA13]. plus [GK11b]. ply [ABR10, ZTW^{+}19]. PML [MSFS18, DL19, HMZ18, MTV13]. POD [DA18b, DFP^{+}13, DNZ^{+}13, FMP19]. POD-assisted [FMP19]. Pohozaev [GTC18, QCT17, Luo19]. Poincaré [BX10, LG12, XD10]. Poincaré-type [BX10]. point [AE12a, AW11, AN12, AIB10, AJ12, AR10d, AER12, AHF10, AH11b, AKS11, APS12, Bac14c, BQ15, BQ17, BJQS18, BJ19, BZ10, Bai11b, BZZ^{+}10, Ber12, BS12b, Bic11, BGP13, BLyS18, CB11a, CB11b, CB11c, CM16a, CMR17, Cao19, CAY12, Cha11b, CCG18, CBS18, Che15a, CTZ17, CAP10, CSW11b, CM10b, ÇA14, ĆSCD11, Das12, DGK10, DYWL19, Dra11a, EM19, ESN10, EE18, FZ14, FWZ16, GRS12, HHY13, HLS11, HH11, HL11b, Hua10b, HWXC16, HWXC18b, JZ11, KAG11, KPR10, Kar10a, KT11a, KL12a, Kim11, KK10b, LCW10, LW11a, LM17, LM18b, LL19b, LZ11e, LZ11d, LZ15a, LZ16b, LZ18b, LC11b, LLL12, Liu12, LT11, MTM11, Mia18, NS11, PPC13, Pop11, RH15, RA11a, RI12, SD15a, SK11a, SMK18, SR17b, SZ12a, SWS19, Sha10, She12, SBJ15, SS16b, SR10a, SCK11, SZ14, SKH12]. point [THC^{+}18, Tha19a, TB10, Tur10, Ver08, Ver12, WZH10, WWA11, qXjH11, XY14, YBC11, Yan18c, YLK10, Yas12, YLH12, Yun13, ZC10, ZJ10b, ZBF11, Zha11d, ZZ18a, Zha14, ZL16, ZL10b, ZZ14, ZYW17, sHC11]. point-collocation [EE18]. point-weighting [CTZ17]. points [Ahm10, ABR10, Ant10, CS11d, Eba11, HR14, HH10a, IN10, KKS10, Mar11a, MAK12, Pir11, Qiu12, SLK12, XLY10, XW18, YK17]. Pointset [DK18]. Poiseuille [LYN11, MA10b, RSV11, SPP18, WCQ^{+}19]. Poisson [LY17, VMC^{+}14, AZ17, AM10c, Bao16, CTG17, CT17a, DLZ17, DLZ19, DO11, DLC19, Fio14, GTC18, Gha17, KKBR19, KL16b, LS17a, LC16, MLL16, MT19b, RCM11, hRWH18, She18b, SWL15, WZM^{+}16, Ye17a]. Poissonian [Liu16e]. poker [Coo10]. polar [BM13a, BJPT16, LY10a, TLR17, WV16]. polarity [LHH10]. polarized [AHP^{+}14]. pole [MCP13]. policies [BO10]. policy [BKP11, CL12a, CWDL17, DRK10, HKHK13, RPTD10, VCM11, ZZ17]. pollutant [MC10b, RAZ19]. pollution [DGOZ13, ODAZ15, RZL11, ZGD13]. Pólya [NP12, WLS10]. Pólya-type [NP12]. Polychotomous [LLSW10]. polygon [CCK18]. polygonal [AOW18, ABSV18, JLL18, LL12b, SWL19, Vac18, VHPVNXW18, WW14a, Wei14, Wei17]. polygroup [YZGW10]. polygroups [WYG11]. Polyharmonic [Seg19]. polyhedra [CDN14, Wac10]. polyhedral [BN16, CBBE16, WW14a, dVDR18]. polyhedron [HMF^{+}19]. polymer [EF14, FG18, GGH18, LDS10, Lin14, STDLM19, ZH11]. polymer-solvent [STDLM19]. polymeric [ZR16]. polymers [CD12]. Polynomial [DCL17, KP10a, PKTH13, BKY10, Boy16, BS17, DE10, EE10, Gem16, GK11b, HA18b, Ikh11, JLCS10, LMW10, LY11a, Liu16a, Lu12, LvdVX18, McN12, MP12, MW14, PZ11a, PW18, RS14a, yS10, Was13, WO18, WFL11a, WY15, WZ11b, YSS11a, YSS11b, ZM14, ZHL12, ZMFL18]. polynomial-and-log [Boy16]. Polynomial-based [PKTH13]. Polynomials [CM11a, YMDZ10, APRM11, Bay19, Boy16, Büy10, CMGR11, CDD12, Dua11, ESL11, Gal11a, GTL16, KK14a, KW11, KAS11a, KYR15, Kim10, LY11a, LS11d, Mah10b, MBH11, MS11a, MRR11, MB10a, Oru17, Öza11, OSS10, Pap15, RT10, SKPW14, TÇA11, VT11, VSI12, WJ11]. polyps [ZJZ^{+}11]. polytopes [Gue13, IBB10, Wac11]. polytree [VHPVNXW18]. polytree-based [VHPVNXW18]. polytropic [Zho16]. Poncelet [CGM10]. pools [GIMZ14]. POOT [HCL12b]. Popoviciu [Yan10b]. Population [VV14, ALMLM14, AB16, Gal12, GS15a, Hu19, LZ11b, MCL15, MW13, MW16, OAKR16, PP10, TNF11, Wri13, YKKS10]. populations [AF13, CN16b, PLKCC13, VDV13]. Pore [SCA14, YLL^{+}14]. Pore-scale [SCA14]. poro [RDE^{+}17]. poro-elasticity [RDE^{+}17]. poroelastic [AKSW19]. poroelasticity [CR18a, KVV14, SR17c, SS13]. Porosity [PHM^{+}19, XY16]. porosity-permeability [XY16]. Porous [Li18e, AM18, AGPR19, AP19b, ASMM11, BASW18, BSK11, BV10, BKNR19, CIN^{+}18, CP15a, CJ12, CLB14, DVY14, EAAS18, EAMA19, EEBM10, FLZ14b, Far11, GQF^{+}10, GZR^{+}13, GGGR17, GDM13, GF19, HWyL11, IK16, KMT19, KNT12, KRP12, KBCS16, KRBS18, LR15, Liu13, LZP^{+}19, MCB10, MCR11, MM19, MMA12, MS10c, NPR10, OHMAK18, PZL^{+}18, PCO16, Pan17, RY10, RHMA18, RBB12, Ran15, REHA11, RTL19, RJGS^{+}19, SSS16, SSHH^{+}18, SD19, SRRP18, TR14, Tim14, TRL19, UABK16, XY16, XZL^{+}11, YYK16, YB13, YKKS10, YLL^{+}14, ZSLZ19, ZD18]. Portfolio [WZG19, BKM11, BKK11, cFpClC13, Jum10, LS11c]. portfolios [Che11b]. posed [CH11a, CDY11, HO19, LP10, Liu14, MZES12, ZT16a]. posedness [AKV11, AS10c, BTB18, CCR16, Dua18, FAHZ17, LL16c, PZA19, Pu12, QY17, QaY18, RA19, Tha19b, Via15, WD16, YZ19]. position [AY18]. positioning [LL12c]. Positive [BDS10, CHS11, Fer11, GCR^{+}18, HL11b, HGW11, JWX14, KLL19, Kon16, LS17a, Li10e, LS11a, LY11c, MM10a, PC11b, Su12, YY10b, Yan11c, Yan11d, Yan12b, YL10b, Zha10, ZZ11a, ZLG18, Amb19, BS12a, BC12, BB12, BKL^{+}19, BR13b, CGY10a, ÇT12, CC19, CM13b, CO19, CP16c, DWY15, DLS18a, DSL18, FIS18b, GJ19, Goo11b, HY10, HLWX11, HYCP11, HM15, HH10b, JK11c, KD10, KB15, Kim11, KM11, LW19a, LL19a, LLZ10, LCW10, LYZ11, LF11c, LZ12a, LY14, Li18b, LZ11e, LZ11d, LWKK10, LZKU11, LLT16a, LLT16b, LS10d, MX10, MW17, Mah11, MS11a, MPLR18, SYY13, Sta11, WSG10, WZH10, WWB13, WLM13, WS17, WS10b, Wu16, XLD11a, XLD11b, Yan10d, ZSH11, ZL10a, ZCH12, ZT16b, ZY15c]. positive-definite [BKL^{+}19, WLM13]. Positivity [CLM14, YDW15, CP15c, MS11b]. Positivity-preserving [CLM14]. possibilistic [LYLX11]. post [DZ17, YW11b, KB10b]. post-buckling [YW11b]. post-processing [DZ17]. postbuckling [KHF^{+}19]. posterior [FdOdSS17]. posteriori [AV14, Bac14a, BPC17, BO18a, BWL18, BV17, BS15c, CCKP15, CCZ18, CGO19, DVY14, FOS19, GRBT16, GGS16, GS18, GOS18, GSZ14, Hou15, Kim14, KT15, KKJ15, LCH19, MRR18, OY19, PS18, PL17, RDE^{+}17, RCRV14, RA11b, WZ18b, ZY11, GG18]. posteriori-based [DVY14]. Postprocessing [Cer18]. potato [RMS12]. potential [AHHM19, Bao16, BDGS13, CRG16, CTG17, CHZ19, CT17b, DL14, DLS18b, EPP18, FJ19, FDG^{+}17, GTC18, HLY17b, JPB11, KVW18, LL19a, LD11b, Li18b, LC16, PC11b, PLT17, SPL19, SGK18, SBM13, Was13, YL16, YW19, YM13, ZKWW17]. potentials [Ali15, SLKK19]. Poussin [Kig10]. powder [RMK19]. Power [CSCM13, HMM12, VBW10a, VBW10b, Ala10, AM14a, BZZ^{+}10, CCJ10, CCJV11, mCfX10, Ciz12, D’A18a, DGA18, DS18a, GIMZ14, HM10, KSMT11, LLH10, LL12d, PTL13, QGGL13, SMBY10, TS11b, VPR11, YLH12, ZY10a, Zha19c, ZT18a]. power-law [AM14a, SMBY10, TS11b, VPR11, Zha19c]. powers [Mah11, dS15]. PPHSS [HWXC16]. PPS [LM18b]. Practical [RWZ13, YÇG12, HKS14, LZL^{+}18, LA11, fNS11, PMM17, WW11a, ZPS^{+}12]. Practice [KKK12]. practices [SXB^{+}12]. Pre [ZD15, XS10]. Pre-asymptotic [ZD15]. pre-matroids [XS10]. precipitation [BP11a]. Precise [FLDZ12, MHH11, YLS12, ST12]. Preconditioned [BJRF19, HS18, HWXC19, LHY18, YLC16, CMR17, CWDL17, CL17b, FZ14, FWZ16, KLK15, KM14, Mia18, WZ17b, XWY17, XWY18, ZYW15]. preconditioner [CWWY15, DGR18, Du12, GSR14, LZ16b, LZ18b, LJJ17, LHL^{+}14b, MY13, Raf12, RTB14, SWS19, SS18a, WWB13, Yan18c, ZZ18a]. preconditioners [BD16, BF16, BK15, Cao19, CHK^{+}18, CBS18, DdSF13, DDMQ19, DHMU16, HH11, HM17d, HMZ18, HWXC18b, HK17, KM18, KCL14b, LZ11c, LXZ18, LZ15a, LZZ18, SBJ15, SS16b, WN18, ZLS13, ZYW17]. Preconditioning [LZ16c, PSS18, AEF15, BJ19, BDHR18, CNP14, DM19a, GGR19, HP19a, KK13a, Pul16, RC17a, Tan18, Tom13, ZT13, ZZ15a]. predation [KALAS11]. Predator [ZY13, ADS14, BP13, CSW11a, Che16, DGT18, GOGYL^{+}11, JLWX18, KL19, LG10, Li13, Li14, LW15, LWN15, LWD15, Li17a, LX17, LLY18a, Li18a, LD19, LY11c, Liu16b, LDL10, LPY16, MLY18, MW16, PQB^{+}16, SZ14, SFM15, WL15, WW19a, WX18b, XZZ16, Xu14, YY10a, YL13, Yan18a, YXWL14, YDW15, Zen11, ZZ11a, ZHW14, ZZ16a, ZSW19, ZG18b]. Predator-induced [ZY13]. predator-prey [DGT18, KL19, LX17, LLY18a, Li18a, MLY18, WW19a, Yan18a, ZSW19, ZG18b]. predator-prey-dependent [JLWX18]. predators [CP16a, MY16]. Predicting [CHT11, MSH10, CGJ^{+}14, KSM12, SH18, XWL18]. Prediction [HBS^{+}10, SKCL19, yS10, BEAA11, Che11e, MLSLM15, VJM15, XJLX10, XCM12, YMSL11, ZNWG11, ZLY12]. predictions [HCL12b]. Predictive [RMK19, BKDM13, BL12, CN13, HMM12, MBT^{+}13]. predictor [AGPCC10, KBAF18, WFL11b, ZC11c]. predictor-corrector [KBAF18]. preemptive [LA11]. Preface [Ano14y, Boc19, CK13, CBM10, DGL18, GR12, HPV^{+}18, LNP^{+}13, Zel13, Zho10]. preinvex [LN10]. Preliminary [CZ11b]. premixed [AM13b]. premixing [AM13b]. preordered [JMLF11]. Preparing [ODAZ15]. preprocessing [ZWJ^{+}11]. preproximity [ÇA10a]. prescribed [HA18a]. presence [BTEM19, COR18, EA10, Gol13, Kia18, MCB10, MM18a, MVB^{+}12, PLKCC12, SSK13, VV14, WLT13a, YLS12]. presentation [OSS10]. presented [FES17]. Preservation [AM12c]. preserve [CMGR11, Mah11]. preserving [AH10a, AX11, AS11c, BBC^{+}11, BC11, CLM14, GW15a, HM15, JJ13, KKVS19, MJB18, MDRRV11, MDVM17, TMZ^{+}15, WK13, Wu11b, YXP^{+}13, YDW15, ZSY19]. press [HT12b]. Pressure [Yan10c, BMSS18, DLF^{+}11, GGR15, HMF16, LB18, MMT18, QWGJ15, VMP15, XZL^{+}11, ZZLB18, ZLZG11]. prevention [Cha13, DA18b, LHL15, MKA^{+}10, TYY^{+}12]. preventive [LJYS18]. prey [ADS14, CP16a, CSW11a, Che16, DGT18, GOGYL^{+}11, GM14c, JLWX18, KL19, LG10, Li13, Li14, LW15, LWN15, LWD15, Li16a, Li17a, LX17, LLY18a, Li18a, LD19, LY11c, Liu16b, LDL10, LPY16, MY16, MLY18, MW16, SZ14, SFM15, WL15, WW19a, WX18b, XZZ16, Xu14, YY10a, YL13, Yan18a, YXWL14, YDW15, Zen11, ZZ11a, ZHW14, ZZ16a, ZSW19, ZG18b]. prey-predator [LD19]. prey-stage [Xu14]. prey-taxis [KL19, Li16a, Li18a, WW19a]. price [BM12a, BC16, CGK14, CEJV16, FES17, GSS11, Sca11, Sin16]. prices [CKM12, ZH18b]. Pricing [LW17, Li18d, LPP15, SS19, ZLL18, ACAS11, CXZ15b, CM18c, CJPR10, CJPB10, DNS16, DH17, DH10c, FSRB15, FSB17, GW15b, GLW18, HMV18, HZ16, HC18, HZ18, HLvS18, KKT13, Lee14, LvS15, MZC17, MvS18, Moh15, SL16b, Tha19a, YKA18, ZM16b, ZC11c, ZBL12, ZH18b]. Primal [FLdS14, BQ17, DG13b, Hof18, KLP10, RTT17]. primary [ABK10]. Prime [ZHB11]. primitive [BG15, LH10b, ZZG19]. principal [LZH16, LR17a, ZM13]. principle [AKT12, BBR10a, EK16, Gal10a, Kim17, Liu17, MF11, Mah14, ÜM16, XFY18, XH11b, ZM17a, ZZG19, ZW11d, ZM17b]. principles [Auc18, KK15, ZLWL11]. printed [WA19]. printers [SLCC12]. priori [BP19, BN16, DLS18a, GGS16, GS18, GOS18, Hou15, NB17]. Priority [CL12c, CZ10, LA11]. Priority-based [CL12c]. prismatic [BMAR18, CWHW17]. Privacy [WK13, YXP^{+}13, GVJ13]. privacy-friendly [GVJ13]. Privacy-preserving [WK13, YXP^{+}13]. private [SBA10]. probabilistic [BA16, HpD11, Wan10a, WCZ13]. probabilities [FSM19]. Probability [Coo10, CKN11, HW11, LL13, MD10, PS12b, WW11b, WL12a, WLT13a, ZC11a]. probe [WY11a]. problem [Aba10b, ABB17, AW11, AK12, AA10b, AHO16, AAD17, AT18b, AT19, AS15a, ÁBÁPM11, ADS14, AV14, AS19, ACTB19, AKL18b, AA10c, BZ10, Bai11b, Bai12, BNR10, BLS18, BPS19, BDS17, BCK11, BKMT14, BKMT17, BDF16, BWL18, BMS13, Bog10, Bog11, BD19, BM18d, BRR16, BV17, BZT16, BPKM10, CW19, CL17a, CCRS17, CGHY11, CCY10, Cha18, CLTA11, CLH13, CLCF14, CM19a, CZF10, CyL11, CPT15, Cho17, CCK18, CS11b, CB19b, CM14, CG15, CN16a, CX18, CHS19, CGJ^{+}14, DL10, DGR18, DRK11, De 10, DCG^{+}12, DFGG13, DMPV10, DDM^{+}18, DH16, DZ17, DLQ16, DH10c, ESN10, EG10, EE18, EKZ17, EKE18, FST19, FES17, FLdS14, FPB17, FBL11, FRSC16, FH17, FOS19, FKeT12, Fur13, FIM18, GG18, Gao11, GHC15a, GS18, GOS18, GR13a, GRS18, GHMN16, GM19, GF19, HB19]. problem [HA18a, Haj18c, HY18, HXX19, HO19, Hes14, HSJ15, HSWZ11, HSZ15, HK15, HCZ16, HMZ18, Ibr16, JRZK11, JPS14, JWX^{+}13, JO19, KAG11, KLP17, Kar17, KP10a, KN11, KV10, KLL19, KTA12, KTDT17, KS12a, Kig10, KLP10, KCL14b, KK19a, KA13, KSS13, KFTT13, Koj10, KKL^{+}13, KM13, KKD13, LeT10, LPK15, LX16, LLZ10, LLW10, LYZ11, LL11, LX12a, LWBW13, LTX^{+}13, LLY18b, Li18b, LKCN19, LJ10a, LZ11d, LZ14, LZ19a, LLZ18, LBH^{+}12, Lin10b, LJJ17, LJLY19, LC10d, LCZ11, LLL12, LLY13, LDL^{+}15a, LDL^{+}15b, Liu16a, Liu16d, Liu17, LW18b, LZF18, LLLW18, LCHZ19, LHTL19, LGVS19, LDHH13, LLX11, MA17, Mar12, MM13, MMOJ14, MMT18, Mat19, Mia18, MGB^{+}11, Moh14, MMR10, MMR11, MLGY16, MRR18, Nab19, Nes10, Ngo18, NSYY13, OY19, ODR10, ORR16, Özu15, PCO16, PAE^{+}12, Pir11, Pit12, PA12, QZ16, RSM17]. problem [RES10, RS15, RDE^{+}17, RBTD14, RCRV14, RHD18, RFP11, hRWH18, RGVR17, Sal10, SYZ19, Sca11, SJL^{+}19, SH12a, She12, SKPW14, ySW10, SLW14, SWL15, SBM10, SMBY10, SS18b, Sle13, SS16c, SbX19, ST19, SWW11, SKM11, SZ11, SZ12b, SSL14, SCGW18, TLR17, eT10, TAS11, TB10, Tia17, Tod15, Tod18, THD19, Tro13, TW18, TKBMT17, TNV19, VDB13, WZW10, WZH10, Wan11, tWqLzGkP11, WG11, WL12b, WL15, WZ18b, WZG19, WZ18c, WY15, WH14, WFZ12, WZ11b, WZY13, WCW13, WW14b, Wu18a, XD17, XW19a, XL15, XZR16, Xu11a, YK18, YF10, Yan11d, Yan18a, YJ19, YGR11, Yil19, YYYH19, Yun13, Zak18a, ZSH11, Zha10, ZHZ14, ZHJ14, ZZ15b, ZT15, ZKWW17, ZHC17, ZJ17, Zha17a, ZL18, ZSL19, ZH19, ZL10b, ZT16b, ZG18a, ZMFL18, dBD17]. Problems [Agr10, BCHS18, AEG18, AA11, AN11a, ANP11, AA10a, Akm15, AH11a, AJRWS12, AAH^{+}18, AIIZ10, AEO15, AD16, AA18, AER12, AYH17, AKLS19, AK11, Ant14, AVZ15, jASzZ12, AG10b, AKA11, Auc18, ABK^{+}13, Bac14a, BJRF19, BS12a, BMJ19, BL10, Ban13, BS14a, BS15b, BS15a, BHM19, BDM^{+}19b, BJS15, BDGS13, BCJ19, Bic11, BR16, BGR14, Bol16, BSN13, BT14, BLyS18, BKP11, CHS11, CB11a, CB11b, CB11c, CKSL^{+}14, CGGM19, CM16a, CMR17, Cao19, CAY12, ÇT12, Cha11a, CHBTD14, CCG18, CBS18, CCD10, Che11d, CM11c, Che12b, CHLY15, CCZ18, CLL19, CO19, CKLL10, CM16c, CAP10, CNP14, Cia12, CJ15, CJPB10, CFLX18, CN11, CDN19, CLN^{+}19, DMP18, DGB10b, Dai14, DWY15, DN10, DN18b, Def10a, Def10b, DM12b, DS18b, DGK10, DBEE11, DB12, DCN^{+}18, DGLU18, DBH^{+}14, DYWL19, DLC19, DSL18, DHMU16]. problems [EFK15, EO15, EDC14, EGG16, ED12, EF14, FJB19, FMS19, FZ14, FWZ16, FZ19, FDG^{+}17, FM18, Fer11, FM12b, FR16, FGHZ17, FP18, GLLC19, GCE18, GP11, GGVRB19, qGpWhL11, GSZ14, GMI12, GR19b, Goo10, GKM11, GN19, GR13b, GHCZ18, GS15b, GS19, GNP14, GW12c, GLZ18, Hak14, HSD10, HHY13, HLWX11, HMY15, HYCP11, HKK^{+}16, HZL17, HCF16, HL11b, HRMS10, HT13, HZ11, HVO17, HT16a, HM18a, Hou15, HLY17a, HLSN15, HRHP17, HC14, HWXC16, HM17d, HZP18, HWXC18b, HD19, HK17, IQR16, Ima17, Jal14, Jan10, JM16, JZ11, JK12, KA10a, KK15, KLK15, Kaw15, KP19a, KM18, Kha10, KFYW11, KL12b, Kim11, KW14a, KW14b, KVV14, KRD16, KAS11b, Ku15, KS12b, LP10, LMR19, LLW11, LG17, LC11a, LCH19, LN98, Li10c, Li10b, LCW10, Li10a, LBZL11, LW11a, LW11b, LS11a, LY13, LCK13, LY15, LH16]. problems [LM17, LXZ18, LM18b, LL19b, LS19b, LZ11e, LZ15a, LZ16b, LZ18b, LZ19b, LYC12, LZ17, Lin10a, LBvB^{+}16, LCLL16, LMRS10, LLH11a, LHF11, LXP11, LLH11b, LS12a, LS12b, Liu12, LW18a, LW18c, LZ18c, LW19b, LRBA15, LDY11, LvdVX18, Luc10, LLW15, MM10a, MJB18, MD18, MT10, MM16, MT12, MDG19, MV11, Mil18, MP19b, MvS18, MP19c, MN10a, MR15, MS12b, NPD17, NMR15, NXHN14, NCC13, OP14, PGQ16, Pad18, PN10, PC11b, PH13, PKD19, Ped18, PW10, PH19, PBK19, PD17, Pir11, PT11, Pov12a, QCLC17, QWL19, RSWZ10, RFK16, RCH19, RWTW19, RSH18, RF12, RHC15, RR11, RS18b, RGdSRLAJ10, RA11b, SK11a, SSSB11, SRGL13, SR17b, SVY16, SWS19, SWL19, She11, She12, SBJ15, SS16b, She18a, Smo17, SW11, Sta11, SST12, Su12, SZW11, SS13, SI17, TPHD18, TM12, TCM18, TA11]. problems [Tha19a, TC10, TDN19, Tur10, Vac18, VAS^{+}18, VB10a, VAB12, WWW11a, WS11a, WWW11b, WWA11, WZWX11, WSCL11, WZF12a, WZF12b, WmN13, WC15, WHS17, WTM17, WCSW18, WZC^{+}19b, WCCS15, WLZ18b, WLGL10, WQNF12, XFL16, XW14, XZC12, XFY18, XC18, XLD11a, XLD11b, XC16, XHH^{+}19, YY14, Yan10d, YY10b, Yan11c, Yan15, YH15, YLK10, YZY10, YC10b, YNLK10, YCLY15, Yas12, YX11c, YDL11, YSB15, ZT16a, ZN18, ZB19, Zbo19, ZYWZ17, ZJ10b, ZBF11, ZLL11, ZLLF12, ZL14b, ZZWG16, ZYT^{+}16, ZZ18a, ZL19b, ZSS10, Zha14, Zha15b, ZcHS18, ZL16, yZjH12, ZY10b, ZGZ13, ZZ14, ZG16, ZZX16, ZYW17, lZxLhY12, ZS18, ZLGL11, dPLM18, sHC11, uRK11]. procedure [AD15, BKMT14, CD14, HYL10, JPCY13, Mai16, SSS11b, WW11c, YLDL11, Zha18d]. procedures [Tom11]. Proceedings [BCHS18]. Procesi [YT11]. Process [SLCC12, ATO19, ALI11b, CsH10, CPP10, CSSW12, CL12b, Chu11b, DK14, FXC18, GSI19, GCG12, GLW18, HHS^{+}10, KYR11a, KYR11b, KT11b, LL12a, Li18d, Mai10, MP11c, NJV13, RMK19, SSIP19, WvDRG19, WA19, XLK11, XLF12, XX19, YLL^{+}14, HLvS18]. processes [ASB12, AD12, AGD10, BGM19a, CIN^{+}18, CGS12, EF14, FLDZ12, GDM13, IW18, JSGP16, MDVM17, MMS^{+}18, PdlF10, SM19, Tim13, XWN11, YjH18, ZZHF12, ZWMD16]. processing [AA10b, AA10c, DSK^{+}14, DZ17, IC12, OKTR11, PMM17, TYY^{+}12, WSS10, WJWW12, WCZ13, WPH11]. processor [KORR10]. produced [RNB11, RB19]. producing [YLZ17]. Product [YZGW10, AAR11, CHM^{+}10, CQRW11, ET12, GTG11, GM18b, KD11, MV12, VC12, WW10c, XD17, YKRV11, ZHJD13, ZZL18a]. product-form [MV12]. product-sum [AAR11]. product-type [ZHJD13]. production [CCCW10, Chu11b, DRK10, DRK11, DK14, DKG14, MAPS10, QMW18, VB10a]. products [CL12a, Git14, LCLL18, San11, VCM11]. Prof. [DFG^{+}18]. professional [KSM12]. profile [MNJ^{+}13, Mom11, ZLG^{+}10]. profiles [DL18, TNT12, WYL19]. program [BCF^{+}14, WBA^{+}18, ZHQG12]. programmable [HHY^{+}11]. Programming [HH15, Ant10, CQ13, DBH^{+}14, GKS10, HDS11, HTGSH13, Koj10, LLW10, OSZP13, RFP11, RF12, SKK12, THZ^{+}11, WZW10, WLGL10, ZLL12]. progress [PZ11a]. projected [Mai16]. Projection [KEHB18, VMP15, Zha17a, BS16b, CGHY11, CD16, FSB17, HD14a, HZM11, JMHF13, KN12, LY15, Liu14, LL15, MKL11, MW14, PLKC16, QZ16, QHW11, SK11a, Von19, WY15, qXjH11, XY14]. Projection-based [KEHB18]. projection-proximal [qXjH11]. projections [ASA16]. projective [JW11, PC12, SH11]. projectors [AAB^{+}13, BT10, DNS15]. projects [BKK11]. prolate [LDHH13]. prolongation [Hu15]. promising [BGPP11]. proof [BCF^{+}14, DC10, ZY17b, ZWY19]. Proofs [YLLN16]. propagated [ABDKD12]. propagating [CWH13, KSG11, MTV13, PLMS14]. Propagation [LMP13, AGT19, Abd18b, AM13b, AM18, BP11a, CBB15, DW18b, EO15, GM18a, GGK18, GGL13, HB19, Ikh11, KB10b, LR17b, MNPD15, MR15, PD17, Ray17, SIL19, WZHW13, XCZQ16, ZZW15]. proper [DA18b, FSZS18]. propertie [Wan13a]. properties [AZ15, AS10a, ASMEE11a, AX11, ADGG13, BM11b, BL17, BCSCB^{+}15, BBD10, CNV14, Che11a, DK12, EAA10, GW15a, HLY12a, HA11, Ili12, JTC^{+}10a, Jia11, KHUO12, MP10a, Mah10a, Mah10b, Mah11, MC10a, NODA11, PKK12, Par17a, QZ11, SK10a, SL16c, SST19, SGQ12, TO11, VLJH18, WS10a, WC10a, WZXL11, WYG12, WSH12, XMWH10, ZJ12, ZGD14]. property [AAR11, AIB10, Che19a, GRS12, LW11c, Med12, MAK12, TKH10, THY^{+}10, XDL12, YLB16, Yil10]. property-based [LW11c]. property-oriented [Med12]. proportional [CW10b, LW17]. Proposing [KK12]. propositions [ABFGZ11]. prospect [DD13]. protection [JCZZ13, LD19, Yan18a]. protein [NJV13]. PROTEUS [FIVM17, FIMV18]. protocol [CHY12, CJP15, GLM^{+}11, HPC12, MKA^{+}10, NCL13, QCYL12, SPLHCB14, VBK13, CJP12]. protocols [KHIB12, LYM12, OO12]. proton [SCA14]. prototypes [NEB14]. provable [LWHY10]. Provably [FSH10, WMSH11, CPL11]. proximal [Hua10b, Ver08, Ver12, qXjH11, XY14]. proximal-point [XY14]. proximity [AH10c, BA11]. proxy [GH12a, LWHY10]. PRP [TNP17]. Pseudo [Abb11, Hes18, CsH10, CZN12, CT10a, DZ21, GDZ11, HGW18, HS13, HSJ15, JLF17, KS10a, Liu11b, MEAMHHV18, NT17, SBEB10, SH10, SLW18, Tha19b, ZL11, ZcHS18]. pseudo-contractions [CsH10, KS10a]. pseudo-contractive [GDZ11]. pseudo-homogeneous [MEAMHHV18]. pseudo-parabolic [DZ21, HGW18, JLF17, SLW18, Tha19b]. Pseudo-spectral [Hes18, HS13, HSJ15, SBEB10, ZcHS18]. pseudo-triangular [Liu11b]. pseudocontractions [CS11c]. pseudocontractive [ZRC11, sHC11]. pseudoconvex [XL11]. pseudodifferential [PT15]. pseudomonotone [AK11]. pseudopotential [WGY^{+}18]. pseudospectral [Ali15, HLX18, Jav11, SD12a]. pseudostress [GGS16]. pseudostress-based [GGS16]. pseudowheels [KW12]. psi [Che11a]. PSM [ZcHS18]. PSO [DWI^{+}12, JS12a, PSD^{+}13, ST12]. PSP [LT15a]. PSP-splines [LT15a]. PSPS [LT15a]. PSS [FWZ16, LZ15a, LZ18b, Yan18c]. PSS-based [LZ15a]. public [Lee11b, WH10, WMZW11]. public-key [Lee11b, WH10]. pull [NGG12, RMA10]. pull-in [NGG12, RMA10]. Pullback [PP14, YL14, ZZ16c]. pulsatile [MDL18]. pulsating [ASFM15]. Pulse [MBH11, KB10b, LS11b]. pulsed [WHLC11, YMDZ10]. pulses [BL12, Dan12]. pump [AdSSS19]. punctual [Yaz11, YAS^{+}11]. purchasing [DH10c]. pure [BM19, CKR10, ZT18a]. Purposeful [AAP12]. push [KSPP11]. put [CW14]. puts [ZC11c]. puttable [ZLL18]. pyramid [CW15b, TKH10].

QFD [WC11b]. QoS [DFS11, WL11a]. Quad [BP18, SCGW18]. Quadratic [SZW11, WH10, APRM11, BDK^{+}11, Cer18, CD10, CM19a, Cie11, Dar11, Dej11, DBH^{+}14, DP15, EGG16, HA18a, Haj18c, HTGSH13, ILP^{+}11, KC11, Kes10, Lee11b, MTM11, Sal11, TCM18, WTM17, You11, ZZ10a]. Quadrature [AuIK11, BC15, MT19a, Bis14, CR13, HL18b, JW15, LH16, LSZ11, MD15, PCS13, RSM17, SNH10, SD15a, Von19, WY18a, WV15, XG10, YW11b, ZMM18, ZW11a]. quadratures [SSSB11]. quadric [Dia17]. quadrilateral [ATH19, KHF^{+}19, WN18, ZL19b, ZMFL18]. quadrilaterals [Li15]. Quadruple [KL12a]. Qualitative [BBDS11, MW16, AÇT11, FdOP17, MP10a]. quality [DRK11, KYW^{+}18, LLX^{+}10]. quantification [BCPS15, BK18, WY15]. Quantitative [AM10c, KC19, BM10b, DB10, FdOP17, Wu10]. quantities [AK16, JQSS12]. quantity [Chu11b, KGJ11]. quantum [Baa13, COR18, FGL10, Iom18, JKS19, LJJ17, Lóp19, MT12, SPL19, TTC14, ZSAN18]. quarantine [SG11a]. Quartic [BM13b, DL11, Ahn10, GKLR11, GMI12, HKT11, SCV10, ST16]. Quasi [DGLS19, NCC13, SL16c, WMW13, YYYH19, Zha19a, Abb10, AM15, ABCR10, ADK10, BGIN13, BNTT14, Bor11, Çak11b, CS10c, CS11d, CT10a, FIW13, GWZ11, GW15a, HLCY12, KKS10, KA10b, KKLJ11, LT13, LX16, LHY11, MYZ12, Mai10, MS15, NHH13, NZ14, NJV13, PT11, QHW11, QL10, SID15, SSR11, ST16, She11, SH10, SW11, SW16b, SJHC14, WXF10, WH18, Wei17, WY18b, Zha11d, Zha18e, vdW14]. quasi- [QHW11]. Quasi-Chebyshev [WMW13]. quasi-compact [WH18]. quasi-contractive [Zha11d]. quasi-convex [ADK10]. quasi-distribution [SID15]. quasi-equilibrium [PT11, SW11]. quasi-geostrophic [FIW13, WY18b, Zha18e]. quasi-Hausdorff [SSR11]. quasi-interpolant [GW15a, ST16]. quasi-interpolants [BGIN13]. quasi-interpolation [GWZ11, WXF10, Wei17]. quasi-linear [KKLJ11, MS15, vdW14]. quasi-minimal [AM15, SJHC14]. Quasi-Monte [DGLS19]. quasi-Newton [LHY11]. quasi-Newtonian [LX16]. quasi-nonexpansive [CS10c, CS11d, KKS10, KA10b, Mai10, She11]. Quasi-optimal [NCC13, BNTT14, SW16b]. quasi-pseudo-contractive [SH10]. quasi-pseudo-monotone [CT10a]. quasi-sliding [HLCY12]. quasi-static [LT13, NHH13]. quasi-stationary [NJV13]. quasi-uniform [Abb10]. Quasi-Variational [YYYH19, ABCR10, NZ14]. quasi-variational-like [QL10]. quasicontraction [AH11b]. quasicontractions [KPR10]. quasiequilibria [AKV11]. quasiequilibrium [yZjH12]. quasigeostrophic [MS17]. quasilinear [BWL18, CW19, CCX13, CF16b, CTC17, CT17b, Cov13, DB11, DS18b, GALO18, HX14, KPK18b, Kon16, LY14, LL14b, MY10, QMW18, She18a, SC16, Wan15, Wan16b, WZ15, XLT17, YX11a, ZW16a, ZW16b, ZCY11, dSSV17]. Quasilinearization [WG11, DMD10]. quasioptimization [AKV11]. quasisecant [LW13]. quasistatic [AKLS19, CXMO19]. quasivariational [CLF10]. quaternion [MJB18, SWC11, WZ10, ZMLZ16]. quaternionic [JZJ18, LJJ17, LJLY19]. qubit [OVV^{+}16]. Quenching [XX19, CPP10, Pad18]. queries [HpD11]. query [IC12, LKK12, YXP^{+}13]. query-by-region [LKK12]. questions [ZFC11]. queue [AM11, AGD10, LA11, LXYT11, WY11b]. queueing [BA11, CTD10, HH10a, PdlF10]. quickest [Lin10a]. Quintic [Moh15, SKM11, AZ10, Che18, ML19b]. quotient [BMJ19].

R [DBH^{+}14, EM19]. R&D [BKK11]. R-linear [DBH^{+}14]. Rabinowitz [She18b, ZZ15b]. Radau [MV10]. RADAU. [MV10]. Radial [BASW18, Cov13, HT12a, HT16b, MvS18, Saj14, SL16b, Ber16, BL14, Boy10, BM13b, CM13b, DHQ11, DO11, DAM14, DM15, DM16, GL10, GK18, HT13, KKT13, KRP12, Kon16, KY11, LW19a, LDW11, LL19b, MG15, MFSL19, PT15, RKP12, RFK16, RKF18, SYY13, Tha19a, TMDTTC16, ZTZ16b, ZC16, Zha18d, ZcHS18]. radial-stiffened [MG15]. radiation [KZ16, LHZ^{+}11, LZZ11a, MCB10, MM18a, NJ16, Pal13]. radiative [BJRF19, GH16, GRS18, KB10a, PDN19]. radiative-conductive [GRS18]. radiative-convective [PDN19]. radii [BK13]. radio [CJP12, CJP15, KP10b]. radio-frequency [CJP12, CJP15]. radiographs [LWWY12]. radius [AA10a, LL12e, NNAS11c, THGG14, ZYS10]. radius-dependent [THGG14]. Radon [Amb12, AM13c, GZ14, XD10]. rainbow [LLL11]. Ramanujan [KK10c]. Random [CJCV10, MX15, VBCJ10, AHOP18, BEAA11, BNTT14, CCJ10, CCJV11, CCNT16, CPT15, CJRR11, CRRS11, CM10c, DRK11, DWZ16, DGLU18, FLDZ12, GGL13, GM18b, JKMS12, LLH14, LG17, LRCG16, Mil18, MNT15, MT19b, Nie10, Ryl15b, SWOF19, SM17, Sun11, TKBMT17, WL17a, WZ18a, WY11b, XLK11, YLS12, YL18b, Zha18h]. Randomized [ASA16, FSH10]. randomly [CBB15]. range [MS12a, Ser19]. rank [AEH19, BK14, CS14b, EHL^{+}14, KL12b, MHL11, RWW18, WWXW19, WZ10, Zha15a]. rank-one [KL12b, MHL11]. rank-two [RWW18]. Rankine [ZZHF12]. Ranking [NM11a, KSKK11, Li10a, Sae11, SSS^{+}11a, WYY11, WYK10, ZHJZ11, ZLY12]. ranks [Als10]. Raphson [SG10a, SL16c]. Rapid [OBCG19, LYL12, MSTB17, Ngo18]. rarefaction [LYN11]. rate [AF13, AR17, CCRS17, CLM12, CL12c, CXZ15a, CCCW10, DRK10, DD10, GCR^{+}18, GGGR13, GD16, HW19a, HZ18, KKLJ11, KPK18a, LBW11, LXYT11, LW11d, Ngo18, XGH17a, XGH17b, ZHW^{+}18]. rates [AM12c, BGH14, BC17, GZN19, HKJ14, JPK17, LZ12c, MT19b, Par15]. ratings [WC11b]. ratio [BSK11, CZY11, CY14b, Ebr11, GOT19, JS12b, Li10a, Li17a, SSHH^{+}18, ST15, SZ14, XA13, ZZ11a, ZZ16b, ZG18b]. ratio-dependent [Li17a, SZ14, ZZ11a, ZG18b]. Rational [LMZK16, Men18, QRMH18, RMY19, Wac10, YGS^{+}16, YS16b, ZDZY17, BC12, BAO^{+}12, DE10, EZRR10, GMB12, LTX10, Pan11, PTZ19, TUT11, YTZ17, YTL^{+}18]. rationality [CEJV16]. ratios [RTRR18, SGK18]. Raviart [CLH13]. ray [CX16, Abd18b]. Rayleigh [AHV10, BMJ19, LK13, RGVR17, Zak18a]. RBF [MFSL19, AuIK17, Bay19, BHJ14, DA18b, FLP13, GCE18, HMV18, JSGP16, LZJY16, QZF19, ZN18]. RBFs [DM16]. RBIE [OP14]. RC [LKL^{+}15]. RC4 [CM13a]. re [CD12, GH12a, RPTD10]. re-assembling [RPTD10]. re-encryption [GH12a]. re-identification [CD12]. reacting [BKR10, RSV11]. Reaction [LN98, Li10c, WFY17, AD15, AD19a, Akm15, AKL18a, AO18, AK19, AAA12, AB16, Ara18, AV14, ABN18, BDM^{+}19b, CBBE16, CL19, CL17a, CG14, CYM13, CFLX18, CCM14, DMP18, DS18b, DHGF17, EKE18, FM18, FMSV17, GD10a, Gao15, Gao17, GM14c, GML17b, HNK13, HTGSH13, HY13, HPR19, HHNLGC18, HM17a, KBAF18, KW14a, KW14b, KAA19, LCM14, LZC13, Li18c, LHL15, LDL^{+}15a, LZ15b, LDL^{+}15b, LWSL19, LZP^{+}19, LZG19, LGVS19, MGTH16, MY16, MF18b, MF18a, MBS17, MPY16, Özu15, QZF19, RKF18, She16, SOJC10, Tam16, TM18, TTM19, TF17, VAS^{+}18, WCB13, WW11c, WHS18, WCLD18, XL15, XZZ16, XZ18, XX19, YP10, YDW15, YSB15, ZWZ16, ZSW19, ZZW15, ZZ16c, ZZ17, ZLW19]. reaction-advection-diffusion [Tam16]. reaction-cross-diffusion [FMSV17]. Reaction-diffusion [WFY17]. reactions [CS13, Kei13]. reactive [BMS13, MC11, MCR11, VPR11]. reactor [CM13b, CL16a]. reactors [IL13]. reactors/bioreactors [IL13]. reader [WNC12]. Real [ACAS11, BK14, CWW15, DSK^{+}14, GS12, Kup11, Li12b, LWQZ18, MS11a, PC11a, PZ11a, RMS12, TL10b, TD10a]. real-time [DSK^{+}14, PC11a, RMS12]. Realistic [LWHL12, BV10, HB12a]. Realization [KRM^{+}10, KSF14, Wei14]. realizations [GD12]. realizing [KK12, ZPWY12]. reasoning [BMS12]. reassembly [PAE^{+}12]. reattaching [AD19b]. rebates [LZL16]. recall [KS10b]. recently [FKC12]. recidivism [SCSF19]. recipes [Res16]. recipients [KLY16]. reciprocal [Ipe12, QZ11]. Reciprocity [CPP15, SC19a]. recognition [CCL^{+}12, PC11a, TNT12, VKJ13, WPH11]. recommendation [HCHH12]. reconstructed [GZR^{+}13]. reconstructing [HN18]. Reconstruction [HLY17b, ZSZ18, CL19, CZY13, CX16, CCM14, GZ14, HH18b, KVW18, LK14, LLG^{+}11, LZJY16, dCMdSGTdC^{+}16, RDE^{+}17]. reconstructions [CM19b, JCF19]. record [McN12, MP12]. recovering [RAD13]. Recovery [SG16a, VAK^{+}19, WLW16, WLW17, WW10c, ZXW13]. rectangles [DRZ10]. rectangular [AYY12, CY14b, DGH17, FBTS19, HCT12, HMWZ16, KTA12, MMFT^{+}19, Mis14, MM18c, Nes10, PGW19, SH18, WCH13, ZW11a, ZWL11, ZPGW16]. recurrence [Dua11, JYL16, RT10, WLGL10, XC11b]. Recursive [KPS10a, KPS10b, LXYT11, ECY11, FRSC16, Ham10, HSBL11, KPP13, LTX10, LL12d]. recursively [EMR10]. recycling [DKG14, MKG13]. Red [KK14b]. Reduced [CWHW17, Che15b, DGH17, DNZ^{+}13, IQR16, NMR15, PGQ16, Zha15a, Akm15, BWZ16, BGH14, CJN19, DDMQ19, DFP^{+}13, FW18, Gup11, KLL10, LS17b, MPY16, MDW13]. reduced-order [Akm15, BWZ16, DFP^{+}13]. Reduced-rank [Zha15a]. Reducing [MNPD15, DEFP11]. Reduction [Sed13, AD11d, CM19a, CLL11, DESV18, DIS19, ES18, HTGSH13, HKS19b, HKW15, LML11, LRV13, MSQ^{+}11, MK18, Med12, MBJ16, RFP11, RF12, SB19, SSK13, VC12, WJ11, Yan11d, YZ12, YMLL18]. reductions [NHIN16, WY16]. redundant [YMM12]. reference [DVMS13, Her14, OSA13, SJL^{+}19]. refinable [LT15a]. refined [BG15, FMPR15, GSZ14, JW19a, JVMF19, JK18, Krn12, RS14a, WmN13]. refinement [AKL18b, DESV18, DVY14, EFK15, LR13, SLW11, Tod13, VFM19, WLS10]. Refinements [ADK10, HX10, Mor11b, Yan10b, FWW14, KK14b]. refiners [McN12, MP12]. reflected [Mai16]. reflection [BJRF19, Ban13, BY11]. reflector [LHZ^{+}11]. reflexive [SCC12b, XHM14, YqS16]. reformulation [DG13b]. refrigerated [GWR^{+}18]. refuge [GM14c]. regime [CEJV16, DNS16, ECJ16, HC18, Lee14, MS18, QCG15, YKA18, ZM16b, ZBL12]. regime-switching [DNS16, Lee14, ZBL12]. regimes [LBZL11]. region [AA10a, Ber16, DNZ^{+}13, LKK12, LW11c, LLH11a, LLH11b, ZM14, ZLL12]. Regional [LL13]. regions [ACD^{+}11, ANP19, Kup11]. registration [Bar17, JZR15, LGHR16, LXZ13, PZJ^{+}16, PZAR19, ZCY16]. regression [Bra13, FL11b, JWX^{+}13, KOPS13, LTL^{+}12, PATA11, QZY11, SW10]. Regular [CPP15, MW11, SJN10a, ZZT11]. Regularity [CYL17, FZ17, GFZ16, JJC11, JLL18, LL12b, Yan19, Yao10, Ye15, Zha18f, AT17, BWL18, CM15, CCNT16, CL15, Cia16, CJK17, CDN14, EG18, FAHZ17, FSZ17, FSZ18, GLR13, GGR15, Liu18c, LB18, Ma18b, MAH18, TDN19, Wen18a, WY18b, Wen18b, XZ17, Ye16, Ye17b, YWT18, YQ18, ZZ16b, ZaY17, ZZ18b, Zha18g, ZWY19]. Regularization [ACTB19, APTZ19, BCD^{+}16, BKMT17, CCR16, GKS17, Gos10, HCL12a, KJK18, KM15, KTDT17, KLRW12, LP10, LHL14a, PGDL18, PB13, TKBMT17, TKHL18, WL12b, WLXG18, Waz11b, XY14, YF10, yYsZyYL13]. Regularized [ZLW18, BGF15, Boy16, CCG18, CM11c, FKF13, FL11b, KBA11, LZ16b, LZ17, NZ14, Ros12, SBJ15, WQRZ14]. regularized-stabilized [FKF13]. regularizers [Liu15b]. regularly [ÇHK11, KMT10]. reinforced [Kia16, LHM11, RYK13, Sza15]. reinforcement [FHH13, PHM^{+}19, ZR18b]. reinitialisation [AMGC19]. reinitialization [KWAS16, TS16]. Reissner [CCN14]. relapse [ZLY17]. related [AKS10, ANR11, BCSCB^{+}15, BFG11, Che11a, DPZ13, DRS11, KT11a, KYR15, LS11d, MS11a, MN11c, NNAS11a, Noo11, ODR10, QH11, RSB14, RGVR17, SGQ12, VT11, WO10, YA11]. Relating [Med12, NODA11]. relation [BL10, CM18a, Gal10a, GC19b, HVA10, JYL16, LZCL18, YG11]. relational [GWL11]. Relations [Tia11, AH10c, BS10a, BL11, BCSCB^{+}15, JMST11, LFJ12, PKK12, RRC11, RES10, RT10, RGdSRLAJ10, SKST10, SYL10, XY16, YG11]. Relationship [BCF10, DT11]. relationships [AX11]. Relative [BKT12, BZK12, LY10a]. relatively [CGHY11, LjHO10, Nil11, ZLL11]. Relativistic [Gal10b, Gal10c, QY13]. relativity [Gal10a, Gal10d, Gal11b, Gal11c, Gal11d]. Relaxation [SZZ11, ABL15, BOT14, FVVS16, GT15, GT16, KM10, KGM11, LZZC12, LSC17, MG15, OHK^{+}19, RSS10, SKTC15, SKTC19, ZSW15, ZHW^{+}18, Zha13]. relaxation/retardation [RSS10]. relaxed [Hua10b, SS17, Ver08, Ver12, Yan18c, YLK10, Zha19b]. reliability [CCDL10, DLT12, HH10a, KLTS11, Lin10a, LLW15, fNS11]. reliable [Abd18a, DFM15, EGG16, HM18a, KTDT17].

RELS [WCD10]. reluctance [MK18]. relying [BP18, Gos10]. remaining [MSH10]. Remap [KKVS19, PBM19]. remark [BS18a, DFG19, Hua10b]. Remarks [Eba14, LC10b, LB18, RR19]. Remediation [ID16]. remodeling [NCV^{+}18]. remoteness [AH10c]. removable [SJS^{+}11]. removal [DCKY15, LX15b]. Removing [HKHK13]. renal [KLY16]. rendering [LWHL12]. renewing [VCM11]. repair [VCM11]. repairable [Tod12]. replacement [VCM11]. replacement-repair [VCM11]. representation [BAO^{+}12, Dej11, MWWL11, RMS10, RHD18, XKH10]. Representations [Ana11b, jASzZ12, DPM15, DSB19, Was13]. represented [fLcJ10, MTN19]. Reprint [GT16]. reproducing [Arq18, DA18c, WSCL11]. reproduction [AF13]. repulsion [Wan16b, ZMH16]. request [KK13b]. request-based-revealing [KK13b]. requirements [KX12]. resampling [HLC11]. rescheduling [OPDC12]. rescue [GW12b]. research [GZW^{+}18, Sen12]. reservoirs [FLWJ11]. Residual [BDGG14, GCDG17, ÖKJR19, Wei17, ABT19, BPC17, CJN19, Haj18a, HLTL17, zLYmL18, MZES12, Naz13, OC14, RH15, SNEP19, Śmi11, SJHC14, VAK^{+}19]. residual-based [BPC17, RH15, VAK^{+}19]. residual-free [CJN19]. resilient [CUK12]. resist [Lin14]. resistive [AHHM19]. resolution [CAC14, DB10, LGHR16, LXZ13, TZXP11, THGG14, WZ16, dBD17]. resolved [GWR^{+}18, NDC^{+}19]. resolvent [BK11a]. resolvents [AD12]. resolving [Boy16, DSWB18]. resonance [Bai11b, ILP^{+}11, LLL12, Yua18, ZBF11, ZLW18, ZCT18]. resonances [ACE17, Hei10a]. Resonant [GZZ^{+}16, LZM18, ZM17a, KST10, WZ11b, WZMY18, ZM17b]. resonators [Hei10a].

Resource [BPKM10, ALMLM14, qGpWhL11, OPDC12, RR11]. resource-leveling [qGpWhL11]. respect [ABR10, MP11c, WLHZ14]. respiratory [IMD11]. Response [GB18, Bae10, BCC14, BKR10, CP16a, DDLM13, IMD11, JLWX18, KLY16, LLL13, LG10, Li13, Li14, LWD15, LKL^{+}15, LY11c, MPMTV15, RS13, RSP18, XZL^{+}11, YX16, YXWL14, YZMA18, ZL14a, ZLC11b]. responses [SM14]. restart [KL12b]. restarted [JW10, WNTW19]. restarting [ZG19]. Restoration [LHLH15, Bai19, BLS17, KJK18, KSZ18, LGHR16, LZ17, LHL14a, Liu16e, LZG19, LYSZ19, PGDL18, ZT16a, Zha19b]. restoring [HM10, LX10a, UCK16, YZM^{+}19]. restrained [CLM11]. restraint [HL11c, ZTW^{+}19]. restricted [Chu10, SWC11]. restriction [MR19]. restrictions [LGG12, ZLZ11]. restrictively [WZ17b]. result [DLS18b, EGSHR10, FH16, FIW17, LYZ17, NODA11, She18b, Yu17]. resulting [Fia15]. results [AKS10, AEO15, AR10d, AKRT14, AKS11, APS12, Bai11a, BK11a, Cha11a, Che12c, Cia16, CM10c, DN18a, DG10a, DQ10, DHMU16, DRS11, GRS12, HL18a, HSBL11, JL11a, Jia12, KR11, KSJ12, KYR15, LD13a, LC12a, LS19b, LZ19b, LH10a, LL10b, LL12e, Lü14, Luc10, MP11b, MP16, MSA12, MV11, Mos10, NUH12, PB11, RA11a, Sha10, She18a, eT10, TB10, Tim14, VRD11, WAZ11a, WS10b, WLL12, XDH16, Yak11, ZBF11, ZW11b, ZZ18c, dSAC11]. retardation [Lóp19, RSS10]. retarded [KMT10, WC10b, Yüz12c]. retinal [MAN^{+}15]. RETRACTED [KYR11b, SK10b]. Retraction [JL20, JFS20, Pen11, SK12]. retractions [PB12]. retrial [AM11, CTD10, GD11, WY11b]. Retrieval [CJP12, CJP15, LKK12, xLlFwWL12, PDHL12, SBA10]. reutilization [PPC13]. revealing [KK13b]. reversal [BRROP19, SR16]. reverse [ZS11b]. reversibility [SAU11]. reversible [CYM13]. reversible-irreversible [CYM13]. reverting [BM12a, TC16]. Reviewing [AaC19]. reviews [LYX11]. revised [RA18]. revisited [PL10a, Sle13]. Revisiting [KKG10, WW10c]. revolution [PLW^{+}18]. reward [ZZXY12]. rewiring [LJYS18]. rework [CCCW10, Chu11b]. Reynolds [BS16b, Far11, GHL18, HBK^{+}19]. RFID [CJP15, CJP12, QCYL12, SPLHCB14, TZMZ12]. RH [PS12b]. RH-conservative [PS12b]. Rham [Oh15]. rheology [DKM17, ST14]. rib [LWWY12]. Riccati [BJS15, Ers16, Gen10, HVR18, KAJ11, LBW11, PIAH10, SH11, Yüz12b]. Riccati-type [Ers16]. Richardson [DZW16, DLZ17, FHZ10, FHZ13, GJX18, ZDF^{+}14, ZGD14]. Ricker [BR13b]. Riemann [Abd11, HKK^{+}16, LEN10, LS12c, Liu12, SK11c, TLR17]. Riesz [fDxZ12, Els10, LLFT17, ZLJ^{+}18]. right [AT18a, CKN11, KM12, Mok11, NM11a, Raf12]. right-hand [AT18a]. right-looking [Raf12]. rigid [DVM12, MGN^{+}16, PTH^{+}16, YGR11]. rigid-body [YGR11]. Rigorous [BPR18, DW18a]. Ring [WMZW11, EHO^{+}12, PBS12, SKdA11, Tso13, Yaz11, YAS^{+}11]. ring-deficient [EHO^{+}12]. rings [AKT10, AS11d, DZK10, DZY11, ÖI12]. Riordan [Wan10b]. ripple [MK18]. rising [ST15]. Risk [ZZXY12, CANA19, KC19, WC11a]. Risk-reward [ZZXY12]. risks [CWW15]. Ritz [AHV10, JW10, LPML19, RAD13, RS14a]. rivers [LZ19c]. RLW [AO13, KBA11, Li16b]. RN [Amb19, MPLR18, ZJ17, Zha18h, ZT18a]. road [VMAVGCM19]. roadpricing [GVJ13]. roaming [SCKH10]. Robin [DA18c, Din13, DB12, Liu13, WRY18]. robot [KBGC12, Kuo16, LH12b, YC12]. robotic [TT12]. robots [SCC^{+}12a, YLC12]. Robust [BV11, BR12b, FM18, FR16, FH17, Hua11, JZ12, Kun12, LZT11, LLC13, LZZ18, LCYC12, Lin10b, LLZ12, LCC13, MMRN12, Raf12, RTB14, ST19, Tom13, TMLF19, VMO10, WS12, AP19a, AHO16, BGF15, BS14b, CNSV17, CHBTD14, CJN19, CMS10, Hu15, HYL10, KJA10, KKL^{+}13, LZ11a, LC12a, LZY12, LW10, MSA12, PC17, SZDO10, WZY13]. rock [BV10]. Rofe [Dos10]. Rogue [QTW^{+}18, WTLS18, ZCT18, CZ18, DTYZ18, LM19a, LYC^{+}19, ML19a, WTYZ17, YTD^{+}18, YHC18, Zha18i]. Role [Grm13, BP13, KGM11, WHLC11]. roof [PBS12, ZD12]. room [JMDA13]. Root [PZ11b, Gem16, Ikh11, McN12, MP12, PZ11a, ZHB11]. Root-finding [PZ11b, Gem16, PZ11a]. root-refiners [McN12, MP12]. roots [BG10a, Boy16, GS12, KG11, LCN10, PPD10, ZM14]. Rose [CZMZ11]. Rosenau [AO13, CdR18, Li16b]. Rosenberg [CJMS10]. Rosenbrock [LO16, MT18]. Rossby [ZYSY17]. Rössler [BBDS11, MHM11]. rotary [Che12c]. rotated [ZZ15a]. rotating [CNH17, CBB15, MSH10, MMA12, ZLZG11]. rotation [GC19b, Mah14, NNAS11b, NNAS11c, Wan16b, WCQ^{+}19]. rotational [ADGL14, NWZ11, YCW^{+}14]. rotationally [HMY15]. Rothe [KQ11]. rotor [MH11]. rotor-bearing [MH11]. rouge [CTSX16]. rough [CP10, Gao12, LXL12, LSZ16, MZQ11, PB13, SH12b, WC10a, ZW11c]. roughness [CD12, LYN11, ZFZ10a]. Roumen [Sen12]. Round [Ikh11, LJJ11, LLSS13]. Round-off [Ikh11]. rounds [KLL10]. router [SSESG12]. routing [KHIB12, MKA^{+}10, SRS11]. row [Ji14]. RPIM [LL19b]. RTk [Ben17]. rule [LZH16, LR17a, LSZ11, QZY11, SWC11, WD12]. rule-based [WD12]. rule-of-thumb [QZY11]. rules [AuIK11, HL18b, KOPS13, PCS13, Pap15, Som13, SAR18, WV15, XG10]. run [CM12b, FER15]. run-time [CM12b]. Runge [Boy10, FHZ13, KMS10, KAS11b, MVKK14, MBJ16, SLL17, TFS11, Tsi11, ZGD14]. runs [MP11b, MP16]. rupture [LZ10]. Ruscheweyh [Lup11, NA11].

S

[Abd18b, EM19, KKAM11, Pan17, HSMG12]. S-boxes [HSMG12]. S-H [KKAM11]. SaaS [KLH^{+}12]. SABR [LSV18]. SABR/LIBOR [LSV18]. Saddle [Ant10, BJQS18, BJ19, Bac14c, BQ15, BQ17, BLyS18, CM16a, CMR17, Cao19, CCG18, CBS18, Che15a, DYWL19, FZ14, FWZ16, HH11, HWXC16, HWXC18b, LM17, LM18b, LZ15a, LZ16b, LZ18b, Mia18, SR17b, SWS19, SBJ15, SS16b, Yan18c, Yun13, ZZ18a, ZL16, ZZ14, ZYW17]. saddle-point [DYWL19, Yan18c, ZYW17]. safe [KAK^{+}12]. Saha [Abd18b]. Saint [ZTR11]. Saint-Venant [ZTR11]. Salagean [Lup11]. Saleh [EM19]. sales [SLCC12]. salesman [LYC12, tWqLzGkP11, ZGZ13]. salesmen [San11]. Salié [ZP10]. Salt [DCKY15]. Samarskii [BCK11]. same [Ikh11]. Sample [YYL11, CHL18, HW11]. Sample-weighted [YYL11]. sampled [LCYC12, THY^{+}10]. sampled-data [LCYC12, THY^{+}10]. Sampling [GV11b, HN18, Chu12a, Li12b]. sandwich [ASMEE11b]. Sasa [SR17a]. satellite [UKAL10]. satisfying [MW13, Tsi11]. Satsuma [LWW19a]. saturated [GQF^{+}10, KMT19, MCR11, REHA11, RRAK19, SSS16, TRL19, UABK16, Yan12a]. saturated-unsaturated [SSS16]. saturating [RHMA18]. saturation [CNSV17, XMW10, Zho13]. Sayed [DR12]. SBFEM [LLL16]. Scalable [Kan15, ALLQ13]. scalar [Fer12, Pal12, SLKK19]. scalarization [AHF10]. scale [AFGL10, BMS19, BJS15, CCY10, CChL14, CMT12, DGOZ13, FGHZ14, FLH10, FL11a, GW12b, GK19, LC12a, LHY11, LY13, LCQF19, Liu16a, LFZ19b, MS17, OPDC12, RJGS^{+}19, SCA14, SM14, TAPA^{+}17, THY^{+}10, TRL19, WLYX13, WL17b, XBHN16, YWW^{+}12, Zha14, Zha15b, ZLC^{+}11a]. scaled [DM15, Śmi11]. scales [Ana10, Ana11b, BG10c, Can11d, ÇT12, DGK10, Dos10, FL10, GGEB12, HY10, HEP10, HZ11, Hon10, LCW10, Li10d, Li11a, LXP11, MP11a, MT11, MZB10, PW11, Rah11a, Rah11b, Sar10, TD10b, WSG10, WZ11a, YX11a, Yas12, Zen11, ZFZ10a, ZFZ10b, ZHJ11, ZGW11, ZQ11b]. scaling [LS10a, Mac12b, XM15]. Scattered [ZL13, HH10b, LW19a]. scatterers [HMY15, Liu18b]. Scattering [HCT12, BJRF19, BT14, CHS18, Col18, CDS15, GN19, HMY15, Hei10a, HMZ18, KK19a, LZ18a, LSZ16, MDG19, NJ16, Par17a, SBB^{+}18]. Schauder [Yil10]. Scheduling [WSS10, AK12, AA10b, ÁBÁPM11, AA10c, DM12a, FLWJ11, KX12, Lee11c, LLW11, LCC12, LBH^{+}12, MMR10, MMR11, RT11, RR11, Wan08, WW10a, WWW11a, WWW11b, YY10c, YH11a, YX11c]. Scheme [CM11a, YMDZ10, ABT19, AZ17, Ara18, AGPCC10, AKA11, BN14a, BSL11, BMSS18, Bra10, BGM19b, CB11c, CFN11, CBBE16, CG13, CTZ17, CV14, CKRW19, CPL11, CCM14, DL19, Def10a, Def10b, DM15, DM10, DGT18, DFW^{+}18, DH18, DZ17, DLT10, DL11, Dun18, EO14, EAMA19, FSH10, FM19, FM11, Fis18a, FIS18b, GMS15, GS15b, GDM13, GA10, HMF^{+}19, HD14b, HLL^{+}15, HD16, HLSN16, HD19, HKW15, IC12, KKT13, KO11a, KS10a, KK13b, KA10b, KSMT11, KSS13, KS15b, Lee11a, LS19a, LJJ11, LF11b, LLML15, Li16b, LW17, LZPZ19, LZZ19, LWHY10, LC10c, LCLL16, LH12b, LZB15, LLYL19, Ma19, MD18, Mah14, MDBCF16, MP19c, MMH11, MM18d, PZZ18, PCO16, PC17, PBM19, PW18, PRS18, QZ16, QX19, QCS^{+}19, RAZ19, RZ16, RR18, RS15, RQ18, RTRR18, RRO17, SNMA12, ST15, Saj12]. scheme [SZA^{+}18, Sha12b, She12, SWL16, SLW14, SR10a, zSdZ10, SSL14, SR18, WZXL11, WMZW11, WW14a, WV14, WV16, WZM^{+}16, WH16, WLL^{+}18, WBZY18, WRY18, WMP^{+}19, WZC^{+}19a, WSCW16, WX18a, WCLD18, WCZ^{+}19, XGH17b, XW18, YZMZ16, YWK^{+}10, YT13, YT18b, YK17, ZYWZ17, ZPWY12, ZZL^{+}18b, ZÖXL^{+}19, ZC11c]. schemes [AD16, AM12c, AY12, AA13, ADGG13, ADGL14, BZ18, BS18a, BDR19, BKNR19, Bra16, BMH19, CGH14, CYP16, CLM12, CLM14, sCYhX18, CS18, Cie13, CG14, DNS16, DL16, DZ18, Dub13, ED12, FDXW11, hGzS15, GGT14, GL16, Gur13, HB19, Ima17, JYYL16, KA11, KVV14, KPG18, KRBS18, KGM11, LCM14, LZ16a, LX18, LLG10, MP19b, MHM11, NH15, OF16, PS16, Por18, Rhe10, SZC^{+}18, STDLM19, Tan18, Wan16a, YL13, YDW15, ZMM18, ZH15a, ZQ14a]. Schiff [AKMUH17, CM18b]. Schinzel [CBKR10]. Scholes [Bho14a, Bho14b, CHXL18, CXZ15b, CWY19, Dav17, DH17, HG18b, Jum10, LC13, Moh15, WY19a, ZLTY16, ZH18b]. Schrödinger [Ali15, AS11b, Bao16, BZ18, BC10, BJPT16, CDW11, CF16b, CTC17, CTG17, CT17a, CHZ19, Cov13, DM16, DZ16, FJ19, FZ18, GL10, GTC18, GMP18, GL17b, HP17, HA16b, HXL11, Hu18, HTWS15, HXS^{+}15, HLX18, Iom18, JZJ18, JKS19, KMS10, KK13a, LS17a, LL19a, LK18, LRZ18, LT15b, LTJ^{+}16, LC16, LLT16a, LLT16b, Lü14, LL14b, Luo18, MKHC11, MI16, PZA19, PJ17, PLT17, QCT17, She18b, SC16, SMBY10, SBM13, zSdZ10, TCHW19, VMC^{+}14, Wan15, WHTZ16, WH16, WZKY12, WZ15, XLT17, YL16, Ye17a, YTC^{+}18, YW19, ZDLC14, ZTZ16b, ZC16, ZZM17, ZH18a, ZLG18, ZW11d]. Schultz [Pet14]. Schur [EMRS12]. Schwarz [Cal19, DdSF13, DGR18, FMPR15, KZ16, LDHH13]. Schwarzian [Asl10]. Science [GGAVGG19, KLK13, LAM^{+}16, YLLK14, DELK13, He11, HLK10, KFYW11, LLR^{+}19, PHWM10, ZLG^{+}11]. Scientific [GIM15, FGHZ13, FGHZ14, LCQF19]. scores [Mok11]. SCR [CI18, CHY19]. SCR-based [CHY19]. scrambling [LH10b]. scrap [CCCW10]. screening [LSW16, WD12]. screw [HHS^{+}17]. SCTP [LJJ11]. SDE [HW11, WZG19]. SDEs [ZZWG16]. SE [RAZ19]. search [EMRS12, HSD10, LW11c, Mai16, MO14, OMS10, SBKS12, VB10a, tWqLzGkP11, WZ17b, WQNF12]. Searching [YC10b, AAR11, CD14, CDP16, CM13a]. Secant [CAH11, Cvi11, EGGS^{+}12, WKG10]. secant-like [WKG10]. Second [BGM19b, CHS18, GGT14, KKVS19, MA10b, PN10, WLL^{+}18, ZM16b, AAZ10, AJ10, AJ11, AJ12, ALI11b, BM11a, Bac14b, BD11a, BD11b, BQ17, BDS10, BS16b, BBBM16, BMH19, BS15c, Can11d, CM11c, CT11b, CLTA11, CWQJ12, DGL12, DWS19, Don10a, DFW^{+}18, ED11a, FM11, FN14, FZL^{+}18, GY15, GMI12, HSD10, HH17, HEP10, HLL^{+}15, HF10, Hua10a, KO11a, Lee15, LL11, LZKU11, LCZ11, LX15b, LLLW18, uHS12, MM18a, MSG11, MBJ16, MB10b, NPD17, NZ16, OCNG12, ÖZ11, PC17, RSDR11, RCH19, RR14, SNMA12, SL12, Sal16, SBvdV13, Set12, Sou12, SKM11, TJ10, THC^{+}18, TMDTTC16, Tod15, VRD11, WS11a, WRW13, WTM17, WW19b, WLZ18b, WHS12, XZC12, YX11a, Yan15, YJ19, YZY10, Zak18a, ZA10, ZYWZ17, ZFZQ10, ZJ10b, ZGW11, ZQ11b, ZL11, ZLL12, ZHL12].

Second-invariant-preserving [KKVS19]. Second-kind [CHS18]. Second-order [BGM19b, WLL^{+}18, ZM16b, AAZ10, ALI11b, Bac14b, BD11b, BDS10, BMH19, Can11d, CM11c, CT11b, DGL12, ED11a, FZL^{+}18, GMI12, HSD10, HH17, HLL^{+}15, KO11a, Lee15, LL11, LCZ11, LLLW18, MBJ16, MB10b, SBvdV13, SKM11, TJ10, THC^{+}18, TMDTTC16, Tod15, WTM17, Yan15, ZA10, ZYWZ17, ZFZQ10, ZGW11, ZQ11b, ZL11, ZLL12]. secondary [DFM15]. Secret [KK13b, OO10, ZPWY12]. section [LZ12b, HPV^{+}18]. sections [Ahn10, WY18a]. sector [ATH18, CCH^{+}12, MG15]. sector-based [CCH^{+}12]. Secure [SPCS13, dlPBTW13, CPL11, FSH10, HPY10, KRM^{+}10, KLL10, LYM12, LLX^{+}10, OO12, OYXK12, PKTH13, UKAL10, WMSH11, yYqWqZC13]. Security [AJJAD^{+}10, LLSS13, SPCS13, Ge10, GH12a, KKG10, KX12, LHHZ12, LWHY10, LC10c, NCL13, Ögü13, PHWM10, QCYL12, Zha11b]. sediment [CWW19]. sediments [JQG14]. seepage [FJB19, KP19a, SVY16]. Segel [FL13b, FHA16]. segment [FdOdSS17]. segmentation [CB19a, FJWW16, FQLC18, HCL11, LK11, LTL^{+}12, PZAR19, QXLL11, XJYL17, YZWW14, yYsZyYL13, jZsQdLmG19]. segmentations [HT16a]. Segmented [DE11]. segregation [AB16, Ara18]. Sehgal [PRR10]. seismic [BPM12, LKL^{+}15, Por18]. seizures [HPR19]. selectable [GLM^{+}11]. selected [PKD19]. selecting [Che12a]. selection [BKM11, BKK11, CSSW12, DFS11, DMPV10, HLC11, HCL11, KJK18, KBDC12, KLP10, LS11c, LHL14a, MLY11, PTL13, WZ17a, WZG19]. selective [KÖC^{+}18, SSIP19, SM19, BGM19a]. selector [QZY11]. Self [BGGCGRSP16, YS19b, Bur13, FOX11, FSCG11, HP10, HM19, JWX^{+}13, KLMV12, TBP19, YMHL18, Yu11b, pZ10, Zha17b, Zha17a, ZLY^{+}13]. Self-adaptive [BGGCGRSP16, FOX11, HM19, Zha17a, ZLY^{+}13]. self-adjointness [Zha17b]. self-consistent [YMHL18, Yu11b]. self-force [TBP19]. self-normalized [HP10, pZ10]. Self-Optimizing [YS19b]. self-organizing [Bur13, JWX^{+}13]. self-shrinking [FSCG11]. self-similar [KLMV12]. selfadjoint [AD10a, AD11a, Dra11a, Dra11b]. selfish [CQLX11]. SEM [ADZ19]. Semantic [JTCC11, Ogi12, BL11]. semantics [WYY11]. Semi [CH17, DYH11, LM17, LXL14, MPMTV15, TC16, YTL^{+}18, ZZ14, AJAR18, BG15, CM16a, CWHW17, CCG18, CLC16, De 10, DFM15, DBH^{+}14, EO15, GGGR13, HKJ14, HHY13, JKB11, LL14a, Lee15, Lee17, LM19b, LRH13, MWWL11, Mia18, PLW^{+}18, QRMH18, SRGL13, TA11, UKI11, WLYX13, XZ10, YH19, ZSZ17, ZL16]. Semi-analytical [MPMTV15, AJAR18, LL14a, Lee17, TA11, UKI11]. Semi-convergence [LM17, ZZ14, CM16a, CCG18, LM19b, Mia18, ZSZ17, ZL16]. Semi-discrete [CH17, HHY13]. Semi-foldon [DYH11]. semi-implicit [DFM15, XZ10, YH19]. semi-infinite [CLC16, De 10]. semi-inverse [JKB11]. semi-Lagrange [EO15]. Semi-Lagrangian [TC16, BG15]. semi-linear [HKJ14, Lee15, LRH13]. semi-monotonic [DBH^{+}14]. Semi-rational [YTL^{+}18, QRMH18]. Semi-structured [LXL14, CWHW17, GGGR13, SRGL13]. semi-supervised [MWWL11, WLYX13]. semiconductor [FP19, HDS11, HHS^{+}10, MDW13]. semicontinuity [AKV11, CLF10, SSL11, WL17a]. semidefinite [DP15]. semidiscrete [NSYY13]. semigroup [Buo11, CLCF14, Pir11, SKdA11]. Semigroups [SJN10b, Ali11a, BR12c, JMLF11, SDM10, SK10a, SJN10a, Yan11a]. semihypergroups [AMD10, JMLF11]. Semilinear [ANP19, Bra16, CL17a, CZN12, CCD10, CC17, HLW19, JPK17, JFC14, KB15, Li18c, LX15a, MT18, MT19b, Oua12, SRM11a, SS16c, Tia17, TNHK19, Wu18b, XZC12, XDH16, ZH19, lZxLhY12]. semilocal [CAH11]. Semimartingale [Dun11]. semiparametric [LLG^{+}11]. semipositone [DGK10, LS11a, ZJ10b]. Semirigid [KP19b]. semiring [BMS12]. semiring-based [BMS12]. semismooth [XX17]. Senator [LX10a]. sense [GDZ11, Set12, ZRC11, sHC11]. Sensing [BMP15]. sensitive [San11]. sensitivities [GSS11]. Sensitivity [ADY12, WCB13, WLHZ14, AD19b, BNP18, DGOZ13, JLL19, LGL^{+}14, ODAZ15, SJL^{+}19, WLT13a, WPL16, Wan18, ZMWH18]. sensor [CQLX11, CL12c, CMS10, Chu12a, HB12a, HCL12b, JCWZ16, PHPK12, SLXC11, ST12, WCB13, YXP^{+}13, YDK^{+}12, gZnZpZbD12, dlPBTW13]. Sensorless [LC12b]. sensors [LX12b]. sentiments [FPW^{+}11]. separable [LL17, XJLX10]. separated [Dol11, STS19, YL18a]. separating [AD19b, PXT10]. separating-reattaching [AD19b]. separation [IK12, LD11a, TAPA^{+}17, YXX11, ZS18]. SephadexTM [LCA^{+}17]. September [Ano18-74, Ano18-56, Ano19-53]. septimal [Che18]. sequel [KAA19]. sequence [AES11, ÁBÁPM11, BD16, CM12a, DÇ12, ECY11, FKDN15, FBB10, Ham10, Haz11, KB10c, Lee11c, MN10c, MN11c, MKPS11, NNL13, Sön11, TÇ11, TD10a, YT12]. sequence-dependent [ÁBÁPM11]. sequences [Bor11, Çak11b, Çan11a, ÇHK11, Çan11b, CM10c, DG10a, DGOZ13, Hak14, KS11, LTX10, LFJ11, LSJ12, MP11b, MP16, Mor10c, Mor10d, MME10, XC11b]. Sequential [AN12, Fur13, NKA18]. sequestration [TZ18]. serial [CKK^{+}10].

Series [RY11, Tur10, Ala10, AM13c, CCJ10, CCJV11, CHT11, CS10b, ED11b, Gos10, GR13b, HZ18, LS11d, MWWL11, Mor10a, PW11, SMBY10, yS10, WHG11, WS11b, YMSL11, dPRVRB13]. series-form [HZ18]. server [LXYT11, WY11b, WMSH11]. server-aided [WMSH11]. servers [GD11]. service [BL11, Cha13, CTD10, DFS11, KK12, LLX^{+}10, MKA^{+}10, TYY^{+}12, YC12]. services [KC12]. servicing [VC12]. servo [BMJ10]. set [AMGC19, AvdW13, BS10a, BBR10a, BDF16, BR12c, CMS10, CChL14, CP10, CCP19, CAC14, DLQ16, FHH13, GXZ10, GXZ11, HMF^{+}19, HRMS10, HRMS12, Hon10, JPS10, JLP10a, JY11, KR11, KWAS16, LKS10, LC10b, LCC12, PKK12, PZAR19, Pop13, ST15, SPT17, TS16, WD10, WG11, YLY^{+}09, YG11, ZW11c, vdW14]. set-up [LCC12]. set-valued [AvdW13, CMS10, HRMS10, HRMS12, JPS10, KR11, LKS10, WD10, vdW14]. sets [AKT10, ATZ11, ASN11, ACD^{+}11, BS11a, BL11, BK10, CCBSRFRM11, CT10a, DDD10, Dol11, FLLF10, Gal11a, GV11b, GN11, HDT11, HLY12a, JTC^{+}10b, JTC^{+}10a, Jia11, JTCC11, JMST11, KST10, LMW10, LZH12, LXL12, LWBW13, LZYW13, MSQ^{+}11, MS10a, MZQ11, MN17, Pop14, SA11a, gShYL10, TK11, Tia11, WYG12, Wan13a, XGXZ10, XZC12, XMWH10, YLJ12, ZJ10a, ZZT11]. Setting [LYLX11]. settling [CV14, DFM15]. setup [ÁBÁPM11, Lee11c]. seventh [FM19, RG18]. seventh-order [FM19]. Several [AB18, HR15, Zha18g, LXK11, MJ10, MRR11, Ped18, RSS10, WLW16]. severity [HSBL11]. SFDD [JS12a]. SFOPDES [GGAVRC^{+}19]. SGL [YLF19]. SGS [AD19b]. Shadow [XLZW11]. shadows [JK11c]. Shafer [KBDC12]. shaft [Che12c]. Shallow [GZD18, HIS19, KK19b, LMPE18, LLG10, LZ19c, Men18, MDBCF16, PLMS14, RAZ19, RRO17, SZA^{+}18, WBN18, WFL11b, YLG10, ZÖXL^{+}19, ZQ14b, ZDZY17]. shallow-water [MDBCF16]. Shamanskii [HM19]. Shamanskii-like [HM19]. Shan [KM10]. Shannon [BPM12]. Shape [APT11, DRZ10, ZB19, CHL18, CPT15, DO11, DNS18, FJB19, FKDN15, GW15a, KP19c, LK14, LH19, Luh12, Pop14, SR10b, Tha19a, WvDRG19, Wu11b, YLLN16]. shaped [CNR10, She16]. shapes [FKDN15, WY18a]. share [LY11a, LSM11, ZPWY12]. shared [WKP^{+}14, WY19b]. sharing [CZ10, OO10, OO12, ZY10a, ZX11, ZPWY12]. Sharp [Cia16, GH12b, XL10, Ciz12, KT15, LG12, Liu10]. sharpening [MA16]. sharply [MSTB17]. shear [ATH18, BKR10, FZBF10, HHS^{+}17, HZLM10, HWY14, Kia19, LGC^{+}17, MSV18, MJWD19, MLSLM15, Sea15, VGC^{+}15, Wan14, WD16, Yan19]. shear-dependent [HZLM10, Wan14]. shear-thinning [BKR10, HHS^{+}17, MLSLM15]. sheet [CYZZ18, Das12, FG18, KWFY11, RRP16, Sah11]. sheet-driven [Sah11]. sheets [GF16, KHF^{+}19, RY10, WL17b]. shell [CRA19, DRS11, Hak14, LGC^{+}17, MTN19, MJWD19, PBS12]. shell-like [DRS11]. shells [GR15, LPML19, PLW^{+}18, VĆV11, ZTW^{+}19]. Shepard [CDD12]. shield [YLDL11]. shift [CM16a, CMR17, Cao19, HWXC18b, SR17b, SS16b]. shift-splitting [CM16a, CMR17, Cao19, HWXC18b, SR17b, SS16b]. Shifted [DHY19, GHM^{+}14, LHL^{+}14b, WmN13, WNTW19, ZT16a, ZG19]. Shishkin [DN18b]. shock [Ban13, BSS18, BMS13, EK16, ILV^{+}19, XZL^{+}11]. shocks [GM18b]. shooting [SM14]. shop [AK12, DM12a]. Short [FLWJ11, ABB17, BMY13, CPL11, HCL12b, VMC^{+}14, WC10a, XH11b, ZPWY12]. Short-term [FLWJ11, HCL12b]. Shortest [Gao11, YDK^{+}12, TMMASG10]. shrinkage [WC11a]. shrinking [CGHY11, Das12, FSCG11, KWFY11, KN12, RRP16]. shrinking/stretching [KWFY11]. SHSS [LM19b, ZM16a]. shunt [Ala10]. SIAM [HPV^{+}18]. Sibanda [Pan17]. side [KM12, KLL10, STS19]. side-channel [KLL10]. sided [FZL^{+}18, MY13, SJPS11, ZD12]. sidedly [CST14]. sides [AT18a, CDG15, NM11a]. Sign [LC10d, Bao16, CTC17, CTG17, CT17a, CHZ19, CT17b, DL14, DLS18b, HY10, HY18, LL19a, LL11, LY14, Li18b, LC16, Su12, ZT18a]. sign-changing [Bao16, CTC17, CTG17, CT17a, CHZ19, CT17b, DL14, DLS18b, HY10, HY18, LL19a, LL11, LY14, Li18b, LC16, ZT18a]. signal [BPM12, JLL19, ZT19, ZHJD13, ZXW13]. signal-dependent [JLL19]. Signaling [SLXC11, KPG13, XBHN16]. signals [CM10a, DB10, HCT12]. signature [CPL11, FSH10, LWZG10, Tso13, WMZW11]. signatures [WMSH11]. signcryption [HPY10, LHHZ12, LWHY10]. Signorini [AD16, LY15, Zha17a]. signs [KK12]. silica [SKCL19]. silico [CANA19]. silicon [SKH12]. similar [KLMV12]. Similarities [OAY11]. Similarity [KKK16, Zha19c, BL11, CYZZ18, FPW^{+}11, KKK14, KKK15, KK17, KTK18, Lam12, VP11, WK13, XZL^{+}11]. Simple [CUK12, KKJ15, AP19a, AA15, AAP12, Che11c, Che14, FGB19, Gur13, HD14b, HL11c, LK15, LMW10, Liu18b, Moh14, WCH18, WHS18, XY15, ZZ19, Zha15b, Zhe11]. Simpler [ZG19, ZWY19]. Simplest [CJ18a, HMWZ16, JKK12, YS16a, ZSAN18, ZSAN19]. simplex [WZW10]. simplexes [Von19]. simplices [KK14b, RR19]. simplicial [Tod13]. simplified [GFZ16, XWL18, YF10]. Simplifying [NCV^{+}18, XA13, Tsi11]. Simply [KBS11, ZWL11]. Simpson [SSO10]. SIMTHESys [IBG12]. simulate [DFJS10, RR14]. Simulated [dCMdSGTdC^{+}16, Che12a, VB10a, ZLY^{+}13]. Simulating [ALHZMC^{+}19, AML^{+}14, BDPM12, LR17b, YNS^{+}14]. Simulation [ARESH18, AM10b, BHK16, BR13a, DM10, GK19, HLSN15, JMDA13, KRAS19, MM19, PC14, QWGJ15, ST14, ZR16, AM13b, ASSV18, BCCZ18, BKL14, BGM19a, BRFH16, CZY11, CLB14, CCY18, CGS12, CJPB10, DRT^{+}15, DGL12, DM16, DSK^{+}14, DSL11, FSZS18, FER15, GQF^{+}10, GLW13, GZR^{+}13, GW12b, GSY10, GIMZ14, HBK^{+}19, HHS^{+}17, HS18, HPR19, HTL10, HLL^{+}15, HKW15, Ima17, JK12, KP18, KORR10, LK13, LZZC12, LCLL16, LYN11, LDL10, MNJ^{+}13, MMFT^{+}19, MJ14, MNPD15, MD15, MDW13, NKM16, NNL13, NDC^{+}19, Özd18, RMK19, RTRR18, RCG15, SSM^{+}17, SK14a, Sag10, SSIP19, SRDD17, SG11b, SGK18, SNEP19, SKTD13, SKFG11, TTT10, THC^{+}18, TY13, UMLF13, Wan19a, WWH12, YSW16, YZ15, YGH11, YLL^{+}14, YCW^{+}14, ZY15b, ZZL19, ZZC13a, ZZC13b]. Simulations [AR17, AVV18, ADD^{+}15, BGP13, CIN^{+}18, CYP16, CM15, Cos18, DSWB18, DTR19, DFM15, FP19, GGAV18, GGGR17, GSD^{+}19, HH17, HMSC10, HIS19, ILP14, JK10, JK11a, JGK13, JO19, KHWK10, KVR11, Kup10, Kup14, LDW11, LLML10, LHM11, MMS^{+}18, MUB^{+}16, MLG17, OHK^{+}19, PTH^{+}16, RKD18, Rua19, SSHH^{+}18, SL16a, SSK13, SZC^{+}18, Tro13, WWLL13, WL13b, WBN18, WS16, WCZ^{+}19, XFH19, YLF19, ZLS19, ZLHF19]. simulator [LWHL12]. Simultaneous [CL19, DIS19, HLI14, Ikh11, LW19b, MO14, XC11a]. simultaneously [Ned12, WQRZ14, ZL12b]. Sinc [MN11a, AEG18]. Sinc-collocation [MN11a]. sinc-Galerkin [AEG18]. sine [KTDT17, LZWC16, MPfTX18, WL17a, ZWG11]. sine-Gordon [KTDT17, LZWC16, MPfTX18, WL17a]. Single [BTEM19, DM19b, Wan08, WW10a, WWW11a, WWW11b, YH11a, ALHZMC^{+}19, BRFH16, BO10, BSN13, CNH17, DJD18, Gal12, JL12, KGJ11, Kia18, Lee11c, LLW11, LYLX11, LXYT11, MR19, MMR10, MMR11, OHK^{+}19, PP10, SYO12, TY13, WHC12, WS16, WL17b, XWY17, YY10c, YX11c, ZQ14b]. single- [BRFH16, OHK^{+}19]. single-buyer [BO10, KGJ11]. single-eyed [CNH17]. single-layered [WL17b]. Single-machine [Wan08, YH11a, Lee11c, LLW11, YY10c, YX11c]. single-pass [WHC12]. Single-phase [BTEM19, DJD18, WS16, ZQ14b]. single-stage [JL12]. single-step [XWY17]. single-vendor [BO10]. single-walled [Kia18]. singletons [AD11c]. Singular [BCF10, LCP16, WCCS15, AR10a, ABJ11, AM10c, AD10b, BS12a, Boy16, CHS11, CM16a, CCG18, CCX13, CCCW16, CR13, CG15, DYWL19, DLS18b, Elb15, Erg19, FID14, FZ14, FSHZ11, Fer12, GMI12, HLWX11, HS11a, HKI12, HH11, HX14, HWXC16, Ibr11, Jal14, JW18a, KA10a, KN11, KS12b, LW11a, LFC16, LS16, LM17, LM18b, Li19b, LM19b, LZ15a, Mia18, MS12b, NB17, Ols10, Oru19, QX19, QWL19, QCG15, SC19a, SR17b, SS16b, Sta11, Su12, WCH18, Wei10a, XLD11a, XLD11b, YCLY15, Yüz11, ZSZ17, ZJ10b, Zha10, ZLC^{+}14, ZL16, ZMWH18, ZZ14, ZCY11, dSSV17]. singularities [Cho17, CCK18, Dos18, FRSW11, KL16b, PPC13, RSWZ10, WS17]. Singularity [YMM12, CK15, DBS12, NB17]. singularly [AHO16, BZT16, CG14, CJ15, CX18, DN18b, FM18, FR16, FH17, HO19, KAG11, KSMN11, KK14c, LN98, Li10c, LGVS19, MN10a, RS12b, Tur10, WSCL11, WZY13]. sinh [FYYT11, LZWC16, LS16, SH11, Wei10b]. sinh-cosh-Gordon [FYYT11, SH11]. sinh-Gordon [LZWC16, Wei10b]. sink [EA10]. Sintered [FSM19]. sintering [WA19]. sinusoidal [HHM12]. SIP [CCK12]. SIPG [BCG17]. SIR [Aki17, Gur13, LJYS18, MMOJ14, Sun10, ZK16]. SIRH [ZLY17]. SIRS [XMW10]. SIS [CHS19, LRV13, XL15]. Sit [RIW12]. Sit-to-Stand [RIW12]. site [LLL13]. sites [LYX11]. Six [Boy10, BJLZ12, YMM12]. six-DOF [BJLZ12]. sixteenth [GK11b]. sixteenth-order [GK11b]. sixth [AH11a, DLZ17, LJ10a, WX18a]. sixth-order [AH11a, DLZ17, WX18a]. size [ALMLM14, EAAS18, GLW13, GGGR17, GD16, HL12, HHM12, RMK19, SSPL10, WAG^{+}14, WHLC11]. size-structured [ALMLM14]. sizes [AM12c, SM19]. sizing [CCY10, DRK11, DH10c, WW10c, ZLG^{+}10]. skew [DH10b, JW15, SWC11, TL10b, WLD13, WLM13, WHW11, XCS18, ZYW15]. skew-Hamiltonian [XCS18]. skew-Hermitian [WLD13, WLM13, ZYW15, ZS16]. skew-symmetric [TL10b, WHW11]. skewness [BKM11]. Sleep [SSM12]. Sleep-based [SSM12]. slender [FXCC18]. Slepian [Gos10]. slice [IS12]. slider [Far11]. Sliding [HGN^{+}10, CZMZ11, CL12b, FDB13, FKC12, HLCY12, Kuo11, LZY12, PC12, YWHC11]. sliding-mode [CL12b, Kuo11]. slightly [AP19b, ZR18b]. Slip [Das12, DSM18, GB16, HKS19a, IBSS11, LTT13, MC10a, RR14, Sah11, UKA15, UABK16]. slip/temperature [RR14]. slipping [STS19]. slope [RAW^{+}16, VFM19]. sloping [FER15]. sloshing [WCH13]. slotted [TZMZ12]. slow [BG13]. slowly [Çak11b, LSM11]. Smagorinsky [GWR^{+}18]. small [Bao16, BO18b, CHM18, Gur13, HAESLB14, JMDA13, KJA10, KALAS11, KYW^{+}18, PQBK17, Sha14, WH11b, WZY13, WW15, YLY12, Ye15, ZX11, ZH19]. small-gain [KJA10]. Smarandache [AS11a]. smoking [EZM12]. smooth [AM10c, CsH10, CS11c, DSL11, DHMU16, HH18b, JZL18, KV17a, KV17b, KBAF18, LZ19a, SDH13, Seg19, YJ19, YN16, ZZ16a, ZZL18a]. Smoothed [ATO19, WRW^{+}19, FM11, LL19b, SPP18]. smoother [ACC18]. smoothers [Che15a]. smoothing [CM11c, LGC^{+}17, MZES12, NZ14, ZXW13]. SMS [KRM^{+}10]. Sobol [DGOZ13]. Sobolev [BKT11, BK12a, DSL18, FWW14, HGHA19, HCF16, HR19, LL16a, LS19b, LLZ18, RSB14, She18a, SW16a, SC19b, WS17, ZLL17]. soccer [TT12]. social [DW18b]. Soft [AKT10, Ali11a, AS11d, BS10a, ÇE10a, ÇE10b, ÇKE11, MZQ11, SA11b, THB12, WYG11, YKD11, ZJ10a, ASN11, BS11a, CD12, FLLF10, GXZ10, GXZ11, AB11, HLY12a, HA11, JTC^{+}10b, JTC^{+}10a, Jia11, JTCC11, JLP10a, JY11, LZH12, LFJ12, MSQ^{+}11, MS10a, MS10b, Min11, ÖI12, PKK12, SA11a, SN11, TK11, WYG12, Wan13a, XGXZ10, XMWH10, YLY^{+}09, Yan11a, YG11, YJZ11, YLJ12]. Software [HH10a, SBB^{+}18, CCDL10, JPP12, KLH^{+}12, LZL^{+}18, LYL12, NEB14, PLT^{+}19, Rob14, ZHQG12]. soil [GQF^{+}10, LLL16]. soils [WC11a, YLDL11]. solar [BZZ^{+}10, SKH12]. Solid [WKBR18, BPZ19, CPP10, CM19b, GLLC19, HAESLB14, HLSN15, JNJ^{+}11, LL19b, MMT18, SK14a, SRDD17, WAG^{+}14, ZLC^{+}11a]. solid-solid [CPP10]. solids [FM17, RNQ13, RNQ16]. Solitary [AKMUH17, DTYZ18, KK19c, Cha11c, Chu11a, ES17, FER15, GZW^{+}18, KK19b, Kuo18, LD11b, LGZ19, LGH^{+}11, zLZ19, Man18, NTR15, RNB11, Sea15, SIL19, SMYK19, TZ15, Tao18, WTYZ17, WF18, Wan19a, WL16, YTD^{+}18, YYLW19, ZLL14]. solitary-wave [TZ15]. Soliton [BC10, JPB11, KK17, KT18b, YYC11, ZTSC16, Bis10, Elb15, GTZ19, HLLM19, HTY^{+}19, HTWS15, KTK18, Kuo18, LMMF17, LMDL11, LTSW16, LL16d, Liu18a, LZM18, zLZ19, LWW19a, MCF18, ML19a, MGS^{+}14, OVV^{+}16, Osm18, XX10, YTD^{+}18, ZZ11b, ZCZ17, ZDM11, ZTZJ14]. solitonary [NNWAS11]. solitonic [LWL11]. Solitons [CTSX16, ZSAN18, AGT19, CDW11, Che18, Elb15, QTW^{+}18, TTG16, WW18a, YHC18, ZCT18, ZM17b]. solute [Kup14]. Solution [EG10, GVSP12, HA18b, RMA10, AB10a, AM13a, AJRWS12, AaC19, AIIZ10, AJ12, AS11b, Alo11, AKL18b, BMAR18, BS16a, BGRS11, BD16, BDF16, Bis14, Bis10, BOT14, BT14, BHJ14, BL17, BN14b, BK10, BPF13, CCJ10, CCJV11, CM11a, CNR10, CP15b, CB11d, CLF10, CYZZ18, CHL18, CJK17, DWY15, DLZ17, DD19a, DAM14, DNP15, DFGG13, DM12b, EMQ18, EM19, EZRR10, EOM11, ESL11, FR15, FLZ14b, FJC16, GMP18, Gha18, GY11, GALO18, Goo11b, GR13b, GW15b, HZ18, HG16, HVR18, HHNLGC18, HB12b, HSZ15, HT12a, HQ19, HXL11, HLZD11, HGW11, HXS^{+}15, HM15, HM17c, HT19, IBSS11, JKN10, JFC14, KB19, Kat11, KV10, KP19a, KPR13, KBCS16, KCC^{+}13, KRD16, KAS11b, KY11, KS12b, LS10a, LS11b, LGG12, LCK13, LSZ16, Li18e, LT18, Li19a, LZ11d, LZ14]. solution [LY11d, Liu14, Liu16a, LLT16a, LLT16b, LCW17, LLP19, LWW19a, LSV18, LLL16, LMP13, LXYT11, MX10, MDW11, MWL18, MDRRV11, MGY11, MO14, Mil18, MN10a, Moh14, NT17, Oru17, ORR16, Pan11, PMM17, PAE^{+}12, PJ17, Pop13, Pop14, RKP12, RY11, RNB11, RBB12, Ran15, RCM11, RC17b, RZZ19, SD10b, SD11a, SMK18, Sat11, STS19, SYI12, SSH15, SBS12, SA16, SKPW14, SMBY10, uIAA15, SWC11, SKM11, SK14b, TZ18, TAS11, TNHK19, Tur10, UKA15, VA10, VA11, VDB13, WYN12, WLT13b, WW19a, WAW15, XZC12, XCS18, XLT17, YYK16, YC11, YJ19, YYLW19, YN16, YM13, ZA10, ZTR11, ZWC10, Zha11c, ZS11b, ZLTY16, ZYSY17, ZH18a, ZWLZ18, yZjH12, ZTW^{+}19, Zho19, ZBL12]. Solutions [ArEM10, CA10b, HKT11, Kes10, LZGZ11, LLX11, MLL16, Abb11, ARESH18, ABL15, AK16, Ade16, Ade17, AJS19, ABJ11, AP10, AÇT11, AA15, AAEG17, AJAR18, AJ10, AJ11, AA18, Amb19, AKLS19, AT17, ACE17, Arq18, ASY^{+}11, AET19, AKMUH17, ALI11b, BZ10, BS12a, BK12a, BM10a, BC12, BNR10, BH10, Bao16, BM18a, BR18, BAO^{+}12, BHM12, BK14, BDS10, BB12, BC10, BKY10, BML11, Bog11, BDO11, CHS11, CW15a, CP15a, CWW19, CW19, ÇT12, Cha11c, CANK11, CZN11, CZN12, CC19, CCD10, CTA12, CCX13, CM13b, CF16b, CTC17, CTG17, CT17a, CT18, CM18b, CO19, CHZ19, jC11, CyL11, CZ15, CZ17, CT17b, CS12, CAP10, Chu11a, Chu11c, CP16c, CC17, Chu18, CCP19, CJCV10, Cov13, CZ18, DLS18a, DCL17, DGA18, DGZ13, DQ10, DSR10, DLD10, Din13, DL14, DD16, DLC19, DSL18, Dua18]. solutions [DDK11, EGAA19, EK16, ESN10, ES17, ES18, Elb15, Erg19, EG18, FKKS11, FYYT11, FSTN18, FST19, FW18, FSHZ11, Fer11, FL13a, GZZ^{+}16, Gao17, GTC18, GJ19, GK13, Gen11a, Gep16, GTL16, GSY10, GL17b, Goo10, Goo11a, GK18, GZD18, GMZ15, GTZ19, Gup11, GH13b, Haj18a, HA18a, Haj18b, HY10, HY18, HL18a, HLWX11, HC16, HLTL17, HLLM19, HGHA19, HKI12, HYCP11, HKK^{+}16, HTM18, HL11a, HL11b, HM10, HRMS10, HM17b, HX14, HTWS15, HYC18, Ibr11, IMS19, JKB11, JKK12, Jal14, JRZK11, JPB11, JPK18, JQSS12, JL11b, JWX14, JPS10, JS11, JLTB12, JD12, JZ11, JLF17, JKS18, JKS19, Jum10, KSG11, KLL19, KK14a, Kim11, KL16b, Kon16, KL19, Ku15, KKK14, KKK15, KKK16, KTK17, KK17, KTK18, KT18a, KT18b, KK19b, KK19c, Kuo18, KMT10, KM11, LK15, LSCG16, LS17a, LL19a, LLZ10, LCW10]. solutions [Li10e, LYZ11, LD11b, LF11c, LDL11, LJ11, LS11a, LL11, LZ12a, Li13, LY14, LL16a, Li16a, LMMF17, LX17, LLY18b, Li18c, Li18b, LZL19, LM19a, LCW19, LGZ19, LJ10a, LZ11e, LWL11, LLZ18, LL10a, LCLL16, LJLY19, LKU10, LWKK10, LJ10b, LC10d, LMDL11, LZKU11, LY11c, LC16, LTSW16, LL16d, Liu16c, Liu18a, LW18a, LCLL18, LZM18, LYC^{+}19, LW19b, LZ19c, zLZ19, zLYLQ19, LWW19a, LDL10, LSX13, LL14b, LMZK16, LXY19, Luo19, MM10a, MP10a, MP11a, MPZ11, MZL13, MW17, MF18a, MYZ18, uHS12, MCKM12, Man18, MZM18, MM18a, MGS^{+}14, MPfTX18, Men18, MAST18, MLZ^{+}16, MV17, MS12a, MPLR18, MY10, MI16, MM12, MK17, MAH18, MCN10, Nab19, NKM15, NTR15, NNWAS11, OZF19, Osm18, Ouy11, ÖZ11, Pad18, PDM11, PRR18, Par11a, PC11b, Pen11, PTZ19, PD11, Pet15, Pir11, PT11, PMA17, PP12, PLT17, QRMH18, QHT16, QCT17, QTW^{+}18, RSM17, RHMA18, RSDR11]. solutions [RFK16, Ray16, RG18, RWW18, RMY19, RÖ10, RNQ13, RNQ16, SR17a, SH11, San12, SAIZ15, Sca11, SZ12a, SJL^{+}19, Sea11, Sea15, SIL19, SVY16, Sha18, SMDI18, SNDK18, SNSK19, SW12, SBM13, SW11, SYY13, Sta11, SZ11, Su12, SZW11, SLL12a, SZ12b, SLL12b, TN11, TJ10, TZWM10, TZ15, TCM18, TUT11, TNF11, Tol12, TTX^{+}16, UABK16, ÜM16, Üns18, VLFS12, WSG10, WZH10, WZF16, Wan16b, WS17, WF18, WWD18, WY19a, Wan19b, WO18, Wei10b, WHW11, WFDW10a, WFDW10b, WLDL11a, WLDL11b, Wu11a, WZ15, WW15, Wu16, WZMY18, WX18b, XX10, XZR16, XYXZ16, XC18, XY10, XLD11a, XLD11b, XHA13, XC17, XZ19, YY15, YTD^{+}18, Yan10d, YY10b, Yan11c, Yan11d, Yan12b, YL16, YGS^{+}16, YM17, YGS17, YKC11, YCLY15, Yao16, YÖ10, Ye17a, YMDZ10, YLG10, YL10b, YMLL18, YTC^{+}18, YW19, YMHL18, YqS16, YS16b, YZS18, YTS^{+}17, YTL^{+}18, YHC18, YSS11a]. solutions [YSS11b, Yüz12b, Yüz12c, ZSAN18, ZA15, Zen11, ZSH11, Zha18a, ZFZQ10, ZJ10b, Zha10, ZLZ11, ZZ11b, ZS11b, ZD11, ZZ11a, ZL11, ZTC14, ZLL14, ZZ15b, ZTZ15, ZTZ16b, ZMLZ16, ZCZ17, ZXZ17, ZJ17, bZM17, ZM17a, ZDZY17, ZZLB18, ZLT18, ZLG18, ZDB19, ZSL19, Zha19c, ZL10a, ZDM11, ZCH12, qZM17a, ZTZJ14, ZTSC16, ZLZG11, ZKW15, ZW16b, ZT16b, ZT18a, ZJ10c, ZWW13, Zho16, ZZ18c, ZP18c, ZSD10, ZCY11, ZLW19, ZMA10]. Solvability [Bai11b, Liu12, Ols10, LJX12, Wan14, WF17]. solve [BDM19a, GL10, HHGA19, Li16b, LLW15, MP19c, SD12a, SD11b, Tom11, XHM14, XW19b, YK17]. solved [Kar17, TLR17]. solvent [CM19a, STDLM19]. Solver [GGAVRC^{+}19, AP19a, BKL^{+}19, Cos18, DTR19, DB12, Du12, FIVM17, FIMV18, Fia15, FXCC18, FP18, HMY15, KKL^{+}13, LN19, Mar16, Ngo18, PÁAP^{+}15, PPC13, QKR19, VMC^{+}14, WFL11a, XZ18, YSW16, YLC16, YW14, ZS16]. solvers [ABK^{+}13, BJQS18, BG10a, DGR18, JRB15, JJ19, LMR19, LCN10, LYC15, PPC15, ST19, SST12, THGG14, WKP^{+}14, ZZL19, ZGL14]. Solving [BBBM16, CCJV11, CRRS11, DGB10b, DT16, GDM13, Haj18c, HLY12b, Lep11b, LLW10, LH12a, LLH11a, LLH11b, LGVS19, PT15, QAA^{+}16, SNH10, SNMA12, SBM10, WNTW19, XL15, YKA18, AD15, ADZ19, AD19a, eMA18, AEH19, Abd18a, AIA13, AO10a, AR10a, ATUC15, AJY13, AER12, AT18a, BM11a, BJRF19, BS11b, BDS15, BGPP11, BE11, BG11, BXKZ11, BX14, BP18, Bra10, BMP15, BPF13, CH11a, CY14a, CAY12, CW10b, CDY11, CM11b, CWDL17, CM17, CW18, CM19a, DZW16, DM12a, DL10, DN18a, DCL17, DM15, DA18c, DHY19, DBS12, DLC19, DZ16, DJZ18, DH10c, Eba11, ECJ16, ESBR10, Els10, EE18, EKE18, FJB19, FT10, FPB17, FH11, Fia15, FIS18b, GGLP15, Gen10, GR13a, GH15, GHR10, GR19b, GHC^{+}15b, GM19, GM11, GGO16, HXX19, HZL17, HA16a, HSWZ11, HHG14, HSMT19]. solving [HL18b, HM14, HWXC16, HWXC18a, HWXC19, HT16b, IB11, JKK11, JGSS10, JYF^{+}11, JA11, JNBK13, JKS12, JJ13, JJ15, JL19, JL20, JWX^{+}13, JYYL16, KAG11, KKAM11, KAJ11, KL12b, KMRN12, LC11a, LHY11, LLJK10, LYZ11, LH16, LBJ10, LBW11, LHL12b, LLLC14, Liu15a, LWZ16, LW18a, LW18b, LFZ19b, LW13, LDY11, LM18c, MN11a, MTM11, MKR12, MHHC18, MBHV10, MZES12, MHM11, Moh15, MSG11, Mor13, MBJ16, NLA19, NUNAS11, OZF19, OP14, Oru19, eOS18, PC17, Pap15, PIAH10, PLR15, PW10, RKF18, RS15, RWTW19, Ros12, SD10a, SK11a, SYG11, SSSB11, SVY16, SG14, SS17, SWL16, SPST18, TM17, TNP17, TA11, TTZW18, TS16, TC10, TTC14, UKI11, VAB12, WD10, WKG10, WSCL11, WKS13, WWB13, WZ17b, WHS17, WTM17, WHS18, WLZ18b, WH11b, WFL11b, cW11, qXjH11, XY15, XWY18, XH11b, YXS10, YX11b, YC10b]. solving [YNLK10, Yao16, YP10, YT18b, Yus09, Yüz11, Yüz12a, ZB19, ZSAN19, Zha11a, ZY17a, Zha19a, ZYZ11, ZÖXL^{+}19, ZH15b, ZQ14a, ZT18b]. SOMA [SZDO10]. Some [ANP11, AS10a, AX11, AS11c, APTZ19, BCB11, BD11c, ÇT11a, ÇHK11, Cha11b, Che11a, Chu11c, DG10a, DÇ12, EAA10, ET12, FM12a, Gao17, GHT^{+}15, GRS12, Gro19, Haz11, HA11, Jic10, KR11, KK15, KM18, KKBR19, KB10c, KKK15, KTK17, KK19b, KHUO12, LD13a, LLW11, LY10b, Li10d, LCN10, Li11a, LC11b, LSJ12, LG12, LL10b, LSZ11, LZJ12, LS11d, Luc10, MZJ11, MM18a, NNWAS11, NODA11, PKK12, RA11a, RR19, RI12, RGdSRLAJ10, Sön11, SGQ12, TÇ11, TÇ12, TO11, VRD11, Wan10a, WS10a, WAZ11a, WSH12, XW10, Yak11, YL10a, YNLK10, YX11c, YWL17, ZW11b, ZJ12, Abb11, AM12a, AvdW13, Arq18, BCFQ19, BJS15, BL17, BBD10, CJ18a, CJ15, CM10c, DGZ13, DE10, DYX11, Din10, Dra11b, EBENEA10, EBENF10, EGGS^{+}12, Gav12, GFZ16, HH16, JZ13, Jum10, Kes10, KA11, Kir10b, KT18a, LWBW13]. some [LZ11e, LTX10, LFJ11, LSV18, MP10a, MB11, MDVM17, MV11, MB10a, MN10b, Mor10d, MN10c, MN11c, MKPS11, Noo10, NA11, NUH12, Pop11, Pop14, QH11, Rah11b, RRC11, RRGTV10, RGHZ15, SG11b, TN11, Von11, WC10a, WS11b, XC11b, YS17, ZWMD16]. Soret [ABV11]. Soret-induced [ABV11]. sorting [DHQ11, PN16]. Sound [ANN10, GC19a, HTL10, IK12, Par17a, Par17b, TTT10]. sound-hard [Par17a, Par17b]. Source [BRROP19, CZF10, AM13a, ABL15, ANN10, CL19, CV14, CLJ11, CCP19, DS18b, EA10, FL13b, FIM18, GD11, HKJ14, HC16, HH18a, HP13, HMP^{+}15, HHM12, IK12, KVW18, KPK18b, LY11b, LMZ18, Liu17, LWSL19, MA17, PC17, RCRV14, RZZ19, SBM10, SS18b, SS16c, SG16a, TKHL18, WQRZ14, WW14b, YF10, YXX11, YMHL18, ZSQ^{+}18, ZL19a, ZKW15, ZW16a, ZY15c]. sources [CM18a, CH19, FT15, NPR10, PXT10, SRG16, Wu16, Yu11b, ZZ10b]. Space [BSY^{+}19, EEBM10, FID14, KV17a, KV17b, ZFLM19, AJS19, AES11, AZ15, AK19, Arq18, Bac14a, BL18, BTEM19, BBR10a, BP13, BZ18, Bis14, BGH14, BGM19b, BMH19, Buo11, CWWY15, CLA19, CyL11, CR19, COR18, CELY18, DDMQ19, DVMS13, DA18d, fDxZ12, DLT12, DZO^{+}19, DH16, DJZ18, EGAA19, FZL^{+}18, FNW18, FOS19, Fur13, GLR13, GGAV18, GLP18, GH14, GM14b, GD12, GM19, GABC16, GF19, Her19, HD14b, HZP18, HLX18, HLY17b, IW18, JW18b, JW19a, JL18, JKS18, Jum10, KK13a, KBAF18, KS15a, LMR19, Li12b, LLFT17, LGZ19, LK18, LN19, LHL12b, LZG13, LWZ16, LFAL19, Lóp19, LDG19, MBS17, MP11c, MS15, MY13, Nab19, NHH13, NH15, NCC13, Ols10, Pad18, PS16, Pov12b, RAZ19, RMS10, RZ16, SD11a, SZA^{+}18, Sha12b, SGK18, SOK19, SND19, SI17, TD10a, THD19, Tru19, TT12, UMY11]. space [WQRZ14, XA13, XLK11, XSYL19, XaZH19, YZ15, Ye19, YDW15, YSX^{+}19, Zha11d, ZLJ^{+}18, ZCLW19, ZZL19, ZLL17, ZLZ10, ZY17d]. space-dependent [HLY17b, WQRZ14]. space-fractional [AK19, BMH19, FZL^{+}18, HLX18, IW18, JW18b, KBAF18, LK18, MY13, SI17, YSX^{+}19, ZLZ10]. space-momentum [KS15a]. Space-time [BSY^{+}19, DH16, FNW18, GABC16, LWZ16, Sha12b, YDW15, ZLL17]. space-time-fractional [Pov12b]. spacecraft [DLWW12, KW12, LLZ12]. spaces [Abb10, AKT12, AE12a, ABRL18, AD10a, AD11a, AIB10, AR10d, AD11c, AHF10, AH11b, AvdW13, AKS11, ASV11, APS12, BKT11, BHM12, Ben17, BM12b, CsH10, CB11b, CDG15, CDG16, CGY11, ÇA10a, CANK11, CZN11, CR19, CSW11b, CS11d, CS11b, CS11c, CM10b, CSN11, ĆSCD11, Deb12, DÇ12, Dra10, Dra11a, Dra11b, DGH17, FL14, FLDZ12, FRZ15, FBB10, Gao12, Ge10, GKK11, GLZ18, HLWX11, HLS11, Haz11, HCF16, HS11b, HA11, IN10, KPR10, KR11, Kar10a, KK10a, KK11b, KC11, KKS10, KA10b, Kha10, KYR11a, KYR11b, KA11, KN12, KB10c, KT11b, KLMV12, Lad16, Lee11a, LS10b, LjHO10, Li17b, LC11b, LZ12b, LLX11, MP10b, Mai10, Min11, MN10c, MME10, MN11c, MKPS11, MAK12, NS11, Nil11, PRR10, PT11, PCM12, QL10, Qiu12, QY17, QaY18, RK10, RA11a, RR19]. spaces [RI12, RA19, RGdSRLAJ10, SP10, SCV10, SK11a, ŞGY11, SLK12, Sal10, SN11, Sha10, SS11c, SH10, SCK11, SW11, Sön11, SCC12b, Via15, WD10, YBC11, Yao10, Yil10, ZLL11, Zha18e, Zha18f, vdW14]. spacewise [RAD13]. spacewise-coefficients [RAD13]. spacing [FES^{+}19, GGGR17]. Spalart [PLR15]. spam [CDFP12]. span [Nab19]. spanning [MW11, RR19]. SParC [ADD^{+}15]. spark [JS12b]. Sparse [BF16, CR19, ASA16, ADD^{+}15, CX16, DGLU18, Git14, HLX18, KM14, KSS13, LCLL16, LSV18, RZZ19, SW16b, TJQS13, THGG14, WY19b, ZS16, ZFY^{+}19, ZZL18a, ZXW13, ZGL14]. sparse-view [CX16]. sparsity [LZ11c, PSS18]. Spatial [FL13b, GV11a, Hu19, Zha18b, AB16, Ara18, CLTA11, CM18c, FSM19, JHW15, LY11b, WYL19, YZMA18, ZH15a, ZSLZ19]. Spatial-temporal [Hu19]. spatially [CL19, CRG16, Che18, GWR^{+}18, GMS18, GLL14, LHL14a, WCCS15, WLZ18b]. spatially-distributed [CL19]. spatially-varying [CRG16]. Spatio [ZZF18, ZLMZ18, AET19, VJM15]. Spatio-temporal [ZZF18, ZLMZ18, AET19, VJM15]. Spatiotemporal [MLY18, SZ14, Zho13, GM14c, YG17]. SPDE [MT19b]. speaker [Hon12]. Special [CBM10, LZL^{+}13, PHWM10, SPCS13, VBW10a, VBW10b, YWL^{+}11b, ZMLZ16, ZFLM19, ADK10, DG13a, Din10, Gal11b, Kir10a, Kir10b, LZGZ11, Lup11, MTN19, MJ10, QY13, RRGTV10, SKST10, TN11, ZWLZ18, Zho16]. Species [YZAX10, BFF^{+}11, Che16, CKRW19, DD19a, FHA16, Gal12, KL16a, Li16a, LMZ18, MW17, PP10, PQB^{+}16, PQBK17, TM18, TZG10, VPR11, ZT19, ZL10a]. Specific [Bra13, Ebr11]. specifications [CFRS10]. Specifying [ZMG10]. Spectral [AS15a, CLC16, CSS10, KDU15, LvdVX18, PGQ16, UMLF13, VLJH18, YCHW18, YQWZ19, ZLHF19, ADZ19, AR10a, AB10b, APRM11, AK19, AT18b, BG14, Boy16, DNS15, DB10, DBEE11, DB12, DHY19, DNS18, DZ16, DHMU16, FID14, GHC15a, GB16, GY11, GSY10, GML17b, HG16, HA16a, HS13, HSJ15, Hes18, HC14, HZP18, HK17, Ipe12, JL18, KB10a, LL14a, Lee17, LCA^{+}17, LK18, LL12e, LZ15b, LFZ19b, PFDG17, PGF18, SD15a, SD18, SBEB10, SD11b, SL18b, SK14b, Tha19b, TXL19, THGG14, ZYS10, ZLJ^{+}18, ZCLW19, ZcHS18, ZJZ18, ZY11, ZZX16, ZC17, ZT18b]. Spectral-element [UMLF13, ZC17]. spectral-Galerkin [AB10b, DB12, DHY19]. Spectrally [AHF16, SMF17, BL14]. Spectrally-accurate [SMF17]. Spectrum [BCF10, AES11, CCL^{+}12, FBB10, LSM10, LL10b]. speech [YZ12, ZJZ^{+}11]. speed [Ebr11, LMP13, LC12b, QCS^{+}19, RR14]. SPH [FL13a, GHT^{+}15, HMF16, LVF^{+}16, TLR17, VLJH18]. sphere [AA18, HR14, LRCG16, Mic17, PT15, RSP18, SKFG11]. spheres [GT15, GT16, PL17]. spherical [BJLZ12, BOY12, CDM10, GVSP12, GH16, JNJ^{+}11, PL17]. spheroid [LDHH13]. spheroids [MKS13]. Spike [DHQ11]. spill [WHD14]. Spin [AHP^{+}14, GTZ19, LM19a]. Spin-polarized [AHP^{+}14]. spiral [HSC17]. Spline [Gha17, Cer18, DSB19, GW15a, GMI11, GMI12, JK18, LXL14, MP10b, MD15, MM18d, Moh15, Pet15, PLT^{+}19, QX19, SHH16, SKM11, SYW11, WZH18, YZX18]. Spline-based [Gha17]. splines [CQRW11, HH15, KB19, LD13b, LT15a, MRS15, PZJ^{+}16, PZAR19, PL17, Seg19, ST16, Sou11, TWLYÖ10, VAK^{+}19]. Split [WH16, CAY12, CX16, DMZ10, JZJ18, WH18]. Split-step [WH16, DMZ10, WH18]. Splitting [AO18, KVV14, SLL17, AM10a, BCF13, BKR^{+}19, BKNR19, CM16a, CMR17, Cao19, CR18a, CWY19, Del13, EO14, FZ19, Gal12, GS15a, GTG11, GR19a, GSD^{+}19, HMF16, HJ13, HWXC18b, KM14, LM17, LWQZ18, MKHC11, MPY16, Oan13, PW10, RWTW19, Saj12, SR17b, SS16b, SS14b, SLM16, WLD13, WWB13, WLM13, WHW11, ZT13, ZT16a, ZYW15, ZS16, ZZ18a, ZYW17, ZQ14b]. splitting-based [CR18a]. splitting-differentiation [GS15a]. splittings [Lad16, LFZ19a, Mis14, MM18c]. SPN [LLSS13]. spotty [ÖŞ11]. spreading [GGK18]. spring [GBG11, KJ11]. Spurious [GOT19, ACE17, BCD^{+}16, MNPD15]. SQP [ZLL12]. SQP-filter [ZLL12]. squall [GZW^{+}18]. Square [CZN11, CCJ10, CCJV11, CCKY12, CJCV10, FGL10, HSK11, RMB^{+}14, SRRP18, TNP17, XH11a, ZHB11]. Square-mean [CZN11]. square-wave-driven [CCKY12]. Squares [BDGG14, GCDG17, KDU15, RCM11, BQ15, BQ17, BJQS18, BJ19, BG14, BC17, CCKP15, DM16, Dia17, GB16, HA18a, HS13, Hes14, HSJ15, Hes18, HM17c, HK17, Ku18, LC10a, LR14, LGG12, LL16b, LZ17, LJJ17, LJLY19, LY11d, Liu14, MLGY16, PFDG17, PGF18, PATA11, TCM18, TJQS13, TBP19, WLT13b, YD12, ZLZ11, ZMLZ16, ZY17a, ZWLZ18, ZL13]. squeeze [CJ12]. squeezing [IIHI10, IIHuI10, MMA12]. squirrel [DGGBTRJF12]. Srivastava [KSJ12]. SST [LZPZ19]. ST [SWS19]. stabilised [LMR19]. Stability [AR09, AR10b, AKSW19, BDB12, CW15a, Chu10, Das15, DD19a, FF14, FZ18, GKK11, Hon10, HTV13, HW19b, JC12, KC11, LCP10, Li17a, MG16, MMFT^{+}19, RMM11, Saj12, Sea14, SSC19, SYW11, WYLZ10, WZ11a, XXH18, YG17, YH19, ZH19, ZLC11b, ZZW15, ZZ17, AM15, AEDL14, Bai11a, BV11, BG19, BTB18, Bis14, BY11, BK13, BK12b, CBKR10, Cie11, CHS19, CFF15, DMZ10, DYQM14, FDB13, FLZ14a, GV11a, GL16, HY13, HM17a, Hua11, HT16b, JZ12, JZL18, KP13, KG14, KLP17, KJA10, KPG18, KGM11, LC12a, LPK15, LHW11, LW15, LWD15, LL16b, LCC13, Luo18, LPY16, MAB19, MSA12, MTV13, MW14, MZ11, Ngo18, Pal12, PCM12, RSV11, Rhe10, SP10, SCV10, SMF10, She16, SI10, Sun10, VB10b, WWG10, WW11a, WZF12a, WLL12, WCW13, XMW10, XH11a, YÇG12, YY10a, YX16, YS19a, YW11a]. stability [YLLN16, ZHW14, ZMWH18, yZjH12, ZGD14, dPLM18]. Stabilization [HH18a, LCQL17, Lin12, BL12, BHH16, CKSL^{+}14, CBBE16, CW17, KEHB18, LMZ18, OSZP13, SZDO10, ZTR11, ZT19]. stabilizations [DM18]. Stabilized [MVKK14, MPY16, TS16, BPS19, BS16b, CBS18, FKF13, FLdS14, FGHZ17, GLP18, GGH18, HZM11, LLML15, LHL18, MFSL19, RC18, WC15, YZ15, ZHJ14]. stabilizing [HK10, Nes10]. Stable [BMH19, Dub13, FXC18, FLP13, SZGG11, AvB16, Ahn12, AKMS18, BR12d, DGTC13, Dun18, FHZ10, GS15b, HP17, HA16b, HKW15, KKT13, LS19a, LLJK10, LLYL19, MM18d, NCC13, RFK16, RKF18, SS14a, SCC^{+}12a, STDLM19, Wei12b, ZLC^{+}14]. stage [Amo15, BP13, CCL^{+}12, CHT11, HHY^{+}11, JL12, LL12c, MLY11, MM18c, SFM15, Xu14, YG17, YD12, Zen11]. stage-based [HHY^{+}11]. stage-structured [BP13, Zen11]. Staggered [KNT12, CDL17, ED12, KG14, KCL14a, KCL14b, KCL16, RS18b, SWL16, ZP18a]. stagnation [BGP13, Das12]. stagnation-point [BGP13]. stairs [KHIB12]. Stancu [Büy10, Gal11a, Mah10b]. Stand [RIW12]. standard [MMH11, MHM11, RMM11, SMM19, TLR17, XGH17b]. standing [DL14, FZ18, Luo18]. star [CNR10]. star-like-shaped [CNR10]. starlike [RS12a, Sok11, SWW11, YL10a]. Starlikeness [Mos10, SK11d]. Start [YZ10a]. Start-up [YZ10a]. State [BKR11, TMSO12, ZHV19, AdAS11, CC19, CTC17, CT17a, CMS10, DNS16, DLT12, FGHZ17, Gao15, GTC18, GRS18, Gro19, GD12, GS19, HP19a, HC14, Kia19, KLTS11, KAK^{+}12, Kup10, Kup11, Lan12, LSCG16, LLZ11, LL13, LWN15, LT18, Li18b, LCYC12, LLT16a, LLT16b, LLP19, Luo19, MCL^{+}13, MN11b, Pen11, PB11, QHT16, RMS10, RFP11, RF12, TMO13, VCM11, VV14, WZ15, XLT17, YY11a, YW19, YKA18, YTZ17, ZHJ14, ZTZ15, ZSQ^{+}18, ZLMZ18, ZT18a, ZZX16, Zho19, ZBL12, dAS18, dSAC11]. state-dependent [AdAS11, DNS16, Pen11, dAS18, dSAC11]. state-space [DLT12]. States [HPV^{+}18, AF13, CL16a, Das15, GCR^{+}18, JLWX18, MCL15, PTL13, Wu18a, ZTZ16a, ZHS^{+}19, ZMWH18]. Static [ZHQG12, CCCW16, JAJ18, KLTS11, LT13, LX18, MMS^{+}18, NHH13, PÁAP^{+}15, TAS11]. station [Mok11]. Stationary [LD19, WL15, CW15a, CGGM19, Cho17, CN16a, CGO19, Ers16, FIW13, FMSV17, HKP17, HZM11, JFC14, KL19, LFZ19a, Liu16b, NCC13, NJV13, SK14b, WH14, Yan18a, YH19, ZHY14, ZT15]. Statistical [AD10b, BA11, DK12, KD11, MHH11, CCJV11, Che11e, DD10, Fio14, JL16, KD10, LZF18, VDV13, WZG19, yXpYxZT11]. statistics [BEAA11]. stator [LC12b]. Steady [DKM17, LWN15, SL12, ZZC13a, BMS13, CL16a, Das15, GCR^{+}18, HP19a, HLSN15, Ima17, JLWX18, Mom11, PD11, TH19, YY11a, Yan18c, YT13, YTZ17, ZHJ14, ZSQ^{+}18, ZLMZ18, ZMWH18]. Steady-state [LWN15, YY11a, YTZ17, ZHJ14, ZSQ^{+}18]. steady-states [JLWX18]. steel [CPP10]. steels [SSIP19]. steep [YL16, YW19]. Stefan [HSZ15, UABK16]. Steffensen [SV11]. Steffensen-type [SV11]. Stein [XW19b]. Steklov [MRR18, RA11b]. stem [Ben12]. Stencil [CCHG17]. Stencil-based [CCHG17]. stencils [FLP13]. stenosed [SR10b]. stenosis [SR10b, SRV10, ZAK18b]. stented [WSC16]. Step [MBJ16, ATUC15, ABCR10, AM12c, CYM13, DN18b, DZW^{+}15, DMZ10, Don10b, FLH10, FL11a, GK11a, GK11b, GGT14, HWXC18a, KRCJ11, KYR11a, KYR11b, KSMN11, LCK17, Mai16, MR19, MS15, Mor13, OBAAD10, SM19, WH16, WH18, XWY17]. Stephenson [HT12b]. Stephenson-I [HT12b]. stepped [LPML19]. stepping [AK19, SZ17, SZC^{+}18, YH19]. steps [WZHW13]. Stepwise [GGAVRC^{+}19]. sterile [ADL12, DD13]. Stewartson [HLTL17, KSG11, QRMH18, THfL17, WLXZ18]. Stieltjes [BR12a, LS12c, SK11c]. stiff [ABL15, AJY13, KAG11, SSSB11, ZR18b]. stiffened [MG15, PBS12]. stiffness [BCG17, MM13, PH13]. stimulus [PLKCC12]. Stirling [Mor10a]. Stochastic [BNTT14, BCPS15, BM19, GHL18, JHW15, KPK18b, KKK12, LHW11, Lin11, LDG19, MP11c, ZZG19, aZW17, ASSV18, BM12a, BS10b, BH10, BH14, BY11, BR13b, BKP11, BK12b, CFN11, CZN11, CLS19, CXZ15a, CKM12, CCCW10, DRK11, DWZ16, DMZ10, DGLU18, EBENF10, EHL^{+}14, FES17, Gao17, Git14, HDS11, HZ18, HP19b, HA18b, JS11, KPR13, KMRN12, KSM12, LD11a, LW17, Li18c, LQMW18, LX12b, LH10a, LSD10, LLH11a, LLH11b, LW12a, LW12b, LW11d, MKR12, MV12, MP19a, MS12a, MB10b, Pal12, PDM11, PZL^{+}18, PW18, PH19, Pul16, SSA12, SYZ19, SP12, VB10b, WS12, WL17a, WZ18a, WLXG18, WW19c, Wan19b, WWH12, XZ10, XH11a, XSYL19, XaZH19, YL18b, ZZ19, ZLW18, ZWMD16, Zha18h, ZC11a, ZC11c, aZ18]. Stock [CXZ15a, FXC18, LWJ10, DRK10, FES17, LX10b, San11, Sin16, WTC^{+}12]. stock-dependent [DRK10]. Stokes [AT17, GOS18, AS19, BT15, BS16b, BO18a, Bis14, BNP18, BC17, BPF13, CLTA11, Cho17, CH17, CB19b, DDMQ19, DM19a, DSM18, DNZ^{+}13, DZ17, DLQ16, Dua18, Ers16, FZ17, FH11, Fis18a, FRZ15, GGLP15, GRBT16, GB16, GH18, HKP17, HKS19a, HS13, Hes14, HSJ15, HQ19, HZM11, ILS13, JRZK11, JPS14, JJH16, JYYL16, KLP10, Kim14, KCL14b, KLRW12, LC10a, LL12b, LR14, LHL18, LHL12b, Liu16c, MM16, MLGY16, NMR15, Ngo18, PGQ16, PLKC16, PGW19, QZM17b, QM19, RBTD14, RTV17, RA19, SMM19, SYI12, Sha14, SZ17, SHH16, SR18, TDM13, TW18, WC15, WPL16, Wan16a, Wan18, WCSW18, WMP^{+}19, WH14, XSYL19, Yan18b, Yan18c, YJ19, YT13, YTZ17, YT18b, Zak18a, ZZ16b, ZL18, Zha18g, ZWY19, ZL19b, ZP18a, ZP18b, ZP18c]. Stokes- [BNP18]. Stokes/Darcy [HQ19, JJH16]. stopping [DVY14]. storage [CWHW17, JFS14, JFS20, KPS10a, KPS10b, SH18]. STR [BRROP19]. strain [KHF^{+}19, NHH13, NH15, RC18, TPHD18, ZHW^{+}18]. strains [ED12, HAESLB14, SCvdV^{+}19, TCM15, YWH14]. Strang [Del13, EO14]. strategies [Ala10, BDPM12, Boy10, BK16, CCHG17, CM13a, FBL11, FMP19, Ögü13, PM13, SJS^{+}11]. strategy [AKL18b, BDHR18, CCH^{+}12, CSSW12, CL17b, DPBL16, DWZ13, DESV18, DDM^{+}18, GYTD12, HD14a, HCL12b, JS12b, LDS10, LTX^{+}13, LW11c, Raf12, RTB14, RC17a, SLXC11, Tan18, Tod13, VC12]. stratification [GIMZ14]. Stratified [Abb10, GZW^{+}18, Liu18b, Sea15]. stream [LWC13, OVV^{+}16, QGGL13, YTZ17]. streamfunction [FIW13, YT13, YT18b]. streamfunction-velocity [YT13, YT18b]. streams [FKC12, MHH11, WCZ13, ZCSG13]. Strength [SLMZ12, SJS^{+}10, ZFC11]. strengthened [LKL^{+}15, Pul16]. Stress [LTT13, RDE^{+}17, BKR^{+}19, BKNR19, CKL18, DB15, Dos18, FZBF10, GGVRB19, HWY14, JPCY13, KL16b, KKVS19, LS16, MSV18, MMT18, Pov19, VGC^{+}15, Yoo17]. stress-assisted [GGVRB19]. stress-based [DB15, JPCY13, Yoo17]. stress-pressure [MMT18]. stresses [DVMS13, DD19b, ÖKJR19, Pov12b, Pov19, SNEP19]. stretched [DDK11, Gen11b]. stretching [BHZJ19, Bog11, CYZZ18, EA10, KWFY11, MMA12, Pal13, RY10, RRP16, Sah11, VPR11]. stretching/shrinking [RRP16]. strict [CsH10, CS11c, GDZ11, KS10a]. strictly [LZ11d, sHC11]. stride [NHIN16]. strike [LTT13]. strike-slip [LTT13]. string [CFLM19, TS11b]. strip [AB18]. stripe [ZCT18]. Strong [CsH10, CB11b, CS11c, KS10a, KK10a, KK11b, KA11, Lee11a, LjHO10, Liu17, Pir11, She11, Wu11a, ZLL11, BTB18, BD11c, CHM^{+}10, CS10c, EZRR10, FST19, GN12, GDZ11, Hua10b, KKLJ11, Kup14, LLL11, Liu16c, MT19b, NCL13, Sun11, YZS18, YM13, ZZ15b]. stronger [WHS12]. strongly [BGPP11, CGY10a, FT10, KRD16, LG10, LXZ18, LFJ11, Liu11a, Nil11, PT15, SOS11, SCC12b, WL17a, WZ18a, YL10a, YASK10, ZZ10b]. Structural [FLZ14a, CFdM^{+}18, FL14, FMP19, KSD^{+}19, Tom11, XC16, YLC18]. Structure [SAU11, AvdW13, BK10, COR18, HB19, HZ10, HM15, JSEM13, JJ13, JL17b, KEHB18, LX16, LMS13, LLD10, LZGZ11, LXL12, LLSS13, LYY12, MJB18, Mac12a, MDVM17, MZM19, MUB^{+}16, Rhe10, RHC15, RZZ19, SLK12, SCSF19, SFM15, TK11, WLYX13, Xu14, XDL12, YG17, YLF19, YZGW10, YLJ12, YLL^{+}14, Yoo17, Yu11a, ZPWY12]. structure-adjustable [LLD10]. structure-preserving [MJB18, MDVM17]. structure-preserving-doubling [HM15]. Structured [BK13, ADY12, AF13, AVV18, Aki17, ALMLM14, BP13, BX14, BFS15, CWHW17, CN16b, EHO^{+}12, GMS18, GGGR13, GGO16, HG16, HCL12b, Hu19, JK10, JZ12, LXL14, LJJ17, LJLY19, LCC13, SRGL13, ZK16, Zen11]. structures [ASN11, BI12, BK12b, DCN^{+}18, FMGR19, GHT^{+}15, JKK10, JLK11, LZ18a, LS16, LGC^{+}17, LKL^{+}15, MJWD19, WLXG18, YC10a, YLS12]. students [WNC12, WTC^{+}12]. studies [AM12a, BV10, DGOZ13, PTP14, PATA11, SL16a]. Study [ALHZMC^{+}19, BM13a, CZL17, Far11, KRCJ11, KM13, WAG^{+}14, AB10a, AÖ10b, AW11, AY18, AM14a, Amo18, ABSV18, AGK15, BM18a, BK18, BM18d, BNP18, CCR16, CD19, CGK14, CJRR11, CS13, DA16, DNR13, EAMA19, FdOP17, FMGR19, GGM^{+}13, GSZ11, HB19, HZLM10, JRB15, JJ19, JAJ18, KRP12, KAK^{+}12, KSMN11, LBZL11, LJK^{+}19, LNP^{+}12, LLG10, LG13, LTJ^{+}16, MC11, MB17, Odi10, OTiSY16, Par17b, RS12b, SSS16, SPT17, SDM10, ST16, uIAH10, TMDTTC16, WSC16, WZKY12, XWN11, XX19, YB13, YWW^{+}12, YCS19, ZMG10, ZCSG13, ZZX^{+}14b, ZYZ^{+}17, ZHS^{+}19, ZD12]. Studying [HHM12, LX12b, ZGD14, BBR10b, RS13]. Sturm [jASzZ12, CKLL10, KA13, YY10b]. style [HCHH12, LZL16]. sub [AM13a, ADZ19, BMS19, CBBE16, GH14, ML19b, Nes10, PS16, SR15, WW16, WW18b, WY18a]. sub-diffusion [AM13a, ADZ19, GH14, PS16, WW16, WW18b]. sub-equation [SR15]. sub-grid [Nes10]. sub-mesh [CBBE16]. sub-parametric [WY18a]. sub-quintic [ML19b]. sub-scale [BMS19]. subbetweennesses [RdSSS11]. subclass [AS10a, MG11, SRM11b, XL10]. Subclasses [ORD11, AX11, Las10, Noo10, ODR10, SK11b]. subcritical [SKFG11, ZT18a]. subdifferential [LMZ17]. Subdiffusion [Iom18, AD15, LZC11, LZLL18, Luk11, Yil19]. subdivided [YA11]. Subdivision [KLCD16, PXXZ16, SR10a, WZXL11, ZB19]. subdomains [DCN^{+}18]. subgradient [ZM18]. Subgraph [Bog10]. subgraphs [MW11]. Subgrid [ZHJ14, ILS13, MS17, Sag10]. subgroups [hYxLL10]. subject [Arq18, DP15, FZBF10, KYA15, LGG12, ZLZ11]. subjected [KLTS11]. sublattices [jHlXZ11]. subliminal [LWZG10]. sublinear [Bao16, KM11, YL16]. submatrix [HA18a, TCM18]. submerged [LFC16, VHPVNXW18]. subnear [ÖI12]. subnear-rings [ÖI12]. Subordination [AKS10, AS10b, XL11, AS11c, BB18, KSJ12]. subordinators [AKMS18]. subsequential [TÇ11]. subsimplices [KK14b]. subsonic [FLZ14a, MS18]. subspace [ACC18, BJRF19, BS11b, BJS15, BOT14, CL16b, DGR18, DNS15, GK13, LSW16, Liu14, Liu15a, Par17b, SL18b, WmN13]. subspaces [GH13a]. substitution [HDS11]. substrates [DVM12]. substructures [AS11d]. substructuring [LP12, OTiSY16]. subsurface [MNJ^{+}13, RRAK19]. success [MP11b, MP16]. successful [AM11]. succession [HHS^{+}10]. successive [Bic11, DH11a, HWXC19]. suction [SRRP18]. suction/injection [SRRP18]. Sufficiency [Ant14, ZLLF12]. Sufficient [AEDL14, KC10, SOS11, Swa10, AD12, CAH11, LC10b, SRM11b, TÇ11, YL10a, ZFZ10b]. Sugeno [Ahn12]. suitable [Bra13, Mok11]. sulfate [DEFP11]. sum [AAR11, WSS10, WS11b, WZ10]. sum-of-actual-processing-time-based [WSS10]. Summability [Ehr18, Bor10, ÇT11a, Çan11c, EÇ10, ET12, Sar11, Sav10, TÇ12]. Summation [KAS11a, Rey12]. summations [Xue13]. sums [CS10b, Cvi10, HP10, MS11b, Sun11, YLS12, YA11, pZ10]. Sumudu [KK11a]. sup [Kim14]. super [CB19a, CT17a, Eba11, IW18, LGHR16, Zha14]. super-cubic [CT17a]. super-diffusion [IW18]. super-diffusive [CB19a]. super-future [Eba11]. super-large [Zha14]. super-resolution [LGHR16]. supercloseness [LZ18c]. supercomputers [ODAZ15, ZGL14]. superconductivity [FSZ18]. Superconvergence [Bac14a, CLH13, Füh18, HZM11, JMHF13, Ku18, LNW19, NSYY13, SW16a, WHTZ16, ZS13, ZY19, BQS16, HLY17a, SW17, SY18, ZZL^{+}18b]. superconvergent [Bac14b, DLT10]. supercritical [Wan15]. superdiffusion [LZC11]. superhydrophobic [CLL11]. superlinear [Bag17, CT18, HEP10, LS11a, LL16a, Wu18a]. superordination [AS10b, AS11c, KSJ12]. superordinations [Lup11]. superposition [MF11, ÜM16, ZM17a, ZM17b]. superpositions [PL10b]. supervised [MWWL11, SLYY13, WLYX13]. SUPG [NNK13]. supplementary [JPP12]. supplier [CW10a, SSS^{+}11a]. suppliers [Che12a]. supply [CSSW12, Chu11b, Chu11c, Chu12b, KGJ11, ZSY14]. Support [GAVOF11, LW12c, CLW11, JWX^{+}13, TJQS13, XKH10]. supported [NJV13, ZWL11]. supporting [DA12]. suppression [CRXL15, Dan12, FDB13, LWWY12, LYY12]. sure [HP10, KPR13, LFJ11]. Surface [EKE18, Lin14, AM15, AA18, AML^{+}14, ABR^{+}14, AM17, CD12, CCY18, CDM10, EA10, GB18, GZN19, ID10, IS12, JK10, JK11a, JGK13, JVMF19, KVW18, Kia19, LHZ^{+}11, Pal13, PGDL18, RRAK19, RB19, SH18, ST18, SG16a, SK19, TTT10, VPR11, WL13b, WY18b, XFY18, ZB19]. Surface-volume [EKE18]. Surfaces [KLCD16, AM10b, AD11d, AM14b, BN16, CLL11, DSL11, FMSV17, HH18b, HT12a, HWW13, JW05, LXL14, MT19a, SLKK19, UMY11, WZ16, XSLS11, XFH19]. surgery [LWHL12]. surgical [FBL11]. Surpassing [WL11c]. surrogate [PKD19]. surround [GW12a]. surveillance [LYM12]. survey [AD10a, AK10, AD11a, DS18a, KY11, McN12, MP12]. survivability [ABDKD12, PHPK12]. survival [GCG12]. suspended [THC^{+}18, WCQ^{+}19]. suspension [MX15, SRV10, VGK^{+}16]. suspensions [WLA18]. sustainable [BDPM12]. Sustaining [DGTC13]. SVEIR [DYQM14]. SVM [ALLH11, ALLQ13]. swaps [HC18, HZ18]. swarm [BBO10, DH10c, FLWJ11, HSZ15, JKK11, LCW12, Laz10, LTX^{+}13, TTC14, XLZW11]. swarming [CS11a]. swaying [SH18]. sweep [KZ16, RWTW19]. sweeping [ALI11b]. swell [WC11a]. Swift [KKAM11, Lee17, PP14]. swirl [AdSSS19, Bis14, ZaY17, Zha18g]. swirling [AD19b]. switched [LCYC12, MK18, MSA12, XH11a]. switching [CY14b, CEJV16, DGTC13, DNS16, ECJ16, HC18, HTV13, Lee14, LJJ11, LS11c, YKA18, ZM16b, ZBL12]. Sylvester [WLDL11a, AEH19, BR18, BK14, CM17, Haj18a, HSMT19, HM17b, HM14, HM17c, HM18b, KM14, LHY18, LM18c, TM17, WLM13, WFDW10a, WFDW10b, WLDL11b, XHM14, Zha15a, ZY17a, ZY17b]. Sylvester-conjugate [WLDL11a, BR18, HM17b, HM17c, HM18b, TM17, WFDW10a, WFDW10b, WLDL11b]. Sylvester-transpose [CM17, HM14, XHM14]. Symbolic [Ade17, EMR10, JL15b, JJ15, KOPS13, LSX13, LMZK16, MWWL11, Pan11, XX10, Zha18i]. Symmetric [ÇT12, GY15, HY10, YTF10, AH10b, ABR10, AEDL14, Bac14c, BKL^{+}19, CM16b, CW18, DH11a, DH10b, Dra10, DIS19, EO14, FHS18, Gem16, GKS10, GJ10, Haj18a, Haj18b, HMY15, HKK^{+}16, HVR18, HDT11, HMWZ16, HWXC18a, HWXC19, KL12b, KSS13, LM19b, LWQZ18, MHL11, Noo10, Pap15, SS18a, TL10b, TM12, WmN13, WHW11, WV15, XY17, XWY17, XWY18, XC16, ZCW15, ZM16a, ZSH11, dCM12]. symmetrical [TLR17]. symmetries [zLYmL18, LXY19, XC17]. Symmetry [WY16, Zed10, Zha17b, Abd18b, Ade17, HLTL17, Kim10, KT18b, KK19b, KK19c, MDRRV11, MDL18, Ray17, RRP16, SR17a, Sin16, SSK13, WF17, WLXZ18]. symmetry-preserving [MDRRV11]. symplectic [DGT18, Dos12, HXS^{+}15, SB19, ZC16, Zha18d]. SYN [JS12a]. Synchronization [AAA12, LZC13, xYsHjL11, AsNAd10, BR12b, FDXW11, HPR19, HGN^{+}10, JW11, Kun12, Kuo11, MMMG12, MRS^{+}12, PC12, WL10, ZFC11]. synchronization/desynchronization [HPR19]. synchronizations [ZGL14]. synchronous [HLCY12, PBK19, ZL14b]. synergetic [LH12b]. synovial [BKR10]. syntax [MBH16]. synthesis [BDM19a, OSZP13]. synthesized [LLZ12]. synthetically [DSWB18]. System [GGAVRC^{+}19, Lin10a, OC10, PFBL10, SBEB10, Ala10, ABCR10, AKL18a, AZ10, ABN18, ACTB19, Bae10, BJLZ12, BS16a, Bao16, BMJ10, BP11b, BCC14, BG19, BTB18, BF11, BD11d, BG11, Bra13, BLyS18, BHH16, BPF13, CGY10b, CR18a, CZY13, CN13, CJ12, CB11d, CP16a, CT11b, CSW11a, CM13b, CM16b, CF16b, CTC17, CHZ19, CKLL10, CTD10, Chu11c, CFB11, CSU13, Cov13, DGB10a, DYH11, DGTC13, DGL12, DVMS13, DSR10, DL14, DLWW12, Du12, DL18, ES17, Elb15, FZ17, FSZ17, FSTN18, FWFL11, FSHZ11, GV18, GW12a, Gao15, GMB12, GM14c, Guo12, GH10, GH13b, HKP17, HC16, HKS19a, HL11b, Hes14, HGN^{+}10, HT12b, HCHH12, HY11, HDHW11, HTWS15, HMZ18, IMD11, JKB11, JNBK13, JMB10, JLL19, KSPP11, KK12, KPK18a, KL19, KPG13, Ku15, Ku18, KKK15, KKK16, KK17, KTK18, KCK19, KPL11, KYY12]. system [Kup14, LS17a, LL19a, LG10, LZY12, Li13, Li14, LY14, Li17a, LY17, LLY18a, Li18a, LHL18, Li19a, LD19, LYY12, Lin12, Liu14, LHL15, LC16, LTSW16, Liu16b, LS17b, LZG19, zLYLQ19, LV11, LW13, LDL10, LLZ12, MCL15, MY16, MG16, Ma18a, Ma18b, MS18, ML19a, MBS17, MM16, MAB19, MH11, MGS^{+}14, MS15, MK17, PC11a, PTL13, PMA17, QL10, QMW18, RMS10, RC17a, RGVR17, SMDI18, SNDK18, SCC^{+}12a, SW11, SSS11b, SLL12b, TZWM10, Tan17, THfL17, TF17, TS11a, TTZW18, TZMZ12, TTC14, VMC^{+}14, Via15, WHS11, WhJxLwW11, WZ11a, Wan12, WWZ12, WL15, WPL16, WZF16, Wan16b, WY16, Wan18, WWD18, Wei12b, WZKY12, WW10c, Wu10, WPH11, WZ11b, WTC^{+}12, XZZ16, XZ17, XCM12, XDL12, YY10a, Yan10d, YW10, Yan11c, xYsHjL11, Yan12b, Yan17, Yan18a, YZX18, YWHC11, YLH12, Ye17a, YGH11, YZAX10, Yu17, YSB15, Yus09, Yüz12c, ZA10, Zed10, ZA15, Zen11]. system [ZM16a, ZZ11b, ZBF11, ZW11b, ZTZ15, ZZ16a, ZTZ16a, ZS16, ZaY17, ZZ18b, ZLG18, Zha18f, ZLMZ18, ZT19, ZWY19, ZSL19, ZFY^{+}19, ZL10a, ZNWG11, ZYZ11, Zha18h, ZW16a, ZMH16, ZMWH18, ZZ10b, Zho16, ZG18b, uRK11]. systems [Abd18a, AIA13, AHV10, ATUC15, ABDKD12, AsNAd10, AEDL14, Alo11, AO18, AAA12, AT18a, AB16, Ara18, ACE17, BM11a, Bac14c, BKT12, BZK12, BR12a, BDM19a, BA11, BDS15, BB12, BD16, BGPP11, BE11, BE12, BKDM13, BWZ16, BDR19, BN14b, BK13, BK12b, CFN11, CS16, CI18, CKW13, CZY13, CW17, CC19, CD10, Che11c, CCH^{+}12, CTA12, Che15a, CTG17, CT17a, CT18, CW18, CM19a, CMS10, CS11a, CAP10, CG14, CFB11, CMT12, Dan12, DH11a, DB11, DYX11, DM10, DGT18, Dos12, FDB13, FL14, FDXW11, FJP18, Fia15, FMSV17, FL10, GD10a, Gal11b, GBG11, GTC18, Gar13, GD11, GKLR11, Gon13, Goo11b, GD12, GHM^{+}14, GHC^{+}15b, GPV11, Gur13, GN11, HK10, HmZ11, HH11, HM17a, HTV13, HXL11, HLZ15, HWXC18a, HWXC19, HWY14, IW18, JKK11, JW11, JKS12, JJ13]. systems [JJ15, JL19, JL20, JCWZ16, JZ12, JC12, JKS19, JKK10, KPS17, KAA19, KST10, Kun12, Kuo11, LX10a, LCW12, Lam13, LZ11a, Lan12, LC12a, LHY11, LCP10, LD11a, LLC13, LX17, LM19b, LZ16c, LCYC12, LLH10, LCT12, Lin12, LSD10, LB11, LS12a, LWQZ18, LCC13, Lü14, LL14b, LRTV10, LZ12c, LHL^{+}14b, MP10a, MP11a, MPZ11, Mac12b, MAPS10, MV10, MZES12, MZ11, MP19a, MM18c, MHM11, MRS^{+}12, MB17, MM12, MN17, MP19d, NHIN16, NN13, Odi10, OBAAD10, Ogi12, Ögü13, Ols10, OPDC12, OSZP13, PC12, Pan11, PS12a, PdlF10, PV12, Pir11, PT11, PCM12, QAA^{+}16, QCYL12, QCT17, RES10, Ram11, RRAK19, RA12, RS13, RR11, SMF10, SYG11, SRM11a, SPS^{+}13, SG14, She12, She18b, SS18a, SD19, SZGG11, SM14, SLL12a, SJHC14, TN11, TJ10, TZZ11, Tan18, TNP17, TZG10, TNF11, TAPA^{+}17, TL12, TÇA12, TMSO12]. systems [THY^{+}10, Vel15, VMO10, VDV13, WSL10, WCD10, WYD10, WW11c, WJ11, WS12, WWB13, WN18, WNTW19, Was13, Wei10a, WFL11a, Wei12a, WAW15, WMW13, Wri13, WX18b, XY15, XY17, XWY17, XWY18, XDH16, XH11a, Xue13, YÇG12, YHZY11, Yan12a, Yan18c, YGR11, YD12, YKA18, YW11a, hYlLL11, YWW^{+}12, YSS11b, Yüz12a, ZZ15a, ZSZ17, ZMG10, ZDL11, ZHJ11, Zha11b, Zha18i, ZJZ^{+}11, ZG19, ZWW13, ZZ18c, ZDV13, ZMA10]. Szabo [SZC^{+}18]. Szasz [VSI12, Mah10a]. Szego [MKS13, ODR10, WLS10].

T [AEDL14, DDLM13, EOM11, MGY11]. T-cell [EOM11]. T-cells [DDLM13, EOM11]. T-hyperbolic [AEDL14]. table [CL12b, MNJ^{+}13]. Tableau [CM14]. tabular [Kup11]. tag [CJP15, SPLHCB14, CJP12]. Taguchi [TL12]. Taguchi-chaos [TL12]. TAIEX [LCCC10]. tailored [HHY13]. taint [ZHQG12]. Taiwan [HHS^{+}10, SLCC12]. Takagi [Ahn12]. taking [CD12, SST19]. Talbot [KBDC12]. talent [HHS^{+}10]. tandem [LA11, UMLF13]. tangent [Ana11a, Cvi11, WhJxLwW11]. tangential [TG14, Tao18]. tanh [Ade17]. tank [WCH13, WFC16]. tanks [DFM15, SH18]. tanlock [EKS10]. Tanner [LG10, Zha18b, ZH15b]. TAPENADE [ZKBE16]. tardiness [LYLX11]. tardy [MMR10, MMR11]. target [HDHL11, WYL19, WPH11]. targeting [SSM^{+}17]. targets [LHM11]. Tarski [JK11b]. Task [WWZ12, DZS10, KX12, OSZP13]. tau [Boy16, SD11a, Zak18a, DE11, ED11a, GY11, VA11]. Tauberian [ÇT11a, ÇHK11, EÇ10, ET12, TÇ12]. tax [ZZ17]. taxis [KL19, Li16a, Li18a, WW19a]. Taylor [Che14, Dra11a, ED11b, FL13a, GR13b, HH18a, HW19a, HLZD11, KLP17, KO13, KSMN11, LK13, LPK15]. TDG [BS14a]. tearing [Hof18]. Technical [WC11b, cFpClC13]. technique [BR12a, BGGCGRSP16, BCD^{+}16, BMP15, CDP16, CJ15, DA18d, DMRS18, DD13, Eba11, FR15, FL14, HLT12, HYS^{+}14, JZR15, JCF19, LX10a, LD11a, LGC^{+}17, LSV18, LDY11, Mil18, MBKK10, fNS11, NGG12, RY11, SD12a, SWS19, SZ12b, TA11, WWZ12, WZ17b, WFL11b, ZXW13, ZD18]. techniques [BA11, BGIN13, Che11e, DM16, DCRL13, DHMU16, JZJ18, KNIF13, Lin14, LB11, MZES12, MHL11, PKD19, PL10a, SH11, SL16c, ZZ15a, ZLY12]. technological [GM18b]. Technologies [LY12a]. technology [ADL12]. telegraph [HA16a, LS10a, Pov12b, SPST18, YK17, ZLY14, ZT18b]. telegraph-diffusion [ZLY14]. Temperature [BCSCB^{+}15, CPP10, AM18, ATH19, BKE18, BMS13, BBL19, CL19, GRW14, Her14, OSA13, Özd18, RR14, SZA^{+}18, SCA14, SKK12, ZC11a]. temperature-composition [Özd18]. temperature-dependent [ATH19, Her14, OSA13]. Temperature-related [BCSCB^{+}15]. temperatures [FM17, RNQ16, ZCH14]. Tempered [GLW18, AKMS18, CWDL17, DA18d, DZ18]. temporal [AET19, DGA18, Hu19, LLG10, LR13, VJM15, ZZF18, ZLMZ18]. temporary [XMW10]. tennis [KSM12]. tension [GZN19, IS12, JVMF19, ZR18b]. tensor [EHL^{+}14, Gal11c, Git14, JW18a, KKVS19, LZ19b, MMRN12, Pov19, Sha12a, SK14b, WPL16, XD17, XW19b, ZLY14, ZHW^{+}18, ZZL18a]. tensor-valued [WPL16]. tensors [JBBL17, LKCN19, LZ19b, Loh16]. tent [GMS15]. term [AM13a, ARESH18, Bag17, BKR11, BTB18, BDS10, BSY^{+}19, CLA19, CZF10, CLJ11, DSL18, FLZ14b, FH16, FLWJ11, HH18a, HP13, HCL12b, JGSS10, JA11, JLTB12, KLL19, KPK18a, LP12, LD11a, Liu12, LYZ17, Liu17, LLLW18, LFAL19, MLZ^{+}16, QZM17b, RZZ19, SC19a, SBM10, SLL12b, Tia17, TY16, WSCW16, yYsZyYL13, yZjM10, ZL19a, ZLA17, ZZL^{+}18b, ZWH^{+}19, ZKW15]. terminal [fNS11]. terminals [RT11]. termination [LWJ10]. terms [ABL15, ANP19, GGR15, Gen11a, HKJ14, HC16, IS12, KNZ19, KPK18b, LSW10, MT10, PC17, PC11b, PL10b, QY17, WZ11a, Was13, WGY^{+}18, YSB15, ZL10a]. Tesla [AdSSS19]. Tesla-type [AdSSS19]. tessellation [FJWW16]. tessellations [WJWW12, XY16]. test [BGH14, CCN14, CHBTD14, DGH17, HK15, NCC13, XLZW11, YW11a]. testbed [KHIB12]. tested [AM15]. testing [CCDL10, Dun18]. tests [BK13, HAESLB14, HLB14, WHG11]. tetragamma [Che11a]. tetrahedra [MW14]. Tetrahedral [ILV^{+}19, CG13, GLLC19]. text [Hon12, MLY11]. text-independent [Hon12]. texts [GA10]. textured [ST18]. TFBSs [fLcJ10]. TFLHOWA [LS10c]. th [LM19a, HLWX11, HL11a, HL10, Yan11d]. th-order [HLWX11, Yan11d]. Thai [TNT12, WNC12]. their [BMRA10, Chu10, DSA10, FF14, GK11a, HPR19, HLY12a, HSMG12, JTC^{+}10a, Jia11, KK10c, LCWZ18, MP10b, MTV13, MS11b, RMY19, RA12, Sal10, Sal11, SK10a, Tia11, THH12, WYG12, Wan13a, XSLS11, XY16, XMWH10, Xue13, YA11]. them [TTG16]. themes [Boy16]. theorem [Çan11b, Çan11c, EÇ10, HP10, LZ12b, SZ12a, WFL11a, WHS12, Zha11d, ZZG19]. theorems [AE12a, AIB10, ASMEE11b, BD11a, BD11b, Ber12, CsH10, CB11b, ÇT11a, ÇHK11, Cha11b, CSW11b, CS10c, ĆSCD11, ET12, GDZ11, HLS11, HRMS12, KPR10, Kar10a, KT11a, KL12a, KK10b, KT11b, Lee11a, LjHO10, LFJ11, LC11b, LSJ12, LZ12b, LT11, MZLF10, NS11, NP12, PRR10, Pop11, RI12, She11, SCK11, XL10, YBC11, ZRC11, ZC10, ZQ11b, ZLL11, sHC11]. Theoretical [AM12a, CC15, GMP18, HLY16, PMM17, Sag10, TC18, XLF12, YWL^{+}11a, ZWL11, ZC11b]. Theories [Pov12b, BC15, BKK12, KM13, WW11b]. Theory [BCF10, DGL12, GW12a, GSPK15, GTG11, JMST11, KKK12, VBCJ10, ZSHL11, AA11, ABFGZ11, AHF10, AH11b, ATH18, BMS13, ÇE10a, ÇE10b, CG15, Col18, CDM10, CFB11, DD19b, EAEH18, FG18, GF16, GMS15, GH10, JZ11, JLP10a, JK11c, KBDC12, KHF^{+}19, KKK16, LCW12, Li11a, LLH10, LJJ17, LS10d, MTN19, MT11, MJWD19, PC11a, PW11, Pov19, SKPW14, SS13, WZF12b, WZF16, WZWS11, WWH12, Yao10, YZMA18, gZnZpZbD12, ZYZ^{+}16, ZYZC18, ZW11c]. theory-based [LLH10]. Thermal [ATH19, HLT12, LLML10, OAKR16, RYK13, SM19, AML^{+}14, BKE18, BSK11, BK15, BHH16, CNH17, CJ18b, DZW^{+}15, DVY14, GSI19, GIMZ14, HKS19b, Hei10b, HNPS13, JMDA13, JLD19, JFS14, JFS20, KWFY11, Kia18, KHF^{+}19, KÖC^{+}18, LWC13, MM18a, MS10c, Pal13, PZL^{+}18, Pov12b, SST19, TL18, VMO10, WWLL13]. thermal-acoustic [BK15]. thermally [HSK11]. thermistor [AT18b, AT19]. Thermo [HAESLB14, BCSCB^{+}15, FG18, GR10, GF16, HLNZ19, NH15, SSIP19, SCBCB^{+}13, SCBCB^{+}17, WY15, YWH14]. thermo-elastic [GF16]. Thermo-elasto-plastic [HAESLB14]. thermo-hydrodynamics [GR10]. thermo-hyperelastic [WY15]. thermo-hyperelasticity [YWH14]. thermo-magneto-electro-elastic [BCSCB^{+}15, SCBCB^{+}13, SCBCB^{+}17]. thermo-mechanical [FG18, SSIP19]. thermo-viscoelasticity [NH15]. thermochronology [GHMN16]. Thermodynamic [BB18, Ebr11, Gro19]. thermodynamics [CCF13, Grm13]. thermoelastic [CBB15, FM17, Her14, Kim17, MTN19, OSA13, RNQ13, RNQ16, SEM13]. thermoelasticity [CCCW16, ED12, KVV14, KM13, ZKR^{+}12]. thermomechanical [HLB14, JM16]. thermoviscoelastic [BDS17]. theta [BM12a]. thick [JW10, STS19, ZW11a]. thick-restarted [JW10]. thickness [GHL18, HD14a]. Thiele [SVY16]. Thien [ZH15b]. thin [ANP19, BPR18, BPZ19, DD19b, GHCZ18, GR15, IS12, JW15, JK12, KEHB18, KWFY11, KS12b, LS16, Lin14, MTN19, Mom11, STS19, YLY12]. Things [gZnZpZbD12, XCM12]. think [HN10]. thinning [BKR10, HHS^{+}17, HWY14, MLSLM15, Yan19]. Third [CSN11, AÇT11, ATH18, BG10a, FKKS11, GK11a, GGEB12, KRCJ11, KS12b, Li10e, LZZ19, LWKK10, LXK11, MC11, Par11a, RHMA18, XW18, YY10b]. third-grade [MC11]. Third-order [CSN11, ATH18, BG10a, Li10e, XW18, YY10b]. third-step [GK11a]. Thomas [CLH13]. thread [SKG^{+}11]. Three [GQF^{+}10, Ima17, JL11b, Kup14, LS16, MMA12, SSHH^{+}18, Sea16, SL18a, WY18a, AN12, AuIA17, BHZJ19, BPS19, BGP13, CB19a, CLA19, DM18, DA16, DLS14, DFW^{+}18, DH18, Dua18, DHMU16, Erg19, FGB19, GGR15, GSPK15, GHCZ18, GDM13, He16, HP17, HVO17, HWyL11, JK12, KL16a, KNT12, KSMN11, LP12, LK13, LDL11, Li16a, LMY19, LNW19, LZ11e, LZ11d, LSC17, LLYL19, LDHH13, MW17, Ma18a, MM18d, MS15, PDN19, PQB^{+}16, PQBK17, Pas14, RKA^{+}18, SK14a, SMF17, SBM13, TH19, TB10, TY13, Tom11, UMY11, UKI11, VFM19, WPL16, WHS17, WMP^{+}19, WLZ18b, XW14, YMM12, YK18, YLZ17, YCLY15, ZJ10b, ZBF11, ZZX^{+}14a, ZSW15, ZHC17, Zha18f, ZCLW19, ZHW^{+}11]. Three-dimensional [GQF^{+}10, Kup14, LS16, MMA12, SSHH^{+}18, Sea16, WY18a, AuIA17, BHZJ19, BGP13, DA16, DFW^{+}18, DH18, Dua18, GGR15, GHCZ18, KNT12, LK13, LMY19, LSC17, LDHH13, Ma18a, RKA^{+}18, SK14a, SMF17, UMY11, VFM19, WPL16, WHS17, WMP^{+}19, YLZ17, YCLY15, ZZX^{+}14a, ZSW15]. three-field [BPS19]. three-level [He16, MM18d]. three-phase [GDM13, TY13]. three-point [AN12, LZ11d, TB10, ZJ10b, ZBF11]. three-species [KL16a, Li16a, MW17, PQB^{+}16]. three-step [KSMN11]. three-wave [LDL11]. Threshold [HL18a, XDH16, Lin10a, Lin11, LRV13, OO10, RIW12, WZ17a, ZZ16a, ZPWY12]. thresholding [LTL^{+}12]. thrust [LLZ12]. Thue [HKT11]. thumb [QZY11]. Ti [GSI19]. tibia [DRT^{+}15]. tiered [YXP^{+}13, YGH11]. ties [LKL^{+}15]. tight [XSLS11]. Tikhonov [BKMT17, SM10, YF10]. tilted [HHM12]. Time [CLT^{+}13, CMT12, DL18, LCWZ18, LLLW18, MWL18, PK19, SBvdV13, SP12, XaZH19, ZCH14, ZFLM18, ZFLM19, AAR11, ARESH18, ABT19, AJS19, AK18, AKL18a, AK19, AJT19, AT19, Ana10, Ana11b, AHOP18, AHJM18, AR17, Arq18, AA10c, BJQS18, Bag17, BKR11, BS16a, BL18, BO10, BMM12a, BS16b, BO18a, BQS16, Bis10, BKDM13, BG10c, BDHR18, BBO10, BKR^{+}19, BOT14, BGM19b, BRROP19, BSY^{+}19, Can11d, Cao19, CFRS10, CHXL18, ÇT12, CXZ15b, CLC16, CJ18a, CLA19, CHT11, CR19, CTM^{+}13, CGK14, CELY18, CM14, CM12b, DH17, DVMS13, DSK^{+}14, DZ18, DFM15, DGT18, fDxZ11, DGK10, DZO^{+}19, DLS14, Dos10, DH16, Dra11a, DHGF17, DJZ18, EGAA19, EAAS18, FID14, FSZ18, FSRB15, FRSW11, FZBF10, FL10, FNW18, FIM18, GD10a, Gal11e, GLP18, GP11, GMP18, GGEB12, GM14b, GGT14, GABC16]. time [Guo12, HB19, HY10, HK10, HGSL18, HXX19, HW19a, HG18a, HP13, HEP10, Her19, HA16b, HZ11, Hon10, HA18b, HHGA19, HGW11, HZP18, HLI14, IK12, JKN10, JPK18, JW19a, JLTB12, JL18, JC12, JLF17, JKS18, Jum10, KKT13, KLP17, KVW18, Kar18, KKAM11, KYA15, KK14c, KCK19, KBK19, KM10, KGM11, LMR19, LHD18, Lee11c, LC12a, LR14, LPK15, LCW10, Li10d, LDW11, Li11a, LHW11, LZZC12, LLML15, LCK17, LLY18b, Li19a, LGZ19, LSC17, LK18, LZ19a, LZZ18, LCYC12, Lin10a, Lin11, LN19, LW10, LXP11, LH12b, LHL12b, LDL^{+}15a, LX15a, LDL^{+}15b, LWZ16, Liu17, LW18c, LW19b, LFAL19, Lóp19, LLL16, LXY19, LHTL19, Luc10, Luk11, LZ12c, MP11a, MF18b, MS18, MDVM17, MAB19, MHHC18, Mar12, MMT18, MT11, MSA12, MDG19, MR19, MZB10, MWWL11, MLZ^{+}16, MV17, MLG17, MM18d, MBJ16, MAH18, Mur08, Nab19, NHH13]. time [NH15, Ngo18, yN11, NWZ11, Oan13, OY19, OHK^{+}19, Pad18, PC11a, PW11, PS16, PQB^{+}16, PD17, PCM12, Pov12a, Pov12b, Pov19, QXG13, QZ16, QZF19, RAZ19, RSWZ10, Rah11a, Rah11b, RS18a, Ray18, RG18, RSB14, RMS12, RZZ19, SD15a, SC19a, SEY12, SZA^{+}18, Sam19, STC18, Sar10, SR16, SZ17, SNSK19, SL18a, Sha12b, She16, SZL^{+}17, SY18, SLM18, SOK19, SS18b, SG16b, SS16c, SG16a, SM19, SZC^{+}18, SI10, yS10, SKTC15, SKTC19, SRG16, SLW18, SW19, TF17, TZ13, Tia17, TCHW19, THD19, Tro13, TKHL18, TDN19, TNHK19, TD10b, UMLF13, VAK^{+}19, WSL10, WSG10, WSS10, WHS11, WWW11a, WWW11b, WJ11, WZ11a, WL12b, WYN12, WS12, WV14, WHTZ16, WW18b, WLA18, WLZ^{+}18a, WLL^{+}18, WRY18, WZC^{+}19a, WZG19, WZKY12, WZ18c, WLGL10, WL17b, XW19a, XZZ16, XWH16, XSYL19, XDL12, YY14, YY15, YF10, YY10c, YMSL11, YX11a]. time [YCHW18, YQWZ19, YH19, YS19a, Yas12, YZAX10, YDW15, YSX^{+}19, Zen11, ZFZ10a, ZFZ10b, ZHJ11, ZGW11, ZQ11b, ZLPM13, ZSW15, ZLTY16, ZHW^{+}18, ZCLW19, ZSL19, ZL19a, ZSW19, ZSLZ19, ZY19, ZLL17, ZLA17, ZZL^{+}18b, ZWH^{+}19, ZJZ18, ZLS13, ZY17d, ZY15c, ZG16, ZP18b, ZP18c, ZY13, aZ18]. time-decay [JLF17]. time-delay [BKDM13, LCYC12]. time-delayed [ZSW19]. Time-dependent [MWL18, ABT19, BO18a, Bis10, CM14, FSZ18, FSRB15, FZBF10, HP13, HLI14, KVW18, KYA15, LR14, LLML15, LZ19a, Oan13, OY19, QZ16, RZZ19, SD15a, SLM18, SG16b, TCHW19, Tro13, WWW11a, WWW11b, WHTZ16, YF10, YS19a, ZL19a]. Time-domain [SP12]. time-efficient [CGK14]. Time-fractional [LCWZ18, PK19, ZCH14, ZFLM18, ARESH18, AJS19, AK18, AJT19, Arq18, CHXL18, CXZ15b, CJ18a, DH17, GMP18, HXX19, JLTB12, Kar18, KKAM11, KCK19, LHD18, LLY18b, LDL^{+}15a, LDL^{+}15b, Liu17, LLLW18, LHTL19, Luc10, MAB19, MV17, MAH18, Pov12a, Pov19, RS18a, RG18, SC19a, SEY12, SS18b, SS16c, TKHL18, TNHK19, WL12b, WLZ^{+}18a, WRY18, WZC^{+}19a, WZKY12, WZ18c, XW19a, YY14, YY15, YCHW18, ZL19a, ZY19, ZLA17, ZWH^{+}19, ZP18b, ZP18c, aZ18]. time-harmonic [BDHR18, Cao19, DLS14, MDG19, PD17, VAK^{+}19, ZLS13]. Time-integration [SBvdV13]. time-invariant [Guo12, JC12]. time-periodic [LZZ18]. time-scales [ÇT12]. Time-space [XaZH19, BSY^{+}19, CLA19, LN19, LFAL19, Nab19, Pad18, PS16, THD19, ZCLW19]. time-step [GGT14]. time-stepping [SZC^{+}18, YH19]. time-variant [IK12, WSL10]. time-varying [BO10, BBO10, DVMS13, DHGF17, HW19a, HGW11, KLP17, LZ12c, SI10, WYN12, WS12, WZG19]. Time-weighted [DL18]. times [AJS19, AA10b, ÁBÁPM11, AA10c, BB10a, CL19, LCC12, RSS10]. Timoshenko [BMAR18, LMR14, TAA14, VLFS12]. Timoshenko-like [BMAR18]. tin [BO18b]. Tinkerbell [dSCM12]. TiO [ZAK18b]. tip [WY11a]. tissue [IS12]. tissues [Mag10]. TLA [ZMG10]. Toda [SR17a]. Toeplitz [BDS15, JKS12, JSEM13, JJ13, JYL16, JL17b, JL17a, LZ17, LWQZ18, WLW16, WLW17]. Tokamak [NKM16, NKM15]. tolerable [Pop13]. tolerance [PA15]. tolerant [ĆCM10, ZPS^{+}12]. Tomography [KY10, dCMdSGTdC^{+}16, BOY12, CX16, DRT^{+}15, GZR^{+}13, GR19b, GKS17, HKS19b, LLG^{+}11]. tool [BBR10b, JPP12, WCB13]. toolbox [JRA^{+}18, RAW^{+}16, FRAK15]. tools [ADD^{+}15, KTDT17]. top [HpD11, LLSW10, PQB^{+}16, WYY11]. top- [WYY11]. top-down [LLSW10]. topic [xLlFwWL12]. topography [MDBCF16, SZA^{+}18]. Topological [HM12, LXL12, TK11, AvdW13, BK10, DB15, DYX11, HA11, Kha10, LS10b, Min11, PDHL12, SN11, SW11]. topologies [LEP11a]. Topology [AGDP19, AdSSS19, DLQ16, HWY14, LYJ15, Tod12, ABSV18, ÇKE11, FMP19, JPCY13, KP19c, LWR16, LGL^{+}14, PH13, PM10, SSM12, VHPVNXW18, Yoo17, gZnZpZbD12]. tornado [GC19b]. tornado-like [GC19b]. torque [MK18]. Torrey [CLA19, LFAL19, QLT^{+}18]. Tortorelli [MPS18]. torus [JQSS12]. Total [ASSV18, CLM11, YZM^{+}19, AA10c, GKS17, LYLX11, LX15b, Liu16e, RZL11, YY10c, Zha19b]. Towered [YT11]. Toxicity [Li11b, CANA19]. toxin [DKG14, YLZ17, ZL14a]. toxin-determined [ZL14a]. trace [CW15b, HKS14, KSS13, MW14]. TraceMIN [KSS13]. Tracer [LPLR19]. track [JSGP16]. tracker [BZZ^{+}10, THY^{+}10]. Tracking [LL12c, FGB19, HDHL11, HCL12b, MZB10, PC14, VMO10, YLH12, YLC12, ZHQG12]. trade [Chu11c, Chu12b, SRS11]. trade-off [SRS11]. trading [WTC^{+}12]. traditional [BKL^{+}19, WHS17]. traffic [BDPM12, BKK12, DNP15, LKLP12, Sal16, SKTD13, WL12a]. trajectory [HW11]. transaction [CJPR10, LW17, LC13, TYY^{+}12]. transcoding [CUK12]. transcritical [PLMS14]. Transdermal [FdOP17]. transfer [BJRF19, BHZJ19, BB15, EA10, EUTS18, GRS18, GD10b, GK18, HLSN15, Ima17, KB10a, LZZ11a, LZP^{+}19, MSV18, MC10a, MM18a, MS10c, NPR10, OP14, Pal13, PCO16, PvdM13, PB11, RJGS^{+}19, Sah11, SL12, SRRP18, UKA15, YL18a, ZYT^{+}16, ZJB19, ZWG11]. transfers [HSMY12, JMNZ19]. transform [ArEM10, APRM11, Alo11, Amb12, AM13c, BG10c, BM13b, DB10, GZ14, Gup11, Hol11, ILP^{+}11, KK11a, KRCJ11, KW11, LWQZ18, LZS12, MZC17, MZM19, NDT11, OBAAD10, Ost11, Rah11a, Sat11, TS11a]. transformation [AB10a, BGM19a, BMY13, CZ15, DD10, EM19, LS16, LM19a, Liu11a, LSX13, MR17, SMK18, SSIP19, SI17, WY11a, WF17, WW18a, Wei10b]. Transformations [Din10, LWL11, EAAED10, EAA10, Gal10c, Gal10d, KK17, KTK18, Lam12, zLYmL18, SB19, WF18]. Transforms [Ehr18, KLMV12, RS12a]. Transient [BDGBM12, MC10b, BDGS13, BGRV15, BPS18, BJPT16, DLS14, FXCC18, GRS18, Ima17, MC10a, QWL19, WCH13]. transients [TMSO12]. transit [ZNWG11]. Transition [Liu16b, GSD^{+}19, LZ12c]. transitional [Özd18]. Transitive [BS11a]. translation [NJV13]. Translational [JKZ11, NWZ11]. translational/rotational [NWZ11]. transmissible [MW16]. transmission [AS15a, jASzZ12, BD19, BM18c, CCRS17, CL12c, CRXL15, CPT15, FH17, FP18, HK15, LZ11b, LHL12a, PM13, SG11a, SW19, WBZY18, ZWZ16, ZTZJ14, ZLG^{+}10]. transmissions [YZMZ16]. transmissivity [BCPS15]. transmittal [Gal10a, Gal11b, Gal11c, Gal11d]. transmitted [ABLS15]. transparent [JXZ^{+}10, KSF14, XD17]. transplant [KLY16]. Transport [RAZ19, AHP^{+}14, ASFM15, ABK10, BNR10, BDF16, CRG16, CDM10, CN16b, EN11, EAMA19, GQF^{+}10, GH16, JSGP16, KK11a, KKSK18, KS15b, LPLR19, NCC13, PZL^{+}18, SSSB11, ySW10, ySGL^{+}10, Sul16, Tim14, Tri11, TRL19, VGC^{+}15, YYK16, YZMA18]. transport-based [Sul16]. transportation [Chu12b, WWH12]. transpose [CM17, HM14, TM17, XHM14]. transverse [AM18, NZ16, TAA14]. transversely [FR15, ZR18c]. trapezoidal [AH10a, ADK10, BBC^{+}11, BC11, KSKK11, LR17a]. trapping [CYP16]. Traub [EGSHR10]. Traveling [BL18, CS16, FYYT11, LX17, Tia19, VDV13, WX18b, ZLY17, ABL15, AAEG17, jC11, DGA18, GMZ15, HBS^{+}10, KS12b, LZY12, LYC12, MS10c, QTW^{+}18, She16, tWqLzGkP11, XXH18, ZGZ13]. Travelling [JKK12, YW10, RG18, YGS17]. traversing [DVMS13]. treating [Gen11b, KA10a]. treatment [BJPT16, DFM15, GS11a, SSM^{+}17, ZHV19]. treatments [Yan13]. tree [EHO^{+}12, FNZ13, KCL12, LL17, LLSW10, WHC12, XDL12, ZZ19]. tree-structured [EHO^{+}12]. Trees [Ili10, BDK^{+}11, GDF12, LHH10, Mar11a, Pas14, QZ11, RdSSS11, RT10, SXB^{+}12, TFS11]. Trefftz [SWL19, TMCM19, Wei14]. Trefftz-like [Wei14]. trellis [Sha12b]. Trends [CKN11]. Tresca [BPZ19]. Tri [LT15a]. Tri-PSPS [LT15a]. triangle [Pap15, SLW11, ZWH^{+}19]. triangles [Erv12, MW14]. Triangular [WHW11, AG11, BMRA10, CBS18, CLH13, Dej11, GLLC19, HMZ18, Li10a, LH12a, LT15a, Liu11b, SRGL13, SBJ15, Zha18c, ZYW17]. triangulated [XSLS11]. trickle [IL13]. trickle-bed [IL13]. Tricomi [DA18c]. Tricomi-type [DA18c]. tricubic [WZH18]. tricyclic [Dol11]. tridiagonal [Abd18a, Gem16, JSEM13, JL15a, JL15b, JL19, JL20, Mar16, dS15]. Trigonometric [LWQZ18, BM12b, Cvi10, GW15a, MS11b, ÜM16]. triharmonic [AEG18]. trim [SG10a]. trimmed [GR15]. trinomials [CKK^{+}10]. trip [LJJ11]. Triple [WZH10, AM18, BFSO12, Sön11, TL10a]. Trivariate [DSB19, GK11b]. trivial [BDO11, SLKK19]. tropical [Lit13]. Truncated [KLCD16, WZH18, CChL14, MCB10, MLG17, PZJ^{+}16, PZAR19]. Truncation [EM19, SMK18, BGGCGRSP16, LY11b]. Trust [Ber16, DNZ^{+}13, SPCS13, AFGL10, AA10a, LLH11a, LLH11b, ZLL12]. Trust-region [Ber16]. trusted [WD13]. Trusting [BCF^{+}14]. truth [Wu10]. Tsallis [BPM12, MBKK10]. Tsanev [Sen12]. tube [GK18]. tubes [AVR17, Tri11]. Tucker [GN12]. tumor [APT11, CB19a, CGJ^{+}14, EE18, HBS^{+}10, HW19b, KFTT13, Özu15, SSM^{+}17, WHLC11, ZHC17]. tumorous [NCV^{+}18]. tumour [AWJH19, Amo18]. tuning [dSCM12]. tunneling [COR18, YLDL11]. tuple [SD12b, Zha11b]. turbid [LTLL12]. turbine [DGGBTRJF12, KDG11, WXYW11]. turbulence [DSWB18, Dun18, FVVS16, GWR^{+}18, JL16, KNIF13, MM16, PLR15, WAG^{+}14, WTSS10, YL18a, YCW^{+}14]. Turbulent [GUK13, Cos18, DdSF13, De 10, GLW13, GSD^{+}19, GJX18, HBK^{+}19, HS18, JQG14, KP18, LDS10, LZZC12, LZPZ19, MS17, NDC^{+}19, PZZ18, PLR15, SKTC15, SKTC19].

Turing [BGL^{+}15, DA16, SD11b, SZ14, TH13]. turn [yN11]. turning [KAG11]. TVD [HLL^{+}15]. TVL1 [LHL14a]. TVL1-based [LHL14a]. twin [KAG11, SD12b]. Twisted [Sim10, WY18a]. Two [AGT19, AA10c, BMRA10, BWL18, CP15a, mCfX10, CLN^{+}19, DM12a, DM16, FMS19, FDG^{+}17, hGzS15, HH11, HCZ16, Hu18, HM15, JJ15, LZ16b, Liu11b, MM18c, Pop11, QM19, RG18, RS13, ySW10, SJPS11, TCHW19, UBF11, WC17, WLW17, YD12, YGH11, ZN18, ZT15, ZLS13, ZYW17, AM13a, ATO19, AHHM19, AJRWS12, Ali15, AA10b, AdSSS19, AER12, And12, AVZ15, ADS14, ASSV18, AGD10, ABK^{+}13, BJLZ12, BN14a, BTEM19, BKL14, BBBM16, Bic11, BSZ16, Bis14, BMSS18, BM18d, BV17, BZT16, BSY^{+}19, BPF13, Cal19, CST14, CP16a, CQ11, CLW11, CLB14, Che14, CCCW16, CS18, CCZ18, CR18b, CHT11, CKRW19, CS11d, CTD10, Chu12b, Col14, CYM13, CX18, CDS15, Def10a, Def10b, DM15, DTR19, DB10, DLS14, DSVS15, DDM^{+}18, Dos12, Dra11a, DIS19, ES18, Elb11, FR15, FLZ14a, FLH10]. two [FZL^{+}18, FM17, FAIV10, Fio14, FIW13, FIM18, GBG11, GK11a, GH14, GMB12, GOT19, GL16, GML17b, GC19a, HGHA19, HY15, HPC12, HP17, HA16a, HSZ15, HD16, Hu19, HA10, HWXC18a, ID10, JJH16, JW05, JL18, KG14, KVJB15, KV17b, KN11, KRCJ11, KK14a, KYO10, KYR11a, KYR11b, KO11b, KLP10, KCL14b, Koj10, KPG18, Kuo11, KYY12, KGM11, LMPE18, LCM14, LA11, Lep11b, LD11b, LZH16, LR17a, LCS18, LZ19a, LNKU12, LXL14, LMZ18, Liu10, LHL12b, LRH13, LDL^{+}15b, LWZ16, LCHZ19, zLYLQ19, LFAL19, LGVS19, MP10a, MP11a, MPZ11, MCL15, MY16, MA17, MZB10, MNJ^{+}13, MLY11, MS15, MY13, MM19, MSFS18, NWZ11, fNS11, Oru19, PS12a, PD11, Por18, PLKCC13, QX19, Qiu12, RZ16, RQ18, RWW18, RWTW19, RFP11, Rus16, SSS16, SSHH^{+}18, SPT17, SK11b, Saj12, San12, SS16a, Ser19, SHH16, SKCL19, SZL^{+}17]. two [SY18, SPST18, uIAA15, SZC^{+}18, SRV10, SYW11, Sul16, TR14, TTG16, TM18, TS11a, TAPA^{+}17, TDM13, TRL19, Tur10, UMLF13, VC12, VMP15, WZ11a, WWLL13, WV14, Wan14, WV16, WW16, WY16, WH16, WCH18, WHS18, WBZY18, WZC^{+}19a, WO18, Wei12b, WFZ12, WZY13, WS16, WCLD18, XG10, XM15, XZ18, XZ11, XX19, XZL10, YGS17, YJ19, Yan19, YQWZ19, YXP^{+}13, YCLY15, Yil19, YYC11, YS16b, YSX^{+}19, YK17, ZCW15, ZZ15a, ZLS19, ZCSG13, ZDLC14, ZTC14, ZLL14, ZC16, ZR16, ZLJ^{+}18, Zha18d, ZT19, ZCLW19, ZJB19, ZL10a, ZYZ11, ZLA17, ZLC^{+}11a, ZJZ18, ZSD10, ZBL12, ZQ14b, ZR18b]. two- [YCLY15]. Two-aircraft [YGH11]. two-by-two [ZZ15a]. two-component [DB10, YYC11]. Two-dimensional [BMRA10, RG18, AM13a, AHHM19, AJRWS12, Ali15, AdSSS19, BSY^{+}19, BPF13, CLB14, CCCW16, CS18, Def10b, DM15, DSVS15, DDM^{+}18, Dos12, FR15, FLZ14a, Fio14, hGzS15, GH14, GL16, GML17b, HSZ15, Hu18, HA10, JL18, KG14, KK14a, KCL14b, KPG18, LCM14, Lep11b, LZH16, LR17a, LCS18, LWZ16, LFAL19, LGVS19, MP10a, MP11a, MPZ11, MM19, MSFS18, Oru19, PD11, Por18, Saj12, SZL^{+}17, SY18, SPST18, uIAA15, SZC^{+}18, TS11a, TCHW19, VC12, WW16, WY16, WCH18, WHS18, WZC^{+}19a, WCLD18, XZ18, XZ11, XX19, YGS17, Yil19, YSX^{+}19, YK17, ZCSG13, ZDLC14, ZTC14, ZC16, ZLJ^{+}18, Zha18d, ZYZ11, ZLA17, ZJZ18, ZSD10]. two-directional [Bis14]. two-fluid [RQ18]. Two-grid [BWL18, HCZ16, Hu18, QM19, TCHW19, WC17, BN14a, CCZ18, CR18b, HD16, JJH16, LZ19a, LRH13, LDL^{+}15b, WBZY18, WFZ12]. two-handed [LNKU12]. Two-layer [AGT19, zLYLQ19]. two-layer-liquid [WO18]. two-layered [SRV10]. Two-level [ZLS13, AVZ15, Chu12b, FIW13, LHL12b, MS15, WCLD18]. Two-machine [AA10c, AA10b]. two-mass-spring [GBG11]. two-material [ZLS19]. two-mode [BJLZ12]. two-parameter [FIM18]. two-party [HPC12]. two-patch [KVJB15, KV17b]. Two-phase [CP15a, ATO19, ABK^{+}13, BTEM19, BSZ16, BV17, Col14, DTR19, GOT19, LMPE18, SSS16, SSHH^{+}18, SPT17, Ser19, Sul16, TR14, TDM13, Wan14, ZR16, ZLC^{+}11a, ZQ14b]. two-point [AER12, Bic11, Tur10]. two-prey [ADS14]. two-relaxation-time [KGM11]. two-scale [TAPA^{+}17, TRL19]. Two-sided [SJPS11, FZL^{+}18, MY13]. two-sidedly [CST14]. two-solitary [LD11b]. two-species [CKRW19, LMZ18, ZT19]. Two-stage [MM18c, YD12, CHT11, MLY11]. two-state [ZBL12]. two-step [CYM13, FLH10, HWXC18a]. two-sweep [RWTW19]. two-terminal [fNS11]. two-tiered [YXP^{+}13]. two-variable [GK11a]. two-way [Koj10]. Type [LN98, SSR11, AR10a, AD10a, AD11a, AÖ10b, AP10, ADK10, AdSSS19, AAA12, Ana11b, AR17, AGPCC10, ANR11, Bae10, BKT11, BK12a, BV11, Ban13, BM11b, BM18a, BCK11, BX10, BD11c, BF11, BML11, BHJ14, CMGR11, CGY10b, CC19, CQRW11, CTG17, CS18, CO19, CZ15, CZ17, CS12, CT10a, CCP19, CD16, CJCV10, Cov13, DLS18a, DN18a, DN10, Dar11, DA18c, DK12, DDM^{+}18, Dra11a, Dra11b, DFS14, DSL18, Ers16, EKE18, FKF13, FLH10, FL11a, FM12a, FXCC18, FP18, Fur13, Gao17, GTC18, Gav12, GKK11, GHM^{+}14, HY18, HGSL18, HLLM19, HmZ11, HCF16, HSBL11, HN10, JJC11, JPS10, JH10, KP13, KPR13, KC11, KKBR19, Kha10, KYR15, KJ11, KM11, Lee11a, LSCG16, Li10c, LLW10, LD11b, LZYW13, LL16a, LX17, Li18b, LZL19, LCW19, LD19, LSZ11, LRV13, LTSW16, LMZ17, LYZ17, LW18b, LCLL18, LYC^{+}19]. type [LLP19, LPP15, Lu11, LS11d, Lü14, Luo19, LPY16, MPLR18, MI16, MN17, Mor10d, NZ14, NP12, Nil11, NUH12, eOS18, ÖAK11, PCS13, PS12a, Par15, Par17a, PRR10, QXG13, QHT16, QZM17b, QM19, RKW12, Sal10, SSO10, Sar10, SS11b, Set12, SV11, SC19b, TL10a, TZZ11, TÇA12, THD11, THH12, TTX^{+}16, TNV19, TD10b, Uze10, VA10, VSI12, VAB12, Von11, WL11b, WW15, Wu18a, WX18b, XZR16, XC18, XWH16, XZL10, YY14, Yan11b, YL13, YM17, YGS17, Yu17, YZ19, Yüz12b, ZS11a, ZHJ11, ZJ17, Zha18b, ZDM11, ZHJD13, ZT16b, ZWW13, Zhu10, Li13, MBKK10, Üns18]. type-2 [LZYW13]. Type-II [Li13]. types [BR13b, GD10a, ID10, JW05, JLP10b, XM15, YZ10b].

UBIAS [Ogi12, Ogi12]. ubiquitous

[AJJAD^{+}10]. UBM [Hon12]. UCM [RY10]. UHF [QCYL12]. Uhlenbeck [Li18d]. Ulam [Cie11, WZF12a]. ULT [LM17, ZL16]. ultra [AY12, Füh18, VMC^{+}14]. ultra-parabolic [AY12]. ultra-short [VMC^{+}14]. ultra-weak [Füh18]. ultraslow [YLC18]. Ultrasonic [CLW11, HB19]. Ultrasound [LYSZ19, FQLC18, GGL13, Ray16, jZsQdLmG19]. unbalance [PFBL10]. Unbounded [SZ11, BKR11, FMS19, GH14, LWBW13, WL17a, YL14, YLG17, YL18b]. uncertain [AEG11, BV11, BR12b, CZMZ11, Gao11, HGN^{+}10, LZ11a, LZT11, LBH^{+}12, LCYC12, Lin12, LZ12c, PG10, WYY11, YWHC11]. uncertainties [FPB17, HLCY12, JZ12, LC12a, LYY12, SP12, WLT13a]. Uncertainty [BCPS15, LZYW13, SH12b, WY15, FSM19, Gor13, HDS11, WS12, WLT13b, ZHT11, dPRVRB13]. Unconditional [GL16, GH18, SW17, YS19a, LLYL19]. Unconditionally [SS14a, AHP^{+}14, HP17, LLJK10]. unconfined [KP19a, SVY16]. unconstrained [AA10a, FLH10, GYTD12, HSD10, KL12b, LFZ19b, MHL11]. unconventional [DPM15]. uncountably [LWKK10, LZKU11]. under-determined [TNP17]. under-resolved [NDC^{+}19]. underlying [ZH18b]. underwater [TTT10]. Unfitted [FVZ18, WC15, dPLM18]. Uniaxial [ZR18b]. Unidirectional [BB18, Sza15]. Unified [FL11b, KSO16, LZB15, Öza11, Ala10, AS19, FDXW11, HL18b, Kun12, LPML19, OSS10, WSCW16, XC11b]. Uniform [FSTN18, FSZ18, FST19, LZ18c, Abb10, AM12c, CJ15, DE10, HKJ14, HO19, JK10, KNZ19, KK14c, LPML19, MPGW19, MN10a, Oru19, SLK12, SKG^{+}11, SKFG11, SR18, TAA14, Tri11, XZ11, yZjM10, ZCSG13]. uniformization [ZWC10]. Uniformly [CG14, LN98, Li10c, AKS10, BBD10, CsH10, CS11c, MS18, NPR10, Noo11, SK11d, Yao10]. unilateral [EKZ17]. unintuitive [SLKK19]. Unique [Che19a, SWC11]. Uniqueness [CP15a, LY11a, LWD15, QYL10, WRY18, WCW13, ZX11, BK10, CM11a, CM15, DQ10, FKeT12, GRS18, Goo11a, JS11, LJ11, LZ12a, LZ11d, LLL12, Liu16c, Luc10, MS12a, MY10, Ouy11, Qiu12, SW12, SLL12b, TNHK19, WAW15, WW14b, XLY10, YWL17, ZSH11, ZDB19, ZLW19]. Unit [Gal10d, DSK^{+}14, DH10c, JFS14, JFS20, LYZ11, LSM11]. Unit-free [Gal10d]. unitary [OVV^{+}16]. united [Wan12]. units [BGRS11, OKTR11]. unity [BASW18, CDP16, GCE18, KdLK19, SL16b]. univalent [KHUO12]. univariate [WXF10]. Universal [Tso13, TL10a, WZCC10, Zhe11]. Unknown [CZY13, AsNAd10, HN18, HP13, SBM10, TNT12, THY^{+}10, TKHL18]. unknowns [KLP10]. unlabeled [MHH11]. unmatched [HLCY12]. unreduced [GH13a]. unreliable [GD11, WY11b]. unsaturated [BKNR19, KMT19, SSS16]. unsaturated-saturated [KMT19]. unstable [PP12]. Unsteady [BPZ19, BKR10, LZZ11b, MCR11, ZWG11, AV14, BHJ14, DZW16, DM19a, EA10, FKKS11, HH17, HHS^{+}17, KNZ19, MC11, MLZ^{+}16, Pal13, RKF18, RS14b, SYI12, SZZ11, YYK16, Yan18b, YT18b, ZAK18b, ZTC14, ZH15a, ZZLB18, ZLZG11]. unstirred [WWG10]. Unstructured [LFAL19, ZQ14a, BDR19, DFP^{+}13, JZ12, JCF19, LPLR19, LZZ19, LZJY16, PZZ18, TBP19, WZH18, ZÖXL^{+}19]. Unstructured-mesh [LFAL19]. Unsymmetric [BS17, LHL^{+}14b]. updated [RWW18, WSM^{+}19]. updates [LHL^{+}14b]. updating [DWZ13, KH13, MHL11, XC16]. UPFC [AR10c]. Upper [HGSL18, Ipe12, Pul16, ZY15c, AJ12, LT18, SB19, San12, SRG16, WL17a, WZ10, ZS11b]. Upscaling [PLKCC13, HJ13, ILS13]. uptake [IS12, XBHN16]. upwind [CW17, sCYhX18, CS18, CKRW19, JYYL16, LHL15, YT18b]. urban [BDPM12, KAK^{+}12, VMAVGCM19]. Urysohn [SNH10]. use [ABFGZ11, GZ10, Ikh11, KK10c, MTN19, OO10, SSSB11, SD12a, SDH13, ZZWG16]. used [HZ16]. useful [Li11a, MSH10]. user [LYX11, SBA10, TXZ^{+}10]. user-private [SBA10]. users [KAK^{+}12]. Using [DFJS10, FHH13, HSZ15, LCW12, LCA^{+}17, WSCL11, AAR11, ATO19, AHV10, AMGC19, AR10a, AÖ10b, ASA16, AB10b, AJAR18, ArEM10, AM14a, Alo11, ATH18, AJS14, BOY12, BPM12, BRFH16, BAO^{+}12, BMM12a, BBBM16, BBO10, BL12, BE18, Bri10, CKSL^{+}14, CKL18, CDFP12, CD14, Cer18, CKK^{+}10, Cha18, CCL^{+}12, CL12b, Che12c, Che15a, CHT11, CUK12, CLJ11, CFLX18, CAC14, DGB10b, DRT^{+}15, DPM15, DM15, DA16, Del13, DMPV10, DB12, DGH17, ES18, EAAS18, FJB19, FKC12, FL13a, FLWJ11, GLLC19, GVSP12, Gep16, GHT^{+}15, GTL16, Git14, GMI11, GMAM12, GGR19, GHCZ18, HTGSH13, HH18b, HKI12, Haz11, HJ13, HGJP19, HJD15, HMM12, HA18b, HYS^{+}14, HH10a, HD19, HPY10, IK12, JKB11, JKK12, JK11a, JCWZ16, KAG11, KRCJ11, KMS19, KW11, KSMT11, KPP13, KW14a]. using [KMS15, KL16b, KHF^{+}19, KVR11, KSMN11, KKK16, KK17, KW12, KORR10, LGHR16, LS10a, LMR19, Laz10, LDS10, LKLP12, LL12a, LKK12, LPLR19, LSW16, LLW10, LK11, LS11b, LZZC12, LY15, LS16, LW17, LYY12, LL12c, Lin14, LLG10, LGH^{+}11, LTL^{+}12, LX15b, LZJY16, LCLL18, LvS15, Lu11, Lup11, MAN^{+}15, MCKM12, MPY16, MNJ^{+}13, MNPD15, MG11, MD15, MBKK10, MN10a, MM19, MAK12, NKM15, NH15, NGG12, OHK^{+}19, ODAZ15, PC14, Pap15, PZJ^{+}16, PZAR19, PJ17, PT15, PM10, PDHL12, QLT^{+}18, QWGJ15, RMS10, RMA10, RSL^{+}18, RES10, RIW12, Ray18, RSB14, RMS12, RG11, RCG15, RR11, ST15, SH11, SS11a, Sea11, SL18b, ST18, SSM12, Sin16, SL16c, SSP13, SM19, Ste16, SW10, TR14, TMMASG10, TTT10, TY13, TBP19, VP11, VMO10, VBK13, WJ11, WAG^{+}14, WLA18, WKBR18, WY15, WO10, WTC^{+}12]. using [WCW13, WL17b, XFL16, XY14, XLZW11, XXG10, YYK16, Yan12a, YWHC11, YMDZ10, YASK10, yYsZyYL13, YW11b, YLLN16, YK17, ZHB11, ZM14, ZTR11, ZSAN18, ZSAN19, ZW11a, ZS11b, Zha18d, ZL13, ZPGW16]. utility [KC12]. Utilization [SZDO10]. utilizing [QXLL11, ZZHF12]. Uzawa [CM16b, FP18, LM18b, Yun13]. Uzawa-type [FP18].

V [McN12, She16, Amb12, AM13c, GSI19, ZFLM19]. V-line [Amb12, AM13c]. V-shaped [She16]. vacation [LXYT11]. vacations [WY11b]. vaccination [DYQM14, KKL16, XXH18, YZ18]. vaccine [SG11a]. vacuum [RJGS^{+}19, Wu11a]. Vague [HLY12a, XMWH10, YJZ11]. Vakhnenko [jC11, LMMF17]. valent [AX11, EAA10, YL10a]. valently [PA12]. validating [SXB^{+}12]. Validation [HHNLGC18, SNEP19, GSI19, GSPK15, MS13, MLSLM15, MBT^{+}13, OAKR16]. validity [AaC19, RZL11]. Vallée [Kig10]. Valuation [Ciu11, CWW15, CW14, CXZ15a, CWDL17, DT17, ELS11, FXC18, LZL16]. value [Aba10b, AEG18, AA11, AN11a, ANP11, AW11, AH11a, AAH^{+}18, AIIZ10, AA18, AER12, ASSV18, BZ10, Bai11b, Bai12, BS12a, BBC^{+}11, Bic11, CHS11, ÇT12, Cha11a, CP16b, CCD10, Che11d, Che12b, CLCF14, CX18, CN11, DGB10b, DL10, DN10, DGK10, DBEE11, DB12, ESN10, Fer11, FM12b, GCE18, GMI12, Goo10, GKM11, GR13b, GW12c, GF19, HLWX11, HYCP11, HL11b, Jan10, JZ11, KA10a, KP10a, Kaw15, KM12, KFYW11, Kim11, KAS11b, LC11a, LN98, Li10c, LLZ10, LCW10, LYZ11, LW11a, LW11b, LS11a, LL11, LYLX11, LCK13, LZH16, LR17a, LJ10a, LZ11e, LZ11d, LZ14, LHF11, LCZ11, LLL12, LS12a, LS12b, Liu12, Luc10, MM10a, MN10a, MS12b, PC11b, Pit12, Pov12a, RS15, RCH19, SKPW14, SKM11, Sta11, SZ11, Su12, SZW11, SZ12b, TB10, TKBMT17, TDN19, VAB12, WZH10, WWA11, Wan11, WG11, WZF12b, WQRZ14, WHS12]. value [XZC12, XLD11a, XLD11b, YK18, YY10b, Yan11c, Yan11d, YZY10, YC10b, Yas12, ZSH11, Zha18a, ZJ10b, ZY10a, Zha10, ZBF11, ZLC^{+}14, ZSS10, ZL10b, lZxLhY12, dPLM18, uRK11]. valued [AD11b, AvdW13, BC11, BBR10a, ÇA10a, CMS10, CS11d, FLLF10, FRZ15, HRMS10, HRMS12, JPS10, JTC^{+}10a, Jia11, JMST11, JY11, KR11, KK10b, LKS10, RI12, gShYL10, WD10, WG11, WYG12, Wan13a, WL13a, WPL16, Wu10, YLY^{+}09, vdW14]. Valuing [JYK16]. valve [KEHB18].

Vandermonde [Res16]. vanilla [SL16b]. vanished [AS11b]. vanishing [FN14]. vapor [Kup11, Kup14, Laz10]. vapor-liquid [Laz10]. VAR [DNZ^{+}13, Che11b, CS14b]. variability [RS12b]. Variable [KK13a, WN18, AJRWS12, BB10a, BZ18, CL17a, CTSX16, CLTA11, CLB14, Dem10a, Dem10b, Ebr11, ES17, ES18, EA10, EUTS18, FZL^{+}18, FRZ15, Gal10c, GK11a, HH18a, HP17, HHGA19, IMS19, JL18, KSZ18, KÖ10, KWFY11, KYR15, Lam12, LY11b, LCS15, LLML15, LWL11, LYY12, LN19, LZ19c, zLYLQ19, LXYT11, MDW11, Mac12a, MCB10, MC10a, MA10b, MCR11, MHHC18, MM18a, MAST18, MV11, MM18d, NLA19, NCV^{+}18, Osm18, RS14b, RA19, SD10b, SYG11, SZA^{+}18, Sha18, SNDK18, SRG16, VMFF18, WW18a, WLZ^{+}18a, WZC^{+}19a, WLXZ18, Wu16, WTLS18, XZ18, YTC^{+}18, ZYG10, ZLPM13, ZDLC14, ZZLB18, ZWH^{+}19, ZW11d, ZY15c]. Variable-coefficient [Dem10b, CTSX16, ES17, IMS19, Lam12, LWL11, WLXZ18, WTLS18, YTC^{+}18]. variable-exponent [MAST18, Sha18]. variable-order [BZ18, CLTA11, CLB14, HHGA19, JL18]. variables [ABFGZ11, CTM^{+}13, FLDZ12, LLZ11, LL13, Liu10, MRR11, Nie10, RÖ10, Sun11, TMSO12, XZL10]. variably [DM15]. variance [BKM11, HZ18, JMNZ19, LS11c, MDW13, WLHZ14]. variance-reduced [MDW13]. variant [IK12, KG11, LHY18, LZ17, LBJ10, SS18a, SJHC14, WSL10, Zha19b]. Variants [EGSHR10, LZ18b, Yun13, FL13a, LZ16b]. variation [GKS17, KS11, LX15b, Liu16e, TD10a, Wei10a, YZM^{+}19, Zha19b]. Variational [Agr10, AMA14, Auc18, CTS19, DRD12, EK16, Elb15, GH15, Kar18, LCZ11, Naw11, PTP14, PMA17, SEY12, YMDZ10, YYYH19, Aba10b, AM14a, ABCR10, AT11, Ant14, BL10, BLS18, BHM19, Bar17, BMM12b, BPG10, BKY10, BXKZ11, BX14, BS15c, CB11a, CB11b, CB11c, CY14a, CGY10a, CKW13, CW10b, CC11, CW14, CAP10, CT10a, CMT12, DSZ18, DWS19, Els10, Far11, Gen10, Gen11b, GALO18, HSWZ11, HZ11, HR15, JJC11, JPS10, KA10a, KFYW11, KTA12, KH18, LGHR16, LKS10, Lu11, Mai16, MT10, MT11, MT12, Mat19, NLA19, NZ14, NUNAS11, NN12, eOS18, PN10, Ped18, QL10, QZM17b, QM19, RGdSRLAJ10, Sea11, She11, She12, SMH18, SS10, UCK16, WD10, cW11, qXjH11, XJLX10, YXS10, YC11, YX11b, YLK10, YDL11, Yua18, ZCW15, Zha11a, ZLLF12, ZHZ14, ZLW18, ZLWL11, ZY10b, ZW11d]. variational [dCM12]. variational-hemivariational [BHM19, SMH18]. variational-like [RGdSRLAJ10]. variational-type [JPS10]. variations [FXCC18, NUH12, OMT12, PAT13]. various [BC15, BK16, KC11, LBZL11, SCV10]. varying [BO10, BBO10, CRG16, CBB15, DVMS13, DHGF17, HW19a, HGW11, KLP17, KMT10, LZ12c, OÖ11, SMC10, SI10, WYN12, WS12, WZG19, WCCS15, Xia11]. vcBK [ZZ11b]. Vector [AHHM19, AH11b, BOY12, CEF^{+}13, CLF10, CLW11, DGGBTRJF12, GAVOF11, HLY12b, JPS10, JWX^{+}13, Kha10, LZ11b, LSJ12, LW12c, LLZ12, Mar12, OVV^{+}16, PT11, RH15, SW11, TJQS13, yZjH12, ZQ14b]. vector-borne [LZ11b]. Vector-potential [AHHM19]. vectorial [RGdSRLAJ10]. Vectorized [KB13]. vectors [GAVOF11, JW10, RS14a]. vegetation [RB19]. vehicle [GWR^{+}18, KPL11, LLWZ11, RT11]. vehicular [CHY12]. veins [MAN^{+}15]. velocimetry [RMB^{+}14]. velocities [Gal10b, Gal10c]. Velocity [GSZ11, BMJ10, BS18b, CT10b, De 10, Fer12, GOT19, HL18b, Liu18c, NWZ11, PLKC16, RS14b, RR14, VMP15, Yan19, YLC12, YT13, YWT18, YT18b, ZSW15, ZZ16b]. velocity-components [PLKC16]. velocity-two [VMP15]. Venant [ZTR11]. Vendor [DMPV10, BO10, KGJ11]. ventricular [AY18]. Verification [PBS12, CZ11b, WMSH11, YWL^{+}11a]. verified [MMS^{+}18]. verifying [ZMG10]. Versatile [GMAM12]. version [Haj18a, SHH16, WHS12]. versions [AT18a, JKS19]. versus [BOT14]. Vertex [CG13, ATZ11, CBBE16, CCM14, DYX11]. vertex-based [CBBE16]. Vertex-centroid [CG13]. vertex-degree [DYX11]. vertical [ABV11, Das12, DGM14, GDM13, MCB10, MS10c, NPR10, REHA11, RRP16, SPH10, ST18]. vertices [LSM10]. Very [CT17c, CM14]. vesicles [DVM12]. vessel [BNR10]. vessels [HZLM10, MDL18, YLF19]. via [AD15, AA15, AHF10, Ant10, AKMUH17, BK11a, BMAR18, BR12b, BGM19a, CP16b, Che11c, CSSW12, CTG17, Dan12, DAM14, Far11, GTG11, Haj18a, Haj18c, HT18, HGHA19, HLSN15, HLY12b, JZ11, JK18, Krn12, Kun12, LK15, LZC13, LTT13, LFZ19a, LCYC12, Lin11, Lin12, fLcJ10, LLH11a, LLH11b, zLYmL18, LLSW10, MR17, ÖAK11, PB12, PW18, QCT17, RSS10, SSAM11, SHM13, SIL19, SDH13, Set12, SS19, Sou11, SZ12b, SK14b, Wan10a, WZWX11, WZF16, WY16, WFL11a, Wen18a, XFH19, XC11b, Yan11d, Yan19, YLK10, Yaz11, YAS^{+}11]. viability [MGW11]. vibrating [CNR10, GHC15a, SSS11b]. vibration [BC15, DD19b, JW15, Kia16, KYA15, LPML19, PLW^{+}18, PHM^{+}19, SSH15, TAA14, WY18a, WL17b, ZTW^{+}19]. vibrations [CKR10, JZ13, RSS10, TS11b]. video [CUK12, HDHL11, LYM12]. view [CX16, GH10, XZL^{+}11]. viewpoint [LS10d]. VIKOR [PCK13]. viral [WYL19, XGH17a, XGH17b]. Virtual [Rus16, WRW13, AAB^{+}13, ABSV18, Cal19, DM18, TMLF19, WTC^{+}12, CM16c, MRR18, Vac18, dVDR18]. virtualization [JXZ^{+}10, JCZZ13]. virus [EOM11, SW19, TTM19, WZXS17, YX16, YZMZ16]. virus-to-cell [YZMZ16]. viscoelastic [BP19, BTB18, CRA19, CYZZ18, FH16, GGL13, HC16, HW19a, KLP17, KDG11, KPK18b, MLZ^{+}16, NB11, QaY18, RSS10, RS13, RSP18, SPH10, Sha18, SP12, Tia17, TY16, WSW11, Wan13b, XC13, XJ18, XHH^{+}19, YZ10a, ZH15b]. viscoelasticity [NHH13, NH15]. viscometer [HAESLB14, HLB14]. viscoplasticity [FKF13]. viscosities [WD16]. Viscosity [KN12, CB11c, CYL17, CZ11b, CL15, DNR13, EUTS18, GRBT16, GC19a, KAG11, KWFY11, LL16c, Ma18a, Ma18b, MCB10, Mai10, MA10b, MCR11, Naz13, RTV17, SD12a, TLR17, Wan14, YZS18, ZZ18b, ZH19]. Viscous [ySGL^{+}10, eMA18, DSL11, GZN19, GGO16, PTH^{+}16, ST14, Wan14, YSW16, Ye15, Ye19]. Vision [KPL11, Hua12, RMS12, YC12]. Vision-based [KPL11]. visualization [LMPG13, RSL^{+}18]. Visualizing [SLKK19]. visually [WNC12]. vital [KK12]. vivo [CSU13]. Vlasov [GSY10, MB10b, RÖ10]. vocal [ZJZ^{+}11]. VOF [JGK13, MLSLM15]. VOF/FVM [MLSLM15]. voids [SWL19]. Voigt [AT17, BN14a, BP19, KLRW12]. volatility [BM12a, BS10b, CKM12, HZ18, LW17, LL17, SYZ19, WY19a, ZZ19, ZC11c]. volcanic [BCF10]. voltage [AR10c, BPS18, ZLG^{+}10]. Volterra [ABJ11, AD12, ACTB19, BMRA10, BR12a, BGPP11, BE12, BK13, CS16, CSW11a, DSR10, GKLR11, HKI12, JWX14, KMRN12, MG16, MTM11, MBH11, MKR12, MBHV10, SNMA12, SYG11, TS11a, Tom11, Via15, YSS11a, dAS18]. Volume [CDS15, FIS18b, LK15, Pas14, ABLS15, BMSS18, CG13, CWW15, CS18, CCZ18, Che19b, CV14, CM14, CCM14, CSU13, CT17c, DPBL16, DRK11, EFK15, EKE18, GY15, GR10, GGO16, HMF^{+}19, HLZM16, HZM11, IS14, JW19a, JYYL16, KRBS18, LLFT17, LW17, LHL18, LHL15, LZJY16, MGTH16, MZM19, MM19, NSYY13, RQ18, SD11b, SLCC12, Tam16, ZL19b, ZSY19, ZQ14a, ZZL15]. volume-filling [HLZM16, IS14, MGTH16]. volumetric [FGPP17, Zho16]. Voronoi [FJWW16, WJWW12, XY16]. Vortex [Che18, GC19a, WFC16]. vortices [CNH17]. vorticity [eMA18, CRG16, RCG15, Wen18a, YTZ17, ZZ16b, ZaY17]. Vries [YYC11, Asl10, Dem10b, ES18, KT18a, LGZ19]. vs [BKL^{+}19, OC10, PMM17]. vulnerability [ZHQG12, ZHJZ11]. vulnerable [JYK16].

W [EMRS12]. WALE [WTSS10]. walk [Mil18]. Walker [EMRS12]. walks [AHOP18, JKMS12]. wall [AD19b, Bog11, CDW11, HBK^{+}19, HWY14, MJ14, MMA12, PAE^{+}12, SRDD17, YL18a]. walled [Kia18]. walls [BNR10, BSK11, RKA^{+}18, STS19]. Walrasian [Sca11]. warranty [CL12a, VCM11, VC12]. waste [MC10b, ZZHF12]. wastewater [DFM15]. Water [GZD18, BASW18, CRXL15, Dem10a, Dem10b, EGAA19, GZR^{+}13, HIS19, KHWK10, KK19b, Laz10, LF11b, LZ16a, LFC16, LZ19c, MZM19, MNJ^{+}13, Men18, MDBCF16, PLMS14, RKA^{+}18, RS18a, RMY19, RRO17, SZA^{+}18, ySGL^{+}10, TTT10, WBN18, WFL11b, YGS^{+}16, YLG10, ZÖXL^{+}19, ZDZY17]. Water-like [ZDZY17]. waterborne [ZWZ16]. waterline [LL17]. Wave [CBB15, XCZQ16, ARESH18, Abd18b, ABL15, Ade16, AJS19, AAEG17, AKMUH17, Bac14b, BCH^{+}18, BL17, CNR10, CM18a, CLC16, jC11, CCKY12, CZ15, Chu11a, CZ18, DGA18, DA18d, DL14, DSVS15, DC15, EGAA19, ES17, ES18, FYYT11, FH16, FIW17, GM18a, GZZ^{+}16, GH15, GSY10, GMI11, GS11b, GR19b, GZD18, GMZ15, GML17b, HKJ14, HB19, Hal13, HC16, HH18a, HW19a, HLL13, HP19b, HNPS13, HSC17, HTY^{+}19, HLY12b, HLY17b, IMS19, JGSS10, JA11, JKK12, JPK17, JPK18, JQSS12, KBA11, KSG11, KKLJ11, KPK18b, KSF14, KS12b, KBK19, KK19b, KK19c, LD11b, LDL11, LZ16a, LFC16, LX17, LZL19, LM19a, LGZ19, LCLL16, LX15a, LYC^{+}19, LZ19c, LWW19a, LMP13, MWL18, ML19a, Man18, Men18, ML19b, MTAS17, MAST18, MNPD15, MNT15, MR15, PD17, Por18, Pov12a, QTW^{+}18, RZ16, RAD13, Ray16, Ray17, RG18, RMY19, Ros12]. wave [SC19a, SAIZ15, Sea15, SIL19, SS18b, SRG16, SMYK19, TZ15, TG14, TTX^{+}16, TDN19, UKI11, VKJ13, WF18, WZ18a, WWD18, Wan19a, WZ18c, WZMY18, WX18b, XW19a, XZL^{+}11, YTD^{+}18, YW10, YC11, YGS17, YCHW18, YL14, YLG17, YHC18, yZjM10, ZLL14, ZM17a, ZCT18, Zha18d, ZBFC19, ZWH^{+}19, ZD15, dVLV18]. wave-guides [KSF14]. wave-like [GH15, SAIZ15]. wave-particle [GSY10]. wave-soliton [HTY^{+}19]. wave/diffusion [JLTB12]. waveform [BOT14]. Wavefront [HLZM16, MR15]. wavefront-based [MR15]. wavefronts [DMP18]. waveguide [Liu18b, Osm18, ZR18a]. waveguides [ZSZ18]. Wavelet [BPM12, KV18, SB14, AuIA17, BP11a, BBR10b, DB10, GM14a, HHGA19, LWZ16, Liu16d, LZS12, Oru19, eOS18, RG18, SNMA12, SKH12, VAB12, XCZQ16]. wavelet-based [BBR10b]. wavelet-Galerkin [SNMA12]. wavelets [AuIK11, BE12, Cer18, HGHA19, JYF^{+}11, Lep11b, MKL11, uIAH10]. wavenumbers [SS14b]. waves [BL18, CJK18, CTSX16, CWH13, Dem10a, Dem10b, DLS14, DTYZ18, Dub13, EGAA19, EK16, FER15, FZ18, GZW^{+}18, HS12, Her14, HTY^{+}19, HYC18, LM18a, LGH^{+}11, LWW19b, Luo18, MS10c, OSA13, PTZ19, PJ17, QTW^{+}18, RS18a, RRO17, Sea14, Sea16, THfL17, Tao18, Tia19, VDV13, WTYZ17, WO18, WL16, WTLS18, XXH18, YGS^{+}16, YTF10, YT11, ZYSY17, Zha18i, ZLY17]. wavy [MS10c]. way [HN10, Koj10, KH13, Tso13]. Wazewska [HGW11, WYN12]. Wazwaz [DR12]. Weak [BHM12, CS10c, EMRS12, FLZ14b, GR15, HMY18, JW15, SLM18, ZP18c, AaC19, ASV11, AET19, BSS18, CP15a, CWWY15, ĆSCD11, FH16, Füh18, GDZ11, HW19a, HX14, ID16, JMLF11, LXZ18, LNW19, Liu16c, MWY17, PLT17, RK10, SLK12, SNDK18, SR17c, TW18, WW14a, Wan16b, WHS18, YLG10, YWT18, YLLN16, ZYWZ17, Zha18a]. weak-strong [Liu16c]. weakened [VLFS12, YLLN16]. weaker [ZL14b]. Weakly [BGL^{+}15, Dem10b, HMF16, AKS11, APS12, BR16, FID14, FL13a, HLS11, HKI12, ML19b, QX19, Sea14, SS11c, SM17, ZCW15, ZM18]. wealth [WZG19]. wear [JO19]. weather [Li18d, TC16]. web [KRM^{+}10, BL11, WTC^{+}12]. Web-based [WTC^{+}12]. Weierstrass [AD10b]. Weight [HCL11, LY14, ÖŞ11, PSD^{+}13, Som13]. Weighted [AH10a, AT18a, CZN12, CDN14, Dra10, GDM13, NDT11, TMZ^{+}15, ZL11, AZ15, BMS12, ÇT11a, CLT^{+}13, CTP10, DÇ12, DS18b, DL18, ED11a, FJWW16, Fur13, HR19, HH15, LC10a, LY10a, LYLX11, LTX^{+}13, LX18, LCC12, LZ17, MF18b, MF18a, Mok11, MN10c, MKPS11, RSB14, hRWH18, Saj12, Sar10, Sha12b, SSP13, SYY13, Sun11, TS14, TÇ12, WC11b, WJWW12, WL13a, WS17, WL10, Yan11b, YYL11, gZnZpZbD12, ZL13]. weighting [CTZ17, GK11a, GK11b, LCW12, Sae11, SYO12]. weights [FM19, RR19]. Weiszfeld [KV10]. weldlines [ZOZZ12]. Well [AKV11, AS10c, BTB18, BT14, CCR16, Dua18, LX18, QY17, FAHZ17, FGL10, LL16c, MDBCF16, PZA19, Pu12, QaY18, RA19, Tha19b, Via15, WD16, YL16, YW19, YZ19]. Well-balanced [LX18, MDBCF16]. Well-conditioned [BT14]. Well-posedness [AKV11, AS10c, BTB18, CCR16, Dua18, QY17, FAHZ17, LL16c, PZA19, Pu12, QaY18, RA19, Tha19b, Via15, WD16, YZ19]. Wendroff [BKZ17]. WENO [FM19, HS18, RR18, RQ18, XW18, ZQ14a]. WENO-Z [XW18]. Werner [CM11b]. wheel [LSM10]. which [Boy16, CMGR11, Mah11]. while [MTV13, San11]. white [GB18, MX15, SM17, XSYL19, SS19]. Whitham [AO10a, GS11c, Lu11, SEY12]. Whitney [RR19]. whole [GM14b, LDG19, Tru19]. whose [CSS10, LY11a, Set12]. wide [Ser19]. wide-range [Ser19]. wideband [CC15]. width [Ray18, SIL19]. Wiener [CRRS11, LY17, LHH10, MP11c, QZ11]. Wigner [LY17, Lóp19]. Wilkinson [ÇA14]. Willmore [VĆC10, VĆV11]. WiMAX [CCK12, KKG10]. wind [DGGBTRJF12, GC19b]. winding [ZM14]. window [FKC12, IC12]. windows [RMK19, LLL13]. windup [Yan12a]. Wire [SISH12, DDK11, Gen11b]. wireless [CQLX11, CL12c, Chu12a, DA12, IC12, KSPP11, KCL12, MKA^{+}10, PHPK12, Sha12a, SLXC11, SCKH10, YXP^{+}13, YDK^{+}12, gZnZpZbD12]. wise [Dan12, OAKR16]. Wishart [YjH18]. within [CL12a, DSWB18, FT15, HMP^{+}15, KB10a, LPLR19, QKR19, WTSS10, YZMZ16]. within-host [YZMZ16]. without [AKV11, BL12, CDG15, CT17a, Col14, DG13b, GZN19, KYW^{+}18, Kim14, LvS15, MR19, NB17, PP12, QY17, She18b, TS16, WZ17b, Wu18a, YLC12, ZH11, ZZ15b]. WKI [Yu11b]. Wolfe [HSD10]. word [TNT12]. worker [MAPS10]. workflow [DRT^{+}15]. working [Ebr11]. workload [Chu12a, YW14]. workstations [Fia15]. world [gZnZpZbD12]. wormhole [LR17b]. wormlike [KL16a]. worms [SJS^{+}11]. WR [Yan17]. WR-HSS [Yan17]. Wright [ASMEE11a]. Wronskian [MGS^{+}14, YTS^{+}17]. WSGD [LZLL18]. WSGD-approximation [LZLL18]. WSNs [SRS11].

X [LY17, CX16]. X-ray [CX16]. xenobiotics [BDB12]. Xeon [CCHG17].

Yang [DT16, DD16, RWW18]. Yield [ĆCM10, HDS11]. YTSF [LD11b]. Yu [SR17a].

Z [XW18]. Zabolotskaya [KKK14]. Zadeh [gShYL10]. Zagreb [DT11]. Zakharov [GABC16, Her19, zLZ19, SR15, Che19a, DGA18, DS18a, GML17a, KK19c, Kuo18, MS18, PMA17, Ray18, Sea14, Sea16, Wan19a, YN16, Yu17, ZTZJ14, ZTSC16]. Zassenhaus [GTG11]. zeons [SS11a]. zero [AJ12, LMW10, Ma19, MCP13, WHC12]. zerodimensional [LMW10]. zero-pole [MCP13]. zero-tree [WHC12]. zeros [GS12, Ikh11, Ned12]. zeroth [Lóp19]. Zeta [CS10b, SJPS11, Cvi10]. Zhabotinskii [LV11]. ZK [Elb15, GZW^{+}18, RS18a, YYLW19, ZSAN18]. ZK-Burgers [YYLW19]. zone [GD16, LD19, Yan18a].

References

Ahookhosh:2010:NTR

[AA10a] Masoud Ahookhosh and Keyvan Amini. A nonmonotone trust region method with adaptive radius for unconstrained optimization problems. Computers and Mathematics with Applications, 60(3):411–422, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003056.

Allahverdi:2010:HTM

[AA10b] Ali Allahverdi and Harun Aydilek. Heuristics for the two-machine flowshop scheduling problem to minimize maximum lateness with bounded processing times. Computers and Mathematics with Applications, 60 (5):1374–1384, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004335.

Aydilek:2010:TMF

[AA10c] Harun Aydilek and Ali Allahverdi. Two-machine flowshop scheduling problem with bounded processing times to minimize total completion time. Computers and Mathematics with Applications, 59(2):684–693, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900724X.

Agarwal:2011:ETA

[AA11] Ravi P. Agarwal and Bashir Ahmad. Existence theory for antiperiodic boundary value problems of fractional differential equations and inclusions. Computers and Mathematics with Applications, 62(3):1200–1214, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001490.

Ashyralyev:2013:CDS

[AA13] Allaberen Ashyralyev and Deniz Agirseven. On convergence of difference schemes for delay parabolic equations. Computers and Mathematics with Applications, 66(7):1232–1244, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004483.

Al-Amr:2015:ESG

[AA15] Mohammed O. Al-Amr.

Exact solutions of the generalized (2 + 1)- dimensional nonlinear evolution equations via the modified simple equation method. Computers and Mathematics with Applications, 69(5):390– 397, March 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114006087.

See comment [LK15].

Alves:2018:MFS

[AA18] Carlos J. S. Alves and Pedro R. S. Antunes. The method of fundamental solutions applied to boundary value problems on the surface of a sphere. Computers and Mathematics with Applications, 75(7):2365–2373, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307782.

Ambrosio:2012:SCC

[AAA12] B. Ambrosio and M. A. Aziz-Alaoui. Synchronization and control of coupled reaction– diffusion systems of the FitzHugh–Nagumo type. Computers and Mathematics with Applications, 64(5):934–943, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000776.

Ahmad:2013:EPV

[AAB^{+}13] B. Ahmad, A. Alsaedi, F. Brezzi, L. D. Marini, and A. Russo. Equivalent projectors for virtual element methods. Computers and Mathematics with Applications, 66(3):376–391, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003179.

Al-arydah:2019:RMV

[AaC19] Mo’tassem Al-arydah and Thomas Carraro. Reviewing the mathematical validity of a fuel cell cathode model. Existence of weak bounded solution. Computers and Mathematics with Applications, 77(6):1425–1436, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304188.

Almeida:2017:FEM

[AAD17] Rui M. P. Almeida, Stanislav N. Antontsev, and José C. M. Duque.

On the finite element method for a nonlocal degenerate parabolic problem. Computers and Mathematics with Applications, 73(8):1724–1740, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300822.

Al-Amr:2017:NET

[AAEG17] Mohammed O. Al-Amr and Shoukry El-Ganaini. New exact traveling wave solutions of the (4 + 1)- dimensional Fokas equation. Computers and Mathematics with Applications, 74(6):1274–1287, September 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117303681.

Aleroev:2018:BVP

[AAH^{+}18] Temirkhan S. Aleroev, Hedi T. Aleroeva, Jianfei Huang, Mikhail V. Tamm, Yifa Tang, and Yue Zhao. Boundary value problems of fractional Fokker–Planck equations. Computers and Mathematics with Applications, 73(6):959– 969, March 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116303704.

Angelova:2012:PMP

[AAP12] Maria Angelova, Krassimir Atanassov, and Tania Pencheva. Purposeful model parameters genesis in simple genetic algorithms. Computers and Mathematics with Applications, 64(3):221–228, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000685.

Aamir:2011:LTC

[AAR11] Khalid Mahmood Aamir, Mujahid Abbas, and Stojan Radenović. A logarithmic time complexity algorithm for pattern searching using productsum property. Computers and Mathematics with Applications, 62 (5):2162–2168, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005554.

Agarwal:2010:IOC

[AAZ10] Ravi P. Agarwal, Douglas R. Anderson, and Agacik Zafer. Interval oscillation criteria for second-order forced delay dynamic equations with mixed nonlinearities. Computers and Mathematics with Applications, 59(2):977–993, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006749.

Abazari:2010:NSS

[AB10a] Reza Abazari and A. Borhanifar. Numerical study of the solution of the Burgers and coupled Burgers equations by a differential transformation method. Computers and Mathematics with Applications, 59(8):2711–2722, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000611.

Akyildiz:2010:BEC

[AB10b] F. Talay Akyildiz and Hamid Bellout. Brinkman equation for a corrugated pipe using a spectral-Galerkin method. Computers and Mathematics with Applications, 59(8): 2443–2451, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109004775.

Gunduz:2011:IFS

[AB11] Cigdem Gunduz (Aras)

and Sadi Bayramov. Intuitionistic fuzzy soft modules. Computers and Mathematics with Applications, 62(6):2480–2486, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005906.

Arakelyan:2016:NAG

[AB16] Avetik Arakelyan and Rafayel Barkhudaryan. A numerical approach for a general class of the spatial segregation of reaction– diffusion systems arising in population dynamics. Computers and Mathematics with Applications, 72(11):2823–2838, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630565X.

Afshar:2018:SEC

[AB18] Habib Afshar and Rasul Bagheri. Several embedded cracks in a functionally graded piezoelectric strip under dynamic loading. Computers and Mathematics with Applications, 76(3):534–550, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302499.

Abassy:2010:IAD

[Aba10a] Tamer A. Abassy. Improved Adomian decomposition method. Computers and Mathematics with Applications, 59(1): 42–54, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003885.

Abassy:2010:MVI

[Aba10b] Tamer A. Abassy. Modified variational iteration method (nonlinear homogeneous initial value problem). Computers and Mathematics with Applications, 59(2):912–918, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006919.

Angel-Bello:2011:HAS

[ÁBÁPM11] Francisco Ángel-Bello, Ada Álvarez, Joaquín Pacheco, and Iris Martínez. A heuristic approach for a scheduling problem with periodic maintenance and sequencedependent setup times. Computers and Mathematics with Applications, 61(4):797–808, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009338.

Abbas:2010:SFQ

[Abb10] S. E. Abbas. Stratified (L, M)-fuzzy quasiuniform spaces. Computers and Mathematics with Applications, 59(8):3014– 3026, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211000129X.

Abbas:2011:PAA

[Abb11] Syed Abbas. Pseudo almost automorphic solutions of some nonlinear integro-differential equations. Computers and Mathematics with Applications, 62(5):2259–2272, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005670.

Acosta:2017:SFI

[ABB17] Gabriel Acosta, Francisco M. Bersetche, and Juan Pablo Borthagaray. A short FE implementation for a 2d homogeneous Dirichlet problem of a fractional Laplacian. Computers and Mathematics with Applications, 74(4):784–816, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303310.

Alimohammady:2010:NPF

[ABCR10] M. Alimohammady, J. Balooee, Y. J. Cho, and M. Roohi. New perturbed finite step iterative algorithms for a system of extended generalized nonlinear mixed quasi-variational inclusions. Computers and Mathematics with Applications, 60(11):2953– 2970, December 2010.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110007522.

Abdeljawad:2011:RCF

[Abd11] Thabet Abdeljawad. On Riemann and Caputo fractional differences. Computers and Mathematics with Applications, 62(3):1602–1611, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100188X.

AbderramanMarrero:2018:RGA

[Abd18a] J. Abderramán Marrero. A reliable Givens– LU approach for solving opposite-bordered tridiagonal linear systems. Computers and Mathematics with Applications, 76(10):2409–2420, November 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118304644.

Abdulwahhab:2018:CPC

[Abd18b] Muhammad Alim Abdulwahhab. Comment on the paper “On conservation laws by Lie symmetry analysis for (2 + 1)- dimensional Bogoyavlensky– Konopelchenko equation in wave propagation” by S. Saha Ray. Computers and Mathematics with Applications, 75(12):4300–4304, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301664.

Al-Begain:2012:GSA

[ABDKD12] Khalid Al-Begain, Alexander Dudin, Valentina Klimenok, and Sergey Dudin. Generalized survivability analysis of systems with propagated failures. Computers and Mathematics with Applications, 64 (12):3777–3791, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001782.

Alcalde:2011:ULV

[ABFGZ11] C. Alcalde, A. Burusco, R. Fuentes-González, and I. Zubia. The use of linguistic variables and fuzzy propositions in the L-fuzzy concept theory.

Computers and Mathematics with Applications, 62(8):3111–3122, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006870.

Almeida:2012:MAC

[ABH12] Luís Almeida, Patrizia Bagnerini, and Abderrahmane Habbal. Modeling actin cable contraction. Computers and Mathematics with Applications, 64(3):310–321, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001666.

Aghajani:2011:ESC

[ABJ11] Asadollah Aghajani, Józef Banaś, and Yaghoub Jalilian. Existence of solutions for a class of nonlinear Volterra singular integral equations. Computers and Mathematics with Applications, 62(3):1215–1227, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111002124.

Asadzadeh:2010:ITI

[ABK10] M. Asadzadeh, A. Brahme, and J. Kempe. Ion transport in inhomogeneous media based on the bipartition model for primary ions. Computers and Mathematics with Applications, 60(8):2445–2459, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006140.

Axelsson:2013:NCE

[ABK^{+}13] O. Axelsson, P. Boyanova, M. Kronbichler, M. Neytcheva, and X. Wu. Numerical and computational efficiency of solvers for two-phase problems. Computers and Mathematics with Applications, 65(3):301–314, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004191.

Amoddeo:2010:MMP

[ABL10] A. Amoddeo, R. Barberi, and G. Lombardo. Moving mesh partial differential equations to describe nematic order dynamics. Computers and Mathematics with Applications, 60(8):2239–2252, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000578X.

Abreu:2015:NMT

[ABL15] Eduardo Abreu, Abel Bustos, and Wanderson Lambert. Non-monotonic traveling wave and computational solutions for gas dynamics Euler equations with stiff relaxation source terms. Computers and Mathematics with Applications, 70 (9):2155–2176, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003375.

Anaya:2015:CFV

[ABLS15] Verónica Anaya, Mostafa Bendahmane, Michel Langlais, and Mauricio Sepúlveda. A convergent finite volume method for a model of indirectly transmitted diseases with nonlocal cross-diffusion. Computers and Mathematics with Applications, 70(2):132– 157, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002023.

Abbas:2011:DAF

[ABM11] Syed Abbas, Malay Banerjee, and Shaher Momani. Dynamical analysis of fractional-order modified logistic model. Computers and Mathematics with Applications, 62(3):1098–1104, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002355.

Atabaigi:2018:BAE

[ABN18] Ali Atabaigi, Ali Barati, and Hamed Norouzi. Bifurcation analysis of an enzyme-catalyzed reaction– diffusion system. Computers and Mathematics with Applications, 75(12):4361–4377, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301718.

Ali:2010:MFR

[ABR10] Rosihan M. Ali, Abeer O.

Badghaish, and V. Ravichandran. Multivalent functions with respect to n-ply points and symmetric conjugate points. Computers and Mathematics with Applications, 60(11):2926–2935, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007509.

Anderl:2014:FSL

[ABR^{+}14] Daniela Anderl, Simon Bogner, Cornelia Rauh, Ulrich Rüde, and Antonio Delgado. Free surface lattice Boltzmann with enhanced bubble model. Computers and Mathematics with Applications, 67(2):331–339, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003726.

Adjerid:2018:HDI

[ABRL18] Slimane Adjerid, Mohamed Ben-Romdhane, and Tao Lin. Higher degree immersed finite element spaces constructed according to the actual interface. Computers and Mathematics with Applications, 75(6):1868–1881, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306582.

Antonietti:2018:VEM

[ABSV18] P. F. Antonietti, M. Bruggi, S. Scacchi, and M. Verani. On the virtual element method for topology optimization on polygonal meshes: A numerical study. Computers and Mathematics with Applications, 74 (5):1091–1109, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303309.

Abgrall:2019:HOR

[ABT19] Rémi Abgrall, Paola Bacigaluppi, and Svetlana Tokareva. Highorder residual distribution scheme for the timedependent Euler equations of fluid dynamics. Computers and Mathematics with Applications, 78(2):274–297, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302712.

Alloui:2011:DDS

[ABV11] Z. Alloui, H. Beji, and P. Vasseur. Doublediffusive and Soret-induced convection of a micropolar fluid in a vertical channel. Computers and Mathematics with Applications, 62(2):725– 736, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004640.

Andalaft-Chacur:2011:ROP

[ACAS11] A. Andalaft-Chacur, M. Montaz Ali, and J. González Salazar. Real options pricing by the finite element method. Computers and Mathematics with Applications, 61(9):2863– 2873, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002331.

Aulisa:2018:IMA

[ACC18] Eugenio Aulisa, Sara Calandrini, and Giacomo Capodaglio. An improved multigrid algorithm for n-irregular meshes with subspace correction smoother. Computers and Mathematics with Applications, 76(3):620–632, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302633.

Allili:2011:DCR

[ACD^{+}11] Madjid Allili, David Corriveau, Sara Derivière, Marc Ethier, and Tomasz Kaczynski. Detecting critical regions in multidimensional data sets. Computers and Mathematics with Applications, 61(2):499–512, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009016.

Araujo-Cabarcas:2017:SSF

[ACE17] Juan Carlos Araujo-Cabarcas and Christian Engström. On spurious solutions in finite element approximations of resonances in open systems. Computers and Mathematics with Applications, 74 (10):2385–2402, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304376.

Aktas:2011:QBS

[AÇT11] Mustafa Fahri Aktas, Devrim Çakmak, and Aydin Tiryaki. On the qualitative behaviors of solutions of third order nonlinear differential equations. Computers and Mathematics with Applications, 62(4):2029–2036, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005256.

Au:2019:RBP

[ACTB19] Vo Van Au, Nguyen Huu Can, Nguyen Huy Tuan, and Tran Thanh Binh. Regularization of a backward problem for a Lotka–Volterra competition system. Computers and Mathematics with Applications, 78(3):765–785, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301191.

Agarwal:2010:SJT

[AD10a] R. P. Agarwal and S. S. Dragomir. A survey of Jensen type inequalities for functions of selfadjoint operators in Hilbert spaces. Computers and Mathematics with Applications, 59(12):3785– 3812, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002853.

See corrigendum [AD11a].

Anastassiou:2010:SAD

[AD10b] George A. Anastassiou and Oktay Duman. Statistical L_{p}-approximation by double Gauss–Weierstrass singular integral operators. Computers and Mathematics with Applications, 59(6):1985–1999, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007482.

Agarwal:2011:CSJ

[AD11a] R. P. Agarwal and S. S. Dragomir. Corrigendum to “A survey of Jensen type inequalities for functions of selfadjoint operators in Hilbert spaces” [Comput. Math. Appl. 59 (2010) 3785–3812]. Computers and Mathematics with Applications, 61(9): 2931, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001714.

See [AD10a].

Akram:2011:IVF

[AD11b] Muhammad Akram and Wieslaw A. Dudek. Intervalvalued fuzzy graphs. Computers and Mathematics with Applications, 61(2):289–299, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008539.

Ameri:2011:FHS

[AD11c] R. Ameri and O. Dehghan. Fuzzy hypervector spaces based on fuzzy singletons. Computers and Mathematics with Applications, 61(10):2933– 2943, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111001416.

Aphirukmatakun:2011:MDE

[AD11d] Chanon Aphirukmatakun and Natasha Dejdumrong. Multiple degree elevation and constrained multiple degree reduction for DP curves and surfaces. Computers and Mathematics with Applications, 61(8):2296–2299, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007492.

Appleby:2012:NSC

[AD12] John A. D. Appleby and John A. Daniels. Necessary and sufficient conditions for periodic decaying resolvents in linear discrete convolution Volterra equations and applications to ARCH(∞) processes. Computers and Mathematics with Applications, 64(7):2312–2325, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002842.

Anguelov:2014:B

[AD14] Roumen Anguelov and Yves Dumont. BIOMATH 2013. Computers and Mathematics with Applications, 68(9):903–904, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002739.

Abbaszadeh:2015:MNP

[AD15] Mostafa Abbaszadeh and Mehdi Dehghan. A meshless numerical procedure for solving fractional reaction subdiffusion model via a new combination of alternating direction implicit (ADI) approach and interpolating element free Galerkin (EFG) method. Computers and Mathematics with Applications, 70(10):2493–2512, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004411.

Alnashri:2016:GSS

[AD16] Yahya Alnashri and Jérôme Droniou. Gradient schemes for the Signorini and the obstacle problems, and application to hybrid mimetic mixed methods. Computers and Mathematics with Applications, 72 (11):2788–2807, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305600.

Abbaszadeh:2019:AMF

[AD19a] Mostafa Abbaszadeh and Mehdi Dehghan. Analysis of mixed finite element method (MFEM) for solving the generalized fractional reaction–diffusion equation on nonrectangular domains. Computers and Mathematics with Applications, 78 (5):1531–1547, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301658.

Arya:2019:EGS

[AD19b] Nitish Arya and Ashoke De. Effect of grid sensitivity on the performance of wall adapting SGS models for LES of swirling and separatingreattaching flows. Computers and Mathematics with Applications, 78 (6):2035–2051, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301634.

Abdullah:2011:IFI

[ADA11] Saleem Abdullah, Bijan Davvaz, and Muhammad Aslam. (α, β)- intuitionistic fuzzy ideals of hemirings. Computers and Mathematics with Applications, 62(8):3077–3090, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006845.

Agarwal:2011:FID

[AdAS11] Ravi P. Agarwal, Bruno de Andrade, and Giovana Siracusa. On fractional integro-differential equations with statedependent delay. Computers and Mathematics with Applications, 62(3):1143–1149, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001192.

Aprovitola:2015:SEL

[ADD^{+}15] Andrea Aprovitola, Pasqua D’Ambra, Filippo M. Denaro, Daniela di Serafino, and Salvatore Filippone. SParC–LES: Enabling large eddy simulations with parallel sparse matrix computation tools. Computers and Mathematics with Applications, 70 (11):2688–2700, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003144.

Adem:2016:GDD

[Ade16] Abdullahi Rashid Adem. The generalized (1 + 1)- dimensional and (2 + 1)-dimensional Itô equations: Multiple Expfunction algorithm and multiple wave solutions. Computers and Mathematics with Applications, 71(6):1248–1258, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300475.

Adem:2017:SCE

[Ade17] Abdullahi Rashid Adem. Symbolic computation on exact solutions of a coupled Kadomtsev– Petviashvili equation: Lie symmetry analysis and extended tanh method. Computers and Mathematics with Applications, 74(8):1897–1902, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304078.

Augier:2013:LLB

[ADGG13] Adeline Augier, François Dubois, Loïc Gouarin, and Benjamin Graille. Linear lattice Boltzmann schemes for acoustic: Parameter choices and isotropy properties. Computers and Mathematics with Applications, 65(6):845–863, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004890.

Augier:2014:RIL

[ADGL14] Adeline Augier, François Dubois, Benjamin Graille, and Pierre Lallemand. On rotational invariance of lattice Boltzmann schemes. Computers and Mathematics with Applications, 67(2):239–255, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300374X.

Abreu:2018:HOC

[ADGS18] Eduardo Abreu, Ciro Díaz, Juan Galvis, and Marcus Sarkis. On high-order conservative finite element methods. Computers and Mathematics with Applications, 75(6):1852–1867, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306685.

Alomari:2010:RHT

[ADK10] M. Alomari, M. Darus, and U. S. Kirmaci. Refinements of Hadamardtype inequalities for quasiconvex functions with applications to trapezoidal formula and to special means. Computers and Mathematics with Applications, 59(1):225–232, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S089812210900532X.

Anguelov:2012:MMS

[ADL12] Roumen Anguelov, Yves Dumont, and Jean Lubuma. Mathematical modeling of sterile insect technology for control of anopheles mosquito. Computers and Mathematics with Applications, 64(3):374– 389, August 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112001939.

Apreutesei:2014:OCP

[ADS14] N. Apreutesei, G. Dimitriu, and R. Strugariu. An optimal control problem for a two-prey and one-predator model with diffusion. Computers and Mathematics with Applications, 67(12):2127– 2143, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000911.

Alonso:2019:TOB

[AdSSS19] Diego Hayashi Alonso, Luís Fernando Nogueira de Sá, Juan Sergio Romero Saenz, and Emílio Carlos Nelli Silva. Topology optimization based on a two-dimensional swirl flow model of Tesla-type pump devices. Computers and Mathematics with Applications, 77(9):2499–2533, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307338.

Ackleh:2012:SAS

[ADY12] Azmy S. Ackleh, Keng Deng, and Xing Yang. Sensitivity analysis for a structured juvenile-adult model. Computers and Mathematics with Applications, 64(3):190–200, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011096.

Abbaszadeh:2019:ADI

[ADZ19]  Mostafa  Abbaszadeh,

Mehdi Dehghan, and Yong Zhou. Alternating direction implicitspectral element method (ADI–SEM) for solving multi-dimensional generalized modified anomalous sub-diffusion equation. Computers and Mathematics with Applications, 78(5):1772–1792, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303281.

Abkar:2012:FPC

[AE12a] Ali Abkar and Mohammad Eslamian. Fixed point and convergence theorems for different classes of generalized nonexpansive mappings in CAT(0) spaces. Computers and Mathematics with Applications, 64(4):643– 650, August 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111011461.

Atici:2012:GID

[AE12b] Ferhan M. Atici and Paul W. Eloe. Gronwall’s inequality on discrete fractional calculus. Computers and Mathematics with Applications, 64(10):3193–3200, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010091.

Aloev:2014:SCS

[AEDL14] R. D. Aloev, Z. K. Eshkuvatov, Sh. O. Davlatov, and N. M. A. Nik Long. Sufficient condition of stability of finite element method for symmetric t-hyperbolic systems with constant coefficients. Computers and Mathematics with Applications, 68(10):1194–1204, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114004118.

Alexandrov:2015:TMC

[AEF15] Vassil Alexandrov and Oscar A. Esquivel-Flores. Towards Monte Carlo preconditioning approach and hybrid Monte Carlo algorithms for matrix computations. Computers and Mathematics with Applications, 70 (11):2709–2718, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004149.

Alwan:2011:CEO

[AEG11] Saba Alwan and Awad El-Gohary. Chaos, estimation and optimal control of habitat destruction model with uncertain parameters. Computers and Mathematics with Applications, 62 (11):4089–4099, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008364.

Abdrabou:2018:SGM

[AEG18] Amgad Abdrabou and Mohamed El-Gamel. On the sinc-Galerkin method for triharmonic boundaryvalue problems. Computers and Mathematics with Applications, 76(3):520–533, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302487.

Aboelenen:2018:HON

[AEH18] Tarek Aboelenen and H. M. El-Hawary. A high-order nodal discontinuous Galerkin method for a linearized fractional Cahn–Hilliard equation. Computers and Mathematics with Applications, 73(6):1197–1217, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304102.

Abdaoui:2019:AEB

[AEH19] I. Abdaoui, L. Elbouyahyaoui, and M. Heyouni. An alternative extended block Arnoldi method for solving low-rank Sylvester equations. Computers and Mathematics with Applications, 78 (8):2817–2830, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302445.

Allaoui:2015:ERC

[AEO15] Mostafa Allaoui, Abdelrachid El Amrouss, and Anass Ourraoui. Existence results for a class of p(x)-Laplacian problems in R^{N}. Computers and Mathematics with Applications, 69(7):582– 591, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000516.

Aly:2012:AAD

[AER12] Emad H. Aly, Abdelhalim Ebaid, and Randolph Rach. Advances in the Adomian decomposition method for solving twopoint nonlinear boundary value problems with Neumann boundary conditions. Computers and Mathematics with Applications, 63(6):1056–1065, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010534.

Akhmedov:2011:FSO

[AES11] A. M. Akhmedov and S. R. El-Shabrawy. On the fine spectrum of the operator Δ_{a,b} over the sequence space c. Computers and Mathematics with Applications, 61(10):2994–3002, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002641.

Ayouch:2019:GWS

[AET19] Chahid Ayouch, El-Hassan Essoufi, and Mouhcine Tilioua. Global weak solutions to a spatio-temporal fractional Landau–Lifshitz–Bloch equation. Computers and Mathematics with Applications, 77(5):1347–1357, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306692.

Ackleh:2013:NRR

[AF13] Azmy S. Ackleh and József Z. Farkas. On the net reproduction rate of continuous structured populations with distributed states at birth. Computers and Mathematics with Applications, 66(9):1685–1694, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002113.

Agudo:2010:SBT

[AFGL10] Isaac Agudo, Carmen Fernandez-Gago, and Javier Lopez. A scale based trust model for multi-context environments. Computers and Mathematics with Applications, 60(2):209–216, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122110001185.

Ali:2010:ELM

[AG10a] M. M. Ali and M. Golalikhani. An electromagnetismlike method for nonlinearly constrained global optimization. Computers and Mathematics with Applications, 60(8):2279–2285, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005821.

Ashyralyev:2010:FDM

[AG10b] Allaberen Ashyralyev and Okan Gercek. Finite difference method for multipoint nonlocal ellipticparabolic problems. Computers and Mathematics with Applications, 60(7):2043–2052, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005365.

Aminossadati:2011:ENC

[AG11] S. M. Aminossadati and B. Ghasemi. Enhanced natural convection in an isosceles triangular enclosure filled with a nanofluid. Computers and Mathematics with Applications, 61(7):1739–1753, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000873.

Arunachalam:2010:FQM

[AGD10] Viswanathan Arunachalam, Vandana Gupta, and S. Dharmaraja. A fluid queue modulated by two independent birth-death processes. Computers and Mathematics with Applications, 60(8):2433–2444, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006036.

Allaire:2019:TOM

[AGDP19] Grégoire Allaire, Perle Geoffroy-Donders, and Olivier Pantz. Topology optimization of modulated and oriented periodic microstructures by the homogenization method. Computers and Mathematics with Applications, 78(7):2197–2229, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304255.

Ayadi:2015:IMP

[AGH^{+}15] Mekki Ayadi, Asma Gdhami, Abderrahmane Habbal, Maroua Mokni, and Boutheina Yahyaoui. Improving the mechanical performances of a multilayered plate with the orientations of its layers of fibers. Computers and Mathematics with Applications, 70(8):1817–1829, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003776.

Atanassov:2015:EAP

[AGK15] E. Atanassov, T. Gurov, and A. Karaivanova. Energy aware performance study for a class of computationally intensive Monte Carlo algorithms. Computers and Mathematics with Applications, 70(11):2719–2725, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003491.

Arenas:2010:NNS

[AGPCC10] Abraham J. Arenas, Gilberto González-Parra, and Benito M. Chen-Charpentier. A nonstandard numerical scheme of predictor–corrector type for epidemic models. Computers and Mathematics with Applications, 59(12):3740–3749, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002671.

Arraras:2019:GMM

[AGPR19] A. Arrarás, F. J. Gaspar, L. Portero, and C. Rodrigo. Geometric multigrid methods for Darcy–Forchheimer flow in fractured porous media. Computers and Mathematics with Applications, 78(9):3139–3151, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302470.

Agrawal:2010:GVP

[Agr10] Om Prakash Agrawal. Generalized variational problems and Euler– Lagrange equations. Computers and Mathematics with Applications, 59(5): 1852–1864, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812210900563X.

Abdel-Gawad:2019:TLF

[AGT19] H. I. Abdel-Gawad and M. Tantawy. Two-layer fluid formation and propagation of periodic solitons induced by (3 + 1)- dimensional KP equation. Computers and Mathematics with Applications, 78(6):2011–2017, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301567.

Anguelov:2014:BBA

[AGU14] Roumen Anguelov, Salisu M. Garba, and Salisu Usaini. Backward bifurcation analysis of epidemiological model with partial immunity. Computers and Mathematics with Applications, 68(9):931–940, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122114002715.

Abbasbandy:2010:WTA

[AH10a] S. Abbasbandy and T. Hajjari. Weighted trapezoidal approximationpreserving cores of a fuzzy number. Computers and Mathematics with Applications, 59(9):3066–3077, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001355.

Ahmad:2010:MMS

[AH10b] I. Ahmad and Z. Husain. Multiobjective mixed symmetric duality involving cones. Computers and Mathematics with Applications, 59(1):319–326, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109004799.

Aouchiche:2010:NGR

[AH10c] M. Aouchiche and P. Hansen. Nordhaus–Gaddum relations for proximity and remoteness in graphs. Computers and Mathematics with Applications, 59(8):2827–2835, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000921.

Al-Hayani:2011:ADM

[AH11a] Waleed Al-Hayani. Adomian decomposition method with Green’s function for sixth-order boundary value problems. Computers and Mathematics with Applications, 61(6):1567– 1575, March 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111000423.

Amini-Harandi:2011:FPT

[AH11b] A. Amini-Harandi. Fixed point theory for generalized quasicontraction maps in vector modular spaces. Computers and Mathematics with Applications, 61(7):1891–1897, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001076.

Amini-Harandi:2010:FPT

[AHF10] A. Amini-Harandi and M. Fakhar. Fixed point theory in cone metric spaces obtained via the scalarization method. Computers and Mathematics with Applications, 59(11):3529–3534, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000235X.

Abtahi:2016:SAA

[AHF16] A. Abtahi, M. Z. Hossain, and J. M. Floryan. Spectrally accurate algorithm for analysis of convection in corrugated conduits. Computers and Mathematics with Applications, 72(10):2636–2659, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305314.

Adler:2019:VPF

[AHHM19] James H. Adler, Yunhui He, Xiaozhe Hu, and Scott P. MacLachlan. Vector-potential finiteelement formulations for two-dimensional resistive magnetohydrodynamics. Computers and Mathematics with Applications, 77(2):476–493, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305716.

Angstmann:2018:ITF

[AHJM18] C. N. Angstmann, B. I.

Henry, B. A. Jacobs, and A. V. McGann. Integrablization of time fractional PDEs. Computers and Mathematics with Applications, 73(6):1053–1062, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306812.

Ahmed:2010:CFP

[Ahm10] M. A. Ahmed. Common fixed points of hybrid maps and an application. Computers and Mathematics with Applications, 60(7):1888–1894, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004980.

Ahn:2010:ACS

[Ahn10] Young Joon Ahn. Approximation of conic sections by curvature continuous quartic Bézier curves. Computers and Mathematics with Applications, 60(7):1986–1993, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005249.

Ahn:2012:ESL

[Ahn12] Choon Ki Ahn. Exponential H_{∞} stable learning method for Takagi– Sugeno fuzzy delayed neural networks: a convex optimization approach. Computers and Mathematics with Applications, 63(5):887–895, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010376.

Allendes:2016:RNM

[AHO16] Alejandro Allendes, Erwin Hernández, and Enrique Otárola. A robust numerical method for a control problem involving singularly perturbed equations. Computers and Mathematics with Applications, 72(4):974–991, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630342X.

Angstmann:2018:GME

[AHOP18] C. N. Angstmann, B. I.

Henry, and I. Ortega-Piwonka. Generalized master equations and fractional Fokker–Planck equations from continuous time random walks with arbitrary initial conditions. Computers and Mathematics with Applications, 73(6):1315–1324, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306320.

Abert:2014:SPT

[AHP^{+}14] Claas Abert, Gino Hrkac, Marcus Page, Dirk Praetorius, Michele Ruggeri, and Dieter Suess. Spinpolarized transport in ferromagnetic multilayers: an unconditionally convergent FEM integrator. Computers and Mathematics with Applications, 68(6):639–654, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122114003162.

Adams:2010:HNE

[AHV10] Jay L. Adams, Tom T. Hartley, and Robert J. Veillette. Hankel-norm estimation for fractionalorder systems using the Rayleigh–Ritz method. Computers and Mathematics with Applications, 59(5):1773–1781, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005550.

Abdollahi:2013:ICA

[AIA13] Mahdi Abdollahi, Ayaz Isazadeh, and Davoud Abdollahi. Imperialist competitive algorithm for solving systems of nonlinear equations. Computers and Mathematics with Applications, 65(12):1894–1908, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002241.

Ali:2010:CFP

[AIB10] Javid Ali, M. Imdad, and D. Bahuguna. Common fixed point theorems in Menger spaces with common property (e.a). Computers and Mathematics with Applications, 60(12):3152–3159, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007996.

Ali:2010:SMB

[AIIZ10] Javed Ali, S. Islam, Sirajul Islam, and Gul Zaman. The solution of multipoint boundary value problems by the optimal homotopy asymptotic method. Computers and Mathematics with Applications, 59(6):2000–2006, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007494.

Almenar:2010:NBS

[AJ10] Pedro Almenar and Lucas Jódar. New bounds for the solutions of second order linear differential equations. Computers and Mathematics with Applications, 59(1):468– 485, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003903.

Almenar:2011:ABS

[AJ11] Pedro Almenar and Lucas Jódar. Asymptotic behaviour of the solutions of second order functional differential equations. Computers and Mathematics with Applications, 62(1):297–309, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003981.

Almenar:2012:UBD

[AJ12] Pedro Almenar and Lucas Jódar. An upper bound for the distance between a zero and a critical point of a solution of a second order linear differential equation. Computers and Mathematics with Applications, 63(1):310–317, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010030.

Al-Jawary:2018:ANS

[AJAR18] Majeed Ahmed Al-Jawary, Mustafa Mahmood Azeez, and Ghassan Hasan Radhi.

Analytical and numerical solutions for the nonlinear Burgers and advection– diffusion equations by using a semi-analytical iterative method. Computers and Mathematics with Applications, 76(1):155– 171, July 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118302086.

Al-Jaroodi:2010:SMA

[AJJAD^{+}10] Jameela Al-Jaroodi, Imad Jawhar, Alyaziyah Al-Dhaheri, Fatmah Al-Abdouli, and Nader Mohamed. Security middleware approaches and issues for ubiquitous applications. Computers and Mathematics with Applications, 60(2):187– 197, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000209.

Al-Jawary:2012:BEF

[AJRWS12] M. A. Al-Jawary, J. Ravnik, L. C. Wrobel, and L. Skerget. Boundary element formulations for the numerical solution of two-dimensional diffusion problems with variable coefficients. Computers and Mathematics with Applications, 64(8):2695–2711, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005135.

Azwadi:2014:EMC

[AJS14] Che Sidik Nor Azwadi, Leila Jahanshaloo, and Arman Safdari. The effect of mixed convection on particle laden flow analysis in a cavity using a lattice Boltzmann method. Computers and Mathematics with Applications, 67(1):52–61, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006408.

Agarwal:2019:NGS

[AJS19] Ravi P. Agarwal, Mohamed Jleli, and Bessem Samet. Nonexistence of global solutions for a time-fractional damped wave equation in a ktimes halved space. Computers and Mathematics with Applications, 78 (5):1608–1620, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300380.

Ammi:2019:FEA

[AJT19] Moulay Rchid Sidi Ammi, Ismail Jamiai, and Delfim F. M. Torres. A finite element approximation for a class of Caputo timefractional diffusion equations. Computers and Mathematics with Applications, 78(5):1334–1344, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302949.

Akinfenwa:2013:CBB

[AJY13] O. A. Akinfenwa, S. N.

Jator, and N. M. Yao. Continuous block backward differentiation formula for solving stiff ordinary differential equations. Computers and Mathematics with Applications, 65(7):996–1005, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000564.

Agarwal:2010:SOI

[AK10] Ravi P. Agarwal and Fatma Karakoç. A survey on oscillation of impulsive delay differential equations. Computers and Mathematics with Applications, 60(6):1648–1685, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000461X.

Anh:2011:OAA

[AK11] Pham Ngoc Anh and Jong Kyu Kim. Outer approximation algorithms for pseudomonotone equilibrium problems. Computers and Mathematics with Applications, 61(9): 2588–2595, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001532.

Ahmad:2012:MBA

[AK12] Farooq Ahmad and Sher Afzal Khan. Module-based architecture for a periodic job-shop scheduling problem. Computers and Mathematics with Applications, 64(1):1–10, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001459.

Adem:2016:CQS

[AK16] Abdullahi Rashid Adem and Chaudry Masood Khalique. Conserved quantities and solutions of a (2 + 1)-dimensional Haragus–Courcelle–Il’ichev model. Computers and Mathematics with Applications, 71(5):1129–1136, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300177.

Check title: accents lost! Akbulut:2018:AEM

[AK18] Arzu Akbulut and Melike Kaplan. Auxiliary equation method for time-fractional differential equations with conformable derivative. Computers and Mathematics with Applications, 75(3):876–882, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306648.

Alzahrani:2019:HOT

[AK19] S. S. Alzahrani and A. Q. M. Khaliq. Highorder time stepping Fourier spectral method for multidimensional space-fractional reaction–diffusion equations. Computers and Mathematics with Applications, 77(3):615–630, February 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118305819.

Ashyralyev:2011:ODS

[AKA11] Allaberen Ashyralyev, Mehmet Emir Koksal, and Ravi P. Agarwal. An operator-difference scheme for abstract Cauchy problems. Computers and Mathematics with Applications, 61(7):1855–1872, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001003.

Akimenko:2017:ASS

[Aki17] Vitalii Akimenko. An age-structured SIR epidemic model with fixed incubation period of infection. Computers and Mathematics with Applications, 73(7):1485–1504, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300524.

Alsaedi:2018:GEA

[AKL18a] Ahmed Alsaedi, Mokhtar Kirane, and Rafika Lassoued. Global existence and asymptotic behavior for a time fractional reaction–diffusion system. Computers and Mathematics with Applications, 73(6):951–958, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302644.

Aulisa:2018:AMR

[AKL18b] E. Aulisa, G. Ke, and S.-Y. Lee. An adaptive mesh refinement strategy for finite element solution of the elliptic problem. Computers and Mathematics with Applications, 76(2):224–244, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302098.

Andrews:2019:MSE

[AKLS19] Kevin T. Andrews, Kenneth L. Kuttler, Ji Li, and Meir Shillor. Measurable solutions for elliptic inclusions and quasistatic problems. Computers and Mathematics with Applications, 77(11):2869– 2882, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305376.

Akman:2015:LIR

[Akm15] Tugba Akman. Local improvements to reducedorder approximations of optimal control problems governed by diffusionconvection-reaction equation. Computers and Mathematics with Applications, 70(2):104–131, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001984.

Alrawashdeh:2018:AIT

[AKMS18] Mahmoud S. Alrawashdeh, James F. Kelly, Mark M. Meerschaert, and Hans-Peter Scheffler. Applications of inverse tempered stable subordinators. Computers and Mathematics with Applications, 73(6):892–905, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630428X.

Ayub:2017:SPW

[AKMUH17] Kamran Ayub, M. Yaqub Khan, and Qazi Mahmood-Ul-Hassan. Solitary and periodic wave solutions of Calogero– Bogoyavlenskii–Schiff equation via exp-function methods. Computers and Mathematics with Applications, 74 (12):3231–3241, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305138.

Anguraj:2014:NER

[AKRT14] A. Anguraj, P. Karthikeyan, M. Rivero, and J. J. Trujillo. On new existence results for fractional integro-differential equations with impulsive and integral conditions. Computers and Mathematics with Applications, 66(12):2587–2594, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000643.

Al-Kharsani:2010:SRH

[AKS10] H. A. Al-Kharsani and A. Sofo. Subordination results on harmonic kuniformly convex mappings and related classes. Computers and Mathematics with Applications, 59(12):3718–3726, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002658.

Aydi:2011:CFP

[AKS11] Hassen Aydi, Erdal Karapınar, and Wasfi Shatanawi. Coupled fixed point results for (ψ, ϕ)-weakly contractive condition in ordered partial metric spaces. Computers and Mathematics with Applications, 62 (12):4449–4460, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008856.

Almani:2019:SME

[AKSW19] T. Almani, K. Kumar, G. Singh, and M. F.

Wheeler. Stability of multirate explicit coupling of geomechanics with flow in a poroelastic medium. Computers and Mathematics with Applications, 78(8):2682–2699, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302056.

Acar:2010:SSS

[AKT10] Ummahan Acar, Fatih Koyuncu, and Bekir Tanay. Soft sets and soft rings. Computers and Mathematics with Applications, 59(11):3458– 3463, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002233.

Abdeljawad:2012:GCP

[AKT12] Thabet Abdeljawad, Erdal Karapinar, and Kenan Tas. A generalized contraction principle with control functions on partial metric spaces. Computers and Mathematics with Applications, 63(3):716–719, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010157.

Anh:2011:WPS

[AKV11] Lam Quoc Anh, Phan Quoc Khanh, and Dang Thi My Van. Well-posedness without semicontinuity for parametric quasiequilibria and quasioptimization. Computers and Mathematics with Applications, 62(4):2045–2057, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100527X.

Alasooly:2010:CBO

[Ala10] Hedaya Alasooly. Comparison between optimal control strategies applied to a system with unified power flow controller, shunt converter and series converter. Computers and Mathematics with Applications, 60(4):954–975, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001616.

Aguirre-Lopez:2019:SFA

[ALHZMC^{+}19] Mario A. Aguirre-López, Filiberto Hueyotl-Zahuantitla, Javier Morales-Castillo, Gerardo J. Escalera Santos, and F.-Javier Almaguer. Simulating the flow around a baseball: Study of a 2Dcylinder with a single bump. Computers and Mathematics with Applications, 78(9):3105–3116, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302500.

Ali:2011:SIS

[Ali11a] Muhammad Irfan Ali.

Soft ideals and soft filters of soft ordered semigroups. Computers and Mathematics with Applications, 62(9):3396–3403, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007279.

Azzam-Laouir:2011:ESS

[ALI11b] Dalila Azzam-Laouir and Sabrina Izza. Existence of solutions for second-order perturbed nonconvex sweeping process. Computers and Mathematics with Applications, 62(4):1736–1744, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004834.

Alici:2015:HPM

[Ali15] H. Alici. The Hermite pseudospectral method for the two-dimensional Schrödinger equation with nonseparable potentials. Computers and Mathematics with Applications, 69(6):466–476, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000048.

Alham:2011:MBD

[ALLH11] Nasullah Khalid Alham, Maozhen Li, Yang Liu, and Suhel Hammoud. A MapReduce-based distributed SVM algorithm for automatic image annotation. Computers and Mathematics with Applications, 62(7):2801–2811, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006171.

Alham:2013:MBD

[ALLQ13] Nasullah Khalid Alham, Maozhen Li, Yang Liu, and Man Qi. A MapReduce-based distributed SVM ensemble for scalable image classification and annotation. Computers and Mathematics with Applications, 66(10):1920–1934, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004434.

Angulo:2014:AEI

[ALMLM14] O. Angulo, J. C. López-Marcos, and M. A. López-Marcos. Analysis of an efficient integrator for a size-structured population model with a dynamical resource. Computers and Mathematics with Applications, 68(9):941–961, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001667.

Alomari:2011:NAS

[Alo11] A. K. Alomari. A new analytic solution for fractional chaotic dynamical systems using the differential transform method. Computers and Mathematics with Applications, 61(9):2528–2534, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001428.

Alsuwaiyel:2010:COP

[Als10] M. H. Alsuwaiyel. On computing an optimal permutation of ranks for multiselection. Computers and Mathematics with Applications, 60(8):2200–2203, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005328.

Acosta:2010:MBO

[AM10a] Carlos D. Acosta and Carlos E. Mejía. A mollification based operator splitting method for convection diffusion equations. Computers and Mathematics with Applications, 59(4):1397–1408, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812210900741X.

Alamyane:2010:SFC

[AM10b] Ahmed A. Alamyane and A. A. Mohamad. Simulation of forced convection in a channel with extended surfaces by the lattice Boltzmann method. Computers and Mathematics with Applications, 59(7):2421–2430, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006415.

Anastassiou:2010:QAF

[AM10c] George A. Anastassiou and Razvan A. Mezei. Quantitative approximation by fractional smooth Poisson Cauchy singular operators. Computers and Mathematics with Applications, 60(1):122– 133, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000310X.

Amador:2011:ASB

[AM11] J. Amador and P. Moreno.

Analysis of the successful and blocked events in the Geo/Geo/c retrial queue. Computers and Mathematics with Applications, 61(9):2667– 2682, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001672.

Alt:2012:TCS

[AM12a] Rene Alt and Svetoslav Markov. Theoretical and computational studies of some bioreactor models. Computers and Mathematics with Applications, 64(3):350–360, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200171X.

Anguelov:2012:B

[AM12b] Roumen Anguelov and Svetoslav Markov. BIOMATH 2011. Computers and Mathematics with Applications, 64(3):161–162, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004488.

Appleby:2012:PGR

[AM12c] John A. D. Appleby and Michael J. Mc-Carthy. Preservation of the growth rates of delay differential equations by Euler schemes with non-uniform step sizes. Computers and Mathematics with Applications, 64(7):2251–2261, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001095.

Abbaszadeh:2013:FOC

[AM13a] Mostafa Abbaszadeh and Akbar Mohebbi. A fourth-order compact solution of the two-dimensional modified anomalous fractional sub-diffusion equation with a nonlinear source term. Computers and Mathematics with Applications, 66 (8):1345–1359, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004914.

Al-Malki:2013:NSI

[AM13b]  Faisal Al-Malki.  Nu-

merical simulation of the influence of partial premixing on the propagation of partially premixed flames. Computers and Mathematics with Applications, 66(3):279–288, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003003.

Ambartsoumian:2013:SFI

[AM13c] Gaik Ambartsoumian and Sunghwan Moon. A series formula for inversion of the V-line Radon transform in a disc. Computers and Mathematics with Applications, 66 (9):1567–1572, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000692.

Anguelov:2013:B

[AM13d] Roumen Anguelov and Svetoslav Markov. BIOMATH 2012. Computers and Mathematics with Applications, 66(9):1533, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005671.

Ali:2014:HFA

[AM14a] Iftikhar Ali and Nadeem A.

Malik. Hilfer fractional advection-diffusion equations with power-law initial condition; a numerical study using variational iteration method. Computers and Mathematics with Applications, 68(10):1161–1179, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004131.

Arnal:2014:GHS

[AM14b] A. Arnal and J. Monterde. Generating harmonic surfaces for interactive design. Computers and Mathematics with Applications, 67(10): 1914–1924, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114001473.

Ahmad:2015:NSQ

[AM15] Daud Ahmad and Bilal Masud. Near-stability of a quasi-minimal surface indicated through a tested curvature algorithm. Computers and Mathematics with Applications, 69(10):1242– 1262, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001200.

Arnal:2017:EBC

[AM17] A. Arnal and J. Monterde. Explicit Bézier control net of a PDE surface. Computers and Mathematics with Applications, 73(3):483–493, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306848.

Al-Malki:2018:ITT

[AM18] Faisal Al-Malki. Influence of transverse temperature gradient on the propagation of triple flames in porous channels. Computers and Mathematics with Applications, 76(2):406–418, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302323.

Akkouche:2014:OCP

[AMA14] Abderrahmane Akkouche, Ahmed Maidi, and Mohamed Aidene. Optimal control of partial differential equations based on the variational iteration method. Computers and Mathematics with Applications, 68(5):622–631, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003137.

Ambartsoumian:2012:IVL

[Amb12] Gaik Ambartsoumian. Inversion of the V-line Radon transform in a disc and its applications in imaging. Computers and Mathematics with Applications, 64(3):260–265, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000806.

Ambrosio:2019:MCP

[Amb19] Vincenzo Ambrosio. On the multiplicity and concentration of positive solutions for a p-fractional Choquard equation in RN. Computers and Mathematics with Applications, 78(8):2593–2617, October 15, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301993.

Anvariyeh:2010:PAS

[AMD10] S. M. Anvariyeh, S. Mirvakili, and B. Davvaz. Pawlak’s approximations in Γ-semihypergroups. Computers and Mathematics with Applications, 60(1):45–53, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002993.

Adams:2019:PLS

[AMGC19] Thomas Adams, Nicholas McLeish, Stefano Giani, and William M. Coombs. A parabolic level set reinitialisation method using a discontinuous Galerkin discretisation. Computers and Mathematics with Applications, 78(9):2944–2960, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300628.

Ammer:2014:SFE

[AML^{+}14] Regina Ammer, Matthias Markl, Ulric Ljungblad, Carolin Körner, and Ulrich Rüde. Simulating fast electron beam melting with a parallel thermal free surface lattice Boltzmann method. Computers and Mathematics with Applications, 67(2):318–330, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005944.

Amoddeo:2015:AGM

[Amo15] Antonino Amoddeo. Adaptive grid modelling for cancer cells in the early stage of invasion. Computers and Mathematics with Applications, 69(7): 610–619, April 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115000590.

Amoddeo:2018:MMS

[Amo18] Antonino Amoddeo. A moving mesh study for diffusion induced effects in avascular tumour growth. Computers and Mathematics with Applications, 75(7):2508–2519, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307873.

Ahmad:2011:APF

[AN11a] Bashir Ahmad and Juan J. Nieto. Anti-periodic fractional boundary value problems. Computers and Mathematics with Applications, 62(3):1150–1156, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001209.

Ameri:2011:FH

[AN11b] R. Ameri and T. Nozari. Fuzzy hyperalgebras. Computers and Mathematics with Applications, 61(2): 149–154, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006334.

Ahmad:2012:SFD

[AN12] Bashir Ahmad and Juan J. Nieto. Sequential fractional differential equations with three-point boundary conditions. Computers and Mathematics with Applications, 64 (10):3046–3052, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001617.

Anastassiou:2010:FNF

[Ana10] George A. Anastassiou.

Foundations of nabla fractional calculus on time scales and inequalities. Computers and Mathematics with Applications, 59(12):3750–3762, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002683.

Anastassiou:2011:MHT

[Ana11a] George A. Anastassiou.

Multivariate hyperbolic tangent neural network approximation. Computers and Mathematics with Applications, 61(4):809–821, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000934X.

Anastassiou:2011:ROT

[Ana11b] George A. Anastassiou.

Representations and Ostrowski type inequalities on time scales. Computers and Mathematics with Applications, 62 (10):3933–3958, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008248.

Anastassiou:2012:FNN

[Ana12] George A. Anastassiou. Fractional neural network approximation. Computers and Mathematics with Applications, 64 (6):1655–1676, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000296.

Andres:2012:PTI

[And12]  Jan Andres.  Period

two implies chaos for a class of multivalued maps: a naive approach. Computers and Mathematics with Applications, 64(7):2160–2165, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010261.

Ayub:2010:SDI

[ANN10] M. Ayub, A. Naeem, and Rab Nawaz. Sound due to an impulsive line source. Computers and Mathematics with Applications, 60(12):3123–3129, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007960.

Anonymous:2010:A

[Ano10a] Anonymous. Acknowledgements. Computers and Mathematics with Applications, 60(4):I, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005146.

Anonymous:2010:EBa

[Ano10b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 59(1):iii–v, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007664.

Anonymous:2010:EBb

[Ano10c] Anonymous. Editorial Board. Computers and Mathematics with Applications, 59(2):iii–v, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007846.

Anonymous:2010:EBc

[Ano10d]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 59(3):iii–v, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000325.

Anonymous:2010:EBd

[Ano10e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 59(4):iii–v, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000696.

Anonymous:2010:EBe

[Ano10f]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 59(5):iii–v, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001021.

Anonymous:2010:EBf

[Ano10g]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 59(6):iii–v, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001446.

Anonymous:2010:EBg

[Ano10h]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 59(7):iii–v, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001781.

Anonymous:2010:EBh

[Ano10i]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 59(8):iii–v, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002087.

Anonymous:2010:EBi

[Ano10j]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 59(9):iii– v, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002762.

Anonymous:2010:EBj

[Ano10k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 59(10):iii–v, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003172.

Anonymous:2010:EBk

[Ano10l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 59(11):iii–v, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003354.

Anonymous:2010:EBl

[Ano10m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 59(12):iii–v, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003524.

Anonymous:2010:EBm

[Ano10n]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 60(1):iii– v, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003755.

Anonymous:2010:EBn

[Ano10o]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 60(2):iii– v, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004025.

Anonymous:2010:EBo

[Ano10p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 60(3):iii–v, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004682.

Anonymous:2010:EBp

[Ano10q] Anonymous. Editorial Board. Computers and Mathematics with Applications, 60(4):iii–v, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005079.

Anonymous:2010:EBq

[Ano10r]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 60(5):iii–v, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005560.

Anonymous:2010:EBr

[Ano10s]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 60(6):iii–v, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006085.

Anonymous:2010:EBs

[Ano10t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 60(7):iii–v, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006425.

Anonymous:2010:EBt

[Ano10u] Anonymous. Editorial Board. Computers and Mathematics with Applications, 60(8):iii–v, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007303.

Anonymous:2010:EBu

[Ano10v]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 60(9):iii–v, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110007674.

Anonymous:2010:EBv

[Ano10w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 60(10):iii–v, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008138.

Anonymous:2010:EBw

[Ano10x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 60(11):iii–v, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008436.

Anonymous:2010:EBx

[Ano10y] Anonymous. Editorial Board. Computers and Mathematics with Applications, 60(12):iii–v, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008771.

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 61(1):iii–v, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008862.

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 61(2):iii–v, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009776.

Anonymous:2011:EBc

[Ano11c]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 61(3):iii–v, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000277.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Computers and Mathematics with Applications, 61(4):v–vii, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000563.

Anonymous:2011:EBe

[Ano11e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 61(5):iii–v, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000782.

Anonymous:2011:EBf

[Ano11f]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 61(6):iii–v, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001283.

Anonymous:2011:EBg

[Ano11g]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 61(7):iii–v, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002008.

Anonymous:2011:EBh

[Ano11h]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 61(8):iii–v, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002501.

Anonymous:2011:EBi

[Ano11i]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 61(9):iii– v, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003105.

Anonymous:2011:EBj

[Ano11j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 61(10):iii–v, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111003804.

Anonymous:2011:EBk

[Ano11k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 61(11):iii–v, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004160.

Anonymous:2011:EBl

[Ano11l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 61(12):iii–v, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004536.

Anonymous:2011:EBm

[Ano11m]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 62(1):iii– v, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004913.

Anonymous:2011:EBn

[Ano11n]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 62(2):iii– v, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005463.

Anonymous:2011:EBo

[Ano11o] Anonymous. Editorial Board. Computers and Mathematics with Applications, 62(3):iii–v, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006055.

Anonymous:2011:EBp

[Ano11p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 62(4):iii–v, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006559.

Anonymous:2011:EBq

[Ano11q]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 62(5):iii–v, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100705X.

Anonymous:2011:EBr

[Ano11r]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 62(6):iii–v, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007358.

Anonymous:2011:EBs

[Ano11s] Anonymous. Editorial Board. Computers and Mathematics with Applications, 62(7):iii–v, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007668.

Anonymous:2011:EBt

[Ano11t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 62(8):iii–v, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008182.

Anonymous:2011:EBu

[Ano11u]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 62(9):iii–v, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008595.

Anonymous:2011:EBv

[Ano11v] Anonymous. Editorial Board. Computers and Mathematics with Applications, 62(10):iii–v, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009205.

Anonymous:2011:EBw

[Ano11w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 62(11):iii–v, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009539.

Anonymous:2011:EBx

[Ano11x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 62(12):iii–v, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009916.

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(1):iii–v, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010637.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(2):iii–v, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011175.

Anonymous:2012:EBc

[Ano12c]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 63(3):iii–v, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000065.

Anonymous:2012:EBd

[Ano12d]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 63(4):iii–v, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000612.

Anonymous:2012:EBe

[Ano12e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 63(5):iii–v, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000971.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(6):iii–iv, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001514.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(7):iii–iv, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001988.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(8):iii–iv, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003070.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(9):iii–iv, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003288.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(10):iii–iv, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003604.

Anonymous:2012:EBk

[Ano12k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(11):iii–iv, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200377X.

Anonymous:2012:EBl

[Ano12l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 63(12):iii–iv, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003926.

Anonymous:2012:EBm

[Ano12m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(1):iii–iv, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004075.

Anonymous:2012:EBn

[Ano12n] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(2):iii–iv, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004257.

Anonymous:2012:EBo

[Ano12o] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(3):iii–iv, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004543.

Anonymous:2012:EBp

[Ano12p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(4):iii–iv, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004786.

Anonymous:2012:EBq

[Ano12q] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(5):iii–iv, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005007.

Anonymous:2012:EBr

[Ano12r] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(6):iii–iv, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005238.

Anonymous:2012:EBs

[Ano12s] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(7):iii–iv, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112005524.

Anonymous:2012:EBt

[Ano12t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(8):iii–iv, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005767.

Anonymous:2012:EBu

[Ano12u] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(9):iii–iv, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005901.

Anonymous:2012:EBv

[Ano12v] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(10):iii–iv, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006128.

Anonymous:2012:EBw

[Ano12w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(11):iii–iv, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006372.

Anonymous:2012:EBx

[Ano12x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 64(12):iii–iv, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006621.

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(1):iii–iv, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006943.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(2):iii–iv, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112007110.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(3):iii–iv, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000072.

Anonymous:2013:EBd

[Ano13d] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(4):iii–iv, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000461.

Anonymous:2013:EBe

[Ano13e] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(5):iii–iv, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000813.

Anonymous:2013:EBf

[Ano13f]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 65(6):iii–v, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001089.

Anonymous:2013:EBg

[Ano13g]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 65(7):iii–v, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001338.

Anonymous:2013:EBh

[Ano13h]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 65(8):iii–v, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122113001776.

Anonymous:2013:EBi

[Ano13i]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 65(9):iii– v, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002009.

Anonymous:2013:EBj

[Ano13j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(10):iii–v, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002368.

Anonymous:2013:EBk

[Ano13k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(11):iii–v, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002629.

Anonymous:2013:EBl

[Ano13l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 65(12):iii–v, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002988.

Anonymous:2013:EBm

[Ano13m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 66(1):iii–v, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003398.

Anonymous:2013:EBn

[Ano13n] Anonymous. Editorial Board. Computers and Mathematics with Applications, 66(2):iii–v, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300357X.

Anonymous:2013:EBo

[Ano13o]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 66(3):iii–v, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003805.

Anonymous:2013:EBp

[Ano13p]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 66(4):iii–v, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004100.

Anonymous:2013:EBq

[Ano13q]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 66(5):iii–v, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004562.

Anonymous:2013:EBr

[Ano13r] Anonymous. Editorial Board. Computers and Mathematics with Applications, 66(6):iii–v, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004987.

Anonymous:2013:EBs

[Ano13s] Anonymous. Editorial Board. Computers and Mathematics with Applications, 66(7):iii–v, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005294.

Anonymous:2013:EBt

[Ano13t]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 66(8):iii–v, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300552X.

Anonymous:2013:EBu

[Ano13u]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 66(9):iii–v, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005725.

Anonymous:2013:EBv

[Ano13v] Anonymous. Editorial Board. Computers and Mathematics with Applications, 66(10):iii–v, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006007.

Anonymous:2013:EBw

[Ano13w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 66(11):iii–v, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006317.

Anonymous:2014:EBa

[Ano14a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 66(12):iii–v, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006470.

Anonymous:2014:EBb

[Ano14b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 67(1):iii–v, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006779.

Anonymous:2014:EBc

[Ano14c]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(2):iii–v, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006986.

Anonymous:2014:EBd

[Ano14d]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(3):iii–v, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400008X.

Anonymous:2014:EBe

[Ano14e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(4):iii–v, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000315.

Anonymous:2014:EBf

[Ano14f]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(5):iii–v, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000534.

Anonymous:2014:EBg

[Ano14g]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(6):iii–v, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400073X.

Anonymous:2014:EBh

[Ano14h]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(7):iii–v, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122114001114.

Anonymous:2014:EBi

[Ano14i]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(8):iii– v, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001217.

Anonymous:2014:EBj

[Ano14j]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 67(9):iii– v, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400145X.

Anonymous:2014:EBk

[Ano14k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 67(10):iii–v, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001928.

Anonymous:2014:EBl

[Ano14l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 67(11):iii–v, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002223.

Anonymous:2014:EBm

[Ano14m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 67(12):iii–v, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002387.

Anonymous:2014:EBn

[Ano14n] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(1–2):iii–v, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002545.

Anonymous:2014:EBo

[Ano14o] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(3):iii–v, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002983.

Anonymous:2014:EBp

[Ano14p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(4):iii–v, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003216.

Anonymous:2014:EBq

[Ano14q]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 68(5):iii–v, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003538.

Anonymous:2014:EBr

[Ano14r]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 68(6):iii–v, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114003939.

Anonymous:2014:EBs

[Ano14s] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(7):iii–v, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114004192.

Anonymous:2014:EBt

[Ano14t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(8):iii–v, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114004441.

Anonymous:2014:EBu

[Ano14u]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 68(9):iii–v, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122114004854.

Anonymous:2014:EBv

[Ano14v] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(10):iii–v, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005148.

Anonymous:2014:EBw

[Ano14w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(11):iii–v, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005380.

Anonymous:2014:EBx

[Ano14x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 68(12):iii–v, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005720.

Anonymous:2014:P

[Ano14y]  Anonymous.  Preface.

Computers and Mathematics with Applications, 68(12):2149–2150, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005677.

Anonymous:2015:EBa

[Ano15a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 69(1):iii–v, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114006002.

Anonymous:2015:EBb

[Ano15b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 69(2):iii–v, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114006142.

Anonymous:2015:EBc

[Ano15c]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 69(3):iii–v, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000127.

Anonymous:2015:EBd

[Ano15d]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 69(4):iii–v, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000255.

Anonymous:2015:EBe

[Ano15e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 69(5):iii–v, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000425.

Anonymous:2015:EBf

[Ano15f]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 69(6):iii–v, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000772.

Anonymous:2015:EBg

[Ano15g]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 69(7):iii–v, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001121.

Anonymous:2015:EBh

[Ano15h]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 69(8):iii–v, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001327.

Anonymous:2015:EBi

[Ano15i]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 69(9):iii– v, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001510.

Anonymous:2015:EBj

[Ano15j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 69(10):iii–v, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001625.

Anonymous:2015:EBk

[Ano15k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 69(11):iii–v, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001881.

Anonymous:2015:EBl

[Ano15l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 69(12):iii–v, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002242.

Anonymous:2015:EBm

[Ano15m]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 70(1):iii– v, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002473.

Anonymous:2015:EBn

[Ano15n]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 70(2):iii– v, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500262X.

Anonymous:2015:EBo

[Ano15o] Anonymous. Editorial Board. Computers and Mathematics with Applications, 70(3):iii–v, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002783.

Anonymous:2015:EBp

[Ano15p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 70(4):iii–v, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003260.

Anonymous:2015:EBq

[Ano15q]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 70(5):iii–v, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003648.

Anonymous:2015:EBr

[Ano15r]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 70(6):iii–v, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003934.

Anonymous:2015:EBs

[Ano15s] Anonymous. Editorial Board. Computers and Mathematics with Applications, 70(7):iii–v, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004289.

Anonymous:2015:EBt

[Ano15t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 70(8):iii–v, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004526.

Anonymous:2015:EBu

[Ano15u]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 70(9):iii–v, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004733.

Anonymous:2015:EBv

[Ano15v] Anonymous. Editorial Board. Computers and Mathematics with Applications, 70(10):iii–v, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500499X.

Anonymous:2015:EBw

[Ano15w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 70(11):iii–v, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005246.

Anonymous:2015:EBx

[Ano15x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 70(12):iii–v, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005489.

Anonymous:2016:A

[Ano16a] Anonymous. ASCAM-2015. Computers and Mathematics with Applications, 71(11):2139, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302462.

Anonymous:2016:EBa

[Ano16b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 71(1):iii–v, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115006148.

Anonymous:2016:EBb

[Ano16c] Anonymous. Editorial Board. Computers and Mathematics with Applications, 71(2):iii–v, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300050.

Anonymous:2016:EBc

[Ano16d]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 71(3):iii–v, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300347.

Anonymous:2016:EBd

[Ano16e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 71(4):iii–v, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S089812211630061X.

Anonymous:2016:EBe

[Ano16f]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 71(5):iii–v, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300967.

Anonymous:2016:EBf

[Ano16g]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 71(6):iii–v, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301092.

Anonymous:2016:EBg

[Ano16h]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 71(7):iii–v, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301377.

Anonymous:2016:EBh

[Ano16i]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 71(8):iii–v, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301547.

Anonymous:2016:EBi

[Ano16j]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 71(9):iii– v, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301791.

Anonymous:2016:EBj

[Ano16k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 71(10):iii–v, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302188.

Anonymous:2016:EBk

[Ano16l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 71(11):iii–v, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302516.

Anonymous:2016:EBl

[Ano16m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 71(12):iii–v, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302735.

Anonymous:2016:EBm

[Ano16n]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 72(1):iii– v, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303054.

Anonymous:2016:EBn

[Ano16o]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 72(2):iii– v, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303200.

Anonymous:2016:EBo

[Ano16p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 72(3):iii–v, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303868.

Anonymous:2016:EBp

[Ano16q] Anonymous. Editorial Board. Computers and Mathematics with Applications, 72(4):iii–v, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304175.

Anonymous:2016:EBq

[Ano16r]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 72(5):iii–v, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122116304497.

Anonymous:2016:EBr

[Ano16s]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 72(6):iii–v, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304941.

Anonymous:2016:EBs

[Ano16t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 72(7):iii–v, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305193.

Anonymous:2016:EBt

[Ano16u] Anonymous. Editorial Board. Computers and Mathematics with Applications, 72(8):iii–v, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630548X.

Anonymous:2016:EBu

[Ano16v]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 72(9):iii–v, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305752.

Anonymous:2016:EBv

[Ano16w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 72(10):iii–v, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211630596X.

Anonymous:2016:EBw

[Ano16x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 72(11):iii–v, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306174.

Anonymous:2016:EBx

[Ano16y] Anonymous. Editorial Board. Computers and Mathematics with Applications, 72(12):iii–v, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306447.

Anonymous:2017:EBa

[Ano17a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 73(1):iii–v, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630699X.

Anonymous:2017:EBb

[Ano17b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 73(2):iii–v, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300202.

Anonymous:2017:EBc

[Ano17c]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 73(3):iii–v, February 1, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300354.

Anonymous:2017:EBd

[Ano17d]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 73(4):iii–v, February 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300597.

Anonymous:2017:EBe

[Ano17e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 73(5):iii–v, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300895.

Anonymous:2017:EBf

[Ano17f]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 73(7):iii–v, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S089812211730158X.

Anonymous:2017:EBg

[Ano17g] Anonymous. Editorial Board. Computers and Mathematics with Applications, 73(8):iii–v, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301906.

Anonymous:2017:EBh

[Ano17h]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 73(9):iii–v, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730216X.

Anonymous:2017:EBi

[Ano17i] Anonymous. Editorial Board. Computers and Mathematics with Applications, 73(10):iii–v, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302407.

Anonymous:2017:EBj

[Ano17j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 73(11):iii–v, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302699.

Anonymous:2017:EBk

[Ano17k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 73(12):iii–v, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302894.

Anonymous:2017:EBl

[Ano17l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 74(2):iii–v, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303401.

Anonymous:2017:EBm

[Ano17m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 74(3):iii–v, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303905.

Anonymous:2017:EBn

[Ano17n] Anonymous. Editorial Board. Computers and Mathematics with Applications, 74(4):iii–v, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304467.

Anonymous:2017:EBo

[Ano17o]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 74(6):iii–v, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305400.

Anonymous:2017:EBp

[Ano17p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 74(8):iii–v, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306028.

Anonymous:2017:EBq

[Ano17q]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 74(10):iii–v, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306739.

Anonymous:2017:EBr

[Ano17r] Anonymous. Editorial Board. Computers and Mathematics with Applications, 74(12):iii–v, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307393.

Anonymous:2018:EBa

[Ano18a]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 73(6):iii–v, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301293.

Anonymous:2018:EBb

[Ano18b]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 74(1):iii–v, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303152.

Anonymous:2018:EBc

[Ano18c]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 74(5):iii–v, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304935.

Anonymous:2018:EBd

[Ano18d] Anonymous. Editorial Board. Computers and Mathematics with Applications, 74(7):iii–v, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305734.

Anonymous:2018:EBe

[Ano18e]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 74(9):iii–v, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306405.

Anonymous:2018:EBf

[Ano18f] Anonymous. Editorial Board. Computers and Mathematics with Applications, 74(11):iii–v, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307174.

Anonymous:2018:EBg

[Ano18g] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(1):i, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307976.

Anonymous:2018:EBh

[Ano18h] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(1):ii, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307988.

Anonymous:2018:EBi

[Ano18i] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(2):i, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300178.

Anonymous:2018:EBj

[Ano18j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(2):ii, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830018X.

Anonymous:2018:EBk

[Ano18k] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(3):i, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300294.

Anonymous:2018:EBl

[Ano18l] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(3):ii, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300300.

Anonymous:2018:EBm

[Ano18m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(4):i, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300683.

Anonymous:2018:EBn

[Ano18n] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(4):ii, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122118300695.

Anonymous:2018:EBo

[Ano18o] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(5):i, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300944.

Anonymous:2018:EBp

[Ano18p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(5):ii, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300956.

Anonymous:2018:EBq

[Ano18q]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 75(6): i, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301196.

Anonymous:2018:EBr

[Ano18r]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 75(6): ii, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301202.

Anonymous:2018:EBs

[Ano18s] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(7):i, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301470.

Anonymous:2018:EBt

[Ano18t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(7):ii, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301482.

Anonymous:2018:EBu

[Ano18u] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(8):i, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301834.

Anonymous:2018:EBv

[Ano18v] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(8):ii, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301846.

Anonymous:2018:EBw

[Ano18w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(9):i, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301974.

Anonymous:2018:EBx

[Ano18x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(9):ii, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301986.

Anonymous:2018:EBy

[Ano18y] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(10):i, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302232.

Anonymous:2018:EBz

[Ano18z] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(10):ii, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302244.

Anonymous:2018:EBaa

[Ano18-27]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 75(11):i, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122118302414.

Anonymous:2018:EBab

[Ano18-28]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 75(11):ii, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302426.

Anonymous:2018:EBac

[Ano18-29] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(12):i, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302578.

Anonymous:2018:EBad

[Ano18-30] Anonymous. Editorial Board. Computers and Mathematics with Applications, 75(12):ii, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830258X.

Anonymous:2018:EBae

[Ano18-31] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(1):i, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302797.

Anonymous:2018:EBaf

[Ano18-32] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(1):ii, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302803.

Anonymous:2018:EBag

[Ano18-33] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(2):i, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303006.

Anonymous:2018:EBah

[Ano18-34] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(2):ii, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303018.

Anonymous:2018:EBai

[Ano18-35] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(3):i, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303213.

Anonymous:2018:EBaj

[Ano18-36] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(3):ii, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303225.

Anonymous:2018:EBak

[Ano18-37] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(4):i, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303997.

Anonymous:2018:EBal

[Ano18-38] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(4):ii, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304000.

Anonymous:2018:EBam

[Ano18-39] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(5):i, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304553.

Anonymous:2018:EBan

[Ano18-40] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(5):ii, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122118304565.

Anonymous:2018:EBao

[Ano18-41] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(6):i, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304929.

Anonymous:2018:EBap

[Ano18-42] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(6):ii, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304930.

Anonymous:2018:EBaq

[Ano18-43] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(7):i, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305145.

Anonymous:2018:EBar

[Ano18-44] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(7):ii, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305157.

Anonymous:2018:EBas

[Ano18-45] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(8):i, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305595.

Anonymous:2018:EBat

[Ano18-46] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(8):ii, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305601.

Anonymous:2018:EBau

[Ano18-47] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(9):i, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305881.

Anonymous:2018:EBav

[Ano18-48] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(9):ii, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305893.

Anonymous:2018:EBaw

[Ano18-49] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(10):i, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306023.

Anonymous:2018:EBax

[Ano18-50] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(10):ii, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306035.

Anonymous:2018:EBay

[Ano18-51] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(11–12):i, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306473.

Anonymous:2018:EBaz

[Ano18-52] Anonymous. Editorial Board. Computers and Mathematics with Applications, 76(11–12):ii, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306485.

Anonymous:2018:PJe

[Ano18-53] Anonymous. Pages 1– 214 (1 July 2018). Computers and Mathematics with Applications, 76(1): ??, July 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PJa

[Ano18-54] Anonymous. Pages 1–334 (1 January 2018). Computers and Mathematics with Applications, 75(1): ??, January 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PFb

[Ano18-55] Anonymous. Pages 1067– 1482 (15 February 2018). Computers and Mathematics with Applications, 75(4):??, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PSb

[Ano18-56]  Anonymous.  Pages

1261–1554 (15 September 2018). Computers and Mathematics with Applications, 76(6):??, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PMa

[Ano18-57] Anonymous. Pages 1483– 1850 (1 March 2018). Computers and Mathematics with Applications, 75(5):??, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:POa

[Ano18-58] Anonymous. Pages 1555– 1826 (1 October 2018). Computers and Mathematics with Applications, 76(7):??, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:POb

[Ano18-59] Anonymous. Pages 1827– 2034 (15 October 2018). Computers and Mathematics with Applications, 76(8):??, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PNa

[Ano18-60] Anonymous. Pages 2035– 2314 (1 November 2018). Computers and Mathematics with Applications, 76(9):??, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PJf

[Ano18-61] Anonymous. Pages 215– 450 (15 July 2018). Computers and Mathematics with Applications, 76(2): ??, July 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PAa

[Ano18-62] Anonymous. Pages 2193– 2606 (1 April 2018). Computers and Mathematics with Applications, 75(7): ??, April 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PNb

[Ano18-63] Anonymous. Pages 2315– 2534 (15 November 2018). Computers and Mathematics with Applications, 76(10):??, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PD

[Ano18-64] Anonymous. Pages 2535– 2766 (1 December 2018). Computers and Mathematics with Applications, 76(11–12):??, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PAb

[Ano18-65] Anonymous. Pages 2607– 3066 (15 April 2018). Computers and Mathematics with Applications, 75(8):??, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PMb

[Ano18-66] Anonymous. Pages 3067– 3452 (1 May 2018). Computers and Mathematics with Applications, 75(9): ??, May 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PJb

[Ano18-67] Anonymous. Pages 335– 720 (15 January 2018). Computers and Mathematics with Applications, 75(2):??, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PMc

[Ano18-68] Anonymous. Pages 3453– 3824 (15 May 2018). Computers and Mathematics with Applications, 75(10):??, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PJc

[Ano18-69] Anonymous. Pages 3825– 4200 (1 June 2018). Computers and Mathematics with Applications, 75(11): ??, June 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PJd

[Ano18-70] Anonymous. Pages 4201– 4540 (15 June 2018). Computers and Mathematics with Applications, 75(12):??, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PAc

[Ano18-71] Anonymous. Pages 451– 672 (1 August 2018). Computers and Mathematics with Applications, 76(3):??, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PAd

[Ano18-72] Anonymous. Pages 673– 966 (15 August 2018). Computers and Mathematics with Applications, 76(4):??, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PFa

[Ano18-73] Anonymous. Pages 721– 1066 (1 February 2018). Computers and Mathematics with Applications, 75(3):??, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2018:PSa

[Ano18-74] Anonymous. Pages 967– 1260 (1 September 2018). Computers and Mathematics with Applications, 76(5):??, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:EBa

[Ano19a] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(1):i, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830717X.

Anonymous:2019:EBb

[Ano19b] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(1):ii, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307181.

Anonymous:2019:EBc

[Ano19c] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(2):i, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307429.

Anonymous:2019:EBd

[Ano19d] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(2):ii, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307430.

Anonymous:2019:EBe

[Ano19e] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(3):i, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300264.

Anonymous:2019:EBf

[Ano19f] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(3):ii, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122119300276.

Anonymous:2019:EBg

[Ano19g] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(4):i, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211930046X.

Anonymous:2019:EBh

[Ano19h] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(4):ii, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300471.

Anonymous:2019:EBi

[Ano19i] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(5):i, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300653.

Anonymous:2019:EBj

[Ano19j] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(5):ii, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300665.

Anonymous:2019:EBk

[Ano19k]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 77(6): i, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300896.

Anonymous:2019:EBl

[Ano19l]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 77(6): ii, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300902.

Anonymous:2019:EBm

[Ano19m] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(7):i, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301154.

Anonymous:2019:EBn

[Ano19n] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(7):ii, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301166.

Anonymous:2019:EBo

[Ano19o] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(8):i, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301348.

Anonymous:2019:EBp

[Ano19p] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(8):ii, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930135X.

Anonymous:2019:EBq

[Ano19q] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(9):i, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301725.

Anonymous:2019:EBr

[Ano19r] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(9):ii, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301737.

Anonymous:2019:EBs

[Ano19s] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(10):i, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122119301956.

Anonymous:2019:EBt

[Ano19t] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(10):ii, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301968.

Anonymous:2019:EBu

[Ano19u]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 77(11):i, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302251.

Anonymous:2019:EBv

[Ano19v]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 77(11):ii, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302263.

Anonymous:2019:EBw

[Ano19w] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(12):i, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302378.

Anonymous:2019:EBx

[Ano19x] Anonymous. Editorial Board. Computers and Mathematics with Applications, 77(12):ii, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930238X.

Anonymous:2019:EBy

[Ano19y] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(1):i, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302536.

Anonymous:2019:EBz

[Ano19z] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(1):ii, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302548.

Anonymous:2019:EBaa

[Ano19-27]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 78(2): i, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302639.

Anonymous:2019:EBab

[Ano19-28]  Anonymous.  Edito-

rial Board. Computers and Mathematics with Applications, 78(2): ii, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302640.

Anonymous:2019:EBac

[Ano19-29] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(3):i, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302986.

Anonymous:2019:EBad

[Ano19-30] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(3):ii, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302998.

Anonymous:2019:EBae

[Ano19-31] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(4):i, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303578.

Anonymous:2019:EBaf

[Ano19-32] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(4):ii, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S089812211930358X.

Anonymous:2019:EBag

[Ano19-33] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(5):i, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119303694.

Anonymous:2019:EBah

[Ano19-34] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(5):ii, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119303700.

Anonymous:2019:EBai

[Ano19-35] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(6):i, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119303992.

Anonymous:2019:EBaj

[Ano19-36] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(6):ii, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304006.

Anonymous:2019:EBak

[Ano19-37] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(7):i, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304201.

Anonymous:2019:EBal

[Ano19-38] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(7):ii, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304213.

Anonymous:2019:EBam

[Ano19-39] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(8):i, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930447X.

Anonymous:2019:EBan

[Ano19-40] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(8):ii, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304481.

Anonymous:2019:EBao

[Ano19-41] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(9):i, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304729.

Anonymous:2019:EBap

[Ano19-42] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(9):ii, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304730.

Anonymous:2019:EBaq

[Ano19-43] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(10):i, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304882.

Anonymous:2019:EBar

[Ano19-44] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(10):ii, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304894.

Anonymous:2019:EBas

[Ano19-45] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(11):i, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122119305140.

Anonymous:2019:EBat

[Ano19-46] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(11):ii, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119305152.

Anonymous:2019:EBau

[Ano19-47] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(12):i, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119305292.

Anonymous:2019:EBav

[Ano19-48] Anonymous. Editorial Board. Computers and Mathematics with Applications, 78(12):ii, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119305309.

Anonymous:2019:PJd

[Ano19-49] Anonymous. Pages 1– 256 (1 July 2019). Computers and Mathematics with Applications, 78(1): ??, July 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PJa

[Ano19-50] Anonymous. Pages 1–310 (1 January 2019). Computers and Mathematics with Applications, 77(1): ??, January 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PMa

[Ano19-51] Anonymous. Pages 1233– 1422 (1 March 2019). Computers and Mathematics with Applications, 77(5):??, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PAa

[Ano19-52] Anonymous. Pages 1799– 2028 (1 April 2019). Computers and Mathematics with Applications, 77(7): ??, April 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PS

[Ano19-53]  Anonymous.  Pages

1801–2166 (15 September 2019). Computers and Mathematics with Applications, 78(6):??, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PAb

[Ano19-54] Anonymous. Pages 2029– 2262 (15 April 2019). Computers and Mathematics with Applications, 77(8):??, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PMb

[Ano19-55] Anonymous. Pages 2263– 2584 (1 May 2019). Computers and Mathematics with Applications, 77(9): ??, May 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PO

[Ano19-56] Anonymous. Pages 2429– 2846 (15 October 2019). Computers and Mathematics with Applications, 78(8):??, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PMc

[Ano19-57] Anonymous. Pages 2585– 2866 (15 May 2019). Computers and Mathematics with Applications, 77(10):??, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PJc

[Ano19-58] Anonymous. Pages 3043– 3280 (15 June 2019). Computers and Mathematics with Applications, 77(12):??, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PJb

[Ano19-59] Anonymous. Pages 311– 614 (15 January 2019). Computers and Mathematics with Applications, 77(2):??, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PN

[Ano19-60] Anonymous. Pages 3237– 3470 (15 November 2019). Computers and Mathematics with Applications, 78(10):??, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PDa

[Ano19-61] Anonymous. Pages 3471– 3712 (1 December 2019). Computers and Mathematics with Applications, 78(11):??, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PDb

[Ano19-62] Anonymous. Pages 3713– 3956 (15 December 2019). Computers and Mathematics with Applications, 78(12):??, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PFa

[Ano19-63] Anonymous. Pages 615– 906 (1 February 2019). Computers and Mathematics with Applications, 77(3):??, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PAc

[Ano19-64] Anonymous. Pages 707– 1050 (1 August 2019). Computers and Mathematics with Applications, 78(3):??, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Anonymous:2019:PFb

[Ano19-65] Anonymous. Pages 907– 1232 (15 February 2019). Computers and Mathematics with Applications, 77(4):??, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Ahmad:2011:SBV

[ANP11] Bashir Ahmad, Juan J.

Nieto, and Johnatan Pimentel. Some boundary value problems of fractional differential equations and inclusions. Computers and Mathematics with Applications, 62(3):1238–1250, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001210.

Arrieta:2019:SEE

[ANP19] José M. Arrieta, Ariadne Nogueira, and Marcone C. Pereira. Semilinear elliptic equations in thin regions with terms concentrating on oscillatory boundaries. Computers and Mathematics with Applications, 77(2):536–554, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305765.

Arif:2011:CAF

[ANR11] Muhammad Arif, Khalida Inayat Noor, and Mohsan Raza. On a class of analytic functions related with generalized bazilevic type functions. Computers and Mathematics with Applications, 61(9):2456–2462, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100112X.

Antczak:2010:SPC

[Ant10] Tadeusz Antczak. Saddle points criteria in nondifferentiable multiobjective programming with V -invex functions via an η-approximation method. Computers and Mathematics with Applications, 60(9):2689–2700, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006991.

Antczak:2014:CSD

[Ant14] Tadeusz Antczak. Comments on “Sufficiency and duality for multiobjective variational control problems with G-invexity” computers and mathematics with applications 63, 838–850 (2012). Computers and Mathematics with Applications, 66(12):2595–2596, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005932.

See [ZLLF12].

Abidi:2010:HAM

[AO10a] Fayçal Abidi and Khaled Omrani. The homotopy analysis method for solving the Fornberg– Whitham equation and comparison with Adomian’s decomposition method. Computers and Mathematics with Applications, 59(8):2743–2750, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000647.

Agirseven:2010:ASF

[AÖ10b] Deniz Agirseven and Turgut Özis. An analytical study for Fisher type equations by using homotopy perturbation method. Computers and Mathematics with Applications, 60(3):602–609, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003469.

Atouani:2013:GFE

[AO13] Noureddine Atouani and Khaled Omrani. Galerkin finite element method for the Rosenau–RLW equation. Computers and Mathematics with Applications, 66(3):289–303, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300309X.

Altmann:2018:SMC

[AO18] R. Altmann and A. Ostermann. Splitting methods for constrained diffusionreaction systems. Computers and Mathematics with Applications, 74(5):962–976, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301232.

Anand:2018:NBF

[AOW18] Akash Anand, Jeffrey S.

Ovall, and Steffen Weißer. A Nyström-based finite element method on polygonal elements. Computers and Mathematics with Applications, 75(11):3971–3986, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301366.

Aksoy:2010:NPI

[AP10] Yigit Aksoy and Mehmet Pakdemirli. New perturbationiteration solutions for Bratu-type equations. Computers and Mathematics with Applications, 59(8):2802–2808, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000854.

Abreu:2019:FRS

[AP19a] Eduardo Abreu and John Pérez. A fast, robust, and simple Lagrangian– Eulerian solver for balance laws and applications. Computers and Mathematics with Applications, 77(9):2310–2336, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307119.

Arraras:2019:MFM

[AP19b] A. Arrarás and L. Portero. Multipoint flux mixed finite element methods for slightly compressible flow in porous media. Computers and Mathematics with Applications, 77(6):1437–1452, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830364X.

Alfaro:2011:OPA

[APRM11] Manuel Alfaro, Ana Peña, M. Luisa Rezola, and Francisco Marcellán. Orthogonal polynomials associated with an inverse quadratic spectral transform. Computers and Mathematics with Applications, 61(4):888–900, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009429.

Aydi:2012:CFP

[APS12] Hassen Aydi, Mihai Postolache, and Wasfi Shatanawi. Coupled fixed point results for (ψ, ϕ)- weakly contractive mappings in ordered G-metric spaces. Computers and Mathematics with Applications, 63(1):298–309, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010029.

Agnelli:2011:SOT

[APT11] J. P. Agnelli, C. Padra, and C. V. Turner. Shape optimization for tumor location. Computers and Mathematics with Applications, 62(11):4068– 4081, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111008339.

Au:2019:SRM

[APTZ19] Vo Van Au, Nguyen Duc Phuong, Nguyen Huy Tuan, and Yong Zhou. Some regularization methods for a class of nonlinear fractional evolution equations. Computers and Mathematics with Applications, 78(5):1752–1771, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119303189.

Arshad:2018:AMM

[APwS18] Muhammad Arshad, Eun-Jae Park, and Dong wook Shin. Analysis of multiscale mortar mixed approximation of nonlinear elliptic equations. Computers and Mathematics with Applications, 75(2):401–418, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305941.

Adivar:2009:SPD

[AR09] Murat Adivar and Youssef N. Raffoul. Stability and periodicity in dynamic delay equations. Computers and Mathematics with Applications, 58(2):264– 272, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109002235.

See note [AR10b].

Adibi:2010:UML

[AR10a] H. Adibi and A. M. Rismani. On using a modified Legendre-spectral method for solving singular IVPs of Lane–Emden type. Computers and Mathematics with Applications, 60(7):2126–2130, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005481.

Adivar:2010:NSP

[AR10b] Murat Adivar and Youssef N.

Raffoul. A note on “Stability and periodicity in dynamic delay equations” [Comput. Math. Appl. 58 (2009) 264– 272]. Computers and Mathematics with Applications, 59(10):3351– 3354, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002014.

See [AR09].

Alasooly:2010:OCU

[AR10c] Hedaya Alasooly and Mohammed Redha. Optimal control of UPFC for load flow control and voltage flicker elimination and current harmonics elimination. Computers and Mathematics with Applications, 60(4):926–943, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001549.

Altun:2010:OCM

[AR10d] Ishak Altun and Vladimir Rakocević. Ordered cone metric spaces and fixed point results. Computers and Mathematics with Applications, 60 (5):1145–1151, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003962.

Anwar:2017:SFR

[AR17] Muhammad Shoaib Anwar and Amer Rasheed. Simulations of a fractional rate type nanofluid flow with non-integer Caputo time derivatives. Computers and Mathematics with Applications, 74(10):2485–2502, November 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0898122117304704.

Arakelyan:2018:CFD

[Ara18] Avetik Arakelyan. Convergence of the finite difference scheme for a general class of the spatial segregation of reaction– diffusion systems. Computers and Mathematics with Applications, 75(12):4232–4240, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301603.

Al-rabtah:2010:SFO

[ArEM10] Adel Al-rabtah, Vedat Suat Ertürk, and Shaher Momani. Solutions of a fractional oscillator by using differential transform method. Computers and Mathematics with Applications, 59(3):1356–1362, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004118.

Abdel-Rehim:2018:SAS

[ARESH18] E. A. Abdel-Rehim, A. M. A. El-Sayed, and A. S. Hashem. Simulation of the approximate solutions of the time-fractional multiterm wave equations. Computers and Mathematics with Applications, 73(6):1134–1154, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303510.

Al-Refai:2013:CBB

[ARK13] Mohammed Al-Refai and Nikos I. Kavallaris. On computation of bounds of the bifurcation parameter for a non-local elliptic equation with increasing nonlinearity. Computers and Mathematics with Applications, 66(4):512–524, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003222.

Arqub:2018:FRK

[Arq18] Omar Abu Arqub. Fitted reproducing kernel Hilbert space method for the solutions of some certain classes of timefractional partial differential equations subject to initial and Neumann boundary conditions. Computers and Mathematics with Applications, 73(6):1243–1261, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306642.

Aouf:2010:SPC

[AS10a] M. K. Aouf and T. M.

Seoudy. Some properties of a certain subclass of multivalent analytic functions involving the Liu–Owa operator. Computers and Mathematics with Applications, 60(6):1525–1535, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004505.

Aouf:2010:SSC

[AS10b] M. K. Aouf and T. M.

Seoudy. Subordination and superordination of a certain integral operator on meromorphic functions. Computers and Mathematics with Applications, 59(12):3669– 3678, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002580.

Ashyralyev:2010:WPP

[AS10c] Allaberen Ashyralyev and Yasar Sözen. Wellposedness of parabolic differential and difference equations. Computers and Mathematics with Applications, 60(3):792– 802, August 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003834.

Ahadpanah:2011:SHB

[AS11a] A. Ahadpanah and A. Borumand Saeid. Smarandache hyper BCC-algebra. Computers and Mathematics with Applications, 61(9):2490–2497, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001179.

Alolyan:2011:FHO

[AS11b] Ibraheem Alolyan and T. E. Simos. A family of high-order multistep methods with vanished phase-lag and its derivatives for the numerical solution of the Schrödinger equation. Computers and Mathematics with Applications, 62(10):3756–3774, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007930.

Aouf:2011:SPS

[AS11c] M. K. Aouf and T. M.

Seoudy. Some preserving subordination and superordination of analytic functions involving the Liu–Owa integral operator. Computers and Mathematics with Applications, 62(9):3575–3580, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007607.

Atagun:2011:SSR

[AS11d] Akin Osman Atagün and Aslihan Sezgin. Soft substructures of rings, fields and modules. Computers and Mathematics with Applications, 61(3):592–601, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110009107.

An:2015:SAT

[AS15a] Jing An and Jie Shen. Spectral approximation to a transmission eigenvalue problem and its applications to an inverse problem. Computers and Mathematics with Applications, 69(10):1132– 1143, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000978.

Axelsson:2015:CNM

[AS15b] Owe Axelsson and Stanislav Sysala. Continuation Newton methods. Computers and Mathematics with Applications, 70 (11):2621–2637, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003818.

Armentano:2019:UMF

[AS19] María Gabriela Armentano and María Lorena Stockdale. A unified mixed finite element approximations of the Stokes–Darcy coupled problem. Computers and Mathematics with Applications, 77(9):2568–2584, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307302.

Aizenbud:2016:RLD

[ASA16] Yariv Aizenbud, Gil Shabat, and Amir Averbuch. Randomized LU decomposition using sparse projections. Computers and Mathematics with Applications, 72(9):2525–2534, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305260.

Angelov:2012:BEO

[ASB12] Angel G. Angelov and Maroussia Slavtchova-Bojkova. Bayesian estimation of the offspring mean in branching processes: Application to infectious disease data. Computers and Mathematics with Applications, 64(3):229–235, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112000703.

Afrouzi:2015:LBA

[ASFM15] H. Hassanzadeh Afrouzi, K. Sedighi, M. Farhadi, and A. Moshfegh. Lattice Boltzmann analysis of micro-particles transport in pulsating obstructed channel flow. Computers and Mathematics with Applications, 70(5):1136–1151, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003430.

Aslan:2010:EFA

[Asl10] Ismail Aslan. The Expfunction approach to the Schwarzian Korteweg–de Vries equation. Computers and Mathematics with Applications, 59(8):2896– 2900, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001161.

Aslan:2011:CAE

[Asl11]  Ismail Aslan.  Com-

ment on: “Application of Exp-function method for (3 + 1)-dimensional nonlinear evolution equations” [Comput. Math. Appl. 56 (2008) 1451– 1456]. Computers and Mathematics with Applications, 61(6):1700–1703, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100071X.

See [BB08].

Aouf:2011:IPC

[ASMEE11a] M. K. Aouf, A. Shamandy, A. O. Mostafa, and F. Z. El-Emam. Inclusion properties of certain classes of meromorphic functions associated with the Wright generalized hypergeometric function. Computers and Mathematics with Applications, 61(5):1419–1424, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000174.

Aouf:2011:STM

[ASMEE11b] M. K. Aouf, A. Shamandy, A. O. Mostafa, and F. Z. El-Emam. On sandwich theorems for multivalent functions involving a generalized differential operator. Computers and Mathematics with Applications, 61(9):2578– 2587, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001507.

Awad:2011:CIC

[ASMM11] F. G. Awad, P. Sibanda, S. S. Motsa, and O. D. Makinde. Convection from an inverted cone in a porous medium with cross-diffusion effects. Computers and Mathematics with Applications, 61(5):1431–1441, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000204.

See comment [Pan17].

Ali:2011:ASS

[ASN11] Muhammad Irfan Ali, Muhammad Shabir, and Munazza Naz. Algebraic structures of soft sets associated with new operations. Computers and Mathematics with Applications, 61(9):2647–2654, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122111001635.

Al-sawalha:2010:AAS

[AsNAd10] M. Mossa Al-sawalha, M. S. M. Noorani, and M. M. Al-dlalah. Adaptive anti-synchronization of chaotic systems with fully unknown parameters. Computers and Mathematics with Applications, 59(10):3234– 3244, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001707.

Arregui:2018:TVA

[ASSV18] Iñigo Arregui, Beatriz Salvador, Daniel Sevcovic, and Carlos Vázquez. Total value adjustment for European options with two stochastic factors. Mathematical model, analysis and numerical simulation. Computers and Mathematics with Applications, 76(4):725–740, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302748.

Aydi:2011:GWC

[ASV11] Hassen Aydi, Wasfi Shatanawi, and Calogero Vetro. On generalized weak Gcontraction mapping in G-metric spaces. Computers and Mathematics with Applications, 62 (11):4222–4229, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008716.

Askari:2011:APS

[ASY^{+}11] H. Askari, Z. Saadatnia, D. Younesian, A. Yildirim, and M. Kalami-Yazdi. Approximate periodic solutions for the Helmholtz– Duffing equation. Computers and Mathematics with Applications, 62 (10):3894–3901, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008108.

Almeida:2011:FVC

[AT11] Ricardo Almeida and Delfim F. M. Torres. Fractional variational calculus for nondifferentiable functions. Computers and Mathematics with Applications, 61(10):3097– 3104, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002781.

Anh:2017:RCS

[AT17] Cung The Anh and Pham Thi Trang. On the regularity and convergence of solutions to the 3D Navier–Stokes– Voigt equations. Computers and Mathematics with Applications, 73(4):601–615, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306940.

Amini:2018:WFV

[AT18a] S. Amini and F. Toutounian. Weighted and flexible versions of block CMRH method for solving nonsymmetric linear systems with multiple right-hand sides. Computers and Mathematics with Applications, 76(8):2011–2021, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118304164.

Ammi:2018:GSM

[AT18b] Moulay Rchid Sidi Ammi and Delfim F. M. Torres. Galerkin spectral method for the fractional nonlocal thermistor problem. Computers and Mathematics with Applications, 73(6):1077–1086, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303133.

Ammi:2019:OCN

[AT19] Moulay Rchid Sidi Ammi and Delfim F. M. Torres. Optimal control of a nonlocal thermistor problem with ABC fractional time derivatives. Computers and Mathematics with Applications, 78(5):1507–1516, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301683.

Ansari:2018:PSB

[ATH18] Reza Ansari, Jalal Torabi, and Ramtin Hassani. Inplane and shear buckling analysis of FG– CNTRC annular sector plates based on the thirdorder shear deformation theory using a numerical approach. Computers and Mathematics with Applications, 75(2):486–502, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305850.

Ansari:2019:TBA

[ATH19] R. Ansari, J. Torabi, and R. Hassani. Thermal buckling analysis of temperature-dependent FG–CNTRC quadrilateral plates. Computers and Mathematics with Applications, 77(5):1294–1311, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830662X.

Abdolahzadeh:2019:MPT

[ATO19] Mohsen Abdolahzadeh, Ali Tayebi, and Pourya Omidvar. Mixing process of two-phase non-Newtonian fluids in 2D using smoothed particle hydrodynamics. Computers and Mathematics with Applications, 78(1):110– 122, July 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300938.

Ahmad:2015:HOM

[ATUC15] Fayyaz Ahmad, Emran Tohidi, Malik Zaka Ullah, and Juan A. Carrasco. Higher order multistep Jarratt-like method for solving systems of nonlinear equations: Application to PDEs and ODEs. Computers and Mathematics with Applications, 70(4):624–636, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002357.

Adams:2011:DMF

[ATZ11] Sarah Spence Adams, Denise Sakai Troxell, and S. Luke Zinnen. Dynamic monopolies and feedback vertex sets in hexagonal grids. Computers and Mathematics with Applications, 62 (11):4049–4057, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100825X.

Auchmuty:2018:VPA

[Auc18] Giles Auchmuty. Variational principles for advection–diffusion problems. Computers and Mathematics with Applications, 75(6):1882–1886, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305862.

Aziz:2017:HWC

[AuIA17] Imran Aziz, Siraj ul Islam, and Muhammad Asif. Haar wavelet collocation method for threedimensional elliptic partial differential equations. Computers and Mathematics with Applications, 73(9):2023–2034, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730113X.

Aziz:2011:QRN

[AuIK11] Imran Aziz, Siraj ul Islam, and Wajid Khan. Quadrature rules for numerical integration based on Haar wavelets and hybrid functions. Computers and Mathematics with Applications, 61(9):2770– 2781, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001957.

Ahmad:2017:LRM

[AuIK17] Imtiaz Ahmad, Siraj ul Islam, and Abdul Q. M. Khaliq. Local RBF method for multi-dimensional partial differential equations. Computers and Mathematics with Applications, 74(2):292–324, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302572.

Araya:2014:PEE

[AV14] Rodolfo Araya and Pablo Venegas. An a posteriori error estimator for an unsteady advectiondiffusion-reaction problem. Computers and Mathematics with Applications, 66(12):2456– 2476, January 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0898122113005890.

Abdelmalik:2016:ESD

[AvB16] M. R. A. Abdelmalik and E. H. van Brummelen. An entropy stable discontinuous Galerkin finiteelement moment method for the Boltzmann equation. Computers and Mathematics with Applications, 72(8):1988–1999, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302899.

Anguelov:2013:ATS

[AvdW13] Roumen Anguelov and Jan Harm van der Walt. Algebraic and topological structure of some spaces of set-valued maps. Computers and Mathematics with Applications, 66(9):1643–1654, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001958.

Abdulhameed:2017:MEF

[AVR17] M. Abdulhameed, D. Vieru, and R. Roslan. Magnetohydrodynamic electroosmotic flow of Maxwell fluids with Caputo-fabrizio derivatives through circular tubes. Computers and Mathematics with Applications, 74 (10):2503–2519, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304698.

Aissa:2018:TGA

[AVV18] Mohamed Aissa, Tom Verstraete, and Cornelis Vuik. Toward a GPUaware comparison of explicit and implicit CFD simulations on structured meshes. Computers and Mathematics with Applications, 74(1):201–217, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301438.

Antonietti:2015:TLM

[AVZ15] Paola F. Antonietti, Marco Verani, and Ludmil Zikatanov. A two-level method for mimetic finite difference discretizations of elliptic problems. Computers and Mathematics with Applications, 70 (11):2674–2687, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002965.

Ahmad:2011:SIF

[AW11] Bashir Ahmad and Guotao Wang. A study of an impulsive four-point nonlocal boundary value problem of nonlinear fractional differential equations. Computers and Mathematics with Applications, 62(3):1341–1349, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003348.

Alrehaili:2019:ENA

[AWJH19] A. H. Alrehaili, M. A.

Walkley, P. K. Jimack, and M. E. Hubbard. An efficient numerical algorithm for a multiphase tumour model. Computers and Mathematics with Applications, 78(8):2734–2745, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302159.

Aouf:2011:SIR

[AX11] M. K. Aouf and N.-Eng Xu. Some inclusion relationships and integralpreserving properties of certain subclasses of pvalent meromorphic functions. Computers and Mathematics with Applications, 61(3):642–650, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009156.

Ashyralyev:2012:MCN

[AY12] Allaberen Ashyralyev and Serhat Yilmaz. Modified Crank–Nicolson difference schemes for ultraparabolic equations. Computers and Mathematics with Applications, 64(8):2756–2764, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005457.

Ahn:2018:SPE

[AY18] Chi Young Ahn and Sangwoon Yun. A study on the 3D position estimation of ventricular borders extracted from 2D echocardiography data.

Computers and Mathematics with Applications, 75(4):1143–1158, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306831.

An:2017:LDG

[AYH17] Na An, Xijun Yu, and Chaobao Huang. Local discontinuous Galerkin methods for parabolic interface problems with homogeneous and nonhomogeneous jump conditions. Computers and Mathematics with Applications, 74(10):2572– 2598, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117304601.

Ay:2012:ALB

[AYY12] Chyung Ay, Chao-Wang Young, and Chuen-Fu Young. Application of lattice Boltzmann method to the fluid analysis in a rectangular microchannel. Computers and Mathematics with Applications, 64(5):1065–1083, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002301.

Asheghi:2010:BLC

[AZ10] Rasoul Asheghi and Hamid R. Z. Zangeneh. Bifurcations of limit cycles for a quintic Hamiltonian system with a double cuspidal loop. Computers and Mathematics with Applications, 59(4):1409–1418, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007913.

Al-Zhour:2015:EGP

[AZ15] Zeyad Al-Zhour. Extension and generalization properties of the weighted Minkowski inverse in a Minkowski space for an arbitrary matrix. Computers and Mathematics with Applications, 70(5):954–961, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003016.

Abide:2017:MDC

[AZ17] Stéphane Abide and Belkacem Zeghmati. Multigrid defect correction and fourth-order compact scheme for Poisson’s equation. Computers and Mathematics with Applications, 73(7):1433–1444, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300469.

Zou:2018:GFE

[aZ18]  Guang an Zou.  A

Galerkin finite element method for time-fractional stochastic heat equation. Computers and Mathematics with Applications, 75(11):4135–4150, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301548.

Akimenko:2013:IPE

[AZB13] V. V. Akimenko, Yu. V.

Zahorodnii, and A. L. Boyko. Identification of parameters of evolutionary model of monocyclic cells aggregation with the hop plants example. Computers and Mathematics with Applications, 66(9):1547–1553, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300151X.

Agarwal:2010:EFN

[AZH10] R. P. Agarwal, Yong Zhou, and Yunyun He. Existence of fractional neutral functional differential equations. Computers and Mathematics with Applications, 59(3):1095–1100, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812210900354X.

Zou:2017:SBE

[aZW17] Guang an Zou and Bo Wang. Stochastic Burgers’ equation with fractional derivative driven by multiplicative noise. Computers and Mathematics with Applications, 74 (12):3195–3208, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305151.

Bareche:2011:STN

[BA11] Aïcha Bareche and Djamil Aïssani. Statistical techniques for a numerical evaluation of the proximity of G/G/1 and G/M/1 queueing systems. Computers and Mathematics with Applications, 61(5):1296–1304, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009624.

Bernal:2016:MLA

[BA16] Francisco Bernal and Juan A. Acebrón. A multigrid-like algorithm for probabilistic domain decomposition. Computers and Mathematics with Applications, 72(7):1790–1810, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304357.

Baaquie:2013:FMQ

[Baa13] Belal E. Baaquie. Financial modeling and quantum mathematics. Computers and Mathematics with Applications, 65(10):1665–1673, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000540.

Babovsky:2014:DKM

[Bab14] Hans Babovsky. Discrete kinetic models in the fluid dynamic limit. Computers and Mathematics with Applications, 67(2):256–271, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004240.

Baccouch:2014:SPE

[Bac14a] Mahboub Baccouch. Superconvergence and a posteriori error estimates for the LDG method for convection–diffusion problems in one space dimension. Computers and Mathematics with Applications, 67(5):1130–1153, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113007086.

Baccouch:2014:SLD

[Bac14b] Mahboub Baccouch. A superconvergent local discontinuous Galerkin method for the secondorder wave equation on Cartesian grids. Computers and Mathematics with Applications, 68 (10):1250–1278, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004271.

Bacuta:2014:CMA

[Bac14c] Constantin Bacuta. Cascadic multilevel algorithms for symmetric saddle point systems. Computers and Mathematics with Applications, 67(10):1905–1913, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001485.

Baek:2010:FCS

[Bae10] Hunki Baek. A food chain system with Holling type IV functional response and impulsive perturbations. Computers and Mathematics with Applications, 60(5):1152–1163, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003974.

Baghaei:2017:LBB

[Bag17] Khadijeh Baghaei. Lower bounds for the blowup time in a superlinear hyperbolic equation with linear damping term. Computers and Mathematics with Applications, 73(4):560–564, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306782.

Bai:2011:NRC

[Bai11a] Chuanzhi Bai. New results concerning the exponential stability of delayed neural networks with impulses. Computers and Mathematics with Applications, 62(7):2719–2726, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005347.

Bai:2011:SCF

[Bai11b] Zhanbing Bai. Solvability for a class of fractional m-point boundary value problem at resonance. Computers and Mathematics with Applications, 62(3):1292–1302, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001544.

Bai:2012:EIC

[Bai12] Zhanbing Bai. Eigenvalue intervals for a class of fractional boundary value problem. Computers and Mathematics with Applications, 64 (10):3253–3257, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000144.

Bai:2019:NNA

[Bai19]  Lufeng Bai.  A new

nonconvex approach for image restoration with gamma noise. Computers and Mathematics with Applications, 77(10):2627–2639, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300045.

Banks:2013:NCG

[Ban13] J. W. Banks. A note on the convergence of Godunov type methods for shock reflection problems. Computers and Mathematics with Applications, 66(1):19–23, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002472.

Belendez:2012:ASN

[BAO^{+}12] A. Beléndez, E. Arribas, M. Ortuño, S. Gallego, A. Márquez, and I. Pascual. Approximate solutions for the nonlinear pendulum equation using a rational harmonic representation. Computers and Mathematics with Applications, 64(6):1602–1611, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200017X.

Bao:2016:IMS

[Bao16]  Gui Bao.  Infinitely

many small solutions for a sublinear Schrödinger– Poisson system with sign-changing potential. Computers and Mathematics with Applications, 71(10):2082–2088, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301742.

Barcelos:2017:LLA

[Bar17] Celia A. Z. Barcelos.

LAVIR — locally adaptive variational image registration. Computers and Mathematics with Applications, 73(7):1403–1413, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306836.

Ben-Ahmed:2018:RBF

[BASW18] E. Ben-Ahmed, M. Sadik, and M. Wakrim. Radial basis function partition of unity method for modelling water flow in porous media. Computers and Mathematics with Applications, 75(8):2925–2941, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300385.

Bayona:2019:IRF

[Bay19] Víctor Bayona. An insight into RBF–FD approximations augmented with polynomials. Computers and Mathematics with Applications, 77(9):2337–2353, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307272.

Boz:2008:AEF

[BB08] Ahmet Boz and Ahmet Bekir. Application of Exp-function method for (3 + 1)-dimensional nonlinear evolution equations. Computers and Mathematics with Applications, 56(5):1451–1456, September 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122108001703.

See [Asl11].

Benchohra:2010:IFD

[BB10a] Mouffak Benchohra and Farida Berhoun. Impulsive fractional differential equations with variable times. Computers and Mathematics with Applications, 59(3):1245–1252, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003575.

Berezansky:2010:OEI

[BB10b] Leonid Berezansky and Elena Braverman. Oscillation of equations with an infinite distributed delay. Computers and Mathematics with Applications, 60(9):2583–2593, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006553.

Berezansky:2012:EPS

[BB12] Leonid Berezansky and Elena Braverman. On the existence of positive solutions for systems of differential equations with a distributed delay. Computers and Mathematics with Applications, 63(7):1256–1265, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211101128X.

Beckers:2015:LHT

[BB15] Pierre Beckers and Benoit Beckers. A 66 line heat transfer finite element code to highlight the dual approach. Computers and Mathematics with Applications, 70 (10):2401–2413, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004216.

Bazhlekova:2018:UFF

[BB18] Emilia Bazhlekova and Ivan Bazhlekov. Unidirectional flows of fractional Jeffreys’ fluids: Thermodynamic constraints and subordination. Computers and Mathematics with Applications, 73(6):1363–1376, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306800.

Bhowmik:2016:STD

[BBBM16] Samir Kumar Bhowmik, Rabah Belbaki, Tahar Zamene Boulmezaoud, and Samy Mziou. Solving two dimensional second order elliptic equations in exterior domains using the inverted finite elements method. Computers and Mathematics with Applications, 72(9):2315–2333, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304849.

Ban:2011:AFN

[BBC^{+}11] A. Ban, A. Brândas, L. Coroianu, C. Negrutiu, and O. Nica. Approximations of fuzzy numbers by trapezoidal fuzzy numbers preserving the ambiguity and value. Computers and Mathematics with Applications, 61(5):1379–1401, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000101.

Breaz:2010:CPS

[BBD10] Nicoleta Breaz, Daniel Breaz, and Maslina Darus. Convexity properties for some general integral operators on uniformly analytic functions classes. Computers and Mathematics with Applications, 60(12):3105–3107, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110007911.

Barrio:2011:QNA

[BBDS11] R. Barrio, F. Blesa, A. Dena, and S. Serrano. Qualitative and numerical analysis of the Rössler model: Bifurcations of equilibria. Computers and Mathematics with Applications, 62 (11):4140–4150, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008455.

Barboteu:2019:E

[BBH^{+}19] Mikaël Barboteu, Robert Brouzet, Weimin Han, Stanislaw Migorski, and Meir Shillor. Editorial. Computers and Mathematics with Applications, 77(11):2867– 2868, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305686.

Bonetti:2019:DII

[BBL19] Elena Bonetti, Giovanna Bonfanti, and Frédéric Lebon. Derivation of imperfect interface models coupling damage and temperature. Computers and Mathematics with Applications, 77(11):2906– 2916, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830539X.

Borm:2018:ABA

[BBM18] S. Börm, C. Börst, and J. M. Melenk. An analysis of a butterfly algorithm. Computers and Mathematics with Applications, 74(9):2125–2143, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303085.

Boonyaritdachochai:2010:OCM

[BBO10] Panida Boonyaritdachochai, Chanwit Boonchuay, and Weerakorn Ongsakul. Optimal congestion management in an electricity market using particle swarm optimization with time-varying acceleration coefficients. Computers and Mathematics with Applications, 60(4):1068–1077, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002543.

Beg:2010:CPS

[BBR10a] Ismat Beg, Asma Rashid Butt, and S. Radojević. The contraction principle for set valued mappings on a metric space with a graph. Computers and Mathematics with Applications, 60(5):1214–1219, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004153.

Benitez:2010:WBT

[BBR10b] R. Benítez, V. J. Bolós, and M. E. Ramírez. A wavelet-based tool for studying non-periodicity. Computers and Mathematics with Applications, 60(3):634–641, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003597.

Bi:2010:SSH

[BC10] Jinbo Bi and Dengyuan Chen. Soliton solutions of higher-order generalized derivative nonlinear Schrödinger equation. Computers and Mathematics with Applications, 60(7):1881–1887, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004979.

Ban:2011:DTF

[BC11] Adrian I. Ban and Lucian C. Coroianu. Discontinuity of the trapezoidal fuzzy number-valued operators preserving core. Computers and Mathematics with Applications, 62(8):3103–3110, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006869.

Balibrea:2012:EPS

[BC12] Francisco Balibrea and Antonio Cascales. Eventually positive solutions in rational difference equations. Computers and Mathematics with Applications, 64(7):2275– 2281, October 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112001794.

Behera:2015:ADQ

[BC15] Laxmi Behera and S. Chakraverty. Application of differential quadrature method in free vibration analysis of nanobeams based on various nonlocal theories. Computers and Mathematics with Applications, 69(12):1444– 1462, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001790.

Ballestra:2016:FNM

[BC16] Luca Vincenzo Ballestra and Liliana Cecere. A fast numerical method to price American options under the Bates model. Computers and Mathematics with Applications, 72(5):1305–1319, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211630373X.

Bringmann:2017:HAL

[BC17] P. Bringmann and C. Carstensen. h-adaptive least-squares finite element methods for the 2D Stokes equations of any order with optimal convergence rates. Computers and Mathematics with Applications, 74(8):1923–1939, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300986.

Berriochoa:2011:SIH

[BCB11] E. Berriochoa, A. Cachafeiro, and E. Martínez Brey. Some improvements to the Hermite–Fejér interpolation on the circle and bounded interval. Computers and Mathematics with Applications, 61(4):1228–1240, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009909.

Bellouquid:2014:AAN

[BCC14] A. Bellouquid and M. Ch-Chaoui. Asymptotic analysis of a nonlinear integro-differential system modeling the immune response. Computers and Mathematics with Applications, 68(9):905–914, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002119.

Bai:2018:IIF

[BCCZ18] Jinwei Bai, Yong Cao, Yuchuan Chu, and Xu Zhang. An improved immersed finite element particle-in-cell method for plasma simulation. Computers and Mathematics with Applications, 75(6):1887– 1899, March 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304790.

Belliard:2016:AAR

[BCD^{+}16] M. Belliard, M. Chandesris, J. Dumas, Y. Gorsse, D. Jamet, C. Josserand, and B. Mathieu. An analysis and an affordable regularization technique for the spurious force oscillations in the context of direct-forcing immersed boundary methods. Computers and Mathematics with Applications, 71(5):1089–1113, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300207.

Bozzo:2010:RBS

[BCF10] Enrico Bozzo, Roberto Carniel, and Dario Fasino. Relationship between singular spectrum analysis and Fourier analysis: Theory and application to the monitoring of volcanic activity. Computers and Mathematics with Applications, 60(3):812–820, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003858.

Batkai:2013:OSN

[BCF13] András Bátkai, Petra Csomós, and Bálint Farkas. Operator splitting for nonautonomous delay equations. Computers and Mathematics with Applications, 65(3):315–324, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003823.

Boldo:2014:TCM

[BCF^{+}14] Sylvie Boldo, François Clément, Jean-Christophe Filliâtre, Micaela Mayero, Guillaume Melquiond, and Pierre Weis. Trusting computations: a mechanized proof from partial differential equations to actual program. Computers and Mathematics with Applications, 68(3):325–352, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002636.

Bazarra:2019:NAS

[BCFQ19] N. Bazarra, M. I. M. Copetti, J. R. Fernández, and R. Quintanilla. Numerical analysis of some dual-phase-lag models. Computers and Mathematics with Applications, 77(2):407–426, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830556X.

Bird:2017:FNM

[BCG17] R. E. Bird, W. M. Coombs, and S. Giani. Fast native-MATLAB stiffness assembly for SIPG linear elasticity. Computers and Mathematics with Applications, 74(12):3209–3230, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730514X.

Bai:2018:MIF

[BCH^{+}18] Jinwei Bai, Yong Cao, Xiaoming He, Hongyan Liu, and Xiaofeng Yang. Modeling and an immersed finite element method for an interface wave equation. Computers and Mathematics with Applications, 76(7):1625–1638, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303808.

Bacuta:2018:PIC

[BCHS18] Constantin Bacuta, Fioralba Cakoni, Houssem Haddar, and Jiguang Sun. Proceedings of the International Conference on Computational Mathematics and Inverse Problems honoring Peter Monk. Computers and Mathematics with Applications, 74(11):2639, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305485.

Beshley:2019:AMB

[BCJ19] Andriy Beshley, Roman Chapko, and B. Tomas Johansson. On the alternating method and boundary-domain integrals for elliptic Cauchy problems. Computers and Mathematics with Applications, 78 (11):3514–3526, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302883.

Berdyshev:2011:SIT

[BCK11] A. S. Berdyshev, A. Cabada, and B. J. Kadirkulov. The Samarskii–Ionkin type problem for the fourth order parabolic equation with fractional differential operator. Computers and Mathematics with Applications, 62 (10):3884–3893, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008066.

Berrone:2015:UQD

[BCPS15] S. Berrone, C. Canuto, S. Pieraccini, and S. Scialò. Uncertainty quantification in discrete fracture network models: Stochastic fracture transmissivity. Computers and Mathematics with Applications, 70(4):603–623, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002369.

Bravo-Castillero:2015:TRE

[BCSCB^{+}15] Julian Bravo-Castillero, Lázaro M. Sixto-Camacho, Renald Brenner, Raúl Guinovart-Díaz, Leslie D. Pérez-Fernández, Reinaldo Rodríguez-Ramos, and Federico J. Sabina. Temperaturerelated effective properties and exact relations for thermo-magneto-electro-elastic fibrous composites. Computers and Mathematics with Applications, 69(9):980–996, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001005.

Baculikova:2011:OTSa

[BD11a] B. Baculíková and J. Dzurina. Oscillation theorems for second order neutral differential equations. Computers and Mathematics with Applications, 61(1):94–99, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008254.

Baculikova:2011:OTSb

[BD11b] B. Baculíková and J. Dzurina. Oscillation theorems for second-order nonlinear neutral differential equations. Computers and Mathematics with Applications, 62 (12):4472–4478, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008881.

Bi:2011:SST

[BD11c] Hui Bi and Shusen Ding. Some strong (p, q)-type inequalities for the homotopy operator. Computers and Mathematics with Applications, 62(4):1780–1789, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004998.

Bianca:2011:MGM

[BD11d] C. Bianca and M. Delitala. On the modelling of genetic mutations and immune system competition. Computers and Mathematics with Applications, 61(9):2362–2375, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000411.

Bertaccini:2016:IPS

[BD16] Daniele Bertaccini and Fabio Durastante. Interpolating preconditioners for the solution of sequence of linear systems. Computers and Mathematics with Applications, 72(4):1118–1130, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303558.

Bonnet:2019:ECM

[BD19] Marc Bonnet and Edouard Demaldent. The eddy current model as a lowfrequency, high-conductivity asymptotic form of the Maxwell transmission problem. Computers and Mathematics with Applications, 77(8):2145–2161, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306989.

Borisov:2012:SAB

[BDB12] Milen Borisov, Neli Dimitrova, and Venko Beschkov. Stability analysis of a bioreactor model for biodegradation of xenobiotics. Computers and Mathematics with Applications, 64(3):361–373, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001927.

Bezchlebova:2016:DGM

[BDF16] Eva Bezchlebová, Vít Dolejsí, and Miloslav Feistauer. Discontinuous Galerkin method for the solution of a transport level-set problem. Computers and Mathematics with Applications, 72(3):455–480, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302267.

Bhalekar:2011:FBE

[BDGBM11] Sachin Bhalekar, Varsha Daftardar-Gejji, Dumitru Baleanu, and Richard Magin. Fractional Bloch equation with delay. Computers and Mathematics with Applications, 61(5):1355–1365, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100006X.

Bhalekar:2012:TCF

[BDGBM12] Sachin Bhalekar, Varsha Daftardar-Gejji, Dumitru Baleanu, and Richard Magin. Transient chaos in fractional Bloch equations. Computers and Mathematics with Applications, 64(10):3367– 3376, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112000909.

Bochev:2014:MRL

[BDGG14] Pavel Bochev, Leszek Demkowicz, Jay Gopalakrishnan, and Max Gunzburger. Minimum residual and least squares finite element methods. Computers and Mathematics with Applications, 68(11):1479, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005331.

Bermudez:2013:FEA

[BDGS13] A. Bermúdez, O. Domínguez, D. Gómez, and P. Salgado. Finite element approximation of nonlinear transient magnetic problems involving periodic potential drop excitations. Computers and Mathematics with Applications, 65(8):1200–1219, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001417.

Bonazzoli:2018:EEI

[BDHR18] Marcella Bonazzoli, Victorita Dolean, Frédéric Hecht, and Francesca Rapetti. An example of explicit implementation strategy and preconditioning for the high order edge finite elements applied to the timeharmonic Maxwell’s equations. Computers and Mathematics with Applications, 75(5):1498–1514, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307277.

Bokal:2011:CQE

[BDK^{+}11] Drago Bokal, Matt De-Vos, Sandi Klavzar, Aki Mimoto, and Arne Ø. Mooers. Computing quadratic entropy in evolutionary trees. Computers and Mathematics with Applications, 62 (10):3821–3828, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100798X.

Barboteu:2019:NAC

[BDM19a] Mikaël Barboteu, Nacera Djehaf, and Matthieu Martel. Numerically accurate code synthesis for Gauss pivoting method to solve linear systems coming from mechanics. Computers and Mathematics with Applications, 77(11):2883–2893, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304401.

Bartos:2019:GOA

[BDM^{+}19b] Ondrej Bartos, Vít Dolejsí, Georg May, Ajay Rangarajan, and Filip Roskovec. A goal-oriented anisotropic hp-mesh adaptation method for linear convection– diffusion-reaction problems. Computers and Mathematics with Applications, 78(9):2973–2993, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301774.

Bonanno:2011:NTS

[BDO11] Gabriele Bonanno, Beatrice Di Bella, and Donal O’Regan. Non-trivial solutions for nonlinear fourth-order elastic beam equations. Computers and Mathematics with Applications, 62(4):1862–1869, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005086.

Berrone:2012:CTM

[BDPM12] Stefano Berrone, Francesca De Santi, Sandra Pieraccini, and Massimo Marro. Coupling traffic models on networks and urban dispersion models for simulating sustainable mobility strategies. Computers and Mathematics with Applications, 64 (6):1975–1991, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002593.

Boscheri:2019:FSM

[BDR19] Walter Boscheri, Michael Dumbser, and Maurizio Righetti. FORCE schemes on moving unstructured meshes for hyperbolic systems. Computers and Mathematics with Applications, 78(2): 362–380, July 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118305182.

Berezansky:2010:PSS

[BDS10] Leonid Berezansky, Josef Diblík, and Zdenek Smarda. Positive solutions of second-order delay differential equations with a damping term. Computers and Mathematics with Applications, 60 (5):1332–1342, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004281.

Belhaj:2015:FAS

[BDS15] Skander Belhaj, Marwa Dridi, and Ahmed Salam. A fast algorithm for solving banded Toeplitz systems. Computers and Mathematics with Applications, 70(12):2958– 2967, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004939.

Bartosz:2017:NAD

[BDS17] Krzysztof Bartosz, David Danan, and Pawel Szafraniec. Numerical analysis of a dynamic bilateral thermoviscoelastic contact problem with nonmonotone friction law. Computers and Mathematics with Applications, 73(5):727–746, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300032.

Biazar:2011:NHP

[BE11] Jafar Biazar and Mostafa Eslami. A new homotopy perturbation method for solving systems of partial differential equations. Computers and Mathematics with Applications, 62(1):225–234, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003737.

Biazar:2012:CWA

[BE12] J. Biazar and H. Ebrahimi.

Chebyshev wavelets approach for nonlinear systems of Volterra integral equations. Computers and Mathematics with Applications, 63(3):608–616, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008376.

Brezina:2018:FAI

[BE18] Jan Brezina and Pavel Exner. Fast algorithms for intersection of non-matching grids using Plücker coordinates. Computers and Mathematics with Applications, 74(1):174–187, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300792.

Barakat:2011:EPI

[BEAA11] H. M. Barakat, Magdy E. El-Adll, and Amany E. Aly. Exact prediction intervals for future exponential lifetime based on random generalized order statistics. Computers and Mathematics with Applications, 61(5):1366–1378, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000071.

Bencheva:2012:CMH

[Ben12] Gergana Bencheva. Computer modelling of haematopoietic stem cells migration. Computers and Mathematics with Applications, 64(3):337–349, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001708.

Bentley:2017:ECC

[Ben17] Alistair Bentley. Explicit construction of computational bases for RTk and BDMk spaces in R^{3}. Computers and Mathematics with Applications, 73(7):1421–1432, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300470.

Berinde:2012:CCP

[Ber12] Vasile Berinde. Coupled coincidence point theorems for mixed monotone nonlinear operators. Computers and Mathematics with Applications, 64(6):1770–1777, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001265.

Bernal:2016:TRM

[Ber16] Francisco Bernal. Trustregion methods for nonlinear elliptic equations with radial basis functions. Computers and Mathematics with Applications, 72(7):1743–1763, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304060.

Bianca:2011:BDM

[BF11] Carlo Bianca and Luisa Fermo. Bifurcation diagrams for the moments of a kinetic type model of keloid-immune system competition. Computers and Mathematics with Applications, 61(2):277– 288, January 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008527.

Bertaccini:2016:SAI

[BF16] Daniele Bertaccini and Salvatore Filippone. Sparse approximate inverse preconditioners on high performance GPU platforms. Computers and Mathematics with Applications, 71(3):693–711, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005763.

Buyuktahtakin:2011:DMC

[BFF^{+}11] I. Esra Büyüktahtakin, Zhuo Feng, George Frisvold, Ferenc Szidarovszky, and Aaryn Olsson. A dynamic model of controlling invasive species. Computers and Mathematics with Applications, 62 (9):3326–3333, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007000.

Bretto:2011:NGR

[BFG11] Alain Bretto, Alain Faisant, and Luc Gillibert. New graphs related to (p, 6) and (p, 8)-cages. Computers and Mathematics with Applications, 62 (6):2472–2479, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005876.

Budinski:2015:LBM

[BFS15] Ljubomir Budinski, Julius Fabian, and Matija Stipić. Lattice Boltzmann method for groundwater flow in non-orthogonal structured lattices. Computers and Mathematics with Applications, 70 (10):2601–2615, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004678.

Bremner:2012:JTD

[BFSO12] Murray R. Bremner, Raúl Felipe, and Juana Sánchez-Ortega. Jordan triple disystems. Computers and Mathematics with Applications, 63(6):1039–1055, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010510.

Biazar:2010:NTO

[BG10a] Jafar Biazar and Behzad Ghanbari. A new thirdorder family of nonlinear solvers for multiple roots. Computers and Mathematics with Applications, 59(10):3315– 3319, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001926.

Black:2010:NIO

[BG10b] Kelly Black and John B. Geddes. Noise-induced oscillations in an actively mode-locked laser. Computers and Mathematics with Applications, 60(1): 1–13, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002877.

Bohner:2010:LTI

[BG10c] Martin Bohner and Gusein Sh. Guseinov. The Laplace transform on isolated time scales. Computers and Mathematics with Applications, 60 (6):1536–1547, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004517.

Biazar:2011:NIH

[BG11] Jafar Biazar and Behzad Ghanbari. Notes on “An improvement to homotopy perturbation method for solving system of linear equations”. Computers and Mathematics with Applications, 61(6):1704, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000125.

See [Yus09].

Bykov:2013:FSI

[BG13] V. Bykov and V. Gol’dshtein.

Fast and slow invariant manifolds in chemical kinetics. Computers and Mathematics with Applications, 65(10):1502– 1515, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000709.

Bochev:2014:SML

[BG14] Pavel Bochev and Marc Gerritsma. A spectral mimetic least-squares method. Computers and Mathematics with Applications, 68(11):1480– 1502, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114004623.

Barros:2015:GSL

[BG15] Saulo R. M. Barros and Claudia I. Garcia. A global semi-Lagrangian model for the adiabatic primitive equations on locally refined grids. Computers and Mathematics with Applications, 69(8):725–742, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122115000930.

Benaissa:2019:ASL

[BG19] Abbes Benaissa and Soumia Gaouar. Asymptotic stability for the Lamé system with fractional boundary damping. Computers and Mathematics with Applications, 77(5):1331–1346, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306643.

Barki:2015:ERE

[BGF15] Hichem Barki, Gaël Guennebaud, and Sebti Foufou. Exact, robust, and efficient regularized Booleans on general 3D meshes. Computers and Mathematics with Applications, 70(6):1235–1254, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003028.

Barrio-Garrido:2016:SAF

[BGGCGRSP16] R. M. Barrio-Garrido, L. E. Garcia-Castillo, I. Gomez-Revuelto, and M. Salazar-Palma. Selfadaptive hp finite element method with iterative mesh truncation technique accelerated with adaptive cross approximation. Computers and Mathematics with Applications, 71(10):1911– 1932, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300840.

Bouma:2014:CRD

[BGH14] Timaeus Bouma, Jay Gopalakrishnan, and Ammar Harb. Convergence rates of the DPG method with reduced test space degree. Computers and Mathematics with Applications, 68 (11):1550–1561, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003642.

Barrera:2013:CTM

[BGIN13] D. Barrera, A. Guessab, M. J. Ibáñez, and O. Nouisser. Construction techniques for multivariate modified quasi-interpolants with high approximation order. Computers and Mathematics with Applications, 65(1):29–41, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006074.

Bozzini:2015:WNA

[BGL^{+}15] B. Bozzini, G. Gambino, D. Lacitignola, S. Lupo, M. Sammartino, and I. Sgura. Weakly nonlinear analysis of Turing patterns in a morphochemical model for metal growth. Computers and Mathematics with Applications, 70(8):1948–1969, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003880.

Bartel:2019:TSS

[BGM19a] Thorsten Bartel, Isabelle Guschke, and Andreas Menzel. Towards the simulation of Selective Laser Melting processes via phase transformation models. Computers and Mathematics with Applications, 78(7):2267–2281, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118304528.

Breil:2019:SOE

[BGM19b] Jérôme Breil, Gabriel Georges, and Pierre-Henri Maire. Secondorder extension in space and time for a 3D cell-centered Lagrangian scheme. Computers and Mathematics with Applications, 78(2):381–401, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303389.

Borrelli:2013:NST

[BGP13] Alessandra Borrelli, Giulia Giantesio, and Maria Cristina Patria. Numerical simulations of three-dimensional MHD stagnation-point flow of a micropolar fluid. Computers and Mathematics with Applications, 66(4):472–489, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003453.

Biazar:2011:HHP

[BGPP11] Jafar Biazar, Behzad Ghanbari, Mehdi Gholami Porshokouhi, and Mohammad Gholami Porshokouhi. He’s homotopy perturbation method: a strongly promising method for solving nonlinear systems of the mixed Volterra–Fredholm integral equations. Computers and Mathematics with Applications, 61(4):1016–1023, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009569.

Boffi:2014:MFI

[BGR14] Daniele Boffi, Lucia Gastaldi, and Michele Ruggeri. Mixed formulation for interface problems with distributed Lagrange multiplier. Computers and Mathematics with Applications, 68 (12):2151–2166, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003381.

Bayon:2011:ACA

[BGRS11] L. Bayón, J. M. Grau, M. M. Ruiz, and P. M. Suárez. Algorithm for calculating the analytic solution for economic dispatch with multiple fuel units. Computers and Mathematics with Applications, 62(5):2225–2234, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005621.

Bermudez:2015:NAT

[BGRV15] Alfredo Bermúdez, Dolores Gómez, Rodolfo Rodríguez, and Pablo Venegas. Numerical analysis of a transient nonlinear axisymmetric eddy current model. Computers and Mathematics with Applications, 70(8):1984–2005, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003867.

Bao:2010:EMS

[BH10] Jianhai Bao and Zhenting Hou. Existence of mild solutions to stochastic neutral partial functional differential equations with non-Lipschitz coefficients. Computers and Mathematics with Applications, 59(1):207–214, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005707.

Boufoussi:2011:FDE

[BH11] B. Boufoussi and S. Hajji. Functional differential equations driven by a fractional Brownian motion. Computers and Mathematics with Applications, 62(2):746– 754, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004664.

Bihlo:2014:PSM

[BH14] Alexander Bihlo and Ronald D. Haynes. Parallel stochastic methods for PDE based grid generation. Computers and Mathematics with Applications, 68(8):804–820, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003356.

Burns:2016:FST

[BHH16] John A. Burns, Xiaoming He, and Weiwei Hu. Feedback stabilization of a thermal fluid system with mixed boundary control. Computers and Mathematics with Applications, 71(11):2170– 2191, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000158.

Bouhamidi:2014:MRM

[BHJ14] A. Bouhamidi, M. Hached, and K. Jbilou. A meshless RBF method for computing a numerical solution of unsteady burgers’-type equations. Computers and Mathematics with Applications, 68(3):238–256, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400248X.

Bae:2016:SIF

[BHK16] Hyeong-Ohk Bae, Seung-Yeal Ha, and Yongsik Kim. Simulation of interaction of flocking particles and an incompressible fluid. Computers and Mathematics with Applications, 71(10):2020– 2033, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301663.

Benchohra:2012:WSH

[BHM12] Mouffak Benchohra, Johnny Henderson, and Fatima-Zohra Mostefai. Weak solutions for hyperbolic partial fractional differential inclusions in Banach spaces. Computers and Mathematics with Applications, 64 (10):3101–3107, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011114.

Barboteu:2019:NAV

[BHM19] Mikaël Barboteu, Weimin Han, and Stanisław Migórski. On numerical approximation of a variationalhemivariational inequality modeling contact problems for locking materials. Computers and Mathematics with Applications, 77(11):2894– 2905, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830422X.

Bhowmik:2014:CFE

[Bho14a] Samir Kumar Bhowmik. Corrigendum to “Fast and efficient numerical methods for an extended Black–Scholes model” [Comput. Math. Appl. (2014) 636–654]. Computers and Mathematics with Applications, 67(10): 2027, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001540.

See [Bho14b].

Bhowmik:2014:FEN

[Bho14b] Samir Kumar Bhowmik. Fast and efficient numerical methods for an extended Black–Scholes model. Computers and Mathematics with Applications, 67(3):636–654, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006895.

See corrigendum [Bho14a].

Bai:2019:FHM

[BHZJ19] Yu Bai, Lamei Huo, Yan Zhang, and Yuehua Jiang. Flow, heat and mass transfer of three-dimensional fractional Maxwell fluid over a bidirectional stretching plate with fractional Fourier’s law and fractional Fick’s law. Computers and Mathematics with Applications, 78(8):2831–2846, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302433.

Baskin:2012:FEE

[BI12] E. Baskin and A. Iomin.

Fractional electrostatic equations in fractal composite structures. Computers and Mathematics with Applications, 64 (10):3302–3309, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001046.

Bica:2011:NMS

[Bic11] Alexandru Mihai Bica.

The numerical method of successive interpolations for two-point boundary value problems with deviating argument. Computers and Mathematics with Applications, 62 (10):3829–3843, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008005.

Biswas:2010:SSE

[Bis10] Anjan Biswas. 1-soliton solution of the K(m, n) equation with generalized evolution and timedependent damping and dispersion. Computers and Mathematics with Applications, 59(8):2536– 2540, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000246.

Bistrian:2014:SPN

[Bis14] Diana A. Bistrian. A solution of the parabolized Navier–Stokes stability model in discrete space by two-directional differential quadrature and application to swirl intense flows. Computers and Mathematics with Applications, 68(3):197–211, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002107.

Bacuta:2019:SPL

[BJ19] Constantin Bacuta and Jacob Jacavage. Saddle point least squares preconditioning of mixed methods. Computers and Mathematics with Applications, 77(5):1396–1407, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306667.

Bai:2012:TMS

[BJLZ12] Kun Bai, Jingjing Ji, Kok-Meng Lee, and Shuyou Zhang. A twomode six-DOF motion system based on a balljoint-like spherical motor for haptic applications. Computers and Mathematics with Applications, 64(5):978–987, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001629.

Bian:2016:ABT

[BJPT16] Lei Bian, Songsong Ji, Gang Pang, and Shaoqiang Tang. Accurate boundary treatment for transient Schrödinger equation under polar coordinates. Computers and Mathematics with Applications, 71(2):479–488, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005866.

Bacuta:2018:SPL

[BJQS18] Constantin Bacuta, Jacob Jacavage, Klajdi Qirko, and Francisco-Javier Sayas. Saddle point least squares iterative solvers for the time harmonic Maxwell equations. Computers and Mathematics with Applications, 74 (11):2915–2928, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730528X.

Badri:2019:PKS

[BJRF19] M. A. Badri, P. Jolivet, B. Rousseau, and Y. Favennec. Preconditioned Krylov subspace methods for solving radiative transfer problems with scattering and reflection. Computers and Mathematics with Applications, 77(6):1453–1465, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305534.

Bentbib:2015:SKS

[BJS15] A. Bentbib, K. Jbilou, and E. M. Sadek. On some Krylov subspace based methods for largescale nonsymmetric algebraic Riccati problems. Computers and Mathematics with Applications, 70(10):2555–2565, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004654.

Bugajewska:2010:EUT

[BK10] Daria Bugajewska and Piotr Kasprzak. On the existence, uniqueness and topological structure of solution sets to a certain fractional differential equation. Computers and Mathematics with Applications, 59(3):1108–1116, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004180.

Balachandran:2011:ERF

[BK11a] K. Balachandran and S. Kiruthika. Existence results for fractional integrodifferential equations with nonlocal condition via resolvent operators. Computers and Mathematics with Applications, 62(3):1350–1358, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003890.

Basar:2011:ACG

[BK11b] Feyzi Basar and Murat Kirisçi. Almost convergence and generalized difference matrix. Computers and Mathematics with Applications, 61(3):602–611, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009119.

Balachandran:2012:ESA

[BK12a] K. Balachandran and S. Kiruthika. Existence of solutions of abstract fractional integrodifferential equations of Sobolev type. Computers and Mathematics with Applications, 64 (10):3406–3413, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011072.

Buckwar:2012:NND

[BK12b] E. Buckwar and C. Kelly.

Non-normal drift structures and linear stability analysis of numerical methods for systems of stochastic differential equations. Computers and Mathematics with Applications, 64(7):2282–2293, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001848.

Braverman:2013:SSR

[BK13] E. Braverman and I. M.

Karabash. Structured stability radii and exponential stability tests for Volterra difference systems. Computers and Mathematics with Applications, 66(11):2259– 2280, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122113003908.

Benner:2014:CRL

[BK14] Peter Benner and Patrick Kürschner. Computing real low-rank solutions of Sylvester equations by the factored ADI method. Computers and Mathematics with Applications, 67(9):1656–1672, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001278.

Brennan:2015:FEA

[BK15] Brian Brennan and Robert C. Kirby. Finite element approximation and preconditioners for a coupled thermal-acoustic model. Computers and Mathematics with Applications, 70(10):2342–2354, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004186.

Braverman:2016:CCM

[BK16] E. Braverman and Md. Kamrujjaman. Competitivecooperative models with various diffusion strategies. Computers and Mathematics with Applications, 72(3):653–662, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302851.

Biswas:2018:MED

[BK18] Sougata Biswas and Jiten C. Kalita. Moffatt eddies in the driven cavity: a quantification study by an HOC approach. Computers and Mathematics with Applications, 76(3):471–487, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302360.

Bobal:2013:APC

[BKDM13] Vladimir Bobal, Marek Kubalcik, Petr Dostal, and Jakub Matejicek. Adaptive predictive control of time-delay systems. Computers and Mathematics with Applications, 66(2):165–176, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000655.

Bagheri:2018:ATB

[BKE18] H. Bagheri, Y. Kiani, and M. R. Eslami. Asymmetric thermal buckling of temperature dependent annular FGM plates on a partial elastic foundation. Computers and Mathematics with Applications, 75(5):1566–1581, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307472.

Bhattacharyya:2011:FRP

[BKK11] Rupak Bhattacharyya, Pankaj Kumar, and Samarjit Kar. Fuzzy R&D portfolio selection of interdependent projects. Computers and Mathematics with Applications, 62(10):3857– 3870, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008042.

Borsche:2012:CMP

[BKK12] Raul Borsche, Mark Kimathi, and Axel Klar. A class of multi-phase traffic theories for microscopic, kinetic and continuum traffic models. Computers and Mathematics with Applications, 64(9):2939–2953, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005639.

Baroudi:2014:EIC

[BKL14] Lina Baroudi, Masahiro Kawaji, and Taehun Lee. Effects of initial conditions on the simulation of inertial coalescence of two drops. Computers and Mathematics with Applications, 67(2):282–289, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002551.

Bohacek:2019:GSS

[BKL^{+}19] Jan Bohacek, Abdellah Kharicha, Andreas Ludwig, Menghuai Wu, Tobias Holzmann, and Ebrahim Karimi-Sibaki. A GPU solver for symmetric positive-definite matrices vs. traditional codes. Computers and Mathematics with Applications, 78(9):2933–2943, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301105.

Bhattacharyya:2011:FMV

[BKM11] Rupak Bhattacharyya, Samarjit Kar, and Dwijesh Dutta Majumder. Fuzzy mean-variance-skewness portfolio selection models by interval analysis. Computers and Mathematics with Applications, 61(1):126–137, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008291.

Berntsson:2014:AAP

[BKMT14] F. Berntsson, V. A. Kozlov, L. Mpinganzima, and B. O. Turesson. An accelerated alternating procedure for the Cauchy problem for the Helmholtz equation. Computers and Mathematics with Applications, 68(1–2):44–60, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001837.

Borsche:2016:MFM

[BKMT16] R. Borsche, A. Klar, A. Meurer, and O. Tse. Mean field models for interacting ellipsoidal particles. Computers and Mathematics with Applications, 72(3):704–719, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302978.

Berntsson:2017:ITR

[BKMT17] F. Berntsson, V. A. Kozlov, L. Mpinganzima, and B. O. Turesson. Iterative Tikhonov regularization for the Cauchy problem for the Helmholtz equation. Computers and Mathematics with Applications, 73(1):163–172, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306113.

Both:2019:AAF

[BKNR19] Jakub Wiktor Both, Kundan Kumar, Jan Martin Nordbotten, and Florin Adrian Radu. Anderson accelerated fixedstress splitting schemes for consolidation of unsaturated porous media. Computers and Mathematics with Applications, 77(6):1479–1502, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304048.

Bruni:2011:MPS

[BKP11] C. Bruni, G. Koch, and F. Papa. A measurement policy in stochastic linear filtering problems. Computers and Mathematics with Applications, 61(3):546–566, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009053.

Bridges:2010:MRS

[BKR10] Craig Bridges, Satish Karra, and K. R. Rajagopal. On modeling the response of the synovial fluid: Unsteady flow of a shear-thinning, chemically-reacting fluid mixture. Computers and Mathematics with Applications, 60(8):2333–2349, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005912.

Balasubramaniam:2011:SEF

[BKR11]  P.  Balasubramaniam,

M. Kalpana, and R. Rakkiyappan. State estimation for fuzzy cellular neural networks with time delay in the leakage term, discrete and unbounded distributed delays. Computers and Mathematics with Applications, 62 (10):3959–3972, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008261.

Borregales:2019:PPT

[BKR^{+}19] Manuel Borregales, Kundan Kumar, Florin Adrian Radu, Carmen Rodrigo, and Francisco José Gaspar. A partially parallelin-time fixed-stress splitting method for Biot’s consolidation model. Computers and Mathematics with Applications, 77(6): 1466–1478, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118305091.

Balachandran:2011:FIE

[BKT11] K. Balachandran, S. Kiruthika, and J. J. Trujillo. On fractional impulsive equations of Sobolev type with nonlocal condition in Banach spaces. Computers and Mathematics with Applications, 62(3):1157–1165, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001830.

Balachandran:2012:RCFa

[BKT12] K. Balachandran, J. Kokila, and J. J. Trujillo. Relative controllability of fractional dynamical systems with multiple delays in control. Computers and Mathematics with Applications, 64 (10):3037–3045, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000922.

Bildik:2010:CLP

[BKY10] Necdet Bildik, Ali Konuralp, and Salih Yalçinbas. Comparison of Legendre polynomial approximation and variational iteration method for the solutions of general linear Fredholm integrodifferential equations.

Computers and Mathematics with Applications, 59(6):1909–1917, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004131.

Burger:2017:ALW

[BKZ17] Raimund Bürger, Sudarshan Kumar Kenettinkara, and David Zorío. Approximate Lax–Wendroff discontinuous Galerkin methods for hyperbolic conservation laws. Computers and Mathematics with Applications, 74 (6):1288–1310, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730367X.

Balaj:2010:MVR

[BL10] Mircea Balaj and Dinh The Luc. On mixed variational relation problems. Computers and Mathematics with Applications, 60(9):2712–2722, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110007121.

Bai:2011:FSS

[BL11] Li Bai and Min Liu. Fuzzy sets and similarity relations for Semantic Web service matching. Computers and Mathematics with Applications, 61(8):2281–2286, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007467.

Braverman:2012:SEU

[BL12] Elena Braverman and Eduardo Liz. On stabilization of equilibria using predictive control with and without pulses. Computers and Mathematics with Applications, 64(7):2192–2201, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000235.

Bettebghor:2014:ORB

[BL14] Dimitri Bettebghor and François-Henri Leroy. Overlapping radial basis function interpolants for spectrally accurate approximation of functions of eigenvalues with application to buckling of composite plates. Computers and Mathematics with Applications, 67(10):1816–1836, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001539.

Boyadjiev:2017:MDF

[BL17] Lyubomir Boyadjiev and Yuri Luchko. Multidimensional α-fractional diffusion-wave equation and some properties of its fundamental solution. Computers and Mathematics with Applications, 73(12):2561–2572, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301827.

Bao:2018:TWE

[BL18] Xiongxiong Bao and Jia Liu. Traveling waves for epidemic models with nonlocal dispersal in time and space periodic habitats. Computers and Mathematics with Applications, 75(7):2404–2413, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307800.

Bettahar:2017:PBE

[BLS17] Salim Bettahar, Patrick Lambert, and Amine Boudghene Stambouli. PDE-based efficient method for colour image restoration. Computers and Mathematics with Applications, 74(3):577–590, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303000.

Banz:2018:HOF

[BLS18] Lothar Banz, Bishnu P. Lamichhane, and Ernst P. Stephan. Higher order FEM for the obstacle problem of the p-Laplacian — a variational inequality approach. Computers and Mathematics with Applications, 76(7):1639–1660, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830381X.

Brenner:2018:MMS

[BLyS18] Susanne C. Brenner, Hengguang Li, and Li yeng Sung. Multigrid methods for saddle point problems: Oseen system. Computers and Mathematics with Applications, 74 (9):2056–2067, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303644.

Baleanu:2010:GES

[BM10a] Dumitru Baleanu and Octavian G. Mustafa. On the global existence of solutions to a class of fractional differential equations. Computers and Mathematics with Applications, 59(5):1835–1841, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005616.

Bardaro:2010:QAF

[BM10b] Carlo Bardaro and Ilaria Mantellini. A quantitative asymptotic formula for a general class of discrete operators. Computers and Mathematics with Applications, 60 (10):2859–2870, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007406.

Babolian:2011:NMS

[BM11a] E. Babolian and M. Mordad. A numerical method for solving systems of linear and nonlinear integral equations of the second kind by hat basis functions. Computers and Mathematics with Applications, 62(1):187– 198, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003695.

Bardaro:2011:APL

[BM11b] Carlo Bardaro and Ilaria Mantellini. Approximation properties for linear combinations of moment type operators. Computers and Mathematics with Applications, 62 (5):2304–2313, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005712.

Baduraliya:2012:EMA

[BM12a] Chaminda H. Baduraliya and Xuerong Mao. The Euler–Maruyama approximation for the asset price in the mean-reverting-theta stochastic volatility model. Computers and Mathematics with Applications, 64(7):2209–2223, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000478.

Brilleaud:2012:MHT

[BM12b] Martine Brilleaud and Marie-Laurence Mazure. Mixed hyperbolic/trigonometric spaces for design. Computers and Mathematics with Applications, 64(8):2459–2477, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200418X.

Bhattacharyya:2013:SLI

[BM13a] A. Bhattacharyya and B. Mukhopadhyay. Study of linear isotropic micropolar plate in an asymptotic approach. Computers and Mathematics with Applications, 66(6):1047–1057, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004288.

Boyd:2013:QGI

[BM13b] John P. Boyd and Philip W. McCauley. Quartic Gaussian and inverse-quartic Gaussian radial basis functions: the importance of a nonnegative Fourier transform. Computers and Mathematics with Applications, 65(1):75–88, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006323.

Batwa:2018:SLT

[BM18a] Sumayah Batwa and Wen-Xiu Ma. A study of lump-type and interaction solutions to a (3 + 1)-dimensional Jimbo– Miwa-like equation. Computers and Mathematics with Applications, 76(7):1576–1582, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118303730.

Bergmann:2018:FFB

[BM18b] Ronny Bergmann and Dennis Merkert. A framework for FFTbased homogenization on anisotropic lattices. Computers and Mathematics with Applications, 76(1):125–140, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302062.

Boubendir:2018:NOD

[BM18c] Yassine Boubendir and Dawid Midura. Nonoverlapping domain decomposition algorithm based on modified transmission conditions for the Helmholtz equation. Computers and Mathematics with Applications, 75(6):1900–1911, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730456X.

Bouchlaghem:2018:ANS

[BM18d] Mohammed Bouchlaghem and El Bekkaye Mermri. Analysis and numerical study of a mixed formulation of a two membranes problem. Computers and Mathematics with Applications, 76 (6):1284–1300, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303468.

Brzezniak:2019:SLL

[BM19] Zdzisław Brzeźniak and Utpal Manna. Stochastic Landau–Lifshitz–Gilbert equation with anisotropy energy driven by pure jump noise. Computers and Mathematics with Applications, 77(6):1503– 1512, March 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118304279.

Baleanu:2011:AIO

[BMA11] Dumitru Băleanu, Octavian G. Mustafa, and Ravi P. Agarwal. Asymptotic integration of (1 + α)-order fractional differential equations. Computers and Mathematics with Applications, 62(3):1492–1500, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001738.

Balduzzi:2018:NPT

[BMAR18] Giuseppe Balduzzi, Simone Morganti, Ferdinando Auricchio, and Alessandro Reali. Nonprismatic Timoshenkolike beam model: Numerical solution via isogeometric collocation. Computers and Mathematics with Applications, 74(7):1531–1541, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302560.

Baleanu:2013:FDA

[BMC13] Dumitru Baleanu, J. A. Tenreiro Machado, and Wen Chen. Fractional differentiation and its applications I. Computers and Mathematics with Applications, 66(5):575, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003714.

Bu:2019:SSO

[BMH19] Linlin Bu, Liquan Mei, and Yan Hou. Stable second-order schemes for the space-fractional Cahn–Hilliard and Allen– Cahn equations. Computers and Mathematics with Applications, 78 (11):3485–3500, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302792.

Barbosa:2010:EFO

[BMJ10] Ramiro S. Barbosa, J. A. Tenreiro Machado, and Isabel S. Jesus. Effect of fractional orders in the velocity control of a servo system. Computers and Mathematics with Applications, 59(5):1679–1686, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005446.

Bai:2019:MRQ

[BMJ19] Zhong-Zhi Bai, Cun-Qiang Miao, and Shuai Jian. On multistep Rayleigh quotient iterations for Hermitian eigenvalue problems. Computers and Mathematics with Applications, 77(9):2396–2406, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307235.

Bin-Mohsin:2011:MFS

[BML11] B. Bin-Mohsin and D. Lesnic.

The method of fundamental solutions for Helmholtz-type equations in composite materials. Computers and Mathematics with Applications, 62(12):4377–4390, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008704.

Bera:2012:IMF

[BMM12a] U. K. Bera, M. K. Maiti, and M. Maiti. Inventory model with fuzzy lead-time and dynamic demand over finite time horizon using a multiobjective genetic algorithm. Computers and Mathematics with Applications, 64(6):1822–1838, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S089812211200185X.

Berkani:2012:OCB

[BMM12b] Samira Berkani, Farida Manseur, and Ahmed Maidi. Optimal control based on the variational iteration method. Computers and Mathematics with Applications, 64(4):604–610, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011370.

Brugiapaglia:2015:CSN

[BMP15] S. Brugiapaglia, S. Micheletti, and S. Perotto. Compressed solving: a numerical approximation technique for elliptic PDEs based on compressed sensing. Computers and Mathematics with Applications, 70(6):1306–1335, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003508.

Babolian:2010:TDT

[BMRA10] E. Babolian, K. Maleknejad, M. Roodaki, and H. Almasieh. Twodimensional triangular functions and their applications to nonlinear 2D Volterra–Fredholm integral equations. Computers and Mathematics with Applications, 60 (6):1711–1722, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004773.

Bistarelli:2012:SBF

[BMS12] Stefano Bistarelli, Fabio Martinelli, and Francesco Santini. A semiringbased framework for the deduction/abduction reasoning in access control with weighted credentials. Computers and Mathematics with Applications, 64(4):447–462, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010728.

Bisi:2013:MTF

[BMS13] M. Bisi, G. Martalò, and G. Spiga. Multitemperature fluid-dynamic model equations from kinetic theory in a reactive gas: the steady shock problem. Computers and Mathematics with Applications, 66(8):1403–1417, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005099.

Barlow:2019:COF

[BMS19] Andrew Barlow, Nathaniel Morgan, and Mikhail Shashkov. Constrained optimization framework for interface-aware subscale dynamics discrete closure model for multimaterial cells in Lagrangian cell-centered hydrodynamics. Computers and Mathematics with Applications, 78(2):541– 564, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303390.

Bouadjila:2018:NAF

[BMSS18] Khaled Bouadjila, Abdelhafid Mokrane, Ali Samir Saad, and Mazen Saad. Numerical analysis of a finite volume scheme for two incompressible phase flow with dynamic capillary pressure. Computers and Mathematics with Applications, 75(10):3614–3631, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301020.

Basile:2012:NMC

[BMTV12] M. Basile, E. Messina, W. Themistoclakis, and A. Vecchio. A numerical method for a class of nonlinear integro-differential equations on the half line. Computers and Mathematics with Applications, 64(7):2354–2363, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004373.

Biloti:2013:SNG

[BMY13] R. Biloti, L. C. Matioli, and Jinyun Yuan. A short note on a generalization of the Givens transformation. Computers and Mathematics with Applications, 66(1):56– 61, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002228.

Bajpai:2014:TGF

[BN14a] S. Bajpai and N. Nataraj.

On a two-grid finite element scheme combined with Crank–Nicolson method for the equations of motion arising in the Kelvin– Voigt model. Computers and Mathematics with Applications, 68 (12):2277–2291, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003290.

Boyanova:2014:ENS

[BN14b] P. Boyanova and M. Neytcheva. Efficient numerical solution of discrete multicomponent Cahn–Hilliard systems. Computers and Mathematics with Applications, 67(1):106–121, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006251.

Bespalov:2016:PEA

[BN16] A. Bespalov and S. Nicaise. A priori error analysis of the BEM with graded meshes for the electric field integral equation on polyhedral surfaces. Computers and Mathematics with Applications, 71(8):1636–1644, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301298.

Breckling:2018:SSN

[BNP18] Sean Breckling, Monika Neda, and Fran Pahlevani. A sensitivity study of the Navier– Stokes-α model. Computers and Mathematics with Applications, 75(2):666–689, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306168.

Balsim:2010:HSM

[BNR10] Igor Balsim, Mathew A.

Neimark, and David S. Rumschitzki. Harmonic solutions of a mixed boundary problem arising in the modeling of macromolecular transport into vessel walls. Computers and Mathematics with Applications, 59(6):1897–1908, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122109002429.

Beck:2014:CQO

[BNTT14] Joakim Beck, Fabio Nobile, Lorenzo Tamellini, and Raúl Tempone. Convergence of quasi-optimal stochastic Galerkin methods for a class of PDES with random coefficients. Computers and Mathematics with Applications, 67(4):732–751, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001569.

Benkherouf:2010:OIP

[BO10] Lakdere Benkherouf and Mohamed Omar. Optimal integrated policies for a single-vendor singlebuyer time-varying demand model. Computers and Mathematics with Applications, 60(7):2066–2077, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005390.

Bernardi:2018:PEA

[BO18a] Christine Bernardi and Ajmia Younes Orfi. A posteriori error analysis of the fully discretized time-dependent coupled Darcy and Stokes equations. Computers and Mathematics with Applications, 76(2):340–360, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302190.

Braun:2018:MDM

[BO18b] Moritz Braun and Kingsley O. Obodo. Multidomain muffin tin finite element density functional calculations for small molecules. Computers and Mathematics with Applications, 74(1):35– 44, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306745.

Bochev:2019:P

[Boc19] Pavel Bochev. Preface.

Computers and Mathematics with Applications, 78(2):257, July 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122119302603.

Bogdanovic:2010:IFG

[Bog10] Milena Bogdanović. An ILP formulation and genetic algorithm for the maximum degreebounded connected subgraph problem. Computers and Mathematics with Applications, 59(9):3029– 3038, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001306.

Bognar:2011:ASB

[Bog11] Gabriella Bognár. Analytic solutions to the boundary layer problem over a stretching wall. Computers and Mathematics with Applications, 61(8):2256–2261, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000725X.

Boltuc:2016:EBP

[Bol16] Agnieszka Boltuć. Elastoplastic boundary problems in PIES comparing to BEM and FEM. Computers and Mathematics with Applications, 72(9):2343–2363, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304795.

Bor:2010:ANS

[Bor10] Hüseyin Bor. On absolute Nörlund summability factors. Computers and Mathematics with Applications, 60(7):2031–2034, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000533X.

Bor:2011:NAQ

[Bor11] Hüseyin Bor. A new application of δ-quasi-monotone and almost increasing sequences. Computers and Mathematics with Applications, 61(9): 2899–2902, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111002409.

Botchev:2014:IAT

[BOT14] M. A. Botchev, I. V.

Oseledets, and E. E. Tyrtyshnikov. Iterative across-time solution of linear differential equations: Krylov subspace versus waveform relaxation. Computers and Mathematics with Applications, 67(12):2088– 2098, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001163.

Boyd:2010:SSD

[Boy10]  John P. Boyd.  Six

strategies for defeating the Runge phenomenon in Gaussian radial basis functions on a finite interval. Computers and Mathematics with Applications, 60 (12):3108–3122, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007947.

Balandin:2012:VTU

[BOY12] A. L. Balandin, Y. Ono, and S. You. 3D vector tomography using vector spherical harmonics decomposition. Computers and Mathematics with Applications, 63(10):1433–1441, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002544.

Boyd:2016:FTC

[Boy16]  John P. Boyd.  Five

themes in Chebyshev spectral methods applied to the regularized Charney eigenproblem: Extra numerical boundary conditions, a boundary-layer-resolving change of coordinate, parameterizing a curve which is singular at an endpoint, extending the tau method to log-and-polynomials and finding the roots of a polynomialand-log approximation. Computers and Mathematics with Applications, 71(6):1227–1241, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300141.

Barros:2011:CAH

[BP11a] Saulo R. M. Barros and Pedro S. Peixoto. Computational aspects of harmonic wavelet Galerkin methods and an application to a precipitation front propagation model. Computers and Mathematics with Applications, 61(4):1217–1227, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009892.

Bashier:2011:FNM

[BP11b] Eihab B. M. Bashier and Kailash C. Patidar. A fitted numerical method for a system of partial delay differential equations. Computers and Mathematics with Applications, 61(6):1475–1492, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000859X.

Bhattacharyya:2013:RSS

[BP13] Joydeb Bhattacharyya and Samares Pal. The role of space in stagestructured cannibalism with harvesting of an adult predator. Computers and Mathematics with Applications, 66(3):339–355, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003131.

Boret:2018:IFS

[BP18] Saúl E. Buitrago Boret and Oswaldo J. Jiménez P. Integrated framework for solving the convection diffusion equation on 2D quad mesh relying on internal boundaries. Computers and Mathematics with Applications, 74(1):218–228, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301402.

Bajpai:2019:PEE

[BP19] Saumya Bajpai and Ambit K. Pany. A priori error estimates of fully discrete finite element Galerkin method for Kelvin–Voigt viscoelastic fluid flow model. Computers and Mathematics with Applications, 78 (12):3872–3895, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303219.

Bakry:2017:NAR

[BPC17] Marc Bakry, Sébastien Pernet, and Francis Collino. A new accurate residual-based a posteriori error indicator for the BEM in 2D-acoustics. Computers and Mathematics with Applications, 73(12):2501–2514, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301785.

Bustamante:2013:GMC

[BPF13] C. A. Bustamante, H. Power, and W. F. Florez. A global meshless collocation particular solution method for solving the two-dimensional Navier–Stokes system of equations. Computers and Mathematics with Applications, 65(12):1939–1955, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002204.

Biazar:2010:EGI

[BPG10] J. Biazar, M. Gholami Porshokuhi, and B. Ghanbari. Extracting a general iterative method from an Adomian decomposition method and comparing it to the variational iteration method. Computers and Mathematics with Applications, 59(2):622–628, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007305.

Bruni:2010:RMN

[BPKM10] Carlo Bruni, Francesco Delli Priscoli, Giorgio Koch, and Ilaria Marchetti. Resource management in network dynamics: an optimal approach to the admission control problem. Computers and Mathematics with Applications, 59(1):305–318, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004842.

Beenamol:2012:WBS

[BPM12] M. Beenamol, S. Prabavathy, and J. Mohanalin. Wavelet based seismic signal de-noising using Shannon and Tsallis entropy. Computers and Mathematics with Applications, 64(11):3580–3593, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112005846.

Banas:2014:NIG

[BPM14] Krzysztof Banaś, Przemyslaw Plaszewski, and Pawel Maciol. Numerical integration on GPUs for higher order finite elements. Computers and Mathematics with Applications, 67(6):1319–1344, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000455.

Benes:2018:RDA

[BPR18] Michal Benes, Igor Pazanin, and Marko Radulović. Rigorous derivation of the asymptotic model describing a nonsteady micropolar fluid flow through a thin pipe. Computers and Mathematics with Applications, 76(9):2035–2060, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830419X.

Bermudez:2018:MNA

[BPS18] A. Bermúdez, M. Piñeiro, and P. Salgado. Mathematical and numerical analysis of a transient magnetic model with voltage drop excitations. Computers and Mathematics with Applications, 76(11–12):2710–2727, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304802.

Banz:2019:FST

[BPS19] Lothar Banz, Jan Petsche, and Andreas Schröder. hp-FEM for a stabilized three-field formulation of the biharmonic problem. Computers and Mathematics with Applications, 77(9):2463–2488, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307351.

Brzdek:2011:NNL

[BPX11] Janusz Brzdek, Dorian Popa, and Bing Xu. Note on nonstability of the linear functional equation of higher order. Computers and Mathematics with Applications, 62 (6):2648–2657, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006705.

Boukrouche:2019:UMF

[BPZ19] M. Boukrouche, L. Paoli, and F. Ziane. Unsteady micropolar fluid flow in a thin domain with Tresca fluid-solid interface law. Computers and Mathematics with Applications, 77(11):2917– 2932, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305042.

Bacuta:2015:SPL

[BQ15] Constantin Bacuta and Klajdi Qirko. A saddle point least squares approach to mixed methods. Computers and Mathematics with Applications, 70(12):2920–2932, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004824.

Bacuta:2017:SPL

[BQ17] Constantin Bacuta and Klajdi Qirko. A saddle point least squares approach for primal mixed formulations of second order PDEs. Computers and Mathematics with Applications, 73(2):173–186, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306319.

Bi:2016:OEE

[BQS16] Hui Bi, Chengeng Qian, and Yang Sun. The optimal error estimate and superconvergence of the local discontinuous Galerkin methods for onedimensional linear fifth order time dependent equations. Computers and Mathematics with Applications, 72(3):687– 703, August 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211630298X.

Banas:2012:TVS

[BR12a] Józef Banaś and Beata Rzepka. The technique of Volterra–Stieltjes integral equations in the application to infinite systems of nonlinear integral equations of fractional orders. Computers and Mathematics with Applications, 64(10):3108–3116, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002118.

Barajas-Ramirez:2012:RSC

[BR12b] J. G. Barajas-Ramírez. Robust synchronization of a class of uncertain complex networks via discontinuous control. Computers and Mathematics with Applications, 64(5):956–964, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001150.

Blanco:2012:ESN

[BR12c] V. Blanco and J. C. Rosales. On the enumeration of the set of numerical semigroups with fixed Frobenius number. Computers and Mathematics with Applications, 63(7):1204–1211, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010893.

Braverman:2012:DEA

[BR12d] E. Braverman and A. Rodkina. On difference equations with asymptotically stable 2-cycles perturbed by a decaying noise. Computers and Mathematics with Applications, 64(7):2224–2232, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000788.

Bogner:2013:SFB

[BR13a] Simon Bogner and Ulrich Rüde. Simulation of floating bodies with the lattice Boltzmann method. Computers and Mathematics with Applications, 65(6):901–913, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006451.

Braverman:2013:DER

[BR13b] E. Braverman and A. Rodkina. Difference equations of Ricker and logistic types under bounded stochastic perturbations with positive mean. Computers and Mathematics with Applications, 66 (11):2281–2294, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003891.

Bochev:2016:OBA

[BR16] Pavel Bochev and Denis Ridzal. Optimizationbased additive decomposition of weakly coercive problems with applications. Computers and Mathematics with Applications, 71(11):2140– 2154, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115006008.

Bayoumi:2018:FIH

[BR18] Ahmed M. E. Bayoumi and Mohamed A. Ramadan. Finite iterative Hermitian R-conjugate solutions of the generalized coupled Sylvesterconjugate matrix equations. Computers and Mathematics with Applications, 75(9):3367–3378, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118300762.

Bratsos:2010:FON

[Bra10] A. G. Bratsos. A fourthorder numerical scheme for solving the modified Burgers equation. Computers and Mathematics with Applications, 60 (5):1393–1400, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004359.

Brandejsky:2013:SMG

[Bra13] Tomas Brandejsky. Specific modification of a GPA–ES evolutionary system suitable for deterministic chaos regression. Computers and Mathematics with Applications, 66(2):106–112, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300028X.

Bradji:2016:ACO

[Bra16]  Abdallah Bradji.  An

analysis for the convergence order of gradient schemes for semilinear parabolic equations. Computers and Mathematics with Applications, 72(5):1287–1304, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303637.

Brechtken:2014:GCA

[Bre14] Stefan Brechtken. GPU and CPU acceleration of a class of kinetic lattice group models. Computers and Mathematics with Applications, 67(2):452–461, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004057.

Begmohammadi:2016:NSS

[BRFH16] A. Begmohammadi, M. H. Rahimian, M. Farhadzadeh, and M. Abbasi Hatani. Numerical simulation of single- and multi-mode film boiling using lattice Boltzmann method. Computers and Mathematics with Applications, 71(9):1861–1874, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122116300876.

Bribiesca:2010:CEN

[Bri10] Ernesto Bribiesca. Computation of the Euler number using the contact perimeter. Computers and Mathematics with Applications, 60 (5):1364–1373, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004323.

Boureanu:2016:BPN

[BRR16] Maria-Magdalena Boureanu, Vicenţiu Rădulescu, and Dušan Repovš. On a p(·)-biharmonic problem with no-flux boundary condition. Computers and Mathematics with Applications, 72(9):2505–2515, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305296.

Brevis:2019:STR

[BRROP19] Ignacio Brevis, Ángel Rodríguez-Rozas, Jaime H. Ortega, and David Pardo. Source time reversal (STR) method for linear elasticity. Computers and Mathematics with Applications, 77(5):1358–1375, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306680.

Bermudo:2011:CHC

[BRS11] Sergio Bermudo, José M. Rodríguez, and José M. Sigarreta. Computing the hyperbolicity constant. Computers and Mathematics with Applications, 62(12):4592–4595, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009059.

Babitha:2010:SSR

[BS10a] K. V. Babitha and J. J. Sunil. Soft set relations and functions. Computers and Mathematics with Applications, 60(7):1840–1849, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004906.

Ballestra:2010:EAO

[BS10b] Luca Vincenzo Ballestra and Carlo Sgarra. The evaluation of American options in a stochastic volatility model with jumps: an efficient finite element approach. Computers and Mathematics with Applications, 60(6):1571–1590, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004542.

Babitha:2011:TCO

[BS11a] K. V. Babitha and Jacob John Sunil. Transitive closures and orderings on soft sets. Computers and Mathematics with Applications, 62 (5):2235–2239, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005645.

Beik:2011:GKS

[BS11b] Fatemeh Panjeh Ali Beik and Davod Khojasteh Salkuyeh. On the global Krylov subspace methods for solving general coupled matrix equations. Computers and Mathematics with Applications, 62(12):4605–4613, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009072.

Bai:2012:EMP

[BS12a] Zhanbing Bai and Weichen Sun. Existence and multiplicity of positive solutions for singular fractional boundary value problems. Computers and Mathematics with Applications, 63(9):1369– 1381, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011497.

Berzig:2012:ECF

[BS12b] Maher Berzig and Bessem Samet. An extension of coupled fixed point’s concept in higher dimension and applications. Computers and Mathematics with Applications, 63(8):1319–1334, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000284.

Banz:2014:AIT

[BS14a] Lothar Banz and Ernst P.

Stephan. hp-adaptive IPDG/TDG–FEM for parabolic obstacle problems. Computers and Mathematics with Applications, 67(4):712–731, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001545.

Broersen:2014:RPG

[BS14b] Dirk Broersen and Rob Stevenson. A robust Petrov–Galerkin discretisation of convection– diffusion equations. Computers and Mathematics with Applications, 68 (11):1605–1618, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002806.

Banz:2015:BBF

[BS15a] Lothar Banz and Andreas Schröder. Biorthogonal basis functions in hp-adaptive FEM for elliptic obstacle problems. Computers and Mathematics with Applications, 70(8):1721–1742, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003454.

Banz:2015:ABF

[BS15b] Lothar Banz and Ernst P. Stephan. On hp-adaptive BEM for frictional contact problems in linear elasticity. Computers and Mathematics with Applications, 69(7):559–581, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000528.

Burg:2015:PEC

[BS15c] Markus Bürg and Andreas Schröder. A posteriori error control of hpfinite elements for variational inequalities of the first and second kind. Computers and Mathematics with Applications, 70(12):2783–2802, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004101.

Bao:2016:BBT

[BS16a] Aiguo Bao and Xianfa Song. Bounds for the blowup time of the solution to a parabolic system with nonlocal factors in nonlinearities. Computers and Mathematics with Applications, 71(3):723–729, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005970.

Bermejo:2016:SOT

[BS16b] R. Bermejo and L. Saavedra. A second order in time local projection stabilized Lagrange– Galerkin method for Navier–Stokes equations at high Reynolds numbers. Computers and Mathematics with Applications, 72(4):820–845, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302802.

Byfut:2017:UML

[BS17] Andreas Byfut and Andreas Schröder. Unsymmetric multi-level hanging nodes and anisotropic polynomial degrees in H^{1}-conforming higherorder finite element methods. Computers and Mathematics with Applications, 73(9):2092–2150, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301086.

Boffi:2018:RFE

[BS18a] Daniele Boffi and Rolf Stenberg. A remark on finite element schemes for nearly incompressible elasticity. Computers and Mathematics with Applications, 74 (9):2047–2055, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303541.

Brechtken:2018:NHO

[BS18b] Stefan Brechtken and Thomas Sasse. Normal, high order discrete velocity models of the Boltzmann equation. Computers and Mathematics with Applications, 75(2):503–519, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122117305874.

Bhuvaneswari:2011:EAR

[BSK11] M. Bhuvaneswari, S. Sivasankaran, and Y. J. Kim. Effect of aspect ratio on convection in a porous enclosure with partially active thermal walls. Computers and Mathematics with Applications, 62 (10):3844–3856, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008017.

Bettahar:2011:NSE

[BSL11] S. Bettahar, A. Boudghene Stambouli, and P. Lambert. Numerical scheme for efficient colour image denoising. Computers and Mathematics with Applications, 61(9):2903– 2913, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002422.

Borges:2013:EGM

[BSN13] Helyane Bronoski Borges, Carlos N. Silla, Jr., and Júlio Cesar Nievola. An evaluation of globalmodel hierarchical classification algorithms for hierarchical classification problems with single path of labels. Computers and Mathematics with Applications, 66 (10):1991–2002, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300432X.

Bira:2018:CCS

[BSS18] B. Bira, T. Raja Sekhar, and G. P. Raja Sekhar. Collision of characteristic shock with weak discontinuity in nonideal magnetogasdynamics. Computers and Mathematics with Applications, 75(11):3873–3883, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301159.

Bu:2019:STF

[BSY^{+}19] Weiping Bu, Shi Shu, Xiaoqiang Yue, Aiguo Xiao, and Wei Zeng. Space-time finite element method for the multi-term time-space fractional diffusion equation on a two-dimensional domain. Computers and Mathematics with Applications, 78(5):1367–1379, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306862.

Bira:2016:ALG

[BSZ16] B. Bira, T. Raja Sekhar, and D. Zeidan. Application of Lie groups to compressible model of two-phase flows. Computers and Mathematics with Applications, 71(1): 46–56, January 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115005118.

Baksalary:2010:FOP

[BT10] Oskar Maria Baksalary and Götz Trenkler. Functions of orthogonal projectors involving the Moore–Penrose inverse. Computers and Mathematics with Applications, 59(2):764–778, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007160.

Baksalary:2011:MD

[BT11] Oskar Maria Baksalary and Götz Trenkler. On the matrix difference I − A. Computers and Mathematics with Applications, 62(5):2279–2288, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005694.

Boubendir:2014:WCB

[BT14] Yassine Boubendir and Catalin Turc. Wellconditioned boundary integral equation formulations for the solution of high-frequency electromagnetic scattering problems. Computers and Mathematics with Applications, 67(10):1772– 1805, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001564.

Benes:2015:CNS

[BT15] Michal Benes and Jakub Tichý. On coupled Navier–Stokes and energy equations in exteriorlike domains. Computers and Mathematics with Applications, 70 (12):2867–2882, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004666.

Beniani:2018:WPE

[BTB18] Abderrahmane Beniani, Noureddine Taouaf, and Abbes Benaissa. Wellposedness and exponential stability for coupled Lamé system with viscoelastic term and strong damping. Computers and Mathematics with Applications, 75(12):4397–4404, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301731.

Barnoon:2019:EGA

[BTEM19] Pouya Barnoon, Davood Toghraie, Farshad Eslami, and Babak Mehmandoust. Entropy generation analysis of different nanofluid flows in the space between two concentric horizontal pipes in the presence of magnetic field: Single-phase and two-phase approaches. Computers and Mathematics with Applications, 77(3):662–692, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305923.

Buong:2011:HII

[Buo11]  Nguyen Buong.  Hy-

brid Ishikawa iterative methods for a nonexpansive semigroup in Hilbert space. Computers and Mathematics with Applications, 61(9):2546–2554, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001465.

Burguillo:2013:PCC

[Bur13] Juan C. Burguillo. Playing with complexity: From cellular evolutionary algorithms with coalitions to self-organizing maps. Computers and Mathematics with Applications, 66(2):201–212, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000370.

Buyukyazici:2010:ASC

[Büy10]  Ibrahim  Büyükyazici.

Approximation by Stancu– Chlodowsky polynomials. Computers and Mathematics with Applications, 59(1):274–282, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005100.

Boek:2010:LBS

[BV10] Edo S. Boek and Maddalena Venturoli. Lattice-Boltzmann studies of fluid flow in porous media with realistic rock geometries. Computers and Mathematics with Applications, 59(7):2305–2314, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006427.

Balasubramaniam:2011:RSU

[BV11] P. Balasubramaniam and V. Vembarasan. Robust stability of uncertain fuzzy BAM neural networks of neutral-type with Markovian jumping parameters and impulses. Computers and Mathematics with Applications, 62(4):1838–1861, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005062.

Bozorgnia:2017:FAT

[BV17] Farid Bozorgnia and Jan Valdman. A FEM approximation of a twophase obstacle problem and its a posteriori error estimate. Computers and Mathematics with Applications, 73(3):419–432, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306691.

Bi:2018:TGF

[BWL18] Chunjia Bi, Cheng Wang, and Yanping Lin. Twogrid finite element method and its a posteriori error estimates for a nonmonotone quasilinear elliptic problem under minimal regularity of data. Computers and Mathematics with Applications, 76(1):98–112, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118302049.

Borggaard:2016:GOR

[BWZ16] Jeff Borggaard, Zhu Wang, and Lizette Zietsman. A goal-oriented reduced-order modeling approach for nonlinear systems. Computers and Mathematics with Applications, 71(11):2155– 2169, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630027X.

Bi:2010:PTI

[BX10] Hui Bi and Yuming Xing.

Poincaré-type inequalities with L^{p}(log L)^{α}-norms for Green’s operator. Computers and Mathematics with Applications, 60(10):2764–2770, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007157.

Bnouhachem:2014:ILA

[BX14] Abdellah Bnouhachem and M. H. Xu. An inexact LQP alternating direction method for solving a class of structured variational inequalities. Computers and Mathematics with Applications, 67(3):671–680, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006913.

Bnouhachem:2011:NAD

[BXKZ11] Abdellah Bnouhachem, M. H. Xu, Mohamed Khalfaoui, and Sheng Zhaohan. A new alternating direction method for solving variational inequalities. Computers and Mathematics with Applications, 62(2):626– 634, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100441X.

Bo:2011:NSD

[BY11] Lijun Bo and Chenggui Yuan. A note on stability in distribution of Markovmodulated stochastic differential equations with reflection. Computers and Mathematics with Applications, 61(10):3010– 3016, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122111002677.

Bai:2010:ESF

[BZ10] Zhanbing Bai and Yinghan Zhang. The existence of solutions for a fractional multi-point boundary value problem. Computers and Mathematics with Applications, 60(8):2364–2372, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005948.

Bhrawy:2018:HAN

[BZ18] A. H. Bhrawy and M. A.

Zaky. Highly accurate numerical schemes for multi-dimensional space variable-order fractional Schrödinger equations. Computers and Mathematics with Applications, 73(6):1100–1117, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306368.

Balachandran:2012:RCFb

[BZK12] K. Balachandran, Yong Zhou, and J. Kokila. Relative controllability of fractional dynamical systems with distributed delays in control. Computers and Mathematics with Applications, 64 (10):3201–3209, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010960.

Bog:2015:NCH

[BZKR15] Tino Bog, Nils Zander, Stefan Kollmannsberger, and Ernst Rank. Normal contact with high order finite elements and a fictitious contact material. Computers and Mathematics with Applications, 70(7):1370–1390, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002011.

Zhang:2017:MLK

[bZM17] Jian bing Zhang and Wen-Xiu Ma. Mixed lump-kink solutions to the BKP equation. Computers and Mathematics with Applications, 74(3):591–596, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302997.

Brdar:2016:SPP

[BZT16] Mirjana Brdar, Helena Zarin, and Ljiljana Teofanov. A singularly perturbed problem with two parameters in two dimensions on graded meshes. Computers and Mathematics with Applications, 72(10):2582–2603, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305338.

Benmessaoud:2010:NAM

[BZZ^{+}10] M. T. Benmessaoud, F. Z.

Zerhouni, M. Zegrar, A. Boudghene Stambouli, and M. Tioursi. New approach modeling and a maximum power point tracker method for solar cells. Computers and Mathematics with Applications, 60(4):1124–1134, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002610.

Cetkin:2010:LVD

[ÇA10a] Vildan Çetkin and Halis Aygün. Lattice valued double fuzzy preproximity spaces. Computers and Mathematics with Applications, 60(3):849– 864, August 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110003895.

Cevikel:2010:SFM

[CA10b] Adem C. Cevikel and Mehmet Ahlatçioğlu. Solutions for fuzzy matrix games. Computers and Mathematics with Applications, 60(3):399–410, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002919.

Cibikdiken:2014:CMM

[ÇA14] Ali Osman Çibikdiken and Kemal Aydin. Computation of the monodromy matrix in floating point arithmetic with the Wilkinson model. Computers and Mathematics with Applications, 67(5):1186–1194, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122113006706.

Cumsille:2014:NMB

[CAC14] Patricio Cumsille, Juan A.

Asenjo, and Carlos Conca. A novel model for biofilm growth and its resolution by using the hybrid immersed interface-level set method. Computers and Mathematics with Applications, 67(1):34–51, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006202.

Cho:2011:ESS

[CAH11] Yeol Je Cho, Ioannis K. Argyros, and Saïd Hilout. Extended sufficient semilocal convergence for the secant method. Computers and Mathematics with Applications, 62(2):599–610, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100438X.

Cakalli:2011:NKC

[Çak11a] Hüseyin Çakalli. New kinds of continuities. Computers and Mathematics with Applications, 61(4):960–965, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009491.

Cakalli:2011:QSO

[Çak11b]  Hüseyin Çakallı.  On

Δ-quasi-slowly oscillating sequences. Computers and Mathematics with Applications, 62 (9):3567–3574, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007590.

Cakalli:2011:C

[Çak11c]  Hüseyin Çakallı.  On

G-continuity. Computers and Mathematics with Applications, 61(2):313– 318, January 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008552.

Calvo:2019:OSM

[Cal19] Juan G. Calvo. An overlapping Schwarz method for virtual element discretizations in two dimensions. Computers and Mathematics with Applications, 77 (4):1163–1177, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306400.

Canak:2011:MS

[Çan11a]  İbrahim Çanak.  On

(C, 1) means of sequences. Computers and Mathematics with Applications, 62(9):3446–3448, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007449.

Canak:2011:TCG

[Çan11b] Ibrahim Çanak. A theorem for convergence of generator sequences. Computers and Mathematics with Applications, 61(2):408–411, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008667.

Canak:2011:TCS

[Çan11c]  Ibrahim Çanak.  A

theorem on the Cesàro summability method. Computers and Mathematics with Applications, 61(4):1162–1166, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009703.

Candan:2011:OSO

[Can11d]  T. Candan.  Oscilla-

tion of second-order nonlinear neutral dynamic equations on time scales with distributed deviating arguments. Computers and Mathematics with Applications, 62 (11):4118–4125, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008431.

Chaudhry:2019:SMR

[CANA19] Qasim Ali Chaudhry, Amna Abbas, Ayesha Noor, and Muqaddas Asif. In silico modeling for the risk assessment of toxicity in cells. Computers and Mathematics with Applications, 77(6):1541–1548, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305108.

Chang:2011:GSF

[CANK11] Yong-Kui Chang, M. Mallika Arjunan, G. M. N’Guérékata, and V. Kavitha. On global solutions to fractional functional differential equations with infinite delay in Fréchet spaces. Computers and Mathematics with Applications, 62(3):1228–1237, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100191X.

Cao:2019:GCS

[Cao19]  Yang Cao.  A gen-

eral class of shift-splitting preconditioners for non-Hermitian saddle point problems with applications to time-harmonic eddy current models. Computers and Mathematics with Applications, 77(4):1124–1143, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306436.

Cho:2010:AMC

[CAP10] Yeol Je Cho, Ioannis K. Argyros, and Narin Petrot. Approximation methods for common solutions of generalized equilibrium, systems of nonlinear variational inequalities and fixed point problems. Computers and Mathematics with Applications, 60(8):2292–2301, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005857.

Ceng:2012:EMS

[CAY12] L.-C. Ceng, Q. H. Ansari, and J.-C. Yao. An extragradient method for solving split feasibility and fixed point problems. Computers and Mathematics with Applications, 64(4):633–642, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211101145X.

Cardenas-Barron:2010:EMD

[CB10] Leopoldo Eduardo Cárdenas-Barrón. An easy method to derive EOQ and EPQ inventory models with backorders. Computers and Mathematics with Applications, 59(2):948– 952, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109006774.

Cai:2011:HAG

[CB11a] Gang Cai and Shangquan Bu. Hybrid algorithm for generalized mixed equilibrium problems and variational inequality problems and fixed point problems. Computers and Mathematics with Applications, 62 (12):4772–4782, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009448.

Cai:2011:SCT

[CB11b] Gang Cai and Shangquan Bu. Strong convergence theorems based on a new modified extragradient method for variational inequality problems and fixed point problems in Banach spaces. Computers and Mathematics with Applications, 62(6):2567–2579, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006316.

Cai:2011:VAS

[CB11c] Gang Cai and Shangquan Bu. A viscosity approximation scheme for finite mixed equilibrium problems and variational inequality problems and fixed point problems. Computers and Mathematics with Applications, 62(1):440–454, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004238.

Charef:2011:ASL

[CB11d] Abdelfatah Charef and Djamel Boucherma. Analytical solution of the linear fractional system of commensurate order. Computers and Mathematics with Applications, 62(12):4415–4428, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008819.

Chandra:2019:MFA

[CB19a] Saroj Kumar Chandra and Manish Kumar Bajpai. Mesh free alternate directional implicit method based three dimensional super-diffusive model for benign brain tumor segmentation. Computers and Mathematics with Applications, 77(12):3212–3223, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930077X.

Cioncolini:2019:MMF

[CB19b] Andrea Cioncolini and Daniele Boffi. The MINI mixed finite element for the Stokes problem: an experimental investigation. Computers and Mathematics with Applications, 77(9):2432–2446, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307260.

Choudhury:2015:WPR

[CBB15] M. Choudhury, U. Basu, and R. K. Bhattacharyya. Wave propagation in a rotating randomly varying granular generalized thermoelastic medium. Computers and Mathematics with Applications, 70(12):2803–2821, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004459.

Cantin:2016:VBS

[CBBE16] Pierre Cantin, Jérôme Bonelle, Erik Burman, and Alexandre Ern. A vertex-based scheme on polyhedral meshes for advection-reaction equations with sub-mesh stabilization. Computers and Mathematics with Applications, 72(9):2057–2071, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304436.

Charifi:2010:SPG

[CBKR10] Ahmed Charifi, Belaid Bouikhalene, Samir Kabbaj, and John Michael Rassias. On the stability of a Pexiderized Gołab–Schinzel equation. Computers and Mathematics with Applications, 59(9):3193–3202, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001653.

Chen:2010:PSI

[CBM10] Wen Chen, Dumitru Baleanu, and J. A. Tenreiro Machado. Preface: Special issue of computers and mathematics with applications on fractional differentiation and its applications. Computers and Mathematics with Applications, 59(5):1585, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005331.

Chaparpordi:2018:BTP

[CBS18] Seyyed Hassan Azizi Chaparpordi, Fatemeh Panjeh Ali Beik, and Davod Khojasteh Salkuyeh. Block triangular preconditioners for stabilized saddle point problems with nonsymmetric (1, 1)-block. Computers and Mathematics with Applications, 76 (6):1544–1553, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118303717.

Chen:2011:NAB

[CC11] Guanghua Chen and Guangming Chen. A numerical algorithm based on a variational iterative approximation for the discrete Hamilton– Jacobi–Bellman (HJB) equation. Computers and Mathematics with Applications, 61(4):901–907, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009430.

Chaillat:2015:WFM

[CC15] Stéphanie Chaillat and Francis Collino. A wideband Fast Multipole Method for the Helmholtz kernel: Theoretical developments. Computers and Mathematics with Applications, 70(4):660–678, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002424.

Chung:2017:NCB

[CC17] Soon-Yeong Chung and Min-Jun Choi. A new condition for blow-up solutions to discrete semilinear heat equations on networks. Computers and Mathematics with Applications, 74 (12):2929–2939, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304595.

Che:2019:EAB

[CC19] Guofeng Che and Haibo Chen. Existence and asymptotic behavior of positive ground state solutions for coupled nonlinear fractional Kirchhofftype systems. Computers and Mathematics with Applications, 77(1):173–188, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305327.

Chalco-Cano:2011:ACF

[CCBSRFRM11] Y. Chalco-Cano, A. D.

Báez-Sánchez, H. Román-Flores, and M. A. Rojas-Medar. On the approximation of compact fuzzy sets. Computers and Mathematics with Applications, 61(2):412–420, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008679.

Chiu:2010:OFP

[CCCW10] Yuan-Shyi Peter Chiu, Kuang-Ku Chen, Feng-Tsung Cheng, and Mei-Fang Wu. Optimization of the finite production rate model with scrap, rework and stochastic machine breakdown. Computers and Mathematics with Applications, 59(2):919–932, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006907.

Chen:2016:SBM

[CCCW16] Bin Chen, Wen Chen, Alexander H. D. Cheng, and Xing Wei. The singular boundary method for two-dimensional static thermoelasticity analysis. Computers and Mathematics with Applications, 72(11):2716–2730, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122116305399.

Chen:2010:AAM

[CCD10] Anping Chen, Fulai Chen, and Siqing Deng. On almost automorphic mild solutions for fractional semilinear initial value problems. Computers and Mathematics with Applications, 59(3):1318–1325, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004192.

Cai:2010:MMS

[CCDL10] Kai-Yuan Cai, Ping Cao, Zhao Dong, and Ke Liu. Mathematical modeling of software reliability testing with imperfect debugging. Computers and Mathematics with Applications, 59(10):3245– 3285, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001719.

Chumley:2013:BT

[CCF13] T. Chumley, S. Cook, and R. Feres. From billiards to thermodynamics. Computers and Mathematics with Applications, 65(10):1596– 1613, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005706.

Chiandussi:2012:CMO

[CCFV12] G. Chiandussi, M. Codegone, S. Ferrero, and F. E. Varesio. Comparison of multi-objective optimization methodologies for engineering applications. Computers and Mathematics with Applications, 63(5):912–942, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010406.

Chao:2018:SCR

[CCG18] Zhen Chao, Guoliang Chen, and Ye Guo. On the semi-convergence of regularized HSS iteration methods for singular saddle point problems. Computers and Mathematics with Applications, 76(2):438–450, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118302359.

Chen:2012:ACE

[CCH^{+}12] Jong-Shin Chen, Hsing-Chung Chen, Yung-Fa Huang, Ching-Chuan Wei, and Kai-Cheng Chuang. An adaptive capacity enhancement strategy for sectorbased cellular systems. Computers and Mathematics with Applications, 64(5):1462–1472, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002982.

Cebrian:2017:CMS

[CCHG17] Juan M. Cebrián, José M.

Cecilia, Mario Hernández, and José M. García. Code modernization strategies to 3-D stencilbased applications on Intel Xeon Phi: KNC and KNL. Computers and Mathematics with Applications, 74 (10):2557–2571, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304613.

Cheng:2014:AST

[CChL14] Wanyou Cheng, Zixin Chen, and Dong hui Li. An active set truncated Newton method for large-scale bound constrained optimization. Computers and Mathematics with Applications, 67(5):1016–1023, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000212.

Calbo:2010:MSP

[CCJ10] G. Calbo, J.-C. Cortés, and L. Jódar. Mean square power series solution of random linear differential equations. Computers and Mathematics with Applications, 59(1):559–572, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003770.

Casaban:2011:NAC

[CCJP11] M.-C. Casabán, R. Company, L. Jódar, and J.- R. Pintos. Numerical analysis and computing of a non-arbitrage liquidity model with observable parameters for derivatives. Computers and Mathematics with Applications, 61(8):1951–1956, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005729.

Calbo:2011:SRL

[CCJV11] G. Calbo, J.-C. Cortés, L. Jódar, and L. Villafuerte. Solving the random Legendre differential equation: Mean square power series solution and its statistical functions. Computers and Mathematics with Applications, 61(9):2782– 2792, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002082.

Cheng:2012:EHS

[CCK12] Yu-Hsin Cheng, Fu-Min Chang, and Shang-Juh Kao. Efficient hierarchical SIP mobility management for WiMAX networks. Computers and Mathematics with Applications, 64(5):1522–1531, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003434.

Choi:2018:FEM

[CCK18] Hyung Jun Choi, Woocheol Choi, and Youngwoo Koh. A finite element method for elliptic optimal control problem on a non-convex polygon with corner singularities. Computers and Mathematics with Applications, 75(1):45–58, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305217.

Cai:2015:AEP

[CCKP15] Zhiqiang Cai, Varis Carey, JaEun Ku, and Eun-Jae Park. Asymptotically exact a posteriori error estimators for first-order div least-squares methods in local and global L_{2} norm. Computers and Mathematics with Applications, 70(4):648–659, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002333.

Cheng:2012:NLF

[CCKY12] Chun-An Cheng, Hung-Liang Cheng, Chen-Wei Ku, and Fu-Li Yang. A novel low-frequency square-wave-driven electronic ballast for automotive HID lamps. Computers and Mathematics with Applications, 64 (5):1409–1419, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002933.

Chen:2012:ARF

[CCL^{+}12] Wen-Ping Chen, Song-Shyong Chen, Chun-Cheng Lin, Ya-Zhung Chen, and Wen-Chih Lin. Automatic recognition of frog calls using a multi-stage average spectrum. Computers and Mathematics with Applications, 64(5):1270–1281, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002763.

Ciric:2010:YAP

[ĆCM10] Vladimir Ćirić, Aleksandar Cvetković, and Ivan Milentijević. Yield analysis of partial defect tolerant bit-plane array. Computers and Mathematics with Applications, 59(1):98–107, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006555.

Costa:2014:NCV

[CCM14] Ricardo Costa, Stéphane Clain, and Gaspar J. Machado. New cell-vertex reconstruction for finite volume scheme: Application to the convection– diffusion-reaction equation. Computers and Mathematics with Applications, 68(10):1229– 1249, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114004076.

Calo:2014:ADP

[CCN14] Victor M. Calo, Nathaniel O. Collier, and Antti H. Niemi. Analysis of the discontinuous Petrov– Galerkin method with optimal test functions for the Reissner–Mindlin plate bending model. Computers and Mathematics with Applications, 66(12):2570–2586, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004409.

Castrillon-Candas:2016:ARC

[CCNT16] Julio E. Castrillón-Candás, Fabio Nobile, and Raúl F. Tempone. Analytic regularity and collocation approximation for elliptic PDEs with random domain deformations. Computers and Mathematics with Applications, 71(6):1173–1197, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000080.

Chung:2019:CSF

[CCP19] Soon-Yeong Chung, Min-Jun Choi, and Jea-Hyun Park. On the critical set for Fujita type blow-up of solutions to the discrete Laplacian parabolic equations with nonlinear source on networks. Computers and Mathematics with Applications, 78(6):1838–1850, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300847.

Cao:2016:WPN

[CCR16] Yanzhao Cao, Song Chen, and Leo G. Rebholz. Well-posedness and a numerical study of a regularization model with adaptive nonlinear filtering for incompressible fluid flow. Computers and Mathematics with Applications, 71(11):2192– 2205, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005805.

Cavalcanti:2017:GDR

[CCRS17] Marcelo M. Cavalcanti, Wellington J. Corrêa, Carole Rosier, and Flávio R. Dias Silva. General decay rate estimates and numerical analysis for a transmission problem with locally distributed nonlinear damping. Computers and Mathematics with Applications, 73(10):2293–2318, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122117301748.

Chen:2014:LBM

[CCSZ14] Junhui Chen, Zhenhua Chai, Baochang Shi, and Wenhuan Zhang. Lattice Boltzmann method for filtering and contour detection of the natural images. Computers and Mathematics with Applications, 68(3):257–268, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002491.

Chen:2013:ENS

[CCX13] Caisheng Chen, Lin Chen, and Zonghu Xiu. Existence of nontrivial solutions for singular quasilinear elliptic equations on R^{N}. Computers and Mathematics with Applications, 65(12):1909–1919, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300223X.

Chang:2010:IDA

[CCY10] Dong-Shang Chang, Fu-Chiao Chyr, and Fu-Chiang Yang. Incorporating a database approach into the largescale multi-level lot sizing problem. Computers and Mathematics with Applications, 60(9):2536–2547, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006012.

Chen:2018:ALB

[CCY18] Ying-Xuan Chen, Shing-Cheng Chang, and Wen-Bin Young. Application of lattice Boltzmann method in free surface flow simulation of micro injection molding. Computers and Mathematics with Applications, 75(7):2374–2386, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307770.

Chen:2018:PEE

[CCZ18] Chuanjun Chen, Yanping Chen, and Xin Zhao. A posteriori error estimates of two-grid finite volume element methods for nonlinear elliptic problems. Computers and Mathematics with Applications, 75(5):1756–1766, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307617.

Chen:2010:LCB

[CD10] Xingwu Chen and Zhengdong Du. Limit cycles bifurcate from centers of discontinuous quadratic systems. Computers and Mathematics with Applications, 59(12):3836– 3848, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002907.

Chen:2012:MPR

[CD12] Zhaoyu Chen and Stefan Diebels. Modelling and parameter reidentification of nanoindentation of soft polymers taking into account effects of surface roughness. Computers and Mathematics with Applications, 64(9):2775–2786, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003501.

Cavoretto:2014:MIA

[CD14] Roberto Cavoretto and Alessandra De Rossi. A meshless interpolation algorithm using a cellbased searching procedure. Computers and Mathematics with Applications, 67(5):1024–1038, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000194.

Chupin:2016:BPM

[CD16] Laurent Chupin and Thierry Dubois. A biprojection method for Bingham type flows. Computers and Mathematics with Applications, 72(5):1263–1286, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303583.

Castelli:2019:NSM

[CD19] G. F. Castelli and W. Dörfler. The numerical study of a microscale model for lithium-ion batteries. Computers and Mathematics with Applications, 77(6):1527–1540, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304991.

Costabile:2012:EAO

[CDD12] F. A. Costabile, F. Dell’Accio, and F. Di Tommaso. Enhancing the approximation order of local Shepard operators by Hermite polynomials. Computers and Mathematics with Applications, 64 (11):3641–3655, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006050.

Castiglione:2012:ACC

[CDFP12] Aniello Castiglione, Alfredo De Santis, Ugo Fiore, and Francesco Palmieri. An asynchronous covert channel using spam. Computers and Mathematics with Applications, 63(2):437– 447, January 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111006432.

Calle:2015:ICA

[CDG15] Jorge L. Díaz Calle, Philippe R. B. Devloo, and Sônia M. Gomes. Implementation of continuous hp-adaptive finite element spaces without limitations on hanging sides and distribution of approximation orders. Computers and Mathematics with Applications, 70(5):1051–1069, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003193.

Carstensen:2016:BSF

[CDG16] C. Carstensen, L. Demkowicz, and J. Gopalakrishnan. Breaking spaces and forms for the DPG method and applications including Maxwell equations. Computers and Mathematics with Applications, 72(3):494–522, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302620.

Chung:2017:DGM

[CDL17] Eric T. Chung, Jie Du, and Chi Yeung Lam. Discontinuous Galerkin methods with staggered hybridization for linear elastodynamics. Computers and Mathematics with Applications, 74 (6):1198–1214, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303516.

Coppa:2010:LTT

[CDM10] G. Coppa, A. D’Angola, and R. Mulas. Linear transport theory for particles moving on a spherical surface. Computers and Mathematics with Applications, 59(1):486– 492, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109003897.

Cecchi:2012:ODE

[CDM12] Mariella Cecchi, Zuzana Doslá, and Mauro Marini. On oscillation of difference equations with bounded phi-Laplacian. Computers and Mathematics with Applications, 64(7):2176–2184, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010790.

Costabel:2014:WAR

[CDN14] Martin Costabel, Monique Dauge, and Serge Nicaise. Weighted analytic regularity in polyhedra. Computers and Mathematics with Applications, 67 (4):807–817, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001582.

Creuse:2019:AGC

[CDN19] Emmanuel Creusé, Patrick Dular, and Serge Nicaise. About the gauge conditions arising in finite element magnetostatic problems. Computers and Mathematics with Applications, 77(6):1563–1582, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303547.

Cavoretto:2016:ECP

[CDP16] R. Cavoretto, A. De Rossi, and E. Perracchione. Efficient computation of partition of unity interpolants through a blockbased searching technique. Computers and Mathematics with Applications, 71(12):2568– 2584, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302024.

Coclite:2018:CMR

[CdR18] Giuseppe Maria Coclite and Lorenzo di Ruvo. On the convergence of the modified Rosenau and the modified Benjamin– Bona–Mahony equations. Computers and Mathematics with Applications, 74(5):899–919, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300700.

Costabel:2015:VIE

[CDS15] Martin Costabel, Eric Darrigrand, and Hamdi Sakly. Volume integral equations for electromagnetic scattering in two dimensions. Computers and Mathematics with Applications, 70(8):2087–2101, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004058.

Chen:2011:DWA

[CDW11] Jun-Lang Chen, Chao-Qing Dai, and Xiao-Gang Wang. Domain wall arrays, fronts, and bright and dark solitons in a generalized derivative nonlinear Schrödinger equation. Computers and Mathematics with Applications, 62(2):620–625, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004408.

Chen:2011:MAA

[CDY11] Zhongying Chen, Shengpei Ding, and Hongqi Yang. Multilevel augmentation algorithms based on fast collocation methods for solving ill-posed integral equations. Computers and Mathematics with Applications, 62(4):2071–2082, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005359.

Cagman:2010:CSM

[ÇE10a] Naim Çagman and Serdar Enginoğlu. Corrigendum to “Soft matrix theory and its decision making” [Comput. Math. Appl. 59 (2010) 3308–3314]. Computers and Mathematics with Applications, 60(9):2723, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004232.

See [ÇE10b].

Cagman:2010:SMT

[ÇE10b] Naim Çagman and Serdar Enginoğlu. Soft matrix theory and its decision making. Computers and Mathematics with Applications, 59(10):3308– 3314, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001914.

See corrigendum [ÇE10a].

Cavazza:2013:CPH

[CEF^{+}13] N. Cavazza, M. Ethier, P. Frosini, T. Kaczynski, and C. Landi. Comparison of persistent homologies for vector functions: From continuous to discrete and back.

Computers and Mathematics with Applications, 66(4):560–573, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003696.

Company:2016:CAO

[CEJV16] Rafael Company, Vera Egorova, Lucas Jódar, and Carlos Vázquez. Computing American option price under regime switching with rationality parameter. Computers and Mathematics with Applications, 72(3):741– 754, August 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116302942.

Chung:2018:GMF

[CELY18] Eric T. Chung, Yalchin Efendiev, Wing Tat Leung, and Shuai Ye. Generalized multiscale finite element methods for space–time heterogeneous parabolic equations. Computers and Mathematics with Applications, 76(2):419–437, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302347.

Chan:2014:DPG

[CEQ14] Jesse Chan, John A.

Evans, and Weifeng Qiu. A dual Petrov–Galerkin finite element method for the convection–diffusion equation. Computers and Mathematics with Applications, 68 (11):1513–1529, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003149.

Cerna:2018:PGM

[Cer18] Dana Cerná. Postprocessing Galerkin method using quadratic spline wavelets and its efficiency. Computers and Mathematics with Applications, 75(9):3186–3200, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300567.

Chaumont-Frelet:2016:HOM

[CF16a] Théophile Chaumont-Frelet. On high order methods for the heterogeneous Helmholtz equation. Computers and Mathematics with Applications, 72(9):2203–2225, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304801.

Chen:2016:IMS

[CF16b] Caisheng Chen and Shuyan Fu. Infinitely many solutions to quasilinear Schrödinger system in R^{N}. Computers and Mathematics with Applications, 71(7):1417–1424, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300517.

Coscia:2011:MTL

[CFB11] V. Coscia, L. Fermo, and N. Bellomo. On the mathematical theory of living systems II: the interplay between mathematics and system biology. Computers and Mathematics with Applications, 62(10):3902–3911, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211100811X.

Chiarelli:2018:CHO

[CFdM^{+}18] L. R. Chiarelli, F. G.

Fumes, E. A. Barros de Moraes, G. A. Haveroth, J. L. Boldrini, and M. L. Bittencourt. Comparison of high order finite element and discontinuous Galerkin methods for phase field equations: Application to structural damage. Computers and Mathematics with Applications, 74(7):1542–1564, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302821.

Csomos:2015:NSN

[CFF15] Petra Csomós, István Faragó, and Imre Fekete. Numerical stability for nonlinear evolution equations. Computers and Mathematics with Applications, 70(11):2752– 2761, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115002680.

Coclite:2019:ADM

[CFLM19] G. M. Coclite, G. Florio, M. Ligabò, and F. Maddalena. Adhesion and debonding in a model of elastic string. Computers and Mathematics with Applications, 78 (6):1897–1909, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301452.

Constantinou:2018:FEA

[CFLX18] P. Constantinou, S. Franz, L. Ludwig, and C. Xenophontos. Finite element approximation of reaction– diffusion problems using an exponentially graded mesh. Computers and Mathematics with Applications, 76 (10):2523–2534, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304772.

Calzolari:2011:NFS

[CFN11] Antonella Calzolari, Patrick Florchinger, and Giovanna Nappo. Nonlinear filtering for stochastic systems with fixed delay: Approximation by a modified Milstein scheme. Computers and Mathematics with Applications, 61(9):2498–2509, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001222.

Fu:2013:AGA

[cFpClC13] Tak chung Fu, Chi pang Chung, and Fu lai Chung. Adopting genetic algorithms for technical analysis and portfolio management. Computers and Mathematics with Applications, 66 (10):1743–1757, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004938.

Carstensen:2014:AA

[CFPP14] C. Carstensen, M. Feischl, M. Page, and D. Praetorius. Axioms of adaptivity. Computers and Mathematics with Applications, 67(6):1195–1253, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006822.

Castillo:2010:DCF

[CFRS10] Fernando J. Castillo, V. Feliu, R. Rivas, and L. Sánchez. Design of a class of fractional controllers from frequency specifications with guaranteed time domain behavior. Computers and Mathematics with Applications, 59(5):1656–1666, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005422.

Ciarlet:2017:AEF

[CFS17] Patrick Ciarlet, Jr., Sonia Fliss, and Christian Stohrer. On the approximation of electromagnetic fields by edge finite elements. Part 2: A heterogeneous multiscale method for Maxwell’s equations. Computers and Mathematics with Applications, 73(9):1900–1919, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301220.

Chandrashekar:2013:VCF

[CG13] Praveen Chandrashekar and Ashish Garg. Vertexcentroid finite volume scheme on tetrahedral grids for conservation laws. Computers and Mathematics with Applications, 65(1):58–74, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006311.

Clavero:2014:UCA

[CG14] C. Clavero and J. L. Gracia. Uniformly convergent additive finite difference schemes for singularly perturbed parabolic reaction–diffusion systems. Computers and Mathematics with Applications, 67(3):655–670, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006901.

Cohen:2015:SPP

[CG15] Yosef Cohen and Gonzalo Galiano. On a singular perturbation problem arising in the theory of evolutionary distributions. Computers and Mathematics with Applications, 69(3):145–156, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005318.

Cangiani:2019:ADG

[CGGM19] A. Cangiani, E. H. Georgoulis, S. Giani, and S. Metcalfe. hpadaptive discontinuous Galerkin methods for non-stationary convection– diffusion problems. Computers and Mathematics with Applications, 78 (9):3090–3104, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302007.

Carstensen:2014:DHD

[CGH14] Carsten Carstensen, Dietmar Gallistl, and Jun Hu. A discrete Helmholtz decomposition with Morley finite element functions and the optimality of adaptive finite element schemes. Computers and Mathematics with Applications, 68 (12):2167–2181, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S089812211400337X.

Carstensen:2014:LOD

[CGHW14] C. Carstensen, D. Gallistl, F. Hellwig, and L. Weggler. Low-order dPG–FEM for an elliptic PDE. Computers and Mathematics with Applications, 68 (11):1503–1512, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004611.

Ceng:2011:HSP

[CGHY11] Lu-Chuan Ceng, Sy-Ming Guu, H.-Y. Hu, and Jen-Chih Yao. Hybrid shrinking projection method for a generalized equilibrium problem, a maximal monotone operator and a countable family of relatively nonexpansive mappings. Computers and Mathematics with Applications, 61(9):2468– 2479, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001143.

Crossen:2014:EEA

[CGJ^{+}14] E. Crossen, M. S. Gockenbach, B. Jadamba, A. A. Khan, and B. Winkler. An equation error approach for the elasticity imaging inverse problem for predicting tumor location. Computers and Mathematics with Applications, 67(1):122–135, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006184.

Chiarella:2014:CST

[CGK14] Carl Chiarella, Susanne Griebsch, and Boda Kang. A comparative study on time-efficient methods to price compound options in the Heston model. Computers and Mathematics with Applications, 67(6):1254– 1270, April 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114000200.

Cima:2010:PM

[CGM10] Anna Cima, Armengol Gasull, and Víctor Mañosa. On Poncelet’s maps. Computers and Mathematics with Applications, 60(5):1457–1464, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004414.

Colmenares:2019:PEA

[CGO19] Eligio Colmenares, Gabriel N.

Gatica, and Ricardo Oyarzúa. A posteriori error analysis of an augmented fully-mixed formulation for the stationary Boussinesq model. Computers and Mathematics with Applications, 77(3):693–714, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305960.

Cimrak:2012:MSP

[CGS12] I. Cimrák, M. Gusenbauer, and T. Schrefl. Modelling and simulation of processes in microfluidic devices for biomedical applications. Computers and Mathematics with Applications, 64(3):278–288, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000831.

Ceng:2010:GIM

[CGY10a] Lu-Chuan Ceng, Sy-Ming Guu, and Jen-Chih Yao. A general iterative method with strongly positive operators for general variational inequalities. Computers and Mathematics with Applications, 59(4):1441–1452, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007378.

Ceng:2010:IAG

[CGY10b] Lu-Chuan Ceng, Sy-Ming Guu, and Jen-Chih Yao. Iterative algorithms for a general system of generalized nonlinear mixed composite-type equilibria. Computers and Mathematics with Applications, 60(3):890–905, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003925.

Ceng:2011:GCI

[CGY11] Lu-Chuan Ceng, Sy-Ming Guu, and Jen-Chih Yao. A general composite iterative algorithm for nonexpansive mappings in Hilbert spaces. Computers and Mathematics with Applications, 61(9):2447– 2455, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001118.

Cao:2011:CAH

[CH11a] Li Cao and Bo Han.

Convergence analysis of the homotopy perturbation method for solving nonlinear ill-posed operator equations. Computers and Mathematics with Applications, 61(8):2058– 2061, April 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211000653X.

Choi:2011:AOH

[CH11b] Junesang Choi and Anvar Hasanov. Applications of the operator H(α, β) to the Humbert double hypergeometric functions. Computers and Mathematics with Applications, 61(3):663–671, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S089812211000917X.

Chrysafinos:2017:SDE

[CH17] Konstantinos Chrysafinos and L. Steven Hou. Semidiscrete error estimates of the evolutionary Stokes equations with inhomogeneous Dirichlet boundary data. Computers and Mathematics with Applications, 73(8):1684–1696, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300779.

Casati:2019:CFE

[CH19] D. Casati and R. Hiptmair. Coupling finite elements and auxiliary sources. Computers and Mathematics with Applications, 77(6):1513–1526, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830511X.

Chai:2011:ERB

[Cha11a]  Guoqing Chai.  Exis-

tence results for boundary value problems of nonlinear fractional differential equations. Computers and Mathematics with Applications, 62 (5):2374–2382, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005797.

Chandok:2011:SCF

[Cha11b] Sumit Chandok. Some common fixed point theorems for generalized nonlinear contractive mappings. Computers and Mathematics with Applications, 62(10):3692– 3699, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111007772.

Chang:2011:EFM

[Cha11c] Jin-Rong Chang. The Exp-function method and generalized solitary solutions. Computers and Mathematics with Applications, 61(8):2081–2084, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006620.

See comments [NTR15].

Chang:2013:DLP

[Cha13] Hangbae Chang. The design of leakage prevention service for industry databases. Computers and Mathematics with Applications, 65(9):1369– 1377, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000466.

Chaudhary:2018:FEA

[Cha18]  Sudhakar  Chaudhary.

Finite element analysis of nonlocal coupled parabolic problem using Newton’s method. Computers and Mathematics with Applications, 75(3):981–1003, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306922.

Chan:2014:RDM

[CHBTD14] Jesse Chan, Norbert Heuer, Tan Bui-Thanh, and Leszek Demkowicz. A robust DPG method for convection-dominated diffusion problems II: Adjoint boundary conditions and mesh-dependent test norms. Computers and Mathematics with Applications, 67(4):771–795, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003751.

Chen:2011:SPF

[Che11a] Chao-Ping Chen. Some properties of functions related to the gamma, psi and tetragamma functions. Computers and Mathematics with Applications, 62(9):3389–3395, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007267.

Chen:2011:AVI

[Che11b] Fen-Ying Chen. Analytical VaR for international portfolios with common jumps. Computers and Mathematics with Applications, 62(8):3066–3076, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100681X.

Chen:2011:CCH

[Che11c] Guoxin Chen. Controlling chaotic and hyperchaotic systems via a simple adaptive feedback controller. Computers and Mathematics with Applications, 61(8):2031–2034, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006358.

Chen:2011:FMA

[Che11d] Jian Chen. Fast multilevel augmentation methods for nonlinear boundary value problems. Computers and Mathematics with Applications, 61(3):612–619, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009120.

Chen:2011:BPF

[Che11e] Mu-Yen Chen. Bankruptcy prediction in firms with statistical and intelligent techniques and a comparison of evolutionary computation approaches. Computers and Mathematics with Applications, 62(12):4514–4524, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008947.

Che:2012:CSS

[Che12a]  Z. H. Che.  Cluster-

ing and selecting suppliers based on simulated annealing algorithms. Computers and Mathematics with Applications, 63(1):228–238, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009801.

Chen:2012:CDF

[Che12b]  Fulai Chen.  Coinci-

dence degree and fractional boundary value problems with impulses. Computers and Mathematics with Applications, 64(10):3444–3455, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001368.

Chen:2012:EDM

[Che12c] Shin-Yong Chen. An equivalent direct modeling of a rotary shaft with hot-fit components using contact element modal analysis results. Computers and Mathematics with Applications, 64 (5):1093–1099, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002325.

Chen:2014:SCF

[Che14]  Long Chen.  A sim-

ple construction of a Fortin operator for the two dimensional Taylor– Hood element. Computers and Mathematics with Applications, 68 (10):1368–1373, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004519.

Chen:2015:MMS

[Che15a]  Long Chen.  Multi-

grid methods for saddle point systems using constrained smoothers. Computers and Mathematics with Applications, 70(12):2854–2866, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122115004605.

Chen:2015:RBD

[Che15b] Yanlai Chen. Reduced basis decomposition: a certified and fast lossy data compression algorithm. Computers and Mathematics with Applications, 70(10):2566–2574, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004630.

Chen:2016:DPP

[Che16] Yujuan Chen. Dynamics of prey–predator nspecies models with density dependent diffusion. Computers and Mathematics with Applications, 72(6):1727–1742, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304394.

Chen:2018:VMC

[Che18] Yi-Xiang Chen. Vortex and multipole coupled solitons in the spatially modulated cubicquintic-septimal nonlinear material. Computers and Mathematics with Applications, 76 (9):2119–2128, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304280.

Chen:2019:UCP

[Che19a] Mo Chen. Unique continuation property for the Zakharov–kuznetsov equation. Computers and Mathematics with Applications, 77(5):1273–1281, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306527.

Chen:2019:DFV

[Che19b] Shuangshuang Chen. A discontinuous finite volume method for a coupled fracture model. Computers and Mathematics with Applications, 78 (10):3429–3449, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302779.

Cliffe:2014:ADG

[CHH14] K. A. Cliffe, E. J. C.

Hall, and P. Houston. hp-adaptive discontinuous Galerkin methods for bifurcation phenomena in open flows. Computers and Mathematics with Applications, 67(4):796– 806, March 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113005919.

Canak:2011:STTb

[ÇHK11] Ibrahim Çanak, Ferhat Hasekiler, and Duygu Kebapci. Some Tauberian theorems for regularly generated sequences. Computers and Mathematics with Applications, 62(12):4486–4491, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008911.

Cermak:2018:CDF

[CHK^{+}18] Martin Cermák, Václav Hapla, Jakub Kruzík, Alexandros Markopoulos, and Alena Vasatová. Comparison of different FETI preconditioners for elastoplasticity. Computers and Mathematics with Applications, 74(1):96–109, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300135.

Chen:2018:SSA

[CHL18] Wen Chen, Yongxing Hong, and Ji Lin. The sample solution approach for determination of the optimal shape parameter in the multiquadric function of the Kansa method. Computers and Mathematics with Applications, 75(8):2942–2954, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300397.

Chen:2015:MMF

[CHLY15] Yanping Chen, Yunqing Huang, Wenbin Liu, and Ningning Yan. A mixed multiscale finite element method for convex optimal control problems with oscillating coefficients. Computers and Mathematics with Applications, 70(4):297–313, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S089812211500125X.

Caceres:2010:GHN

[CHM^{+}10] J. Cáceres, C. Hernando, M. Mora, I. M. Pelayo, and M. L. Puertas. On the geodetic and the hull numbers in strong product graphs. Computers and Mathematics with Applications, 60 (11):3020–3031, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007765.

Cakoni:2018:ISP

[CHM18] Fioralba Cakoni, Isaac Harris, and Shari Moskow. The imaging of small perturbations in an anisotropic media. Computers and Mathematics with Applications, 74(11):2769– 2783, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211730408X.

Choi:2017:EAF

[Cho17] Hyung Jun Choi. An error analysis of the finite element method overcoming corner singularities for the stationary Stokes problem. Computers and Mathematics with Applications, 73(3):450–464, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630671X.

Caballero:2011:PSC

[CHS11] J. Caballero, J. Harjani, and K. Sadarangani. Positive solutions for a class of singular fractional boundary value problems. Computers and Mathematics with Applications, 62(3):1325–1332, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003002.

Claeys:2018:SKB

[CHS18] Xavier Claeys, Ralf Hiptmair, and Elke Spindler. Second-kind boundary integral equations for electromagnetic scattering at composite objects. Computers and Mathematics with Applications, 74 (11):2650–2670, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305060.

Coronel:2019:NES

[CHS19] Aníbal Coronel, Fernando Huancas, and Mauricio Sepúlveda. A note on the existence and stability of an inverse problem for a SIS model. Computers and Mathematics with Applications, 77(12):3186–3194, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300616.

Cheng:2011:PDO

[CHT11] Ching-Hsue Cheng, Sue-Fen Huang, and Hia-Jong Teoh. Predicting daily ozone concentration maxima using fuzzy time series based on a twostage linguistic partition method. Computers and Mathematics with Applications, 62(4):2016–2028, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005244.

Chung:2010:SFE

[Chu10] Jae-Young Chung. Stability of functional equations on restricted domains in a group and their asymptotic behaviors. Computers and Mathematics with Applications, 60(9):2653–2665, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006887.

Chun:2011:NSW

[Chu11a] Changbum Chun. New solitary wave solutions to nonlinear evolution equations by the Exp-function method. Computers and Mathematics with Applications, 61(8):2107–2110, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000667X.

Chung:2011:EPQ

[Chu11b] Kun-Jen Chung. The economic production quantity with rework process in supply chain management. Computers and Mathematics with Applications, 62(6):2547–2550, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005979.

Chung:2011:SIA

[Chu11c] Kun-Jen Chung. Some improved algorithms to locate the optimal solutions for exponentially deteriorating items under trade credit financing in a supply chain system. Computers and Mathematics with Applications, 61(9):2353–2361, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009594.

Chu:2012:EBS

[Chu12a] Edward T.-H. Chu. Energybalanced sampling workload allocation in wireless sensor networks. Computers and Mathematics with Applications, 64 (5):1376–1389, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200288X.

Chung:2012:IIM

[Chu12b] Kun-Jen Chung. The integrated inventory model with the transportation cost and two-level trade credit in supply chain management. Computers and Mathematics with Applications, 64 (6):2011–2033, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003203.

Chung:2018:DSD

[Chu18] Soon-Yeong Chung. Dichotomy of solutions to discrete p-Laplace equations and p-Laplace parabolic equations. Computers and Mathematics with Applications, 75(8):2915–2924, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300373.

Cen:2018:NAT

[CHXL18] Zhongdi Cen, Jian Huang, Aimin Xu, and Anbo Le. Numerical approximation of a timefractional Black–Scholes equation. Computers and Mathematics with Applications, 75(8):2874–2887, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300269.

Chen:2012:IPP

[CHY12] Yuh-Shyan Chen, Chih-Shun Hsu, and Wei-Han Yi. An IP passing protocol for vehicular ad hoc networks with network fragmentation. Computers and Mathematics with Applications, 63(2):407–426, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006377.

Chen:2019:SBE

[CHY19] Yaoyao Chen, Yunqing Huang, and Nianyu Yi. A SCR-based error estimation and adaptive finite element method for the Allen–Cahn equation. Computers and Mathematics with Applications, 78(1):204–223, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300963.

Chen:2019:EMS

[CHZ19] Jianhua Chen, Xianjiu Huang, and Chuanxi Zhu. Existence of multiple solutions for nonhomogeneous Schrödinger– Kirchhoff system involving the fractional p-Laplacian with signchanging potential. Computers and Mathematics with Applications, 77(10):2725–2739, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300094.

Capetillo:2018:MIM

[CI18] Azael Capetillo and Fernando Ibarra. Multiphase injector modelling for automotive SCR systems: A full factorial design of experiment and optimization. Computers and Mathematics with Applications, 74(1):188–200, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300652.

Ciarlet:2012:CAD

[Cia12] Patrick Ciarlet, Jr. T - coercivity: Application to the discretization of Helmholtz-like problems. Computers and Mathematics with Applications, 64(1):22–34, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001599.

Ciarlet:2016:AEF

[Cia16] Patrick Ciarlet, Jr. On the approximation of electromagnetic fields by edge finite elements.Part 1: Sharp interpolation results for low-regularity fields. Computers and Mathematics with Applications, 71(1):85–104, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005155.

Cieplinski:2011:GHU

[Cie11] Krzysztof Ciepliński. On the generalized Hyers– Ulam stability of multiquadratic mappings. Computers and Mathematics with Applications, 62 (9):3418–3426, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007309.

Cieslinski:2013:LEM

[Cie13] Jan L. Cieśliński. Locally exact modifications of numerical schemes. Computers and Mathematics with Applications, 65(12):1920–1938, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002216.

Calo:2018:CES

[CIN^{+}18] Victor M. Calo, Oleg Iliev, Suzana P. Nunes, Galina Printsypar, and Meixia Shi. Cellelement simulations to optimize the performance of osmotic processes in porous membranes. Computers and Mathematics with Applications, 76(2):361–376, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302293.

Ciurlia:2011:VEC

[Ciu11] Pierangelo Ciurlia. Valuation of European continuousinstallment options. Computers and Mathematics with Applications, 62 (6):2518–2534, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005955.

Cizmesija:2012:NSD

[Ciz12] Aleksandra Cizmesija. A new sharp double inequality for generalized Heronian, harmonic and power means. Computers and Mathematics with Applications, 64(4):664–671, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011527.

Chang-Jian:2012:BCA

[CJ12]  Cai-Wan  Chang-Jian.

Bifurcation and chaos analysis of the porous squeeze film damper mounted gear-bearing system. Computers and Mathematics with Applications, 64(5):798–812, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010820.

Clavero:2015:AUC

[CJ15] C. Clavero and J. C.

Jorge. Another uniform convergence analysis technique of some numerical methods for parabolic singularly perturbed problems. Computers and Mathematics with Applications, 70(3): 222–235, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115001753.

Chen:2018:SEM

[CJ18a] Cheng Chen and Yao-Lin Jiang. Simplest equation method for some timefractional partial differential equations with conformable derivative. Computers and Mathematics with Applications, 75(8):2978–2988, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300415.

Chen:2018:PEN

[CJ18b] S. Chen and X. Y. Jiang. Parameters estimation for a new anomalous thermal diffusion model in layered media. Computers and Mathematics with Applications, 73(6):1172–1181, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305661.

Cortes:2010:RAT

[CJCV10] J.-C. Cortés, L. Jódar, F. Camacho, and L. Villafuerte. Random Airy type differential equations: Mean square exact and numerical solutions. Computers and Mathematics with Applications, 60(5):1237–1244, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004189.

Ciarlet:2017:DDM

[CJK17] P. Ciarlet, E. Jamelot, and F. D. Kpadonou. Domain decomposition methods for the diffusion equation with lowregularity solution. Computers and Mathematics with Applications, 74 (10):2369–2384, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304340.

Cassier:2018:MMD

[CJK18] Maxence Cassier, Patrick Joly, and Maryna Kachanovska. Mathematical models for dispersive electromagnetic waves: An overview. Computers and Mathematics with Applications, 74(11):2792–2830, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304546.

Cristea:2010:ERH

[CJMS10] Irina Cristea, Morteza Jafarpour, Seyed S. Mousavi, and Ali Soleymani. Enumeration of Rosenberg hypergroups. Computers and Mathematics with Applications, 60 (10):2753–2763, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007133.

Chen:2019:REE

[CJN19] Yanlai Chen, Jiahua Jiang, and Akil Narayan. A robust error estimator and a residual-free error indicator for reduced basis methods. Computers and Mathematics with Applications, 77(7):1963–1979, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306850.

Cho:2012:CBF

[CJP12] Jung-Sik Cho, Young-Sik Jeong, and Sang Oh Park. Consideration on the brute-force attack cost and retrieval cost: a hashbased radio-frequency identification (RFID) Tag Mutual Authentication Protocol. Computers and Mathematics with Applications, 69(1):58–65, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001393.

See cryptanalysis [SPLHCB14].

Cho:2015:CBF

[CJP15] Jung-Sik Cho, Young-Sik Jeong, and Sang Oh Park. Consideration on the brute-force attack cost and retrieval cost: a hashbased radio-frequency identification (RFID) tag mutual authentication protocol. Computers and Mathematics with Applications, 69(1):58–65, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001393.

Company:2010:NAS

[CJPB10] R. Company, L. Jódar, E. Ponsoda, and C. Ballester. Numerical analysis and simulation of option pricing problems modeling illiquid markets. Computers and Mathematics with Applications, 59(8):2964–2975, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001239.

Company:2010:COP

[CJPR10] R. Company, L. Jódar, J.-R. Pintos, and M.- D. Roselló. Computing option pricing models under transaction costs. Computers and Mathematics with Applications, 59(2):651–662, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007275.

Cortes:2011:CSN

[CJRR11] J.-C. Cortés, L. Jódar, J.-V. Romero, and M.- D. Roselló. A comparative study of the numerical approximation of the random Airy differential equation. Computers and Mathematics with Applications, 62(9):3411–3417, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007292.

Chan:2013:P

[CK13] David Chan and Candace Kent. Preface. Computers and Mathematics with Applications, 66(11):2155, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006263.

Choi:2015:FFE

[CK15] Hyung Jun Choi and Jae Ryong Kweon. The Fourier finite element method for the corner singularity expansion of the heat equation. Computers and Mathematics with Applications, 69(1):13–30, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400532X.

Cagman:2011:ST

[ÇKE11] Naim Çagman, Serkan Karatas, and Serdar Enginoğlu. Soft topology. Computers and Mathematics with Applications, 62(1):351–358, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004044.

Chang:2010:NBS

[CKK^{+}10] Nam Su Chang, Tae Hyun Kim, Chang Han Kim, Dong-Guk Han, and Jongin Lim. A new bit-serial multiplier over GF(p^{m}) using irreducible trinomials. Computers and Mathematics with Applications, 60(2):355–361, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000076.

Cai:2018:EEF

[CKL18] Zhiqiang Cai, Seokchan Kim, and Hyung-Chun Lee. Error estimate of a finite element method using stress intensity factor. Computers and Mathematics with Applications, 76(10):2402–2408, November 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118304619.

Cheng:2010:DEP

[CKLL10] Y. H. Cheng, S. Y.

Kung, C. K. Law, and W. C. Lian. The dual eigenvalue problems for the Sturm–Liouville system. Computers and Mathematics with Applications, 60(9):2556–2563, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006322.

Chiarella:2012:EBO

[CKM12] Carl Chiarella, Boda Kang, and Gunter H. Meyer. The evaluation of barrier option prices under stochastic volatility. Computers and Mathematics with Applications, 64(6):2034–2048, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003215.

Castro:2011:DNP

[CKMR11] Aline Castro, Sandi Klavzar, Michel Mollard, and Yoomi Rho. On the domination number and the 2-packing number of Fibonacci cubes and Lucas cubes. Computers and Mathematics with Applications, 61(9):2655– 2660, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001647.

Cha:2011:TPF

[CKN11] Kyung Joon Cha, Hee Sik Kim, and J. Neggers. Trends, probability functions and fuzzy right ideals for d-algebras. Computers and Mathematics with Applications, 62(8):2988–2994, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006687.

Cveticanin:2010:AMV

[CKR10] L. Cveticanin, I. Kovacic, and Z. Rakaric.

Asymptotic methods for vibrations of the pure non-integer order oscillator. Computers and Mathematics with Applications, 60(9):2616–2628, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006735.

Chertock:2019:AMM

[CKRW19] Alina Chertock, Alexander Kurganov, Mario Ricchiuto, and Tong Wu. Adaptive moving mesh upwind scheme for the two-species chemotaxis model. Computers and Mathematics with Applications, 77(12):3172–3185, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300513.

Cai:2014:NSC

[CKSL^{+}14] Q. Cai, S. Kollmannsberger, E. Sala-Lardies, A. Huerta, and E. Rank. On the natural stabilization of convection dominated problems using high order Bubnov– Galerkin finite elements. Computers and Mathematics with Applications, 66(12):2545–2558, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005609.

Ceng:2013:GSV

[CKW13] Lu-Chuan Ceng, Zhao-Rong Kong, and Ching-Feng Wen. On general systems of variational inequalities. Computers and Mathematics with Applications, 66(8):1514–1532, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005105.

Cai:2011:AMF

[CL11] Xu-Chu Cai and Jun-Fang Liu. Application of the modified frequency formulation to a nonlinear oscillator. Computers and Mathematics with Applications, 61(8):2237–2240, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000711X.

Chang:2012:OMP

[CL12a] Wen Liang Chang and Jyh-Horng Lin. Optimal maintenance policy and length of extended warranty within the life cycle of products. Computers and Mathematics with Applications, 63(1):144–150, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009679.

Chen:2012:CPA

[CL12b] Hung-Yi Chen and Jin-Wei Liang. Control of a 3D piezo-actuating table by using an adaptive sliding-mode controller for a drilling process. Computers and Mathematics with Applications, 64(5):1226–1234, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002714.

Chen:2012:PBT

[CL12c] Young-Long Chen and Hung-Pin Lai. Prioritybased transmission rate control with a fuzzy logical controller in wireless multimedia sensor networks. Computers and Mathematics with Applications, 64(5):688–698, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008029.

Cheng:2015:GRM

[CL15] Jianfeng Cheng and Yujun Liu. Global regularity of the 2D magnetic micropolar fluid flows with mixed partial viscosity. Computers and Mathematics with Applications, 70(1):66– 72, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002072.

Chen:2016:SSN

[CL16a] Ruipeng Chen and Xiaoya Li. The steady states of a non-cooperative model arising in reactor dynamics. Computers and Mathematics with Applications, 72(3):594–602, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122116302826.

Cong:2016:BKS

[CL16b] Yuhao Cong and Dongping Li. Block Krylov subspace methods for approximating the linear combination of φfunctions arising in exponential integrators. Computers and Mathematics with Applications, 72(4):846–855, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302991.

Castillo:2017:SPP

[CL17a] Ricardo Castillo and Miguel Loayza. A semilinear parabolic problem with variable reaction on a general domain. Computers and Mathematics with Applications, 74(3):351–359, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302511.

Chou:2017:FAM

[CL17b] Lot-Kei Chou and Siu-Long Lei. Fast ADI method for high dimensional fractional diffusion equations in conservative form with preconditioned strategy. Computers and Mathematics with Applications, 73(3):385–403, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306666.

Cao:2019:SRS

[CL19] K. Cao and D. Lesnic. Simultaneous reconstruction of the spatiallydistributed reaction coefficient, initial temperature and heat source from temperature measurements at different times. Computers and Mathematics with Applications, 78(10):3237–3249, November 15, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122119304638.

Chen:2019:FAD

[CLA19] Ruige Chen, Fawang Liu, and Vo Anh. A fractional alternating-direction implicit method for a multiterm time-space fractional Bloch–Torrey equations in three dimensions. Computers and Mathematics with Applications, 78(5):1261–1273, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306886.

Chen:2014:NSN

[CLB14] S. Chen, F. Liu, and K. Burrage. Numerical simulation of a new two-dimensional variableorder fractional percolation equation in nonhomogeneous porous media. Computers and Mathematics with Applications, 67(9):1673–1681, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001266.

Chen:2016:SMT

[CLC16] Hu Chen, Shujuan Lü, and Wenping Chen. Spectral methods for the time fractional diffusion-wave equation in a semi-infinite channel. Computers and Mathematics with Applications, 71(9):1818–1830, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300785.

Chen:2014:IVP

[CLCF14] Pengyu Chen, Yongxiang Li, Qiyu Chen, and Binhua Feng. On the initial value problem of fractional evolution equations with noncompact semigroup. Computers and Mathematics with Applications, 67(5):1108–1115, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000030.

Chen:2010:SSP

[CLF10] C. R. Chen, S. J. Li, and Z. M. Fang. On the solution semicontinuity to a parametric generalized vector quasivariational inequality. Computers and Mathematics with Applications, 60(8):2417–2425, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006000.

Chen:2013:STR

[CLH13] Yanping Chen, Zuliang Lu, and Yunqing Huang. Superconvergence of triangular Raviart–Thomas mixed finite element methods for a bilinear constrained optimal control problem. Computers and Mathematics with Applications, 66 (8):1498–1513, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005130.

Chi:2011:NIS

[CLJ11] Guangsheng Chi, Gongsheng Li, and Xianzheng Jia. Numerical inversions of a source term in the FADE with a Dirichlet boundary condition using final observations. Computers and Mathematics with Applications, 62(4):1619–1626, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001155.

Cui:2011:NID

[CLL11] Jing Cui, Weizhong Li, and Wei-Haur Lam. Numerical investigation on drag reduction with superhydrophobic surfaces by lattice-Boltzmann method. Computers and Mathematics with Applications, 61(12):3678–3689, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005298.

Chen:2019:BEF

[CLL19] Linchong Chen, Xin Liu, and Xiaolin Li. The boundary element-free method for 2D interior and exterior Helmholtz problems. Computers and Mathematics with Applications, 77 (3):846–864, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306199.

Chen:2011:TRD

[CLM11] Xing Chen, Juan Liu, and Jixiang Meng. Total restrained domination in graphs. Computers and Mathematics with Applications, 62(8):2892–2898, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006341.

Chapwanya:2012:EKE

[CLM12] Michael Chapwanya, Jean M.-S. Lubuma, and Ronald E. Mickens. From enzyme kinetics to epidemiological models with Michaelis– Menten contact rate: Design of nonstandard finite difference schemes. Computers and Mathematics with Applications, 64(3):201–213, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011291.

Chapwanya:2014:PPN

[CLM14] Michael Chapwanya, Jean M.-S. Lubuma, and Ronald E. Mickens. Positivitypreserving nonstandard finite difference schemes for cross-diffusion equations in biosciences. Computers and Mathematics with Applications, 68 (9):1071–1082, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001813.

Cheng:2018:LBJ

[CLMM18] Guanghui Cheng, Shanman Li, Jifei Miao, and Eric Moreau. LU-based Jacobi-like algorithms for non-orthogonal joint diagonalization. Computers and Mathematics with Applications, 76(1):113– 124, July 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118302050.

Creuse:2019:TGE

[CLN^{+}19] E. Creusé, Y. Le Menach, S. Nicaise, F. Piriou, and R. Tittarelli. Two guaranteed equilibrated error estimators for harmonic formulations in eddy current problems. Computers and Mathematics with Applications, 77(6):1549–1562, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304723.

Chang:2019:DSD

[CLS19] Qingquan Chang, Dandan Li, and Chunyou Sun. Dynamics for a stochastic degenerate parabolic equation. Computers and Mathematics with Applications, 77(9):2407–2431, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307211.

Cao:2013:TDA

[CLT^{+}13] Lili Cao, Yan Li, Guohui Tian, Baodong Liu, and YangQuan Chen. Time domain analysis of the fractional order weighted distributed parameter Maxwell model. Computers and Mathematics with Applications, 66(5):813–823, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006876.

Chen:2011:NMF

[CLTA11] Chang-Ming Chen, F. Liu, I. Turner, and V. Anh. Numerical methods with fourth-order spatial accuracy for variable-order nonlinear Stokes’ first problem for a heated generalized second grade fluid. Computers and Mathematics with Applications, 62(3):971–986, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002288.

Chen:2011:UIC

[CLW11] Weishi Chen, Tiejun Liu, and Baofa Wang. Ultrasonic image classification based on support vector machine with two independent component features. Computers and Mathematics with Applications, 62(7):2696–2703, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005311.

Caboussat:2010:NAE

[CM10a] A. Caboussat and G. K. Miers. Numerical approximation of electromagnetic signals arising in the evaluation of geological formations. Computers and Mathematics with Applications, 59(1):338–351, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812210900457X.

Choudhury:2010:PCC

[CM10b] Binayak S. Choudhury and N. Metiya. The point of coincidence and common fixed point for a pair of mappings in cone metric spaces. Computers and Mathematics with Applications, 60(6):1686–1695, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004621.

Cureg:2010:NRS

[CM10c] Edgardo Cureg and Arunava Mukherjea. Numerical results on some generalized random Fibonacci sequences. Computers and Mathematics with Applications, 59(1):233–246, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005136.

Caldera:2011:NUB

[CM11a] C. R. Caldera and A. Milgram. Notes on uniqueness of the Boubaker Polynomials Expansion Scheme (BPES) solution in the case of the Klein–Gordon equation. Computers and Mathematics with Applications, 62(1):536–538, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111002914.

Chen:2011:NMK

[CM11b] Liang Chen and Yanfang Ma. A new modified King–Werner method for solving nonlinear equations. Computers and Mathematics with Applications, 62 (10):3700–3705, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007802.

Chen:2011:MSR

[CM11c] Linjie Chen and Changfeng Ma. A modified smoothing and regularized Newton method for monotone second-order cone complementarity problems. Computers and Mathematics with Applications, 61(5):1407–1418, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000149.

Chen:2012:NSC

[CM12a] Chao-Ping Chen and Cristinel Mortici. New sequence converging towards the Euler–Mascheroni constant. Computers and Mathematics with Applications, 64(4):391–398, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002793.

Costa:2012:GAD

[CM12b] Gabriele Costa and Ilaria Matteucci. Gate automatadriven run-time enforcement. Computers and Mathematics with Applications, 63(2):518–524, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100719X.

Chen:2013:NSS

[CM13a] Jiageng Chen and Atsuko Miyaji. Novel strategies for searching RC4 key collisions. Computers and Mathematics with Applications, 66(1):81– 90, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001867.

Chen:2013:GBP

[CM13b] Ruipeng Chen and Ruyun Ma. Global bifurcation of positive radial solutions for an elliptic system in reactor dynamics. Computers and Mathematics with Applications, 65(8):1119–1128, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000680.

Clain:2014:VHO

[CM14] S. Clain and G. J. Machado. A very high-order finite volume method for the timedependent convection– diffusion problem with Butcher Tableau extension. Computers and Mathematics with Applications, 68(10):1292– 1311, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114004337.

Castillo:2015:DDM

[CM15] María Emilia Castillo and Pedro Morin. On a dissolution-diffusion model. existence, uniqueness, regularity and simulations. Computers and Mathematics with Applications, 70(8):1887–1905, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003727.

Cao:2016:SCG

[CM16a] Yang Cao and Shu-Xin Miao. On semiconvergence of the generalized shift-splitting iteration method for singular nonsymmetric saddle point problems. Computers and Mathematics with Applications, 71(7):1503–1511, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300815.

Chen:2016:AUI

[CM16b] Cai-Rong Chen and Chang-Feng Ma. AOR–Uzawa iterative method for a class of complex symmetric linear system of equations. Computers and Mathematics with Applications, 72(9):2462–2472, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305077.

Chinosi:2016:VEM

[CM16c] Claudia Chinosi and L. Donatella Marini. Virtual Element Method for fourth order problems: L^{2}-estimates. Computers and Mathematics with Applications, 72(8):1959–1967, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300438.

Chen:2017:MCI

[CM17] Cai-Rong Chen and Chang-Feng Ma. A matrix CRS iterative method for solving a class of coupled Sylvester-transpose matrix equations. Computers and Mathematics with Applications, 74 (6):1223–1231, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303498.

Carcione:2018:RBS

[CM18a] José M. Carcione and Francesco Mainardi. On the relation between sources and initial conditions for the wave and diffusion equations. Computers and Mathematics with Applications, 73(6):906–913, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301973.

Chen:2018:LSG

[CM18b] Shou-Ting Chen and Wen-Xiu Ma. Lump solutions of a generalized Calogero–Bogoyavlenskii– Schiff equation. Computers and Mathematics with Applications, 76(7):1680–1685, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303845.

Chen:2018:NMP

[CM18c] Yong Chen and Jingtang Ma. Numerical methods for a partial differential equation with spatial delay arising in option pricing under hardto-borrow model. Computers and Mathematics with Applications, 76 (9):2129–2140, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118304292.

Chen:2019:ACR

[CM19a] Cairong Chen and Changfeng Ma. An accelerated cyclic-reduction-based solvent method for solving quadratic eigenvalue problem of gyroscopic systems. Computers and Mathematics with Applications, 77(10):2585– 2595, May 15, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307387.

Chiravalle:2019:LCC

[CM19b] Vincent P. Chiravalle and Nathaniel R. Morgan. A 3D Lagrangian cellcentered hydrodynamic method with higherorder reconstructions for gas and solid dynamics. Computers and Mathematics with Applications, 78(2):298–317, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303353.

Cardenas-Morales:2011:BTO

[CMGR11]  D.  Cárdenas-Morales,

P. Garrancho, and I. Rasa.

Bernstein-type operators which preserve polynomials. Computers and Mathematics with Applications, 62(1):158–163, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100366X.

Caceres:2011:TNF

[CMMP11] José Cáceres, Alberto Márquez, María Morales, and María Luz Puertas. Towards a new framework for domination. Computers and Mathematics with Applications, 62(1):44– 50, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003397.

Cao:2017:PGS

[CMR17] Yang Cao, Shu-Xin Miao, and Zhi-Ru Ren. On preconditioned generalized shift-splitting iteration methods for saddle point problems. Computers and Mathematics with Applications, 74(4):859–872, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117303462.

Cheng:2010:DRS

[CMS10] Teddy M. Cheng, Veerachai Malyavej, and Andrey V. Savkin. Decentralized robust set-valued state estimation in networked multiple sensor systems. Computers and Mathematics with Applications, 59(8):2636–2646, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000544.

Cresson:2012:TSD

[CMT12] Jacky Cresson, Agnieszka B. Malinowska, and Delfim F. M. Torres. Time scale differential, integral, and variational embeddings of Lagrangian systems. Computers and Mathematics with Applications, 64(7):2294–2301, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002088.

Costabile:2011:CCM

[CN11] F. Costabile and A. Napoli. A class of collocation methods for numerical integration of initial value problems. Computers and Mathematics with Applications, 62(8):3221–3235, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006997.

Chalupa:2013:MMP

[CN13] Petr Chalupa and Jakub Novák. Modeling and model predictive control of a nonlinear hydraulic system. Computers and Mathematics with Applications, 66(2):155–164, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000382.

Colmenares:2016:DMF

[CN16a] Eligio Colmenares and Michael Neilan. Dualmixed finite element methods for the stationary Boussinesq problem. Computers and Mathematics with Applications, 72(7):1828–1850, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116304655.

Coyle:2016:HOD

[CN16b] Joe Coyle and Nilima Nigam. High-order discontinuous Galerkin methods for a class of transport equations with structured populations. Computers and Mathematics with Applications, 72(3):768–784, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302929.

Castano:2017:DVS

[CNH17] D. Castaño, M. C. Navarro, and H. Herrero. Double vortices and single-eyed vortices in a rotating cylinder under thermal gradients. Computers and Mathematics with Applications, 73(10):2238–2257, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301724.

Cho:2014:DDP

[CNP14] Sungmin Cho, S. V. Nepomnyaschikh, and Eun-Jae Park. Domain decomposition preconditioning for elliptic problems with jumps in coefficients. Computers and Mathematics with Applications, 68 (12):2292–2313, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004039.

Caratelli:2010:FSW

[CNR10] Diego Caratelli, Pierpaolo Natalini, and Paolo E. Ricci. Fourier solution of the wave equation for a star-like-shaped vibrating membrane. Computers and Mathematics with Applications, 59(1):176–184, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900580X.

Canuto:2017:RSH

[CNSV17] Claudio Canuto, Ricardo H. Nochetto, Rob Stevenson, and Marco Verani. On p-robust saturation for hp-AFEM. Computers and Mathematics with Applications, 73(9):2004–2022, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301141.

Canuto:2014:COP

[CNV14] C. Canuto, R. H. Nochetto, and M. Verani. Contraction and optimality properties of adaptive Legendre–Galerkin methods: the one-dimensional case. Computers and Mathematics with Applications, 67(4):752–770, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003477.

Cecot:2015:HOF

[CO15] W. Cecot and M. Oleksy. High order FEM for multigrid homogenization. Computers and Mathematics with Applications, 70(7):1391–1400, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003107.

Chen:2019:EBB

[CO19] Bin Chen and Zeng Qi Ou. Existence and bifurcation behavior of positive solutions for a class of Kirchhoff-type problems. Computers and Mathematics with Applications, 77(10):2859–2866, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300422.

Collado:2014:HCL

[Col14] Francisco J. Collado.

Hyperbolic conservation laws for continuous twophase flow without mass exchange. Computers and Mathematics with Applications, 67(8):1622–1630, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000856.

Colton:2018:PMI

[Col18]  David Colton.  Peter

Monk and inverse scattering theory. Computers and Mathematics with Applications, 74 (11):2640–2644, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117300676.

Cooke:2010:PMB

[Coo10] Charlie H. Cooke. Probability models for blackjack poker. Computers and Mathematics with Applications, 59(1):108– 114, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006543.

Chichurin:2018:MQT

[COR18] Alexander V. Chichurin, Elena M. Ovsiyuk, and Viktor M. Red’kov. Modeling the quantum tunneling effect for a particle with intrinsic structure in presence of external magnetic field in the Lobachevsky space. Computers and Mathematics with Applications, 75(5):1550–1565, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307332.

Costa:2018:FBF

[Cos18] Pedro Costa. A FFTbased finite-difference solver for massivelyparallel direct numerical simulations of turbulent flows. Computers and Mathematics with Applications, 76(8):1853–1862, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830405X.

Covei:2013:RSQ

[Cov13] Dragos-Patru Covei. Radial solutions for a quasilinear elliptic system of Schrödinger type. Computers and Mathematics with Applications, 65(8):1187–1193, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001211.

Chitcharoen:2010:NMF

[CP10] Doungrat Chitcharoen and Puntip Pattaraintakorn. Novel matrix forms of rough set flow graphs with applications to data integration. Computers and Mathematics with Applications, 60 (10):2880–2897, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007534.

Cao:2015:TPP

[CP15a] X. Cao and I. S. Pop. Two-phase porous media flows with dynamic capillary effects and hysteresis: Uniqueness of weak solutions. Computers and Mathematics with Applications, 69(7):688– 695, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000620.

Chadha:2015:EAS

[CP15b] Alka Chadha and D. N. Pandey. Existence and approximation of solution to neutral fractional differential equation with nonlocal conditions. Computers and Mathematics with Applications, 69(9):893–908, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000498.

Chung:2015:EPE

[CP15c] Soon-Yeong Chung and Jea-Hyun Park. Extinction and positivity for the evolution p-Laplacian equations with absorption on networks. Computers and Mathematics with Applications, 69(3):223–234, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005951.

Chatterjee:2016:ICB

[CP16a] Anal Chatterjee and Samares Pal. Interspecies competition between prey and two different predators with Holling IV functional response in diffusive system. Computers and Mathematics with Applications, 71(2):615–632, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005908.

Chehab:2016:PMF

[CP16b] Jean-Paul Chehab and Madalina Petcu. Parallel matrix function evaluation via initial value ODE modeling. Computers and Mathematics with Applications, 72(1):76– 91, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302292.

Chung:2016:CCN

[CP16c] Soon-Yeong Chung and Jea-Hyun Park. A complete characterization of nonlinear absorption for the evolution p-Laplacian equations to have positive or extinctive solutions. Computers and Mathematics with Applications, 71(8):1624–1635, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211630075X.

Choi:2011:NPS

[CPL11] Kyu Young Choi, Jong Hwan Park, and Dong Hoon Lee. A new provably secure certificateless short signature scheme. Computers and Mathematics with Applications, 61(7):1760–1768, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000897.

Carlone:2010:FEA

[CPP10] P. Carlone, G. S. Palazzo, and R. Pasquino. Finite element analysis of the steel quenching process: Temperature field and solid-solid phase change. Computers and Mathematics with Applications, 59(1):585–594, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900371X.

Caruso:2015:EAI

[CPP15] N. Caruso, M. Portapila, and H. Power. An efficient and accurate implementation of the localized regular dual reciprocity method. Computers and Mathematics with Applications, 69(11):1342– 1366, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001431.

Chernov:2015:SCB

[CPT15] Alexey Chernov, Duong Pham, and Thanh Tran. A shape calculus based method for a transmission problem with a random interface. Computers and Mathematics with Applications, 70(7):1401–1424, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003077.

Chen:2011:LMC

[CQ11] Xufeng Chen and Sen Qin. On-line machine covering on two machines with local migration. Computers and Mathematics with Applications, 62(5):2336–2341, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005748.

Chen:2013:GLC

[CQ13] Zhongwen Chen and Songqiang Qiu. Global and local convergence of a penalty-free method for nonlinear programming. Computers and Mathematics with Applications, 65(4):589–608, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200630X.

Chen:2011:IMS

[CQLX11] Zhide Chen, Yihui Qiu, Jingjing Liu, and Li Xu. Incentive mechanism for selfish nodes in wireless sensor networks based on evolutionary game. Computers and Mathematics with Applications, 62 (9):3378–3388, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007255.

Chen:2011:BSB

[CQRW11] Qiuhui Chen, Tao Qian, Guangbin Ren, and Yi Wang. B-splines of Blaschke product type. Computers and Mathematics with Applications, 62(10):3669–3681, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007577.

Chernov:2013:NQH

[CR13] Alexey Chernov and Anne Reinarz. Numerical quadrature for highdimensional singular integrals over parallelotopes. Computers and Mathematics with Applications, 66(7):1213–1231, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004458.

Chaabane:2018:SBF

[CR18a] Nabil Chaabane and Béatrice Rivière. A splitting-based finite element method for the Biot poroelasticity system. Computers and Mathematics with Applications, 75(7):2328–2337, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307721.

Chen:2018:TGD

[CR18b] Shuangshuang Chen and Hongxing Rui. A twogrid decoupled algorithm for fracture models. Computers and Mathematics with Applications, 76 (5):1161–1173, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303298.

Chernov:2019:SGA

[CR19] Alexey Chernov and Anne Reinarz. Sparse grid approximation spaces for space–time boundary integral formulations of the heat equation. Computers and Mathematics with Applications, 78 (11):3605–3619, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304626.

Castineira:2019:JVF

[CRA19] G. Castiñeira and Á. Rodríguez-Arós. On the justification of viscoelastic flexural shell equations. Computers and Mathematics with Applications, 77(11):2933– 2942, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304887.

Chen:2016:EPV

[CRG16] Qingshan Chen, Todd Ringler, and Peter R. Gent. Extending a potential vorticity transport eddy closure to include a spatially-varying coefficient. Computers and Mathematics with Applications, 71(11):2206– 2217, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115006094.

Cortes:2011:SRD

[CRRS11] J.-C. Cortés, J.-V. Romero, M.-D. Roselló, and C. Santamaría. Solving random diffusion models with nonlinear perturbations by the Wiener–Hermite expansion method. Computers and Mathematics with Applications, 61(8): 1946–1950, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005493.

Chen:2015:OBC

[CRXL15] Tehuan Chen, Zhigang Ren, Chao Xu, and Ryan Loxton. Optimal boundary control for water hammer suppression in fluid transmission pipelines. Computers and Mathematics with Applications, 69(4):275–290, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005525.

Caballero:2010:FCI

[CS10a] J. Caballero and K. Sadarangani.

Fritz Carlson’s inequality for fuzzy integrals. Computers and Mathematics with Applications, 59(8):2763–2767, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000805.

Choi:2010:MSA

[CS10b] Junesang Choi and H. M.

Srivastava. Mathieu series and associated sums involving the zeta functions. Computers and Mathematics with Applications, 59(2):861–867, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900707X.

Cholamjiak:2010:WSC

[CS10c] Watcharaporn Cholamjiak and Suthep Suantai. Weak and strong convergence theorems for a finite family of generalized asymptotically quasi-nonexpansive mappings. Computers and Mathematics with Applications, 60(7):1917–1923, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000502X.

Cheng:2011:DCM

[CS11a] Teddy M. Cheng and Andrey V. Savkin. Decentralized control of multiagent systems for swarming with a given geometric pattern. Computers and Mathematics with Applications, 61(4):731–744, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008953.

Cholamjiak:2011:EIM

[CS11b] Prasit Cholamjiak and Suthep Suantai. Existence and iteration for a mixed equilibrium problem and a countable family of nonexpansive mappings in Banach spaces. Computers and Mathematics with Applications, 61(9):2725–2733, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001817.

Cholamjiak:2011:SCC

[CS11c] Prasit Cholamjiak and Suthep Suantai. Strong convergence for a countable family of strict pseudocontractions in quniformly smooth Banach spaces. Computers and Mathematics with Applications, 62(2):787– 796, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004767.

Cholamjiak:2011:ACF

[CS11d] Watcharaporn Cholamjiak and Suthep Suantai. Approximation of common fixed points of two quasi-nonexpansive multi-valued maps in Banach spaces. Computers and Mathematics with Applications, 61(4):941–949, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009478.

Chidume:2012:ASG

[CS12] C. E. Chidume and Y. Shehu. Approximation of solutions of generalized equations of Hammerstein type. Computers and Mathematics with Applications, 63(5):966– 974, March 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111010443.

Csorgo:2013:NAS

[CS13] Gábor Csörgő and Péter L. Simon. Numerical and analytical study of bifurcations in a model of electrochemical reactions in fuel cells. Computers and Mathematics with Applications, 65(3):325–337, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003690.

Chernov:2014:HOF

[CS14a] Alexey Chernov and Christoph Schwab. Highorder finite element approximation for partial differential equations. Computers and Mathematics with Applications, 67(4):709–711, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000248.

Cioaca:2014:LRA

[CS14b] Alexandru Cioaca and Adrian Sandu. Lowrank approximations for computing observation impact in 4D-Var data assimilation. Computers and Mathematics with Applications, 67(12): 2112–2126, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114000650.

Cao:2016:TCF

[CS16] Meiling Cao and Weijie Sheng. Traveling curved fronts of bistable Lotka– Volterra competitiondiffusion systems in R^{3}. Computers and Mathematics with Applications, 71(6):1270–1286, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300451.

Chen:2018:GTU

[CS18] Leitao Chen and Laura Schaefer. Godunov-type upwind flux schemes of the two-dimensional finite volume discrete Boltzmann method. Computers and Mathematics with Applications, 75(9):3105–3126, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300506.

Ciric:2011:CFP

[ĆSCD11] Ljubomir Ćirić, Bessem Samet, Nenad Cakić, and Boško Damjanović. Coincidence and fixed point theorems for generalized (ψ, ϕ)-weak nonlinear contraction in ordered K-metric spaces. Computers and Mathematics with Applications, 62(9):3305–3316, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006365.

Cho:2013:PDA

[CSCM13] Jaeik Cho, Setiawan Soekamtoputra, Ken Choi, and Jongsub Moon. Power dissipation and area comparison of 512-bit and 1024-bit key AES. Computers and Mathematics with Applications, 65(9):1378–1383, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000454.

Cai:2010:SCT

[CsH10] Gang Cai and Chang song Hu. Strong convergence theorems of a general iterative process for a finite family of λ_{i}-strict pseudo-contractions in quniformly smooth Banach spaces. Computers and Mathematics with Applications, 59(1):149–160, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005872.

Chun:2011:TOF

[CSN11] Changbum Chun, Pantelimon Stanica, and Beny Neta. Third-order family of methods in Banach spaces. Computers and Mathematics with Applications, 61(6):1665–1675, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000514.

Cvetkovic:2010:SDG

[CSS10] Dragos Cvetković, Slobodan K. Simić, and Zoran Stanić. Spectral determination of graphs whose components are paths and cycles. Computers and Mathematics with Applications, 59(12):3849– 3857, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002920.

Chen:2012:BSS

[CSSW12] Chiau-Ching Chen, Hsu-Shih Shih, Huan-Jyh Shyur, and Kun-Shan Wu. A business strategy selection of green supply chain management via an analytic network process. Computers and Mathematics with Applications, 64(8):2544–2557, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004439.

Chamoun:2014:CAC

[CST14] Georges Chamoun, Mazen Saad, and Raafat Talhouk. A coupled anisotropic chemotaxis-fluid model: the case of two-sidedly degenerate diffusion. Computers and Mathematics with Applications, 68 (9):1052–1070, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001679.

Coudiere:2013:AFV

[CSU13] Yves Coudière, Mazen Saad, and Alexandre Uzureau. Analysis of a finite volume method for a bone growth system in vivo. Computers and Mathematics with Applications, 66(9):1581–1594, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000941.

Chen:2011:NHB

[CSW11a] Shanshan Chen, Junping Shi, and Junjie Wei. A note on Hopf bifurcations in a delayed diffusive Lotka–Volterra predator– prey system. Computers and Mathematics with Applications, 62 (5):2240–2245, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005657.

Cho:2011:CFP

[CSW11b] Yeol Je Cho, Reza Saadati, and Shenghua Wang. Common fixed point theorems on generalized distance in ordered cone metric spaces. Computers and Mathematics with Applications, 61(4):1254–1260, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000095.

Chen:2010:ADM

[CSZK10] Wen Chen, Hongguang Sun, Xiaodi Zhang, and Dean Korosak. Anomalous diffusion modeling by fractal and fractional derivatives. Computers and Mathematics with Applications, 59(5):1754– 1758, March 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005525.

Chowdhury:2010:GBQ

[CT10a] Mohammad S. R. Chowdhury and Kok-Keong Tan. Generalized biquasi-variational inequalities for quasi-pseudo-monotone type I operators on non-compact sets. Computers and Mathematics with Applications, 60(3):423–431, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000307X.

Citterio:2010:DOC

[CT10b] Maurizio Citterio and Rodolfo Talamo. Damped oscillators: a continuous model for velocity dependent drag. Computers and Mathematics with Applications, 59(1):352–359, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004453.

Canak:2011:STTa

[ÇT11a] Ibrahim Çanak and Ümit Totur. Some Tauberian theorems for the weighted mean methods of summability. Computers and Mathematics with Applications, 62(6):2609–2615, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006419.

Chen:2011:NEH

[CT11b] Peng Chen and X. H. Tang. New existence of homoclinic orbits for a second-order Hamiltonian system. Computers and Mathematics with Applications, 62(1):131– 141, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003634.

Cetin:2012:SPS

[ÇT12] Erbil Çetin and F. Serap Topal. Symmetric positive solutions of fourth order boundary value problems for an increasing homeomorphism and homomorphism on timescales. Computers and Mathematics with Applications, 63(3):669–678, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010054.

Chen:2017:GSS

[CT17a] Sitong Chen and Xianhua Tang. Ground state sign-changing solutions for asymptotically cubic or super-cubic Schrödinger–Poisson systems without compact condition. Computers and Mathematics with Applications, 74(3):446–458, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302626.

Cheng:2017:HES

[CT17b] Bitao Cheng and Xianhua Tang. High energy solutions of modified quasilinear fourthorder elliptic equations with sign-changing potential. Computers and Mathematics with Applications, 73(1):27–36, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305867.

Coudiere:2017:VHO

[CT17c] Yves Coudière and Rodolphe Turpault. Very high order finite volume methods for cardiac electrophysiology. Computers and Mathematics with Applications, 74(4):684–700, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117303012.

Chen:2018:IMS

[CT18] Sitong Chen and Xianhua Tang. Infinitely many solutions and least energy solutions for Klein– Gordon–Maxwell systems with general superlinear nonlinearity. Computers and Mathematics with Applications, 75(9):3358–3366, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300774.

Chen:2012:IMH

[CTA12] Peng Chen, X. H. Tang, and Ravi P. Agarwal. Infinitely many homoclinic solutions for nonautonomous p(t)-Laplacian Hamiltonian systems. Computers and Mathematics with Applications, 63(4):751–763, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010224.

Chen:2017:EGS

[CTC17] Jianhua Chen, Xianhua Tang, and Bitao Cheng.

Existence of ground state sign-changing solutions for a class of generalized quasilinear Schrödinger– Maxwell system in R^{3}. Computers and Mathematics with Applications, 74(3):466–481, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302596.

Choudhury:2010:BAR

[CTD10] Gautam Choudhury, Lotfi Tadj, and Kandarpa Deka. A batch arrival retrial queueing system with two phases of service and service interruption. Computers and Mathematics with Applications, 59(1):437–450, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900399X.

Chen:2017:EMS

[CTG17] Jianhua Chen, Xianhua Tang, and Zu Gao. Existence of multiple solutions for modified Schrödinger– Kirchhoff–Poisson type systems via perturbation method with signchanging potential. Computers and Mathematics with Applications, 73(3):505–519, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306770.

Chetoui:2013:NCM

[CTM^{+}13] Manel Chetoui, Magalie Thomassin, Rachid Malti, Mohamed Aoun, Slaheddine Najar, Mohamed Naceur Abdelkrim, and Alain Oustaloup. New consistent methods for order and coefficient estimation of continuous-time errors-in-variables fractional models. Computers and Mathematics with Applications, 66(5):860–872, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002873.

Capizzano:2010:NAM

[CTP10] Stefano Serra Capizzano and Cristina Tablino-Possio. A note on algebraic multigrid methods for the discrete weighted Laplacian. Computers and Mathematics with Applications, 60 (5):1290–1298, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004220.

Cleja-Tigoiu:2019:VIC

[CTS19] S. Cleja-Tigoiu and N. E.

Stoicuta. Variational inequality in classical plasticity. Applications to Armstrong-frederick elasto-plastic model. Computers and Mathematics with Applications, 77(11): 2953–2970, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118305030.

Chai:2016:SRW

[CTSX16] Jun Chai, Bo Tian, Wen-Rong Sun, and Xi-Yang Xie. Solitons and rouge waves for a generalized (3 + 1)-dimensional variablecoefficient Kadomtsev– Petviashvili equation in fluid mechanics. Computers and Mathematics with Applications, 71(10): 2060–2068, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116301638.

Cheng:2017:DMF

[CTZ17] Dongsheng Cheng, Xu Tan, and Taishan Zeng. A dispersion minimizing finite difference scheme for the Helmholtz equation based on point-weighting. Computers and Mathematics with Applications, 73(11):2345–2359, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302262.

Cheng:2012:SDE

[CUK12] Kwee-Li Cheng, Naofumi Uchihara, and Hiroyuki Kasai. Simple drift error-resilient H.264/AVC encoder for fast video transcoding using DCT coefficients. Computers and Mathematics with Applications, 64(5):1420–1430, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002945.

Chernogorova:2014:FVD

[CV14] T. Chernogorova and L. Vulkov. A finite volume difference scheme for a model of settling particle dispersion from an elevated source in an open-channel flow. Computers and Mathematics with Applications, 67(12):2099–2111, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001357.

Cveticanin:2011:ONE

[Cve11] L. Cveticanin. Oscillators with nonlinear elastic and damping forces. Computers and Mathematics with Applications, 62(4):1745–1757, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004846.

Cvijovic:2010:ETS

[Cvi10] Djurdje Cvijović. Exponential and trigonometric sums associated with the Lerch zeta and Legendre chi functions. Computers and Mathematics with Applications, 59(4):1484–1490, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000489.

Cvijovic:2011:HOT

[Cvi11] Djurdje Cvijović. Higherorder tangent and secant numbers. Computers and Mathematics with Applications, 62(4):1879–1886, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005116.

Che:2010:HAS

[CW10a] Z. H. Che and H. S. Wang. A hybrid approach for supplier cluster analysis. Computers and Mathematics with Applications, 59(2):745–763, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007172.

Chen:2010:VIM

[CW10b] Xumei Chen and Linjun Wang. The variational iteration method for solving a neutral functionaldifferential equation with proportional delays. Computers and Mathematics with Applications, 59(8):

2696–2702, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000593.

Wu:2011:FVI

[cW11] Guo cheng Wu. A fractional variational iteration method for solving fractional nonlinear differential equations. Computers and Mathematics with Applications, 61(8):2186–2190, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006966.

Chen:2014:PMF

[CW14] Wen Chen and Song Wang. A penalty method for a fractional order parabolic variational inequality governing American put option valuation. Computers and Mathematics with Applications, 67(1):77–90, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006196.

Cai:2015:SHB

[CW15a] Yongli Cai and Weiming Wang. Stability and Hopf bifurcation of the stationary solutions to an epidemic model with crossdiffusion. Computers and Mathematics with Applications, 70(8):1906–1920, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003715.

Chan:2015:FET

[CW15b] Jesse Chan and T. Warburton. hp-finite element trace inequalities for the pyramid. Computers and Mathematics with Applications, 69(6):510–517, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000358.

Chan:2017:PSM

[CW17] Jesse Chan and T. Warburton. On the penalty stabilization mechanism for upwind discontinuous Galerkin formulations of first order hyperbolic systems. Computers and Mathematics with Applications, 74 (12):3099–3110, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304819.

Chen:2018:MND

[CW18] Min-Hong Chen and Qing-Biao Wu. On modified Newton–DGPMHSS method for solving nonlinear systems with complex symmetric Jacobian matrices. Computers and Mathematics with Applications, 76(1):45–57, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302013.

Cao:2019:EMS

[CW19] Xiaofei Cao and Shu Wen. Existence and multiplicity of solutions for a quasilinear nonlocal problem. Computers and Mathematics with Applications, 78(8):2575–2583, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301919.

Chen:2017:FPP

[CWDL17] Xu Chen, Wenfei Wang, Deng Ding, and Siu-Long Lei. A fast preconditioned policy iteration method for solving the tempered fractional HJB equation governing American options valuation. Computers and Mathematics with Applications, 73(9):1932–1944, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301190.

Chen:2013:LBM

[CWH13] Poting Chen, Chen-Hao Wang, and Jeng-Rong Ho. A lattice Boltzmann model for electromagnetic waves propagating in a onedimensional dispersive medium. Computers and Mathematics with Applications, 65(6):961–973, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300031X.

Chan:2017:RSN

[CWHW17] Jesse Chan, Zheng Wang, Russell J. Hewett, and T. Warburton. Reduced storage nodal discontinuous Galerkin methods on semi-structured prismatic meshes. Computers and Mathematics with Applications, 73(5):775–793, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300305.

Chu:2012:BCE

[CWQJ12] Yu-Ming Chu, Miao-Kun Wang, Song-Liang Qiu, and Yue-Ping Jiang. Bounds for complete elliptic integrals of the second kind with applications. Computers and Mathematics with Applications, 63(7):1177–1184, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010935.

Chang:2015:FFV

[CWW15] Shuhua Chang, Jing Wang, and Xinyu Wang. A fitted finite volume method for real option valuation of risks in climate change. Computers and Mathematics with Applications, 70 (5):1198–1219, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003387.

Cao:2019:BSD

[CWW19] Qian Cao, Jianhua Wu, and Yan’e Wang. Bifurcation solutions in the diffusive minimal sediment. Computers and Mathematics with Applications, 77(3):888–906, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306333.

Chen:2015:ASM

[CWWY15] Long Chen, Junping Wang, Yanqiu Wang, and Xiu Ye. An auxiliary space multigrid preconditioner for the weak Galerkin method. Computers and Mathematics with Applications, 70(4):330–344, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001972.

Chen:2019:NOS

[CWY19] Chris Chen, Zeqi Wang, and Yue Yang. A new operator splitting method for American options under fractional Black– Scholes models. Computers and Mathematics with Applications, 77(8):2130–2144, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306990.

Chen:2016:NLS

[CX16] Chong Chen and Guoliang Xu. A new linearized split Bregman iterative algorithm for image reconstruction in sparse-view Xray computed tomography. Computers and Mathematics with Applications, 71(8):1537–1559, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000067.

Constantinou:2018:FEM

[CX18] P. Constantinou and C. Xenophontos. An hp finite element method for a 4th order singularly perturbed boundary value problem in two dimensions. Computers and Mathematics with Applications, 74(7):1565–1575, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300755.

Cheng:2019:EEQ

[CXMO19] Xiaoliang Cheng, Qichang Xiao, Stanisław Migórski, and Anna Ochal. Error estimate for quasistatic history-dependent contact model. Computers and Mathematics with Applications, 77(11): 2943–2952, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830484X.

Chen:2015:SLV

[CXZ15a] Wenting Chen, Liangbin Xu, and Song-Ping Zhu. Stock loan valuation under a stochastic interest rate model. Computers and Mathematics with Applications, 70(8):1757–1771, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003545.

Chen:2015:APD

[CXZ15b] Wenting Chen, Xiang Xu, and Song-Ping Zhu. Analytically pricing double barrier options based on a time-fractional Black– Scholes equation. Computers and Mathematics with Applications, 69(12): 1407–1419, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211500142X.

Cao:2014:AOC

[CY14a] Weidong Cao and Danping Yang. Adaptive optimal control approximation for solving a fourth-order elliptic variational inequality. Computers and Mathematics with Applications, 66(12):2517–2531, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005853.

Chen:2014:MAS

[CY14b] Nan Chen and Huidan Yu. Mechanism of axis switching in low aspect-ratio rectangular jets. Computers and Mathematics with Applications, 67(2):437–444, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001703.

Chen:2019:EIC

[CY19] Jixin Chen and Danping Yang. Explicit/ implicit and Crank– Nicolson domain decomposition methods for parabolic partial differential equations. Computers and Mathematics with Applications, 77(7):1841–1863, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306734.

Cheng:2011:ESG

[CyL11] Bo Cheng and San yang Liu. Existence of solutions for generalized equilibrium problem in g-convex space. Computers and Mathematics with Applications, 62 (9):3404–3410, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007280.

Chen:2017:RCI

[CYL17] Liying Chen, Haifang Yu, and Yang Liu. Regularity criteria for the incompressible magnetohydrodynamic flows with density-dependent viscosity coefficient. Computers and Mathematics with Applications, 74 (12):2950–2957, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304716.

Constales:2013:ICC

[CYM13] D. Constales, G. S.

Yablonsky, and G. B. Marin. Intersections and coincidences in chemical kinetics: Linear two-step reversible-irreversible reaction mechanism. Computers and Mathematics with Applications, 65(10): 1614–1624, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112006517.

Cartalade:2016:LBS

[CYP16] Alain Cartalade, Amina Younsi, and Mathis Plapp. Lattice Boltzmann simulations of 3D crystal growth: Numerical schemes for a phase-field model with anti-trapping current. Computers and Mathematics with Applications, 71(9):1784–1798, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300839.

Chen:2010:FDE

[CYS10] Wen Chen, Linjuan Ye, and Hongguang Sun. Fractional diffusion equations by the Kansa method. Computers and Mathematics with Applications, 59(5):1614–1620, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005379.

Chen:2018:LGS

[CYZZ18] Xuehui Chen, Yifan Ye, Xinru Zhang, and Liancun Zheng. Lie-group similarity solution and analysis for fractional viscoelastic MHD fluid over a stretching sheet. Computers and Mathematics with Applications, 75(8):3002–3011, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300440.

Chen:2010:NSF

[CZ10] Zhenmin Chen and Chi Zhang. Network sharing fairness function with consideration of priority levels. Computers and Mathematics with Applications, 60(9):2548–2555, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006188.

Chen:2011:AFF

[CZ11a] Fulai Chen and Yong Zhou. Attractivity of fractional functional differential equations. Computers and Mathematics with Applications, 62(3):1359–1369, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002252.

Cheng:2011:VCA

[CZ11b] Yongguang Cheng and Hui Zhang. A viscosity counteracting approach in the lattice Boltzmann BGK model for low viscosity flow: Preliminary verification. Computers and Mathematics with Applications, 61(12):3690–3702, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000903X.

Cheng:2015:MWS

[CZ15] Li Cheng and Yi Zhang. Multiple wave solutions and auto-Bäcklund transformation for the (3 + 1)-dimensional generalized B-type Kadomtsev– Petviashvili equation. Computers and Mathematics with Applications, 70(5):765–775, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002734.

Cheng:2017:GTD

[CZ17] Li Cheng and Yi Zhang. Grammian-type determinant solutions to generalized KP and BKP equations. Computers and Mathematics with Applications, 74(4):727–735, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303097.

Cui:2018:MRW

[CZ18] Wenying Cui and Zhaqilao. Multiple rogue wave and breather solutions for the (3 + 1)-dimensional KPI equation. Computers and Mathematics with Applications, 76 (5):1099–1107, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303250.

Cheng:2010:STI

[CZF10] Wei Cheng, Ling-Ling Zhao, and Chu-Li Fu. Source term identification for an axisymmetric inverse heat conduction problem. Computers and Mathematics with Applications, 59(1):142–148, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005914.

Chen:2017:SFN

[CZL17] Pengyu Chen, Xuping Zhang, and Yongxiang Li. Study on fractional non-autonomous evolution equations with delay. Computers and Mathematics with Applications, 73(5):794–803, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300299.

Chen:2011:NCS

[CZMZ11] Di-Yi Chen, Wei-Li Zhao, Xiao-Yi Ma, and Run-Fan Zhang. No-chattering sliding mode control chaos in Hindmarsh– Rose neurons with uncertain parameters. Computers and Mathematics with Applications, 61(10): 3161–3171, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111002975.

Chang:2011:SMA

[CZN11] Yong-Kui Chang, Zhi-Han Zhao, and G. M. N’Guérékata. Squaremean almost automorphic mild solutions to non-autonomous stochastic differential equations in Hilbert spaces. Computers and Mathematics with Applications, 61(2):384–391, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008631.

Chang:2012:WPA

[CZN12] Yong-Kui Chang, Rui Zhang, and G. M. N’Guérékata. Weighted pseudo almost automorphic mild solutions to semilinear fractional differential equations. Computers and Mathematics with Applications, 64 (10):3160–3170, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001642.

Chen:2011:LBS

[CZY11] Xiao-Peng Chen, Cheng-Wen Zhong, and Xu-Long Yuan. Lattice Boltzmann simulation of cavitating bubble growth with large density ratio. Computers and Mathematics with Applications, 61(12):3577– 3584, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110004943.

Chadli:2013:UIO

[CZY13] M. Chadli, I. Zelinka, and T. Youssef. Unknown inputs observer design for fuzzy systems with application to chaotic system reconstruction. Computers and Mathematics with Applications, 66(2):147–154, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000357.

Dung:2012:MFS

[DA12] Le The Dung and Beongku An. A modeling framework for supporting and evaluating performance of multi-hop paths in mobile ad-hoc wireless networks. Computers and Mathematics with Applications, 64(5):1197–1205, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002684.

Dehghan:2016:NST

[DA16] Mehdi Dehghan and Mostafa Abbaszadeh. Numerical study of threedimensional Turing patterns using a meshless method based on moving kriging element free Galerkin (EFG) approach. Computers and Mathematics with Applications, 72(3):427–454, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302310.

DAbbicco:2018:CED

[D’A18a] M. D’Abbicco. The critical exponent for the dissipative plate equation with power nonlinearity. Computers and Mathematics with Applications, 74(5):1006–1014, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301244.

Dehghan:2018:CPO

[DA18b] Mehdi Dehghan and Mostafa Abbaszadeh. A combination of proper orthogonal decompositiondiscrete empirical interpolation method (POD– DEIM) and meshless local RBF–DQ approach for prevention of groundwater contamination. Computers and Mathematics with Applications, 75(4):1390–1412, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307125.

Dehghan:2018:EFG

[DA18c] Mehdi Dehghan and Mostafa Abbaszadeh. Element free Galerkin approach based on the reproducing kernel particle method for solving 2D fractional Tricomitype equation with Robin boundary condition. Computers and Mathematics with Applications, 73(6):1270–1285, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211630637X.

Dehghan:2018:FDF

[DA18d] Mehdi Dehghan and Mostafa Abbaszadeh. A finite difference/finite element technique with error estimate for space fractional tempered diffusionwave equation. Computers and Mathematics with Applications, 75(8):2903–2914, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300361.

Dai:2014:EPP

[Dai14] Limei Dai. Exterior problems of parabolic Monge– Ampère equations for n = 2. Computers and Mathematics with Applications, 67(8):1497–1506, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000807.

Dehghan:2014:NSN

[DAM14] Mehdi Dehghan, Mostafa Abbaszadeh, and Akbar Mohebbi. The numerical solution of nonlinear high dimensional generalized Benjamin–Bona– Mahony–Burgers equation via the meshless method of radial basis functions. Computers and Mathematics with Applications, 68(3):212–237, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002314.

Danca:2012:CSP

[Dan12] Marius-F. Danca. Chaos suppression via periodic pulses in a class of piecewise continuous systems. Computers and Mathematics with Applications, 64(5):849–855, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011436.

Darwish:2011:PQF

[Dar11] Mohamed Abdalla Darwish. On a perturbed quadratic fractional integral equation of Abel type. Computers and Mathematics with Applications, 61(2):182–190, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008230.

Das:2012:SEM

[Das12] K. Das. Slip effects on MHD mixed convection stagnation point flow of a micropolar fluid towards a shrinking vertical sheet. Computers and Mathematics with Applications, 63(1):255–267, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009849.

Dassios:2015:SBS

[Das15] Ioannis K. Dassios. Stability of basic steady states of networks in bounded domains. Computers and Mathematics with Applications, 70 (9):2177–2196, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500379X.

deAndrade:2018:EVE

[dAS18] Bruno de Andrade and Giovana Siracusa. On evolutionary Volterra equations with state-dependent delay. Computers and Mathematics with Applications, 75(4):1181–1190, February 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117306971.

Davvaz:2010:FKH

[Dav10] B. Davvaz. Fuzzy Krasner (m, n)-hyperrings. Computers and Mathematics with Applications, 59(12):3879–3891, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002968.

David:2017:CBS

[Dav17] Claire David. Control of the Black–Scholes equation. Computers and Mathematics with Applications, 73(7):1566–1575, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300731.

Dinc:2010:FWT

[DB10] Erdal Dinç and Dumitru Baleanu. Fractional wavelet transform for the quantitative spectral resolution of the composite signals of the active compounds in a two-component mixture. Computers and Mathematics with Applications, 59(5):1701–1708, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005471.

Debbouche:2011:CFE

[DB11] Amar Debbouche and Dumitru Baleanu. Controllability of fractional evolution nonlocal impulsive quasilinear delay integro-differential systems. Computers and Mathematics with Applications, 62(3):1442–1450, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002380.

Doha:2012:EDS

[DB12] E. H. Doha and A. H.

Bhrawy. An efficient direct solver for multidimensional elliptic Robin boundary value problems using a Legendre spectral-Galerkin method. Computers and Mathematics with Applications, 64(4): 558–571, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111011060.

Delgado:2015:TDS

[DB15] Gabriel Delgado and Marc Bonnet. The topological derivative of stress-based cost functionals in anisotropic elasticity. Computers and Mathematics with Applications, 69(10):1144– 1166, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001054.

deBuhan:2017:NRE

[dBD17] M. de Buhan and M. Darbas. Numerical resolution of an electromagnetic inverse medium problem at fixed frequency. Computers and Mathematics with Applications, 74 (12):3111–3128, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304807.

Doha:2011:CSM

[DBEE11] E. H. Doha, A. H. Bhrawy, and S. S. Ezz-Eldien. A Chebyshev spectral method based on operational matrix for initial and boundary value problems of fractional order. Computers and Mathematics with Applications, 62(5):2364–2373, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122111005785.

Dostal:2014:RLC

[DBH^{+}14] Zdenek Dostál, Tomás Brzobohatý, David Horák, Tomás Kozubek, and Petr Vodstrcil. On r-linear convergence of semi-monotonic inexact augmented Lagrangians for bound and equality constrained quadratic programming problems with application. Computers and Mathematics with Applications, 67(3):515–526, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006652.

Dosiyev:2012:HAB

[DBS12] A. A. Dosiyev, S. C. Buranay, and D. Subasi. The highly accurate blockgrid method in solving Laplace’s equation for nonanalytic boundary condition with corner singularity. Computers and Mathematics with Applications, 64(4):616–632, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111011394.

Du:2010:CPE

[DC10] Juan Du and Minggen Cui. Constructive proof of existence for a class of fourth-order nonlinear BVPs. Computers and Mathematics with Applications, 59(2):903–911, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006920.

Demiriz:2012:SNP

[DÇ12] Serkan Demiriz and Celal Çakan. Some new paranormed difference sequence spaces and weighted core. Computers and Mathematics with Applications, 64 (6):1726–1739, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000715.

Dong:2015:HFM

[DC15] Qiaoli Dong and Liqun Cao. The hole-filling method and the multiscale computation for the wave equations in perforated domains. Computers and Mathematics with Applications, 70(8):1743–1756, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003442.

Deng:2012:NPH

[DCG^{+}12] Wu Deng, Rong Chen, Jian Gao, Yingjie Song, and Junjie Xu. A novel parallel hybrid intelligence optimization algorithm for a function approximation problem. Computers and Mathematics with Applications, 63(1):325–336, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211101008X.

Dong:2015:SPN

[DCKY15] Fangfang Dong, Yunmei Chen, De-Xing Kong, and Bailin Yang. Salt and pepper noise removal based on an approximation of l_{0} norm. Computers and Mathematics with Applications, 70(5):789–804, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002710.

Dangal:2017:PPS

[DCL17] Thir Dangal, C. S. Chen, and Ji Lin. Polynomial particular solutions for solving elliptic partial differential equations. Computers and Mathematics with Applications, 73(1):60–70, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306071.

daCruz:2012:SVC

[dCM12] Artur M. C. Brito da Cruz and Natália Martins. The q-symmetric variational calculus. Computers and Mathematics with Applications, 64(7):2241–2250, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001083.

Martins:2016:ISA

[dCMdSGTdC^{+}16] Thiago de Castro Martins, Marcos de Sales Guerra Tsuzuki, Erick Dario León Bueno de Camargo, Raul Gonzalez Lima, Fernando Silva de Moura, and Marcelo Brito Passos Amato. Interval simulated annealing applied to electrical impedance tomography image reconstruction with fast objective function evaluation. Computers and Mathematics with Applications, 72 (5):1230–1243, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303534.

Dong:2018:MCM

[DCN^{+}18] Hao Dong, Junzhi Cui, Yufeng Nie, Zihao Yang, and Zhiqiang Yang. Multiscale computational method for heat conduction problems of composite structures with diverse periodic configurations in different subdomains. Computers and Mathematics with Applications, 76(11– 12):2549–2565, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304875.

Duan:2013:ADM

[DCRL13] Jun-Sheng Duan, Temuer Chaolu, Randolph Rach, and Lei Lu. The Adomian decomposition method with convergence acceleration techniques for nonlinear fractional differential equations. Computers and Mathematics with Applications, 66(5):728–736, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000369.

Dirik:2010:FDM

[DD10] Fadime Dirik and Kamil Demirci. Four-dimensional matrix transformation and the rate of Astatistical convergence of continuous functions. Computers and Mathematics with Applications, 59(8):2976–2981, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001240.

Dufourd:2013:IEF

[DD13] Claire Dufourd and Yves Dumont. Impact of environmental factors on mosquito dispersal in the prospect of sterile insect technique control.

Computers and Mathematics with Applications, 66(9):1695–1715, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002307.

Dong:2016:CCS

[DD16] Qixiang Dong and Jiu Ding. Complete commuting solutions of the Yang–Baxter-like matrix equation for diagonalizable matrices. Computers and Mathematics with Applications, 72(1):194– 201, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302401.

DeSouza:2019:SCN

[DD19a] M. M. De Souza and A. L. De Bortoli. Stability and convergence of the numerical solution for the species equation of a model for PEMFCs. Computers and Mathematics with Applications, 77(4):1197–1215, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118306631.

Dimitrienko:2019:CSE

[DD19b] Yu. I. Dimitrienko and I. D. Dimitrienko. Computations of stresses and energy dissipation in composite thin laminates with the asymptotic vibration theory. Computers and Mathematics with Applications, 78(8):2541–2559, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301889.

DeLoof:2010:CLE

[DDD10] K. De Loof, B. De Baets, and H. De Meyer. Counting linear extension majority cycles in partially ordered sets on up to 13 elements. Computers and Mathematics with Applications, 59(4):1541–1547, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007883.

Durmaz:2011:ASN

[DDK11] Seher Durmaz, Sezgin Altay Demirbag, and Metin Orhan Kaya. Approximate solutions for nonlinear oscillation of a mass attached to a stretched elastic wire. Computers and Mathematics with Applications, 61(3):578–585, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009089.

Delitala:2013:MMI

[DDLM13] Marcello Delitala, Umberto Dianzani, Tommaso Lorenzi, and Matteo Melensi. A mathematical model for immune and autoimmune response mediated by Tcells. Computers and Mathematics with Applications, 66(6):1010–1023, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300401X.

Donatelli:2018:FBB

[DDM^{+}18] Marco Donatelli, Ali Dorostkar, Mariarosa Mazza, Maya Neytcheva, and Stefano Serra-Capizzano. Function-based block multigrid strategy for a twodimensional linear elasticitytype problem. Computers and Mathematics with Applications, 74 (5):1015–1028, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303292.

DalSanto:2019:MSR

[DDMQ19] N. Dal Santo, S. Deparis, A. Manzoni, and A. Quarteroni. Multi space reduced basis preconditioners for parametrized Stokes equations. Computers and Mathematics with Applications, 77(6): 1583–1604, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305480.

DAmbra:2013:PAP

[DdSF13] Pasqua D’Ambra, Daniela di Serafino, and Salvatore Filippone. Performance analysis of parallel Schwarz preconditioners in the LES of turbulent channel flows. Computers and Mathematics with Applications, 65(3):352–361, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004865.

DeChant:2010:SID

[De 10] Lawrence J. De Chant.

A semi-infinite domain eigenvalue problem describing turbulent velocity fluctuations. Computers and Mathematics with Applications, 60 (5):1177–1183, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004098.

Dehghan:2010:BUP

[DE10] Mehdi Dehghan and M. R. Eslahchi. Best uniform polynomial approximation of some rational functions. Computers and Mathematics with Applications, 59(1):382–390, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004283.

DaSilva:2011:STA

[DE11] Carmen Da Silva and René Escalante. Segmented Tau approximation for a forwardbackward functional differential equation. Computers and Mathematics with Applications, 62 (12):4582–4591, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009047.

Debnath:2012:LIC

[Deb12] Pradip Debnath. Lacunary ideal convergence in intuitionistic fuzzy normed linear spaces. Computers and Mathematics with Applications, 63(3):708–715, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010145.

Defterli:2010:CNS

[Def10a] Ozlem Defterli. Corrigendum to “A numerical scheme for two dimensional optimal control problems with memory effect” [Comput. Math. Appl. 59 (2010) 1630–1636]. Computers and Mathematics with Applications, 59(8), April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001604.

See [Def10b].

Defterli:2010:NST

[Def10b] Ozlem Defterli. A numerical scheme for twodimensional optimal control problems with memory effect. Computers and Mathematics with Applications, 59(5):1630–1636, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005392.

DAcunto:2011:DMS

[DEFP11] B. D’Acunto, G. Esposito, L. Frunzo, and F. Pirozzi. Dynamic modeling of sulfate reducing biofilms. Computers and Mathematics with Applications, 62(6):2601–2608, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006390.

Dejdumrong:2011:NBB

[Dej11] Natasha Dejdumrong. A new bivariate basis representation for Bézierbased triangular patches with quadratic complexity. Computers and Mathematics with Applications, 61(8):2292–2295, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007480.

Dellar:2013:IDL

[Del13] Paul J. Dellar. An interpretation and derivation of the lattice Boltzmann method using Strang splitting. Computers and Mathematics with Applications, 65(2):129–141, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007206.

Derksen:2013:MME

[DELK13] Jos Derksen, Dmitry Eskin, Li-Shi Luo, and Manfred Krafczyk. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 65(2):127–128, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112007067.

Demiray:2010:HOP

[Dem10a] Hilmi Demiray. Higher order perturbation expansion of waves in water of variable depth. Computers and Mathematics with Applications, 59(1):298–304, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004994.

Demiray:2010:WNW

[Dem10b] Hilmi Demiray. Weakly nonlinear waves in water of variable depth: Variable-coefficient Korteweg– de Vries equation. Computers and Mathematics with Applications, 60 (6):1747–1755, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004803.

Daniel:2018:AHR

[DESV18] Patrik Daniel, Alexandre Ern, Iain Smears, and Martin Vohralík. An adaptive hp-refinement strategy with computable guaranteed bound on the error reduction factor. Computers and Mathematics with Applications, 76(5):967–983, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303109.

Deutsch:2010:CFD

[Deu10] Jesse Ira Deutsch. Conjectures on the fundamental domain of the Hilbert modular group. Computers and Mathematics with Applications, 59(2):700–705, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007226.

Demkowicz:2018:PIB

[DFG^{+}18] Leszek Demkowicz, Rick Falk, Benqi Guo, Michael Vogelius, and Zhimin Zhang. Prof. Ivo Babuška. Computers and Mathematics with Applications, 74(9):2045–2046, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305503.

Devloo:2019:RCD

[DFG19] Philippe R. B. Devloo, Agnaldo M. Farias, and Sônia M. Gomes. A remark concerning divergence accuracy order for H (div)-conforming finite element flux approximations. Computers and Mathematics with Applications, 77(7):1864–1872, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306722.

Devloo:2013:ACC

[DFGG13] Philippe R. B. Devloo, Agnaldo M. Farias, Sônia M. Gomes, and João L. Gonçalves. Application of a combined continuous-discontinuous Galerkin finite element method for the solution of the girkmann problem. Computers and Mathematics with Applications, 65(11):1786– 1794, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001673.

Dupuy:2010:UCH

[DFJS10] Pablo M. Dupuy, Maria Fernandino, Hugo A. Jakobsen, and Hallvard F. Svendsen. Using Cahn–Hilliard mobility to simulate coalescence dynamics. Computers and Mathematics with Applications, 59(7):2246–2259, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006373.

Diehl:2015:FRS

[DFM15] S. Diehl, S. Farås, and G. Mauritsson. Fast reliable simulations of secondary settling tanks in wastewater treatment with semi-implicit time discretization. Computers and Mathematics with Applications, 70(4):459– 477, August 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002163.

Du:2013:PRO

[DFP^{+}13] J. Du, F. Fang, C. C. Pain, I. M. Navon, J. Zhu, and D. A. Ham. POD reduced-order unstructured mesh modeling applied to 2D and 3D fluid flow. Computers and Mathematics with Applications, 65(3):362–379, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004397.

Di:2011:LMD

[DFS11] Xiaofeng Di, Yushun Fan, and Yimin Shen. Local martingale difference approach for service selection with dynamic QoS. Computers and Mathematics with Applications, 61(9):2638– 2646, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001623.

Du:2014:BCG

[DFS14] Lei Du, Yasunori Futamura, and Tetsuya Sakurai. Block conjugate gradient type methods for the approximation of bilinear form C^{H}A^{−1}B. Computers and Mathematics with Applications, 66(12):2446–2455, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005907.

Dong:2018:CAN

[DFW^{+}18] Lixiu Dong, Wenqiang Feng, Cheng Wang, Steven M. Wise, and Zhengru Zhang. Convergence analysis and numerical implementation of a second order numerical scheme for the three-dimensional phase field crystal equation. Computers and Mathematics with Applications, 75(6):1912–1928, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304297.

Das:2010:SFR

[DG10a] Pratulananda Das and Sanjoy Kr. Ghosal. Some further results on I-Cauchy sequences and condition (AP). Computers and Mathematics with Applications, 59(8):2597– 2600, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000490.

Ding:2010:BGE

[DG10b] Juntang Ding and Bao-Zhu Guo. Blow-up and global existence for nonlinear parabolic equations with Neumann boundary conditions. Computers and Mathematics with Applications, 60(3):670–679, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003640.

DElia:2013:FLO

[DG13a] Marta D’Elia and Max Gunzburger. The fractional Laplacian operator on bounded domains as a special case of the nonlocal diffusion operator. Computers and Mathematics with Applications, 66(7):1245–1260, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004707.

Demkowicz:2013:PDM

[DG13b] L. Demkowicz and J. Gopalakrishnan. A primal DPG method without a first-order reformulation. Computers and Mathematics with Applications, 66(6):1058–1064, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122113004343.

Das:2018:BET

[DGA18] Amiya Das, Niladri Ghosh, and Khusboo Ansari. Bifurcation and exact traveling wave solutions for dual power Zakharov– Kuznetsov–Burgers equation with fractional temporal evolution. Computers and Mathematics with Applications, 75(1):59– 69, January 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305357.

Daftardar-Gejji:2010:CFO

[DGB10a] Varsha Daftardar-Gejji and Sachin Bhalekar. Chaos in fractional ordered Liu system. Computers and Mathematics with Applications, 59(3):1117–1127, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812210900426X.

Daftardar-Gejji:2010:SFB

[DGB10b] Varsha Daftardar-Gejji and Sachin Bhalekar. Solving fractional boundary value problems with Dirichlet boundary conditions using a new iterative method. Computers and Mathematics with Applications, 59(5):1801–1809, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005586.

Dominguez-Garcia:2012:IVC

[DGGBTRJF12] José Luis Domínguez-García, Oriol Gomis-Bellmunt, Lluís Trilla-Romero, and Adrià Junyent-Ferré. Indirect vector control of a squirrel cage induction generator wind turbine. Computers and Mathematics with Applications, 64(2):102–114, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000314.

Drake:2017:RTS

[DGH17] Dow Drake, Jay Gopalakrishnan, and Ammar Harb. Reduced test spaces for DPG methods using rectangular elements. Computers and Mathematics with Applications, 74 (8):1955–1963, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304042.

Dogan:2010:HOS

[DGK10] Abdulkadir Dogan, John R. Graef, and Lingju Kong. Higher order semipositone multi-point boundary value problems on time scales. Computers and Mathematics with Applications, 60(1):23– 35, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000297X.

Darbas:2012:EBC

[DGL12] M. Darbas, O. Goubet, and S. Lohrengel. Exact boundary controllability of the second-order Maxwell system: Theory and numerical simulation. Computers and Mathematics with Applications, 63(7):1212–1237, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011023.

Dieci:2018:P

[DGL18] Luca Dieci, Nicola Guglielmi, and Luciano Lopez. Preface. Computers and Mathematics with Applications, 74(5):881, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304285.

Dick:2019:HOQ

[DGLS19] Josef Dick, Robert N.

Gantner, Quoc T. Le Gia, and Christoph Schwab. Higher order quasi-Monte Carlo integration for Bayesian PDE inversion. Computers and Mathematics with Applications, 77(1):144–172, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305315.

Dong:2018:MSK

[DGLU18] Zhaonan Dong, Emmanuil H. Georgoulis, Jeremy Levesley, and Fuat Usta. A multilevel sparse kernel-based stochastic collocation finite element method for elliptic problems with random coefficients.

Computers and Mathematics with Applications, 76(8):1950–1965, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304127.

Donat:2014:IEM

[DGM14] R. Donat, F. Guerrero, and P. Mulet. Implicitexplicit methods for models for vertical equilibrium multiphase flow. Computers and Mathematics with Applications, 68(3):363–383, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400265X.

Dimov:2013:AAM

[DGOZ13] I. Dimov, R. Georgieva, Tz. Ostromsky, and Z. Zlatev. Advanced algorithms for multidimensional sensitivity studies of large-scale air pollution models based on Sobol sequences. Computers and Mathematics with Applications, 65(3):338–351, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004725.

Darrigrand:2018:CKS

[DGR18] Eric Darrigrand, Nabil Gmati, and Rania Rais. Convergence of Krylov subspace solvers with Schwarz preconditioner for the exterior Maxwell problem. Computers and Mathematics with Applications, 74(11):2691– 2709, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117305199.

Diele:2018:NAF

[DGT18] Fasma Diele, Marcus Garvie, and Catalin Trenchea. Numerical analysis of a first-order in time implicit-symplectic scheme for predatorprey systems. Computers and Mathematics with Applications, 74(5):948–961, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302614.

Danca:2013:SSD

[DGTC13] Marius-F. Danca, Roberto Garrappa, Wallace K. S. Tang, and Guanrong Chen. Sustaining stable dynamics of a fractionalorder chaotic financial system by parameter switching. Computers and Mathematics with Applications, 66(5):702–716, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000588.

DeLeenheer:2013:NEN

[DGZ13] Patrick De Leenheer, Jay Gopalakrishnan, and Erica Zuhr. Nonnegativity of exact and numerical solutions of some chemotactic models. Computers and Mathematics with Applications, 66(3):356–375, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003167.

Dang:2010:CGU

[DH10a] Jr-Fong Dang and I-Hsuan Hong. The Cournot game under a fuzzy decision environment. Computers and Mathematics with Applications, 59(9):3099–3109, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001501.

Dehghan:2010:MES

[DH10b] Mehdi Dehghan and Masoud Hajarian. Matrix equations over (R, S)- symmetric and (R, S)- skew symmetric matrices. Computers and Mathematics with Applications, 59(11):3583– 3594, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002415.

Dye:2010:PSO

[DH10c] Chung-Yuan Dye and Tsu-Pang Hsieh. A particle swarm optimization for solving joint pricing and lot-sizing problem with fluctuating demand and unit purchasing cost. Computers and Mathematics with Applications, 60(7):1895–1907, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004992.

Darvishi:2011:MSS

[DH11a] M. T. Darvishi and P. Hessari. A modified symmetric successive overrelaxation method for augmented systems. Computers and Mathematics with Applications, 61(10):3128–3135, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002835.

Dehghan:2011:IAB

[DH11b] Mehdi Dehghan and Hossein Hosseinzadeh. Improvement of the accuracy in boundary element method based on high-order discretization. Computers and Mathematics with Applications, 62(12):4461–4471, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100887X.

Dou:2016:FKB

[DH16] F. F. Dou and Y. C. Hon. Fundamental kernel-based method for backward space-time fractional diffusion problem. Computers and Mathematics with Applications, 71(1):356–367, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005660.

DeStaelen:2017:NPD

[DH17] R. H. De Staelen and A. S. Hendy. Numerically pricing double barrier options in a time-fractional Black– Scholes model. Computers and Mathematics with Applications, 74 (6):1166–1175, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730353X.

Dong:2018:OCA

[DH18] Xiaojing Dong and Yinnian He. Optimal convergence analysis of Crank–Nicolson extrapolation scheme for the three-dimensional incompressible magnetohydrodynamics. Computers and Mathematics with Applications, 76(11– 12):2678–2700, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304863.

Duan:2017:PAR

[DHGF17] Lian Duan, Lihong Huang, Zhenyuan Guo, and Xianwen Fang. Periodic attractor for reaction– diffusion high-order Hopfield neural networks with time-varying delays. Computers and Mathematics with Applications, 73(2):233–245, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306277.

Dutt:2016:SEM

[DHMU16] P. Dutt, A. Husain, A. S. Vasudeva Murthy, and C. S. Upadhyay. h– p spectral element methods for three dimensional elliptic problems on nonsmooth domains, Part III: Error estimates, preconditioners, computational techniques and numerical results. Computers and Mathematics with Applications, 71(9):1745– 1771, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122116000134.

Dai:2011:SSB

[DHQ11] Min Dai, Su-Rui Hu, and Jia Qi. Spike sorting based on radial basis function network with overlap decomposition. Computers and Mathematics with Applications, 62(7):2736–2742, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005529.

Doha:2019:SJS

[DHY19] E. H. Doha, R. M.

Hafez, and Y. H. Youssri. Shifted Jacobi spectral-Galerkin method for solving hyperbolic partial differential equations. Computers and Mathematics with Applications, 78(3):889–904, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301300.

Diao:2017:CNL

[Dia17] Huai-An Diao. On condition numbers for least squares with quadric inequality constraint. Computers and Mathematics with Applications, 73(4):616–627, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300123.

DAmbrosio:2012:PMM

[DIJ12] Raffaele D’Ambrosio, Giuseppe Izzo, and Zdzislaw Jackiewicz. Perturbed MEBDF methods. Computers and Mathematics with Applications, 63(4):851–861, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010339.

Ding:2010:TBS

[Din10] Feng Ding. Transformations between some special matrices. Computers and Mathematics with Applications, 59(8):2676– 2695, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000581.

Ding:2013:GBS

[Din13] Juntang Ding. Global and blow-up solutions for nonlinear parabolic equations with Robin boundary conditions. Computers and Mathematics with Applications, 65(11): 1808–1822, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211300165X.

Du:2019:SBR

[DIS19] Lei Du, Akira Imakura, and Tetsuya Sakurai. Simultaneous band reduction of two symmetric matrices. Computers and Mathematics with Applications, 77(8):2207–2220, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306953.

Dzikowski:2018:DAL

[DJD18] Michal Dzikowski, Lukasz Jasinski, and Marcin Dabrowski. Depthaveraged Lattice Boltzmann and Finite Element methods for singlephase flows in fractures with obstacles. Computers and Mathematics with Applications, 75(10):3453–3470, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830083X.

Duo:2018:FAS

[DJZ18] Siwei Duo, Lili Ju, and Yanzhi Zhang. A fast algorithm for solving the space–time fractional diffusion equation. Computers and Mathematics with Applications, 75(6):1929–1941, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302298.

Dogru:2012:SAP

[DK12] Ogün Dogru and Kadir Kanat. Statistical approximation properties of King-type modification of Lupas operators. Computers and Mathematics with Applications, 64(4):511–517, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010881.

Dimitrova:2014:MBO

[DK14] Neli Dimitrova and Mikhail Krastanov. Model-based optimization of biogas production in an anaerobic biodegradation process. Computers and Mathematics with Applications, 68(9):986–993, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400159X.

Doss:2018:FPM

[DK18] L. Jones Tarcius Doss and N. Kousalya. Finite pointset method for biharmonic equations. Computers and Mathematics with Applications, 75(10):3756–3785, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830110X.

Dong:2014:MMN

[DKG14] Xiaoyang Dong, Hristo V.

Kojouharov, and James P. Grover. Mathematical models of nutrient recycling and toxin production in a gradostat. Computers and Mathematics with Applications, 68(9):972–985, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000972.

Dudek:2017:SFG

[DKM17] Sylwia Dudek, Piotr Kalita, and Stanislaw Migórski. Steady flow of generalized Newtonian fluid with multivalued rheology and nonmonotone friction law. Computers and Mathematics with Applications, 74(8):1813–1825, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303966.

Duster:2015:H

[DKRY15] Alexander Düster, Stefan Kollmannsberger, Alessandro Reali, and Zohar Yosibash. HOFEIM 2014. Computers and Mathematics with Applications, 70(7):1369, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003594.

Dang:2010:IMS

[DL10] Quang A. Dang and Vu Thai Luan. Iterative method for solving a nonlinear fourth order boundary value problem. Computers and Mathematics with Applications, 60(1):112–121, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003081.

Dubois:2011:QPA

[DL11] François Dubois and Pierre Lallemand. Quartic parameters for acoustic applications of lattice Boltzmann scheme. Computers and Mathematics with Applications, 61(12):3404–3416, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000162.

Ding:2014:IMS

[DL14] Ling Ding and Lin Li.

Infinitely many standing wave solutions for the nonlinear Klein–Gordon– Maxwell system with sign-changing potential. Computers and Mathematics with Applications, 68(5):589–595, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002843.

DeRossi:2016:CML

[DL16] Alessandro De Rossi and Emmanuel Lévêque. Central-moment lattice Boltzmann schemes with fixed and moving immersed boundaries. Computers and Mathematics with Applications, 72 (6):1616–1628, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304278.

Du:2018:TWB

[DL18] Yini Du and Bingchen Liu. Time-weighted blowup profiles in a nonlinear parabolic system with Fujita exponent. Computers and Mathematics with Applications, 76 (5):1034–1055, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303158.

Dastour:2019:FOO

[DL19] Hatef Dastour and Wenyuan Liao. A fourth-order optimal finite difference scheme for the Helmholtz equation with PML. Computers and Mathematics with Applications, 78(6):2147–2165, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302676.

Dou:2019:MPS

[DLC19] Fangfang Dou, Yanshan Liu, and C. S. Chen. The method of particular solutions for solving nonlinear Poisson problems. Computers and Mathematics with Applications, 77(2):501–513, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830573X.

Ding:2010:ISM

[DLD10] Jie Ding, Yanjun Liu, and Feng Ding. Iterative solutions to matrix equations of the form A_{i}XB_{i} = F_{i}. Computers and Mathematics with Applications, 59(11):3500–3507, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002300.

Dupuy:2011:MHP

[DLF^{+}11] Pablo M. Dupuy, Yi Lin, Maria Fernandino, Hugo A. Jakobsen, and Hallvard F. Svendsen. Modelling of high pressure binary droplet collisions. Computers and Mathematics with Applications, 61(12):3564–3576, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000413X.

delaPiedra:2013:SEL

[dlPBTW13] Antonio de la Piedra, An Braeken, Abdellah Touhafi, and Karel Wouters. Secure event logging in sensor networks. Computers and Mathematics with Applications, 65(5):762–773, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004646.

Duan:2016:TOI

[DLQ16] Xianbao Duan, Feifei Li, and Xinqiang Qin.

Topology optimization of incompressible Navier– Stokes problem by level set based adaptive mesh method. Computers and Mathematics with Applications, 72(4):1131–1141, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303662.

Dominguez:2014:NFC

[DLS14] Víctor Domínguez, Sijiang L. Lu, and Francisco-Javier Sayas. A Nyström flavored Calderón calculus of order three for two dimensional waves, time-harmonic and transient. Computers and Mathematics with Applications, 67(1):217–236, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006615.

Dai:2018:PBP

[DLS18a] Qiuyi Dai, Enhao Lan, and Feilin Shi. A priori bounds for positive solutions of Kirchhoff type equations. Computers and Mathematics with Applications, 76 (6):1525–1534, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303699.

Duan:2018:EER

[DLS18b] Yu Duan, Jia-Feng Liao, and Xin Sun. An exact estimate result for singular p-Laplacian equations with sign-changing potential. Computers and Mathematics with Applications, 76 (8):2022–2034, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304176.

Dubois:2010:SLB

[DLT10] François Dubois, Pierre Lallemand, and Mahdi Tekitek. On a superconvergent lattice Boltzmann boundary scheme. Computers and Mathematics with Applications, 59(7):2141–2149, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006439.

Distefano:2012:IDR

[DLT12] Salvatore Distefano, Francesco Longo, and Kishor S. Trivedi. Investigating dynamic reliability and availability through statespace models. Computers and Mathematics with Applications, 64 (12):3701–3716, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001630.

Dong:2012:FIC

[DLWW12] Su-Jun Dong, Yun-Ze Li, Jin Wang, and Jun Wang. Fuzzy incremental control algorithm of loop heat pipe cooling system for spacecraft applications. Computers and Mathematics with Applications, 64(5):877–886, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000405.

Dai:2017:ESO

[DLZ17] Ruxin Dai, Pengpeng Lin, and Jun Zhang. An efficient sixth-order solution for anisotropic Poisson equation with completed Richardson extrapolation and multiscale multigrid method. Computers and Mathematics with Applications, 73(8):1865–1877, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300998.

Dai:2019:EAM

[DLZ19] Ruxin Dai, Pengpeng Lin, and Jun Zhang. An EX-CMG accelerated multiscale multigrid computation for 3D Poisson equation. Computers and Mathematics with Applications, 77(8):2051–2060, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307223.

Deu:2010:SFD

[DM10] J.-F. Deü and D. Matignon. Simulation of fractionally damped mechanical systems by means of a newmark-diffusive scheme. Computers and Mathematics with Applications, 59(5):1745–1753, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122109005513.

Dalfard:2012:TMH

[DM12a] Vahid Majazi Dalfard and Ghorbanali Mohammadi. Two meta-heuristic algorithms for solving multiobjective flexible job-shop scheduling with parallel machine and maintenance constraints. Computers and Mathematics with Applications, 64 (6):2111–2117, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003471.

DiCostanzo:2012:AAS

[DM12b] E. Di Costanzo and A. Marasco. Approximate analytic solution of the Dirichlet problems for Laplace’s equation in planar domains by a perturbation method. Computers and Mathematics with Applications, 63(1):60–67, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009485.

Dehghan:2015:MVS

[DM15] Mehdi Dehghan and Vahid Mohammadi. The method of variably scaled radial kernels for solving two-dimensional magnetohydrodynamic (MHD) equations using two discretizations: the Crank– Nicolson scheme and the method of lines (MOL). Computers and Mathematics with Applications, 70(10):2292–2315, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004113.

Dehghan:2016:TNM

[DM16] Mehdi Dehghan and Vahid Mohammadi. Two numerical meshless techniques based on radial basis functions (RBFs) and the method of generalized moving least squares (GMLS) for simulation of coupled Klein–Gordon– Schrödinger (KGS) equations. Computers and Mathematics with Applications, 71(4):892–921, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500601X.

Dassi:2018:EHO

[DM18] F. Dassi and L. Mascotto. Exploring highorder three dimensional virtual elements: Bases and stabilizations. Computers and Mathematics with Applications, 75(9):3379–3401, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300786.

Dessole:2019:FII

[DM19a] M. Dessole and F. Marcuzzi. Fully iterative ILU preconditioning of the unsteady Navier–Stokes equations for GPGPU. Computers and Mathematics with Applications, 77(4):907–927, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306345.

Dong:2019:SIB

[DM19b] Fangfang Dong and Qianting Ma. Single image blind deblurring based on the fractional-order differential. Computers and Mathematics with Applications, 78 (6):1960–1977, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301580.

Devi:2010:GQF

[DMD10] J. Vasundhara Devi, F. A. McRae, and Z. Drici. Generalized quasilinearization for fractional differential equations. Computers and Mathematics with Applications, 59(3):1057–1062, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003514.

DAmbrosio:2018:ANM

[DMP18] Raffaele D’Ambrosio, Martina Moccaldi, and Beatrice Paternoster. Adapted numerical methods for advection-reaction–diffusion problems generating periodic wavefronts. Computers and Mathematics with Applications, 74 (5):1029–1042, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122117302547.

Diaz-Madronero:2010:VSP

[DMPV10] Manuel Díaz-Madroñero, David Peidro, and Pandian Vasant. Vendor selection problem by using an interactive fuzzy multi-objective approach with modified S-curve membership functions. Computers and Mathematics with Applications, 60(4):1038–1048, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002506.

Dolejsi:2018:AHM

[DMRS18] Vít Dolejsí, Georg May, Filip Roskovec, and Pavel Solin. Anisotropic hpmesh optimization technique based on the continuous mesh and error models. Computers and Mathematics with Applications, 74(1):45–63, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306861.

Dorville:2011:OCN

[DMV11] René Dorville, Gisèle M.

Mophou, and Vincent S. Valmorin. Optimal control of a nonhomogeneous Dirichlet boundary fractional diffusion equation. Computers and Mathematics with Applications, 62(3):1472–1481, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001775.

Ding:2010:CSS

[DMZ10] Xiaohua Ding, Qiang Ma, and Lei Zhang. Convergence and stability of the split-step θ-method for stochastic differential equations. Computers and Mathematics with Applications, 60 (5):1310–1321, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004256.

Darwish:2010:IBV

[DN10] Mohamed Abdalla Darwish and S. K. Ntouyas. On initial and boundary value problems for fractional order mixed type functional differential inclusions. Computers and Mathematics with Applications, 59(3):1253–1265, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003563.

Dang:2018:ERI

[DN18a] Quang A. Dang and Thanh Huong Nguyen. Existence results and iterative method for solving a nonlinear biharmonic equation of Kirchhoff type. Computers and Mathematics with Applications, 76(1):11–22, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301925.

Das:2018:HOC

[DN18b] Abhishek Das and Srinivasan Natesan. Higherorder convergence with fractional-step method for singularly perturbed 2D parabolic convection– diffusion problems on Shishkin mesh. Computers and Mathematics with Applications, 75(7):2387–2403, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307769.

Delis:2015:MTF

[DNP15] A. I. Delis, I. K. Nikolos, and M. Papageorgiou. Macroscopic traffic flow modeling with adaptive cruise control: Development and numerical solution. Computers and Mathematics with Applications, 70(8):1921–1947, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003703.

Dunca:2013:MNS

[DNR13] Argus A. Dunca, Monika Neda, and Leo G. Rebholz. A mathematical and numerical study of a filtering-based multiscale fluid model with nonlinear eddy viscosity. Computers and Mathematics with Applications, 66(6):917–933, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300388X.

Demyanko:2015:ISB

[DNS15] Kirill V. Demyanko, Yuri M. Nechepurenko, and Miloud Sadkane. Inverse subspace biiteration and bi-Newton methods for computing spectral projectors. Computers and Mathematics with Applications, 69(7): 592–600, April 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115000486.

Dang:2016:NSP

[DNS16] Duy-Minh Dang, Duy Nguyen, and Granville Sewell. Numerical schemes for pricing Asian options under state-dependent regime-switching jumpdiffusion models. Computers and Mathematics with Applications, 71(1): 443–458, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115005854.

Dominguez:2018:CFE

[DNS18] Sebastian Dominguez, Nilima Nigam, and Bobak Shahriari. A combined finite element and Bayesian optimization framework for shape optimization in spectral geometry. Computers and Mathematics with Applications, 74 (11):2874–2896, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305461.

Du:2013:ROM

[DNZ^{+}13] Juan Du, I. M. Navon, Jiang Zhu, Fangxin Fang, and A. K. Alekseev. Reduced order modeling based on POD of a parabolized Navier– Stokes equations model II: Trust region POD 4D VAR data assimilation. Computers and Mathematics with Applications, 65(3):380–394, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004312.

Davydov:2011:OSP

[DO11] Oleg Davydov and Dang Thi Oanh. On the optimal shape parameter for Gaussian radial basis function finite difference approximation of the Poisson equation. Computers and Mathematics with Applications, 62 (5):2143–2161, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005177.

Dolati:2011:NIS

[Dol11] Ardeshir Dolati. On the number of independent sets in cycle-separated tricyclic graphs. Computers and Mathematics with Applications, 61(6):1542– 1546, March 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111000381.

Dong:2010:OBS

[Don10a] Jiu-Gang Dong. Oscillation behavior of second order nonlinear neutral differential equations with deviating arguments. Computers and Mathematics with Applications, 59(12):3710– 3717, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002646.

Dong:2010:NSL

[Don10b]  Yunda Dong.  New

step lengths in conjugate gradient methods.

Computers and Mathematics with Applications, 60(3):563–571, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003433.

Dosly:2010:RBF

[Dos10]  Ondrej Doslý.  Rofe–

Beketov’s formula on time scales. Computers and Mathematics with Applications, 60(8):2382–2386, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005973.

Dosly:2012:OCC

[Dos12] Ondrej Doslý. Oscillation and conjugacy criteria for two-dimensional symplectic difference systems. Computers and Mathematics with Applications, 64(7):2202–2208, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000247.

Dosiyev:2018:MHE

[Dos18] A. A. Dosiyev. A method of harmonic extension for computing the generalized stress intensity factors for Laplace’s equation with singularities. Computers and Mathematics with Applications, 75(5):1767–1777, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307605.

Dostal:2015:MQF

[DP15] Zdenek Dostál and Lukás Pospísil. Minimizing quadratic functions with semidefinite Hessian subject to bound constraints. Computers and Mathematics with Applications, 70(8):2014–2028, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003843.

DElia:2016:CSN

[DPBL16] Marta D’Elia, Mauro Perego, Pavel Bochev, and David Littlewood. A coupling strategy for nonlocal and local diffusion models with mixed volume constraints and boundary conditions. Computers and Mathematics with Applications, 71(11): 2218–2230, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500574X.

dePrenter:2018:NSP

[dPLM18] Frits de Prenter, Christoph Lehrenfeld, and André Massing. A note on the stability parameter in Nitsche’s method for unfitted boundary value problems. Computers and Mathematics with Applications, 75(12):4322–4336, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301688.

Darrigrand:2015:GOA

[DPM15] Vincent Darrigrand, David Pardo, and Ignacio Muga. Goal-oriented adaptivity using unconventional error representations for the 1D Helmholtz equation. Computers and Mathematics with Applications, 69(9):964–979, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122115001017.

dePereda:2013:GCM

[dPRVRB13] Diego de Pereda, Sergio Romero-Vivo, Beatriz Ricarte, and Jorge Bondia. Guaranteed computation methods for compartmental in-series models under uncertainty. Computers and Mathematics with Applications, 66 (9):1595–1605, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001600.

DiPaola:2013:FDE

[DPZ13] Mario Di Paola, Francesco Paolo Pinnola, and Massimiliano Zingales. Fractional differential equations and related exact mechanical models. Computers and Mathematics with Applications, 66(5):608–620, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001648.

Deng:2010:NUR

[DQ10] Jiqin Deng and Hailiang Qu. New uniqueness results of solutions for fractional differential equations with infinite delay. Computers and Mathematics with Applications, 60(8):2253–2259, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005791.

Duan:2012:HON

[DR12] Jun-Sheng Duan and Randolph Rach. Higherorder numeric Wazwaz– El-Sayed modified Adomian decomposition algorithms. Computers and Mathematics with Applications, 63(11):1557– 1568, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002556.

Dragomir:2010:WGM

[Dra10] S. S. Dragomir. Weighted f-Gini mean difference for convex and symmetric functions in linear spaces. Computers and Mathematics with Applications, 60(3):734–743, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211000369X.

Dragomir:2011:ATD

[Dra11a] S. S. Dragomir. Approximating n-time differentiable functions of selfadjoint operators in Hilbert spaces by two point Taylor type expansion. Computers and Mathematics with Applications, 61(10):2958– 2970, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002410.

Dragomir:2011:OTI

[Dra11b] S. S. Dragomir. Ostrowski’s type inequalities for some classes of continuous functions of selfadjoint operators in Hilbert spaces. Computers and Mathematics with Applications, 62 (12):4439–4448, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008844.

Devi:2012:VLM

[DRD12] J. Vasundhara Devi, F. A. Mc Rae, and Z. Drici. Variational Lyapunov method for fractional differential equations. Computers and Mathematics with Applications, 64(10):2982– 2989, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112000910.

Das:2010:IPP

[DRK10] Debasis Das, Arindam Roy, and Samarjit Kar. Improving production policy for a deteriorating item under permissible delay in payments with stock-dependent demand rate. Computers and Mathematics with Applications, 60(7):1973–1985, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005237.

Das:2011:VFE

[DRK11] Debasis Das, Arindam Roy, and Samarjit Kar. A volume flexible economic production lotsizing problem with imperfect quality and random machine failure in fuzzy-stochastic environment. Computers and Mathematics with Applications, 61(9):2388–2400, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001015.

Dziok:2011:CRC

[DRS11] Jacek Dziok, Ravinder Krishna Raina, and Janusz Sokól. Certain results for a class of convex functions related to a shell-like curve connected with Fibonacci numbers. Computers and Mathematics with Applications, 61(9):2605–2613, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100157X.

Dahmen:2015:AWB

[DRT^{+}15] T. Dahmen, M. Roland, T. Tjardes, B. Bouillon, P. Slusallek, and S. Diebels. An automated workflow for the biomechanical simulation of a tibia with implant using computed tomography and the finite element method. Computers and Mathematics with Applications, 70(5):903–916, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002953.

Drazic:2010:SEO

[DRZ10] Slobodan Drazić, Nebojsa Ralević, and Jovisa Zunić. Shape elongation from optimal encasing rectangles. Computers and Mathematics with Applications, 60(7):2035–2042, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005353.

daSilva:2015:IPA

[dS15] João Lita da Silva. Integer powers of antitridiagonal matrices of the form antitridiag_{n}(a, 0, b), a, b ∈ R. Computers and Mathematics with Applications, 69(11):1313– 1328, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001212.

daSilva:2016:APP

[dS16] João Lita da Silva. On anti-pentadiagonal persymmetric Hankel matrices with perturbed corners. Computers and Mathematics with Applications, 72(3):415–426, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302048.

Das:2018:DSD

[DS18a] Amiya Das and Asit Saha. Dynamical survey of the dual power Zakharov–Kuznetsov–Burgers equation with external periodic perturbation. Computers and Mathematics with Applications, 76(5):1174–1183, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303304.

Ding:2018:BAQ

[DS18b] Juntang Ding and Xuhui Shen. Blow-up analysis in quasilinear reaction– diffusion problems with weighted nonlocal source. Computers and Mathematics with Applications, 75(4):1288–1301, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307095.

Dudek:2009:FIH

[DSA09] W. A. Dudek, M. Shabir, and M. Irfan Ali. (α, β)- fuzzy ideals of hemirings. Computers and Mathematics with Applications, 58(2):310–321, July 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109002636.

See note [Jun10].

Dudek:2010:CHT

[DSA10] Wieslaw A. Dudek, Muhammad Shabir, and Rukhshanda Anjum. Characterizations of hemirings by their h-ideals. Computers and Mathematics with Applications, 59(9):3167– 3179, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000163X.

dosSantos:2011:ERF

[dSAC11] José Paulo Carvalho dos Santos, M. Mallika Arjunan, and Claudio Cuevas. Existence results for fractional neutral integro-differential equations with statedependent delay. Computers and Mathematics with Applications, 62(3):1275–1283, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002112.

Dokken:2019:TSR

[DSB19] Tor Dokken, Vibeke Skytt, and Oliver Barrowclough. Trivariate spline representations for computer aided design and additive manufacturing. Computers and Mathematics with Applications, 78(7):2168–2182, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304358.

Coelho:2012:FAA

[dSCM12] Leandro dos Santos Coelho and Viviana Cocco Mariani. Firefly algorithm approach based on chaotic Tinkerbell map applied to multivariable PID controller tuning. Computers and Mathematics with Applications, 64(8):2371–2382, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003975.

Delbosc:2014:OIL

[DSK^{+}14] N. Delbosc, J. L. Summers, A. I. Khan, N. Kapur, and C. J. Noakes. Optimized implementation of the lattice Boltzmann method on a graphics processing unit towards real-time fluid simulation. Computers and Mathematics with Applications, 67(2):462–475, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006068.

Ding:2011:NSV

[DSL11] Lixia Ding, Weiping Shi, and Hongwen Luo. Numerical simulation of viscous flow over nonsmooth surfaces. Computers and Mathematics with Applications, 61(12): 3703–3710, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002841.

Duan:2018:MPS

[DSL18] Yu Duan, Xin Sun, and Jia-Feng Liao. Multiplicity of positive solutions for a class of critical Sobolev exponent problems involving Kirchhofftype nonlocal term. Computers and Mathematics with Applications, 75(12):4427–4437, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301779.

Djoko:2018:IMS

[DSM18] J. K. Djoko, V. Konlack Socgnia, and M. Mbehou. Iterative methods for Stokes flow under nonlinear slip boundary condition coupled with the heat equation. Computers and Mathematics with Applications, 76(11– 12):2613–2634, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304838.

Diblik:2010:APS

[DSR10] Josef Diblík, Ewa Schmeidel, and Miroslava Ruzicková. Asymptotically periodic solutions of Volterra system of difference equations. Computers and Mathematics with Applications, 59(8):2854– 2867, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000957.

deSouza:2017:NSQ

[dSSV17] Manassés de Souza, Uberlandio B. Severo, and Gilberto F. Vieira. On a nonhomogeneous and singular quasilinear equation involving critical growth in R^{2}. Computers and Mathematics with Applications, 74(3):513–531, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302808.

Dominguez:2015:FDC

[DSVS15] Víctor Domínguez, Tonatiuh Sánchez-Vizuet, and Francisco-Javier Sayas. A fully discrete Calderón calculus for the twodimensional elastic wave equation. Computers and Mathematics with Applications, 69(7):620–635, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000589.

DeNayer:2018:EIM

[DSWB18] G. De Nayer, S. Schmidt, J. N. Wood, and M. Breuer. Enhanced injection method for synthetically generated turbulence within the flow domain of eddy-resolving simulations. Computers and Mathematics with Applications, 75(7):2338–2355, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307757.

Ding:2018:CAE

[DSZ18] Rui Ding, Quan Shen, and Zhengcheng Zhu. Convergence analysis and error estimates of the element-free Galerkin method for a class of parabolic evolutionary variational inequalities. Computers and Mathematics with Applications, 75(1):22–32, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305230.

Das:2011:RBE

[DT11] Kinkar Ch. Das and N. Trinajstić. Relationship between the eccentric connectivity index and Zagreb indices. Computers and Mathematics with Applications, 62(4):1758–1764, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004858.

Ding:2016:SYB

[DT16] Jiu Ding and Haiyan Tian. Solving the Yang– Baxter-like matrix equation for a class of elementary matrices. Computers and Mathematics with Applications, 72 (6):1541–1548, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304072.

Dilloo:2017:HOF

[DT17] Mehzabeen Jumanah Dilloo and Désiré Yannick Tangman. A highorder finite difference method for option valuation. Computers and Mathematics with Applications, 74(4):652–670, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730295X.

Devolder:2019:ANS

[DTR19] Brecht Devolder, Peter Troch, and Pieter Rauwoens. Accelerated numerical simulations of a heaving floating body by coupling a motion solver with a two-phase fluid solver. Computers and Mathematics with Applications, 77(6):1605–1625, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304966.

Dong:2018:SWH

[DTYZ18] Min-Jie Dong, Shou-Fu Tian, Xue-Wei Yan, and Li Zou. Solitary waves, homoclinic breather waves and rogue waves of the (3 + 1)- dimensional Hirota bilinear equation. Computers and Mathematics with Applications, 75(3):957–964, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306958.

Du:2012:NPI

[Du12] Kui Du. A new preconditioner for the interface system arising in a fast Helmholtz solver. Computers and Mathematics with Applications, 63(4):794–806, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010285.

Duan:2011:NRA

[Dua11] Jun-Sheng Duan. New recurrence algorithms for the nonclassic Adomian polynomials. Computers and Mathematics with Applications, 62(8):2961–2977, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006493.

Duan:2018:WPD

[Dua18]  Ning Duan.  Well-

posedness and decay of solutions for threedimensional generalized Navier–Stokes equations. Computers and Mathematics with Applications, 76(5):1026–1033, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303146.

Dubois:2013:SLB

[Dub13] François Dubois. Stable lattice Boltzmann schemes with a dual entropy approach for monodimensional nonlinear waves. Computers and Mathematics with Applications, 65(2):142–159, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005718.

Dung:2011:SAF

[Dun11] Nguyen Tien Dung. Semimartingale approximation of fractional Brownian motion and its applications. Computers and Mathematics with Applications, 61(7):1844–1854, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100099X.

Dunca:2018:NAT

[Dun18] Argus A. Dunca. Numerical analysis and testing of a stable and convergent finite element scheme for approximate deconvolution turbulence models. Computers and Mathematics with Applications, 75(2):690–702, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306156.

Denton:2010:FII

[DV10] Z. Denton and A. S. Vatsala. Fractional integral inequalities and applications. Computers and Mathematics with Applications, 59(3):1087–1094, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003538.

daVeiga:2018:HOV

[dVDR18] L. Beirão da Veiga, F. Dassi, and A. Russo. High-order Virtual Element Method on polyhedral meshes. Computers and Mathematics with Applications, 74 (5):1110–1122, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301839.

daVeiga:2018:MFD

[dVLV18] L. Beirão da Veiga, L. Lopez, and G. Vacca. Mimetic finite difference methods for Hamiltonian wave equations in 2D. Computers and Mathematics with Applications, 74(5):1123–1141, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303279.

Djondjorov:2012:DIV

[DVM12] Peter A. Djondjorov, Vassil M. Vassilev, and Ivaïlo M. Mladenov. Deformation of injected vesicles adhering onto flat rigid substrates. Computers and Mathematics with Applications, 64(3):214–220, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000545.

Dehestani:2013:CSM

[DVMS13] M. Dehestani, A. Vafai, M. Mofid, and F. Szidarovszky. Computation of the stresses in a moving reference system in a half-space due to a traversing time-varying concentrated load. Computers and Mathematics with Applications, 65(11): 1849–1862, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001612.

DiPietro:2014:PBF

[DVY14] Daniele A. Di Pietro, Martin Vohralík, and Soleiman Yousef. An a posteriori-based, fully adaptive algorithm with adaptive stopping criteria and mesh refinement for thermal multiphase compositional flows in porous media. Computers and Mathematics with Applications, 68 (12):2331–2347, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400368X.

Dlotko:2018:RCA

[DW18a] Paweł Dłotko and Thomas Wanner. Rigorous cubical approximation and persistent homology of continuous functions. Computers and Mathematics with Applications, 75(5):1648–1666, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307538.

Du:2018:PDE

[DW18b] Bo Du and Haiyan Wang.

Partial differential equation modeling of malware propagation in social networks with mixed delays. Computers and Mathematics with Applications, 75(10):3537–3548, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300889.

Dong:2012:OBC

[DWI^{+}12] Na Dong, Chun-Ho Wu, Wai-Hung Ip, Zeng-Qiang Chen, Ching-Yuen Chan, and Kai-Leung Yung. An opposition-based chaotic GA/PSO hybrid algorithm and its application in circle detection. Computers and Mathematics with Applications, 64(6):1886–1902, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002453.

Ding:2019:CAE

[DWS19] Rui Ding, Yu Wang, and Quan Shen. Convergence analysis and error estimates of the element-free Galerkin method for the second kind of elliptic variational inequalities. Computers and Mathematics with Applications, 78(8):2584–2592, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301907.

Dai:2015:GBP

[DWY15] Guowei Dai, Haiyan Wang, and Bianxia Yang. Global bifurcation and positive solution for a class of fully nonlinear problems. Computers and Mathematics with Applications, 69(8):771–776, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122115000875.

Dai:2013:FHA

[DWZ13] Ruxin Dai, Yin Wang, and Jun Zhang. Fast and high accuracy multiscale multigrid method with multiple coarse grid updating strategy for the 3D convection–diffusion equation. Computers and Mathematics with Applications, 66(4):542–559, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003738.

Dexter:2016:ECB

[DWZ16] Nick C. Dexter, Clayton G. Webster, and Guannan Zhang. Explicit cost bounds of stochastic Galerkin approximations for parameterized PDEs with random coefficients. Computers and Mathematics with Applications, 71(11):2231– 2256, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005738.

Dai:2011:SFF

[DYH11] Chao-Qing Dai, Qin Yang, and Bing-Chuanlong Huang. Semi-foldon fission and fusion in the (2+1)-dimensional higher order Broer–Kaup system. Computers and Mathematics with Applications, 61(7):1724–1727, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000642.

Duan:2014:GSS

[DYQM14] Xichao Duan, Sanling Yuan, Zhipeng Qiu, and Junling Ma. Global stability of an SVEIR epidemic model with ages of vaccination and latency. Computers and Mathematics with Applications, 68(3):288–308, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002612.

Dou:2019:CAM

[DYWL19] Yan Dou, Ai-Li Yang, Yu-Jiang Wu, and Zhao-Zheng Liang. Convergence analysis of modified PGSS methods for singular saddle-point problems. Computers and Mathematics with Applications, 77(1):93–104, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305285.

Deng:2011:GMS

[DYX11] Hanyuan Deng, Jianguang Yang, and Fangli Xia. A general modeling of some vertexdegree based topological indices in benzenoid systems and phenylenes. Computers and Mathematics with Applications, 61(10):3017–3023, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002689.

Duo:2016:MCF

[DZ16] Siwei Duo and Yanzhi Zhang. Mass-conservative Fourier spectral methods for solving the fractional nonlinear Schrödinger equation. Computers and Mathematics with Applications, 71(11):2257– 2271, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115006203.

Du:2017:LPF

[DZ17] Guangzhi Du and Liyun Zuo. Local and parallel finite element postprocessing scheme for the Stokes problem. Computers and Mathematics with Applications, 73(1):129–140, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306095.

Deng:2018:NST

[DZ18] W. H. Deng and Z. J. Zhang. Numerical schemes of the time tempered fractional Feynman–Kac equation. Computers and Mathematics with Applications, 73(6):1063–1076, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306885.

Ding:2021:CBD

[DZ21] Hang Ding and Jun Zhou. Comments on “Blow-up and decay for a class of pseudo-parabolic p-Laplacian equation with logarithmic nonlinearity”

[Comput. Math. Appl. 75(2) (2018) 459–469]. Computers and Mathematics with Applications, 84(??):144–147, February 15, 2021. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122120304703.

See [HGW18].

Davvaz:2010:FHR

[DZK10] Bijan Davvaz, Jianming Zhan, and Kyung Ho Kim. Fuzzy Γ-hypernear-rings. Computers and Mathematics with Applications, 59(8):2846–2853, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000945.

Dohr:2019:PST

[DZO^{+}19] Stefan Dohr, Jan Zapletal, Günther Of, Michal Merta, and Michal Kravcenko. A parallel space– time boundary element method for the heat equation. Computers and Mathematics with Applications, 78(9):2852–2866, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307296.

Davendra:2010:CDE

[DZS10] Donald Davendra, Ivan Zelinka, and Roman Senkerik. Chaos driven evolutionary algorithms for the task of PID control. Computers and Mathematics with Applications, 60(4):1088–1104, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002567.

Deng:2015:FST

[DZW^{+}15] Lin Deng, Yun Zhang, Yanwei Wen, Bin Shan, and Huamin Zhou. A fractional-step thermal lattice Boltzmann model for high Peclet number flow. Computers and Mathematics with Applications, 70(5):1152–1161, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003417.

Dai:2016:HOA

[DZW16] Ruxin Dai, Jun Zhang, and Yin Wang. Higher order ADI method with completed Richardson extrapolation for solving unsteady convection– diffusion equations. Computers and Mathematics with Applications, 71(1): 431–442, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115005751.

Davvaz:2011:FIR

[DZY11] Bijan Davvaz, Jianming Zhan, and Yunqiang Yin. Fuzzy H_{v}-ideals in Γ-H_{v}- rings. Computers and Mathematics with Applications, 61(3):690–698, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009211.

Elbashbeshy:2010:HTU

[EA10] Elsayed M. A. Elbashbeshy and Dalia A. Aldawody. Heat transfer over an unsteady stretching surface with variable heat flux in the presence of a heat source or sink. Computers and Mathematics with Applications, 60(10):2806–2811, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007212.

El-Ashwah:2010:SPC

[EAA10] R. M. El-Ashwah and M. K. Aouf. Some properties of certain classes of meromorphically p-valent functions involving extended multiplier transformations. Computers and Mathematics with Applications, 59(6):2111–2120, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007755.

El-Ashwah:2010:CMF

[EAAED10] R. M. El-Ashwah, M. K. Aouf, and S. M. El-Deeb. On a class of multivalent functions defined by an extended multiplier transformations. Computers and Mathematics with Applications, 60(3):623–628, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003573.

Espinoza-Andaluz:2018:CTD

[EAAS18] Mayken Espinoza-Andaluz, Martin Andersson, and Bengt Sundén. Computational time and domain size analysis of porous media flows using the lattice Boltzmann method. Computers and Mathematics with Applications, 74(1):26–34, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306721.

Elmoussaoui:2018:DKT

[EAEH18] A. Elmoussaoui, P. Argoul, M. El Rhabi, and A. Hakim. Discrete kinetic theory for 2D modeling of a moving crowd: Application to the evacuation of a non-connected bounded domain. Computers and Mathematics with Applications, 75(4):1159–1180, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306818.

Espinoza-Andaluz:2019:CSB

[EAMA19] Mayken Espinoza-Andaluz, Ayrton Moyón, and Martin Andersson. A comparative study between D2Q9 and D2Q5 lattice Boltzmann scheme for mass transport phenomena in porous media. Computers and Mathematics with Applications, 78(9):2886–2896, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930080X.

Ebadi:2011:CMM

[Eba11] Moosa Ebadi. A class of multistep methods based on a super-future points technique for solving IVPs. Computers and Mathematics with Applications, 61(11):3288– 3297, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003270.

Ebaid:2014:RHP

[Eba14] Abdelhalim Ebaid. Remarks on the homotopy perturbation method for the peristaltic flow of Jeffrey fluid with nanoparticles in an asymmetric channel. Computers and Mathematics with Applications, 68(3):77– 85, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002016.

El-Borai:2010:SFE

[EBENEA10] Mahmoud M. El-Borai, Khairia El-Said El-Nadi, and Eman G. El-Akabawy. On some fractional evolution equations. Computers and Mathematics with Applications, 59(3):1352–1355, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003587.

El-Borai:2010:SFS

[EBENF10] Mahmoud M. El-Borai, Khairia El-Said El-Nadi, and Hoda A. Fouad. On some fractional stochastic delay differential equations. Computers and Mathematics with Applications, 59(3):1165–1170, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003551.

Ebrahimi:2011:TMP

[Ebr11] Rahim Ebrahimi. Thermodynamic modeling of performance of a Miller cycle with engine speed and variable specific heat ratio of working fluid. Computers and Mathematics with Applications, 62(5):2169–2176, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005566.

Erdem:2010:TTS

[EÇ10] Yılmaz Erdem and İbrahim Çanak. A Tauberian theorem for (A)(C, α) summability. Computers and Mathematics with Applications, 60 (11):2920–2925, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007443.

Egorova:2016:NEN

[ECJ16] Vera N. Egorova, Rafael Company, and Lucas Jódar. A new efficient numerical method for solving American option under regime switching model. Computers and Mathematics with Applications, 71(1):224–237, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005623.

Erdogan:2011:DRS

[ECY11] M. Emre Erdogan, Cengiz Cinar, and Ibrahim Yalcinkaya. On the dynamics of the recursive sequence x_{n+1} = \frac{x_{n−1}}{β+γx_{n−2}^{2}x_{n−4}+γx_{n−2}x_{n−4}^{2}} . Computers and Mathematics with Applications, 61(3):533–537, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009028.

El-Daou:2011:EWL

[ED11a] Mohamed K. El-Daou.

Exponentially weighted Legendre–Gauss Tau methods for linear secondorder differential equations. Computers and Mathematics with Applications, 62(1):51–64, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003488.

Eslahchi:2011:ATS

[ED11b] M. R. Eslahchi and Mehdi Dehghan. Application of Taylor series in obtaining the orthogonal operational matrix. Computers and Mathematics with Applications, 61(9):2596– 2604, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001556.

Erbts:2012:ASC

[ED12] Patrick Erbts and Alexander Düster. Accelerated staggered coupling schemes for problems of thermoelasticity at finite strains. Computers and Mathematics with Applications, 64(8):2408–2430, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004002.

Ellis:2014:LCD

[EDC14] Truman Ellis, Leszek Demkowicz, and Jesse Chan. Locally conservative discontinuous Petrov–Galerkin finite elements for fluid problems. Computers and Mathematics with Applications, 68(11):1530–1549, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003113.

Esmaeili:2010:FPB

[EE10] Mostafa Esmaeili and Morteza Esmaeili. A Fibonacci-polynomial based coding method with error detection and correction. Computers and Mathematics with Applications, 60(10):2738–2752, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006760.

Esmaili:2018:AFP

[EE18] Sakine Esmaili and M. R.

Eslahchi. Application of fixed point-collocation method for solving an optimal control problem of a parabolic-hyperbolic free boundary problem modeling the growth of tumor with drug application. Computers and Mathematics with Applications, 75(7):2193–2216, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307058.

Ezzat:2010:SAH

[EEBM10] M. A. Ezzat, A. A. El-Bary, and M. M. Morsey. Space approach to the hydromagnetic flow of a dusty fluid through a porous medium. Computers and Mathematics with Applications, 59(8):2868–2879, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000969.

Escobedo:2014:FBP

[EF14] Ramón Escobedo and Luis A. Fernández. Free boundary problems and optimal control of axisymmetric polymer crystallization processes. Computers and Mathematics with Applications, 68(1– 2):27–43, July 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001849.

Ebrahimnejad:2015:ARM

[EFK15] M. Ebrahimnejad, N. Fallah, and A. R. Khoei. Adaptive refinement in the meshless finite volume method for elasticity problems. Computers and Mathematics with Applications, 69(12):1420– 1443, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001406.

Esmaeilpour:2010:SJH

[EG10] M. Esmaeilpour and D. D. Ganji. Solution of the Jeffery–Hamel flow problem by optimal homotopy asymptotic method. Computers and Mathematics with Applications, 59(11):3405– 3411, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002002.

Ern:2018:AEF

[EG18] Alexandre Ern and Jean-Luc Guermond. Analysis of the edge finite element approximation of the Maxwell equations with low regularity solutions. Computers and Mathematics with Applications, 75(3):918–932, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730665X.

El-Ganaini:2019:NAW

[EGAA19] Shoukry El-Ganaini and Mohammed O. Al-Amr. New abundant wave solutions of the conformable space–time fractional (4+1)-dimensional Fokas equation in water waves. Computers and Mathematics with Applications, 78 (6):2094–2106, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301816.

Engstrom:2016:ERF

[EGG16] Christian Engström, Stefano Giani, and Luka Grubisić. Efficient and reliable hp–FEM estimates for quadratic eigenvalue problems and photonic crystal applications. Computers and Mathematics with Applications, 72(4):952–973, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303339.

Ezquerro:2012:AES

[EGGS^{+}12] J. A. Ezquerro, A. Grau, M. Grau-Sánchez, M. A. Hernández, and M. Noguera. Analysing the efficiency of some modifications of the secant method. Computers and Mathematics with Applications, 64 (6):2066–2073, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003239.

Ezquerro:2010:VCT

[EGSHR10] J. A. Ezquerro, M. Grau-Sánchez, M. A. Hernández, and N. Romero. Variants of a classic Traub’s result. Computers and Mathematics with Applications, 60(11):2899–2908, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007145.

Espig:2014:ELR

[EHL^{+}14] Mike Espig, Wolfgang Hackbusch, Alexander Litvinenko, Hermann G. Matthies, and Philipp Wähnert. Efficient lowrank approximation of the stochastic Galerkin matrix in tensor formats. Computers and Mathematics with Applications, 67(4):818–829, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006244.

Ellabaan:2012:TSC

[EHO^{+}12] M. M. Ellabaan, S. D. Handoko, Y. S. Ong, C. K. Kwoh, S. A. Bahnassy, F. M. Elassawy, and H. Y. Man. A tree-structured covalentbond-driven molecular memetic algorithm for optimization of ringdeficient molecules. Computers and Mathematics with Applications, 64 (12):3792–3804, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001800.

Ehrenmark:2018:SEC

[Ehr18] Ulf Ehrenmark. Summability experiments with a class of divergent inverse Kontorovich–Lebedev transforms. Computers and Mathematics with Applications, 76(1):141–154, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302074.

El-Khazali:2013:FOC

[EK13] Reyad El-Khazali. Fractionalorder P I^{λ}D^{μ} controller design. Computers and Mathematics with Applications, 66(5):639–646, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001247.

El-Kalaawy:2016:VPC

[EK16] O. H. El-Kalaawy. Variational principle, conservation laws and exact solutions for dust ion acoustic shock waves modeling modified Burger equation. Computers and Mathematics with Applications, 72(4):1031–1041, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303455.

Evans:2018:AFC

[EKE18] Ryan M. Evans, Udita N.

Katugampola, and David A. Edwards. Applications of fractional calculus in solving Abel-type integral equations: Surfacevolume reaction problem. Computers and Mathematics with Applications, 73(6):1346–1362, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306769.

El-Khazali:2010:DDF

[EKS10] Reyad El-Khazali and M. H. B. M. Shariff. Double-delay fractional and integer-order tanlock loops. Computers and Mathematics with Applications, 59(5):1874–1884, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005653.

Essoufi:2017:ADM

[EKZ17] E.-H. Essoufi, J. Koko, and A. Zafrar. Alternating direction method of multiplier for a unilateral contact problem in electro-elastostatics. Computers and Mathematics with Applications, 73(8):1789–1802, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301062.

Elboree:2011:JEF

[Elb11] Mohammed K. Elboree.

The Jacobi elliptic function method and its application for two component BKP hierarchy equations. Computers and Mathematics with Applications, 62(12):4402–4414, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008790.

Elboree:2015:VAS

[Elb15] Mohammed K. Elboree.

Variational approach, soliton solutions and singular solitons for new coupled ZK system. Computers and Mathematics with Applications, 70(5):934–941, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003041.

Elsaid:2010:VIM

[Els10] A. Elsaid. The variational iteration method for solving Riesz fractional partial differential equations. Computers and Mathematics with Applications, 60(7):1940–1947, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005195.

Elliott:2011:CFO

[ELS11] Robert J. Elliott, Chuin Ching Liew, and Tak Kuen Siu. Characteristic functions and option valuation in a Markov chain market. Computers and Mathematics with Applications, 62(1):65–74, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003531.

Esterhazy:2014:ADH

[EM14] S. Esterhazy and J. M. Melenk. An analysis of discretizations of the Helmholtz equation in L^{2} and in negative norms. Computers and Mathematics with Applications, 67(4):830–853, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006123.

El-Mistikawy:2019:CTM

[EM19] Tarek M. A. El-Mistikawy.

Comment on “Truncation method with point transformation for exact solution of Liouville Bratu Gelfand equation” by R. Saleh, S. M. Mabrouk, M. Kassem, Computers and Mathematics with Applications 76(5) (2018) 1219–1227. Computers and Mathematics with Applications, 78(6):1887–1888, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930149X.

See [SMK18].

Abd-el-Malek:2018:LGM

[eMA18] Mina B. Abd el Malek and Amr M. Amin. Lie group method for solving viscous barotropic vorticity equation in ocean climate models. Computers and Mathematics with Applications, 75(4):1443–1461, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117307307.

Egidi:2018:IEM

[EMQ18] Nadaniela Egidi, Pierluigi Maponi, and Michela Quadrini. An integral equation method for the numerical solution of the Burgers equation. Computers and Mathematics with Applications, 76(1):35–44, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301949.

El-Mikkawy:2010:SAI

[EMR10] Moawwad El-Mikkawy and El-Desouky Rahmo. Symbolic algorithm for inverting cyclic pentadiagonal matrices recursively — derivation and implementation. Computers and Mathematics with Applications, 59(4):1386–1396, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007792.

Eliahou:2012:WSN

[EMRS12] S. Eliahou, J. M. Marín, M. P. Revuelta, and M. I.

Sanz. Weak Schur numbers and the search for G. W. Walker’s lost partitions. Computers and Mathematics with Applications, 63(1):175–182, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009722.

El-Nabulsi:2011:FBT

[EN11] Rami Ahmad El-Nabulsi.

The fractional Boltzmann transport equation. Computers and Mathematics with Applications, 62(3):1568–1575, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001921.

Einkemmer:2014:ASS

[EO14] Lukas Einkemmer and Alexander Ostermann. An almost symmetric Strang splitting scheme for nonlinear evolution equations. Computers and Mathematics with Applications, 67(12): 2144–2157, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122114000996.

Einkemmer:2015:EPS

[EO15] Lukas Einkemmer and Alexander Ostermann. On the error propagation of semi-Lagrange and Fourier methods for advection problems. Computers and Mathematics with Applications, 69(3):170–179, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005902.

Erturk:2011:ASF

[EOM11] Vedat Suat Ertürk, Zaid M. Odibat, and Shaher Momani. An approximate solution of a fractional order differential equation model of human T-cell lymphotropic virus I (HTLV-I) infection of CD4^{+} T-cells. Computers and Mathematics with Applications, 62(3):996–1002, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002719.

Oshagh:2018:AWC

[eOS18] M. Khaksar e Oshagh and M. Shamsi. An adaptive wavelet collocation method for solving optimal control of elliptic variational inequalities of the obstacle type. Computers and Mathematics with Applications, 75(2):470–485, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305898.

Erdozain:2018:AME

[EPP18] A. Erdozain, V. Péron, and D. Pardo. Asymptotic models for the electric potential across a highly conductive casing. Computers and Mathematics with Applications, 76(8):1975–2000, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304140.

Ergashev:2019:FSM

[Erg19] T. G. Ergashev. On fundamental solutions for multidimensional Helmholtz equation with three singular coefficients. Computers and Mathematics with Applications, 77(1): 69–76, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305261.

Ershkov:2016:NSR

[Ers16] Sergey V. Ershkov. Nonstationary Riccati-type flows for incompressible 3D Navier–Stokes equations. Computers and Mathematics with Applications, 71(7):1392–1404, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300542.

Ervin:2012:CBT

[Erv12] V. J. Ervin. Computational bases for RT_{k} and BDM_{k} on triangles. Computers and Mathematics with Applications, 64(8):2765–2774, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005469.

Eriksson:2010:CCD

[ES10] Robert Eriksson and Lennart Söder. Coordinated control design of multiple HVDC links based on model identification. Computers and Mathematics with Applications, 60(4):944–953, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001562.

Elliott:2011:MAD

[ES11] Robert J. Elliott and Tak Kuen Siu. An mary detection approach for asset allocation. Computers and Mathematics with Applications, 62(4):2083–2094, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005360.

El-Shiekh:2017:PSW

[ES17] Rehab M. El-Shiekh. Periodic and solitary wave solutions for a generalized variable-coefficient Boiti– Leon–Pempinlli system. Computers and Mathematics with Applications, 73(7):1414–1420, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300287.

El-Shiekh:2018:JEW

[ES18] Rehab M. El-Shiekh. Jacobi elliptic wave solutions for two variable coefficients cylindrical Korteweg–de Vries models arising in dusty plasmas by using direct reduction method. Computers and Mathematics with Applications, 75(5):1676–1684, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307575.

El-Sayed:2010:ADM

[ESBR10] A. M. A. El-Sayed, S. H. Behiry, and W. E. Raslan. Adomian’s decomposition method for solving an intermediate fractional advectiondispersion equation. Computers and Mathematics with Applications, 59(5): 1759–1765, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005537.

Esmaeili:2011:NSF

[ESL11] Shahrokh Esmaeili, M. Shamsi, and Yury Luchko. Numerical solution of fractional differential equations with a collocation method based on Müntz polynomials. Computers and Mathematics with Applications, 62(3):918– 929, August 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111003245.

El-Shahed:2010:NSN

[ESN10] Moustafa El-Shahed and Juan J. Nieto. Nontrivial solutions for a nonlinear multi-point boundary value problem of fractional order. Computers and Mathematics with Applications, 59(11): 3438–3443, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110002208.

Tatar:2010:ERE

[eT10] Nasser eddine Tatar.

Existence results for an evolution problem with fractional nonlocal conditions. Computers and Mathematics with Applications, 60 (11):2971–2982, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007546.

Erdem:2012:STT

[ET12] Yilmaz Erdem and Ümit Totur. Some Tauberian theorems for the product method of Borel and Cesàro summability. Computers and Mathematics with Applications, 64(9):2871–2876, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003719.

Evcin:2018:DOP

[EUTS18] Cansu Evcin, Ömür Ugur, and Münevver Tezer-Sezgin. Determining the optimal parameters for the MHD flow and heat transfer with variable viscosity and Hall effect. Computers and Mathematics with Applications, 76(6):1338–1355, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118303511.

Erturk:2012:NAM

[EZM12] Vedat Suat Ertürk, Gul Zaman, and Shaher Momani. A numeric-analytic method for approximating a giving up smoking model containing fractional derivatives. Computers and Mathematics with Applications, 64 (10):3065–3074, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001162.

Elias-Zuniga:2010:SSN

[EZRR10] Alex Elías-Zúñiga, Ciro A.

Rodríguez, and Oscar Martínez Romero. On the solution of strong nonlinear oscillators by applying a rational elliptic balance method. Computers and Mathematics with Applications, 60(5):1409–1420, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004372.

Fan:2017:GWP

[FAHZ17] Jishan Fan, Ahmed Alsaedi, Tasawar Hayat, and Yong Zhou. Global well-posedness and regularity criteria for epitaxial growth models. Computers and Mathematics with Applications, 74(3):459–465, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302602.

Fetecau:2010:OMO

[FAIV10] Corina Fetecau, W. Akhtar, M. A. Imran, and D. Vieru. On the oscillating motion of an Oldroyd-b fluid between two infinite circular cylinders. Computers and Mathematics with Applications, 59(8):2836–2845, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000933.

Faraz:2011:SER

[Far11]  Naeem Faraz.  Study

of the effects of the Reynolds number on circular porous slider via variational iteration algorithm — II. Computers and Mathematics with Applications, 61(8):1991–1994, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000622X.

Furkan:2010:FSO

[FBB10] Hasan Furkan, Hüseyin Bilgiç, and Feyzi Basar. On the fine spectrum of the operator B(r, s, t) over the sequence spaces ℓ_{p} and bv_{p}, (1 < p < ∞). Computers and Mathematics with Applications, 60(7):2141–2152, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005511.

Fermo:2011:ASS

[FBL11] L. Fermo, N. Bellomo, and D. B. Lumenta. Assessment of surgical strategies for addressing keloids: an optimization problem. Computers and Mathematics with Applications, 62(6):2417–2423, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005827.

Fendoglu:2019:MFR

[FBTS19] Hande Fendoglu, Canan Bozkaya, and Münevver Tezer-Sezgin. MHD flow in a rectangular duct with a perturbed boundary. Computers and Mathematics with Applications, 77(2):374–388, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305522.

Fan:2012:DCN

[FCZ12] Zhengping Fan, Guanrong Chen, and Yunong Zhang. Differentiating complex network models: an engineering perspective. Computers and Mathematics with Applications, 64(5):840–848, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011345.

Faieghi:2013:NSS

[FDB13] Mohammad Reza Faieghi, Hadi Delavari, and Dumitru Baleanu. A note on stability of sliding mode dynamics in suppression of fractional-order chaotic systems. Computers and Mathematics with Applications, 66(5):832–837, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006761.

Farias:2017:TDM

[FDG^{+}17] Agnaldo M. Farias, Philippe R. B. Devloo, Sônia M. Gomes, Denise de Siqueira, and Douglas A. Castro. Two dimensional mixed finite element approximations for elliptic problems with enhanced accuracy for the potential and flux divergence. Computers and Mathematics with Applications, 74 (12):3283–3295, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305059.

Ferreira:2017:MAD

[FdOdSS17] J. A. Ferreira, Paula de Oliveira, P. M. da Silva, and R. Silva. Mathematics of aging: Diseases of the posterior segment of the eye. Computers and Mathematics with Applications, 73(1):11–26, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305843.

Ferreira:2017:TIQ

[FdOP17] J. A. Ferreira, P. de Oliveira, and G. Pena. Transdermal iontophoresis — a quantitative and qualitative study. Computers and Mathematics with Applications, 74 (10):2231–2242, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304170.

Feng:2011:DLS

[FDXW11] Jianwen Feng, Anding Dai, Chen Xu, and Jingyi Wang. Designing lag synchronization schemes for unified chaotic systems. Computers and Mathematics with Applications, 61(8):2123–2128, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006772.

Ding:2011:NIF

[fDxZ11] Heng fei Ding and Yu xin Zhang. Notes on “Implicit finite difference approximation for time fractional diffusion equations” [Comput. Math. Appl. 56 (2008) 1138– 1145]. Computers and Mathematics with Applications, 61(9):2924–2928, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001520.

See [Mur08].

Ding:2012:NNM

[fDxZ12] Heng fei Ding and Yu xin Zhang. New numerical methods for the Riesz space fractional partial differential equations. Computers and Mathematics with Applications, 63(7):1135–1146, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010832.

Ferreira:2011:PSC

[Fer11] Rui A. C. Ferreira.

Positive solutions for a class of boundary value problems with fractional q-differences. Computers and Mathematics with Applications, 61(2):367–373, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008618.

Ferreira:2012:FDA

[Fer12] Lucas C. F. Ferreira. A family of dissipative active scalar equations with singular velocity and measure initial data. Computers and Mathematics with Applications, 64(10):3292–3301, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200106X.

Farhadi:2015:NSN

[FER15] Asghar Farhadi, Homayoun Emdad, and Ebrahim Goshtasbi Rad. On the numerical simulation of the nonbreaking solitary waves run up on sloping beaches. Computers and Mathematics with Applications, 70(9):2270–2281, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004150.

Farhadi:2017:DNM

[FES17] A. Farhadi, G. H. Erjaee, and M. Salehi. Derivation of a new Merton’s optimal problem presented by fractional stochastic stock price and its applications. Computers and Mathematics with Applications, 73(9):2066–2075, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301104.

Fucik:2019:ONS

[FES^{+}19] Radek Fučík, Pavel Eichler, Robert Straka, Petr Pauš, Jakub Klinkovský, and Tomáš Oberhuber. On optimal node spacing for immersed boundary-lattice Boltzmann method in 2D and 3D. Computers and Mathematics with Applications, 77 (4):1144–1162, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306424.

Fekete:2014:SCT

[FF14] Imre Fekete and István Faragó. Stability concepts and their applications. Computers and Mathematics with Applications, 67(12):2158– 2170, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000959.

Fornberg:2015:FGD

[FF15] Bengt Fornberg and Natasha Flyer. Fast generation of 2-D node distributions for meshfree PDE discretizations. Computers and Mathematics with Applications, 69(7):531–544, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000334.

Far:2018:LDT

[FG18] M. N. Sadraee Far and M. E. Golmakani. Large deflection of thermomechanical loaded bilayer orthotropic graphene sheet in/on polymer matrix based on nonlocal elasticity theory. Computers and Mathematics with Applications, 76 (9):2061–2089, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304206.

Fakhari:2019:SPF

[FGB19] Abbas Fakhari, Martin Geier, and Diogo Bolster. A simple phasefield model for interface tracking in three dimensions. Computers and Mathematics with Applications, 78(4):1154–1165, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304758.

Farago:2013:ENM

[FGHZ13] István Faragó, Krassimir Georgiev, Ágnes Havasi, and Zahari Zlatev. Efficient numerical methods for scientific applications: Introduction. Computers and Mathematics with Applications, 65(3):297–300, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000023.

Farago:2014:EAL

[FGHZ14] István Faragó, Krassimir Georgiev, Ágnes Havasi, and Zahari Zlatev. Efficient algorithms for large scale scientific computations: Introduction. Computers and Mathematics with Applications, 67(12):2085–2087, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002478.

Fu:2017:SMF

[FGHZ17] Hongfei Fu, Hui Guo, Jian Hou, and Jiansong Zhang. A stabilized mixed finite element approximation of bilinear state optimal control problems. Computers and Mathematics with Applications, 74 (6):1246–1261, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303589.

Fojon:2010:QSW

[FGL10] O. Fojón, M. Gadella, and L. P. Lara. The quantum square well with moving boundaries: a numerical analysis. Computers and Mathematics with Applications, 59(2):964–976, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006750.

Fortes:2017:FHG

[FGPP17] M. A. Fortes, P. González, A. Palomares, and M. Pasadas. Filling holes with geometric and volumetric constraints. Computers and Mathematics with Applications, 74(4):671–683, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302985.

Fang:2017:EMD

[FGY^{+}17] Hong Fang, Chunye Gong, Caihui Yu, Changwan Min, Xing Zhang, Jie Liu, and Liquan Xiao. Efficient mesh deformation based on Cartesian background mesh. Computers and Mathematics with Applications, 73(1):71–86, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630606X.

Feng:2011:MHP

[FH11] Xinlong Feng and Yinnian He. Modified homotopy perturbation method for solving the Stokes equations. Computers and Mathematics with Applications, 61(8): 2262–2266, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007388.

Ferhat:2016:GEE

[FH16] Mohamed Ferhat and Ali Hakem. Global existence and energy decay result for a weak viscoelastic wave equations with a dynamic boundary and nonlinear delay term. Computers and Mathematics with Applications, 71(3):779–804, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115006070.

Fuhrer:2017:RCD

[FH17] Thomas Führer and Norbert Heuer. Robust coupling of DPG and BEM for a singularly perturbed transmission problem. Computers and Mathematics with Applications, 74(8):1940–1954, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305284.

Fu:2016:IGM

[FHA16] Shengmao Fu, Guangjian Huang, and Badradeen Adam. Instability in a generalized multi-species Keller–Segel chemotaxis model. Computers and Mathematics with Applications, 72(9):2280–2288, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304734.

Fard:2013:URL

[FHH13] Seyed Mehdi Hazrati Fard, Ali Hamzeh, and Sattar Hashemi. Using reinforcement learning to find an optimal set of features. Computers and Mathematics with Applications, 66 (10):1892–1904, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004495.

Fu:2018:NSC

[FHS18] Zhixing Fu, Norbert Heuer, and Francisco-Javier Sayas. A nonsymmetric coupling of boundary elements with the hybridizable discontinuous Galerkin method. Computers and Mathematics with Applications, 74(11):2752–2768, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305278.

Farago:2010:EIS

[FHZ10] István Faragó, Ágnes Havasi, and Zahari Zlatev. Efficient implementation of stable Richardson extrapolation algorithms. Computers and Mathematics with Applications, 60(8):2309–2325, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005894.

Farago:2013:CDI

[FHZ13] István Faragó, Ágnes Havasi, and Zahari Zlatev. The convergence of diagonally implicit Runge– Kutta methods combined with Richardson extrapolation. Computers and Mathematics with Applications, 65(3):395–401, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003653.

Fialko:2015:PDS

[Fia15] S. Fialko. Parallel direct solver for solving systems of linear equations resulting from finite element method on multi-core desktops and workstations. Computers and Mathematics with Applications, 70 (12):2968–2987, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004927.

Fakhar-Izadi:2014:STS

[FID14] Farhad Fakhar-Izadi and Mehdi Dehghan. Space– time spectral method for a weakly singular parabolic partial integro-differential equation on irregular domains. Computers and Mathematics with Applications, 67(10):1884– 1904, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001497.

Furati:2018:ISP

[FIM18] Khaled M. Furati, Olaniyi S. Iyiola, and Kassem Mustapha. An inverse source problem for a two-parameter anomalous diffusion with local time datum. Computers and Mathematics with Applications, 73(6):1008–1015, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303686.

Falagkaris:2018:PCI

[FIMV18] E. J. Falagkaris, D. M. Ingram, K. Markakis, and I. M. Viola. PRO-TEUS: a coupled iterative force-correction immersed-boundary cascaded lattice Boltzmann solver for moving and deformable boundary applications. Computers and Mathematics with Applications, 75 (4):1330–1354, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307071.

Fiori:2014:TDP

[Fio14]  S. Fiori.  A two-

dimensional Poisson equation formulation of nonparametric statistical nonlinear modeling. Computers and Mathematics with Applications, 67(5):1171– 1185, March 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211300672X.

Fishelov:2018:NFO

[Fis18a]  D. Fishelov.  A new

fourth-order compact scheme for the Navier–Stokes equations in irregular domains. Computers and Mathematics with Applications, 74(1):6–25, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305910.

Foucher:2018:CPN

[FIS18b] Françoise Foucher, Moustafa Ibrahim, and Mazen Saad. Convergence of a positive nonlinear Control Volume Finite Element scheme for solving an anisotropic degenerate breast cancer development model. Computers and Mathematics with Applications, 76(3):551–578, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302517.

Falagkaris:2017:PCI

[FIVM17] E. J. Falagkaris, D. M. Ingram, I. M. Viola, and K. Markakis. PRO-TEUS: A coupled iterative force-correction immersed-boundary multidomain cascaded lattice Boltzmann solver. Computers and Mathematics with Applications, 74 (10):2348–2368, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304339.

Foster:2013:TLF

[FIW13] Erich L. Foster, Traian Iliescu, and David R. Wells. A two-level finite element discretization of the streamfunction formulation of the stationary quasi-geostrophic equations of the ocean.

Computers and Mathematics with Applications, 66(7):1261–1271, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004768.

Fino:2017:BRN

[FIW17] A. Z. Fino, H. Ibrahim, and A. Wehbe. A blow-up result for a nonlinear damped wave equation in exterior domain: The critical case. Computers and Mathematics with Applications, 73(11):2415–2420, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302043.

Fang:2019:FSE

[FJ19] Fei Fang and Chao Ji. On a fractional Schrödinger equation with periodic potential. Computers and Mathematics with Applications, 78(5):1517–1530, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301695.

Fallah:2019:DKM

[FJB19] Alireza Fallah, Ehsan Jabbari, and Reza Babaee. Development of the Kansa method for solving seepage problems using a new algorithm for the shape parameter optimization. Computers and Mathematics with Applications, 77(3):815–829, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306187.

Fan:2016:PEF

[FJC16] Wenping Fan, Xiaoyun Jiang, and Shanzhen Chen. Parameter estimation for the fractional fractal diffusion model based on its numerical solution. Computers and Mathematics with Applications, 71(2):642–651, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005982.

Ferreira:2018:ACH

[FJP18] J. A. Ferreira, D. Jordão, and L. Pinto. Approximating coupled hyperbolicparabolic systems arising in enhanced drug delivery. Computers and Mathematics with Applications, 76(1):81–97, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302037.

Fan:2016:FEW

[FJWW16] Xiaochuan Fan, Lili Ju, Xiaoqiang Wang, and Song Wang. A fuzzy edge-weighted centroidal Voronoi tessellation model for image segmentation. Computers and Mathematics with Applications, 71(11): 2272–2284, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115005180.

Farzanyar:2012:MFM

[FKC12] Zahra Farzanyar, Mohammadreza Kangavari, and Nick Cercone. Max– FISM: Mining (recently) maximal frequent itemsets over data streams using the sliding window model. Computers and Mathematics with Applications, 64(6):1706–1718, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000557.

Fuentes:2015:OEH

[FKDN15] Federico Fuentes, Brendan Keith, Leszek Demkowicz, and Sriram Nagaraj. Orientation embedded high order shape functions for the exact sequence elements of all shapes. Computers and Mathematics with Applications, 70(4):353–458, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002084.

Furati:2012:EUP

[FKeT12] K. M. Furati, M. D. Kassim, and N. e. Tatar. Existence and uniqueness for a problem involving Hilfer fractional derivative. Computers and Mathematics with Applications, 64(6):1616–1626, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000193.

Faria:2013:RSM

[FKF13] C. O. Faria and J. Karam-Filho. A regularizedstabilized mixed finite element formulation for viscoplasticity of Bingham type. Computers and Mathematics with Applications, 66(6):975–995, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003994.

Fakhar:2011:CAS

[FKKS11] K. Fakhar, A. H. Kara, I. Khan, and M. Sajid. On the computation of analytical solutions of an unsteady magnetohydrodynamics flow of a third grade fluid with Hall effects. Computers and Mathematics with Applications, 61(4):980–987, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000951X.

Fu:2010:ONC

[FL10] Sheng-Chen Fu and Ming-Li Lin. Oscillation and nonoscillation criteria for linear dynamic systems on time scales. Computers and Mathematics with Applications, 59(8):2552–2565, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000026X.

Farid:2011:IMS

[FL11a] Mahboubeh Farid and Wah June Leong. An improved multi-step gradienttype method for large scale optimization. Computers and Mathematics with Applications, 61(11): 3312–3318, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003312.

Feng:2011:UAC

[FL11b] Yun-Long Feng and Shao-Gao Lv. Unified approach to coefficientbased regularized regression. Computers and Mathematics with Applications, 62(1):506–515, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004329.

Fredini:2013:EWC

[FL13a] Pablo S. Rojas Fredini and Alejandro C.

Limache. Evaluation of weakly compressible SPH variants using derived analytical solutions of Taylor–Couette flows. Computers and Mathematics with Applications, 66(3):304–317, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003106.

Fu:2013:SPF

[FL13b] Shengmao Fu and Ji Liu.

Spatial pattern formation in the Keller–Segel model with a logistic source. Computers and Mathematics with Applications, 66(3):403–417, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003192.

Fan:2014:MIT

[FL14] Hongxia Fan and Yongxiang Li. Monotone iterative technique for the elastic systems with structural damping in Banach spaces. Computers and Mathematics with Applications, 68(3):384– 391, August 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002685.

Liu:2010:DRM

[fLcJ10] Li fang Liu and Li cheng Jiao. Detection of overrepresented motifs corresponding to known TFBSs via motif clustering and matching. Computers and Mathematics with Applications, 59(2):779–786, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007159.

Faria:2014:PSH

[FLdS14] Cristiane O. Faria, Abimael F. D. Loula, and Antônio J. B. dos Santos. Primal stabilized hybrid and DG finite element methods for the linear elasticity problem. Computers and Mathematics with Applications, 68(4):486–507, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002752.

Fu:2012:PAL

[FLDZ12] Ke-Ang Fu, Jie Li, Ya-Juan Dong, and Hui Zhou. Precise asymptotics for the linear processes generated by associated random variables in Hilbert spaces. Computers and Mathematics with Applications, 64 (6):1937–1943, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002519.

Farid:2010:NTS

[FLH10] Mahboubeh Farid, Wah June Leong, and Malik Abu Hassan. A new two-step gradient-type method for large-scale unconstrained optimization. Computers and Mathematics with Applications, 59(10):3301–3307, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001902.

Feng:2010:ALS

[FLLF10] Feng Feng, Yongming Li, and Violeta Leoreanu-Fotea. Application of level soft sets in decision making based on interval-valued fuzzy soft sets. Computers and Mathematics with Applications, 60(6):1756–1767, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004815.

Fornberg:2013:SCG

[FLP13] Bengt Fornberg, Erik Lehto, and Collin Powell. Stable calculation of Gaussian-based RBF–FD stencils. Computers and Mathematics with Applications, 65(4):627–637, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006529.

Fu:2011:STS

[FLWJ11] Xiang Fu, Anqiang Li, Liping Wang, and Changming Ji. Short-term scheduling of cascade reservoirs using an immune algorithm-based particle swarm optimization. Computers and Mathematics with Applications, 62(6):2463–2471, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005864.

Fan:2014:SSS

[FLZ14a] Mingshu Fan, Shan Li, and Lei Zhang. Structural stability of subsonic irrotational flows in two-dimensional infinitely long nozzles. Computers and Mathematics with Applications, 68(1–2):61– 66, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400176X.

Fan:2014:WSE

[FLZ14b] Mingshu Fan, Shan Li, and Lei Zhang. Weak solution of the equation for a fractional porous medium with a forcing term. Computers and Mathematics with Applications, 67(1):145–150, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005920.

Fatehi:2011:EES

[FM11] R. Fatehi and M. T. Manzari. Error estimation in smoothed particle hydrodynamics and a new scheme for second derivatives. Computers and Mathematics with Applications, 61(2):482–498, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009004.

Feng:2012:SGO

[FM12a] Qinghua Feng and Fanwei Meng. Some generalized Ostrowski-Grüss type integral inequalities. Computers and Mathematics with Applications, 63(3):652–659, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009837.

Ford:2012:DOE

[FM12b] N. J. Ford and M. L. Morgado. Distributed order equations as boundary value problems. Computers and Mathematics with Applications, 64 (10):2973–2981, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000740.

Fernandez:2017:MTS

[FM17] José R. Fernández and Maria Masid. A mixture of thermoelastic solids with two temperatures. Computers and Mathematics with Applications, 73(9):1886–1899, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301049.

Faustmann:2018:REC

[FM18] M. Faustmann and J. M.

Melenk. Robust exponential convergence of hp-FEM in balanced norms for singularly perturbed reaction–diffusion problems: Corner domains. Computers and Mathematics with Applications, 74(7):1576–1589, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301773.

Fardipour:2019:MSO

[FM19] Kaveh Fardipour and Kamyar Mansour. A modified seventh-order WENO scheme with new nonlinear weights for hyperbolic conservation laws. Computers and Mathematics with Applications, 78 (12):3748–3769, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303098.

Fogerty:2019:CSM

[FMGR19] Shane Fogerty, Matt Martineau, Rao Garimella, and Robert Robey. A comparative study of multi-material data structures for computational physics applications. Computers and Mathematics with Applications, 78(2): 565–581, July 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118303341.

Ferro:2019:PAS

[FMP19] Nicola Ferro, Stefano Micheletti, and Simona Perotto. POD-assisted strategies for structural topology optimization. Computers and Mathematics with Applications, 77(10):2804–2820, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122119300318.

Fuhrer:2015:OAS

[FMPR15] T. Führer, J. M. Melenk, D. Praetorius, and A. Rieder. Optimal additive Schwarz methods for the hp-BEM: the hypersingular integral operator in 3D on locally refined meshes. Computers and Mathematics with Applications, 70(7):1583–1605, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003119.

Falletta:2019:TBI

[FMS19] S. Falletta, G. Monegato, and L. Scuderi. Two boundary integral equation methods for linear elastodynamics problems on unbounded domains. Computers and Mathematics with Applications, 78(12):3841–3861, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303207.

Frittelli:2017:LFE

[FMSV17] Massimo Frittelli, Anotida Madzvamuse, Ivonne Sgura, and Chandrasekhar Venkataraman. Lumped finite elements for reactioncross-diffusion systems on stationary surfaces. Computers and Mathematics with Applications, 74 (12):3008–3023, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730473X.

Feng:2014:FEA

[FN14] Xiaobing Feng and Michael Neilan. Finite element approximations of general fully nonlinear second order elliptic partial differential equations based on the vanishing moment method. Computers and Mathematics with Applications, 68 (12):2182–2204, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003411.

Niu:2011:PBA

[fNS11] Yi feng Niu and Fang-Ming Shao. A practical bounding algorithm for computing two-terminal reliability based on decomposition technique.

Computers and Mathematics with Applications, 61(8):2241–2246, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007194.

Fu:2018:DCF

[FNW18] Hongfei Fu, Michael K.

Ng, and Hong Wang. A divide-and-conquer fast finite difference method for space-time fractional partial differential equation. Computers and Mathematics with Applications, 73(6):1233–1242, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306526.

Fang:2013:IMA

[FNZ13] Min Fang, WenKe Niu, and XiaoSong Zhang. An improved multipleattractor cellular automata classifier with a tree frame based on CART. Computers and Mathematics with Applications, 66 (10):1836–1844, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004653.

Fuica:2019:PEA

[FOS19] Francisco Fuica, Enrique Otárola, and Abner J. Salgado. An a posteriori error analysis of an elliptic optimal control problem in measure space. Computers and Mathematics with Applications, 77(10):2659–2675, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300021.

Fu:2011:SAD

[FOX11] Hongjie Fu, Dantong Ouyang, and Jiaming Xu. A self-adaptive differential evolution algorithm for binary CSPs. Computers and Mathematics with Applications, 62(7):2712–2718, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005335.

Fuhrer:2018:LUT

[FP18] Thomas Führer and Dirk Praetorius. A linear Uzawa-type FEM– BEM solver for nonlinear transmission problems. Computers and Mathematics with Applications, 75(8):2678–2697, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830004X.

Farrell:2019:NDB

[FP19] Patricio Farrell and Dirk Peschka. Nonlinear diffusion, boundary layers and nonsmoothness: Analysis of challenges in drift-diffusion semiconductor simulations. Computers and Mathematics with Applications, 78 (12):3731–3747, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303104.

Faverjon:2017:IBC

[FPB17] B. Faverjon, B. Puig, and T. N. Baranger. Identification of boundary conditions by solving Cauchy problem in linear elasticity with material uncertainties. Computers and Mathematics with Applications, 73(3):494–504, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306824.

Feng:2011:NAC

[FPW^{+}11] Shi Feng, Jun Pang, Daling Wang, Ge Yu, Feng Yang, and Dongping Xu. A novel approach for clustering sentiments in Chinese blogs based on graph similarity. Computers and Mathematics with Applications, 62(7):2770–2778, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006146.

Fang:2018:ACM

[FQLC18] Lingling Fang, Tianshuang Qiu, Yin Liu, and Chaofeng Chen. Active contour model driven by global and local intensity information for ultrasound image segmentation. Computers and Mathematics with Applications, 75(12):4286–4299, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301652.

Fam:2015:EMT

[FR15] George S. A. Fam and Youssef F. Rashed. An efficient meshless technique for the solution of transversely isotropic twodimensional piezoelectricity. Computers and Mathematics with Applications, 69(5):438–454, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500019X.

Franz:2016:REE

[FR16] Sebastian Franz and Hans-Görg Roos. Robust error estimation in energy and balanced norms for singularly perturbed fourth order problems. Computers and Mathematics with Applications, 72(1):233–247, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302449.

Frasin:2011:NGI

[Fra11] B. A. Frasin. New general integral operator.

Computers and Mathematics with Applications, 62(11):4272–4276, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008807.

Frank:2015:FMG

[FRAK15] Florian Frank, Balthasar Reuter, Vadym Aizinger, and Peter Knabner. FES-TUNG: A MATLAB/ GNU Octave toolbox for the discontinuous Galerkin method, Part I: Diffusion operator. Computers and Mathematics with Applications, 70(1): 11–46, July 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115001820.

Fortes:2016:IFR

[FRSC16] M. A. Fortes, M. Raydan, and A. M. Sajo-Castelli. Inverse-free recursive multiresolution algorithms for a data approximation problem. Computers and Mathematics with Applications, 72(4):1177–1187, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303741.

Feichtinger:2011:PBO

[FRSW11] Anna Feichtinger, Irena Rachunková, Svatoslav Stanek, and Ewa Weinmüller. Periodic BVPs in ODEs with time singularities. Computers and Mathematics with Applications, 62(4):2058–2070, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005281.

Fu:2015:HDV

[FRZ15] Yongqiang Fu, Vicentiu D. Radulescu, and Binlin Zhang. Hodge decomposition of variable exponent spaces of Clifford-valued functions and applications to Dirac and Stokes equations. Computers and Mathematics with Applications, 70(4):691–704, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002552.

Farnoosh:2017:EFN

[FSB17] Rahman Farnoosh, Amirhossein Sobhani, and Mohammad Hossein Beheshti. Efficient and fast numerical method for pricing discrete double barrier option by projection method. Computers and Mathematics with Applications, 73(7):1539–1545, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300494.

Fuster-Sabater:2011:AGS

[FSCG11] A. Fúster-Sabater and P. Caballero-Gil. Analysis of the generalized self-shrinking generator. Computers and Mathematics with Applications, 61(4):871–880, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009405.

Fan:2010:PSR

[FSH10] Chun-I Fan, Wei-Zhe Sun, and Vincent Shi-Ming Huang. Provably secure randomized blind signature scheme based on bilinear pairing. Computers and Mathematics with Applications, 60(2):285–293, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000043X.

Feng:2011:ESS

[FSHZ11] Wenquan Feng, Shurong Sun, Zhenlai Han, and Yige Zhao. Existence of solutions for a singular system of nonlinear fractional differential equations. Computers and Mathematics with Applications, 62(3):1370–1378, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002392.

Faes:2019:HSU

[FSM19] Matthias Faes, Ghosh Dastidar Sabyasachi, and David Moens. Hybrid spatial uncertainty analysis for the estimation of imprecise failure probabilities in Laser Sintered PA-12 parts. Computers and Mathematics with Applications, 78(7):2395–2406, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304826.

Farnoosh:2015:NMD

[FSRB15] R. Farnoosh, Amirhossein Sobhani, Hamidreza Rezazadeh, and Mohammad Hossein Beheshti. Numerical method for discrete double barrier option pricing with timedependent parameters. Computers and Mathematics with Applications, 70(8):2006–2013, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003855.

Fan:2019:UGS

[FST19] Jishan Fan, Jianzhu Sun, and Tong Tang. Uniform global strong solutions of the 2D densitydependent incompressible magnetic Bénard problem in a bounded domain. Computers and Mathematics with Applications, 77(2):494–500, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122118305728.

Fan:2018:UGS

[FSTN18] Jishan Fan, Jianzhu Sun, Tong Tang, and Gen Nakamura. Uniform global solutions of the 3D compressible MHD system in a bounded domain. Computers and Mathematics with Applications, 76(11–12):2758–2766, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305388.

Fan:2017:RCG

[FSZ17] Jishan Fan, Bessem Samet, and Yong Zhou. A regularity criterion for a generalized Hall– MHD system. Computers and Mathematics with Applications, 74 (10):2438–2443, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304418.

Fan:2018:URT

[FSZ18] Jishan Fan, Bessem Samet, and Yong Zhou. Uniform regularity for a 3D time-dependent Ginzburg–Landau model in superconductivity. Computers and Mathematics with Applications, 75(9):3244–3248, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300609.

Fareed:2018:IPO

[FSZS18] Hiba Fareed, John R. Singler, Yangwen Zhang, and Jiguang Shen. Incremental proper orthogonal decomposition for PDE simulation data. Computers and Mathematics with Applications, 75(6):1942–1960, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305606.

Farzaneh:2010:GEM

[FT10] Yadollah Farzaneh and Ali Akbarzadeh Tootoonchi. Global error minimization method for solving strongly nonlinear oscillator differential equations. Computers and Mathematics with Applications, 59(8):2887–2895, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000115X.

Fikioris:2015:CIO

[FT15] G. Fikioris and N. L.

Tsitsas. On convergence and inherent oscillations within computational methods employing fictitious sources. Computers and Mathematics with Applications, 69(7):636– 649, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000565.

Fu:2019:HOH

[Fu19] Guosheng Fu. A highorder HDG method for the Biot’s consolidation model. Computers and Mathematics with Applications, 77(1):237–252, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305418.

Fuhrer:2018:SDM

[Füh18] Thomas Führer. Superconvergence in a DPG method for an ultraweak formulation. Computers and Mathematics with Applications, 75(5):1705–1718, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307551.

Furati:2013:CTP

[Fur13] Khaled M. Furati. A Cauchy-type problem involving a weighted sequential derivative in the space of integrable functions. Computers and Mathematics with Applications, 66(5):883–891, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000746.

Flint:2016:BMR

[FVVS16] Christopher Flint, George Vahala, Linda Vahala, and Min Soe. A 9-bit multiple relaxation lattice Boltzmann magnetohydrodynamic algorithm for 2D turbulence. Computers and Mathematics with Applications, 72(2):394– 403, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122115004228.

Formaggia:2018:UEF

[FVZ18] Luca Formaggia, Christian Vergara, and Stefano Zonca. Unfitted extended finite elements for composite grids. Computers and Mathematics with Applications, 76(4):893–904, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302967.

Fang:2018:ISD

[FW18] Tao Fang and Yun-Hu Wang. Interaction solutions for a dimensionally reduced Hirota bilinear equation. Computers and Mathematics with Applications, 76 (6):1476–1485, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303651.

Fang:2011:PIA

[FWFL11] Haipeng Fang, Juan Wang, Enmin Feng, and Zhijun Li. Parameter identification and application of a distributed parameter coupled system with a movable inner boundary. Computers and Mathematics with Applications, 62 (11):4015–4020, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008030.

Fankhauser:2014:AFB

[FWW14] Thomas Fankhauser, Thomas P. Wihler, and Marcel Wirz. The hp-adaptive FEM based on continuous Sobolev embeddings: isotropic refinements. Computers and Mathematics with Applications, 67(4):854– 868, March 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113003465.

Fan:2016:CPG

[FWZ16] Hong-Tao Fan, Xin Wang, and Bing Zheng. A class of preconditioned generalized local PSS iteration methods for non-Hermitian saddle point problems. Computers and Mathematics with Applications, 72(4):1188–1204, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303728.

Fan:2018:SLV

[FXC18] Congyin Fan, Kaili Xiang, and Shanzhen Chen. Stock loan valuation based on the finite moment log-stable process. Computers and Mathematics with Applications, 75(2):374–387, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305631.

Fu:2018:BTM

[FXCC18] Zhuo-Jia Fu, Qiang Xi, Wen Chen, and Alexander H.-D. Cheng. A boundary-type meshless solver for transient heat conduction analysis of slender functionally graded materials with exponential variations. Computers and Mathematics with Applications, 76(4):760–773, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302852.

Fan:2011:TWS

[FYYT11] Xinghua Fan, Shouxiang Yang, Jiuli Yin, and Lixin Tian. Traveling wave solutions to the (n + 1)-dimensional sinh-cosh-Gordon equation. Computers and Mathematics with Applications, 61(3):699–707, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009223.

Fan:2014:PGI

[FZ14] Hong-Tao Fan and Bing Zheng. A preconditioned GLHSS iteration method for non-Hermitian singular saddle point problems. Computers and Mathematics with Applications, 67(3):614–626, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006871.

Fan:2017:RCD

[FZ17] Jishan Fan and Yong Zhou. Regularity criteria for the 3D densitydependent incompressible Maxwell–Navier– Stokes system. Computers and Mathematics with Applications, 73(11):2421–2425, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302055.

Feng:2018:SSW

[FZ18] Binhua Feng and Honghong Zhang. Stability of standing waves for the fractional Schrödinger– Choquard equation. Computers and Mathematics with Applications, 75(7):2499–2507, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307885.

Fang:2019:MBM

[FZ19] Xi-Ming Fang and Zhi-Wei Zhu. The modulusbased matrix double splitting iteration method for linear complementarity problems. Computers and Mathematics with Applications, 78 (11):3633–3643, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122119303153.

Fetecau:2010:EBF

[FZBF10] C. Fetecau, J. Zierep, R. Bohning, and Corina Fetecau. On the energetic balance for the flow of an Oldroyd-b fluid due to a flat plate subject to a time-dependent shear stress. Computers and Mathematics with Applications, 60(1):74– 82, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003020.

Feng:2018:FSO

[FZL^{+}18] L. B. Feng, P. Zhuang, F. Liu, I. Turner, V. Anh, and J. Li. A fast secondorder accurate method for a two-sided spacefractional diffusion equation with variable coefficients. Computers and Mathematics with Applications, 73(6):1155–1171, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211630339X.

Gupta:2010:FDC

[GA10] Ashutosh Gupta and Suneeta Agarwal. A fast dynamic compression scheme for natural language texts. Computers and Mathematics with Applications, 60 (12):3139–3151, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007984.

Guner:2016:DMD

[GABC16] Ozkan Guner, Esin Aksoy, Ahmet Bekir, and Adem C. Cevikel. Different methods for (3 + 1)- dimensional space-time fractional modified KdV– Zakharov–Kuznetsov equation. Computers and Mathematics with Applications, 71(6):1259–1269, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300463.

See comments [Her19].

Garcia-Alonso:2013:MOE

[GACMO13] Carlos R. García-Alonso, Pilar Campoy-Muñoz, and Melania Salazar Ordoñez. A multiobjective evolutionary algorithm for enhancing Bayesian networks hybrid-based modeling. Computers and Mathematics with Applications, 66(10):1971–1980, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300059X.

Galperin:2010:ITP

[Gal10a]  E. A. Galperin.  In-

formation transmittal, principle of relativity and mass-energy relation. Computers and Mathematics with Applications, 59(8):2490–2509, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007950.

Galperin:2010:RLO

[Gal10b] E. A. Galperin. Relativistic limits of observable velocities. Computers and Mathematics with Applications, 59(8):2478–2489, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007949.

Galperin:2010:RTV

[Gal10c] E. A. Galperin. Relativistic transformations at variable velocities. Computers and Mathematics with Applications, 59(2):886–897, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006944.

Galperin:2010:UFT

[Gal10d] E. A. Galperin. Unit-free transformations of linear relativity. Computers and Mathematics with Applications, 59(1):365–375, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004428.

Gal:2011:ACS

[Gal11a] Sorin G. Gal. Approximation in compact sets by q-Stancu–Faber polynomials, q > 1. Computers and Mathematics with Applications, 61(10):3003– 3009, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002665.

Galperin:2011:HSI

[Gal11b] E. A. Galperin. Hamiltonian systems, information transmittal and special relativity. Computers and Mathematics with Applications, 61(2):211–235, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008333.

Galperin:2011:ITN

[Gal11c]  E. A. Galperin.  In-

formation transmittal, Newton’s law of gravitation, and tensor approach to general relativity. Computers and Mathematics with Applications, 62(2):709–724, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004639.

Galperin:2011:ITR

[Gal11d] E. A. Galperin. Information transmittal, relativity and gravitation. Computers and Mathematics with Applications, 61(6):1517–1535, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000368.

Galperin:2011:LTD

[Gal11e] E. A. Galperin. Left time derivatives in mathematics, mechanics and control of motion. Computers and Mathematics with Applications, 62 (12):4742–4757, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009424.

Galiano:2012:CDP

[Gal12] Gonzalo Galiano. On a cross-diffusion population model deduced from mutation and splitting of a single species. Computers and Mathematics with Applications, 64 (6):1927–1936, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002507.

Gonzalez-Andrade:2018:MOA

[GALO18] Sergio González-Andrade and Sofía López-Ordóñez. A multigrid optimization algorithm for the numerical solution of quasilinear variational inequalities involving the p-Laplacian. Computers and Mathematics with Applications, 75(4):1107–1127, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306855.

Gao:2011:SPP

[Gao11]  Yuan Gao.  Shortest

path problem with uncertain arc lengths. Computers and Mathematics with Applications, 62 (6):2591–2600, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100633X.

Gao:2012:BCD

[Gao12] Guilian Gao. Boundedness for commutators of n-dimensional rough Hardy operators on Morrey–Herz spaces. Computers and Mathematics with Applications, 64(4):544–549, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122111011011.

Gao:2015:NCC

[Gao15] Peng Gao. Null controllability with constraints on the state for the reaction–diffusion system. Computers and Mathematics with Applications, 70(5):776–788, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002722.

Gao:2017:SPT

[Gao17] Peng Gao. Some periodic type solutions for stochastic reaction– diffusion equation with cubic nonlinearities. Computers and Mathematics with Applications, 74 (10):2281–2297, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304224.

Garrappa:2013:FAE

[Gar13] Roberto Garrappa. A family of Adams exponential integrators for fractional linear systems. Computers and Mathematics with Applications, 66(5):717–727, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000394.

Gavrea:2012:ISI

[Gav12] Bogdan Gavrea. Improvement of some inequalities of Chebysev–Grüss type. Computers and Mathematics with Applications, 64(6):2003–2010, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003197.

Gonzalez-Abril:2011:SVM

[GAVOF11] L. Gonzalez-Abril, F. Velasco, J. A. Ortega, and L. Franco. Support vector machines for classification of input vectors with different metrics. Computers and Mathematics with Applications, 61(9):2874– 2878, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002343.

Gerritsma:2016:SML

[GB16] Marc Gerritsma and Pavel Bochev. A spectral mimetic least-squares method for the Stokes equations with no-slip boundary condition. Computers and Mathematics with Applications, 71(11): 2285–2300, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116300293.

Ge:2018:RCM

[GB18] Gen Ge and Zhe Bo.

Response of a cantilever model with a surface crack under basal white noise excitation. Computers and Mathematics with Applications, 76(11– 12):2728–2743, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305017.

Ganji:2011:AAT

[GBG11] S. S. Ganji, A. Barari, and D. D. Ganji. Approximate analysis of two-mass-spring systems and buckling of a column. Computers and Mathematics with Applications, 61(4):1088–1095, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009648.

Guo:2019:SGT

[GC19a] Ru-Qian Guo and Xiao-Peng Chen. Sound generation by two dimensional vortex pair motion and the influence of viscosity. Computers and Mathematics with Applications, 78(8):2761–2771, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302305.

Guo:2019:ILI

[GC19b] Xixiong Guo and Jun Cao. An IB–LBM investigation into the aerodynamic coefficients in relation to the rotation intensity of a tornado-like wind. Computers and Mathematics with Applications, 78(4):1206–1226, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304084.

Gerritsma:2017:MRL

[GCDG17] Marc Gerritsma, Carsten Carstensen, Leszek Demkowicz, and Jay Gopalakrishnan. Minimum residual and least squares finite element methods II. Computers and Mathematics with Applications, 74(8):1922, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304388.

Garmanjani:2018:RPU

[GCE18] G. Garmanjani, R. Cavoretto, and M. Esmaeilbeigi. A RBF partition of unity collocation method based on finite difference for initialboundary value problems. Computers and Mathematics with Applications, 75(11):4066–4090, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301433.

Gomez-Corral:2012:NBD

[GCG12] A. Gómez-Corral and M. López García. On the number of births and deaths during an extinction cycle, and the survival of a certain individual in a competition process. Computers and Mathematics with Applications, 64(3):236–259, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200079X.

Gelisken:2010:ABP

[GÇK10] Ali Gelisken, Cengiz Çinar, and Abdullah Selçuk Kurbanli. On the asymptotic behavior and periodic nature of a difference equation with maximum. Computers and Mathematics with Applications, 59(2):898–902, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006932.

Gao:2018:PSS

[GCR^{+}18] Xiaoyan Gao, Yongli Cai, Feng Rao, Shengmao Fu, and Weiming Wang. Positive steady states in an epidemic model with nonlinear incidence rate. Computers and Mathematics with Applications, 75(2):424–443, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305928.

Gafiychuk:2010:MMD

[GD10a] V. Gafiychuk and B. Datsko.

Mathematical modeling of different types of instabilities in time fractional reaction–diffusion systems. Computers and Mathematics with Applications, 59(3):1101–1107, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003496.

Gokaltun:2010:LBC

[GD10b] S. Gokaltun and G. S. Dulikravich. Lattice Boltzmann computations of incompressible laminar flow and heat transfer in a constricted channel. Computers and Mathematics with Applications, 59(7):2431–2441, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006348.

Gharbi:2011:AAA

[GD11] Nawel Gharbi and Claude Dutheillet. An algorithmic approach for analysis of finite-source retrial systems with unreliable servers. Computers and Mathematics with Applications, 62(6):2535–2546, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005967.

Gu:2012:OSS

[GD12] Ya Gu and Ruifeng Ding. Observable state space realizations for multivariable systems. Computers and Mathematics with Applications, 63(9):1389– 1399, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200034X.

Gupta:2016:EZS

[GD16] Varun Gupta and C. Armando Duarte. On the enrichment zone size for optimal convergence rate of the Generalized/ Extended Finite Element Method. Computers and Mathematics with Applications, 72(3):481–493, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211630236X.

Guillon:2012:NMB

[GDF12] Thomas Guillon, Yves Dumont, and Thierry Fourcaud. Numerical methods for the biomechanics of growing trees. Computers and Mathematics with Applications, 64(3):289–309, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001654.

Guerrero:2013:SMD

[GDM13] F. Guerrero, R. Donat, and P. Mulet. Solving a model for 1-D, three-phase flow vertical equilibrium processes in a homogeneous porous medium by means of a weighted essentially non oscillatory numerical scheme. Computers and Mathematics with Applications, 66(7):1284–1298, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004781.

Guo:2011:NWS

[GDZ11] Yan-Ni Guo, Qiao-Li Dong, and Zhi-Fei Zhang. Notes on weak and strong convergence theorems for a finite family of asymptotically strict pseudocontractive mappings in the intermediate sense. Computers and Mathematics with Applications, 62(4):2132–2141, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005001.

Ge:2010:ACA

[Ge10] Xun Ge. An application of covering approximation spaces on network security. Computers and Mathematics with Applications, 60(5):1191–1199, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004116.

Gemignani:2016:APR

[Gem16] L. Gemignani. Accurate polynomial root-finding methods for symmetric tridiagonal matrix eigenproblems. Computers and Mathematics with Applications, 72(4):992–1001, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303418.

Geng:2010:MVI

[Gen10] Fazhan Geng. A modified variational iteration method for solving Riccati differential equations. Computers and Mathematics with Applications, 60(7):1868–1872, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004931.

Geng:2011:NSD

[Gen11a] Fazhan Geng. Numerical solutions of Duffing equations involving both integral and non-integral forcing terms. Computers and Mathematics with Applications, 61(8):1935–1938, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005456.

Geng:2011:PVI

[Gen11b] Fazhan Geng. A piecewise variational iteration method for treating a nonlinear oscillator of a mass attached to a stretched elastic wire. Computers and Mathematics with Applications, 62(4):1641–1644, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003920.

Gepreel:2016:ESN

[Gep16] Khaled A. Gepreel. Exact solutions for nonlinear integral member of Kadomtsev–Petviashvili hierarchy differential equations using the modified (w/g)-expansion method. Computers and Mathematics with Applications, 72(9):2072–2083, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630459X.

Check title: garbled text?? Gandham:2014:GAA

[GEZ14] Rajesh Gandham, Kenneth Esler, and Yongpeng Zhang. A GPU accelerated aggregation algebraic multigrid method. Computers and Mathematics with Applications, 68(10):1151–1160, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004143.

Golmakani:2016:NTE

[GF16] M. E. Golmakani and M. N. Sadraee Far. Nonlinear thermo-elastic bending behavior of graphene sheets embedded in an elastic medium based on nonlocal elasticity theory. Computers and Mathematics with Applications, 72(3):785–805, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303546.

Guo:2019:IBV

[GF19] Xiaoyi Guo and Zunwei Fu. An initial and boundary value problem of fractional Jeffreys’ fluid in a porous half space. Computers and Mathematics with Applications, 78(6):1801–1810, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005635.

Gu:2016:RCS

[GFZ16] Weijiang Gu, Jishan Fan, and Yong Zhou. Regularity criteria for some simplified non-isothermal models for nematic liquid crystals. Computers and Mathematics with Applications, 72 (12):2839–2853, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305648.

Gaddam:2018:IDB

[GG18] Sharat Gaddam and Thirupathi Gudi. Inhomogeneous Dirichlet boundary condition in the a posteriori error control of the obstacle problem. Computers and Mathematics with Applications, 75(7):2311–2327, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307733.

Galan-Garcia:2018:PDE

[GGAV18] José L. Galán-García and Gabriel Aguilera-Venegas. Partial differential equations based simulations in multiple space dimensions. Computers and Mathematics with Applications, 74(1):1–5, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730264X.

Galan-Garcia:2019:APD

[GGAVGG19] José L. Galán-García, Gabriel Aguilera-Venegas, and María Á. Galán-García. Applications of partial differential equations in science and engineering. Computers and Mathematics with Applications, 78(9):2847–2851, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930433X.

Galan-Garcia:2019:SSF

[GGAVRC^{+}19] José Luis Galán-García, Gabriel Aguilera-Venegas, Pedro Rodríguez-Cielos, Yolanda Padilla-Domínguez, and María Ángeles Galán-García. SFOPDES: a stepwise first order partial differential equations solver with a computer algebra system. Computers and Mathematics with Applications, 78 (9):3152–3164, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302731.

Grace:2012:OTO

[GGEB12] Said R. Grace, John R. Graef, and Mohamed A. El-Beltagy. On the oscillation of third order neutral delay dynamic equations on time scales. Computers and Mathematics with Applications, 63(4):775–782, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211101025X.

Gmeiner:2013:OMC

[GGGR13] B. Gmeiner, T. Gradl, F. Gaspar, and U. Rüde. Optimization of the multigridconvergence rate on semistructured meshes by local Fourier analysis. Computers and Mathematics with Applications, 65(4):694–711, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200702X.

Gil:2017:OCA

[GGGR17] A. Gil, J. P. G. Galache, C. Godenschwager, and U. Rüde. Optimum configuration for accurate simulations of chaotic porous media with Lattice Boltzmann Methods considering boundary conditions, lattice spacing and domain size. Computers and Mathematics with Applications, 73(12):2515–2528, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301797.

Ge:2018:FDS

[GGH18] Zhihao Ge, Zhen Guan, and Yinnian He. Fully discrete stabilized multiphysics finite element method for the polymer gel model. Computers and Mathematics with Applications, 76(2):393–405, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302311.

Gerardo-Giorda:2018:CMM

[GGK18] Luca Gerardo-Giorda and Julia M. Kroos. A computational multiscale model of cortical spreading depression propagation. Computers and Mathematics with Applications, 74(5):1076–1090, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303024.

Gilbert:2013:VMR

[GGL13] Robert P. Gilbert, Philippe Guyenne, and Jing Li. A viscoelastic model for random ultrasound propagation in cancellous bone. Computers and Mathematics with Applications, 66(6):943–964, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003970.

Ganzha:2015:APM

[GGLP15] Maria Ganzha, Krassimir Georgiev, Ivan Lirkov, and Marcin Paprzycki. An application of partition method for solving 3D Stokes equation. Computers and Mathematics with Applications, 70(11):2762–2772, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002060.

Gadzhev:2013:NSA

[GGM^{+}13] Georgi K. Gadzhev, Kostadin G. Ganev, Nikolay G. Miloshev, Dimiter E. Syrakov, and Maria Prodanova. Numerical study of the atmospheric composition in Bulgaria. Computers and Mathematics with Applications, 65(3):402–422, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004683.

Guzik:2016:HPF

[GGO16] S. M. Guzik, X. Gao, and C. Olschanowsky. A high-performance finitevolume algorithm for solving partial differential equations governing compressible viscous flows on structured grids. Computers and Mathematics with Applications, 72 (9):2098–2118, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304588.

Gala:2015:RCT

[GGR15] Sadek Gala, Zhengguang Guo, and Maria Alessandra Ragusa. A regularity criterion for the three-dimensional MHD equations in terms of one directional derivative of the pressure. Computers and Mathematics with Applications, 70 (12):3057–3061, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500509X.

Gossye:2019:EMH

[GGR19] Michiel Gossye, Dries Vande Ginste, and Hendrik Rogier. Electromagnetic modeling of high magnetic contrast media using Calderón preconditioning. Computers and Mathematics with Applications, 77(6):1626–1638, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304711.

Gatica:2016:PPE

[GGS16] Gabriel N. Gatica, Luis F. Gatica, and Filánder A. Sequeira. A priori and a posteriori error analyses of a pseudostressbased mixed formulation for linear elasticity. Computers and Mathematics with Applications, 71(2):585–614, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005775.

Guillen-Gonzalez:2014:SOS

[GGT14] Francisco Guillén-González and Giordano Tierra. Second order schemes and time-step adaptivity for Allen–Cahn and Cahn–Hilliard models. Computers and Mathematics with Applications, 68(8):821–846, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003320.

Gatica:2019:FAF

[GGVRB19] Gabriel N. Gatica, Bryan Gomez-Vargas, and Ricardo Ruiz-Baier. Formulation and analysis of fully-mixed methods for stress-assisted diffusion problems. Computers and Mathematics with Applications, 77(5):1312–1330, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306618.

Gyori:2010:NVT

[GH10] I. Győri and L. Horváth. A new view of the l^{p}- theory for a system of higher order difference equations. Computers and Mathematics with Applications, 59(8):2918– 2932, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001197.

Guo:2012:EBP

[GH12a] Lifeng Guo and Lei Hu. Efficient bidirectional proxy re-encryption with direct chosen-ciphertext security. Computers and Mathematics with Applications, 63(1):151–157, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009680.

Gyori:2012:SAP

[GH12b] István Győri and László Horváth. Sharp algebraic periodicity conditions for linear higher order difference equations. Computers and Mathematics with Applications, 64(7):2262–2274, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001320.

Galantai:2013:PIS

[GH13a] A. Galántai and C. J.

Hegedüs. Perturbations of invariant subspaces of unreduced Hessenberg matrices. Computers and Mathematics with Applications, 65(3):423–434, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003379.

Gyori:2013:EPS

[GH13b] István Győri and László Horváth. Existence of periodic solutions in a linear higher order system of difference equations. Computers and Mathematics with Applications, 66(11):2239–2250, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003933.

Ghaffari:2014:OAB

[GH14] Rezvan Ghaffari and S. Mohammad Hosseini. Obtaining artificial boundary conditions for fractional sub-diffusion equation on space two-dimensional unbounded domains. Computers and Mathematics with Applications, 68(1– 2):13–26, July 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001874.

Ghaneai:2015:VIM

[GH15] H. Ghaneai and M. M. Hosseini. Variational iteration method with an auxiliary parameter for solving wave-like and heat-like equations in large domains. Computers and Mathematics with Applications, 69(5):363– 373, March 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005513.

Garrett:2016:ESH

[GH16] C. Kristopher Garrett and Cory D. Hauck.

On the eigenstructure of spherical harmonic equations for radiative transport. Computers and Mathematics with Applications, 72(2):264–270, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002850.

Guo:2018:UCO

[GH18] Yingwen Guo and Yinnian He. Unconditional convergence and optimal L^{2} error estimates of the Crank– Nicolson extrapolation FEM for the nonstationary Navier–Stokes equations. Computers and Mathematics with Applications, 75(1):134–152, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305266.

Ghasemi:2017:SBD

[Gha17]  M. Ghasemi.  Spline-

based DQM for multidimensional PDEs: Application to biharmonic and Poisson equations in 2D and 3D. Computers and Mathematics with Applications, 73(7):1576–1592, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730072X.

Ghasemi:2018:NSH

[Gha18] M. Ghasemi. On the numerical solution of high order multi-dimensional elliptic PDEs. Computers and Mathematics with Applications, 76 (5):1228–1245, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303419.

Gao:2015:ISP

[GHC15a] Qin Gao, Zhengda Huang, and Xiaoliang Cheng. Inverse spectral problem for the density of a vibrating elastic membrane. Computers and Mathematics with Applications, 70(5):980–993, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003053.

Gu:2015:HIA

[GHC^{+}15b] Xian-Ming Gu, Ting-Zhu Huang, Bruno Carpentieri, Liang Li, and Chun Wen. A hybridized iterative algorithm of the BiCORSTAB and GPBi-COR methods for solving non-Hermitian linear systems. Computers and Mathematics with Applications, 70 (12):3019–3031, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005076.

Gu:2018:ATD

[GHCZ18] Yan Gu, Xiaoqiao He, Wen Chen, and Chuanzeng Zhang. Analysis of threedimensional anisotropic heat conduction problems on thin domains using an advanced boundary element method. Computers and Mathematics with Applications, 75(1):33–44, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305229.

Gustafsson:2018:SGA

[GHL18] Tom Gustafsson, Harri Hakula, and Matti Leinonen. Stochastic Galerkin approximation of the Reynolds equation with irregular film thickness. Computers and Mathematics with Applications, 74(7):1590–1606, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303607.

Gu:2014:BTM

[GHM^{+}14] Xian-Ming Gu, Ting-Zhu Huang, Jing Meng, Tomohiro Sogabe, Hou-Biao Li, and Liang Li. BiCRtype methods for families of shifted linear systems. Computers and Mathematics with Applications, 68(7):746–758, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003472.

Glotov:2016:ICP

[GHMN16] Dmitry Glotov, Willis E. Hames, A. J. Meir, and Sedar Ngoma. An integral constrained parabolic problem with applications in thermochronology. Computers and Mathematics with Applications, 71(11):2301– 2312, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630013X.

Golev:2010:AAS

[GHR10] A. Golev, S. Hristova, and A. Rahnev. An algorithm for approximate solving of differential equations with “Maxima”. Computers and Mathematics with Applications, 60 (10):2771–2778, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007169.

Ghadampour:2015:SNA

[GHT^{+}15] Z. Ghadampour, M. R.

Hashemi, N. Talebbeydokhti, S. P. Neill, and A. H. Nikseresht. Some numerical aspects of modelling flow around hydraulic structures using incompressible SPH. Computers and Mathematics with Applications, 69(12):1470–1483, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115001704.

Georgiev:2015:NMS

[GIM15] Krassimir Georgiev, Oleg Iliev, and Peter Minev. Numerical methods for scientific computations and advanced applications. Computers and Mathematics with Applications, 70(11):2619– 2620, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005192.

Gornak:2014:FAS

[GIMZ14] T. Gornak, O. Iliev, P. Minev, and A. Zemitis. A fast algorithm for 3D simulation of thermal stratification in containment pools of nuclear power plants. Computers and Mathematics with Applications, 67(12):2228– 2239, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001345.

Gittelson:2014:HOM

[Git14] Claude Jeffrey Gittelson. High-order methods as an alternative to using sparse tensor products for stochastic Galerkin FEM. Computers and Mathematics with Applications, 67(4):888–898, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003702.

Gupta:2010:MHO

[GJ10] S. K. Gupta and Anurag Jayswal. Multiobjective higher-order symmetric duality involving generalized cone-invex functions. Computers and Mathematics with Applications, 60(12):3187–3192, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008023.

Guo:2012:IFF

[GJ12] Tian Liang Guo and Wei Jiang. Impulsive fractional functional differential equations. Computers and Mathematics with Applications, 64 (10):3414–3424, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111011102.

Gasinski:2019:MPS

[GJ19] Leszek Gasiński and João R. Santos Júnior. Multiplicity of positive solutions for an equation with degenerate nonlocal diffusion. Computers and Mathematics with Applications, 78(1):136–143, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301038.

Gunzburger:2018:AAF

[GJX18] Max Gunzburger, Nan Jiang, and Feifei Xu. Analysis and approximation of a fractional Laplacian-based closure model for turbulent flows and its connection to Richardson pair dispersion. Computers and Mathematics with Applications, 75(6):1973–2001, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303838.

Geum:2011:BFE

[GK11a] Young Hee Geum and Young Ik Kim. A biparametric family of eighthorder methods with their third-step weighting function decomposed into a one-variable linear fraction and a two-variable generic function. Computers and Mathematics with Applications, 61(3):708–714, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009259.

Geum:2011:FOS

[GK11b] Young Hee Geum and Young Ik Kim. A family of optimal sixteenthorder multipoint methods with a linear fraction plus a trivariate polynomial as the fourth-step weighting function. Computers and Mathematics with Applications, 61(11):3278– 3287, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002938.

Gazizov:2013:CES

[GK13] R. K. Gazizov and A. A.

Kasatkin. Construction of exact solutions for fractional order differential equations by the invariant subspace method. Computers and Mathematics with Applications, 66(5):576–584, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002915.

Grabski:2016:ACF

[GK16] Jakub Krzysztof Grabski and Jan Adam Kolodziej. Analysis of Carreau fluid flow between corrugated plates. Computers and Mathematics with Applications, 72(6):1501–1514, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303960.

Grabski:2018:LFF

[GK18] Jakub Krzysztof Grabski and Jan Adam Kołodziej. Laminar fluid flow and heat transfer in an internally corrugated tube by means of the method of fundamental solutions and radial basis functions. Computers and Mathematics with Applications, 75(4):1413–1433, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307113.

Griebel:2019:SMS

[GK19] M. Griebel and M. Klitz.

Simulation of micronscale drop impact. Computers and Mathematics with Applications, 78 (9):3027–3043, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301762.

Gordji:2011:SCJ

[GKK11] M. Eshaghi Gordji, H. Khodaei, and M. Kamyar. Stability of Cauchy– Jensen type functional equation in generalized fuzzy normed spaces. Computers and Mathematics with Applications, 62(8):2950–2960, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100647X.

Gine:2011:LCL

[GKLR11] Jaume Giné, Zhibek Kadyrsizova, Yirong Liu, and Valery G. Romanovski. Linearizability conditions for Lotka–Volterra planar complex quartic systems having homogeneous nonlinearities. Computers and Mathematics with Applications, 61(4):1190–1201, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009855.

Graef:2011:HOB

[GKM11] John R. Graef, Lingju Kong, and Feliz M. Minhós. Higher order boundary value problems with ϕ-Laplacian and functional boundary conditions. Computers and Mathematics with Applications, 61(2):236–249, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008345.

Gupta:2010:HOC

[GKS10] S. K. Gupta, N. Kailey, and M. K. Sharma. Higher-order (F, α, ρ, d)- convexity and symmetric duality in multiobjective programming. Computers and Mathematics with Applications, 60(8):2373–2381, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000595X.

Gonzalez:2017:IAT

[GKS17] Gerardo González, Ville Kolehmainen, and Aku Seppänen. Isotropic and anisotropic total variation regularization in electrical impedance tomography. Computers and Mathematics with Applications, 74(3):564–576, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302833.

Gadella:2010:AMS

[GL10] M. Gadella and L. P.

Lara. An algebraic method to solve the radial Schrödinger equation. Computers and Mathematics with Applications, 60(9):2701–2711, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007091.

Guo:2016:USA

[GL16] Geyang Guo and Shujuan Lü. Unconditional stability of alternating difference schemes with intrinsic parallelism for two-dimensional fourthorder diffusion equation. Computers and Mathematics with Applications, 71(10):1944–1959, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301201.

Ganesan:2017:GFE

[GL17a] Sashikumaar Ganesan and Shangerganesh Lingeshwaran. Galerkin finite element method for cancer invasion mathematical model. Computers and Mathematics with Applications, 73(12):2603–2617, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302274.

Gong:2017:ESA

[GL17b] Yanping Gong and Shulin Liang. Existence of solutions for asymptotically periodic fractional Schrödinger equation.

Computers and Mathematics with Applications, 74(12):3175–3182, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305175.

Guo:2014:HBS

[GLL14] Gaihui Guo, Bingfang Li, and Xiaolin Lin. Hopf bifurcation in spatially homogeneous and inhomogeneous autocatalysis models. Computers and Mathematics with Applications, 67(1):151–163, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005087.

Gao:2019:NED

[GLLC19] Xiao-Wei Gao, Hua-Yu Liu, Jun Lv, and Miao Cui. A novel element differential method for solid mechanical problems using isoparametric triangular and tetrahedral elements. Computers and Mathematics with Applications, 78 (11):3563–3585, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302895.

Guo:2011:EDA

[GLM^{+}11] Hua Guo, Zhoujun Li, Yi Mu, Fan Zhang, Chuankun Wu, and Jikai Teng. An efficient dynamic authenticated key exchange protocol with selectable identities. Computers and Mathematics with Applications, 61(9):2518–2527, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001404.

Garg:2018:CCI

[GLP18] Deepak Garg, Antonella Longo, and Paolo Papale. Computation of compressible and incompressible flows with a space–time stabilized finite element method. Computers and Mathematics with Applications, 75(12):4272–4285, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301640.

Gala:2013:LIR

[GLR13] Sadek Gala, Qiao Liu, and Maria Alessandra Ragusa. Logarithmically improved regularity criterion for the nematic liquid crystal flows in Ḃ_{∞,∞}^{−1} space. Computers and Mathematics with Applications, 65(11):1738– 1745, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001909.

Gao:2013:LBSa

[GLW13] Hui Gao, Hui Li, and Lian-Ping Wang. Lattice Boltzmann simulation of turbulent flow laden with finite-size particles. Computers and Mathematics with Applications, 65(2):194–210, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005074.

Guo:2018:TFD

[GLW18] Xu Guo, Yutian Li, and Hong Wang. Tempered fractional diffusion equations for pricing multi-asset options under CGMYe process.

Computers and Mathematics with Applications, 76(6):1500–1514, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303675.

Guo:2018:NIF

[GLZ18] Ruchi Guo, Tao Lin, and Xu Zhang. Nonconforming immersed finite element spaces for elliptic interface problems. Computers and Mathematics with Applications, 75(6):2002–2016, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306983.

Gurefe:2011:EFM

[GM11] Yusuf Gurefe and Emine Misirli. Exp-function method for solving nonlinear evolution equations with higher order nonlinearity. Computers and Mathematics with Applications, 61(8):2025–2030, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110006346.

Goyal:2014:FAD

[GM14a] Kavita Goyal and Mani Mehra. A fast adaptive diffusion wavelet method for Burger’s equation. Computers and Mathematics with Applications, 68(4):568–577, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002661.

Grunfeld:2014:TSD

[GM14b] C. P. Grünfeld and D. Marinescu. On a time and space discretized approximation of the Boltzmann equation in the whole space. Computers and Mathematics with Applications, 68 (10):1393–1408, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004556.

Guin:2014:EPR

[GM14c] Lakshmi Narayan Guin and Prashanta Kumar Mandal. Effect of prey refuge on spatiotemporal dynamics of the reaction– diffusion system. Computers and Mathematics with Applications, 68 (10):1325–1340, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004313.

Ganesh:2018:HOF

[GM18a] M. Ganesh and C. Morgenstern. High-order FEM domain decomposition models for highfrequency wave propagation in heterogeneous media. Computers and Mathematics with Applications, 75(6):1961–1972, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306995.

Gorajski:2018:ETS

[GM18b] Mariusz Górajski and Dominika Machowska. The effects of technological shocks in an optimal goodwill model with a random product life cycle. Computers and Mathematics with Applications, 76(4):905–922, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303067.

Gudi:2019:CDG

[GM19] Thirupathi Gudi and Papri Majumder. Conforming and discontinuous Galerkin FEM in space for solving parabolic obstacle problem. Computers and Mathematics with Applications, 78 (12):3896–3915, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303256.

Gomez-Miralles:2012:VIF

[GMAM12] Luis Gómez-Miralles and Joan Arnedo-Moreno. Versatile iPad forensic acquisition using the Apple camera connection kit. Computers and Mathematics with Applications, 63(2):544–553, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008315.

Gleska:2012:ABS

[GMB12] Alina Gleska and Ewa Magnucka-Blandzi. On the asymptotic behavior of a system of two rational difference equations. Computers and Mathematics with Applications, 64(7):2345–2353, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004324.

Goh:2011:NMU

[GMI11] Joan Goh, Ahmad Abd. Majid, and Ahmad Izani Md. Ismail. Numerical method using cubic B-spline for the heat and wave equation. Computers and Mathematics with Applications, 62 (12):4492–4498, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008923.

Goh:2012:QSS

[GMI12] Joan Goh, Ahmad Abd. Majid, and Ahmad Izani Md. Ismail. A quartic Bspline for second-order singular boundary value problems. Computers and Mathematics with Applications, 64(2):115– 120, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000326.

Gao:2017:GFE

[GML17a] Yali Gao, Liquan Mei, and Rui Li. Galerkin finite element methods for the generalized Klein–Gordon–Zakharov equations. Computers and Mathematics with Applications, 74 (10):2466–2484, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304571.

Guo:2017:EGS

[GML17b] Shimin Guo, Liquan Mei, and Ying Li. An efficient Galerkin spectral method for two-dimensional fractional nonlinear reaction– diffusion-wave equation. Computers and Mathematics with Applications, 74(10):2449–2465, November 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211730439X.

Garrappa:2018:TFS

[GMP18] Roberto Garrappa, Igor Moret, and Marina Popolizio. On the time-fractional Schrödinger equation: Theoretical analysis and numerical solution by matrix Mittag-Leffler functions. Computers and Mathematics with Applications, 74(5):977–992, September 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306605.

Gopalakrishnan:2015:TPS

[GMS15] Jay Gopalakrishnan, Peter Monk, and Paulina Sepúlveda. A tent pitching scheme motivated by Friedrichs theory. Computers and Mathematics with Applications, 70 (5):1114–1135, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003351.

Garvie:2018:FEA

[GMS18] Marcus R. Garvie, Jeff Morgan, and Vandana Sharma. Finite element approximation of a spatially structured metapopulation PDE model.

Computers and Mathematics with Applications, 74(5):934–947, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306939.

Guo:2015:CEM

[GMZ15] Shimin Guo, Liquan Mei, and Yubin Zhou. The compound (\frac{G′}{G})- expansion method and double non-traveling wave solutions of (2 + 1)- dimensional nonlinear partial differential equations. Computers and Mathematics with Applications, 69(8):804–816, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000693.

Guseinov:2011:AAC

[GN11] Khalik G. Guseinov and Ali S. Nazlipinar. An algorithm for approximate calculation of the attainable sets of the nonlinear control systems with integral constraint on controls. Computers and Mathematics with Applications, 62(4):1887–1895, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005128.

Golestani:2012:CSK

[GN12] M. Golestani and S. Nobakhtian. Convexificators and strong Kuhn–Tucker conditions. Computers and Mathematics with Applications, 64(4):550–557, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011035.

Grote:2019:AEM

[GN19] Marcus J. Grote and Uri Nahum. Adaptive eigenspace for multiparameter inverse scattering problems. Computers and Mathematics with Applications, 77(12):3264–3280, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300732.

Gudi:2014:IPM

[GNP14] Thirupathi Gudi, Neela Nataraj, and Kamana Porwal. An interior penalty method for distributed optimal control problems governed by the biharmonic operator. Computers and Mathematics with Applications, 68(12):2205–2221, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004040.

Gonzalez-Olivares:2011:CDA

[GOGYL^{+}11] Eduardo González-Olivares, Betsabé González-Yañez, Jaime Mena Lorca, Alejandro Rojas-Palma, and José D. Flores. Consequences of double Allee effect on the number of limit cycles in a predator– prey model. Computers and Mathematics with Applications, 62 (9):3449–3463, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007450.

Golse:2013:BEE

[Gol13] François Golse. From the Boltzmann equation to the Euler equations in the presence of boundaries. Computers and Mathematics with Applications, 65(6):815–830, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200123X.

Goncalves:2013:LCG

[Gon13] M. L. N. Gonçalves. Local convergence of the Gauss–Newton method for injective-overdetermined systems of equations under a majorant condition. Computers and Mathematics with Applications, 66(4):490–499, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003210.

Goodrich:2010:CSD

[Goo10] Christopher S. Goodrich. Continuity of solutions to discrete fractional initial value problems. Computers and Mathematics with Applications, 59(11):3489–3499, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002294.

Goodrich:2011:EUS

[Goo11a] Christopher S. Goodrich.

Existence and uniqueness of solutions to a fractional difference equation with nonlocal conditions. Computers and Mathematics with Applications, 61(2):191–202, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000831X.

Goodrich:2011:EPS

[Goo11b] Christopher S. Goodrich.

Existence of a positive solution to systems of differential equations of fractional order. Computers and Mathematics with Applications, 62(3):1251–1268, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001386.

Gorban:2013:MMA

[Gor13] A. N. Gorban. Maxallent: Maximizers of all entropies and uncertainty of uncertainty. Computers and Mathematics with Applications, 65(10):1438–1456, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000217.

Gosse:2010:EBL

[Gos10] Laurent Gosse. Effective band-limited extrapolation relying on Slepian series and ℓ^{1} regularization. Computers and Mathematics with Applications, 60(5):1259–1279, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004207.

Gatica:2018:PPEb

[GOS18] Luis F. Gatica, Ricardo Oyarzúa, and Nestor Sánchez. A priori and a posteriori error analysis of an augmented mixed-FEM for the Navier– Stokes–Brinkman problem. Computers and Mathematics with Applications, 75(7):2420–2444, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307927.

Gong:2019:SVC

[GOT19] Jiaming Gong, Nobuyuki Oshima, and Yutaka Tabe. Spurious velocity from the cutoff and magnification equation in free energy-based LBM for two-phase flow with a large density ratio. Computers and Mathematics with Applications, 78(4):1166–1181, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305028.

Garrappa:2011:GET

[GP11] Roberto Garrappa and Marina Popolizio. Generalized exponential time differencing methods for fractional order problems. Computers and Mathematics with Applications, 62(3):876–890, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003579.

Garg:2019:EFE

[GP19] Vikram V. Garg and Serge Prudhomme. Enhanced functional evaluation for the finite element penalty method. Computers and Mathematics with Applications, 78 (12):3821–3840, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303049.

Guirao:2011:MDC

[GPV11] Juan L. G. Guirao, Fernando L. Pelayo, and Jose C. Valverde. Modeling the dynamics of concurrent computing systems. Computers and Mathematics with Applications, 61(5):1402–1406, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000137.

See notes [PV12].

Gao:2010:TDM

[GQF^{+}10] Hui Gao, Charmaine Q. Qiu, Dimin Fan, Yan Jin, and Lian-Ping Wang. Three-dimensional microscale flow simulation and colloid transport modeling in saturated soil porous media. Computers and Mathematics with Applications, 59(7):2271–2289, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006464.

Ghasemi:2010:FVL

[GR10] J. Ghasemi and S. E.

Razavi. On the finitevolume lattice Boltzmann modeling of thermohydrodynamics. Computers and Mathematics with Applications, 60 (5):1135–1144, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003603.

Gyori:2012:P

[GR12] István Győri and David W.

Reynolds. Preface. Computers and Mathematics with Applications, 64(7):2159, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005470.

Ghandehari:2013:NMS

[GR13a] Mohammad Ali Mohebbi Ghandehari and Mojtaba Ranjbar. A numerical method for solving a fractional partial differential equation through converting it into an NLP problem. Computers and Mathematics with Applications, 65(7):975–982, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000205.

Groza:2013:TSM

[GR13b] Ghiocel Groza and Mohsen Razzaghi. A Taylor series method for the solution of the linear initialboundary-value problems for partial differential equations. Computers and Mathematics with Applications, 66(7):1329–1343, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004859.

Guo:2015:WDB

[GR15] Yujie Guo and Martin Ruess. Weak Dirichlet boundary conditions for trimmed thin isogeometric shells. Computers and Mathematics with Applications, 70(7):1425–1440, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002989.

Gidey:2019:OSM

[GR19a] H. H. Gidey and B. D.

Reddy. Operator-splitting methods for the 2D convective Cahn–Hilliard equation. Computers and Mathematics with Applications, 77 (12):3128–3153, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300537.

Goncharsky:2019:MSC

[GR19b] Alexander V. Goncharsky and Sergey Y. Romanov. A method of solving the coefficient inverse problems of wave tomography. Computers and Mathematics with Applications, 77(4):967–980, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306308.

Gatica:2016:PEA

[GRBT16] Gabriel N. Gatica, Ricardo Ruiz-Baier, and Giordano Tierra. A posteriori error analysis of an augmented mixed method for the Navier–Stokes equations with nonlinear viscosity. Computers and Mathematics with Applications, 72(9):2289–2310, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304862.

Grmela:2013:RTM

[Grm13] Miroslav Grmela. Role of thermodynamics in multiscale physics. Computers and Mathematics with Applications, 65(10): 1457–1470, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112006803.

Grove:2019:SCT

[Gro19] John W. Grove. Some comments on thermodynamic consistency for equilibrium mixture equations of state. Computers and Mathematics with Applications, 78(2):582– 597, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211830141X.

Ghorbanian:2012:SOF

[GRS12] V. Ghorbanian, Sh. Rezapour, and N. Shahzad. Some ordered fixed point results and the property (p). Computers and Mathematics with Applications, 63(9):1361–1368, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011424.

Ghattassi:2018:EUT

[GRS18] Mohamed Ghattassi, Jean Rodolphe Roche, and Didier Schmitt. Existence and uniqueness of a transient state for the coupled radiativeconductive heat transfer problem. Computers and Mathematics with Applications, 75(11):3918–3928, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301317.

Gasser:2014:OCT

[GRW14] Ingenuin Gasser, Martin Rybicki, and Winnifried Wollner. Optimal control of the temperature in a catalytic converter. Computers and Mathematics with Applications, 67(8):1521–1544, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000686.

Golbabai:2011:ATD

[GS11a] A. Golbabai and K. Sayevand. Analytical treatment of differential equations with fractional coordinate derivatives. Computers and Mathematics with Applications, 62(3):1003–1012, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002100.

Golbabai:2011:FCN

[GS11b] A. Golbabai and K. Sayevand. Fractional calculus — a new approach to the analysis of generalized fourth-order diffusionwave equations. Computers and Mathematics with Applications, 61(8):2227– 2231, April 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S089812211000708X.

Gupta:2011:HPM

[GS11c] Praveen Kumar Gupta and Mithilesh Singh. Homotopy perturbation method for fractional Fornberg–Whitham equation. Computers and Mathematics with Applications, 61(2):250–254, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008357.

Gil:2012:CRZ

[GS12] Amparo Gil and Javier Segura. Computing the real zeros of cylinder functions and the roots of the equation xC′_{ν}(x) + γC_{ν}(x) = 0. Computers and Mathematics with Applications, 64(1):11– 21, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001460.

Galiano:2015:ASD

[GS15a] Gonzalo Galiano and Virginia Selgas. Analysis of a splitting-differentiation population model leading to cross-diffusion. Computers and Mathematics with Applications, 70 (12):2933–2945, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004873.

Guan:2015:SNM

[GS15b] Hongbo Guan and Dongyang Shi. A stable nonconforming mixed finite element scheme for elliptic optimal control problems. Computers and Mathematics with Applications, 70(3):236–243, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001741.

Gatica:2018:PPEa

[GS18] Luis F. Gatica and Filánder A. Sequeira. A priori and a posteriori error analyses of an HDG method for the Brinkman problem. Computers and Mathematics with Applications, 75(4):1191–1212, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211730696X.

Guan:2019:ENO

[GS19] Hongbo Guan and Dongyang Shi. An efficient NFEM for optimal control problems governed by a bilinear state equation. Computers and Mathematics with Applications, 77(7):1821–1827, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306709.

Grinstein:2019:EOS

[GSD^{+}19] F. F. Grinstein, J. A.

Saenz, J. C. Dolence, T. O. Masser, R. M. Rauenzahn, and M. M. Francois. Effects of operator splitting and low Mach-number correction in turbulent mixing transition simulations. Computers and Mathematics with Applications, 78(2):437–458, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302700.

Shang:2010:DFS

[gShYL10] You guang Shang, Xue hai Yuan, and E. Stanley Lee. The n-dimensional fuzzy sets and Zadeh fuzzy sets based on the finite valued fuzzy sets. Computers and Mathematics with Applications, 60(3):442–463, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003263.

Galati:2019:EVN

[GSI19] Manuela Galati, Anders Snis, and Luca Iuliano. Experimental validation of a numerical thermal model of the EBM process for Ti_{6}Al_{4}V. Computers and Mathematics with Applications, 78(7):2417–2427, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303857.

Geier:2015:CLB

[GSPK15] Martin Geier, Martin Schönherr, Andrea Pasquali, and Manfred Krafczyk. The cumulant lattice Boltzmann equation in three dimensions: Theory and validation. Computers and Mathematics with Applications, 70(4):507– 547, August 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115002126.

Gonzalez:2014:GOD

[GSR14] Luis González, Antonio Suárez, and Eduardo Rodríguez. A generalization of the optimal diagonal approximate inverse preconditioner. Computers and Mathematics with Applications, 66(12):2433–2445, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006111.

Guerra:2011:OPS

[GSS11] Maria Letizia Guerra, Laerte Sorini, and Luciano Stefanini. Option price sensitivities through fuzzy numbers. Computers and Mathematics with Applications, 61(3):515–526, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008965.

Gibelli:2010:IEW

[GSY10] L. Gibelli, B. D. Shizgal, and A. W. Yau. Ion energization by wave-particle interactions: Comparison of spectral and particle simulation solutions of the Vlasov equation. Computers and Mathematics with Applications, 59(8):2566–2581, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000386.

Guo:2011:VIM

[GSZ11] Zhaoli Guo, Baochang Shi, and Chuguang Zheng. Velocity inversion of micro cylindrical Couette flow: a lattice Boltzmann study. Computers and Mathematics with Applications, 61(12):3519– 3527, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000441.

Giani:2014:PEE

[GSZ14] Stefano Giani, Dominik Schötzau, and Liang Zhu. An a-posteriori error estimate for hp-adaptive DG methods for convection– diffusion problems on anisotropically refined meshes. Computers and Mathematics with Applications, 67(4):869–887, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200644X.

Gupta:2015:CRP

[GT15] Vinay Kumar Gupta and Manuel Torrilhon. Comparison of relaxation phenomena in binary gas-mixtures of Maxwell molecules and hard spheres. Computers and Mathematics with Applications, 70(1):73– 88, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002096.

Gupta:2016:RCR

[GT16] Vinay Kumar Gupta and Manuel Torrilhon. Reprint of: Comparison of relaxation phenomena in binary gas-mixtures of Maxwell molecules and hard spheres. Computers and Mathematics with Applications, 72(2):271– 287, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302000.

Gao:2018:EGS

[GTC18] Zu Gao, Xianhua Tang, and Sitong Chen. Existence of ground state solutions of Nehari– Pohozaev type for fractional Schrödinger–Poisson systems with a general potential. Computers and Mathematics with Applications, 75(2):614–631, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306181.

Geiser:2011:HOO

[GTG11] Jürgen Geiser, Gamze Tanoğlu, and Nurcan Gücüyenen. Higher order operator splitting methods via Zassenhaus product formula: Theory and applications. Computers and Mathematics with Applications, 62(4):1994–2015, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005232.

Ghimire:2016:NSE

[GTL16] B. Khatri Ghimire, H. Y.

Tian, and A. R. Lamichhane. Numerical solutions of elliptic partial differential equations using Chebyshev polynomials. Computers and Mathematics with Applications, 72(4):1042–1054, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303443.

Guo:2019:ISS

[GTZ19] Ding Guo, Shou-Fu Tian, and Tian-Tian Zhang. Integrability, soliton solutions and modulation instability analysis of a (2+ 1)-dimensional nonlinear Heisenberg ferromagnetic spin chain equation. Computers and Mathematics with Applications, 77(3):770–778, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830614X.

Guessab:2013:GBC

[Gue13] Allal Guessab. Generalized barycentric coordinates and approximations of convex functions on arbitrary convex polytopes. Computers and Mathematics with Applications, 66(6):1120–1136, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004422.

Geller:2013:TJC

[GUK13] S. Geller, S. Uphoff, and M. Krafczyk. Turbulent jet computations based on MRT and cascaded lattice Boltzmann models. Computers and Mathematics with Applications, 65(12):1956–1966, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002198.

Guo:2012:COI

[Guo12] Tian Liang Guo. Controllability and observability of impulsive fractional linear time-invariant system. Computers and Mathematics with Applications, 64(10):3171–3182, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122112001344.

Gupta:2011:AAS

[Gup11] Praveen Kumar Gupta.

Approximate analytical solutions of fractional Benney–Lin equation by reduced differential transform method and the homotopy perturbation method. Computers and Mathematics with Applications, 61(9):2829–2842, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002203.

Gurski:2013:SCN

[Gur13] K. F. Gurski. A simple construction of nonstandard finite-difference schemes for small nonlinear systems applied to SIR models. Computers and Mathematics with Applications, 66 (11):2165–2177, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005142.

Gonzalez:2011:SAM

[GV11a] Eduardo González and Marcelo J. Villena. Spatial attrition modeling: stability conditions for a 2D + t FD formulation. Computers and Mathematics with Applications, 61(11):3246– 3257, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002963.

Guan:2011:SAS

[GV11b] Yun Guan and Jan Verschelde. Sampling algebraic sets in local intrinsic coordinates. Computers and Mathematics with Applications, 62 (10):3706–3721, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007814.

Galiano:2018:CDS

[GV18] Gonzalo Galiano and Julián Velasco. On a cross-diffusion system arising in image denoising. Computers and Mathematics with Applications, 76 (5):984–996, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303110.

Garcia:2013:CBP

[GVJ13] Flavio D. Garcia, Eric R.

Verheul, and Bart Jacobs. Cell-based privacyfriendly roadpricing. Computers and Mathematics with Applications, 65 (5):774–785, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112004853.

Garate:2012:SHE

[GVSP12] G. Garate, E. G. Vadillo, J. Santamaria, and D. Pardo. Solution of the 3D-Helmholtz equation in exterior domains using spherical harmonic decomposition. Computers and Mathematics with Applications, 64(8):2520–2543, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004403.

Gao:2012:TSS

[GW12a] Chaobang Gao and Jiajin Wen. Theory of surround system and associated inequalities. Computers and Mathematics with Applications, 63(12):1621– 1640, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002428.

Gelenbe:2012:LSS

[GW12b] Erol Gelenbe and Fang-Jing Wu. Large scale simulation for human evacuation and rescue. Computers and Mathematics with Applications, 64 (12):3869–3880, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002611.

Guo:2012:IPF

[GW12c] Tian Liang Guo and Jiang Wei. Impulsive problems for fractional differential equations with boundary value conditions. Computers and Mathematics with Applications, 64 (10):3281–3291, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001204.

Gao:2015:AOS

[GW15a] Wenwu Gao and Zongmin Wu. Approximation orders and shape preserving properties of the multiquadric trigonometric B-spline quasiinterpolant. Computers and Mathematics with Applications, 69(7):696– 707, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000619.

Guo:2015:NSN

[GW15b] Jianqiang Guo and Wansheng Wang. On the numerical solution of nonlinear option pricing equation in illiquid markets. Computers and Mathematics with Applications, 69(2):117–133, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005835.

Guu:2011:MOO

[GWL11] Sy-Ming Guu, Yan-Kuen Wu, and E. S. Lee. Multiobjective optimization with a max-t-norm fuzzy relational equation constraint. Computers and Mathematics with Applications, 61(6):1559–1566, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100040X.

Gaedtke:2018:ALB

[GWR^{+}18] Maximilian Gaedtke, Simon Wachter, Matthias Rädle, Hermann Nirschl, and Mathias J. Krause. Application of a lattice Boltzmann method combined with a Smagorinsky turbulence model to spatially resolved heat flux inside a refrigerated vehicle. Computers and Mathematics with Applications, 76 (10):2315–2329, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830436X.

Gao:2011:AMQ

[GWZ11] Qinjiao Gao, Zongmin Wu, and Shenggang Zhang. Applying multiquadric quasi-interpolation for boundary detection. Computers and Mathematics with Applications, 62(12):4356–4361, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008509.

Gong:2010:BSS

[GXZ10] Ke Gong, Zhi Xiao, and Xia Zhang. The bijective soft set with its operations. Computers and Mathematics with Applications, 60(8):2270–2278, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000581X.

Gong:2011:CED

[GXZ11] Ke Gong, Zhi Xiao, and Xia Zhang. Corrigendum to “Exclusive disjunctive soft set” [Comput. Math. Appl. 59 (2010) 2128–2137]. Computers and Mathematics with Applications, 61(7):1906, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000915.

See [XGXZ10].

Ghoreishi:2011:EST

[GY11] F. Ghoreishi and S. Yazdani. An extension of the spectral Tau method for numerical solution of multi-order fractional differential equations with convergence analysis. Computers and Mathematics with Applications, 61(1):30–43, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008060.

Gorban:2013:GC

[GY13] A. N. Gorban and G. S. Yablonsky. Grasping complexity. Computers and Mathematics with Applications, 65(10): 1421–1426, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113002411.

Gan:2015:SFV

[GY15] Xiao-Ting Gan and Jun-Feng Yin. Symmetric finite volume element approximations of second order linear hyperbolic integro-differential equations. Computers and Mathematics with Applications, 70 (10):2589–2600, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004599.

Gao:2011:NCF

[GYH11] Changliang Gao, Yongjian Yang, and Boshun Han. A new class of filled functions with one parameter for global optimization. Computers and Mathematics with Applications, 62(6):2393–2403, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003944.

Gandomi:2012:CES

[GYTD12] Amir Hossein Gandomi, Xin-She Yang, Siamak Talatahari, and Suash Deb. Coupled eagle strategy and differential evolution for unconstrained and constrained global optimization. Computers and Mathematics with Applications, 63(1):191–200, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100976X.

Grahovac:2010:MHM

[GZ10] N. M. Grahovac and M. M. Zigić. Modelling of the hamstring muscle group by use of fractional derivatives. Computers and Mathematics with Applications, 59(5):1695–1700, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812210900546X.

Gouia-Zarrad:2014:ARF

[GZ14] Rim Gouia-Zarrad. Analytical reconstruction formula for n-dimensional conical Radon transform. Computers and Mathematics with Applications, 68(9):1016–1023, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001795.

Gu:2018:DBI

[GZD18] Jiayue Gu, Yong Zhang, and Huanhe Dong. Dynamic behaviors of interaction solutions of (3+1)- dimensional shallow water wave equation. Computers and Mathematics with Applications, 76 (6):1408–1419, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303584.

Guo:2019:DRV

[GZN19] Boling Guo, Lan Zeng, and Guoxi Ni. Decay rates for the viscous incompressible MHD equations with and without surface tension. Computers and Mathematics with Applications, 77(12):3224–3249, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300768.

Zhang:2012:NCA

[gZnZpZbD12] De gan Zhang, Ya nan Zhu, Chen peng Zhao, and Wen bo Dai. A new constructing approach for a weighted topology of wireless sensor networks based on local-world theory for the Internet of Things (IOT). Computers and Mathematics with Applications, 64 (5):1044–1055, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002283.

Gao:2013:LBSb

[GZR^{+}13] Yuan Gao, Xiaoxian Zhang, Pratp Rama, Rui Chen, Hossein Ostadi, and Kyle Jiang. Lattice Boltzmann simulation of water and gas flow in porous gas diffusion layers in fuel cells reconstructed from microtomography. Computers and Mathematics with Applications, 65(6):891– 900, March 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200538X.

Guo:2018:NZI

[GZW^{+}18] Min Guo, Yu Zhang, Man Wang, Yaodeng Chen, and Hongwei Yang. A new ZK–ILW equation for algebraic gravity solitary waves in finite depth stratified atmosphere and the research of squall lines formation mechanism. Computers and Mathematics with Applications, 75(10):3589–3603, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301007.

Gao:2016:RBM

[GZZ^{+}16] Li-Na Gao, Xue-Ying Zhao, Yao-Yao Zi, Jun Yu, and Xing Lü. Resonant behavior of multiple wave solutions to a Hirota bilinear equation. Computers and Mathematics with Applications, 72(5):1225–1229, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303406.

Huang:2010:MLC

[HA10] Pengzhan Huang and Abdurishit Abduwali. The modified local Crank– Nicolson method for one- and two-dimensional Burgers’ equations. Computers and Mathematics with Applications, 59(8): 2452–2463, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109006567.

Hussain:2011:SPS

[HA11] Sabir Hussain and Bashir Ahmad. Some properties of soft topological spaces.

Computers and Mathematics with Applications, 62(11):4058–4067, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008297.

Hesameddini:2016:NSG

[HA16a] Esmail Hesameddini and Elham Asadolahifard. A new spectral Galerkin method for solving the two dimensional hyperbolic telegraph equation. Computers and Mathematics with Applications, 72(7):1926–1942, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304576.

Hicdurmaz:2016:SNM

[HA16b] Betul Hicdurmaz and Allaberen Ashyralyev. A stable numerical method for multidimensional time fractional Schrödinger equations. Computers and Mathematics with Applications, 72 (6):1703–1713, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304412.

Hajarian:2018:LSS

[HA18a] Masoud Hajarian and Hassan Abbas. Least squares solutions of quadratic inverse eigenvalue problem with partially bisymmetric matrices under prescribed submatrix constraints. Computers and Mathematics with Applications, 76(6):1458–1475, September 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118303626.

Hosseini:2018:SSN

[HA18b] S. M. Hosseini and Zohreh Asgari. Solution of stochastic nonlinear time fractional PDEs using polynomial chaos expansion combined with an exponential integrator. Computers and Mathematics with Applications, 73(6):997–1007, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304230.

Heuze:2014:BTBa

[HAESLB14] Thomas Heuzé, Hussein Amin-El-Sayed, Jean-Baptiste Leblond, and Jean-Michel Bergheau. Benchmark tests based on the Couette viscometer — II: Thermo-elasto-plastic solid behaviour in small and large strains. Computers and Mathematics with Applications, 67(8):1482–1496, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000819.

Haidar:2010:GNI

[Hai10] Nassar H. S. Haidar. A geometric note on integration. Computers and Mathematics with Applications, 59(9):3130–3136, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001550.

Hajarian:2018:CSS

[Haj18a] Masoud Hajarian. Computing symmetric solutions of general Sylvester matrix equations via Lanczos version of biconjugate residual algorithm. Computers and Mathematics with Applications, 76(4):686–700, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302724.

Hajarian:2018:PCD

[Haj18b] Masoud Hajarian. Periodic conjugate direction algorithm for symmetric periodic solutions of general coupled periodic matrix equations. Computers and Mathematics with Applications, 75(11):4151–4178, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830155X.

Hajarian:2018:SCQ

[Haj18c] Masoud Hajarian. Solving constrained quadratic inverse eigenvalue problem via conjugate direction method. Computers and Mathematics with Applications, 76 (10):2384–2401, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118304607.

Hakula:2014:BLM

[Hak14]  Harri Hakula.  hp-

boundary layer mesh sequences with applications to shell problems. Computers and Mathematics with Applications, 67(4):899–917, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001594.

Haltmeier:2013:ICM

[Hal13] Markus Haltmeier. Inversion of circular means and the wave equation on convex planar domains. Computers and Mathematics with Applications, 65(7):1025–1036, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000667.

Hamza:2010:RS

[Ham10] Alaa E. Hamza. On the recursive sequence x_{n+1} = (α−βx_{n−k})/g(x_{n}, x_{n−1}, …, x_{n−k+1}). Computers and Mathematics with Applications, 60(7):2170–2177, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005687.

Hazarika:2011:SNS

[Haz11] Bipan Hazarika. Some new sequence spaces of fuzzy numbers defined by Orlicz functions using a fuzzy metric. Computers and Mathematics with Applications, 61(9):2762– 2769, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001945.

Hernandez:2012:DAC

[HB12a] Hugo Hernández and Christian Blum. Distributed ant colony optimization for minimum energy broadcasting in sensor networks with realistic antennas. Computers and Mathematics with Applications, 64 (12):3683–3700, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001605.

Herzallah:2012:EPM

[HB12b] Mohamed A. E. Herzallah and Dumitru Baleanu. Existence of a periodic mild solution for a nonlinear fractional differential equation. Computers and Mathematics with Applications, 64 (10):3059–3064, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211101131X.

Hai:2019:NSC

[HB19] Bhuiyan Shameem Mahmood Ebna Hai and Markus Bause. Numerical study and comparison of alternative time discretization schemes for an ultrasonic guided wave propagation problem coupled with fluid-structure interaction. Computers and Mathematics with Applications, 78 (9):2867–2885, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300306.

Heubes:2014:EAB

[HBE14] Daniel Heubes, Andreas Bartel, and Matthias Ehrhardt. Exact artificial boundary conditions for a lattice Boltzmann method. Computers and Mathematics with Applications, 67(11):2041– 2054, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001801.

Heubes:2015:COD

[HBE15] Daniel Heubes, Andreas Bartel, and Matthias Ehrhardt. Concept for a one-dimensional discrete artificial boundary condition for the lattice Boltzmann method. Computers and Mathematics with Applications, 70 (10):2316–2330, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004095.

Haussmann:2019:LES

[HBK^{+}19] Marc Haussmann, Alejandro Claro Barreto, Gislain Lipeme Kouyi, Nicolas Rivière, Hermann Nirschl, and Mathias J. Krause. Largeeddy simulation coupled with wall models for turbulent channel flows at high Reynolds numbers with a lattice Boltzmann method — application to Coriolis mass flowmeter. Computers and Mathematics with Applications, 78 (10):3285–3302, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302494.

Hatzikirou:2010:PTF

[HBS^{+}10] H. Hatzikirou, L. Brusch, C. Schaller, M. Simon, and A. Deutsch. Prediction of traveling front behavior in a latticegas cellular automaton model for tumor invasion. Computers and Mathematics with Applications, 59(7):2326–2339, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006294.

Huang:2014:EES

[HC14] Fenglin Huang and Yanping Chen. Error estimates for spectral approximation of elliptic control problems with integral state and control constraints. Computers and Mathematics with Applications, 68(8):789–803, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002855.

Hao:2016:GDS

[HC16] Jianghao Hao and Li Cai.

General decay of solutions for a system of coupled viscoelastic wave equations with source terms. Computers and Mathematics with Applications, 72(5):1418–1425, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304126.

He:2018:MCB

[HC18] Xin-Jiang He and Wenting Chen. A Monte-Carlo based approach for pricing credit default swaps with regime switching. Computers and Mathematics with Applications, 76(7):1758–1766, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303924.

Heidarkhani:2016:PKT

[HCF16] Shapour Heidarkhani, Giuseppe Caristi, and Massimiliano Ferrara. Perturbed Kirchhoff-type Neumann problems in Orlicz–Sobolev spaces. Computers and Mathematics with Applications, 71(10):2008–2019, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301481.

Hsu:2012:PAM

[HCHH12] Chia-Cheng Hsu, Hsin-Chin Chen, Kuo-Kuang Huang, and Yueh-Min Huang. A personalized auxiliary material recommendation system based on learning style on Facebook applying an artificial bee colony algorithm. Computers and Mathematics with Applications, 64(5):1506–1513, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003161.

Hung:2011:WSM

[HCL11] Wen-Liang Hung, Yen-Chang Chang, and E. Stanley Lee. Weight selection in W –K-means algorithm with an application in color image segmentation. Computers and Mathematics with Applications, 62(2):668– 676, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004469.

Hao:2012:HRM

[HCL12a] Dinh Nho Hào, La Huu Chuong, and D. Lesnic. Heuristic regularization methods for numerical differentiation. Computers and Mathematics with Applications, 63(4):816–826, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010303.

Hsu:2012:PEO

[HCL12b] Jenq-Muh Hsu, Chao-Chun Chen, and Chia-Chi Li. POOT: an efficient object tracking strategy based on short-term optimistic predictions for face-structured sensor networks. Computers and Mathematics with Applications, 63(2):391–406, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005888.

Huang:2018:NAL

[HCLL18] Yunqing Huang, Meng Chen, Jichun Li, and Yanping Lin. Numerical analysis of a leapfrog ADI–FDTD method for Maxwell’s equations in lossy media. Computers and Mathematics with Applications, 76(4):938–956, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303080.

Ho:2012:SSM

[HCT12] Ching-Yen Ho, Bor-Chyuan Chen, and Yu-Hsiang Tsai. Scattering signals of monochromatic light incident on a rectangular microchannel. Computers and Mathematics with Applications, 64(5):1514–1521, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003422.

Chen:2011:NCN

[hCTM11] Guo hua Chen, Zhao-Ling Tao, and Jin-Zhong Min. Notes on a conservative nonlinear oscillator. Computers and Mathematics with Applications, 61(8):2120–2122, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006711.

Hu:2016:TGM

[HCZ16] Hanzhang Hu, Yanping Chen, and Jie Zhou. Twogrid method for miscible displacement problem by mixed finite element methods and finite element method of characteristics. Computers and Mathematics with Applications, 72 (11):2694–2715, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305041.

Helfen:2014:CHC

[HD14a] Cécile E. Helfen and Stefan Diebels. Computational homogenisation of composite plates: Consideration of the thickness change with a modified projection strategy. Computers and Mathematics with Applications, 67(5):1116–1129, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300713X.

Hosseinzadeh:2014:SAS

[HD14b] Hossein Hosseinzadeh and Mehdi Dehghan. A simple and accurate scheme based on complex space C to calculate boundary integrals of 2D boundary elements method. Computers and Mathematics with Applications, 68(4):531–542, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002727.

Hou:2016:ELP

[HD16] Yanren Hou and GuangZhi Du. An expandable local and parallel two-grid finite element scheme. Computers and Mathematics with Applications, 71(12):2541–2556, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301699.

Hubrich:2019:NIN

[HD19] S. Hubrich and A. Düster.

Numerical integration for nonlinear problems of the finite cell method using an adaptive scheme based on moment fitting. Computers and Mathematics with Applications, 77(7):1983–1997, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306837.

Han:2011:EPF

[HDHL11] Hua Han, Yong-Sheng Ding, Kuang-Rong Hao, and Xiao Liang. An evolutionary particle filter with the immune genetic algorithm for intelligent video target tracking. Computers and Mathematics with Applications, 62(7):2685–2695, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100530X.

Hu:2011:PHS

[HDHW11] Yingjing Hu, Lixia Duan, Yi Huang, and Yuhai Wu. On the perturbations of a Hamiltonian system. Computers and Mathematics with Applications, 61(4):773–782, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009296.

Han:2011:SDP

[HDS11] Guanghua Han, Ming Dong, and Xiaofeng Shao. A stochastic dynamic programming approachbased yield management with substitution and uncertainty in semiconductor manufacturing. Computers and Mathematics with Applications, 61(4):1241–1253, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000083.

Hladik:2011:CAE

[HDT11] Milan Hladík, David Daney, and Elias Tsigaridas. Characterizing and approximating eigenvalue sets of symmetric interval matrices. Computers and Mathematics with Applications, 62(8):3152–3163, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006912.

He:2011:AMN

[He11] Ji-Huan He. Asymptotic methods: the next frontier towards nonlinear science. Computers and Mathematics with Applications, 61(8):1907–1908, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002458.

He:2016:NCT

[He16] Dongdong He. On the L^{∞}-norm convergence of a three-level linearly implicit finite difference method for the extended Fisher–Kolmogorov equation in both 1D and 2D. Computers and Mathematics with Applications, 71(12):2594– 2607, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302073.

Heider:2010:CSR

[Hei10a] P. Heider. Computation of scattering resonances for dielectric resonators. Computers and Mathematics with Applications, 60(6):1620–1632, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000458X.

Heino:2010:LBF

[Hei10b] Pekka Heino. Lattice-Boltzmann finite-difference model with optical phonons for nanoscale thermal conduction. Computers and Mathematics with Applications, 59(7):2351– 2359, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006452.

Hassan:2010:OSO

[HEP10] Taher S. Hassan, Lynn Erbe, and Allan Peterson. Oscillation of second order superlinear dynamic equations with damping on time scales. Computers and Mathematics with Applications, 59(1):550–558, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003782.

Herzallah:2014:CEF

[Her14] Mohamed A. E. Herzallah. Comments on “Effect of fractional parameter on plane waves of generalized magnetothermoelastic diffusion with reference temperaturedependent elastic medium” [Comput. Math. Appl. 65 (2013) 1104–1118]. Computers and Mathematics with Applications, 67 (5):997–998, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114000224.

See [OSA13].

Herzallah:2019:CDM

[Her19] Mohamed A. E. Herzallah. Comments on “Different methods for (3+1)- dimensional space–time fractional modified KdV-Zakharov–Kuznetsov equation” [Comput. Math. Appl. 71(6) (2016) 1259– 1269]. Computers and Mathematics with Applications, 77(1):66–68, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830525X.

See [GABC16].

Hessari:2014:FOS

[Hes14] Peyman Hessari. First order system least squares method for the interface problem of the Stokes equations. Computers and Mathematics with Applications, 68(3):309– 324, August 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002624.

Hessari:2018:PSL

[Hes18] Peyman Hessari. Pseudospectral least squares method for linear elasticity. Computers and Mathematics with Applications, 76(6):1356–1371, September 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303523.

Huang:2010:FOS

[HF10] Mugen Huang and Weizhen Feng. Forced oscillations for second order delay differential equations with impulses. Computers and Mathematics with Applications, 59(1):18–30, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109002272.

Hejranfar:2016:SDL

[HG16] Kazem Hejranfar and Ali Ghaffarian. A spectral difference lattice Boltzmann method for solution of inviscid compressible flows on structured grids. Computers and Mathematics with Applications, 72(5):1341–1368, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303753.

Haq:2018:ENA

[HG18a] Sirajul Haq and Abdul Ghafoor. An efficient numerical algorithm for multi-dimensional time dependent partial differential equations. Computers and Mathematics with Applications, 75(8):2723–2734, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300087.

Hu:2018:HOM

[HG18b] Jinhao Hu and Siqing Gan. High order method for Black–Scholes PDE. Computers and Mathematics with Applications, 75(7):2259–2270, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307654.

Haq:2019:NST

[HGHA19] Sirajul Haq, Abdul Ghafoor, Manzoor Hussain, and Shamsul Arifeen. Numerical solutions of two dimensional Sobolev and generalized Benjamin– Bona–Mahony–Burgers equations via Haar wavelets. Computers and Mathematics with Applications, 77(2):565–575, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305789.

He:2019:CDC

[HGJP19] Linwei He, Matthew Gilbert, Thomas Johnson, and Tom Pritchard.

Conceptual design of AM components using layout and geometry optimization. Computers and Mathematics with Applications, 78(7):2308–2324, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303778.

Hosseinnia:2010:SMS

[HGN^{+}10] S. H. Hosseinnia, R. Ghaderi, A. Ranjbar N., M. Mahmoudian, and S. Momani. Sliding mode synchronization of an uncertain fractional order chaotic system. Computers and Mathematics with Applications, 59(5):1637–1643, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005409.

Han:2018:ULB

[HGSL18] Yuzhu Han, Wenjie Gao, Zhe Sun, and Haixia Li. Upper and lower bounds of blow-up time to a parabolic type Kirchhoff equation with arbitrary initial energy. Computers and Mathematics with Applications, 76 (10):2477–2483, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304693.

Huang:2011:PAP

[HGW11] Zuda Huang, Shuhua Gong, and Lijuan Wang. Positive almost periodic solution for a class of Lasota–Wazewska model with multiple time-varying delays. Computers and Mathematics with Applications, 61(4):755–760, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009247.

He:2018:BDC

[HGW18] Yijun He, Huaihong Gao, and Hua Wang. Blowup and decay for a class of pseudo-parabolic p-Laplacian equation with logarithmic nonlinearity. Computers and Mathematics with Applications, 75(2):459–469, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/

pii/S0898122117305904.

See comments [DZ21].

Gao:2015:TAD

[hGzS15] Guang hua Gao and Zhi zhong Sun. Two alternating direction implicit difference schemes with the extrapolation method for the two-dimensional distributed-order differential equations. Computers and Mathematics with Applications, 69(9):926– 948, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000905.

Huang:2010:SRA

[HH10a] Chin-Yu Huang and Tsui-Ying Hung. Software reliability analysis and assessment using queueing models with multiple change-points. Computers and Mathematics with Applications, 60(7):2015–2030, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005316.

Hussain:2010:PSD

[HH10b] Malik Zawwar Hussain and Maria Hussain. C^{1} positive scattered data interpolation. Computers and Mathematics with Applications, 59(1):457– 467, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109003927.

He:2011:TAP

[HH11] Jun He and Ting-Zhu Huang. Two augmentation preconditioners for nonsymmetric and indefinite saddle point linear systems with singular (1, 1) blocks. Computers and Mathematics with Applications, 62(1):87– 92, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003555.

Hollig:2015:PFE

[HH15] Klaus Höllig and Jörg Hörner. Programming finite element methods with weighted B-splines. Computers and Mathematics with Applications, 70(7):1441–1456, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122115000863.

Hasanov:2016:CAS

[HH16] Vejdi I. Hasanov and Sevdzhan A. Hakkaev. Convergence analysis of some iterative methods for a nonlinear matrix equation. Computers and Mathematics with Applications, 72(4):1164–1176, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303674.

Hashemabadi:2017:ISO

[HH17] M. Hashemabadi and M. Hadidoolabi. Implicit second-order CUSP gridless method for unsteady moving boundary simulations. Computers and Mathematics with Applications, 74(4):842–858, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303322.

Hao:2018:SWE

[HH18a] Jianghao Hao and Yaxin Hou. Stabilization for wave equation of variable coefficients with Balakrishnan–Taylor damping and source term. Computers and Mathematics with Applications, 76(9):2235–2245, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304437.

Hashemian:2018:IFF

[HH18b] Ali Hashemian and Seyed Farhad Hosseini. An integrated fitting and fairing approach for object reconstruction using smooth NURBS curves and surfaces. Computers and Mathematics with Applications, 76(7):1555–1575, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303729.

Heydari:2014:ECM

[HHG14] M. H. Heydari, M. R. Hooshmandasl, and F. M. Maalek Ghaini. An efficient computational method for solving fractional biharmonic equation. Computers and Mathematics with Applications, 68(3):269– 287, August 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114002600.

Hosseininia:2019:WMS

[HHGA19] M. Hosseininia, M. H.

Heydari, F. M. Maalek Ghaini, and Z. Avazzadeh. A wavelet method to solve nonlinear variable-order time fractional 2D Klein– Gordon equation. Computers and Mathematics with Applications, 78 (12):3713–3730, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303116.

Hussain:2012:SEL

[HHM12] Salam Hadi Hussain, Ahmed Kadhim Hussein, and Rehab Noor Mohammed. Studying the effects of a longitudinal magnetic field and discrete isoflux heat source size on natural convection inside a tilted sinusoidal corrugated enclosure. Computers and Mathematics with Applications, 64(4):476–488, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111010777.

Herrera-Hernandez:2018:VNS

[HHNLGC18] E. C. Herrera-Hernández, M. Núñez-López, and J. A. González-Calderón. Validation of numerical solution of diffusive part in a reaction–diffusion model. Computers and Mathematics with Applications, 74(1):143–156, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300433.

Hor:2010:ETM

[HHS^{+}10] F. C. Hor, Liang-Chih Huang, Hsu-Shih Shih, Yen-Hua Lee, and E. Stanley Lee. Establishing talent management for company’s succession planning through analytic network process: Application to an MNC semiconductor company in Taiwan. Computers and Mathematics with Applications, 60(3):528–540, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000341X.

He:2017:NSU

[HHS^{+}17] Qiaolin He, Jinyang Huang, Xiaoding Shi, Xiao-Ping Wang, and Chao Bi. Numerical simulation of 2D unsteady shear-thinning non-Newtonian incompressible fluid in screw extruder with fictitious domain method. Computers and Mathematics with Applications, 73(1):109–121, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306125.

He:2011:MAS

[HHY^{+}11] Anping He, William N. N.

Hung, Guowu Yang, Jinzhao Wu, and Lian Li. Mathematical analysis of stage-based programmable logic controller. Computers and Mathematics with Applications, 61(7):1769–1785, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000903.

Han:2013:SDT

[HHY13] Houde Han, Zhongyi Huang, and Wenjun Ying.

A semi-discrete tailored finite point method for a class of anisotropic diffusion problems. Computers and Mathematics with Applications, 65(11): 1760–1774, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113001697.

Herty:2019:FAS

[HIS19] Michael Herty, Nouh Izem, and Mohammed Seaid. Fast and accurate simulations of shallow water equations in large networks. Computers and Mathematics with Applications, 78 (6):2107–2126, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301804.

He:2013:UMU

[HJ13] Xinguang He and Lijian Jiang. An upscaling method using coefficient splitting and its applications to elliptic PDEs. Computers and Mathematics with Applications, 65(4):712–730, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112007031.

Heinze:2015:NHH

[HJD15] Stephan Heinze, Meysam Joulaian, and Alexander Düster. Numerical homogenization of hybrid metal foams using the finite cell method. Computers and Mathematics with Applications, 70(7):1501–1517, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002321.

Hamamci:2010:CAS

[HK10] Serdar Ethem Hamamci and Muhammet Koksal. Calculation of all stabilizing fractional-order PD controllers for integrating time delay systems. Computers and Mathematics with Applications, 59(5):1621–1629, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005380.

Heuer:2015:DMO

[HK15] Norbert Heuer and Michael Karkulik. DPG method with optimal test functions for a transmission problem. Computers and Mathematics with Applications, 70(5):1070–1081, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003181.

Husain:2017:LSS

[HK17] Akhlaq Husain and Arbaz Khan. Least-squares spectral element preconditioners for fourth order elliptic problems. Computers and Mathematics with Applications, 74(3):482–503, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302638.

Hur:2013:REC

[HKHK13] Junbeom Hur, Dongyoung Koo, Seong Oun Hwang, and Kyungtae Kang. Removing escrow from ciphertext policy attribute-based encryption. Computers and Mathematics with Applications, 65(9):1310–1317, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001198.

Hashmi:2012:NSW

[HKI12] M. S. Hashmi, N. Khan, and S. Iqbal. Numerical solutions of weakly singular Volterra integral equations using the optimal homotopy asymptotic method. Computers and Mathematics with Applications, 64 (6):1567–1574, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011564.

Ha:2014:GEU

[HKJ14] Tae Gab Ha, Daewook Kim, and Il Hyo Jung. Global existence and uniform decay rates for the semi-linear wave equation with damping and source terms. Computers and Mathematics with Applications, 67(3):692–707, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122113007062.

He:2016:RHP

[HKK^{+}16] Fuli He, Min Ku, Uwe Kähler, Frank Sommen, and Swanhild Bernstein. Riemann–Hilbert problems for null-solutions to iterated generalized Cauchy–Riemann equations in axially symmetric domains. Computers and Mathematics with Applications, 71(10):1990– 2000, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301468.

Hong:2013:HOE

[HKKK13] Deukjo Hong, Dong-Chan Kim, Woo-Hwan Kim, and Jongsung Kim. Higher order eTCR hash functions. Computers and Mathematics with Applications, 65(9):1396–1402, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000430.

Han:2017:IDS

[HKP17] Joo Hyeong Han, Jae Ryong Kweon, and Minje Park. Interior discontinuity for a stationary compressible Stokes system with inflow datum. Computers and Mathematics with Applications, 74 (10):2321–2329, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304194.

Heuer:2014:NDT

[HKS14] Norbert Heuer, Michael Karkulik, and Francisco-Javier Sayas. Note on discontinuous trace approximation in the practical DPG method. Computers and Mathematics with Applications, 68 (11):1562–1568, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003125.

Haslinger:2019:SSL

[HKS19a] Jaroslav Haslinger, Radek Kucera, and Václav Sátek. Stokes system with local Coulomb’s slip boundary conditions: Analysis of discretized models and implementation. Computers and Mathematics with Applications, 77(6):1655–1667, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302384.

Havelka:2019:DRT

[HKS19b] Jan Havelka, Anna Kucerová, and Jan Sýkora. Dimensionality reduction in thermal tomography. Computers and Mathematics with Applications, 78(9):3077–3089, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302275.

He:2011:SCQ

[HKT11] Bo He, Omar Kihel, and Alain Togbé. Solutions of a class of quartic Thue inequalities. Computers and Mathematics with Applications, 61(9):2914– 2923, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002434.

Hyun:2015:ESM

[HKW15] Jaeyub Hyun, Junghwan Kook, and Semyung Wang. Efficient and stable model reduction scheme for the numerical simulation of broadband acoustic metamaterials. Computers and Mathematics with Applications, 69(8):876–892, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000061.

Huang:2010:CAT

[HL10] Yufei Huang and Bolian Liu. On a conjecture about the k th lower multiexponent. Computers and Mathematics with Applications, 60(1):36– 44, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002981.

Henderson:2011:BOD

[HL11a] Johnny Henderson and Xueyan Liu. BVP’s with odd differences of gaps in boundary conditions for n th order ODE’s by matching solutions. Computers and Mathematics with Applications, 62(10):3722–3728, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007826.

Henderson:2011:PSS

[HL11b] Johnny Henderson and Rodica Luca. Positive solutions for a system of higher-order multi-point boundary value problems. Computers and Mathematics with Applications, 62(10):3920–3932, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008133.

Huang:2011:SMD

[HL11c] Yong Huang and Qi-Zhi Luo. A simple method to determine the critical buckling loads for axially inhomogeneous beams with elastic restraint. Computers and Mathematics with Applications, 61(9):2510–2517, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001362.

He:2012:CBS

[HL12] Ji-Huan He and Yong Liu. Control of bubble size and bubble number in bubble electrospinning. Computers and Mathematics with Applications, 64(5):1033–1035, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200226X.

Han:2018:TRE

[HL18a] Yuzhu Han and Qingwei Li. Threshold results for the existence of global and blow-up solutions to Kirchhoff equations with arbitrary initial energy. Computers and Mathematics with Applications, 75(9):3283–3297, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300634.

Hu:2018:IDD

[HL18b] Wen-Qiang Hu and Zhi-Hui Li. Investigation on different discrete velocity quadrature rules in gas-kinetic unified algorithm solving Boltzmann model equation. Computers and Mathematics with Applications, 75(11):4179–4200, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301561.

Heuze:2014:BTBb

[HLB14] Thomas Heuzé, Jean-Baptiste Leblond, and Jean-Michel Bergheau. Benchmark tests based on the Couette viscometer — I: Laminar flow of incompressible fluids with inertia effects and thermomechanical coupling. Computers and Mathematics with Applications, 67(10):1925– 1937, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001400.

Hung:2011:BBR

[HLC11] Wen-Liang Hung, E. Stanley Lee, and Shun-Chin Chuang. Balanced bootstrap resampling method for neural model selection. Computers and Mathematics with Applications, 62 (12):4576–4581, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122111009035.

Huang:2012:DQS

[HLCY12] Cheng-Fang Huang, Teh-Lu Liao, Chih-Yung Chen, and Jun-Juh Yan. The design of quasisliding mode control for a permanent magnet synchronous motor with unmatched uncertainties. Computers and Mathematics with Applications, 64(5):1036–1043, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002271.

Hussein:2014:SDT

[HLI14] M. S. Hussein, D. Lesnic, and M. I. Ivanchov. Simultaneous determination of time-dependent coefficients in the heat equation. Computers and Mathematics with Applications, 67(5):1065–1091, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000169.

Hoekstra:2010:MME

[HLK10] Alfons Hoekstra, Li-Shi Luo, and Manfred Krafczyk. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 59(7):2139–2140, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007287.

He:2013:AMF

[HLL13] Siriguleng He, Hong Li, and Yang Liu. Analysis of mixed finite element methods for fourth-order wave equations. Computers and Mathematics with Applications, 65(1):1–16, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006037.

Hou:2015:NEC

[HLL^{+}15] Jingming Hou, Qiuhua Liang, Zhanbin Li, Shifeng Wang, and Reinhard Hinkelmann. Numerical error control for secondorder explicit TVD scheme with limiters in advection simulation. Computers and Mathematics with Applications, 70 (9):2197–2209, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004010.

Hao:2019:PAS

[HLLM19] Xiazhi Hao, Yinping Liu, Zhibin Li, and Wen-Xiu Ma. Painlevé analysis, soliton solutions and lump-type solutions of the (3 + 1)-dimensional generalized KP equation. Computers and Mathematics with Applications, 77(3):724–730, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305947.

Hu:2019:LBM

[HLNZ19] Yang Hu, Decai Li, Xiaodong Niu, and Yanjuan Zhang. Lattice Boltzmann model for the axisymmetric electrothermo-convection. Computers and Mathematics with Applications, 78(1): 55–65, July 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122119300744.

Harjani:2011:FPT

[HLS11] J. Harjani, B. López, and K. Sadarangani. Fixed point theorems for weakly C-contractive mappings in ordered metric spaces. Computers and Mathematics with Applications, 61(4):790–796, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009326.

Hu:2015:SSF

[HLSN15] Yang Hu, Decai Li, Shi Shu, and Xiaodong Niu. Simulation of steady fluid-solid conjugate heat transfer problems via immersed boundary-lattice Boltzmann method. Computers and Mathematics with Applications, 70 (9):2227–2237, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004034.

Hu:2016:LBF

[HLSN16] Yang Hu, Decai Li, Shi Shu, and Xiaodong Niu. Lattice Boltzmann flux scheme for the convection–diffusion equation and its applications. Computers and Mathematics with Applications, 72(1):48–63, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302139.

Hu:2012:TDM

[HLT12] Jhen-Jia Hu, Hui-Chieh Li, and Hung-Ming Tai. Thermal distribution monitoring of the container data center by a fast infrared image fusion technique. Computers and Mathematics with Applications, 64(5):1484–1494, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003148.

Hao:2017:RSE

[HLTL17] Xiazhi Hao, Yinping Liu, Xiaoyan Tang, and Zhibin Li. The residual symmetry and exact solutions of the Davey– Stewartson III equation. Computers and Mathematics with Applications, 73(11):2404–2414, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302092.

Hook:2018:KFF

[HLvS18] Lars Josef Höök, Gustav Ludvigsson, and Lina von Sydow. The Kolmogorov forward fractional partial differential equation for the CGMYprocess with applications in option pricing. Computers and Mathematics with Applications, 76 (10):2330–2344, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304486.

Huang:2019:ACF

[HLW19] Yong Huang, Zhenhai Liu, and Ching-Feng Wen. Approximate controllability for fractional semilinear parabolic equations. Computers and Mathematics with Applications, 77(11):2971– 2979, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304218.

Hao:2011:MPS

[HLWX11] Xinan Hao, Lishan Liu, Yonghong Wu, and Naiwei Xu. Multiple positive solutions for singular n th-order nonlocal boundary value problems in Banach spaces. Computers and Mathematics with Applications, 61(7):1880–1890, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001039.

Huang:2018:FPM

[HLX18] Yunqing Huang, Xueyang Li, and Aiguo Xiao. Fourier pseudospectral method on generalized sparse grids for the spacefractional Schrödinger equation. Computers and Mathematics with Applications, 75 (12):4241–4255, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301627.

Huang:2012:NVS

[HLY12a] Xiaokun Huang, Hongjie Li, and Yunqiang Yin. Notes on “Vague soft sets and their properties”. Computers and Mathematics with Applications, 64(6):2153–2157, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000119.

See [XMWH10].

Huang:2012:SMM

[HLY12b] Yunqing Huang, Jichun Li, and Wei Yang. Solving metamaterial Maxwell’s equations via a vector wave integro-differential equation. Computers and Mathematics with Applications, 63(12):1597– 1606, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002404.

Huang:2016:TNA

[HLY16] Yunqing Huang, Jichun Li, and Wei Yang. Theoretical and numerical analysis of a nonlocal dispersion model for light interaction with metallic nanostructures. Computers and Mathematics with Applications, 72(4):921–932, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303352.

Hou:2017:EES

[HLY17a] Tianliang Hou, Chunmei Liu, and Yin Yang. Error estimates and superconvergence of a mixed finite element method for elliptic optimal control problems. Computers and Mathematics with Applications, 74(4):714–726, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303103.

Hussein:2017:RSD

[HLY17b] S. O. Hussein, D. Lesnic, and M. Yamamoto. Reconstruction of spacedependent potential and/or damping coefficients in the wave equation. Computers and Mathematics with Applications, 74 (6):1435–1454, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303784.

Huang:2015:OCF

[HLZ15] Yong Huang, Zhenhai Liu, and Biao Zeng. Optimal control of feedback control systems governed by hemivariational inequalities. Computers and Mathematics with Applications, 70(8):2125–2136, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004083.

Huang:2011:ASF

[HLZD11] Li Huang, Xian-Fang Li, Yulin Zhao, and Xiang-Yang Duan. Approximate solution of fractional integro-differential equations by Taylor expansion method. Computers and Mathematics with Applications, 62(3):1127–1134, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001891.

Han:2016:WIV

[HLZM16] Yazhou Han, Zhongfang Li, Shutao Zhang, and Manjun Ma. Wavefront invasion for a volumefilling chemotaxis model with logistic growth.

Computers and Mathematics with Applications, 71(2):471–478, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005702.

Herisanu:2010:AAS

[HM10] N. Herisanu and V. Marinca. Accurate analytical solutions to oscillators with discontinuities and fractional-power restoring force by means of the optimal homotopy asymptotic method. Computers and Mathematics with Applications, 60(6):1607–1615, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004566.

Hoskova-Mayerova:2012:TH

[HM12] Sárka Hosková-Mayerová.

Topological hypergroupoids. Computers and Mathematics with Applications, 64(9):2845–2849, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112003665.

Huang:2014:MCG

[HM14] Na Huang and Changfeng Ma. The modified conjugate gradient methods for solving a class of generalized coupled Sylvestertranspose matrix equations. Computers and Mathematics with Applications, 67(8):1545–1558, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000613.

Huang:2015:TSP

[HM15] Na Huang and Chang-Feng Ma. Two structurepreserving-doubling like algorithms for obtaining the positive definite solution to a class of nonlinear matrix equation. Computers and Mathematics with Applications, 69(6):494–502, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000206.

Hinow:2017:LSD

[HM17a] Peter Hinow and Maya Mincheva. Linear stability of delayed reaction– diffusion systems. Computers and Mathematics with Applications, 73(2):226–232, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630623X.

Hu:2017:MNH

[HM17b] Jing-Jing Hu and Chang-Feng Ma. Minimumnorm Hamiltonian solutions of a class of generalized Sylvesterconjugate matrix equations. Computers and Mathematics with Applications, 73(5):747–764, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300068.

Huang:2017:LSG

[HM17c] Bao-Hua Huang and Chang-Feng Ma. On the least squares generalized Hamiltonian solution of generalized coupled Sylvester-conjugate matrix equations. Computers and Mathematics with Applications, 74(3):532–555, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302791.

Huang:2017:AIB

[HM17d] Na Huang and Chang-Feng Ma. Analysis on inexact block diagonal preconditioners for elliptic PDE-constrained optimization problems. Computers and Mathematics with Applications, 74 (10):2423–2437, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304352.

Holm:2018:FRE

[HM18a] Bärbel Holm and Svetlana Matculevich. Fully reliable error control for first-order evolutionary problems. Computers and Mathematics with Applications, 75(4):1302–1329, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307083.

Huang:2018:GBI

[HM18b] Bao-Hua Huang and Chang-Feng Ma. Gradientbased iterative algorithms for generalized coupled Sylvester-conjugate matrix equations. Computers and Mathematics with Applications, 75(7):2295–2310, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307745.

Huang:2019:SLS

[HM19] Baohua Huang and Changfeng Ma. A Shamanskii-like self-adaptive Levenberg– Marquardt method for nonlinear equations. Computers and Mathematics with Applications, 77(2):357–373, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305510.

Hashemi:2016:EPS

[HMF16] M. R. Hashemi, M. T.

Manzari, and R. Fatehi. Evaluation of a pressure splitting formulation for weakly compressible SPH: Fluid flow around periodic array of cylinders. Computers and Mathematics with Applications, 71(3):758–778, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115006021.

Hahn:2019:III

[HMF^{+}19] Jooyoung Hahn, Karol Mikula, Peter Frolkovic, Matej Medl’a, and Branislav Basara. Iterative inflowimplicit outflow-explicit finite volume scheme for level-set equations on polyhedron meshes. Computers and Mathematics with Applications, 77(6):1639–1654, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303572.

Hooshmand:2012:PFM

[HMM12] A. Hooshmand, H. A. Malki, and J. Mohammadpour. Power flow management of microgrid networks using model predictive control. Computers and Mathematics with Applications, 64(5):869–876, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000387.

Helfer:2015:IOS

[HMP^{+}15] Thomas Helfer, Bruno Michel, Jean-Michel Proix, Maxime Salvo, Jérôme Sercombe, and Michel Casella. Introducing the open-source mfront code generator: Application to mechanical behaviours and material knowledge management within the PLEIADES fuel element modelling platform. Computers and Mathematics with Applications, 70(5):994–1023, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003132.

Henrich:2010:DGC

[HMSC10] O. Henrich, D. Marenduzzo, K. Stratford, and M. E. Cates. Domain growth in cholesteric blue phases: Hybrid lattice Boltzmann simulations. Computers and Mathematics with Applications, 59(7):2360–2369, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006324.

Haghi:2018:RFM

[HMV18] Majid Haghi, Reza Mollapourasl, and Michèle Vanmaele. An RBF–FD method for pricing American options under jumpdiffusion models. Computers and Mathematics with Applications, 76 (10):2434–2459, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304668.

Hu:2016:SNM

[HMWZ16] Jun Hu, Hongying Man, Jianye Wang, and Shangyou Zhang. The simplest nonconforming mixed finite element method for linear elasticity in the symmetric formulation on n-rectangular grids. Computers and Mathematics with Applications, 71(7):1317–1336, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300190.

Hao:2015:EHA

[HMY15] S. Hao, P. G. Martinsson, and P. Young. An efficient and highly accurate solver for multibody acoustic scattering problems involving rotationally symmetric scatterers. Computers and Mathematics with Applications, 69(4):304–318, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005823.

Hu:2018:WGM

[HMY18] Xiaozhe Hu, Lin Mu, and Xiu Ye. Weak Galerkin method for the Biot’s consolidation model. Computers and Mathematics with Applications, 75(6):2017–2030, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304303.

He:2011:DAL

[HmZ11] Xiaofei He and Qi ming Zhang. A discrete analogue of Lyapunovtype inequalities for nonlinear difference systems. Computers and Mathematics with Applications, 62(2):677–684, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004470.

Huang:2018:ABD

[HMZ18] Na Huang, Chang-Feng Ma, and Jun Zou. Analysis on block diagonal and triangular preconditioners for a PML system of an electromagnetic scattering problem. Computers and Mathematics with Applications, 74 (11):2856–2873, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304558.

Huy:2010:NWT

[HN10] Vu Nhat Huy and Quô’c-Anh Ngô. A new way to think about Ostrowskilike type inequalities. Computers and Mathematics with Applications, 59(9):3045–3052, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110001331.

Haddar:2018:SMR

[HN18] Houssem Haddar and Thi-Phong Nguyen. Sampling methods for reconstructing the geometry of a local perturbation in unknown periodic layers. Computers and Mathematics with Applications, 74(11):2831–2855, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304327.

Hankins:2013:MNB

[HNK13] Michael J. Hankins, Timea Nagy, and István Z. Kiss. Methodology for a nullcline-based model from direct experiments: Applications to electrochemical reaction models. Computers and Mathematics with Applications, 65(10):1633– 1644, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006773.

Holm:2013:DFA

[HNPS13] Sverre Holm, Sven Peter Näsholm, Fabrice Prieur, and Ralph Sinkus. Deriving fractional acoustic wave equations from mechanical and thermal constitutive equations. Computers and Mathematics with Applications, 66(5):621–629, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001557.

Henderson:2010:IDI

[HO10] Johnny Henderson and Abdelghani Ouahab. Impulsive differential inclusions with fractional order. Computers and Mathematics with Applications, 59(3):1191–1226, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003484.

Hegarty:2019:PUN

[HO19] A. F. Hegarty and E. O’Riordan. A parameter-uniform numerical method for a singularly perturbed convection– diffusion problem posed on an annulus. Computers and Mathematics with Applications, 78 (10):3329–3344, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930269X.

Hofer:2018:PCD

[Hof18] Christoph Hofer. Parallelization of continuous and discontinuous Galerkin dual-primal isogeometric tearing and interconnecting methods. Computers and Mathematics with Applications, 74(7):1607–1625, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304091.

Holm:2011:LTD

[Hol11] Michael T. Holm. The Laplace transform in discrete fractional calculus. Computers and Mathematics with Applications, 62(3):1591–1601, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003208.

Hong:2010:SCS

[Hon10] Shihuang Hong. Stability criteria for set dynamic equations on time scales. Computers and Mathematics with Applications, 59(11):3444– 3457, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002221.

Hong:2012:HUA

[Hon12] Wei-Tyng Hong. HCRF– UBM approach for textindependent speaker identification. Computers and Mathematics with Applications, 64(5):1120–1127, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002350.

Hotar:2013:FGI

[Hot13] Vlastimil Hotar. Fractal geometry for industrial data evaluation. Computers and Mathematics with Applications, 66(2):113–121, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000321.

Hou:2015:PPE

[Hou15] Tianliang Hou. A priori and a posteriori error estimates of H^{1}- Galerkin mixed finite element methods for elliptic optimal control problems. Computers and Mathematics with Applications, 70(10):2542–2554, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004587.

Huang:2010:ASC

[HP10] Sai-Hua Huang and Tian-Xiao Pang. An almost sure central limit theorem for self-normalized partial sums. Computers and Mathematics with Applications, 60(9):2639–2644, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006784.

Hasanov:2013:IUT

[HP13] Alemdar Hasanov and Burhan Pektas. Identification of an unknown time-dependent heat source term from overspecified Dirichlet boundary data by conjugate gradient method.

Computers and Mathematics with Applications, 65(1):42–57, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200627X.

He:2017:USL

[HP17] Dongdong He and Kejia Pan. An unconditionally stable linearized CCD– ADI method for generalized nonlinear Schrödinger equations with variable coefficients in two and three dimensions. Computers and Mathematics with Applications, 73(11):2360–2374, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302304.

Hajabdollahi:2019:ILM

[HP19a] Farzaneh Hajabdollahi and Kannan N. Premnath. Improving the low Mach number steady state convergence of the cascaded lattice Boltzmann method by preconditioning. Computers and Mathematics with Applications, 78(4):1115–1130, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300664.

He:2019:ACC

[HP19b] Jia Wei He and Li Peng.

Approximate controllability for a class of fractional stochastic wave equations. Computers and Mathematics with Applications, 78(5):1463–1476, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300331.

He:2012:ECT

[HPC12] Debiao He, Sahadeo Padhye, and Jianhua Chen. An efficient certificateless two-party authenticated key agreement protocol. Computers and Mathematics with Applications, 64(6):1914–1926, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002490.

He:2011:TQP

[HpD11] Ming He and Yong ping Du. P -top-k queries in a probabilistic framework from information extraction models. Computers and Mathematics with Applications, 62(7):2755–2769, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005542.

Hemami:2019:NSR

[HPR19]  Mohammad  Hemami,

Kourosh Parand, and Jamal Amani Rad. Numerical simulation of reaction–diffusion neural dynamics models and their synchronization/desynchronization: Application to epileptic seizures. Computers and Mathematics with Applications, 78 (11):3644–3677, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303141.

Hungria:2018:HME

[HPS18] Allan Hungria, Daniele Prada, and Francisco-Javier Sayas. HDG methods for elastodynamics. Computers and Mathematics with Applications, 74(11):2671–2690, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305084.

He:2018:PAM

[HPV^{+}18] Xiaoming He, Stephen Pankavich, Erik Van Vleck, Zhu Wang, and Xiu Ye. Preface of 2nd Annual Meeting of SIAM Central States Section. Computers and Mathematics with Applications, 75(6):1851, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830021X.

Hur:2010:CCS

[HPY10] Junbeom Hur, Chanil Park, and Hyunsoo Yoon. Chosen ciphertext secure authenticated group communication using identitybased signcryption. Computers and Mathematics with Applications, 60(2): 362–375, July 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000167.

Hou:2019:SCS

[HQ19] Yanren Hou and Yi Qin. On the solution of coupled Stokes/Darcy model with Beavers–Joseph interface condition. Computers and Mathematics with Applications, 77(1): 50–65, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305236.

Holhos:2014:OEA

[HR14] Adrian Holhos and Daniela Rosca. An octahedral equal area partition of the sphere and near optimal configurations of points. Computers and Mathematics with Applications, 67(5):1092–1107, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000133.

Hintermuller:2015:SPF

[HR15] M. Hintermüller and J. Rasch. Several pathfollowing methods for a class of gradient constrained variational inequalities. Computers and Mathematics with Applications, 69(10):1045– 1067, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005872.

Heltai:2019:EEW

[HR19] Luca Heltai and Nella Rotundo. Error estimates in weighted Sobolev norms for finite element immersed interface methods. Computers and Mathematics with Applications, 78 (11):3586–3604, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302925.

Hu:2017:CEC

[HRHP17] Hongling Hu, Zhengyong Ren, Dongdong He, and Kejia Pan. On the convergence of an extrapolation cascadic multigrid method for elliptic problems. Computers and Mathematics with Applications, 74(4):759–771, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303280.

Hernandez:2010:SSV

[HRMS10] Elvira Hernández, Luis Rodríguez-Marín, and Miguel Sama. On solutions of set-valued optimization problems. Computers and Mathematics with Applications, 60 (5):1401–1408, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004360.

Hernandez:2012:AHB

[HRMS12] Elvira Hernández, Luis Rodríguez-Marín, and Miguel Sama. About Hahn–Banach extension theorems and applications to set-valued optimization. Computers and Mathematics with Applications, 64(6):1778–1788, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001277.

Ruan:2018:OPW

[hRWH18] Qi hua Ruan, Weihua Wang, and Qin Huang. An overdetermined problem for a weighted Poisson’s equation. Computers and Mathematics with Applications, 75(9):3139–3146, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830052X.

Hasegawa:2011:AMH

[HS11a] Takemitsu Hasegawa and Hiroshi Sugiura. An approximation method for high-order fractional derivatives of algebraically singular functions. Computers and Mathematics with Applications, 62(3): 930–937, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111002732.

Hussain:2011:KMC

[HS11b] N. Hussain and M. H.

Shah. KKM mappings in cone b-metric spaces. Computers and Mathematics with Applications, 62(4):1677–1684, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100472X.

Helal:2012:BFI

[HS12] M. A. Helal and A. R. Seadawy. Benjamin– Feir instability in nonlinear dispersive waves. Computers and Mathematics with Applications, 64(11):3557–3568, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005810.

Hessari:2013:LSP

[HS13] Peyman Hessari and Byeong-Chun Shin. The least-squares pseudo-spectral method for Navier–Stokes equations. Computers and Mathematics with Applications, 66 (3):318–329, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003118.

Hejranfar:2018:PWF

[HS18] Kazem Hejranfar and Mohammad Hossein Saadat. Preconditioned WENO finite-difference lattice Boltzmann method for simulation of incompressible turbulent flows. Computers and Mathematics with Applications, 76 (6):1427–1446, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303602.

Hou:2011:FRR

[HSBL11] Wen Hou, Lixin Song, Zhenhua Bao, and He Liu. Further results of recursive evaluation for compound distribution with the severity distribution of mixed type. Computers and Mathematics with Applications, 62(1):261– 271, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003956.

Hou:2017:LBB

[HSC17] Zhimin Hou, Baochang Shi, and Zhenhua Chai. A lattice Boltzmann based local feedback control approach for spiral wave. Computers and Mathematics with Applications, 74 (10):2330–2340, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117304273.

Han:2010:NSO

[HSD10] Xue Han, Wenyu Sun, and Chuangyin Dang. Nonmonotone secondorder Wolfe’s line search method for unconstrained optimization problems. Computers and Mathematics with Applications, 60(9):2517–2525, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005651.

Hessari:2015:ALS

[HSJ15] Peyman Hessari, Beyong-Chun Shin, and Bongsoo Jang. Analysis of least squares pseudospectral method for the interface problem of the Navier–Stokes equations. Computers and Mathematics with Applications, 69(8):838–851, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000577.

Hakeem:2011:NCS

[HSK11] A. K. Abdul Hakeem, S. Saravanan, and P. Kandaswamy. Natural convection in a square cavity due to thermally active plates for different boundary conditions. Computers and Mathematics with Applications, 62(1):491– 496, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004287.

Hussain:2012:CLJ

[HSMG12] Iqtadar Hussain, Tariq Shah, Hasan Mahmood, and Muhammad Asif Gondal. Construction of S_{8} Liu J S-boxes and their applications. Computers and Mathematics with Applications, 64(8):2450–2458, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004166.

Heyouni:2019:GHB

[HSMT19] Mohammed Heyouni, Farid Saberi-Movahed, and Azita Tajaddini. On global Hessenberg based methods for solving Sylvester matrix equations. Computers and Mathematics with Applications, 77(1): 77–92, January 1, 2019.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305273.

Han:2012:EOT

[HSMY12] Jinguang Han, Willy Susilo, Yi Mu, and Jun Yan. Efficient oblivious transfers with access control. Computers and Mathematics with Applications, 63(4):827–837, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010315.

Hung:2012:EMA

[HSS^{+}12] Shih-Hao Hung, Chi-Sheng Shih, Jeng-Peng Shieh, Chen-Pang Lee, and Yi-Hsiang Huang. Executing mobile applications on the cloud: Framework and issues. Computers and Mathematics with Applications, 63(2):573–587, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009084.

Hetmaniok:2011:CAD

[HSWZ11] Edyta Hetmaniok, Damian Slota, Roman Witula, and Adam Zielonka. Comparison of the Adomian decomposition method and the variational iteration method in solving the moving boundary problem. Computers and Mathematics with Applications, 61(8):1931–1934, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005420.

Hetmaniok:2015:USI

[HSZ15] Edyta Hetmaniok, Damian Slota, and Adam Zielonka. Using the swarm intelligence algorithms in solution of the twodimensional inverse Stefan problem. Computers and Mathematics with Applications, 69(4):347–361, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114006221.

Heuer:2012:RBF

[HT12a] Norbert Heuer and Thanh Tran. Radial basis functions for the solution of hypersingular operators on open surfaces. Computers and Mathematics with Applications, 63(11):1504–1518, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200243X.

Hsieh:2012:ODN

[HT12b] Wen-Hsiang Hsieh and Chia-Heng Tsai. Optimum design of a novel press system with Stephensoni mechanism. Computers and Mathematics with Applications, 64(5):897–907, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000739.

Heuer:2013:MMD

[HT13] Norbert Heuer and Thanh Tran. A mixed method for Dirichlet problems with radial basis functions. Computers and Mathematics with Applications, 66(10):2045–2055, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122113005245.

Hofer:2016:DGI

[HT16a] Christoph Hofer and Ioannis Toulopoulos. Discontinuous Galerkin isogeometric analysis of elliptic problems on segmentations with nonmatching interfaces. Computers and Mathematics with Applications, 72(7):1811–1827, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304448.

Huber:2016:RBF

[HT16b] Sarah E. Huber and Manfred R. Trummer. Radial basis functions for solving differential equations: Ill-conditioned matrices and numerical stability. Computers and Mathematics with Applications, 71(1):319–327, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005568.

Hanbay:2018:NAC

[HT18] Kazim Hanbay and Muhammed Fatih Talu. A novel active contour model for medical images via the Hessian matrix and eigenvalues. Computers and Mathematics with Applications, 75(9):3081–3104, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830049X.

Huang:2019:MES

[HT19] Shuibo Huang and Qiaoyu Tian. Marcinkiewicz estimates for solution to fractional elliptic Laplacian equation. Computers and Mathematics with Applications, 78(5):1732–1738, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302482.

Hannemann-Tamas:2013:MCR

[HTGSH13] Ralf Hannemann-Tamás, Attila Gábor, Gábor Szederkényi, and Katalin M. Hangos. Model complexity reduction of chemical reaction networks using mixed-integer quadratic programming. Computers and Mathematics with Applications, 65(10): 1575–1595, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112006852.

Hiraishi:2010:NSS

[HTL10] Masayuki Hiraishi, Michihisa Tsutahara, and R. C. K. Leung. Numerical simulation of sound generation in a mixing layer by the finite difference lattice Boltzmann method. Computers and Mathematics with Applications, 59(7):2403–2410, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006476.

He:2018:NIS

[HTM18] Chunhua He, Yaning Tang, and Jinli Ma. New interaction solutions for the (3 + 1)-dimensional Jimbo–Miwa equation. Computers and Mathematics with Applications, 76(9):2141–2147, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304309.

HosseinNia:2013:SFO

[HTV13] S. Hassan HosseinNia, Inés Tejado, and Blas M. Vinagre. Stability of fractional order switching systems. Computers and Mathematics with Applications, 66(5):585–596, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002903.

Huang:2015:BSS

[HTWS15] Zhi-Ruo Huang, Bo Tian, Yun-Po Wang, and Ya Sun. Bright soliton solutions and collisions for a (3+1)- dimensional coupled nonlinear Schrödinger system in optical-fiber communication. Computers and Mathematics with Applications, 69(12):1383– 1389, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001030.

Hu:2019:DBW

[HTY^{+}19] Cong-Cong Hu, Bo Tian, Hui-Min Yin, Chen-Rong Zhang, and Ze Zhang. Dark breather waves, dark lump waves and lump wave-soliton interactions for a (3 + 1)- dimensional generalized Kadomtsev–Petviashvili equation in a fluid. Computers and Mathematics with Applications, 78(1):166–177, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301002.

Hu:2015:RPO

[Hu15] Jun Hu. A robust prolongation operator for non-nested finite element methods. Computers and Mathematics with Applications, 69(3):235–246, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400594X.

Hu:2018:TGM

[Hu18]  Hanzhang Hu.  Two-

grid method for twodimensional nonlinear Schrödinger equation by mixed finite element method. Computers and Mathematics with Applications, 75(3):900–917, February 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117306661.

Hu:2019:STP

[Hu19]  Wenjie Hu.  Spatial-

temporal patterns of a two age structured population model with spatial non-locality. Computers and Mathematics with Applications, 78(1):123– 135, July 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211930104X.

Huang:2010:OCS

[Hua10a] Mugen Huang. Oscillation criteria for second order nonlinear dynamic equations with impulses. Computers and Mathematics with Applications, 59(1):31–41, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109002107.

Huang:2010:RSC

[Hua10b] Zhenyu Huang. A remark on the strong convergence of the overrelaxed proximal point algorithm. Computers and Mathematics with Applications, 60(6):1616–1619, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004578.

Huang:2011:RSD

[Hua11] Tingwen Huang. Robust stability of delayed fuzzy Cohen–Grossberg neural networks. Computers and Mathematics with Applications, 61(8):2247–2250, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007236.

Huang:2012:DCA

[Hua12] Kuo-Yi Huang. Detection and classification of areca nuts with machine vision. Computers and Mathematics with Applications, 64(5):739–746, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010248.

Hadi-Vencheh:2010:RBF

[HVA10] A. Hadi-Vencheh and M. Allame. On the relation between a fuzzy number and its centroid. Computers and Mathematics with Applications, 59(11):3578– 3582, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002403.

Hoang:2017:LOF

[HVO17] Thi-Thao-Phuong Hoang, Duc Cam Hai Vo, and Thanh Hai Ong. A low-order finite element method for three dimensional linear elasticity problems with general meshes. Computers and Mathematics with Applications, 74 (6):1379–1398, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303711.

Hernandez-Veron:2018:ELA

[HVR18] M. A. Hernández-Verón and N. Romero. Existence, localization and approximation of solution of symmetric algebraic Riccati equations. Computers and Mathematics with Applications, 76(1):187–203, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302128.

Hu:2011:EPD

[HW11] Guixin Hu and Ke Wang.

The estimation of probability distribution of SDE by only one sample trajectory. Computers and Mathematics with Applications, 62(4):1798–1806, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005025.

Houston:2016:AEM

[HW16] Paul Houston and Thomas P.

Wihler. Adaptive energy minimisation for hp-finite element methods. Computers and Mathematics with Applications, 71(4):977–990, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000055.

Hao:2019:GDR

[HW19a] Jianghao Hao and Fei Wang. General decay rate for weak viscoelastic wave equation with Balakrishnan– Taylor damping and timevarying delay. Computers and Mathematics with Applications, 78(8):2632–2640, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302081.

Hou:2019:SAM

[HW19b] Xiumei Hou and Junde Wu. Stability analysis of a multi-layer tumor model with free boundary. Computers and Mathematics with Applications, 77(1):199–208, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305340.

Hsu:2015:IGM

[HWH^{+}15] Ming-Chen Hsu, Chenglong Wang, Austin J. Herrema, Dominik Schillinger, Anindya Ghoshal, and Yuri Bazilevs. An interactive geometry modeling and parametric design platform for isogeometric analysis. Computers and Mathematics with Applications, 70(7):1481–1500, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001716.

Huang:2013:CPB

[HWW13] Wei-Xian Huang, Hua-Jing-Ling Wu, and Guo-Jin Wang. Constructing PDE-based surfaces bounded by geodesics or lines of curvature. Computers and Mathematics with Applications, 65(4):673–681, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006797.

Huang:2016:IPI

[HWXC16] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. Improved PPHSS iterative methods for solving nonsingular and singular saddle point problems. Computers and Mathematics with Applications, 72(1):92–109, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302280.

Huang:2018:ETS

[HWXC18a] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. An efficient two-step iterative method for solving a class of complex symmetric linear systems. Computers and Mathematics with Applications, 75(7):2473–2498, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307897.

Huang:2018:PGS

[HWXC18b] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. Parameterized generalized shift-splitting preconditioners for nonsymmetric saddle point problems. Computers and Mathematics with Applications, 75(2):349–373, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305643.

Huang:2019:PAG

[HWXC19] Zheng-Ge Huang, Li-Gong Wang, Zhong Xu, and Jing-Jing Cui. Preconditioned accelerated generalized successive overrelaxation method for solving complex symmetric linear systems. Computers and Mathematics with Applications, 77(7):1902–1916, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306771.

Hyun:2014:TOS

[HWY14] Jaeyub Hyun, Semyung Wang, and Sung Yang. Topology optimization of the shear thinning non-Newtonian fluidic systems for minimizing wall shear stress. Computers and Mathematics with Applications, 67(5):1154–1170, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113007074.

Huang:2011:ETL

[HWyL11] Haibo Huang, Lei Wang, and Xi yun Lu. Evaluation of three lattice Boltzmann models for multiphase flows in porous media. Computers and Mathematics with Applications, 61(12):3606– 3617, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004487.

Hu:2010:RAB

[HX10] Zhicheng Hu and Aimin Xu. Refinements of Aczél and Bellman’s inequalities. Computers and Mathematics with Applications, 59(9):3078–3083, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001367.

He:2011:FH

[HX11] Pengfei He and Xiaolong Xin. Fuzzy hyperlattices. Computers and Mathematics with Applications, 62(12):4682–4690, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009357.

Huang:2014:EMW

[HX14] Jincheng Huang and Zonghu Xiu. Existence and multiplicity of weak solutions for a singular quasilinear elliptic equation. Computers and Mathematics with Applications, 67(8):1450–1460, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000868.

Hu:2011:GSC

[HXL11] Jiaqian Hu, Jie Xin, and Hong Lu. The global solution for a class of systems of fractional nonlinear Schrödinger equations with periodic boundary condition. Computers and Mathematics with Applications, 62(3):1510–1521, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004378.

Huang:2015:NSS

[HXS^{+}15] Zhixiang Huang, Jie Xu, Bingbing Sun, Bo Wu, and Xianliang Wu. A new solution of Schrödinger equation based on symplectic algorithm. Computers and Mathematics with Applications, 69(11): 1303–1312, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115000929.

Han:2019:FLM

[HXX19] Yaozong Han, Xiangtuan Xiong, and Xuemin Xue. A fractional Landweber method for solving backward time-fractional diffusion problem. Computers and Mathematics with Applications, 78(1):81– 91, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300859.

Hamal:2010:SPS

[HY10] Nuket Aykut Hamal and Fulya Yoruk. Symmetric positive solutions of fourth order integral BVP for an increasing homeomorphism and homomorphism with sign-changing nonlinearity on time scales. Computers and Mathematics with Applications, 59(11):3603– 3611, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002452.

Hu:2011:HDE

[HY11] Chunping Hu and Xuefeng Yan. A hybrid differential evolution algorithm integrated with an ant system and its application. Computers and Mathematics with Applications, 62(1):32– 43, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100335X.

Hattaf:2013:GSR

[HY13] Khalid Hattaf and Noura Yousfi. Global stability for reaction–diffusion equations in biology. Computers and Mathematics with Applications, 66(8):1488–1497, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300518X.

Hattaf:2015:GHM

[HY15] Khalid Hattaf and Noura Yousfi. A generalized HBV model with diffusion and two delays. Computers and Mathematics with Applications, 69(1):31–40, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122114005549.

Hattaf:2016:NMD

[HY16] Khalid Hattaf and Noura Yousfi. A numerical method for delayed partial differential equations describing infectious diseases. Computers and Mathematics with Applications, 72 (11):2741–2750, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305363.

Han:2018:SCS

[HY18] Wei Han and Jiangyan Yao. The sign-changing solutions for a class of p-Laplacian Kirchhoff type problem in bounded domains. Computers and Mathematics with Applications, 76(7):1779–1790, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303948.

Huang:2018:LWI

[HYC18] Lili Huang, Yunfei Yue, and Yong Chen. Localized waves and interaction solutions to a (3 + 1)-dimensional generalized KP equation. Computers and Mathematics with Applications, 76(4):831–844, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302918.

He:2011:EMP

[HYCP11] Tieshan He, Fengjian Yang, Chuanyong Chen, and Shiguo Peng. Existence and multiplicity of positive solutions for nonlinear boundary value problems with a parameter. Computers and Mathematics with Applications, 61(11):3355– 3363, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003403.

Hung:2010:RCP

[HYL10] Wen-Liang Hung, Miin-Shen Yang, and E. Stanley Lee. A robust clustering procedure for fuzzy data. Computers and Mathematics with Applications, 60(1):151–165, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000324X.

Yuan:2011:CGF

[hYlLL11] Xue hai Yuan, Zeng liang Liu, and E. Stanley Lee. Center-of-gravity fuzzy systems based on normal fuzzy implications. Computers and Mathematics with Applications, 61(9):2879–2898, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002379.

Hu:2014:IME

[HYS^{+}14] Yang Hu, Haizhuan Yuan, Shi Shu, Xiaodong Niu, and Mingjun Li. An improved momentum exchanged-based immersed boundary-lattice Boltzmann method by using an iterative technique. Computers and Mathematics with Applications, 68(3):140–155, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002065.

Yuan:2010:DIF

[hYxLL10] Xue hai Yuan, Hong xing Li, and E. Stanley Lee. On the definition of the intuitionistic fuzzy subgroups. Computers and Mathematics with Applications, 59(9):3117–3129, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001525.

Hao:2010:LBB

[HZ10] Jian Hao and Luoding Zhu. A lattice Boltzmann based implicit immersed boundary method for fluid-structure interaction. Computers and Mathematics with Applications, 59(1):185–193, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005781.

Hilscher:2011:FOC

[HZ11] Roman Simon Hilscher and Vera Zeidan. First order conditions for generalized variational problems over time scales. Computers and Mathematics with Applications, 62(9):3490–3503, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007498.

He:2016:AFU

[HZ16] Xin-Jiang He and Song-Ping Zhu. An alternative form used to calibrate the Heston option pricing model. Computers and Mathematics with Applications, 71(9):1831– 1842, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300773.

He:2018:SFS

[HZ18] Xin-Jiang He and Song-Ping Zhu. A seriesform solution for pricing variance and volatility swaps with stochastic volatility and stochastic interest rate. Computers and Mathematics with Applications, 76 (9):2223–2234, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304425.

He:2017:HAC

[HZL17] Kejing He, Xiaoqiang Zhou, and Qian Lin. High accuracy complete elliptic integrals for solving the Hertzian elliptical contact problems. Computers and Mathematics with Applications, 73(1):122–128, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306101.

Hundertmark-Zauskova:2010:NSS

[HZLM10] A. Hundertmark-Zausková and M. Lukácová-Medvid’ová. Numerical study of sheardependent non-Newtonian fluids in compliant vessels. Computers and Mathematics with Applications, 60(3):572–590, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003445.

Huang:2011:SPS

[HZM11] Pengzhan Huang, Tong Zhang, and Xiaoling Ma. Superconvergence by L^{2}- projection for a stabilized finite volume method for the stationary Navier– Stokes equations. Computers and Mathematics with Applications, 62 (11):4249–4257, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008765.

Huang:2018:EES

[HZP18] Fenglin Huang, Zhong Zheng, and Yucheng Peng. Error estimates of the space–time spectral method for parabolic control problems. Computers and Mathematics with Applications, 75(2):335–348, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305667.

Ilic:2010:CNI

[IB10] Aleksandar Ilić and Milan Basić. On the chromatic number of integral circulant graphs. Computers and Mathematics with Applications, 60(1):144– 150, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003238.

Ibis:2011:NCM

[IB11] Birol Ibis and Mustafa Bayram. Numerical comparison of methods for solving fractional differential-algebraic equations (FDAEs). Computers and Mathematics with Applications, 62(8):3270–3278, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007164.

Imran:2010:FCP

[IBB10] Muhammad Imran, Syed Ahtsham Ul Haq Bokhary, and A. Q. Baig. On families of convex polytopes with constant metric dimension. Computers and Mathematics with Applications, 60(9):2629–2638, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006759.

Iacono:2012:SMM

[IBG12] M. Iacono, E. Barbierato, and M. Gribaudo. The SIMTHESys multiformalism modeling framework. Computers and Mathematics with Applications, 64(12):3828–3839, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002143.

Ibrahim:2011:HSN

[Ibr11] Rabha W. Ibrahim. On holomorphic solutions for nonlinear singular fractional differential equations. Computers and Mathematics with Applications, 62(3):1084–1090, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003385.

Ibrahima:2016:TOF

[Ibr16] Dione Ibrahima. Towards optimal finite element error estimates for the penalized Dirichlet problem in a domain with curved boundary. Computers and Mathematics with Applications, 71(1):76–84, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005143.

Islam:2011:OSF

[IBSS11] S. Islam, Z. Bano, I. Siddique, and A. M. Siddiqui. The optimal solution for the flow of a fourth-grade fluid with partial slip. Computers and Mathematics with Applications, 61(6):1507–1516, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000198.

Im:2012:MML

[IC12] Seokjin Im and Jin-Tak Choi. MLAIN: Multi-leveled air indexing scheme in non-flat wireless data broadcast for efficient window query processing. Computers and Mathematics with Applications, 64(5):1242–1251, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002738.

Itsariyawanich:2010:ECE

[ID10] Kan Itsariyawanich and Natasha Dejdumrong. Erratum to: “Conversion and evaluation for two types of parametric surface constructed by NTP bases” [Comput. Math. Appl. 49(2) (2005) 321– 329]. Computers and Mathematics with Applications, 59(1):595–601, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005884.

See [JW05].

Ilati:2016:RCG

[ID16] Mohammad Ilati and Mehdi Dehghan. Remediation of contaminated groundwater by meshless local weak forms. Computers and Mathematics with Applications, 72(9):2408–2416, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305119.

Idrees:2010:AOH

[IIHI10] M. Idrees, S. Islam, Sirajul Haq, and Sirajul Islam. Application of the optimal homotopy asymptotic method to squeezing flow. Computers and Mathematics with Applications, 59(12):3858–3866, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002944.

See corrigendum [IIHuI10].

Idrees:2010:CAO

[IIHuI10] M. Idrees, S. Islam, Sirajul Haq, and Siraj ul Islam. Corrigendum to “Application of the optimal homotopy asymptotic method to squeezing flow” [Comput. Math. Appl. 59 (2010) 3858– 3866]. Computers and Mathematics with Applications, 60(9):2724, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007054.

See [IIHI10].

Innami:2012:NBE

[IK12] Satoshi Innami and Hiroyuki Kasai. NMF-based environmental sound source separation using timevariant gain features. Computers and Mathematics with Applications, 64(5):1333–1342, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112002829.

Ibragimov:2016:EMF

[IK16] Akif Ibragimov and Thinh T. Kieu. An expanded mixed finite element method for generalized Forchheimer flows in porous media. Computers and Mathematics with Applications, 72(6):1467–1483, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303613.

Ikhile:2011:RBD

[Ikh11] M. N. O. Ikhile. The root and Bell’s disk iteration methods are of the same error propagation characteristics in the simultaneous determination of the zeros of a polynomial, part II: Round-off error analysis by use of interval arithmetic. Computers and Mathematics with Applications, 61(11):3191– 3217, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002860.

Iliuta:2013:DFB

[IL13] Ion Iliuta and Faical Larachi. Dynamics of fines/bacterial cells accumulation in tricklebed reactors/bioreactors — multiscale modeling framework. Computers and Mathematics with Applications, 65(10): 1698–1718, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113000278.

Ilic:2010:TML

[Ili10] Aleksandar Ilić. Trees with minimal Laplacian coefficients. Computers and Mathematics with Applications, 59(8):2776– 2783, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000829.

Ilic:2012:EPA

[Ili12]  Aleksandar Ilić.  On

the extremal properties of the average eccentricity. Computers and Mathematics with Applications, 64(9):2877–2885, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003720.

Irarrazaval:2011:FFT

[ILP^{+}11] Pablo Irarrazaval, Carlos Lizama, Vicente Parot, Carlos Sing-Long, and Cristian Tejos. The fractional Fourier transform and quadratic field magnetic resonance imaging. Computers and Mathematics with Applications, 62(3):1576–1590, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001799.

Iliev:2014:MAF

[ILP14] O. Iliev, Z. Lakdawala, and G. Printsypar. On a multiscale approach for filter efficiency simulations. Computers and Mathematics with Applications, 67(12):2171– 2184, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000935.

Iliev:2013:NSU

[ILS13] O. Iliev, Z. Lakdawala, and V. Starikovicius.

On a numerical subgrid upscaling algorithm for Stokes–Brinkman equations. Computers and Mathematics with Applications, 65(3):435–448, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004014.

Ibanez:2019:TMA

[ILV^{+}19] D. A. Ibanez, E. Love, T. E. Voth, J. R. Overfelt, N. V. Roberts, and G. A. Hansen. Tetrahedral mesh adaptation for Lagrangian shock hydrodynamics. Computers and Mathematics with Applications, 78(2):402– 416, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303377.

Imani:2017:TDL

[Ima17] Gholamreza Imani. Three dimensional lattice Boltzmann simulation of steady and transient finned natural convection problems with evaluation of different forcing and conjugate heat transfer schemes. Computers and Mathematics with Applications, 74(6):1362–1378, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303723.

Ionescu:2011:FOI

[IMD11] Clara Ionescu, Jose Tenreiro Machado, and Robin De Keyser. Fractionalorder impulse response of the respiratory system. Computers and Mathematics with Applications, 62(3):845–854, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003221.

Ilhan:2019:LWS

[IMS19] Onur Alp Ilhan, Jalil Manafian, and Mohammad Shahriari. Lump wave solutions and the interaction phenomenon for a variable-coefficient Kadomtsev–Petviashvili equation. Computers and Mathematics with Applications, 78 (8):2429–2448, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122119301798.

Izadi:2010:FPC

[IN10] Zahra Izadi and Kourosh Nourouzi. Fixed points of correspondences defined on cone metric spaces. Computers and Mathematics with Applications, 60(3):653–659, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003627.

Iomin:2018:SCQ

[Iom18] A. Iomin. Subdiffusion in classical and quantum nonlinear Schrödinger equations with disorder. Computers and Mathematics with Applications, 73(6):914–930, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303431.

Ipek:2012:UBC

[Ipe12]  Ahmet Ipek.  Upper

bounds for the condition numbers of the GCD and the reciprocal GCD matrices in spectral norm. Computers and Mathematics with Applications, 63(3):645–651, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009825.

Iapichino:2016:RBM

[IQR16] Laura Iapichino, Alfio Quarteroni, and Gianluigi Rozza. Reduced basis method and domain decomposition for elliptic problems in networks and complex parametrized geometries. Computers and Mathematics with Applications, 71(1):408–430, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005696.

Ivanova:2012:FOU

[IS12] R. Ivanova and G. Simeonov. A formula for the oxygen uptake of thin tissue slice in terms of its surface oxygen tension. Computers and Mathematics with Applications, 64(3):322–336, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122112001691.

Ibrahim:2014:ECV

[IS14] Moustafa Ibrahim and Mazen Saad. On the efficacy of a control volume finite element method for the capture of patterns for a volume-filling chemotaxis model. Computers and Mathematics with Applications, 68 (9):1032–1051, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001333.

Iyiola:2018:EIM

[IW18] O. S. Iyiola and B. A.

Wade. Exponential integrator methods for systems of non-linear spacefractional models with super-diffusion processes in pattern formation. Computers and Mathematics with Applications, 75(10):3719–3736, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301081.

Izsak:2015:ENE

[Izs15] Ferenc Izsák. Energy norm error estimates for averaged discontinuous Galerkin methods: Multidimensional case. Computers and Mathematics with Applications, 70(4):705–725, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002941.

Jafari:2011:ASM

[JA11] M. A. Jafari and A. Aminataei. An algorithm for solving multi-term diffusion-wave equations of fractional order. Computers and Mathematics with Applications, 62(3):1091–1097, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100229X.

Jones:2018:SMF

[JAJ18] G. Sosa Jones, J. Arteaga, and O. Jiménez. A study of mimetic and finite difference methods for the static diffusion equation. Computers and Mathematics with Applications, 76(3):633–648, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302645.

Jalilian:2014:EMS

[Jal14] Y. Jalilian. On the existence and multiplicity of solutions for a class of singular elliptic problems. Computers and Mathematics with Applications, 68(6):664–680, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003332.

Jankowski:2010:FOF

[Jan10] Tadeusz Jankowski. Firstorder functional difference equations with nonlinear boundary value problems. Computers and Mathematics with Applications, 59(6):1937–1943, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900738X.

Ao:2012:MRS

[jASzZ12] Ji jun Ao, Jiong Sun, and Mao zhu Zhang. Matrix representations of Sturm–Liouville problems with transmission conditions. Computers and Mathematics with Applications, 63(8):1335–1348, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000727.

Javidi:2011:MCP

[Jav11] M. Javidi. A modified Chebyshev pseudospectral DD algorithm for the GBH equation. Computers and Mathematics with Applications, 62 (9):3366–3377, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007243.

Jaworska:2013:ICN

[Jaw13] Irena Jaworska. On the ill-conditioning in the new higher order multipoint method. Computers and Mathematics with Applications, 66(3):238–249, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300254X.

Jin:2017:GIT

[JBBL17] Hongwei Jin, Minru Bai, Julio Benítez, and Xiaoji Liu. The generalized inverses of tensors and an application to linear models. Computers and Mathematics with Applications, 74(3):385–397, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302481.

Cheng:2011:CTW

[jC11] Yan jun Cheng. Classification of traveling wave solutions to the Vakhnenko equations. Computers and Mathematics with Applications, 62(10):3987–3996, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008418.

Jiao:2012:SFO

[JC12] Zhuang Jiao and Yang Quan Chen. Stability of fractional-order linear time-invariant systems with multiple noncommensurate orders. Computers and Mathematics with Applications, 64 (10):3053–3058, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008789.

Jibben:2019:PFT

[JCF19] Z. Jibben, N. N. Carlson, and M. M. François. A paraboloid fitting technique for calculating curvature from piecewiselinear interface reconstructions on 3D unstructured meshes. Computers and Mathematics with Applications, 78(2):643– 653, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305194.

Jiang:2016:EDO

[JCWZ16] Zhengxian Jiang, Baotong Cui, Wei Wu, and Bo Zhuang. Event-driven observer-based control for distributed parameter systems using mobile sensor and actuator. Computers and Mathematics with Applications, 72(12):2854–2864, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305673.

Jin:2013:CFG

[JCZZ13] Hai Jin, Ge Cheng, Deqing Zou, and Xinwen Zhang. Cherub: Finegrained application protection with on-demand virtualization. Computers and Mathematics with Applications, 65(9):1326– 1338, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001149.

Jiang:2012:NSF

[JD12] Yao-Lin Jiang and Xiao-Li Ding. Nonnegative solutions of fractional functional differential equations. Computers and Mathematics with Applications, 63(5):896–904, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010388.

Ji:2014:ABI

[JFC14] Ruihong Ji, Mingshu Fan, and Hui Chen. Asymptotic behavior at infinity of the stationary solution to a semilinear heat equation. Computers and Mathematics with Applications, 67(6):1289–1292, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000406.

Jourabian:2014:EMP

[JFS14]  Mahmoud  Jourabian,

Mousa Farhadi, and Kurosh Sedighi. On the expedited melting of phase change material (PCM) through dispersion of nanoparticles in the thermal storage unit. Computers and Mathematics with Applications, 67(7):1358–1372, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000625.

See retraction notice [JFS20].

Jourabian:2020:RNE

[JFS20]  Mahmoud  Jourabian,

Mousa Farhadi, and Kurosh Sedighi. Retraction notice to “On the expedited melting of phase change material (PCM) through dispersion of nanoparticles in the thermal storage unit” [Comput. Math. Appl. 67(7) (2014) 1358– 1372]. Computers and Mathematics with Applications, 80(12):3193, December 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122120304144.

See [JFS14].

Janssen:2013:ENL

[JGK13] Christian F. Janßen, Stephan T. Grilli, and Manfred Krafczyk. On enhanced non-linear free surface flow simulations with a hybrid LBM–VOF model. Computers and Mathematics with Applications, 65(2):211–229, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004026.

Jafari:2010:HAM

[JGSS10] H. Jafari, A. Golbabai, S. Seifi, and K. Sayevand. Homotopy analysis method for solving multi-term linear and nonlinear diffusion-wave equations of fractional order. Computers and Mathematics with Applications, 59(3):1337–1344, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003952.

Jiang:2010:IHT

[JH10] Weidong Jiang and Yun Hua. On improvement of a Hilbert-type inequality. Computers and Mathematics with Applications, 60(3):629–633, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003585.

Hua:2011:GCF

[jHlXZ11] Xiu juan Hua, Xiao long Xin, and Xi Zhu. Generalized (convex) fuzzy sublattices. Computers and Mathematics with Applications, 62(2):699– 708, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004627.

Jiang:2015:SEI

[JHW15] Fengze Jiang, Chengming Huang, and Xiaojie Wang. Stochastic exponential integrator for finite element spatial discretization of stochastic elastic equation. Computers and Mathematics with Applications, 69(8):817– 827, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000656.

Ji:2014:COG

[Ji14]  Jun Ji.  Computing

the outer and group inverses through elementary row operations. Computers and Mathematics with Applications, 68(6):655–663, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003344.

Jiang:2011:CIV

[Jia11] Yuncheng Jiang. Corrigendum to “Intervalvalued intuitionistic fuzzy soft sets and their properties” [Comput. Math. Appl. 60 (2010) 906– 918]. Computers and Mathematics with Applications, 61(10):3179, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002884.

See [JTC^{+}10a, WYG12].

Jiang:2012:ERF

[Jia12] Heping Jiang. Existence results for fractional order functional differential equations with impulse. Computers and Mathematics with Applications, 64(10):3477–3483, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001812.

Jichang:2010:SNE

[Jic10] Kuang Jichang. Some new extensions on H-G-A inequality. Computers and Mathematics with Applications, 60(8):2204–2205, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005699.

Jia:2013:SPM

[JJ13] Jiteng Jia and Yaolin Jiang. A structure preserving matrix factorization for solving general periodic pentadiagonal Toeplitz linear systems. Computers and Mathematics with Applications, 66(6):965–974, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003982.

Jia:2015:TSA

[JJ15] Jiteng Jia and Yaolin Jiang. Two symbolic algorithms for solving general periodic pentadiagonal linear systems. Computers and Mathematics with Applications, 69(9):1020– 1029, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001042.

Jha:2019:SSN

[JJ19] Abhinav Jha and Volker John. A study of solvers for nonlinear AFC discretizations of convection–diffusion equations. Computers and Mathematics with Applications, 78(9):3117–3138, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302287.

Jeong:2011:RNV

[JJC11] Jin-Mun Jeong, Eun-Young Ju, and Su-Jin Cheon. Regularity for nonlinear variational inequalities of hyperbolic type. Computers and Mathematics with Applications, 62 (11):4230–4237, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008728.

Jia:2016:MTG

[JJH16] Hongen Jia, Huiyong Jia, and Yunqing Huang. A modified two-grid decoupling method for the mixed Navier–Stokes/ Darcy model. Computers and Mathematics with Applications, 72(4):1142–1152, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303650.

Janssen:2010:LBA

[JK10] Christian Janssen and Manfred Krafczyk. A lattice Boltzmann approach for free-surface-flow simulations on nonuniform block-structured grids. Computers and Mathematics with Applications, 59(7):2215–2235, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006488.

Janssen:2011:FSF

[JK11a] Christian Janßen and Manfred Krafczyk. Free surface flow simulations on GPGPUs using the LBM. Computers and Mathematics with Applications, 61(12):3549– 3563, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001684.

Jun:2011:FGT

[JK11b] Young Bae Jun and Min Su Kang. Fuzzifications of generalized Tarski filters in Tarski algebras. Computers and Mathematics with Applications, 61(1):1–7, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008035.

Jun:2011:FPI

[JK11c] Young Bae Jun and Min Su Kang. Fuzzy positive implicative ideals of BCK-algebras based on the theory of falling shadows. Computers and Mathematics with Applications, 61(1):62–67, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008199.

Joshi:2012:MMC

[JK12] Pratibha Joshi and Manoj Kumar. Mathematical model and computer simulation of three dimensional thin film elliptic interface problems. Computers and Mathematics with Applications, 63(1):25–35, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009308.

Juttler:2018:CAR

[JK18] Bert Jüttler and Stefan K. Kleiss. Coupling adaptively refined multipatch spline discretizations via boundary compatibility. Computers and Mathematics with Applications, 74(7):1626–1647, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302316.

Jabbari:2011:ESC

[JKB11] A. Jabbari, H. Kheiri, and A. Bekir. Exact solutions of the coupled Higgs equation and the Maccari system using He’s semi-inverse method and (\frac{G′}{G})-expansion method. Computers and Mathematics with Applications, 62(5):2177–2186, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005578.

Jun:2010:FSH

[JKK10] Young Bae Jun, Min Su Kang, and Hee Sik Kim. Fuzzy structures of hyper-MV-deductive systems in hyper-MV-algebras. Computers and Mathematics with Applications, 59(8):2982–2989, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110001252.

Jaberipour:2011:PSA

[JKK11] Majid Jaberipour, Esmaile Khorram, and Behrooz Karimi. Particle swarm algorithm for solving systems of nonlinear equations. Computers and Mathematics with Applications, 62(2):566– 576, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004299.

Jafari:2012:TWS

[JKK12] H. Jafari, N. Kadkhoda, and C. M. Khalique. Travelling wave solutions of nonlinear evolution equations using the simplest equation method. Computers and Mathematics with Applications, 64(6):2084–2088, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003409.

Jurlewicz:2012:FGE

[JKMS12] A. Jurlewicz, P. Kern, M. M. Meerschaert, and H.-P. Scheffler. Fractional governing equations for coupled random walks. Computers and Mathematics with Applications, 64(10):3021–3036, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008741.

Jangveladze:2010:LTA

[JKN10] Temur Jangveladze, Zurab Kiguradze, and Beny Neta. Large time asymptotic and numerical solution of a nonlinear diffusion model with memory. Computers and Mathematics with Applications, 59(1):254–273, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005112.

Jia:2012:NAS

[JKS12] Jiteng Jia, Qiongxiang Kong, and Tomohiro Sogabe. A new algorithm for solving nearly pentadiagonal Toeplitz linear systems. Computers and Mathematics with Applications, 63(7):1238–1243, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211101100X.

Jleli:2018:NGS

[JKS18] Mohamed Jleli, Mokhtar Kirane, and Bessem Samet. Nonexistence of global solutions for a class of nonlocal in time and space nonlinear evolution equations. Computers and Mathematics with Applications, 75(8):2698–2709, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300051.

Jleli:2019:AGS

[JKS19] Mohamed Jleli, Mokhtar Kirane, and Bessem Samet. On the absence of global solutions for quantum versions of Schrödinger equations and systems. Computers and Mathematics with Applications, 77(3):740–751, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305972.

Jamil:2011:TFO

[JKZ11] M. Jamil, N. A. Khan, and A. A. Zafar. Translational flows of an Oldroyd-b fluid with fractional derivatives. Computers and Mathematics with Applications, 62(3):1540–1553, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002690.

Ji:2011:ERI

[JL11a] Shaochun Ji and Gang Li.

Existence results for impulsive differential inclusions with nonlocal conditions. Computers and Mathematics with Applications, 62(4):1908–1915, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005141.

Jia:2011:TNS

[JL11b] Mei Jia and Xiping Liu.

Three nonnegative solutions for fractional differential equations with integral boundary conditions. Computers and Mathematics with Applications, 62(3):1405–1412, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001787.

Jayabal:2012:EES

[JL12] Raymond J. Jayabal and Chiew Tong Lau. On the exact expression for single-stage CSMA… CA’s idle period distribution. Computers and Mathematics with Applications, 64(5):1343–1351, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002830.

Jia:2015:IDG

[JL15a] Jiteng Jia and Sumei Li. On the inverse and determinant of general bordered tridiagonal matrices. Computers and Mathematics with Applications, 69(6):503–509, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500036X.

Jia:2015:SAI

[JL15b] Jiteng Jia and Sumei Li. Symbolic algorithms for the inverses of general k-tridiagonal matrices. Computers and Mathematics with Applications, 70(12):3032–3042, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005131.

Jiang:2016:AMT

[JL16] Nan Jiang and William Layton. Algorithms and models for turbulence not at statistical equilibrium. Computers and Mathematics with Applications, 71(11):2352– 2372, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004861.

Jia:2017:ENA

[JL17a] Jiteng Jia and Sumei Li.

An efficient numerical algorithm for the determinant of a cyclic pentadiagonal Toeplitz matrix. Computers and Mathematics with Applications, 74(12):2992–2999, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117304649.

Jia:2017:DCP

[JL17b] Jiteng Jia and Sumei Li. On determinants of cyclic pentadiagonal matrices with Toeplitz structure. Computers and Mathematics with Applications, 73(2):304–309, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306630.

Jiang:2018:STS

[JL18] Wei Jiang and Hui Li. A space–time spectral collocation method for the two-dimensional variableorder fractional percolation equations. Computers and Mathematics with Applications, 75(10):3508–3520, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300865.

Jia:2019:NAN

[JL19] Jiteng Jia and Sumei Li. New algorithms for numerically solving a class of bordered tridiagonal systems of linear equations. Computers and Mathematics with Applications, 78(1):144–151, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301026.

See retraction notice [JL20].

Jia:2020:RNN

[JL20] Jiteng Jia and Sumei Li. Retraction notice to “New algorithms for numerically solving a class of bordered tridiagonal systems of linear equations” [Comput. Math. Appl. 78 (1) (2019) 144–151]. Computers and Mathematics with Applications, 80(6):1744, September 15, 2020. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122120303278.

See [JL19].

Jiang:2010:AEP

[JLCS10] Hao Jiang, Shengguo Li, Lizhi Cheng, and Fang Su. Accurate evaluation of a polynomial and its derivative in Bernstein form. Computers and Mathematics with Applications, 60(3):744–755, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003706.

Jiang:2019:LBM

[JLD19] Jibing Jiang, Dinggen Li, and Ruzhen Dou. A lattice Boltzmann modeling and analysis of the thermal convection in a lithium-ion battery. Computers and Mathematics with Applications, 77(10):2695–2706, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300070.

Jin:2017:GET

[JLF17] Lingyu Jin, Lang Li, and Shaomei Fang. The global existence and timedecay for the solutions of the fractional pseudoparabolic equation. Computers and Mathematics with Applications, 73(10):2221–2232, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122117301670.

Jun:2011:CSA

[JLK11] Young Bae Jun, Kyoung Ja Lee, and Min Su Kang. Cubic structures applied to ideals of BCIalgebras. Computers and Mathematics with Applications, 62(9):3334–3342, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007152.

Jin:2018:REO

[JLL18] Guanghao Jin, Young-Ju Lee, and Hengguang Li. Regularity estimates and optimal finite element methods in W_{p}^{1} on polygonal domains. Computers and Mathematics with Applications, 74(9):2089–2105, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300809.

Jiang:2019:GEA

[JLL19] Kerui Jiang, Zhi Ling, and Zuhan Liu. Global existence and asymptotic behavior of the fractional chemotaxis system with signal-dependent sensitivity. Computers and Mathematics with Applications, 78 (10):3450–3470, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302810.

Jun:2010:FSS

[JLP10a] Young Bae Jun, Kyoung Ja Lee, and Chul Hwan Park. Fuzzy soft set theory applied to BCK/ BCI-algebras. Computers and Mathematics with Applications, 59(9):3180– 3192, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001641.

Jun:2010:NTF

[JLP10b] Young Bae Jun, Kyoung Ja Lee, and Chul Hwan Park. New types of fuzzy ideals in BCK/BCIalgebras. Computers and Mathematics with Applications, 60(3):771–785, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110003810.

Jiang:2012:ASM

[JLTB12] H. Jiang, F. Liu, I. Turner, and K. Burrage. Analytical solutions for the multi-term timefractional diffusion-wave/ diffusion equations in a finite domain. Computers and Mathematics with Applications, 64 (10):3377–3388, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001678.

Jia:2018:AES

[JLWX18] Yunfeng Jia, Bimei Luo, Jianhua Wu, and Hong-Kun Xu. Analysis on the existence of the steadystates for an ecologicalmathematical model with predator-prey-dependent functional response. Computers and Mathematics with Applications, 76(7):1767–1778, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303936.

Jaskowiec:2015:EIA

[JM15] J. Jaśkowiec and S. Milewski. The effective interface approach for coupling of the FE and meshless FD methods and applying essential boundary conditions. Computers and Mathematics with Applications, 70(5):962–979, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003065.

Jaskowiec:2016:CFE

[JM16] J. Jaśkowiec and S. Milewski. Coupling finite element method with meshless finite difference method in thermomechanical problems. Computers and Mathematics with Applications, 72(9):2259–2279, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304746.

Jesus:2010:CHD

[JMB10] Isabel S. Jesus, J. A. Tenreiro Machado, and Ramiro S. Barbosa. Control of a heat diffusion system through a fractional order nonlinear algorithm.

Computers and Mathematics with Applications, 59(5):1687–1694, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005458.

Jia:2013:SMT

[JMDA13] Xinli Jia, John B. McLaughlin, Jos Derksen, and Goodarz Ahmadi. Simulation of a mannequin’s thermal plume in a small room. Computers and Mathematics with Applications, 65(2):287–295, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005372.

Jari:2013:SDG

[JMHF13] Rabeea Jari, Lin Mu, Anna Harris, and Lynn Fox. Superconvergence for discontinuous Galerkin finite element methods by L^{2}-projection methods. Computers and Mathematics with Applications, 65(4):665–672, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112006748.

Jafarpour:2011:CSC

[JMLF11] Morteza Jafarpour, Seyed S. Mousavi, and V. Leoreanu-Fotea. A class of semihypergroups connected to preordered weak Γ-semigroups. Computers and Mathematics with Applications, 62(8):2944–2949, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006468.

Jin:2019:ICF

[JMNZ19] Zhengmeng Jin, Lihua Min, Michael K. Ng, and Minling Zheng. Image colorization by fusion of color transfers based on DFT and variance features. Computers and Mathematics with Applications, 77(9):2553–2567, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307314.

Jimenez:2011:LVA

[JMST11] Jorge Jiménez, Susana Montes, Branimir Seselja, and Andreja Tepavcević.

Lattice-valued approach to closed sets under fuzzy relations: Theory and applications. Computers and Mathematics with Applications, 62 (10):3729–3740, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007899.

Jog:2014:MFE

[JN14] C. S. Jog and Arup Nandy. Mixed finite elements for electromagnetic analysis. Computers and Mathematics with Applications, 68(8):887–902, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003666.

Jafari:2013:NAS

[JNBK13] H. Jafari, M. Nazari, D. Baleanu, and C. M. Khalique. A new approach for solving a system of fractional partial differential equations. Computers and Mathematics with Applications, 66(5):838–843, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200675X.

Jalaal:2011:HPM

[JNJ^{+}11] M. Jalaal, M. G. Nejad, P. Jalili, M. Esmaeilpour, H. Bararnia, E. Ghasemi, Soheil Soleimani, D. D. Ganji, and S. M. Moghimi. Homotopy perturbation method for motion of a spherical solid particle in plane Couette fluid flow. Computers and Mathematics with Applications, 61(8):2267–2270, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000739X.

Jureczka:2019:NAS

[JO19] Michal Jureczka and Anna Ochal. Numerical analysis and simulations of contact problem with wear. Computers and Mathematics with Applications, 77(11):2980– 2988, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830470X.

Jawad:2011:SSN

[JPB11] Anwar Ja’afar Mohamad Jawad, Marko D. Petković, and Anjan Biswas. Soliton solutions for nonlinear Calaogero–Degasperis and potential Kadomtsev– Petviashvili equations. Computers and Mathematics with Applications, 62(6):2621–2628, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100650X.

Jeong:2013:TSB

[JPCY13] Seung Hyun Jeong, Seon Ho Park, Dong-Hoon Choi, and Gil Ho Yoon. Toward a stress-based topology optimization procedure with indirect calculation of internal finite element information. Computers and Mathematics with Applications, 66(6):1065–1081, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004355.

Jeong:2017:EDR

[JPK17] Jin-Mun Jeong, Jong Yeoul Park, and Yong Han Kang. Energy decay rates for the semilinear wave equation with memory boundary condition and acoustic boundary conditions. Computers and Mathematics with Applications, 73(9):1975–1986, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301177.

Jeong:2018:GNS

[JPK18] Jin-Mun Jeong, Jong-Yeoul Park, and Yong Han Kang. Global nonexistence of solutions for a nonlinear wave equation with time delay and acoustic boundary conditions. Computers and Mathematics with Applications, 76(3):661–671, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302669.

Jereva:2012:DEM

[JPP12] Dessislava Jereva, Ilza Pajeva, and Tania Pencheva. Data extraction module — a supplementary tool for the AM-MOS ProtLig software package. Computers and Mathematics with Applications, 64(3):266–271, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000818.

Jiang:2010:ESG

[JPS10] Si-Jia Jiang, Li-Ping Pang, and Jie Shen. Existence of solutions of generalized vector variational-type inequalities with set-valued mappings. Computers and Mathematics with Applications, 59(4):1453–1461, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007408.

Jeon:2014:HFE

[JPS14] Y. Jeon, E.-J. Park, and D. Sheen. A hybridized finite element method for the Stokes problem. Computers and Mathematics with Applications, 68(12):2222–2232, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003654.

Jinfeng:2014:LBM

[JQG14] Zhang Jinfeng, Zhang Qinghe, and Qiao Guangquan. A lattice Boltzmann model for the nonequilibrium flocculation of cohesive sediments in turbulent flow. Computers and Mathematics with Applications, 67(2):381–392, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002290.

Jhangeer:2012:ASC

[JQSS12] Adil Jhangeer, M. N. Qureshi, S. Sial, and S. Sharif. Analytic solutions and conserved quantities of wave equation on torus. Computers and Mathematics with Applications, 64(6):1627–1635, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000211.

Jaust:2018:FAG

[JRA^{+}18] Alexander Jaust, Balthasar Reuter, Vadym Aizinger, Jochen Schütz, and Peter Knabner. FES-TUNG: a MATLAB/ GNU Octave toolbox for the discontinuous Galerkin method. Part III: Hybridized discontinuous Galerkin (HDG) formulation. Computers and Mathematics with Applications, 75(12):4505–4533, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301895.

Jain:2015:ISI

[JRB15] Subit K. Jain, Rajendra K. Ray, and Arnav Bhavsar. Iterative solvers for image denoising with diffusion models: a comparative study. Computers and Mathematics with Applications, 70(3):191–211, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001789.

Jablan:2012:MCK

[JRSZ12] Slavik Jablan, Ljiljana Radović, Radmila Sazdanović, and Ana Zeković. Mirror-curves and knot mosaics. Computers and Mathematics with Applications, 64(4):527–543, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010984.

Jamil:2011:NEA

[JRZK11] M. Jamil, A. Rauf, A. A. Zafar, and N. A. Khan. New exact analytical solutions for Stokes’ first problem of Maxwell fluid with fractional derivative approach. Computers and Mathematics with Applications, 62(3):1013–1023, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100174X.

Jiang:2011:NEU

[JS11] Feng Jiang and Yi Shen. A note on the existence and uniqueness of mild solutions to neutral stochastic partial functional differential equations with non-Lipschitz coefficients. Computers and Mathematics with Applications, 61(6):1590–1594, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000447.

Jamali:2012:PSD

[JS12a] Shahram Jamali and Gholam Shaker. PSO– SFDD: Defense against SYN flooding DoS attacks by employing PSO algorithm. Computers and Mathematics with Applications, 63(1):214–221, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009783.

Jansri:2012:EMF

[JS12b] Anurak Jansri and Pitikhate Sooraksa. Enhanced model and fuzzy strategy of air to fuel ratio control for spark ignition engines. Computers and Mathematics with Applications, 64(5):922–933, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000764.

Jia:2013:ITM

[JSEM13] Jiteng Jia, Tomohiro Sogabe, and Moawwad El-Mikkawy. Inversion of k-tridiagonal matrices with Toeplitz structure. Computers and Mathematics with Applications, 65(1):116–125, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006475.

Jackson:2016:ARF

[JSGP16] S. J. Jackson, D. Stevens, D. Giddings, and H. Power. An adaptive RBF finite collocation approach to track transport processes across moving fronts. Computers and Mathematics with Applications, 71(1):278–300, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005581.

Jiang:2010:IVI

[JTC^{+}10a] Yuncheng Jiang, Yong Tang, Qimai Chen, Hai Liu, and Jianchao Tang. Interval-valued intuitionistic fuzzy soft sets and their properties. Computers and Mathematics with Applications, 60(3):906–918, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003949.

See corrigenda [Jia11, Wan13a] and notes [WYG12].

Jiang:2010:ESS

[JTC^{+}10b] Yuncheng Jiang, Yong Tang, Qimai Chen, Ju Wang, and Suqin Tang. Extending soft sets with description logics. Computers and Mathematics with Applications, 59(6):2087– 2096, March 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109007615.

Jiang:2011:SOM

[JTCC11] Yuncheng Jiang, Yong Tang, Qimai Chen, and Zhanmao Cao. Semantic operations of multiple soft sets under conflict. Computers and Mathematics with Applications, 62(4):1923–1939, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005165.

Jumarie:2010:DSS

[Jum10] Guy Jumarie. Derivation and solutions of some fractional Black– Scholes equations in coarse-grained space and time. application to Merton’s optimal portfolio. Computers and Mathematics with Applications, 59(3):1142–1164, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003599.

Jun:2010:NFI

[Jun10] Young Bae Jun. Note on “(α, β)-fuzzy ideals of hemirings”. Computers and Mathematics with Applications, 59(8):2582– 2586, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000398.

See [DSA09].

Jibben:2019:MST

[JVMF19] Z. Jibben, J. Velechovsky, T. Masser, and M. M. François. Modeling surface tension in compressible flow on an adaptively refined mesh. Computers and Mathematics with Applications, 78(2):504– 516, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305248.

Jiang:2005:CET

[JW05] Su-Rong Jiang and Guo-Jin Wang. Conversion and evaluation for two types of parametric surfaces constructed by NTP bases. Computers and Mathematics with Applications, 49 (2–3):321–329, January/ February 2005. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122105000362.

See erratum [ID10].

Jiang:2010:TRB

[JW10] Wei Jiang and Gang Wu. A thick-restarted block Arnoldi algorithm with modified Ritz vectors for large eigenproblems. Computers and Mathematics with Applications, 60(3):873–889, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003913.

Jia:2011:CCH

[JW11] Nuo Jia and Tao Wang.

Chaos control and hybrid projective synchronization for a class of new chaotic systems. Computers and Mathematics with Applications, 62 (12):4783–4795, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100945X.

Jin:2015:WFQ

[JW15] Chunhua Jin and Xinwei Wang. Weak form quadrature element method for accurate free vibration analysis of thin skew plates. Computers and Mathematics with Applications, 70(8):2074–2086, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004009.

Ji:2018:DIE

[JW18a] Jun Ji and Yimin Wei. The Drazin inverse of an even-order tensor and its application to singular tensor equations. Computers and Mathematics with Applications, 75(9):3402–3413, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300798.

Jia:2018:FFD

[JW18b] Jinhong Jia and Hong Wang. A fast finite difference method for distributed-order spacefractional partial differential equations on convex domains. Computers and Mathematics with Applications, 75(6):2031–2043, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305515.

Jia:2019:FFV

[JW19a] Jinhong Jia and Hong Wang. A fast finite volume method for conservative space–time fractional diffusion equations discretized on space–time locally refined meshes. Computers and Mathematics with Applications, 78(5):1345–1356, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302019.

Jin:2019:ACM

[JW19b] Ri Jin and Guirong Weng.

Active contour model based on improved fuzzy c-means algorithm and adaptive functions. Computers and Mathematics with Applications, 78 (11):3678–3691, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930313X.

Jiang:2013:CSO

[JWX^{+}13] Mingfeng Jiang, Yaming Wang, Ling Xia, Feng Liu, Shanshan Jiang, and Wenqing Huang. The combination of selforganizing feature maps and support vector regression for solving the inverse ECG problem. Computers and Mathematics with Applications, 66(10):1981–1990, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005610.

Jia:2014:PSL

[JWX14] Yunfeng Jia, Jianhua Wu, and Hong-Kun Xu. Positive solutions of a Lotka–Volterra competition model with crossdiffusion. Computers and Mathematics with Applications, 68 (10):1220–1228, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004088.

Jin:2010:GTF

[JXZ^{+}10] Hai Jin, Guofu Xiang, Deqing Zou, Feng Zhao, Min Li, and Chen Yu. A guest-transparent file integrity monitoring method in virtualization environment. Computers and Mathematics with Applications, 60(2):256– 266, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000180.

Jun:2011:NPC

[JY11] Young Bae Jun and Xibei Yang. A note on the paper “Combination of interval-valued fuzzy set and soft set” [Comput. Math. Appl. 58 (2009) 521–527]. Computers and Mathematics with Applications, 61(5):1468–1470, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/

                                        pii/S0898122111000034.

See [YLY^{+}09].

Jafari:2011:ALW

[JYF^{+}11] H. Jafari, S. A. Yousefi, M. A. Firoozjaee, S. Momani, and C. M. Khalique. Application of Legendre wavelets for solving fractional differential equations. Computers and Mathematics with Applications, 62(3):1038–1045, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003257.

Jeon:2016:VVG

[JYK16] Junkee Jeon, Ji-Hun Yoon, and Myungjoo Kang. Valuing vulnerable geometric Asian options. Computers and Mathematics with Applications, 71(2):676–691, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115006069.

Jia:2016:HRR

[JYL16] Jiteng Jia, Boting Yang, and Sumei Li. On a homogeneous recurrence relation for the determinants of general pentadiagonal Toeplitz matrices. Computers and Mathematics with Applications, 71(4):1036–1044, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300232.

Jiang:2016:HCU

[JYYL16] Zhen-Hua Jiang, Chao Yan, Jian Yu, and Yansu Li. Hybrid central-upwind finite volume schemes for solving the Euler and Navier– Stokes equations. Computers and Mathematics with Applications, 72 (9):2241–2258, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630476X.

Jiao:2011:ESC

[JZ11] Feng Jiao and Yong Zhou.

Existence of solutions for a class of fractional boundary value problems via critical point theory. Computers and Mathematics with Applications, 62(3):1181–1199, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002653.

Jiao:2012:RSF

[JZ12] Zhuang Jiao and Yisheng Zhong. Robust stability for fractional-order systems with structured and unstructured uncertainties. Computers and Mathematics with Applications, 64 (10):3258–3266, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002167.

Jasinski:2013:SAF

[JZ13] Maciej Jasinski and Grzegorz Zboinski. On some hp-adaptive finite element method for natural vibrations. Computers and Mathematics with Applications, 66 (11):2376–2399, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005385.

Jomo:2018:PML

[JZE^{+}18] John N. Jomo, Nils Zander, Mohamed Elhaddad, Ali Özcan, Stefan Kollmannsberger, Ralf-Peter Mundani, and Ernst Rank. Parallelization of the multi-level hp-adaptive finite cell method. Computers and Mathematics with Applications, 74(1):126–142, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300147.

Jiang:2018:ATS

[JZJ18] Tongsong Jiang, Zhaozhong Zhang, and Ziwu Jiang. Algebraic techniques for Schrödinger equations in split quaternionic mechanics. Computers and Mathematics with Applications, 75(7):2217–2222, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307691.

Jin:2018:HSF

[JZL18] Jiaojiao Jin, Tong Zhang, and Jian Li. H^{2}- stability of the first order Galerkin method for the Boussinesq equations with smooth and non-smooth initial data. Computers and Mathematics with Applications, 75(1):248–288, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730562X.

Jia:2015:NDM

[JZR15] Yue Jia, Yongjie Zhang, and Timon Rabczuk. A novel dynamic multilevel technique for image registration. Computers and Mathematics with Applications, 69(9):909–925, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000632.

Zong:2019:AUI

[jZsQdLmG19] Jing jing Zong, Tian shuang Qiu, Wei dong Li, and Dong mei Guo. Automatic ultrasound image segmentation based on local entropy and active contour model. Computers and Mathematics with Applications, 78(3):929–943, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122119301476.

Kanth:2010:HVI

[KA10a] A. S. V. Ravi Kanth and K. Aruna. He’s variational iteration method for treating nonlinear singular boundary value problems. Computers and Mathematics with Applications, 60(3):821–829, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000386X.

Khan:2010:CGI

[KA10b] Abdul Rahim Khan and M. A. Ahmed. Convergence of a general iterative scheme for a finite family of asymptotically quasi-nonexpansive mappings in convex metric spaces and applications. Computers and Mathematics with Applications, 59(8):2990–2995, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001264.

Khan:2011:SCS

[KA11] Safeer Hussain Khan and Mujahid Abbas. Strong and Δ-convergence of some iterative schemes in CAT(0) spaces. Computers and Mathematics with Applications, 61(1):109– 116, January 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110008278.

Klimek:2013:FSL

[KA13] M. Klimek and O. P. Agrawal. Fractional Sturm–Liouville problem. Computers and Mathematics with Applications, 66(5):795–812, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112007225.

Kirane:2019:SRD

[KAA19] Mokhtar Kirane, Ahmed Alsaedi, and Bashir Ahmad. On systems of reaction–diffusion equations with a balance law: the sequel. Computers and Mathematics with Applications, 78 (5):1244–1260, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118305856.

Kadalbajoo:2011:CMU

[KAG11] Mohan K. Kadalbajoo, Puneet Arora, and Vikas Gupta. Collocation method using artificial viscosity for solving stiff singularly perturbed turning point problem having twin boundary layers. Computers and Mathematics with Applications, 61(6):1595–1607, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000459.

Khan:2011:EAS

[KAJ11] Najeeb Alam Khan, Asmat Ara, and Muhammad Jamil. An efficient approach for solving the Riccati equation with fractional orders. Computers and Mathematics with Applications, 61(9):2683–2689, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001696.

Ko:2012:SDS

[KAK^{+}12] Hoon Ko, Kyungjin An, Taihoon Kim, Goreti Marreiros, Zita Vale, and Jongmyoung Choi. A study on dynamic state information (DSI) around users for safe urban life. Computers and Mathematics with Applications, 63(2):554–563, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008510.

Khan:2011:BAS

[KALAS11] Q. J. A. Khan, M. Al-Lawatia, and M. H. M. Al-Senaidi. Breeding adjustment of small mammals to avoid predation. Computers and Mathematics with Applications, 62(12):4337–4355, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008327.

Kang:2015:SIP

[Kan15]  K. S. Kang.  Scal-

able implementation of the parallel multigrid method on massively parallel computers. Computers and Mathematics with Applications, 70 (11):2701–2708, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003582.

Karapinar:2010:CFP

[Kar10a] Erdal Karapinar. Couple fixed point theorems for nonlinear contractions in cone metric spaces. Computers and Mathematics with Applications, 59(12):3656– 3668, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000252X.

Karatas:2010:GBH

[Kar10b] Ramazan Karatas. Global behavior of a higher order difference equation. Computers and Mathematics with Applications, 60(3):830–839, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003871.

Karafiat:2017:CEP

[Kar17] Andrzej Karafiat. Convergence estimates for a plane elasticity problem solved by the Galerkin boundary integral formulation with NURBS. Computers and Mathematics with Applications, 74(2):336–349, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302523.

Karkulik:2018:VFT

[Kar18] Michael Karkulik. Variational formulation of time-fractional parabolic equations. Computers and Mathematics with Applications, 75 (11):3929–3938, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301329.

Khan:2011:SFG

[KAS11a] Subuhi Khan and Mustafa Walid Al-Saad. Summation formulae for Gould–Hopper generalized Hermite polynomials. Computers and Mathematics with Applications, 61(6):1536–1541, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211100037X.

Kosti:2011:COE

[KAS11b] A. A. Kosti, Z. A. Anastassi, and T. E. Simos. Construction of an optimized explicit Runge– Kutta–Nyström method for the numerical solution of oscillatory initial value problems. Computers and Mathematics with Applications, 61(11):3381– 3390, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100349X.

Katsikadelis:2011:BNS

[Kat11] John T. Katsikadelis. The BEM for numerical solution of partial fractional differential equations. Computers and Mathematics with Applications, 62(3):891–901, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002707.

Kawai:2015:EHB

[Kaw15] Reiichiro Kawai. Explicit hard bounding functions for boundary value problems for elliptic partial differential equations. Computers and Mathematics with Applications, 70(12):2822–2837, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004642.

Kadem:2010:FRT

[KB10a] Abdelouhab Kadem and Dumitru Baleanu. Fractional radiative transfer equation within Chebyshev spectral approach. Computers and Mathematics with Applications, 59(5):1865–1873, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005641.

Kano:2010:APF

[KB10b] Patrick O. Kano and Moysey Brio. Application of Post’s formula to optical pulse propagation in dispersive media. Computers and Mathematics with Applications, 59(2):629–650, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122109007299.

Kirisci:2010:SNS

[KB10c] Murat Kirisçi and Feyzi Basar. Some new sequence spaces derived by the domain of generalized difference matrix. Computers and Mathematics with Applications, 60(5):1299–1309, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004244.

Kruzel:2013:VOI

[KB13] Filip Kruzel and Krzysztof Banaś. Vectorized OpenCL implementation of numerical integration for higher order finite elements. Computers and Mathematics with Applications, 66(10):2030–2044, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300521X.

Khelghati:2015:BPN

[KB15] Ali Khelghati and Khadijeh Baghaei. Blow-up phenomena for a nonlocal semilinear parabolic equation with positive initial energy. Computers and Mathematics with Applications, 70(5):896–902, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002898.

Karakoc:2019:GFE

[KB19] Seydi Battal Gazi Karakoc and Samir Kumar Bhowmik. Galerkin finite element solution for Benjamin– Bona–Mahony–Burgers equation with cubic Bsplines. Computers and Mathematics with Applications, 77(7):1917–1932, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830676X.

Kabir:2011:AEM

[KBA11] M. M. Kabir, A. Borhanifar, and R. Abazari. Application of (\frac{G′}{G})- expansion method to regularized long wave (RLW) equation. Computers and Mathematics with Applications, 61(8):2044–2047, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006486.

Khaliq:2018:LIP

[KBAF18] A. Q. M. Khaliq, T. A.

Biala, S. S. Alzahrani, and K. M. Furati. Linearly implicit predictorcorrector methods for space-fractional reaction– diffusion equations with non-smooth initial data. Computers and Mathematics with Applications, 75(8):2629–2657, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300026.

Khochemane:2016:PSM

[KBCS16] H. E. Khochemane, H. Boutabia, M. Chau, and P. Spiteri. Parallel solution of the modified porous medium equation. Computers and Mathematics with Applications, 71(4):931–948, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000110.

Kano:2012:DSE

[KBDC12] Patrick O. Kano, Moysey Brio, Paul Dostert, and Jon Cain. Dempster– Shafer evidential theory for the automated selection of parameters for Talbot’s method contours and application to matrix exponentiation. Computers and Mathematics with Applications, 63(11):1519– 1535, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002465.

Kang:2012:DMC

[KBGC12] Rongjie Kang, David T. Branson, Emanuele Guglielmino, and Darwin G. Caldwell. Dynamic modeling and control of an octopus inspired multiple continuum arm robot. Computers and Mathematics with Applications, 64(5):1004–1016, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002234.

Kumar:2019:MLC

[KBK19] Alpesh Kumar, Akanksha Bhardwaj, and B. V. Rathish Kumar. A meshless local collocation method for time fractional diffusion wave equation. Computers and Mathematics with Applications, 78 (6):1851–1861, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930152X.

Kanwar:2011:SCF

[KBS11] V. Kanwar, Ramandeep Behl, and Kapil K. Sharma. Simply constructed family of a Ostrowski’s method with optimal order of convergence. Computers and Mathematics with Applications, 62 (11):4021–4027, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008078.

Kim:2010:SCC

[KC10] In Hwa Kim and Nak Eun Cho. Sufficient conditions for Carathéodory functions. Computers and Mathematics with Applications, 59(6):2067–2073, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122109007597.

Kenary:2011:SMA

[KC11] Hassan Azadi Kenary and Yeol Je Cho. Stability of mixed additivequadratic Jensen type functional equation in various spaces. Computers and Mathematics with Applications, 61(9):2704– 2724, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001763.

Ke:2012:MNA

[KC12] Chih-Kun Ke and Young-Long Chen. A message negotiation approach to e-services by utility function and multi-criteria decision analysis. Computers and Mathematics with Applications, 64 (5):1056–1064, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002295.

Khalid:2019:QAC

[KC19] Sumera Khalid and Qasim Ali Chaudhry. Quantitative analysis of cancer risk assessment in a mammalian cell with the inclusion of mitochondria. Computers and Mathematics with Applications, 78(8):2449–2467, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301786.

Khodja:2013:PSA

[KCC^{+}13] Lilia Ziane Khodja, Ming Chau, Raphaël Couturier, Jacques Bahi, and Pierre Spitéri. Parallel solution of American option derivatives on GPU clusters. Computers and Mathematics with Applications, 65(11):1830– 1848, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001624.

Kumar:2019:FEA

[KCK19] Dileep Kumar, Sudhakar Chaudhary, and V. V. K. Srinivas Kumar. Finite element analysis for coupled time-fractional nonlinear diffusion system. Computers and Mathematics with Applications, 78(6):1919–1936, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301610.

Kumar:2012:NMD

[KCL12] Neeraj Kumar, Naveen Chilamkurti, and Jong-Hyouk Lee. A novel minimum delay maximum flow multicast algorithm to construct a multicast tree in wireless mesh networks. Computers and Mathematics with Applications, 63(2):481–491, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006444.

Kim:2014:BAC

[KCL14a] Hyea Hyun Kim, Eric T. Chung, and Chak Shing Lee. A BDDC algorithm for a class of staggered discontinuous Galerkin methods. Computers and Mathematics with Applications, 67(7):1373–1389, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000480.

Kim:2014:FDP

[KCL14b] Hyea Hyun Kim, Eric T.

Chung, and Chak Shing Lee. FETI–DP preconditioners for a staggered discontinuous Galerkin formulation of the twodimensional Stokes problem. Computers and Mathematics with Applications, 68(12):2233– 2250, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114004647.

Kim:2016:MFC

[KCL16] Hyea Hyun Kim, Eric T.

Chung, and Chi Yeung Lam. Mortar formulation for a class of staggered discontinuous Galerkin methods. Computers and Mathematics with Applications, 71(8):1568–1585, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630089X.

Karakus:2010:ESC

[KD10] Sevda Karakuş and Kamil Demirci. Equi-statistical σ-convergence of positive linear operators. Computers and Mathematics with Applications, 60(8):2212–2218, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005730.

Karakus:2011:SAM

[KD11] Sevda Karakuş and Kamil Demirci. Statistical σ approximation to maxproduct operators. Computers and Mathematics with Applications, 61(4):1024–1031, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009570.

Kazanci:2012:MFL

[KD12] O. Kazanci and B. Davvaz. More on fuzzy lattices. Computers and Mathematics with Applications, 64(9):2917–2925, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003872.

Khani:2011:AIC

[KDG11] F. Khani, M. T. Darvishi, and Rama Subba Reddy Gorla. Analytical investigation for cooling turbine disks with a non-Newtonian viscoelastic fluid. Computers and Mathematics with Applications, 61(7):1728–1738, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100068X.

Kuzmin:2019:PUA

[KdLK19] Dmitri Kuzmin, Manuel Quezada de Luna, and Christopher E. Kees. A partition of unity approach to adaptivity and limiting in continuous finite element methods. Computers and Mathematics with Applications, 78(3):944–957, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301464.

Khan:2015:NLS

[KDU15] Arbaz Khan, Pravir Dutt, and Chandra Shekhar Upadhyay. Nonconforming least-squares spectral element method for European options. Computers and Mathematics with Applications, 70(1):47– 65, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500200X.

Kamensky:2018:PBS

[KEHB18] David Kamensky, John A. Evans, Ming-Chen Hsu, and Yuri Bazilevs. Projectionbased stabilization of interface Lagrange multipliers in immersogeometric fluid-thin structure interaction analysis, with application to heart valve modeling. Computers and Mathematics with Applications, 74 (9):2068–2088, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304236.

Keil:2013:CMH

[Kei13] Frerich J. Keil. Complexities in modeling of heterogeneous catalytic reactions. Computers and Mathematics with Applications, 65(10):1674– 1697, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006840.

Keskin:2010:SSQ

[Kes10]  Refik Keskin.  Solu-

tions of some quadratic Diophantine equations. Computers and Mathematics with Applications, 60(8):2225–2230, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005766.

Knopoff:2013:AMT

[KFTT13] D. A. Knopoff, D. R.

Fernández, G. A. Torres, and C. V. Turner. Adjoint method for a tumor growth PDE-constrained optimization problem. Computers and Mathematics with Applications, 66(6):1104–1119, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004380.

Khan:2011:FVI

[KFYW11] Yasir Khan, Naeem Faraz, Ahmet Yildirim, and Qingbiao Wu. Fractional variational iteration method for fractional initial-boundary value problems arising in the application of nonlinear science. Computers and Mathematics with Applications, 62(5):2273–2278, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005682.

Kim:2011:COV

[KG11] Young Ik Kim and Young Hee Geum. A cubic-order variant of Newton’s method for finding multiple roots of nonlinear equations. Computers and Mathematics with Applications, 62(4):1634–1640, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003725.

Kalita:2014:GTD

[KG14] Jiten C. Kalita and Bidyut B. Gogoi. Global two-dimensional stability of the staggered cavity flow with an HOC approach. Computers and Mathematics with Applications, 67(3):569–590, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122113006718.

Kamali:2011:MOQ

[KGJ11] Amir Kamali, S. M.

T. Fatemi Ghomi, and F. Jolai. A multiobjective quantity discount and joint optimization model for coordination of a singlebuyer multi-vendor supply chain. Computers and Mathematics with Applications, 62(8):3251–3269, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007139.

Kuzmin:2011:RKP

[KGM11] A. Kuzmin, I. Ginzburg, and A. A. Mohamad. The role of the kinetic parameter in the stability of two-relaxation-time advection-diffusion lattice Boltzmann schemes. Computers and Mathematics with Applications, 61(12):3417–3442, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005286.

Kong:2013:WUD

[KH13] Xiangyin Kong and Zhengda Huang. A way of updating the density function for the design of the drum. Computers and Mathematics with Applications, 66(1):62–80, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002186.

Kolasinski:2018:NFV

[KH18] Avary Kolasinski and Weizhang Huang. A new functional for variational mesh generation and adaptation based on equidistribution and alignment conditions. Computers and Mathematics with Applications, 75(6):2044–2058, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304017.

Khan:2010:GVC

[Kha10] Suhel A. Khan. Generalized vector complementaritytype problems in topological vector spaces. Computers and Mathematics with Applications, 59(11):3595–3602, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002440.

Kolahchi:2019:NMM

[KHF^{+}19] Reza Kolahchi, Hadi Hosseini, Mohammad Hosein Fakhar, Reza Taherifar, and Maryam Mahmoudi. A numerical method for magnetohygro-thermal postbuckling analysis of defective quadrilateral graphene sheets using higher order nonlocal strain gradient theory with different movable boundary conditions. Computers and Mathematics with Applications, 78 (6):2018–2034, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301671.

Kulla:2012:PCO

[KHIB12] Elis Kulla, Masahiro Hiyama, Makoto Ikeda, and Leonard Barolli. Performance comparison of OLSR and BAT-MAN routing protocols by a MANET testbed in stairs environment.

Computers and Mathematics with Applications, 63(2):339–349, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100589X.

Kuroki:2012:SPC

[KHUO12] Kazuo Kuroki, Toshio Hayami, Neslihan Uyanik, and Shigeyoshi Owa. Some properties for a certain class concerned with univalent functions. Computers and Mathematics with Applications, 63(10):1425–1432, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002532.

Kotsalis:2010:NPM

[KHWK10] E. M. Kotsalis, I. Hanasaki, J. H. Walther, and P. Koumoutsakos. Nonperiodic molecular dynamics simulations of coarse grained lipid bilayer in water. Computers and Mathematics with Applications, 59(7):2370– 2373, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109006403.

Kiani:2016:FVF

[Kia16] Y. Kiani. Free vibration of functionally graded carbon nanotube reinforced composite plates integrated with piezoelectric layers. Computers and Mathematics with Applications, 72 (9):2433–2449, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305090.

Kiani:2018:NFD

[Kia18] Keivan Kiani. Nonlocal free dynamic analysis of periodic arrays of single-walled carbon nanotubes in the presence of longitudinal thermal and magnetic fields. Computers and Mathematics with Applications, 75(11):3849–3872, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301147.

Kiani:2019:DFI

[Kia19] Keivan Kiani. Divergence and flutter instabilities of nanobeams in moving state accounting for surface and shear effects. Computers and Mathematics with Applications, 77(10):2764–2785, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300136.

Kiguradze:2010:VPP

[Kig10] Tariel Kiguradze. The Vallée–Poussin problem for higher order nonlinear hyperbolic equations. Computers and Mathematics with Applications, 59(2):994–1002, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006737.

Kim:2010:ISB

[Kim10] Dae San Kim. Identities of symmetry for q-Bernoulli polynomials. Computers and Mathematics with Applications, 60(8):2350–2359, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110005924.

Kim:2011:EIP

[Kim11] Chan-Gyun Kim. Existence and iteration of positive solutions for multipoint boundary value problems on a half-line. Computers and Mathematics with Applications, 61(7):1898–1905, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100109X.

Kim:2014:FCP

[Kim14] Kwang-Yeon Kim. Fully computable a posteriori error estimates for the Stokes equation without the global inf-sup constant. Computers and Mathematics with Applications, 67(3):681–691, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006925.

Kim:2015:DGM

[Kim15]  Mi-Young Kim.  A

discontinuous Galerkin method with Lagrange multiplier for hyperbolic conservation laws with boundary conditions. Computers and Mathematics with Applications, 70(4):488–506, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500214X.

Kim:2017:EFH

[Kim17] Jinkyu Kim. Extended framework of Hamilton’s principle for thermoelastic continua. Computers and Mathematics with Applications, 73(7):1505–1523, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300512.

Kim:2018:HOD

[Kim18] Mi-Young Kim. High order DG–DGLM method for hyperbolic conservation laws. Computers and Mathematics with Applications, 75(12):4458–4489, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301792.

Kiryakova:2010:MIM

[Kir10a] V. Kiryakova. The multiindex Mittag-Leffler functions as an important class of special functions of fractional calculus. Computers and Mathematics with Applications, 59(5):1885–1895, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005665.

Kiryakova:2010:SFF

[Kir10b] Virginia Kiryakova. The special functions of fractional calculus as generalized fractional calculus operators of some basic functions. Computers and Mathematics with Applications, 59(3):1128–1141, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003526.

Kim:2011:ABN

[KJ11] Daewook Kim and Il Hyo Jung. Asymptotic behavior of a nonlinear Kirchhoff type equation with spring boundary conditions. Computers and Mathematics with Applications, 62(8):3004–3014, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006742.

Karafyllis:2010:NER

[KJA10] Iasson Karafyllis, Zhong-Ping Jiang, and George Athanasiou. Nash equilibrium and robust stability in dynamic games: a small-gain perspective. Computers and Mathematics with Applications, 60(11):2936–2952, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007510.

Kang:2018:HOR

[KJK18] Myeongmin Kang, Miyoun Jung, and Myungjoo Kang. Higher-order regularization based image restoration with automatic regularization parameter selection. Computers and Mathematics with Applications, 76(1): 58–80, July 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118302025.

Katchang:2010:SCM

[KK10a] Phayap Katchang and Poom Kumam. Strong convergence of the modified Ishikawa iterative method for infinitely many nonexpansive mappings in Banach spaces. Computers and Mathematics with Applications, 59(4):1473–1483, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000477.

See corrigendum [KK11b].

Kiran:2010:FPT

[KK10b] Quanita Kiran and Tayyab Kamran. Fixed point theorems for generalized contractive multi-valued maps. Computers and Mathematics with Applications, 59(12):3813– 3823, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002865.

Konstantinou:2010:RCI

[KK10c] Elisavet Konstantinou and Aristides Kontogeorgis. Ramanujan’s class invariants and their use in elliptic curve cryptography. Computers and Mathematics with Applications, 59(8):2901–2917, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001173.

Kadem:2011:NTE

[KK11a] Abdelouahab Kadem and Adem Kiliçman. Note on transport equation and fractional Sumudu transform. Computers and Mathematics with Applications, 62(8):2995–3003, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006729.

Katchang:2011:CSC

[KK11b] Phayap Katchang and Poom Kumam. Corrigendum to “Strong convergence of the modified Ishikawa iterative method for infinitely many nonexpansive mappings in Banach spaces” [Comput. Math. Appl. 59 (2010) 1473–1483]. Computers and Mathematics with Applications, 61(1):148, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007935.

See [KK10a].

Kim:2012:HSS

[KK12] Tae-Woong Kim and Hee-Cheol Kim. A healthcare system as a service in the context of vital signs: Proposing a framework for realizing a model. Computers and Mathematics with Applications, 64(5):1324–1332, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002817.

Karatson:2013:VPC

[KK13a] J. Karátson and B. Kovács.

Variable preconditioning in complex Hilbert space and its application to the nonlinear Schrödinger equation. Computers and Mathematics with Applications, 65(3):449–459, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003707.

Kawai:2013:SHS

[KK13b] Yutaka Kawai and Noboru Kunihiro. Secret handshake scheme with requestbased-revealing. Computers and Mathematics with Applications, 65(5):786– 798, March 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112004919.

Khalil:2014:NMB

[KK14a] Hammad Khalil and Rahmat Ali Khan. A new method based on Legendre polynomials for solutions of the fractional two-dimensional heat conduction equation. Computers and Mathematics with Applications, 67(10):1938– 1953, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400131X.

Korotov:2014:RRS

[KK14b] Sergey Korotov and Michal Krízek. Red refinements of simplices into congruent subsimplices. Computers and Mathematics with Applications, 67(12):2199–2204, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000662.

Kumar:2014:HOP

[KK14c] Sunil Kumar and Mukesh Kumar. High order parameter-uniform discretization for singularly perturbed parabolic partial differential equations with time delay. Computers and Mathematics with Applications, 68 (10):1355–1367, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004520.

Karatson:2015:SDM

[KK15] J. Karátson and S. Korotov. Some discrete maximum principles arising for nonlinear elliptic finite element problems. Computers and Mathematics with Applications, 70(11):2732–2741, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500334X.

Kumar:2017:SSK

[KK17] Mukesh Kumar and Raj Kumar. Soliton solutions of KD system using similarity transformations method. Computers and Mathematics with Applications, 73(4):701–712, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300020.

Klibanov:2019:CDC

[KK19a] Michael V. Klibanov and Aleksandr E. Kolesov. Convexification of a 3-d coefficient inverse scattering problem. Computers and Mathematics with Applications, 77(6):1681– 1702, March 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118301512.

Kumar:2019:SNP

[KK19b] Dharmendra Kumar and Sachin Kumar. Some new periodic solitary wave solutions of (3 + 1)- dimensional generalized shallow water wave equation by Lie symmetry approach. Computers and Mathematics with Applications, 78(3):857–877, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301269.

Kumar:2019:SWS

[KK19c] Sachin Kumar and Dharmendra Kumar. Solitary wave solutions of (3 + 1)-dimensional extended Zakharov–kuznetsov equation by Lie symmetry approach. Computers and Mathematics with Applications, 77(8):2096–2113, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307016.

Khan:2011:AMS

[KKAM11] Najeeb Alam Khan, Nasir-Uddin Khan, Muhammad Ayaz, and Amir Mahmood. Analytical methods for solving the time-fractional Swift– Hohenberg (S-H) equation. Computers and Mathematics with Applications, 61(8):2182–2185, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006954.

Khalil:2019:SNO

[KKBR19] Hammad Khalil, Rahmat Ali Khan, Dumitru Baleanu, and Mohammad Mehdi Rashidi. Some new operational matrices and its application to fractional order Poisson equations with integral type boundary constrains. Computers and Mathematics with Applications, 78 (6):1826–1837, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301924.

Kim:2010:IMD

[KKC^{+}10] H. J. Kim, C. Kim, Y. Choi, S. Wang, and X. Zhang. Improved modification direction methods. Computers and Mathematics with Applications, 60(2):319–325, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000179.

Kutev:2013:GEC

[KKD13] N. Kutev, N. Kolkovska, and M. Dimova. Global existence of Cauchy problem for Boussinesq paradigm equation. Computers and Mathematics with Applications, 65(3):500–511, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004877.

Kambourakis:2010:RWM

[KKG10] Georgios Kambourakis, Elisavet Konstantinou, and Stefanos Gritzalis. Revisiting WiMAX MBS security. Computers and Mathematics with Applications, 60(2):217–223, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000052.

Kumar:2015:SPE

[KKJ15] Mukesh Kumar, Trond Kvamsdal, and Kjetil André Johannessen. Simple a posteriori error estimators in adaptive isogeometric analysis. Computers and Mathematics with Applications, 70(7):1555–1582, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115002862.

Kolodziej:2012:TPS

[KKK12] Joanna Kolodziej, William J. Knottenbelt, and Samee U. Khan. Theory and practice of stochastic modelling. Computers and Mathematics with Applications, 64(12):3657, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006578.

Kumar:2014:SSZ

[KKK14] Mukesh Kumar, Raj Kumar, and Anshu Kumar. On similarity solutions of Zabolotskaya–Khokhlov equation. Computers and Mathematics with Applications, 68(4):454–463, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002818.

Kumar:2015:SMS

[KKK15] Mukesh Kumar, Raj Kumar, and Anshu Kumar. Some more similarity solutions of the (2 + 1)-dimensional BLP system. Computers and Mathematics with Applications, 70(3):212–221, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001777.

Kumar:2016:SSK

[KKK16] Mukesh Kumar, Anshu Kumar, and Raj Kumar. Similarity solutions of the Konopelchenko– Dubrovsky system using Lie group theory. Computers and Mathematics with Applications, 71(10):2051–2059, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630164X.

Kollmann:2013:RFE

[KKL^{+}13] M. Kollmann, M. Kolmbauer, U. Langer, M. Wolfmayr, and W. Zulehner. A robust finite element solver for a multiharmonic parabolic optimal control problem. Computers and Mathematics with Applications, 65(3):469–486, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112004427.

Kim:2016:COC

[KKL16] Jungeun Kim, Hee-Dae Kwon, and Jeehyun Lee. Constrained optimal control applied to vaccination for influenza. Computers and Mathematics with Applications, 71(11):2313–2329, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211600002X.

Kim:2011:EDR

[KKLJ11] Daewook Kim, Yong Han Kang, Mi Jin Lee, and Il Hyo Jung. Energy decay rate for a quasilinear wave equation with localized strong dissipation. Computers and Mathematics with Applications, 62(1):164–172, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003671.

Kettapun:2010:NAM

[KKS10] Atichart Kettapun, Amnuay Kananthai, and Suthep Suantai. A new approximation method for common fixed points of a finite family of asymptotically quasi-nonexpansive mappings in Banach spaces. Computers and Mathematics with Applications, 60(5):1430–1439, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004396.

Krejci:2018:NAC

[KKSK18] Tomás Krejcí, Jaroslav Kruis, Michal Sejnoha, and Tomás Koudelka. Numerical analysis of coupled heat and moisture transport in masonry. Computers and Mathematics with Applications, 74(1):229–248, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301840.

Kadalbajoo:2013:ARB

[KKT13] Mohan K. Kadalbajoo, Alpesh Kumar, and Lok Pati Tripathi. Application of radial basis function with l-stable Padé time marching scheme for pricing exotic option. Computers and Mathematics with Applications, 66(4):500–511, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003490.

Klima:2019:SIP

[KKVS19] Matej Klíma, Milan Kucharík, Jan Velechovský, and Mikhail Shashkov. Secondinvariant-preserving remap of the 2D deviatoric stress tensor in ALE methods. Computers and Mathematics with Applications, 78(2):654–669, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303365.

Karapinar:2012:QFP

[KL12a] Erdal Karapinar and Nguyen Van Luong. Quadruple fixed point theorems for nonlinear contractions. Computers and Mathematics with Applications, 64(6):1839–1848, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001861.

Khiyabani:2012:PNS

[KL12b] Farzin Modarres Khiyabani and Wah June Leong. On the performance of a new symmetric rank-one method with restart for solving unconstrained optimization problems. Computers and Mathematics with Applications, 64(6):2141–2152, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200449X.

Kang:2016:TSM

[KL16a] Hunseok Kang and Young-Ju Lee. A three-species model for wormlike micellar fluids. Computers and Mathematics with Applications, 71(7):1349–1363, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300566.

Kim:2016:FEM

[KL16b] Seokchan Kim and Hyung-Chun Lee. A finite element method for computing accurate solutions for Poisson equations with corner singularities using the stress intensity factor. Computers and Mathematics with Applications, 71(11):2330– 2337, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500591X.

Kong:2019:BBS

[KL19] Lei Kong and Fengjiao Lu. Bifurcation branch of stationary solutions in a general predatorprey system with preytaxis. Computers and Mathematics with Applications, 78(1):191–203, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300975.

Kang:2016:THL

[KLCD16] Hongmei Kang, Xin Li, Falai Chen, and Jiansong Deng. Truncated hierarchical loop subdivision surfaces and application in isogeometric analysis. Computers and Mathematics with Applications, 72(8):2041–2055, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122116303777.

Kim:2012:IMD

[KLH^{+}12] Wonil Kim, Joo Hwan Lee, Chuleui Hong, Changhee Han, Hanku Lee, and Bongshik Jang. An innovative method for data and software integration in SaaS. Computers and Mathematics with Applications, 64(5):1252–1258, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200274X.

Klimek:2010:AEF

[Kli10] Małgorzata Klimek. On analogues of exponential functions for antisymmetric fractional derivatives. Computers and Mathematics with Applications, 59(5):1709–1717, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005483.

Kuznik:2013:MME

[KLK13] Frédéric Kuznik, Li-Shi Luo, and Manfred Krafczyk. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 65(6):813– 814, March 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211300103X.

Kardani:2015:AGA

[KLK15] Omid Kardani, Andrei Lyamin, and Kristian Krabbenhøft. Application of a GPU-accelerated hybrid preconditioned conjugate gradient approach for large 3D problems in computational geomechanics. Computers and Mathematics with Applications, 69(10):1114– 1131, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500098X.

Kim:2010:RMR

[KLL10] Jongsung Kim, Yuseop Lee, and Sangjin Lee. DES with any reduced masked rounds is not secure against side-channel attacks. Computers and Mathematics with Applications, 60(2):347–354, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000222.

Ke:2019:PSC

[KLL19] Xiao-Feng Ke, Jiu Liu, and Jia-Feng Liao. Positive solutions for a critical p-Laplacian problem with a Kirchhoff term. Computers and Mathematics with Applications, 77(9):2279–2290, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307132.

Kunze:2012:GFT

[KLMV12] H. Kunze, D. La Torre, F. Mendivil, and E. R. Vrscay. Generalized fractal transforms and selfsimilar objects in cone metric spaces. Computers and Mathematics with Applications, 64 (6):1761–1769, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001253.

Kim:2010:SPU

[KLP10] Hyea Hyun Kim, Chang-Ock Lee, and Eun-Hee Park. On the selection of primal unknowns for a FETI–DP formulation of the Stokes problem in two dimensions. Computers and Mathematics with Applications, 60(12):3047–3057, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007789.

Kang:2017:ASV

[KLP17] Jum-Ran Kang, Mi Jin Lee, and Sun Hye Park. Asymptotic stability of a viscoelastic problem with Balakrishnan–Taylor damping and time-varying delay. Computers and Mathematics with Applications, 74(6):1506–1515, September 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117303814.

Kuberry:2012:NAV

[KLRW12] Paul Kuberry, Adam Larios, Leo G. Rebholz, and Nicholas E. Wilson. Numerical approximation of the Voigt regularization for incompressible Navier–Stokes and magnetohydrodynamic flows. Computers and Mathematics with Applications, 64(8):2647–2662, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004944.

Kimiaeifar:2011:AHA

[KLTS11] A. Kimiaeifar, E. Lund, O. T. Thomsen, and J. D. Sørensen. Application of the homotopy analysis method to determine the analytical limit state functions and reliability index for large deflection of a cantilever beam subjected to static co-planar loading. Computers and Mathematics with Applications, 62(12):4646–4655, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100914X.

Kwon:2016:FCI

[KLY16] Hee-Dae Kwon, Jeehyun Lee, and Myoungho Yoon. Feedback control of the immune response of renal transplant recipients with inequality constraints. Computers and Mathematics with Applications, 71(11):2338– 2351, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500588X.

Kuzmin:2010:MMR

[KM10] A. Kuzmin and A. A. Mohamad. Multirange multirelaxation time Shan– Chen model with extended equilibrium. Computers and Mathematics with Applications, 59(7): 2260–2270, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006300.

Kusano:2011:ABP

[KM11] Takasi Kusano and Jelena Manojlović. Asymptotic behavior of positive solutions of sublinear differential equations of Emden– Fowler type. Computers and Mathematics with Applications, 62(2):551– 565, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100407X.

Ketabchi:2012:EMO

[KM12] Saeed Ketabchi and Hossein Moosaei. An efficient method for optimal correcting of absolute value equations by minimal changes in the right hand side. Computers and Mathematics with Applications, 64(6):1882–1885, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002209.

Kothari:2013:SPF

[KM13] Shweta Kothari and Santwana Mukhopadhyay. Study of a problem of functionally graded hollow disk under different thermoelasticity theories — an analysis of phase-lag effects. Computers and Mathematics with Applications, 66(7):1306–1321, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300480X.

Ke:2014:PNS

[KM14] Yi-Fen Ke and Chang-Feng Ma. A preconditioned nested splitting conjugate gradient iterative method for the large sparse generalized Sylvester equation. Computers and Mathematics with Applications, 68 (10):1409–1420, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400457X.

Karkulik:2015:LHO

[KM15] M. Karkulik and J. M. Melenk. Local high-order regularization and applications to hp-methods. Computers and Mathematics with Applications, 70(7):1606–1639, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003120.

Ke:2018:SPE

[KM18] Yi-Fen Ke and Chang-Feng Ma. Some preconditioners for elliptic PDEconstrained optimization problems. Computers and Mathematics with Applications, 75(8):2795–2813, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300130.

Khodabin:2012:NAS

[KMRN12] M. Khodabin, K. Maleknejad, M. Rostami, and M. Nouri. Numerical approach for solving stochastic Volterra– Fredholm integral equations by stochastic operational matrix. Computers and Mathematics with Applications, 64 (6):1903–1913, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002477.

Kalogiratou:2010:NMR

[KMS10] Z. Kalogiratou, Th.

Monovasilis, and T. E. Simos. New modified Runge–Kutta–Nyström methods for the numerical integration of the Schrödinger equation. Computers and Mathematics with Applications, 60(6):1639–1647, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004608.

Kim:2015:HMD

[KMS15] Jibum Kim, Brian J.

Miller, and Suzanne M. Shontz. A hybrid mesh deformation algorithm using anisotropic PDEs and multiobjective mesh optimization. Computers and Mathematics with Applications, 70(8):1830–1851, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003764.

Kergassner:2019:MAM

[KMS19] Andreas Kergaßner, Julia Mergheim, and Paul Steinmann. Modeling of additively manufactured materials using gradientenhanced crystal plasticity. Computers and Mathematics with Applications, 78(7):2338–2350, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302840.

Kusano:2010:ERV

[KMT10] Takaŝi Kusano, Jelena Manojlović, and Tomoyuki Tanigawa. Existence of regularly varying solutions with nonzero indices of half-linear differential equations with retarded arguments. Computers and Mathematics with Applications, 59(1): 411–425, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109004179.

Kacur:2019:NMH

[KMT19] Jozef Kacur, Patrik Mihala, and Michal Tóth. Numerical modeling of heat exchange and unsaturatedsaturated flow in porous media. Computers and Mathematics with Applications, 77(6):1668–1680, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830333X.

Karimov:2011:DPE

[KN11] E. T. Karimov and J. J.

Nieto. The Dirichlet problem for a 3D elliptic equation with two singular coefficients. Computers and Mathematics with Applications, 62(1):214– 224, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003713.

Kimura:2012:VAS

[KN12] Yasunori Kimura and Kazuhide Nakajo. Viscosity approximations by the shrinking projection method in Hilbert spaces. Computers and Mathematics with Applications, 63(9):1400–1408, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000533.

Kareem:2013:HAF

[KNIF13] Waleed Abdel Kareem, Tamer Nabil, Seiichiro Izawa, and Yu Fukunishi. Harmonic analysis filtering techniques for forced and decaying homogeneous isotropic turbulence. Computers and Mathematics with Applications, 65(7):1059–1085, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000722.

Karasozen:2012:SGT

[KNT12] Bülent Karasözen, Andrew D. Nemtsev, and Vyacheslav G. Tsybulin. Staggered grids for three-dimensional convection of a multicomponent fluid in a porous medium. Computers and Mathematics with Applications, 64(6):1740–1751, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001216.

Khazaeli:2019:IME

[KNZ19] Reza Khazaeli, Mohammad Reza Nazari, and Abed Zadehgol. Introducing a modified exact difference method for incorporating unsteady and non-uniform force terms in lattice kinetic models. Computers and Mathematics with Applications, 78(3):723–740, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301087.

Karpuz:2010:FOC

[KÖ10] Basak Karpuz and Özkan Öcalan. Further oscillation criteria for partial difference equations with variable coefficients. Computers and Mathematics with Applications, 59(1):55–63, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006695.

Kadri:2011:SOA

[KO11a] Tlili Kadri and Khaled Omrani. A secondorder accurate difference scheme for an extended Fisher–Kolmogorov equation. Computers and Mathematics with Applications, 61(2):451–459, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008710.

Khiari:2011:FDD

[KO11b] Noomen Khiari and Khaled Omrani. Finite difference discretization of the extended Fisher– Kolmogorov equation in two dimensions. Computers and Mathematics with Applications, 62 (11):4151–4160, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008467.

Koskela:2013:ETM

[KO13] Antti Koskela and Alexander Ostermann. Exponential Taylor methods: Analysis and implementation. Computers and Mathematics with Applications, 65(3):487–499, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004348.

Kollmannsberger:2018:HCM

[KÖC^{+}18] S. Kollmannsberger, A. Özcan, M. Carraturo, N. Zander, and E. Rank. A hierarchical computational model for moving thermal loads and phase changes with applications to selective laser melting. Computers and Mathematics with Applications, 75(5):1483–1497, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307289.

Kojic:2010:ILP

[Koj10] Jelena Kojić. Integer linear programming model for multidimensional twoway number partitioning problem. Computers and Mathematics with Applications, 60(8):2302–2308, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005882.

Kong:2016:PRS

[Kon16] Lingju Kong. Positive radial solutions for quasilinear biharmonic equations. Computers and Mathematics with Applications, 72(12):2878–2886, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305697.

Kromer:2013:TND

[KOPS13] P. Krömer, S. Owais, J. Platos, and V. Snásel. Towards new directions of data mining by evolutionary fuzzy rules and symbolic regression. Computers and Mathematics with Applications, 66(2):190–200, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001284.

Kuznik:2010:LBF

[KORR10] Frédéric Kuznik, Christian Obrecht, Gilles Rusaouen, and Jean-Jacques Roux. LBM based flow simulation using GPU computing processor. Computers and Mathematics with Applications, 59(7):2380–2392, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006361.

Karatsompanis:2010:PAN

[KP10a] I. Karatsompanis and Panos K. Palamides. Polynomial approximation to a non-local boundary value problem. Computers and Mathematics with Applications, 60 (12):3058–3071, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007856.

Kola:2010:ILB

[KP10b] Srinivasa Rao Kola and Pratima Panigrahi. An improved lower bound for the radio k-chromatic number of the hypercube Q_{n}. Computers and Mathematics with Applications, 60(7):2131–2140, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000550X.

Kadiev:2013:ESI

[KP13] Ramazan Kadiev and Arcady Ponosov. Exponential stability of Itô-type linear functional difference equations. Computers and Mathematics with Applications, 66 (11):2295–2306, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003878.

Kajzer:2018:AED

[KP18] Adam Kajzer and Jacek Pozorski. Application of the Entropically Damped Artificial Compressibility model to direct numerical simulation of turbulent channel flow. Computers and Mathematics with Applications, 76 (5):997–1013, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303122.

Kazemzadeh-Parsi:2019:IAS

[KP19a] Mohammad Javad Kazemzadeh-Parsi. Isogeometric analysis in solution of unconfined seepage problems. Computers and Mathematics with Applications, 78(1):66–80, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300793.

Khludnev:2019:SIE

[KP19b] Alexander Khludnev and Tatiana Popova. Semirigid inclusions in elastic bodies: Mechanical interplay and optimal control. Computers and Mathematics with Applications, 77(1):253–262, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830542X.

Kindermann:2019:OST

[KP19c] Stefan Kindermann and Stepán Papácek. Optimization of the shape (and topology) of the initial conditions for diffusion parameter identification. Computers and Mathematics with Applications, 77 (12):3102–3116, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300550.

Kourdis:2013:AAA

[KPG13] Panayotis D. Kourdis, Athanasia G. Palasantza, and Dimitris A. Goussis. Algorithmic asymptotic analysis of the NF–κB signaling system. Computers and Mathematics with Applications, 65(10):1516– 1534, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006505.

Konangi:2018:NSA

[KPG18] Santosh Konangi, Nikhil K. Palakurthi, and Urmila Ghia. von Neumann stability analysis of firstorder accurate discretization schemes for onedimensional (1D) and two-dimensional (2D) fluid flow equations. Computers and Mathematics with Applications, 75(2):643–665, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730620X.

Kim:2018:EDR

[KPK18a] Daewook Kim, Jong Yeoul Park, and Yong Han Kang. Energy decay rate for a von Karman system with a boundary nonlinear delay term. Computers and Mathematics with Applications, 75(9):3269–3282, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300622.

Kim:2018:SQV

[KPK18b] Sangil Kim, Jong-Yeoul Park, and Yong Han Kang. Stochastic quasilinear viscoelastic wave equation with degenerate damping and source terms. Computers and Mathematics with Applications, 75(11):3987– 3994, June 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118301378.

Kuo:2011:VBV

[KPL11] Ying-Che Kuo, Neng-Sheng Pai, and Yen-Feng Li. Vision-based vehicle detection for a driver assistance system. Computers and Mathematics with Applications, 61(8):2096–2100, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006656.

Kim:2013:NRE

[KPP13] Young-Rok Kim, Sang-Young Park, and Chandeok Park. Non-recursive estimation using a batch filter based on particle filtering. Computers and Mathematics with Applications, 66 (10):1905–1919, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004471.

Kadelburg:2010:CFP

[KPR10] Zoran Kadelburg, Mirjana Pavlović, and Stojan Radenović. Common fixed point theorems for ordered contractions and quasicontractions in ordered cone metric spaces. Computers and Mathematics with Applications, 59(9):3148– 3159, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001586.

Kelly:2013:ASI

[KPR13] Cónall Kelly, Peter Palmer, and Alexandra Rodkina. Almost sure instability of the equilibrium solution of a Milsteintype stochastic difference equation. Computers and Mathematics with Applications, 66 (11):2220–2230, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003957.

Kaliyappan:2010:CRF

[KPS10a] M. Kaliyappan, S. Ponnusamy, and S. Sundar. Corrigendum to “Recursive formulation of the matrix Padé approximation in packed storage” [Comput. Math. Appl. 59 (2010) 1532– 1540]. Computers and Mathematics with Applications, 59(8):3027, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001318.

See [KPS10b].

Kaliyappan:2010:RFM

[KPS10b] M. Kaliyappan, S. Ponnusamy, and S. Sundar. Recursive formulation of the matrix Padé approximation in packed storage. Computers and Mathematics with Applications, 59(4):1532–1540, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007457.

See corrigendum [KPS10a].

Kim:2017:HOD

[KPS17] Mi-Young Kim, Eun-Jae Park, and Jaemin Shin. High-order discontinuous Galerkin methods with Lagrange multiplier for hyperbolic systems of conservation laws. Computers and Mathematics with Applications, 73(9):1945–1974, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301189.

Kropinski:2011:FIE

[KQ11] Mary Catherine A. Kropinski and Bryan D. Quaife. Fast integral equation methods for Rothe’s method applied to the isotropic heat equation. Computers and Mathematics with Applications, 61(9):2436–2446, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001106.

Kadelburg:2011:SRS

[KR11] Zoran Kadelburg and Stojan Radenović. Some results on set-valued contractions in abstract metric spaces. Computers and Mathematics with Applications, 62(1):342– 350, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004032.

Kollmannsberger:2019:SAM

[KRAS19] Stefan Kollmannsberger, Ernst Rank, Ferdinando Auricchio, and Paul Steinmann. Simulation for additive manufacturing. Computers and Mathematics with Applications, 78(7):2167, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119304171.

Kumar:2018:MDF

[KRBS18] Sarvesh Kumar, Ricardo Ruiz-Baier, and Ruchi Sandilya. Mixed and discontinuous finite volume element schemes for the optimal control of immiscible flow in porous media. Computers and Mathematics with Applications, 76(4):923–937, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303079.

Keimanesh:2011:STG

[KRCJ11] M. Keimanesh, M. M. Rashidi, Ali J. Chamkha, and R. Jafari. Study of a third grade non-Newtonian fluid flow between two parallel plates using the multistep differential transform method. Computers and Mathematics with Applications, 62(8):2871–2891, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006298.

Konig:2016:FCF

[KRD16] Marcel König, Lars Radtke, and Alexander Düster. A flexible C++ framework for the partitioned solution of strongly coupled multifield problems. Computers and Mathematics with Applications, 72(7):1764–1789, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304369.

Kilbas:2010:KFE

[KRGS^{+}10] Anatoly A. Kilbas, Luis Rodríguez-Germá, Megumi Saigo, R. K. Saxena, and J. J. Trujillo. The Krätzel function and evaluation of integrals. Computers and Mathematics with Applications, 59(5):1790–1800, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005574.

Khan:2010:RCB

[KRM^{+}10] Zeeshan Shafi Khan, Khalid Rashid, Fahad Bin Muhaya, Qutbuddin, and Aneel Rahim. Realization of call-back authentication (CBA) for secure web to cellular phone SMS communication. Computers and Mathematics with Applications, 60(2):198– 208, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000118.

Krnic:2012:RDH

[Krn12]  Mario Krnić.  A re-

fined discrete Hilbert inequality obtained via the Hermite–Hadamard inequality. Computers and Mathematics with Applications, 63(12):1587– 1596, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002192.

Kazem:2012:MMN

[KRP12] S. Kazem, J. A. Rad, and K. Parand. A meshless method on non-Fickian flows with mixing length growth in porous media based on radial basis functions: a comparative study. Computers and Mathematics with Applications, 64(4):399–412, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100928X.

Kangtunyakarn:2010:SCN

[KS10a] Atid Kangtunyakarn and Suthep Suantai. Strong convergence of a new iterative scheme for a finite family of strict pseudocontractions. Computers and Mathematics with Applications, 60(3):680– 694, August 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110003652.

Kumar:2010:PRA

[KS10b] Somesh Kumar and Manu Pratap Singh. Pattern recall analysis of the Hopfield neural network with a genetic algorithm. Computers and Mathematics with Applications, 60(4):1049–1057, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002518.

Karakaya:2011:ABV

[KS11] Vatan Karakaya and Ekrem Savas. On almost p-bounded variation of lacunary sequences. Computers and Mathematics with Applications, 61(6):1502–1506, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000150.

Khuri:2012:BLP

[KS12a] S. A. Khuri and A. Sayfy. The boundary layer problem: a fourth-order adaptive collocation approach. Computers and Mathematics with Applications, 64(6):2089–2099, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003410.

Kumar:2012:PPA

[KS12b] Manoj Kumar and Neelima Singh. Phase plane analysis and traveling wave solution of third order nonlinear singular problems arising in thin film evolution. Computers and Mathematics with Applications, 64 (9):2886–2895, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003847.

Kitzler:2015:HOS

[KS15a] G. Kitzler and J. Schöberl.

A high order spacemomentum discontinuous Galerkin method for the Boltzmann equation. Computers and Mathematics with Applications, 70(7):1539–1554, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002977.

Kumar:2015:HOM

[KS15b] Santosh Kumar and Paramjeet Singh. Higherorder MUSCL scheme for transport equation originating in a neuronal model. Computers and Mathematics with Applications, 70 (12):2838–2853, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004629.

Kanarska:2019:AMD

[KSD^{+}19] Yuliya Kanarska, Samuel Schofield, Timothy Dunn, Benjamin Liu, and Charles Noble. Advanced multidomain method for multiphase flow interaction with Lagrangian structural meshes. Computers and Mathematics with Applications, 78(2):598– 610, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306060.

Klindworth:2014:NRD

[KSF14] Dirk Klindworth, Kersten Schmidt, and Sonia Fliss. Numerical realization of Dirichletto-Neumann transparent boundary conditions for photonic crystal waveguides. Computers and Mathematics with Applications, 67(4):918–943, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001570.

Kavitha:2011:EPD

[KSG11] L. Kavitha, B. Srividya, and D. Gopi. Exact propagating dromionlike localized wave solutions of generalized (2 + 1)-dimensional Davey– Stewartson equations. Computers and Mathematics with Applications, 62(12):4691–4707, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009369.

Kavitha:2012:DSS

[KSJ12] S. Kavitha, S. Sivasubramanian, and R. Jayasankar. Differential subordination and superordination results for Cho– Kwon–Srivastava operator. Computers and Mathematics with Applications, 64(6):1789–1803, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001307.

Kumar:2011:NAR

[KSKK11] Amit Kumar, Pushpinder Singh, Amarpreet Kaur, and Parmpreet Kaur. A new approach for ranking nonnormal p-norm trapezoidal fuzzy numbers. Computers and Mathematics with Applications, 61(4):881–887, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009417.

Knottenbelt:2012:COS

[KSM12] William J. Knottenbelt, Demetris Spanias, and Agnieszka M. Madurska. A common-opponent stochastic model for predicting the outcome of professional tennis matches. Computers and Mathematics with Applications, 64(12):3820–3827, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002106.

Kumar:2011:NSS

[KSMN11] B. V. Rathish Kumar, Vivek Sangwan, S. V. S. S. N. V. G. K. Murthy, and Mohit Nigam. A numerical study of singularly perturbed generalized Burgers–Huxley equation using three-step Taylor– Galerkin method. Computers and Mathematics with Applications, 62(2): 776–786, July 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0898122111004755.

Khoshkhoo:2011:EPC

[KSMT11] H. Khoshkhoo, S. H. H.

Sadeghi, R. Moini, and H. A. Talebi. An efficient power control scheme for electric arc furnaces using online estimation of flexible cable inductance. Computers and Mathematics with Applications, 62(12):4391–4401, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100873X.

Krasnopolsky:2016:UGB

[KSO16] B. Krasnopolsky, A. Starostin, and A. A. Osiptsov. Unified graph-based multifluid model for gas-liquid pipeline flows. Computers and Mathematics with Applications, 72 (5):1244–1262, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303522.

Kakali:2011:NHB

[KSPP11] Vasiliki L. Kakali, Panagiotis G. Sarigiannidis, Georgios I. Papadimitriou, and Andreas S. Pomportsis. A novel HMM-based learning framework for improving dynamic wireless push system performance. Computers and Mathematics with Applications, 62(1):474–485, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004263.

Klinvex:2013:PIT

[KSS13] A. Klinvex, F. Saied, and A. Sameh. Parallel implementations of the trace minimization scheme TraceMIN for the sparse symmetric eigenvalue problem. Computers and Mathematics with Applications, 65(3):460–468, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004415.

Klavzar:2010:MCR

[KST10] Sandi Klavzar, Khaled Salem, and Andrej Taranenko. Maximum cardinality resonant sets and maximal alternating sets of hexagonal systems. Computers and Mathematics with Applications, 59(1):506–513, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900385X.

Karami:2018:VEN

[KSZ18] F. Karami, K. Sadik, and L. Ziad. A variable exponent nonlocal p(x)-Laplacian equation for image restoration. Computers and Mathematics with Applications, 75(2):534–546, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305977.

Karapinar:2011:GCR

[KT11a] Erdal Karapınar and Kenan Taş. Generalized (C)-conditions and related fixed point theorems. Computers and Mathematics with Applications, 61(11):3370– 3380, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003361.

Kiziltunc:2011:CTN

[KT11b] Hukmi Kiziltunc and Seyit Temir. Convergence theorems by a new iteration process for a finite family of nonself asymptotically nonexpansive mappings with errors in Banach spaces. Computers and Mathematics with Applications, 61(9):2480– 2489, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001167.

Kleiss:2015:GSP

[KT15] Stefan K. Kleiss and Satyendra K. Tomar. Guaranteed and sharp a posteriori error estimates in isogeometric analysis. Computers and Mathematics with Applications, 70(3):167–190, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001807.

Kumar:2018:SIS

[KT18a] Mukesh Kumar and Dig Vijay Tanwar. On some invariant solutions of (2 + 1)-dimensional Korteweg– de Vries equations. Computers and Mathematics with Applications, 76(11– 12):2535–2548, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304796.

Kumar:2018:SSB

[KT18b] Mukesh Kumar and Atul Kumar Tiwari. Soliton solutions of BLMP equation by Lie symmetry approach. Computers and Mathematics with Applications, 75(4):1434–1442, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307320.

Khan:2012:AVM

[KTA12] Y. Khan, P. Tiwari, and R. Ali. Application of variational methods to a rectangular clamped plate problem. Computers and Mathematics with Applications, 63(4):862–869, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010340.

Khoa:2017:CPC

[KTDT17] Vo Anh Khoa, Mai Thanh Nhat Truong, Nguyen Ho Minh Duy, and Nguyen Huy Tuan. The Cauchy problem of coupled elliptic sine-Gordon equations with noise: Analysis of a general kernel-based regularization and reliable tools of computing. Computers and Mathematics with Applications, 73(1):141–162, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306083.

Krause:2013:ABF

[KTH13] Mathias J. Krause, Gudrun Thäter, and Vincent Heuveline. Adjointbased fluid flow control and optimisation with lattice Boltzmann methods. Computers and Mathematics with Applications, 65(6):945–960, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005421.

Kumar:2017:SMS

[KTK17] Mukesh Kumar, Atul Kumar Tiwari, and Raj Kumar. Some more solutions of Kadomtsev– Petviashvili equation. Computers and Mathematics with Applications, 74(10):2599–2607, November 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117304637.

Kumar:2018:CFS

[KTK18] Mukesh Kumar, Dig Vijay Tanwar, and Raj Kumar. On closed form solutions of (2+1)-breaking soliton system by similarity transformations method. Computers and Mathematics with Applications, 75(1):218–234, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305539.

Ku:2015:NSN

[Ku15]  JaEun Ku.  Numeri-

cal solutions for nonlinear elliptic problems based on first-order system. Computers and Mathematics with Applications, 69(7):601–609, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000474.

Ku:2018:SLS

[Ku18] JaEun Ku. Superconvergence of least-squares methods for a coupled system of elliptic equations. Computers and Mathematics with Applications, 75(6):2059–2070, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307009.

Kuntanapreeda:2012:RSF

[Kun12] Suwat Kuntanapreeda. Robust synchronization of fractional-order unified chaotic systems via linear control. Computers and Mathematics with Applications, 63(1):183–190, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009734.

Kuo:2011:DFS

[Kuo11] Chao-Lin Kuo. Design of a fuzzy sliding-mode synchronization controller for two different chaos systems. Computers and Mathematics with Applications, 61(8):2090–2095, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006644.

Kuo:2016:MMA

[Kuo16] Yong-Lin Kuo. Mathematical modeling and analysis of the delta robot with flexible links. Computers and Mathematics with Applications, 71(10):1973–1989, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630147X.

Kuo:2018:NES

[Kuo18]  Chun-Ku Kuo.  The

new exact solitary and multi-soliton solutions for the (2 + 1)-dimensional Zakharov–kuznetsov equation. Computers and Mathematics with Applications, 75(8):2851–2857, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300245.

Kupershtokh:2010:CNI

[Kup10] Alexander L. Kupershtokh. Criterion of numerical instability of liquid state in LBE simulations. Computers and Mathematics with Applications, 59(7):2236–2245, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812210900649X.

Kupershtokh:2011:LBE

[Kup11] Alexander L. Kupershtokh. A lattice Boltzmann equation method for real fluids with the equation of state known in tabular form only in regions of liquid and vapor phases. Computers and Mathematics with Applications, 61(12):3537– 3548, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004463.

Kupershtokh:2014:TDL

[Kup14] A. L. Kupershtokh. Threedimensional LBE simulations of a decay of liquid dielectrics with a solute gas into the system of gas-vapor channels under the action of strong electric fields. Computers and Mathematics with Applications, 67(2):340–349, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005373.

Katz:2010:WAS

[KV10] I. Norman Katz and Steven R. Vogl. A Weiszfeld algorithm for the solution of an asymmetric extension of the generalized Fermat location problem. Computers and Mathematics with Applications, 59(1):399– 410, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109004234.

Kapl:2017:SCS

[KV17a] Mario Kapl and Vito Vitrih. Space of C^{2}-smooth geometrically continuous isogeometric functions on planar multi-patch geometries: Dimension and numerical experiments. Computers and Mathematics with Applications, 73(10):2319–2338, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301736.

Kapl:2017:SSG

[KV17b] Mario Kapl and Vito Vitrih. Space of C^{2}-smooth geometrically continuous isogeometric functions on two-patch geometries. Computers and Mathematics with Applications, 73(1):37–59, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305855.

Kumar:2018:WBI

[KV18] K. Harish Kumar and V. Antony Vijesh. Wavelet based iterative methods for a class of 2D-partial integro differential equations. Computers and Mathematics with Applications, 75(1):187–198, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305564.

Kapl:2015:IAG

[KVJB15] Mario Kapl, Vito Vitrih, Bert Jüttler, and Katharina Birner. Isogeometric analysis with geometrically continuous functions on two-patch geometries. Computers and Mathematics with Applications, 70(7):1518–1538, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500173X.

Kruger:2011:EAS

[KVR11] T. Krüger, F. Varnik, and D. Raabe. Efficient and accurate simulations of deformable particles immersed in a fluid using a combined immersed boundary lattice Boltzmann finite element method. Computers and Mathematics with Applications, 61(12):3485– 3505, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002476.

Kolesov:2014:SSP

[KVV14] A. E. Kolesov, P. N.

Vabishchevich, and M. V. Vasilyeva. Splitting schemes for poroelasticity and thermoelasticity problems. Computers and Mathematics with Applications, 67(12):2185– 2198, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000674.

Kang:2018:RTD

[KVW18] T. Kang, K. Van Bockstal, and R. Wang. The reconstruction of a timedependent source from a surface measurement for full Maxwell’s equations by means of the potential field method. Computers and Mathematics with Applications, 75(3):764–786, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306521.

Khan:2011:HPT

[KW11] Yasir Khan and Qingbiao Wu. Homotopy perturbation transform method for nonlinear equations using He’s polynomials. Computers and Mathematics with Applications, 61(8):1963–1967, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122110005869.

Kuo:2012:OLC

[KW12] Yong-Lin Kuo and Tsung-Liang Wu. Open-loop and closed-loop attitude dynamics and controls of miniature spacecraft using pseudowheels. Computers and Mathematics with Applications, 64 (5):1282–1290, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002775.

Kim:2014:CDG

[KW14a] Mi-Young Kim and Mary F.

Wheeler. Coupling discontinuous Galerkin discretizations using mortar finite elements for advection-diffusion-reaction problems. Computers and Mathematics with Applications, 67(1):181–198, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006561.

Kim:2014:MDG

[KW14b] Mi-Young Kim and Mary F.

Wheeler. A multiscale discontinuous Galerkin method for convection– diffusion-reaction problems. Computers and Mathematics with Applications, 68(12):2251– 2261, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114003678.

Karakus:2016:GAL

[KWAS16] A. Karakus, T. Warburton, M. H. Aksel, and C. Sert. A GPU accelerated level set reinitialization for an adaptive discontinuous Galerkin method. Computers and Mathematics with Applications, 72(3):755–767, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302930.

Khan:2011:EVV

[KWFY11] Yasir Khan, Qingbiao Wu, Naeem Faraz, and Ahmet Yildirim. The effects of variable viscosity and thermal conductivity on a thin film flow over a shrinking/stretching sheet. Computers and Mathematics with Applications, 61(11):3391– 3399, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003567.

Kang:2013:OGB

[KWPK13] Ah Reum Kang, Jiyoung Woo, Juyong Park, and Huy Kang Kim. Online game bot detection based on party-play log analysis. Computers and Mathematics with Applications, 65(9):1384–1395, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000442.

Kolodziej:2012:ITA

[KX12] Joanna Kolodziej and Fatos Xhafa. Integration of task abortion and security requirements in GA-based metaheuristics for independent batch grid scheduling. Computers and Mathematics with Applications, 63(2):350–364, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100592X.

Kwon:2010:BEF

[KY10] Kiwoon Kwon and Birsen Yazici. Born expansion and Fréchet derivatives in nonlinear diffuse optical tomography. Computers and Mathematics with Applications, 59(11):3377–3397, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006129.

Kumar:2011:MPR

[KY11] Manoj Kumar and Neha Yadav. Multilayer perceptrons and radial basis function neural network methods for the solution of differential equations: a survey. Computers and Mathematics with Applications, 62 (10):3796–3811, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007966.

Kucuk:2015:OPC

[KYA15] Ismail Kucuk, Kenan Yildirim, and Sarp Adali. Optimal piezoelectric control of a plate subject to time-dependent boundary moments and forcing function for vibration damping. Computers and Mathematics with Applications, 69(4):291–303, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005562.

Khalifeh:2010:AAG

[KYAA10] M. H. Khalifeh, H. Yousefi-Azari, and A. R. Ashrafi. Another aspect of graph invariants depending on the path metric and an application in nanoscience. Computers and Mathematics with Applications, 60(8):2460–2468, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006164.

Khan:2010:CIA

[KYO10] Safeer Hussain Khan, Isa Yildirim, and Murat Ozdemir. Convergence of an implicit algorithm for two families of nonexpansive mappings. Computers and Mathematics with Applications, 59(9):3084– 3091, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001380.

Khan:2011:EOS

[KYR11a] Safeer Hussain Khan, Isa Yildirim, and B. E. Rhoades. Erratum to “A one-step iterative process for two multivalued nonexpansive mappings in Banach spaces” [Comput. Math. Appl. 61 (2011) 3172–3178]. Computers and Mathematics with Applications, 62(9):3648, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007516.

See [KYR11b].

Khan:2011:ROS

[KYR11b] Safeer Hussain Khan, Isa Yildirim, and B. E. Rhoades. RETRACTED: a one-step iterative process for two multivalued nonexpansive mappings in Banach spaces. Computers and Mathematics with Applications, 61(10):3172–3178, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/

pii/S0898122111002987.

See erratum [KYR11a].

Khan:2015:CRV

[KYR15] Subuhi Khan, Ghazala Yasmin, and Mumtaz Riyasat. Certain results for the 2-variable Apostol type and related polynomials. Computers and Mathematics with Applications, 69(11):1367– 1382, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001418.

Khan:2018:HQM

[KYW^{+}18] Dawar Khan, Dong-Ming Yan, Yiqun Wang, Kaimo Hu, Juntao Ye, and Xiaopeng Zhang. Highquality 2D mesh generation without obtuse and small angles. Computers and Mathematics with Applications, 75(2):582–595, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306211.

Kuo:2012:FPD

[KYY12] Ying-Che Kuo, Zih-Yi Yang, and Chih-Hung Yen. Fast pedestrian detection system with a two layer cascade of classifiers. Computers and Mathematics with Applications, 64(5):1311–1323, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002805.

Kim:2016:ODS

[KZ16] Seungil Kim and Hui Zhang. Optimized double sweep Schwarz method by complete radiation boundary conditions. Computers and Mathematics with Applications, 72 (6):1573–1589, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304266.

Lekadir:2011:EBP

[LA11] Ouiza Lekadir and Djamil Aïssani. Error bounds on practical approximation for two tandem queue with blocking and non-preemptive priority. Computers and Mathematics with Applications, 61(7):1810–1822, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000952.

Ladics:2016:COS

[Lad16] Tamás Ladics. Convergence of operator splittings for locally Lipschitzcontinuous operators in Banach spaces. Computers and Mathematics with Applications, 71(1): 57–75, January 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115005106.

Lambers:2012:ADV

[Lam12] James V. Lambers. Approximate diagonalization of variable-coefficient differential operators through similarity transformations. Computers and Mathematics with Applications, 64(8):2575–2593, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200466X.

Lampe:2013:ODI

[Lam13] Richard E. Lampe. Orbit distributions of iterated function systems with finitely many forms.

Computers and Mathematics with Applications, 66(11):2211–2219, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003969.

Lee:2016:MME

[LAM^{+}16] Taehun Lee, Yiannis Andreopoulos, Jeffrey F. Morris, Li-Shi Luo, and Manfred Krafczyk. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 72(2): 249–250, July 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116303157.

Lan:2012:ILC

[Lan12]  Yong-Hong Lan.  It-

erative learning control with initial state learning for fractional order nonlinear systems. Computers and Mathematics with Applications, 64 (10):3210–3216, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211200291X.

Lashin:2010:CSM

[Las10] A. Y. Lashin. On certain subclasses of meromorphic functions associated with certain integral operators. Computers and Mathematics with Applications, 59(1):524–531, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003824.

Lazzus:2010:OAC

[Laz10] Juan A. Lazzús. Optimization of activity coefficient models to describe vapor-liquid equilibrium of (alcohol + water) mixtures using a particle swarm algorithm. Computers and Mathematics with Applications, 60(8):2260–2269, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005808.

Liu:2011:NTP

[LB11] F. Liu and K. Burrage. Novel techniques in parameter estimation for fractional dynamical models arising from biological systems. Computers and Mathematics with Applications, 62(3):822–833, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001519.

Li:2012:IDD

[LB12] Xiaodi Li and Martin Bohner. An impulsive delay differential inequality and applications. Computers and Mathematics with Applications, 64 (6):1875–1881, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002180.

Liu:2018:RPR

[LB18] Longshen Liu and Meng Bai. Remarks on pressure regularity criterion for the 3D Boussinesq equations. Computers and Mathematics with Applications, 76(7):1661–1668, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118303821.

Liefooghe:2012:OBO

[LBH^{+}12] Arnaud Liefooghe, Matthieu Basseur, Jérémie Humeau, Laetitia Jourdan, and El-Ghazali Talbi. On optimizing a bi-objective flowshop scheduling problem in an uncertain environment. Computers and Mathematics with Applications, 64 (12):3747–3762, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001769.

Lin:2010:CAV

[LBJ10] Yiqin Lin, Liang Bao, and Xianzheng Jia. Convergence analysis of a variant of the Newton method for solving nonlinear equations. Computers and Mathematics with Applications, 59(6):2121–2127, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007767.

Lycett-Brown:2014:MCL

[LBL14] Daniel Lycett-Brown and Kai H. Luo. Multiphase cascaded lattice Boltzmann method. Computers and Mathematics with Applications, 67(2):350–362, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005403.

Lin:2016:FEF

[LBvB^{+}16] Chao-An Lin, Yuri Bazilevs, Harald van Brummelen, Kenji Takizawa, and Jong-Shinn Wu. Finite elements in flow problems 2015. Computers and Mathematics with Applications, 72(8):1957–1958, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305405.

Lin:2011:CRI

[LBW11] Yiqin Lin, Liang Bao, and Qinghua Wu. On the convergence rate of an iterative method for solving nonsymmetric algebraic Riccati equations. Computers and Mathematics with Applications, 62(11):4178–4184, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008492.

Li:2011:GKN

[LBZL11] Zhi-Hui Li, Lin Bi, Han-Xin Zhang, and Lin Li. Gas-kinetic numerical study of complex flow problems covering various flow regimes. Computers and Mathematics with Applications, 61(12):3653–3667, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008369.

Lee:2010:NWL

[LC10a] Hsueh-Chen Lee and Tsu-Fen Chen. A nonlinear weighted leastsquares finite element method for Stokes equations. Computers and Mathematics with Applications, 59(1):215–224, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005677.

Li:2010:RSS

[LC10b] Jing-Wen Li and Sui Sun Cheng. Remarks on a set of sufficient conditions for global attractivity in a model of hematopoiesis. Computers and Mathematics with Applications, 59(8):2751–2755, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000787.

Lin:2010:SMS

[LC10c] Huang Lin and Zhenfu Cao. On the security of metering scheme. Computers and Mathematics with Applications, 60(2): 272–275, July 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000192.

Liu:2010:SCS

[LC10d] Jingkun Liu and Jianqing Chen. Sign changing solutions and multiple solutions of an elliptic eigenvalue problem with constraint in H^{1}(R^{N}). Computers and Mathematics with Applications, 59(8):3005–3013, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110001288.

Lei:2011:NMS

[LC11a] Lu Lei and Temuer Chaolu. A new method for solving boundary value problems for partial differential equations. Computers and Mathematics with Applications, 61(8):2164–2167, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006875.

Lin:2011:SNF

[LC11b] Lai-Jiu Lin and Chih-Sheng Chuang. Some new fixed point theorems of generalized nonlinear contractive multivalued maps in complete metric spaces. Computers and Mathematics with Applications, 62(9):3555–3566, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007565.

Lee:2012:FRR

[LC12a] Chien-Hua Lee and Cheng-Yi Chen. Further results for robust stability of homogeneous large-scale bilinear systems with time delays and uncertainties. Computers and Mathematics with Applications, 64(5):1532–1544, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003446.

Luo:2012:SSF

[LC12b] Yung-Chang Luo and Wei-Xian Chen. Sensorless stator field orientation controlled induction motor drive with a fuzzy speed controller. Computers and Mathematics with Applications, 64(5):1206–1216, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002696.

Liu:2013:CFA

[LC13] Hsuan-Ku Liu and Jui-Jane Chang. A closedform approximation for the fractional Black– Scholes model with transaction costs. Computers and Mathematics with Applications, 65(11): 1719–1726, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113001934.

Liu:2016:MSN

[LC16] Hongliang Liu and Haibo Chen. Multiple solutions for a nonlinear Schrödinger–Poisson system with sign-changing potential. Computers and Mathematics with Applications, 71(7):1405–1416, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300529.

Liang:2017:USC

[LCA^{+}17] Yingjie Liang, Wen Chen, Belinda S. Akpa, Thomas Neuberger, Andrew G. Webb, and Richard L. Magin. Using spectral and cumulative spectral entropy to classify anomalous diffusion in SephadexTM gels. Computers and Mathematics with Applications, 73(5):765–774, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300056.

Liao:2012:IHP

[LCC12] Ching-Jong Liao, Chien-Wen Chao, and Liang-Chuan Chen. An improved heuristic for parallel machine weighted flowtime scheduling with family set-up times. Computers and Mathematics with Applications, 63(1):110– 117, January 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009643.

Lu:2013:RAS

[LCC13] Jun-Guo Lu, Yangquan Chen, and Weidong Chen. Robust asymptotical stability of fractionalorder linear systems with structured perturbations. Computers and Mathematics with Applications, 66(5):873–882, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001521.

Liu:2010:AEB

[LCCC10] Jing-Wei Liu, Tai-Liang Chen, Ching-Hsue Cheng, and Yao-Hsien Chen. Adaptive-expectation based multi-attribute FTS model for forecasting TAIEX.

Computers and Mathematics with Applications, 59(2):795–802, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007135.

Leng:2019:EPE

[LCH19] Haitao Leng, Yanping Chen, and Yunqing Huang. Equivalent a posteriori error estimates for elliptic optimal control problems with L^{1}-control cost. Computers and Mathematics with Applications, 77(2):342–356, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305509.

Liu:2019:ETG

[LCHZ19] Shang Liu, Yanping Chen, Yunqing Huang, and Jie Zhou. An efficient two grid method for miscible displacement problem approximated by mixed finite element methods. Computers and Mathematics with Applications, 77(3):752–764, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306084.

Li:2013:MSH

[LCK13] Ming Li, C. S. Chen, and A. Karageorghis. The MFS for the solution of harmonic boundary value problems with non-harmonic boundary conditions. Computers and Mathematics with Applications, 66 (11):2400–2424, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005452.

Li:2017:CEA

[LCK17] Yibao Li, Yongho Choi, and Junseok Kim. Computationally efficient adaptive time step method for the Cahn–Hilliard equation. Computers and Mathematics with Applications, 73(8):1855–1864, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301001.

Lin:2016:FSM

[LCLL16] Ji Lin, C. S. Chen, Chein-Shan Liu, and Jun Lu. Fast simulation of multi-dimensional wave problems by the sparse scheme of the method of fundamental solutions. Computers and Mathematics with Applications, 72(3):555–567, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630284X.

Liu:2018:PSP

[LCLL18] Xiao-Yan Liu, C. S.

Chen, Wen Li, and Ming Li. Particular solutions of products of Helmholtz-type equations using the matern function. Computers and Mathematics with Applications, 75(9):3158–3171, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300543.

Lapinska-Chrzczonowicz:2014:EDS

[LCM14] Magdalena Lapinska-Chrzczonowicz and Piotr Matus. Exact difference schemes for a two-dimensional convection–diffusion-reaction equation. Computers and Mathematics with Applications, 67(12):2205– 2217, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000637.

Li:2010:SFOa

[LCN10] S. G. Li, L. Z. Cheng, and B. Neta. Some fourth-order nonlinear solvers with closed formulae for multiple roots. Computers and Mathematics with Applications, 59(1):126–135, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900618X.

Li:2010:SFOb

[LCP10] Yan Li, YangQuan Chen, and Igor Podlubny. Stability of fractional-order nonlinear dynamic systems: Lyapunov direct method and generalized Mittag-Leffler stability. Computers and Mathematics with Applications, 59(5):1810–1821, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005598.

Li:2016:SBM

[LCP16] Weiwei Li, Wen Chen, and Guofei Pang. Singular boundary method for acoustic eigenanalysis. Computers and Mathematics with Applications, 72(3):663–674, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302917.

Li:2019:MMA

[LCQF19] Junpu Li, Wen Chen, Qing-Hua Qin, and Zhuojia Fu. A modified multilevel algorithm for large-scale scientific and engineering computing. Computers and Mathematics with Applications, 77(8):2061–2076, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307041.

Li:2017:SLO

[LCQL17] Zhenzhen Li, Shaochun Chen, Shuanghong Qu, and Minghao Li. Stabilization of low-order mixed finite elements for the plane elasticity equations. Computers and Mathematics with Applications, 73(3):363–373, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306629.

Li:2015:LBM

[LCS15] Qianhuan Li, Zhenhua Chai, and Baochang Shi. Lattice Boltzmann model for a class of convection–diffusion equations with variable coefficients. Computers and Mathematics with Applications, 70(4):548– 561, August 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115002199.

Li:2018:LBM

[LCS18] Qianhuan Li, Zhenhua Chai, and Baochang Shi. Lattice Boltzmann models for two-dimensional coupled Burgers’ equations. Computers and Mathematics with Applications, 75(3):864–875, February 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117306612.

Lin:2012:LCA

[LCT12] Kuang-Hao Lin, Wei-Hao Chiu, and Jan-Dong Tseng. Low-complexity architecture of carrier frequency offset estimation and compensation for body area network systems. Computers and Mathematics with Applications, 64(5):1400–1408, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002908.

Li:2010:MPS

[LCW10] Peiluan Li, Haibo Chen, and Yusen Wu. Multiple positive solutions of n-point boundary value problems for p-Laplacian impulsive dynamic equations on time scales. Computers and Mathematics with Applications, 60 (9):2572–2582, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006516.

Lai:2012:UPS

[LCW12] Uei-Dar Lai, Kuei-Hsiang Chao, and Meng-Hui Wang. Using a particle swarm method to optimize the weighting in extension theory for the detection of islanding in photovoltaic systems. Computers and Mathematics with Applications, 64(5):1441–1449, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002969.

Liu:2017:HGM

[LCW17] Zhengguang Liu, Aijie Cheng, and Hong Wang. An hp-Galerkin method with fast solution for linear peridynamic models in one dimension. Computers and Mathematics with Applications, 73(7):1546–1565, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300743.

Li:2019:LTS

[LCW19] Qiang Li, Temuer Chaolu, and Yun-Hu Wang. Lumptype solutions and lump solutions for the (2 + 1)-dimensional generalized Bogoyavlensky– Konopelchenko equation. Computers and Mathematics with Applications, 77(8):2077–2085, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830703X.

Liu:2018:TFA

[LCWZ18] Huan Liu, Aijie Cheng, Hong Wang, and Jia Zhao. Time-fractional Allen–Cahn and Cahn– Hilliard phase-field models and their numerical investigation. Computers and Mathematics with Applications, 76(8):1876–1892, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304073.

Lien:2012:RDD

[LCYC12] Chang-Hua Lien, Jenq-Der Chen, Ker-Wei Yu, and Long-Yeu Chung. Robust delay-dependent H_{∞} control for uncertain switched time-delay systems via sampleddata state feedback input. Computers and Mathematics with Applications, 64(5):1187–1196, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002672.

Liu:2011:VMS

[LCZ11] Zhisu Liu, Haibo Chen, and Tiejun Zhou. Variational methods to the second-order impulsive differential equation with Dirichlet boundary value problem. Computers and Mathematics with Applications, 61(6):1687–1699, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000708.

Li:2011:MLS

[LD11a] Junhong Li and Feng Ding. Maximum likelihood stochastic gradient estimation for Hammerstein systems with colored noise based on the key term separation technique. Computers and Mathematics with Applications, 62 (11):4170–4177, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008480.

Li:2011:EPC

[LD11b] Zitian Li and Zhengde Dai. Exact periodic crosskink wave solutions and breather type of twosolitary wave solutions for the (3 + 1)-dimensional potential–YTSF equation. Computers and Mathematics with Applications, 61(8):1939–1945, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000547X.

Lallemand:2013:SRE

[LD13a] Pierre Lallemand and François Dubois. Some results on energy-conserving lattice Boltzmann models. Computers and Mathematics with Applications, 65(6):831–844, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006530.

Lei:2013:AEB

[LD13b] Yunwen Lei and Lixin Ding. Approximation and estimation bounds for free knot splines. Computers and Mathematics with Applications, 65(7):1006–1024, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000606.

Li:2018:EFG

[LD18] Xiaolin Li and Haiyun Dong. The elementfree Galerkin method for the nonlinear p-Laplacian equation. Computers and Mathematics with Applications, 75(7):2549–2560, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307824.

Li:2019:SPP

[LD19] Shanbing Li and Yaying Dong. Stationary patterns of a prey-predator system with a protection zone and crossdiffusion of fractional type. Computers and Mathematics with Applications, 77(7):1873–1887, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306795.

Lv:2019:SNC

[LDG19] Guangying Lv, Jinqiao Duan, and Hongjun Gao. Stochastic nonlocal conservation laws on whole space. Computers and Mathematics with Applications, 77(7):1945–1962, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306801.

Luo:2013:SAA

[LDHH13] Xuqiong Luo, Qikui Du, Hongying Huang, and Tianshu He. A Schwarz alternating algorithm for a three-dimensional exterior harmonic problem with prolate spheroid boundary. Computers and Mathematics with Applications, 65(8):1129–1139, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000965.

Lu:2010:NSP

[LDL10] Chun Lu, Xiaohua Ding, and Mingzhu Liu. The numerical simulation of periodic solutions for a predator–prey system. Computers and Mathematics with Applications, 59(2):868–879, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007068.

Li:2011:ETW

[LDL11] Zitian Li, Zhengde Dai, and Jun Liu. Exact three-wave solutions for the (3 + 1)-dimensional Jimbo–Miwa equation. Computers and Mathematics with Applications, 61(8):2062–2066, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006541.

Liu:2015:FDF

[LDL^{+}15a] Yang Liu, Yanwei Du, Hong Li, Siriguleng He, and Wei Gao. Finite difference/finite element method for a nonlinear time-fractional fourthorder reaction–diffusion problem. Computers and Mathematics with Applications, 70(4):573–591, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002382.

Liu:2015:TGM

[LDL^{+}15b] Yang Liu, Yanwei Du, Hong Li, Jichun Li, and Siriguleng He. A twogrid mixed finite element method for a nonlinear fourth-order reaction– diffusion problem with time-fractional derivative. Computers and Mathematics with Applications, 70(10):2474–2492, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004423.

Lee:2010:MMS

[LDS10] Jonghoon Lee, Burkhard Dünweg, and Jörg Schumacher. Multiscale modelling strategy using the lattice Boltzmann method for polymer dynamics in a turbulent flow. Computers and Mathematics with Applications, 59(7):2374–2379, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122109006397.

Li:2011:NAP

[LDW11] Can Li, Weihua Deng, and Yujiang Wu. Numerical analysis and physical simulations for the time fractional radial diffusion equation. Computers and Mathematics with Applications, 62(3):1024–1037, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100321X.

Lotfi:2011:NTS

[LDY11] A. Lotfi, Mehdi Dehghan, and S. A. Yousefi. A numerical technique for solving fractional optimal control problems. Computers and Mathematics with Applications, 62(3):1055–1067, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002070.

Lee:2011:SCT

[Lee11a] Byung-Soo Lee. Strong convergence theorems with a Noor-type iterative scheme in convex metric spaces. Computers and Mathematics with Applications, 61(11):3218– 3225, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100304X.

Lee:2011:CQC

[Lee11b] Moon Sung Lee. Cryptanalysis of a quadratic compact knapsack publickey cryptosystem. Computers and Mathematics with Applications, 62 (9):3614–3621, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007838.

Lee:2011:NSM

[Lee11c]  Wen-Chiung Lee.  A

note on single-machine scheduling with general learning effect and past-sequence-dependent setup time. Computers and Mathematics with Applications, 62(4):2095–2100, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005384.

Lee:2014:FOP

[Lee14] Younhee Lee. Financial options pricing with regime-switching jumpdiffusions. Computers and Mathematics with Applications, 68(3):392–404, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002764.

Lee:2015:NMS

[Lee15] Eunjung Lee. Newton-LL^{∗} method for the second-order semi-linear elliptic partial differential equations. Computers and Mathematics with Applications, 69(10): 1031–1044, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005501.

Lee:2016:HOM

[Lee16] Hyun Geun Lee. Highorder and mass conservative methods for the conservative Allen–Cahn equation. Computers and Mathematics with Applications, 72(3):620–631, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302796.

Lee:2017:SAF

[Lee17]  Hyun Geun Lee.  A

semi-analytical Fourier spectral method for the Swift–Hohenberg equation. Computers and Mathematics with Applications, 74(8):1885–1896, October 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211730411X.

LeMehaute:2010:RCF

[LEN10] A. Le Méhauté, A. El Kaabouchi, and L. Nivanen. Riemann’s conjecture and a fractional derivative. Computers and Mathematics with Applications, 59(5):1610– 1613, March 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109005367.

Lendak:2011:ACT

[LEP11a] Imre I. Lendak, Aleksandar M. Erdeljan, and Dragan S. Popović. Algorithm for cataloging topologies in the common information model (CIM). Computers and Mathematics with Applications, 61(3):715–721, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009260.

Lepik:2011:SPA

[Lep11b]  Ü. Lepik.  Solving

PDEs with the aid of two-dimensional Haar wavelets. Computers and Mathematics with Applications, 61(7):1873–1879, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001027.

Laskri:2010:CEO

[LeT10] Yamina Laskri and Nasser eddine Tatar. The critical exponent for an ordinary fractional differential problem. Computers and Mathematics with Applications, 59(3):1266–1270, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003939.

Leoreanu-Fotea:2011:AHF

[LF11a] Violeta Leoreanu-Fotea.

Approximations in hypergroups and fuzzy hypergroups. Computers and Mathematics with Applications, 61(9):2734–2741, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001829.

Li:2011:GKB

[LF11b] Qibing Li and Song Fu. A gas-kinetic BGK scheme for gas-water flow. Computers and Mathematics with Applications, 61(12):3639–3652, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007959.

Li:2011:EPP

[LF11c] Yongxiang Li and Hongxia Fan. Existence of positive periodic solutions for higher-order ordinary differential equations. Computers and Mathematics with Applications, 62(4):1715–1722, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004810.

Liu:2019:UMG

[LFAL19] Fawang Liu, Libo Feng, Vo Anh, and Jing Li. Unstructured-mesh Galerkin finite element method for the two-dimensional multi-term time-space fractional Bloch–Torrey equations on irregular convex domains. Computers and Mathematics with Applications, 78 (5):1637–1650, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300124.

Li:2016:NIO

[LFC16] Junpu Li, Zhuojia Fu, and Wen Chen. Numerical investigation on the obliquely incident water wave passing through the submerged breakwater by singular boundary method. Computers and Mathematics with Applications, 71(1):381–390, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122115005684.

Lin:2011:ASL

[LFJ11] Fuming Lin, Yu Fu, and Yingying Jiang. Almost sure limit theorems for the maxima of some strongly dependent Gaussian sequences. Computers and Mathematics with Applications, 62(2):635– 640, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004421.

Liu:2012:NGS

[LFJ12] Xiaoyan Liu, Feng Feng, and Young Bae Jun. A note on generalized soft equal relations. Computers and Mathematics with Applications, 64(4):572–578, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011084.

Li:2019:ESI

[LFZ19a] Rong Li, Hong-Tao Fan, and Bing Zheng. An effective stationary iterative method via double splittings of matrices. Computers and Mathematics with Applications, 77(4):981–990, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306291.

Liu:2019:SCG

[LFZ19b] J. K. Liu, Y. M. Feng, and L. M. Zou. A spectral conjugate gradient method for solving large-scale unconstrained optimization. Computers and Mathematics with Applications, 77(3):731–739, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305832.

Li:2010:SCP

[LG10] Jianjun Li and Wenjie Gao. A strongly coupled predator–prey system with modified Holling– Tanner functional response. Computers and Mathematics with Applications, 60(7):1908–1916, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110005006.

Ling:2012:SLP

[LG12] Yi Ling and Bao Gejun.

Some local Poincaré inequalities for the composition of the sharp maximal operator and the Green’s operator. Computers and Mathematics with Applications, 63(3):720–727, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010169.

Liu:2013:LBS

[LG13] Xiuliang Liu and Zhaoli Guo. A lattice Boltzmann study of gas flows in a long micro-channel. Computers and Mathematics with Applications, 65(2):186–193, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000617.

Lee:2017:CAR

[LG17] Hyung-Chun Lee and Max D. Gunzburger. Comparison of approaches for random PDE optimization problems based on different matching functionals. Computers and Mathematics with Applications, 73(8):1657– 1672, April 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117300688.

Li:2017:HGS

[LGC^{+}17] W. Li, Z. X. Gong, Y. B. Chai, C. Cheng, T. Y. Li, Q. F. Zhang, and M. S. Wang. Hybrid gradient smoothing technique with discrete shear gap method for shell structures. Computers and Mathematics with Applications, 74(8):1826–1855, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304054.

Li:2012:HLS

[LGG12] Ying Li, Yan Gao, and Wenbin Guo. A Hermitian least squares solution of the matrix equation AXB = C subject to inequality restrictions. Computers and Mathematics with Applications, 64(6):1752–1760, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001241.

Liu:2011:NII

[LGH^{+}11] Hongwei Liu, Mohamed S.

Ghidaoui, Zhenhua Huang, Zhida Yuan, and Jun Wang. Numerical investigation of the interactions between solitary waves and pile breakwaters using BGK-based methods. Computers and Mathematics with Applications, 61(12):3668– 3677, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004268.

Laghrib:2016:MFS

[LGHR16] Amine Laghrib, Abdelghani Ghazdali, Abdelilah Hakim, and Said Raghay. A multi-frame super-resolution using diffusion registration and a nonlocal variational image restoration. Computers and Mathematics with Applications, 72 (9):2535–2548, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122116305259.

Liu:2014:DAS

[LGL^{+}14] Geng Liu, Martin Geier, Zhenyu Liu, Manfred Krafczyk, and Tao Chen. Discrete adjoint sensitivity analysis for fluid flow topology optimization based on the generalized lattice Boltzmann method. Computers and Mathematics with Applications, 68 (10):1374–1392, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004507.

Lukyanenko:2019:SCI

[LGVS19] D. V. Lukyanenko, V. B. Grigorev, V. T. Volkov, and M. A. Shishlenin. Solving of the coefficient inverse problem for a nonlinear singularly perturbed two-dimensional reaction–diffusion equation with the location of moving front data. Computers and Mathematics with Applications, 77(5):1245–1254, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118306552.

Li:2019:NSW

[LGZ19] Chao Li, Qilong Guo, and Meimei Zhao. New solitary wave solutions of (2+1)-dimensional space– time fractional Burgers equation and Korteweg– de Vries equation. Computers and Mathematics with Applications, 77(8):2255–2262, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307053.

Lin:2010:ERI

[LH10a] Aihong Lin and Lanying Hu. Existence results for impulsive neutral stochastic functional integro-differential inclusions with nonlocal initial conditions. Computers and Mathematics with Applications, 59(1):64–73, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006683.

Liu:2010:SIP

[LH10b] Bolian Liu and Yufei Huang. The scrambling index of primitive digraphs. Computers and Mathematics with Applications, 60(3):706–721, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003676.

Li:2012:SCM

[LH12a] Deng-Feng Li and Fang-Xuan Hong. Solving constrained matrix games with payoffs of triangular fuzzy numbers. Computers and Mathematics with Applications, 64(4):432–446, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010522.

Liu:2012:FTS

[LH12b] Chi-Hua Liu and Ming-Ying Hsiao. A finite time synergetic control scheme for robot manipulators. Computers and Mathematics with Applications, 64(5):1163–1169, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122112002635.

Li:2016:HAQ

[LH16] Hu Li and Jin Huang.

High-accuracy quadrature methods for solving boundary integral equations of axisymmetric elasticity problems. Computers and Mathematics with Applications, 71(1):459–469, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005799.

Li:2019:PFM

[LH19] Futuan Li and Xianliang Hu. A phase-field method for shape optimization of incompressible flows. Computers and Mathematics with Applications, 77(4):1029–1041, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306266.

Lashkarian:2018:CLD

[LHD18] Elham Lashkarian, S. Reza Hejazi, and Elham Dastranj. Conservation laws of (3 + α)-dimensional time-fractional diffusion equation. Computers and Mathematics with Applications, 75 (3):740–754, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306235.

Liu:2011:IBV

[LHF11] Zhenhai Liu, Jiangfeng Han, and Lijing Fang. Integral boundary value problems for first order integro-differential equations with impulsive integral conditions. Computers and Mathematics with Applications, 61(10):3035–3043, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002744.

Liu:2010:WPI

[LHH10] Bolian Liu, Huoquan Hou, and Yufei Huang. On the Wiener polarity index of trees with maximum degree or given number of leaves. Computers and Mathematics with Applications, 60(7):2053–2057, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005377.

Li:2012:CBS

[LHHZ12] Jiguo Li, Xinyi Huang, Meixue Hong, and Yichen Zhang. Certificate-based signcryption with enhanced security features. Computers and Mathematics with Applications, 64(6):1587–1601, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000168.

Lee:2012:MEM

[LHL12a] Hau-Wei Lee, Shih-Hsiang Hsu, and Chien-Hung Liu. A mathematic error model of a linear moving mechanism based on the transmission medium of a diffraction grating. Computers and Mathematics with Applications, 64(5):813–821, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010996.

Liu:2012:TLM

[LHL12b] Qingfang Liu, Yanren Hou, and Qingchang Liu. A two-level method in time and space for solving the Navier–Stokes equations based on Newton iteration. Computers and Mathematics with Applications, 64 (11):3569–3579, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005822.

Liu:2014:HOT

[LHL14a] Gang Liu, Ting-Zhu Huang, and Jun Liu. High-order TVL1-based images restoration and spatially adapted regularization parameter selection. Computers and Mathematics with Applications, 67(10):2015– 2026, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001655.

Luo:2014:EPU

[LHL^{+}14b] Wei-Hua Luo, Ting-Zhu Huang, Liang Li, Yong Zhang, and Xian-Ming Gu. Efficient preconditioner updates for unsymmetric shifted linear systems. Computers and Mathematics with Applications, 67(9):1643–1655, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400128X.

Liu:2015:CNA

[LHL15] Wei Liu, Jian Huang, and Xiaohan Long. Coupled nonlinear advectiondiffusion-reaction system for prevention of groundwater contamination by modified upwind finite volume element method. Computers and Mathematics with Applications, 69(6):477–493, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000036.

Li:2018:SFV

[LHL18] Yi Li, Yanren Hou, and Rui Li. A stabilized finite volume method for the evolutionary Stokes– Darcy system. Computers and Mathematics with Applications, 75(2):596–613, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306193.

Liu:2015:RBC

[LHLH15] Jun Liu, Ting-Zhu Huang, Xiao-Guang Lv, and Jie Huang. Restoration of blurred color images with impulse noise. Computers and Mathematics with Applications, 70 (6):1255–1265, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003156.

Liu:2011:NSO

[LHM11] Yan Liu, Fenglei Huang, and Aie Ma. Numerical simulations of oblique penetration into reinforced concrete targets. Computers and Mathematics with Applications, 61(8):2168–2171, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006929.

Luc:2019:BPI

[LHTL19] Nguyen Hoang Luc, Le Nhat Huynh, Nguyen Huy Tuan, and Le Dinh Long. On a backward problem for inhomogeneous timefractional diffusion equations. Computers and Mathematics with Applications, 78(5):1317–1333, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300823.

Li:2011:SSI

[LHW11] Chuandong Li, Wenfeng Hu, and Sichao Wu. Stochastic stability of impulsive BAM neural networks with time delays. Computers and Mathematics with Applications, 61(8):2313–2316, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007790.

Leong:2011:MFQ

[LHY11] Wah June Leong, Malik Abu Hassan, and Muhammad Waziri Yusuf. A matrix-free quasi-Newton method for solving large-scale nonlinear systems. Computers and Mathematics with Applications, 62(5):2354–2363, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005773.

Li:2018:PHI

[LHY18] Xu Li, Hai-Feng Huo, and Ai-Li Yang. Preconditioned HSS iteration method and its non-alternating variant for continuous Sylvester equations. Computers and Mathematics with Applications, 75 (4):1095–1106, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306867.

Liu:2011:CRC

[LHZ^{+}11] Yan Liu, Fenglei Huang, Qingming Zhang, Junming Yuan, and Tao Dong. Calculations of radiation characteristics of reflector antennas with surface deformation and perforation. Computers and Mathematics with Applications, 61(8):2349–2352, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122110007868.

Li:2010:RRM

[Li10a] Deng-Feng Li. A ratio ranking method of triangular intuitionistic fuzzy numbers and its application to MADM problems. Computers and Mathematics with Applications, 60(6):1557–1570, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004530.

Li:2010:MFE

[Li10b] Jian Li. A multiscale finite element method for optimal control problems governed by the elliptic homogenization equations. Computers and Mathematics with Applications, 60(3):390–398, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002889.

Li:2010:CUC

[Li10c] Jichun Li. Corrigendum to “Uniformly convergent finite element methods for singularly perturbed elliptic boundary value problems I: Reaction– diffusion type” [Comput. Math. Appl. 35 (3) (1998) 57–70]. Computers and Mathematics with Applications, 59(10):3374– 3376, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002026.

See [LN98].

Li:2010:SDI

[Li10d]  Wei Nian Li.  Some

delay integral inequalities on time scales. Computers and Mathematics with Applications, 59(6):1929–1936, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007366.

Li:2010:PPS

[Li10e]  Yongxiang Li.  Posi-

tive periodic solutions for fully third-order ordinary differential equations. Computers and Mathematics with Applications, 59(11):3464– 3471, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110002245.

Li:2011:SII

[Li11a] Wei Nian Li. Some integral inequalities useful in the theory of certain partial dynamic equations on time scales. Computers and Mathematics with Applications, 61(7):1754–1759, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000885.

Li:2011:TIP

[Li11b] Ya Li. Toxicity impact on a plant-herbivore model with disease in herbivores. Computers and Mathematics with Applications, 62(6):2671–2680, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006754.

Li:2012:IIG

[Li12a] Tuo Li. Invariant intervals and global behavior of a higher order difference equation. Computers and Mathematics with Applications, 64(6):1719–1725, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000569.

Li:2012:MEM

[Li12b]  Yaohang Li.  MOM-

CMC: an efficient Monte Carlo method for multiobjective sampling over real parameter space. Computers and Mathematics with Applications, 64(11):3542–3556, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200569X.

Li:2013:GES

[Li13] Chenglin Li. Global existence of solutions to a cross-diffusion predator– prey system with Holling Type-II functional response. Computers and Mathematics with Applications, 65(8):1152–1162, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000990.

Li:2014:ABM

[Li14]  Haixia Li.  Asymp-

totic behavior and multiplicity for a diffusive Leslie–Gower predator– prey system with crowley-Martin functional response. Computers and Mathematics with Applications, 68(7):693–705, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003368.

Li:2015:NFN

[Li15] Youai Li. A new family of nonconforming finite elements on quadrilaterals. Computers and Mathematics with Applications, 70(4):637–647, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002345.

Li:2016:GEC

[Li16a] Chenglin Li. Global existence of classical solutions to the crossdiffusion three-species model with prey-taxis. Computers and Mathematics with Applications, 72(5):1394–1401, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303790.

Li:2016:NAF

[Li16b] Shuguang Li. Numerical analysis for fourth-order compact conservative difference scheme to solve the 3D Rosenau–RLW equation. Computers and Mathematics with Applications, 72(9):2388–2407, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305120.

Li:2017:SBR

[Li17a]  Chenglin Li.  Stabil-

ity and bifurcation of a ratio-dependent prey– predator system with cross-diffusion. Computers and Mathematics with Applications, 73(4):565–575, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306903.

Li:2017:CLE

[Li17b]  Kexue Li.  A char-

acteristic of local existence for nonlinear fractional heat equations in Lebesgue spaces. Computers and Mathematics with Applications, 73(4):653–665, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300081.

Li:2018:GBC

[Li18a] Chenglin Li. On global bifurcation for a crossdiffusion predator-prey system with prey-taxis. Computers and Mathematics with Applications, 76(5):1014–1025, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303134.

Li:2018:EPG

[Li18b] Hong-Ying Li. Existence of positive ground state solutions for a critical Kirchhoff type problem with sign-changing potential. Computers and Mathematics with Applications, 75(8):2858–2873, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300257.

Li:2018:BSS

[Li18c]  Kexue Li.  Blow-up

of solutions for semilinear stochastic delayed reaction–diffusion equations with Lévy noise. Computers and Mathematics with Applications, 75(2):388–400, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305849.

Li:2018:PWD

[Li18d] Peng Li. Pricing weather derivatives with partial differential equations of the Ornstein– Uhlenbeck process. Computers and Mathematics with Applications, 75(3):1044–1059, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306880.

Li:2018:AMA

[Li18e] Xianping Li. Anisotropic mesh adaptation for finite element solution of Anisotropic Porous Medium Equation. Computers and Mathematics with Applications, 75(6): 2086–2099, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117304832.

Li:2019:LTB

[Li19a]  Bin Li.  Long time

behavior of the solution to a parabolic– elliptic system. Computers and Mathematics with Applications, 78 (10):3345–3362, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302688.

Li:2019:FSB

[Li19b] Weiwei Li. A fast singular boundary method for 3D Helmholtz equation. Computers and Mathematics with Applications, 77(2):525–535, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305753.

Lin:2010:SRQ

[Lin10a]  Yi-Kuei Lin.  System

reliability for quickest path problems under time threshold and budget. Computers and Mathematics with Applications, 60(8):2326–2332, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005900.

Lin:2010:REP

[Lin10b] Yih-Lon Lin. Robust estimation of parameter for fractal inverse problem. Computers and Mathematics with Applications, 60(7):2099–2108, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005432.

Lin:2011:SFN

[Lin11] Yi-Kuei Lin. Stochastic flow networks via multiple paths under time threshold and budget constraint. Computers and Mathematics with Applications, 62(6):2629–2638, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006651.

Lin:2012:SUF

[Lin12] Kuo-Jung Lin. Stabilization of uncertain fuzzy control systems via a new descriptor system approach. Computers and Mathematics with Applications, 64(5):1170–1178, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002647.

Lin:2014:SIT

[Lin14] Ming-Che Lin. Surface instability of thin polymer resist films with phase change effects on coating flow using numerical approximation techniques. Computers and Mathematics with Applications, 68(8):847–858, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003630.

Litvinov:2013:ITM

[Lit13] G. L. Litvinov. Idempotent and tropical mathematics; complexity of algorithms and interval analysis. Computers and Mathematics with Applications, 65(10):1483– 1496, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005834.

Liu:2010:GTS

[Liu10] Zheng Liu. A generalization of two sharp inequalities in two independent variables. Computers and Mathematics with Applications, 59(8):2809–2814, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000866.

Liu:2011:ADT

[Liu11a] Hsuan-Ku Liu. Application of a differential transformation method to strongly nonlinear damped q-difference equations. Computers and Mathematics with Applications, 61(9):2555–2561, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001477.

Liu:2011:TCP

[Liu11b]  Hua-Wen Liu.  Two

classes of pseudo-triangular norms and fuzzy implications. Computers and Mathematics with Applications, 61(4):783–789, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009302.

Liu:2012:SMP

[Liu12]  Yuji Liu.  Solvability

of multi-point boundary value problems for multiple term Riemann– Liouville fractional differential equations. Computers and Mathematics with Applications, 64(4):413– 431, August 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111010479.

Liu:2013:BPN

[Liu13] Yan Liu. Blow-up phenomena for the nonlinear nonlocal porous medium equation under Robin boundary condition. Computers and Mathematics with Applications, 66 (10):2092–2095, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005191.

Liu:2014:MPS

[Liu14] Chein-Shan Liu. A maximal projection solution of ill-posed linear system in a column subspace, better than the least squares solution. Computers and Mathematics with Applications, 67(10):1998– 2014, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001709.

Liu:2015:DOI

[Liu15a]  Chein-Shan Liu.  A

double optimal iterative algorithm in an affine Krylov subspace for solving nonlinear algebraic equations. Computers and Mathematics with Applications, 70 (10):2376–2400, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004162.

Liu:2015:EAH

[Liu15b] Xinwu Liu. Efficient algorithms for hybrid regularizers based image denoising and deblurring. Computers and Mathematics with Applications, 69(7):675–687, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000644.

Liu:2016:FMS

[Liu16a]  Chein-Shan Liu.  A

fast multiple-scale polynomial solution for the inverse Cauchy problem of elasticity in an arbitrary plane domain. Computers and Mathematics with Applications, 72(4):1205–1224, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303765.

Liu:2016:TSP

[Liu16b] Pan-Ping Liu. Transition from stationary patterns to no-stationary patterns in a predator–prey system. Computers and Mathematics with Applications, 71(7):1512–1518, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300803.

Liu:2016:WSU

[Liu16c] Qiao Liu. On weakstrong uniqueness of solutions to the generalized incompressible Navier–Stokes equations. Computers and Mathematics with Applications, 72(3):675–686, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303005.

Liu:2016:WMH

[Liu16d]  Tao Liu.  A wavelet

multiscale-homotopy method for the parameter identification problem of partial differential equations. Computers and Mathematics with Applications, 71(7):1519–1523, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300906.

Liu:2016:ALM

[Liu16e] Xinwu Liu. Augmented Lagrangian method for total generalized variation based Poissonian image restoration. Computers and Mathematics with Applications, 71(8):1694– 1705, April 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116301213.

Liu:2017:SMP

[Liu17] Yikan Liu. Strong maximum principle for multiterm time-fractional diffusion equations and its application to an inverse source problem. Computers and Mathematics with Applications, 73(1):96–108, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306046.

Liu:2018:DPS

[Liu18a] Jian-Guo Liu. Doubleperiodic soliton solutions for the (3 + 1)- dimensional Boiti–Leon– Manna–Pempinelli equation in incompressible fluid. Computers and Mathematics with Applications, 75(10):3604– 3613, May 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301019.

Liu:2018:SMD

[Liu18b]  Keji Liu.  A sim-

ple method for detecting scatterers in a stratified ocean waveguide. Computers and Mathematics with Applications, 76(7):1791–1802, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830395X.

Liu:2018:MER

[Liu18c] Qiao Liu. On 3D MHD equations with regularity in one directional derivative of the velocity. Computers and Mathematics with Applications, 76 (10):2375–2383, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304516.

Liang:2010:ASP

[LJ10a] Songxin Liang and David J. Jeffrey. Approximate solutions to a parameterized sixth order boundary value problem. Computers and Mathematics with Applications, 59(1):247– 253, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109005124.

Liu:2010:MSF

[LJ10b] Xiping Liu and Mei Jia. Multiple solutions for fractional differential equations with nonlinear boundary conditions. Computers and Mathematics with Applications, 59(8):2880–2886, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000970.

Li:2011:EUM

[LJ11] Kexue Li and Junxiong Jia. Existence and uniqueness of mild solutions for abstract delay fractional differential equations. Computers and Mathematics with Applications, 62(3):1398–1404, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001374.

Li:2010:SCT

[LjHO10] Xi Li, Nan jing Huang, and Donal O’Regan. Strong convergence theorems for relatively nonexpansive mappings in Banach spaces with applications. Computers and Mathematics with Applications, 60(5):1322–1331, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000427X.

Leu:2011:PSS

[LJJ11] Fang-Yie Leu, Fenq-Lin Jenq, and Fuu-Cheng Jiang. A path switching scheme for SCTP based on round trip delays. Computers and Mathematics with Applications, 62(9):3504–3523, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007504.

Ling:2017:LAS

[LJJ17] Si-Tao Ling, Zhi-Gang Jia, and Tong-Song Jiang. LSQR algorithm with structured preconditioner for the least squares problem in quaternionic quantum theory. Computers and Mathematics with Applications, 73(10):2208–2220, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301682.

Li:2019:CSD

[LJK^{+}19] Yibao Li, Darae Jeong, Hyundong Kim, Chaeyoung Lee, and Junseok Kim. Comparison study on the different dynamics between the Allen–Cahn and the Cahn–Hilliard equations. Computers and Mathematics with Applications, 77(2):311–322, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305467.

Ling:2019:MLA

[LJLY19] Si-Tao Ling, Zhi-Gang Jia, Xin Lu, and Bing Yang. Matrix LSQR algorithm for structured solutions to quaternionic least squares problem. Computers and Mathematics with Applications, 77(3):830–845, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306205.

Li:2013:CNM

[LJSK13] Yibao Li, Darae Jeong, Jaemin Shin, and Junseok Kim. A conservative numerical method for the Cahn–Hilliard equation with Dirichlet boundary conditions in complex domains. Computers and Mathematics with Applications, 65(1):102–115, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006463.

Liu:2012:SFD

[LJX12] Xiping Liu, Mei Jia, and Xiufen Xiang. On the solvability of a fractional differential equation model involving the p-Laplacian operator. Computers and Mathematics with Applications, 64(10):3267–3275, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002064.

Li:2018:NMS

[LJYS18] Jing Li, Zhen Jin, Yuan Yuan, and Gui-Quan Sun. A non-Markovian SIR network model with fixed infectious period and preventive rewiring. Computers and Mathematics with Applications, 75(11):3884–3902, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301160.

Li:2011:MIS

[LK11] Yibao Li and Junseok Kim. Multiphase image segmentation using a phase-field model. Computers and Mathematics with Applications, 62(2): 737–745, July 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111004652.

Lee:2013:NST

[LK13] Hyun Geun Lee and Junseok Kim. Numerical simulation of the threedimensional Rayleigh– Taylor instability. Computers and Mathematics with Applications, 66 (8):1466–1474, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005166.

Lee:2014:CNP

[LK14] Ju-Hyun Lee and Sungkwon Kang. Complex nonlinear parameter estimation (CNPE) and obstacle shape reconstruction. Computers and Mathematics with Applications, 67(8):1631–1642, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000820.

Latif:2015:CES

[LK15] M. S. Abdel Latif and A. H. Abdel Kader. Comment on: “Exact solutions of the generalized (2 + 1)-dimensional nonlinear evolution equations via the modified simple equation method” [Computers & Mathematics with Applications Volume 69, Issue 5, March 2015, Pages 390–397]. Computers and Mathematics with Applications, 70 (10):2616–2617, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004460.

See [AA15].

Liang:2018:EFS

[LK18] Xiao Liang and Abdul Q. M. Khaliq. An efficient Fourier spectral exponential time differencing method for the space-fractional nonlinear Schrödinger equations. Computers and Mathematics with Applications, 75(12):4438–4457, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301780.

Li:2019:EPT

[LKCN19] Wen Li, Rihuan Ke, Wai-Ki Ching, and Michael K. Ng. A C-eigenvalue problem for tensors with applications to higher-order multivariate Markov chains. Computers and Mathematics with Applications, 78(3):1008–1025, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301415.

Lee:2012:EOI

[LKK12] Yong-Hwan Lee, Bonam Kim, and Heung-Jun Kim. Efficient object identification and localization for image retrieval using query-by-region. Computers and Mathematics with Applications, 63(2):511–517, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006821.

Liolios:2015:CAS

[LKL^{+}15] Angelos Liolios, Athanasios Karabinis, Asterios Liolios, Stefan Radev, Krassimir Georgiev, and Ivan Georgiev. A computational approach for the seismic damage response under multiple earthquakes excitations of adjacent RC structures strengthened by ties. Computers and Mathematics with Applications, 70(11):2742–2751, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122115003806.

Lee:2012:DDA

[LKLP12] Sang Min Lee, Dong Seong Kim, Je Hak Lee, and Jong Sou Park. Detection of DDoS attacks using optimized traffic matrix. Computers and Mathematics with Applications, 63(2):501–510, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006833.

Lee:2010:FNS

[LKS10] Byung Soo Lee, M. Firdosh Khan, and Salahuddin. Fuzzy nonlinear setvalued variational inclusions. Computers and Mathematics with Applications, 60(6):1768–1775, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004827.

Liu:2010:EBN

[LKU10] Zeqing Liu, Shin Min Kang, and Jeong Sheok Ume. Existence of bounded nonoscillatory solutions of first-order nonlinear neutral delay differential equations. Computers and Mathematics with Applications, 59(11):3535–3547, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002361.

Lin:2010:DBE

[LL10a] Bin Lin and Kaitai Li. The (1 + 3)-dimensional Burgers equation and its comparative solutions. Computers and Mathematics with Applications, 60(12):3082–3087, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007881.

Liu:2010:SRL

[LL10b] Muhuo Liu and Bolian Liu. Some results on the Laplacian spectrum. Computers and Mathematics with Applications, 59(11):3612–3616, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002488.

Li:2011:SCS

[LL11] Haitao Li and Yansheng Liu. On sign-changing solutions for a secondorder integral boundary value problem. Computers and Mathematics with Applications, 62(2):651– 656, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004445.

Lee:2012:EDF

[LL12a] Kok-Meng Lee and Chih-Hsing Liu. Explicit dynamic finite element analysis of an automated grasping process using highly damped compliant fingers. Computers and Mathematics with Applications, 64(5):965–977, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001587.

Lee:2012:ASE

[LL12b] Young-Ju Lee and Hengguang Li. Axisymmetric Stokes equations in polygonal domains: Regularity and finite element approximations. Computers and Mathematics with Applications, 64 (11):3500–3521, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005640.

Lin:2012:TCB

[LL12c] Chih-Jer Lin and Po-Ting Lin. Tracking control of a biaxial piezo-actuated positioning stage using generalized Duhem model. Computers and Mathematics with Applications, 64(5):766–787, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010583.

Lin:2012:DBR

[LL12d] Hsiung-Cheng Lin and Liang-Yih Liu. DFTbased recursive groupharmonic energy distribution approach for power interharmonic identification. Computers and Mathematics with Applications, 64(5):1128–1139, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002362.

Liu:2012:NMN

[LL12e] Muhuo Liu and Bolian Liu. New method and new results on the order of spectral radius. Computers and Mathematics with Applications, 63(3):679–686, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010078.

Li:2013:RIE

[LL13] Luyi Li and Zhenzhou Lu. Regional importance effect analysis of the input variables on failure probability and its state dependent parameter estimation. Computers and Mathematics with Applications, 66 (10):2075–2091, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005427.

Lee:2014:SAF

[LL14a] Hyun Geun Lee and June-Yub Lee. A semianalytical Fourier spectral method for the Allen–Cahn equation. Computers and Mathematics with Applications, 68(3):174–184, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002089.

Lu:2014:MSC

[LL14b] Dengfeng Lü and Qiong Liu. Multiplicity of solutions for a class of quasilinear Schrödinger systems in R^{N}. Computers and Mathematics with Applications, 66(12):2532–2544, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005786.

Liu:2015:PMC

[LL15] J. K. Liu and S. J. Li. A projection method for convex constrained monotone nonlinear equations with applications. Computers and Mathematics with Applications, 70 (10):2442–2453, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004447.

Li:2016:EMS

[LL16a] Hong-Ying Li and Jia-Feng Liao. Existence and multiplicity of solutions for a superlinear Kirchhoff-type equations with critical Sobolev exponent in R^{N}. Computers and Mathematics with Applications, 72 (12):2900–2907, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305880.

Li:2016:SML

[LL16b] Xiaolin Li and Shuling Li. On the stability of the moving least squares approximation and the element-free Galerkin method. Computers and Mathematics with Applications, 72 (6):1515–1531, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630400X.

Lin:2016:GWP

[LL16c] Hongxia Lin and Shan Li.

Global well-posedness for the 2\frac{1}{2}D incompressible magneto-micropolar fluid equations with mixed partial viscosity. Computers and Mathematics with Applications, 72(4):1066–1075, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303601.

Liu:2016:NMS

[LL16d] Na Liu and Yansheng Liu. New multi-soliton solutions of a (3 + 1)-dimensional nonlinear evolution equation. Computers and Mathematics with Applications, 71(8):1645–1654, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301286.

Lok:2017:WTS

[LL17] U. Hou Lok and Yuh-Dauh Lyuu. The waterline tree for separable local-volatility models. Computers and Mathematics with Applications, 73(4):537–559, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122116306794.

Lei:2019:MPS

[LL19a] Chun-Yu Lei and Gao-Sheng Liu. Multiple positive solutions for a Schrödinger–Newton system with sign-changing potential. Computers and Mathematics with Applications, 77 (3):631–640, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305820.

Li:2019:EFS

[LL19b] Y. Li and G. R. Liu. An element-free smoothed radial point interpolation method (EFS–RPIM) for 2D and 3D solid mechanics problems. Computers and Mathematics with Applications, 77(2):441–465, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305650.

Li:2013:RDC

[LLC13] Jun Li, Jun-Guo Lu, and Yangquan Chen. Robust decentralized control of perturbed fractionalorder linear interconnected systems. Computers and Mathematics with Applications, 66(5):844–859, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004021.

Liao:2016:IDC

[LLCG16] Xiangke Liao, Shengguo Li, Lizhi Cheng, and Ming Gu. An improved divide-and-conquer algorithm for the banded matrices with narrow bandwidths. Computers and Mathematics with Applications, 71(10):1933– 1943, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301249.

Li:2010:NOL

[LLD10] Guohu Li, Min Liu, and Mingyu Dong. A new online learning algorithm for structure-adjustable extreme learning machine. Computers and Mathematics with Applications, 60(3):377–389, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001999.

Li:2017:NFV

[LLFT17] J. Li, F. Liu, L. Feng, and I. Turner. A novel finite volume method for the Riesz space distributedorder diffusion equation. Computers and Mathematics with Applications, 74(4):772–783, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303061.

Liu:2010:NST

[LLG10] Hongwei Liu, Man Yue Lam, and Mohamed S. Ghidaoui. A numerical study of temporal shallow mixing layers using BGK-based schemes. Computers and Mathematics with Applications, 59(7):2393–2402, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900635X.

Lei:2011:IRA

[LLG^{+}11] J. Lei, S. Liu, H. H. Guo, Z. H. Li, J. T. Li, and Z. X. Han. An image reconstruction algorithm based on the semiparametric model for electrical capacitance tomography. Computers and Mathematics with Applications, 61(9):2843–2853, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002239.

Lin:2010:OOT

[LLH10] Shieh-Shing Lin, Ch’i-Hsin Lin, and Shih-Cheng Horng. An ordinal optimization theory-based algorithm for large distributed power systems. Computers and Mathematics with Applications, 59(10):3361–3373, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002178.

Liu:2011:CSE

[LLH11a] Hongwei Liu, Xiangli Li, and Yakui Huang. Corrigendum to “Solving equations via the trust region and its application to a class of stochastic linear complementarity problems” [Comput. Math. Appl. 61 (2011) 1646– 1664]. Computers and Mathematics with Applications, 61(11):3400, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100352X.

See [LLH11b].

Liu:2011:SET

[LLH11b] Hongwei Liu, Xiangli Li, and Yakui Huang. Solving equations via the trust region and its application to a class of stochastic linear complementarity problems. Computers and Mathematics with Applications, 61(6):1646–1664, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000502.

See [LLH11a].

Lee:2014:OBD

[LLH14] Jangwoon Lee, Jeehyun Lee, and Yoongu Hwang. An optimization based domain decomposition method for PDEs with random inputs. Computers and Mathematics with Applications, 68 (12):2262–2276, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114003393.

Li:2010:USH

[LLJK10] Yibao Li, Hyun Geun Lee, Darae Jeong, and Junseok Kim. An unconditionally stable hybrid numerical method for solving the Allen–Cahn equation. Computers and Mathematics with Applications, 60(6):1591–1606, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004554.

Li:2011:SRC

[LLL11] Hengzhe Li, Xueliang Li, and Sujuan Liu. The (strong) rainbow connection numbers of Cayley graphs on Abelian groups. Computers and Mathematics with Applications, 62(11):4082–4088, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008340.

Liu:2012:EUP

[LLL12] Bingmei Liu, Junling Li, and Lishan Liu. Existence and uniqueness for an m-point boundary value problem at resonance on infinite intervals. Computers and Mathematics with Applications, 64(6):1677–1690, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000338.

Lee:2013:SIM

[LLL13] Keungi Lee, Changhoon Lee, and Sangjin Lee. Onsite investigation methodology for incident response in Windows environments. Computers and Mathematics with Applications, 65(9):1413– 1420, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000399.

Lu:2016:TDS

[LLL16] Shan Lu, Jun Liu, and Gao Lin. A time domain solution for complex multilayered soil model with circular inhomogeneity by the SBFEM.

Computers and Mathematics with Applications, 71(2):652–675, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005969.

Liu:2014:CMC

[LLLC14] Xiao-Yan Liu, Wen Li, Ming Li, and C. S. Chen. Circulant matrix and conformal mapping for solving partial differential equations. Computers and Mathematics with Applications, 68(3):67– 76, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001862.

Liu:2018:TSO

[LLLW18] Nan Liu, Yang Liu, Hong Li, and Jinfeng Wang. Time second-order finite difference/finite element algorithm for nonlinear time-fractional diffusion problem with fourthorder derivative term. Computers and Mathematics with Applications, 75(10):3521–3536, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300877.

Liu:2010:TBC

[LLML10] Chih-Hao Liu, Kuen-Hau Lin, Hao-Chueh Mai, and Chao-An Lin. Thermal boundary conditions for thermal lattice Boltzmann simulations. Computers and Mathematics with Applications, 59(7):2178–2193, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006312.

Li:2015:MSF

[LLML15] Ying Li, Jian Li, Liquan Mei, and Yingping Li. Mixed stabilized finite element methods based on backward difference/ Adams–Bashforth scheme for the time-dependent variable density incompressible flows. Computers and Mathematics with Applications, 70 (10):2575–2588, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004617.

Liu:2019:GSS

[LLP19] Jiu Liu, Jia-Feng Liao, and Hui-Lan Pan. Ground state solution on a nonautonomous Kirchhoff type equation. Computers and Mathematics with Applications, 78(3):878–888, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301282.

Lee:2019:MME

[LLR^{+}19] Cunbiao Lee, Tiegang Liu, Timothy Reis, Baolin Tian, Manfred Krafczyk, and Li-Shi Luo. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 78(4):1051–1052, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302950.

Li:2013:SEM

[LLSS13] Ruilin Li, Chao Li, Jinshu Su, and Bing Sun. Security evaluation of MISTY structure with SPN round function. Computers and Mathematics with Applications, 65(9):1264–1279, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001435.

Lu:2010:PKF

[LLSW10] Zhao Lu, Lily Rui Liang, Gangbing Song, and Shufang Wang. Polychotomous kernel Fisher discriminant via top-down induction of binary tree. Computers and Mathematics with Applications, 60(3):511–519, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003305.

Liu:2016:PGSa

[LLT16a] Jiu Liu, Jia-Feng Liao, and Chun-Lei Tang. A positive ground state solution for a class of asymptotically periodic Schrödinger equations. Computers and Mathematics with Applications, 71(4):965–976, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000079.

Liu:2016:PGSb

[LLT16b] Jiu Liu, Jia-Feng Liao, and Chun-Lei Tang. A positive ground state solution for a class of asymptotically periodic Schrödinger equations with critical exponent. Computers and Mathematics with Applications, 72(7):1851–1864, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304643.

Li:2010:STB

[LLW10] Minqiang Li, Dan Lin, and Shouyang Wang. Solving a type of biobjective bilevel programming problem using NSGA– II. Computers and Mathematics with Applications, 59(2):706–715, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007214.

Lee:2011:SSM

[LLW11] Wen-Chiung Lee, Peng-Jen Lai, and Chin-Chia Wu. Some single-machine and flowshop scheduling problems with a nonlinear deterioration function. Computers and Mathematics with Applications, 62(6):2487–2496, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005918.

Lv:2015:NLF

[LLW15] Zhaoyan Lv, Zhenzhou Lu, and Pan Wang. A new learning function for kriging and its applications to solve reliability problems in engineering. Computers and Mathematics with Applications, 70(5):1182–1197, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003399.

Li:2011:NFD

[LLWZ11] Demin Li, Qiuran Li, Jiacun Wang, and Zhitao Zhang. Navigation function design for backbone connectivity in vehicle ad hoc networks. Computers and Mathematics with Applications, 61(8):2067–2070, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006565.

Liu:2010:ISM

[LLX^{+}10] Yuhua Liu, Yuling Li, Naixue Xiong, Jong Hyuk Park, and Yang Sun Lee. The incentive secure mechanism based on quality of service in P2P network. Computers and Mathematics with Applications, 60(2):224– 233, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000106.

Lv:2011:SCP

[LLX11] Zhi-Wei Lv, Jin Liang, and Ti-Jun Xiao. Solutions to the Cauchy problem for differential equations in Banach spaces with fractional order. Computers and Mathematics with Applications, 62(3):1303–1311, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003282.

Lin:2010:OCE

[LLY10] Xiaojie Lin, Wenbin Liu, and Yuanhong Yu. Oscillation criteria for evenorder half-linear distributed delay differential equations with damping. Computers and Mathematics with Applications, 60(8):2206–2211, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005705.

Liu:2013:BPP

[LLY13] Yan Liu, Shiguang Luo, and Yunhua Ye. Blowup phenomena for a parabolic problem with a gradient nonlinearity under nonlinear boundary conditions. Computers and Mathematics with Applications, 65(8):1194– 1199, April 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113001223.

Li:2018:DDP

[LLY18a] Yanfeng Li, Haicheng Liu, and Ruizhi Yang. A delayed diffusive predatorprey system with predator cannibalism. Computers and Mathematics with Applications, 75(4):1355–1367, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730706X.

Li:2018:ASD

[LLY18b] Zhiyuan Li, Yuri Luchko, and Masahiro Yamamoto. Analyticity of solutions to a distributed order time-fractional diffusion equation and its application to an inverse problem. Computers and Mathematics with Applications, 73(6):1041–1052, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303625.

Long:2019:USC

[LLYL19] Jianmin Long, Chaojun Luo, Qian Yu, and Yibao Li. An unconditional stable compact fourth-order finite difference scheme for three dimensional Allen–Cahn equation. Computers and Mathematics with Applications, 77(4):1042–1054, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122118306254.

Li:2010:EPS

[LLZ10] C. F. Li, X. N. Luo, and Yong Zhou. Existence of positive solutions of the boundary value problem for nonlinear fractional differential equations. Computers and Mathematics with Applications, 59(3):1363–1375, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003915.

Li:2011:IAM

[LLZ11] Luyi Li, Zhenzhou Lu, and Changcong Zhou. Importance analysis for models with correlated input variables by the state dependent parameters method. Computers and Mathematics with Applications, 62 (12):4547–4556, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008984.

Lu:2012:RSC

[LLZ12] Hao Lu, Yunhua Li, and Chenglin Zhu. Robust synthesized control of electromechanical actuator for thrust vector system in spacecraft. Computers and Mathematics with Applications, 64(5):699–708, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100890X.

Liao:2018:EMS

[LLZ18] Jia-Feng Liao, Hong-Ying Li, and Peng Zhang. Existence and multiplicity of solutions for a nonlocal problem with critical Sobolev exponent. Computers and Mathematics with Applications, 75(3):787–797, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306600.

Li:2017:SCA

[LM17] Jing-Tao Li and Chang-Feng Ma. Semi-convergence analysis of parameterized ULT splitting iteration methods for singular saddle point problems. Computers and Mathematics with Applications, 73(10):2285–2292, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730175X.

Li:2018:MLW

[LM18a] Bang-Qing Li and Yu-Lan Ma. Multiple-lump waves for a (3 + 1)- dimensional Boiti–Leon– Manna–Pempinelli equation arising from incompressible fluid. Computers and Mathematics with Applications, 76(1):204– 214, July 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211830213X.

Li:2018:UPI

[LM18b] Cheng-Liang Li and Chang-Feng Ma. The Uzawa–PPS iteration methods for nonsingular and singular non-Hermitian saddle point problems. Computers and Mathematics with Applications, 75(2):703–720, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306259.

Lv:2018:BMS

[LM18c] Chang-Qing Lv and Chang-Feng Ma. BCR method for solving generalized coupled Sylvester equations over centrosymmetric or anti-centrosymmetric matrix. Computers and Mathematics with Applications, 75(1):70–88, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305333.

Li:2019:LPD

[LM19a] Bang-Qing Li and Yu-Lan Ma. Lax pair, Darboux transformation and N-th-order rogue wave solutions for a (2 + 1)-dimensional Heisenberg ferromagnetic spin chain equation. Computers and Mathematics with Applications, 77(2):514–524, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305741.

Li:2019:SCP

[LM19b] Cheng-Liang Li and Chang-Feng Ma. On semi-convergence of parameterized SHSS method for a class of singular complex symmetric linear systems. Computers and Mathematics with Applications, 77(2):466–475, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305704.

Liu:2011:AMS

[LMDL11] Jun Liu, Gui Mu, Zhengde Dai, and Xi Liu. Analytic multi-soliton solutions of the generalized Burgers equation. Computers and Mathematics with Applications, 61(8):1995–1999, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006231.

Li:2011:HKR

[LML11] Jinhai Li, Changlin Mei, and Yuejin Lv. A heuristic knowledge-reduction method for decision formal contexts. Computers and Mathematics with Applications, 61(4):1096–1106, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000965X.

Lieberman:2019:HOL

[LMLB19] Evan J. Lieberman, Nathaniel R. Morgan, Darby J. Luscher, and Donald E. Burton. A higher-order Lagrangian discontinuous Galerkin hydrodynamic method for elasticplastic flows. Computers and Mathematics with Applications, 78(2):318– 334, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304383.

Li:2017:LSS

[LMMF17] Bang-Qing Li, Yu-Lan Ma, Li-Po Mo, and Ying-Ying Fu. The N-loop soliton solutions for (2 + 1)- dimensional Vakhnenko equation. Computers and Mathematics with Applications, 74(3):504–512, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730281X.

Langer:2015:MGI

[LMMT15] Ulrich Langer, Angelos Mantzaflaris, Stephen E.

Moore, and Ioannis Toulopoulos. Mesh grading in isogeometric analysis. Computers and Mathematics with Applications, 70(7):1685–1700, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001066.

Luchko:2013:PSM

[LMP13] Yuri Luchko, Francesco Mainardi, and Yuriy Povstenko. Propagation speed of the maximum of the fundamental solution to the fractional diffusion-wave equation. Computers and Mathematics with Applications, 66(5):774–784, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000229.

LaRocca:2018:DBE

[LMPE18] Michele La Rocca, Andrea Montessori, Pietro Prestininzi, and Lakshmanan Elango. A discrete Boltzmann equation model for two-phase shallow granular flows. Computers and Mathematics with Applications, 75(8):2814–2824, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300142.

Lopes:2013:FDM

[LMPG13] António M. Lopes, J. A. Tenreiro Machado, C. M. A. Pinto, and A. M. S. F. Galhano. Fractional dynamics and MDS visualization of earthquake phenomena. Computers and Mathematics with Applications, 66(5):647–658, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000953.

Lepe:2014:LFF

[LMR14] Felipe Lepe, David Mora, and Rodolfo Rodríguez. Locking-free finite element method for a bending moment formulation of Timoshenko beams. Computers and Mathematics with Applications, 68(3):118–131, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122114002041.

Langer:2019:GEB

[LMR19] Ulrich Langer, Svetlana Matculevich, and Sergey Repin. Guaranteed error bounds and local indicators for adaptive solvers using stabilised space– time IgA approximations to parabolic problems. Computers and Mathematics with Applications, 78(8):2641–2671, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930207X.

Litvinchev:2010:LHC

[LMRS10] Igor Litvinchev, Miguel Mata, Socorro Rangel, and Jania Saucedo. Lagrangian heuristic for a class of the generalized assignment problems. Computers and Mathematics with Applications, 60(4):1115–1123, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002609.

Leonenko:2013:CSF

[LMS13] Nikolai N. Leonenko, Mark M. Meerschaert, and Alla Sikorskii. Correlation structure of fractional Pearson diffusions. Computers and Mathematics with Applications, 66(5):737–745, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000266.

Li:2010:DPS

[LMW10] Xiaoliang Li, Chenqi Mou, and Dongming Wang. Decomposing polynomial sets into simple sets over finite fields: the zero-dimensional case. Computers and Mathematics with Applications, 60(11):2983–2997, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000756X.

Li:2019:GBT

[LMY19] Dan Li, Chunlai Mu, and Hong Yi. Global boundedness in a threedimensional chemotaxishaptotaxis model. Computers and Mathematics with Applications, 77(9):2447–2462, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307284.

Liu:2017:OFC

[LMZ17] Zhenhai Liu, Stanislaw Migórski, and Biao Zeng. Optimal feedback control and controllability for hyperbolic evolution inclusions of Clarke’s subdifferential type. Computers and Mathematics with Applications, 74 (12):3183–3194, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305163.

Lin:2018:NAT

[LMZ18] Ke Lin, Chunlai Mu, and Hua Zhong. A new approach toward stabilization in a two-species chemotaxis model with logistic source. Computers and Mathematics with Applications, 75(3):837–849, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306557.

Lu:2016:RSE

[LMZK16] Xing Lü, Wen-Xiu Ma, Yuan Zhou, and Chaudry Masood Khalique. Rational solutions to an extended Kadomtsev– Petviashvili-like equation with symbolic computation. Computers and Mathematics with Applications, 71(8):1560–1567, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300712.

Li:1998:UCF

[LN98] J. Li and I. M. Navon. Uniformly convergent finite element methods for singularly perturbed elliptic boundary value problems i: Reaction–diffusion type. Computers and Mathematics with Applications, 35(3):57–70, February 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122197002794.

See corrigendum [Li10c].

Li:2010:CPF

[LN10] Jueyou Li and Muhammad Aslam Noor. On characterizations of preinvex fuzzy mappings. Computers and Mathematics with Applications, 59(2):933–940, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006798.

Lin:2019:FSM

[LN19] Xue-Lei Lin and Michael K.

Ng. A fast solver for multidimensional time-space fractional diffusion equation with variable coefficients. Computers and Mathematics with Applications, 78(5):1477–1489, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930210X.

Lin:2012:MTH

[LNKU12] Jiaqing Lin, Hiroaki Nishino, Tsuneo Kagawa, and Kouichi Utsumiya. A method of two-handed gesture interactions with applications based on commodity devices. Computers and Mathematics with Applications, 63(2):448–457, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006274.

Lilkova:2012:MSM

[LNP^{+}12] Elena Lilkova, Genoveva Nacheva, Peicho Petkov, Petko Petkov, Stoyan Markov, Nevena Ilieva, and Leandar Litov. Metadynamics study of mutant human interferon-gamma forms. Computers and Mathematics with Applications, 64(3):272–277, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200082X.

Lopez:2013:P

[LNP^{+}13] Javier Lopez, Svetla Nikova, Andreas Pashalidis, Günther Pernul, and Bart Preneel. Preface. Computers and Mathematics with Applications, 65(5):747, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000795.

Li:2019:SNG

[LNW19] Dan Li, Yufeng Nie, and Chunmei Wang. Superconvergence of numerical gradient for weak Galerkin finite element methods on nonuniform Cartesian partitions in three dimensions. Computers and Mathematics with Applications, 78(3):905–928, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301294.

Luan:2016:PER

[LO16] Vu Thai Luan and Alexander Ostermann. Parallel exponential Rosenbrock methods. Computers and Mathematics with Applications, 71(5):1137– 1150, March 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116300165.

Lohmann:2016:EAC

[Loh16] Christoph Lohmann. Efficient algorithms for constraining orientation tensors in Galerkin methods for the Fokker–Planck equation. Computers and Mathematics with Applications, 71(5):1059–1073, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211600016X.

Lopez:2019:DZR

[Lóp19] José Luis López. Derivation of the zeroth retardation order Boltzmann equation from nonstandard space–time Wigner quantum kinetics. Computers and Mathematics with Applications, 78(8):2512–2524, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301853.

Landi:2010:ILM

[LP10] G. Landi and E. Loli Piccolomini. An iterative Lagrange method for the regularization of discrete ill-posed inverse problems. Computers and Mathematics with Applications, 60(6):1723–1738, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004785.

Lee:2012:DIS

[LP12] Chang-Ock Lee and Eun-Hee Park. A dual iterative substructuring method with a penalty term in three dimensions. Computers and Mathematics with Applications, 64(9):2787–2805, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003513.

Lee:2015:ASP

[LPK15] Mi Jin Lee, Jong Yeoul Park, and Yong Han Kang. Asymptotic stability of a problem with Balakrishnan– Taylor damping and a time delay. Computers and Mathematics with Applications, 70(4):478– 487, August 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115002151.

Lee:2019:TTW

[LPLR19] D. Lee, M. Petersen, R. Lowrie, and T. Ringler. Tracer transport within an unstructured grid ocean model using characteristic discontinuous Galerkin advection. Computers and Mathematics with Applications, 78(2):

611–622, July 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305364.

Li:2019:JRM

[LPML19] Haichao Li, Fuzhen Pang, Xuhong Miao, and Yuhui Li. Jacobi–Ritz method for free vibration analysis of uniform and stepped circular cylindrical shells with arbitrary boundary conditions: a unified formulation. Computers and Mathematics with Applications, 77(2):427–440, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305649.

Lorig:2015:PAE

[LPP15] Matthew Lorig, Stefano Pagliarani, and Andrea Pascucci. Pricing approximations and error estimates for local Lévytype models with default. Computers and Mathematics with Applications, 69(10):1189– 1219, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001182.

Lv:2016:HBG

[LPY16] Yunfei Lv, Yongzhen Pei, and Rong Yuan. Hopf bifurcation and global stability of a diffusive Gausetype predator–prey models. Computers and Mathematics with Applications, 72(10):2620–2635, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630534X.

Li:2018:DGM

[LQMW18] Chen Li, Ruibin Qin, Ju Ming, and Zhongming Wang. A discontinuous Galerkin method for stochastic Cahn–Hilliard equations. Computers and Mathematics with Applications, 75(6):2100– 2114, March 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117303346.

Love:2013:CFD

[LR13] E. Love and W. J. Rider.

On the convergence of finite difference methods for PDE under temporal refinement. Computers and Mathematics with Applications, 66(1):33– 40, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002265.

Lee:2014:LSA

[LR14] Hyesuk Lee and Kelsey Rife. Least squares approach for the timedependent nonlinear Stokes– Darcy flow. Computers and Mathematics with Applications, 67(10): 1806–1815, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114001552.

Li:2015:MFE

[LR15] Xindong Li and Hongxing Rui. A MCC finite element approximation of incompressible miscible displacement in porous media. Computers and Mathematics with Applications, 70(5):750–764, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002412.

Li:2017:EET

[LR17a] Jin Li and Hongxing Rui. Error expansion of trapezoidal rule for certain two-dimensional Cauchy principal value integrals. Computers and Mathematics with Applications, 74 (10):2608–2637, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305886.

Li:2017:CBC

[LR17b] Xiaoli Li and Hongxing Rui. Characteristic block-centered finite difference method for simulating incompressible wormhole propagation. Computers and Mathematics with Applications, 73(10):2171–2190, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301426.

Lopez-Rincon:2015:NIP

[LRBA15] Alejandro Lopez-Rincon, Mostafa Bendahmane, and Bedr’Eddine Ainseba. On 3D numerical inverse problems for the bidomain model in electrocardiology. Computers and Mathematics with Applications, 69(4):255–274, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004714.

Lozano:2016:EAG

[LRCG16] Elias Lozano, Deane Roehl, Waldemar Celes, and Marcelo Gattass. An efficient algorithm to generate random sphere packs in arbitrary domains. Computers and Mathematics with Applications, 71(8):1586–1601, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300864.

Liu:2013:TGA

[LRH13] Wei Liu, Hongxing Rui, and Fengzhu Hu. A twogrid algorithm for expanded mixed finite element approximations of semi-linear elliptic equations. Computers and Mathematics with Applications, 66(3):392–402, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003180.

Luchko:2010:FMN

[LRTV10] Yury F. Luchko, Margarita Rivero, Juan J. Trujillo, and M. Pilar Velasco. Fractional models, non-locality, and complex systems. Computers and Mathematics with Applications, 59(3):1048–1056, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003502.

Liu:2013:SMH

[LRV13] Maoxing Liu, Gergely Röst, and Gabriella Vas. SIS model on homogeneous networks with threshold type delayed contact reduction. Computers and Mathematics with Applications, 66 (9):1534–1546, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001016.

Liang:2018:FSK

[LRZ18] Sihua Liang, Dusan Repovs, and Binlin Zhang.

On the fractional Schrödinger– Kirchhoff equations with electromagnetic fields and critical nonlinearity. Computers and Mathematics with Applications, 75(5):1778–1794, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307599.

Lakestani:2010:NST

[LS10a] Mehrdad Lakestani and Behzad Nemati Saray. Numerical solution of telegraph equation using interpolating scaling functions. Computers and Mathematics with Applications, 60(7):1964–1972, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005225.

Li:2010:MFC

[LS10b] Hong-Yan Li and Fu-Gui Shi. Measures of fuzzy compactness in Lfuzzy topological spaces. Computers and Mathematics with Applications, 59(2):941–947, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006786.

Liu:2010:MAD

[LS10c] Peide Liu and Yu Su.

The multiple-attribute decision making method based on the TFLHOWA operator. Computers and Mathematics with Applications, 60(9):2609–2615, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006723.

Lv:2010:LTV

[LS10d] Shaogao Lv and Lei Shi.

Learning theory viewpoint of approximation by positive linear operators. Computers and Mathematics with Applications, 60(12):3177–3186, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008011.

Li:2011:PSS

[LS11a] Hongyu Li and Jingxian Sun. Positive solutions of superlinear semipositone nonlinear boundary value problems. Computers and Mathematics with Applications, 61(9):2806– 2815, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002148.

Li:2011:NSF

[LS11b] Yuanlu Li and Ning Sun. Numerical solution of fractional differential equations using the generalized block pulse operational matrix. Computers and Mathematics with Applications, 62(3):1046–1054, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001842.

Li:2011:OPS

[LS11c] Zhicheng Li and Huisheng Shu. Optimal portfolio selection with liability management and Markov switching under constrained variance. Computers and Mathematics with Applications, 61(8):2271–2277, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211000742X.

Lu:2011:SSI

[LS11d] Da-Qian Lu and H. M.

Srivastava. Some series identities involving the generalized Apostol type and related polynomials. Computers and Mathematics with Applications, 62(9):3591–3602, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007784.

Liu:2012:NBVa

[LS12a] Zhenhai Liu and Jihua Sun. Nonlinear boundary value problems of fractional differential systems. Computers and Mathematics with Applications, 64(4):463–475, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010753.

Liu:2012:NBVb

[LS12b] Zhenhai Liu and Jihua Sun. Nonlinear boundary value problems of fractional functional integrodifferential equations. Computers and Mathematics with Applications, 64(10):3228–3234, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200140X.

Liu:2012:ICO

[LS12c] Zhi Liu and Yuangong Sun. Interval criteria for oscillation of a forced impulsive differential equation with Riemann–Stieltjes integral. Computers and Mathematics with Applications, 63(12):1577– 1586, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002155.

Li:2016:TDS

[LS16] Xiaochao Li and Yu Su. Three-dimensional stress analysis of thin structures using a boundary element method with sinh transformation for nearly singular integrals. Computers and Mathematics with Applications, 72(11):2773–2787, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305594.

Lei:2017:PSS

[LS17a] Chun-Yu Lei and Hong-Min Suo. Positive solutions for a Schrödinger– Poisson system involving concave-convex nonlinearities. Computers and Mathematics with Applications, 74 (6):1516–1524, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303772.

Liu:2017:BCF

[LS17b] Wei Liu and Zhengjia Sun. A block-centered finite difference method for reduced fracture model in Karst aquifer system. Computers and Mathematics with Applications, 74(6):1455–1470, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303760.

Lee:2019:ESC

[LS19a] Seunggyu Lee and Jaemin Shin. Energy stable compact scheme for Cahn– Hilliard equation with periodic boundary condition. Computers and Mathematics with Applications, 77(1):189–198, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305339.

Li:2019:ERB

[LS19b] Pan-Li Li and Hong-Rui Sun. Existence results and bifurcation for nonlocal fractional problems with critical Sobolev exponent. Computers and Mathematics with Applications, 78(5):1720–1731, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302032.

Liang:2017:EPF

[LSC17] H. Liang, B. C. Shi, and Z. H. Chai. An efficient phase-field-based multiple-relaxation-time lattice Boltzmann model for three-dimensional multiphase flows. Computers and Mathematics with Applications, 73(7):1524–1538, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300500.

Lei:2016:GSS

[LSCG16] Chun-Yu Lei, Hong-Min Suo, Chang-Mu Chu, and Liu-Tao Guo. On ground state solutions for a Kirchhoff type equation with critical growth. Computers and Mathematics with Applications, 72(3):729–740, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302954.

Liu:2010:CSG

[LSD10] Yanjun Liu, Jie Sheng, and Ruifeng Ding. Convergence of stochastic gradient estimation algorithm for multivariable ARX-like systems. Computers and Mathematics with Applications, 59(8):2615–2627, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000520.

Lin:2012:SDL

[LSJ12] Fuming Lin, Daimin Shi, and Yingying Jiang. Some distributional limit theorems for the maxima of Gaussian vector sequences. Computers and Mathematics with Applications, 64(8):2497–2506, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200435X.

Lin:2010:LSC

[LSM10] Yuanqing Lin, Jinlong Shu, and Yao Meng. Laplacian spectrum characterization of extensions of vertices of wheel graphs and multi-fan graphs. Computers and Mathematics with Applications, 60(7):2003–2008, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005274.

Liu:2011:MFU

[LSM11] Huifang Liu, Daochun Sun, and Zhiqiang Mao. Meromorphic functions in the unit disc that share slowly growing functions in an angular domain.

Computers and Mathematics with Applications, 62(12):4539–4546, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008972.

Li:2017:OND

[LSS17] Jichun Li, Cengke Shi, and Chi-Wang Shu. Optimal non-dissipative discontinuous Galerkin methods for Maxwell’s equations in Drude metamaterials. Computers and Mathematics with Applications, 73(8):1760–1780, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300974.

Lopez-Salas:2018:PFS

[LSV18] José G. López-Salas and Carlos Vázquez. PDE formulation of some SABR/ LIBOR market models and its numerical solution with a sparse grid combination technique. Computers and Mathematics with Applications, 75(5):1616–1634, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307502.

Li:2010:PDD

[LSW10] Jing Li, Jiebao Sun, and Boying Wu. Periodic doubly degenerate parabolic equation with nonlocal terms. Computers and Mathematics with Applications, 60(3):490–500, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003287.

Lewis:2016:GFA

[LSW16] Allison Lewis, Ralph Smith, and Brian Williams. Gradient free active subspace construction using Morris screening elementary effects. Computers and Mathematics with Applications, 72 (6):1603–1615, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304242.

Lu:2013:CES

[LSX13] Zhuosheng Lü, Jianzhong Su, and Fuding Xie. Construction of exact solutions to the Jimbo– Miwa equation through Bäcklund transformation and symbolic computation. Computers and Mathematics with Applications, 65(4):648–656, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006554.

Liu:2011:SNE

[LSZ11] Wenjun Liu, Yanan Sun, and Qilin Zhang. Some new error inequalities for a generalized quadrature rule of open type. Computers and Mathematics with Applications, 62 (5):2218–2224, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100561X.

Li:2016:NSS

[LSZ16] Jianliang Li, Guanying Sun, and Ruming Zhang. The numerical solution of scattering by infinite rough interfaces based on the integral equation method. Computers and Mathematics with Applications, 71(7):1491–1502, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300852.

Luong:2011:CFP

[LT11] Nguyen Van Luong and Nguyen Xuan Thuan. Coupled fixed point theorems for mixed monotone mappings and an application to integral equations. Computers and Mathematics with Applications, 62(11):4238–4248, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008753.

Le:2013:CFE

[LT13] Kim-Ngan Le and Thanh Tran. A convergent finite element approximation for the quasistatic Maxwell–Landau– Lifshitz–Gilbert equations. Computers and Mathematics with Applications, 66(8):1389–1402, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004902.

Li:2015:MRT

[LT15a] Qingde Li and Jie Tian.

Multilevel refinable triangular PSP-splines (Tri-PSPS). Computers and Mathematics with Applications, 70(8):1781–1798, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003521.

Lin:2015:APA

[LT15b] Xiaoyan Lin and Xianhua Tang. An asymptotically periodic and asymptotically linear Schrödinger equation with indefinite linear part. Computers and Mathematics with Applications, 70(4):726– 736, August 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115002990.

Li:2018:EGS

[LT18] Gui-Dong Li and Chun-Lei Tang. Existence of a ground state solution for Choquard equation with the upper critical exponent. Computers and Mathematics with Applications, 76(11– 12):2635–2647, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304784.

Liu:2016:ASD

[LTJ^{+}16] De-Yin Liu, Bo Tian, Yan Jiang, Xi-Yang Xie, and Xiao-Yu Wu. Analytic study on a (2 + 1)-dimensional nonlinear Schrödinger equation in the Heisenberg ferromagnetism. Computers and Mathematics with Applications, 71(10):2001– 2007, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301493.

Liu:2012:PMS

[LTL^{+}12] Chen-Chung Liu, Chung-Yen Tsai, Jui Liu, Chun-Yuan Yu, and Shyr-Shen Yu. A pectoral muscle segmentation algorithm for digital mammograms using Otsu thresholding and multiple regression analysis. Computers and Mathematics with Applications, 64(5):1100–1107, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122112002337.

Liao:2012:DTI

[LTLL12] Yu-Yi Liao, Shen-Chuan Tai, Jzau-Sheng Lin, and Ping-Jui Liu. Degradation of turbid images based on the adaptive logarithmic algorithm. Computers and Mathematics with Applications, 64 (5):1259–1269, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002751.

Liu:2016:MTS

[LTSW16] Lei Liu, Bo Tian, Wen-Rong Sun, and Xiao-Yu Wu. Mixed-type soliton solutions for the N-coupled Hirota system in an optical fiber. Computers and Mathematics with Applications, 72(4):807–819, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302668.

Li:2013:SFA

[LTT13] Xian-Fang Li, Guo-Jin Tang, and Bao-Qing Tang. Stress field around a strike-slip fault in orthotropic elastic layers via a hypersingular integral equation. Computers and Mathematics with Applications, 66 (11):2317–2326, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005440.

Liao:2010:GFS

[LTX10] Maoxin Liao, Xianhua Tang, and Changjin Xu. General form of some rational recursive sequences. Computers and Mathematics with Applications, 59(1):360–364, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004441.

Li:2013:KBH

[LTX^{+}13] Ziqiang Li, Zhuojun Tian, Yanfang Xie, Rong Huang, and Jiyang Tan. A knowledge-based heuristic particle swarm optimization approach with the adjustment strategy for the weighted circle packing problem. Computers and Mathematics with Applications, 66 (10):1758–1769, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004926.

Lu:2011:AAF

[Lu11] Junfeng Lu. An analytical approach to the Fornberg–Whitham type equations by using the variational iteration method. Computers and Mathematics with Applications, 61(8):2010–2013, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006267.

Lu:2012:PAC

[Lu12] Lizheng Lu. On polynomial approximation of circular arcs and helices. Computers and Mathematics with Applications, 63(7):1192–1196, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010911.

Lu:2014:EMR

[Lü14] Dengfeng Lü. Existence and multiplicity results for perturbed Kirchhofftype Schrödinger systems in R^{3}. Computers and Mathematics with Applications, 68 (10):1180–1193, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400412X.

Luchko:2010:SUE

[Luc10]  Yury Luchko.  Some

uniqueness and existence results for the initialboundary-value problems for the generalized timefractional diffusion equation. Computers and Mathematics with Applications, 59(5):1766–1772, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005549.

Luh:2012:SPG

[Luh12] Lin-Tian Luh. The shape parameter in the Gaussian function. Computers and Mathematics with Applications, 63(3):687–694, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010121.

Lukashchuk:2011:EPF

[Luk11] S. Yu. Lukashchuk. Estimation of parameters in fractional subdiffusion equations by the time integral characteristics method. Computers and Mathematics with Applications, 62(3):834–844, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002215.

Luo:2018:ESS

[Luo18]  Xiao Luo.  Existence

and stability of standing waves for a planar gauged nonlinear Schrödinger equation. Computers and Mathematics with Applications, 76(11–12):2701–2709, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305078.

Luo:2019:GSS

[Luo19] Huxiao Luo. Ground state solutions of Pohožaev type for fractional Choquard equations with general nonlinearities. Computers and Mathematics with Applications, 77(3):877–887, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306217.

Lupas:2011:SDS

[Lup11] Alina Alb Lupas. On special differential superordinations using a generalized Salagean operator and Ruscheweyh derivative. Computers and Mathematics with Applications, 61(4):1048–1058, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009600.

Llibre:2011:NIB

[LV11] Jaume Llibre and Clàudia Valls. On the C^{1} non-integrability of the Belousov–Zhabotinskii system. Computers and Mathematics with Applications, 62(5):2342–2348, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211100575X.

Lu:2018:SAP

[LvdVX18] Zhongjie Lu, J. J. W.

van der Vegt, and Yan Xu. Spectral approximation for polynomial eigenvalue problems. Computers and Mathematics with Applications, 76 (5):1184–1197, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303316.

Leroy:2016:NOB

[LVF^{+}16] A. Leroy, D. Violeau, M. Ferrand, L. Fratter, and A. Joly. A new open boundary formulation for incompressible SPH. Computers and Mathematics with Applications, 72(9):2417–2432, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305107.

Lotstedt:2015:NOP

[LvS15] Per Lötstedt and Lina von Sydow. Numerical option pricing without oscillations using flux limiters. Computers and Mathematics with Applications, 70(1):1–10, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001728.

Liu:2010:CIE

[LW10] Hongjun Liu and Xingyuan Wang. Color image encryption based on onetime keys and robust chaotic maps. Computers and Mathematics with Applications, 59(10): 3320–3327, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110001938.

Li:2011:NMN

[LW11a] X. Y. Li and B. Y. Wu. A novel method for nonlinear singular fourth order four-point boundary value problems. Computers and Mathematics with Applications, 62(1):27– 31, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003300.

Li:2011:PBV

[LW11b] X. Y. Li and B. Y.

Wu. Periodic boundary value problems for neutral multi-pantograph equations. Computers and Mathematics with Applications, 61(8):1983–1986, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000619X.

Lin:2011:EPB

[LW11c] Yih-Lon Lin and Ming-Sheng Wu. An edge property-based neighborhood region search strategy for fractal image compression. Computers and Mathematics with Applications, 62(1):310– 318, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003993.

Lyuu:2011:CCB

[LW11d] Yuh-Dauh Lyuu and Chuan-Ju Wang. On the construction and complexity of the bivariate lattice with stochastic interest rate models. Computers and Mathematics with Applications, 61(4):1107–1121, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009661.

Liu:2012:CSL

[LW12a] Meng Liu and Ke Wang. Corrigendum to “On a stochastic logistic equation with impulsive perturbations” [Comput. Math. Appl. 63 (2012) 871–886]. Computers and Mathematics with Applications, 64(6):2158, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004440.

See [LW12b].

Liu:2012:SLE

[LW12b] Meng Liu and Ke Wang. On a stochastic logistic equation with impulsive perturbations. Computers and Mathematics with Applications, 63(5):871– 886, March 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111009692.

See corrigendum [LW12a].

Lo:2012:SVM

[LW12c] Chien-Shun Lo and Chuin-Mu Wang. Support vector machine for breast MR image classification. Computers and Mathematics with Applications, 64(5):1153–1162, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002386.

Long:2013:QMS

[LW13] Qiang Long and Changzhi Wu. A quasisecant method for solving a system of nonsmooth equations. Computers and Mathematics with Applications, 66(4):419–431, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003209.

Li:2015:HBG

[LW15] Yan Li and Mingxin Wang. Hopf bifurcation and global stability of a delayed predator– prey model with prey harvesting. Computers and Mathematics with Applications, 69(5):398–410, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500005X.

Li:2017:PEO

[LW17] Wen Li and Song Wang. Pricing European options with proportional transaction costs and stochastic volatility using a penalty approach and a finite volume scheme. Computers and Mathematics with Applications, 73(11):2454–2469, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301980.

Liu:2018:EMF

[LW18a] Chein-Shan Liu and Fajie Wang. An energy method of fundamental solutions for solving the inverse Cauchy problems of the Laplace equation. Computers and Mathematics with Applications, 75(12):4405–4413, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301743.

Liu:2018:MMS

[LW18b] Chein-Shan Liu and Fajie Wang. A meshless method for solving the nonlinear inverse Cauchy problem of elliptic type equation in a doublyconnected domain. Computers and Mathematics with Applications, 76(8):1837–1852, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304036.

Liu:2018:ETD

[LW18c] Jun Liu and Zhu Wang.

Efficient time domain decomposition algorithms for parabolic PDE-constrained optimization problems. Computers and Mathematics with Applications, 75(6):2115–2133, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305655.

LeBorne:2019:DDM

[LW19a] Sabine Le Borne and Michael Wende. Domain decomposition methods in scattered data interpolation with conditionally positive definite radial basis functions. Computers and Mathematics with Applications, 77(4):1178–1196, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306394.

Liu:2019:BTE

[LW19b] Bingchen Liu and Guicheng Wu. Blow-up time estimates and simultaneous blow-up of solutions in nonlinear diffusion problems. Computers and Mathematics with Applications, 77(2):597–614, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305807.

Li:2013:DPN

[LWBW13] Guanfeng Li, Yong Wang, Gejun Bao, and Tingting Wang. The Dirichlet problem of nonhomogeneous A-harmonic equations in unbounded open sets and some estimates. Computers and Mathematics with Applications, 65(11):1795– 1807, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001661.

Luo:2013:DFI

[LWC13] Zhaoming Luo, Guangjun Wang, and Hong Chen. Decentralized fuzzy inference method for estimating thermal boundary condition of a heated cylinder normal to a laminar air stream. Computers and Mathematics with Applications, 66 (10):1869–1878, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004513.

Li:2015:USP

[LWD15] Shanbing Li, Jianhua Wu, and Yaying Dong. Uniqueness and stability of a predator–prey model with c-m functional response. Computers and Mathematics with Applications, 69(10):1080– 1095, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001029.

Liu:2012:RRO

[LWHL12] Xue-Mei Liu, Sheng-Min Wang, Ai-Min Hao, and Huan Liu. Realistic rendering of organ for surgery simulator. Computers and Mathematics with Applications, 64(5):721–728, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010108.

Lin:2010:EPS

[LWHY10] Han-Yu Lin, Tzong-Sun Wu, Shih-Kun Huang, and Yi-Shiung Yeh. Efficient proxy signcryption scheme with provable CCA and CMA security. Computers and Mathematics with Applications, 60(7):1850–1858, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004918.

Liang:2010:SLA

[LWJ10] Zongxia Liang, Weiming Wu, and Shuqing Jiang. Stock loan with automatic termination clause, cap and margin. Computers and Mathematics with Applications, 60 (12):3160–3176, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000800X.

Liu:2010:EUM

[LWKK10] Zeqing Liu, Lili Wang, Gang Il Kim, and Shin Min Kang. Existence of uncountably many bounded positive solutions for a third order nonlinear neutral delay difference equation. Computers and Mathematics with Applications, 60(8):2399–2416, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005997.

Liang:2011:TMS

[LWL11] Yueqian Liang, Guangmei Wei, and Xiaonan Li. Transformations and multi-solitonic solutions for a generalized variable-coefficient Kadomtsev–Petviashvili equation. Computers and Mathematics with Applications, 61(11):3268– 3277, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100294X.

Liu:2014:AFE

[LWL14] Wei Liu, Zhifeng Wang, and Jin Li. Anisotropic finite element approximation for a coupled continuum pipe-flow/Darcy model in Karst aquifers. Computers and Mathematics with Applications, 68(3):86–100, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002028.

Li:2015:SSB

[LWN15] Shanbing Li, Jianhua Wu, and Hua Nie. Steadystate bifurcation and Hopf bifurcation for a diffusive Leslie–Gower predator–prey model. Computers and Mathematics with Applications, 70(12):3043–3056, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500512X.

Liu:2018:TTS

[LWQZ18] Zhongyun Liu, Nianci Wu, Xiaorong Qin, and Yulin Zhang. Trigonometric transform splitting methods for real symmetric Toeplitz systems. Computers and Mathematics with Applications, 75(8):2782–2794, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300129.

Laniewski-Wollk:2016:ALB

[LWR16] L. Laniewski-Wollk and J. Rokicki. Adjoint lattice Boltzmann for topology optimization on multi-GPU architecture. Computers and Mathematics with Applications, 71(3):833–848, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115006215.

Liu:2019:BER

[LWSL19] Bingchen Liu, Guicheng Wu, Xizheng Sun, and Fengjie Li. Blow-up estimate in a reaction– diffusion equation with nonlinear nonlocal flux and source. Computers and Mathematics with Applications, 78 (6):1862–1877, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301518.

Liu:2019:SSL

[LWW19a] Yaqing Liu, Xiao-Yong Wen, and Deng-Shan Wang. The N-soliton solution and localized wave interaction solutions of the (2 + 1)- dimensional generalized Hirota–Satsuma–Ito equation. Computers and Mathematics with Applications, 77(4):947–966, February 15, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306321.

Liu:2019:NIP

[LWW19b] Yaqing Liu, Xiao-Yong Wen, and Deng-Shan Wang. Novel interaction phenomena of localized waves in the generalized (3 + 1)-dimensional KP equation. Computers and Mathematics with Applications, 78(1):1–19, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122119301245.

Lee:2012:NBR

[LWWY12] Jiann-Shu Lee, Jing-Wein Wang, Hsing-Hsien Wu, and Ming-Zheng Yuan. A nonparametric-based rib suppression method for chest radiographs. Computers and Mathematics with Applications, 64 (5):1390–1399, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002891.

Liu:2016:STF

[LWZ16] Xiaojing Liu, Jizeng Wang, and Youhe Zhou. A space-time fully decoupled wavelet Galerkin method for solving twodimensional Burgers’ equations. Computers and Mathematics with Applications, 72(12):2908– 2919, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116305879.

Lin:2010:DSM

[LWZG10] Dai-Rui Lin, Chih-I Wang, Zhi-Kai Zhang, and D. J. Guan. A digital signature with multiple subliminal channels and its applications. Computers and Mathematics with Applications, 60(2):276– 284, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000012X.

Lai:2010:AGS

[LX10a] S. K. Lai and Y. Xiang. Application of a generalized Senator– Bapat perturbation technique to nonlinear dynamical systems with an irrational restoring force. Computers and Mathematics with Applications, 60(7):2078–2086, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005407.

Liu:2010:CSL

[LX10b] Guangying Liu and Yongqing Xu. Capped stock loans. Computers and Mathematics with Applications, 59(11):3548–3558, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110002373.

Li:2012:FBH

[LX12a] Ming Li and Xiangtuan Xiong. On a fractional backward heat conduction problem: Application to deblurring. Computers and Mathematics with Applications, 64(8):2594–2602, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004695.

Liang:2012:SSC

[LX12b] Xiannuan Liang and Yang Xiao. Studying the stochastic capturing of moving intruders by mobile sensors. Computers and Mathematics with Applications, 64(8):2431–2449, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004154.

Liu:2015:GEF

[LX15a] Gongwei Liu and Suxia Xia. Global existence and finite time blow up for a class of semilinear wave equations on R^{N}. Computers and Mathematics with Applications, 70(6):1345–1356, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003569.

Liu:2015:EMN

[LX15b] Pengfei Liu and Liang Xiao. Efficient multiplicative noise removal method using isotropic second order total variation. Computers and Mathematics with Applications, 70(8):2029–2048, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003831.

Lee:2016:FDE

[LX16] Hyesuk Lee and Shuhan Xu. Fully discrete error estimation for a quasi-Newtonian fluid-structure interaction problem. Computers and Mathematics with Applications, 71(11): 2373–2388, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115005921.

Li:2017:TWS

[LX17] Huiru Li and Haibin Xiao. Traveling wave solutions for diffusive predator-prey type systems with nonlinear density dependence. Computers and Mathematics with Applications, 74 (10):2221–2230, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304066.

Li:2018:WBF

[LX18] Gang Li and Yulong Xing. Well-balanced finite difference weighted essentially non-oscillatory schemes for the Euler equations with static gravitational fields. Computers and Mathematics with Applications, 75(6): 2071–2085, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117306636.

Liu:2011:FGG

[LXF11] Hongbing Liu, Shengwu Xiong, and Zhixiang Fang. FL–GrCCA: a granular computing classification algorithm based on fuzzy lattices.

Computers and Mathematics with Applications, 61(1):138–147, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008308.

Liu:2011:BOC

[LXK11] Zeqing Liu, Yuguang Xu, and Shin Min Kang. Bounded oscillation criteria for certain third order nonlinear difference equations with several delays and advances. Computers and Mathematics with Applications, 61(4):1145–1161, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009697.

Li:2012:TSG

[LXL12] Zhaowen Li, Tusheng Xie, and Qingguo Li. Topological structure of generalized rough sets. Computers and Mathematics with Applications, 63(6):1066–1071, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122111010546.

Lin:2014:SSS

[LXL14] Hongwei Lin, Yunyang Xiong, and Hongwei Liao. Semi-structured B-spline for blending two B-spline surfaces. Computers and Mathematics with Applications, 68(7):706–718, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003319.

Liu:2011:NID

[LXP11] G. Liu, X. Xiang, and Y. Peng. Nonlinear integro-differential equations and optimal control problems on time scales. Computers and Mathematics with Applications, 61(2):155–169, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007923.

Lu:2019:ALS

[LXY19] Changna Lu, Luoyan Xie, and Hongwei Yang. Analysis of Lie symmetries with conservation laws and solutions for the generalized (3 + 1)- dimensional time fractional Camassa–Holm– Kadomtsev–Petviashvili equation. Computers and Mathematics with Applications, 77 (12):3154–3171, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300525.

Luo:2011:RSQ

[LXYT11] Chuanyi Luo, Kaili Xiang, Miaomiao Yu, and Yinghui Tang. Recursive solution of queue length distribution for Geo/G/1 queue with single server vacation and variable input rate. Computers and Mathematics with Applications, 61(9):2401–2411, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001040.

Leng:2013:MII

[LXZ13] Juelin Leng, Guoliang Xu, and Yongjie Zhang. Medical image interpolation based on multiresolution registration. Computers and Mathematics with Applications, 66(1):1–18, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002538.

Li:2018:BPW

[LXZ18] Binjie Li, Xiaoping Xie, and Shiquan Zhang. BPS preconditioners for a weak Galerkin finite element method for 2D diffusion problems with strongly discontinuous coefficients. Computers and Mathematics with Applications, 76(4):701–724, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302736.

Li:2010:RPB

[LY10a] Hanyu Li and Hu Yang.

Relative perturbation bounds for weighted polar decomposition. Computers and Mathematics with Applications, 59(2):853– 860, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109007081.

Li:2010:SCC

[LY10b] Yuan-Chuan Li and Cheh-Chih Yeh. Some characterizations of convex functions. Computers and Mathematics with Applications, 59(1):327–337, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109004659.

Li:2011:UMF

[LY11a] Xiao-Min Li and Hong-Xun Yi. Uniqueness of meromorphic functions whose certain nonlinear differential polynomials share a polynomial. Computers and Mathematics with Applications, 62(2):539–550, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002926.

Li:2011:TMI

[LY11b] Xiao-Xiao Li and Fan Yang. The truncation method for identifying the heat source dependent on a spatial variable. Computers and Mathematics with Applications, 62(6):2497–2505, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005931.

Liu:2011:PPS

[LY11c] Guirong Liu and Jurang Yan. Positive periodic solutions of neutral predator–prey model with Beddington– DeAngelis functional response. Computers and Mathematics with Applications, 61(8):2317–2322, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007807.

Liu:2011:EGC

[LY11d] Xifu Liu and Hu Yang.

An expression of the general common leastsquares solution to a pair of matrix equations with applications. Computers and Mathematics with Applications, 61(10):3071–3078, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002768.

Li:2012:ATC

[LY12a] Tzuu-Hseng S. Li and Her-Terng Yau. Advanced technologies in computer, consumer and control. Computers and Mathematics with Applications, 64(5):687, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004208.

Li:2012:MOB

[LY12b] Xiangtao Li and Minghao Yin. Multi-operator based biogeography based optimization with mutation for global numerical optimization. Computers and Mathematics with Applications, 64 (9):2833–2844, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003550.

Li:2013:AFE

[LY13] Hao Li and Yidu Yang. The adaptive finite element method based on multi-scale discretizations for eigenvalue problems. Computers and Mathematics with Applications, 65(7):1086–1102, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000734.

Li:2014:MPS

[LY14] Qin Li and Zuodong Yang. Multiple positive solutions for a quasilinear elliptic system with critical exponent and sign-changing weight. Computers and Mathematics with Applications, 67(10):1848– 1863, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001515.

Li:2015:MPI

[LY15] Xiaolin Li and Chunjun Yu. Meshless projection iterative analysis of Signorini problems using a boundary element-free method. Computers and Mathematics with Applications, 70(5):869–882, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002916.

Li:2017:WPX

[LY17] Bin Li and Han Yang. The Wigner–Poisson–Xα system in Wiener algebra. Computers and Mathematics with Applications, 73(9):1987–1995, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301165.

Liao:2012:EAT

[LYC12] Yen-Far Liao, Dun-Han Yau, and Chieh-Li Chen. Evolutionary algorithm to traveling salesman problems. Computers and Mathematics with Applications, 64(5):788–797, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211101073X.

Liu:2015:AAM

[LYC15] Hui Liu, Bo Yang, and Zhangxin Chen. Accelerating algebraic multigrid solvers on NVIDIA GPUs. Computers and Mathematics with Applications, 70(5):1162–1181, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003405.

Liu:2019:ARW

[LYC^{+}19] Jiangen Liu, Xiaojun Yang, Menghong Cheng, Yiying Feng, and Yaodong Wang. Abound rogue wave type solutions to the extended (3 + 1)- dimensional Jimbo–Miwa equation. Computers and Mathematics with Applications, 78 (6):1947–1959, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301592.

Lee:2015:TOC

[LYJ15] Jong Wook Lee, Gil Ho Yoon, and Seung Hyun Jeong. Topology optimization considering fatigue life in the frequency domain. Computers and Mathematics with Applications, 70(8):1852–1877, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003740.

Lin:2012:LBM

[LYL12] Lendy Lin, Weipang Yang, and Jyhjong Lin. A layer-based method for rapid software development. Computers and Mathematics with Applications, 64(5):1364–1375, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002878.

Li:2011:SDD

[LYLX11] Jinquan Li, Xuehai Yuan, E. S. Lee, and Dehua Xu. Setting due dates to minimize the total weighted possibilistic mean value of the weighted earlinesstardiness costs on a single machine. Computers and Mathematics with Applications, 62 (11):4126–4139, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008443.

Lee:2012:SFS

[LYM12] Kyungroul Lee, Kangbin Yim, and Mohammad A. Mikki. A secure framework of the surveillance video network integrating heterogeneous video formats and protocols. Computers and Mathematics with Applications, 63(2):525–535, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007218.

Liu:2011:MDP

[LYN11] Chaofeng Liu, Jidong Yang, and Yushan Ni. A multiplicative decomposition of Poiseuille number on rarefaction and roughness by lattice Boltzmann simulation. Computers and Mathematics with Applications, 61(12):3528– 3536, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002191.

Li:2012:EPA

[LYS12a] Hanyu Li, Hu Yang, and Hua Shao. Erratum to “Perturbation analysis for the hyperbolic QR factorization” [Comput. Math. Appl. 63 (2012) 1607– 1620]. Computers and Mathematics with Applications, 64(4):686, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004452.

See [LYS12b].

Li:2012:PAH

[LYS12b] Hanyu Li, Hu Yang, and Hua Shao. Perturbation analysis for the hyperbolic QR factorization. Computers and Mathematics with Applications, 63(12):1607– 1620, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002416.

See erratum [LYS12a].

Lu:2019:UIR

[LYSZ19] Jian Lu, Hanmei Yang, Lixin Shen, and Yuru Zou. Ultrasound image restoration based on a learned dictionary and a higher-order MRF. Computers and Mathematics with Applications, 77(4):991–1009, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830628X.

Liu:2011:AEU

[LYX11] Wei Liu, Hualiang Yan, and Jianguo Xiao. Automatically extracting user reviews from forum sites. Computers and Mathematics with Applications, 62(7):2779–2792, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006158.

Lin:2012:CSC

[LYY12] Chih-Jer Lin, Shyi-Kae Yang, and Her-Terng Yau. Chaos suppression control of a coronary artery system with uncertainties by using variable structure control. Computers and Mathematics with Applications, 64(5):988–995, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200212X.

Li:2011:BMS

[LYZ11] Zhao-Xiang Li, Zhong-Hua Yang, and Hai-Long Zhu. A bifurcation method for solving multiple positive solutions to the boundary value problem of the Henon equation on a unit disk. Computers and Mathematics with Applications, 62(10):3775–3784, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007942.

Liu:2017:RCE

[LYZ17] Jiu Liu, Zhongquan Yan, and Zhibo Zheng. A result on a class of elliptic equations involving Kirchhoff type nonlocal term. Computers and Mathematics with Applications, 73(2):355–361, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630668X.

Li:2010:EGG

[LZ10] Yinkui Li and Shenggui Zhang. Extremal graphs with given order and the rupture degree. Computers and Mathematics with Applications, 60(6):1706–1710, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110004761.

Lan:2011:LBR

[LZ11a] Yong-Hong Lan and Yong Zhou. LMI-based robust control of fractionalorder uncertain linear systems. Computers and Mathematics with Applications, 62(3):1460–1471, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001805.

Lashari:2011:GDV

[LZ11b] Abid Ali Lashari and Gul Zaman. Global dynamics of vector-borne diseases with horizontal transmission in host population. Computers and Mathematics with Applications, 61(4):745–754, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009235.

Lee:2011:FAI

[LZ11c] Eun-Joo Lee and Jun Zhang. Factored approximate inverse preconditioners with dynamic sparsity patterns. Computers and Mathematics with Applications, 62(1): 235–242, July 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111003749.

Liang:2011:EUS

[LZ11d] Sihua Liang and Jihui Zhang. Existence and uniqueness of strictly nondecreasing and positive solution for a fractional three-point boundary value problem. Computers and Mathematics with Applications, 62(3):1333–1340, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002367.

Liang:2011:ETP

[LZ11e] Sihua Liang and Jihui Zhang. Existence of three positive solutions of m-point boundary value problems for some nonlinear fractional differential equations on an infinite interval. Computers and Mathematics with Applications, 61(11):3343– 3354, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003051.

Li:2012:EUM

[LZ12a] Shunyong Li and Xiaoqin Zhang. Existence and uniqueness of monotone positive solutions for an elastic beam equation with nonlinear boundary conditions. Computers and Mathematics with Applications, 63(9):1355– 1360, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011369.

Lu:2012:STA

[LZ12b] Haishu Lu and Jihui Zhang. A section theorem with applications to coincidence theorems and minimax inequalities in F W C-spaces. Computers and Mathematics with Applications, 64(4):579–588, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011321.

Luo:2012:PAP

[LZ12c] Mengzhuo Luo and Shouming Zhong. Passivity analysis and passification of uncertain Markovian jump systems with partially known transition rates and modedependent interval timevarying delays. Computers and Mathematics with Applications, 63(7):1266– 1278, April 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111011412.

Liang:2014:GSI

[LZ14] Chuangchuang Liang and Kaijun Zhang. Global solution of the initial boundary value problem to a hyperbolic nonlocal MEMS equation. Computers and Mathematics with Applications, 67(3):549–554, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006688.

Liang:2015:PBC

[LZ15a] Zhao-Zheng Liang and Guo-Feng Zhang. On PSS-based constraint preconditioners for singular nonsymmetric saddle point problems. Computers and Mathematics with Applications, 69(6):455– 465, March 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115000073.

Liu:2015:SGM

[LZ15b] Bochao Liu and Chengjian Zhang. A spectral Galerkin method for nonlinear delay convection– diffusion-reaction equations. Computers and Mathematics with Applications, 69(8):709–724, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000942.

Li:2016:ENS

[LZ16a] Can Li and Shan Zhao.

Efficient numerical schemes for fractional water wave models. Computers and Mathematics with Applications, 71(1):238–254, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005611.

Liang:2016:TNV

[LZ16b] Zhao-Zheng Liang and Guo-Feng Zhang. Two new variants of the HSS preconditioner for regularized saddle point problems. Computers and Mathematics with Applications, 72(3):603–619, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302814.

Liao:2016:PCL

[LZ16c] Li-Dan Liao and Guo-Feng Zhang. Preconditioning of complex linear systems from the Helmholtz equation. Computers and Mathematics with Applications, 72(9):2473–2485, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305065.

Liao:2017:NVH

[LZ17] Li-Dan Liao and Guo-Feng Zhang. New variant of the HSS iteration method for weighted Toeplitz regularized leastsquares problems from image restoration. Computers and Mathematics with Applications, 73(11):2482–2499, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302018.

Lechleiter:2018:NPA

[LZ18a] Armin Lechleiter and Ruming Zhang. Nonperiodic acoustic and electromagnetic, scattering from periodic structures in 3D. Computers and Mathematics with Applications, 74 (11):2723–2738, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305345.

Liang:2018:VDP

[LZ18b] Zhao-Zheng Liang and Guo-Feng Zhang. Variants of the deteriorated PSS preconditioner for saddle point problems. Computers and Mathematics with Applications, 75(8):3024–3046, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300464.

Liu:2018:USG

[LZ18c] Xiaowei Liu and Jin Zhang. Uniform supercloseness of Galerkin finite element method for convection–diffusion problems with characteristic layers. Computers and Mathematics with Applications, 75(2):444–458, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305916.

Liang:2019:PTG

[LZ19a] Hongxia Liang and Tong Zhang. Parallel two-grid finite element method for the time-dependent natural convection problem with non-smooth initial data. Computers and Mathematics with Applications, 77(8):2221–2241, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306941.

Liang:2019:FRM

[LZ19b] Maolin Liang and Bing Zheng. Further results on Moore–Penrose inverses of tensors with application to tensor nearness problems. Computers and Mathematics with Applications, 77(5):1282–1293, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306515.

Liu:2019:BWS

[LZ19c] Jian-Guo Liu and Wen-Hui Zhu. Breather wave solutions for the generalized shallow water wave equation with variable coefficients in the atmosphere, rivers, lakes and oceans. Computers and Mathematics with Applications, 78(3):848–856, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301270.

Liu:2012:NMA

[LZB12] F. Liu, P. Zhuang, and K. Burrage. Numerical methods and analysis for a class of fractional advectiondispersion models. Computers and Mathematics with Applications, 64 (10):2990–3007, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000302.

Liu:2015:UGK

[LZB15] Sha Liu, Chengwen Zhong, and Jie Bai. Unified gas-kinetic scheme for microchannel and nanochannel flows. Computers and Mathematics with Applications, 69(1): 41–57, January 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005537.

Li:2011:NAN

[LZC11] Changpin Li, Zhengang Zhao, and YangQuan Chen. Numerical approximation of nonlinear fractional differential equations with subdiffusion and superdiffusion. Computers and Mathematics with Applications, 62(3):855–875, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001441.

Li:2013:SDR

[LZC13] Junmin Li, Weiyuan Zhang, and Minglai Chen.

Synchronization of delayed reaction–diffusion neural networks via an adaptive learning control approach. Computers and Mathematics with Applications, 65(11):1775– 1785, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001685.

Liu:2018:ADC

[LZCL18] Lin Liu, Liancun Zheng, Yanping Chen, and Fawang Liu. Anomalous diffusion in comb model with fractional dualphase-lag constitutive relation. Computers and Mathematics with Applications, 76(2):245–256, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302141.

Li:2017:CMF

[LZD17] Xiangli Li, Wen Zhang, and Xiaoliang Dong. A class of modified FR conjugate gradient method and applications to non-negative matrix factorization. Computers and Mathematics with Applications, 73(2):270–276, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306344.

Liu:2018:OIP

[LZF18] Limin Liu, Lin Zhang, and Shiqi Fan. Optimal investment problem under non-extensive statistical mechanics. Computers and Mathematics with Applications, 75(10):3549–3557, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300890.

Liu:2013:LIC

[LZG13] Qiao Liu, Pingan Zhang, and Sadek Gala. Logarithmically improved criteria for the 3D nematic liquid crystal flows in the Morrey–Campanato space. Computers and Mathematics with Applications, 66(11):2327– 2334, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005592.

Liu:2019:FRD

[LZG19] Qiang Liu, Zhiguang Zhang, and Zhichang Guo. On a fractional reaction–diffusion system applied to image decomposition and restoration. Computers and Mathematics with Applications, 78(5):1739–1751, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302937.

Li:2011:SSS

[LZGZ11] Ying Li, Fengxia Zhang, Wenbin Guo, and Jianli Zhao. Solutions with special structure to the linear matrix equation AX = B. Computers and Mathematics with Applications, 61(2):374–383, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000862X.

Li:2012:FSS

[LZH12] Zhaowen Li, Dingwei Zheng, and Jing Hao. L-fuzzy soft sets based on complete Boolean lattices. Computers and Mathematics with Applications, 64(8):2558–2574, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004658.

Li:2016:EEP

[LZH16] Jin Li, Qingli Zhao, and Hongying Huang. Error expansion of piecewise constant interpolation rule for certain two-dimensional Cauchy principal value integrals. Computers and Mathematics with Applications, 72(9):2119–2142, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304552.

Liu:2012:SGI

[LZJ12] Jianzhong Liu, Xuande Zhang, and Bo Jiang. Some generalizations and improvements of discrete Hardy’s inequality. Computers and Mathematics with Applications, 63(3):601–607, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007875.

Liu:2016:HOF

[LZJY16] Yilang Liu, Weiwei Zhang, Yuewen Jiang, and Zhengyin Ye. A high-order finite volume method on unstructured grids using RBF reconstruction. Computers and Mathematics with Applications, 72(4):1096–1117, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630356X.

Liu:2011:EUM

[LZKU11] Zeqing Liu, Liangshi Zhao, Shin Min Kang, and Jeong Sheok Ume. Existence of uncountably many bounded positive solutions for second order nonlinear neutral delay difference equations. Computers and Mathematics with Applications, 61(9):2535–2545, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001453.

Li:2013:IFS

[LZL^{+}13] Maozhen Li, Xiao-Jun Zeng, Frank H. F. Leung, Shigang Yue, and Lipo Wang. ICNC–FSKD 2012 Special Issue on “Computers & Mathematics in Natural Computation and Data Mining”. Computers and Mathematics with Applications, 66 (10):1727–1728, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005956.

Le:2016:IEA

[LZL16] Nhat-Tan Le, Song-Ping Zhu, and Xiaoping Lu. An integral equation approach for the valuation of American-style downand-out calls with rebates. Computers and Mathematics with Applications, 71(2):544–564, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005817.

Lai:2018:ICA

[LZL^{+}18] Yicong Lai, Yongjie Jessica Zhang, Lei Liu, Xiaodong Wei, Eugene Wang, and Jim Lua. Integrating CAD with Abaqus: A practical isogeometric analysis software platform for industrial applications. Computers and Mathematics with Applications, 74(7):1648–1660, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302079.

Li:2019:EWS

[LZL19] Wei Li, Yan Zhang, and Yinping Liu. Exact wave solutions for a (3 + 1)-dimensional generalized B-type Kadomtsev– Petviashvili equation. Computers and Mathematics with Applications, 77(12):3087–3101, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300574.

Liu:2018:HOL

[LZLL18] Yang Liu, Min Zhang, Hong Li, and Jichun Li. High-order local discontinuous Galerkin method combined with WSGDapproximation for a fractional subdiffusion equation. Computers and Mathematics with Applications, 73(6):1298–1314, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304692.

Liu:2018:RSC

[LZM18] Jiangen Liu, Yufeng Zhang, and Iqbal Muhammad. Resonant soliton and complexiton solutions for (3 + 1)- dimensional Boiti–Leon– Manna–Pempinelli equation. Computers and Mathematics with Applications, 75(11):3939– 3945, June 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301330.

Liu:2019:EFM

[LZP^{+}19] Chunyan Liu, Liancun Zheng, Mingyang Pan, Ping Lin, and Fawang Liu. Effects of fractional mass transfer and chemical reaction on MHD flow in a heterogeneous porous medium. Computers and Mathematics with Applications, 78(8):2618–2631, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302093.

Li:2019:GKS

[LZPZ19] Ji Li, Chengwen Zhong, Dongxin Pan, and Congshan Zhuo. A gas-kinetic scheme coupled with SST model for turbulent flows. Computers and Mathematics with Applications, 78(4):1227–1242, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305144.

Lu:2012:MLE

[LZS12] Huimin Lu, Lifeng Zhang, and Seiichi Serikawa. Maximum local energy: an effective approach for multisensor image fusion in beyond wavelet transform domain. Computers and Mathematics with Applications, 64 (5):996–1003, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002222.

Li:2011:ROF

[LZT11] Shukai Li, Jianxiong Zhang, and Wansheng Tang. Robust H_{∞} output feedback control for uncertain complex delayed dynamical networks. Computers and Mathematics with Applications, 62(1):497–505, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004317.

Li:2016:AAE

[LZWC16] Xiaolin Li, Shougui Zhang, Yan Wang, and Hao Chen. Analysis and application of the element-free Galerkin method for nonlinear sine-Gordon and generalized sinh-Gordon equations. Computers and Mathematics with Applications, 71(8):1655– 1678, April 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116301237.

Zhu:2012:ACC

[lZxLhY12] Hai long Zhu, Zhao xiang Li, and Zhong hua Yang. Analysis and computation for a class of semilinear elliptic boundary value problems. Computers and Mathematics with Applications, 64(8):2735–2743, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005366.

Liu:2011:OEO

[LZY11] Shouhua Liu, Quanxin Zhang, and Yuanhong Yu. Oscillation of even-order half-linear functional differential equations with damping. Computers and Mathematics with Applications, 61(8):2191–2196, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006978.

Li:2012:DRS

[LZY12] Yunhua Li, Qi Zheng, and Liman Yang. Design of robust sliding mode control with disturbance observer for multi-axis coordinated traveling system. Computers and Mathematics with Applications, 64(5):759–765, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010364.

Li:2013:UDM

[LZYW13] Chengdong Li, Guiqing Zhang, Jianqiang Yi, and Ming Wang. Uncertainty degree and modeling of interval type-2 fuzzy sets: Definition, method and application. Computers and Mathematics with Applications, 66(10):1822–1835, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004665.

Liu:2011:MFH

[LZZ11a] Yaqing Liu, Liancun Zheng, and Xinxin Zhang. MHD flow and heat transfer of a generalized Burgers’ fluid due to an exponential accelerating plate with the effect of radiation. Computers and Mathematics with Applications, 62(8):3123–3131, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006882.

Liu:2011:UMC

[LZZ11b] Yaqing Liu, Liancun Zheng, and Xinxin Zhang. Unsteady MHD Couette flow of a generalized Oldroyd-b fluid with fractional derivative. Computers and Mathematics with Applications, 61(2):443–450, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008709.

Liao:2018:RPO

[LZZ18] Li-Dan Liao, Guo-Feng Zhang, and Lei Zhang. Robust preconditioners for optimal control with time-periodic parabolic equation. Computers and Mathematics with Applications, 76 (10):2514–2522, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304760.

Li:2019:TOG

[LZZ19] Ji Li, Chengwen Zhong, and Congshan Zhuo. A third order gas-kinetic scheme for unstructured grid. Computers and Mathematics with Applications, 78(1):92–109, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930094X.

Li:2012:NBF

[LZZC12] Kai Li, Chengwen Zhong, Congshan Zhuo, and Jun Cao. Non-body-fitted Cartesian-mesh simulation of highly turbulent flows using multirelaxation-time lattice Boltzmann method. Computers and Mathematics with Applications, 63(10): 1481–1496, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002854.

Ma:2010:CAM

[Ma10a] Jingtang Ma. Convergence analysis of moving Godunov methods for dynamical boundary layers. Computers and Mathematics with Applications, 59(1):80–93, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006580.

Makinde:2010:SLA

[MA10b] O. D. Makinde and A. Aziz. Second law analysis for a variable viscosity plane Poiseuille flow with asymmetric convective cooling. Computers and Mathematics with Applications, 60 (11):3012–3019, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007601.

Majidi:2016:AIS

[MA16] Sahand Majidi and Asghar Afshari. An adaptive interface sharpening methodology for compressible multiphase flows. Computers and Mathematics with Applications, 72(10):2660– 2684, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116305351.

Malik:2017:ISP

[MA17] Salman A. Malik and Sara Aziz. An inverse source problem for a two parameter anomalous diffusion equation with nonlocal boundary conditions. Computers and Mathematics with Applications, 73(12):2548–2560, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117301815.

Ma:2018:GET

[Ma18a] Liangliang Ma. Global existence of three-dimensional incompressible magnetomicropolar system with mixed partial dissipation, magnetic diffusion and angular viscosity. Computers and Mathematics with Applications, 75(1):170–186, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305576.

Ma:2018:GRM

[Ma18b] Liangliang Ma. Global regularity for the 2D magnetic Bénard fluid system with mixed partial viscosity. Computers and Mathematics with Applications, 76 (9):2148–2166, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304310.

Ma:2019:GSC

[Ma19] Wen-Xiu Ma. A generating scheme for conservation laws of discrete zero curvature equations and its application. Computers and Mathematics with Applications, 78 (10):3422–3428, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302755.

Mansouri:2019:AST

[MAB19] Djamel Mansouri, Salem Abdelmalek, and Samir Bendoukha. On the asymptotic stability of the time-fractional lengyel-Epstein system. Computers and Mathematics with Applications, 78 (5):1415–1430, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302135.

Machado:2012:EFO

[Mac12a] J. A. Tenreiro Machado.

The effect of fractional order in variable structure control. Computers and Mathematics with Applications, 64 (10):3340–3350, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001186.

Machado:2012:MSA

[Mac12b] J. Tenreiro Machado. Multidimensional scaling analysis of fractional systems. Computers and Mathematics with Applications, 64 (10):2966–2972, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003458.

Magin:2010:FCM

[Mag10] Richard L. Magin. Fractional calculus models of complex dynamics in biological tissues. Computers and Mathematics with Applications, 59(5):1586–1593, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005343.

Mahmudov:2010:APC

[Mah10a] N. I. Mahmudov. Approximation properties of complex q-Szász–Mirakjan operators in compact disks. Computers and Mathematics with Applications, 60(6):1784–1791, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004840.

Mahmudov:2010:CPI

[Mah10b] N. I. Mahmudov. Convergence properties and iterations for q-Stancu polynomials in compact disks. Computers and Mathematics with Applications, 59(12):3763– 3769, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002816.

Mahmudov:2011:APP

[Mah11] N. I. Mahmudov. Asymptotic properties of powers of linear positive operators which preserve e_{2}. Computers and Mathematics with Applications, 62(12):4568–4575, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100900X.

Mahipal:2014:SDM

[Mah14] J. Mahipal. A scheme with discrete maximum principle and optimized rotation invariance for coherence enhancing diffusion. Computers and Mathematics with Applications, 68(8):859–871, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003617.

Mu:2018:ERS

[MAH18] Jia Mu, Bashir Ahmad, and Shuibo Huang. Existence and regularity of solutions to timefractional diffusion equations. Computers and Mathematics with Applications, 73(6):985–996, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302322.

Mainge:2010:VAP

[Mai10] Paul-Emile Maingé. The viscosity approximation process for quasi-nonexpansive mappings in Hilbert spaces. Computers and Mathematics with Applications, 59(1):74–79, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006592.

Maini:2012:F

[Mai12] Philip K. Maini. Foreword. Computers and Mathematics with Applications, 64(3):163, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004476.

Mainge:2016:NAM

[Mai16] Paul-Emile Maingé. Numerical approach to monotone variational inequalities by a one-step projected reflected gradient method with line-search procedure. Computers and Mathematics with Applications, 72(3):720– 728, August 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116302966.

Mustafa:2012:CFP

[MAK12] Zead Mustafa, Hassen Aydi, and Erdal Karapınar. On common fixed points in G-metric spaces using (E.A) property. Computers and Mathematics with Applications, 64(6):1944–1956, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002568.

Malek:2015:CAB

[MAN^{+}15] Jihene Malek, Ahmad Taher Azar, Boutheina Nasralli, Mehdi Tekari, Heykel Kamoun, and Rached Tourki. Computational analysis of blood flow in the retinal arteries and veins using fundus image. Computers and Mathematics with Applications, 69(2):101–116, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005859.

Manafian:2018:NSW

[Man18] Jalil Manafian. Novel solitary wave solutions for the (3 + 1)-dimensional extended Jimbo–Miwa equations. Computers and Mathematics with Applications, 76 (5):1246–1260, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303420.

Mahdavi:2010:DMM

[MAPS10] Iraj Mahdavi, Amin Aalaei, Mohammad Mahdi Paydar, and Maghsud Solimanpur. Designing a mathematical model for dynamic cellular manufacturing systems considering production planning and worker assignment. Computers and Mathematics with Applications, 60(4):1014–1025, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002336.

Markin:2011:FPG

[Mar11a]  Jack Markin.  Fixed

points for generalized nonexpansive mappings in R-trees. Computers and Mathematics with Applications, 62 (12):4614–4618, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009096.

Marrero:2011:CMM

[Mar11b] Isabel Marrero. A class of meromorphic multivalent functions with negative coefficients defined by a differential operator. Computers and Mathematics with Applications, 62(8):3027–3036, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006778.

Marinkovic:2012:OCV

[Mar12] Boban Marinković. Optimality conditions in a vector continuous-time optimization problem. Computers and Mathematics with Applications, 63(1):318–324, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010066.

Marrero:2016:NSG

[Mar16] J. Abderramán Marrero. A numerical solver for general bordered tridiagonal matrix equations. Computers and Mathematics with Applications, 72(11):2731–2740, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305375.

Mejia:2011:NIN

[MAS11] Carlos E. Mejía, Carlos D. Acosta, and Katerine I. Saleme. Numerical identification of a nonlinear diffusion coefficient by discrete mollification. Computers and Mathematics with Applications, 62(5):2187–2199, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100558X.

Messaoudi:2018:DSN

[MAST18] Salim A. Messaoudi, Jamal H. Al-Smail, and Ala A. Talahmeh. Decay for solutions of a nonlinear damped wave equation with variableexponent nonlinearities. Computers and Mathematics with Applications, 76(8):1863–1875, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304061.

Matei:2019:MHV

[Mat19]  Andaluzia Matei.  A

mixed hemivariationalvariational problem and applications. Computers and Mathematics with Applications, 77(11): 2989–3000, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118305005.

Mirevski:2010:SFG

[MB10a] S. P. Mirevski and L. Boyadjiev. On some fractional generalizations of the Laguerre polynomials and the Kummer function. Computers and Mathematics with Applications, 59(3):1271–1277, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004106.

Muthukumar:2010:ACS

[MB10b] P. Muthukumar and P. Balasubramaniam. Approximate controllability of second-order damped McKean–Vlasov stochastic evolution equations. Computers and Mathematics with Applications, 60(10):2788–2796, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007182.

Macarie:2011:OCS

[MB11] Vasile Marius Macarie and Daniel Breaz. The order of convexity of some general integral operators. Computers and Mathematics with Applications, 62 (12):4667–4673, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009291.

Monica:2017:ASO

[MB17] Stefania Monica and Federico Bergenti. An analytic study of opinion dynamics in multiagent systems. Computers and Mathematics with Applications, 73(10):2272–2284, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301700.

Maleknejad:2011:HLP

[MBH11] K. Maleknejad, B. Basirat, and E. Hashemizadeh. Hybrid Legendre polynomials and block–pulse functions approach for nonlinear Volterra–Fredholm integro-differential equations. Computers and Mathematics with Applications, 61(9):2821–2828, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002185.

Maier:2016:LGH

[MBH16] Matthias Maier, Mauro Bardelloni, and Luca Heltai. LinearOperator — a generic, highlevel expression syntax for linear algebra. Computers and Mathematics with Applications, 72(1): 1–24, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630205X.

Masouri:2010:EIM

[MBHV10] Z. Masouri, E. Babolian, and S. Hatamzadeh-Varmazyar. An expansioniterative method for numerically solving Volterra integral equation of the first kind. Computers and Mathematics with Applications, 59(4):1491–1499, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007342.

Moreta:2016:AOR

[MBJ16] M. J. Moreta, B. Bujanda, and J. C. Jorge. Avoiding the order reduction when solving secondorder in time PDEs with fractional step Runge– Kutta–Nyström methods. Computers and Mathematics with Applications, 71(7):1425–1447, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300694.

Mohanalin:2010:NAM

[MBKK10] Mohanalin, Beenamol, Prem Kumar Kalra, and Nirmal Kumar. A novel automatic microcalcification detection technique using Tsallis entropy & a Type II fuzzy index. Computers and Mathematics with Applications, 60(8):2426–2432, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110006024.

Mach:2017:NGF

[MBS17] Jan Mach, Michal Benes, and Pavel Strachota. Nonlinear Galerkin finite element method applied to the system of reaction–diffusion equations in one space dimension. Computers and Mathematics with Applications, 73(9):2053–2065, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301116.

Morrison:2013:DPM

[MBT^{+}13] Rebecca E. Morrison, Corey M. Bryant, Gabriel Terejanu, Serge Prudhomme, and Kenji Miki. Data partition methodology for validation of predictive models. Computers and Mathematics with Applications, 66 (10):2114–2125, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005476.

Makinde:2010:MTF

[MC10a] O. D. Makinde and T. Chinyoka. MHD transient flows and heat transfer of dusty fluid in a channel with variable physical properties and Navier slip condition. Computers and Mathematics with Applications, 60(3):660–669, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003639.

Makinde:2010:TAP

[MC10b] O. D. Makinde and T. Chinyoka. Transient analysis of pollutant dispersion in a cylindrical pipe with a nonlinear waste discharge concentration. Computers and Mathematics with Applications, 60(3):642–652, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003615.

Makinde:2011:NSU

[MC11] O. D. Makinde and T. Chinyoka. Numerical study of unsteady hydromagnetic generalized Couette flow of a reactive third-grade fluid with asymmetric convective cooling. Computers and Mathematics with Applications, 61(4):1167–1179, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009715.

Mahdy:2010:DDC

[MCB10] A. Mahdy, A. J. Chamkha, and Yousef Baba. Doublediffusive convection with variable viscosity from a vertical truncated cone in porous media in the presence of magnetic field and radiation effects. Computers and Mathematics with Applications, 59(12):3867– 3878, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002956.

Ma:2018:LLS

[MCF18] Zhengyi Ma, Junchao Chen, and Jinxi Fei. Lump and line soliton pairs to a (2 + 1)-dimensional integrable Kadomtsev–Petviashvili equation. Computers and Mathematics with Applications, 76 (5):1130–1138, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303274.

Chu:2010:TOD

[mCfX10] Yu ming Chu and Wei feng Xia. Two optimal double inequalities between power mean and logarithmic mean. Computers and Mathematics with Applications, 60(1): 83–89, July 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110003032.

Malik:2012:ESB

[MCKM12] Anand Malik, Fakir Chand, Hitender Kumar, and S. C. Mishra. Exact solutions of the Bogoyavlenskii equation using the multiple (\frac{G′}{G})- expansion method. Computers and Mathematics with Applications, 64 (9):2850–2859, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003677.

Maryshev:2013:ASM

[MCL^{+}13] B. Maryshev, A. Cartalade, C. Latrille, M. Joelson, and M. C. Néel. Adjoint state method for fractional diffusion: Parameter identification. Computers and Mathematics with Applications, 66(5):630–638, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001442.

Ma:2015:CSE

[MCL15] Ruyun Ma, Ruipeng Chen, and Yanqiong Lu. Coexistence states of an elliptic system modeling a population with two age groups. Computers and Mathematics with Applications, 69(10):1263– 1271, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001194.

Muslim:2010:ASF

[MCN10] M. Muslim, Carlos Conca, and A. K. Nandakumaran. Approximation of solutions to fractional integral equation. Computers and Mathematics with Applications, 59(3):1236–1244, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003940.

McNamee:2012:CEP

[McN12] J. M. McNamee. Corrigendum to “Efficient polynomial root-refiners: a survey and new record efficiency estimates”, by J. M. McNamee and V. Y. Pan [Comput. Math. Appl. 63 (2012) 239– 254]. Computers and Mathematics with Applications, 63(8):1354, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001101.

See [MP12].

Maione:2013:OZP

[MCP13] G. Maione, R. Caponetto, and A. Pisano. Optimization of zero-pole interlacing for indirect discrete approximations of noninteger order operators. Computers and Mathematics with Applications, 66(5):746–754, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000242.

Machado:2011:EAD

[MCQ11] J. A. Tenreiro Machado, António C. Costa, and Maria Dulce Quelhas. Entropy analysis of the DNA code dynamics in human chromosomes. Computers and Mathematics with Applications, 62(3):1612–1617, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001568.

Makinde:2011:UFR

[MCR11] O. D. Makinde, T. Chinyoka, and L. Rundora. Unsteady flow of a reactive variable viscosity non-Newtonian fluid through a porous saturated medium with asymmetric convective boundary conditions. Computers and Mathematics with Applications, 62 (9):3343–3352, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100722X.

Mustapha:2010:GLE

[MD10] Hussein Mustapha and Roussos Dimitrakopoulos. Generalized Laguerre expansions of multivariate probability densities with moments. Computers and Mathematics with Applications, 60(7):2178–2189, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005717.

Mittal:2015:NSH

[MD15] R. C. Mittal and Sumita Dahiya. Numerical simulation on hyperbolic diffusion equations using modified cubic B-spline differential quadrature methods. Computers and Mathematics with Applications, 70(5):737–749, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002035.

Macias-Diaz:2018:BES

[MD18] J. E. Macías-Díaz. A bounded and efficient scheme for multidimensional problems with anomalous convection and diffusion. Computers and Mathematics with Applications, 75(11):3995–4011, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830138X.

Michel-Dansac:2016:WBS

[MDBCF16] Victor Michel-Dansac, Christophe Berthon, Stéphane Clain, and Françoise Foucher. A well-balanced scheme for the shallowwater equations with topography. Computers and Mathematics with Applications, 72(3):568–593, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302838.

Mattesi:2019:HOA

[MDG19] V. Mattesi, M. Darbas, and C. Geuzaine. A highorder absorbing boundary condition for 2D time-harmonic elastodynamic scattering problems. Computers and Mathematics with Applications, 77(6):1703–1721, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302761.

Mitsotakis:2018:AND

[MDL18] Dimitrios Mitsotakis, Denys Dutykh, and Qian Li. Asymptotic nonlinear and dispersive pulsatile flow in elastic vessels with cylindrical symmetry. Computers and Mathematics with Applications, 75(11):4022–4047, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301408.

Macias-Diaz:2011:NSG

[MDRRV11] J. E. Macías-Díaz, J. Ruiz-Ramírez, and J. Villa. The numerical solution of a generalized Burgers– Huxley equation through a conditionally bounded and symmetry-preserving method. Computers and Mathematics with Applications, 61(11):3330– 3342, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003233.

Macias-Diaz:2017:SPM

[MDVM17] J. E. Macías-Díaz and J. Villa-Morales. A structure-preserving method for the distribution of the first hitting time to a moving boundary for some Gaussian processes. Computers and Mathematics with Applications, 74(8):1799–1812, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303978.

Ma:2011:ESK

[MDW11] Hongcai Ma, Aiping Deng, and Yan Wang. Exact solution of a KdV equation with variable coefficients. Computers and Mathematics with Applications, 61(8):2278–2280, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007455.

Muscato:2013:VRE

[MDW13] Orazio Muscato, Vincenza Di Stefano, and Wolfgang Wagner. A variance-reduced electrothermal Monte Carlo method for semiconductor device simulation.

Computers and Mathematics with Applications, 65(3):520–527, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003185.

Molina-Espinosa:2018:NMP

[MEAMHHV18] L. Molina-Espinosa, C. G.

Aguilar-Madera, E. C. Herrera-Hernández, and C. Verde. Numerical modeling of pseudohomogeneous fluid flow in a pipe with leaks. Computers and Mathematics with Applications, 74(1):64–73, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306733.

Medina:2012:RAR

[Med12] Jesús Medina. Relating attribute reduction in formal, object-oriented and property-oriented concept lattices. Computers and Mathematics with Applications, 64(6):1992–2002, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112002921.

Meng:2018:RSG

[Men18] Xiang-Hua Meng. Rational solutions in Grammian form for the (3 + 1)-dimensional generalized shallow water wave equation. Computers and Mathematics with Applications, 75(12):4534–4539, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301901.

Ma:2011:LSP

[MF11] Wen-Xiu Ma and Engui Fan. Linear superposition principle applying to Hirota bilinear equations. Computers and Mathematics with Applications, 61(4):950–959, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000948X.

Ma:2018:BPS

[MF18a] Lingwei Ma and Zhong Bo Fang. Blow-up phenomena of solutions for a reaction–diffusion equation with weighted exponential nonlinearity. Computers and Mathematics with Applications, 75(8):2735–2745, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300099.

Ma:2018:BBT

[MF18b] Lingwei Ma and Zhong Bo Fang. Bounds for blowup time of a reaction– diffusion equation with weighted gradient nonlinearity. Computers and Mathematics with Applications, 76(3):508–519, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302451.

Mishra:2019:SRB

[MFSL19] Pankaj K. Mishra, Gregory E. Fasshauer, Mrinal K. Sen, and Leevan Ling. A stabilized radial basis-finite difference (RBF–FD) method with hybrid kernels. Computers and Mathematics with Applications, 77(9):2354–2368, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307259.

Mishra:2011:CIS

[MG11] A. K. Mishra and P. Gochhayat. A coefficient inequality for a subclass of the Carathéodory functions defined using conical domains. Computers and Mathematics with Applications, 61(9):2816–2820, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100215X.

Mehrabian:2015:NBA

[MG15] M. Mehrabian and M. E. Golmakani. Nonlinear bending analysis of radial-stiffened annular laminated sector plates with dynamic relaxation method. Computers and Mathematics with Applications, 69(10):1272– 1302, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001261.

Ma:2016:SBD

[MG16] Li Ma and Shangjiang Guo. Stability and bifurcation in a diffusive Lotka–Volterra system with delay. Computers and Mathematics with Applications, 72(1):147– 177, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302425.

Moghimi:2011:HPM

[MGB^{+}11] S. M. Moghimi, D. D.

Ganji, H. Bararnia, M. Hosseini, and M. Jalaal. Homotopy perturbation method for nonlinear MHD Jeffery–Hamel problem. Computers and Mathematics with Applications, 61(8):2213–2216, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007042.

Malvandi:2016:MPC

[MGN^{+}16] Amir Malvandi, Amirmahdi Ghasemi, Rasoul Nikbakhti, Amirreza Ghasemi, and Faraz Hedayati. Modeling and parallel computation of the nonlinear interaction of rigid bodies with incompressible multi-phase flow. Computers and Mathematics with Applications, 72(4):1055–1065, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303509.

Meng:2014:MSD

[MGS^{+}14] Gao-Qing Meng, Yi-Tian Gao, Yu-Hao Sun, Yi Qin, and Xin Yu. Multi-soliton and double Wronskian solutions of a (2 + 1)-dimensional modified Heisenberg ferromagnetic system. Computers and Mathematics with Applications, 66(12):2559–2569, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005154.

Ma:2016:CDP

[MGTH16] Manjun Ma, Meiyan Gao, Changqing Tong, and Yazhou Han. Chemotaxisdriven pattern formation for a reaction–diffusionchemotaxis model with volume-filling effect. Computers and Mathematics with Applications, 72 (5):1320–1340, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303716.

Mozyrska:2011:NCV

[MGW11] Dorota Mozyrska, Ewa Girejko, and Małgorzata Wyrwas. A necessary condition of viability for fractional differential equations with initialization. Computers and Mathematics with Applications, 62(9):3642–3647, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007851.

Merdan:2011:NSM

[MGY11] Mehmet Merdan, Ahmet Gökdogan, and Ahmet Yildirim. On the numerical solution of the model for HIV infection of CD4^{+} T cells. Computers and Mathematics with Applications, 62(1):118– 123, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003610.

Marinca:2011:NDA

[MH11] V. Marinca and N. Herisanu.

Nonlinear dynamic analysis of an electrical machine rotor-bearing system by the optimal homotopy perturbation method. Computers and Mathematics with Applications, 61(8):2019–2024, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006309.

Mozafari:2011:PSA

[MHH11] Niloofar Mozafari, Sattar Hashemi, and Ali Hamzeh. A precise statistical approach for concept change detection in unlabeled data streams. Computers and Mathematics with Applications, 62(4):1655–1669, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004688.

Mardani:2018:MMS

[MHHC18] A. Mardani, M. R. Hooshmandasl, M. H. Heydari, and C. Cattani. A meshless method for solving the time fractional advection–diffusion equation with variable coefficients. Computers and Mathematics with Applications, 75(1):122–133, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305308.

Modarres:2011:SRO

[MHL11] Farzin Modarres, Malik Abu Hassan, and Wah June Leong. A symmetric rank-one method based on extra updating techniques for unconstrained optimization. Computers and Mathematics with Applications, 62(1):392–400, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004202.

Moaddy:2011:NSFb

[MHM11] K. Moaddy, I. Hashim, and S. Momani. Nonstandard finite difference schemes for solving fractional-order Rössler chaotic and hyperchaotic systems. Computers and Mathematics with Applications, 62(3):1068–1074, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111002227.

Moawad:2016:EFC

[MI16] S. M. Moawad and D. A. Ibrahim. Elementary functions for constructing exact solutions to nonlinear partial differential equations with applications to nonlinear Schrödinger type and MHD equations. Computers and Mathematics with Applications, 72 (9):2151–2166, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304679.

Miao:2018:SCP

[Mia18] Shu-Xin Miao. On the semi-convergence of preconditioned GLHSS iteration method for non-Hermitian singular saddle point problem. Computers and Mathematics with Applications, 75(2):419–423, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730593X.

Michaels:2017:EIC

[Mic17] T. Michaels. Equidistributed icosahedral configurations on the sphere. Computers and Mathematics with Applications, 74(4):605–612, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302286.

Milewski:2018:CMF

[Mil18] Sławomir Milewski. Combination of the meshless finite difference approach with the Monte Carlo random walk technique for solution of elliptic problems. Computers and Mathematics with Applications, 76(4):854–876, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302931.

Min:2011:NST

[Min11]  Won Keun Min.  A

note on soft topological spaces. Computers and Mathematics with Applications, 62(9):3524–3528, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007528.

Mishra:2014:NSR

[Mis14] Debasisha Mishra. Nonnegative splittings for rectangular matrices. Computers and Mathematics with Applications, 67(1): 136–144, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300624X.

Masjed-Jamei:2010:MIS

[MJ10] Mohammad Masjed-Jamei. A main inequality for several special functions. Computers and Mathematics with Applications, 60(5):1280–1289, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004219.

Ming:2014:LBS

[MJ14] Cheng Ming and Lou Jing. Lattice Boltzmann simulation of a drop impact on a moving wall with a liquid film. Computers and Mathematics with Applications, 67(2):307–317, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004215.

Ma:2018:SPJ

[MJB18] Ru-Ru Ma, Zhi-Gang Jia, and Zheng-Jian Bai. A structure-preserving Jacobi algorithm for quaternion Hermitian eigenvalue problems. Computers and Mathematics with Applications, 75(3):809–820, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306570.

Mellouli:2019:MIA

[MJWD19] H. Mellouli, H. Jrad, M. Wali, and F. Dammak. Meshless implementation of arbitrary 3Dshell structures based on a modified first order shear deformation theory. Computers and Mathematics with Applications, 77(1):34–49, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118305200.

Motsepa:2017:CLS

[MK17] Tanki Motsepa and Chaudry Masood Khalique. Conservation laws and solutions of a generalized coupled (2 + 1)-dimensional Burgers system. Computers and Mathematics with Applications, 74 (6):1333–1339, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303632.

Marcsa:2018:DCT

[MK18] Dániel Marcsa and Miklós Kuczmann. Design and control for torque ripple reduction of a 3phase switched reluctance motor. Computers and Mathematics with Applications, 74(1):89–95, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730010X.

Misra:2010:ALR

[MKA^{+}10] Sudip Misra, P. Venkata Krishna, Kiran Isaac Abraham, Navin Sasikumar, and S. Fredun. An adaptive learning routing protocol for the prevention of distributed denial of service attacks in wireless mesh networks. Computers and Mathematics with Applications, 60(2):294–306, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000088.

Martines:2013:NRA

[MKG13] Ian P. Martines, Hristo V.

Kojouharov, and James P. Grover. Nutrient recycling and allelopathy in a gradostat. Computers and Mathematics with Applications, 66 (9):1613–1626, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000977.

Ma:2011:HOC

[MKHC11] Yuanping Ma, Linghua Kong, Jialin Hong, and Ying Cao. High-order compact splitting multisymplectic method for the coupled nonlinear Schrödinger equations. Computers and Mathematics with Applications, 61(2):319–333, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008564.

Maleknejad:2011:CCN

[MKL11] K. Maleknejad, A. Khademi, and T. Lotfi. Convergence and condition number of multi-projection operators by Legendre wavelets. Computers and Mathematics with Applications, 62(9):3538–3550, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007541.

Mursaleen:2011:MNM

[MKPS11] M. Mursaleen, V. Karakaya, H. Polat, and N. Simsek. Measure of noncompactness of matrix operators on some difference sequence spaces of weighted means. Computers and Mathematics with Applications, 62(2):814– 820, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004792.

Maleknejad:2012:NMS

[MKR12] K. Maleknejad, M. Khodabin, and M. Rostami. A numerical method for solving m-dimensional stochastic Itô–Volterra integral equations by stochastic operational matrix. Computers and Mathematics with Applications, 63(1):133–143, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009667.

Morais:2013:GHS

[MKS13] J. Morais, K. I. Kou, and W. Sprößig. Generalized holomorphic Szegő kernel in 3D spheroids. Computers and Mathematics with Applications, 65(4):576–588, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006293.

Ma:2019:IBS

[ML19a] Yu-Lan Ma and Bang-Qing Li. Interactions between soliton and rogue wave for a (2 + 1)- dimensional generalized breaking soliton system: Hidden rogue wave and hidden soliton. Computers and Mathematics with Applications, 78(3):827–839, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930121X.

Meng:2019:EAW

[ML19b] Fengjuan Meng and Cuncai Liu. Exponential attractors for weakly damped wave equation with sub-quintic nonlinearity. Computers and Mathematics with Applications, 78(3):1026–1036, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301403.

Modave:2017:PML

[MLG17] Axel Modave, Jonathan Lambrechts, and Christophe Geuzaine. Perfectly matched layers for convex truncated domains with discontinuous Galerkin time domain simulations. Computers and Mathematics with Applications, 73(4):684–700, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300044.

Monnesland:2016:LSF

[MLGY16] Irene Sonja Monnesland, Eunjung Lee, Max Gunzburger, and Ryeongkyung Yoon. A least-squares finite element method for a nonlinear Stokes problem in glaciology. Computers and Mathematics with Applications, 71(11):2421–2431, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005167.

Ma:2016:SDI

[MLL16] Zongwei Ma, Rui Li, and Xing Lin. Solutions to the 2-dimensional isothermal Euler–Poisson equations with a cosmological constant. Computers and Mathematics with Applications, 71(3):730–736, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005957.

Moraga:2015:VFP

[MLSLM15] Nelson O. Moraga, Luis A.

Lemus, Mario A. Saavedra, and Roberto A. Lemus-Mondaca. VOF/ FVM prediction and experimental validation for shear-thinning fluid column collapse. Computers and Mathematics with Applications, 69(2):89– 100, January 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114005860.

Meng:2011:TSF

[MLY11] Jiana Meng, Hongfei Lin, and Yuhai Yu. A twostage feature selection method for text categorization. Computers and Mathematics with Applications, 62(7):2793–2800, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100616X.

Ma:2018:SDI

[MLY18] Zhan-Ping Ma, Jie Liu, and Jia-Long Yue. Spatiotemporal dynamics induced by delay and diffusion in a predatorprey model with mutual interference among the predator. Computers and Mathematics with Applications, 75(10):3488–3507, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300853.

Ming:2016:ASM

[MLZ^{+}16] Chunying Ming, Fawang Liu, Liancun Zheng, Ian Turner, and Vo Anh. Analytical solutions of multiterm time fractional differential equations and application to unsteady flows of generalized viscoelastic fluid. Computers and Mathematics with Applications, 72 (9):2084–2097, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304667.

Ma:2010:PSN

[MM10a] Ruyun Ma and Huili Ma.

Positive solutions for nonlinear discrete periodic boundary value problems. Computers and Mathematics with Applications, 59(1):136–141, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005938.

Mursaleen:2010:CBC

[MM10b] M. Mursaleen and S. A. Mohiuddine. On σconservative and boundedly σ-conservative fourdimensional matrices. Computers and Mathematics with Applications, 59(2):880–885, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006956.

Mishra:2011:MFD

[MM11] S. S. Mishra and P. P. Mishra. A (Q, R) model for fuzzified deterioration under cobweb phenomenon and permissible delay in payment. Computers and Mathematics with Applications, 61(4):921–932, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009454.

Morchalo:2012:BSD

[MM12] Jarosław Morchalo and Małgorzata Migda. Boundedness of solutions of difference systems with delays. Computers and Mathematics with Applications, 64(7):2233–2240, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001071.

Marinov:2013:IPE

[MM13] Tchavdar T. Marinov and Rossitza S. Marinova. An inverse problem for estimation of bending stiffness in Kirchhoff–Love plates. Computers and Mathematics with Applications, 65(3):512–519, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004920.

Manservisi:2016:OCP

[MM16] Sandro Manservisi and Filippo Menghini. Optimal control problems for the Navier–Stokes system coupled with the k–ω turbulence model. Computers and Mathematics with Applications, 71(11):2389–2406, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500485X.

Marinca:2018:SES

[MM18a] Bogdan Marinca and Vasile Marinca. Some exact solutions for MHD flow and heat transfer to modified second grade fluid with variable thermal conductivity in the presence of thermal radiation and heat generation/absorption. Computers and Mathematics with Applications, 76 (6):1515–1524, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303687.

Martelli:2018:NMI

[MM18b] P.-W. Martelli and S. M. Mefire. Numerical modeling and investigations of 3D devices with ferroelectric layer fully embedded in a paraelectric environment. Computers and Mathematics with Applications, 76(3):451–470, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302335.

Mishra:2018:TSI

[MM18c] Nachiketa Mishra and Debasisha Mishra. Twostage iterations based on composite splittings for rectangular linear systems. Computers and Mathematics with Applications, 75(8):2746–2756, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300105.

Moghaddam:2018:STL

[MM18d] B. P. Moghaddam and J. A. T. Machado. A stable three-level explicit spline finite difference scheme for a class of nonlinear time variable order fractional partial differential equations. Computers and Mathematics with Applications, 73(6):1262–1269, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304023.

Moshiri:2019:SMC

[MM19] M. Moshiri and M. T.

Manzari. Simulation of multi-component multiphase fluid flow in twodimensional anisotropic heterogeneous porous media using high-order control volume distributed methods. Computers and Mathematics with Applications, 78 (10):3303–3328, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302597.

Munawar:2012:TDS

[MMA12] Sufian Munawar, Ahmer Mehmood, and Asif Ali. Three-dimensional squeezing flow in a rotating channel of lower stretching porous wall. Computers and Mathematics with Applications, 64(6):1575–1586, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000132.

Mursaleen:2010:ICD

[MME10] M. Mursaleen, S. A. Mohiuddine, and Osama H. H. Edely. On the ideal convergence of double sequences in intuitionistic fuzzy normed spaces. Computers and Mathematics with Applications, 59(2):603–611, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007329.

Mil-Martinez:2019:SAN

[MMFT^{+}19] R. Mil-Martínez, V. H.

Ferrer, M. Turcio, F. López-Serrano, J. A. Ortega, and R. O. Vargas. Stability analysis and numerical simulation of gravitactic bioconvection in a rectangular cavity. Computers and Mathematics with Applications, 77(1):222–236, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305406.

Moaddy:2011:NSFa

[MMH11] K. Moaddy, S. Momani, and I. Hashim. The nonstandard finite difference scheme for linear fractional PDEs in fluid mechanics. Computers and Mathematics with Applications, 61(4):1209–1216, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110009880.

Matin:2017:EFE

[MMHGM17] Rastin Matin, Marek Krzysztof Misztal, Anier Hernández-García, and Joachim Mathiesen. Evaluation of the finite element lattice Boltzmann method for binary fluid flows. Computers and Mathematics with Applications, 74(2):281–291, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302584.

Mata-Machuca:2012:ASC

[MMMG12] Juan L. Mata-Machuca and Rafael Martínez-Guerra. Asymptotic synchronization of the Colpitts oscillator. Computers and Mathematics with Applications, 63(6):1072– 1078, March 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010558.

Marinov:2014:IPC

[MMOJ14] Tchavdar T. Marinov, Rossitza S. Marinova, Joe Omojola, and Michael Jackson. Inverse problem for coefficient identification in SIR epidemic models. Computers and Mathematics with Applications, 67(12):2218– 2227, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000601.

Molaee:2010:MME

[MMR10] Ehsan Molaee, Ghasem Moslehi, and Mohammad Reisi. Minimizing maximum earliness and number of tardy jobs in the single machine scheduling problem. Computers and Mathematics with Applications, 60(11):2909–2919, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007431.

Molaee:2011:MME

[MMR11] Ehsan Molaee, Ghasem Moslehi, and Mohammad Reisi. Minimizing maximum earliness and number of tardy jobs in the single machine scheduling problem with availability constraint. Computers and Mathematics with Applications, 62(9):3622–3641, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100784X.

MacDonald:2012:RAC

[MMRN12] Craig S. MacDonald, John A. Mackenzie, Alison Ramage, and Christopher J. P. Newton. Robust adaptive computation of a one-dimensional Q-tensor model of nematic liquid crystals. Computers and Mathematics with Applications, 64(11):3627–3640, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006049.

Maier:2018:FFS

[MMS^{+}18] Marie-Luise Maier, Stefanie Milles, Sebastian Schuhmann, Gisela Guthausen, Hermann Nirschl, and Mathias J. Krause. Fluid flow simulations verified by measurements to investigate adsorption processes in a static mixer. Computers and Mathematics with Applications, 76(11–12):2744–2757, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830498X.

Marquez:2018:FEA

[MMT18] Antonio Márquez, Salim Meddahi, and Thanh Tran. Frequency-explicit asymptotic error estimates for a stresspressure formulation of a time harmonic fluidsolid interaction problem. Computers and Mathematics with Applications, 76(9):2090–2109, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304243.

Mohapatra:2010:PUN

[MN10a] Jugal Mohapatra and Srinivasan Natesan. Parameteruniform numerical method for global solution and global normalized flux of singularly perturbed boundary value problems using grid equidistribution. Computers and Mathematics with Applications, 60(7):1924–1939, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005183.

Mophou:2010:SCA

[MN10b] G. M. Mophou and G. M. N’Guérékata. On some classes of almost automorphic functions and applications to fractional differential equations. Computers and Mathematics with Applications, 59(3):1310–1317, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003472.

Mursaleen:2010:AHM

[MN10c] M. Mursaleen and Abdullah K. Noman. Applications of the Hausdorff measure of noncompactness in some sequence spaces of weighted means. Computers and Mathematics with Applications, 60(5):1245–1258, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004190.

Maleknejad:2011:ASC

[MN11a] K. Maleknejad and K. Nedaiasl. Application of sinc-collocation method for solving a class of nonlinear Fredholm integral equations. Computers and Mathematics with Applications, 62(8):3292–3303, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007188.

Mophou:2011:OCFb

[MN11b] Gisèle M. Mophou and Gaston M. N’Guérékata. Optimal control of a fractional diffusion equation with state constraints. Computers and Mathematics with Applications, 62(3):1413–1426, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003476.

Mursaleen:2011:GMS

[MN11c] M. Mursaleen and Abdullah K. Noman. On generalized means and some related sequence spaces. Computers and Mathematics with Applications, 61(4):988–999, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009521.

Mohammadi:2012:CAA

[MN12] M. Saleh Mohammadi and Ali Naserasadi. A criticism of the ACW algorithm. Computers and Mathematics with Applications, 64(8):2616–2620, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004889.

Morosanu:2017:ISA

[MN17] Gheorghe Morosanu and Mihai Nechita. Invariant sets and attractors for Hanusse-type chemical systems with diffusions. Computers and Mathematics with Applications, 73(8):1815–1823, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301037.

Mehdinejadiani:2013:MMS

[MNJ^{+}13] Behrouz Mehdinejadiani, Abd Ali Naseri, Hossein Jafari, Afshin Ghanbarzadeh, and Dumitru Baleanu. A mathematical model for simulation of a water table profile between two parallel subsurface drains using fractional derivatives. Computers and Mathematics with Applications, 66(5):785–794, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000199.

Mirbagheri:2015:RSO

[MNPD15] Y. Mirbagheri, H. Nahvi, J. Parvizian, and A. Düster. Reducing spurious oscillations in discontinuous wave propagation simulation using high-order finite elements. Computers and Mathematics with Applications, 70(7):1640–1658, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003089.

Motamed:2015:ACE

[MNT15] Mohammad Motamed, Fabio Nobile, and Raúl Tempone. Analysis and computation of the elastic wave equation with random coefficients. Computers and Mathematics with Applications, 70 (10):2454–2473, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004435.

Milewski:2014:SOA

[MO14] Slawomir Milewski and Janusz Orkisz. In search of optimal acceleration approach to iterative solution methods of simultaneous algebraic equations. Computers and Mathematics with Applications, 68(3):101–117, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400203X.

Mohsen:2014:SSB

[Moh14] A. Mohsen. A simple solution of the Bratu problem. Computers and Mathematics with Applications, 67(1):26–33, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300610X.

Mohammadi:2015:QBS

[Moh15] Reza Mohammadi. Quintic B-spline collocation approach for solving generalized Black–Scholes equation governing option pricing. Computers and Mathematics with Applications, 69(8):777– 797, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000851.

Mokhtarian:2011:NFW

[Mok11] M. N. Mokhtarian. A new fuzzy weighted average (FWA) method based on left and right scores: an application for determining a suitable location for a gas oil station. Computers and Mathematics with Applications, 61(10):3136– 3145, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002847.

Momoniat:2011:DSF

[Mom11] E. Momoniat. On the determination of the steady film profile for a non-Newtonian thin droplet. Computers and Mathematics with Applications, 62(1):383–391, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004093.

Moore:2018:DGI

[Moo18] Stephen Edward Moore. Discontinuous Galerkin isogeometric analysis for the biharmonic equation. Computers and Mathematics with Applications, 76(4):673–685, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830261X.

Mophou:2011:OCFa

[Mop11] Gisèle. M. Mophou. Optimal control of fractional diffusion equation. Computers and Mathematics with Applications, 61(1):68–78, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008205.

Mortici:2010:ASG

[Mor10a] Cristinel Mortici. The asymptotic series of the generalized Stirling formula. Computers and Mathematics with Applications, 60(3):786–791, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003822.

Mortici:2010:CIA

[Mor10b]  Cristinel Mortici.  A

class of integral approximations for the factorial function. Computers and Mathematics with Applications, 59(6):2053–2058, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007573.

Mortici:2010:NSC

[Mor10c] Cristinel Mortici. On new sequences converging towards the Euler– Mascheroni constant. Computers and Mathematics with Applications, 59(8): 2610–2614, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000519.

Mortici:2010:SEM

[Mor10d] Cristinel Mortici. On some Euler–Mascheroni type sequences. Computers and Mathematics with Applications, 60(7):2009–2014, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005304.

Mortici:2011:IAF

[Mor11a] Cristinel Mortici. Improved asymptotic formulas for the gamma function. Computers and Mathematics with Applications, 61(11):3364– 3369, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003373.

Mortici:2011:RGF

[Mor11b] Cristinel Mortici. Refinements of Gurland’s formula for pi. Computers and Mathematics with Applications, 62 (6):2616–2620, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122111006481.

Moreno:2013:IFO

[Mor13] J. Moreno. An infinite family of one-step iterators for solving nonlinear equations to increase the order of convergence and a new algorithm of global convergence. Computers and Mathematics with Applications, 66(8):1418–1436, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004847.

Mostafa:2010:SCR

[Mos10] A. O. Mostafa. Starlikeness and convexity results for hypergeometric functions. Computers and Mathematics with Applications, 59(8):2821–2826, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000088X.

Ma:2010:SQP

[MP10a] Qing-Hua Ma and Josip Pecarić. On some qualitative properties for solutions of a certain two-dimensional fractional differential systems. Computers and Mathematics with Applications, 59(3):1294–1299, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004246.

Mainar:2010:OBC

[MP10b] E. Mainar and J. M. Peña. Optimal bases for a class of mixed spaces and their associated spline spaces. Computers and Mathematics with Applications, 59(4):1509–1523, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007391.

Ma:2011:BSC

[MP11a] Qing-Hua Ma and Josip Pecarić. The bounds on the solutions of certain two-dimensional delay dynamic systems on time scales. Computers and Mathematics with Applications, 61(8):2158–2163, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110006863.

Makri:2011:SRF

[MP11b] Frosso S. Makri and Zaharias M. Psillakis. On success runs of a fixed length in Bernoulli sequences: Exact and asymptotic results. Computers and Mathematics with Applications, 61(4):761–772, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009284.

See corrigendum [MP16].

Mishura:2011:SDE

[MP11c] Yu. S. Mishura and S. V.

Posashkova. Stochastic differential equations driven by a Wiener process and fractional Brownian motion: Convergence in Besov space with respect to a parameter. Computers and Mathematics with Applications, 62(3):1166–1180, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001180.

McNamee:2012:EPR

[MP12] J. M. McNamee and Victor Y. Pan. Efficient polynomial rootrefiners: a survey and new record efficiency estimates. Computers and Mathematics with Applications, 63(1):239–254, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009813.

See corrigendum [McN12].

Makri:2016:CSR

[MP16] Frosso S. Makri and Zaharias M. Psillakis. Corrigendum to “On success runs of a fixed length in Bernoulli sequences: Exact and asymptotic results” [Comput. Math. Appl. 61 (2011) 761– 772]. Computers and Mathematics with Applications, 72(3):806, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303388.

See [MP11b].

Mertz:2019:NAD

[MP19a] Laurent Mertz and Olivier Pironneau. Numerical analysis of degenerate Kolmogorov equations of constrained stochastic Hamiltonian systems. Computers and Mathematics with Applications, 78(8):2719–2733, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302111.

Milewski:2019:HOM

[MP19b] Sławomir Milewski and Roman Putanowicz. Higher order meshless schemes applied to the finite element method in elliptic problems. Computers and Mathematics with Applications, 77(3):779–802, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306138.

Mitra:2019:MSS

[MP19c] K. Mitra and I. S. Pop.

A modified L-scheme to solve nonlinear diffusion problems. Computers and Mathematics with Applications, 77(6):1722–1738, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305546.

Motreanu:2019:DCS

[MP19d] Dumitru Motreanu and Zijia Peng. Doubly coupled systems of elliptic hemivariational inequalities: Existence and location. Computers and Mathematics with Applications, 77(11):3001– 3009, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304620.

Meng:2018:ASD

[MPfTX18] Gao-Qing Meng, Yu-Song Pan, Hao feng Tan, and Xi-Yang Xie. Analytic solutions for the (2 + 1)-dimensional generalized sine-Gordon equations in nonlinear optics. Computers and Mathematics with Applications, 76(6):1535–1543, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303705.

Min:2019:IDA

[MPGW19] Haoda Min, Cheng Peng, Zhaoli Guo, and Lian-Ping Wang. An inverse design analysis of mesoscopic implementation of non-uniform forcing in MRT lattice Boltzmann models. Computers and Mathematics with Applications, 78(4):1095–1114, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302334.

Miyagaki:2018:MPS

[MPLR18] O. H. Miyagaki, L. C.

Paes-Leme, and B. M. Rodrigues. Multiplicity of positive solutions for the Kirchhoff-type equations with critical exponent in RN. Computers and Mathematics with Applications, 75(9):3201–3212, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300579.

Muga:2015:SAR

[MPMTV15] Ignacio Muga, David Pardo, Paweł J. Matuszyk, and Carlos Torres-Verdín. Semi-analytical response of acoustic logging measurements in frequency domain. Computers and Mathematics with Applications, 70(4):314– 329, August 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001455.

Micheletti:2018:AMA

[MPS18] Stefano Micheletti, Simona Perotto, and Marianna Signorini. Anisotropic mesh adaptation for the generalized Ambrosio– Tortorelli functional with application to brittle fracture. Computers and Mathematics with Applications, 75(6):2134–2152, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730487X.

McLaughlin:2016:SRO

[MPY16] Benjamin McLaughlin, Janet Peterson, and Ming Ye. Stabilized reduced order models for the advection-diffusion-reaction equation using operator splitting. Computers and Mathematics with Applications, 71(11):2407–2420, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300281.

Ma:2011:IIS

[MPZ11] Qing-Hua Ma, Josip Pecarić, and Jian-Mei Zhang. Integral inequalities of systems and the estimate for solutions of certain nonlinear two-dimensional fractional differential systems. Computers and Mathematics with Applications, 61(11):3258– 3267, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002951.

Matucci:2010:NHL

[MR10] S. Matucci and P. Rehák.

Nonoscillation of halflinear dynamic equations. Computers and Mathematics with Applications, 60(5):1421–1429, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004384.

Martins:2014:MMN

[MR14] Nuno F. M. Martins and Magda Rebelo. Meshfree methods for nonhomogeneous Brinkman flows. Computers and Mathematics with Applications, 68(8):872–886, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003629.

Motamed:2015:WBG

[MR15] Mohammad Motamed and Olof Runborg. A wavefront-based Gaussian beam method for computing high frequency wave propagation problems. Computers and Mathematics with Applications, 69(9):949–963, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000887.

Mabrouk:2017:ADB

[MR17] S. M. Mabrouk and A. S. Rashed. Analysis of (3 + 1)-dimensional Boiti-Leon-manna-Pempinelli equation via Lax pair investigation and group transformation method. Computers and Mathematics with Applications, 74(10):2546–2556, November 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117304625.

McGregor:2019:IAD

[MR19] D. A. McGregor and A. C. Robinson. An indirect ALE discretization of single fluid plasma without a fast magnetosonic time step restriction. Computers and Mathematics with Applications, 78(2):417–436, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306072.

Milojevic:2011:JFP

[MRR11] Marija Milojević, Stojan Radenović, and Bozidar Rosić. Jensen’s functional and polynomials in several variables. Computers and Mathematics with Applications, 61(11):3322– 3329, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003324.

Mora:2018:PEE

[MRR18] David Mora, Gonzalo Rivera, and Rodolfo Rodríguez. A posteriori error estimates for a Virtual Element Method for the Steklov eigenvalue problem. Computers and Mathematics with Applications, 74 (9):2172–2190, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730305X.

Moaddy:2012:FOM

[MRS^{+}12] K. Moaddy, A. G. Radwan, K. N. Salama, S. Momani, and I. Hashim. The fractional-order modeling and synchronization of electrically coupled neuron systems. Computers and Mathematics with Applications, 64 (10):3329–3339, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000156.

Manni:2015:ICM

[MRS15] Carla Manni, Alessandro Reali, and Hendrik Speleers. Isogeometric collocation methods with generalized Bsplines. Computers and Mathematics with Applications, 70(7):1659–1675, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001443.

Majumdar:2010:GFS

[MS10a] Pinaki Majumdar and S. K. Samanta. Generalised fuzzy soft sets. Computers and Mathematics with Applications, 59(4):1425–1432, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007536.

Majumdar:2010:SM

[MS10b] Pinaki Majumdar and S. K. Samanta. On soft mappings. Computers and Mathematics with Applications, 60 (9):2666–2672, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006899.

Muthuraj:2010:MCH

[MS10c] R. Muthuraj and S. Srinivas. Mixed convective heat and mass transfer in a vertical wavy channel with traveling thermal waves and porous medium. Computers and Mathematics with Applications, 59(11):3516– 3528, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002348.

Malik:2011:DFC

[MS11a] Pradeep Malik and A. Swaminathan. Derivatives of a finite class of orthogonal polynomials defined on the positive real line related to F -distribution. Computers and Mathematics with Applications, 61(4):1180–1189, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009727.

Mondal:2011:PCT

[MS11b] Saiful R. Mondal and A. Swaminathan. On the positivity of certain trigonometric sums and their applications. Computers and Mathematics with Applications, 62 (10):3871–3883, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008054.

Mishura:2012:MSD

[MS12a] Yuliya Mishura and Georgiy Shevchenko. Mixed stochastic differential equations with long-range dependence: Existence, uniqueness and convergence of solutions. Computers and Mathematics with Applications, 64 (10):3217–3227, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002660.

Motsa:2012:LMN

[MS12b] S. S. Motsa and P. Sibanda.

A linearisation method for non-linear singular boundary value problems. Computers and Mathematics with Applications, 63(7):1197–1203, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211101090X.

MBarka:2013:CVB

[MS13] Moez Ben MBarka and Julien P. Stern. Certification validation: Back to the past. Computers and Mathematics with Applications, 65(5):799–811, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000187.

Mohanty:2015:NHA

[MS15] R. K. Mohanty and Nikita Setia. A new high accuracy two-level implicit off-step discretization for the system of three space dimensional quasi-linear parabolic partial differential equations. Computers and Mathematics with Applications, 69(10):1096– 1113, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000991.

Maulik:2017:NDF

[MS17] Romit Maulik and Omer San. A novel dynamic framework for subgrid scale parametrization of mesoscale eddies in quasigeostrophic turbulent flows. Computers and Mathematics with Applications, 74(3):420–445, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730247X.

Ma:2018:UOA

[MS18] Ying Ma and Chunmei Su. A uniformly and optimally accurate multiscale time integrator method for the Klein– Gordon–Zakharov system in the subsonic limit regime. Computers and Mathematics with Applications, 76(3):602–619, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302542.

Mathiyalagan:2012:NRE

[MSA12] K. Mathiyalagan, R. Sakthivel, and S. Marshal Anthoni. New robust exponential stability results for discrete-time switched fuzzy neural networks with time delays. Computers and Mathematics with Applications, 64(9):2926–2938, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005433.

Moura:2018:CPD

[MSFS18] A. S. Moura, E. J. Silva, W. G. Facco, and R. R. Saldanha. CFS– PML–DEC formulation in two-dimensional convex and non-convex domains. Computers and Mathematics with Applications, 76(1):172–178, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302104.

Molabahrami:2011:AMS

[MSG11] A. Molabahrami, A. Shidfar, and A. Ghyasi. An analytical method for solving linear Fredholm fuzzy integral equations of the second kind. Computers and Mathematics with Applications, 61(9):2754–2761, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001866.

Mahamad:2010:PRU

[MSH10] Abd Kadir Mahamad, Sharifah Saon, and Takashi Hiyama. Predicting remaining useful life of rotating machinery based artificial neural network. Computers and Mathematics with Applications, 60(4):1078–1087, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002555.

Ma:2011:NEN

[MSQ^{+}11] Xiuqin Ma, Norrozila Sulaiman, Hongwu Qin, Tutut Herawan, and Jasni Mohamad Zain. A new efficient normal parameter reduction algorithm of soft sets. Computers and Mathematics with Applications, 62(2):588–598, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004366.

Mbarki:2017:NRA

[MSTB17] Zouhair Mbarki, Hassene Seddik, Sondes Tebini, and Ezzedine Ben Braiek. A new rapid auto-adapting diffusion function for adaptive anisotropic image denoising and sharply conserved edges. Computers and Mathematics with Applications, 74(8):1751–1768, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303747.

Mahsud:2018:NCF

[MSV18] Yasir Mahsud, Nehad Ali Shah, and Dumitru Vieru. Natural convection flows and heat transfer with exponential memory of a Maxwell fluid with damped shear stress. Computers and Mathematics with Applications, 76(9):2246–2261, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304449.

Morvant:2018:CEA

[MSW18] Angelique Morvant, Ethan Seal, and Shawn W. Walker. A coupled Ericksen/Allen–Cahn model for liquid crystal droplets. Computers and Mathematics with Applications, 75(11):4048–4065, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301421.

Mossaiby:2017:OIH

[MSZG17] F. Mossaiby, A. Shojaei, M. Zaccariotto, and U. Galvanetto. OpenCL implementation of a high performance 3D peridynamic model on graphics accelerators. Computers and Mathematics with Applications, 74(8):1856–1870, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304030.

Malinowska:2010:GNB

[MT10] Agnieszka B. Malinowska and Delfim F. M. Torres. Generalized natural boundary conditions for fractional variational problems in terms of the Caputo derivative. Computers and Mathematics with Applications, 59(9):3110–3116, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001513.

Martins:2011:GVT

[MT11] Natália Martins and Delfim F. M. Torres. Generalizing the variational theory on time scales to include the delta indefinite integral. Computers and Mathematics with Applications, 61(9):2424– 2435, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001088.

Martins:2012:HOI

[MT12] Natália Martins and Delfim F. M. Torres. Higher-order infinite horizon variational problems in discrete quantum calculus. Computers and Mathematics with Applications, 64(7):2166–2175, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010492.

Mukam:2018:NER

[MT18] Jean Daniel Mukam and Antoine Tambue. A note on exponential Rosenbrock–Euler method for the finite element discretization of a semilinear parabolic partial differential equation.

Computers and Mathematics with Applications, 76(7):1719–1738, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303900.

Manson:2019:QPG

[MT19a] Nicole Manson and Johannes Tausch. Quadrature for parabolic Galerkin BEM with moving surfaces. Computers and Mathematics with Applications, 77(1):1–14, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830508X.

Mukam:2019:OSC

[MT19b] Jean Daniel Mukam and Antoine Tambue. Optimal strong convergence rates of numerical methods for semilinear parabolic SPDE driven by Gaussian noise and Poisson random measure. Computers and Mathematics with Applications, 77(10):2786–2803, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211930032X.

Messaoudi:2017:NDW

[MTAS17] Salim A. Messaoudi, Ala A. Talahmeh, and Jamal H. Al-Smail. Nonlinear damped wave equation: Existence and blowup. Computers and Mathematics with Applications, 74(12):3024–3041, December 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304777.

Maleknejad:2011:FPM

[MTM11] K. Maleknejad, P. Torabi, and R. Mollapourasl. Fixed point method for solving nonlinear quadratic Volterra integral equations. Computers and Mathematics with Applications, 62(6):2555–2566, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006304.

Marchuk:2019:NUM

[MTN19] Mykhaylo Marchuk, Roman Tuchapskyy, and Dmytro Nespliak. Numerical use of {m, n}- approximation method thermoelastic anisotropic thin shell theory equations represented in a special form. Computers and Mathematics with Applications, 77(10):2740–2763, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300082.

Matuszyk:2013:BPM

[MTV13] Pawel Jerzy Matuszyk and Carlos Torres-Verdín. Back-propagating modes in elastic logging-while-drilling collars and their effect on PML stability. Computers and Mathematics with Applications, 66(11):2335–2343, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005634.

Mehl:2016:PCN

[MUB^{+}16] Miriam Mehl, Benjamin Uekermann, Hester Bijl, David Blom, Bernhard Gatzhammer, and Alexander van Zuijlen. Parallel coupling numerics for partitioned fluidstructure interaction simulations. Computers and Mathematics with Applications, 71(4):869–891, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005933.

Murio:2008:IFD

[Mur08] Diego A. Murio. Implicit finite difference approximation for time fractional diffusion equations. Computers and Mathematics with Applications, 56(4):1138–1145, August 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122108001156.

See [fDxZ11].

Martin-Vaquero:2010:ORI

[MV10] J. Martín-Vaquero. A 17th-order Radau IIA method for package RADAU. applications in mechanical systems. Computers and Mathematics with Applications, 59(8):2464– 2472, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007925.

Mihailescu:2011:MRS

[MV11] Mihai Mihailescu and Csaba Varga. Multiplicity results for some elliptic problems with nonlinear boundary conditions involving variable exponents. Computers and Mathematics with Applications, 62(9):3464–3471, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007462.

Marin:2012:APF

[MV12] Andrea Marin and Maria Grazia Vigliotti. Algorithmic product-form approximations of interacting stochastic models. Computers and Mathematics with Applications, 64 (12):3852–3868, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002398.

Mirza:2017:FSA

[MV17] Itrat Abbas Mirza and Dumitru Vieru. Fundamental solutions to advection-diffusion equation with time-fractional Caputo–Fabrizio derivative. Computers and Mathematics with Applications, 73(1):1–10, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305387.

Muszynski:2012:CAE

[MVB^{+}12] Jakub Muszyński, Sébastien Varrette, Pascal Bouvry, Franciszek Seredyński, and Samee U. Khan. Convergence analysis of evolutionary algorithms in the presence of crashfaults and cheaters. Computers and Mathematics with Applications, 64 (12):3805–3819, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200209X.

Martin-Vaquero:2014:SER

[MVKK14] J. Martín-Vaquero, A. Q. M. Khaliq, and B. Kleefeld. Stabilized explicit Runge– Kutta methods for multiasset American options. Computers and Mathematics with Applications, 67(6):1293–1308, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400042X.

Milovanovic:2018:RBF

[MvS18] Slobodan Milovanović and Lina von Sydow. Radial basis function generated finite differences for option pricing problems. Computers and Mathematics with Applications, 75(4):1462–1481, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307290.

Mrozek:2010:CHA

[MW10] Marian Mrozek and Thomas Wanner. Coreduction homology algorithm for inclusions and persistent homology. Computers and Mathematics with Applications, 60 (10):2812–2833, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007224.

Mane:2011:RCB

[MW11] S. A. Mane and B. N.

Waphare. Regular connected bipancyclic spanning subgraphs of hypercubes. Computers and Mathematics with Applications, 62(9):3551–3554, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007553.

Mickens:2013:NDI

[MW13] Ronald E. Mickens and Talitha M. Washington. NSFD discretizations of interacting population models satisfying conservation laws. Computers and Mathematics with Applications, 66 (11):2307–2316, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003866.

Melenk:2014:SBT

[MW14] J. M. Melenk and T. Wurzer. On the stability of the boundary trace of the polynomial L^{2}-projection on triangles and tetrahedra. Computers and Mathematics with Applications, 67(4):944–965, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122113007128.

Min:2016:QAD

[MW16] Na Min and Mingxin Wang. Qualitative analysis for a diffusive predator–prey model with a transmissible disease in the prey population. Computers and Mathematics with Applications, 72(6):1670–1689, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304333.

Ma:2017:BPS

[MW17] Zhan-Ping Ma and Yu-Xia Wang. Bifurcation of positive solutions for a three-species food chain model with diffusion. Computers and Mathematics with Applications, 74(12):3271–3282, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305072.

Ma:2018:TDA

[MWL18] Qiaozhen Ma, Jing Wang, and Tingting Liu. Timedependent asymptotic behavior of the solution for wave equations with linear memory. Computers and Mathematics with Applications, 76 (6):1372–1387, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303559.

Meng:2011:GBS

[MWWL11] Jun Meng, LiXia Wu, XiuKun Wang, and TsauYoung Lin. Granulationbased symbolic representation of time series and semi-supervised classification. Computers and Mathematics with Applications, 62(9):3581–3590, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007619.

Mu:2017:EIW

[MWY17] Lin Mu, Junping Wang, and Xiu Ye. Effective implementation of the weak Galerkin finite element methods for the biharmonic equation. Computers and Mathematics with Applications, 74 (6):1215–1222, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303504.

Ma:2010:EPS

[MX10] Ruyun Ma and Youji Xu.

Existence of positive solution for nonlinear fourthorder difference equations. Computers and Mathematics with Applications, 59(12):3770– 3777, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002828.

Ma:2015:RAE

[MX15] Qiaozhen Ma and Ling Xu. Random attractors for the extensible suspension bridge equation with white noise. Computers and Mathematics with Applications, 70(12):2895–2903, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004812.

Mo:2010:BAB

[MY10] Jing Mo and Zuodong Yang. Boundary asymptotic behavior and uniqueness of large solutions to quasilinear elliptic equations. Computers and Mathematics with Applications, 59(6):2007–2017, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007500.

Moroney:2013:BPT

[MY13] Timothy Moroney and Qianqian Yang. A banded preconditioner for the two-sided, nonlinear space-fractional diffusion equation. Computers and Mathematics with Applications, 66(5):659–667, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000783.

Ma:2016:CEC

[MY16] Zhan-Ping Ma and Jia-Long Yue. Competitive exclusion and coexistence of a delayed reaction– diffusion system modeling two predators competing for one prey. Computers and Mathematics with Applications, 71(9):1799– 1817, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300797.

Ma:2012:CIQ

[MYZ12] Xueling Ma, Yunqiang Yin, and Jianming Zhan. Characterizations of hintra- and h-quasi-hemiregular hemirings. Computers and Mathematics with Applications, 63(4):783– 793, February 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111010273.

Ma:2018:DIS

[MYZ18] Wen-Xiu Ma, Xuelin Yong, and Hai-Qiang Zhang. Diversity of interaction solutions to the (2 + 1)-dimensional Ito equation. Computers and Mathematics with Applications, 75(1):289–295, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305618.

Ma:2010:CND

[MZ10] Wen-Xiu Ma and Zuo-Nong Zhu. Constructing nonlinear discrete integrable Hamiltonian couplings. Computers and Mathematics with Applications, 60(9):2601–2608, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006607.

Mert:2011:DSC

[MZ11] R. Mert and A. Zafer. On disconjugacy and stability criteria for discrete Hamiltonian systems. Computers and Mathematics with Applications, 62(8):3015–3026, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006766.

Meerschaert:2010:PTF

[MZB10] Mark M. Meerschaert, Yong Zhang, and Boris Baeumer. Particle tracking for fractional diffusion with two time scales. Computers and Mathematics with Applications, 59(3):1078–1086, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003460.

Ma:2017:HLT

[MZC17] Jingtang Ma, Zhiqiang Zhou, and Zhenyu Cui. Hybrid Laplace transform and finite difference methods for pricing American options under complex models. Computers and Mathematics with Applications, 74(3):369–384, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302493.

Matinfar:2012:GIR

[MZES12] M. Matinfar, H. Zareamoghaddam, M. Eslami, and M. Saeidy. GMRES implementations and residual smoothing techniques for solving ill-posed linear systems. Computers and Mathematics with Applications, 63(1):1–13, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007905.

Ma:2011:SKF

[MZJ11] Xueling Ma, Jianming Zhan, and Young Bae Jun. Some kinds of (ϵ_{γ}, ϵ_{γ} ∨ q_{δ})-fuzzy ideals of BCI-algebras. Computers and Mathematics with Applications, 61(4):1005–1015, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009545.

Ma:2013:ASA

[MZL13] Yutian Ma, Fengrong Zhang, and Changpin Li. The asymptotics of the solutions to the anomalous diffusion equations. Computers and Mathematics with Applications, 66(5):682–692, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300062X.

Ma:2010:FIT

[MZLF10] Xueling Ma, Jianming Zhan, and V. Leoreanu-Fotea. On (fuzzy) isomorphism theorems of Γ-hyperrings. Computers and Mathematics with Applications, 60 (9):2594–2600, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110006590.

Manukure:2018:LSD

[MZM18] Solomon Manukure, Yuan Zhou, and Wen-Xiu Ma. Lump solutions to a (2 + 1)-dimensional extended KP equation. Computers and Mathematics with Applications, 75(7):2414–2419, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307939.

Mahmoodi:2019:LTF

[MZM19] R. Mahmoodi, A. Zolfaghari, and A. Minuchehr. Laplace transform finite volume modeling of water hammer along fluid-structure interaction. Computers and Mathematics with Applications, 77(10): 2821–2832, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122119300379.

Meng:2011:SRF

[MZQ11] Dan Meng, Xiaohong Zhang, and Keyun Qin. Soft rough fuzzy sets and soft fuzzy rough sets. Computers and Mathematics with Applications, 62(12):4635–4645, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009138.

Noor:2011:SAR

[NA11] Khalida Inayat Noor and Muhammad Arif. On some applications of Ruscheweyh derivative. Computers and Mathematics with Applications, 62(12):4726–4732, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009394.

Nabti:2019:LSB

[Nab19] Abderrazak Nabti. Life span of blowing-up solutions to the Cauchy problem for a time-space fractional diffusion equation. Computers and Mathematics with Applications, 78(5):1302–1316, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830631X.

Nawaz:2011:VIM

[Naw11]  Yasir Nawaz.  Varia-

tional iteration method and homotopy perturbation method for fourthorder fractional integrodifferential equations. Computers and Mathematics with Applications, 61(8):2330–2341, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007832.

Nazarov:2013:CRB

[Naz13] Murtazo Nazarov. Convergence of a residual based artificial viscosity finite element method. Computers and Mathematics with Applications, 65(4):616–626, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006499.

Nerantzaki:2011:AIA

[NB11] M. S. Nerantzaki and N. G. Babouskos. Analysis of inhomogeneous anisotropic viscoelastic bodies described by multiparameter fractional differential constitutive models. Computers and Mathematics with Applications, 62(3):945–960, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003919.

Noormohammadi:2017:CES

[NB17] N. Noormohammadi and B. Boroomand. Construction of equilibrated singular basis functions without a priori knowledge of analytical singularity order. Computers and Mathematics with Applications, 73(7):1611–1626, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300706.

Niemi:2013:ASD

[NCC13] Antti H. Niemi, Nathaniel O. Collier, and Victor M. Calo. Automatically stable discontinuous Petrov– Galerkin methods for stationary transport problems: Quasi-optimal test space norm. Computers and Mathematics with Applications, 66 (10):2096–2113, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004446.

Ni:2013:EIB

[NCL13] Liang Ni, Gongliang Chen, and Jianhua Li. Escrowable identity-based authenticated key agreement protocol with strong security. Computers and Mathematics with Applications, 65(9):1339–1349, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200051X.

Neto:2018:SBT

[NCV^{+}18] Joana P. Neto, Rui Moura Coelho, Duarte Valério, Susana Vinga, Dominik Sierociuk, Wiktor Malesza, Michal Macias, and Andrzej Dzieliński. Simplifying biochemical tumorous bone remodeling models through variable order derivatives. Computers and Mathematics with Applications, 75(9):3147–3157, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300531.

Nigro:2019:LDD

[NDC^{+}19] A. Nigro, C. De Bartolo, A. Crivellini, M. Franciolini, A. Colombo, and F. Bassi. A lowdissipation DG method for the under-resolved simulation of low Mach number turbulent flows. Computers and Mathematics with Applications, 77(6):1739–1755, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305698.

Nhan:2011:WNI

[NDT11] Nguyen Du Vi Nhan, Dinh Thanh Duc, and Vu Kim Tuan. Weighted norm inequalities for a nonlinear transform. Computers and Mathematics with Applications, 61(4):832–839, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009363.

Neumann:2014:HMC

[NEB14] Philipp Neumann, Wolfgang Eckhardt, and Hans-Joachim Bungartz. Hybrid molecular-continuum methods: From prototypes to coupling software. Computers and Mathematics with Applications, 67(2):272–281, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004252.

Nedzhibov:2012:DFI

[Ned12] Gyurhan H. Nedzhibov.

A derivative-free iterative method for simultaneously computing an arbitrary number of zeros of nonlinear equations. Computers and Mathematics with Applications, 63(7):1185–1191, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010923.

Nesliturk:2010:CSS

[Nes10]  A. I. Nesliturk.  On

the choice of stabilizing sub-grid for convection– diffusion problem on rectangular grids. Computers and Mathematics with Applications, 59(12): 3687–3699, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002622.

Noghrehabadi:2012:NAE

[NGG12] Aminreza Noghrehabadi, Mohammad Ghalambaz, and Afshin Ghanbarzadeh. A new approach to the electrostatic pull-in instability of nanocantilever actuators using the ADMpadé technique. Computers and Mathematics with Applications, 64 (9):2806–2815, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003537.

Naadimuthu:2010:FCC

[NGL10] G. Naadimuthu, P. Gultom, and E. S. Lee. Fuzzy clustering in cell formation with multiple attributes. Computers and Mathematics with Applications, 59(9):3137–3147, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001574.

Ngondiep:2018:LTS

[Ngo18] Eric Ngondiep. Long time stability and convergence rate of Mac-Cormack rapid solver method for nonstationary Stokes–Darcy problem. Computers and Mathematics with Applications, 75(10):3663–3684, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301056.

Netz:2015:MFE

[NH15] Torben Netz and Stefan Hartmann. A monolithic finite element approach using high-order schemes in time and space applied to finite strain thermoviscoelasticity. Computers and Mathematics with Applications, 70(7):1457–1480, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001571.

Netz:2013:HOQ

[NHH13] Torben Netz, Ahmad-Wahadj Hamkar, and Stefan Hartmann. Highorder quasi-static finite element computations in space and time with application to finite strain viscoelasticity. Computers and Mathematics with Applications, 66(4):441–459, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003349.

Nagata:2016:ECC

[NHIN16] Munehiro Nagata, Masatsugu Hada, Masashi Iwasaki, and Yoshimasa Nakamura. Eigenvalue clustering of coefficient matrices in the iterative stride reductions for linear systems. Computers and Mathematics with Applications, 71(1):349– 355, January 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005659.

Niezgoda:2010:NBM

[Nie10] Marek Niezgoda. New bounds for moments of continuous random variables. Computers and Mathematics with Applications, 60 (12):3130–3138, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122110007972.

Nilsrakoo:2011:HTI

[Nil11]  Weerayuth  Nilsrakoo.

Halpern-type iterations for strongly relatively nonexpansive mappings in Banach spaces. Computers and Mathematics with Applications, 62 (12):4656–4666, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009151.

Nandy:2016:AFE

[NJ16] Arup Nandy and C. S.

Jog. An amplitude finite element formulation for electromagnetic radiation and scattering. Computers and Mathematics with Applications, 71(7):1364–1391, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300554.

Nikolova:2013:DFQ

[NJV13] Elena Nikolova, Ivan Jordanov, and Nikolay K. Vitanov. Dynamical features of the quasistationary microRNAmediated protein translation process supported by eIF4F translation initiation factors. Computers and Mathematics with Applications, 66 (9):1716–1725, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002289.

Nikazad:2018:NFS

[NKA18] Touraj Nikazad, Mehdi Karimpour, and Mokhtar Abbasi. Notes on flexible sequential block iterative methods. Computers and Mathematics with Applications, 76 (6):1321–1332, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303481.

Nath:2015:CFB

[NKM15] D. Nath, M. S. Kalra, and P. Munshi. Computation of fixed boundary tokamak equilibria using a method based on approximate particular solutions. Computers and Mathematics with Applications, 70(5):1220–1233, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500348X.

Nath:2016:AMM

[NKM16] D. Nath, M. S. Kalra, and P. Munshi. Application of MFS–MPS to the current-hole simulation in a tokamak. Computers and Mathematics with Applications, 71(8):1706–1721, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301195.

Nadeem:2019:MLV

[NLA19]  Muhammad  Nadeem,

Fengquan Li, and Hijaz Ahmad. Modified Laplace variational iteration method for solving fourth-order parabolic partial differential equation with variable coefficients. Computers and Mathematics with Applications, 78 (6):2052–2062, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122119301841.

Nejad:2011:RFN

[NM11a] Ali Mahmodi Nejad and Mashaallah Mashinchi. Ranking fuzzy numbers based on the areas on the left and the right sides of fuzzy number. Computers and Mathematics with Applications, 61(2):431–442, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008692.

Noor:2011:NCA

[NM11b] Khalida Inayat Noor and Sarfraz Nawaz Malik. On a new class of analytic functions associated with conic domain. Computers and Mathematics with Applications, 62(1):367– 375, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004068.

Noor:2011:CIF

[NM11c] Khalida Inayat Noor and Sarfraz Nawaz Malik. On coefficient inequalities of functions associated with conic domains. Computers and Mathematics with Applications, 62(5):2209–2217, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005608.

Negri:2015:RBA

[NMR15] Federico Negri, Andrea Manzoni, and Gianluigi Rozza. Reduced basis approximation of parametrized optimal flow control problems for the Stokes equations. Computers and Mathematics with Applications, 69(4):319–336, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114006075.

Noor:2012:GEB

[NN12] Muhammad Aslam Noor and Khalida Inayat Noor. General equilibrium bifunction variational inequalities. Computers and Mathematics with Applications, 64 (11):3522–3526, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005676.

Noack:2013:HME

[NN13] Bernd R. Noack and Robert K. Niven. A hierarchy of maximum entropy closures for Galerkin systems of incompressible flows. Computers and Mathematics with Applications, 65(10):1558– 1574, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001259.

Noor:2011:ACO

[NNAS11a] Khalida Inayat Noor, Muhammad Aslam Noor, and Eisa Al-Said. Applications of certain operators to the classes of analytic functions related with generalized conic domains. Computers and Mathematics with Applications, 62 (11):4194–4206, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008674.

Noor:2011:AFB

[NNAS11b] Khalida Inayat Noor, Muhammad Aslam Noor, and Eisa Al-Said. On analytic functions of bounded boundary rotation of complex order. Computers and Mathematics with Applications, 62(4):2112–2125, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005402.

Noor:2011:CAF

[NNAS11c] Khalida Inayat Noor, Muhammad Aslam Noor, and Eisa A. Al-Said. On certain analytic functions with bounded radius rotation. Computers and Mathematics with Applications, 61(10):2987– 2993, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100263X.

Nasu:2013:SFE

[NNK13] Shoichi Nasu, Kazuya Nojima, and Mutsuto Kawahara. SUPG finite element method for adiabatic flows. Computers and Mathematics with Applications, 66(3):250–268, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300285X.

Ng:2013:PSF

[NNL13] K. C. Ng, Y. L. Ng, and W. H. Lam. Particle simulation and flow sequence on drainage of liquid particles. Computers and Mathematics with Applications, 66(8):1437–1451, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005129.

Nah:2014:MDL

[NNR14] Kyeongah Nah, Yukihiko Nakata, and Gergely Röst. Malaria dynamics with long incubation period in hosts. Computers and Mathematics with Applications, 68(9):915–930, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001825.

Noor:2011:SNS

[NNWAS11] Muhammad Aslam Noor, Khalida Inayat Noor, Asif Waheed, and Eisa A. Al-Said. Some new solitonary solutions of the modified Benjamin– Bona–Mahony equation. Computers and Mathematics with Applications, 62(4):2126–2131, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005414.

Nunokawa:2011:SPA

[NODA11] Mamoru Nunokawa, Shigeyoshi Owa, Emel Yavuz Duman, and Melike Aydogan. Some properties of analytic functions relating to the Miller and Mocanu result. Computers and Mathematics with Applications, 61(5):1291–1295, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009557.

Noor:2010:SSF

[Noo10] Khalida Inayat Noor. On some subclasses of mfold symmetric analytic functions. Computers and Mathematics with Applications, 60(1):14– 22, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002932.

Noor:2011:GUC

[Noo11] Khalida Inayat Noor. On a generalization of uniformly convex and related functions. Computers and Mathematics with Applications, 61(1):117–125, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000828X.

Niezgoda:2012:HLP

[NP12] Marek Niezgoda and Josip Pecarić. Hardy– Littlewood–Pólya-type theorems for invex functions. Computers and Mathematics with Applications, 64(4):518–526, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010947.

Nagaraj:2017:CDF

[NPD17] Sriram Nagaraj, Socratis Petrides, and Leszek F. Demkowicz. Construction of DPG Fortin operators for second order problems. Computers and Mathematics with Applications, 74(8):1964–1980, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303358.

Nath:2010:CHM

[NPR10] P. Raveendra Nath, P. M. V.

Prasad, and D. R. V. Prasada Rao. Computational hydromagnetic mixed convective heat and mass transfer through a porous medium in a non-uniformly heated vertical channel with heat sources and dissipation. Computers and Mathematics with Applications, 59(2):803–811, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007111.

Nashine:2011:CCF

[NS11] Hemant Kumar Nashine and Wasfi Shatanawi. Coupled common fixed point theorems for a pair of commuting mappings in partially ordered complete metric spaces. Computers and Mathematics with Applications, 62(4):1984–1993, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005220.

Nie:2013:SAE

[NSYY13] Cunyun Nie, Shi Shu, Haiyuan Yu, and Yuyue Yang. Superconvergence and asymptotic expansion for semidiscrete bilinear finite volume element approximation of the parabolic problem. Computers and Mathematics with Applications, 66(1):91–104, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001405.

Nhan:2017:GSB

[NT17] Le Cong Nhan and Le Xuan Truong. Global solution and blow-up for a class of pseudo p-Laplacian evolution equations with logarithmic nonlinearity. Computers and Mathematics with Applications, 73(9):2076–2091, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301098.

Navickas:2015:CEF

[NTR15] Z. Navickas, T. Telksnys, and M. Ragulskis. Comments on “The Exp-function method and generalized solitary solutions”. Computers and Mathematics with Applications, 69(8):798–803, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500070X.

See [Cha11c].

Noor:2012:SIT

[NUH12] Khalida Inayat Noor and Wasim Ul-Haq. On some implication type results involving generalized bounded Mocanu variations. Computers and Mathematics with Applications, 63(10): 1456–1461, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211200260X.

Noor:2011:IMS

[NUNAS11] Muhammad Aslam Noor, Saleem Ullah, Khalida Inayat Noor, and Eisa Al-Said. Iterative methods for solving extended general mixed variational inequalities. Computers and Mathematics with Applications, 62(2):804– 813, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004780.

Nie:2011:LTD

[NWZ11] Deming Nie, Yuming Wang, and Kai Zhang. Long-time decay of the translational/rotational velocity autocorrelation function for colloidal particles in two dimensions. Computers and Mathematics with Applications, 61(8):2152–2157, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006851.

Nguyen-Xuan:2014:IAE

[NXHN14] H. Nguyen-Xuan, T. Hoang, and V. P. Nguyen. An isogeometric analysis for elliptic homogenization problems. Computers and Mathematics with Applications, 67(9):1722–1741, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000029.

Ni:2014:RSN

[NZ14] Tie Ni and Jun Zhai.

A regularized smoothing Newton-type algorithm for quasi-variational inequalities. Computers and Mathematics with Applications, 68 (10):1312–1324, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004325.

Nguyen:2016:SOD

[NZ16] Duc Duy Nguyen and Shan Zhao. A second order dispersive FDTD algorithm for transverse electric Maxwell’s equations with complex interfaces. Computers and Mathematics with Applications, 71(4):1010–1035, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116000183.

Ozdemir:2011:HHT

[ÖAK11] M. Emin Özdemir, Merve Avci, and Havva Kavurmaci. Hermite–Hadamardtype inequalities via (α, m)-convexity. Computers and Mathematics with Applications, 61(9): 2614–2620, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001581.

Obrecht:2016:TLW

[OAKR16] Christian Obrecht, Pietro Asinari, Frédéric Kuznik, and Jean-Jacques Roux. Thermal link-wise artificial compressibility method: GPU implementation and validation of a double-population model. Computers and Mathematics with Applications, 72(2):375–385, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002576.

Oanh:2013:SMB

[Oan13] Nguyen Thi Ngoc Oanh. A splitting method for a backward parabolic equation with time-dependent coefficients. Computers and Mathematics with Applications, 65(1):17– 28, January 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112006062.

Ohwada:2011:ACM

[OAY11] Taku Ohwada, Pietro Asinari, and Daisuke Yabusaki. Artificial compressibility method and lattice Boltzmann method: Similarities and differences. Computers and Mathematics with Applications, 61(12): 3461–3474, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110005961.

Odibat:2010:MSD

[OBAAD10] Zaid M. Odibat, Cyrille Bertelle, M. A. Aziz-Alaoui, and Gérard H. E. Duchamp. A multistep differential transform method and application to non-chaotic or chaotic systems. Computers and Mathematics with Applications, 59(4):1462–1472, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007354.

OSullivan:2019:RNL

[OBCG19] S. O’Sullivan, R. E. Bird, W. M. Coombs, and S. Giani. Rapid nonlinear finite element analysis of continuous and discontinuous Galerkin methods in MATLAB. Computers and Mathematics with Applications, 78(9):3007–3026, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301312.

Ortigueira:2010:SIC

[OC10] Manuel D. Ortigueira and Fernando J. Coito. System initial conditions vs derivative initial conditions. Computers and Mathematics with Applications, 59(5):1782–1789, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005562.

Oleksy:2014:ECH

[OC14] M. Oleksy and W. Cecot. Estimation of computational homogenization error by explicit residual method. Computers and Mathematics with Applications, 66(12):2504–2516, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005865.

Orozco-Castaneda:2012:GBB

[OCNG12] Johanna Marcela Orozco-Castañeda, Daya K. Nagar, and Arjun K. Gupta. Generalized bivariate beta distributions involving Appell’s hypergeometric function of the second kind. Computers and Mathematics with Applications, 64(8):2507–2519, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004361.

Ostromsky:2015:PID

[ODAZ15] Tzvetan Ostromsky, Ivan Dimov, Vassil Alexandrov, and Zahari Zlatev. Preparing input data for sensitivity analysis of an air pollution model by using highperformance supercomputers and algorithms. Computers and Mathematics with Applications, 70(11):2773–2782, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003557.

Odibat:2010:ASL

[Odi10] Zaid M. Odibat. Analytic study on linear systems of fractional differential equations. Computers and Mathematics with Applications, 59(3):1171–1183, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109004155.

Orhan:2010:FSP

[ODR10] Halit Orhan, Erhan Deniz, and Dorina Raducanu. The Fekete–Szegő problem for subclasses of analytic functions defined by a differential operator related to conic domains. Computers and Mathematics with Applications, 59(1):283–295, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005082.

Otte:2016:DAL

[OF16] Philipp Otte and Martin Frank. Derivation and analysis of lattice Boltzmann schemes for the linearized Euler equations. Computers and Mathematics with Applications, 72(2):311–327, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005726.

Ogiela:2012:SAC

[Ogi12] Lidia Ogiela. Semantic analysis in cognitive UBIAS & E-UBIAS systems. Computers and Mathematics with Applications, 63(2):378–390, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005992.

Ogut:2013:CDS

[Ögü13] Hulisi Ögüt. The configuration and detection strategies for information security systems. Computers and Mathematics with Applications, 65(9):1234–1253, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004142.

Oh:2015:RCA

[Oh15] Minah Oh. de Rham complexes arising from Fourier finite element methods in axisymmetric domains. Computers and Mathematics with Applications, 70(8):2063–2073, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003995.

Osaki:2019:NSF

[OHK^{+}19] Susumu Osaki, Kosuke Hayashi, Hidehito Kimura, Takeshi Seta, Eiji Kohmura, and Akio Tomiyama. Numerical simulations of flows in cerebral aneurysms using the lattice Boltzmann method with single- and multiplerelaxation time collision models. Computers and Mathematics with Applications, 78(8):2746–2760, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302299.

Obembe:2018:MMB

[OHMAK18] Abiola D. Obembe, M. Enamul Hossain, Kassem Mustapha, and Sidqi A. Abu-Khamsin. A modified memory-based mathematical model describing fluid flow in porous media. Computers and Mathematics with Applications, 73(6):1385–1402, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306393.

Ozturk:2012:FSS

[ÖI12] Mehmet Ali Öztürk and Ebubekir Inan. Fuzzy soft subnear-rings and (∈, ∈ ∨q)-fuzzy soft subnearrings. Computers and Mathematics with Applications, 63(3):617–628, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009746.

Ozcan:2019:RSM

[ÖKJR19] A. Özcan, S. Kollmannsberger, J. Jomo, and E. Rank. Residual stresses in metal deposition modeling: Discretizations of higher order. Computers and Mathematics with Applications, 78(7):2247–2266, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306242.

Obrecht:2011:NAL

[OKTR11] Christian Obrecht, Frédéric Kuznik, Bernard Tourancheau, and Jean-Jacques Roux. A new approach to the lattice Boltzmann method for graphics processing units. Computers and Mathematics with Applications, 61(12): 3628–3638, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211000091X.

Obrecht:2013:EGI

[OKTR13a] Christian Obrecht, Frédéric Kuznik, Bernard Tourancheau, and Jean-Jacques Roux. Efficient GPU implementation of the linearly interpolated bounce-back boundary condition. Computers and Mathematics with Applications, 65 (6):936–944, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112004130.

Obrecht:2013:MGI

[OKTR13b] Christian Obrecht, Frédéric Kuznik, Bernard Tourancheau, and Jean-Jacques Roux. Multi-GPU implementation of the lattice Boltzmann method. Computers and Mathematics with Applications, 65(2):252– 261, January 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111001064.

Olszowy:2010:SIS

[Ols10] Leszek Olszowy. Solvability of infinite systems of singular integral equations in Fréchet space of continuous functions. Computers and Mathematics with Applications, 59(8):2794–2801, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000842.

Oftadeh:2010:NMH

[OMS10] R. Oftadeh, M. J. Mahjoob, and M. Shariatpanahi. A novel meta-heuristic optimization algorithm inspired by group hunting of animals: Hunting search. Computers and Mathematics with Applications, 60(7):2087–2098, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005419.

Odzijewicz:2012:GFC

[OMT12] Tatiana Odzijewicz, Agnieszka B. Malinowska, and Delfim F. M. Torres. Generalized fractional calculus with applications to the calculus of variations. Computers and Mathematics with Applications, 64 (10):3351–3366, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001058.

Ogiela:2010:UML

[OO10] Marek R. Ogiela and Urszula Ogiela. The use of mathematical linguistic methods in creating secret sharing threshold algorithms. Computers and Mathematics with Applications, 60(2):267– 271, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001379.

Ostrovska:2011:NEB

[OÖ11] Sofiya Ostrovska and Ahmet Yaşar Özban. The norm estimates of the q-Bernstein operators for varying q > 1. Computers and Mathematics with Applications, 62 (12):4758–4771, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009436.

Ogiela:2012:LPS

[OO12] Marek R. Ogiela and Urszula Ogiela. Linguistic protocols for secure information management and sharing. Computers and Mathematics with Applications, 63(2):564–572, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009023.

Ooi:2014:EHB

[OP14] Ean Hin Ooi and Viktor Popov. An efficient hybrid BEM–RBIE method for solving conjugate heat transfer problems. Computers and Mathematics with Applications, 66(12):2489–2503, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005877.

Olteanu:2012:DRA

[OPDC12] Alexandra Olteanu, Florin Pop, Ciprian Dobre, and Valentin Cristea. A dynamic rescheduling algorithm for resource management in large scale dependable distributed systems. Computers and Mathematics with Applications, 63(9):1409–1423, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001915.

Orhan:2011:SMM

[ORD11] Halit Orhan, Dorina Raducanu, and Erhan Deniz. Subclasses of meromorphically multivalent functions defined by a differential operator. Computers and Mathematics with Applications, 61(4):966–979, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009508.

Ossandon:2016:NNS

[ORR16] Sebastián Ossandón, Camilo Reyes, and Carlos M. Reyes. Neural network solution for an inverse problem associated with the Dirichlet eigenvalues of the anisotropic Laplace operator. Computers and Mathematics with Applications, 72(4):1153–1163, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303698.

Oruc:2017:NAB

[Oru17] Ömer Oruç. A new algorithm based on Lucas polynomials for approximate solution of 1D and 2D nonlinear generalized Benjamin– Bona–Mahony–Burgers equation. Computers and Mathematics with Applications, 74 (12):3042–3057, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304753.

Oruc:2019:NUH

[Oru19] Ömer Oruç. A nonuniform Haar wavelet method for numerically solving two-dimensional convection-dominated equations and two-dimensional near singular elliptic equations. Computers and Mathematics with Applications, 77(7):1799– 1820, April 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306710.

Ozen:2011:MIS

[ÖŞ11] Mehmet Özen and Vedat Şiap. The MacWilliams identity for m-spotty weight enumerators of linear codes over finite fields. Computers and Mathematics with Applications, 61(4):1000–1004, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009533.

Othman:2013:EFP

[OSA13] M. I. A. Othman, N. Sarkar, and Sarhan. Y. Atwa. Effect of fractional parameter on plane waves of generalized magnetothermoelastic diffusion with reference temperaturedependent elastic medium. Computers and Mathematics with Applications, 65(7):1103–1118, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000771.

See comments [Her14].

Osman:2018:MSS

[Osm18] M. S. Osman. On multisoliton solutions for the (2+1)-dimensional breaking soliton equation with variable coefficients in a graded-index waveguide. Computers and Mathematics with Applications, 75(1):1–6, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305254.

Ozden:2010:UPG

[OSS10] Hacer Ozden, Yilmaz Simsek, and H. M. Srivastava. A unified presentation of the generating functions of the generalized Bernoulli, Euler and Genocchi polynomials. Computers and Mathematics with Applications, 60(10):2779–2787, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007170.

Ostrovska:2011:LT

[Ost11] Sofiya Ostrovska. On the Lupaş q-transform. Computers and Mathematics with Applications, 61(3):527–532, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008977.

Oplatkova:2013:APT

[OSZP13] Zuzana Kominkova Oplatkova, Roman Senkerik, Ivan Zelinka, and Michal Pluhacek. Analytic programming in the task of evolutionary synthesis of a controller for high order oscillations stabilization of discrete chaotic systems. Computers and Mathematics with Applications, 66(2):177–189, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001004.

Ogino:2016:NSI

[OTiSY16] Masao Ogino, Amane Takei, Shin ichiro Sugimoto, and Shinobu Yoshimura. A numerical study of iterative substructuring method for finite element analysis of high frequency electromagnetic fields. Computers and Mathematics with Applications, 72(8):2020–2027, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302097.

Ouahab:2012:FSD

[Oua12] Abdelghani Ouahab. Fractional semilinear differential inclusions. Computers and Mathematics with Applications, 64 (10):3235–3252, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002441.

Ouyang:2011:EUS

[Ouy11]  Zigen Ouyang.  Ex-

istence and uniqueness of the solutions for a class of nonlinear fractional order partial differential equations with delay. Computers and Mathematics with Applications, 61(4):860–870, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009399.

Oganesov:2016:BDG

[OVV^{+}16] Armen Oganesov, George Vahala, Linda Vahala, Jeffrey Yepez, and Min Soe. Benchmarking the Dirac-generated unitary lattice qubit collisionstream algorithm for 1D vector Manakov soliton collisions. Computers and Mathematics with Applications, 72(2):386– 393, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002874.

Orfi:2019:PEE

[OY19] Ajmia Younes Orfi and Driss Yakoubi. A posteriori error estimates of finite element method for the time-dependent Darcy problem in an axisymmetric domain. Computers and Mathematics with Applications, 77(10):2833–2853, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300409.

Ogiela:2012:TCC

[OYXK12] Marek R. Ogiela, Ilsun You, Fatos Xhafa, and Hoon Ko. Towards context, cognitive, and secure computing. Computers and Mathematics with Applications, 63(2):337–338, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011126.

Ozbekler:2011:OSS

[ÖZ11] A. Özbekler and A. Zafer.

Oscillation of solutions of second order mixed nonlinear differential equations under impulsive perturbations. Computers and Mathematics with Applications, 61(4):933–940, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009466.

Ozarslan:2011:UAB

[Öza11] M. Ali Özarslan. Unified Apostol–Bernoulli, Euler and Genocchi polynomials. Computers and Mathematics with Applications, 62(6):2452–2462, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005852.

Ozdemir:2018:MTC

[Özd18] I. Bedii Özdemir. A modification to temperaturecomposition pdf method and its application to the simulation of a transitional bluff-body flame. Computers and Mathematics with Applications, 75(7):2574–2592, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307940.

Oh:2019:AMF

[OZF19] Jaeyoun Oh, Huiqing Zhu, and Zhuojia Fu. An adaptive method of fundamental solutions for solving the Laplace equation. Computers and Mathematics with Applications, 77(7):1828–1840, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306746.

Ozugurlu:2015:NNA

[Özu15] E. Özuğurlu. A note on the numerical approach for the reaction– diffusion problem to model the density of the tumor growth dynamics. Computers and Mathematics with Applications, 69(12):1504– 1517, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001996.

Prajapat:2012:MPC

[PA12] J. K. Prajapat and M. K.

Aouf. Majorization problem for certain class of pvalently analytic function defined by generalized fractional differintegral operator. Computers and Mathematics with Applications, 63(1):42–47, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009412.

Pauli:2015:DOM

[PA15] Stefan Pauli and Peter Arbenz. Determining optimal multilevel Monte Carlo parameters with application to fault tolerance. Computers and Mathematics with Applications, 70(11):2638–2651, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003466.

Pardo:2015:IEL

[PÁAP^{+}15] D. Pardo, J. Álvarez-Aramberri, M. Paszynski, L. Dalcin, and V. M. Calo. Impact of elementlevel static condensation on iterative solver performance. Computers and Mathematics with Applications, 70 (10):2331–2341, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004198.

Padgett:2018:QST

[Pad18] Joshua L. Padgett. The quenching of solutions to time-space fractional Kawarada problems. Computers and Mathematics with Applications, 76(7):1583–1592, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303742.

Papaodysseus:2012:ESP

[PAE^{+}12] Constantin Papaodysseus, Dimitris Arabadjis, Michalis Exarhos, Panayiotis Rousopoulos, Solomon Zannos, Michail Panagopoulos, and Lena Papazoglou-Manioudaki. Efficient solution to the 3D problem of automatic wall paintings reassembly. Computers and Mathematics with Applications, 64(8):2712–2734, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005147.

Palmer:2012:ADI

[Pal12] Peter Palmer. Application of a discrete Itô formula to determine stability (instability) of the equilibrium of a scalar linear stochastic difference equation. Computers and Mathematics with Applications, 64(7):2302–2311, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002179.

Pal:2013:HCM

[Pal13] Dulal Pal. Hall current and MHD effects on heat transfer over an unsteady stretching permeable surface with thermal radiation. Computers and Mathematics with Applications, 66(7):1161–1180, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004379.

Pan:2011:NOS

[Pan11] Victor Y. Pan. Nearly optimal solution of rational linear systems of equations with symbolic lifting and numerical initialization. Computers and Mathematics with Applications, 62(4):1685–1706, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004743.

Pantokratoras:2017:CP

[Pan17] Asterios Pantokratoras.

Comment on the paper “Convection from an inverted cone in a porous medium with cross-diffusion effects”, F. G. Awad, P. Sibanda, S. S. Motsa, O. D. Makinde, [Comput. Math. Appl. 61 (2011) 1431–1441]. Computers and Mathematics with Applications, 74(8):1779–1781, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304005.

See [ASMM11].

Papanicolopulos:2015:CMD

[Pap15] Stefanos-Aldo Papanicolopulos. Computation of moderate-degree fully-symmetric cubature rules on the triangle using symmetric polynomials and algebraic solving. Computers and Mathematics with Applications, 69(7):650–666, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S089812211500067X.

Parhi:2011:NOS

[Par11a] N. Parhi. Non-oscillation of solutions of difference equations of third order. Computers and Mathematics with Applications, 62(10):3812–3820, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007978.

Paris:2011:DAL

[Par11b] R. B. Paris. The discrete analogue of Laplace’s method. Computers and Mathematics with Applications, 61(10):3024– 3034, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002720.

Park:2015:ADR

[Par15] Sun-Hye Park. Arbitrary decay rates of energy for a von Kármán equation of memory type. Computers and Mathematics with Applications, 70(8):1878–1886, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003739.

Park:2017:CPM

[Par17a] Won-Kwang Park. Certain properties of a MUSIC-type imaging functional in inverse scattering from an open sound-hard arc. Computers and Mathematics with Applications, 74 (6):1232–1245, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303577.

Park:2017:NSS

[Par17b] Won-Kwang Park. A novel study on subspace migration for imaging of a sound-hard arc. Computers and Mathematics with Applications, 74 (12):3000–3007, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304741.

Park:2018:GDK

[Par18]  Sun-Hye Park.  Gen-

eral decay of a von karman plate equation with memory on the boundary. Computers and Mathematics with Applications, 75(9):3067–3080, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300488.

Paszynska:2014:VNA

[Pas14] Anna Paszyńska. Volume and neighbors algorithm for finding elimination trees for three dimensional h-adaptive grids. Computers and Mathematics with Applications, 68(10):1467–1478, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400460X.

Pooseh:2012:AFI

[PAT12] Shakoor Pooseh, Ricardo Almeida, and Delfim F. M. Torres. Approximation of fractional integrals by means of derivatives. Computers and Mathematics with Applications, 64(10):3090–3100, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122112000892.

Pooseh:2013:DDM

[PAT13] Shakoor Pooseh, Ricardo Almeida, and Delfim F. M. Torres. Discrete direct methods in the fractional calculus of variations. Computers and Mathematics with Applications, 66(5):668–676, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000758.

Pourahmad:2011:FLR

[PATA11] Saeedeh Pourahmad, Seyyed Mohammad Taghi Ayatollahi, S. Mahmoud Taheri, and Zahra Habib Agahi. Fuzzy logistic regression based on the least squares approach with application in clinical studies. Computers and Mathematics with Applications, 62 (9):3353–3365, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007231.

Petkovic:2011:FRP

[PB11] Marko D. Petković and Milan Basić. Further results on the perfect state transfer in integral circulant graphs. Computers and Mathematics with Applications, 61(2):300–312, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008540.

Palmeira:2012:EFL

[PB12] E. S. Palmeira and B. C. Bedregal. Extension of fuzzy logic operators defined on bounded lattices via retractions. Computers and Mathematics with Applications, 63(6):1026–1038, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010509.

Prasad:2013:FRB

[PB13] Yamuna Prasad and K. K. Biswas. Fuzzy rough based regularization in generalized multiple kernel learning. Computers and Mathematics with Applications, 66 (10):1770–1781, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004896.

Peterson:2019:ESP

[PBK19] Kara Peterson, Pavel Bochev, and Paul Kuberry. Explicit synchronous partitioned algorithms for interface problems based on Lagrange multipliers. Computers and Mathematics with Applications, 78(2): 459–482, July 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118305637.

Paulin:2019:ICC

[PBM19] C. Paulin, J.-P. Braeunig, and R. Motte. Isentropic correction for collocated Lagrange–Remap scheme. Computers and Mathematics with Applications, 78(2):623–642, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303638.

Pitkaranta:2012:DRV

[PBS12] Juhani Pitkäranta, Ivo Babuska, and Barna Szabó. The dome and the ring: Verification of an old mathematical model for the design of a stiffened shell roof. Computers and Mathematics with Applications, 64(1):48– 72, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002131.

Pai:2011:ESR

[PC11a] Neng-Sheng Pai and Shih-Ping Chang. An embedded system for real-time facial expression recognition based on the extension theory. Computers and Mathematics with Applications, 61(8):2101–2106, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006668.

Park:2011:PSD

[PC11b] Jea-Hyun Park and Soon-Yeong Chung. Positive solutions for discrete boundary value problems involving the p-Laplacian with potential terms. Computers and Mathematics with Applications, 61(1):17–29, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008059.

Pai:2012:DIF

[PC12] Neng-Sheng Pai and Shih-Ping Chang. Design and implementation of fuzzy sliding mode controllers for generalized projective synchronization of chaos horizontal platform systems. Computers and Mathematics with Applications, 64(5):709–720, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009011.

Pan:2014:SBD

[PC14] Kuo-Long Pan and Zhi-Jen Chen. Simulation of bubble dynamics in a microchannel using a front-tracking method. Computers and Mathematics with Applications, 67(2):290–306, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002514.

Pandit:2017:RHO

[PC17] Swapan K. Pandit and Anirban Chattopadhyay. A robust higher order compact scheme for solving general second order partial differential equation with derivative source terms on nonuniform curvilinear meshes. Computers and Mathematics with Applications, 74(6):1414–1434, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303796.

Park:2013:EVM

[PCK13] Jin Han Park, Hyun Ju Cho, and Young Chel Kwun. Extension of the VIKOR method to dynamic intuitionistic fuzzy multiple attribute decision making. Computers and Mathematics with Applications, 65(4):731–744, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112007043.

Popa:2012:ESL

[PCM12] Ioan-Lucian Popa, Traian Ceausu, and Mihail Megan. On exponential stability for linear discrete-time systems in Banach spaces. Computers and Mathematics with Applications, 63(11):1497–1503, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000375.

Pandit:2016:FOC

[PCO16] Swapan K. Pandit, Anirban Chattopadhyay, and Hakan F. Oztop. A fourth order compact scheme for heat transfer problem in porous media. Computers and Mathematics with Applications, 71(3):805–832, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115006057.

Pang:2013:GJT

[PCS13] Guofei Pang, Wen Chen, and K. Y. Sze. Gauss– Jacobi-type quadrature rules for fractional directional integrals. Computers and Mathematics with Applications, 66(5):597–607, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002277.

Perumal:2011:MSS

[PD11] D. Arumuga Perumal and Anoop K. Dass. Multiplicity of steady solutions in two-dimensional liddriven cavity flows by lattice Boltzmann method. Computers and Mathematics with Applications, 61(12):3711–3721, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002427.

Petrides:2017:ADM

[PD17] Socratis Petrides and Leszek F. Demkowicz. An adaptive DPG method for high frequency timeharmonic wave propagation problems. Computers and Mathematics with Applications, 74(8):1999–2017, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304029.

Pu:2012:IFE

[PDHL12] Yu-Chi Pu, Wei-Chang Du, Chien-Hsiang Huang, and Chen-Kuo Lai. Invariant feature extraction for 3D model retrieval: an adaptive approach using Euclidean and topological metrics. Computers and Mathematics with Applications, 64(5):1217–1225, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002702.

Pardo:2010:FMD

[PdlF10] María José Pardo and David de la Fuente. Fuzzy Markovian decision processes: Application to queueing systems. Computers and Mathematics with Applications, 60(9):2526–2535, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005675.

Pan:2011:ASS

[PDM11] Xiao-Jun Pan, Chao-Qing Dai, and Lu-Feng Mo. Analytical solutions for the stochastic Gardner equation. Computers and Mathematics with Applications, 61(8):2138–2141, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006826.

Parmananda:2019:NAT

[PDN19] Mukul Parmananda, Amaresh Dalal, and Ganesh Natarajan. Numerical appraisal of three low Mach number algorithms for radiativeconvective flows in enclosures. Computers and Mathematics with Applications, 77(8):2162–2181, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306977.

Pedregal:2018:DAC

[Ped18] Pablo Pedregal. A direct algorithm for constrained variational problems in several dimensions. Computers and Mathematics with Applications, 75(1):105–121, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730531X.

Peng:2011:RNA

[Pen11] Lequn Peng. Retraction notice to “Asymptotic behavior of solutions to a differential equation with state-dependent delay” [Comput. Math. Appl. 57 (2009) 1511– 1514]. Computers and Mathematics with Applications, 61(2):513, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009181.

Perugia:2018:PMC

[Per18]  Ilaria Perugia.  Pe-

ter Monk’s contributions to numerical analysis and Maxwell’s equations. Computers and Mathematics with Applications, 74(11):2645–2649, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301852.

Pestov:2013:NCH

[Pes13] Vladimir Pestov. Is the k–NN classifier in high dimensions affected by the curse of dimensionality? Computers and Mathematics with Applications, 65(10):1427– 1437, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006426.

Petras:2011:MNA

[Pet11] Ivo Petrás. Modeling and numerical analysis of fractional-order Bloch equations. Computers and Mathematics with Applications, 61(2):341– 356, January 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008588.

Petkovic:2014:GSI

[Pet14] Marko D. Petković. Generalized schultz iterative methods for the computation of outer inverses. Computers and Mathematics with Applications, 67(10):1837– 1847, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001527.

Peters:2015:GSF

[Pet15]  Jörg Peters.  General

spline filters for discontinuous Galerkin solutions. Computers and Mathematics with Applications, 70(5):1046–1050, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500320X.

Paz:2010:SUF

[PFBL10] Martín C. Rodríguez Paz, Renato G. Ferraz, Arturo Suman Bretas, and Roberto Chouhy Leborgne. System unbalance and fault impedance effect on faulted distribution networks. Computers and Mathematics with Applications, 60(4):1105–1114, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002579.

Park:2017:LSS

[PFDG17] Keunsoo Park, Maria Fernandino, Carlos A. Dorao, and Marc Gerritsma. The least-squares spectral element method for phase-field models for isothermal fluid mixture. Computers and Mathematics with Applications, 74(8):1981–1998, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304182.

Pineda:2010:MCG

[PG10] M. Delgado Pineda and E. A. Galperin. MAPLE code for the gamma algorithm for global optimization of uncertain functions in economy and finance. Computers and Mathematics with Applications, 59(8):2951–2963, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001227.

Pang:2018:IRB

[PGDL18] Zhi-Feng Pang, Li-Zhen Guo, Yuping Duan, and Jian Lu. Image restoration based on the minimized surface regularization. Computers and Mathematics with Applications, 76(8):1893–1905, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304085.

Park:2018:CCA

[PGF18] Keunsoo Park, Marc Gerritsma, and Maria Fernandino. C^{1} continuous h-adaptive leastsquares spectral element method for phase-field models. Computers and Mathematics with Applications, 75(5):1582–1594, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307526.

Pacciarini:2016:SBD

[PGQ16] Paolo Pacciarini, Paola Gervasio, and Alfio Quarteroni. Spectral based discontinuous Galerkin reduced basis element method for parametrized Stokes problems. Computers and Mathematics with Applications, 72(8):1977–1987, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300268.

Peng:2019:LBM

[PGW19] Cheng Peng, Zhaoli Guo, and Lian-Ping Wang. A lattice-BGK model for the Navier–Stokes equations based on a rectangular grid. Computers and Mathematics with Applications, 78(4):1076–1094, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302656.

Park:2013:AAB

[PH13] Ji-Yong Park and Seog-Young Han. Application of artificial bee colony algorithm to topology optimization for dynamic stiffness problems. Computers and Mathematics with Applications, 66 (10):1879–1891, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004501.

Peng:2019:NBP

[PH19] Li Peng and Yunqing Huang. On nonlocal backward problems for fractional stochastic diffusion equations. Computers and Mathematics with Applications, 78 (5):1450–1462, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122119300343.

Pourjabari:2019:EPF

[PHM^{+}19] Amin Pourjabari, Zanyar Esmailpoor Hajilak, Alireza Mohammadi, Mostafa Habibi, and Hamed Safarpour. Effect of porosity on free and forced vibration characteristics of the GPL reinforcement composite nanostructures. Computers and Mathematics with Applications, 77(10):2608–2626, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307399.

Parvin:2012:SMW

[PHPK12] Sazia Parvin, Farookh Khadeer Hussain, Jong Sou Park, and Dong Seong Kim. A survivability model in wireless sensor networks. Computers and Mathematics with Applications, 64(12):3666–3682, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001411.

Park:2010:SIC

[PHWM10] Jong Hyuk Park, Sajid Hussain, Guilin Wang, and Yi Mu. Special issue of computers and mathematics with applications on “Advances in cryptography, security and applications for future computer science”. Computers and Mathematics with Applications, 60(2): 175, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002695.

Peinado:2010:ABA

[PIAH10] J. Peinado, J. Ibáñez, E. Arias, and V. Hernández. Adams–Bashforth and Adams–Moulton methods for solving differential Riccati equations. Computers and Mathematics with Applications, 60(11):3032–3045, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007777.

Piri:2011:SCM

[Pir11] Hossein Piri. Strong convergence for a minimization problem on solutions of systems of equilibrium problems and common fixed points of an infinite family and semigroup of nonexpansive mappings. Computers and Mathematics with Applications, 61(9):2562– 2577, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001489.

Pituk:2012:LBV

[Pit12] Mihály Pituk. A limit boundary value problem for a nonlinear difference equation. Computers and Mathematics with Applications, 64(7):2364–2369, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004385.

Perez-Jorda:2017:FSS

[PJ17] José M. Pérez-Jordá. Fast solution of Schrödinger’s equation using linear combinations of plane waves. Computers and Mathematics with Applications, 74(12):3318– 3327, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117305187.

Povstenko:2019:TFH

[PK19] Yuriy Povstenko and Tamara Kyrylych. Timefractional heat conduction in an infinite plane containing an external crack under heat flux loading. Computers and Mathematics with Applications, 78(5):1386–1395, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300501.

Pavlicek:2019:ACS

[PKD19] Karel Pavlícek, Václav Kotlan, and Ivo Dolezel. Applicability and comparison of surrogate techniques for modeling of selected heating problems. Computers and Mathematics with Applications, 78(9):2897–2910, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300811.

Park:2012:SPE

[PKK12] Jin Han Park, Oe Hyeon Kim, and Young Chel Kwun. Some properties of equivalence soft set relations. Computers and Mathematics with Applications, 63(6):1079–1088, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211101056X.

Piao:2013:PBK

[PKTH13] Yanji Piao, JongUk Kim, Usman Tariq, and Manpyo Hong. Polynomialbased key management for secure intra-group and inter-group communication. Computers and Mathematics with Applications, 65(9):1300–1309, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001228.

Pearson:2010:KRA

[PL10a] Jane M. Pearson and Noel G. Lloyd. Kukles revisited: Advances in computing techniques. Computers and Mathematics with Applications, 60(10):2797–2805, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110007200.

Puninagool:2010:CTS

[PL10b] Wattapong Puninagool and Sorasak Leeratanavalee. Complexity of terms, superpositions, and generalized hypersubstitutions. Computers and Mathematics with Applications, 59(2):1038–1045, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109004143.

Pham:2017:PEE

[PL17] Duong Thanh Pham and Tung Le. A posteriori error estimation for the Laplace–Beltrami equation on spheres with spherical splines. Computers and Mathematics with Applications, 74 (10):2298–2320, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304200.

Pan:2016:AVC

[PLKC16] Xiaomin Pan, Changhoon Lee, Kyoungyoun Kim, and Jung-Il Choi. Analysis of velocity-components decoupled projection method for the incompressible Navier–Stokes equations. Computers and Mathematics with Applications, 71(8):1722–1743, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301183.

Prieto-Langarica:2012:DCA

[PLKCC12] Alicia Prieto-Langarica, Hristo V. Kojouharov, and Benito M. Chen-Charpentier. Discrete and continuous approaches to modeling cell movement in the presence of a foreign stimulus. Computers and Mathematics with Applications, 64(3):167–174, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010418.

Prieto-Langarica:2013:UDC

[PLKCC13] Alicia Prieto-Langarica, Hristo V. Kojouharov, and Benito M. Chen-Charpentier. Upscaling from discrete to continuous mathematical models of two interacting populations. Computers and Mathematics with Applications, 66(9):1606–1612, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001028.

Prestininzi:2014:GKM

[PLMS14] Pietro Prestininzi, Michele La Rocca, Andrea Montessori, and Giampiero Sciortino. A gas-kinetic model for 2D transcritical shallow water flows propagating over dry bed. Computers and Mathematics with Applications, 68(4):439–453, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002831.

Pellerin:2015:ISA

[PLR15] Nicolas Pellerin, Sébastien Leclaire, and Marcelo Reggio. An implementation of the Spalart– Allmaras turbulence model in a multi-domain lattice Boltzmann method for solving turbulent airfoil flows. Computers and Mathematics with Applications, 70 (12):3001–3018, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004897.

Pu:2017:EWS

[PLT17] Yang Pu, Jiu Liu, and Chun-Lei Tang. Existence of weak solutions for a class of fractional Schrödinger equations with periodic potential. Computers and Mathematics with Applications, 73(3):465–482, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306757.

Pew:2019:PAE

[PLT^{+}19] Jack Pew, Zhi Li, Connor Tannahill, Paul Muir, and Graeme Fairweather. Performance analysis of error-control B-spline Gaussian collocation software for PDEs. Computers and Mathematics with Applications, 77(7):1888–1901, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306783.

Pang:2018:SAM

[PLW^{+}18] Fuzhen Pang, Haichao Li, Xueren Wang, Xuhong Miao, and Shuo Li. A semi analytical method for the free vibration of doubly-curved shells of revolution. Computers and Mathematics with Applications, 75(9):3249–3268, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300610.

Pingen:2010:ODN

[PM10] Georg Pingen and Kurt Maute. Optimal design for non-Newtonian flows using a topology optimization approach. Computers and Mathematics with Applications, 59(7):2340–2350, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006336.

Pinto:2013:FMM

[PM13] Carla M. A. Pinto and J. A. Tenreiro Machado. Fractional model for malaria transmission under control strategies.

Computers and Mathematics with Applications, 66(5):908–916, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006785.

Porogo:2017:VAE

[PMA17] O. P. Porogo, B. Muatjetjeja, and A. R. Adem. Variational approach and exact solutions for a generalized coupled Zakharov– Kuznetsov system. Computers and Mathematics with Applications, 73(5):864–872, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730041X.

Papacek:2017:CFF

[PMM17] Stepán Papácek, Benn Macdonald, and Ctirad Matonoha. Closedform formulas vs. PDE based numerical solution for the FRAP data processing: Theoretical and practical comparison. Computers and Mathematics with Applications, 73(8):1673–1683, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300767.

Padhan:2010:SOD

[PN10] S. K. Padhan and C. Nahak. Second order duality for the variational problems under ρ − (η, θ)-invexity. Computers and Mathematics with Applications, 60 (12):3072–3081, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000787X.

Pazouki:2016:NIM

[PN16] Arman Pazouki and Dan Negrut. Numerical investigation of microfluidic sorting of microtissues. Computers and Mathematics with Applications, 72(2):251–263, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004885.

Popescu:2011:TGS

[Pop11] Ovidiu Popescu. Two generalizations of some fixed point theorems. Computers and Mathematics with Applications, 62(10):3912–3919, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008121.

Popova:2013:IEP

[Pop13] Evgenija D. Popova. Inner estimation of the parametric tolerable solution set. Computers and Mathematics with Applications, 66(9):1655–1665, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001946.

Popova:2014:IES

[Pop14] Evgenija D. Popova. Improved enclosure for some parametric solution sets with linear shape. Computers and Mathematics with Applications, 68(9):994–1005, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001588.

Portillo:2018:NCE

[Por18] A. M. Portillo. Near conserving energy numerical schemes for twodimensional coupled seismic wave equations. Computers and Mathematics with Applications, 75(3):1016–1037, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306909.

Povstenko:2012:NBV

[Pov12a] Yuriy Povstenko. Neumann boundary-value problems for a timefractional diffusion-wave equation in a half-plane. Computers and Mathematics with Applications, 64(10):3183–3192, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001897.

Povstenko:2012:TTS

[Pov12b] Yuriy Povstenko. Theories of thermal stresses based on space-time-fractional telegraph equations. Computers and Mathematics with Applications, 64(10):3321– 3328, November 2012.

CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112000879.

Povstenko:2019:GTD

[Pov19] Yuriy Povstenko. Generalized theory of diffusive stresses associated with the time-fractional diffusion equation and nonlocal constitutive equations for the stress tensor. Computers and Mathematics with Applications, 78(6):1819–1825, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300888.

Pamuk:2010:HHP

[PP10] Serdal Pamuk and Nevin Pamuk. He’s homotopy perturbation method for continuous population models for single and interacting species. Computers and Mathematics with Applications, 59(2): 612–621, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109007317.

Preda:2012:ABS

[PP12] Ciprian Preda and Petre Preda. On the asymptotic behavior of the solutions of autonomous equations without unstable invariant manifolds. Computers and Mathematics with Applications, 64(1):35– 47, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001721.

Park:2014:PAN

[PP14] Sun Hye Park and Jong Yeoul Park. Pullback attractor for a non-autonomous modified Swift–Hohenberg equation. Computers and Mathematics with Applications, 67(3):542–548, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006676.

Paszynski:2013:DSR

[PPC13] Maciej Paszynski, David Pardo, and Victor M. Calo. A direct solver with reutilization of LU factorizations for h-adaptive finite element grids with point singularities. Computers and Mathematics with Applications, 65(8):

1140–1151, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113000989.

Paszynski:2015:DSP

[PPC15] Maciej Paszynski, David Pardo, and Victor M. Calo. Direct solvers performance on h-adapted grids. Computers and Mathematics with Applications, 70(3):282–295, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002187.

Petkovic:2010:AGI

[PPD10] M. S. Petković, L. D.

Petković, and J. Dzunić. Accelerating generators of iterative methods for finding multiple roots of nonlinear equations. Computers and Mathematics with Applications, 59(8):2784–2793, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000830.

Parshad:2016:LTD

[PQB^{+}16] Rana D. Parshad, Emmanuel Quansah, Kelly Black, Ranjit Kumar Upadhyay, S. K. Tiwari, and Nitu Kumari. Long time dynamics of a three-species food chain model with Allee effect in the top predator. Computers and Mathematics with Applications, 71(2):503–528, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005830.

Parshad:2017:SDB

[PQBK17] Rana D. Parshad, Emmanuel Quansah, Matthew A. Beauregard, and Said Kouachi. On “small” data blow-up in a three species food chain model. Computers and Mathematics with Applications, 73(4):576–587, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306897.

Panigrahi:2011:OAB

[PR11] S. Panigrahi and P. Rami Reddy. On oscillatory and asymptotic behavior of fourth order nonlinear neutral delay dynamic equations. Computers and Mathematics with Applications, 62 (11):4258–4271, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008777.

Pavlovic:2010:AMS

[PRR10] Mirjana Pavlović, Stojan Radenović, and Slobodan Radojević. Abstract metric spaces and Sehgal– Guseman-type theorems. Computers and Mathematics with Applications, 60(3):865–872, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003901.

Papageorgiou:2018:PSC

[PRR18] Nikolaos S. Papageorgiou, Vicentiu D. Radulescu, and Dusan D. Repovs. Periodic solutions for a class of evolution inclusions. Computers and Mathematics with Applications, 75(8):3047–3065, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300476.

Praetorius:2018:CIE

[PRS18] Dirk Praetorius, Michele Ruggeri, and Bernhard Stiftner. Convergence of an implicit-explicit midpoint scheme for computational micromagnetics. Computers and Mathematics with Applications, 75(5):1719–1738, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730754X.

Papaschinopoulos:2012:DTE

[PS12a] G. Papaschinopoulos and C. J. Schinas. On the dynamics of two exponential type systems of difference equations. Computers and Mathematics with Applications, 64(7):2326–2334, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003380.

Patterson:2012:RCM

[PS12b] Richard F. Patterson and Ekrem Savas. RHconservative matrix characterization of P -convergence in probability. Computers and Mathematics with Applications, 63(6):1020– 1025, March 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111009333.

Pang:2016:FOF

[PS16] Hong-Kui Pang and Hai-Wei Sun. Fourth order finite difference schemes for time-space fractional sub-diffusion equations. Computers and Mathematics with Applications, 71(6):1287–1302, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300530.

Petsche:2018:PEC

[PS18] Jan Petsche and Andreas Schröder. A posteriori error control and adaptivity of hp-finite elements for mixed and mixed-hybrid methods. Computers and Mathematics with Applications, 74(7):1661–1674, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303474.

Pluhacek:2013:BPC

[PSD^{+}13] Michal Pluhacek, Roman Senkerik, Donald Davendra, Zuzana Kominkova Oplatkova, and Ivan Zelinka. On the behavior and performance of chaos driven PSO algorithm with inertia weight. Computers and Mathematics with Applications, 66(2):122–134, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000333.

Papaschinopoulos:2010:MDE

[PSP10]  G.  Papaschinopoulos,

G. Stefanidou, and K. B. Papadopoulos. On a modification of a discrete epidemic model. Computers and Mathematics with Applications, 59(11):3559–3569, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002385.

Porcelli:2018:PPC

[PSS18] Margherita Porcelli, Valeria Simoncini, and Martin Stoll. Preconditioning PDE-constrained optimization with L^{1}-sparsity and control constraints. Computers and Mathematics with Applications, 74(5):1059–1075, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302778.

Plubtieng:2011:ESS

[PT11] Somyot Plubtieng and Tipphawan Thammathiwat. Existence of solutions of systems of generalized implicit vector quasi-equilibrium problems in G-convex spaces. Computers and Mathematics with Applications, 62(1):124–130, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111003622.

Pham:2015:SNS

[PT15] D. T. Pham and T. Tran.

Solving non-strongly elliptic pseudodifferential equations on a sphere using radial basis functions. Computers and Mathematics with Applications, 70(8):1970–1983, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003879.

Peng:2016:IIB

[PTH^{+}16] Cheng Peng, Yihua Teng, Brian Hwang, Zhaoli Guo, and Lian-Ping Wang. Implementation issues and benchmarking of lattice Boltzmann method for moving rigid particle simulations in a viscous flow. Computers and Mathematics with Applications, 72(2):349– 374, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500406X.

Pisica:2013:FSF

[PTL13] Ioana Pisica, Gareth Taylor, and Laurentiu Lipan. Feature selection filter for classification of power system operating states. Computers and Mathematics with Applications, 66(10):1795–1807, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004689.

Penenko:2014:VAE

[PTP14] V. V. Penenko, E. A.

Tsvetova, and A. V. Penenko. Variational approach and Euler’s integrating factors for environmental studies. Computers and Mathematics with Applications, 67(12):2240–2256, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001576.

Peng:2019:BWR

[PTZ19] Wei-Qi Peng, Shou-Fu Tian, and Tian-Tian Zhang. Breather waves and rational solutions in the (3 + 1)-dimensional Boiti– Leon–Manna–Pempinelli equation. Computers and Mathematics with Applications, 77 (3):715–723, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305959.

Paramanathan:2010:FIK

[PU10] P. Paramanathan and R. Uthayakumar. Fractal interpolation on the Koch curve. Computers and Mathematics with Applications, 59(10):3229– 3233, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001689.

Pu:2012:LWP

[Pu12]  Xueke Pu.  On the

local well-posedness for the fractional Landau– Lifshitz–Gilbert equation. Computers and Mathematics with Applications, 64(10):3276–3280, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000855.

Pultarova:2016:HPS

[Pul16] Ivana Pultarová. Hierarchical preconditioning for the stochastic Galerkin method: Upper bounds to the strengthened CBS constants. Computers and Mathematics with Applications, 71(4):949–964, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000092.

Pelayo:2012:NMD

[PV12] Fernando L. Pelayo and Jose C. Valverde. Notes on “Modeling the dynamics of concurrent computing systems”. Computers and Mathematics with Applications, 64(4):661– 663, August 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111011515.

See [GPV11].

Pelevic:2013:NIH

[PvdM13] Nikola Pelevic and Theo van der Meer. Numerical investigation of heat transfer enhancement by carbon nano fibers deposited on a flat plate. Computers and Mathematics with Applications, 65(6):914–923, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000576.

Peng:2010:PPS

[PW10] Zheng Peng and Donghua Wu. A partial parallel splitting augmented Lagrangian method for solving constrained matrix optimization problems. Computers and Mathematics with Applications, 60(6):1515–1524, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004499.

Pang:2011:FST

[PW11] Lisha Pang and Ke Wang. Function series theory of time scales. Computers and Mathematics with Applications, 62 (9):3427–3437, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007425.

Peng:2018:GSL

[PW18] Wei Peng and Hongxia Wang. A general scheme for log-determinant computation of matrices via stochastic polynomial approximation. Computers and Mathematics with Applications, 75(4):1259–1271, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307010.

Peng:2010:NBM

[PXT10] Dezhong Peng, Yong Xiang, and Hieu Trinh. A new blind method for separating M + 1 sources from M mixtures. Computers and Mathematics with Applications, 60(7):1829–1839, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004876.

Pan:2016:IAB

[PXXZ16] Qing Pan, Guoliang Xu, Gang Xu, and Yongjie Zhang. Isogeometric analysis based on extended Catmull–Clark subdivision. Computers and Mathematics with Applications, 71(1):105–119, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500543X.

Zang:2010:KCM

[pZ10] Qing pei Zang. A kind of complete moment convergence for self-normalized sums. Computers and Mathematics with Applications, 60(6):1803–1809, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004888.

Pan:2011:NPR

[PZ11a] Victor Y. Pan and Ai-Long Zheng. New progress in real and complex polynomial rootfinding. Computers and Mathematics with Applications, 61(5):1305–1334, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009867.

Pan:2011:RFE

[PZ11b] Victor Y. Pan and Ai-Long Zheng. Root-finding by expansion with independent constraints. Computers and Mathematics with Applications, 62(8):3164–3182, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006924.

Peng:2019:WPF

[PZA19] Li Peng, Yong Zhou, and Bashir Ahmad. The wellposedness for fractional nonlinear Schrödinger equations. Computers and Mathematics with Applications, 77(7):1998– 2005, April 1, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118306904.

Pawar:2019:JIS

[PZAR19] Aishwarya Pawar, Yongjie Jessica Zhang, Cosmin Anitescu, and Timon Rabczuk. Joint image segmentation and registration based on a dynamic level set approach using truncated hierarchical B-splines. Computers and Mathematics with Applications, 78 (10):3250–3267, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302421.

Pawar:2016:AFB

[PZJ^{+}16] Aishwarya Pawar, Yongjie Zhang, Yue Jia, Xiaodong Wei, Timon Rabczuk, Chiu Ling Chan, and Cosmin Anitescu. Adaptive FEM-based nonrigid image registration using truncated hierarchical Bsplines. Computers and Mathematics with Applications, 72(8):2028–2040, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302887.

Pan:2018:SMT

[PZL^{+}18] Mingyang Pan, Liancun Zheng, Chunyan Liu, Fawang Liu, Ping Lin, and Goong Chen. A stochastic model for thermal transport of nanofluid in porous media: Derivation and applications. Computers and Mathematics with Applications, 75(4):1226–1236, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306806.

Pan:2018:IGK

[PZZ18] Dongxin Pan, Chengwen Zhong, and Congshan Zhuo. An implicit gas-kinetic scheme for turbulent flow on unstructured hybrid mesh. Computers and Mathematics with Applications, 75(11):3825–3848, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301135.

Qasim:2016:SSN

[QAA^{+}16] Sabahat Qasim, Zulifqar Ali, Fayyaz Ahmad, S. Serra-Capizzano, Malik Zaka Ullah, and Arshad Mahmood. Solving systems of nonlinear equations when the nonlinearity is expensive. Computers and Mathematics with Applications, 71(7):1464–1478, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300724.

Qiu:2018:WPI

[QaY18] Hua Qiu and Zheng an Yao. On the wellposedness of the ideal incompressible viscoelastic flow in the critical Besov spaces. Computers and Mathematics with Applications, 76(2):257–275, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302153.

Qu:2015:FMA

[QCG15] Wenzhen Qu, Wen Chen, and Yan Gu. Fast multipole accelerated singular boundary method for the 3D Helmholtz equation in low frequency regime. Computers and Mathematics with Applications, 70(4):679–690, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002400.

Qin:2017:CGN

[QCLC17] Fangfang Qin, Jinru Chen, Zhilin Li, and Mingchao Cai. A Cartesian grid nonconforming immersed finite element method for planar elasticity interface problems. Computers and Mathematics with Applications, 73(3):404–418, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306654.

Qu:2019:NAS

[QCS^{+}19] Feng Qu, Jiaojiao Chen, Di Sun, Junqiang Bai, and Chao Yan. A new all-speed flux scheme for the Euler equations. Computers and Mathematics with Applications, 77(4):1216–1231, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306540.

Qin:2017:ENE

[QCT17] Dongdong Qin, Jing Chen, and XianHua Tang. Existence and nonexistence of nontrivial solutions for Schrödinger systems via Nehari– Pohozaev manifold. Computers and Mathematics with Applications, 74 (12):3141–3160, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304881.

Qian:2012:ANS

[QCYL12] Zhuzhong Qian, Ce Chen, Ilsun You, and Sanglu Lu. ACSP: a novel security protocol against counting attack for UHF RFID systems. Computers and Mathematics with Applications, 63(2):492–500, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006936.

Qu:2013:DPA

[QGGL13] Bo Qu, Dawu Gu, Zheng Guo, and Junrong Liu. Differential power analysis of stream ciphers with LFSRs. Computers and Mathematics with Applications, 65(9):1291– 1299, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001381.

Geng:2011:IAC

[qGpWhL11] Jin qiang Geng, Li ping Weng, and Si hong Liu. An improved ant colony optimization algorithm for nonlinear resource-leveling problems. Computers and Mathematics with Applications, 61(8):2300–2305, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007558.

Qiang:2011:GHS

[QH11] Hua Qiang and Zhicheng Hu. Generalizations of Hölder’s and some related inequalities. Computers and Mathematics with Applications, 61(2):392– 396, January 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110008643.

Qin:2016:GSS

[QHT16] Dongdong Qin, Yubo He, and Xianhua Tang. Ground state solutions for Kirchhoff type equations with asymptotically 4-linear nonlinearity. Computers and Mathematics with Applications, 71(7):1524–1536, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300918.

Qin:2011:CHP

[QHW11] Xiaolong Qin, Shuechin Huang, and Tianze Wang. On the convergence of hybrid projection algorithms for asymptotically quasi-ϕ-nonexpansive mappings. Computers and Mathematics with Applications, 61(4):851–859, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009387.

Qiu:2012:EUC

[Qiu12]  Zheyong Qiu.  Exis-

tence and uniqueness of common fixed points for two multivalued operators in ordered metric spaces. Computers and Mathematics with Applications, 63(8):1279–1286, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010765.

Jiang:2012:IFA

[qJhY12] Shen qing Jiang and Cong hua Yan. L-intuitionistic fuzzy σ-algebras. Computers and Mathematics with Applications, 64 (6):1849–1865, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001873.

Qi:2019:ICI

[QKR19] Jiaxing Qi, Harald Klimach, and Sabine Roller.

Implementation of the compact interpolation within the octree based Lattice Boltzmann solver Musubi. Computers and Mathematics with Applications, 78(4):1131–1141, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303571.

Qiu:2010:NSG

[QL10] Yang-Qing Qiu and Li-Wei Liu. A new system of generalized quasi-variational-like inclusion in Hilbert spaces. Computers and Mathematics with Applications, 59(1):1–8, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005045.

Qin:2018:MAD

[QLT^{+}18] Shanlin Qin, Fawang Liu, Ian W. Turner, Qianqian Yang, and Qiang Yu. Modelling anomalous diffusion using fractional Bloch–Torrey equations on approximate irregular domains. Computers and Mathematics with Applications, 75(1):7–21, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305242.

Qiu:2019:TGM

[QM19] Hailong Qiu and Liquan Mei. Two-grid MFEAs for the incompressible Stokes type variational inequality with damping. Computers and Mathematics with Applications, 78(8):2772–2788, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302342.

Qiu:2018:BHD

[QMW18] Shuyan Qiu, Chunlai Mu, and Liangchen Wang. Boundedness in the higherdimensional quasilinear chemotaxis-growth system with indirect attractant production. Computers and Mathematics with Applications, 75(9):3213–3223, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300580.

Qian:2018:RSR

[QRMH18] Chao Qian, Jiguang Rao, Dumitru Mihalache, and Jingsong He. Rational and semi-rational solutions of the y-nonlocal Davey– Stewartson I equation. Computers and Mathematics with Applications, 75(9):3317–3330, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300658.

Qin:2018:RWB

[QTW^{+}18] Chun-Yan Qin, Shou-Fu Tian, Xiu-Bin Wang, Tian-Tian Zhang, and Jin Li. Rogue waves, brightdark solitons and traveling wave solutions of the (3 + 1)-dimensional generalized Kadomtsev– Petviashvili equation. Computers and Mathematics with Applications, 75(12):4221–4231, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301597.

Qiu:2015:SEB

[QWGJ15] Ruo-Fan Qiu, An-Lin Wang, Qi-Wei Gong, and Tao Jiang. Simulation of expanding bubble through a hole in a channel driven by pressure using lattice Boltzmann method. Computers and Mathematics with Applications, 70(3):244–253, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002047.

Qiu:2019:MSB

[QWL19] Lin Qiu, Fajie Wang, and Ji Lin. A meshless singular boundary method for transient heat conduction problems in layered materials. Computers and Mathematics with Applications, 78 (11):3544–3562, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302901.

Qiao:2019:BAO

[QX19] Leijie Qiao and Da Xu. BDF ADI orthogonal spline collocation scheme for the fractional integrodifferential equation with two weakly singular kernels. Computers and Mathematics with Applications, 78(12):3807– 3820, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122119303050.

Qi:2013:CTT

[QXG13] Hai-Tao Qi, Huan-Ying Xu, and Xin-Wei Guo. The Cattaneo-type time fractional heat conduction equation for laser heating. Computers and Mathematics with Applications, 66(5):824–831, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006827.

See notes [XWH16].

Xia:2011:IHP

[qXjH11] Fu quan Xia and Nan jing Huang. An inexact hybrid projection-proximal point algorithm for solving generalized mixed variational inequalities. Computers and Mathematics with Applications, 62(12):4596–4604, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111009060.

Qin:2011:ISB

[QXLL11] Kun Qin, Kai Xu, Feilong Liu, and Deyi Li. Image segmentation based on histogram analysis utilizing the cloud model. Computers and Mathematics with Applications, 62(7):2824–2833, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006195.

Qamar:2013:ADG

[QY13] Shamsul Qamar and Muhammad Yousaf. Application of a discontinuous Galerkin finite element method to special relativistic hydrodynamic models. Computers and Mathematics with Applications, 65(8):1220–1232, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001430.

Qiu:2017:WPD

[QY17] Hua Qiu and Zheng’an Yao. Well-posedness for density-dependent Boussinesq equations without dissipation terms in Besov spaces. Computers and Mathematics with Applications, 73(9):1920–1931, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301207.

Qi:2010:UPM

[QYL10] Xiao-Guang Qi, Lian-Zhong Yang, and Kai Liu. Uniqueness and periodicity of meromorphic functions concerning the difference operator. Computers and Mathematics with Applications, 60(6):1739–1746, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004797.

Qi:2011:EPR

[QZ11] Xuli Qi and Bo Zhou.

Extremal properties of reciprocal complementary Wiener number of trees. Computers and Mathematics with Applications, 62(1):523–531, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004342.

Qian:2016:EEP

[QZ16] Yanxia Qian and Tong Zhang. On error estimates of the projection method for the timedependent natural convection problem: First order scheme. Computers and Mathematics with Applications, 72 (5):1444–1465, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304059.

Qiao:2019:CIR

[QZF19] Yuanyang Qiao, Jianping Zhao, and Xinlong Feng. A compact integrated RBF method for time fractional convection– diffusion-reaction equations. Computers and Mathematics with Applications, 77(9):2263–2278, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307090.

Zhao:2017:MLK

[qZM17a] Hai qiong Zhao and Wen-Xiu Ma. Mixed lumpkink solutions to the KP equation. Computers and Mathematics with Applications, 74 (6):1399–1405, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303826.

Qiu:2017:MFN

[QZM17b] Hailong Qiu, Yongchao Zhang, and Liquan Mei. A Mixed-FEM for Navier– Stokes type variational inequality with nonlinear damping term. Computers and Mathematics with Applications, 73(10):2191–2207, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301414.

Qin:2011:NCL

[QZY11] Xu Qin, Jiang-She Zhang, and Xiao-Dong Yan. A nonparametric circularlinear multivariate regression model with a ruleof-thumb bandwidth selector. Computers and Mathematics with Applications, 62(8):3048–3055, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006791.

Rezapour:2011:SFP

[RA11a] Sh. Rezapour and P. Amiri. Some fixed point results for multivalued operators in generalized metric spaces. Computers and Mathematics with Applications, 61(9):2661– 2666, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001660.

Russo:2011:PEE

[RA11b] Anahí Dello Russo and Ana E. Alonso. A posteriori error estimates for nonconforming approximations of Steklov eigenvalue problems. Computers and Mathematics with Applications, 62 (11):4100–4117, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100842X.

Roodaki:2012:DBF

[RA12] M. Roodaki and H. Almasieh. Delta basis functions and their applications to systems of integral equations. Computers and Mathematics with Applications, 63(1):100–109, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009631.

Rasheed:2018:NCF

[RA18] Amer Rasheed and Muhammad Shoaib Anwar. Numerical computations of fractional nonlinear Hartmann flow with revised heat flux model. Computers and Mathematics with Applications, 76 (10):2421–2433, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304656.

Ru:2019:GWP

[RA19] Shaolei Ru and Muhammad Zainul Abidin. Global well-posedness of the incompressible fractional Navier–Stokes equations in Fourier–Besov spaces with variable exponents. Computers and Mathematics with Applications, 77(4):1082–1090, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306369.

Rashedi:2013:ARG

[RAD13] Kamal Rashedi, Hojatollah Adibi, and Mehdi Dehghan. Application of the Ritz–Galerkin method for recovering the spacewise-coefficients in the wave equation. Computers and Mathematics with Applications, 65(12):1990–2008, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001922.

Rafiei:2012:CPS

[Raf12]  A. Rafiei.  A com-

plete pivoting strategy for the right-looking robust incomplete factorization preconditioner. Computers and Mathematics with Applications, 64(8):2682–2694, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122112005123.

Rahmat:2011:LTD

[Rah11a] Mohamad Rafi Segi Rahmat. The (q, h)–Laplace transform on discrete time scales. Computers and Mathematics with Applications, 62(1):272– 281, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003968.

Rahmat:2011:SAI

[Rah11b] Mohamad Rafi Segi Rahmat. On some (q, h)- analogues of integral inequalities on discrete time scales. Computers and Mathematics with Applications, 62(4):1790–1797, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005013.

Ramadan:2011:FIS

[Ram11] A. A. Ramadan. Lfuzzy interior systems. Computers and Mathematics with Applications, 62(12):4301–4307, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007759.

Rana:2015:CNS

[Ran15] Puneet Rana. Corrigendum to “Numerical solution for mixed convection boundary layer flow of a nanofluid along an inclined plate embedded in a porous medium” [Comput. Math. Appl. 64 (2012) 2816–2832]. Computers and Mathematics with Applications, 69(12): 1518, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001467.

See [RBB12].

Reuter:2016:FMG

[RAW^{+}16] Balthasar Reuter, Vadym Aizinger, Manuel Wieland, Florian Frank, and Peter Knabner. FESTUNG: a MATLAB/GNU Octave toolbox for the discontinuous Galerkin method, Part II: Advection operator and slope limiting. Computers and Mathematics with Applications, 72(7):1896–1925, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304606.

Ray:2016:NES

[Ray16] S. Saha Ray. New exact solutions of nonlinear fractional acoustic wave equations in ultrasound. Computers and Mathematics with Applications, 71(3):859–868, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000043.

Ray:2017:CLL

[Ray17]  S. Saha Ray.  On

conservation laws by Lie symmetry analysis for (2 + 1)-dimensional Bogoyavlensky-Konopelchenko equation in wave propagation. Computers and Mathematics with Applications, 74(6):1158–1165, September 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117303553.

Ray:2018:IACb

[Ray18] S. Saha Ray. Invariant analysis and conservation laws for the time fractional (2 + 1)- dimensional Zakharov– kuznetsov modified equal width equation using Lie group analysis. Computers and Mathematics with Applications, 76 (9):2110–2118, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304267.

Rabbani:2019:TPS

[RAZ19] Omar Rabbani, Munshoor Ahmed, and Saqib Zia. Transport of pollutant in shallow flows: a space–time CE/SE scheme. Computers and Mathematics with Applications, 77(12):3195– 3211, June 15, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122119300781.

Reznicek:2019:IVD

[RB19] Hynek Reznícek and Ludek Benes. Impact of vegetation on dustiness produced by surface coal mine in North Bohemia. Computers and Mathematics with Applications, 78(9):3175–3186, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302743.

Rana:2012:NSM

[RBB12] Puneet Rana, R. Bhargava, and O. A. Bég. Numerical solution for mixed convection boundary layer flow of a nanofluid along an inclined plate embedded in a porous medium. Computers and Mathematics with Applications, 64 (9):2816–2832, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003549.

See corrigendum [Ran15].

Roberts:2014:DMS

[RBTD14] Nathan V. Roberts, Tan Bui-Thanh, and Leszek Demkowicz. The DPG method for the Stokes problem. Computers and Mathematics with Applications, 67(4):966–995, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113007116.

Roberts:2017:GMP

[RC17a] Nathan V. Roberts and Jesse Chan. A geometric multigrid preconditioning strategy for DPG system matrices. Computers and Mathematics with Applications, 74(8):2018–2043, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304133.

Ru:2017:GSC

[RC17b] Shaolei Ru and Jiecheng Chen. Global solution of the critical Burgers equation in N dimensions. Computers and Mathematics with Applications, 74(2):325–335, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302535.

Ruter:2018:ESE

[RC18] Marcus Olavi Rüter and Jiun-Shyan Chen. An enhanced-strain error estimator for Galerkin meshfree methods based on stabilized conforming nodal integration. Computers and Mathematics with Applications, 74 (9):2144–2171, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304108.

Rossi:2015:NSV

[RCG15] E. Rossi, A. Colagrossi, and G. Graziani. Numerical simulation of 2Dvorticity dynamics using particle methods. Computers and Mathematics with Applications, 69(12):1484–1503, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001832.

Rebollo:2019:PGM

[RCH19] Tomás Chacón Rebollo, Daniel Franco Coronil, and Frédéric Hecht. A Petrov–Galerkin multilayer discretization to second order elliptic boundary value problems. Computers and Mathematics with Applications, 77(12):3068–3086, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122119300586.

Romao:2011:AGL

[RCM11] E. C. Romão, M. D. Campos, and L. F. M. Moura. Application of the Galerkin and least-squares finite element methods in the solution of 3D Poisson and Helmholtz equations. Computers and Mathematics with Applications, 62(11):4288–4299, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008868.

Rodriguez:2014:PEE

[RCRV14] A. Alonso Rodríguez, J. Camaño, R. Rodríguez, and A. Valli. A posteriori error estimates for the problem of electrostatics with a dipole source. Computers and Mathematics with Applications, 68(4):464–485, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002788.

Riedlbeck:2017:SFR

[RDE^{+}17] Rita Riedlbeck, Daniele A.

Di Pietro, Alexandre Ern, Sylvie Granet, and Kyrylo Kazymyrenko. Stress and flux reconstruction in Biot’s poroelasticity problem with application to a posteriori error analysis. Computers and Mathematics with Applications, 73(7):1593–1610, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300718.

Reis:2019:MMI

[RdSRL19] Ruy Freitas Reis, Rodrigo Weber dos Santos, Bernardo Martins Rocha, and Marcelo Lobosco. On the mathematical modeling of inflammatory edema formation. Computers and Mathematics with Applications, 78(9):2994–3006, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301890.

Rautenbach:2011:STH

[RdSSS11] Dieter Rautenbach, Vinícius Fernandes dos Santos, Philipp M. Schäfer, and Jayme L. Szwarcfiter. On subbetweennesses of trees: Hardness, algorithms, and characterizations. Computers and Mathematics with Applications, 62(12):4674–4681, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009321.

Rashad:2011:NCB

[REHA11] A. M. Rashad, M. A. El-Hakiem, and M. M. M. Abdou. Natural convection boundary layer of a non-Newtonian fluid about a permeable vertical cone embedded in a porous medium saturated with a nanofluid. Computers and Mathematics with Applications, 62(8):3140–3151, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006900.

Ramadan:2010:PEA

[RES10] Mohamed A. Ramadan and Ehab A. El-Sayed. Partial eigenvalue assignment problem of high order control systems using orthogonality relations. Computers and Mathematics with Applications, 59(6):1918–1928, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005835.

Respondek:2016:INR

[Res16] Jerzy S. Respondek.

Incremental numerical recipes for the high efficient inversion of the confluent Vandermonde matrices. Computers and Mathematics with Applications, 71(2):489–502, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005842.

Reynolds:2012:ACL

[Rey12] David W. Reynolds. On asymptotic constancy for linear discrete summation equations. Computers and Mathematics with Applications, 64(7):2335–2344, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004129.

Rong:2012:CPB

[RF12] Aiying Rong and José Rui Figueira. Computational performance of basic state reduction based dynamic programming algorithms for bi-objective 0-1 knapsack problems. Computers and Mathematics with Applications, 63(10):1462–1480, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002623.

Rashidinia:2016:SME

[RFK16] J. Rashidinia, G. E. Fasshauer, and M. Khasi. A stable method for the evaluation of Gaussian radial basis function solutions of interpolation and collocation problems. Computers and Mathematics with Applications, 72(1):178– 193, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302413.

Rong:2011:TSR

[RFP11] Aiying Rong, José Rui Figueira, and Margarida Vaz Pato. A two state reduction based dynamic programming algorithm for the bi-objective 0-1 knapsack problem. Computers and Mathematics with Applications, 62(8):2913–2930, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006420.

Ren:2011:HFF

[RG11] Zhong-Fu Ren and Wei-Kui Gui. He’s frequency formulation for nonlinear oscillators using a golden mean location. Computers and Mathematics with Applications, 61(8):1987–1990, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006218.

Ray:2018:TDL

[RG18] S. Saha Ray and A. K.

Gupta. Two-dimensional Legendre wavelet method for travelling wave solutions of time-fractional generalized seventh order KdV equation. Computers and Mathematics with Applications, 73(6):1118–1133, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303972.

Ruiz-Garzon:2010:SRB

[RGdSRLAJ10]  Gabriel  Ruiz-Garzón,

Lucelina Batista dos Santos, Antonio Rufián-Lizana, and Manuel Arana-Jiménez. Some relations between Minty variational-like inequality problems and vectorial optimization problems in Banach spaces. Computers and Mathematics with Applications, 60(9):2679–2688, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000698X.

Rodrigo:2015:FEF

[RGHZ15] C. Rodrigo, F. J. Gaspar, X. Hu, and L. Zikatanov. A finite element framework for some mimetic finite difference discretizations. Computers and Mathematics with Applications, 70(11):2661– 2673, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0898122115003478.

Rueda-Gomez:2017:RBM

[RGVR17] D. A. Rueda-Gómez and E. J. Villamizar-Roa. On the Rayleigh–Bénard– Marangoni system and a related optimal control problem. Computers and Mathematics with Applications, 74 (12):2969–2991, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304674.

Ramiere:2015:IRB

[RH15] Isabelle Ramière and Thomas Helfer. Iterative residual-based vector methods to accelerate fixed point iterations. Computers and Mathematics with Applications, 70(9):2210–2226, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004046.

Roy:2015:BIF

[RHC15] Saswati Roy, Luca Heltai, and Francesco Costanzo. Benchmarking the immersed finite element method for fluid-structure interaction problems. Computers and Mathematics with Applications, 69(10): 1167–1188, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115001078.

Rojas:2018:ERG

[RHD18] S. Rojas, R. Hein, and M. Durán. On an equivalent representation of the Green’s function for the Helmholtz problem in a non-absorbing impedance half-plane. Computers and Mathematics with Applications, 75(11):3903–3917, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301305.

Rheinlander:2010:SSL

[Rhe10] Martin Rheinländer. On the stability structure for lattice Boltzmann schemes. Computers and Mathematics with Applications, 59(7):2150–2167, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122109006282.

Rahman:2018:GES

[RHMA18] S. Rahman, T. Hayat, M. Muneer, and B. Ahmad. Global existence of solutions for MHD third grade flow equations saturating porous medium. Computers and Mathematics with Applications, 76(10):2360–2374, November 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118304504.

Rouzkard:2012:SCF

[RI12] Fayyaz Rouzkard and Mohammad Imdad. Some common fixed point theorems on complex valued metric spaces. Computers and Mathematics with Applications, 64 (6):1866–1874, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001885.

Rasool:2012:MAD

[RIW12] Ghulam Rasool, Kamran Iqbal, and Gannon A. White. Myoelectric activity detection during a sit-to-stand movement using threshold methods. Computers and Mathematics with Applications, 64(5):1473–1483, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002994.

Ross-Jones:2019:CHT

[RJGS^{+}19] Jesse Ross-Jones, Maximilian Gaedtke, Sebastian Sonnick, Matthias Rädle, Hermann Nirschl, and Mathias J. Krause. Conjugate heat transfer through nano scale porous media to optimize vacuum insulation panels with lattice Boltzmann methods. Computers and Mathematics with Applications, 77(1):209–221, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305352.

Radenovic:2010:GWC

[RK10] Stojan Radenović and Zoran Kadelburg. Generalized weak contractions in partially ordered metric spaces. Computers and Mathematics with Applications, 60(6):1776–1783, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004839.

Rahimi:2018:LBM

[RKA^{+}18] Alireza Rahimi, Abbas Kasaeipoor, Ali Amiri, Mohammad Hossein Doranehgard, Emad Hasani Malekshah, and Lioua Kolsi. Lattice Boltzmann method based on Dual-MRT model for three-dimensional natural convection and entropy generation in CuO– water nanofluid filled cuboid enclosure included with discrete active walls. Computers and Mathematics with Applications, 75(5):1795–1813, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307630.

Radtke:2018:IGI

[RKD18] Lars Radtke, Marcel König, and Alexander Düster. The influence of geometric imperfections in cardiovascular FSI simulations. Computers and Mathematics with Applications, 74(7):1675–1689, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730233X.

Rashidinia:2018:SGR

[RKF18] J. Rashidinia, M. Khasi, and G. E. Fasshauer. A stable Gaussian radial basis function method for solving nonlinear unsteady convection–diffusion– reaction equations. Computers and Mathematics with Applications, 75(5):1831–1850, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307708.

Rad:2012:NSN

[RKP12] J. A. Rad, S. Kazem, and K. Parand. A numerical solution of the nonlinear controlled Duffing oscillator by radial basis functions. Computers and Mathematics with Applications, 64(6):2049–2065, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003227.

Radha:2018:DLD

[RKSA18] R. Radha, C. Senthil Kumar, K. Subramanian, and T. Alagesan. Drone like dynamics of dromion pairs in the (2 + 1) AKNS equation. Computers and Mathematics with Applications, 75(7):2356–2364, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307794.

Reinecke:2012:CBF

[RKW12] Philipp Reinecke, Tilman Krauß, and Katinka Wolter. Cluster-based fitting of phase-type distributions to empirical data. Computers and Mathematics with Applications, 64(12):3840–3851, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002210.

Raheem:2017:OCI

[RM17] A. Raheem and Md.

Maqbul. Oscillation criteria for impulsive partial fractional differential equations. Computers and Mathematics with Applications, 73(8):1781–1788, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300950.

Rafiq:2010:SNP

[RMA10] Arif Rafiq, Muhammad Yousaf Malik, and Tariq Abbasi. Solution of nonlinear pull-in behavior in electrostatic microactuators by using He’s homotopy perturbation method. Computers and Mathematics with Applications, 59(8):2723–2733, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000623.

Rajaraman:2014:EPI

[RMB^{+}14] Prathish K. Rajaraman, T. A. Manteuffel, M. Belohlavek, E. McMahon, and Jeffrey J. Heys. Echocardiographic particle imaging velocimetry data assimilation with least square finite element methods. Computers and Mathematics with Applications, 68 (11):1569–1580, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003447.

Rausch:2019:PSP

[RMK19] Alexander M. Rausch, Matthias Markl, and Carolin Körner. Predictive simulation of process windows for powder bed fusion additive manufacturing: Influence of the powder size distribution. Computers and Mathematics with Applications, 78(7):2351–2359, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303535.

Radwan:2011:SNS

[RMM11] A. G. Radwan, K. Moaddy, and Shaher Momani. Stability and non-standard finite difference method of the generalized Chua’s circuit. Computers and Mathematics with Applications, 62(3):961–970, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003506.

Rachid:2010:MFS

[RMS10] Mansouri Rachid, Bettayeb Maamar, and Djennoune Said. Multivariable fractional system approximation with initial conditions using integral state space representation. Computers and Mathematics with Applications, 59(5):1842–1851, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005628.

Razmjooy:2012:RTM

[RMS12] Navid Razmjooy, B. Somayeh Mousavi, and F. Soleymani. A realtime mathematical computer method for potato inspection using machine vision. Computers and Mathematics with Applications, 63(1):268–279, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009850.

Ren:2019:RST

[RMY19] Bo Ren, Wen-Xiu Ma, and Jun Yu. Rational solutions and their interaction solutions of the (2 + 1)-dimensional modified dispersive water wave equation. Computers and Mathematics with Applications, 77(8):2086–2095, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307028.

Ragulskis:2011:SSL

[RNB11] Minvydas Ragulskis, Zenonas Navickas, and Liepa Bikulciene. The solitary solution of the Liouville equation produced by the Exp-function method does not hold for all initial conditions. Computers and Mathematics with Applications, 62(1):376– 382, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004081.

Rivera:2013:DSM

[RNQ13] Jaime E. Muñoz Rivera, Maria Grazia Naso, and Ramon Quintanilla. Decay of solutions for a mixture of thermoelastic one dimensional solids. Computers and Mathematics with Applications, 66(1):41–55, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002253.

Rivera:2016:DSM

[RNQ16] Jaime E. Muñoz Rivera, Maria Grazia Naso, and Ramon Quintanilla. Decay of solutions for a mixture of thermoelastic solids with different temperatures. Computers and Mathematics with Applications, 71(4):991–1009, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116000146.

Rezvan:2010:ISI

[RÖ10] Farshad Rezvan and Teoman Özer. Invariant solutions of integrodifferential Vlasov–Maxwell equations in Lagrangian variables by Lie group analysis. Computers and Mathematics with Applications, 59(11):3412– 3437, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000218X.

Roberts:2014:CSF

[Rob14] Nathan V. Roberts. Camellia: a software framework for discontinuous Petrov– Galerkin methods. Computers and Mathematics with Applications, 68 (11):1581–1604, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004027.

Roshan:2012:PGM

[Ros12] Thoudam Roshan. A Petrov–Galerkin method for solving the generalized regularized long wave (GRLW) equation. Computers and Mathematics with Applications, 63(5):943–956, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211101042X.

Rachaniotis:2010:MOR

[RPTD10] N. P. Rachaniotis, C. P.

Pappis, G. T. Tsoulfas, and T. Dasaklis. A multi-objective reassembling policy model and its implementation in the case of a network of personal computers. Computers and Mathematics with Applications, 59(1):391–398, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004271.

Rehman:2018:HOF

[RQ18] Asad Rehman and Shamsul Qamar. High order finite-volume WENO scheme for five-equation model of compressible two-fluid flow. Computers and Mathematics with Applications, 76(11– 12):2648–2664, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305066.

Rudek:2011:NOD

[RR11] Agnieszka Rudek and Radoslaw Rudek. A note on optimization in deteriorating systems using scheduling problems with the aging effect and resource allocation models. Computers and Mathematics with Applications, 62(4):1870–1878, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005104.

Rooholghdos:2014:ESO

[RR14] Seyed Ali Rooholghdos and Ehsan Roohi. Extension of a second order velocity slip/temperature jump boundary condition to simulate high speed micro/nanoflows. Computers and Mathematics with Applications, 67(11):2029–2040, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001850.

Rathan:2018:MFO

[RR18] Samala Rathan and G. Naga Raju. A modified fifthorder WENO scheme for hyperbolic conservation laws. Computers and Mathematics with Applications, 75(5):1531–1549, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307344.

Rodriguez:2019:SRS

[RR19] Ana Alonso Rodríguez and Francesca Rapetti. Some remarks on spanning families and weights for high order Whitney spaces on simplices. Computers and Mathematics with Applications, 78 (9):2961–2972, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301257.

Reuter:2019:DGM

[RRAK19] Balthasar Reuter, Andreas Rupp, Vadym Aizinger, and Peter Knabner. Discontinuous Galerkin method for coupling hydrostatic free surface flows to saturated subsurface systems. Computers and Mathematics with Applications, 77(9):2291–2309, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307120.

Rakha:2011:SNC

[RRC11] Medhat A. Rakha, Arjun K. Rathie, and Purnima Chopra. On some new contiguous relations for the Gauss hypergeometric function with applications. Computers and Mathematics with Applications, 61(3):620–629, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009132.

Rivero:2010:FOS

[RRGTV10] Margarita Rivero, Luis Rodríguez-Germá, Juan J. Trujillo, and M. Pilar Velasco. Fractional operators and some special functions. Computers and Mathematics with Applications, 59(5):1822–1834, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005604.

Rouatbi:2017:NSM

[RRO17] Asma Rouatbi, Moeiz Rouis, and Khaled Omrani. Numerical scheme for a model of shallow water waves in (2 + 1)-dimensions. Computers and Mathematics with Applications, 74(8):1871–1884, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117304121.

Rosca:2016:LGS

[RRP16] Natalia C. Rosca, Alin V. Rosca, and Ioan Pop. Lie group symmetry method for MHD double-diffusive convection from a permeable vertical stretching/ shrinking sheet. Computers and Mathematics with Applications, 71(8):1679– 1693, April 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301225.

Raghavendar:2012:ITF

[RS12a] K. Raghavendar and A. Swaminathan. Integral transforms of functions to be in certain class defined by the combination of starlike and convex functions. Computers and Mathematics with Applications, 63(8):1296–1304, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011485.

Rai:2012:NSS

[RS12b] Pratima Rai and Kapil K. Sharma. Numerical study of singularly perturbed differential-difference equation arising in the modeling of neuronal variability. Computers and Mathematics with Applications, 63(1):118–132, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009655.

Rossikhin:2013:TAS

[RS13] Yu. A. Rossikhin and M. V. Shitikova. Two approaches for studying the impact response of viscoelastic engineering systems: an overview. Computers and Mathematics with Applications, 66(5):755–773, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000230.

Ravibabu:2014:RRV

[RS14a] Mashetti Ravibabu and Arindama Singh. On refined Ritz vectors and polynomial characterization. Computers and Mathematics with Applications, 67(5):1057–1064, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000170.

Ravnik:2014:IEF

[RS14b] J. Ravnik and L. Skerget. Integral equation formulation of an unsteady diffusion-convection equation with variable coefficient and velocity. Computers and Mathematics with Applications, 66(12):2477–2488, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005889.

Rebenda:2015:CAI

[RS15] Josef Rebenda and Zdenek Smarda. Convergence analysis of an iterative scheme for solving initial value problem for multidimensional partial differential equations. Computers and Mathematics with Applications, 70(8):1772–1780, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003533.

Ray:2018:IACa

[RS18a] S. Saha Ray and S. Sahoo. Invariant analysis and conservation laws of (2 + 1) dimensional time-fractional ZK–BBM equation in gravity water waves. Computers and Mathematics with Applications, 75(7):2271–2279, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307642.

Rui:2018:LFF

[RS18b] Hongxing Rui and Ming Sun. A locking-free finite difference method on staggered grids for linear elasticity problems. Computers and Mathematics with Applications, 76(6):1301–1320, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830347X.

Raza:2014:NAT

[RSB14] Nauman Raza, Sultan Sial, and Asma Rashid Butt. Numerical approximation of time evolution related to Ginzburg– Landau functionals using weighted Sobolev gradients. Computers and Mathematics with Applications, 67(1):210–216, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006627.

Rajovic:2011:PSO

[RSDR11] Miloje Rajović, Rade Stojiljković, Dragan Dimitrovski, and Dragana Radosavljević. Perturbation of solutions of ordinary linear homogeneous differential equations of the second order. Computers and Mathematics with Applications, 62(5):2330–2335, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005736.

Repin:2018:CML

[RSH18] Sergey Repin, Stanislav Sysala, and Jaroslav Haslinger. Computable majorants of the limit load in Hencky’s plasticity problems. Computers and Mathematics with Applications, 75(1):199–217, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305552.

Rahimi:2018:EGA

[RSL^{+}18] Alireza Rahimi, Mohammad Sepehr, Milad Janghorban Lariche, Abbas Kasaeipoor, Emad Hasani Malekshah, and Lioua Kolsi. Entropy generation analysis and heatline visualization of free convection in nanofluid (KKL model-based)-filled cavity including internal active fins using lattice Boltzmann method. Computers and Mathematics with Applications, 75(5):1814–1830, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730771X.

Ragb:2017:IDQ

[RSM17] Ola Ragb, L. F. Seddek, and M. S. Matbuly. Iterative differential quadrature solutions for Bratu problem. Computers and Mathematics with Applications, 74(2):249–257, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302080.

Rossikhin:2018:DRV

[RSP18] Yu.A. Rossikhin, M. V. Shitikova, and I. I. Popov. Dynamic response of a viscoelastic beam impacted by a viscoelastic sphere. Computers and Mathematics with Applications, 73(6):970–984, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211630267X.

Rossikhin:2010:AFV

[RSS10] Yu. A. Rossikhin, M. V. Shitikova, and T. A. Shcheglova. Analysis of free vibrations of a viscoelastic oscillator via the models involving several fractional parameters and relaxation/retardation times. Computers and Mathematics with Applications, 59(5):1727–1744, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005501.

Rafsanjani:2016:EDC

[RSS16] Hossein Khodabakhshi Rafsanjani, Mohammad Hossein Sedaaghi, and Saeid Saryazdi. Efficient diffusion coefficient for image denoising. Computers and Mathematics with Applications, 72(4):893–903, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303376.

Rajagopal:2011:LSH

[RSV11] K. R. Rajagopal, G. Saccomandi, and L. Vergori. Linear stability of Hagen–Poiseuille flow in a chemically reacting fluid. Computers and Mathematics with Applications, 61(2):460–469, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008989.

Rachunkova:2010:NPT

[RSWZ10] Irena Rachunková, Svatoslav Stanek, Ewa Weinmüller, and Michael Zenz. Neumann problems with time singularities. Computers and Mathematics with Applications, 60(3):722– 733, August 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003688.

Robbiano:2010:ARR

[RT10] María Robbiano and Vilmar Trevisan. Applications of recurrence relations for the characteristic polynomials of Bethe trees. Computers and Mathematics with Applications, 59(9):3039– 3044, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000132X.

Rashidi:2011:CIA

[RT11] Hassan Rashidi and Edward P. K. Tsang. A complete and an incomplete algorithm for automated guided vehicle scheduling in container terminals. Computers and Mathematics with Applications, 61(3):630–641, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009144.

Rafiei:2014:CPS

[RTB14] Amin Rafiei, Behnaz Tolue, and Matthias Bollhöfer. Complete pivoting strategy for the leftlooking robust incomplete factorization preconditioner. Computers and Mathematics with Applications, 67(11):2055– 2070, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001734.

Rohan:2019:BDB

[RTL19] Eduard Rohan, Jana Turjanicová, and Vladimír Lukes. The Biot–Darcy– Brinkman model of flow in deformable double porous media; homogenization and numerical modelling. Computers and Mathematics with Applications, 78 (9):3044–3066, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302020.

Rezavand:2018:ISN

[RTRR18] Massoud Rezavand, Mohammad Taeibi-Rahni, and Wolfgang Rauch. An ISPH scheme for numerical simulation of multiphase flows with complex interfaces and high density ratios. Computers and Mathematics with Applications, 75(8):2658–2677, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300038.

Riviere:2017:EAP

[RTT17] Béatrice Rivière, Jun Tan, and Travis Thompson. Error analysis of primal discontinuous Galerkin methods for a mixed formulation of the Biot equations. Computers and Mathematics with Applications, 73(4):666–683, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730007X.

Rodriguez:2017:NMD

[RTV17] J. M. Rodríguez and R. Taboada-Vázquez. A new LES model derived from generalized Navier–Stokes equations with nonlinear viscosity. Computers and Mathematics with Applications, 73(2):294–303, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306538.

Ruas:2019:AEN

[Rua19]  Vitoriano Ruas.  Ac-

curacy enhancement for non-isoparametric finiteelement simulations in curved domains; application to fluid flow. Computers and Mathematics with Applications, 77(6):1756–1769, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302979.

Russo:2016:CID

[Rus16] Alessandro Russo. On the choice of the internal degrees of freedom for the nodal virtual element method in two dimensions. Computers and Mathematics with Applications, 72(8):1968–1976, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116301328.

Ren:2019:GTS

[RWTW19] Huan Ren, Xiang Wang, Xiao-Bin Tang, and Teng Wang. The general two-sweep modulus-based matrix splitting iteration method for solving linear complementarity problems. Computers and Mathematics with Applications, 77(4):1071–1081, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306370.

Ren:2018:CSY

[RWW18] Huan Ren, Xiang Wang, and Teng Wang. Commuting solutions of the Yang–Baxter-like matrix equation for a class of rank-two updated matrices. Computers and Mathematics with Applications, 76 (5):1085–1098, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303183.

Ren:2013:PPK

[RWZ13] Yanli Ren, Shuozhong Wang, and Xinpeng Zhang. Practical parallel key-insulated encryption with multiple helper keys. Computers and Mathematics with Applications, 65(9):1403–1412, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000429.

Raftari:2010:AHP

[RY10] Behrouz Raftari and Ahmet Yildirim. The application of homotopy perturbation method for MHD flows of UCM fluids above porous stretching sheets. Computers and Mathematics with Applications, 59(10):3328– 3337, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000194X.

Raftari:2011:SSN

[RY11] Behrouz Raftari and Ahmet Yildirim. Series solution of a nonlinear ODE arising in magnetohydrodynamic by HPM-padé technique.

Computers and Mathematics with Applications, 61(6):1676–1681, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000630.

Rafiee:2013:TBB

[RYK13] M. Rafiee, Jie Yang, and Siritiwat Kitipornchai. Thermal bifurcation buckling of piezoelectric carbon nanotube reinforced composite beams. Computers and Mathematics with Applications, 66(7):1147–1160, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300429X.

Rylko:2015:EEH

[Ryl15a] Natalia Rylko. Edge effects for heat flux in fibrous composites. Computers and Mathematics with Applications, 70 (10):2283–2291, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003752.

Rylko:2015:FLF

[Ryl15b] Natalia Rylko. Fractal local fields in random composites. Computers and Mathematics with Applications, 69(3):247–254, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005926.

Ran:2016:CDS

[RZ16] Maohua Ran and Chengjian Zhang. Compact difference scheme for a class of fractional-in-space nonlinear damped wave equations in two space dimensions. Computers and Mathematics with Applications, 71(5):1151–1162, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300153.

Rachowicz:2017:AMF

[RZ17] Waldemar Rachowicz and Adam Zdunek. An hadaptive mortar finite element method for finite deformation contact with higher order p extension. Computers and Mathematics with Applications, 73(8):1834–1854, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301013.

Rao:2011:IMA

[RZL11] Congjun Rao, Yong Zhao, and Chuanfeng Li. Incentive mechanism for allocating total permitted pollution discharge capacity and evaluating the validity of free allocation. Computers and Mathematics with Applications, 62(8):3037–3047, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100678X.

Ruan:2019:NST

[RZZ19] Zhousheng Ruan, Sen Zhang, and Wen Zhang. Numerical solution of time-dependent component with sparse structure of source term for a time fractional diffusion equation. Computers and Mathematics with Applications, 77(5):1408–1422, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118306655.

Sezgin:2011:OSS

[SA11a] Aslihan Sezgin and Akin Osman Atagün. On operations of soft sets. Computers and Mathematics with Applications, 61(5):1457–1467, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000356.

Sezgin:2011:SGN

[SA11b] Aslihan Sezgin and Akin Osman Atagün. Soft groups and normalistic soft groups. Computers and Mathematics with Applications, 62(2):685– 698, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004482.

Sharifnia:2016:CAM

[SA16] Mahdi Sharifnia and Alireza Akbarzadeh. A constrained assumed modes method for solution of a new dynamic equation for an axially moving beam. Computers and Mathematics with Applications, 72(9):2167–2180, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304722.

Saeidifar:2011:AWF

[Sae11] A. Saeidifar. Application of weighting functions to the ranking of fuzzy numbers. Computers and Mathematics with Applications, 62(5):2246–2258, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005669.

Sagaut:2010:TAS

[Sag10] Pierre Sagaut. Toward advanced subgrid models for lattice-Boltzmann-based large-eddy simulation: Theoretical formulations. Computers and Mathematics with Applications, 59(7):2194–2199, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006385.

Sahoo:2011:ESS

[Sah11] Bikash Sahoo. Effects of slip on sheet-driven flow and heat transfer of a non-Newtonian fluid past a stretching sheet. Computers and Mathematics with Applications, 61(5):1442–1456, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000228.

Saha:2017:DGK

[Sah17]  Asit Saha.  Dynam-

ics of the generalized KP–MEW–Burgers equation with external periodic perturbation. Computers and Mathematics with Applications, 73(9):1879–1885, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300962.

Sarwar:2015:NOH

[SAIZ15] S. Sarwar, Salem Alkhalaf, S. Iqbal, and M. A. Zahid. A note on optimal homotopy asymptotic method for the solutions of fractional order heat- and wave-like partial differential equations. Computers and Mathematics with Applications, 70(5):942–953, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500303X.

Sajavicius:2012:SWS

[Saj12] Svajunas Sajavicius. Stability of the weighted splitting finite-difference scheme for a two-dimensional parabolic equation with two nonlocal integral conditions. Computers and Mathematics with Applications, 64 (11):3485–3499, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005445.

Sajavicius:2014:RBF

[Saj14] Svajunas Sajavicius. Radial basis function method for a multidimensional linear elliptic equation with nonlocal boundary conditions. Computers and Mathematics with Applications, 67(7):1407– 1420, April 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000261.

Salem:2010:FCA

[Sal10] Hussein A. H. Salem.

On the fractional calculus in abstract spaces and their applications to the Dirichlet-type problem of fractional order. Computers and Mathematics with Applications, 59(3):1278–1293, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003976.

Salem:2011:QIE

[Sal11] Hussein A. H. Salem.

On the quadratic integral equations and their applications. Computers and Mathematics with Applications, 62(8):2931–2943, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006456.

Salazar:2016:NHS

[Sal16]  W. Salazar.  Numer-

ical homogenization of a second order discrete model for traffic flow. Computers and Mathematics with Applications, 71(1):29–45, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005088.

Samet:2019:BPN

[Sam19] Bessem Samet. Blow-up phenomena for a nonlinear time fractional heat equation in an exterior domain. Computers and Mathematics with Applications, 78(5):1380–1385, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305844.

Sana:2011:EMH

[San11] Shib Sankar Sana. An EOQ model of homogeneous products while demand is salesmen’s initiatives and stock sensitive. Computers and Mathematics with Applications, 62(2):577–587, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004305.

Sang:2012:MSN

[San12]  Yanbin Sang.  Multi-

ple solutions for nonlinear operator equations under the condition of two pairs of paralleled lower and upper solutions. Computers and Mathematics with Applications, 63(8):1349–1353, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001319.

Sarikaya:2010:NWO

[Sar10] Mehmet Zeki Sarikaya.

New weighted Ostrowski and cebysev type inequalities on time scales. Computers and Mathematics with Applications, 60 (5):1510–1514, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004475.

Sarigol:2011:CGS

[Sar11] Mehmet Ali Sarigöl.

Characterization of general summability factors and applications. Computers and Mathematics with Applications, 62 (6):2665–2670, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006730.

Szabo:2018:FAD

[SAR18] Barna Szabó, Ricardo Actis, and David Rusk. On the formulation and application of design rules. Computers and Mathematics with Applications, 74(9):2191–2202, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730398X.

Satsanit:2011:SDF

[Sat11] Wanchak Satsanit. On the solution of n-dimensional and the Fourier Bessel transform of the operator ⊗_{B}^{k}. Computers and Mathematics with Applications, 61(10):3105– 3116, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211100280X.

Siap:2011:SRH

[SAU11] Irfan Siap, Hasan Akin, and Selman Uguz. Structure and reversibility of 2D hexagonal cellular automata. Computers and Mathematics with Applications, 62 (11):4161–4169, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008479.

Savas:2010:GSF

[Sav10] Ekrem Savaş. On generalized |A|_{k}-summability factors. Computers and Mathematics with Applications, 59(1):514–518, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003848.

Singh:2014:WOF

[SB14] Bani Singh and Anuj Bhardwaj. Wavelet optimized finite difference mesh for MHD flow in a circular duct. Computers and Mathematics with Applications, 67(8):1582– 1594, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000893.

Salam:2019:UHR

[SB19] Ahmed Salam and Haithem Ben Kahla. An upper Jhessenberg reduction of a matrix through symplectic Householder transformations. Computers and Mathematics with Applications, 78(1):178–190, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300999.

Stokes:2010:OCP

[SBA10] Klara Stokes and Maria Bras-Amorós. Optimal configurations for peerto-peer user-private information retrieval. Computers and Mathematics with Applications, 59(4):1568–1577, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000143.

Scroggs:2018:SFI

[SBB^{+}18] Matthew W. Scroggs, Timo Betcke, Erik Burman, Wojciech Śmigaj, and Elwin van ’t Wout. Software frameworks for integral equations in electromagnetic scattering based on Calderón identities. Computers and Mathematics with Applications, 74 (11):2897–2914, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304789.

Saravi:2010:SLO

[SBEB10] M. Saravi, E. Babolian, R. England, and M. Bromilow. System of linear ordinary differential and differentialalgebraic equations and pseudo-spectral method. Computers and Mathematics with Applications, 59(4):1524–1531, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007895.

Shen:2015:ABT

[SBJ15] Shu-Qian Shen, Wen-Di Bao, and Ling Jian. On augmentation block triangular preconditioners for regularized saddle point problems. Computers and Mathematics with Applications, 69(8):828–837, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115000607.

Schaefer:2012:ABM

[SBKS12] Robert Schaefer, Aleksander Byrski, Joanna Kolodziej, and Maciej Smolka. An agent-based model of hierarchic genetic search. Computers and Mathematics with Applications, 64 (12):3763–3776, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001770.

Shidfar:2010:SIP

[SBM10] A. Shidfar, A. Babaei, and A. Molabahrami. Solving the inverse problem of identifying an unknown source term in a parabolic equation. Computers and Mathematics with Applications, 60 (5):1209–1213, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004141.

Singh:2013:CFS

[SBM13] Ram Mehar Singh, S. B. Bhardwaj, and S. C. Mishra. Closed-form solutions of the Schrödinger equation for a coupled harmonic potential in three dimensions. Computers and Mathematics with Applications, 66(4):537–541, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003507.

Sharifi:2012:FSN

[SBS12] M. Sharifi, D. K. R.

Babajee, and F. Soleymani. Finding the solution of nonlinear equations by a class of optimal methods. Computers and Mathematics with Applications, 63(4):764–774, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010236.

Sarmany:2013:TIM

[SBvdV13] D. Sármány, M. A.

Botchev, and J. J. W. van der Vegt. Timeintegration methods for finite element discretisations of the secondorder Maxwell equation. Computers and Mathematics with Applications, 65(3):528–543, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004713.

Sofonea:2019:BOC

[SbX19] Mircea Sofonea and Yi bin Xiao. Boundary optimal control of a nonsmooth frictionless contact problem. Computers and Mathematics with Applications, 78(1):152–165, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301014.

Sheng:2013:IBG

[SC13] Xingping Sheng and Guoliang Chen. Innovation based on Gaussian elimination to compute generalized inverse A_{T,S}^{(2)}. Computers and Mathematics with Applications, 65(11):1823– 1829, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113001636.

Shi:2016:GQA

[SC16] Hongxia Shi and Haibo Chen. Generalized quasilinear asymptotically periodic Schrödinger equations with critical growth. Computers and Mathematics with Applications, 71(3):849–858, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116000109.

Safari:2019:CIS

[SC19a] Farzaneh Safari and Wen Chen. Coupling of the improved singular boundary method and dual reciprocity method for multi-term time-fractional mixed diffusion-wave equations. Computers and Mathematics with Applications, 78(5):1594–1607, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300690.

Su:2019:FKT

[SC19b] Yu Su and Haibo Chen.

Fractional Kirchhoff-type equation with Hardy– Littlewood–Sobolev critical exponent. Computers and Mathematics with Applications, 78 (6):2063–2082, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211930183X.

Scaramuzzino:2011:HCS

[Sca11] F. Scaramuzzino. Hölderian continuity of solutions of a dynamic Walrasian price equilibrium problem with application. Computers and Mathematics with Applications, 61(7):1800–1809, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000940.

Salomov:2014:PSM

[SCA14] Uktam R. Salomov, Eliodoro Chiavazzo, and Pietro Asinari. Pore-scale modeling of fluid flow through gas diffusion and catalyst layers for high temperature proton exchange membrane (HT-PEM) fuel cells. Computers and Mathematics with Applications, 67(2):393–411, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122113004872.

Sixto-Camacho:2013:AHP

[SCBCB^{+}13] Lázaro M. Sixto-Camacho, Julian Bravo-Castillero, Renald Brenner, Raúl Guinovart-Díaz, Houari Mechkour, Reinaldo Rodríguez-Ramos, and Federico J. Sabina. Asymptotic homogenization of periodic thermo-magneto-electro-elastic heterogeneous media. Computers and Mathematics with Applications, 66 (10):2056–2074, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005221.

See corrigendum [SCBCB^{+}17].

Sixto-Camacho:2017:CAH

[SCBCB^{+}17] Lázaro M. Sixto-Camacho, Julián Bravo-Castillero, Renald Brenner, Raúl Guinovart-Díaz, Houari Mechkour, Reinaldo Rodríguez-Ramos, and Federico J. Sabina. Corrigendum to: “Asymptotic homogenization of periodic thermo-magneto-electro-elastic heterogeneous media” [Comput. Math. Appl. 66 (2013) 2056–2074]. Computers and Mathematics with Applications, 74 (6):1525–1527, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304169.

See [SCBCB^{+}13].

Shieh:2012:GDB

[SCC^{+}12a] Ming-Yuan Shieh, Chien-Sheng Chen, Chen-Hsin Chuang, Yi-Rong Liou, and Jeng-Han Li. Gait detection based stable locomotion control system for biped robots. Computers and Mathematics with Applications, 64 (5):1431–1440, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002957.

Suantai:2012:HIB

[SCC12b] Suthep Suantai, Yeol Je Cho, and Prasit Cholamjiak. Halpern’s iteration for Bregman strongly nonexpansive mappings in reflexive Banach spaces. Computers and Mathematics with Applications, 64(4):489–499, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010819.

Sun:2018:MMQ

[SCGW18] Zhengjia Sun, Jintao Cui, Fuzheng Gao, and Chao Wang. Multigrid methods for a quad–curl problem based on C^{0} interior penalty method. Computers and Mathematics with Applications, 76(9):2192–2211, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304395.

Sintunavarat:2011:CFP

[SCK11] Wutiphol Sintunavarat, Yeol Je Cho, and Poom Kumam. Common fixed point theorems for cdistance in ordered cone metric spaces. Computers and Mathematics with Applications, 62(4):1969–1978, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005207.

Shrestha:2010:KBA

[SCKH10] Anish Prasad Shrestha, Dong-You Choi, Goo Rak Kwon, and Seung-Jo Han. Kerberos based authentication for interdomain roaming in wireless heterogeneous network. Computers and Mathematics with Applications, 60(2):245–255, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000416.

Sanchez:2019:PDE

[SCSF19] Fabio Sanchez, Juan G. Calvo, Esteban Segura, and Zhilan Feng. A partial differential equation model with agestructure and nonlinear recidivism: Conditions for a backward bifurcation and a general numerical implementation. Computers and Mathematics with Applications, 78(12):3916–3930, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303244.

Saadati:2010:SQF

[SCV10] Reza Saadati, Yeol J. Cho, and Javad Vahidi. The stability of the quartic functional equation in various spaces. Computers and Mathematics with Applications, 60(7):1994–2002, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005262.

Setien:2019:EMD

[SCvdV^{+}19] Iñaki Setien, Michele Chiumenti, Sjoerd van der Veen, Maria San Sebastian, Fermín Garciandía, and Alberto Echeverría. Empirical methodology to determine inherent strains in additive manufacturing. Computers and Mathematics with Applications, 78(7):2282–2295, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302839.

Chen:2018:ELM

[sCYhX18] Shu sheng Chen, Chao Yan, and Xing hao Xiang. Effective low-Mach number improvement for upwind schemes. Computers and Mathematics with Applications, 75(10):3737–3755, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301093.

Saadatmandi:2010:NOM

[SD10a] Abbas Saadatmandi and Mehdi Dehghan. A new operational matrix for solving fractionalorder differential equations. Computers and Mathematics with Applications, 59(3):1326–1336, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109004258.

Saadatmandi:2010:NSH

[SD10b] Abbas Saadatmandi and Mehdi Dehghan. Numerical solution of the higher-order linear Fredholm integro-differential-difference equation with variable coefficients. Computers and Mathematics with Applications, 59(8): 2996–3004, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001276.

Saadatmandi:2011:TAS

[SD11a] Abbas Saadatmandi and Mehdi Dehghan. A tau approach for solution of the space fractional diffusion equation. Computers and Mathematics with Applications, 62(3):1135–1142, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003014.

Shakeri:2011:FVS

[SD11b] Fatemeh Shakeri and Mehdi Dehghan. The finite volume spectral element method to solve Turing models in the biological pattern formation. Computers and Mathematics with Applications, 62(12):4322–4336, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008273.

Salehi:2012:ULP

[SD12a] Rezvan Salehi and Mehdi Dehghan. The use of a Legendre pseudospectral viscosity technique to solve a class of nonlinear dynamic Hamilton– Jacobi equations. Computers and Mathematics with Applications, 63(3):629–644, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009758.

Shan:2012:TTD

[SD12b] Erfang Shan and Yanxia Dong. The k-tuple twin domination in generalized de Bruijn and Kautz networks. Computers and Mathematics with Applications, 63(1):222–227, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009795.

Sabouri:2015:EIS

[SD15a] Mania Sabouri and Mehdi Dehghan. An efficient implicit spectral element method for timedependent nonlinear diffusion equations by evaluating integrals at one quadrature point. Computers and Mathematics with Applications, 70 (10):2513–2541, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500440X.

Solis:2015:EMM

[SD15b] Francisco J. Solis and Sandra E. Delgadillo. Evolution of a mathematical model of an aggressive-invasive cancer under chemotherapy. Computers and Mathematics with Applications, 69(7):545–558, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000371.

Sabouri:2018:HMS

[SD18] Mania Sabouri and Mehdi Dehghan. A hk mortar spectral element method for the p-Laplacian equation. Computers and Mathematics with Applications, 76(7):1803–1826, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303961.

Shen:2019:BPP

[SD19] Xuhui Shen and Juntang Ding. Blow-up phenomena in porous medium equation systems with nonlinear boundary conditions. Computers and Mathematics with Applications, 77(12):3250–3263, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300756.

Sedaghatjoo:2013:UCB

[SDH13] Zeinab Sedaghatjoo, Mehdi Dehghan, and Hossein Hosseinzadeh. The use of continuous boundary elements in the boundary elements method for domains with non-smooth boundaries via finite difference approach. Computers and Mathematics with Applications, 65(7): 983–995, April 2013. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113000552.

Schmidt:2015:NCG

[SDH15] Kersten Schmidt, Julien Diaz, and Christian Heier. Non-conforming Galerkin finite element methods for local absorbing boundary conditions of higher order. Computers and Mathematics with Applications, 70(9):2252–2269, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004137.

Sardar:2010:SFI

[SDM10] Sujit Kumar Sardar, Bijan Davvaz, and Samit Kumar Majumder. A study on fuzzy interior ideals of Γ-semigroups. Computers and Mathematics with Applications, 60(1):90– 94, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003044.

Seadawy:2011:NES

[Sea11] A. R. Seadawy. New exact solutions for the KdV equation with higher order nonlinearity by using the variational method. Computers and Mathematics with Applications, 62(10):3741–3755, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007917.

Seadawy:2014:SAZ

[Sea14] A. R. Seadawy. Stability analysis for Zakharov– Kuznetsov equation of weakly nonlinear ionacoustic waves in a plasma. Computers and Mathematics with Applications, 67(1):172–180, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006421.

Seadawy:2015:FSW

[Sea15] Aly R. Seadawy. Fractional solitary wave solutions of the nonlinear higher-order extended KdV equation in a stratified shear flow: Part I. Computers and Mathematics with Applications, 70(4):345–352, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001960.

Seadawy:2016:TDN

[Sea16] Aly R. Seadawy. Threedimensional nonlinear modified Zakharov–Kuznetsov equation of ion-acoustic waves in a magnetized plasma. Computers and Mathematics with Applications, 71(1):201–212, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005374.

Sedaghat:2013:ROP

[Sed13] H. Sedaghat. Reduction of order, periodicity and boundedness in a class of nonlinear, higher order difference equations. Computers and Mathematics with Applications, 66(11):2231–2238, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003945.

Segeth:2019:PSG

[Seg19] Karel Segeth. Polyharmonic splines generated by multivariate smooth interpolation. Computers and Mathematics with Applications, 78 (9):3067–3076, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302160.

Sherief:2013:EBF

[SEM13] Hany H. Sherief and Nasser M. El-Maghraby. Effect of body forces on a 2D generalized thermoelastic long cylinder. Computers and Mathematics with Applications, 66(7):1181–1191, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004392.

Sendov:2012:RTP

[Sen12] Blagovest Sendov. Roumen tsanev — the pioneer in bio-mathematical research in Bulgaria. Computers and Mathematics with Applications, 64(3): 164–166, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004464.

Serezhkin:2019:MMW

[Ser19] A. Serezhkin. Mathematical modeling of widerange compressible twophase flows. Computers and Mathematics with Applications, 78(2):517– 540, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304334.

Set:2012:NIO

[Set12]  Erhan Set.  New in-

equalities of Ostrowski type for mappings whose derivatives are s-convex in the second sense via fractional integrals. Computers and Mathematics with Applications, 63(7):1147–1154, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010789.

Sakar:2012:VIM

[SEY12] Mehmet Giyas Sakar, Fevzi Erdogan, and Ahmet Yildirim. Variational iteration method for the time-fractional Fornberg–Whitham equation. Computers and Mathematics with Applications, 63(9):1382–1388, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000417.

Sun:2015:PFP

[SFM15] Liangliang Sun, Shengmao Fu, and Wenjun Ma. Pattern formation in a predator–prey diffusion model with stage structure for the predator. Computers and Mathematics with Applications, 70(12):2988–3000, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004915.

Sekhar:2010:FBB

[SG10a] D. Chandra Sekhar and R. Ganguli. Fractal boundaries of basin of attraction of Newton– Raphson method in helicopter trim. Computers and Mathematics with Applications, 60 (10):2834–2858, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007364.

Song:2010:AEM

[SG10b] Ming Song and Yuli Ge. Application of the (\frac{G′}{G})-expansion method to (3 + 1)-dimensional nonlinear evolution equations. Computers and Mathematics with Applications, 60(5):1220–1227, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004165.

Safi:2011:MAD

[SG11a] Mohammad A. Safi and Abba B. Gumel. Mathematical analysis of a disease transmission model with quarantine, isolation and an imperfect vaccine. Computers and Mathematics with Applications, 61(10):3044– 3070, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002756.

Shi:2011:LBS

[SG11b] Baochang Shi and Zhaoli Guo. Lattice Boltzmann simulation of some nonlinear convection– diffusion equations. Computers and Mathematics with Applications, 61(12): 3443–3452, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111000691.

Sharma:2014:EFO

[SG14] Janak Raj Sharma and Puneet Gupta. An efficient fifth order method for solving systems of nonlinear equations. Computers and Mathematics with Applications, 67(3):591–601, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006834.

Slodicka:2016:RTD

[SG16a] M. Slodicka and M. Galba. Recovery of a time dependent source from a surface measurement in Maxwell’s equations. Computers and Mathematics with Applications, 71(1):368–380, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005672.

Slodicka:2016:DTD

[SG16b] Marian Slodicka and Michal Galba. Determination of a timedependent convolution kernel from a boundary measurement in nonlinear Maxwell’s equations. Computers and Mathematics with Applications, 72(6):1484–1500, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303649.

Shimizu:2018:MBP

[SGK18] Yuma Shimizu, Hitoshi Gotoh, and Abbas Khayyer. An MPS-based particle method for simulation of multiphase flows characterized by high density ratios by incorporation of space potential particle concept. Computers and Mathematics with Applications, 76 (5):1108–1129, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303262.

Srivastava:2012:SPC

[SGQ12] H. M. Srivastava, Senlin Guo, and Feng Qi. Some properties of a class of functions related to completely monotonic functions. Computers and Mathematics with Applications, 64(6):1649–1654, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000260.

Sahiner:2011:ICC

[ŞGY11] A. Şahiner, M. Gürdal, and T. Yiǧit. Ideal convergence characterization of the completion of linear n-normed spaces. Computers and Mathematics with Applications, 61(3):683–689, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000920X.

Siddig:2018:IDM

[SGZW18] Abdelgader Siddig, Zhichang Guo, Zhenyu Zhou, and Boying Wu. An image denoising model based on a fourth-order nonlinear partial differential equation. Computers and Mathematics with Applications, 76 (5):1056–1074, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830316X.

Shi:2010:MHA

[SH10] Tian Shi and Songnian He. Modified hybrid algorithms for Lipschitz quasi-pseudo-contractive mappings in Hilbert spaces. Computers and Mathematics with Applications, 59(8):2940–2950, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001215.

Salas:2011:NES

[SH11] Alvaro H. Salas and Jairo E. Castillo H. New exact solutions to sinh-cosh-Gordon equation by using techniques based on projective Riccati equations. Computers and Mathematics with Applications, 61(2):470–481, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008990.

Shayanfar:2012:MFA

[SH12a] N. Shayanfar and M. Hadizadeh.

λ-matrix formulation applied to the Hertz contact problem with finite friction. Computers and Mathematics with Applications, 64(8):2478–2483, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004300.

She:2012:UMR

[SH12b] Yanhong She and Xiaoli He. Uncertainty measures for rough formulae in rough logic: an axiomatic approach. Computers and Mathematics with Applications, 63(1):83–93, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009618.

Saghi:2018:MDN

[SH18] Hassan Saghi and Ali Hashemian. Multidimensional NURBS model for predicting maximum free surface oscillation in swaying rectangular storage tanks. Computers and Mathematics with Applications, 76 (10):2496–2513, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304759.

Shatanawi:2010:POC

[Sha10]  W. Shatanawi.  Par-

tially ordered cone metric spaces and coupled fixed point results. Computers and Mathematics with Applications, 60(8):2508–2515, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006589.

Sharma:2012:CMM

[Sha12a] S. K. Sharma. Characterization and modeling of MIMO wireless channels based on correlation tensor. Computers and Mathematics with Applications, 64(2):89– 101, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000272.

Sharma:2012:NWM

[Sha12b]  Sanjay Sharma.  A

novel weighted multilevel space-time trellis coding scheme. Computers and Mathematics with Applications, 63(1):280–287, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111009862.

Shakhmurov:2014:SES

[Sha14] Veli B. Shakhmurov. Stokes equations with small parameters in half plane. Computers and Mathematics with Applications, 67(1):91–105, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300641X.

Shahrouzi:2018:BSN

[Sha18] Mohammad Shahrouzi. On behaviour of solutions for a nonlinear viscoelastic equation with variable-exponent nonlinearities. Computers and Mathematics with Applications, 75(11):3946–3956, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301342.

songHu:2011:CTE

[sHC11] Chang song Hu and Gang Cai. Convergence theorems for equilibrium problems and fixed point problems of a finite family of asymptotically k-strictly pseudocontractive mappings in the intermediate sense. Computers and Mathematics with Applications, 61(1):79–93, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008242.

Shehu:2011:SCT

[She11] Yekini Shehu. Strong convergence theorems for an infinite family of quasinonexpansive mappings and generalized equilibrium problems and variational inequality problems. Computers and Mathematics with Applications, 61(2):357–366, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008606.

Shehu:2012:HIS

[She12] Yekini Shehu. Hybrid iterative scheme for fixed point problem, infinite systems of equilibrium and variational inequality problems. Computers and Mathematics with Applications, 63(6):1089–1103, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010571.

Sheng:2016:MSV

[She16] Wei-Jie Sheng. Multidimensional stability of v-shaped traveling fronts in time periodic bistable reaction–diffusion equations. Computers and Mathematics with Applications, 72(6):1714–1726, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304400.

Shen:2018:ENE

[She18a] Liejun Shen. Existence and non-existence results for quasilinear Kirchhoff problems with the Hardy–Sobolev exponent. Computers and Mathematics with Applications, 76(8):1923–1937, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304103.

Shen:2018:ERF

[She18b]  Liejun Shen.  Exis-

tence result for fractional Schrödinger–Poisson systems involving a Bessel operator without Ambrosetti– Rabinowitz condition. Computers and Mathematics with Applications, 75(1):296–306, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730559X.

Shao:2016:VTL

[SHH16] Xinping Shao, Danfu Han, and Xianliang Hu. A p-version two level spline method for 2D Navier–Stokes equations. Computers and Mathematics with Applications, 71(12):2557–2567, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301870.

Sathinarain:2013:NIP

[SHM13] M. Sathinarain, C. Harley, and E. Momoniat. Numerical investigation of the parabolic mixed derivative diffusion equation via alternating direction implicit methods. Computers and Mathematics with Applications, 66(8):1452–1465, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005208.

Stamova:2010:GES

[SI10] Ivanka M. Stamova and Rajcho Ilarionov. On global exponential stability for impulsive cellular neural networks with time-varying delays. Computers and Mathematics with Applications, 59(11):3508– 3515, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002324.

Szekeres:2017:FDA

[SI17] Béla J. Szekeres and Ferenc Izsák. Finite difference approximation of space-fractional diffusion problems: the matrix transformation method. Computers and Mathematics with Applications, 73(2):261–269, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122116306381.

Sellier:2015:MDQ

[SID15] J. M. Sellier, D. Y.

Ivanova, and I. Dimov. Molecular descriptors and quasi-distribution functions. Computers and Mathematics with Applications, 70 (11):2726–2731, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003363.

Seadawy:2019:APL

[SIL19] Aly R. Seadawy, Mujahid Iqbal, and Dianchen Lu. Applications of propagation of long-wave with dissipation and dispersion in nonlinear media via solitary wave solutions of generalized Kadomtsev–Petviashvili modified equal width dynamical equation. Computers and Mathematics with Applications, 78 (11):3620–3632, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303165.

Simsek:2010:TAF

[Sim10] Yilmaz Simsek. Twisted p-adic (h, q)–L-functions. Computers and Mathematics with Applications, 59(6):2097–2110, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007743.

Sinkala:2016:GAF

[Sin16] Winter Sinkala. On the generation of arbitragefree stock price models using Lie symmetry analysis. Computers and Mathematics with Applications, 72(5):1386–1393, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303807.

Shah:2012:WCA

[SISH12] Rehan Ali Shah, Saeed Islam, A. M. Siddiqui, and T. Haroon. Wire coating analysis with Oldroyd 8constant fluid by optimal homotopy asymptotic method. Computers and Mathematics with Applications, 63(3):695–707, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010133.

Song:2014:CAH

[SJC14] Huailing Song, Lijian Jiang, and Gaojie Chen. Convergence analysis of hybrid expanded mixed finite element method for elliptic equations. Computers and Mathematics with Applications, 68 (10):1205–1219, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400409X.

Sun:2014:QMR

[SJHC14] Dong-Lin Sun, Yan-Fei Jing, Ting-Zhu Huang, and Bruno Carpentieri. A quasi-minimal residual variant of the Bi-CORSTAB method for nonsymmetric linear systems. Computers and Mathematics with Applications, 67(10):1743– 1755, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001746.

Schroeder:2019:RSS

[SJL^{+}19] Philipp W. Schroeder, Volker John, Philip L. Lederer, Christoph Lehrenfeld, Gert Lube, and Joachim Schöberl. On reference solutions and the sensitivity of the 2D Kelvin–Helmholtz instability problem. Computers and Mathematics with Applications, 77(4):1010–1028, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306278.

Shabir:2010:CRS

[SJN10a] Muhammad Shabir, Young Bae Jun, and Yasir Nawaz. Characterizations of regular semigroups by (α, β)- fuzzy ideals. Computers and Mathematics with Applications, 59(1):161– 175, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005823.

Shabir:2010:SCF

[SJN10b] Muhammad Shabir, Young Bae Jun, and Yasir Nawaz. Semigroups characterized by (∈, ∈ ∨q_{k})-fuzzy ideals. Computers and Mathematics with Applications, 60(5):1473–1493, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000444X.

Srivastava:2011:TSI

[SJPS11] H. M. Srivastava, Dragana Jankov, Tibor K. Pogány, and R. K. Saxena. Two-sided inequalities for the extended Hurwitz–Lerch zeta function. Computers and Mathematics with Applications, 62(1):516–522, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004330.

Song:2010:MBO

[SJS^{+}10] Xue Guan Song, Ji Hoon Jung, Hwan Jung Son, Joon Hong Park, Kwon Hee Lee, and Young Chul Park. Metamodel-based optimization of a control arm considering strength and durability performance. Computers and Mathematics with Applications, 60(4):976–980, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001951.

Song:2011:IRD

[SJS^{+}11] Li-Peng Song, Zhen Jin, Gui-Quan Sun, Juan Zhang, and Xie Han. Influence of removable devices on computer worms: Dynamic analysis and control strategies. Computers and Mathematics with Applications, 61(7):1823–1829, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000964.

Shabir:2010:COS

[SK10a] Muhammad Shabir and Asghar Khan. Characterizations of ordered semigroups by the properties of their fuzzy ideals. Computers and Mathematics with Applications, 59(1):539–549, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003800.

Sivasubramanian:2010:RCA

[SK10b] M. Sivasubramanian and S. Kalimuthu. RE-TRACTED: a computer application in mathematics. Computers and Mathematics with Applications, 59(1):296–297, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005070.

Saewan:2011:MHP

[SK11a] Siwaporn Saewan and Poom Kumam. A modified hybrid projection method for solving generalized mixed equilibrium problems and fixed point problems in Banach spaces. Computers and Mathematics with Applications, 62(4):1723–1735, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004822.

Sagsoz:2011:NTN

[SK11b] Fatma Sagsöz and Muhammet Kamali. On neighborhoods of two new subclasses of multivalent functions with negative coefficients. Computers and Mathematics with Applications, 62(4):1772–1779, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004986.

Sun:2011:ICF

[SK11c] Yuangong Sun and Qingkai Kong. Interval criteria for forced oscillation with nonlinearities given by Riemann–Stieltjes integrals. Computers and Mathematics with Applications, 62(1):243–252, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003750.

Szasz:2011:EOS

[SK11d] Róbert Szász and Pál Aurel Kupán. The exact order of starlikeness of uniformly convex functions. Computers and Mathematics with Applications, 62(1):173–186, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003683.

Sivasubramanian:2012:RNC

[SK12] M. Sivasubramanian and S. Kalimuthu. Retraction notice to: “Computer application in mathematics” [Comput. Math. Appl. 59 (1) (2009) 296– 297]. Computers and Mathematics with Applications, 63(9):1424, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003008.

Safdari:2014:LBS

[SK14a] Arman Safdari and Kyung Chun Kim. Lattice Boltzmann simulation of solid particles behavior in a three-dimensional liddriven cavity flow. Computers and Mathematics with Applications, 68(5):606–621, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002879.

Sun:2014:NSHa

[SK14b] Yifei Sun and Mrinal Kumar. Numerical solution of high dimensional stationary Fokker–Planck equations via tensor decomposition and Chebyshev spectral differentiation. Computers and Mathematics with Applications, 67(10):1960– 1977, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001771.

Suchde:2019:MGF

[SK19] Pratik Suchde and Jörg Kuhnert. A meshfree generalized finite difference method for surface PDEs. Computers and Mathematics with Applications, 78(8):2789–2805, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302469.

Sheu:2019:PEC

[SKCL19] Tony W. H. Sheu, C. Y. Kao, Y. W. Chang, and J. H. Li. Prediction of evanescent coupling efficiency in two parallel silica nanowires. Computers and Mathematics with Applications, 78(3):707–722, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301099.

Shah:2011:CCT

[SKdA11] Tariq Shah, Atlas Khan, and Antonio Aparecido de Andrade. Constructions of codes through the semigroup ring B[X; \frac{1}{2^{2}} Z_{0}] and encoding. Computers and Mathematics with Applications, 62(4):1645–1654, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004676.

Stiebler:2011:LBL

[SKFG11] M. Stiebler, M. Krafczyk, S. Freudiger, and M. Geier. Lattice Boltzmann large eddy simulation of subcritical flows around a sphere on non-uniform grids. Computers and Mathematics with Applications, 61(12):3475– 3484, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002264.

Schonherr:2011:MTI

[SKG^{+}11] M. Schönherr, K. Kucher, M. Geier, M. Stiebler, S. Freudiger, and M. Krafczyk. Multi-thread implementations of the lattice Boltzmann method on non-uniform grids for CPUs and GPUs. Computers and Mathematics with Applications, 61(12): 3730–3743, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111002999.

Syafaruddin:2012:FWN

[SKH12] Syafaruddin, Engin Karatepe, and Takashi Hiyama. Fuzzy wavelet network identification of optimum operating point of noncrystalline silicon solar cells. Computers and Mathematics with Applications, 63(1):68–82, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009606.

Shemyakova:2010:CCC

[SKJ10] E. Shemyakova, S. I. Khashin, and D. J. Jeffrey. A conjecture concerning a completely monotonic function. Computers and Mathematics with Applications, 60(5):1360–1363, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004311.

Stanislawska:2012:MGT

[SKK12] Karolina Stanislawska, Krzysztof Krawiec, and Zbigniew W. Kundzewicz. Modeling global temperature changes with genetic programming. Computers and Mathematics with Applications, 64 (12):3717–3728, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001745.

Srivastava:2011:QNS

[SKM11] Pankaj Kumar Srivastava, Manoj Kumar, and R. N. Mohapatra. Quintic nonpolynomial spline method for the solution of a second-order boundary-value problem with engineering applications. Computers and Mathematics with Applications, 62(4):1707–1714, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004809.

Sheshko:2014:AFP

[SKPW14] Michail A. Sheshko, Pawel Karczmarek, Dorota Pylak, and Pawel Wójcik. Application of Faber polynomials to the approximate solution of a generalized boundary value problem of linear conjugation in the theory of analytic functions. Computers and Mathematics with Applications, 67(8):1474–1481, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000832.

Spartalis:2010:HOS

[SKST10] S. Spartalis, M. Konstantinidou-Serafimidou, and A. Taouktsoglou. C-hypergroupoids obtained by special binary relations. Computers and Mathematics with Applications, 59(8):2628– 2635, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000532.

Suga:2015:DMR

[SKTC15] K. Suga, Y. Kuwata, K. Takashima, and R. Chikasue. A D3Q27 multiplerelaxation-time lattice Boltzmann method for turbulent flows. Computers and Mathematics with Applications, 69(6):518– 529, March 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115000346.

See corrigendum [SKTC19].

Suga:2019:CDM

[SKTC19] K. Suga, Y. Kuwata, K. Takashima, and R. Chikasue. Corrigendum to “A D3Q27 multiplerelaxation-time lattice Boltzmann method for turbulent flows” [Comput. Math. Appl. 69(6) (2015) 518–529]. Computers and Mathematics with Applications, 78(1):254–256, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301075.

See [SKTC15].

Soheili:2013:ANS

[SKTD13] Ali R. Soheili, A. Kerayechian, H. R. Tareghian, and N. Davoodi. Adaptive numerical simulation of traffic flow density. Computers and Mathematics with Applications, 66(3):227–237, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113002526.

Scherer:2011:GLM

[SKTH11] Rudolf Scherer, Shyam L. Kalla, Yifa Tang, and Jianfei Huang. The Grünwald–Letnikov method for fractional differential equations. Computers and Mathematics with Applications, 62(3):902– 917, August 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002173.

Sahoo:2012:SHF

[SL12] Bikash Sahoo and Fotini Labropulu. Steady homann flow and heat transfer of an electrically conducting second grade fluid. Computers and Mathematics with Applications, 63(7):1244–1255, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010972.

Seleson:2016:CSM

[SL16a] Pablo Seleson and David J.

Littlewood. Convergence studies in meshfree peridynamic simulations. Computers and Mathematics with Applications, 71(11):2432– 2448, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005891.

Shcherbakov:2016:RBF

[SL16b] Victor Shcherbakov and Elisabeth Larsson. Radial basis function partition of unity methods for pricing vanilla basket options. Computers and Mathematics with Applications, 71(1):185–200, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005386.

Smith:2016:ACP

[SL16c] Matthew R. Smith and Yi-Hsin Lin. Analysis of the convergence properties for a nonlinear implicit equilibrium flux method using quasi Newton–Raphson and BiCGStab techniques. Computers and Mathematics with Applications, 72(8):2008–2019, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302875.

Shao:2018:TTI

[SL18a] Weidong Shao and Jun Li. Three time integration methods for incompressible flows with discontinuous Galerkin Boltzmann method. Computers and Mathematics with Applications, 75(11):4091–4106, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301445.

Sheikholeslami:2018:MFO

[SL18b] Somayyeh Sheikholeslami and James V. Lambers. Modeling of firstorder photobleaching kinetics using Krylov subspace spectral methods. Computers and Mathematics with Applications, 75(6):2153–2172, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122117306673.

Shih:2012:FDS

[SLCC12] Hsu-Shih Shih, E. Stanley Lee, Shun-Hsiang Chuang, and Chiau-Ching Chen. A forecasting decision on the sales volume of printers in Taiwan: an exploitation of the analytic network process. Computers and Mathematics with Applications, 64(6):1545–1556, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011540.

Slemrod:2013:BEH

[Sle13] M. Slemrod. From Boltzmann to Euler: Hilbert’s 6th problem revisited. Computers and Mathematics with Applications, 65(10):1497–1501, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005664.

Sahu:2012:EAF

[SLK12] D. R. Sahu, Zeqing Liu, and Shin Min Kang. Existence and approximation of fixed points of nonlinear mappings in spaces with weak uniform normal structure. Computers and Mathematics with Applications, 64(4):672–685, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011539.

Stockrahm:2019:CDM

[SLKK19] Alex Stockrahm, Valtteri Lahtinen, Jari J. J. Kangas, and P. Robert Kotiuga. Cuts for 3d magnetic scalar potentials: Visualizing unintuitive surfaces arising from trivial knots. Computers and Mathematics with Applications, 78(9):3200–3210, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930286X.

Sun:2012:ESFa

[SLL12a] Jihua Sun, Yiliang Liu, and Guifang Liu. Existence of solutions for fractional differential systems with antiperiodic boundary conditions. Computers and Mathematics with Applications, 64 (6):1557–1566, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011552.

Sun:2012:EUS

[SLL12b] Shurong Sun, Qiuping Li, and Yanan Li. Existence and uniqueness of solutions for a coupled system of multi-term nonlinear fractional differential equations. Computers and Mathematics with Applications, 64 (10):3310–3320, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000867.

Shin:2017:CSR

[SLL17] Jaemin Shin, Hyun Geun Lee, and June-Yub Lee. Convex splitting Runge– Kutta methods for phasefield models. Computers and Mathematics with Applications, 73(11):2388–2403, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117302250.

Srinivasan:2016:MDS

[SLM16] Shriram Srinivasan, Raytcho Lazarov, and Peter Minev. Multiscale direction-splitting algorithms for parabolic equations with highly heterogeneous coefficients. Computers and Mathematics with Applications, 72(6):1641–1654, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304370.

Shields:2018:WGM

[SLM18] Sidney Shields, Jichun Li, and Eric A. Machorro. Weak Galerkin methods for time-dependent Maxwell’s equations. Computers and Mathematics with Applications, 74 (9):2106–2124, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304765.

Sheng:2012:ISP

[SLMZ12] Wanxing Sheng, Yongmei Liu, Xiaoli Meng, and Tianshu Zhang. An improved strength Pareto evolutionary algorithm 2 with application to the optimization of distributed generations. Computers and Mathematics with Applications, 64(5):944–955, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000843.

Srivastava:2011:NRJ

[SLW11] H. M. Srivastava, V. Lokesha, and Yu-Dong Wu. A new refinement of the Janous–Gmeiner inequality for a triangle. Computers and Mathematics with Applications, 62 (5):2349–2353, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005761.

Shi:2014:LOC

[SLW14] Dongyang Shi, Xin Liao, and Lele Wang. The lowest order characteristic mixed finite element scheme for convectiondominated diffusion problem. Computers and Mathematics with Applications, 68(7):759–769, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114003459.

Sun:2018:FTB

[SLW18] Fenglong Sun, Lishan Liu, and Yonghong Wu. Finite time blow-up for a class of parabolic or pseudo-parabolic equations. Computers and Mathematics with Applications, 75(10):3685– 3701, May 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301068.

Shen:2011:SGB

[SLXC11] Shigen Shen, Yuanjie Li, Hongyun Xu, and Qiying Cao. Signaling game based strategy of intrusion detection in wireless sensor networks. Computers and Mathematics with Applications, 62(6):2404–2416, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005815.

Song:2013:FSL

[SLYY13] Xiaoning Song, Zi Liu, Xibei Yang, and Jingyu Yang. A fuzzy supervised learning method with dynamical parameter estimation for nonlinear discriminant analysis. Computers and Mathematics with Applications, 66(10):1782–1794, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004690.

Sun:2011:IID

[SLZ11] Xiaoli Sun, Min Li, and Weiqiang Zhang. An improved image denoising model based on the directed diffusion equation. Computers and Mathematics with Applications, 61(8):2177–2181, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006942.

Stanculescu:2010:NAL

[SM10] Iuliana Stanculescu and Carolina C. Manica. Numerical analysis of Leray– Tikhonov deconvolution models of fluid motion.

Computers and Mathematics with Applications, 60(5):1440–1456, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004402.

Shabir:2011:CHF

[SM11] Muhammad Shabir and Tahir Mahmood. Characterizations of hemirings by (∈, ∈ ∨q_{k})-fuzzy ideals. Computers and Mathematics with Applications, 61(4):1059–1078, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009612.

Stoykov:2014:NCP

[SM14] S. Stoykov and S. Margenov. Numerical computation of periodic responses of nonlinear large-scale systems by shooting method. Computers and Mathematics with Applications, 67(12):2257–2267, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000649.

Shen:2017:ERA

[SM17] Xiaoying Shen and Qiaozhen Ma. Existence of random attractors for weakly dissipative plate equations with memory and additive white noise. Computers and Mathematics with Applications, 73(10):2258–2271, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301712.

Soldner:2019:TMS

[SM19] Dominic Soldner and Julia Mergheim. Thermal modelling of selective beam melting processes using heterogeneous time step sizes. Computers and Mathematics with Applications, 78(7):2183–2196, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302505.

Shidfar:2010:SSC

[SMBY10] A. Shidfar, A. Molabahrami, A. Babaei, and A. Yazdanian. A series solution of the Cauchy problem for the generalized d-dimensional Schrödinger equation with a power-law nonlinearity. Computers and Mathematics with Applications, 59(4):1500–1508, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007470.

Shen:2010:FVC

[SMC10] Si-Lian Shen, Chang-Lin Mei, and Jian-Ling Cui. A fuzzy varying coefficient model and its estimation. Computers and Mathematics with Applications, 60(6):1696–1705, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004633.

Shakeel:2018:CFS

[SMDI18] Muhammad Shakeel, Syed Tauseef Mohyud-Din, and Muhammad Asad Iqbal. Closed form solutions for coupled nonlinear Maccari system. Computers and Mathematics with Applications, 76(4):799–809, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118302888.

Sabatier:2010:LSC

[SMF10] Jocelyn Sabatier, Mathieu Moze, and Christophe Farges. LMI stability conditions for fractional order systems. Computers and Mathematics with Applications, 59(5):1594–1609, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005355.

Sakib:2017:SAI

[SMF17] N. Sakib, A. Mohammadi, and J. M. Floryan. Spectrally-accurate immersed boundary conditions method for threedimensional flows. Computers and Mathematics with Applications, 73(11):2426–2453, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302067.

Sofonea:2018:PMH

[SMH18] Mircea Sofonea, Stanisław Migórski, and Weimin Han. A penalty method for history-dependent variational-hemivariational inequalities. Computers and Mathematics with Applications, 75(7):2561– 2573, April 1, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117307812.

Smietanski:2011:CIG

[Śmi11] Marek J. Śmietański. Convergence of an inexact generalized Newton method with a scaled residual control. Computers and Mathematics with Applications, 61(6):1624–1632, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000484.

Saleh:2018:TMP

[SMK18] R. Saleh, S. M. Mabrouk, and M. Kassem. Truncation method with point transformation for exact solution of Liouville Bratu Gelfand equation. Computers and Mathematics with Applications, 76(5):1219–1227, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303407.

See comment [EM19].

Sayevand:2019:NNS

[SMM19] K. Sayevand, J. Tenreiro Machado, and V. Moradi. A new non-standard finite difference method for analyzing the fractional Navier–Stokes equations. Computers and Mathematics with Applications, 78(5):1681–1694, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307089.

Smolka:2017:DOC

[Smo17] Maciej Smolka. Differentiability of the objective in a class of coefficient inverse problems. Computers and Mathematics with Applications, 73(11):2375–2387, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302328.

Sun:2019:DLS

[SMYK19] Yong-Li Sun, Wen-Xiu Ma, Jian-Ping Yu, and Chaudry Masood Khalique.

Dynamics of lump solitary wave of Kadomtsev– Petviashvili–Boussinesqlike equation. Computers and Mathematics with Applications, 78(3):840–847, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301208.

Shabir:2011:STS

[SN11] Muhammad Shabir and Munazza Naz. On soft topological spaces. Computers and Mathematics with Applications, 61(7):1786–1799, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000927.

Sun:2019:CLD

[SND19] Jing Sun, Daxin Nie, and Weihua Deng. Central local discontinuous Galerkin method for the space fractional diffusion equation. Computers and Mathematics with Applications, 78(5):1274–1287, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300707.

Shangerganesh:2018:EWS

[SNDK18] L. Shangerganesh, N. Nyamoradi, V. N. Deiva Mani, and S. Karthikeyan. On the existence of weak solutions of nonlinear degenerate parabolic system with variable exponents. Computers and Mathematics with Applications, 75(1):322–334, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305679.

Siewert:2019:VML

[SNEP19] Marvin Siewert, Fabian Neugebauer, Jérémy Epp, and Vasily Ploshikhin. Validation of mechanical layer equivalent method for simulation of residual stresses in additive manufactured components. Computers and Mathematics with Applications, 78(7):2407–2416, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304346.

Saberi-Nadjafi:2010:SNI

[SNH10] Jafar Saberi-Nadjafi and Mahdi Heidari. Solving nonlinear integral equations in the Urysohn form by Newton–Kantorovichquadrature method. Computers and Mathematics with Applications, 60(7):2058–2065, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005389.

Saberi-Nadjafi:2012:SVI

[SNMA12] J. Saberi-Nadjafi, M. Mehrabinezhad, and H. Akbari. Solving Volterra integral equations of the second kind by wavelet-Galerkin scheme. Computers and Mathematics with Applications, 63(11):1536– 1547, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002489.

Shangerganesh:2019:FTB

[SNSK19] L. Shangerganesh, N. Nyamoradi, G. Sathishkumar, and S. Karthikeyan. Finitetime blow-up of solutions to a cancer invasion mathematical model with haptotaxis effects.

Computers and Mathematics with Applications, 77(8):2242–2254, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830693X.

Solis:2010:CFD

[SOJC10] Francisco J. Solis, Fausto Ongay, Silvia Jerez, and Marcos Capistran. Convergence for a family of discrete advectionreaction operators. Computers and Mathematics with Applications, 59(1): 499–505, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109003861.

Sokol:2010:CMF

[Sok10] Janusz Sokól. Classes of multivalent functions associated with a convolution operator. Computers and Mathematics with Applications, 60 (5):1343–1350, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004293.

Sokol:2011:CCS

[Sok11] Janusz Sokól. A certain class of starlike functions. Computers and Mathematics with Applications, 62(2):611–619, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004391.

Sin:2019:DEG

[SOK19] Chung-Sik Sin, Hyong-Chol O., and Sang-Mun Kim. Diffusion equations with general nonlocal time and space derivatives. Computers and Mathematics with Applications, 78 (10):3268–3284, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211930241X.

Sommariva:2013:FCF

[Som13] Alvise Sommariva. Fast construction of Fejér and Clenshaw–Curtis rules for general weight functions. Computers and Mathematics with Applications, 65(4):682–693, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200689X.

Sonmez:2011:SNS

[Sön11] Abdulcabbar Sönmez.

Some new sequence spaces derived by the domain of the triple band matrix. Computers and Mathematics with Applications, 62(2):641–650, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004433.

Shiraishi:2011:SCS

[SOS11] Hitoshi Shiraishi, Shigeyoshi Owa, and H. M. Srivastava. Sufficient conditions for strongly Carathéodory functions. Computers and Mathematics with Applications, 62(8):2978–2987, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006675.

Sousa:2011:NAF

[Sou11]  Ercília Sousa.  Nu-

merical approximations for fractional diffusion equations via splines. Computers and Mathematics with Applications, 62(3):938–944, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003026.

Sousa:2012:SOE

[Sou12] Ercília Sousa. A second order explicit finite difference method for the fractional advection diffusion equation. Computers and Mathematics with Applications, 64 (10):3141–3152, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002076.

Saadati:2010:NAF

[SP10] Reza Saadati and Choonkil Park. Non-Archimedian L-fuzzy normed spaces and stability of functional equations. Computers and Mathematics with Applications, 60(8):2488–2496, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006292.

Soize:2012:TDF

[SP12] C. Soize and I. E.

Poloskov. Time-domain formulation in computational dynamics for linear viscoelastic media with model uncertainties and stochastic excitation. Computers and Mathematics with Applications, 64(11):3594–3612, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005858.

Shon:2013:SIA

[SPCS13] Taeshik Shon, James J.

(Jong Hyuk) Park, Lambrinoudakis Costas, and Jean-Marc Seigneur. Special issue on advanced information security for secure and trust computing. Computers and Mathematics with Applications, 65(9):1233, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001727.

Sajid:2010:FDM

[SPH10] M. Sajid, I. Pop, and T. Hayat. Fully developed mixed convection flow of a viscoelastic fluid between permeable parallel vertical plates. Computers and Mathematics with Applications, 59(1):493–498, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003873.

Sandev:2019:CQM

[SPL19] Trifce Sandev, Irina Petreska, and Ervin K. Lenzi. Constrained quantum motion in δ-potential and application of a generalized integral operator. Computers and Mathematics with Applications, 78(5):1695–1704, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306874.

Safkhani:2014:CCA

[SPLHCB14] Masoumeh Safkhani, Pedro Peris-Lopez, Julio Cesar Hernandez-Castro, and Nasour Bagheri. Cryptanalysis of the Cho et al. protocol: a hashbased RFID tag mutual authentication protocol. Journal of Computational and Applied Mathematics, 259 (part B)(??):571–577, March 15, 2014. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0377042713005281.

See [CJP12].

Song:2018:ISP

[SPP18] Baofang Song, Arman Pazouki, and Thorsten Pöschel. Instability of smoothed particle hydrodynamics applied to Poiseuille flows. Computers and Mathematics with Applications, 76 (6):1447–1457, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303614.

Sayama:2013:MCS

[SPS^{+}13] Hiroki Sayama, Irene Pestov, Jeffrey Schmidt, Benjamin James Bush, Chun Wong, Junichi Yamanoi, and Thilo Gross. Modeling complex systems with adaptive networks. Computers and Mathematics with Applications, 65(10):1645– 1664, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112007018.

Singh:2018:ABM

[SPST18] Somveer Singh, Vijay Kumar Patel, Vineet Kumar Singh, and Emran Tohidi. Application of Bernoulli matrix method for solving twodimensional hyperbolic telegraph equations with Dirichlet boundary conditions. Computers and Mathematics with Applications, 75(7):2280–2294, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307666.

Safi:2017:BCT

[SPT17] Mohammad Amin Safi, Nikolaos Prasianakis, and Stefan Turek. Benchmark computations for 3D twophase flows: a coupled lattice Boltzmannlevel set study. Computers and Mathematics with Applications, 73(3):520–536, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211630685X.

Siddiqi:2010:IBF

[SR10a] Shahid S. Siddiqi and Kashif Rehan. Improved binary four point subdivision scheme and new corner cutting scheme. Computers and Mathematics with Applications, 59(8):2647–2657, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000568.

Srivastava:2010:BFT

[SR10b] V. P. Srivastava and Rati Rastogi. Blood flow through a stenosed catheterized artery: Effects of hematocrit and stenosis shape. Computers and Mathematics with Applications, 59(4):1377–1385, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007548.

Saeid:2012:IFA

[SR12] A. Borumand Saeid and A. Rezaei. Intuitionistic (T, S)-fuzzy CIalgebras. Computers and Mathematics with Applications, 63(1):158–166, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009709.

Sahoo:2015:IFS

[SR15] S. Sahoo and S. Saha Ray. Improved fractional sub-equation method for (3 + 1)-dimensional generalized fractional KdV– Zakharov–Kuznetsov equations. Computers and Mathematics with Applications, 70(2):158–166, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002138.

Seidl:2016:ITR

[SR16] Robert Seidl and Ernst Rank. Iterative time reversal based flaw identification. Computers and Mathematics with Applications, 72(4):879–892, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303327.

Sahoo:2017:LSA

[SR17a] S. Sahoo and S. Saha Ray. Lie symmetry analysis and exact solutions of (3+ 1) dimensional Yu–Toda– Sasa–Fukuyama equation in mathematical physics. Computers and Mathematics with Applications, 73(2):253–260, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306332.

Salkuyeh:2017:MGS

[SR17b] Davod Khojasteh Salkuyeh and Maryam Rahimian. A modification of the generalized shift-splitting method for singular saddle point problems. Computers and Mathematics with Applications, 74 (12):2940–2949, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304583.

Sun:2017:CWG

[SR17c] Ming Sun and Hongxing Rui. A coupling of weak Galerkin and mixed finite element methods for poroelasticity. Computers and Mathematics with Applications, 73(5):804–823, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300275.

Sun:2018:MFD

[SR18] Yue Sun and Hongxing Rui. MAC finite difference scheme for Stokes equations with damping on non-uniform grids. Computers and Mathematics with Applications, 75(4):1272–1287, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307101.

Sepahi:2017:AHS

[SRDD17] O. Sepahi, L. Radtke, S. E. Debus, and A. Düster. Anisotropic hierarchic solid finite elements for the simulation of passiveactive arterial wall models. Computers and Mathematics with Applications, 74(12):3058–3079, December 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304856.

Sun:2016:LUB

[SRG16] Lili Sun, Yuxue Ren, and Wenjie Gao. Lower and upper bounds for the blow-up time for nonlinear wave equation with variable sources. Computers and Mathematics with Applications, 71(1):267–277, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005593.

Salinas:2013:ECC

[SRGL13] P. Salinas, C. Rodrigo, F. J. Gaspar, and F. J. Lisbona. An efficient cell-centered multigrid method for problems with discontinuous coefficients on semi-structured triangular grids. Computers and Mathematics with Applications, 65(12):1978–1989, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002101.

Sakthivel:2011:ACS

[SRM11a] R. Sakthivel, Yong Ren, and N. I. Mahmudov. On the approximate controllability of semilinear fractional differential systems. Computers and Mathematics with Applications, 62(3):1451–1459, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003415.

Sivasubramanian:2011:CSC

[SRM11b] S. Sivasubramanian, Thomas Rosy, and K. Muthunagai. Certain sufficient conditions for a subclass of analytic functions involving hohlov operator. Computers and Mathematics with Applications, 62(12):4479–4485, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008893.

Sheremet:2018:MCH

[SRRP18] Mikhail A. Sheremet, Natalia C. Rosca, Alin V. Rosca, and Ioan Pop. Mixed convection heat transfer in a square porous cavity filled with a nanofluid with suction/ injection effect. Computers and Mathematics with Applications, 76(11– 12):2665–2677, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305029.

Shahraki:2011:NAE

[SRS11] Amin Shahraki, Marjan Kuchaki Rafsanjani, and Arsham Borumand Saeid. A new approach for energy and delay trade-off intra-clustering routing in WSNs. Computers and Mathematics with Applications, 62(4):1670–1676, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004718.

Srivastava:2010:TLS

[SRV10] V. P. Srivastava, Rati Rastogi, and Rochana Vishnoi. A two-layered suspension blood flow through an overlapping stenosis. Computers and Mathematics with Applications, 60(3):432–441, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003093.

Soltani:2010:NMV

[SS10] L. Ahmad Soltani and Ahmad Shirzadi. A new modification of the variational iteration method. Computers and Mathematics with Applications, 59(8):2528–2535, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000234.

Schott:2011:CCC

[SS11a] René Schott and G. Stacey Staples. Complexity of counting cycles using zeons. Computers and Mathematics with Applications, 62(4):1828–1837, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005050.

Set:2011:GOG

[SS11b] Erhan Set and Mehmet Zeki Sarikaya. On the generalization of Ostrowski and Grüss type discrete inequalities. Computers and Mathematics with Applications, 62(1):455– 461, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100424X.

Shatanawi:2011:WCC

[SS11c] Wasfi Shatanawi and Bessem Samet. On (ψ, ϕ)- weakly contractive condition in partially ordered metric spaces. Computers and Mathematics with Applications, 62(8):3204–3214, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006961.

Svanadze:2013:MPC

[SS13] Merab Svanadze and Antonio Scalia. Mathematical problems in the coupled linear theory of bone poroelasticity. Computers and Mathematics with Applications, 66 (9):1554–1566, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300076X.

Sharma:2014:USN

[SS14a] Nisha Sharma and Kapil K.

Sharma. Unconditionally stable numerical method for a nonlinear partial integro-differential equation. Computers and Mathematics with Applications, 67(1):62–76, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006226.

Sheng:2014:ESD

[SS14b] Qin Sheng and Hai-Wei Sun. Exponential splitting for n-dimensional paraxial Helmholtz equation with high wavenumbers. Computers and Mathematics with Applications, 68(10):1341– 1354, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114004532.

Sekhar:2016:GCI

[SS16a] T. Raja Sekhar and Purnima Satapathy. Group classification for isothermal drift flux model of two phase flows. Computers and Mathematics with Applications, 72 (5):1436–1443, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304096.

Shen:2016:GSS

[SS16b] Qin-Qin Shen and Quan Shi. Generalized shiftsplitting preconditioners for nonsingular and singular generalized saddle point problems. Computers and Mathematics with Applications, 72(3):632– 641, August 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116302905.

Slodicka:2016:ISP

[SS16c] M. Slodicka and K. Sisková.

An inverse source problem in a semilinear timefractional diffusion equation. Computers and Mathematics with Applications, 72(6):1655–1669, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304345.

Sheng:2017:RGB

[SS17] Xingping Sheng and Weiwei Sun. The relaxed gradient based iterative algorithm for solving matrix equations A_{i}XB_{i} = F_{i}.

Computers and Mathematics with Applications, 74(3):597–604, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302973.

Shen:2018:VHP

[SS18a] Qin-Qin Shen and Quan Shi. A variant of the HSS preconditioner for complex symmetric indefinite linear systems. Computers and Mathematics with Applications, 75(3):850–863, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306545.

Siskova:2018:SIP

[SS18b] K. Sisková and M. Slodicka. A source identification problem in a timefractional wave equation with a dynamical boundary condition. Computers and Mathematics with Applications, 75(12):4337–4354, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211830169X.

Soleymani:2019:PFH

[SS19] Fazlollah Soleymani and Behzad Nemati Saray. Pricing the financial Heston–Hull–White model with arbitrary correlation factors via an adaptive FDM. Computers and Mathematics with Applications, 77(4):1107–1123, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306448.

Sakthivel:2012:ACF

[SSA12] R. Sakthivel, S. Suganya, and S. M. Anthoni. Approximate controllability of fractional stochastic evolution equations. Computers and Mathematics with Applications, 63(3):660–668, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010042.

Sadeghian:2011:CCF

[SSAM11] Hoda Sadeghian, Hassan Salarieh, Aria Alasty, and Ali Meghdari. On the control of chaos via fractional delayed feedback method. Computers and Mathematics with Applications, 62(3):1482–1491, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003907.

Sriburadet:2019:SAC

[SSC19] Sirilak Sriburadet, Yin-Tzer Shih, and C.- S. Chien. Stability analysis and continuation for the coupled Gross–Pitaevskii equations. Computers and Mathematics with Applications, 78(3):807–826, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301221.

Sayed:2012:FRA

[SSESG12] Mostafa S. Sayed, Ahmed Shalaby, Mohamed El-Sayed, and Victor Goulart. Flexible router architecture for network-on-chip. Computers and Mathematics with Applications, 64(5):1301–1310, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002799.

Shahraeeni:2015:ASF

[SSH15] Mehran Shahraeeni, Rezgar Shakeri, and Seyyed Mohammad Hasheminejad. An analytical solution for free and forced vibration of a piezoelectric laminated plate coupled with an acoustic enclosure. Computers and Mathematics with Applications, 69(11):1329– 1341, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001273.

Sadeghi:2018:TDL

[SSHH^{+}18] R. Sadeghi, M. S. Shadloo, M. Hopp-Hirschler, A. Hadjadj, and U. Nieken. Three-dimensional lattice Boltzmann simulations of high density ratio twophase flows in porous media. Computers and Mathematics with Applications, 75(7):2445–2465, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307915.

Schanzel:2019:CTM

[SSIP19] Michael Schänzel, Damir Shakirov, Alexander Ilin, and Vasily Ploshikhin. Coupled thermo-mechanical process simulation method for selective laser melting considering phase transformation steels. Computers and Mathematics with Applications, 78(7):2230–2246, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300434.

Sonday:2013:NDS

[SSK13] Benjamin Sonday, Amit Singer, and Ioannis G. Kevrekidis. Noisy dynamic simulations in the presence of symmetry: Data alignment and model reduction. Computers and Mathematics with Applications, 65(10):1535–1557, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000412.

Syau:2011:CSF

[SSL11] Yu-Ru Syau, Ly-Fie Sugianto, and E. S. Lee.

Continuity and semicontinuity of fuzzy mappings. Computers and Mathematics with Applications, 61(4):1122–1128, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009673.

Sun:2014:NSHb

[SSL14] Ming Sun, Shicang Song, and Zeyi Liu. A nonconforming scheme with high accuracy for the plate bending problem. Computers and Mathematics with Applications, 68(10):1083–1092, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003460.

Shirali:2012:SBT

[SSM12] Mina Shirali, Nasrin Shirali, and Mohamad Reza Meybodi. Sleep-based topology control in the ad hoc networks by using fitness aware learning automata. Computers and Mathematics with Applications, 64(2):137– 146, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001289.

Sabir:2017:MMT

[SSM^{+}17] Muhammad Sabir, Abdullah Shah, Wazir Muhammad, Ijaz Ali, and Peter Bastian. A mathematical model of tumor hypoxia targeting in cancer treatment and its numerical simulation. Computers and Mathematics with Applications, 74 (12):3250–3259, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305114.

Sarikaya:2010:NIS

[SSO10] Mehmet Zeki Sarikaya, Erhan Set, and M. Emin Ozdemir. On new inequalities of Simpson’s type for s-convex functions. Computers and Mathematics with Applications, 60(8):2191–2199, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005250.

Soheili:2013:CWM

[SSP13] Ali R. Soheili, F. Soleymani, and M. D. Petković. On the computation of weighted Moore–Penrose inverse using a high-order matrix method. Computers and Mathematics with Applications, 66 (11):2344–2351, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005580.

Sivakumar:2010:EHL

[SSPL10] V. Sivakumar, S. Sivasankaran, P. Prakash, and Jinho Lee. Effect of heating location and size on mixed convection in liddriven cavities. Computers and Mathematics with Applications, 59(9):3053– 3065, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001343.

Selmanogullari:2011:TJG

[SSR11] T. Selmanogullari, Ekrem Savas, and B. E. Rhoades. Type M for E–J generalized quasi-Hausdorff matrices. Computers and Mathematics with Applications, 61(2):178–181, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008229.

Soroor:2011:AAM

[SSS^{+}11a] Javad Soroor, Sara Sajjadi, S. Sahar Sajjadi, Seyed Naeim Alavi, and Abdolmajid Soheilinia. An advanced adoption model and an algorithm of evaluation agents in automated supplier ranking. Computers and Mathematics with Applications, 62(10):3649–3662, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007140.

Subasi:2011:PGM

[SSS11b] Murat Subasi, Sidika Sule Sener, and Yesim Saraç. A procedure for the Galerkin method for a vibrating system. Computers and Mathematics with Applications, 61(9):2854– 2862, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100232X.

Saad:2016:NSC

[SSS16] Ali S. Saad, Bilal Saad, and Mazen Saad. Numerical study of compositional compressible degenerate two-phase flow in saturated-unsaturated heterogeneous porous media. Computers and Mathematics with Applications, 71(2):565–584, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005787.

Salcedo:2011:UQS

[SSSB11] Javier Salcedo, Bruno Salcedo, and F. J. Sánchez-Bernabe. The use of quadratures for solving convective and highly stiff transport problems. Computers and Mathematics with Applications, 61(3):586–591, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009090.

Strehl:2012:EAF

[SST12] Robert Strehl, Andriy Sokolov, and Stefan Turek. Efficient, accurate and flexible finite element solvers for chemotaxis problems. Computers and Mathematics with Applications, 64(3):175–189, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010959.

Smolyanov:2019:CLI

[SST19] Ivan Smolyanov, Fedor Sarapulov, and Fedor Tarasov. Calculation of linear induction motor features by detailed equivalent circuit method taking into account nonlinear electromagnetic and thermal properties. Computers and Mathematics with Applications, 78(9):3187–3199, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302780.

Sung:2012:DFM

[ST12] Wen-Tsai Sung and Ming-Han Tsai. Data fusion of multi-sensor for IOT precise measurement based on improved PSO algorithms. Computers and Mathematics with Applications, 64(5):1450–1461, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002970.

Shi:2014:SNN

[ST14] Y. Shi and G. H. Tang.

Simulation of Newtonian and non-Newtonian rheology behavior of viscous fingering in channels by the lattice Boltzmann method. Computers and Mathematics with Applications, 68 (10):1279–1291, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004301.

Safi:2015:ECH

[ST15] Mohammad Amin Safi and Stefan Turek. Efficient computations for high density ratio rising bubble flows using a diffused interface, coupled lattice Boltzmannlevel set scheme. Computers and Mathematics with Applications, 70 (6):1290–1305, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115003429.

Serghini:2016:NAS

[ST16] A. Serghini and A. Tijini. New approach to study splines by blossoming method and application to the construction of a bivariate C^{1} quartic quasi-interpolant. Computers and Mathematics with Applications, 71(2):529–543, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005829.

Shi:2018:ICD

[ST18] Y. Shi and G. H. Tang. Investigation of coalesced droplet vertical jumping and horizontal moving on textured surface using the lattice Boltzmann method. Computers and Mathematics with Applications, 75(4):1213–1225, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730682X.

Sogn:2019:RMS

[ST19] Jarle Sogn and Stefan Takacs. Robust multigrid solvers for the biharmonic problem in isogeometric analysis. Computers and Mathematics with Applications, 77(1):105–124, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305297.

Stanek:2011:EPS

[Sta11] Svatoslav Stanek. The existence of positive solutions of singular fractional boundary value problems. Computers and Mathematics with Applications, 62(3):1379–1388, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003518.

Sandev:2018:GDO

[STC18] Trifce Sandev, Zivorad Tomovski, and Bojan Crnkovic. Generalized distributed order diffusion equations with composite time fractional derivative. Computers and Mathematics with Applications, 73(6):1028–1040, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304011.

Strasser:2019:ESL

[STDLM19] Paul J. Strasser, Giordano Tierra, Burkhard Dünweg, and Mária Lukácová-Medvid’ová. Energystable linear schemes for polymer-solvent phase field models. Computers and Mathematics with Applications, 77(1):125–143, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305303.

Stensholt:2016:OOP

[Ste16] Sigvat Stensholt. Oriented ordering parameters for free energy lattice Boltzmann methods using the bounceback boundary condition. Computers and Mathematics with Applications, 72(2):404–413, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003004.

Senel:2019:DSM

[STS19] P. Senel and M. Tezer-Sezgin. DRBEM solution to MHD flow in ducts with thin slipping side walls and separated by conducting thick Hartmann walls. Computers and Mathematics with Applications, 78 (9):3165–3174, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302822.

Su:2012:PSS

[Su12] Xinwei Su. Positive solutions to singular boundary value problems for fractional functional differential equations with changing sign nonlinearity. Computers and Mathematics with Applications, 64(10):3425–3435, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200168X.

Sulman:2016:OMT

[Sul16] Mohamed H. M. Sulman.

Optimal mass transportbased adaptive mesh method for phase-field models of two-phase fluid flows. Computers and Mathematics with Applications, 72(9):2181–2193, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304837.

Sun:2010:GSE

[Sun10] Ruoyan Sun. Global stability of the endemic equilibrium of multigroup SIR models with nonlinear incidence. Computers and Mathematics with Applications, 60(8):2286–2291, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005845.

Sung:2011:SLL

[Sun11] Soo Hak Sung. On the strong law of large numbers for weighted sums of random variables. Computers and Mathematics with Applications, 62 (11):4277–4287, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008820.

Soleymani:2011:OST

[SV11] F. Soleymani and S. Karimi Vanani. Optimal Steffensentype methods with eighth order of convergence. Computers and Mathematics with Applications, 62(12):4619–4626, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009114.

Slodicka:2019:ACM

[SVP^{+}19] Marián Slodicka, Karel Van Bockstal, Iuliu Sorin Pop, Christophe Geuzaine, and Rob H. De Staelen. Advanced COmputational Methods in ENgineering (ACOMEN 2017). Computers and Mathematics with Applications, 77(6):1423–1424, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300860.

Shahrokhabadi:2016:ITC

[SVY16] Shahriar Shahrokhabadi, Farshid Vahedifard, and Shantia Yarahmadian. Integration of Thiele continued fractions and the method of fundamental solutions for solving unconfined seepage problems. Computers and Mathematics with Applications, 71(7):1479–1490, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300748.

Strijov:2010:NRM

[SW10] Vadim Strijov and Gerhard Wilhelm Weber. Nonlinear regression model generation using hyperparameter optimization. Computers and Mathematics with Applications, 60(4):981–988, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001975.

Song:2011:ESS

[SW11] Q. Q. Song and L. S. Wang. The existence of solutions for the system of vector quasiequilibrium problems in topological order spaces. Computers and Mathematics with Applications, 62(4):1979–1983, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005219.

Shu:2012:EUM

[SW12] Xiao-Bao Shu and Qianqian Wang. The existence and uniqueness of mild solutions for fractional differential equations with nonlocal conditions of order 1 < α < 2. Computers and Mathematics with Applications, 64(6):2100–2110, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200346X.

Shi:2016:SAG

[SW16a] Dongyang Shi and Junjun Wang. Superconvergence analysis of an H^{1}–Galerkin mixed finite element method for Sobolev equations. Computers and Mathematics with Applications, 72 (6):1590–1602, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304254.

Stoyanov:2016:DAS

[SW16b] Miroslav K. Stoyanov and Clayton G. Webster. A dynamically adaptive sparse grids method for quasi-optimal interpolation of multidimensional functions. Computers and Mathematics with Applications, 71(11):2449– 2465, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116000031.

Shi:2017:USA

[SW17] Dongyang Shi and Junjun Wang. Unconditional superconvergence analysis of a linearized Galerkin FEM for nonlinear hyperbolic equations. Computers and Mathematics with Applications, 74(4):634–651, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302961.

Sun:2019:DDV

[SW19] Hongquan Sun and Jinliang Wang. Dynamics of a diffusive virus model with general incidence function, cell-to-cell transmission and time delay. Computers and Mathematics with Applications, 77(1):284–301, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305443.

Swaminathan:2010:SCH

[Swa10] A. Swaminathan. Sufficient conditions for hypergeometric functions to be in a certain class of analytic functions. Computers and Mathematics with Applications, 59(4):1578–1583, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000027.

Song:2011:CRU

[SWC11] Guang-Jing Song, Qing-Wen Wang, and Hai-Xia Chang. Cramer rule for the unique solution of restricted matrix equations over the quaternion skew field. Computers and Mathematics with Applications, 61(6):1576–1589, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111000435.

Shi:2015:NFE

[SWL15] Dongyang Shi, Lele Wang, and Xin Liao. Nonconforming finite element analysis for Poisson eigenvalue problem. Computers and Mathematics with Applications, 70(5):835–845, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002849.

Sheu:2016:DSF

[SWL16] Tony W. H. Sheu, Y. C. Wang, and J. H. Li. Development of a 3D staggered FDTD scheme for solving Maxwell’s equations in Drude medium. Computers and Mathematics with Applications, 71(6):1198–1226, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300219.

She:2019:HTP

[SWL19] Ze She, Keyong Wang, and Peichao Li. Hybrid Trefftz polygonal elements for heat conduction problems with inclusions/voids. Computers and Mathematics with Applications, 78 (6):1978–1992, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301579.

Scarabosio:2019:GOA

[SWOF19] Laura Scarabosio, Barbara Wohlmuth, J. Tinsley Oden, and Danial Faghihi. Goaloriented adaptive modeling of random heterogeneous media and modelbased multilevel Monte Carlo methods. Computers and Mathematics with Applications, 78(8):2700–2718, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302123.

Shao:2019:CAL

[SWS19] Xin-Hui Shao, Hong-Yu Wu, and Hai-Long Shen. Combination of augmented Lagrangian technique and ST preconditioner for saddle point problems. Computers and Mathematics with Applications, 77(3):865–876, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306230.

Sokol:2011:CPC

[SWW11] Janusz Sokół and Agnieszka Wiśniowska-Wajnryb. On certain problem in the class of k-starlike functions. Computers and Mathematics with Applications, 62 (12):4733–4741, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009400.

Shi:2012:OTC

[SXB^{+}12] Quan Shi, Yanghua Xiao, Nik Bessis, Yiqi Lu, Yaoliang Chen, and Richard Hill. Optimizing K^{2} trees: a case for validating the maturity of network of practices. Computers and Mathematics with Applications, 63(2):427–436, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122111006353.

Shen:2011:LPG

[SXM11] Ting-Ting Shen, Kang-Zheng Xing, and He-Ping Ma. A Legendre Petrov–Galerkin method for fourth-order differential equations. Computers and Mathematics with Applications, 61(1):8–16, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008047.

Shi:2018:NAS

[SY18] Dongyang Shi and Huaijun Yang. A new approach of superconvergence analysis for twodimensional time fractional diffusion equation. Computers and Mathematics with Applications, 75(8):3012–3023, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300452.

Sahin:2011:CAS

[SYG11] Niyazi Sahin, Suayip Yüzbasi, and Mustafa Gülsu. A collocation approach for solving systems of linear Volterra integral equations with variable coefficients. Computers and Mathematics with Applications, 62(2):755–769, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100469X.

Shah:2012:NSU

[SYI12] Abdullah Shah, Li Yuan, and Shamsul Islam. Numerical solution of unsteady Navier–Stokes equations on curvilinear meshes. Computers and Mathematics with Applications, 63(11):1548–1556, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002520.

Sun:2010:FHB

[SYL10] Kaibiao Sun, Xuehai Yuan, and Hongxing Li. Fuzzy hypergroups based on fuzzy relations. Computers and Mathematics with Applications, 60(3):610–622, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003470.

Seifollahi:2012:NWS

[SYO12] Sattar Seifollahi, John Yearwood, and Bahadorreza Ofoghi. Novel weighting in single hidden layer feedforward neural networks for data classification. Computers and Mathematics with Applications, 64(2):128– 136, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000521.

Su:2011:SCT

[SYW11] Hong Su, Shui-Ping Yang, and Li-Ping Wen. Stability and convergence of the two parameter cubic spline collocation method for delay differential equations. Computers and Mathematics with Applications, 62(6):2580–2590, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006328.

Song:2013:MPR

[SYY13] Huijuan Song, Jingxue Yin, and Ying Yang.

Multiplicity of positive radial solutions for the weighted p-Laplacian in R^{n} \ {0}. Computers and Mathematics with Applications, 66(8):1475–1487, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005117.

Saporito:2019:CSL

[SYZ19] Yuri F. Saporito, Xu Yang, and Jorge P. Zubelli. The calibration of stochastic local-volatility models: an inverse problem perspective. Computers and Mathematics with Applications, 77(12):3054–3067, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300598.

Su:2011:USB

[SZ11] Xinwei Su and Shuqin Zhang. Unbounded solutions to a boundary value problem of fractional order on the halfline. Computers and Mathematics with Applications, 61(4):1079–1087, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009636.

Schmeidel:2012:ADF

[SZ12a] Ewa Schmeidel and Zenon Zbaszyniak. An application of Darbo’s fixed point theorem in the investigation of periodicity of solutions of difference equations. Computers and Mathematics with Applications, 64(7):2185–2191, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010807.

Sun:2012:ESFb

[SZ12b] Hong-Rui Sun and Quan-Guo Zhang. Existence of solutions for a fractional boundary value problem via the mountain pass method and an iterative technique. Computers and Mathematics with Applications, 64(10):3436–3443, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200137X.

Song:2014:SDD

[SZ14] Yongli Song and Xingfu Zou. Spatiotemporal dynamics in a diffusive ratio-dependent predator–prey model near a Hopf–Turing bifurcation point. Computers and Mathematics with Applications, 67(10): 1978–1997, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001758.

Shan:2017:EEP

[SZ17] Li Shan and Yuhong Zhang. Error estimates of the partitioned time stepping method for the evolutionary Stokes– Darcy flows. Computers and Mathematics with Applications, 73(4):713–726, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300019.

Szabo:2015:UFR

[Sza15]  Barna Szabó.  Unidi-

rectional fiber-reinforced composite laminae: Homogenization and localization. Computers and Mathematics with Applications, 70(7):1676–1684, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000681.

Saleem:2018:STC

[SZA^{+}18] M. Rehan Saleem, Saqib Zia, Waqas Ashraf, Ishtiaq Ali, and Shamsul Qamar. The space–time CESE scheme for shallow water equations incorporating variable bottom topography and horizontal temperature gradients. Computers and Mathematics with Applications, 75(3):933–956, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730679X.

Song:2018:ETD

[SZC^{+}18] Fangying Song, Fanhai Zeng, Wei Cai, Wen Chen, and George Em Karniadakis. Efficient two-dimensional simulations of the fractional Szabo equation with different time-stepping schemes. Computers and Mathematics with Applications, 73(6):1286–1297, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306356.

Senkerik:2010:USD

[SZDO10] Roman Senkerik, Ivan Zelinka, Donald Davendra, and Zuzana Oplatkova. Utilization of SOMA and differential evolution for robust stabilization of chaotic logistic equation. Computers and Mathematics with Applications, 60(4):1026–1037, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000249X.

Shi:2011:SAF

[SZGG11] Wuxi Shi, Mu Zhang, Wencheng Guo, and Lijin Guo. Stable adaptive fuzzy control for MIMO nonlinear systems. Computers and Mathematics with Applications, 62(7):2843–2853, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006213.

Shi:2017:HAA

[SZL^{+}17] Z. G. Shi, Y. M. Zhao, F. Liu, Y. F. Tang, F. L. Wang, and Y. H. Shi. High accuracy analysis of an H^{1}-Galerkin mixed finite element method for two-dimensional time fractional diffusion equations. Computers and Mathematics with Applications, 74(8):1903–1914, October 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122117304157.

Sun:2011:LBM

[SZP^{+}11] D. K. Sun, M. F. Zhu, S. Y. Pan, C. R. Yang, and D. Raabe. Lattice Boltzmann modeling of dendritic growth in forced and natural convection. Computers and Mathematics with Applications, 61(12):3585– 3592, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008503.

Sun:2011:QAS

[SZW11] Li Sun, Mingru Zhou, and Guangwa Wang. Quadratic approximation of solutions for boundary value problems with nonlocal boundary conditions. Computers and Mathematics with Applications, 61(6):1547–1558, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000393.

Shen:2011:RMU

[SZZ11] Wensheng Shen, Changjiang Zhang, and Jun Zhang. Relaxation method for unsteady convection–diffusion equations. Computers and Mathematics with Applications, 61(4):908– 920, February 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110009442.

Temimi:2011:SAI

[TA11] H. Temimi and A. R. Ansari. A semi-analytical iterative technique for solving nonlinear problems. Computers and Mathematics with Applications, 61(2):203–210, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110008321.

Torabi:2014:DTV

[TAA14] K. Torabi, H. Afshari, and F. Haji Aboutalebi. A DQEM for transverse vibration analysis of multiple cracked non-uniform Timoshenko beams with general boundary conditions. Computers and Mathematics with Applications, 67(3):527–541, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006664.

Tachev:2011:MBO

[Tac11] Gancho Tachev. On multiplicativity of the Bernstein operator. Computers and Mathematics with Applications, 62(8):3236–3240, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007115.

Tambue:2016:EIF

[Tam16] Antoine Tambue. An exponential integrator for finite volume discretization of a reactionadvection-diffusion equation. Computers and Mathematics with Applications, 71(9):1875–1897, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301171.

Tang:2017:BPP

[Tan17] Gusheng Tang. Blowup phenomena for a parabolic system with gradient nonlinearity under nonlinear boundary conditions. Computers and Mathematics with Applications, 74(3):360–368, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730250X.

Tani:2018:PSL

[Tan18] Mattia Tani. A preconditioning strategy for linear systems arising from nonsymmetric schemes in isogeometric analysis. Computers and Mathematics with Applications, 74(7):1690–1702, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122117303619.

Tao:2018:FNC

[Tao18] Bo Tao. Fully nonlinear capillary-gravity solitary waves under a tangential electric field, Part II: Dynamics. Computers and Mathematics with Applications, 76(4):788–798, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302876.

Thiele:2017:IHS

[TAPA^{+}17] Christopher Thiele, Mauricio Araya-Polo, Faruk O. Alpak, Beatrice Riviere, and Florian Frank. Inexact hierarchical scale separation: A two-scale approach for linear systems from discontinuous Galerkin discretizations. Computers and Mathematics with Applications, 74(8):1769–1778, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303735.

Tarasov:2017:EDF

[Tar17] Vasily E. Tarasov. Exact discretization of fractional Laplacian. Computers and Mathematics with Applications, 73(5):855–863, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300421.

Temimi:2011:ASS

[TAS11] H. Temimi, A. R. Ansari, and A. M. Siddiqui. An approximate solution for the static beam problem and nonlinear integro-differential equations. Computers and Mathematics with Applications, 62(8):3132–3139, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006894.

Tian:2010:ERT

[TB10] Yuansheng Tian and Zhanbing Bai. Existence results for the three-point impulsive boundary value problem involving fractional differential equations. Computers and Mathematics with Applications, 59(8):2601–2609, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000507.

Trask:2019:MSF

[TBP19] Nathaniel Trask, Pavel Bochev, and Mauro Perego. Mitigation of the self-force effect in unstructured PIC codes using generalized moving least squares. Computers and Mathematics with Applications, 78(2):688– 705, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306102.

Tricaud:2010:AMN

[TC10] Christophe Tricaud and YangQuan Chen. An approximate method for numerically solving fractional order optimal control problems of general form. Computers and Mathematics with Applications, 59(5):1644–1655, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005410.

Totur:2011:SSC

[TÇ11] Ümit Totur and Ibrahim Çanak. Some sufficient conditions for subsequential convergence of a sequence. Computers and Mathematics with Applications, 61(3):567–572, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009065.

Totur:2012:SGT

[TÇ12] Ümit Totur and Ibrahim Çanak. Some general Tauberian conditions for the weighted mean summability method. Computers and Mathematics with Applications, 63(5): 999–1006, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111010480.

Tang:2016:SLM

[TC16] Wenguang Tang and Shuhua Chang. A semi-Lagrangian method for the weather options of mean-reverting Brownian motion with jumpdiffusion. Computers and Mathematics with Applications, 71(5):1045–1058, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115006082.

Taranets:2018:TAB

[TC18] Roman M. Taranets and Marina Chugunova. Theoretical aspects of a binary mixture flow. Computers and Mathematics with Applications, 75(11):3957–3970, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301354.

Tasdelen:2011:MEJ

[TÇA11] Fatma Tasdelen, Bayram Çekim, and Rabia Aktas. On a multivariable extension of Jacobi matrix polynomials. Computers and Mathematics with Applications, 61(9):2412– 2423, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001052.

Tiryaki:2012:LTI

[TÇA12] Aydin Tiryaki, Devrim Çakmak, and Mustafa Fahri Aktas. Lyapunov-type inequalities for a certain class of nonlinear systems. Computers and Mathematics with Applications, 64(6):1804–1811, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001332.

Tian:2019:TGM

[TCHW19] Zhikun Tian, Yanping Chen, Yunqing Huang, and Jianyun Wang. Twogrid method for the two-dimensional timedependent Schrödinger equation by the finite element method. Computers and Mathematics with Applications, 77(12):3043–3053, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300604.

Truster:2015:AIA

[TCM15] Timothy J. Truster, Pinlei Chen, and Arif Masud. On the algorithmic and implementational aspects of a discontinuous Galerkin method at finite strains. Computers and Mathematics with Applications, 70(6):1266–1289, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003211.

Tang:2018:IMO

[TCM18] Jia Tang, Linjie Chen, and Changfeng Ma. An iterative method for obtaining the least squares solutions of quadratic inverse eigenvalue problems over generalized Hamiltonian matrix with submatrix constraints. Computers and Mathematics with Applications, 76(7):1608–1624, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303791.

Tripathy:2010:BVD

[TD10a] Binod Chandra Tripathy and Amar Jyoti Dutta. Bounded variation double sequence space of fuzzy real numbers. Computers and Mathematics with Applications, 59(2):1031–1037, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006701.

Tuna:2010:GOO

[TD10b] Adnan Tuna and Durmus Daghan. Generalization of Ostrowski and Ostrowski-Grüss type inequalities on time scales. Computers and Mathematics with Applications, 60(3):803–811, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003846.

Tolke:2013:LBM

[TDM13] Jonas Tölke, Giuseppe De Prisco, and Yaoming Mu. A lattice Boltzmann method for immiscible two-phase Stokes flow with a local collision operator. Computers and Mathematics with Applications, 65(6):864–881, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004178.

Tuan:2019:ERF

[TDN19] Nguyen Huy Tuan, Amar Debbouche, and Tran Bao Ngoc. Existence and regularity of final value problems for time fractional wave equations. Computers and Mathematics with Applications, 78 (5):1396–1414, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306898.

Tan:2018:PLM

[TDXQ18] Wei Tan, ZhengDe Dai, JingLi Xie, and De-Qing Qiu. Parameter limit method and its application in the (4 + 1)-dimensional Fokas equation. Computers and Mathematics with Applications, 75 (12):4214–4220, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301585.

Tao:2017:BPN

[TF17] Xueyan Tao and Zhong Bo Fang. Blow-up phenomena for a nonlinear reaction–diffusion system with time dependent coefficients. Computers and Mathematics with Applications, 74 (10):2520–2528, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304662.

Tsitouras:2011:MRK

[TFS11] Ch. Tsitouras, I. Th. Famelis, and T. E. Simos. On modified Runge– Kutta trees and methods. Computers and Mathematics with Applications, 62(4):2101–2111, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005396.

Tao:2011:CPE

[TG11] Qiang Tao and Hang Gao. Controllability for parabolic equations with nonlinear memory. Computers and Mathematics with Applications, 61(3):538–545, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009041.

Tao:2014:FNC

[TG14] B. Tao and D. L. Guo. Fully nonlinear capillarygravity wave patterns under the tangential electric field. Computers and Mathematics with Applications, 67(3):627–635, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006883.

Toole:2013:TMC

[TH13] Gregory Toole and Monica K. Hurdal. Turing models of cortical folding on exponentially and logistically growing domains. Computers and Mathematics with Applications, 66(9):1627–1642, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001715.

Tang:2019:APF

[TH19] Qili Tang and Yunqing Huang. Analysis of parallel finite element algorithm based on three linearization methods for the steady incompressible MHD flow. Computers and Mathematics with Applications, 78(1):35– 54, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300719.

Thakoor:2019:ASF

[Tha19a] Nawdha Thakoor. Analytical shape functions and derivatives approximation formulas in local radial point interpolation methods with applications to financial option pricing problems. Computers and Mathematics with Applications, 78 (12):3770–3789, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119303086.

Thanh:2019:WPS

[Tha19b] Bui Le Trong Thanh. On the well-posedness of a spectral fractional forward-backward pseudoparabolic equation. Computers and Mathematics with Applications, 77(2):323–333, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305479.

Taba:2012:SN

[THB12] D. Afkhami Taba, A. Hasankhani, and M. Bolurian. Soft nexuses. Computers and Mathematics with Applications, 64(6):1812–1821, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001733.

Tao:2018:OPS

[THC^{+}18] Shi Tao, Qing He, Baiman Chen, Xiaoping Yang, and Simin Huang. Onepoint second-order curved boundary condition for lattice Boltzmann simulation of suspended particles. Computers and Mathematics with Applications, 76(7):1593–1607, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830378X.

Tseng:2011:NHH

[THD11] Kuei-Lin Tseng, Shiow-Ru Hwang, and Sever S. Dragomir. New Hermite– Hadamard-type inequalities for convex functions (II). Computers and Mathematics with Applications, 62(1):401– 418, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004214.

Trong:2019:TSF

[THD19] Dang Duc Trong, Dinh Nguyen Duy Hai, and Nguyen Minh Dien. On a time-space fractional backward diffusion problem with inexact orders. Computers and Mathematics with Applications, 78(5):1572–1593, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301397.

Tang:2017:IWD

[THfL17] Xiao-Yan Tang, Xia-Zhi Hao, and Zu feng Liang. Interacting waves of Davey–Stewartson III system. Computers and Mathematics with Applications, 74(6):1311–1320, September 15, 2017. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303668.

Torregrosa:2014:SMC

[THGG14] A. J. Torregrosa, S. Hoyas, A. Gil, and J. P. G. Galache. A sparse mesh for Compact Finite Difference–Fourier solvers with radius-dependent spectral resolution in circular domains. Computers and Mathematics with Applications, 67(6):1309– 1318, April 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114000443.

Tseng:2012:HTB

[THH12] Kuei-Lin Tseng, Shiow-Ru Hwang, and Kai-Chen Hsu. Hadamard-type and bullen-type inequalities for Lipschitzian functions and their applications. Computers and Mathematics with Applications, 64(4):651–660, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011473.

Tsai:2010:MDL

[THY^{+}10] Jason Sheng-Hong Tsai, Nien-Tsu Hu, Po-Chuan Yang, Shu-Mei Guo, and Leang-San Shieh. Modeling of decentralized linear observer and tracker for a class of unknown interconnected large-scale sampled-data nonlinear systems with closedloop decoupling property. Computers and Mathematics with Applications, 60(3):541–562, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003421.

Tan:2011:IFF

[THZ^{+}11] Qing Tan, Qing He, Weizhong Zhao, Zhongzhi Shi, and E. S. Lee. An improved FCMBP fuzzy clustering method based on evolutionary programming. Computers and Mathematics with Applications, 61(4):1129–1144, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009685.

Tian:2011:RBM

[Tia11]  Yongge Tian.  Rela-

tions between matrix sets generated from linear matrix expressions and their applications. Computers and Mathematics with Applications, 61(6):1493–1501, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009843.

Tian:2017:BBT

[Tia17] Shuying Tian. Bounds for blow-up time in a semilinear parabolic problem with viscoelastic term. Computers and Mathematics with Applications, 74(4):736–743, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303073.

Tian:2019:TWM

[Tia19] Yanling Tian. Traveling waves for monotone or non-monotone equations with nonlocal delays in a cylinder. Computers and Mathematics with Applications, 78(3):958–978, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301440.

Timofte:2013:MAD

[Tim13] Claudia Timofte. Multiscale analysis of diffusion processes in composite media. Computers and Mathematics with Applications, 66(9):1573–1580, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006888.

Timofte:2014:HRI

[Tim14] Claudia Timofte. Homogenization results for ionic transport in periodic porous media. Computers and Mathematics with Applications, 68 (9):1024–1031, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001321.

Tang:2010:EMP

[TJ10] X. H. Tang and Jianchu Jiang. Existence and multiplicity of periodic solutions for a class of secondorder Hamiltonian systems. Computers and Mathematics with Applications, 59(12):3646– 3655, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002282.

Tian:2013:ESL

[TJQS13] Yingjie Tian, Xuchan Ju, Zhiquan Qi, and Yong Shi. Efficient sparse least squares support vector machines for pattern classification. Computers and Mathematics with Applications, 66 (10):1935–1947, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004331.

Tanay:2011:TSF

[TK11] Bekir Tanay and M. Burç Kandemir. Topological structure of fuzzy soft sets. Computers and Mathematics with Applications, 61(10):2952– 2957, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002197.

Tuan:2017:FRF

[TKBMT17] Nguyen Huy Tuan, Mokhtar Kirane, Bandar Bin-Mohsin, and Pham Thi Minh Tam. Filter regularization for final value fractional diffusion problem with deterministic and random noise. Computers and Mathematics with Applications, 74 (6):1340–1361, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303620.

Teng:2010:CPP

[TKH10] Yuan-Hsiang Teng, Tzu-Liang Kung, and Lih-Hsing Hsu. The 3^{∗}- connected property of pyramid networks. Computers and Mathematics with Applications, 60(8):2360–2363, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005936.

Tuan:2018:IRU

[TKHL18] Nguyen Huy Tuan, Mokhtar Kirane, Luu Vu Cam Hoan, and Le Dinh Long. Identification and regularization for unknown source for a timefractional diffusion equation. Computers and Mathematics with Applications, 73(6):931–950, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305582.

Tang:2010:DIU

[TL10a] Yi-Ming Tang and Xiao-Ping Liu. Differently implicational universal triple I method of (1, 2,

2) type. Computers and Mathematics with Applications, 59(6):1965–1984, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007469.

Tang:2010:ANN

[TL10b] Ying Tang and Jianping Li. Another neural network based approach for computing eigenvalues and eigenvectors of real skew-symmetric matrices. Computers and Mathematics with Applications, 60(5):1385–1392, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004347.

Tien:2012:HTC

[TL12] Jia-Ping Tien and Tzuu-Hseng S. Li. Hybrid taguchi-chaos of multilevel immune and the artificial bee colony algorithm for parameter identification of chaotic systems. Computers and Mathematics with Applications, 64(5):1108–1119, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002349.

Taoum:2018:DAH

[TL18] Samer Taoum and Emmanuel Lefrançois. Dual analysis for heat exchange: Application to thermal bridges. Computers and Mathematics with Applications, 75(10):3471–3487, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300841.

Taddei:2017:ASR

[TLR17] L. Taddei, N. Lebaal, and S. Roth. Axissymmetrical Riemann problem solved with standard SPH method. development of a polar formulation with artificial viscosity. Computers and Mathematics with Applications, 74 (12):3161–3174, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122117305035.

Tsitouras:2010:EHO

[TM10] Ch. Tsitouras and Ch. G.

Massouros. On enumeration of hypergroups of order 3. Computers and Mathematics with Applications, 59(1):519–523, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003836.

Tang:2012:ADG

[TM12] Jia Tang and Changfeng Ma. An application of H differentiability to generalized complementarity problems over symmetric cones. Computers and Mathematics with Applications, 63(1):14–24, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009102.

Tang:2017:GCD

[TM17] Jia Tang and Chang-Feng Ma. Generalized conjugate direction method for solving a class of generalized coupled Sylvester-conjugate transpose matrix equations over generalized Hamiltonian matrices. Computers and Mathematics with Applications, 74(12):3303–3317, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305102.

Tang:2018:DBG

[TM18] De Tang and Li Ma. Dynamical behavior of a general reaction–diffusion– advection model for two competing species. Computers and Mathematics with Applications, 75(4):1128–1142, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306843.

Tiba:2019:OPH

[TM19] Dan Tiba and Cornel Marius Murea. Optimization of a plate with holes. Computers and Mathematics with Applications, 77(11):3010– 3020, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118304632.

Tsukerman:2019:TAC

[TMCM19] Igor Tsukerman, Shampy Mansha, Y. D. Chong, and Vadim A. Markel. Trefftz approximations in complex media: Accuracy and applications. Computers and Mathematics with Applications, 77(6):1770–1785, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304978.

Tien:2016:NSC

[TMDTTC16] C. M. T. Tien, N. Mai-Duy, C.-D. Tran, and T. Tran-Cong. A numerical study of compact approximations based on flat integrated radial basis functions for secondorder differential equations. Computers and Mathematics with Applications, 72(9):2364–2387, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630503X.

Touzeau:2019:RMI

[TMLF19] Clément Touzeau, Benoît Magnain, Gilles Lubineau, and Eric Florentin. Robust method for identifying material parameters based on virtual fields in elastodynamics. Computers and Mathematics with Applications, 77(11):3021– 3042, June 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304231.

Tajdin:2010:CFS

[TMMASG10] Ali Tajdin, Iraj Mahdavi, Nezam Mahdavi-Amiri, and Bahram Sadeghpour-Gildeh. Computing a fuzzy shortest path in a network with mixed fuzzy arc lengths using α-cuts. Computers and Mathematics with Applications, 60(4):989–1002, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002270.

Trigeassou:2013:ISA

[TMO13] J. C. Trigeassou, N. Maamri, and A. Oustaloup. The infinite state approach: Origin and necessity. Computers and Mathematics with Applications, 66(5):892–907, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006815.

Trigeassou:2012:SVT

[TMSO12] J. C. Trigeassou, N. Maamri, J. Sabatier, and A. Oustaloup. State variables and transients of fractional order differential systems. Computers and Mathematics with Applications, 64(10):3117–3140, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003173.

Tian:2015:WCP

[TMZ^{+}15] Wei Tian, Tinghuai Ma, Yuhui Zheng, Xin Wang, Yuan Tian, Abdullah Al-Dhelaan, and Mznah Al-Rodhaan. Weighted curvature-preserving PDE image filtering method. Computers and Mathematics with Applications, 70(6):1336–1344, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115003570.

Taghizadeh:2011:NCS

[TN11] N. Taghizadeh and A. Neirameh. New complex solutions for some special nonlinear partial differential systems. Computers and Mathematics with Applications, 62(4):2037–2044, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005268.

Teng:2011:PSG

[TNF11] Zhidong Teng, Linfei Nie, and Xining Fang. The periodic solutions for general periodic impulsive population systems of functional differential equations and its applications. Computers and Mathematics with Applications, 61(9):2690– 2703, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001751.

Tuan:2019:EUM

[TNHK19] Nguyen Huy Tuan, Tran Bao Ngoc, Le Nhat Huynh, and Mokhtar Kirane. Existence and uniqueness of mild solution of timefractional semilinear differential equations with a nonlocal final condition. Computers and Mathematics with Applications, 78(5):1651–1668, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306606.

Tarzanagh:2017:NPC

[TNP17] D. Ataee Tarzanagh, P. Nazari, and M. Reza Peyghami. A nonmonotone PRP conjugate gradient method for solving square and under-determined systems of equations. Computers and Mathematics with Applications, 73(2):339–354, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306587.

TeCho:2012:BBE

[TNT12] Jakkrit TeCho, Cholwich Nattee, and Thanaruk Theeramunkong. Boosting-based ensemble learning with penalty profiles for automatic Thai unknown word recognition. Computers and Mathematics with Applications, 63(6):1117–1134, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011503.

Tuan:2019:BPK

[TNV19] Nguyen Huy Tuan, Danh Hua Quoc Nam, and Thi Minh Nhat Vo. On a backward problem for the Kirchhoff’s model of parabolic type. Computers and Mathematics with Applications, 77(1):15–33, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305054.

Tuneski:2011:SPC

[TO11] Nikola Tuneski and Milutin Obradović. Some properties of certain expressions of analytic functions. Computers and Mathematics with Applications, 62(9):3438–3445, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122111007437.

Todinov:2012:TOR

[Tod12] M. T. Todinov. Topology optimisation of repairable flow networks for a maximum average availability. Computers and Mathematics with Applications, 64(12):3729–3746, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001757.

Todorov:2013:ORS

[Tod13] Todor D. Todorov. The optimal refinement strategy for 3-D simplicial meshes. Computers and Mathematics with Applications, 66(7):1272–1283, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300477X.

Todorov:2015:NPG

[Tod15] Todor D. Todorov. Nonlocal problem for a general second-order elliptic operator. Computers and Mathematics with Applications, 69(5):411–422, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000024.

Todorov:2018:DPN

[Tod18] Todor D. Todorov. Dirichlet problem for a nonlocal p-Laplacian elliptic equation. Computers and Mathematics with Applications, 76 (6):1261–1274, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303432.

Toloo:2012:ASC

[Tol12] Mehdi Toloo. Alternative solutions for classifying inputs and outputs in data envelopment analysis. Computers and Mathematics with Applications, 63(6):1104–1110, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010716.

Tomasiello:2011:NTN

[Tom11] S. Tomasiello. A note on three numerical procedures to solve Volterra integro-differential equations in structural analysis. Computers and Mathematics with Applications, 62(8):3183–3193, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006948.

Tomar:2013:RAM

[Tom13] S. K. Tomar. Robust algebraic multilevel preconditioning in H (curl) and H (div). Computers and Mathematics with Applications, 66(6):1024–1046, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004239.

Taghipour:2018:FCM

[TPHD18] Aliakbar Taghipour, Jamshid Parvizian, Stephan Heinze, and Alexander Düster. The finite cell method for nearly incompressible finite strain plasticity problems with complex geometries. Computers and Mathematics with Applications, 75(9):3298–3316, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118300646.

Taghilou:2014:ITP

[TR14] Mohammad Taghilou and Mohammad Hassan Rahimian. Investigation of twophase flow in porous media using lattice Boltzmann method. Computers and Mathematics with Applications, 67(2):424–436, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004860.

Treanta:2018:HOH

[Tre18] Savin Treanta. Higherorder Hamilton dynamics and Hamilton–Jacobi divergence PDE. Computers and Mathematics with Applications, 75(2):547–560, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305965.

Tripathi:2011:PTF

[Tri11] Dharmendra Tripathi. Peristaltic transport of fractional Maxwell fluids in uniform tubes: Applications in endoscopy.

Computers and Mathematics with Applications, 62(3):1116–1126, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001908.

Turjanicova:2019:HBT

[TRL19] Jana Turjanicová, Eduard Rohan, and Vladimír Lukes. Homogenization based two-scale modelling of ionic transport in fluid saturated deformable porous media. Computers and Mathematics with Applications, 78(9):3211–3235, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302913.

Trouette:2013:LBS

[Tro13] Benoît Trouette. Lattice Boltzmann simulations of a time-dependent natural convection problem. Computers and Mathematics with Applications, 66(8):1360–1371, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122113004756.

Truong:2019:NMF

[Tru19] Le Xuan Truong. The Nehari manifold for fractional p-Laplacian equation with logarithmic nonlinearity on whole space. Computers and Mathematics with Applications, 78(12):3931– 3940, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211930327X.

Tari:2011:DTM

[TS11a] A. Tari and S. Shahmorad. Differential transform method for the system of twodimensional nonlinear Volterra integro-differential equations. Computers and Mathematics with Applications, 61(9):2621– 2629, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001593.

Tomovski:2011:EFF

[TS11b] Zivorad Tomovski and Trifce Sandev. Effects of a fractional friction with power-law memory kernel on string vibrations. Computers and Mathematics with Applications, 62(3):1554–1561, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003452.

Thiagarajan:2014:AWN

[TS14] Vaidyanathan Thiagarajan and Vadim Shapiro. Adaptively weighted numerical integration over arbitrary domains. Computers and Mathematics with Applications, 67(9): 1682–1702, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122114000984.

Toure:2016:SFE

[TS16] Mamadou Kabirou Touré and Azzeddine Soulaïmani. Stabilized finite element methods for solving the level set equation without reinitialization. Computers and Mathematics with Applications, 71(8):1602– 1623, April 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116300827.

Tebini:2016:AAM

[TSB16] Sondes Tebini, Hassene Seddik, and Ezzedine Ben Braiek. An advanced and adaptive mathematical function for an efficient anisotropic image filtering. Computers and Mathematics with Applications, 72(5):1369–1385, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303819.

Tsitouras:2011:RKP

[Tsi11] Ch. Tsitouras. Runge– Kutta pairs of order 5(4) satisfying only the first column simplifying assumption. Computers and Mathematics with Applications, 62(2):770– 775, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004706.

Tso:2013:NWG

[Tso13] Raylin Tso. A new way to generate a ring: Universal ring signature. Computers and Mathematics with Applications, 65(9):1350– 1359, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000491.

Tsai:2012:NCD

[TT12] Shun-Hung Tsai and Yu-Hsiang Tseng. A novel color detection method based on HSL color space for robotic soccer competition. Computers and Mathematics with Applications, 64(5):1291–1300, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002787.

Tohidi:2014:CAB

[TT14] Emran Tohidi and Faezeh Toutounian. Convergence analysis of Bernoulli matrix approach for onedimensional matrix hyperbolic equations of the first order. Computers and Mathematics with Applications, 68(1–2):1– 12, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002004.

Turgut:2014:CQB

[TTC14] Oguz Emrah Turgut, Mert Sinan Turgut, and Mustafa Turhan Coban. Chaotic quantum behaved particle swarm optimization algorithm for solving nonlinear system of equations. Computers and Mathematics with Applications, 68(4):508–530, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002740.

Tang:2016:LSI

[TTG16] Yaning Tang, Siqiao Tao, and Qing Guan. Lump solitons and the interaction phenomena of them for two classes of nonlinear evolution equations. Computers and Mathematics with Applications, 72(9):2334–2342, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304813.

Tang:2019:GDR

[TTM19] Sitian Tang, Zhidong Teng, and Hui Miao. Global dynamics of a reaction–diffusion virus infection model with humoral immunity and nonlinear incidence. Computers and Mathematics with Applications, 78(3):786–806, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301233.

Taheri-Tolgari:2012:IMI

[TTMJ12] Javad Taheri-Tolgari, Abolfazl Mirzazadeh, and Fariborz Jolai. An inventory model for imperfect items under inflationary conditions with considering inspection errors. Computers and Mathematics with Applications, 63(6):1007–1019, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008285.

Tajiri:2010:DSS

[TTT10] Shinsuke Tajiri, Michihisa Tsutahara, and Hisao Tanaka. Direct simulation of sound and underwater sound generated by a water drop hitting a water surface using the finite difference lattice Boltzmann method. Computers and Mathematics with Applications, 59(7):2411–2420, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006270.

Tu:2016:PWS

[TTX^{+}16] Jian-Min Tu, Shou-Fu Tian, Mei-Juan Xu, Pan-Li Ma, and Tian-Tian Zhang. On periodic wave solutions with asymptotic behaviors to a (3 + 1)-dimensional generalized B-type Kadomtsev– Petviashvili equation in fluid dynamics. Computers and Mathematics with Applications, 72 (9):2486–2504, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116305053.

Tian:2018:IMS

[TTZW18] Zhaolu Tian, Maoyi Tian, Yan Zhang, and Pihua Wen. An iteration method for solving the linear system Ax = b. Computers and Mathematics with Applications, 75(8):2710–2722, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300075.

Turkyilmazoglu:2010:SSN

[Tur10] M. Turkyilmazoğlu. Series solution of nonlinear two-point singularly perturbed boundary layer problems. Computers and Mathematics with Applications, 60(7):2109–2114, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005444.

Taskara:2011:PSR

[TUT11] N. Taskara, K. Uslu, and D. T. Tollu. The periodicity and solutions of the rational difference equation with periodic coefficients. Computers and Mathematics with Applications, 62(4):1807–1813, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005037.

Tu:2018:BAS

[TW18] Xuemin Tu and Bin Wang. A BDDC algorithm for the Stokes problem with weak Galerkin discretizations. Computers and Mathematics with Applications, 76(2):377– 392, July 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211830230X.

Taylan:2010:FPE

[TWLYÖ10] Pakize Taylan, Gerhard-Wilhelm Weber, Lian Liu, and Fatma Yerlikaya-Özkurt. On the foundations of parameter estimation for generalized partial linear models with B-splines and continuous optimization. Computers and Mathematics with Applications, 60(1):134– 143, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003111.

Wang:2011:MAB

[tWqLzGkP11] Yu ting Wang, Jun qing Li, Kai zhou Gao, and Quan ke Pan. Memetic algorithm based on improved inver-over operator and Lin–Kernighan local search for the Euclidean traveling salesman problem. Computers and Mathematics with Applications, 62(7):2743–2754, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005530.

Tian:2019:FCS

[TXL19] Yan Tian, Shuhuang Xiang, and Guidong Liu. Fast computation of the spectral differentiation by the fast multipole method. Computers and Mathematics with Applications, 78(1):240–253, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300987.

Takahashi:2010:IUF

[TXZ^{+}10] Daisuke Takahashi, Yang Xiao, Yan Zhang, Periklis Chatzimisios, and Hsiao-Hwa Chen. IEEE 802.11 user fingerprinting and its applications for intrusion detection. Computers and Mathematics with Applications, 60(2):307– 318, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110000131.

Tofighi:2013:NSS

[TY13] Nima Tofighi and Mehmet Yildiz. Numerical simulation of single droplet dynamics in three-phase flows using ISPH. Computers and Mathematics with Applications, 66(4):525–536, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003143.

Truong:2016:CNH

[TY16] Le Xuan Truong and Nguyen Van Y. On a class of nonlinear heat equations with viscoelastic term. Computers and Mathematics with Applications, 72(1):216– 232, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302371.

Tang:2012:EDP

[TYY^{+}12] Feilong Tang, Ilsun You, Shui Yu, Cho-Li Wang, Minyi Guo, and Wenlong Liu. An efficient deadlock prevention approach for service oriented transaction processing. Computers and Mathematics with Applications, 63(2):458–468, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006389.

Tian:2013:HBA

[TZ13] Canrong Tian and Lai Zhang. Hopf bifurcation analysis in a diffusive food-chain model with time delay. Computers and Mathematics with Applications, 66 (10):2139–2153, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005439.

Tang:2015:NEP

[TZ15] Yaning Tang and Weijian Zai. New exact periodic solitary-wave solutions for the (3 + 1)- dimensional generalized KP and BKP equations. Computers and Mathematics with Applications, 70(10):2432–2441, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004472.

Tang:2018:ODM

[TZ18] Wenguang Tang and Shuhua Zhang. Optimal decision model and solution for carbon sequestration by afforestation. Computers and Mathematics with Applications, 76 (10):2484–2495, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304747.

Teng:2010:PCG

[TZG10] Zhidong Teng, Yu Zhang, and Shujing Gao. Permanence criteria for general delayed discrete nonautonomous n-species Kolmogorov systems and its applications. Computers and Mathematics with Applications, 59(2):812– 828, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S089812210900710X.

Tong:2012:BED

[TZMZ12] Qiaoling Tong, Qiao Zhang, Run Min, and Xuecheng Zou. Bayesian estimation in dynamic framed slotted ALOHA algorithm for RFID system. Computers and Mathematics with Applications, 64(5):1179–1186, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002659.

Tang:2010:PSS

[TZWM10] Hengsheng Tang, Zhengqiu Zhang, Zhicheng Wang, and Manjun Ma. Periodic solutions for a system of the first order nonautonomous differential equations. Computers and Mathematics with Applications, 60(7):1948–1958, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005201.

Tian:2011:ICM

[TZXP11] Yan Tian, Kun Zhao, Yiping Xu, and Fuyuan Peng. An image compression method based on the multi-resolution characteristics of BEMD. Computers and Mathematics with Applications, 61(8):2142–2147, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006838.

Tang:2011:LTI

[TZZ11] X. H. Tang, Qi-Ming Zhang, and Meirong Zhang. Lyapunov-type inequalities for the firstorder nonlinear Hamiltonian systems. Computers and Mathematics with Applications, 62 (9):3603–3613, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007796.

Uddin:2016:NSG

[UABK16] M. J. Uddin, Yasser Alginahi, O. Anwar Bég, and M. N. Kabir. Numerical solutions for gyrotactic bioconvection in nanofluid-saturated porous media with Stefan blowing and multiple slip effects. Computers and Mathematics with Applications, 72 (10):2562–2581, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305302.

Ularu:2011:TIO

[UBF11] Nicoleta Ularu, Daniel Breaz, and B. A. Frasin. Two integral operators on the class N (β). Computers and Mathematics with Applications, 62 (6):2551–2554, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006286.

Ullah:2016:NVA

[UCK16] Asmat Ullah, Wen Chen, and Mushtaq Ahmad Khan. A new variational approach for restoring images with multiplicative noise. Computers and Mathematics with Applications, 71(10):2034– 2050, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301651.

Mahmood-ul-Hassan:2012:MAE

[uHS12] Mahmood ul Hassan and A. M. Siddiqui. A modified approach to exact solutions of a general form of a non-Newtonian second grade fluid. Computers and Mathematics with Applications, 63(1):94–99, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100962X.

Siraj-ul-Islam:2015:NST

[uIAA15] Siraj ul Islam, Imran Aziz, and Masood Ahmad. Numerical solution of two-dimensional elliptic PDEs with nonlocal boundary conditions. Computers and Mathematics with Applications, 69(3):180–205, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005896.

Siraj-ul-Islam:2010:CSN

[uIAH10] Siraj ul Islam, Imran Aziz, and Fazal Haq. A comparative study of numerical integration based on Haar wavelets and hybrid functions. Computers and Mathematics with Applications, 59(6):2026–2036, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007524.

Uddin:2015:LGA

[UKA15] M. J. Uddin, M. N. Kabir, and Y. M. Alginahi. Lie group analysis and numerical solution of magnetohydrodynamic free convective slip flow of micropolar fluid over a moving plate with heat transfer. Computers and Mathematics with Applications, 70(5):846–856, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002886.

Usama:2010:CBS

[UKAL10] Muhammad Usama, Muhammad Khurram Khan, Khaled Alghathbar, and Changhoon Lee. Chaosbased secure satellite imagery cryptosystem. Computers and Mathematics with Applications, 60(2):326–337, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000064.

Ugurlu:2011:CTS

[UKI11] Yavuz Ugurlu, Dogan Kaya, and Ibrahim E. Inan. Comparison of three semi-analytical methods for solving (1 + 1)-dimensional dispersive long wave equations. Computers and Mathematics with Applications, 61(5):1278–1290, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009314.

Uzunca:2017:OCC

[UKYK17] Murat Uzunca, Tugba Küçükseyhan, Hamdullah Yücel, and Bülent Karasözen. Optimal control of convective FitzHugh–Nagumo equation. Computers and Mathematics with Applications, 73(9):2151–2169, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301074.

Unsal:2016:LSP

[ÜM16] Ömer Ünsal and Wen-Xiu Ma. Linear superposition principle of hyperbolic and trigonometric function solutions to generalized bilinear equations. Computers and Mathematics with Applications, 71(6):1242–1247, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300487.

Uga:2013:SED

[UMLF13] Kalu Chibueze Uga, Misun Min, Taehun Lee, and Paul F. Fischer. Spectral-element discontinuous Galerkin lattice Boltzmann simulation of flow past two cylinders in tandem with an exponential time integrator. Computers and Mathematics with Applications, 65(2):239–251, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011308.

Ugail:2011:BST

[UMY11] H. Ugail, M. C. Márquez, and A. Yilmaz. On Bézier surfaces in threedimensional Minkowski space. Computers and Mathematics with Applications, 62(8):2899–2912, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006407.

Unsal:2018:CSD

[Üns18] Ömer Ünsal. Complexiton solutions for (3 + 1) dimensional KdV-type equation. Computers and Mathematics with Applications, 75(7):2466–2472, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307903.

urRehman:2011:NBV

[uRK11] Mujeeb ur Rehman and Rahmat Ali Khan. A note on boundary value problems for a coupled system of fractional differential equations. Computers and Mathematics with Applications, 61(9):2630– 2637, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001611.

Uzer:2010:MTC

[Uze10] Ali Uzer. Multiplicative type complex calculus as an alternative to the classical calculus. Computers and Mathematics with Applications, 60(10):2725–2737, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006747.

Vanani:2010:NSD

[VA10] S. Karimi Vanani and A. Aminataei. On the numerical solution of differential equations of Lane–Emden type. Computers and Mathematics with Applications, 59(8):2815–2820, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000878.

Vanani:2011:TAS

[VA11] S. Karimi Vanani and A. Aminataei. Tau approximate solution of fractional partial differential equations. Computers and Mathematics with Applications, 62(3):1075–1083, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111001659.

Venkatesh:2012:LWM

[VAB12] S. G. Venkatesh, S. K. Ayyaswamy, and S. Raja Balachandar. The Legendre wavelet method for solving initial value problems of Bratu-type. Computers and Mathematics with Applications, 63(8):1287–1295, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011400.

Vacca:2018:VEM

[Vac18] Giuseppe Vacca. Virtual Element Methods for hyperbolic problems on polygonal meshes. Computers and Mathematics with Applications, 74(5):882–898, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302103.

Videla:2019:ADR

[VAK^{+}19] Javier Videla, Cosmin Anitescu, Tahsin Khajah, Stéphane P. A. Bordas, and Elena Atroshchenko. h- and p-adaptivity driven by recovery and residualbased error estimators for PHT-splines applied to time-harmonic acoustics. Computers and Mathematics with Applications, 77(9):2369–2395, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307247.

Valli:2018:PFD

[VAS^{+}18] Andrea M. P. Valli, Regina C. Almeida, Isaac P. Santos, Lucia Catabriga, Sandra M. C. Malta, and Alvaro L. G. A. Coutinho. A parameterfree dynamic diffusion method for advection– diffusion–reaction problems. Computers and Mathematics with Applications, 75(1):307–321, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305680.

Vazquez:2016:NDI

[Váz16]  R. Vázquez.  A new

design for the implementation of isogeometric analysis in Octave and Matlab: GeoPDEs 3.0. Computers and Mathematics with Applications, 72(3):523–554, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302681.

Vasant:2010:HPS

[VB10a] P. Vasant and N. Barsoum. Hybrid pattern search and simulated annealing for fuzzy production planning problems. Computers and Mathematics with Applications, 60(4):1058–1067, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002531.

See [VBW10b].

Vidhya:2010:SSP

[VB10b] C. Vidhya and P. Balasubramaniam. On the stability of the stochastic parabolic Itô equation with delay and Markovian jump. Computers and Mathematics with Applications, 60(7):1959–1963, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122110005213.

Villafuerte:2010:RDO

[VBCJ10] L. Villafuerte, C. A.

Braumann, J.-C. Cortés, and L. Jódar. Random differential operational calculus: Theory and applications. Computers and Mathematics with Applications, 59(1):115–125, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006531.

Vijayakumar:2013:CKD

[VBK13] P. Vijayakumar, S. Bose, and A. Kannan. Centralized key distribution protocol using the greatest common divisor method. Computers and Mathematics with Applications, 65(9):1360– 1368, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200048X.

Vasant:2010:CSI

[VBW10a] Pandian Vasant, Nader Barsoum, and Jeffrey Frank Webb. Corrigendum to “Special Issue: PCO’2010 — Gold Coast, Australia 2–4th December, 2010, 3rd Global Conference on Power Control Optimization” [CAMWA 60 (4) (2010) 925–1134]. Computers and Mathematics with Applications, 60:2898, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007613.

See [VBW10b].

Vasant:2010:SIG

[VBW10b] Pandian Vasant, Nader Barsoum, and Jeffrey Frank Webb. Special issue for 3rd Global Conference on Power Control Optimization in Gold Coast, Australia 2–4th December 2010. Computers and Mathematics with Applications, 60(4):925, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002701.

See corrigendum [VBW10a].

Varnosafaderani:2012:TDW

[VC12] Sima Varnosafaderani and Stefanka Chukova. A two-dimensional warranty servicing strategy based on reduction in product failure intensity. Computers and Mathematics with Applications, 63(1):201–213, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009771.

Velimirovic:2010:CWE

[VĆC10] Ljubica S. Velimirović, Marija S. Ćirić, and Milica D. Cvetković. Change of the Willmore energy under infinitesimal bending of membranes. Computers and Mathematics with Applications, 59(12):3679–3686, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002592.

Vahdani:2011:ORR

[VCM11] Hashem Vahdani, Stefanka Chukova, and Hashem Mahlooji. On optimal replacement-repair policy for multi-state deteriorating products under renewing free replacement warranty. Computers and Mathematics with Applications, 61(4):840–850, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009375.

Velimirovic:2011:WES

[VĆV11] Ljubica S. Velimirović, Marija S. Ćirić, and Nikola M. Velimirović. On the Willmore energy of shells under infinitesimal deformations. Computers and Mathematics with Applications, 61(11):3181–3190, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001878.

Volkov:2013:SNP

[VDB13] E. A. Volkov, A. A. Dosiyev, and S. C. Buranay. On the solution of a nonlocal problem. Computers and Mathematics with Applications, 66(3):330–338, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300312X.

Vitanov:2013:TWS

[VDV13] Nikolay K. Vitanov, Zlatinka I. Dimitrova, and Kaloyan N. Vitanov. Traveling waves and statistical distributions connected to systems of interacting populations. Computers and Mathematics with Applications, 66(9):1666–1684, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001892.

vanderWalt:2014:LQL

[vdW14] Jan Harm van der Walt.

Linear and quasi-linear spaces of set-valued maps. Computers and Mathematics with Applications, 68(9):1006–1015, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001308.

Veliov:2015:NAO

[Vel15] V. M. Veliov. Numerical approximations in optimal control of a class of heterogeneous systems. Computers and Mathematics with Applications, 70(11):2652–2660, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002102.

Verma:2008:RPP

[Ver08] Ram U. Verma. The overrelaxed proximal point algorithm based on Hmaximal monotonicity design and applications. Computers and Mathematics with Applications, 55(11):2673–2679, June 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122107007663.

See corrigendum [Ver12].

Verma:2012:CRP

[Ver12] Ram U. Verma. Corrigendum to “The overrelaxed proximal point algorithm based on Hmaximal monotonicity design and applications” [Comput. Math. Appl. 55 (2008) 2673–2679]. Computers and Mathematics with Applications, 64(9):2961–2963, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/

pii/S0898122112005652.

See [Ver08].

Velechovsky:2019:DAS

[VFM19] Jan Velechovsky, Marianne François, and Thomas Masser. Direction-aware slope limiter for threedimensional cubic grids with adaptive mesh refinement. Computers and Mathematics with Applications, 78(2):670– 687, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302943.

Vitillo:2015:ASS

[VGC^{+}15] F. Vitillo, C. Galati, L. Cachon, E. Laroche, and P. Millan. An anisotropic shear stress transport (ASST) model formulation. Computers and Mathematics with Applications, 70 (9):2238–2251, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004022.

Valero:2016:CEN

[VGK^{+}16] J. Valero, A. Giménez, O. V. Kapustyan, P. O. Kasyanov, and J. M.

Amigó. Convergence of equilibria for numerical approximations of a suspension model. Computers and Mathematics with Applications, 72(4):856–878, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303145.

Vu-Huu:2018:PBA

[VHPVNXW18] T. Vu-Huu, P. Phung-Van, H. Nguyen-Xuan, and M. Abdel Wahab. A polytree-based adaptive polygonal finite element method for topology optimization of fluidsubmerged breakwater interaction. Computers and Mathematics with Applications, 76 (5):1198–1218, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303328.

Viana:2015:LWP

[Via15] Arlúcio Viana. Local wellposedness for a Lotka– Volterra system in Besov spaces. Computers and Mathematics with Applications, 69(7):667–674, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000668.

Vlad:2015:BST

[VJM15] Iulian T. Vlad, Pablo Juan, and Jorge Mateu. Bayesian spatio-temporal prediction of cancer dynamics. Computers and Mathematics with Applications, 70(5):857–868, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002928.

Volna:2013:MCB

[VKJ13] Eva Volna, Martin Kotyrba, and Robert Jarusek. Multi-classifier based on Elliott wave’s recognition. Computers and Mathematics with Applications, 66(2):213–225, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000291.

Vadillo:2012:FOS

[VLFS12] G. Vadillo, J. A. Loya, and J. Fernández-Sáez.

First order solutions for the buckling loads of weakened Timoshenko columns. Computers and Mathematics with Applications, 64(8):2395–2407, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003999.

Violeau:2018:SPS

[VLJH18] Damien Violeau, Agnès Leroy, Antoine Joly, and Alexis Hérault. Spectral properties of the SPH Laplacian operator. Computers and Mathematics with Applications, 75(10):3649–3662, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301044.

Vazquez-Mendez:2019:OMU

[VMAVGCM19] M. E. Vázquez-Méndez, L. J. Alvarez-Vázquez, N. García-Chan, and A. Martínez. Optimal management of an urban road network with an environmental perspective. Computers and Mathematics with Applications, 77(6):1786–1797, March 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303456.

Vecil:2014:PDS

[VMC^{+}14] Francesco Vecil, José M.

Mantas, María J. Cáceres, Carlos Sampedro, Andrés Godoy, and Francisco Gámiz. A parallel deterministic solver for the Schrödinger–Poisson– Boltzmann system in ultra-short DG–MOSFETs: Comparison with Monte-Carlo. Computers and Mathematics with Applications, 67(9):1703–1721, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000923.

Vlasiuk:2018:FHD

[VMFF18] O. Vlasiuk, T. Michaels, N. Flyer, and B. Fornberg. Fast high-dimensional node generation with variable density. Computers and Mathematics with Applications, 76(7):1739–1757, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122118303912.

Victor:2010:RPT

[VMO10] Stéphane Victor, Pierre Melchior, and Alain Oustaloup. Robust path tracking using flatness for fractional linear MIMO systems: a thermal application. Computers and Mathematics with Applications, 59(5):1667–1678, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109005434.

Varsakelis:2015:PMT

[VMP15] C. Varsakelis, D. Monsorno, and M. V. Papalexandris. Projection methods for two velocitytwo pressure models for flows of heterogeneous mixtures. Computers and Mathematics with Applications, 70(5):1024–1045, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003090.

Vong:2011:NSO

[Von11] Seak Weng Vong. A note on some Ostrowski-like type inequalities. Computers and Mathematics with Applications, 62(1): 532–535, July 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111004354.

Vondrejc:2019:DGQ

[Von19] Jaroslav Vondrejc. Doublegrid quadrature with interpolation–projection (DoGIP) as a novel discretisation approach: an application to FEM on simplexes. Computers and Mathematics with Applications, 78 (11):3501–3513, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302846.

Valenzuela:2011:NAE

[VP11] O. Valenzuela and M. Pasadas.

A new approach to estimate the interpolation error of fuzzy data using similarity measures of fuzzy numbers. Computers and Mathematics with Applications, 61(6):1633–1645, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000496.

Vajravelu:2011:DCR

[VPR11] K. Vajravelu, K. V. Prasad, and N. S. Prasanna Rao. Diffusion of a chemically reactive species of a power-law fluid past a stretching surface. Computers and Mathematics with Applications, 62(1):93–108, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003580.

Vujakovic:2011:SNR

[VRD11] Jelena Vujaković, Miloje Rajović, and Dragan Dimitrovski. Some new results on a linear equation of the second order. Computers and Mathematics with Applications, 61(7):1837–1843, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000988.

Varma:2012:GSO

[VSI12] Serhan Varma, Sezgin Sucu, and Gürhan Içöz. Generalization of Szasz operators involving Brenke type polynomials. Computers and Mathematics with Applications, 64(2):121–127, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000351.

Varma:2011:BMP

[VT11] Serhan Varma and Fatma Tasdelen. Biorthogonal matrix polynomials related to Jacobi matrix polynomials. Computers and Mathematics with Applications, 62 (10):3663–3668, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007474.

Vitanov:2014:PDP

[VV14] Nikolay K. Vitanov and Kaloyan N. Vitanov. Population dynamics in presence of state dependent fluctuations. Computers and Mathematics with Applications, 68(9):962–971, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001291.

Vondrejc:2014:FBG

[VZM14] Jaroslav Vondrejc, Jan Zeman, and Ivo Marek. An FFT-based Galerkin method for homogenization of periodic media. Computers and Mathematics with Applications, 68(3):156–173, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002077.

Wucher:2019:MMS

[WA19] B. Wucher and L. Arbaoui. Multiscale modeling of the sintering process of printed nanoinks. Computers and Mathematics with Applications, 78(7):2325–2337, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303894.

Wachspress:2010:RBC

[Wac10] E. Wachspress. Rational bases for convex polyhedra. Computers and Mathematics with Applications, 59(6):1953–1956, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007433.

Wachspress:2011:BCP

[Wac11] Eugene L. Wachspress.

Barycentric coordinates for polytopes. Computers and Mathematics with Applications, 61(11): 3319–3321, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111003336.

Wang:2014:SFT

[WAG^{+}14] Lian-Ping Wang, Orlando Ayala, Hui Gao, Charles Andersen, and Kevin L. Mathews. Study of forced turbulence and its modulation by finitesize solid particles using the lattice Boltzmann approach. Computers and Mathematics with Applications, 67(2):363–380, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001880.

Wang:2008:SMS

[Wan08]  Ji-Bo Wang.  Single-

machine scheduling with general learning functions. Computers and Mathematics with Applications, 56(8):1941–1947, October 2008. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812210800271X.

See note [YH11a].

Wang:2010:SIP

[Wan10a] Mingjin Wang. Some inequalities via probabilistic method. Computers and Mathematics with Applications, 59(11): 3481–3488, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002269.

Wang:2010:RAH

[Wan10b] Weiping Wang. Riordan arrays and harmonic number identities. Computers and Mathematics with Applications, 60 (5):1494–1509, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122110004451.

Wang:2011:IBV

[Wan11] Xuhuan Wang. Impulsive boundary value problem for nonlinear differential equations of fractional order. Computers and Mathematics with Applications, 62(5):2383–2391, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005803.

Wang:2012:BND

[Wan12] Cheng-Chi Wang. Bifurcation and nonlinear dynamic analysis of united gas-lubricated bearing system. Computers and Mathematics with Applications, 64(5):729–738, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211101011X.

Wang:2013:CNI

[Wan13a]  Jinyan Wang.  Corri-

gendum to “Notes on: ‘Interval-valued intuitionistic fuzzy soft sets and their properties’ [Comput. Math. Appl. 60 (2010) 906–918]” [Comput. Math. Appl. 64 (2012) 2954–2960]. Computers and Mathematics with Applications, 65(4):745–746, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006864.

See [JTC^{+}10a, WYG12].

Wang:2013:NDE

[Wan13b]  Kun Wang.  A new

discrete EVSS method for the viscoelastic flows. Computers and Mathematics with Applications, 65(4):609–615, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006487.

Wang:2014:SMT

[Wan14] Changjia Wang. On the solvability of models for two-phase flows of viscous incompressible fluid with shear-dependent viscosity. Computers and Mathematics with Applications, 68(3):132–139, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122114002053.

Wang:2015:CQS

[Wan15] Youjun Wang. A class of quasilinear Schrödinger equations with critical or supercritical exponents. Computers and Mathematics with Applications, 70(4):562–572, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002394.

Wang:2016:ISN

[Wan16a] Kun Wang. Iterative schemes for the nonhomogeneous Navier– Stokes equations based on the finite element approximation. Computers and Mathematics with Applications, 71(1):120–132, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005428.

Wang:2016:GBW

[Wan16b] Yilong Wang. Global bounded weak solutions to a degenerate quasilinear attraction-repulsion chemotaxis system with rotation. Computers and Mathematics with Applications, 72(9):2226–2240, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304783.

Wang:2018:BCS

[Wan18] Yilong Wang. Boundedness in a 2D chemotaxis-Stokes system with general sensitivity and nonlinear diffusion. Computers and Mathematics with Applications, 76(4):818–830, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302906.

Wang:2019:NSS

[Wan19a]  Huimin Wang.  Nu-

merical simulation for solitary wave of Klein– Gordon–Zakharov equation based on the lattice Boltzmann model. Computers and Mathematics with Applications, 78 (12):3941–3955, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122119303323.

Wang:2019:HCS

[Wan19b] Xiaohuan Wang. Hölder continuous of the solutions to stochastic nonlocal heat equations. Computers and Mathematics with Applications, 78(3):741–753, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301129.

Washington:2013:NRP

[Was13] Talitha M. Washington. NSFD representations for polynomial terms appearing in the potential functions of 1-dimensional conservative systems. Computers and Mathematics with Applications, 66 (11):2251–2258, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003921.

Wei:2015:DEU

[WAW15] Li Wei, Ravi P. Agarwal, and Patricia J. Y. Wong. Discussion on the existence and uniqueness of solution to nonlinear integro-differential systems. Computers and Mathematics with Applications, 69(5):374–389, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005938.

Wang:2011:SER

[WAZ11a] Guotao Wang, Bashir Ahmad, and Lihong Zhang. Some existence results for impulsive nonlinear fractional differential equations with mixed boundary conditions. Computers and Mathematics with Applications, 62(3):1389–1397, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002902.

Wazwaz:2011:RMF

[Waz11b] Abdul-Majid Wazwaz. The regularization method for Fredholm integral equations of the first kind. Computers and Mathematics with Applications, 61(10):2981– 2986, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002628.

Wilbrandt:2018:PMP

[WBA^{+}18] Ulrich Wilbrandt, Clemens Bartsch, Naveed Ahmed, Najib Alia, Felix Anker, Laura Blank, Alfonso Caiazzo, Sashikumaar Ganesan, Swetlana Giere, Gunar Matthies, Raviteja Meesala, Abdus Shamim, Jagannath Venkatesan, and Volker John. Par-MooN — a modernized program package based on mapped finite elements. Computers and Mathematics with Applications, 74(1):74–88, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306915.

Wittmann:2018:HPS

[WBN18] Roland Wittmann, Hans-Joachim Bungartz, and Philipp Neumann. High performance shallow water kernels for parallel overland flow simulations based on FullSWOF2D. Computers and Mathematics with Applications, 74(1):110–125, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300159.

Wang:2018:TGD

[WBZY18] Shixi Wang, Hai Bi, Yu Zhang, and Yidu Yang. A two-grid discretization scheme of non-conforming finite elements for transmission eigenvalues. Computers and Mathematics with Applications, 75(2):520–533, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305953.

Wang:2010:SNS

[WC10a] Changzhong Wang and Degang Chen. A short note on some properties of rough groups. Computers and Mathematics with Applications, 59(1):431–436, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109004003.

Wang:2010:CRI

[WC10b] Rong-Nian Wang and De-Han Chen. On a class of retarded integrodifferential equations with nonlocal initial conditions. Computers and Mathematics with Applications, 59(12):3700– 3709, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002634.

Wang:2011:NCM

[WC11a] Mingwu Wang and Guangyi Chen. A novel coupling model for risk analysis of swell and shrinkage of expansive soils. Computers and Mathematics with Applications, 62(7):2854–2861, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006225.

Wang:2011:TIR

[WC11b] Ying-Ming Wang and Kwai-Sang Chin. Technical importance ratings in fuzzy QFD by integrating fuzzy normalization and fuzzy weighted average. Computers and Mathematics with Applications, 62(11):4207–4221, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008698.

Wang:2015:NUS

[WC15] Qiuliang Wang and Jinru Chen. A new unfitted stabilized Nitsche’s finite element method for Stokes interface problems. Computers and Mathematics with Applications, 70(5):820–834, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002692.

Wang:2017:TGM

[WC17] Keyan Wang and Yanping Chen. Two-grid mixed finite element method for nonlinear hyperbolic equations. Computers and Mathematics with Applications, 74 (6):1489–1505, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211730370X.

Waeytens:2013:SIA

[WCB13] Julien Waeytens, Patrice Chatellier, and Frédéric Bourquin. Sensitivity of inverse advectiondiffusion-reaction to sensor and control: a low computational cost tool. Computers and Mathematics with Applications, 66(6):1082–1103, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004367.

Wei:2015:SBM

[WCCS15] Xing Wei, Wen Chen, Bin Chen, and Linlin Sun. Singular boundary method for heat conduction problems with certain spatially varying conductivity. Computers and Mathematics with Applications, 69(3):206–222, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005914.

Wang:2010:AMB

[WCD10] Dongqing Wang, Yanyun Chu, and Feng Ding. Auxiliary model-based RELS and MI–ELS algorithm for Hammerstein OEMA systems. Computers and Mathematics with Applications, 59(9):3092– 3098, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001392.

Wu:2013:HFI

[WCH13] Chih-Hua Wu, Bang-Fuh Chen, and Tin-Kan Hung. Hydrodynamic forces induced by transient sloshing in a 3D rectangular tank due to oblique horizontal excitation. Computers and Mathematics with Applications, 65(8):1163–1186, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300120X.

Wang:2018:SEF

[WCH18] Fajie Wang, Wen Chen, and Qingsong Hua. A simple empirical formula of origin intensity factor in singular boundary method for two-dimensional Hausdorff derivative Laplace equations with Dirichlet boundary. Computers and Mathematics with Applications, 76 (5):1075–1084, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303171.

Wu:2018:TLL

[WCLD18] Fengyan Wu, Xiujun Cheng, Dongfang Li, and Jinqiao Duan. A twolevel linearized compact ADI scheme for twodimensional nonlinear reaction–diffusion equations. Computers and Mathematics with Applications, 75(8):2835–2850, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300233.

Wen:2019:LMN

[WCQ^{+}19] Binghai Wen, Hui Chen, Zhangrong Qin, Bing He, and Chaoying Zhang. Lateral migration and nonuniform rotation of suspended ellipse in Poiseuille flow. Computers and Mathematics with Applications, 78(4):1142–1153, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305132.

Wang:2018:CEF

[WCSW18] Hua Wang, Jinru Chen, Pengtao Sun, and Nan Wang. A conforming enriched finite element method for Stokes interface problems. Computers and Mathematics with Applications, 75(12):4256–4271, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118301639.

Wu:2013:USI

[WCW13] Bin Wu, Qun Chen, and Zewen Wang. Uniqueness and stability of an inverse problem for a phase field model using data from one component. Computers and Mathematics with Applications, 66 (10):2126–2138, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005464.

Wang:2013:CEP

[WCZ13] Y. H. Wang, K. Cao, and X. M. Zhang. Complex event processing over distributed probabilistic event streams. Computers and Mathematics with Applications, 66 (10):1808–1821, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004677.

Wu:2019:GAF

[WCZ^{+}19] Jiayang Wu, Yongguang Cheng, Wei Zhou, Chunze Zhang, and Wei Diao. GPU acceleration of FSI simulations by the immersed boundary-lattice Boltzmann coupling scheme. Computers and Mathematics with Applications, 78(4):1194–1205, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305636.

Wang:2010:AOA

[WD10] Zhong Bao Wang and Xie Ping Ding. (H(·, ·), η)- accretive operators with an application for solving set-valued variational inclusions in Banach spaces. Computers and Mathematics with Applications, 59(4):1559–1567, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007330.

Wibowo:2012:FRB

[WD12] Santoso Wibowo and Hepu Deng. A fuzzy rule-based approach for screening international distribution centres. Computers and Mathematics with Applications, 64 (5):1084–1092, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002313.

Winter:2013:HGC

[WD13] Johannes Winter and Kurt Dietrich. A hijacker’s guide to communication interfaces of the trusted platform module. Computers and Mathematics with Applications, 65(5):748–761, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004634.

Wang:2016:LWP

[WD16] Changjia Wang and Qun Dai. Local well-posedness for Boussinesq approximation with shear dependent viscosities in 3D. Computers and Mathematics with Applications, 72(1):131–146, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302437.

Wei:2010:CVF

[Wei10a] Jiang Wei. The constant variation formulae for singular fractional differential systems with delay. Computers and Mathematics with Applications, 59(3):1184–1190, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004295.

Wei:2010:FTM

[Wei10b] Long Wei. A function transformation method and exact solutions to a generalized sinh-Gordon equation. Computers and Mathematics with Applications, 60 (11):3003–3011, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110007595.

Wei:2012:CFC

[Wei12a] Jiang Wei. The controllability of fractional control systems with control delay. Computers and Mathematics with Applications, 64(10):3153–3159, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001903.

Wei:2012:DFD

[Wei12b]  Zhouchao Wei.  De-

layed feedback on the 3-D chaotic system only with two stable nodefoci. Computers and Mathematics with Applications, 63(3):728–738, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010200.

Weisser:2014:AOT

[Wei14] Steffen Weißer. Arbitrary order Trefftz-like basis functions on polygonal meshes and realization in BEM-based FEM. Computers and Mathematics with Applications, 67(7):1390–1406, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000431.

Weisser:2017:RBE

[Wei17] Steffen Weißer. Residual based error estimate and quasi-interpolation on polygonal meshes for high order BEM-based FEM. Computers and Mathematics with Applications, 73(2):187–202, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306307.

Wen:2018:IRC

[Wen18a] Zhihong Wen. An improved regularity criterion for the 3D Hall– MHD equations via the vorticity. Computers and Mathematics with Applications, 75 (3):821–836, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306569.

Wen:2018:NRC

[Wen18b] Zhihong Wen. A new regularity criterion of the 2D MHD equations. Computers and Mathematics with Applications, 76 (6):1333–1337, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830350X.

Wang:2017:NAB

[WF17] Chuanjian Wang and Hui Fang. Non-auto Bäcklund transformation, nonlocal symmetry and CRE solvability for the Bogoyavlenskii– Kadomtsev–Petviashvili equation. Computers and Mathematics with Applications, 74 (12):3296–3302, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305047.

Wang:2018:BBT

[WF18] Chuanjian Wang and Hui Fang. Bilinear Bäcklund transformations, kink periodic solitary wave and lump wave solutions of the Bogoyavlenskii– Kadomtsev–Petviashvili equation. Computers and Mathematics with Applications, 76(1):1–10, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301913.

Wu:2016:DVE

[WFC16] Chih-Hua Wu, Odd Magnus Faltinsen, and Bang-Fuh Chen. Dynamics of vortex evolution in a 2D baffled tank. Computers and Mathematics with Applications, 71(1):1–28, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004903.

Wu:2010:CFS

[WFDW10a] Ai-Guo Wu, Gang Feng, Guang-Ren Duan, and Wei-Jun Wu. Closed-form solutions to Sylvesterconjugate matrix equations. Computers and Mathematics with Applications, 60(1):95–111, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003068.

Wu:2010:ISC

[WFDW10b] Ai-Guo Wu, Gang Feng, Guang-Ren Duan, and Wei-Jun Wu. Iterative solutions to coupled Sylvester-conjugate matrix equations. Computers and Mathematics with Applications, 60(1):54– 66, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003007.

Wei:2011:GBP

[WFL11a] Feifei Wei, Jieqing Feng, and Hongwei Lin. GPUbased parallel solver via the Kantorovich theorem for the nonlinear Bernstein polynomial systems. Computers and Mathematics with Applications, 62(6):2506–2517, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005943.

Wu:2011:EPC

[WFL11b] Zhao-Chun Wu, Jing-Mei Feng, and Jian-Bing Liu. Euler’s predictor–corrector technique for solving the depth-integrated shallow water equations.

Computers and Mathematics with Applications, 61(8):2287–2291, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007479.

Wallace:2017:RDM

[WFY17] M. Wallace, R. Feres, and G. Yablonsky. Reactiondiffusion on metric graphs: From 3D to 1D. Computers and Mathematics with Applications, 73(9):2035–2052, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301128.

Weng:2012:ITK

[WFZ12] Zhifeng Weng, Xinlong Feng, and Shuying Zhai. Investigations on two kinds of two-grid mixed finite element methods for the elliptic eigenvalue problem. Computers and Mathematics with Applications, 64(8):2635–2646, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122112004932.

Wang:2011:QIV

[WG11] Peiguang Wang and Wei Gao. Quasilinearization of an initial value problem for a set valued integro-differential equation. Computers and Mathematics with Applications, 61(8):2111–2115, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006693.

Wu:2018:FOA

[WGY^{+}18] Yongyong Wu, Nan Gui, Xingtuan Yang, Jiyuan Tu, and Shengyao Jiang. Fourth-order analysis of force terms in multiphase pseudopotential lattice Boltzmann model. Computers and Mathematics with Applications, 76(7):1699–1712, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303870.

Wang:2010:QCK

[WH10] Baocang Wang and Yupu Hu. Quadratic compact knapsack public-key cryptosystem. Computers and Mathematics with Applications, 59(1):194–206, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005756.

Wang:2011:LBA

[WH11a] Chen-Hao Wang and Jeng-Rong Ho. A lattice Boltzmann approach for the non-Newtonian effect in the blood flow. Computers and Mathematics with Applications, 62(1):75– 86, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003543.

Wu:2011:DEM

[WH11b] Xionghua Wu and Guofeng Han. Direct expansion method of boundary condition for solving 3D elliptic equations with small parameters in the irregular domain. Computers and Mathematics with Applications, 61(10):2971– 2980, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002446.

Wen:2014:CAN

[WH14] Juan Wen and Yinnian He. Convergence analysis of a new multiscale finite element method for the stationary Navier– Stokes problem. Computers and Mathematics with Applications, 67(1):1–25, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006238.

Wang:2016:SSA

[WH16] Pengde Wang and Chengming Huang. Splitstep alternating direction implicit difference scheme for the fractional Schrödinger equation in two dimensions. Computers and Mathematics with Applications, 71(5):1114–1128, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300189.

Wang:2018:ESS

[WH18] Nan Wang and Chengming Huang. An efficient split-step quasicompact finite difference method for the nonlinear fractional Ginzburg– Landau equations. Computers and Mathematics with Applications, 75(7):2223–2242, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730768X.

Wu:2012:SPP

[WHC12] Bing-Fei Wu, Hao-Yu Huang, and Yen-Lin Chen. The singlepass perceptual embedded zero-tree coding implementation on DSP. Computers and Mathematics with Applications, 64(5):1140–1152, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002374.

Wang:2014:GMH

[WHD14] Hui Wang, Ting-Zhu Huang, and Ying-Qiong Du. A global minimization hybrid active contour model with applications to oil spill images. Computers and Mathematics with Applications, 68(3):353–362, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114002648.

Wen:2011:CTC

[WHG11] Jiajin Wen, Tianyong Han, and Chaobang Gao. Convergence tests on constant Dirichlet series. Computers and Mathematics with Applications, 62(9):3472–3489, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007486.

Wang:2011:DDE

[WhJxLwW11] Jun Wang, Jiu hong Jiang, Li xin Lu, and Zhi wei Wang. Dropping damage evaluation for a tangent nonlinear system with a critical component. Computers and Mathematics with Applications, 61(8):1979–1982, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006176.

Wei:2011:RIT

[WHLC11] Hsiu-Chuan Wei, Shin-Feng Hwang, Jenn-Tsann Lin, and Tze-Jang Chen. The role of initial tumor biomass size in a mathematical model of periodically pulsed chemotherapy. Computers and Mathematics with Applications, 61(10):3117– 3127, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002823.

Wen:2018:CAM

[WHQ^{+}18] Binghai Wen, Bingfang Huang, Zhangrong Qin, Chunlei Wang, and Chaoying Zhang. Contact angle measurement in lattice Boltzmann method. Computers and Mathematics with Applications, 76(7):1686–1698, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303869.

Wang:2011:AND

[WHS11] Zhen Wang, Xia Huang, and Guodong Shi. Analysis of nonlinear dynamics and chaos in a fractional order financial system with time delay. Computers and Mathematics with Applications, 62(3):1531–1539, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003609.

Witula:2012:SVS

[WHS12] Roman Witula, Edyta Hetmaniok, and Damian Slota. A stronger version of the second mean value theorem for integrals. Computers and Mathematics with Applications, 64(6):1612–1615, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000181.

Wang:2017:INT

[WHS17] Liqun Wang, Songming Hou, and Liwei Shi. An improved non-traditional finite element formulation for solving threedimensional elliptic interface problems. Computers and Mathematics with Applications, 73(3):374–384, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306678.

Wang:2018:SWF

[WHS18] Liqun Wang, Songming Hou, and Liwei Shi. A simple weak formulation for solving two-dimensional diffusion equation with local reaction on the interface. Computers and Mathematics with Applications, 75(4):1378–1389, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307034.

Wang:2016:SAF

[WHTZ16] Jianyun Wang, Yunqing Huang, Zhikun Tian, and Jie Zhou. Superconvergence analysis of finite element method for the time-dependent Schrödinger equation. Computers and Mathematics with Applications, 71(10):1960–1972, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301316.

Wen:2011:TSS

[WHW11] Chun Wen, Ting-Zhu Huang, and Chao Wang. Triangular and skewsymmetric splitting method for numerical solutions of Markov chains. Computers and Mathematics with Applications, 62 (11):4039–4048, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008091.

Wang:2017:BCM

[WHZL17] Liangchen Wang, Xuegang Hu, Pan Zheng, and Ling Li. Boundedness in a chemotaxis model with exponentially decaying diffusivity and consumption of chemoattractant. Computers and Mathematics with Applications, 74 (10):2444–2448, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304406.

Wang:2011:MOR

[WJ11] Xiao-Long Wang and Yao-Lin Jiang. Model order reduction methods for coupled systems in the time domain using Laguerre polynomials. Computers and Mathematics with Applications, 62(8):3241–3250, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007127.

Wang:2012:GEW

[WJWW12] Yu Wang, Lili Ju, Desheng Wang, and Xiaoqiang Wang. Generalized edge-weighted centroidal Voronoi tessellations for geometry processing. Computers and Mathematics with Applications, 64(8):2663–2681, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004956.

Wong:2013:PPS

[WK13] Kok-Seng Wong and Myung Ho Kim. Privacypreserving similarity coefficients for binary data. Computers and Mathematics with Applications, 65(9):1280–1290, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001423.

Wassermann:2018:GDN

[WKBR18] Benjamin Wassermann, Stefan Kollmannsberger, Tino Bog, and Ernst Rank. From geometric design to numerical analysis: A direct approach using the Finite Cell Method on Constructive Solid Geometry. Computers and Mathematics with Applications, 74(7):1703–1726, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300780.

Wang:2010:NMS

[WKG10] Xiuhua Wang, Jisheng Kou, and Chuanqing Gu. A new modified secant-like method for solving nonlinear equations. Computers and Mathematics with Applications, 60(6):1633–1638, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004591.

Wozniak:2014:CCE

[WKP^{+}14] M. Woźniak, K. Kuźnik, M. Paszyński, V. M. Calo, and D. Pardo. Computational cost estimates for parallel shared memory isogeometric multi-frontal solvers. Computers and Mathematics with Applications, 67(10):1864– 1883, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001503.

Wang:2013:CIM

[WKS13] Xiuhua Wang, Jisheng Kou, and Dongyang Shi. Convergence of an iterative method for solving a class of nonlinear equations. Computers and Mathematics with Applications, 66(7):1322–1328, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004823.

Wu:2010:ESW

[WL10] Xiangjun Wu and Hongtao Lu. Exponential synchronization of weighted general delay coupled and non-delay coupled dynamical networks. Computers and Mathematics with Applications, 60(8):2476–2487, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006280.

Wang:2011:ABA

[WL11a] Chia-Hung Wang and Hsing Paul Luh. Analysis of bandwidth allocation on end-to-end QoS networks under budget control. Computers and Mathematics with Applications, 62(1):419– 439, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004226.

Wang:2011:MEA

[WL11b] Fu-Kwun Wang and Chih-Wen Lee. Mestimator with asymmetric influence function for estimating the Burr type III parameters with outliers. Computers and Mathematics with Applications, 62(4):1896–1907, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S089812211100513X.

Wang:2011:SFD

[WL11c] Jin-Liang Wang and Hui-Feng Li. Surpassing the fractional derivative: Concept of the memorydependent derivative. Computers and Mathematics with Applications, 62(3):1562–1567, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003294.

Wang:2011:OCE

[WL11d] Yuming Wang and Jianzhong Lin. The oblique collision efficiency of nanoparticles at different angles in Brownian coagulation. Computers and Mathematics with Applications, 61(8):1917–1922, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005018.

Wang:2012:ELP

[WL12a] Chia-Hung Wang and Hsing Paul Luh. Estimating the loss probability under heavy traffic conditions. Computers and Mathematics with Applications, 64(5):1352–1363, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002866.

Wang:2012:DRB

[WL12b] Liyan Wang and Jijun Liu. Data regularization for a backward time-fractional diffusion problem. Computers and Mathematics with Applications, 64 (11):3613–3626, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006025.

Wang:2013:MAD

[WL13a] Weize Wang and Xinwang Liu. The multi-attribute decision making method based on interval-valued intuitionistic fuzzy Einstein hybrid weighted geometric operator. Computers and Mathematics with Applications, 66 (10):1845–1856, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122113004641.

Wardle:2013:FEL

[WL13b] Kent E. Wardle and Taehun Lee. Finite element lattice Boltzmann simulations of free surface flow in a concentric cylinder. Computers and Mathematics with Applications, 65(2):230–238, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007887.

Wang:2015:SPP

[WL15] Yu-Xia Wang and Wan-Tong Li. Stationary problem of a predator– prey system with nonlinear diffusion effects. Computers and Mathematics with Applications, 70(8):2102–2124, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004125.

Whang:2016:FLS

[WL16] Sungim Whang and Sunmi Lee. Fully localized solitary waves for the forced Kadomtsev–Petviashvili equation. Computers and Mathematics with Applications, 72(7):1865–1879, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630462X.

Wang:2017:EUS

[WL17a] Zhaojuan Wang and Yanan Liu. Existence and upper semicontinuity of random attractors for non-autonomous stochastic strongly damped sine-Gordon equation on unbounded domains. Computers and Mathematics with Applications, 73(7):1445–1460, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300457.

Wu:2017:FVA

[WL17b] Chih-Ping Wu and Wei-Chen Li. Free vibration analysis of embedded single-layered nanoplates and graphene sheets by using the multiple time scale method. Computers and Mathematics with Applications, 73(5):838–854, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300445.

Wang:2018:ICT

[WLA18] Duo Wang, Christopher R. Leonardi, and Saiied M. Aminossadati. Improved coupling of time integration and hydrodynamic interaction in particle suspensions using the lattice Boltzmann and discrete element methods. Computers and Mathematics with Applications, 75(7):2593–2606, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300063.

Wong:2011:EDA

[WLC11] Kwok-Wo Wong, Qiuzhen Lin, and Jianyong Chen. Error detection in arithmetic coding with artificial markers. Computers and Mathematics with Applications, 62(1):359– 366, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004056.

Wang:2013:HSH

[WLD13] Xiang Wang, Yan Li, and Lin Dai. On Hermitian and skew-Hermitian splitting iteration methods for the linear matrix equation AXB = C. Computers and Mathematics with Applications, 65(4):657–664, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006566.

Wu:2011:CPS

[WLDL11a] Ai-Guo Wu, Lingling Lv, Guang-Ren Duan, and Wanquan Liu. Corrigendum to “Parametric solutions to Sylvesterconjugate matrix equations” [Comput. Math. Appl. 62 (2011) 3317– 3325]. Computers and Mathematics with Applications, 62(12):4806, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009345.

See [WLDL11b].

Wu:2011:PSSb

[WLDL11b] Ai-Guo Wu, Lingling Lv, Guang-Ren Duan, and Wanquan Liu. Parametric solutions to Sylvesterconjugate matrix equations. Computers and Mathematics with Applications, 62(9):3317–3325, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006973.

See corrigendum [WLDL11a].

Wen:2010:RAC

[WLGL10] Ching-Feng Wen, Yung-Yih Lur, Sy-Ming Guu, and E. Stanley Lee. On a recurrence algorithm for continuous-time linear fractional programming problems. Computers and Mathematics with Applications, 59(2):829–852, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007093.

Wang:2014:SAV

[WLHZ14] Pan Wang, Zhenzhou Lu, Jixiang Hu, and Changcong Zhou. Sensitivity analysis of the variance contributions with respect to the distribution parameters by the kernel function. Computers and Mathematics with Applications, 67(10):1756– 1771, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114001643.

Wei:2012:NSR

[WLL12] Wei Wei, Xuezhu Li, and Xia Li. New stability results for fractional integral equation. Computers and Mathematics with Applications, 64 (10):3468–3476, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001824.

Wang:2018:SOA

[WLL^{+}18] Jinfeng Wang, Tianqi Liu, Hong Li, Yang Liu, and Siriguleng He. Second-order approximation scheme combined with H^{1}-Galerkin MFE method for nonlinear time fractional convection– diffusion equation. Computers and Mathematics with Applications, 73(6):1182–1196, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122116304424.

Wang:2013:PDS

[WLM13] Xiang Wang, Wen-Wei Li, and Liang-Zhi Mao. On positive-definite and skew-Hermitian splitting iteration methods for continuous Sylvester equation AX + XB = C. Computers and Mathematics with Applications, 66(11):2352–2361, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005622.

Wu:2010:ARP

[WLS10] Yu-Dong Wu, V. Lokesha, and H. M. Srivastava. Another refinement of the Pólya–Szegő inequality. Computers and Mathematics with Applications, 60(3):761–770, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003809.

Wang:2013:DBS

[WLT13a] Pan Wang, Zhenzhou Lu, and Zhangchun Tang. A derivative based sensitivity measure of failure probability in the presence of epistemic and aleatory uncertainties. Computers and Mathematics with Applications, 65(1):89–101, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006438.

Wang:2013:IMA

[WLT13b] Pan Wang, Zhenzhou Lu, and Zhangchun Tang. Importance measure analysis with epistemic uncertainty and its moving least squares solution. Computers and Mathematics with Applications, 66(4):460–471, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003489.

Wang:2011:AHE

[WLW^{+}11] Yu-Zhi Wang, Bo Li, Ren-Qing Wang, Jing Su, and Xiao-Xia Rong. Application of the Hurst exponent in ecology. Computers and Mathematics with Applications, 61(8):2129–2131, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006802.

Wang:2016:CSA

[WLW16] Chuan-Long Wang, Chao Li, and Jin Wang. Comparisons of several algorithms for Toeplitz matrix recovery. Computers and Mathematics with Applications, 71(1):133–146, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005416.

Wang:2017:TMA

[WLW17] Chuan-Long Wang, Chao Li, and Jin Wang. Two modified augmented Lagrange multiplier algorithms for Toeplitz matrix compressive recovery. Computers and Mathematics with Applications, 74(8):1915–1921, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304145.

Wang:2018:NRM

[WLXG18] Linjun Wang, Jinwei Liu, Youxiang Xie, and Yuantong Gu. A new regularization method for the dynamic load identification of stochastic structures. Computers and Mathematics with Applications, 76(4):741–759, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830275X.

Wei:2018:LSA

[WLXZ18] Guang-Mei Wei, Ying-Lin Lu, Ya-Qin Xie, and Wen-Xin Zheng. Lie symmetry analysis and conservation law of variable-coefficient Davey–Stewartson equation. Computers and Mathematics with Applications, 75(9):3420–3430, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300816.

Wu:2013:LLI

[WLYX13] Guangchao Wu, Yuhan Li, Xiaowei Yang, and Jianqing Xi. Local learning integrating global structure for large scale semi-supervised classification. Computers and Mathematics with Applications, 66(10):1961– 1970, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113004306.

Wang:2018:NAH

[WLZ^{+}18a] F. L. Wang, F. Liu, Y. M.

Zhao, Y. H. Shi, and Z. G. Shi. A novel approach of high accuracy analysis of anisotropic bilinear finite element for time-fractional diffusion equations with variable coefficient. Computers and Mathematics with Applications, 75(10):3786–3800, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301111.

Wei:2018:SSO

[WLZ18b] Zhihan Wei, Chuan Li, and Shan Zhao. A spatially second order alternating direction implicit (ADI) method for solving three dimensional parabolic interface problems. Computers and Mathematics with Applications, 75(6):2173–2192, March 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303851.

Wang:2013:RSB

[WmN13] Xiang Wang and Yong ming Nie. A refined shifted block inverse-free Krylov subspace method for symmetric generalized eigenvalue problems. Computers and Mathematics with Applications, 66(6):1137–1146, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004719.

Wang:2019:LBS

[WMP^{+}19] Lian-Ping Wang, Haoda Min, Cheng Peng, Nicholas Geneva, and Zhaoli Guo. A lattice-Boltzmann scheme of the Navier–Stokes equation on a threedimensional cuboid lattice. Computers and Mathematics with Applications, 78(4):1053–1075, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303492.

Wu:2011:PSSa

[WMSH11] Wei Wu, Yi Mu, Willy Susilo, and Xinyi Huang. Provably secure serveraided verification signatures. Computers and Mathematics with Applications, 61(7):1705–1723, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000629.

Wen:2013:QCA

[WMW13] Rui-Ping Wen, Guo-Yan Meng, and Chuan-Long Wang. Quasi-Chebyshev accelerated iteration methods based on optimization for linear systems. Computers and Mathematics with Applications, 66(6):934–942, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300391X.

Wang:2011:RSS

[WMZW11] Shangping Wang, Rui Ma, Yaling Zhang, and Xiaofeng Wang. Ring signature scheme based on multivariate public key cryptosystems. Computers and Mathematics with Applications, 62 (10):3973–3979, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008303.

Wang:2018:VVC

[WN18] Huiling Wang and Yufeng Nie. Variable V -cycle multigrid preconditioners for the discrete systems from combined hybrid quadrilateral elements. Computers and Mathematics with Applications, 76(3):649–660, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302657.

Wongkia:2012:MAM

[WNC12] Wararat Wongkia, Kanlaya Naruedomkul, and Nick Cercone. i-Math: Automatic math reader for Thai blind and visually impaired students. Computers and Mathematics with Applications, 64(6):2128–2140, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122112003495.

Wang:2019:SSL

[WNTW19] Rui-Rui Wang, Qiang Niu, Xiao-Bin Tang, and Xiang Wang. Solving shifted linear systems with restarted GM-RES augmented with error approximations. Computers and Mathematics with Applications, 78 (6):1910–1918, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301622.

Wu:2010:FMG

[WO10] Desheng Dash Wu and David L. Olson. Fuzzy multiattribute grey related analysis using DEA. Computers and Mathematics with Applications, 60(1):166–174, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110003251.

Wazwaz:2018:ACM

[WO18] Abdul-Majid Wazwaz and M. S. Osman. Analyzing the combined multi-waves polynomial solutions in a two-layer-liquid medium. Computers and Mathematics with Applications, 76(2):276– 283, July 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118302165.

Wu:2011:FAS

[WPH11] Qing-E Wu, Xue-Min Pang, and Zhen-Yu Han. Fuzzy automata system with application to target recognition based on image processing. Computers and Mathematics with Applications, 61(5):1267–1277, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006905.

Wang:2016:BTD

[WPL16] Yulan Wang, Fengqin Pang, and Huifang Li. Boundedness in a threedimensional chemotaxis-Stokes system with tensorvalued sensitivity. Computers and Mathematics with Applications, 71(3):712–722, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005945.

Wu:2012:HHS

[WQNF12] Bin Wu, Cunhua Qian, Weihong Ni, and Shuhai Fan. Hybrid harmony search and artificial bee colony algorithm for global optimization problems. Computers and Mathematics with Applications, 64(8):2621–2634, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112004907.

Wang:2014:ROM

[WQRZ14] Zewen Wang, Shufang Qiu, Zhousheng Ruan, and Wen Zhang. A regularized optimization method for identifying the space-dependent source and the initial value simultaneously in a parabolic equation. Computers and Mathematics with Applications, 67(7):1345–1357, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000789.

Wright:2013:PSP

[Wri13]  Justin Wright.  Peri-

odic systems of population models and enveloping functions. Computers and Mathematics with Applications, 66 (11):2178–2195, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005075.

Wang:2013:ASF

[WRW13] Yuanquan Wang, Wenqi Ren, and Huaibin Wang. Anisotropic second and fourth order diffusion models based on convolutional virtual electric field for image denoising. Computers and Mathematics with Applications, 66(10):1729–1742, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005415.

Weirather:2019:SPH

[WRW^{+}19] Johannes Weirather, Vladyslav Rozov, Mario Wille, Paul Schuler, Christian Seidel, Nikolaus A. Adams, and Michael F. Zaeh. A smoothed particle hydrodynamics model for laser beam melting of Ni-based alloy 718. Computers and Mathematics with Applications, 78(7):2377–2394, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306175.

Wang:2018:UNS

[WRY18] Jun-Gang Wang, Yu-Hong Ran, and Zhan-Bin Yuan. Uniqueness and numerical scheme for the Robin coefficient identification of the time-fractional diffusion equation. Computers and Mathematics with Applications, 75(11):4107–4114, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301524.

Wang:2010:SMP

[WS10a] Haichao Wang and Erfang Shan. Some matching properties in 4-γ_{×2}- critical graphs. Computers and Mathematics with Applications, 59(2):694– 699, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007238.

Wei:2010:BAB

[WS10b] Gengping Wei and Jianhua Shen. Boundedness and asymptotic behavior results for nonlinear difference equations with positive and negative coefficients. Computers and Mathematics with Applications, 60(8):2469–2475, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006206.

Wang:2011:EPS

[WS11a] Weibing Wang and Jianhua Shen. Eigenvalue problems of second order impulsive differential equations. Computers and Mathematics with Applications, 62(1):142– 150, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003646.

Witula:2011:SSA

[WS11b] Roman Witula and Damian Slota. On the sum of some alternating series. Computers and Mathematics with Applications, 62(6):2658–2664, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006717.

Wang:2012:RCS

[WS12] Cheng Wang and Yi Shen.

Robust H_{∞} control for stochastic systems with nonlinearity, uncertainty and time-varying delay. Computers and Mathematics with Applications, 63(5):985–998, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010467.

Wu:2016:ETM

[WS16] Yuanqing Wu and Shuyu Sun. Equivalence of two models in single-phase multicomponent flow simulations. Computers and Mathematics with Applications, 71(6):1303–1316, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116300505.

Wang:2017:EMP

[WS17] Cong Wang and Yan-Ying Shang. Existence and multiplicity of positive solutions for elliptic equation with critical weighted Hardy– Sobolev exponents and boundary singularities. Computers and Mathematics with Applications, 74(4):701–713, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303036.

Wang:2016:LBS

[WSC16] Lei Wang, Baochang Shi, and Zhenhua Chai. A lattice Boltzmann study of the asymmetry effect on the hemodynamics in stented fusiform aneurysms. Computers and Mathematics with Applications, 71(1):328– 348, January 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005647.

Wang:2011:URK

[WSCL11] Yulan Wang, Lijuan Su, Xuejun Cao, and Xiaona Li. Using reproducing kernel for solving a class of singularly perturbed problems. Computers and Mathematics with Applications, 61(2):421–430, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008680.

Wu:2016:DUG

[WSCW16] Chen Wu, Baochang Shi, Zhenhua Chai, and Peng Wang. Discrete unified gas kinetic scheme with a force term for incompressible fluid flows. Computers and Mathematics with Applications, 71(12):2608–2629, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302061.

Wang:2010:MPS

[WSG10] Da-Bin Wang, Jian-Ping Sun, and Wen Guan. Multiple positive solutions for functional dynamic equations on time scales. Computers and Mathematics with Applications, 59(4):1433–1440, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007780.

Witula:2012:SPI

[WSH12] Roman Witula, Damian Slota, and Edyta Hetmaniok. Some properties of inverses of the full matrices. Computers and Mathematics with Applications, 63(5):905–911, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211101039X.

Wang:2010:ADT

[WSL10] De-Gang Wang, Wen-Yan Song, and Hong-Xing Li. Analysis and design of time-variant fuzzy systems based on dynamic fuzzy inference. Computers and Mathematics with Applications, 60(3):464–489, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003275.

Wu:2019:ULD

[WSM^{+}19] T. Wu, M. Shashkov, N. Morgan, D. Kuzmin, and H. Luo. An updated Lagrangian discontinuous Galerkin hydrodynamic method for gas dynamics. Computers and Mathematics with Applications, 78(2):258– 273, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301767.

Wang:2010:SJE

[WSS10] Ji-Bo Wang, Lin-Hui Sun, and Lin-Yan Sun. Scheduling jobs with an exponential sum-of-actual-processing-time-based learning effect. Computers and Mathematics with Applications, 60(9):2673–2678, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006917.

Wang:2011:FEP

[WSW11] Kun Wang, Yueqiang Shang, and Hongbo Wei. A finite element penalty method for the linearized viscoelastic Oldroyd fluid motion equations. Computers and Mathematics with Applications, 62(4):1814–1827, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005049.

Wu:2012:EST

[WTC^{+}12] Hui-Chi Wu, Chien-Ming Tseng, Po-Chou Chan, Sue-Fen Huang, Wei-Wei Chu, and Yung-Fu Chen. Evaluation of stock trading performance of students using a Web-based virtual stock trading system. Computers and Mathematics with Applications, 64(5):1495–1505, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211200315X.

Wu:2018:RWV

[WTLS18] Xiao-Yu Wu, Bo Tian, Lei Liu, and Yan Sun. Rogue waves for a variablecoefficient Kadomtsev– Petviashvili equation in fluid mechanics. Computers and Mathematics with Applications, 76(2):215– 223, July 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0898122117307848.

Wang:2017:MSO

[WTM17] Xiang Wang, Xiao-Bin Tang, and Liang-Zhi Mao. A modified secondorder Arnoldi method for solving the quadratic eigenvalue problems. Computers and Mathematics with Applications, 73(2):327–338, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306599.

Weickert:2010:IWT

[WTSS10] M. Weickert, G. Teike, O. Schmidt, and M. Sommerfeld. Investigation of the LES WALE turbulence model within the lattice Boltzmann framework. Computers and Mathematics with Applications, 59(7):2200–2214, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900652X.

Wang:2017:SWB

[WTYZ17] Xiu-Bin Wang, Shou-Fu Tian, Hui Yan, and Tian Tian Zhang. On the solitary waves, breather waves and rogue waves to a generalized (3 + 1)- dimensional Kadomtsev– Petviashvili equation. Computers and Mathematics with Applications, 74(3):556–563, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730278X.

Wu:2010:GTD

[Wu10] Hongbo Wu. The generalized truth degree of quantitative logic in the logic system P_{n}^{∗} (nvalued NM-logic system). Computers and Mathematics with Applications, 59(8):2587–2596, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000465.

Wu:2011:SSI

[Wu11a] Hongwei Wu. Strong solutions to the incompressible magnetohydrodynamic equations with vacuum. Computers and Mathematics with Applications, 61(9):2742–2753, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001854.

Wu:2011:NAS

[Wu11b] Jinming Wu. A new approach for shape preserving interpolating curves. Computers and Mathematics with Applications, 61(5):1425–1430, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000186.

Wu:2016:BSL

[Wu16] Xiulan Wu. The blowup of solutions for m-Laplacian equations with variable sources under positive initial energy. Computers and Mathematics with Applications, 72(9):2516–2524, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305272.

Wu:2018:EGS

[Wu18a] Ke Wu. Existence of ground states for a Kirchhoff type problem without 4-superlinear condition.

Computers and Mathematics with Applications, 75(3):755–763, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306533.

Wu:2018:DLE

[Wu18b] Yiting Wu. On-diagonal lower estimate of heat kernels for locally finite graphs and its application to the semilinear heat equations. Computers and Mathematics with Applications, 76(4):810–817, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830289X.

Wang:2014:HOE

[WV14] Zhibo Wang and Seakweng Vong. A high-order exponential ADI scheme for two dimensional time fractional convection– diffusion equations. Computers and Mathematics with Applications, 68(3): 185–196, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.

com/science/article/
pii/S0898122114002090.

Witherden:2015:ISQ

[WV15] F. D. Witherden and P. E. Vincent. On the identification of symmetric quadrature rules for finite element methods. Computers and Mathematics with Applications, 69(10):1232– 1241, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001224.

Wang:2016:CDS

[WV16] Zhibo Wang and Seakweng Vong. A compact difference scheme for a two dimensional nonlinear fractional Klein– Gordon equation in polar coordinates. Computers and Mathematics with Applications, 71(12): 2524–2540, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116301730.

Westbeek:2019:IPS

[WvDRG19] S. Westbeek, J. A. W.

van Dommelen, J. J. C. Remmers, and M. G. D. Geers. Influence of particle shape in the additive manufacturing process for ceramics. Computers and Mathematics with Applications, 78(7):2360–2376, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304590.

Wang:2010:SMM

[WW10a] Ji-Bo Wang and Ming-Zheng Wang. Single machine multiple common due dates scheduling with learning effects. Computers and Mathematics with Applications, 60(11):2998–3002, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007583.

Wang:2010:BNL

[WW10b] Renhong Wang and Shaofan Wang. The Bezout number for linear piecewise algebraic curves. Computers and Mathematics with Applications, 59(2):1019–1030, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122109006713.

Widyadana:2010:RLS

[WW10c] Gede Agus Widyadana and Hui Ming Wee. Revisiting lot sizing for an inventory system with product recovery. Computers and Mathematics with Applications, 59(8):2933–2939, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001203.

Wang:2011:BPS

[WW11a] Peiguang Wang and Meng Wu. ϕ_{0}-boundedness and practical ϕ_{0}-stability of difference equations. Computers and Mathematics with Applications, 62(8):2863–2870, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005633.

Wang:2011:PTI

[WW11b] Qing-Ping Wang and Guo-Jun Wang. The probability theories for IVFSs and IVIFSs. Computers and Mathematics with Applications, 62 (12):4535–4538, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008960.

Wang:2011:HOCb

[WW11c] Yuan-Ming Wang and Jie Wang. A higher-order compact ADI method with monotone iterative procedure for systems of reaction–diffusion equations. Computers and Mathematics with Applications, 62(6):2434–2451, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005840.

Wang:2014:ENS

[WW14a] Chunmei Wang and Junping Wang. An efficient numerical scheme for the biharmonic equation by weak Galerkin finite element methods on polygonal or polyhedral meshes. Computers and Mathematics with Applications, 68(12):2314–2330, December 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122114002326.

Wu:2014:EUI

[WW14b] Bin Wu and Siyuan Wu.

Existence and uniqueness of an inverse source problem for a fractional integrodifferential equation. Computers and Mathematics with Applications, 68(10):1123–1136, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004064.

Wu:2015:IMS

[WW15] Ke Wu and Xian Wu. Infinitely many small energy solutions for a modified Kirchhoff-type equation in R^{N}. Computers and Mathematics with Applications, 70(4):592– 602, August 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115002370.

Wang:2016:CLM

[WW16] Tao Wang and Yuan-Ming Wang. A compact LOD method and its extrapolation for two-dimensional modified anomalous fractional sub-diffusion equations. Computers and Mathematics with Applications, 71(1):147–170, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005404.

Wang:2018:DTN

[WW18a] Xin Wang and Lei Wang. Darboux transformation and nonautonomous solitons for a modified Kadomtsev–Petviashvili equation with variable coefficients. Computers and Mathematics with Applications, 75 (12):4201–4213, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301573.

Wang:2018:CAM

[WW18b] Yuan-Ming Wang and Tao Wang. A compact ADI method and its extrapolation for time fractional sub-diffusion equations with nonhomogeneous Neumann boundary conditions. Computers and Mathematics with Applications, 75(3):721–739, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306247.

Wang:2019:GSD

[WW19a] Jianping Wang and Mingxin Wang. Global solution of a diffusive predator-prey model with prey-taxis. Computers and Mathematics with Applications, 77(10):2676–2694, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930001X.

Wang:2019:EMF

[WW19b] Keyan Wang and Qisheng Wang. Expanded mixed finite element method for second order hyperbolic equations. Computers and Mathematics with Applications, 78(8):2560–2574, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301920.

Wang:2019:ABN

[WW19c] Renhai Wang and Bixiang Wang. Asymptotic behavior of nonautonomous fractional stochastic p-Laplacian equations. Computers and Mathematics with Applications, 78 (11):3527–3543, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302871.

Wang:2011:FOC

[WWA11] Yuan-Ming Wang, Wen-Jia Wu, and Ravi P. Agarwal. A fourthorder compact finite difference method for nonlinear higher-order multipoint boundary value problems. Computers and Mathematics with Applications, 61(11):3226– 3245, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003038.

Wang:2013:NSP

[WWB13] Chuan-Long Wang, Rui-Ping Wen, and Yan-Hong Bai. A new splitting and preconditioner for iteratively solving non-Hermitian positive definite systems.

Computers and Mathematics with Applications, 65(7):1047–1058, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000710.

Wang:2018:ISD

[WWD18] Hui Wang, Yun-Hu Wang, and Huan-He Dong. Interaction solutions of a (2 + 1)-dimensional dispersive long wave system. Computers and Mathematics with Applications, 75(8):2625–2628, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300014.

Wang:2010:CSU

[WWG10] Yan’e Wang, Jianhua Wu, and Gaihui Guo. Coexistence and stability of an unstirred chemostat model with Beddington– DeAngelis function. Computers and Mathematics with Applications, 60(8):2497–2507, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110006310.

Wojtusiak:2012:MLA

[WWH12] Janusz Wojtusiak, Tobias Warden, and Otthein Herzog. Machine learning in agent-based stochastic simulation: Inferential theory and evaluation in transportation logistics. Computers and Mathematics with Applications, 64(12):3658–3665, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001113.

Wang:2013:LBS

[WWLL13] Jia Wang, Donghai Wang, Pierre Lallemand, and Li-Shi Luo. Lattice Boltzmann simulations of thermal convective flows in two dimensions. Computers and Mathematics with Applications, 65(2):262–286, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004671.

Wang:2011:ENS

[WWW11a] Xiao-Yuan Wang, Li-Yan Wang, and Ji-Bo Wang. Erratum to “Notes on ‘Single machine scheduling problems under the effects of nonlinear deterioration and time-dependent learning’ [Math. Comput. Modelling 50 (2009) 401– 406]” [CAMWA 61(5) (2011) 1471–1473]. Computers and Mathematics with Applications, 61(10): 3180, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001702.

See [WWW11b].

Wang:2011:ESM

[WWW11b] Xiao-Yuan Wang, Li-Yan Wang, and Ji-Bo Wang. Erratum to “Single machine scheduling problems under the effects of nonlinear deterioration and time-dependent learning” [math. comput. modelling 50 (2009) 401– 406]. Computers and Mathematics with Applications, 61(5):1471–1473, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000216.

See erratum [WWW11a].

Wang:2014:NHO

[WWW14] Heping Wang, Kai Wang, and Xiaoli Wang. On the norm of the hyperinterpolation operator on the d-dimensional cube. Computers and Mathematics with Applications, 68(5):632–638, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114003150.

Wang:2019:ALR

[WWXW19] Jin Wang, Yan-Ping Wang, Zhi Xu, and Chuan-Long Wang. Accelerated low rank matrix approximate algorithms for matrix completion. Computers and Mathematics with Applications, 77(2):334–341, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305492.

Wang:2012:TOB

[WWZ12] Bo Wang, Yongwei Wu, and Weimin Zheng. Task optimization based on CPU pipeline technique in a multicore system. Computers and Mathematics with Applications, 63(2):536–543, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007929.

Wu:2018:OCS

[WX18a] Tingting Wu and Ruimin Xu. An optimal compact sixth-order finite difference scheme for the Helmholtz equation. Computers and Mathematics with Applications, 75(7):2520–2537, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307861.

Wu:2018:TWS

[WX18b] Yuanyong Wu and Haibin Xiao. Traveling wave solutions for Gause type predator–prey systems with density dependence: a heteroclinic orbit in R^{4}. Computers and Mathematics with Applications, 76(5):1139–1160, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303286.

Wang:2010:KIU

[WXF10] Ren-Hong Wang, Min Xu, and Qin Fang. A kind of improved univariate multiquadric quasiinterpolation operators. Computers and Mathematics with Applications, 59(1):451–456, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003988.

Wang:2011:HOCa

[WXYW11] Shujie Wang, Chao Xu, Peng Yuan, and Yingying Wang. Hydrodynamic optimization of channelling device for hydro turbine based on lattice Boltzmann method. Computers and Mathematics with Applications, 61(12):3722–3729, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002439.

Wang:2011:ADT

[WY11a] Cheng-Chi Wang and Her-Terng Yau. Application of the differential transformation method to bifurcation and chaotic analysis of an AFM probe tip. Computers and Mathematics with Applications, 61(8):1957–1962, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005833.

Wu:2011:MGR

[WY11b] Jinbiao Wu and Xiaoling Yin. An M/G/1 retrial G-queue with nonexhaustive random vacations and an unreliable server. Computers and Mathematics with Applications, 62(5):2314–2329, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005724.

Weiss:2015:UQT

[WY15] Danny Weiss and Zohar Yosibash. Uncertainty quantification for a 1D thermo-hyperelastic coupled problem using polynomial chaos projection and p-FEMs. Computers and Mathematics with Applications, 70(7):1701–1720, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002059.

Wang:2016:SAR

[WY16] Deng-Shan Wang and Yanbin Yin. Symmetry analysis and reductions of the two-dimensional generalized Benney system via geometric approach. Computers and Mathematics with Applications, 71(3):748–757, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115006033.

Wang:2018:TDV

[WY18a] Xinwei Wang and Zhangxian Yuan. Threedimensional vibration analysis of curved and twisted beams with irregular shapes of crosssections by sub-parametric quadrature element method. Computers and Mathematics with Applications, 76(6):1486–1499, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303663.

Wen:2018:LIR

[WY18b] Zhihong Wen and Zhuan Ye. A logarithmically improved regularity criterion for the surface quasi-geostrophic equation. Computers and Mathematics with Applications, 75 (4):1368–1377, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307046.

Wang:2019:ABS

[WY19a] Shu Wang and Fang Yuan. The asymptotic behavior of the solutions of the Black–Scholes equation as volatility σ → 0+. Computers and Mathematics with Applications, 78(3):1037–1050, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301531.

Wu:2019:PSG

[WY19b] Yuanqing Wu and Maoqing Ye. A parallel sparse grid construction algorithm based on the shared memory architecture and its application to flash calculations. Computers and Mathematics with Applications, 77(8):2114–2129, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307004.

Wang:2010:GBI

[WYD10] Dongqing Wang, Guowei Yang, and Ruifeng Ding. Gradient-based iterative parameter estimation for Box–Jenkins systems. Computers and Mathematics with Applications, 60(5):1200–1208, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004128.

Wang:2011:SP

[WYG11] Jinyan Wang, Minghao Yin, and Wenxiang Gu. Soft polygroups. Computers and Mathematics with Applications, 62 (9):3529–3537, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100753X.

Wang:2012:NIV

[WYG12] Jinyan Wang, Minghao Yin, and Wenxiang Gu. Notes on: “Intervalvalued intuitionistic fuzzy soft sets and their properties” [Comput. Math. Appl. 60 (2010) 906– 918]. Computers and Mathematics with Applications, 64(9):2954–2960, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003525.

See [JTC^{+}10a, Jia11, Wan13a].

Wen:2010:NRM

[WYK10] Meilin Wen, Cuilian You, and Rui Kang. A new ranking method to fuzzy data envelopment analysis. Computers and Mathematics with Applications, 59(11):3398– 3404, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001537.

Wang:2019:APV

[WYL19] Xiaoyan Wang, Junyuan Yang, and Xiaofeng Luo. Asymptotical profiles of a viral infection model with multi-target cells and spatial diffusion. Computers and Mathematics with Applications, 77(2):389–406, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305558.

Wan:2010:SCA

[WYLZ10] Min Wan, Zhang Yi, Jian Cheng Lv, and Jiliu Zhou. Stability and chaos analysis for an ICA algorithm. Computers and Mathematics with Applications, 60(6):1810–1827, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000489X.

Wang:2012:PSA

[WYN12] Li Wang, Mei Yu, and Pengcheng Niu. Periodic solution and almost periodic solution of impulsive Lasota–Wazewska model with multiple timevarying delays. Computers and Mathematics with Applications, 64(8):2383–2394, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003987.

Wang:2011:STR

[WYY11] Chonghai Wang, Li Yan Yuan, and Jia-Huai You. On the semantics of top-k ranking for objects with uncertain data. Computers and Mathematics with Applications, 62(7):2812–2823, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006183.

Wu:2010:ELB

[WZ10] Junliang Wu and Yan Zhang. Estimate for the lower bound of rank and the upper bound of eigenvalues norms’ sum of given quaternion matrix. Computers and Mathematics with Applications, 59(9):3160–3166, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001628.

Wang:2011:STT

[WZ11a] Peiguang Wang and Zhengran Zhan. Stability in terms of two measures of dynamic system on time scales. Computers and Mathematics with Applications, 62 (12):4717–4725, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009382.

Wu:2011:GCP

[WZ11b] Yusen Wu and Cui Zhang. The generalized center problem of resonant infinity for a polynomial differential system. Computers and Mathematics with Applications, 62(1):151– 157, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003658.

Wu:2015:EGS

[WZ15] Ke Wu and Fen Zhou. Existence of ground state solutions for a quasilinear Schrödinger equation with critical growth. Computers and Mathematics with Applications, 69(2):81–88, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122114005550.

Wang:2016:AMR

[WZ16] Aizeng Wang and Gang Zhao. An algorithm for multi-resolution analysis of NURBS surfaces. Computers and Mathematics with Applications, 72(3):642–652, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302863.

Wang:2017:BSN

[WZ17a] Hanchao Wang and Likai Zhou. Bandwidth selection of nonparametric threshold estimator in jump-diffusion models. Computers and Mathematics with Applications, 73(2):211–219, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306289.

Wang:2017:DFR

[WZ17b] Jueyu Wang and Detong Zhu. Derivative-free restrictively preconditioned conjugate gradient path method without line search technique for solving linear equality constrained optimization. Computers and Mathematics with Applications, 73(2):277–293, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211630654X.

Wang:2018:FFD

[WZ18a] Zhaojuan Wang and Lingping Zhang. Finite fractal dimension of random attractor for stochastic non-autonomous strongly damped wave equation. Computers and Mathematics with Applications, 75(9):3343–3357, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300750.

Wang:2018:PEA

[WZ18b] Zhoufeng Wang and Yunzhang Zhang. A posteriori error analysis for the conical diffraction problem. Computers and Mathematics with Applications, 74(5):993–1005, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302559.

Wei:2018:BPT

[WZ18c] Ting Wei and Yun Zhang.

The backward problem for a time-fractional diffusion-wave equation in a bounded domain. Computers and Mathematics with Applications, 75(10):3632–3648, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301032.

Wang:2019:NHO

[WZC^{+}19a] Fenling Wang, Yanmin Zhao, Chen Chen, Yabing Wei, and Yifa Tang. A novel high-order approximate scheme for two-dimensional timefractional diffusion equations with variable coefficient. Computers and Mathematics with Applications, 78(5):1288–1301, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306825.

Wang:2019:NEL

[WZC^{+}19b] Qiao Wang, Wei Zhou, Yonggang Cheng, Gang Ma, and Xiaolin Chang. NURBS-enhanced line integration boundary element method for 2D elasticity problems with body forces. Computers and Mathematics with Applications, 77(7):2006–2028, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306928.

Wang:2010:UAC

[WZCC10] Hongchao Wang, Hongke Zhang, Chi-Yuan Chang, and Han-Chieh Chao. A universal access control method based on host identifiers for future Internet. Computers and Mathematics with Applications, 60(2):176– 186, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000040.

Wang:2012:NIP

[WZF12a] JinRong Wang, Yong Zhou, and Michal Fečkan. Nonlinear impulsive problems for fractional differential equations and Ulam stability. Computers and Mathematics with Applications, 64 (10):3389–3405, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001356.

Wang:2012:RDT

[WZF12b] JinRong Wang, Yong Zhou, and Michal Fečkan. On recent developments in the theory of boundary value problems for impulsive fractional differential equations. Computers and Mathematics with Applications, 64 (10):3008–3020, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011357.

Wang:2016:EMS

[WZF16] Liben Wang, Xingyong Zhang, and Hui Fang. Existence and multiplicity of solutions for a class of (ϕ_{1}, ϕ_{2})-Laplacian elliptic system in R^{N} via genus theory. Computers and Mathematics with Applications, 72(1):110– 130, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302279.

Wang:2019:PSP

[WZG19] Jixia Wang, Pan Zhao, and Qinghui Gao. Portfolio selection problem with nonlinear wealth equations under non-extensive statistical mechanics for time-varying SDE. Computers and Mathematics with Applications, 77(2):555–564, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305777.

Wang:2010:TPS

[WZH10] Youyu Wang, Meng Zhao, and Yinping Hu. Triple positive solutions for a multi-point boundary value problem with a onedimensional p-Laplacian. Computers and Mathematics with Applications, 60(6):1792–1802, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110004852.

Wei:2018:THT

[WZH18] Xiaodong Wei, Yongjie Jessica Zhang, and Thomas J. R. Hughes. Truncated hierarchical tricubic C^{0} spline construction on unstructured hexahedral meshes for isogeometric analysis applications. Computers and Mathematics with Applications, 74(9):2203–2220, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304728.

Wittmann:2013:CDP

[WZHW13] Markus Wittmann, Thomas Zeiser, Georg Hager, and Gerhard Wellein. Comparison of different propagation steps for lattice Boltzmann methods. Computers and Mathematics with Applications, 65(6):924–935, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003835.

Wei:2012:NSB

[WZKY12] Leilei Wei, Xindong Zhang, Sunil Kumar, and Ahmet Yildirim. A numerical study based on an implicit fully discrete local discontinuous Galerkin method for the time-fractional coupled Schrödinger system. Computers and Mathematics with Applications, 64(8):2603–2615, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004701.

Wang:2016:EIF

[WZM^{+}16] Hanquan Wang, Yong Zhang, Xiu Ma, Jun Qiu, and Yan Liang. An efficient implementation of fourth-order compact finite difference scheme for Poisson equation with Dirichlet boundary conditions. Computers and Mathematics with Applications, 71(9):1843–1860, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300761.

Wu:2018:CRM

[WZMY18] Pinxia Wu, Yufeng Zhang, Iqbal Muhammad, and Qiqi Yin. Complexiton and resonant multiple wave solutions to the (2 + 1)-dimensional Konopelchenko–Dubrovsky equation. Computers and Mathematics with Applications, 76(4):845–853, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830292X.

Wang:2010:APM

[WZW10] Guangmin Wang, Kejun Zhu, and Zhongping Wan. An approximate programming method based on the simplex method for bilevel programming problem. Computers and Mathematics with Applications, 59(10):3355– 3360, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002038.

Wanyan:2011:PAA

[WZWS11] Xiaoru Wanyan, Damin Zhuang, Hengyang Wei, and Jianshuang Song. Pilot attention allocation model based on fuzzy theory. Computers and Mathematics with Applications, 62(7):2727–2735, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122111005517.

Wang:2011:NPF

[WZWX11] JinRong Wang, Yong Zhou, Wei Wei, and Honglei Xu. Nonlocal problems for fractional integrodifferential equations via fractional operators and optimal controls. Computers and Mathematics with Applications, 62(3):1427–1441, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111001398.

Wang:2011:FPG

[WZXL11] Juan Wang, Hongchan Zheng, Feng Xu, and Dekong Liu. Fractal properties of the generalized Chaikin corner-cutting subdivision scheme. Computers and Mathematics with Applications, 61(8): 2197–2200, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007005.

Wang:2017:DVI

[WZXS17] Shaoli Wang, Jiafang Zhang, Fei Xu, and Xinyu Song. Dynamics of virus infection models with density-dependent diffusion. Computers and Mathematics with Applications, 74 (10):2403–2422, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304364.

Wu:2013:RAM

[WZY13] Yujiang Wu, Na Zhang, and Jinyun Yuan. A robust adaptive method for singularly perturbed convection–diffusion problem with two small parameters. Computers and Mathematics with Applications, 66(6):996–1009, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004008.

Xia:2013:SCD

[XA13] Fei Xia and Richard L.

Axelbaum. Simplifying the complexity of diffusion flames through interpretation in C/O ratio space. Computers and Mathematics with Applications, 65(10):1625– 1632, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113000254.

Xu:2019:TSF

[XaZH19] Pengfei Xu, Guang an Zou, and Jianhua Huang. Time-space fractional stochastic Ginzburg–Landau equation driven by fractional Brownian motion. Computers and Mathematics with Applications, 78(12):3790–3806, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119303074.

Xu:2016:MSM

[XBHN16] Feifei Xu, Robert Bierman, Frank Healy, and Hoa Nguyen. A multiscale model of Escherichia coli chemotaxis from intracellular signaling pathway to motility and nutrient uptake in nutrient gradient and isotropic fluid environments. Computers and Mathematics with Applications, 71(11): 2466–2478, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.

com/science/article/
pii/S0898122115005878.

Xie:2011:ESA

[XC11a] Tingfan Xie and Feilong Cao. The errors of simultaneous approximation of multivariate functions by neural networks. Computers and Mathematics with Applications, 61(10):3146– 3152, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002859.

Xu:2011:UAS

[XC11b] Aimin Xu and Zhongdi Cen. A unified approach to some recurrence sequences via Faà di Bruno’s formula. Computers and Mathematics with Applications, 62(1):253–260, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003932.

Xu:2013:FOM

[XC13] Zheng Xu and Wen Chen. A fractionalorder model on new experiments of linear viscoelastic creep of hami melon. Computers and Mathematics with Applications, 66(5):677–681, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000631.

Xu:2016:IPS

[XC16] Wei-Ru Xu and Guo-Liang Chen. Inverse problems for (R, S)-symmetric matrices in structural dynamic model updating. Computers and Mathematics with Applications, 71(5):1074–1088, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300220.

Xu:2017:SLM

[XC17] Wei-Ru Xu and Guo-Liang Chen. The solutions to linear matrix equations AX = B, Y A = D with kinvolutory symmetries. Computers and Mathematics with Applications, 73(8):1741–1759, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122117300810.

Xie:2018:EMN

[XC18] Weihong Xie and Haibo Chen. Existence and multiplicity of normalized solutions for the nonlinear Kirchhoff type problems. Computers and Mathematics with Applications, 76(3):579–591, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302529.

Xu:2012:IFP

[XCM12] Xiaoli Xu, Tao Chen, and Mamoru Minami. Intelligent fault prediction system based on Internet of Things. Computers and Mathematics with Applications, 64(5):833–839, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011059.

Xu:2018:ABA

[XCS18] Wei-Ru Xu, Guo-Liang Chen, and Xing-Ping Sheng. Analytical best approximate Hermitian and generalized skew-Hamiltonian solution of matrix equation AXAH+ CY CH = F. Computers and Mathematics with Applications, 75(10):3702–3718, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830107X.

Xu:2010:FNB

[XCXW10] Xirong Xu, Yongchang Cao, Jun-Ming Xu, and Yezhou Wu. Feedback numbers of de Bruijn digraphs. Computers and Mathematics with Applications, 59(2):716–723, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007202.

Xue:2016:HPW

[XCZQ16] Xiaofeng Xue, Xuefeng Chen, Xingwu Zhang, and Baijie Qiao. Hermitian plane wavelet finite element method: Wave propagation and load identification. Computers and Mathematics with Applications, 72 (12):2920–2942, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305909.

Xing:2010:PIR

[XD10] Yuming Xing and Shusen Ding. Poincaré inequalities with the Radon measure for differential forms. Computers and Mathematics with Applications, 59(6):1944–1952, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007421.

Xia:2017:TPF

[XD17] Zhi Xia and Kui Du.

A tensor product finite element method for the diffraction grating problem with transparent boundary conditions. Computers and Mathematics with Applications, 73(4):628–639, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300111.

Xie:2016:TRS

[XDH16] Junhui Xie, Qiuyi Dai, and Haiyang He. Threshold results for semilinear parabolic systems. Computers and Mathematics with Applications, 71(12):2585–2593, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302085.

Xue:2012:EIS

[XDL12] Chaogai Xue, Lili Dong, and Junjuan Liu. Enterprise information system structure optimization based on time property with improved immune genetic algorithm and binary tree. Computers and Mathematics with Applications, 63(7):1155–1168, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211101087X.

Xiao:2019:NSC

[XFH19] Xufeng Xiao, Xinlong Feng, and Yinnian He. Numerical simulations for the chemotaxis models on surfaces via a novel characteristic finite element method. Computers and Mathematics with Applications, 78(1):20– 34, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300720.

Xenophontos:2016:FEA

[XFL16] Christos Xenophontos, Sebastian Franz, and Lars Ludwig. Finite element approximation of convection–diffusion problems using an exponentially graded mesh. Computers and Mathematics with Applications, 72(6):1532–1540, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303996.

Xiao:2018:LMF

[XFY18] Xufeng Xiao, Xinlong Feng, and Jinyun Yuan. The lumped mass finite element method for surface parabolic problems: Error estimates and maximum principle. Computers and Mathematics with Applications, 76(3):488–507, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302372.

Xiao:2010:NAC

[XG10] Hong Xiao and Zydrunas Gimbutas. A numerical algorithm for the construction of efficient quadrature rules in two and higher dimensions. Computers and Mathematics with Applications, 59(2):663–676, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007263.

Xu:2017:GDD

[XGH17a] Jinhu Xu, Yan Geng, and Jiangyong Hou. Global dynamics of a diffusive and delayed viral infection model with cellular infection and nonlinear infection rate. Computers and Mathematics with Applications, 73(4):640–652, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300093.

Xu:2017:NSF

[XGH17b] Jinhu Xu, Yan Geng, and Jiangyong Hou. A non-standard finite difference scheme for a delayed and diffusive viral infection model with general nonlinear incidence rate. Computers and Mathematics with Applications, 74(8):1782–1798, October 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303991.

Xiao:2010:EDS

[XGXZ10] Zhi Xiao, Ke Gong, Sisi Xia, and Yan Zou. Exclusive disjunctive soft sets. Computers and Mathematics with Applications, 59(6):2128–2137, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007779.

See corrigendum [GXZ11].

Xu:2011:MSE

[XH11a] Liguang Xu and Danhua He. Mean square exponential stability analysis of impulsive stochastic switched systems with mixed delays. Computers and Mathematics with Applications, 62(1):109– 117, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003592.

Xu:2011:SMP

[XH11b] Yufeng Xu and Zhimin He. The short memory principle for solving Abel differential equation of fractional order. Computers and Mathematics with Applications, 62(12):4796–4805, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009473.

Xu:2013:NAS

[XHA13] Yufeng Xu, Zhimin He, and Om P. Agrawal. Numerical and analytical solutions of new generalized fractional diffusion equation. Computers and Mathematics with Applications, 66 (10):2019–2029, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005233.

Xu:2019:NAH

[XHH^{+}19] Wei Xu, Ziping Huang, Weimin Han, Wenbin Chen, and Cheng Wang.

Numerical analysis of history-dependent hemivariational inequalities and applications to viscoelastic contact problems with normal penetration. Computers and Mathematics with Applications, 77(10):2596– 2607, May 15, 2019. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118307363.

Xie:2014:IMS

[XHM14] Yajun Xie, Na Huang, and Changfeng Ma. Iterative method to solve the generalized coupled Sylvester-transpose linear matrix equations over reflexive or anti-reflexive matrix. Computers and Mathematics with Applications, 67(11):2071– 2084, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001722.

Xiang:2011:LCV

[Xia11] Dao-Hong Xiang. Logistic classification with varying Gaussians. Computers and Mathematics with Applications, 61(2):397– 407, January 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008655.

Xu:2018:CCM

[XJ18] Huanying Xu and Xiaoyun Jiang. Creep constitutive models for viscoelastic materials based on fractional derivatives. Computers and Mathematics with Applications, 73(6):1377–1384, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302450.

Xu:2010:IPC

[XJLX10] M. H. Xu, J. L. Jiang, B. Li, and B. Xu. An improved prediction– correction method for monotone variational inequalities with separable operators. Computers and Mathematics with Applications, 59(6):2074–2086, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007603.

Xu:2017:GLA

[XJYL17] Haiyong Xu, Gangyi Jiang, Mei Yu, and Ting Luo. A global and local active contour model based on dual algorithm for image segmentation. Computers and Mathematics with Applications, 74(6):1471–1488, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303759.

Xu:2010:MRE

[XKH10] Haiyan Xu, D. Marc Kilgour, and Keith W. Hipel. Matrix representation and extension of coalition analysis in group decision support. Computers and Mathematics with Applications, 60(5):1164–1176, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004086.

Xu:2010:SGD

[XL10] Qing-Hua Xu and Tai-Shun Liu. Sharp growth and distortion theorems for a subclass of biholomorphic mappings.

Computers and Mathematics with Applications, 59(12):3778–3784, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000283X.

Xu:2011:SCB

[XL11] Qing-Hua Xu and Tai-Shun Liu. Subordination chains and biholomorphic mappings on bounded balanced pseudoconvex domains. Computers and Mathematics with Applications, 61(7):1830–1836, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000976.

Xiang:2015:SIP

[XL15] Huili Xiang and Bin Liu. Solving the inverse problem of an SIS epidemic reaction–diffusion model by optimal control methods. Computers and Mathematics with Applications, 70(5):805–819, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122115002709.

Xu:2011:CEP

[XLD11a] Xingye Xu, Luanying Lian, and Lokenath Debnath. Corrigendum to “Existence of positive solutions of singular elliptic boundary value problems in a ball” [Comput. Math. Appl. 61 (2011) 1335–1341]. Computers and Mathematics with Applications, 61(9): 2932, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001726.

See [XLD11b].

Xu:2011:EPS

[XLD11b] Yingye Xu, Luanying Lian, and Lokenath Debnath. Existence of positive solutions of singular elliptic boundary value problems in a ball. Computers and Mathematics with Applications, 61(5):1335–1341, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009922.

See corrigendum [XLD11a].

Xu:2012:TME

[XLF12] Lan Xu, Fujuan Liu, and Naeem Faraz. Theoretical model for the electrospinning nanoporous materials process. Computers and Mathematics with Applications, 64 (5):1017–1021, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002246.

Xu:2011:ACR

[XLK11] Yuguang Xu, Zeqing Liu, and Shin Min Kang. Accumulation and control of random errors in the Ishikawa iterative process in arbitrary Banach space. Computers and Mathematics with Applications, 61(8):2217–2220, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007066.

Li:2012:LTB

[xLlFwWL12] Da xiang Li, Jiu lun Fan, Dian wei Wang, and Ying Liu. Latent topic based multi-instance learning method for localized content-based image retrieval. Computers and Mathematics with Applications, 64(4):500–510, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010856.

Xue:2017:GSS

[XLT17] Yan-Fang Xue, Jiu Liu, and Chun-Lei Tang. A ground state solution for an asymptotically periodic quasilinear Schrödinger equation. Computers and Mathematics with Applications, 74(6):1143–1157, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303486.

Xu:2010:FPU

[XLY10] Jun-Feng Xu, Feng Lü, and Hong-Xun Yi. Fixedpoints and uniqueness of meromorphic functions. Computers and Mathematics with Applications, 59(1):9–17, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004520.

Xing:2011:SDU

[XLZW11] Chao Xing, Yanjun Li, Ke Zhang, and Ling Wang. Shadow detecting using particle swarm optimization and the Kolmogorov test. Computers and Mathematics with Applications, 62(7):2704–2711, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005323.

Xie:2015:SCG

[XM15] Ya-Jun Xie and Chang-Feng Ma. The scaling conjugate gradient iterative method for two types of linear matrix equations. Computers and Mathematics with Applications, 70(5):1098–1113, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115003168.

Xu:2010:GSD

[XMW10] Rui Xu, Zhien Ma, and Zhiping Wang. Global stability of a delayed SIRS epidemic model with saturation incidence and temporary immunity. Computers and Mathematics with Applications, 59(9):3211–3221, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001690.

Xu:2010:VSS

[XMWH10] Wei Xu, Jian Ma, Shouyang Wang, and Gang Hao. Vague soft sets and their properties. Computers and Mathematics with Applications, 59(2):787– 794, January 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109007147.

See notes [HLY12a].

Xin:2010:CBF

[XS10] Xiu Xin and Fu-Gui Shi. Categories of bifuzzy pre-matroids. Computers and Mathematics with Applications, 59(4):1548–1558, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000155.

Xia:2011:TOH

[XSLS11] Shugao Xia, Xiquan Shi, Fengshan Liu, and Zhixun Su. The tight orthogonal homotopic bases of closed oriented triangulated surfaces and their computing. Computers and Mathematics with Applications, 61(10):2944– 2951, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002161.

Xu:2019:ATF

[XSYL19] Liyang Xu, Tianlong Shen, Xuejun Yang, and Jiarui Liang. Analysis of time fractional and space nonlocal stochastic incompressible Navier– Stokes equation driven by white noise. Computers and Mathematics with Applications, 78 (5):1669–1680, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307144.

Xu:2011:HAP

[Xu11a]  Lan Xu.  A Hamil-

tonian approach for a plasma physics problem. Computers and Mathematics with Applications, 61(8):1909–1911, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004426.

Xu:2011:GDH

[Xu11b]  Rui Xu.  Global dy-

namics of an HIV-1 infection model with distributed intracellular delays. Computers and Mathematics with Applications, 61(9):2799–2805, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002136.

Xu:2014:DGP

[Xu14] Shenghu Xu. Dynamics of a general prey–predator model with prey-stage structure and diffusive effects. Computers and Mathematics with Applications, 68(3):405–423, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002776.

Xue:2013:ASA

[Xue13] Fei Xue. On asymptotic summations of almost diagonal difference systems and their applications. Computers and Mathematics with Applications, 66(11):2156–2164, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005804.

Xu:2010:SEF

[XW10] Aimin Xu and Chengjing Wang. Some extensions of Faà di Bruno’s formula with divided differences. Computers and Mathematics with Applications, 59(6):2047–2052, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007561.

Xia:2014:GFM

[XW14] Kelin Xia and Guo-Wei Wei. A Galerkin formulation of the MIB method for three dimensional elliptic interface problems. Computers and Mathematics with Applications, 68(7):719–745, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400340X.

Xu:2018:ITO

[XW18] Weizheng Xu and Weiguo Wu. Improvement of third-order WENO-Z scheme at critical points. Computers and Mathematics with Applications, 75(9):3431–3452, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300828.

Xian:2019:DID

[XW19a] J. Xian and T. Wei. Determination of the initial data in a time-fractional diffusion-wave problem by a final time data. Computers and Mathematics with Applications, 78(8):2525–2540, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301877.

Xu:2019:EBB

[XW19b] Xiangjian Xu and Qing-Wen Wang. Extending BiCG and BiCR methods to solve the Stein tensor equation. Computers and Mathematics with Applications, 77(12):3117–3127, June 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300549.

Xu:2016:NCT

[XWH16] Guang-Ying Xu, Jin-Bao Wang, and Zhi Han. Notes on “The Cattaneotype time fractional heat conduction equation for laser heating” [Comput. Math. Appl. 66 (2013) 824–831]. Computers and Mathematics with Applications, 71(10):2132– 2137, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301274.

See [QXG13].

Xu:2018:NMS

[XWL18] Weizheng Xu, Weiguo Wu, and Yongshui Lin. Numerical method and simplified analytical model for predicting the blast load in a partially confined chamber. Computers and Mathematics with Applications, 76(2):284– 314, July 15, 2018. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118302177.

Xu:2011:NSM

[XWN11] Lan Xu, Yue Wu, and Yasir Nawaz. Numerical study of magnetic electrospinning processes. Computers and Mathematics with Applications, 61(8):2116–2119, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000670X.

Xiang:2011:GEN

[XWY11] Zhaoyin Xiang, Ying Wang, and Huizhi Yang. Global existence and nonexistence for degenerate parabolic equations with nonlinear boundary flux. Computers and Mathematics with Applications, 62(8):3056–3065, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006808.

Xiao:2017:ESS

[XWY17] Xiao-Yong Xiao, Xiang Wang, and Hong-Wei Yin. Efficient singlestep preconditioned HSS iteration methods for complex symmetric linear systems. Computers and Mathematics with Applications, 74 (10):2269–2280, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304248.

Xiao:2018:EPN

[XWY18] Xiao-Yong Xiao, Xiang Wang, and Hong-Wei Yin. Efficient preconditioned NHSS iteration methods for solving complex symmetric linear systems. Computers and Mathematics with Applications, 75(1):235–247, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305527.

Xia:2010:ESD

[XX10] Tiecheng Xia and Shouquan Xiong. Exact solutions of (2 + 1)-dimensional Bogoyavlenskii’s breaking soliton equation with symbolic computation. Computers and Mathematics with Applications, 60(3):919–923, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003950.

Xu:2017:SNM

[XX17] Hong-Ru Xu and Shui-Lian Xie. A semismooth Newton method for a kind of HJB equation. Computers and Mathematics with Applications, 73(12):2581–2586, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302031.

Xu:2019:QST

[XX19] Qinwu Xu and Yufeng Xu. Quenching study of two-dimensional fractional reaction–diffusion equation from combustion process. Computers and Mathematics with Applications, 78 (5):1490–1506, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302044.

Xu:2010:DNF

[XXG10] M.-R. Xu, S.-P. Xu, and H.-Y. Guo. Determination of natural frequencies of fluidconveying pipes using homotopy perturbation method. Computers and Mathematics with Applications, 60(3):520–527, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003408.

Xu:2018:STW

[XXH18] Zhiting Xu, Youqing Xu, and Yehui Huang. Stability and traveling waves of a vaccination model with nonlinear incidence. Computers and Mathematics with Applications, 75(2):561–581, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306223.

Xiong:2010:APS

[XY10] Wanmin Xiong and Guangxue Yue. Almost periodic solutions for a class of fourth-order nonlinear differential equations with a deviating argument. Computers and Mathematics with Applications, 60(5):1184–1190, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004104.

Xu:2011:CAC

[XY11] Fang Xu and Zhang Yi.

Continuous attractors of a class of neural networks with a large number of neurons. Computers and Mathematics with Applications, 62 (10):3785–3795, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007954.

Xie:2014:PPP

[XY14] Wei-Si Xie and Yu-Fei Yang. A projection proximal-point algorithm for MR imaging using the hybrid regularization model. Computers and Mathematics with Applications, 67(12):2268– 2278, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000947.

Xiao:2015:SEM

[XY15] Xiaoyong Xiao and Hongwei Yin. A simple and efficient method with high order convergence for solving systems of nonlinear equations. Computers and Mathematics with Applications, 69(10):1220– 1231, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001236.

Xiao:2016:GMP

[XY16] Feng Xiao and Xiaolong Yin. Geometry models of porous media based on Voronoi tessellations and their porosity-permeability relations. Computers and Mathematics with Applications, 72(2):328–348, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004393.

Xiao:2017:EPH

[XY17] Xiao-Yong Xiao and Hong-Wei Yin. Efficient parameterized HSS iteration methods for complex symmetric linear systems. Computers and Mathematics with Applications, 73(1):87–95, January 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306058.

Yang:2011:SBF

[xYsHjL11] Li xin Yang, Wan sheng He, and Xiao jun Liu. Synchronization between a fractional-order system and an integer order system. Computers and Mathematics with Applications, 62 (12):4708–4716, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009370.

Xiazhi:2016:MPF

[XYXZ16] Hao Xiazhi, Liu Yinping, Tang Xiaoyan, and Li Zhibin. A Maple package for finding interaction solutions of nonlinear evolution equations. Computers and Mathematics with Applications, 72(9):2450–2461, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116305089.

Xiao:2010:NCS

[XZ10] Y. Xiao and H. Y. Zhang. A note on convergence of semi-implicit Euler methods for stochastic pantograph equations. Computers and Mathematics with Applications, 59(4):1419–1424, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007901.

Xiong:2011:TDL

[XZ11] Wenjuan Xiong and Junfeng Zhang. A two-dimensional lattice Boltzmann model for uniform channel flows. Computers and Mathematics with Applications, 61(12):3453–3460, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001598.

Xie:2017:RCI

[XZ17] Hongyan Xie and Mingxuan Zhu. A regularity criterion for 3D ideal density-dependent MHD system. Computers and Mathematics with Applications, 73(10):2233–2237, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301669.

Xie:2018:HOM

[XZ18] Jianqiang Xie and Zhiyue Zhang. The high-order multistep ADI solver for two-dimensional nonlinear delayed reaction– diffusion equations with variable coefficients. Computers and Mathematics with Applications, 75(10):3558–3570, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300907.

Xu:2019:GEB

[XZ19] Guangyu Xu and Jun Zhou. Global existence and blow-up of solutions to a class of nonlocal parabolic equations. Computers and Mathematics with Applications, 78(3):979–996, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301439.

Xiao:2012:SSS

[XZC12] Jian-Zhong Xiao, Xing-Hua Zhu, and Rong Cheng. The solution sets for second order semilinear impulsive multivalued boundary value problems. Computers and Mathematics with Applications, 64(2):147– 160, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001290.

Xue:2010:NGO

[XZL10] Qiaoling Xue, Jian Zhu, and Wenjun Liu. A new generalization of Ostrowski-type inequality involving functions of two independent variables. Computers and Mathematics with Applications, 60(8):2219–2224, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110005754.

Xu:2011:DSS

[XZL^{+}11] Aiguo Xu, Guangcai Zhang, Hua Li, Yangjun Ying, and Jianshi Zhu. Dynamical similarity in shock wave response of porous material: From the view of pressure. Computers and Mathematics with Applications, 61(12):3618–3627, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007820.

Xiang:2016:ESB

[XZR16] Mingqi Xiang, Binlin Zhang, and Vicenţiu D. Rădulescu. Existence of solutions for a binonlocal fractional p-Kirchhoff type problem. Computers and Mathematics with Applications, 71(1):255–266, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500560X.

Xiao:2016:IET

[XZZ16] Aiguo Xiao, Gengen Zhang, and Jie Zhou. Implicit-explicit time discretization coupled with finite element methods for delayed predator–prey competition reaction– diffusion system. Computers and Mathematics with Applications, 71(10): 2106–2123, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301717.

Yarahmadi:2011:BEF

[YA11] Zahra Yarahmadi and Ali Reza Ashrafi. The bipartite edge frustration of graphs under subdivided edges and their related sums. Computers and Mathematics with Applications, 62(1):319– 325, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004007.

Yakar:2011:SGC

[Yak11] Ali Yakar. Some generalizations of comparison results for fractional differential equations. Computers and Mathematics with Applications, 62(8):3215–3220, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006985.

Yang:2010:BBH

[Yan10a] Ming-Chien Yang. Bipanconnectivity of balanced hypercubes. Computers and Mathematics with Applications, 60(7):1859–1867, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000492X.

Yang:2010:RGA

[Yan10b]  Wengui Yang.  Re-

finements of generalized Aczél–Popoviciu’s inequality and Bellman’s inequality. Computers and Mathematics with Applications, 59(11):3570– 3577, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002397.

Yang:2010:PCL

[Yan10c]  Zhaoxia Yang.  Pres-

sure condition for lattice Boltzmann methods on domains with curved boundaries. Computers and Mathematics with Applications, 59(7):2168– 2177, April 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006518.

Yang:2010:EPS

[Yan10d]  Zhilin Yang.  Exis-

tence of positive solutions for a system of generalized Lidstone problems. Computers and Mathematics with Applications, 60(3):501–510, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110003299.

Yang:2011:FSS

[Yan11a] Cheng-Fu Yang. Fuzzy soft semigroups and fuzzy soft ideals. Computers and Mathematics with Applications, 61(2):255– 261, January 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008370.

Yang:2011:WCG

[Yan11b]  Wengui Yang.  On

weighted q-Čebyšev–Grüss type inequalities. Computers and Mathematics with Applications, 61(5): 1342–1347, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111000046.

Yang:2011:PSS

[Yan11c] Zhilin Yang. Positive solutions for a system of p-Laplacian boundary value problems. Computers and Mathematics with Applications, 62 (12):4429–4438, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008832.

Yang:2011:PST

[Yan11d] Zhilin Yang. Positive solutions of a 2n th-order boundary value problem involving all derivatives via the order reduction method. Computers and Mathematics with Applications, 61(4):822–831, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009351.

Yang:2012:OBA

[Yan12a] Shyi-Kae Yang. Observerbased anti-windup compensator design for saturated control systems using an LMI approach. Computers and Mathematics with Applications, 64(5):747–758, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010352.

Yang:2012:PSC

[Yan12b]  Wengui Yang.  Posi-

tive solutions for a coupled system of nonlinear fractional differential equations with integral boundary conditions. Computers and Mathematics with Applications, 63(1):288–297, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010017.

Yang:2013:LBO

[Yan13]  Zhaoxia Yang.  Lat-

tice Boltzmann outflow treatments: Convective conditions and others. Computers and Mathematics with Applications, 65(2):160–171, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122112006736.

Yang:2015:CLS

[Yan15]  Chaoxia Yang.  Con-

vergence of a linearized second-order BDF–FEM for nonlinear parabolic interface problems. Computers and Mathematics with Applications, 70(3): 265–281, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115002175.

Yang:2017:CWH

[Yan17]  Xi Yang.  On con-

vergence of the WR-HSS iteration method for a system of linear differential equations. Computers and Mathematics with Applications, 73(11):2470–2481, June 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302006.

Yang:2018:ECD

[Yan18a] Wenbin Yang. Effect of cross-diffusion on the stationary problem of a predator-prey system with a protection zone. Computers and Mathematics with Applications, 76(9):2262–2271, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304450.

Yang:2018:NCI

[Yan18b]  Xi Yang.  Numerical

contour integral methods for unsteady Stokes equations. Computers and Mathematics with Applications, 75 (12):4414–4426, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301755.

Yang:2018:RGP

[Yan18c]  Xi Yang.  A relaxed

generalized-PSS preconditioner for saddle-point linear systems from steady incompressible Navier– Stokes equations. Computers and Mathematics with Applications, 76(8):1906–1922, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304097.

Yang:2019:RCS

[Yan19] Jiaqi Yang. Regularity criteria for 3D shear thinning fluids via two velocity components. Computers and Mathematics with Applications, 77(10):2854–2858, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300410.

Yao:2010:RTU

[Yao10] Fengping Yao. Regularity theory for the uniformly elliptic operators in Orlicz spaces. Computers and Mathematics with Applications, 60 (12):3098–3104, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000790X.

Yao:2016:ILM

[Yao16] Guangming Yao. An improved localized method of approximate particular solutions for solving elliptic PDEs. Computers and Mathematics with Applications, 71(1):171–184, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005398.

Yildirim:2011:ANO

[YAS^{+}11] A. Yildirim, H. Askari, Z. Saadatnia, M. KalamiYazdi, and Y. Khan. Analysis of nonlinear oscillations of a punctual charge in the electric field of a charged ring via a Hamiltonian approach and the energy balance method. Computers and Mathematics with Applications, 62(1):486– 490, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004275.

See corrigendum [Yaz11].

Yaslan:2012:HOP

[Yas12] İsmail Yaslan. Higher order m-point boundary value problems on time scales. Computers and Mathematics with Applications, 63(3):739–750, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010212.

Younesian:2010:FAS

[YASK10] Davood Younesian, Hassan Askari, Zia Saadatnia, and Mohammad KalamiYazdi. Frequency analysis of strongly nonlinear generalized Duffing oscillators using He’s frequency-amplitude formulation and He’s energy balance method. Computers and Mathematics with Applications, 59(9):3222– 3228, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001896.

Yazdi:2011:CAN

[Yaz11] M. Kalami Yazdi. Corrigendum to “Analysis of nonlinear oscillations of a punctual charge in the electric field of a charged ring via a Hamiltonian approach and the energy balance method” [Computers & Mathematics with Applications 62 (2011) 486–490]. Computers and Mathematics with Applications, 62 (6):2681–2682, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006699.

See [YAS^{+}11].

Yang:2013:CSM

[YB13] Jianhui Yang and Edo S. Boek. A comparison study of multicomponent lattice Boltzmann models for flow in porous media applications. Computers and Mathematics with Applications, 65(6):882–890, March 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112006839.

Yang:2011:CCF

[YBC11] Seung-Kab Yang, Jong-Sook Bae, and Seong-Hoon Cho. Coincidence and common fixed and periodic point theorems in cone metric spaces. Computers and Mathematics with Applications, 61(2):170–177, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008217.

Yan:2010:BNM

[YC10a] Lisa Jing Yan and Nick Cercone. Bayesian network modeling for evolutionary genetic structures. Computers and Mathematics with Applications, 59(8):2541–2551, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000258.

Yao:2010:SLV

[YC10b] Huanmin Yao and Minggen Cui. Searching the least value method for solving fourth-order nonlinear boundary value problems. Computers and Mathematics with Applications, 59(2):677–683, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007251.

Yang:2011:COI

[YC11] Gang Yang and Ruyun Chen. Choice of an optimal initial solution for a wave equation in the variational iteration method. Computers and Mathematics with Applications, 61(8):2053–2057, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006528.

Yang:2012:MVB

[YC12] Yang Yang and Qi-Xin Cao. Monocular vision based 6D object localization for service robot’s intelligent grasping. Computers and Mathematics with Applications, 64(5):1235–1241, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002726.

Yakar:2012:PSB

[YÇG12] Coskun Yakar, Muhammed Çiçek, and M. Bayram Gücen. Practical stability, boundedness criteria and Lagrange stability of fuzzy differential systems. Computers and Mathematics with Applications, 64(6):2118–2127, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003483.

Yang:2018:SCM

[YCHW18] Yin Yang, Yanping Chen, Yunqing Huang, and Huayi Wei. Spectral collocation method for the time-fractional diffusionwave equation and convergence analysis. Computers and Mathematics with Applications, 73(6):1218–1232, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304710.

Yao:2015:LMA

[YCLY15] Guangming Yao, C. S.

Chen, Wen Li, and D. L. Young. The localized method of approximated particular solutions for near-singular two- and three-dimensional problems. Computers and Mathematics with Applications, 70(12):2883– 2894, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211500468X.

Yuan:2019:DBD

[YCS19] Xiaolei Yuan, Zhenhua Chai, and Baochang Shi. Dynamic behavior of droplet through a confining orifice:a lattice Boltzmann study. Computers and Mathematics with Applications, 77(10):2640–2658, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119300033.

Yu:2014:GAL

[YCW^{+}14] Huidan (Whitney) Yu, Rou Chen, Hengjie Wang, Zhi Yuan, Ye Zhao, Yiran An, Yousheng Xu, and Luoding Zhu. GPU accelerated lattice Boltzmann simulation for rotational turbulence. Computers and Mathematics with Applications, 67(2):445–451, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005816.

Yao:2012:TSL

[YD12] Guoyu Yao and Ruifeng Ding. Two-stage least squares based iterative identification algorithm for controlled autoregressive moving average (CARMA) systems. Computers and Mathematics with Applications, 63(5):975–984, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

                                        URL http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010455.

Yilmaz:2012:SHM

[YDK^{+}12] Onur Yilmaz, Sercan Demirci, Yagiz Kaymak, Serkan Ergun, and Ahmet Yildirim. Shortest hop multipath algorithm for wireless sensor networks. Computers and Mathematics with Applications, 63(1):48–59, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111009461.

Yousefi:2011:GEL

[YDL11] S. A. Yousefi, Mehdi Dehghan, and A. Lotfi. Generalized Euler–Lagrange equations for fractional variational problems with free boundary conditions. Computers and Mathematics with Applications, 62(3):987–995, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002276.

Yu:2015:PBP

[YDW15] Yanyan Yu, Weihua Deng, and Yujiang Wu. Positivity and boundedness preserving schemes for space-time fractional predator–prey reaction– diffusion model. Computers and Mathematics with Applications, 69(8):743– 759, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000917.

Ye:2015:RCS

[Ye15] Zhuan Ye. Regularity criteria and small data global existence to the generalized viscous Hallmagnetohydrodynamics. Computers and Mathematics with Applications, 70(8):2137–2154, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004071.

Ye:2016:RCB

[Ye16] Zhuan Ye. On the regularity criteria of the 2D Boussinesq equations with partial dissipation. Computers and Mathematics with Applications, 72(7):1880–1895, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122116304618.

Ye:2017:ECB

[Ye17a] Hongyu Ye. The existence and the concentration behavior of normalized solutions for the L^{2}-critical Schrödinger–Poisson system. Computers and Mathematics with Applications, 74(2):266–280, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302341.

Ye:2017:IRC

[Ye17b] Zhuan Ye. An improved regularity criterion for the 2D inviscid Boussinesq equations. Computers and Mathematics with Applications, 74 (12):3095–3098, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304820.

Ye:2019:BLP

[Ye19]  Xia Ye.  Boundary-

layer phenomena for the boundary layer for the heat-conductive incompressible viscous fluids in the half space. Computers and Mathematics with Applications, 78 (6):1889–1896, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301488.

Yang:2010:MST

[YF10] Fan Yang and Chu-Li Fu. The method of simplified Tikhonov regularization for dealing with the inverse timedependent heat source problem. Computers and Mathematics with Applications, 60(5):1228–1236, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004177.

Yang:2011:KCS

[YG11] Hai-Long Yang and Zhi-Lian Guo. Kernels and closures of soft set relations, and soft set relation mappings. Computers and Mathematics with Applications, 61(3):651–662, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110009168.

Yan:2017:SAS

[YG17] Shuling Yan and Shangjiang Guo. Stability analysis of a stage structure model with spatiotemporal delay effect. Computers and Mathematics with Applications, 73(2):310–326, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306617.

Yin:2011:TAF

[YGH11] Yunfei Yin, Guanghong Gong, and Liang Han. Two-aircraft formation flight simulation system based on four-tiered architecture. Computers and Mathematics with Applications, 61(8):2323– 2329, April 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110007819.

Yao:2011:MDR

[YGR11] Wen-Li Yao, Lu-Lu Gao, and Yong-Sheng Ren. Modeling for dynamics of rigid-body systems with friction by linear complementary problem (LCP). Computers and Mathematics with Applications, 61(8):2232–2236, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007108.

Yang:2016:RSD

[YGS^{+}16] Jin-Wei Yang, Yi-Tian Gao, Chuan-Qi Su, Chen Zhao, and Yu-Jie Feng. Rational solutions for a (2 + 1)-dimensional nonlinear model in water waves generated by the Jaulent–Miodek hierarchy. Computers and Mathematics with Applications, 72(11):2685– 2693, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116304771.

Yang:2017:ETW

[YGS17] Xiao-Jun Yang, Feng Gao, and H. M. Srivastava. Exact travelling wave solutions for the local fractional two-dimensional Burgerstype equations. Computers and Mathematics with Applications, 73(2):203–210, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306290.

Yin:2011:NSM

[YH11a] Na Yin and Xue Huang.

A note on “Singlemachine scheduling with general learning functions”. Computers and Mathematics with Applications, 61(9):2929–2930, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002318.

See [Wan08].

Yuan:2011:DEI

[YH11b] Pingzhi Yuan and Yongzhong Hu. On the Diophantine equation x^{2} − kxy + y^{2} + lx = 0, l ∈ {1, 2, 4}. Computers and Mathematics with Applications, 61(3):573–577, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009077.

Yang:2012:CEC

[YH12] Junmin Yang and Maoan Han. Computation of expansion coefficients of Melnikov functions near a nilpotent center. Computers and Mathematics with Applications, 64 (6):1957–1974, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200257X.

Yang:2015:MFE

[YH15] Yidu Yang and Jiayu Han. Multilevel finite element discretizations based on local defect correction for nonsymmetric eigenvalue problems. Computers and Mathematics with Applications, 70(8):1799–1816, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115003788.

Yang:2019:SCS

[YH19] Jinjin Yang and Yinnian He. Stability and convergence of semi-implicit time-stepping algorithm for stationary incompressible magnetohydrodynamics. Computers and Mathematics with Applications, 77(5):1376–1395, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306679.

Yue:2018:SBL

[YHC18] Yunfei Yue, Lili Huang, and Yong Chen. Nsolitons, breathers, lumps and rogue wave solutions to a (3 + 1)-dimensional nonlinear evolution equation. Computers and Mathematics with Applications, 75(7):2538–2548, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730785X.

Yang:2011:NCI

[YHZY11] Zhichun Yang, Tingwen Huang, Linhua Zhang, and Zhiguo Yang. On networked control of impulsive hybrid systems. Computers and Mathematics with Applications, 61(8):2076–2080, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006619.

Yilmaz:2010:SBA

[Yil10] Yilmaz Yilmaz. Schauder bases and approximation property in fuzzy normed spaces. Computers and Mathematics with Applications, 59(6):1957–1964, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109007445.

Yildiz:2019:OCP

[Yil19] Tugba Akman Yildiz. Optimal control problem of the two-dimensional modified anomalous subdiffusion equation with discontinuous Galerkin approximation. Computers and Mathematics with Applications, 78 (6):2127–2146, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119302858.

Yang:2019:NAT

[YJ19] Xiu Yang and Xiaoyun Jiang. Numerical algorithm for two dimensional fractional Stokes’ first problem for a heated generalized second grade fluid with smooth and non-smooth solution. Computers and Mathematics with Applications, 78 (5):1562–1571, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301543.

Yue:2018:FWP

[YjH18] Jia Yue and Nan jing Huang. Fractional Wishart processes and ϵ-fractional Wishart processes with applications. Computers and Mathematics with Applications, 75(8):2955–2977, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300403.

Yin:2011:VSH

[YJZ11] Yunqiang Yin, Young Bae Jun, and Jianming Zhan. Vague soft hemirings. Computers and Mathematics with Applications, 62(1):199–213, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111003701.

Yuzbasi:2017:GLS

[YK17] Suayip Yüzbasi and Murat Karaçayir. A Galerkinlike scheme to solve twodimensional telegraph equation using collocation points in initial and boundary conditions. Computers and Mathematics with Applications, 74(12):3242–3249, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305126.

Yaman:2018:NIE

[YK18] Olha Ivanyshyn Yaman and Rainer Kress. Nonlinear integral equations for Bernoulli’s free boundary value problem in three dimensions. Computers and Mathematics with Applications, 74 (11):2784–2791, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303590.

Yousuf:2018:SCP

[YKA18] M. Yousuf, A. Q. M. Khaliq, and Salah Alrabeei. Solving complex PIDE systems for pricing American option under multistate regime switching jump-diffusion model. Computers and Mathematics with Applications, 75(8):2989–3001, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300427.

Yao:2011:LAM

[YKC11] Guangming Yao, Joseph Kolibal, and C. S. Chen. A localized approach for the method of approximate particular solutions. Computers and Mathematics with Applications, 61(9):2376–2387, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000939.

Yamak:2011:SH

[YKD11] S. Yamak, O. Kazanci, and B. Davvaz. Soft hyperstructure. Computers and Mathematics with Applications, 62(2):797– 803, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004779.

Yiotis:2010:ALB

[YKKS10] A. G. Yiotis, M. E. Kainourgiakis, E. S. Kikkinides, and A. K. Stubos. Application of the lattice-Boltzmann method to the modeling of population blob dynamics in 2D porous domains. Computers and Mathematics with Applications, 59(7):2315–2325, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000453.

Yero:2011:MDC

[YKRV11] I. G. Yero, D. Kuziak, and J. A. Rodríguez-Velázquez. On the metric dimension of corona product graphs. Computers and Mathematics with Applications, 61(9):2793– 2798, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002094.

Yang:2010:SSC

[YL10a] Ding-Gong Yang and Jin-Lin Liu. Some sufficient conditions for pvalent strongly starlike functions. Computers and Mathematics with Applications, 59(6):2018–2025, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007512.

Yin:2010:PDS

[YL10b] Xiaoyan Yin and Sanyang Liu. Positive definite solutions of the matrix equations X ± A^{∗}X^{−q}A = Q(q ≥ 1). Computers and Mathematics with Applications, 59(12):3727– 3739, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000266X.

Yang:2013:DDP

[YL13] Wensheng Yang and Yongqing Li. Dynamics of a diffusive predator– prey model with modified Leslie–Gower and Holling-type III schemes. Computers and Mathematics with Applications, 65(11):1727–1737, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001910.

Yin:2014:PAN

[YL14] Fuqi Yin and Linfang Liu. D-pullback attractor for a non-autonomous wave equation with additive noise on unbounded domains. Computers and Mathematics with Applications, 68(3):424–438, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400279X.

Yang:2016:MCS

[YL16] Liu Yang and Zhisu Liu. Multiplicity and concentration of solutions for fractional Schrödinger equation with sublinear perturbation and steep potential well. Computers and Mathematics with Applications, 72 (6):1629–1640, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304382.

Yang:2018:ITM

[YL18a] X. L. Yang and Y. Liu. An improved k–ω–ϕ– α turbulence model applied to near-wall, separated and impinging jet flows and heat transfer. Computers and Mathematics with Applications, 76(2):315–339, July 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302189.

Yin:2018:FDR

[YL18b] Fuqi Yin and Xueli Li. Fractal dimensions of random attractors for stochastic Benjamin– Bona–Mahony equation on unbounded domains. Computers and Mathematics with Applications, 75(5):1595–1615, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307514.

Yang:2016:LBP

[YLB16] Yidu Yang, Hao Li, and Hai Bi. The lower bound property of the Morley element eigenvalues. Computers and Mathematics with Applications, 72(4):904–920, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116303364.

Yen:2012:ANN

[YLC12] Hui-Min Yen, Tzuu-Hseng S. Li, and Yeong-Chan Chang. Adaptive neural network based tracking control for electrically driven flexiblejoint robots without velocity measurements. Computers and Mathematics with Applications, 64(5):1022–1032, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112002258.

Yang:2016:PGS

[YLC16] Bo Yang, Hui Liu, and Zhangxin Chen. Preconditioned GMRES solver on multiple-GPU architecture. Computers and Mathematics with Applications, 72(4):1076–1095, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303595.

Yang:2018:LSD

[YLC18] Xu Yang, Yingjie Liang, and Wen Chen. A local structural derivative PDE model for ultraslow creep. Computers and Mathematics with Applications, 76(7):1713–1718, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303882.

Yu:2011:PPI

[YLDL11] He Yu, Shouju Li, Hongxia Duan, and Yingxi Liu. A procedure of parameter inversion for a nonlinear constitutive model of soils with shield tunneling. Computers and Mathematics with Applications, 61(8):2005–2009, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006255.

Yan:2019:LSI

[YLF19] Weiwei Yan, Yang Liu, and Bingmei Fu. LBM simulations on the influence of endothelial SGL structure on cell adhesion in the micro-vessels. Computers and Mathematics with Applications, 78(4):1182–1193, August 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303959.

Yin:2010:GEW

[YLG10] Zheng Yin, Shaoyong Lai, and Yunxi Guo. Global existence of weak solutions for a shallow water equation. Computers and Mathematics with Applications, 60(9):2645–2652, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006796.

Yin:2017:BCA

[YLG17] Jinyan Yin, Yangrong Li, and Anhui Gu. Backwards compact attractors and periodic attractors for non-autonomous damped wave equations on an unbounded domain. Computers and Mathematics with Applications, 74(4):744–758, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303048.

Yau:2012:MPP

[YLH12] Her-Terng Yau, Qin-Cheng Liang, and Chin-Tsung Hsieh. Maximum power point tracking and optimal Li-ion battery charging control for photovoltaic charging system. Computers and Mathematics with Applications, 64(5):822–832, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011047.

Yin:2012:ASI

[YLJ12] Yunqiang Yin, Hongjie Li, and Young Bae Jun. On algebraic structure of intuitionistic fuzzy soft sets. Computers and Mathematics with Applications, 64(9):2896–2911, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003859.

Yao:2010:ACE

[YLK10] Yonghong Yao, Yeong-Cheng Liou, and Shin Min Kang. Approach to common elements of variational inequality problems and fixed point problems via a relaxed extragradient method. Computers and Mathematics with Applications, 59(11):3472–3480, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110002257.

Yong:2014:DSI

[YLL^{+}14] Yumei Yong, Xiaojun Lou, Sha Li, Chao Yang, and Xiaolong Yin. Direct simulation of the influence of the pore structure on the diffusion process in porous media. Computers and Mathematics with Applications, 67(2):412–423, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113005397.

Yang:2014:MME

[YLLK14] Jaw-Yen Yang, Chao-An Lin, Li-Shi Luo, and Manfred Krafczyk. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 67(2):237–238, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122113006937.

Yue:2016:PSC

[YLLN16] J. H. Yue, M. Li, G. R.

Liu, and R. P. Niu. Proofs of the stability and convergence of a weakened weak method using PIM shape functions. Computers and Mathematics with Applications, 72(4):933–951, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303340.

Yang:2012:PLD

[YLS12] Yang Yang, Remigijus Leipus, and Jonas Siaulys. Precise large deviations for compound random sums in the presence of dependence structures. Computers and Mathematics with Applications, 64(6):2074–2083, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003392.

Yang:2009:CIV

[YLY^{+}09] Xibei Yang, Tsau Young Lin, Jingyu Yang, Yan Li, and Dongjun Yu.

Combination of intervalvalued fuzzy set and soft set. Computers and Mathematics with Applications, 58(3):521–527, August 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003228.

See note [JY11].

Yan:2012:LBM

[YLY12] Guangwu Yan, Tingting Li, and Xianli Yin. Lattice Boltzmann model for elastic thin plate with small deflection. Computers and Mathematics with Applications, 63(8):1305–1318, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000259.

Yang:2017:DTP

[YLZ17] Ruizhi Yang, Ming Liu, and Chunrui Zhang. A diffusive toxin producing phytoplankton model with maturation delay and three-dimensional patch. Computers and Mathematics with Applications, 73(5):824–837, March 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300263.

Yuan:2013:LES

[YM13] Hongjun Yuan and Qiu Meng. Local existence of strong solution for a class of compressible non-Newtonian fluids with non-Newtonian potential. Computers and Mathematics with Applications, 65(4):563–575, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006281.

Yang:2017:ALT

[YM17] J. Y. Yang and W. X.

Ma. Abundant lump-type solutions of the Jimbo– Miwa equation in (3 + 1)-dimensions. Computers and Mathematics with Applications, 73(2):220–225, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306241.

Yildirim:2010:ASP

[YMDZ10] A. Yildirim, S. T. Mohyud-Din, and D. H. Zhang. Analytical solutions to the pulsed Klein–Gordon equation using modified variational iteration method (MVIM) and Boubaker polynomials expansion scheme (BPES). Computers and Mathematics with Applications, 59(8):2473–2477, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007937.

Yong:2018:LSK

[YMHL18] Xuelin Yong, Wen-Xiu Ma, Yehui Huang, and Yong Liu. Lump solutions to the Kadomtsev– Petviashvili I equation with a self-consistent source. Computers and Mathematics with Applications, 75(9):3414–3419, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300804.

Yin:2018:DES

[YMLL18] Yu-Hang Yin, Wen-Xiu Ma, Jian-Guo Liu, and Xing Lü. Diversity of exact solutions to a (3 + 1)-dimensional nonlinear evolution equation and its reduction. Computers and Mathematics with Applications, 76 (6):1275–1283, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303444.

Yahya:2012:SAS

[YMM12] Samer Yahya, M. Moghavvemi, and Haider A. F. Mohamed. Singularity avoidance of a six degree of freedom three dimensional redundant planar manipulator. Computers and Mathematics with Applications, 64(5):856–868, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011448.

Yang:2011:CBO

[YMSL11] Xiao-Hua Yang, Ying Mei, Dun-Xian She, and Jian-Qiang Li. Chaotic Bayesian optimal prediction method and its application in hydrological time series. Computers and Mathematics with Applications, 61(8):1975– 1978, April 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006152.

Nie:2011:DDT

[yN11] Pu yan Nie. Dynamic discrete-time multi-leader-follower games with leaders in turn. Computers and Mathematics with Applications, 61(8):2039– 2043, April 2011. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006474.

You:2016:GSS

[YN16] Shujun You and Xiaoqi Ning. On global smooth solution for generalized Zakharov equations. Computers and Mathematics with Applications, 72(1):64–75, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302309.

Yao:2010:SNA

[YNLK10] Yonghong Yao, Muhammad Aslam Noor, Yeong-Cheng Liou, and Shin Min Kang. Some new algorithms for solving mixed equilibrium problems. Computers and Mathematics with Applications, 60(5):1351–1359, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000430X.

Yuan:2014:MEB

[YNS^{+}14] Hai-Zhuan Yuan, Xiao-Dong Niu, Shi Shu, Mingjun Li, and Hiroshi Yamaguchi. A momentum exchange-based immersed boundary-lattice Boltzmann method for simulating a flexible filament in an incompressible flow. Computers and Mathematics with Applications, 67(5):1039–1056, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000182.

Yasar:2010:ISC

[YÖ10] Emrullah Yasar and Teoman Özer. Invariant solutions and conservation laws to nonconservative FP equation. Computers and Mathematics with Applications, 59(9):3203– 3210, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001665.

Yoon:2017:BDF

[Yoo17] Gil Ho Yoon. Brittle and ductile failure constraints of stress-based topology optimization method for fluid-structure interactions. Computers and Mathematics with Applications, 74(3):398–419, August 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302468.

Youssef:2011:CQK

[You11] Amr M. Youssef. Cryptanalysis of a quadratic knapsack cryptosystem. Computers and Mathematics with Applications, 61(4):1261–1265, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000113.

Yildirim:2010:AEF

[YP10] Ahmet Yildirim and Zehra Pinar. Application of the exp-function method for solving nonlinear reaction–diffusion equations arising in mathematical biology. Computers and Mathematics with Applications, 60(7):1873–1880, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004967.

Yuan:2018:GRM

[YQ18] Baoquan Yuan and Yuanyuan Qiao. Global regularity for the 2D magnetomicropolar equations with partial and fractional dissipation. Computers and Mathematics with Applications, 76 (10):2345–2359, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304498.

Yu:2016:HAR

[YqS16] Juan Yu and Shu qian Shen. The Hermitian {P, k + 1}-(anti- )reflexive solutions of a linear matrix equation. Computers and Mathematics with Applications, 71(12):2513–2523, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301626.

Yang:2019:SCM

[YQWZ19] Yin Yang, Wanying Qiao, Jindi Wang, and Shangyou Zhang. Spectral collocation methods for nonlinear coupled time fractional Nernst– Planck equations in two dimensions and its convergence analysis. Computers and Mathematics with Applications, 78 (5):1431–1449, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118307107.

Yosibash:2018:H

[YRDR18] Zohar Yosibash, Ernst Rank, Alexander Düster, and Alessandro Reali. HOFEIM 2016. Computers and Mathematics with Applications, 74(7):1529–1530, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304315.

Su:2010:PMC

[yS10] Li yun Su. Prediction of multivariate chaotic time series with local polynomial fitting. Computers and Mathematics with Applications, 59(2):737–744, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007184.

Yildiz:2012:CCA

[YS12] Bahattin Yildiz and Irfan Siap. Cyclic codes over F_{2}[u]/(u^{4} − 1) and applications to DNA codes. Computers and Mathematics with Applications, 63(7):1169–1176, April 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010844.

Yu:2016:MMS

[YS16a] Jianping Yu and Yongli Sun. Modified method of simplest equation and its applications to the Bogoyavlenskii equation. Computers and Mathematics with Applications, 72(7):1943–1955, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304564.

Yu:2016:RST

[YS16b] Jianping Yu and Yongli Sun. Rational solutions to two new KP-like equations. Computers and Mathematics with Applications, 72(6):1556–1572, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304035.

Yu:2017:NGS

[YS17] Jianping Yu and Yongli Sun. A note on the gaussons of some new logarithmic evolution equations. Computers and Mathematics with Applications, 74(2):258–265, July 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117302353.

Yang:2019:USE

[YS19a] Yang Yang and Zhiyong Si. Unconditional stability and error estimates of the modified characteristics FEMs for the time-dependent incompressible MHD equations. Computers and Mathematics with Applications, 77(1):263–283, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305431.

Yun:2019:EGM

[YS19b] Gun Jin Yun and Shen Shang. On the existence of a global minimum in inverse parameters identification by self-optimizing inverse analysis method. Computers and Mathematics with Applications, 77(3):803–814, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306163.

Yucel:2015:DGM

[YSB15] Hamdullah Yücel, Martin Stoll, and Peter Benner. A discontinuous Galerkin method for optimal control problems governed by a system of convection– diffusion PDEs with nonlinear reaction terms. Computers and Mathematics with Applications, 70(10):2414–2431, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004204.

Shi:2010:VFC

[ySGL^{+}10] Xiao yan Shi, Hui Gao, Volha I. Lazouskaya, Qinjun Kang, Yan Jin, and Lian-Ping Wang. Viscous flow and colloid transport near air-water interface in a microchannel. Computers and Mathematics with Applications, 59(7):2290–2304, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006506.

Yuzbasi:2011:BPS

[YSS11a] Suayip Yüzbasi, Niyazi Sahin, and Mehmet Sezer. Bessel polynomial solutions of highorder linear Volterra integro-differential equations. Computers and Mathematics with Applications, 62(4):1940–1956, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122111005189.

Yuzbasi:2011:NSS

[YSS11b] Suayip Yüzbasi, Niyazi Sahin, and Mehmet Sezer. Numerical solutions of systems of linear Fredholm integro-differential equations with Bessel polynomial bases. Computers and Mathematics with Applications, 61(10): 3079–3096, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211100277X.

Shi:2010:TLO

[ySW10] Dong yang Shi and Xiao-Ling Wang. Two low order characteristic finite element methods for a convection-dominated transport problem. Computers and Mathematics with Applications, 59(12): 3630–3639, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110001677.

Yang:2016:ELB

[YSW16] L. M. Yang, C. Shu, and J. Wu. Extension of lattice Boltzmann flux solver for simulation of 3D viscous compressible flows. Computers and Mathematics with Applications, 71(10):2069– 2081, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116301687.

Yue:2019:PTM

[YSX^{+}19] Xiaoqiang Yue, Shi Shu, Xiaowen Xu, Weiping Bu, and Kejia Pan. Parallel-in-time multigrid for space–time finite element approximations of two-dimensional space-fractional diffusion equations. Computers and Mathematics with Applications, 78 (11):3471–3484, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302809.

Yin:2011:TWA

[YT11] Jiuli Yin and Lixin Tian. Towered waves and antiwaves in the generalized Degasperis–Procesi equation. Computers and Mathematics with Applications, 62(2):657–667, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004457.

Yazlik:2012:NGH

[YT12] Yasin Yazlik and Necati Taskara. A note on generalized k-Horadam sequence. Computers and Mathematics with Applications, 63(1):36–41, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100931X.

Yu:2013:CSV

[YT13] P. X. Yu and Zhen F.

Tian. A compact streamfunction-velocity scheme on nonuniform grids for the 2D steady incompressible Navier– Stokes equations. Computers and Mathematics with Applications, 66(7):1192–1212, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004410.

Yacouba:2018:NCN

[YT18a] Simpore Yacouba and Antoine Tambue. Null controllability and numerical method for Crocco equation with incomplete data based on an exponential integrator and finite difference-finite element method. Computers and Mathematics with Applications, 74 (5):1043–1058, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301694.

Yu:2018:UCD

[YT18b] P. X. Yu and Zhen F. Tian. An upwind compact difference scheme for solving the streamfunctionvelocity formulation of the unsteady incompressible Navier–Stokes equation. Computers and Mathematics with Applications, 75(9):3224–3243, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300592.

Yin:2018:NSV

[YTC^{+}18] Hui-Min Yin, Bo Tian, Jun Chai, Lei Liu, and Yan Sun. Numerical solutions of a variable-coefficient nonlinear Schrödinger equation for an inhomogeneous optical fiber. Computers and Mathematics with Applications, 76(8):1827–1836, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303493.

Yan:2018:CSW

[YTD^{+}18] Xue-Wei Yan, Shou-Fu Tian, Min-Jie Dong, Li Zhou, and Tian-Tian Zhang. Characteristics of solitary wave, homoclinic breather wave and rogue wave solutions in a (2 + 1)-dimensional generalized breaking soliton equation. Computers and Mathematics with Applications, 76(1):179–186, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302116.

Yin:2010:SNS

[YTF10] Jiuli Yin, Lixin Tian, and Xinghua Fan. Symmetric and non-symmetric waves in the osmosis k(2, 2) equation. Computers and Mathematics with Applications, 59(8):2756–2762, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000799.

Yuan:2018:SRS

[YTL^{+}18] Yu-Qiang Yuan, Bo Tian, Lei Liu, Han-Peng Chai, and Yan Sun. Semirational solutions for the (3 + 1)-dimensional Kadomtsev–Petviashvili equation in a plasma or fluid. Computers and Mathematics with Applications, 76(11– 12):2566–2574, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304851.

Yuan:2017:WGS

[YTS^{+}17] Yu-Qiang Yuan, Bo Tian, Wen-Rong Sun, Jun Chai, and Lei Liu. Wronskian and Grammian solutions for a (2 + 1)- dimensional Date–Jimbo– Kashiwara–Miwa equation. Computers and Mathematics with Applications, 74(4):873–879, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303565.

Yu:2017:RHO

[YTZ17] P. X. Yu, Z. F. Tian, and Hongjie Zhang. A rational high-order compact difference method for the steady-state stream function-vorticity formulation of the Navier– Stokes equations. Computers and Mathematics with Applications, 73(7):1461–1484, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300548.

Yu:2011:GFK

[Yu11a] Fajun Yu. A generalized fractional KN equation hierarchy and its fractional Hamiltonian structure. Computers and Mathematics with Applications, 62(3):1522–1530, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003464.

Yu:2011:ICG

[Yu11b]  Fajun Yu.  An inte-

grable couplings of G-WKI equations hierarchy with self-consistent sources. Computers and Mathematics with Applications, 61(8):2085–2089, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006632.

Yu:2017:BRT

[Yu17] Liju Yu. Blowup result for a type of generalized Zakharov system. Computers and Mathematics with Applications, 74(6):1406–1413, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303802.

Yuan:2018:IVM

[Yua18] Jianjun Yuan. An improved variational model for denoising magnetic resonance images. Computers and Mathematics with Applications, 76 (9):2212–2222, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304413.

Yun:2013:VUM

[Yun13] Jae Heon Yun. Variants of the Uzawa method for saddle point problem. Computers and Mathematics with Applications, 65(7):1037–1046, April 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000679.

Yusufoglu:2009:IHP

[Yus09] Elçin Yusufoğlu. An improvement to homotopy perturbation method for solving system of linear equations. Computers and Mathematics with Applications, 58 (11–12):2231–2235, December 2009. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109001497.

See notes [BG11].

Yuzbasi:2011:NAS

[Yüz11] Suayip Yüzbasi. A numerical approach for solving the high-order linear singular differentialdifference equations. Computers and Mathematics with Applications, 62 (5):2289–2303, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005700.

Yuzbasi:2012:EAS

[Yüz12a]  Suayip Yüzbasi.  An

efficient algorithm for solving multi-pantograph equation systems. Computers and Mathematics with Applications, 64(4): 589–603, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111011333.

Yuzbasi:2012:NAB

[Yüz12b] Suayip Yüzbasi. A numerical approximation based on the Bessel functions of first kind for solutions of Riccati type differential– difference equations. Computers and Mathematics with Applications, 64 (6):1691–1705, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000363.

Yuzbasi:2012:SSL

[Yüz12c]  Suayip Yüzbasi.  On

the solutions of a system of linear retarded and advanced differential equations by the Bessel collocation approximation. Computers and Mathematics with Applications, 63(10):1442– 1455, May 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112002581.

Yang:2010:TWG

[YW10] Xiping Yang and Yifu Wang. Travelling wave and global attractivity in a competition-diffusion system with nonlocal delays. Computers and Mathematics with Applications, 59(10):3338– 3350, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001963.

Yu:2011:GTI

[YW11a] Y. J. Yu and Z. H.

Wang. A graphical test for the interval stability of fractional-delay systems. Computers and Mathematics with Applications, 62(3):1501–1509, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002306.

Yuan:2011:BPB

[YW11b] Zhangxian Yuan and Xinwei Wang. Buckling and post-buckling analysis of extensible beam-columns by using the differential quadrature method. Computers and Mathematics with Applications, 62(12):4499–4513, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008935.

Yu:2014:PMW

[YW14] Chenhan D. Yu and Weichung Wang. Performance models and workload distribution algorithms for optimizing a hybrid CPU– GPU multifrontal solver. Computers and Mathematics with Applications, 67(7):1421–1437, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400025X.

Yin:2019:ECG

[YW19] Li-Feng Yin and Xing-Ping Wu. Existence and concentration of ground state solutions for critical Schrödinger equation with steep potential well. Computers and Mathematics with Applications, 78(12):3862–3871, December 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303190.

Yosibash:2014:HOF

[YWH14] Zohar Yosibash, Danny Weiss, and Stefan Hartmann. High-order FEMs for thermo-hyperelasticity at finite strains. Computers and Mathematics with Applications, 67(3):477–496, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006573.

Yau:2011:NAC

[YWHC11] Her-Terng Yau, Cheng-Chi Wang, Chin-Tsung Hsieh, and Ching-Chang Cho. Nonlinear analysis and control of the uncertain micro-electro-mechanical system by using a fuzzy sliding mode control design. Computers and Mathematics with Applications, 61(8):1912–1916, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004955.

Yao:2010:IDA

[YWK^{+}10] Lin Yao, Lei Wang, Xiangwei Kong, Guowei Wu, and Feng Xia. An inter-domain authentication scheme for pervasive computing environment. Computers and Mathematics with Applications, 60(2):234–244, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000210.

Yang:2011:HMT

[YWL^{+}11a] Jun-Yuan Yang, Xiao-Yan Wang, Xue-Zhi Li, Feng-Qin Zhang, and Souvik Bhattacharya. An HIV model: Theoretical analysis and experimental verification. Computers and Mathematics with Applications, 61(8):2172–2176, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006930.

Yue:2011:IFS

[YWL^{+}11b] Shigang Yue, Hua-Liang Wei, Maozhen Li, Qilian Liang, and Lipo Wang. ICNC–FSKD 2010 Special Issue on Computers & Mathematics in Natural Computation and Knowledge Discovery. Computers and Mathematics with Applications, 62(7):2683–2684, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005293.

Yuan:2017:SUM

[YWL17] Hailong Yuan, Jianhua Wu, and Yanling Li. Some uniqueness, multiplicity and complete dynamics for a cooperative model. Computers and Mathematics with Applications, 73(12):2587–2602, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302110.

Yu:2018:GRL

[YWT18] Yanghai Yu, Xing Wu, and Yanbin Tang. Global regularity of the 2D liquid crystal equations with weak velocity dissipation. Computers and Mathematics with Applications, 74(5):920–933, September 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306253.

Yuan:2012:JFH

[YWW^{+}12] Yulai Yuan, Yongwei Wu, Qiuping Wang, Guangwen Yang, and Weimin Zheng. Job failures in high performance computing systems: a largescale empirical study. Computers and Mathematics with Applications, 63(2):365–377, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111005980.

Yang:2011:OCS

[YX11a] Qiaoshun Yang and Zhiting Xu. Oscillation criteria for second order quasilinear neutral delay differential equations on time scales. Computers and Mathematics with Applications, 62 (10):3682–3691, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007760.

Yang:2011:CVI

[YX11b] Shui-Ping Yang and Ai-Guo Xiao. Convergence of the variational iteration method for solving multidelay differential equations. Computers and Mathematics with Applications, 61(8):2148–2151, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000684X.

Yin:2011:SSM

[YX11c] Yunqiang Yin and Dehua Xu. Some singlemachine scheduling problems with general effects of learning and deterioration. Computers and Mathematics with Applications, 61(1):100–108, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110008266.

Yang:2016:GSD

[YX16] Yu Yang and Yancong Xu. Global stability of a diffusive and delayed virus dynamics model with Beddington– DeAngelis incidence function and CTL immune response. Computers and Mathematics with Applications, 71(4):922–930, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116000122.

Yao:2013:PPM

[YXP^{+}13] Yonglei Yao, Naixue Xiong, Jong Hyuk Park, Li Ma, and Jingfa Liu. Privacy-preserving max/ min query in two-tiered wireless sensor networks. Computers and Mathematics with Applications, 65(9):1318–1325, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112001174.

Xi:2011:AAI

[yXpYxZT11] Guang yong Xi, Jian ping Yue, Bao xing Zhou, and Pu Tang. Application of an artificial immune algorithm on a statistical model of dam displacement. Computers and Mathematics with Applications, 62 (10):3980–3986, November 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008352.

Yang:2010:CVI

[YXS10] Shuiping Yang, Aiguo Xiao, and Hong Su. Convergence of the variational iteration method for solving multi-order fractional differential equations. Computers and Mathematics with Applications, 60(10):2871– 2879, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110007418.

Yin:2014:PAM

[YXWL14] Hongwei Yin, Xiaoyong Xiao, Xiaoqing Wen, and Kai Liu. Pattern analysis of a modified Leslie–Gower predator– prey model with Crowley– Martin functional response and diffusion.

Computers and Mathematics with Applications, 67(8):1607–1621, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211400087X.

Yang:2011:MCA

[YXX11] Zuyuan Yang, Yong Xiang, and Shengli Xie. Maximum contrast analysis for nonnegative blind source separation. Computers and Mathematics with Applications, 62 (11):3997–4006, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007589.

Yang:2010:IPG

[YXYH10] Laurence T. Yang, Li Xu, Sang-Soo Yeo, and Sajid Hussain. An integrated parallel GNFS algorithm for integer factorization based on Linbox Montgomery block Lanczos method over GF(2). Computers and Mathematics with Applications, 60(2):338–346, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000428.

Yan:2010:PAS

[YY10a] Weiping Yan and Jurang Yan. Periodicity and asymptotic stability of a predator–prey system with infinite delays. Computers and Mathematics with Applications, 60(5):1465–1472, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110004438.

Yang:2010:PST

[YY10b] Chen Yang and Jurang Yan. Positive solutions for third-order Sturm–Liouville boundary value problems with p-Laplacian. Computers and Mathematics with Applications, 59(6):2059– 2066, March 2010. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122109007585.

Yang:2010:MTC

[YY10c] Suh-Jenq Yang and Dar-Li Yang. Minimizing the total completion time in single-machine scheduling with aging/deteriorating effects and deteriorating maintenance activities. Computers and Mathematics with Applications, 60(7):2161–2169, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005663.

Yan:2011:SSL

[YY11a] Bo Yan and Guangwu Yan. A steady-state lattice Boltzmann model for incompressible flows. Computers and Mathematics with Applications, 61(5):1348–1354, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000058.

Yasar:2011:NNC

[YY11b] Elif Yasar and Sibel Yalçin. Neighborhoods of a new class of harmonic multivalent functions. Computers and Mathematics with Applications, 62(1):462–473, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004251.

Yang:2012:MDB

[YY12] Shangming Yang and Mao Ye. Multistability of α-divergence based NMF algorithms. Computers and Mathematics with Applications, 64(2):73– 88, July 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211200020X.

Yan:2014:EKT

[YY14] Liang Yan and Fenglian Yang. Efficient Kansatype MFS algorithm for time-fractional inverse diffusion problems. Computers and Mathematics with Applications, 67(8):1507–1520, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000790.

Yan:2015:MAP

[YY15] Liang Yan and Fenglian Yang. The method of approximate particular solutions for the time-fractional diffusion equation with a nonlocal boundary condition. Computers and Mathematics with Applications, 70(3):254–264, August 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115002114.

Yong:2011:SFF

[YYC11] Xuelin Yong, Yongping Yang, and Degang Chen. Soliton fission and fusion of a new two-component Korteweg–de Vries (KdV) equation. Computers and Mathematics with Applications, 62(4):1765–1771, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100486X.

Yadav:2016:NSU

[YYK16] Neha Yadav, Anupam Yadav, and Joong Hoon Kim. Numerical solution of unsteady advection dispersion equation arising in contaminant transport through porous media using neural networks. Computers and Mathematics with Applications, 72(4):1021–1030, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303467.

Yu:2011:SWC

[YYL11] Jian Yu, Miin-Shen Yang, and E. Stanley Lee. Sample-weighted clustering methods. Computers and Mathematics with Applications, 62 (5):2200–2208, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005591.

Yin:2019:DZB

[YYLW19] Xiaojun Yin, Liangui Yang, Quansheng Liu, and Guorong Wu. (2+1)- dimensional ZK-Burgers equation with the generalized beta effect and its exact solitary solution. Computers and Mathematics with Applications, 77(1):302–310, January 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118305455.

Yang:2013:ECS

[yYqWqZC13] Xiao yuan Yang, Li qiang Wu, Min qing Zhang, and Xiao-Feng Chen. An efficient CCA-secure cryptosystem over ideal lattices from identity-based encryption. Computers and Mathematics with Applications, 65(9):1254– 1263, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001447.

Yu:2013:NAC

[yYsZyYL13] Chun yan Yu, Wei shi Zhang, Ying ying Yu, and Ying Li. A novel active contour model for image segmentation using distance regularization term. Computers and Mathematics with Applications, 65(11):1746– 1759, July 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001879.

Yoshioka:2019:HJB

[YYYH19] Hidekazu Yoshioka, Yuta Yaegashi, Yumi Yoshioka, and Kunihiko Hamagami. Hamilton–Jacobi– Bellman Quasi-Variational Inequality arising in an environmental problem and its numerical discretization. Computers and Mathematics with Applications, 77(8):2182–2206, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306965.

Yang:2010:SFV

[YZ10a] Di Yang and Ke-Qin Zhu. Start-up flow of a viscoelastic fluid in a pipe with a fractional Maxwell’s model. Computers and Mathematics with Applications, 60(8):2231–2238, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005778.

Yin:2010:NTF

[YZ10b] Yunqiang Yin and Jianming Zhan. New types of fuzzy filters of BLalgebras. Computers and Mathematics with Applications, 60(7):2115–2125, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005468.

Yeh:2012:ECR

[YZ12] Cheng-Yu Yeh and Chang-Zhi Zhuo. An efficient complexity reduction algorithm for g.729 speech codec. Computers and Mathematics with Applications, 64(5):887–896, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000697.

Yao:2015:NSH

[YZ15] Qinghe Yao and Hongkun Zhu. Numerical simulation of hydrogen dispersion behaviour in a partially open space by a stabilized balancing domain decomposition method. Computers and Mathematics with Applications, 69(10):1068–1079, May 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000954.

Yang:2018:DDV

[YZ18] Yu Yang and Shengliang Zhang. Dynamics of a diffusive vaccination model with nonlinear incidence. Computers and Mathematics with Applications, 75(12):4355–4360, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301706.

Yuan:2019:GWP

[YZ19] Baoquan Yuan and Qiuyue Zhang. Global well-posedness of the dd Oldroyd-B type models with fractional Laplacian dissipation. Computers and Mathematics with Applications, 77(7):1933–1944, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306758.

Yu:2010:SPD

[YZAX10] Hengguo Yu, Shouming Zhong, Ravi P. Agarwal, and Lianglin Xiong. Species permanence and dynamical behavior analysis of an impulsively controlled ecological system with distributed time delay. Computers and Mathematics with Applications, 59(12):3824– 3835, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110002890.

Yin:2010:PSF

[YZGW10] Minghao Yin, Tingting Zou, Wenxiang Gu, and Jianan Wang. Product structure of the fuzzy n-ary factor polygroup. Computers and Mathematics with Applications, 59(8):2734–2742, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000635.

Zhong:2012:SSM

[yZjH12] Ren you Zhong and Nan jing Huang. On the stability of solution mapping for parametric generalized vector quasiequilibrium problems. Computers and Mathematics with Applications, 63(4):807–815, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010297.

Zhang:2010:GEU

[yZjM10] Zai yun Zhang and Xiu jin Miao. Global existence and uniform decay for wave equation with dissipative term and boundary damping. Computers and Mathematics with Applications, 59(2):1003–1018, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006725.

Yang:2019:TVH

[YZM^{+}19] Jing-Hua Yang, Xi-Le Zhao, Jin-Jin Mei, Si Wang, Tian-Hui Ma, and Ting-Zhu Huang. Total variation and high-order total variation adaptive model for restoring blurred images with Cauchy noise. Computers and Mathematics with Applications, 77(5):1255–1272, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306539.

Yousefi:2018:EPB

[YZMA18] Mostafa Yousefi, A. Zolfaghari, A. Minuchehr, and M. R. Abbassi. Even-parity Boltzmann transport equation applied for response (contributon) flux calculation based on the spatial channel theory. Computers and Mathematics with Applications, 75(12):4378–4396, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830172X.

Yang:2016:NFD

[YZMZ16] Yu Yang, Jinling Zhou, Xinsheng Ma, and Tonghua Zhang. Nonstandard finite difference scheme for a diffusive within-host virus dynamics model with both virus-to-cell and cell-to-cell transmissions. Computers and Mathematics with Applications, 72(4):1013–1020, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303479.

Yu:2018:GSS

[YZS18] Haibo Yu, Peixin Zhang, and Xiujuan Shi. Global strong solutions to the 3D incompressible MHD equations with densitydependent viscosity. Computers and Mathematics with Applications, 75(8):2825–2834, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300221.

Yang:2014:FMI

[YZWW14] Yunyun Yang, Yi Zhao, Boying Wu, and Hongpeng Wang. A fast multiphase image segmentation model for gray images. Computers and Mathematics with Applications, 67(8):1559–1581, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000418.

Yang:2018:OSC

[YZX18] Xuehua Yang, Haixiang Zhang, and Da Xu. Orthogonal spline collocation method for the fourth-order diffusion system. Computers and Mathematics with Applications, 75(9):3172–3185, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300555.

Yin:2010:FH

[YZXW10] Yunqiang Yin, Jianming Zhan, Dehua Xu, and Jiayin Wang. The L-fuzzy hypermodules.

Computers and Mathematics with Applications, 59(2):953–963, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109006762.

Yao:2010:APB

[YZY10] Meiping Yao, Aimin Zhao, and Jurang Yan. Anti-periodic boundary value problems of second order impulsive differential equations. Computers and Mathematics with Applications, 59(12): 3617–3629, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122110000556.

Zarebnia:2010:NSS

[ZA10] M. Zarebnia and M. G. Ali Abadi. Numerical solution of system of nonlinear second-order integro-differential equations. Computers and Mathematics with Applications, 60(3):591–601, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110003457.

Zedan:2015:NCS

[ZA15] Hassan Zedan and Mohamed M. Alshamrani. A novel class of solutions for the (2 + 1)-dimensional higher-order Broer–Kaup system. Computers and Mathematics with Applications, 69(2):67–80, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005227.

Zaky:2018:ITM

[Zak18a] Mahmoud A. Zaky. An improved tau method for the multi-dimensional fractional Rayleigh–Stokes problem for a heated generalized second grade fluid. Computers and Mathematics with Applications, 75(7):2243–2258, April 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307678.

Zaman:2018:NCT

[ZAK18b] A. Zaman, N. Ali, and Nabeela Kousar. Nanoparticles (Cu, TiO_{2}, Al_{2}O_{3}) analysis on unsteady blood flow through an artery with a combination of stenosis and aneurysm. Computers and Mathematics with Applications, 76 (9):2179–2191, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304371.

Zhang:2017:AMS

[ZaY17] Zujin Zhang and Zheng an Yao. 3D axisymmetric MHD system with regularity in the swirl component of the vorticity. Computers and Mathematics with Applications, 73(12):2573–2580, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730202X.

Zapletal:2019:SOS

[ZB19] Jan Zapletal and Jirí Bouchala. Shape optimization and subdivision surface based approach to solving 3D Bernoulli problems. Computers and Mathematics with Applications, 78(9):2911–2932, November 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300835.

Zhang:2011:ERC

[ZBF11] Yinghan Zhang, Zhanbing Bai, and Tingting Feng. Existence results for a coupled system of nonlinear fractional three-point boundary value problems at resonance. Computers and Mathematics with Applications, 61(4):1032–1047, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009582.

Zhang:2019:NPW

[ZBFC19] Runfa Zhang, Sudao Bilige, Tao Fang, and Temuer Chaolu. New periodic wave, cross-kink wave and the interaction phenomenon for the Jimbo–Miwa-like equation. Computers and Mathematics with Applications, 78(3):754–764, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301117.

Zhu:2012:NES

[ZBL12] Song-Ping Zhu, Alexander Badran, and Xiaoping Lu. A new exact solution for pricing European options in a two-state regime-switching economy. Computers and Mathematics with Applications, 64(8):2744–2755, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112005378.

Zboinski:2019:BHM

[Zbo19] G. Zboiński. 3D-based hierarchical models and hpq-approximations for adaptive finite element method of Laplace problems as exemplified by linear dielectricity. Computers and Mathematics with Applications, 78(8):2468–2511, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301865.

Zhai:2010:FPT

[ZC10] Cheng-Bo Zhai and Xiao-Min Cao. Fixed point theorems for τ–ϕ-concave operators and applications. Computers and Mathematics with Applications, 59(1):532–538, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003812.

Zheng:2011:ESF

[ZC11a] Mianbin Zheng and Guohua Chen. An efficient 3D stochastic finite element method for failure probability analysis of high temperature components. Computers and Mathematics with Applications, 62(1):333– 341, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004020.

Zheng:2011:MCC

[ZC11b] Xiaoping Zheng and Yuan Cheng. Modeling cooperative and competitive behaviors in emergency evacuation: a game-theoretical approach. Computers and Mathematics with Applications, 62 (12):4627–4634, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009126.

Zhu:2011:PCS

[ZC11c] Song-Ping Zhu and Wen-Ting Chen. A predictor– corrector scheme based on the ADI method for pricing American puts with stochastic volatility. Computers and Mathematics with Applications, 62(1):1–26, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002811.

Zhang:2016:MSM

[ZC16] Shengliang Zhang and Siyan Chen. A meshless symplectic method for two-dimensional Schrödinger equation with radial basis functions. Computers and Mathematics with Applications, 72(9):2143–2150, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304680.

Zhuang:2017:LGS

[ZC17] Qingqu Zhuang and Lizhen Chen. Legendre– Galerkin spectral-element method for the biharmonic equations and its applications. Computers and Mathematics with Applications, 74 (12):2958–2968, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304686.

Zhao:2012:EPS

[ZCH12] Yulin Zhao, Haibo Chen, and Li Huang. Existence of positive solutions for nonlinear fractional functional differential equation. Computers and Mathematics with Applications, 64 (10):3456–3467, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112001137.

Zhang:2014:TFH

[ZCH14] Wei Zhang, Xing Cai, and Sverre Holm. Timefractional heat equations and negative absolute temperatures. Computers and Mathematics with Applications, 67(1):164– 171, January 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006639.

Zhao:2018:LRB

[ZcHS18] Wei Zhao, Yiu chung Hon, and Martin Stoll. Localized radial basis functions-based pseudospectral method (LRBF– PSM) for nonlocal diffusion problems. Computers and Mathematics with Applications, 75(5):1685–1704, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307563.

Zhang:2019:FDS

[ZCLW19] Jun Zhang, Hu Chen, Shimin Lin, and Jinrong Wang. Finite difference/ spectral approximation for a time-space fractional equation on two and three space dimensions. Computers and Mathematics with Applications, 78 (6):1937–1946, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301609.

Zhang:2013:LBS

[ZCSG13] Wenhuan Zhang, Zhenhua Chai, Baochang Shi, and Zhaoli Guo. Lattice Boltzmann study of flow and mixing characteristics of twodimensional confined impinging streams with uniform and non-uniform inlet jets. Computers and Mathematics with Applications, 65(4):638–647, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112006542.

Zhang:2018:RWP

[ZCT18] Xiaoen Zhang, Yong Chen, and Xiaoyan Tang. Rogue wave and a pair of resonance stripe solitons to KP equation. Computers and Mathematics with Applications, 76(8):1938–1949, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304115.

Zeng:2015:EEW

[ZCW15] Yuping Zeng, Jinru Chen, and Feng Wang. Error estimates of the weakly over-penalized symmetric interior penalty method for two variational inequalities. Computers and Mathematics with Applications, 69(8):760–770, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000899.

Zhu:2011:EMS

[ZCY11] Shenglan Zhu, Caisheng Chen, and Huaping Yao. Existence of multiple solutions for a singular quasilinear elliptic equation in R^{N}. Computers and Mathematics with Applications, 62 (12):4525–4534, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008959.

Zhang:2016:NHO

[ZCY16] Jin Zhang, Ke Chen, and Bo Yu. A novel highorder functional based image registration model with inequality constraint. Computers and Mathematics with Applications, 72(12):2887– 2899, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116305892.

Zhang:2017:BLS

[ZCZ17] Xiaoen Zhang, Yong Chen, and Yong Zhang. Breather, lump and X soliton solutions to nonlocal KP equation. Computers and Mathematics with Applications, 74 (10):2341–2347, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304212.

Zhang:2011:IMS

[ZD11] Dan Zhang and Binxiang Dai. Infinitely many solutions for a class of nonlinear impulsive differential equations with periodic boundary conditions. Computers and Mathematics with Applications, 61(10):3153– 3160, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002896.

Ziapour:2012:NSA

[ZD12] Behrooz M. Ziapour and Resam Dehnavi. A numerical study of the arcroof and the one-sided roof enclosures based on the entropy generation minimization. Computers and Mathematics with Applications, 64 (6):1636–1648, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000223.

Zhu:2015:PAE

[ZD15] Lingxue Zhu and Yu Du. Pre-asymptotic error analysis of hp-interior penalty discontinuous Galerkin methods for the Helmholtz equation with large wave number. Computers and Mathematics with Applications, 70(5):917–933, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500293X.

Zuo:2018:MGT

[ZD18] Liyun Zuo and Guangzhi Du. A multi-grid technique for coupling fluid flow with porous media flow. Computers and Mathematics with Applications, 75(11):4012–4021, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301391.

Zhang:2019:EUM

[ZDB19] Qin Zhang, Xuemei Deng, and Qunyi Bie. Existence and uniqueness of mild solutions to the incompressible nematic liquid crystal flow. Computers and Mathematics with Applications, 77(9):2489–2498, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211830734X.

Zlatev:2014:ARE

[ZDF^{+}14] Zahari Zlatev, Ivan Dimov, István Faragó, Krassimir Georgiev, Ágnes Havasi, and Tzvetan Ostromsky. Application of Richardson extrapolation for multi-dimensional advection equations. Computers and Mathematics with Applications, 67(12): 2279–2293, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S089812211400100X.

Zhang:2011:HGB

[ZDL11] Zhening Zhang, Feng Ding, and Xinggao Liu. Hierarchical gradient based iterative parameter estimation algorithm for multivariable output error moving average systems. Computers and Mathematics with Applications, 61(3):672–682, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009193.

Zhang:2014:ICV

[ZDLC14] L. W. Zhang, Y. J. Deng, K. M. Liew, and Y. M. Cheng. The improved complex variable elementfree Galerkin method for two-dimensional Schrödinger equation. Computers and Mathematics with Applications, 68 (10):1093–1106, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114003423.

Zhao:2011:BTP

[ZDM11] Zhanhui Zhao, Zhengde Dai, and Gui Mu. The breather-type and periodictype soliton solutions for the (2 + 1)-dimensional breaking soliton equation. Computers and Mathematics with Applications, 61(8):2048–2052, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006498.

Zmeskal:2013:EFS

[ZDV13] Oldrich Zmeskal, Petr Dzik, and Michal Vesely. Entropy of fractal systems. Computers and Mathematics with Applications, 66(2):135–146, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000345.

Zhang:2017:RSL

[ZDZY17] Yong Zhang, Huanhe Dong, Xiaoen Zhang, and Hongwei Yang. Rational solutions and lump solutions to the generalized (3 + 1)-dimensional Shallow Water-like equation. Computers and Mathematics with Applications, 73(2):246–252, January 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116306265.

Zedan:2010:SAI

[Zed10] Hassan A. Zedan. Symmetry analysis of an integrable Ito coupled system. Computers and Mathematics with Applications, 60(12):3088–3097, December 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007893.

Zelinka:2013:P

[Zel13]  Ivan Zelinka.  Pref-

ace. Computers and Mathematics with Applications, 66(2):105, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113003593.

Zeng:2011:PSD

[Zen11]  Zhijun Zeng.  Peri-

odic solutions for a delayed predator–prey system with stage-structured predator on time scales. Computers and Mathematics with Applications, 61(11):3298–3311, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111003269.

Zhou:2011:ECS

[ZFC11] Xuan Zhou, Hui Feng, and Shihua Chen. The effect of control strength on the synchronization in pinning control questions. Computers and Mathematics with Applications, 61(8):2014–2018, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006279.

Zhou:2018:AFD

[ZFLM18] Yong Zhou, Michal Feckan, Fawang Liu, and J. A. Tenreiro Machado. Advances in fractional differential equations (IV): Time-fractional PDEs. Computers and Mathematics with Applications, 73(6):873, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306873.

Zhou:2019:SIA

[ZFLM19] Yong Zhou, Michal Feckan, Fawang Liu, and J. A. Tenreiro Machado. Special issue: Advances in fractional differential equations (V): Time–space fractional PDEs. Computers and Mathematics with Applications, 78(5): 1243, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122119303669.

Zhang:2019:GSG

[ZFY^{+}19] Rui Zhang, Xiangchu Feng, Lixia Yang, Lihong Chang, and Xiaolong Zhu. A global sparse gradient based coupled system for image denoising. Computers and Mathematics with Applications, 78(11):3692–3711, December 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119303128.

Zhang:2010:ERE

[ZFZ10a] Jimin Zhang, Meng Fan, and Huaiping Zhu. Existence and roughness of exponential dichotomies of linear dynamic equations on time scales. Computers and Mathematics with Applications, 59(8):2658–2675, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000057X.

Zhang:2010:NSC

[ZFZ10b] Jimin Zhang, Meng Fan, and Huaiping Zhu. Necessary and sufficient criteria for the existence of exponential dichotomy on time scales. Computers and Mathematics with Applications, 60(8):2387–2398, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005985.

Zhang:2010:GES

[ZFZQ10] Dongkai Zhang, Wenli Feng, Yongqiang Zhao, and Jiqing Qiu. Global existence of solutions for fuzzy second-order differential equations under generalized H-differentiability. Computers and Mathematics with Applications, 60(6):1548–1556, September 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122110004529.

Zhou:2014:IAN

[ZG14] Xiuxiang Zhou and Hang Gao. Interior approximate and null controllability of the heat equation with memory. Computers and Mathematics with Applications, 67(3):602–613, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113006846.

Zhou:2016:FEA

[ZG16] Zhaojie Zhou and Wei Gong. Finite element approximation of optimal control problems governed by time fractional diffusion equation. Computers and Mathematics with Applications, 71(1):301–318, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211500557X.

Zhou:2018:FMA

[ZG18a] Qing-Mei Zhou and Bin Ge. The fibering map approach to a nonlocal problem involving p(x)-Laplacian. Computers and Mathematics with Applications, 75(2):632–642, January 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730617X.

Zou:2018:DDP

[ZG18b] Rong Zou and Shangjiang Guo. Dynamics in a diffusive predator-prey system with ratio-dependent predator influence. Computers and Mathematics with Applications, 75(4):1237–1258, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307022.

Zhong:2019:FAS

[ZG19] Hong-Xiu Zhong and Xian-Ming Gu. A flexible and adaptive simpler GMRES with deflated restarting for shifted linear systems. Computers and Mathematics with Applications, 78(3):997–1007, August 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301427.

Zlatev:2013:ICC

[ZGD13] Z. Zlatev, K. Georgiev, and I. Dimov. Influence of climatic changes on pollution levels in the Balkan peninsula. Computers and Mathematics with Applications, 65(3):544–562, February 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004737.

Zlatev:2014:SAS

[ZGD14] Zahari Zlatev, Krassimir Georgiev, and Ivan Dimov. Studying absolute stability properties of the Richardson extrapolation combined with explicit Runge–Kutta methods. Computers and Mathematics with Applications, 67(12):2294– 2307, July 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000960.

Zhu:2014:MSS

[ZGL14] Sheng-Xin Zhu, Tong-Xiang Gu, and Xing-Ping Liu. Minimizing synchronizations in sparse iterative solvers for distributed supercomputers. Computers and Mathematics with Applications, 67(1):199–209, January 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006640.

Zhang:2011:OSO

[ZGW11] Quanxin Zhang, Li Gao, and Lei Wang. Oscillation of second-order nonlinear delay dynamic equations on time scales. Computers and Mathematics with Applications, 61(8):2342–2348, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007844.

Zhou:2013:DBE

[ZGZ13] Aimin Zhou, Feng Gao, and Guixu Zhang. A decomposition based estimation of distribution algorithm for multiobjective traveling salesman problems. Computers and Mathematics with Applications, 66 (10):1857–1868, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211300463X.

Zhang:2011:CNM

[ZH11] Juan Zhang and Huan Huang. A complex network model of polymer melts without growth. Computers and Mathematics with Applications, 61(8):2205–2208, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007029.

Zhang:2015:DSU

[ZH15a] Guo-Dong Zhang and Yinnian He. Decoupled schemes for unsteady MHD equations II: Finite element spatial discretization and numerical implementation. Computers and Mathematics with Applications, 69(12):1390– 1406, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115001248.

Zhou:2015:DAS

[ZH15b] Shaoling Zhou and Lei Hou. Decoupled algorithm for solving Phan-Thien–Tanner viscoelastic fluid by finite element method. Computers and Mathematics with Applications, 69(5):423–437, March 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115000188.

Zhang:2018:EGS

[ZH18a] Hongwei Zhang and Qingying Hu. Existence of the global solution for fractional logarithmic Schrödinger equation. Computers and Mathematics with Applications, 75(1):161–169, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305588.

Zhu:2018:MBS

[ZH18b] Song-Ping Zhu and Xin-Jiang He. A modified Black–Scholes pricing formula for European options with bounded underlying prices. Computers and Mathematics with Applications, 75(5):1635–1647, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307496.

Zhang:2019:SCF

[ZH19] Yarong Zhang and Yinnian He. Stability and convergence of a finite element method for a semilinear elliptical problem with small viscosity. Computers and Mathematics with Applications, 78 (10):3363–3374, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930272X.

Zhang:2010:PSS

[Zha10]  Shuqin Zhang.  Posi-

tive solutions to singular boundary value problem for nonlinear fractional differential equation. Computers and Mathematics with Applications, 59(3):1300–1309, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122109004167.

Zhang:2011:ASG

[Zha11a] Qing-Bang Zhang. An algorithm for solving the general variational inclusion involving Amonotone operators. Computers and Mathematics with Applications, 61(6): 1682–1686, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111000654.

Zhang:2011:MEC

[Zha11b] Shilin Zhang. A model for evaluating computer network security systems with 2-tuple linguistic information. Computers and Mathematics with Applications, 62(4):1916–1922, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005153.

Zhang:2011:ESFa

[Zha11c]  Shuqin Zhang.  Exis-

tence of a solution for the fractional differential equation with nonlinear boundary conditions. Computers and Mathematics with Applications, 61(4):1202–1208, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110009879.

Zhang:2011:FPT

[Zha11d]  Xian Zhang.  Fixed

point theorem of generalized quasi-contractive mapping in cone metric space. Computers and Mathematics with Applications, 62(4):1627–1633, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111002872.

Zhao:2013:ORC

[Zha13] Fuzhang Zhao. Optimal relaxation collisions for lattice Boltzmann methods. Computers and Mathematics with Applications, 65(2):172–185, January 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111004731.

Zhao:2014:FPM

[Zha14] Jian Xun Zhao. A fixedpoint method for a class of super-large scale nonlinear complementarity problems. Computers and Mathematics with Applications, 67(5):999–1015, March 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000236.

Zhang:2015:RRG

[Zha15a] Huamin Zhang. Reducedrank gradient-based algorithms for generalized coupled Sylvester matrix equations and its applications. Computers and Mathematics with Applications, 70(8):2049–2062, October 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500382X.

Zhao:2015:SFM

[Zha15b] Jian-Xun Zhao. A simple and feasible method for a class of large-scale l^{1}- problems. Computers and Mathematics with Applications, 70(5):883–895, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122115002904.

Zhang:2017:PSA

[Zha17a] Shougui Zhang. Projection and self-adaptive projection methods for the Signorini problem with the BEM. Computers and Mathematics with Applications, 74 (6):1262–1273, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303693.

Zhang:2017:CLP

[Zha17b] Zhi-Yong Zhang. Conservation laws of partial differential equations: Symmetry, adjoint symmetry and nonlinear self-adjointness. Computers and Mathematics with Applications, 74 (12):3129–3140, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304868.

Zhan:2018:WSE

[Zha18a] Huashui Zhan. The weak solutions of an evolutionary p(x)-Laplacian equation are controlled by the initial value. Computers and Mathematics with Applications, 76 (9):2272–2285, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304462.

Zhang:2018:SPF

[Zha18b] Jia-Fang Zhang. Spatial patterns of a fractional type cross-diffusion Holling–Tanner model. Computers and Mathematics with Applications, 76(4):957–965, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303092.

Zhang:2018:PBE

[Zha18c]  Shangyou Zhang.  A

P_{4} bubble enriched P_{3} divergence-free finite element on triangular grids. Computers and Mathematics with Applications, 74(11):2710–2722, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730384X.

Zhang:2018:SPT

[Zha18d] Shengliang Zhang. A symplectic procedure for two-dimensional coupled elastic wave equations using radial basis functions interpolation. Computers and Mathematics with Applications, 76 (9):2167–2178, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304322.

Zhang:2018:BCQ

[Zha18e] Zujin Zhang. On the blow-up criterion for the quasi-geostrophic equations in homogeneous Besov spaces. Computers and Mathematics with Applications, 75(3):1038–1043, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306892.

Zhang:2018:RCT

[Zha18f] Zujin Zhang. Regularity criteria for the three dimensional Ericksen– Leslie system in homogeneous Besov spaces. Computers and Mathematics with Applications, 75(3):1060–1065, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306879.

Zhang:2018:SNR

[Zha18g]  Zujin Zhang.  Sev-

eral new regularity criteria for the axisymmetric Navier–Stokes equations with swirl. Computers and Mathematics with Applications, 76 (6):1420–1426, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303596.

Zhao:2018:CRD

[Zha18h] Wenqiang Zhao. Continuity and random dynamics of the nonautonomous stochastic FitzHugh–Nagumo system on RN. Computers and Mathematics with Applications, 75(10):3801–3824, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301123.

Zhaqilao:2018:SCA

[Zha18i] Zhaqilao. A symbolic computation approach to constructing rogue waves with a controllable center in the nonlinear systems. Computers and Mathematics with Applications, 75(9):3331–3342, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118300749.

Zhang:2019:QGB

[Zha19a] Huamin Zhang. Quasi gradient-based inversionfree iterative algorithm for solving a class of the nonlinear matrix equations. Computers and Mathematics with Applications, 77(5):1233–1244, March 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306564.

Zhang:2019:RNP

[Zha19b] Jianjun Zhang. A relaxed Newton–Picard like method for Huber variant of total variation based image restoration. Computers and Mathematics with Applications, 78(1):224–239, July 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300951.

Zhang:2019:SSM

[Zha19c] Zhongxin Zhang. Similarity solutions to the MHD boundary layer equations with a negative parameter for power-law fluids. Computers and Mathematics with Applications, 78(8):2806–2816, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302457.

Zalaket:2011:PFU

[ZHB11] Joseph Zalaket and Joseph Hajj-Boutros. Prime factorization using square root approximation. Computers and Mathematics with Applications, 61(9): 2463–2467, May 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122111001131.

Zhang:2017:BEM

[ZHC17] Yarong Zhang, Yinnian He, and Hongbin Chen. Boundary element method for a free boundary problem modeling three dimensional tumor growth. Computers and Mathematics with Applications, 73(7):1627–1641, April 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730069X.

Zheng:2011:SUA

[Zhe11] Jinling Zheng. A simple universal adaptive feedback controller for chaos and hyperchaos control. Computers and Mathematics with Applications, 61(8):2000–2004, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006243.

Zhang:2011:LTI

[ZHJ11] Qi-Ming Zhang, Xiaofei He, and Jianchu Jiang. On Lyapunovtype inequalities for nonlinear dynamic systems on time scales. Computers and Mathematics with Applications, 62 (11):4028–4038, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100808X.

Zhang:2014:SSD

[ZHJ14] Yunzhang Zhang, Yanren Hou, and Hongen Jia. Subgrid stabilized defect-correction method for a steady-state natural convection problem. Computers and Mathematics with Applications, 67(3):497–514, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113006585.

Zhao:2013:GPT

[ZHJD13] Liang Zhao, Ting-Zhu Huang, Yan-Fei Jing, and Liang-Jian Deng. A generalized product-type Bi-COR method and its application in signal deconvolution. Computers and Mathematics with Applications, 66(8):1372–1388, November 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004884.

Zhao:2011:HRA

[ZHJZ11] Feng Zhao, Heqing Huang, Hai Jin, and Qin Zhang. A hybrid ranking approach to estimate vulnerability for dynamic attacks. Computers and Mathematics with Applications, 62(12):4308– 4321, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111007991.

Zhao:2012:IIM

[ZHL12] Jianping Zhao, Yanren Hou, and Yongfei Li. Immersed interface method for elliptic equations based on a piecewise second order polynomial. Computers and Mathematics with Applications, 63(5):957–965, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010431.

Zhou:2010:P

[Zho10]  Yong Zhou.  Preface.

Computers and Mathematics with Applications, 59(3):1047, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000271.

Zhou:2011:AFD

[Zho11] Yong Zhou. Advances in fractional differential equations II. Computers and Mathematics with Applications, 62(3):821, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006006.

Zhou:2012:AFD

[Zho12] Yong Zhou. Advances in fractional differential equations, III. Computers and Mathematics with Applications, 64(10): 2965, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112006177.

Zhou:2013:SPF

[Zho13] Jun Zhou. Spatiotemporal pattern formation of a diffusive bimolecular model with autocatalysis and saturation law. Computers and Mathematics with Applications, 66(10):2003–2018, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113005178.

Zhou:2016:GEB

[Zho16] Jun Zhou. Global existence and blow-up of solutions for a non-Newton polytropic filtration system with special volumetric moisture content. Computers and Mathematics with Applications, 71(5):1163–1172, March 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300256.

Zhou:2019:GSS

[Zho19] Jun Zhou. Ground state solution for a fourthorder elliptic equation with logarithmic nonlinearity modeling epitaxial growth. Computers and Mathematics with Applications, 78 (6):1878–1886, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301506.

Zhang:2012:SPA

[ZHQG12] Ruoyu Zhang, Shiqiu Huang, Zhengwei Qi, and Haibing Guan. Static program analysis assisted dynamic taint tracking for software vulnerability discovery. Computers and Mathematics with Applications, 63(2):469–480, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100664X.

Zhang:2019:NSG

[ZHS^{+}19] Rongpei Zhang, Zijian Han, Yongyun Shao, Zhen Wang, and Yu Wang. The numerical study for the ground and excited states of fractional Bose–Einstein condensates. Computers and Mathematics with Applications, 78(5):1548–1561, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211930166X.

Zhang:2011:OMC

[ZHT11] Qun Zhang, Xiaoxia Huang, and Leming Tang. Optimal multinational capital budgeting under uncertainty. Computers and Mathematics with Applications, 62 (12):4557–4567, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008996.

Zhu:2010:JTI

[Zhu10]  Ling Zhu.  Jordan

type inequalities involving the Bessel and modified Bessel functions. Computers and Mathematics with Applications, 59(2):724–736, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109007196.

Zeman:2019:HFC

[ZHV19] Michal Zeman, Milan Holec, and Pavel Váchal. HerEOS: a framework for consistent treatment of the equation of state in ALE hydrodynamics. Computers and Mathematics with Applications, 78(2):483–503, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122118306096.

Zhu:2011:IBM

[ZHW^{+}11] Luoding Zhu, Guowei He, Shizhao Wang, Laura Miller, Xing Zhang, Qian You, and Shiaofen Fang. An immersed boundary method based on the lattice Boltzmann approach in three dimensions, with application. Computers and Mathematics with Applications, 61(12):3506–3518, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110001987.

Zhang:2014:PSD

[ZHW14] Xueli Zhang, Yehui Huang, and Peixuan Weng. Permanence and stability of a diffusive predator– prey model with disease in the prey. Computers and Mathematics with Applications, 68 (10):1431–1445, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004593.

Zhang:2018:CSR

[ZHW^{+}18] Wenhuan Zhang, Changsheng Huang, Yihang Wang, Baochang Shi, Shibo Kuang, and Zhenhua Chai. The computation of strain rate tensor in multiple-relaxation-time lattice Boltzmann model. Computers and Mathematics with Applications, 75(8):2888–2902, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830035X.

Zhang:2014:ACI

[ZHY14] Guo-Dong Zhang, Yinnian He, and Di Yang. Analysis of coupling iterations based on the finite element method for stationary magnetohydrodynamics on a general domain. Computers and Mathematics with Applications, 68(7):770–788, October 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114003435.

Zhang:2014:EAF

[ZHZ14] Yunzhang Zhang, Yanren Hou, and Jianping Zhao. Error analysis of a fully discrete finite element variational multiscale method for the natural convection problem. Computers and Mathematics with Applications, 68(4):543–567, August 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002673.

Zhan:2010:SAB

[ZJ10a] Jianming Zhan and Young Bae Jun. Soft BL-algebras based on fuzzy sets. Computers and Mathematics with Applications, 59(6):2037–2046, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900755X.

Zhang:2010:MSS

[ZJ10b] Qiumei Zhang and Daqing Jiang. Multiple solutions to semipositone Dirichlet boundary value problems with singular dependent nonlinearities for second order threepoint differential equations. Computers and Mathematics with Applications, 59(8):2516–2527, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211000009X.

Zhou:2010:EMS

[ZJ10c] Yong Zhou and Feng Jiao. Existence of mild solutions for fractional neutral evolution equations. Computers and Mathematics with Applications, 59(3):1063–1077, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109003964.

Zhang:2012:SPL

[ZJ12] Xiaoming Zhang and Weidong Jiang. Some properties of log-convex function and applications for the exponential function. Computers and Mathematics with Applications, 63(6):1111–1116, March 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111010741.

Zhang:2017:ENS

[ZJ17] Jian Zhang and Yanju Ji. The existence of nontrivial solutions for the critical Kirchhoff type problem in RN. Computers and Mathematics with Applications, 74 (12):3080–3094, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117304844.

Zhang:2019:MFH

[ZJB19] Yan Zhang, Jinxia Jiang, and Yu Bai. MHD flow and heat transfer analysis of fractional Oldroyd-B nanofluid between two coaxial cylinders. Computers and Mathematics with Applications, 78 (10):3408–3421, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302767.

Zhong:2011:ASP

[ZJZ^{+}11] Zhen Zhong, Ting Jiang, Wenshuo Zhang, Haipeng Yao, and Shuifang Xiao. Analyzing speech of patients with vocal polyps based on channel parameters and fuzzy logic systems. Computers and Mathematics with Applications, 62(7):2834–2842, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006201.

Zheng:2018:LFS

[ZJZ18] Rumeng Zheng, Xiaoyun Jiang, and Hui Zhang. L1 Fourier spectral methods for a class of generalized two-dimensional time fractional nonlinear anomalous diffusion equations. Computers and Mathematics with Applications, 75(5):1515–1530, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307319.

Zaman:2016:DAA

[ZK16] Gul Zaman and Asaf Khan. Dynamical aspects of an age-structured SIR endemic model. Computers and Mathematics with Applications, 72 (6):1690–1702, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304291.

Zwicke:2016:AIM

[ZKBE16] Florian Zwicke, Philipp Knechtges, Marek Behr, and Stefanie Elgeti. Automatic implementation of material laws: Jacobian calculation in a finite element code with TAPENADE. Computers and Mathematics with Applications, 72 (11):2808–2822, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305685.

Zander:2012:FCM

[ZKR^{+}12] N. Zander, S. Kollmannsberger, M. Ruess, Z. Yosibash, and E. Rank. The finite cell method for linear thermoelasticity. Computers and Mathematics with Applications, 64(11):3527–3541, December 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112005688.

Zheng:2015:PSH

[ZKW15] Jiashan Zheng, Yuanyuan Ke, and Yifu Wang. Periodic solutions to a heat equation with hysteresis in the source term.

Computers and Mathematics with Applications, 69(2):134–143, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005884.

Zhang:2017:BDI

[ZKWW17] Huai Zhang, Tong Kang, Ran Wang, and Yanfang Wang. Boundary data identification for an electromagnetic problem by means of the potential field method. Computers and Mathematics with Applications, 73(4):588–600, February 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116306927.

Zhao:2010:FPP

[ZL10a] Kaihong Zhao and Yongkun Li. Four positive periodic solutions to two species parasitical system with harvesting terms. Computers and Mathematics with Applications, 59(8):2703–2710, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000060X.

Zhong:2010:NMP

[ZL10b] Wenyong Zhong and Wei Lin. Nonlocal and multiple-point boundary value problem for fractional differential equations. Computers and Mathematics with Applications, 59(3):1345–1351, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812210900412X.

Zhang:2011:WPA

[ZL11] Li-Li Zhang and Hong-Xu Li. Weighted pseudo almost periodic solutions of second-order neutraldelay differential equations with piecewise constant argument. Computers and Mathematics with Applications, 62 (12):4362–4376, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008686.

Zhao:2012:NAG

[ZL12a] Zhengang Zhao and Changpin Li. A numerical approach to the generalized nonlinear fractional Fokker–Planck equation. Computers and Mathematics with Applications, 64(10):3075–3089, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112000880.

Zhou:2012:SNA

[ZL12b] Jituan Zhou and Linzhang Lu. On simultaneously nilpotent antiring matrices. Computers and Mathematics with Applications, 64(9):2912–2916, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122112003860.

Zhou:2013:SNH

[ZL13] Tianhe Zhou and Zhong Li. Scattered noisy Hermite data fitting using an extension of the weighted least squares method. Computers and Mathematics with Applications, 65(12):1967–1977, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122113002174.

Zhang:2014:DDP

[ZL14a] Cun-Hua Zhang and Zhi-Zhen Li. Dynamics in a diffusive plantherbivore model with toxin-determined functional response. Computers and Mathematics with Applications, 67(8):1439– 1449, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400090X.

Zhang:2014:WCM

[ZL14b] Li-Tao Zhang and Jian-Lei Li. The weaker convergence of modulusbased synchronous multisplitting multi-parameters methods for linear complementarity problems. Computers and Mathematics with Applications, 67(10):1954–1959, June 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114001783.

Zheng:2016:SCU

[ZL16] Qingqing Zheng and Linzhang Lu. On semiconvergence of ULT iterative method for the singular saddle point problems. Computers and Mathematics with Applications, 72(6):1549–1555, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304047.

Zhang:2018:GEF

[ZL18] Tao Zhang and Xiaolin Li. A generalized elementfree Galerkin method for Stokes problem. Computers and Mathematics with Applications, 75(9):3127–3138, May 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300518.

Zhang:2019:ITD

[ZL19a] Mengmeng Zhang and Jijun Liu. Identification of a time-dependent source term in a distributedorder time-fractional equation from a nonlocal integral observation. Computers and Mathematics with Applications, 78 (10):3375–3389, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302718.

Zhang:2019:FVM

[ZL19b] Tie Zhang and Zheng Li. A finite volume method for Stokes problems on quadrilateral meshes. Computers and Mathematics with Applications, 77(4):1091–1106, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306412.

Zhao:2017:NMT

[ZLA17] Linlin Zhao, Fawang Liu, and Vo V. Anh. Numerical methods for the two-dimensional multiterm time-fractional diffusion equations. Computers and Mathematics with Applications, 74 (10):2253–2268, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211730425X.

Zheng:2011:DSL

[ZLC^{+}11a] Zhong Zheng, Liangcheng Lin, Zhiwei Chen, Jim Lim, and John Grace. A dual-scale lattice gas automata model for gassolid two-phase flow in bubbling fluidized beds. Computers and Mathematics with Applications, 61(12):3593–3605, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111000472.

Zhu:2011:SHB

[ZLC11b] Huiyan Zhu, Yang Luo, and Meiling Chen. Stability and Hopf bifurcation of a HIV infection model with CTL-response delay. Computers and Mathematics with Applications, 62(8):3091–3102, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111006857.

Zhang:2014:FSA

[ZLC^{+}14] Jieyuan Zhang, Shengguo Li, Lizhi Cheng, Xiangke Liao, and Guangquan Cheng. A fast and stable algorithm for downdating the singular value decomposition. Computers and Mathematics with Applications, 68 (10):1421–1430, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004568.

Ziari:2010:OAS

[ZLG^{+}10] I. Ziari, G. Ledwich, A. Ghosh, D. Cornforth, and M. Wishart. Optimal allocation and sizing of capacitors to minimize the transmission line loss and to improve the voltage profile. Computers and Mathematics with Applications, 60(4):1003–1013, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110002312.

Zheng:2011:MME

[ZLG^{+}11] Chuguang Zheng, Jidong Lu, Zhaoli Guo, Li-Shi Luo, and Manfred Krafczyk. Mesoscopic methods in engineering and science. Computers and Mathematics with Applications, 61(12):3401–3403, June 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111003427.

Zhang:2018:PSN

[ZLG18] Jia-Feng Zhang, Chun-Yu Lei, and Liu-Tao Guo. Positive solutions for a nonlocal Schrödinger– Newton system involving critical nonlinearity. Computers and Mathematics with Applications, 76(8):1966–1974, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304139.

Zou:2011:NMD

[ZLGL11] Dexuan Zou, Haikuan Liu, Liqun Gao, and Steven Li. A novel modified differential evolution algorithm for constrained optimization problems. Computers and Mathematics with Applications, 61(6):1608–1623, March 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111000460.

Zheng:2019:SES

[ZLHF19] Yong-Lai Zheng, Don Liu, Hui-Li Han, and Mohammad Ferdows. Spectral element simulations of interactive particles in a fluid. Computers and Mathematics with Applications, 77(8):2029–2050, April 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300100.

Zhang:2018:CNA

[ZLJ^{+}18] Hui Zhang, Fawang Liu, Xiaoyun Jiang, Fanhai Zeng, and Ian Turner. A Crank–Nicolson ADI Galerkin–Legendre spectral method for the two-dimensional Riesz space distributed-order advection–diffusion equation. Computers and Mathematics with Applications, 76(10):2460– 2476, November 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122118304681.

Zhang:2011:SCT

[ZLL11] Congjun Zhang, Jinlu Li, and Baoqing Liu. Strong convergence theorems for equilibrium problems and relatively nonexpansive mappings in Banach spaces. Computers and Mathematics with Applications, 61(2):262–276, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110008515.

Zhang:2012:TRS

[ZLL12] Xiangsong Zhang, Zhenhua Liu, and Sanyang Liu. A trust region SQP-filter method for nonlinear secondorder cone programming. Computers and Mathematics with Applications, 63(12):1569–1576, June 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112000120.

Zhang:2014:NSW

[ZLL14] Li Zhang, Yezhi Lin, and Yinping Liu. New solitary wave solutions for two nonlinear evolution equations. Computers and Mathematics with Applications, 67(8):1595– 1606, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114000881.

Zhao:2017:AST

[ZLL17] Zhihui Zhao, Hong Li, and Zhendong Luo. Analysis of a space-time continuous Galerkin method for convection-dominated Sobolev equations. Computers and Mathematics with Applications, 73(8):1643–1656, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300536.

Zhu:2018:PPC

[ZLL18] Song-Ping Zhu, Sha Lin, and Xiaoping Lu. Pricing puttable convertible bonds with integral equation approaches. Computers and Mathematics with Applications, 75(8):2757–2781, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300117.

Zhang:2012:SDM

[ZLLF12] Jianke Zhang, Sanyang Liu, Lifeng Li, and Quanxi Feng. Sufficiency and duality for multiobjective variational control problems with Ginvexity. Computers and Mathematics with Applications, 63(4):838–850, February 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111010327.

See comments [Ant14].

Zhang:2018:STD

[ZLMZ18] Tonghua Zhang, Xia Liu, Xinzhu Meng, and Tongqian Zhang. Spatiotemporal dynamics near the steady state of a planktonic system. Computers and Mathematics with Applications, 75(12):4490–4504, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301809.

Zhang:2013:NNM

[ZLPM13] H. Zhang, F. Liu, Mantha S. Phanikumar, and Mark M. Meerschaert. A novel numerical method for the time variable fractional order mobileimmobile advection-dispersion model. Computers and Mathematics with Applications, 66(5):693–701, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113000618.

Zhong:2013:TLA

[ZLS13] Liuqiang Zhong, Chunmei Liu, and Shi Shu. Two-level additive preconditioners for edge element discretizations of time-harmonic Maxwell equations. Computers and Mathematics with Applications, 66 (4):432–440, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113003234.

Zeng:2019:AEB

[ZLS19] Xianyi Zeng, Kangan Li, and Guglielmo Scovazzi. An ALE/embedded boundary method for two-material flow simulations. Computers and Mathematics with Applications, 78(2):335–361, July 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302621.

Zhang:2018:MLS

[ZLT18] Yan Zhang, Yinping Liu, and Xiaoyan Tang. Mlump solutions to a (3 + 1)-dimensional nonlinear evolution equation. Computers and Mathematics with Applications, 76(3):592–601, August 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302530.

Zhang:2016:NST

[ZLTY16] H. Zhang, F. Liu, I. Turner, and Q. Yang. Numerical solution of the time fractional Black–Scholes model governing European options. Computers and Mathematics with Applications, 71(9):1772– 1783, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300499.

Zeng:2010:ANO

[ZLW10] D. Q. Zeng, Y. Y. Lee, and C. K. Wong. Analysis of a nonlinear oscillator with discontinuity. Computers and Mathematics with Applications, 59(8):2510–2515, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110000039.

Zhang:2018:RVD

[ZLW18] Jun Zhang, Haijiao Liu, and Zhihui Wei. Regularized variational dynamic stochastic resonance method for enhancement of dark and low-contrast image. Computers and Mathematics with Applications, 76(4):774–787, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118302864.

Zhu:2019:EUG

[ZLW19] Bo Zhu, Lishan Liu, and Yonghong Wu. Existence and uniqueness of global mild solutions for a class of nonlinear fractional reaction–diffusion equations with delay. Computers and Mathematics with Applications, 78 (6):1811–1818, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116300244.

Zheng:2011:GVP

[ZLWL11] Cheng-Bo Zheng, Bin Liu, Zuo-Jun Wang, and Hong-Shi Lü. Generalized variational principles for micromorphic magnetoelectroelastodynamics. Computers and Mathematics with Applications, 61(8):2201–2204, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007017.

Zhou:2012:EMB

[ZLY12] Ligang Zhou, Kin Keung Lai, and Jerome Yen. Empirical models based on features ranking techniques for corporate financial distress prediction. Computers and Mathematics with Applications, 64(8):2484–2496, October 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112004336.

Zhao:2013:NHD

[ZLY^{+}13] Xinchao Zhao, Wenqiao Lin, Chengchi Yu, Jing Chen, and Shuguang Wang. A new hybrid differential evolution with simulated annealing and self-adaptive immune operation. Computers and Mathematics with Applications, 66 (10):1948–1960, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113004318.

Zhang:2014:CNT

[ZLY14] Wei Zhang, Jiaojie Li, and Yupu Yang. A class of nonlocal tensor telegraph-diffusion equations applied to coherence enhancement. Computers and Mathematics with Applications, 67(8):1461– 1473, May 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114000844.

Zhu:2017:TWN

[ZLY17] Cheng-Cheng Zhu, Wan-Tong Li, and Fei-Ying Yang. Traveling waves in a nonlocal dispersal SIRH model with relapse. Computers and Mathematics with Applications, 73(8):1707–1723, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117300834.

Liu:2019:NVC

[zLYLQ19] Xi zhong Liu, Jun Yu, Zhi-Mei Lou, and Xian-Min Qian. A nonlocal variable coefficient modified KdV equation derived from a two-layer fluid system and its exact solutions. Computers and Mathematics with Applications, 78 (6):2083–2093, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301828.

Liu:2018:NBT

[zLYmL18] Xi zhong Liu, Jun Yu, and Zhi mei Lou. New Bäcklund transformations of the (2 + 1)- dimensional Bogoyavlenskii equation via localization of residual symmetries. Computers and Mathematics with Applications, 76(7):1669–1679, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303833.

Zheng:2010:NFE

[ZLZ10] Yunying Zheng, Changpin Li, and Zhengang Zhao. A note on the finite element method for the space-fractional advection diffusion equation. Computers and Mathematics with Applications, 59(5):1718–1726, March 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109005495.

Zhang:2011:CHL

[ZLZ11] Fengxia Zhang, Ying Li, and Jianli Zhao. Common Hermitian least squares solutions of matrix equations A_{1}XA_{1}^{∗} = B_{1} and A_{2}XA_{2}^{∗} = B_{2} subject to inequality restrictions. Computers and Mathematics with Applications, 62(6):2424–2433, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005839.

Zhu:2018:CCD

[ZLZ18] Min Zhu, Zhigui Lin, and Qunying Zhang. Coexistence of a crossdiffusive dengue fever model in a heterogeneous environment. Computers and Mathematics with Applications, 75(3):1004–1015, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306910.

Liu:2019:NNE

[zLZ19] Hong zhun Liu and Tong Zhang. Notes on “The new exact solitary and multi-soliton solutions for the (2 + 1)- dimensional Zakharov– Kuznetsov equation”. Computers and Mathematics with Applications, 77(7):1980–1982, April 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306849.

Zheng:2011:ESU

[ZLZG11] Liancun Zheng, Chunrui Li, Xinxin Zhang, and Yingtao Gao. Exact solutions for the unsteady rotating flows of a generalized Maxwell fluid with oscillating pressure gradient between coaxial cylinders. Computers and Mathematics with Applications, 62(3):1105–1115, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100143X.

Zinovyev:2013:DCM

[ZM13] Andrei Zinovyev and Evgeny Mirkes. Data complexity measured by principal graphs. Computers and Mathematics with Applications, 65(10): 1471–1482, May 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122112007055.

Zapata:2014:FNR

[ZM14] Juan Luis García Zapata and Juan Carlos Díaz Martín. Finding the number of roots of a polynomial in a plane region using the winding number. Computers and Mathematics with Applications, 67(3):555–568, February 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300669X.

Zeng:2016:PSI

[ZM16a] Min-Li Zeng and Chang-Feng Ma. A parameterized SHSS iteration method for a class of complex symmetric system of linear equations. Computers and Mathematics with Applications, 71(10):2124–2131, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301705.

Zhou:2016:SOL

[ZM16b] Zhiqiang Zhou and Jingtang Ma. Second-order lattice Boltzmann methods for PDEs of Asian option pricing with regime switching. Computers and Mathematics with Applications, 71(7):1448–1463, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116300736.

Zhang:2017:RMW

[ZM17a] Hai-Qiang Zhang and Wen-Xiu Ma. Resonant multiple wave solutions for a (3 + 1)-dimensional nonlinear evolution equation by linear superposition principle. Computers and Mathematics with Applications, 73(10):2339–2343, May 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301761.

Zhou:2017:ALS

[ZM17b] Yuan Zhou and Wen-Xiu Ma. Applications of linear superposition principle to resonant solitons and complexitons. Computers and Mathematics with Applications, 73(8):1697–1706, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300846.

Zeng:2018:ESI

[ZM18] Biao Zeng and Stanisław Migórski. Evolutionary subgradient inclusions with nonlinear weakly continuous operators and applications. Computers and Mathematics with Applications, 75(1):89–104, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122117305321.

Zurigat:2010:AAS

[ZMA10] Mohammad Zurigat, Shaher Momani, and Ahmad Alawneh. Analytical approximate solutions of systems of fractional algebraic-differential equations by homotopy analysis method. Computers and Mathematics with Applications, 59(3):1227–1235, February 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109004209.

Zhou:2018:NPM

[ZMFL18] Xinchen Zhou, Zhaoliang Meng, Xin Fan, and Zhongxuan Luo. Nonconforming polynomial mixed finite element for the Brinkman problem over quadrilateral meshes. Computers and Mathematics with Applications, 76(4):877–892, August 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118302955.

Zhang:2010:SVP

[ZMG10] Hehua Zhang, Stephan Merz, and Ming Gu. Specifying and verifying PLC systems with TLA ^{+}: a case study. Computers and Mathematics with Applications, 60(3):695–705, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003664.

Zheng:2016:BHD

[ZMH16] Pan Zheng, Chunlai Mu, and Xuegang Hu. Boundedness in the higher dimensional attractionrepulsion chemotaxisgrowth system. Computers and Mathematics with Applications, 72 (9):2194–2202, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304825.

Zhang:2016:SLS

[ZMLZ16] Fengxia Zhang, Weisheng Mu, Ying Li, and Jianli Zhao. Special least squares solutions of the quaternion matrix equation AXB + CXD = E. Computers and Mathematics with Applications, 72(5):1426–1435, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304114.

Zapletal:2018:BEQ

[ZMM18] Jan Zapletal, Michal Merta, and Lukás Malý. Boundary element quadrature schemes for multiand many-core architectures. Computers and Mathematics with Applications, 74(1):157–173, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117300482.

Zheng:2018:GAS

[ZMWH18] Pan Zheng, Chunlai Mu, Robert Willie, and Xuegang Hu. Global asymptotic stability of steady states in a chemotaxisgrowth system with singular sensitivity. Computers and Mathematics with Applications, 75(5):1667–1675, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307587.

Zamolo:2018:TAF

[ZN18] Riccardo Zamolo and Enrico Nobile. Two algorithms for fast 2D node generation: Application to RBF meshless discretization of diffusion problems and image halftoning. Computers and Mathematics with Applications, 75(12):4305–4321, June 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301676.

Zhao:2011:NAP

[ZNWG11] Shu-Zhi Zhao, Tong-He Ni, Yang Wang, and Xiang-Tao Gao. A new approach to the prediction of passenger flow in a transit system. Computers and Mathematics with Applications, 61(8):1968–1974, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005870.

Zhao:2019:IMM

[ZÖXL^{+}19] Jiaheng Zhao, Ilhan Özgen-Xian, Dongfang Liang, Tian Wang, and Reinhard Hinkelmann. An improved multislope MUSCL scheme for solving shallow water equations on unstructured grids. Computers and Mathematics with Applications, 77(2):576–596, January 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305790.

Zheng:2012:AMC

[ZOZZ12] Supei Zheng, Jie Ouyang, Zhifeng Zhao, and Ling Zhang. An adaptive method to capture weldlines during the injection mold filling. Computers and Mathematics with Applications, 64 (9):2860–2870, November 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122112003689.

Zhang:2010:DSN

[ZP10] Yong Zhang and Hao Pan. On the divisibility of q-Salié numbers. Computers and Mathematics with Applications, 59(12):3640– 3645, June 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000221X.

Zhao:2018:FCB

[ZP18a] Lina Zhao and Eun-Jae Park. Fully computable bounds for a staggered discontinuous Galerkin method for the Stokes equations. Computers and Mathematics with Applications, 75(11):4115–4134, June 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301536.

Zhou:2018:TFN

[ZP18b] Yong Zhou and Li Peng.

On the time-fractional Navier–Stokes equations. Computers and Mathematics with Applications, 73(6):874–891, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116301675.

Zhou:2018:WST

[ZP18c] Yong Zhou and Li Peng. Weak solutions of the time-fractional Navier– Stokes equations and optimal control. Computers and Mathematics with Applications, 73(6):1016–1027, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116303984.

Zong:2016:DCF

[ZPGW16] Yuan Zong, Cheng Peng, Zhaoli Guo, and Lian-Ping Wang. Designing correct fluid hydrodynamics on a rectangular grid using MRT lattice Boltzmann approach. Computers and Mathematics with Applications, 72(2):288–310, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115002564.

Zhao:2012:TEP

[ZPS^{+}12] Baokang Zhao, Wei Peng, Ziming Song, Jinshu Su, Chunqing Wu, Wanrong Yu, and Qiaolin Hu. Towards efficient and practical network coding in delay tolerant networks. Computers and Mathematics with Applications, 63(2):588–600, January 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111008522.

Zhao:2012:SSS

[ZPWY12] Dawei Zhao, Haipeng Peng, Cong Wang, and Yixian Yang. A secret sharing scheme with a short share realizing the (t, n) threshold and the adversary structure. Computers and Mathematics with Applications, 64(4):611–615, August 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111011382.

Zhang:2011:ACM

[ZQ11a] Hui-Li Zhang and Li-Juan Qin. An ancient Chinese mathematical algorithm and its application to nonlinear oscillators. Computers and Mathematics with Applications, 61(8):2071–2075, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110006577.

Zhang:2011:OTS

[ZQ11b] Quanxin Zhang and Fang Qiu. Oscillation theorems for second-order half-linear delay dynamic equations with damping on time scales. Computers and Mathematics with Applications, 62 (11):4185–4193, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111008662.

Zhu:2014:FVH

[ZQ14a] Jun Zhu and Jianxian Qiu. Finite volume Hermite WENO schemes for solving the Hamilton– Jacobi equations II: Unstructured meshes. Computers and Mathematics with Applications, 68 (10):1137–1150, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114004052.

Zia:2014:KFV

[ZQ14b] Saqib Zia and Shamsul Qamar. A kinetic flux-vector splitting method for single-phase and two-phase shallow flows. Computers and Mathematics with Applications, 67(6):1271–1288, April 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400039X.

Zhang:2016:SMC

[ZR16] Zhen Zhang and Weiqing Ren. Simulation of moving contact lines in twophase polymeric fluids. Computers and Mathematics with Applications, 72(4):1002–1012, August 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116303480.

Zdunek:2018:ILW

[ZR18a] Adam Zdunek and Waldemar Rachowicz. Inhomogeneous lossy waveguide mode analysis. Computers and Mathematics with Applications, 75(3):798–808, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306594.

Zdunek:2018:MFE

[ZR18b] Adam Zdunek and Waldemar Rachowicz. A mixed finite element formulation for slightly compressible finite elasticity with stiff fibre reinforcement. Two fibre families. Uniaxial tension formulation. Computers and Mathematics with Applications, 75(8):2607–2624, April 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117307836.

Zdunek:2018:MHO

[ZR18c] Adam Zdunek and Waldemar Rachowicz. A mixed higher order FEM for fully coupled compressible transversely isotropic finite hyperelasticity. Computers and Mathematics with Applications, 74(7):1727–1750, October 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122117301219.

Zegeye:2011:CTA

[ZRC11] H. Zegeye, M. Robdera, and B. Choudhary. Convergence theorems for asymptotically pseudocontractive mappings in the intermediate sense. Computers and Mathematics with Applications, 62(1):326– 332, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111004019.

Zdunek:2016:FFF

[ZRE16] Adam Zdunek, Waldemar Rachowicz, and Thomas Eriksson. A five-field finite element formulation for nearly inextensible and nearly incompressible finite hyperelasticity. Computers and Mathematics with Applications, 72(1):25–47, July 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302036.

Zegeye:2011:CMT

[ZS11a] H. Zegeye and N. Shahzad.

Convergence of Mann’s type iteration method for generalized asymptotically nonexpansive mappings. Computers and Mathematics with Applications, 62 (11):4007–4014, December 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007863.

Zhang:2011:ESFb

[ZS11b] Shuqin Zhang and Xinwei Su. The existence of a solution for a fractional differential equation with nonlinear boundary conditions considered using upper and lower solutions in reverse order. Computers and Mathematics with Applications, 62(3):1269–1274, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100160X.

Zhang:2013:SGN

[ZS13] Yadong Zhang and Dongyang Shi. Superconvergence of an H^{1}–Galerkin nonconforming mixed finite element method for a parabolic equation. Computers and Mathematics with Applications, 66 (11):2362–2375, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211300566X.

Zheng:2015:PFF

[ZS15] Qianqian Zheng and Jianwei Shen. Pattern formation in the FitzHugh– Nagumo model. Computers and Mathematics with Applications, 70 (5):1082–1097, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500317X.

Zhang:2016:PSL

[ZS16] Zhengyi Zhang and Ahmed H.

Sameh. A parallel sparse linear system solver based on Hermitian/skew-Hermitian splitting. Computers and Mathematics with Applications, 72(8):2000–2007, October 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116304709.

Zieniuk:2018:SBG

[ZS18] Eugeniusz Zieniuk and Krzysztof Szerszeń. A separation of the boundary geometry from the boundary functions in PIES for 3D problems modeled by the Navier– Lamé equation. Computers and Mathematics with Applications, 75(4):1067–1094, February 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117306946.

Zayed:2018:SOS

[ZSAN18] Elsayed M. E. Zayed, Reham M. A. Shohib, and Abdul-Ghani Al-Nowehy. Solitons and other solutions for higher-order NLS equation and quantum ZK equation using the extended simplest equation method. Computers and Mathematics with Applications, 76(9):2286–2303, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304474.

Zayed:2019:SDN

[ZSAN19] Elsayed M. E. Zayed, Reham. M. A. Shohib, and Abdul-Ghani Al-Nowehy. On solving the (3 + 1)-dimensional NLEQZK equation and the (3 + 1)-dimensional NLmZK equation using the extended simplest equation method. Computers and Mathematics with Applications, 78 (10):3390–3407, November 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302706.

Zhu:2010:NST

[ZSD10] Hongqing Zhu, Huazhong Shu, and Meiyu Ding. Numerical solutions of two-dimensional Burgers’ equations by discrete Adomian decomposition method. Computers and Mathematics with Applications, 60(3):840–848, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003883.

Sun:2010:CDS

[zSdZ10] Zhi zhong Sun and Dan dan Zhao. On the L_{∞} convergence of a difference scheme for coupled nonlinear Schrödinger equations. Computers and Mathematics with Applications, 59(10):3286–3300, May 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110001720.

Zhai:2011:EUS

[ZSH11] Chengbo Zhai, Ruipeng Song, and Qianqian Han. The existence and the uniqueness of symmetric positive solutions for a fourth-order boundary value problem. Computers and Mathematics with Applications, 62 (6):2639–2647, September 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111006663.

Zhao:2011:TFH

[ZSHL11] Yige Zhao, Shurong Sun, Zhenlai Han, and Qiuping Li. Theory of fractional hybrid differential equations. Computers and Mathematics with Applications, 62(3):1312–1324, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111001933.

Zhang:2019:GEB

[ZSL19] Quanguo Zhang, Hong-Rui Sun, and Yaning Li. Global existence and blow-up of solutions of the Cauchy problem for a time fractional diffusion system. Computers and Mathematics with Applications, 78(5):1357–1366, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119301385.

Zhang:2019:NTS

[ZSLZ19] Mengchen Zhang, Ming Shen, Fawang Liu, and Hongmei Zhang. A new time and spatial fractional heat conduction model for Maxwell nanofluid in porous medium. Computers and Mathematics with Applications, 78(5):1621–1636, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300112.

Zhang:2018:MAS

[ZSQ^{+}18] Yao-Ming Zhang, Fang-Ling Sun, Wen-Zhen Qu, Yan Gu, and Der-Liang Young. A meshless average source boundary node method for steadystate heat conduction in general anisotropic media. Computers and Mathematics with Applications, 75(5):1739–1755, March 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117307629.

Zhao:2010:IBV

[ZSS10] Yian Zhao, Guangxing Song, and Xiaoyan Sun. Integral boundary value problems with causal operators. Computers and Mathematics with Applications, 59(8):2768–2775, April 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110000817.

Zhang:2015:VVM

[ZSW15] Wenhuan Zhang, Baochang Shi, and Yihang Wang. 14-velocity and 18-velocity multiple-relaxation-time lattice Boltzmann models for three-dimensional incompressible flows. Computers and Mathematics with Applications, 69(9): 997–1019, May 2015. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115000966.

Zhang:2019:NTD

[ZSW19] Liang Zhang, Wan-Xia Shi, and Shuang-Ming Wang. A nonlocal and time-delayed reaction– diffusion eco-epidemiological predator-prey model. Computers and Mathematics with Applications, 77(9):2534–2552, May 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118307326.

Zhang:2014:OCD

[ZSY14] Shuhua Zhang, Shuyu Sun, and Hongtao Yang. Optimal convergence of discontinuous Galerkin methods for continuum modeling of supply chain networks. Computers and Mathematics with Applications, 68(6):681–691, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114003307.

Zhou:2019:PBP

[ZSY19] Huifang Zhou, Zhiqiang Sheng, and Guangwei Yuan. Physical-bound-preserving finite volume methods for the Nagumo equation on distorted meshes. Computers and Mathematics with Applications, 77(4):1055–1070, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306357.

Zeng:2017:SCP

[ZSZ17] Min-Li Zeng, Walker Paul Sevin, and Guo-Feng Zhang. On semi-convergence of the parameterized generalized MHSS method for singular complex linear systems. Computers and Mathematics with Applications, 73(8):1824–1833, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117301025.

Zhang:2018:RPO

[ZSZ18] Ruming Zhang, Jiguang Sun, and Chunxiong Zheng. Reconstruction of a penetrable obstacle in periodic waveguides. Computers and Mathematics with Applications, 74(11):2739–2751, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305205.

Zak:2013:NSC

[ZT13] Mohammad Khorsand Zak and Faezeh Toutounian. Nested splitting conjugate gradient method for matrix equation AXB = C and preconditioning. Computers and Mathematics with Applications, 66(3):269–278, September 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002861.

Zhang:2015:TLP

[ZT15] Tong Zhang and Zhenzhen Tao. Two level penalty finite element methods for the stationary incompressible magnetohydrodynamics problem. Computers and Mathematics with Applications, 70(10):2355– 2375, November 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115004174.

Zak:2016:SNS

[ZT16a] Mohammad Khorsand Zak and Faezeh Toutounian. A shifted nested splitting iterative method with applications to illposed problems and image restoration. Computers and Mathematics with Applications, 71(1):213– 223, January 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122115005362.

Zhong:2016:MPS

[ZT16b] Xiao-Jing Zhong and Chun-Lei Tang. Multiple positive solutions to a Kirchhoff type problem involving a critical nonlinearity. Computers and Mathematics with Applications, 72 (12):2865–2877, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305703.

Zhong:2018:GSS

[ZT18a] Xiao-Jing Zhong and Chun-Lei Tang. Ground state sign-changing solutions for a class of subcritical Choquard equations with a critical pure power nonlinearity in RN. Computers and Mathematics with Applications, 76(1):23–34, July 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118301937.

Zogheib:2018:MHS

[ZT18b] Bashar Zogheib and Emran Tohidi. Modal Hermite spectral collocation method for solving multi-dimensional hyperbolic telegraph equations. Computers and Mathematics with Applications, 75(10):3571–3588, May 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118300919.

Zhang:2019:BST

[ZT19] Qingshan Zhang and Weirun Tao. Boundedness and stabilization in a two-species chemotaxis system with signal absorption. Computers and Mathematics with Applications, 78(8):2672–2681, October 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119302068.

Zhang:2014:LMA

[ZTC14] Xueying Zhang, Haiyan Tian, and Wen Chen. Local method of approximate particular solutions for two-dimensional unsteady Burgers’ equations. Computers and Mathematics with Applications, 66(12):2425– 2432, January 2014. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122113006214.

Zhang:2011:HAM

[ZTH11] Xindong Zhang, Bo Tang, and Yinnian He. Homotopy analysis method for higher-order fractional integro-differential equations. Computers and Mathematics with Applications, 62(8):3194–3203, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100695X.

Zarmehi:2011:NSS

[ZTR11] Fatemeh Zarmehi, Ali Tavakoli, and Majid Rahimpour. On numerical stabilization in the solution of Saint-Venant equations using the finite element method. Computers and Mathematics with Applications, 62(4):1957–1968, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111005190.

Zhen:2016:SSC

[ZTSC16] Hui-Ling Zhen, Bo Tian, Ya Sun, and Jun Chai. Soliton solutions and chaotic motions for the (2 + 1)-dimensional Zakharov equations in a laser-induced plasma. Computers and Mathematics with Applications, 71(7):1337–1348, April 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S089812211630044X.

Zhong:2019:ESF

[ZTW^{+}19] Rui Zhong, Jinyuan Tang, Ailun Wang, Cijun Shuai, and Qingshan Wang. An exact solution for free vibration of cross-ply laminated composite cylindrical shells with elastic restraint ends. Computers and Mathematics with Applications, 77(3):641–661, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118305935.

Zhang:2019:NMM

[ZTY^{+}19] Z. Zhang, Z. J. Tan, X. X. Yao, C. P. Hu, P. Ge, Z. Y. Wan, J. Y. Li, and Q. Wu. Numerical methods for microstructural evolutions in laser additive manufacturing. Computers and Mathematics with Applications, 78(7):2296–2307, October 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118303766.

Zhang:2015:GSS

[ZTZ15] Wen Zhang, Xianhua Tang, and Jian Zhang. Ground state solutions for a diffusion system. Computers and Mathematics with Applications, 69(4):337–346, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114006099.

Zhang:2016:GSD

[ZTZ16a] Jian Zhang, Xianhua Tang, and Wen Zhang. Ground states for diffusion system with periodic and asymptotically periodic nonlinearity. Computers and Mathematics with Applications, 71(2):633–641, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005994.

Zhang:2016:IMR

[ZTZ16b] Wen Zhang, Xianhua Tang, and Jian Zhang. Infinitely many radial and non-radial solutions for a fractional Schrödinger equation. Computers and Mathematics with Applications, 71(3):737–747, February 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115006045.

Zhen:2014:DBS

[ZTZJ14] Hui-Ling Zhen, Bo Tian, Hui Zhong, and Yan Jiang. Dynamic behaviors and soliton solutions of the modified Zakharov– Kuznetsov equation in the electrical transmission line. Computers and Mathematics with Applications, 68(5):579–588, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211400282X.

Zhang:2011:EBA

[ZW11a] Wei Zhang and Xinwei Wang. Elastoplastic buckling analysis of thick rectangular plates by using the differential quadrature method. Computers and Mathematics with Applications, 61(1):44–61, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122110008187.

Zhang:2011:SRD

[ZW11b] Zhixin Zhang and Jiang Wei. Some results of the degenerate fractional differential system with delay. Computers and Mathematics with Applications, 62(3):1284–1291, August 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111002240.

Zhou:2011:CRS

[ZW11c] Lei Zhou and Wei-Zhi Wu. Characterization of rough set approximations in atanassov intuitionistic fuzzy set theory. Computers and Mathematics with Applications, 62(1):282– 296, July 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211100397X.

Zhou:2011:VPC

[ZW11d] Xin-Wei Zhou and Lin Wang. A variational principle for coupled nonlinear Schrödinger equations with variable coefficients and high nonlinearity. Computers and Mathematics with Applications, 61(8):2035–2038, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000636X.

Zheng:2016:BDB

[ZW16a] Jiashan Zheng and Yifu Wang. Boundedness and decay behavior in a higher-dimensional quasilinear chemotaxis system with nonlinear logistic source. Computers and Mathematics with Applications, 72 (10):2604–2619, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305326.

Zheng:2016:BSQ

[ZW16b] Jiashan Zheng and Yifu Wang. Boundedness of solutions to a quasilinear chemotaxis-haptotaxis model. Computers and Mathematics with Applications, 71(9):1898–1909, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

com/science/article/
pii/S0898122116301304.

Zhang:2010:MUM

[ZWC10] Jingwei Zhang, Layne T.

Watson, and Yang Cao. A modified uniformization method for the solution of the chemical master equation. Computers and Mathematics with Applications, 59(1):573–584, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109003745.

Zheng:2011:UFH

[ZWG11] Lian-Cun Zheng, Kang-Ning Wang, and Ying-Tao Gao. Unsteady flow and heat transfer of a generalized Maxwell fluid due to a hyperbolic sine accelerating plate. Computers and Mathematics with Applications, 61(8):2209–2212, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007030.

Zhao:2019:ALT

[ZWH^{+}19] Yanmin Zhao, Fenling Wang, Xiaohan Hu, Zhengguang Shi, and Yifa Tang. Anisotropic linear triangle finite element approximation for multiterm time-fractional mixed diffusion and diffusionwave equations with variable coefficient on 2D bounded domain. Computers and Mathematics with Applications, 78 (5):1705–1719, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118306813.

Zhong:2011:MIP

[ZWJ^{+}11] Ping Zhong, Shile Wang, Ye Jin, Xinxing Tu, and Nian Luo. A method of image preprocessing based on nonlinear diffusion and information extraction. Computers and Mathematics with Applications, 61(8):2132–2137, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006814.

Zhang:2011:TAL

[ZWL11] Qing-Ming Zhang, Xiao-Cheng Wang, and Hua Liu. Theoretical analysis of load-carrying characteristics of a simply supported dual-duct rectangular plate under moving loads. Computers and Mathematics with Applications, 61(8):2306–2312, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007571.

Zhang:2018:EMS

[ZWLZ18] Fengxia Zhang, Musheng Wei, Ying Li, and Jianli Zhao. An efficient method for special least squares solution of the complex matrix equation (AXB, CXD) = (E, F). Computers and Mathematics with Applications, 76(8):2001–2010, October 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118304152.

Zhao:2016:GLN

[ZWMD16] Yanxiang Zhao, Jiakou Wang, Yanping Ma, and Qiang Du. Generalized local and nonlocal master equations for some stochastic processes. Computers and Mathematics with Applications, 71(11):2497–2512, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004848.

Zhou:2013:HES

[ZWW13] Fen Zhou, Ke Wu, and Xian Wu. High energy solutions of systems of Kirchhoff-type equations on R^{N}. Computers and Mathematics with Applications, 66(7):1299–1305, October 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004793.

Zhang:2019:ESP

[ZWY19] Zujin Zhang, Weihua Wang, and Xian Yang. An extension and simpler proof of Berselli– Córdoba’s geometric regularity condition for the Navier–Stokes system. Computers and Mathematics with Applications, 77(3):765–769, February 1, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306151.

Zhang:2016:DRD

[ZWZ16] Liang Zhang, Zhi-Cheng Wang, and Yan Zhang. Dynamics of a reaction– diffusion waterborne pathogen model with direct and indirect transmission. Computers and Mathematics with Applications, 72(1): 202–215, July 2016. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

http://www.sciencedirect.
com/science/article/
pii/S0898122116302395.

Zhang:2011:UMF

[ZX11] Xiao-Bin Zhang and Jun-Feng Xu. Uniqueness of meromorphic functions sharing a small function and its applications. Computers and Mathematics with Applications, 61(3):722–730, February 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110009272.

Zhu:2013:LSS

[ZXW13] Hong Zhu, Yunhai Xiao, and Soon-Yi Wu. Large sparse signal recovery by conjugate gradient algorithm based on smoothing technique. Computers and Mathematics with Applications, 66(1):24– 32, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113002319.

Zhang:2017:EMS

[ZXZ17] Hui Zhang, Junxiang Xu, and Fubao Zhang. Existence and multiplicity of solutions for a generalized Choquard equation. Computers and Mathematics with Applications, 73(8):1803–1814, April 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301050.

Zhang:2010:PEF

[ZY10a] Ji-Long Zhang and Lian-Zhong Yang. A power of an entire function sharing one value with its derivative. Computers and Mathematics with Applications, 60(7):2153–2160, October 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211000564X.

Zhou:2010:VIM

[ZY10b] Xin-Wei Zhou and Li Yao.

The variational iteration method for Cauchy problems. Computers and Mathematics with Applications, 60(3):756–760, August 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003718.

Zhou:2011:IPE

[ZY11] Jianwei Zhou and Danping Yang. An improved a posteriori error estimate for the Galerkin spectral method in one dimension. Computers and Mathematics with Applications, 61(2):334–340, January 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110008576.

Ziyadi:2013:PIM

[ZY13] Najat Ziyadi and Abdul-Aziz Yakubu. Predatorinduced and mating limitation-induced Allee effects in a discrete-time SIMS epidemic model. Computers and Mathematics with Applications, 66(11):2196–2210, December 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122113004835.

Zhang:2015:LBM

[ZY15a] Jianying Zhang and Guangwu Yan. Lattice Boltzmann model for complex Ginzburg–Landau equation in curvilinear coordinates. Computers and Mathematics with Applications, 70 (12):2904–2919, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115004836.

Zhang:2015:LBS

[ZY15b] Jianying Zhang and Guangwu Yan. Lattice Boltzmann simulation of pattern formation under crossdiffusion. Computers and Mathematics with Applications, 69(3):157–169, February 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122114005847.

Zhou:2015:UBE

[ZY15c] Jun Zhou and Di Yang.

Upper bound estimate for the blow-up time of an evolution m-Laplace equation involving variable source and positive initial energy. Computers and Mathematics with Applications, 69(12):1463–1469, June 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001765.

Zhang:2017:CGL

[ZY17a] Huamin Zhang and Hongcai Yin. Conjugate gradient least squares algorithm for solving the generalized coupled Sylvester matrix equations. Computers and Mathematics with Applications, 73(12):2529–2547, June 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301803.

Zhang:2017:NPG

[ZY17b] Huamin Zhang and Hongcai Yin. New proof of the gradient-based iterative algorithm for the Sylvester conjugate matrix equation. Computers and Mathematics with Applications, 74 (12):3260–3270, December 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117305096.

Zhang:2017:AFO

[ZY17c] Xiaojuan Zhang and Wanzhou Ye. An adaptive fourth-order partial differential equation for image denoising. Computers and Mathematics with Applications, 74 (10):2529–2545, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304650.

Zhong:2017:FTB

[ZY17d] Penghong Zhong and Ganshan Yang. Finite time blowup of multidimensional inhomogeneous isotropic Landau– Lifshitz equation on a hyperbolic space. Computers and Mathematics with Applications, 73(3):433–449, February 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122116306708.

Zhang:2019:SAN

[ZY19] Houchao Zhang and Xiaoxia Yang. Superconvergence analysis of nonconforming finite element method for time-fractional nonlinear parabolic equations on anisotropic meshes. Computers and Mathematics with Applications, 77(10):2707–2724, May 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300069.

Zhang:2010:OBE

[ZYG10] Quanxin Zhang, Jurang Yan, and Li Gao. Oscillation behavior of even-order nonlinear neutral differential equations with variable coefficients. Computers and Mathematics with Applications, 59(1):426–430, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004015.

Zhai:2010:LSR

[ZYS10] Mingqing Zhai, Guanglong Yu, and Jinlong Shu. The Laplacian spectral radius of bicyclic graphs with a given girth. Computers and Mathematics with Applications, 59(1):376–381, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122109004416.

Zhang:2017:DRW

[ZYSY17] Ruigang Zhang, Liangui Yang, Jian Song, and Hongli Yang. (2 + 1) dimensional Rossby waves with complete Coriolis force and its solution by homotopy perturbation method. Computers and Mathematics with Applications, 73(9):1996–2003, May 1, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117301153.

Zhang:2016:PIB

[ZYT^{+}16] Hao Zhang, Haizhuan Yuan, F. Xavier Trias, Aibing Yu, Yuanqiang Tan, and Assensi Oliva. Particulate immersed boundary method for complex fluid-particle interaction problems with heat transfer. Computers and Mathematics with Applications, 71(1):391–407, January 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115005714.

Zhang:2015:PPH

[ZYW15] Wei-Hong Zhang, Ai-Li Yang, and Yu-Jiang Wu. Parameterized preconditioned Hermitian and skew-Hermitian splitting iteration method for a class of linear matrix equations. Computers and Mathematics with Applications, 70 (6):1357–1367, September 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S089812211500351X.

Zhou:2017:TMB

[ZYW17] Sheng-Wei Zhou, Ai-Li Yang, and Yu-Jiang Wu. Two modified block-triangular splitting preconditioners for generalized saddle-point problems. Computers and Mathematics with Applications, 74 (6):1176–1197, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117303528.

Zhai:2017:WGF

[ZYWZ17] Qilong Zhai, Xiu Ye, Ruishu Wang, and Ran Zhang. A weak Galerkin finite element scheme with boundary continuity for second-order elliptic problems. Computers and Mathematics with Applications, 74 (10):2243–2252, November 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122117304261.

Zhongxue:2010:NCI

[ZYY10] Lü Zhongxue, Gao Youcai, and Wei Yuxiang. Note on the Carleman’s inequality and Hardy’s inequality. Computers and Mathematics with Applications, 59(1):94–97, January 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122109006579.

Zhao:2011:NNM

[ZYZ11] Guozhong Zhao, Xijun Yu, and Rongpei Zhang. The new numerical method for solving the system of two-dimensional Burgers’ equations. Computers and Mathematics with Applications, 62(8):3279–3291, October 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122111007176.

Zhang:2016:AKT

[ZYZ^{+}16] Liangqi Zhang, Shiliang Yang, Zhong Zeng, Liping Yao, and Jia Wei Chew. Alternative kinetic theory based lattice Boltzmann model for incompressible axisymmetric flows. Computers and Mathematics with Applications, 72(11):2751–2772, December 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116305570.

Zhang:2017:CSA

[ZYZ^{+}17] Liangqi Zhang, Shiliang Yang, Zhong Zeng, Jie Chen, Lingquan Wang, and Jia Wei Chew. A comparative study of the axisymmetric lattice Boltzmann models under the incompressible limit. Computers and Mathematics with Applications, 74(4):817–841, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303334.

Zhang:2018:AIK

[ZYZC18] Liangqi Zhang, Shiliang Yang, Zhong Zeng, and Jia Wei Chew. An alternative implementation of the kinetic theory based axisymmetric lattice Boltzmann model. Computers and Mathematics with Applications, 76(6):1388–1407, September 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122118303560.

Zhao:2010:FOC

[ZZ10a] Yanjuan Zhao and Yuanguo Zhu. Fuzzy optimal control of linear quadratic models. Computers and Mathematics with Applications, 60(1):67–73, July 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110003019.

Zhou:2010:GBD

[ZZ10b] Shuangshuang Zhou and Sining Zheng. Global blow-up in a degenerate and strongly coupled parabolic system with localized sources. Computers and Mathematics with Applications, 60 (9):2564–2571, November 2010. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110006504.

Zhang:2011:PPS

[ZZ11a] Fengqin Zhang and Chongwu Zheng. Positive periodic solutions for the neutral ratio-dependent predator–prey model. Computers and Mathematics with Applications, 61(8):2221–2226, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110007078.

Zhang:2011:EFM

[ZZ11b] Sheng Zhang and Hong-Qing Zhang. An Expfunction method for a new N-soliton solutions with arbitrary functions of a (2 + 1)- dimensional vcBK system. Computers and Mathematics with Applications, 61(8):1923–1930, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122110005341.

Zhou:2014:SCA

[ZZ14] Lijuan Zhou and Naimin Zhang. Semi-convergence analysis of GMSSOR methods for singular saddle point problems. Computers and Mathematics with Applications, 68(5):596–605, September 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114002867.

Zeng:2015:PRB

[ZZ15a] Min-Li Zeng and Guo-Feng Zhang. Parameterized rotated block preconditioning techniques for block two-by-two systems with application to complex linear systems. Computers and Mathematics with Applications, 70(12):2946–2957, December 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115004940.

Zhang:2015:ESS

[ZZ15b] Qihu Zhang and Chunshan Zhao. Existence of strong solutions of a p(x)-Laplacian Dirichlet problem without the Ambrosetti–Rabinowitz condition. Computers and Mathematics with Applications, 69(1):1–12, January 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114005276.

Zhang:2016:DAD

[ZZ16a] Xuebing Zhang and Hongyong Zhao. Dynamics analysis of a delayed diffusive predator–prey system with non-smooth continuous threshold harvesting. Computers and Mathematics with Applications, 72(5):1402–1417, September 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.

                                        com/science/article/
                                        pii/S0898122116303789.

Zhang:2016:RCN

[ZZ16b] Zujin Zhang and Yong Zhou. On regularity criteria for the 3D Navier– Stokes equations involving the ratio of the vorticity and the velocity. Computers and Mathematics with Applications, 72(9):2311–2314, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304850.

Zhu:2016:CPA

[ZZ16c] Kaixuan Zhu and Feng Zhou. Continuity and pullback attractors for a non-autonomous reaction– diffusion equation in R^{N}. Computers and Mathematics with Applications, 71(10):2089–2105, May 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116301729.

Zhu:2017:SRD

[ZZ17] Cheng-Cheng Zhu and Jiang Zhu. Stability of a reaction–diffusion alcohol model with the impact of tax policy. Computers and Mathematics with Applications, 74(4):613–633, August 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117302845.

Zhang:2018:NMS

[ZZ18a] Jianhua Zhang and Jing Zhao. A new matrix splitting preconditioner for generalized saddle point problems. Computers and Mathematics with Applications, 75(1):153–160, January 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117305291.

Zhang:2018:GRN

[ZZ18b] Peixin Zhang and Mingxuan Zhu. Global regularity of 3D nonhomogeneous incompressible magneto-micropolar system with the densitydependent viscosity. Computers and Mathematics with Applications, 76 (9):2304–2314, November 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S089812211830467X.

Zhou:2018:EMR

[ZZ18c] Yong Zhou and Lu Zhang. Existence and multiplicity results of homoclinic solutions for fractional Hamiltonian systems. Computers and Mathematics with Applications, 73(6):1325–1345, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116302346.

Zeng:2019:NST

[ZZ19] Xiang-Chen Zeng and Song-Ping Zhu. A new simple tree approach for the Heston’s stochastic volatility model. Computers and Mathematics with Applications, 78 (6):1993–2010, September 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122119301555.

Zhuo:2013:FMLa

[ZZC13a] Congshan Zhuo, Chengwen Zhong, and Jun Cao. Filter-matrix lattice Boltzmann simulation of lid-driven deepcavity flows, part Isteady flows. Computers and Mathematics with Applications, 65(12):1863–1882, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001429.

Zhuo:2013:FMLb

[ZZC13b] Congshan Zhuo, Chengwen Zhong, and Jun Cao. Filter-matrix lattice Boltzmann simulation of lid-driven deepcavity flows, part II — flow bifurcation. Computers and Mathematics with Applications, 65(12):1883–1893, August 2013. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122113001533.

Zhang:2018:STC

[ZZF18] Xuebing Zhang, Hongyong Zhao, and Zhaosheng Feng. Spatio-temporal complexity of a delayed diffusive model for plant invasion. Computers and Mathematics with Applications, 76(11– 12):2575–2612, December 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118304899.

Zhang:2019:SPE

[ZZG19] Rangrang Zhang, Guoli Zhou, and Boling Guo. Stochastic 2D primitive equations: Central limit theorem and moderate deviation principle. Computers and Mathematics with Applications, 77(4):928–946, February 15, 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122118306229.

Zhang:2012:DMM

[ZZHF12] Jianhua Zhang, Wenfang Zhang, Guolian Hou, and Fang Fang. Dynamic modeling and multivariable control of organic Rankine cycles in waste heat utilizing processes. Computers and Mathematics with Applications, 64(5):908–921, September 2012. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.

com/science/article/
pii/S0898122112000752.

Zhu:2015:IFV

[ZZL15] Ling Zhu, Zhiyue Zhang, and Zhilin Li. An immersed finite volume element method for 2D PDEs with discontinuous coefficients and nonhomogeneous jump conditions. Computers and Mathematics with Applications, 70(2):89–103, July 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122115001819.

Zhao:2018:MSG

[ZZL18a] Yangzhang Zhao, Qi Zhang, and Jeremy Levesley. Multilevel sparse grids collocation for linear partial differential equations, with tensor product smooth basis functions. Computers and Mathematics with Applications, 75(3):883–899, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306624.

Zhao:2018:CSF

[ZZL^{+}18b] Yanmin Zhao, Yadong Zhang, F. Liu, I. Turner, Yifa Tang, and V. Anh. Convergence and superconvergence of a fullydiscrete scheme for multiterm time fractional diffusion equations. Computers and Mathematics with Applications, 73(6):1087–1099, March 15, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122116302632.

Zhang:2019:FIS

[ZZL19] Min Zhang, Guo-Feng Zhang, and Li-Dan Liao. Fast iterative solvers and simulation for the space fractional Ginzburg–Landau equations. Computers and Mathematics with Applications, 78 (5):1793–1800, September 2019. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122119300562.

Zhang:2018:ANS

[ZZLB18] Yan Zhang, Haojie Zhao, Fawang Liu, and Yu Bai. Analytical and numerical solutions of the unsteady 2D flow of MHD fractional Maxwell fluid induced by variable pressure gradient. Computers and Mathematics with Applications, 75(3):965–980, February 1, 2018. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117306934.

Zhang:2017:FSE

[ZZM17] Wen Zhang, Jian Zhang, and Heilong Mi. On fractional Schrödinger equation with periodic and asymptotically periodic conditions. Computers and Mathematics with Applications, 74 (6):1321–1332, September 15, 2017. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122117303656.

Zhong:2011:CCS

[ZZT11] Ting Zhong, Jing-Jing Zhang, and Liang Tang. A class of Cantor sets associated with the regular continued fractions. Computers and Mathematics with Applications, 61(8):2251–2255, April 2011. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122110007248.

Zhu:2015:SBA

[ZZW15] Linhe Zhu, Hongyong Zhao, and Xiaoming Wang. Stability and bifurcation analysis in a delayed reaction–diffusion malware propagation model. Computers and Mathematics with Applications, 69(8):852–875, April 2015. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115000504.

Zhang:2016:NMC

[ZZWG16] Guannan Zhang, Weidong Zhao, Clayton Webster, and Max Gunzburger. Numerical methods for a class of nonlocal diffusion problems with the use of backward SDEs. Computers and Mathematics with Applications, 71(11):2479– 2496, June 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122115005179.

Zhang:2014:ALB

[ZZX^{+}14a] Liangqi Zhang, Zhong Zeng, Haiqiong Xie, Xutang Tao, Yongxiang Zhang, Yiyu Lu, Akira Yoshikawa, and Yoshiyuki Kawazoe. An alternative lattice Boltzmann model for three-dimensional incompressible flow. Computers and Mathematics with Applications, 68 (10):1107–1122, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114003691.

Zhang:2014:CSL

[ZZX^{+}14b] Liangqi Zhang, Zhong Zeng, Haiqiong Xie, Yongxiang Zhang, Yiyu Lu, Akira Yoshikawa, Hiroshi Mizuseki, and Yoshiyuki Kawazoe. A comparative study of lattice Boltzmann models for incompressible flow. Computers and Mathematics with Applications, 68(10):1446–1466, November 2014. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

//www.sciencedirect.
com/science/article/
pii/S0898122114004581.

Zhou:2016:GSA

[ZZX16] Jianwei Zhou, Juan Zhang, and Xiaoqing Xing. Galerkin spectral approximations for optimal control problems governed by the fourth order equation with an integral constraint on state. Computers and Mathematics with Applications, 72 (10):2549–2561, November 2016. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL http:

                                        //www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122116304631.

Zhang:2012:RRM

[ZZXY12] Yong Zhang, Weiguo Zhang, Weijun Xu, and Xingyu Yang. Riskreward models for online leasing of depreciable equipment. Computers and Mathematics with Applications, 63(1):167– 174, January 2012. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/
                                        pii/S0898122111009710.